A Selected Bibliography of Publications by, and about, Albert Einstein

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 April 2024 Version 4.350

Title word cross-reference

#10.2 [Bel98].

$1 [TT70, Duf46]. $10 [Kuh54]. $10.00 [Hea13]. $11.50 [Jam70]. $12.50 [Bla76a, Bla76b]. $138.00 [Wal02b]. $14.95 [Har79d]. $15 [Ano54b]. $15.00 [Jan73]. $19.50 [War96b]. $19.95 [Ste22]. 2 [GB15, SSS^{+}98]. $2.00 [Ano50a]. $2.50 [Kae38]. $2.9 [Bar18b]. $20.00 [Ham11a, Tre88]. $22.50 [Pri99]. $22.95 [For01, Kar05, vB02]. $24 [Eng17b, Ree15d, Woo06]. $24.95 [Bed01a, GRC18]. $25 [Ree17d, Cah90b, Nah01]. $25.00 [Dys15]. $26 [Rus17]. $26.95 [Gor02, Ivr99]. $27 [Kai81]. $27.00 [Sto17]. $27.95 [Hes04, Kar05, vB01]. $28 [Ano45b]. $28.00 [Myr21b]. $29.95 [Kle14a, Sta21, Ste22]. 3 [Ano13c]. 3 + 1 [Alc08]. $3.50 [Hen82b]. $30.00 [Bru03c, Kra02, Sta99b]. $32 [Weg08]. $32.00 [Gor07a, Sch07c]. $35.00 [Par06]. $36.95 [Sta99b]. $38 [PJMS11]. $38.00 [Mar10]. $39 [Kai81]. $39.95 [WM08]. $40.00 [For05]. $42 [Boy88]. $44.95 [Nau17, Hoc93].

$45.00 [Pri99]. $49.95 [Alf09]. $5 [Dac20b]. $51 [Smi02]. $55.00 [Noe11]. $59.50 [Goo85]. $59.95 [Pec04]. $6 [Ano60, Hie68]. $60.00 [For01]. $69 [Gly90]. $69.95 [Hoh13]. $7.50 [Lin60]. $74.50 [Sci80]. $74.95 [Ach13]. $78.00 [Vog20]. $80 [Ree17c]. $85 [Cah90b]. $85.00 [Bru06]. $98 [PJMS11]. = [GB08]. = mc^{2} [For80a]. ^{(1)} [Aga07]. ^{235} [Ree17a]. ^{239} [Ree17a]. ^{3} [Leg04]. A [All90, CM67, Lid67, Tur66, Tur67]. B^{(3)} [Eva95]. • [Ano05b]. · [dB65]. E = [GB06]. E = Mc^{2} [Mer11, Ano79d, Ari14, Bal11b, Bal11c, B^{+}06, Ber98, Bod00, Bod05a, Bod05b, Bou13, CF09, Ein46a, Ein91a, Ein07a, Flo98, Flo05, Flo06, Jan03, KN19, Kra06, LD08, ME14a, Moo02, Oha09, Oha12a, Oha12b, Raf17, Ros12, Vog20, vB05]. e^{+}e^{−} → ΛΛ̄ → π^{−}pπ^{+}p̄ [Tör81]. E_{0} = mc^{2} [Hec11]. \frac{1}{2} [Sha14]. G [Duc11, GAHS96, LHF^{+}98, Mos13a, Mos13b, QPSD13, TLW^{+}10, VT99, GHK^{+}06, GHK^{+}07]. ≫ [Mas02, Seg91]. hν = (1/2)mv^{2} [Kir80]. k ln n! [Ste49]. Λ [Lah20, Nov18, Tur66, Lon38a]. ≪ [Mas02, Seg91]. S [Cos87a]. U^{238} [Tur46].

-Absorption [Ein16m]. -c [Ryc92]. -Matrix [Cos87a]. -Phenomenon [Lon38a].

// [EWJ^{+}25]. //doi.org/10.1088/0031 [Ree16a]. //doi.org/10.1088/0031-8949/91/11/113002 [Ree16a]. /Brochure [Ano79g]. /Celebration [Ano79g]. /Deutsche [Ano79g]. /Einstein [Ano79g]. /Feier [Ano79g]. /German [Ano79g]. /Hubert [Ano79g]. /Max [vBL96a, vBL96b]. /Rath [GRE^{+}01].

0 [Bon94, Dix00, Gor02, Har02b, Kar05, Kra02, Kri08, Mau09, Pec04, Sch07c, Smi02, Sto17, Tur05, Tur06a, War03a, Weg08, Woo06]. 0-19-850594-9 [Har02b, War03a]. 0-19-853907-X [Red84b]. 0-309-08408-3 [Hes04]. 0-309-08998-0 [Kar05]. 0-393-05165-X [Kar05]. 0-395-95563-7 [Nah01]. 0-465-09293-4 [Woo06]. 0-471-11459-6 [Sta99b]. 0-521-26584-3 [Rec86]. 0-521-40386-3 [Hoc93]. 0-571-16744-6 [Bon94]. 0-674-01519-3 [Tur06a]. 0-691-00099-9 [War96b]. 0-691-00699-7 [Mas01, Gor02]. 0-691-05939-X [Hen00b]. 0-691-08299-5 [Kil84b]. 0-691-08407-6 [Hen88b]. 0-7167-3103-7 [Kra02, Tur05]. 0-8047-2824-0 [Smi02]. 0-8176-4263-3 [Pec04]. 0-85224-458-4 [Hen86a]. 0-88135-283-7 [WM08]. 0-88135-288-8 [For05]. 0-918102-15-4 [Tre88]. 000-year [Mao07]. 0031-8949 [Ree16a]. 0456 [HTS^{+}88].

1 [B^{+}79, Bro02, BL07a, Dix00, Gre19a, JNR^{+}07b, Myr21c, Per10, Pes13, Ree05a, Ren21, Shi75, vL19, vL21, vL52, vL56a, vL65]. 1-900405-04-0 [Dix00]. 1.5 [GRE^{+}01]. 100 [Gre15a, JS87]. 100th [Ano79h, Ano79g, B^{+}79, Bro78a, CGT11, EHM^{+}79, Har79c, KRW05, Lam79, LFKW06, Rec02, RH02, SKD^{+}87, She16b, Spe79, Tre79a, Wis79]. 109 [Con05]. 11 [Ein05s, Kil95]. 11.-16 [MK56]. 110th [SS97]. 112 [FW07]. 12.14.1930/Albert [Ano69c]. 125th [Wie06]. 12th [B^{+}79, Sch28]. 13.-14 [Mag95]. 13.80 [Hei53a]. 1346 [LCL^{+}90]. 14 [Bre79, Ein05-40, HY86]. 14.12.1930/Albert [Ano69c]. 14th [Ano93c, Sch28]. 15 [Red84b]. 15.99pb/$69/95hb [Jon01]. 150 [Ros86]. 150th [Jac08]. 16 [ČA99, Ein29h, GHK^{+}96]. 16. [Dix00]. 1663 [Jet77, Jet07]. 17 [Ein05-41, Ein05-37, Ein05-35, Kil98]. 175 [Ein79e, Ein85e]. 18 [Ein05v]. 1800-1905 [Tha77]. 1801 [vS21a, vS21b]. 1887 [Leg15]. 1896 [Hun01]. 19 [Ein79e, Ein85e, Ein05r, Ein05-42, Kra06]. 19-175 [Ein79e, Ein85e]. 1900 [B^{+}01]. 1901 [Enk18]. 1901-1950 [dVNHdVH53]. 1902 [Bal91, Gol87, Hen88a, Ker89b, Loe88a, Mcc87, Pye89, Smi89b, Smi89a, Yan87, Loe88b]. 1903 [Adl03b]. 1905’ [Lil86, Ber06c, Rig05a]. 1908 [Jar08a]. 1909 [Har91, Cah90b, Cah90a, Ram90, Str90]. 1910 [Kle75b]. 1911 [Ein10b, LdB12b, Meh75c, Pye95, Str11a, War95]. 1912 [RS96]. 1913 [Jan07, RA13]. 1914 [Wei12s, Har96a, Har96b, Pye95, Sie97, War96a, Ano99b]. 1915 [Jen01, Kie82, Lin15, Van04, Wei12s, Wei12t, Wei18a, Mat86b, Rei86]. 1916 [Wei12t, Wei12u]. 1916-1955 [Kle70b, Mar73]. 1917 [Pye97, Wal99, Ano99c, Sta00]. 1918 [de 15, dP17, Row00]. 1919 [Bau20, DED20, Edd19, SE79, SW12]. 1919-1925 [SW12]. 1920 [Ber07a, BSKS04a, BSKS04b, BSR^{+}06a, BSR^{+}06b, Jüt21, Lac08, Pla20b, Hol05b, Kev05]. 1920er [Waz09]. 1920s [D’A97, Hol81b, Oha19, Waz09, Nav15, Waz14, Hen16]. 1921 [Ano23a, Bad15, BRS^{+}09a, BRS^{+}09b, Ein21l, EA22, Ein22p, EA23, Pye06, Don70]. 1922 [HY86, LLH^{+}80, Waz10]. 1923 [BIRS12a, BIRS12b, Ros18, Bad15, Ano26a]. 1924 [Mil24]. 1925 [AMS04, BIR^{+}15a, BIR^{+}15b, Ein29h, SW12, Ser88]. 1927 [BIJ^{+}18]. 1928 [Sch28]. 1929 [BRI^{+}21a, HEB^{+}80, HvMW79]. 1930 [Cho10a]. 1930s [Hoc83, Tur06b]. 1931 [Ein31a]. 1932 [Ein31b, For80a, Hei32, Sim06, Ano84a]. 1933 [Ano13a, For80a, HW78, Pye80]. 1933/Chadwick [Ano51a]. 1934-1984 [A^{+}84]. 1935/1936 [Ano51a]. 1938 [Fer38, Fer39, Fer70, Kra88]. 1939 [Ein91b, OSWR86, vMHW85]. 1942 [Hei95, KW93]. 1944 [Roe93]. 1945 [Ano00g, BC95, Fre08, GHK^{+}96, Ham97, Pac72, Pac81, Pau45, Pau47b, SS65, Sza84, WY05, Wei95, YJ95]. 1946 [CSW97b, Roe93, dP94, P.50]. 1948 [Eme48, Ein48g, Hah81, T– D82b]. 1949 [Kri08, Krö53, PvM93, Shi49, vM93]. 1950 [Coh54, Coh54]. 1950s [BDH89, Yam09]. 1952 [vM96]. 1954 [vM99b]. 1954/Mladjenovic [ARK^{+}99]. 1955 [Ano93c, Bre79, Whe80, Whi55b, For70a, Kle73]. 1956 [vM01]. 1958 [Enz05, vM05c]. 1960s [ARK^{+}99, HBRD97]. 1960s/Frenkel [ARK^{+}99]. 1961 [Ano61, Hei61b, Duv64]. 1961/Max [Ano61]. 1961/Preis [Ano61]. 1963 [Wig64b, Wig64c, Wig64a, Wig67a]. 1964 [Tow65]. 1965 [Fey64]. 1966 [MB66]. 1967 [Bet67]. 1969 [Ano69a]. 1970 [Ihd74, Hah73]. 1970s [HBRD97]. 1971 [Ihd74]. 1972 [Meh73b]. 1978 [Ano78, Opp79a]. 1979 [BS81, B^{+}79, EO79, Lam79, NHP^{+}79, Wig79, Wis79, Cao06]. 1980 [Dür80, Sch83b]. 1981 [Ano81d, BCS82]. 1983 [BS83b, Ber84a]. 1984 [Kri01]. 1985 [BPS87, HMW24]. 1986 [Gly90, HS89b]. 1987 [Ano88a, Ano88b]. 1988 [AS91, BB89, EK92, MTVB88, Rig00]. 1988/Green [MTVB88]. 1988/Grünes [MTVB88]. 1989 [HLZJ91, L.89]. 1991 [Est06c, SW94]. 1992 [Ano00g, Tra00]. 1993 [BCEP94]. 1994 [Sto96]. 1995 [Flo08, GRRS99b, RMR96]. 1996 [Ano97a]. 19th [Kra88].

2 [Ach13, Ano98, B.50, Ber07a, BL07b, Eng17b, Gre19b, JRN^{+}07b, Koj08, Pla20b, Ree11a, Ree17d, Rei50, UM86, Ste90]. 2-8 [JRG02]. 2.0 [Ein95b]. 2.6.1920 [Pla20a]. 20 [Ein05n, Ein05-43, Flo06, Hen86a]. 2000 [JRG02]. 2001 [BF02]. 2002 [Bar05f]. 2004 [KK05]. 2005 [Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, BL09, DDDR06]. 2006 [BBL08, Ber06b]. 2007 [Cli09a, JP08]. 2008 [Cli09b, Dys09]. 2010 [HL10]. 2011 [Bad11]. 2013 [Pap15]. 2016 [Sch17a]. 20th [Ano05c, Cea97, Cea98, ECEP11, Meh73b, Sch96a, WSG00, Bre97, Cat98, Gin01, Has04, Kai02, Ren08, Rog06]. 21 [Ano69a, Ein05-36]. 21-year [Eds82, Wad81b]. 211 [JRT^{+}91]. 21st [FJTS05, Neu06, GHS08, GHS18]. 22 [Ein05l, Ein05-32, Ein05p]. 22.50 [Rec86]. 23 [BPS87, Ein05k, Ein05-33, Ein05-34]. 24.70 [Kil84b]. 24.80 [Kle70b, Kle70a]. 24th [Ein53l]. 25 [NHP^{+}79]. 25.00/$35.00hb [Ber00]. 250th [Aĭn77]. 26 [EL05c, EL05d]. 27 [EL05b]. 272 [Fri73]. 28 [Ein05g, EL05a]. 288pp [Jon01]. 29 [BF02, Pos11]. 29th [Edd19]. 2nd [EK92].

3 [Ano13a, BL07d, Cah04, Hea13, Pec04, Ree15d, Ree16c, RSS^{+}07a, Row06a, Row06b, Sta21, Tur06a]. 3-540-90642-8 [Hen86a]. 3-540-90667-3 [Hen86a]. 3-540-90674-6 [Hen86a]. 3-540-90675-4 [Hen86a]. 3-540-90680-0 [Hen86a]. 3-938016-04-3 [Row06a, Row06b]. 3-Minute [Par11]. 3.85 [Bro72]. 30 [LdB12b]. 31.1.1930 [Ano69c]. 315pp [Pec04]. 33 [EH05b, EH05c, Ein05-29]. 34 [Ein05f, Ein05i, Ein05m, Ein05j, Ein05q]. 35 [Ein05t, Ein05-31]. 364pp [Ber00]. 37 [Ein86c, Ein05-30]. 38 [Ein05d, Ein05x, Ein05y, Ein05-28, Ein05-44]. 39 [Ein05e, Kan92b].

4 [Ano11b, BL07c, Ein05u, Mar10, RSS^{+}07b, Woo06]. 4-Vectors [Wal07]. 4.0 [BHJ^{+}00]. 4.0/Buchtipp [BHJ^{+}00]. 4/Pauli [OSWR86]. 40 [ES05a]. 400th [EVB^{+}72]. 410th [Ein53k]. 44 [EF05]. ’45 [Ree06]. 47 [Ein05c]. 49 [Ein05o]. 49/Atmanspacher/Primas [MGG^{+}95]. 4th [Bor50, Ein53n].

5 [Del11a, Kle14a, Lew79, Myr21b]. 50 [Kai81]. 509th [KK05]. 50th [Ano81d, DS83a, Sti05]. 51 [Ein05-38, Jac11]. 55 [BB69, Ein05-27]. 59 [The12]. 5th [Sch28].

6 [Bon94, Rus17]. 60s [Hea64]. 60th [MB66, Wig96b]. 62nd [Bad12]. 65.00/$99.00 [Nav15]. 67th [Ano46a]. 69 [Ein05h]. 6th [JP89].

7 [Del11b, Del12, Gor02, Kra02, Tur05]. 70-tem [vL49a]. 70th [HBD71, Som49b]. 71 [Her03]. 72 [’t 02]. 73 [Leg03]. 76th [Lau55b]. 7th [Sch28, Ano21i].

8 [Coo04, Ein05-39, Myr21a, Van22]. 80th [CS79b, FPLR59, Lin60, Ros79c]. 81 [Nam10]. 88 [BS92]. 88-15-02809-9 [Seg91]. 888-423-1212 [Wac06]. 8949/91/11/113002 [Ree16a]. 8th [Cli09b].

9 [BFB82, Ein05w, Har02b, Hoh13, Nau17, Ree17c, Tho11b, War03a]. 91 [Ree16a]. 93 [Nod34]. 97 [GHK^{+}07]. 978 [Ach13, Ber07a, Del11b, Del11a, Del12, Eng17b, Hea13, Hoh13, Kle14a, Koj08, Kri08, Mar10, Mau09, Myr21b, Myr21a, Myr21c, Nau17, Per10, Pes13, Ree15d, Ree17d, Ree17c, Ren21, Rus17, Sch07c, Sta21, Sto17, Tho11b, Van22, Weg08]. 978-0-12-398498-2 [Ach13]. 978-0-19-874299-9 [Nau17]. 978-0-19-920566-0 [Kri08]. 978-0-226-51951-7 [Del11b]. 978-0-226-77057-4 [Mar10]. 978-0-226-77057-5 [Del11a]. 978-0-230-35487-6 [The12]. 978-0-2620-3961-1 [Ste22]. 978-0-2625-3897-8 [Ste22]. 978-0-521-82834-5 [Hus15]. 978-0-521-88346-7 [Del12]. 978-0-544-80856-0 [Sto17]. 978-0-674-02828-9 [Tho11b]. 978-0-691-12094-2 [Koj08]. 978-0-691-12825-2 [Ber07a]. 978-0-691-13968-5 [Kle14a]. 978-0-691-14174-9 [GRC18]. 978-0-691-17737-3 [Sta21]. 978-0-691-20166-5 [Myr21b]. 978-0-691-20672-1 [Myr21c]. 978-0-691-20739-1 [Vog21]. 978-0-691-20983-8 [Myr21a]. 978-0-7432-6473-0 [Gor07a, Sch07c, Weg08]. 978-0-8018-9359-9 [Mil11a]. 978-0-8129-9898-6 [Rus17]. 978-0-88135-399-0 [Mau09]. 978-0-941232-44-9 [Ree17c]. 978-1-107-01744-3 [Nav15]. 978-1-4008-6577-2 [Eng17b]. 978-1-48-750177-8 [Van22]. 978-1-58988-044-3 [Hea13]. 978-1-63388-223-2 [Ree17d]. 978-3-030-61372-3 [Sho21]. 978-3-0340-1006-1 [Pes13]. 978-3-540-69221-8 [Hen11c]. 978-3-642-27560-9 [Hoh13]. 978-88-8342-530-1 [Per10]. 978-9-81-461270-08 [She15]. 978-981-3237-70-4 [Vog20]. 978-981-4299-37-4 [Gil12]. 978-981-4612-67-8 [She15]. 978-981-4612-68-5 [She15]. 978-981-4618-79-3 [Ree15d]. 978-989-8363-41-1 [Ren21]. 981 [Cah04, Coo04]. 981-238-004-3 [Cah04]. 981-238-233-X [Coo04]. 981-238-234-8 [Coo04]. 9th [JP08].

= [Ano04c, BD92, HLZJ91, MK56, Rol19, Sch09a, Sou19, Uns00]. Ē [EI78a]. ōinaru [Gol98b].

ôlam [Ayn34]. ôlami [Ayn34].

A&E [Wac06]. A-bomb [LRD^{+}91, Asa95, Mal09, Eva53, Hir98, Mac85]. A. [Asi83a, Bok15, Bur97, Ein22c, Ein23f, Ein86d, Fer81, Fer82c, Fra37, Gil84b, HH10, Hei83a, Hel84, How97a, Isa85, Lak85, Lar84b, Lov86, Met95, Mur25, Nav94, Pal83a, Pip83, Pip97, Red84a, Rus43, Sta07c, Těš12, Vie29, War84, Whi83b, Wig96j, PHvMW93]. A./Kraus [GRE^{+}01]. a.c. [LLSS07]. a.o [HvMW79, vM93, vM96, vM99b, vM01, vM05c]. Aachen [B^{+}01]. Aachener [B^{+}01]. Aant [WM08, Wid07]. år [Sch50b]. Aarau [Hun15]. åsikter [EI38d]. Abandoned [Bar15, OMNM14]. Abbildungen [HH14]. ABC [Rus31, Rus58, Rus69, Rus85b, Rus97, Rus09]. Abdominal [CG90, Cer11]. Abdus [Alf09, Bus12a, Fra08, Gha82]. Abel [Hel14]. Abendländer [Sch17c]. Abenteuer [EI38f, EI49c, N.38]. abenteuerliche [Bod03]. Aberration [Ein14f, Ram22]. Abhandlung [EL08a, Ein14f, Ein15b, Ein16n, Ein20d, Ein22b, Ein25b, Ein05c, EL05b, Ein05-38]. Abhandlungen [BSW22, Bor63a, LEM13, LME23]. abhängig [Ein05b, Ein05v]. Abiko [Dar04b]. ability [Sti13c]. Able [Ano29d]. Ableitung [Ein17c]. Ablenkung [vS21a, vS21b]. Abner [CHS97]. Abolition [BEI^{+}55b, RE63, Whi55a]. Above [Rol15, Rol19, Rol71]. Abraham [Ber00, Ein05-28, For84b, Gla00, Gol84a, Pei83, Smo07a, Wal02b, Ein12g, Ein05e, Ein05-28, Kle84, Red84b, Ren07c]. Abrahams [Ein12b, Ein05e, Ein12b, Nad02]. Abroad [Jof62]. Abscheidung [HS44]. Abschlusses [Pau25]. absence [vD21]. Absent [Rad08]. Absent-minded [Rad08]. absolu [CE79a]. absolut [Fra52]. Absolute [Dor78, ES13, ES05a, EF05, Föp07, For80a, Fri07, Gru57, Kil80, Sha99, Ano19a, Ano84a, CE79b, CE79a, Deb29, Deb79, EF13, FK90, Fra52, Her77, Jac08]. Absolutely [Fay10]. Absoluten [Her77, Jac08, ES13, EF13, ES05a, EF05]. Absorption [BJ41, Ein16m, Cha11, Ein06f, Ein05-43]. Abstraction [Nee79, Pye13]. Abteilung [BDR93]. Abwehr [Ein21v]. abyss [RB04]. Acad [LG97]. Academia [Ein29h, You79, Buc80]. Academic [Bla76a, Bla76b, Jam70, Wal02b, Arp98, Ber69b, Van04, Ano69a]. Academies [SG82]. Academy [Bar05f, Ein33d, Fri12, Hil39, Pla20a, Pla20b, Sim06, BEM12, Ano33e, Ano80b, Bro79a, Dir80a, Egg02, Ein21h, LMOS15, Sim06]. acausal [JP55]. accelerate [Hom06]. Accelerated [HF05]. Accelerates [Gon96e]. Accelerating [Kir02, Kir13, Sch12e, Kra07a, Rie12]. Acceleration [Dic81, GM00, HEB^{+}80, Raf17, Per12]. Acceleration/Brosche [HEB^{+}80]. accelerator [Ano51b]. Acceptance [ES13, Wie06]. Accepted [Ano19b, Bru99, Del11a]. accident [McK03b]. Accidents [Mah14, Mah15, Mey07]. acclaimed [Gla00, PP07, PP06]. accommodation [HW12]. Accorded [Man86]. according [Ano97b, Ein16m, EG23, HH10, Hei53a, Hen91, Pat85, Thi46]. Account [Ein17g, Lat88, Abr03, But23, Hau54, KV11, Nic87, Ros02a, Smy45c, Smy45d, Smy45e]. accounts [Cam09b, Dar21, Sha87]. accuracy [Bra37, Gré19c, Mac93]. Accusations [Sha92]. Accused [Ste95b]. achieve [Koc01]. Achievement [Hay91, Jam06, SR11, Whi67]. Achievements [Ano84c, Fer82b, Gol84b, RN16, SS97, Woo80b]. Achieving [Hym12]. Achtzehn [ABH^{+}55, ABH^{+}87]. acknowledgement [LHZ^{+}23]. Acoustics [Ano79g, Ste98]. Acrobat [BHJ^{+}00]. Across [Hei74b, Ano00a, Ano00b, CZPR17, Cra00a, Goo00, GP11, Pat00, Pat01b, Pea06, PBM^{+}12, Vea01]. Act [Dow80, Lau55a, PA21, Sha73b, Ein10e, Gol83a]. Acta [Sch09a]. actes [MK56, Sto96]. acting [TN03]. actinide [KWMS05]. Action [Ano21d, Fra14, HOO88, Pet10, Sau05, Sau07b, Tib13, WF49, Ano10b, Ano15j, CPDH^{+}87, Ein36e, Lea07, Mus15b, Ach13]. active [HS39a, PB06]. Actives [PB06]. activist [Bes93]. activities [Tur06b]. actors [Kai94]. Acts [Gru97, Sta07a, Des72, Dür62, Dür80, Dür06]. Actually [NHG89, Lee06, Sch61]. actuel [Cun22a, Ein12c, EK53]. ad [Pal99, Ray92, Seg86a, VR06]. Adaptation [For71b, For80b, MC80, Wei15c]. Added [McC84]. Addenda [Duk49]. Addendum [Aga07, Ein25a, Ein05y]. addicního [Žáč12]. Adding [HG06a]. Addison [Kai81]. Addison-Wesley [Kai81]. Addition [Wal02b, Ein21t, Kin95, Lie78, Ung07, Žáč12]. additional [ES46a, Ein14k]. Address [Ano30a, Ano33a, CCD^{+}52, Ein32b, Wig96a, Wig96d, Abi00a, Ein21h, Ein30g, vL55b]. addressed [Ein48e]. addresses [JDB^{+}97]. addresses/Looking [JDB^{+}97]. adds [Hak07]. Adi [Pye80]. adjustment [Gré19c]. Adler [Ein18g, Ein18g]. ADLON [JDB^{+}97, JDB^{+}97]. ADLON/in [JDB^{+}97]. ADLON/kurzgefaßt/Auf [JDB^{+}97]. administration [Rol47]. Adobe [BHJ^{+}00]. Adolf [Kac96]. AdP [Ein05c, Ein05d, Ein05-40, Ein05g, EL05b, Ein05e, Ein05h, Ein05f, Ein05l, Ein05i, Ein05m, Ein05k, Ein05j, EL05a, Ein05q, ES05a, Ein05-41, Ein05t, EL05c, Ein05-31, Ein05-37, Ein05u, Ein05-38, Ein05o, Ein05-32, EL05d, Ein05w, Ein05x, Ein05-36, Ein05-33, Ein05-35, Ein05r, EF05, Ein05y, Ein05-27, Ein05n, Ein05p, Ein05-28, Ein05-34, EH05b, EH05c, Ein05-42, Ein05-30, Ein05s, Ein05v, Ein05-43, Ein05-29, Ein05-44, Ein05-39]. Adressen/Blick [JDB^{+}97]. Adrien [KW87]. Advance [Ano34b, Van10]. Advanced [Fle33, Reg87, Sha73b, Myr21b, Myr21c, She16b, Vog21]. Advancement [Ein50k, Hol86, Hol98b]. Advances [Ano23a, MTVB88, NJP92, Hei93]. Advent [Ein51a, Tho13, Pye85]. adventure [EI39c, EI58, EI65a, HB01, MB11, Ren21]. Adventurer [Sta03b]. Adventures [BFB82, Mah17, SS01, Ula76, vH62, CK^{+}04, EI38f, FL88, FL06, Kee04, FLH85]. adventurous [Bod03]. Adviser [Eck90]. Advising [Her00, vB02]. Advisors [Dys76, Yor89]. Advisory [FR49a, Uni96]. AE [Ein79e, Ein85e]. Aegean [Pap15]. Aera [NS88]. Aerodynamics [Ste98]. aesthetic [Cur82, Eng05a]. aesthetics [Cha87]. Aether [Ein18c, Ein22k, Ein24j, Ein30h, Hei10, Nav21, Swe13, Gol83b, Lan21, Ein20b, Ein21k]. Aether-drift [Swe13]. aether-equations [Lan21]. Aethergleichungen [Lan21]. Aetherzustandes [Ein95]. af [RRK^{+}64]. Affair [Gol99, Van22, Uda94, Uda98, Lit91, McC77]. Affairs [GR63, GR65, Ano06f, Rab63]. Affect [Ken13]. affects [ACT23]. Affine [Car07a, Ein23j, Ein23i]. affinen [Ein23j]. Affirm [ACU^{+}54]. affirmative [Mon58]. Afoul [Bro15]. Africa [AA02]. Africa-related [AA02]. African [AE06, AA02, Cle08, Dur08]. Africans [Hec12]. After [AE06, BS81, Dir60b, Ein79c, Kae31, Kol77, Pow15, Rob39a, Rob39b, Sch21, Ski04, Wyd84, Ano15a, BL14, Cap05, Dac11, Ein19c, Els95, For73, FG20, GGW94, Hel80, Kae48, Ber84a, LSW12, Lin15, McC68, Pal98, Pra06, Sch03a, Wal79]. Aftereffects [Fre85]. Aftermath [Her85b, Kri08, Hen07]. AG [Sho21]. AGA [SR11]. Again [Ano50c, Ein54c, Eis06, Abr12a, Kle85]. Against [Bel90, Foc53a, Fox65, Fre08, G.33, Kri99, ABH^{+}55, ABH^{+}87, Bäu83, Cra00b, Ein33b, Hau54, Hau74, Hol96b, Hol00a, Ian93, IRW31, Jer02, Kra20b, Lit09, Lyn33, Mal08, Mal12, Mel79b, Urb22, Wal97, Wal04a, de 15, dP17, vB73]. Agashe [Aga07]. Age [Ano49a, Bru77, CP15, Dix00, GR63, GR65, Joh06, JW11, Lan48, SL93, Stu18, Acz09, Ano65a, Bjø19, Bus12c, Car10, Cha05c, CG01, Els78, EE72, Fer03, Ham08, Her79a, Kra88, Lec86, Lui11, Mag09, McK84, MP99b, NKY15, Nel14, New88, New89, Rab63, Rob58, Sal95, Say85, SS01, Boy85, Boy94, Cre15, ČA99, ES68, Erc55, Gil08, Gil09, Hah79, Her93, Mas95, Pri12b, Roc03, SFW91, Swa92, Swa95, And81]. Aged [Sha14]. agencies [Gru05d]. Agenda [Ano12c]. Agent [Ove13]. Ages [Jam70, Lan70]. Aggregation [WL12]. Agile [Mar96b]. Aging [Lui11, Diz05]. agissent [Ein10e]. Agitation [ES13, ES05a]. Agnon [Ano07p, BL07a]. Ago [Ano05b, Kov01, Gib91a, Poo90, Spo16]. Agog [Ano19b]. agreed [Gio13b]. Agreement [Res50, Ein33i]. A’h [New56, New60, Jon89]. A’h-mose [New56, New60, Jon89]. Ahead [Ano97c, Ein48f]. Ahmed [Str04]. ai [EI65b, EI13, Seg83]. Aichelburg [Pye80]. Aikenhead [Dow80]. Aim [Max93c, Max93b]. Aim-oriented [Max93c, Max93b]. Ainshutain [Gol98b]. Ainshutain-no-ōinaru [Gol98b]. Air [Ano29d, Mei12, TM77, TM90]. Aires [Nir25]. Ajax [Lan00]. Ajnstajn [Inf79a, Mag95, Mil10a, Mil12a]. Ajnstajn-Marić [Mag95]. Ajnstajna [Pop98, Pop05, Sre29, T– D69, T– D95]. Ajnstajnom [Mil10a, Mil12a]. Ajnstajnove [Mar95b]. Akad [Pla20a]. Akademie [Jüt21, Pla20b, Sim06, Tre66, Ein21h]. Akademie-Vorträge [Sim06]. Akademiebibliothek [Egg02]. Akademische [Ano69a]. akausaler [JP52]. Akte [Gru97, Ber90a]. Akten [Des72, Dür80, Dür62]. aktiver [HS39a]. Akustik [Ano79g]. Akzeptanz [Wie06]. al [Bat03, Bjø20a, Bjø20c, Bla86, Mar11c]. Alamogordo [GHK^{+}96, GHK^{+}96, LW71]. Alamos [Car04a, Her21, Mai99, Ree16a, Ree16b, Ree17c, Ree17b, Roe93, Sei99a, Ano45a, Ano86, Ano96a, BHB80, BFB82, Bed01b, Bet91, Bro60, BS97, Bus12c, Cha74, Chu60, CC74, Chu98b, CC98, Chu01, Con05, Fis85, HTS83, HHMW93, HB05, Hun07, Kan97, LE84, Mar00, Mas95, Nes14, Pet72, Rho93, Ric93, Ros02a, Rot92, Rot97, Sei95, Ser43, SR92, SG11, Sza92a, Sza92b, WB01, WS88, WS08, WA03, Woh76]. Alan [Sha14]. Albert [Adl03a, Ano39d, Ano39e, Ano55d, Ano60, Ano69c, Ano79h, Ano79g, Ano84a, Ano84c, Ano93c, Ano99b, Ano99c, Ano00b, Ano05o, Ano13a, Arm38, B.50, Bac89b, Bad15, Bai89, Bal91, Bal96, B^{+}79, Ben99, Ber07a, Ber90a, Bes04, Bin06, BHJ^{+}00, Bla50, Bon94, Bor56, Bor63c, Bre08, Bri66, Bro66, Bro72, Bro39, BW91, Buc93, Büh98a, Bus13a, C.50, C.51b, Cah90b, Cah90a, CR07, Cao06, Col68, Cou52, Deb29, Egg02, Ehl05a, Ein91g, ER13, Fel11, For70b, For70a, For80a, For84b, FF97, Gea94, GKP12a, GKP12b, Gli84, Gol84a, Gol84b, Gol87, Gru05d, Grü90, Gut15, Hal55, Ham11a, Har91, Har96a, Har96b, HY86, Hei53a, Hei75a, Hen11a, Hen82b, Hen88a, Hen88b, Hen90, Her63b, HC94a, HC96a, HC04, Hof08c, Hol05b, Hor00, Hou65, Ill89c, Inf50b, Ing05, Joe19]. Albert [Joh11b, Jun05a, Kae38, Kar05, Ker89b, Kev05, Kie82, Kil84b, Kil95, Kil98, Kle70b, Kle70a, Kle73, Kle75a, Kle84, Kle75b, Kon15, Kra25, Krs04, Kuh54, L.58, Lac08, Lam79, Lau62, Leg15, Len39, Loe88a, Loe88b, LS05, Mal65, Mar73, Mar11b, Mar11c, Mcc87, Mcc55, Mic68, Mil11a, Müh11, N.38, Nuv11, Pau61b, Pei83, Pol83, Pop98, Pop05, Pye89, Pye93, Pye95, Pye97, Pye06, Qui40, Rai46, Ram79a, Ram81, Ram90, Red84b, Rei50, Rei51, Ren97b, RD05, Ren05e, Rob48, Roo39, Row00, Row06a, Row06b, RSPS49, Rus17, Sar35, SE79, Sch79b, Sch17c, Sch78, Sch05e, Sch12f, See65b, See56b, Set11, Sha75, Sha54, She16a, She16b, Sie97, Sie07, Smi89b, Smi89a, Smo07a, Sol59, Spe79, Spe62, Spe67, Spe68, Spi96, Sta00, Sta95c, Ste90, Sti05]. Albert [Str90, Swa02, Tay21, Tou05, T– D82a, T– D83, T– D85, T– D88, T– D93, Tre66, Tre21, Wal99, War95, War96a, War96b, Wie06, Wig96g, Woo28, Wre66, Wue05, Yan87, Zau05, vB05, vL49a, Ano79a, Abb05, AS79, Ald93, All05, AH60, A^{+}09, AH96, And15, And17, Ano33b, Ano50a, Ano55a, Ano55b, Ano57b, Ano59, Ano66, Ano67, Ano69c, Ano75, Ano79b, Ano79a, Ano79c, Ano79f, Ano79g, Ano80a, Ano81a, Ano83, Ano87a, Ano00a, Ano04a, Ano05c, Ano05e, Ano05g, Ano05f, Ano05h, Ano05k, Ano05i, Ano05j, Ano05l, Ano05d, Ano05n, Ano08a, Ano11b, Ano13a, Ano15b, Ano16a, Ano18a, Are05, ABM79, Arr07, Ash11, Ash12, Asi83a, Aud06, Bac89b, Bac89a, Bac05, Bad83a, Bad15, Bai07a, Bak97, Bal05, Ban12, Bar80a, B^{+}06]. Albert [Bäu83, B^{+}79, Bel96a, Bel98, BM09, Ben01, Ber07a, Ber74a, Ber92a, Ber82b, Ber94, BF78, Ber79b, Ber96a, Ber96b, Bič79c, B^{+}89, Bje02, Blu22, BR55, Bor55a, Bor56, Bor63c, BB69, BB71a, BB71b, BB05, Bra15, Bra17, Bra02, Bre79, Bre08, Bri05d, Bro73a, Bro79b, Bru09, BRI^{+}21a, BRI^{+}21b, BW91, BW92, Buc93, Büh98a, Bur94, Bur97, Cal02, CL05, CH98, Cas94a, Cas94b, Cau91, Cer11, Cha79, Cha05b, CD07, Cle06, Cob07, Coh79, CG90, Col09, Col57, Col93, Col00b, Cou79a, Cou79b, Cra00a, Cro05, Cug06, Cwi87, Dad78, DL15, Deb79, Ded85, Del13, Den84, DSMH85, Die05, Din93, Diz05, Don97, Duf81, DH79, DH81, DHR13, Dur05a, Dür79, Dür82, EH05a, Egg02, Ehl79, Ehl05a, Ein24f, Ein31m, Ein36c, Ein39a, Ein55a, Ein55f, EHM^{+}79, Ein85c]. Albert [vM90, Ein91b, Ein95a, Ein97, ECS02, Ein05-46, Ein05a, Ein05-45, Ein05b, EKS^{+}08, Ein22, Elb13, ER13, Fal09, FLN13, Fan06, Far05a, Fau79, Fer05a, Fer81, Fer82c, Fer82b, Fer83, Fin05, Flü74, Fok55, Fol17, Föl93, Föl97, Föl98, For80a, For63, FF97, Fre06, Fre15a, Fre85, Fre19, Fri79a, dFV04, Gal99, Gal07, Gal79, Gan97, Gar39, Gar40, Gar41, Gar42, Gar77, Gen98, Ger79, Gil84a, Gil84b, Gin15, Gin22, Gla07, GKP12a, GKP12b, GC96a, Goe03, Goe15, Gol83a, Gol84b, Gol04, Gol15, GE97, Goo00, Goo75, Got80, Gra99, Gre03a, Gre04a, Gri79a, Gri79b, Gri81, Gri93, GG05b, Gru05d, Grü90, Gun15, GP11, Gut94, Gut15, Gwy02, Hal55, Ham11b, Ham85, Ham73, Ham86, HD14, Han79c, HH16, Har91, Har96a]. Albert [Has04, Has09, HY86, Hav77, Haw07, Hay91, Hei83a, Hei49a, Hei55, Hei74c, Hei75a, Hei83b, Hel80, Hel84, Hen11b, Hen82a, Hen90, HG05a, Her66a, Her78, Her79a, Her33, Her63a, Her63b, Her67, Her74, Her75, Her79b, Her11, HC93, HC94a, HC94d, HC94b, HC94c, HC95, HC96a, HC96b, HC96c, HC98, HC04, HD75, Hof97a, Hof49a, HD72, HD73, Hof82a, Hof08c, Hol81b, HE82, Hol00b, HG06a, Hoo77, Hor79b, Hor79a, How94a, How97a, How05a, Hu05, HL08b, HL12b, Hun10, Hun01, Hun15, Ill79, Ill89a, Ill89c, Ill06, Inf50a, Inf50b, Inf53a, Inf53b, Inf54, Inf55a, Inf55b, Inf56a, Inf56b, Inf57b, Inf57a, Inf58, Inf60, Inf61b, Inf61a, Inf63, Inf68, Inf74, Inf79a, Inf79b, Inf84, Inf92, Ing05, Ire89, Jay05, Joe19, Jof55, Jon89, Jor69, Jun08]. Albert [Kai81, Kak04, Kak05, Kal35, Kan92a, Kan94a, Kan96a, Kan92b, Kan94b, Kap80, Kap81, KK89, Kap91b, Ker89a, Kie15, KWGH97, Kil80, Kil84a, KH05, KT79, Kle82a, Kle05, Kle10, Kle16, Kle75b, Kli62a, Kli62b, KS12, Kni16, Kol56, Kol58, Kol72, Kol74, Kon15, Kor79, Kos98, Kox06, Kra25, Kra14a, KWM57, Krö53, Kru94, KK09, Kur79, Kuz79, Kva79, Lak85, Lam79, Lan55, Lan59, Lan65, Lan67b, Lan67a, Lan74c, LA88, Lar84a, Lar84b, Lau62, Lau15, LLH^{+}80, Lev15, Lew79, Li07, Lin15, LFKW06, Lov86, Lut05, Man86, Man55, MH12, MO13, MFHR06, Mas05a, Mas02, McC15, McP95, McP97, Mcc78, Meh87, Mel78, Mel79a, Mel79b, Mel82, ME14b, MFO14, Met95, Mic78, Mil10a, Mil12a, Mil13a, Mil15, Mil75, Mil81, Mil92, Mil93, Mil98, Mil11b]. Albert [Mil49a, Min84, Min89, Mor96, MRS17, Mor79b, MNOW00, Muk05, Mur25, Nag50, Nav94, Nee79, Nef11, Nev92, New56, New60, NJ05, Nov92, OK81, Opp39, Opp65, Opp66, Opp79b, Opp79c, Oxl03, Pai82, Pal83a, Pal83b, Pan05, Par04, Pat00, Pat01a, Pat01b, Pát79a, Pat79b, Pat93, Pau58, Pau61b, Pau94a, Pau94c, Pea49, Pea06, Pei55, PS79, Pfl99, Phi74, Pip83, Pip97, PP95, Pol83, Pol91, Pop03, Pop57, PL17, Prz86, Prz11, Pye80, Pye95, Pye97, Pye06, RMI21a, RMI21b, RF55, RF79, Rad08, Rai75, RR82, Rav97, Red84a, Rei34, Rei01, Rei30, Rei31, Rei07, RS92, Ren97b, RCD98, Ren05c, Ren05b, RD05, Rin03, RS07e, Rol10, Ron07, Ros82a, Ros98, Ros05, RW07, Ros18, Row00, Row06d, Rya06, Sal99, San77, SE79, Sch02a, Sch05a]. Albert [Sch02b, Sch49b, Sch49c, Sch49d, Sch49a, Sch51a, Sch51b, Sch55, Sch70, Sch79a, Sch83a, Sch84a, Sch12c, Sch50a, Sch81a, Sch93a, Sch87b, Sch17c, Sch95a, Sch96b, Sch84b, Sch88a, Sch88b, Sch92b, Sch82b, Sch05f, Sch05e, Sch12f, Sch03d, Sea79, See52, See53, See54, See55, See56a, See56b, See60, See61, See64, See80, See86, See91, SL05a, Sei99b, Sha92, Sha63, Sha73a, Sha79a, She16a, Shi49, Sig92, Sim80, Sim05, Sim06, Sin01a, Sin10a, Smi04, Smi05b, Smi07a, Sol87, Sol56, Sol59, Som49b, Som49a, Som49c, Som55a, Som55b, Spe79, Spe72, Spo16, Sre29, Sri94, Sta95b, SL91, Sta92b, Sta95c, SL05b, Sta06, SE90, Ste05a, Ste05b, Sti05, Sti13c, Str94, Str82, Sug89, SE45, Tal82, Tan66, Tau79, Tee02, Tel78, Tho00, Tho83, Tho05]. Albert [Tho88b, Tob94, Ton97, Tou14, T– D69, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D95, T– D07, T– D12, Tre79a, Tul79, Uni00, Vai89, Val54a, Val55, Van07a, Van96, Van79, VFIS63, Vea01, Viz79, VW86, Vrá57, Wad05, Wad07, War84, Wei12a, Wei04, Wei12e, Wei12b, Wei12f, Wei12d, Wei12c, Wei12g, Wei12w, Wei13e, Wes05, Wet80, Whe79a, Whe80, Whe91, Whi05, Whi83b, Whi55b, Wie06, Wig49, Wig96g, Wig96l, Wil84, Win69, WE24, WE87, Wis79, WKH99b, WKH99c, WKH99a, Won05, Woo80a, Woo80b, Wue05, Yan06, Yan13a, Zal91, Zau05, Zei99, Zho85, Zim84, Zwe80, dB49b, dB49c, dB55a, dB55b, van06, van12, vD15, vD17, vB05, vL55a, vL55b, vL55c, vL57, vL59, vM05b, vM09, vM12]. ‘Albert [Lil86, Pye80, Ano84b, Bus12b, C.51a, Coh51, Cus82, Fri73, Gin82, Hei74a, Jeh51, Lig97, Med76, Moy81, Noe93, Oha79, Ram79b, Ram79c, Ram79a, Ram83, Sch92c, See65a, Smi51, Wig79]. Alberta [Bič79a, Pop98, Pop05, T– D69, T– D95, Hav77, HC95, Isr73, Jof55, Rus90, Sre29, T– D07]. Alberto [Ein31m, Kol58, T– D12]. Albertom [Mil10a, Mil12a]. Albertu [Pát79a]. Albertus [Sch17c, Sch17c]. Albie [Tou14]. Albrecht [BL09, Lig97]. Albuquerque [Hel11]. Alchemist [Kri19c]. Alchemy [Scl89, Hab02]. äldre [EI38d]. Aleksander [Tam14]. Alexander [Jos95a, Ein17a, TFC93, Wei03, Jos95a]. Alfonso [Fel82b]. Alfred [Adl03a, BCS82, Goo85, Hel83, Hop07, Sei84, Van22, Bai06, Ber62, EF69, Fan91, Fan93, Fre13, Hal59, KWB^{+}02, Lar08, Lin99, LB02, Lun74, Mau09, Mei54, Pau42, Pau47a, Sta89b, Stê60, Str83, Wil74, Wor10]. Algebra [KN08, Kin95, NP88]. Algebraic [EK54]. Algorithm [HG06a]. algorithmic [Van96]. Alice [GRC18]. Alive [Koe95]. alkali [Leg01, Leg03]. alkaline [HS39c]. All-Embracing [Lau55a, Hon04]. alla [EI65b, EI13]. Allan [Ste22]. alle [Cap01]. allegedly [Ben99]. Allein [de 15, dP17]. Allen [Rob48]. aller [Bac05]. alles [Whe68b, Ing05]. allgemein [Ano22a, Ein18d]. Allgemeine [Ein18b, vL23, Ein16j, Ein17g, Ein18n, Ein20o, Ein25e, Ein54i, Ein88b, vI11b, Wey18, EG27a, EG27b]. Allgemeinen [Jun05a, Row06a, Row06b, Ein04, Ein14g, Ein15j, Ein15e, Ein15g, Ein16h, Ein17e, Ein18e, Ein18m, Ein19e, Ein19i, Ein21t, Ein23k, Ein23b, Ein30n, Ein31o, Ein05-40, Ein05o, Ein05-27, Foc53b, Hen90, Isa85, Kra21, Kra25, Pye93, uJR15, Sch17, Tre66, Voi86, Wey23, Wue05]. Allgemeinene [Ein16b]. Allgemeiner [HS44]. Allgemeinverständlich [Rei41]. alliance [Bla02, She75]. Allied [Irv67b, Log96]. allies [Row06c, Ber93]. Allocation [AKD23]. Allvar [MH12]. almanac [Cal04]. Almost [Ren07c, Oli99]. alone [Kle81, de 15, dP17]. Alperovitz [Ham97, Ove96, VB96, You95, VB96]. Alpha [Bar02a, Bar02b, Sei03a]. already [Bre08, Dor78]. Also [Ano29d, Ano41, Die88]. Alsos [Gou48, Ano01b, Egl97, Gou47b, Gou47c, Gou48, Gou83, Gou96, Pas69, Pas70]. alt [Hei49a]. Általános [Ein22a]. alte [RCD98, Sti05]. Alten [Jun08]. Alteration [Ano25a]. Alterations [AH96]. Altered [Lau50]. alternatief [Gui09]. Alternative [Ano21b, Grü73b, Nav21, Voi86, Gui09, Kle74, Lon84, Lon03, Ska94]. Alternatives [Mos18, GMG^{+}77]. Althoff [Som05a, Som05b, Som05a]. Alunni [Del13]. Alva [Fel82b]. Always [Pal95]. Am [Her03, AH93, Ano69a, Bre05, Ein20b, Ein21c, EP06, Fla15, GHK^{+}96, HY86, Jüt21, Lam79, ME14b, Pla20a, Pla20b, RK86, Tre66, WvLuAE18, Wig64a]. Amaldi [Bat03, RRR07]. Amateur [Cra79]. ambiguous [BSPKN03].

Ambition [Jon08, Kea18]. Ambitionen [Sta98d]. ambitions [Hym12, Sta98d]. Ambitious [ESA15]. Ambr [Jüt21]. America [Ano00b, AK97, Ano00a, Bal86, Bla02, Bra12, Cra00a, D’A93, Ein21g, Ein22, Ful84b, Ful84a, GPSR01, Goo00, GM06, Ill06, Isa09, Jac11, Jun14, LML88, LM95, McK03b, Mis85, Moo17, MT98, Ndi07, Nic87, Pat00, Pat01b, Pea06, Say85, Smi65, Smi70, Smi07a, Sop88, Tak95, Tho88a, Vea01, Wac06, Wel99, Wel00, Wil16b, Woo80a]. American [Ham97, Hei53a, Ove96, Pas10, Ree17c, VB96, Vis94, You95, Alp65, Alp85, Alp94, ATW^{+}95, Asa95, Bel96a, BBM78, Bet63a, BS05, Boy85, Boy94, Bro21, CHW91, Cas05c, Cre81, Ein32d, Ein33i, Ein41f, EM49, EBU^{+}52, FH11, Fis92, Fra89, Fra04, FK68, Gal93, Gil62, Gon96a, Gon96c, Gon00, Goo03a, HCD^{+}78, Hec16, Her82b, Hol81b, How50, LE96, Mon58, Ric06a, Rol89, Ros79b, Val06, Win93, Win99, Zac97, Zac99a, Ano05b, MKBF30a, MKBF30b, Rom91]. americana [Val06]. Americans [Amb97, Eic94, Gro49]. amerikanische [Ein21g]. amerikanischen [Ein33i, Hei53a]. Amherst [Bar05f, Ree17d]. Amid [Sta19]. Amis [Ami04, Die95]. Among [Ein40d, Whi55a, Cre81, Hub29, Med94, Mer79]. Amount [Ano39a]. Ampère [Har02b, War03a, Ein15f, EdH15a, EdH16a, EdH15b, EdH15c, EdH16b, Dar00, Ein15c, EdH15c, Ein16e, MP74]. Ampèreschen [Ein15c, Ein15f, EdH15a, EdH15b, Ein16e]. AMS [Fre18]. Amsterdam [Hea64]. Amy [Rob48]. Ana [Ren21]. Analog [Ein12d]. Analogies [Gin15]. Analogue [Cun20b]. Analogy [Dor71, Whi83a]. analyse [Pia31]. Analysed [Gin82]. Analyses [Bal91, Bal96, Bro83, Col68, Don70, Hum51, Lam07, Mat86b, Mat86a, P.50, Ram82, Ram83, Spe62, Spe67, Spe68, Tat64, Wal99]. Analysis [Ao04, Ano23a, Bar98a, DFB22, Hay91, AKD23, Bin06, Cha11, CZPR17, CD23, EG80a, FLN13, GW10, LR14, MSYW12, MTY79, OM14, OM22, Rol47, SBB16, UKKP04, ZZ23, ZXLX14]. Analytical [FGPP06]. analytiques [Pia31]. analyzing [TTB20]. anarchy [Fol17]. anatomy [Cra01, Gar93]. Ancient [Ano14a, Buc13, Sac81]. and/or [Mil12b]. andere [LS05]. Anderson [Ein23a, Ano51a, Bad72, Ein23a, Ein24b, RDF^{+}08]. Andover [AS91, Gly90, HS89b]. André [Hum51]. Andrei [Bro79b]. Andrew [Fre08, Hen86a, Kai05a, Nau17]. Anecdotes [BFB82]. Aneurysm [CG90, Cer11]. aneurysms [MNOW00]. Anfassen [Sic09]. Angeles [Bed01a]. Anger [Alf09, Bus12a, Fra08]. anglaise [CE79a]. angle [Lam21]. Anglo [Goo03a]. Angriff [Bre05]. Angrist [Mea21]. angry [MP99b]. Angst [Ein17b]. Angst-Traum [Ein17b]. Angular [GM00]. Anhang [Rei41]. anisotropies [Smo07c]. Anlaß [NHP^{+}79, Tre66]. anlässlich [Ren05e, Hal55]. Anmerkungen [Han18]. Ann [Row00, Smi02, Bro10b]. Anna [Cah90b, Har91, Hen88b, Pye89, Pye95, War96b]. Annahme [ES13, ES05a]. Annalen [AP65, Ren05d, Ren05f]. Anne [Kil95, Sie97, War96b]. Annette [Pye97]. anni [CDS01]. annihilation [Kee11]. anniversaire [Ano80a, Sto96]. Anniversaries [Kha20, Kis82]. Anniversary [Ano81d, Ano93a, Ano95a, Ein24k, Ein53k, OJCP^{+}47, Rot95, She16b, Wis79, Aĭn77, Ano67, Ano80a, BRVF84, CGT11, DS83a, EVB^{+}72, KK05, LFKW06, NHP^{+}79, Sti05, Sto96]. annotated [WK12]. annotations [She16a]. Announce [Ano23e]. Announced [Ano39g, Bet63b, Spo16]. Annual [Ano93c, Ein53l, Mir94, Pin00a, Wig65, WK12]. Annus [GG05b, Mas05a, RR06, Jan06]. anomalies [Chr22, Hof82b, Sta06]. anomalous [Mes83]. años [Paz10]. Anregungen [Stu87]. ans [A^{+}84]. Ansätze [Hen92a]. anschauliche [Ein16d]. Anschauung [Stu14, Stu23]. Anschauungen [Ein09b, Ein09c]. Anschluss [Pes13, Sie10]. Anschütz [LS05]. Anschütz-Kaempfe [LS05]. Ansprachen [WvLuAE18]. Answer [Boh62, How93a, Ano69c, Ein14k, Ein20h, Ein12a, Ein14b, Ein15b, Ein24b, Ein05c]. Answering [Mea21]. answers [Wol00]. anthology [Hen96, The95, Hie84]. Anthropic [BTM84, BT86, BT88]. anthropique [BTM84]. Anti [CHR96, Fol96, Goe93, Mir11, Yam09]. Anti-Einstein [Goe93]. anti-nuclear [Yam09]. Anti-Realism [CHR96, Fol96]. anticipated [TC94]. anticipates [Obe16]. Anticipations [RS49]. Antimatter [Clo10, QN08]. Antiproton [Kae13a]. antirelativitätstheoretische [Ein20h]. Antirelativity [Ein20h]. Antisemitismus [Ein20a]. Antoinette [Ber94]. Anton [Fis10, Fis12, Fis10, Fis12, Kle75b]. Antoon [Kox88]. Antrittsrede [Ein14a, EP06]. Antwort [Ano69c, Ein14k, Ein20h, Ein12a, Ein14b, Ein15b, Ein24b, Ein05c, Ein05d]. Anwendung [EH10b, Ein12a, Ein15b, Ein15g, Ein21e, Ein05c, Ein05d, EH05c]. Anwesenheit [Bos24b, Ein25a]. anxiety [Win93, Win99]. Any [GB89, HS14]. Anyagokra [von90]. anymore [Coo83]. Anything [CF12, Gol87]. Anziehung [Eöt91]. Aortic [CG90, Cer11, MNOW00]. Aortic-Aneurysm [CG90]. aos [Seg87]. apart [Kai05b]. Apfel [Büh98b]. aphorisms [Ein31d]. Apologetics [BC95]. Apology [Gen23c]. Aporie [Mas99]. apparatus [Coo83]. Apparent [Hum36, Ank15]. Apparently [Ein14h, HS14]. Appeal [BEI^{+}55a, BEI^{+}55b, Ein35a, Res50, RE63, Pye82b]. Appearance [XMY^{+}09]. appeared [Ein19b]. Appendice [Ein53a]. Appendix [Jan14a, Gru05c, Rei41, Ano82b, Ano82a, Ein53n, Krs91]. apple [Büh98b, SS15]. Applicable [Ano15d]. Application [BC82, Ein15g, Ein21e, Ein05c, EH05c, Hei47b, Hei47a, Lem49a, ST06, Bad09, Car89, EH10b, Ein15b, Ein05d, Iro04, Mou19, NP88, Sta16, Ein12a]. Applications [Ano23a, EWJ^{+}25, HEB^{+}80, Mur25, Par79, Per00, Sch17b, YC18, Alf01, Ano15a, GH07, RDB^{+}09, Wei72, Wei05b, Whi22, Whi04b, Pia31]. Applications/Lehmann [HEB^{+}80]. Applied [BZ10, Hay91, Par79]. Apply [Lev01b]. applying [Koc01]. Appointed [Ano34b]. appointment [Hav77]. Appraisal [Grü73b, HR02, Kan94b, RW02, Tha77, How76, Sri94]. appraised [Che13]. Appreciation [Ein82a, Wig96b, Bra29, Hol03a, Zim84]. apprentice [Jak13]. Approach [HBN00, Kap91b, Kra05, PS03, Tor82, Tre79b, Vec17, Bon64, EE59, EE90, FGMvH14, FBD05, Hag08, Kau85, Mit87, Nav91, RdO07, RdO16, Sar96, Sau06, Bri18]. approaches [FGPP06, Hen92a, Hsu00]. Appropriate [CC67]. Appropriation [Leg15]. approve [Ano04a]. Approximate [Ein16l]. Approximation [Gás54]. après [Ano15a]. April [B^{+}79, Bre79, BSKS04a, BSKS04b, Ein29h, Hol05b, Kev05, WvLuAE18, Whe80, BIR^{+}15a, BIR^{+}15b, Eme48, LLH^{+}80, Roe93, Van04]. Apriori [Rei20, Rei20]. Aquinas [Kil70, Wal61, Wal72]. Äquivalentgesetzes [Ein05-30, Ein05y, Ein12f, Ein12i, Ein12h]. Äquivalenz [Gio13a, Gio13a]. Äquivalenzhypothese [Ein16n, Ein05-38]. Äquivalenzprinzip [Tre71]. Ära [BS92]. Arab [Zia94]. Arabs [EK44]. Arbeit [Ein07b, Ein09a, EL09, Ein11a, Ein11e, Ein12f, Ein15c, EdH15b, Ein18k, Ein22c, Ein22d, Ein23b, Ein23f, Ein05g, Ein05h, Ein05l, Ein05i, Ein05m, EL05a, Ein05y, vI11a, vI11b, Sch05c]. Arbeiten [ARK^{+}99, Ein11f, Ein05j, Hei46, vM90]. Arbeitsgemeinschaft [Ano79g]. Arbitrary [MSS22]. Arbor [GW83]. arc [Dar06a]. Arch [JR15]. Arch. [Ein86c]. archetypal [JP55]. archetypischer [JP52]. Archimedes [Pic08]. architect [Pau42, Pau47a]. Architecture [Ham97, Ove96, VB96, You95, ATW^{+}95, Egg95, Gra99]. Architektur [Egg95]. Archival [Ano13b, Wid07, Elz06]. archive [Wal81, Eks99, Ros12]. Archives [Ano04b, Bar05f, Bar06, DH79, For05, ABC^{+}06, Ano03a, Bat03, DH81, DHR13, MFO14, Cra02b, For80a, Hav80, Ein03d, Moy81, Wig79]. archivi [Bat03]. Archivu [Hav80]. Area [Ein07a, Kan92b, Rot92, Rot97, Jac11]. Areas [Ano15d, Ein16d]. Argentina [Ein08b, GO08, Ort95, Pye86]. Arguing [Mar04a]. Argument [Dor71, How93a, Rös95, Wei15b, Ein12b, Ein05e, Gio13a, Hni98, Joh19, Kra07b, Nor06, Sta16, Wal77b, Wal92d, Wal06e, Wal15, Wei13g]. Argumente [ES13, ES05a]. Arguments [ES13, ES05a, Gio13a, Sla13, McH07, Mol74, Pat85, PS10]. Argyris [MGG^{+}95]. Argyris/Faust [MGG^{+}95]. arising [AAA^{+}09]. Aristarchus [Van85]. Aristid [Erm89]. Aristole [vW02, vWR04]. Aristoteles [Bar94b, vW02, vWR04]. Aristotle [Bar94b, Cro07, Dar06a, Mau90, Slo11, Whi43a, Whi43b]. arm [Cle06]. Armageddon [Kap91a]. armaments [Bro92]. armchair [GS59, Nat05]. Armin [For70b, Win69]. Arms [Ein50d, Hol81a, Yor95, Bus49, Bus68, Bus85, Dre93, Dre07, Far13b, HGS87, HY15, Hof10, Man64, Man68a, Man70, Man03, McM05, Rho07, She87, Yor70]. Army [Jon85]. Arnold [Ein05-46, For70b, Jon01, Sch10a, Win69, Sch10a, Sch87b, Ton97]. Arons [Ein19h, Ein19h]. Arrhenius [Cra84a, Kra13]. Arsenal [D’A93]. Arsenals [Rho07]. Art [GHS08, GHS18, Kov01, Pep04c, AMM99, Bar99, Ein59d, FF95, Foe11, HKS10, Hof03, Lei08, Sch00a, Cob07, dGK11]. artful [Bar95, Bar05b]. Arthur [Adl03a, Bla50, Coh51, Cus82, Duf46, For80a, Jeh51, Kai81, Ram79b, Ram83, Smi51, McC91, Mel82, Pal83b]. article [Dar04b, Ein28d, Ein47f]. Articles [Bič89a, Din65, GR63, GR65, Ano01a, Hen13, How50, LHDH21, LLH22, LLSH23, LR14, LHZ^{+}23, San15, YCL20]. Artificial [Fer39, Fer70, Gin59]. artificiality [Gio16]. artificielle [A^{+}84]. Artist [Rad08]. Artistic [Ein21c]. Artists [Mic78]. Arts [Bad11, Kie82, Pye13].

Artur [Jos95a]. Arzamas [ČA99, Dix00]. Arzamas-16 [ČA99]. Arzamas-16. [Dix00]. Ascent [Bro73b]. Asian [Abr09, Lor08]. Ask [Whi55a, Mor79a]. Asked [Lau55b]. Askew [Ano19b, Kae13b]. Aspden [For71a]. aspect [Sch44, Sch45, Sch62a, Sch67, Sch92a, Ano23a]. Aspects [LSF94, Rey25, ST06, Cle17, Fla15, Hen92a, Lan37, Meh75c, Möl84]. Aspekte [Cle17, Fla15, Möl84]. Aspekten [Hen92a]. Asperger [Jam06]. Assembly [Ein20m, HHMW93, Ein18l, Ein47c]. Assert [Ano45a]. Assessing [BLR15, Gré19c]. Assessment [Scl89, Whi04a, Wil17, BBD^{+}11, RN16, vD21]. Asset [Com46a]. Assimilation [Ein20a]. assist [Wei12i]. Associate [Ano50c, Cah90b]. Association [Ano79g, SG82, BC71]. Assuming [ES05a]. Assumptions [MNT^{+}00]. astonish [RR17]. astonishing [Lin15]. astounding [Haw11]. Astride [How98]. Astron [Sch09a]. Astronomer [Bro23, Cre83]. Astronomers [Fle18, Cre81, MCS09]. Astronomical [Ano07a, Kra17, Sch05b, dSR16a, dSR16b, dSR17, Per23b]. Astronomie [Ein15g]. Astronomy [Ein15g, GT80, Myr21b, Myr21c, PD15, RSS^{+}07a, Wil89, Cul80, Dea13, FF91, Gia03, Haw02c, Hei05, Hen97, Lin04b, Nor94, Smi82, Top07a, Smi02]. Astrophysical [Ben05, Bla10, BBCPL09, Sta07d]. astrophysicist [Cha83]. Astrophysics [BTM84, Bič00, BL14, Cho10b, CHR90, Har79a, KN08, RBB03, Tau63, Egg95, Bet03, CF11, Gou10, Gou13, Isr73, Kos03, Lon13, RR82, Sch96c]. Astrophysik [Egg95]. astrophysique [BTM84, Gou10]. asua [Inf57a, Inf60, Inf61a]. asymmetric [EK54]. asymmetries [Byr81]. Asymmetry [HG05b, FP16, HG06b]. Asymptotic [Wil05a, Gro05]. ate [EI39b]. Athens [Dys80]. Äther [Ein30h, Ein30b, Ein18c, Ein20c, Ein20b, Ein24j]. Äther- [Ein30b]. Atiyah [JW06]. Atlantic [Man86]. Atlas [Kar05, Chr22]. Atom [Ano39b, Ano39a, Ano39g, Ano41, Ano45a, Bad83b, Boh65b, Cas79, CM02b, DCTW02, Dro20, Duf46, Far09, Foe23, Fus88, Hau54, HK69, Kae45, Lau45, Lau51, Lau55b, Pau25, Rod19, Ser77, dCENdCA64, Ano51b, Aue95, BRH^{+}09, Boh85b, Bru03b, CHW91, Cat04, Cha09, CSD^{+}07, DB86, For82a, For84a, Fus90, HS44, Hau74, Hen00a, JDB^{+}97, Ket02, Lau88, Mos87a, Mos87b, Uda94, Uda98, Wal92b, Wil87, Win93, Win99, vCH18]. atomaren [Sal95]. Atomarten [HS44, Hah42]. Atombombe [Com56a, Rho88a, RT90, Thi46, Aue95, HG92]. atombomben [Jak13, RE90]. Atome [Hen11c, Rec09, Rec10, Ano23c, Boh23]. atomes [Per13]. Atomfallen [JDB^{+}97]. Atomfallen/kurzgefaßt/Max [JDB^{+}97]. Atomic [Ada07, Alp65, Alp66, Alp85, Alp94, And81, Ano39c, Ano39a, Ano45d, Ano50b, Ano94a, Ano95a, Ano08b, Ano16c, ARK^{+}99, Bad83b, Bag10, Ber76b, Ber93, Ber95, Bla46, Boh49, Boh58, Boh10, BMW89, BBSR69, BM77, Bur45, CP15, Can00, Cas17, Cau91, Com46a, Com46b, Com56b, CR08, Cra84a, D’A93, Duf46, Ein45a, Ein47a, Ein48b, Ein49b, Erm89, FR13, Fie07, FH11, For01, Gol82, Gow64b, GR63, GR65, Hal97, Ham97, HG92, HSD34a, Hei47b, Hei47a, Her00, Her21, HF98, HA62, HD69, HDHA89, HAD90, Hoh13, Hol81a, Hun07, Joh46, Kel04, Klo97, Kra79, KZS99, Lan48, Lan94, Lau51, Lau55a, Lef95, Lin60, Man5 a, Man5 b, Mah09, Mah14, Mah15, Mah17, Mal08, Mal12, Mar96a, Mei12, MM58, Mor57, Nef11, Nye99, OO01, Olw97]. Atomic [Ove96, Pal00, Pal98, Pei97, Pos11, Ree06, Rho86, Rho88a, Rho88b, RT90, RE90, RA00, Rho12, RK39, Ros79b, SK82a, SK82b, Sch65, Sch54, Scl89, Sig78, Smi58, Smy45b, Smy45a, Smy48, Smy76a, Smy78, Smy89, SM08, Str68b, Swi45, Szi45, Tur03, Tur46, VB96, Wal17, WHF^{+}96, You95, vB02, Abr98, Abr09, AK97, ATW^{+}95, Amr59, AK02, Ano65a, Ano81c, Bai33c, Bak68, Bak76b, Bak76a, Bal86, Bet91, Bla48, Bla49a, Bla49b, Car10, Com56a, Dee08, Del09a, Els78, Els95, EE72, Fei61, Fei66, Fer61, FV13, Flü39, FPLR59, GS50, Gon00, Gor09, Gos94, Gos99, Gos01, Gou48, Gow64a, GA79, Gow93, Gro67a, Gro67b, Gro68, Gro00, Hab02, Hei46, HW49, Hei69a, Hei72, HFN06, Her80, Her82a, Her88, Her79a, Her82b, HW08, Hol94]. atomic [Irv67a, Irv67b, Irv83, Jak13, Jay90, Joh17, Jon85, Jun58, Jun82, Kel06a, Kel07, Kel09, Kha12, Kra88, Kra15a, KMN^{+}97, Kur06, Lau46, Lau59, Lau72, LML88, LW71, Log96, Mai99, MW46a, MW07, McK03a, McK84, Mel00, Nac79, NS88, NS96, Nel14, Olw04, Pur63, Ree15c, Rho88c, Rob01, Rob58, Rol47, Rus90, Sal95, SS01, Sei95, SR92, She75, Sla60, Smy45c, Smy45d, Smy45e, Spa58, Swa92, Swa95, Tak95, Thi46, TM77, TM90, Van03, Wal95, Wal97, Wal04a, Wel06, Wil16b, Bed01a, Boy85, Boy94, Cre15, Erc55, For82a, For84a, For86, Fre08, Hac94, Hah79, HH89, Kie13, Mas95, Mog02, Pri12b, Roc03, Ros02a, Smy46, SFW91, Swa92, Swa95, Lit09]. atomic-bomb [Dee08]. atomica [Pye86]. Atómico [Mar96a, Pye86]. atomique [Gou48]. Atomism [Ein15a, Rüg88, Kox14, Yos95]. Atomismus [Yos95]. Atomistik [Ein15a]. Atomkeller [Fec90]. Atomkeller-Museum [Fec90]. Atomkerne [Flü39, HW49]. Atomkernenergie [Hei46]. Atomnaëiìa [Smy46]. atomnoi [Smy46]. atomowa [RA00]. Atomphysik [BS92, NS88, Thi46, Hei69a, Hei72]. Atoms [Ano33h, Ano45b, Boh23, CR12, Fer54, Fer57, Fer82a, Fer87, HH89, Nor06, Per23a, Per90, Sal19, Seg95, Tre75b, Ano23c, Bai33c, Cha09, Che13, CT98, FS01, Hah42, Ket02, Lau59, Lau97, Phi98, Rec10, Sla53, Tow65, Per13]. Atomtheorie [HSD34a, Wil36]. Atomwaffen [ABH^{+}55, ABH^{+}87]. Atomzeitalter [Kra88, Erc55]. Atomzeitalters [Hah79]. atoombom [Hoe18, vC16]. ATP [All02]. attachment [Ryc92]. Attack [Bre05, Hau74, Bie03]. Attacks [Ano35]. Attempt [Zim84, Sch19]. attempted [Bro92]. Attempts [JN14, Per23b, Sil30]. Attends [Swa02]. Attention [Gri93, Ken54]. Attitude [Gri81, Hen92b]. Attitudes [CD07]. Attraction [BR21, EJN93, vS21a, vS21b]. attractive [GSPH00]. Attraktion [vS21a, vS21b]. attributed [Hen13]. attribution [Pol05]. ATV [Ano11b]. ATV-4 [Ano11b]. Au-dessus [Rol15, Rol71, Rol19]. auch [JDB^{+}97, KRZ^{+}05]. Auction [Bar18c, Ano95c]. Audacious [Pye82a, Gib91a]. Aufbau [Ein19j]. Aufbaus [Cun22a]. Aufbruch [MGG^{+}95]. Aufrauf [ENFB14]. Aufrechterhaltung [Ein28e]. Aufsatz [Rit09b, Ein11c]. Aufsätze [Her76c, Pau61a, SW12, Jos95b]. Aufstieg [Flü74]. auftretenden [Hah42]. Aufwind [JDB^{+}97]. Aufwind/Trägheitsfusion/Erstes [JDB^{+}97]. aufzufassen [Whe68b]. Auge [Bre08]. Augengläser [Ein52e]. Auger [Tat64]. August [BCS82, BC95, BB89, Ein48g, Ein91b, GRRS99b, HW55, Ham97, Isr73, JP89, Pac72, Pac81, WY05, Ano21k, Gor07b, Jan07, Uda94, Uda98, Wei95, YJ95]. aula [GdS15]. Aung [Joh11a]. auroral [Kra13]. Ausbreitung [Ein11i, Ein05-31]. Auseinandersetzung [Ein12b, Ein05e]. Ausgabe [Hei53a]. ausgeübten [EL08c, EL05d]. Ausgewählte [Bor63a, Egg02]. Ausgezeichnete [Sch12b, Sch12a]. Ausnutzung [Hei46]. Auspices [Ano22a, LdB12a, LdB12b, Smy45b, Smy45c, Smy45d, Smy45e, Smy76a, Smy78, Smy89]. Äußerungen [Kru94]. Ausstellung [Rec02, RD05, Ren05e, Spe79]. Austin [Hen82b, Lew79]. Australia [BB89, CK^{+}04, Hom06, Kee04, Poc98]. Australian [Cha23]. Austrian [Ano93b, McL99]. Auswahl [Hah81, Sch49a, Sch51a, Sch55, Sch79a, Sch83a]. Author [Del11a, YCL20, Ein23h, Fla15]. Authority [Ano45a, Sch02a]. authors [Bes04, IRW31, Ren05c, Ren05b, vB73]. authorship [Fie15, KKKM05]. Autobigraphical [Ram79b]. autobiográficas [Sch84a]. autobiographic [HOO88]. Autobiographical [Bla50, Ein49a, Ein79d, Ein91c, For80a, Hei35, Sch79b, GR20b, Sch35c, Ein56a, Sch92a]. autobiographies [Ber74d]. Autobiographische [Ein56a]. Autobiography [Hoh13, vD17, Hah62, Hah66, Hah67, Hah70, HM11, Inf42, Inf80, Inf06, Kha12, Man13, PGH84, PGH96]. autonomy [HW12]. Autor [Fla15]. Autoren [IRW31, Ren05c, Ren05b]. Autorenkollektiv [Spe68]. autorovou [Ein23h]. autour [Com95]. autres [Gis12, Kle05, Ano05o]. aux [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Pat85, Per39, Seg84, Wei15c]. av [Gar41, Ves82]. Available [Ano90b, Hor79b, Flü39, Hen88b]. avant [Kie94]. avant-garde [Kie94]. Avantgarde [Kie94]. aventura [EI65a]. Avert [Hec16, Ree17c, Res50]. avez [Mon12]. aviation [Pia31]. Avogadro [Mil13]. Avoid [RFP^{+}45]. avoiding [GBE95]. Await [Jam30]. Awaited [Lau50]. Awakening [Mah09]. Award [Ano11a, Ano33c, Ano45b, Ano54a, LW34, Pin00a, CJT22, Ano21c, Ano54b, Bet63b, Ein48i, How21]. Awarded [Ano12, Ano50b, Ano56, Dys65, Ano79g]. Awarding [KUW82]. Awards [Ano63a, Eng05b, Adl12, CT15]. Away [Cam18, HM90, Smi21, BCD93]. ax [Ein06d]. Axel [Per09, LB02]. Axes [Bro83]. Axiomatic [Cor07]. Axiomatics [Mil99]. Axiomatik [Rei24]. Axiomatization [Cor04a, RR69, Rei24]. Axis [Hen00a]. axisymmetric [Bre89]. ayer [Ano25b]. Aynshtayns [Sch27]. Azarkh [Dix00]. Azcona [M08].

B [Ber79b, Con17, Fri73, Her82b, Her89b, Her93, Ken25, Mur25, Nag50, Rav97, Rob38, Rus43, vM05c, Ano11a, Ger78, Gib91a, Pos11, Sta02]. B-29 [Pos11]. B. [Sch95a, Sch96b]. ‘Baby [DC10]. Babylonians [Tay49]. Bach [Ein34m]. Bachelier [Dim93]. Bacher [Car08]. Back [Ano31a, JR07, dB36, DS01, Kai94, Mei64a, Mei64b, Mof10, Wei18a].

Background [HW07b, How92, How97b, Löw78, Tha77, Bel96b, CT15, Gon96a, How76, Smo07c, TTB20]. backlash [HS14]. Backs [Ano54c]. Bacteriology [SB82]. Bad [Van07c]. Baer [AH60, Ein26c]. Baerschen [Ein26c]. bag [Jak13, Joh14]. Bai [LR02]. Balance [GM00, Kol59, vE10]. Balancing [PA21]. Balashov [Pec04]. Balibar [Mat86a]. Ball [Ano39a]. Balleinrubin [BSS^{+}17]. Baltimore [Ham11a, Mil11a]. Ban [Bad95, Div78, Whi55a, Tau12]. Bananas [Smi97]. Band [ARK^{+}99, Eis00, Fle18, HvMW79, MGG^{+}95, PvM93, vL19, vL21, vL23, vL52, vL56a, vL65, vMHW85, vM93, vM96, vM99b, vM01, vM05c, Cas07, Kro01]. Banesh [For80a, Hei74a, Moy81, Ram79c, Wig79]. Banff [Isr73]. Bang [FH10, Lev07a, Ber84a, Mus15b, Ost15, SN93, Wal77a, Wal92a, Bar05g, Cle09, Haw88, HBN00, Kra96b, Kra92, Ler91, Mat07, Mit02, Pee14, Sch96c, Sch99b, Seg11, Sin04b, Sta95c, Wil13a]. Bangs [DM96]. Banquet [Ano81d, ACT23]. banquets [Söd10]. Bar [UM92]. Bar-Hillel [UM92]. Barak [Wac06]. barber [Wol00]. Bardeen [Hes04, HD02]. Bare [Ano39a]. barium [HS39a]. Bariumisotope [HS39a]. Barn [Ves82]. Barren [Ein39b]. Barrie [SE90]. barrières [Dat38]. barriers [Dat38]. Barth [Jüt21]. Bartlett [Ein35f]. Base [Ein14c, Mer01, Jac11]. Based [Ano23a, GRRS99b, HS89b, AS91, Bro19, Bro20b, BDH89, CZPR17, DCL23, EG14, Ein95b, EK92, FP40c, LLSH23, RN16]. Basel [Pec04, Ros86]. Bases [Ein11f, Ano23a]. Basic [Ano05m, Fow03, Roh87, Woo06, Che10, vM90, Ein05g, EL05a, EL05c, Sch35a, Ein15g, Ein23d]. Basics [Dat15, Fis11, PFM17, JDB^{+}97]. Basingstoke [The12]. Basis [Bha49, Ein13e, Ein05o, Zeh07, Bor22a, Ein14c, Ein14g, Ein14l, Kop21]. Basque [EI00]. Bass [OSWR86]. Bath [Bro79a]. Bathtub [Ree17c]. Battle [Bow87a, Fer14c, Kae45, Nor22b, Sus08, Ein44d]. Battled [Sau16, Hal15]. battlefield [Bru03b, Bus12c, Gro12]. Battles [Ham97, LE96]. Bau [Ano23c, Boh23, MTVB88]. Baumgarte [Pul11]. BBC [Mos16a]. Bd [HEB^{+}80]. ,,von [Som05a]. Be [Ano45b, Ano45a, Far03a, Lev01b, RFP^{+}45, Ros79a, Ano19b, Ano35, Cao07, Car12, Car15, CL19, Del11a, Die88, Ein18j, Ein19b, Ein20g, EM31b, EPR35, Ein35f, Enk18, Enz02, Far02, Flü39, Gio13a, Gow88, Gro62, Gro83, HMW24, Hen13, Jak13, Kie15, Mas74, Mau96, Mor95b, Sch19, Sch15a, vC05b]. Beach [Hag05, Mau09, Ros86, WM08, GW93, LRL15, LRL17, OV86, Bot13, Lan00]. Beam [Bus13a, Wei12m, Kni13, Nac73, Nac77]. beantworten [Ano69c]. Beaten [Fey05]. Beautiful [Far03a, GM04, Eng02, Far02]. Beauty [Dir79a, Pep04c, Cha87, Goo02, Mil01, Sch04c]. Became [Bar15, Mis85, Ein21u, Gor07b, Pow02, Ree15c, Sch19, Spo16]. Because [Mau96]. Bechthold [dP94]. Beck [Cah90b, Har91, Hen88b, Mcc87, Pye89, Pye95, Ste90, War96b]. Become [Ano20d, Ein39b, Ein54c, Kae31, Ano15f, GW10]. becomes [Bra17]. Becoming [Čap71a, Dor95]. BECs [Fri09]. Bedeutung [APWB95, MGG^{+}95, Sch15, Hah48]. Bedrohung [Sal95]. Bedside [Rei82]. beeinflusst [Ank15]. Been [Ano20f, Com46b, Kae31, Opp79a, Bre08, Dor78].

Beerhofmann [Ber94]. bees [RMI21a]. Before [Bor66, Bru82, Cle09, How21, Jan77, KK83, Mac86, Pre05, RMI21b, Thr17, Wyd84, Bal11c, BBM78, Cap05, Gea90, Gio20, Han79a, MSYW12, Mes83, Nik12, Pet72, Pla20a, RMI21a, Ros11, Sch03a]. begabten [Ste89]. Began [FW67, Kae48, Mas95, Dys16]. begann [Bre05]. Begegnungen [Hei75a, Jof67, RS13]. begets [Tol23]. begin [Bre05]. beginner [Doh06]. beginners [Ano05-42, SM79, Sch81b, SM92, Con83]. Beginning [Hel09a, Hoo77, MCW47, Rec86, Ven12, Wei12s, Ano86, Ano96a, Gef14, Gin04, Lum11, Mar95a, RP09, Sei03a]. Beginnings [Ber87b, Gon96d, Kat04, Rei86, Cra84b, Mat86b]. Begleitbuch [Rec02]. Begriffes [Ein28e]. begrunden [Ein20g]. Begründer [Hah79, Kle17b]. Begründung [BHJ62, Ein12f, Ein12i, Ein12h, Ein05y, Ein05-30]. BEH [Eng14]. Behandlung [Ang13]. behave [Ket02]. Behavior [Ein05q, Ein11a, Ein11g, Ein05i, Lin15]. Behind [Ano96a, Bar05f, Bru03c, Fri01, Kae38, Kra02, Smi58, Tur05, Ber05, Bri05c, Elz06, Gre96, How97b, Jak13, JJ81, LS92, LS94, LHZ^{+}23, Pic08, Wal06a, Wal02a, WM08, Wid07]. Behring [Enk18, Enk18]. beiden [Her66a]. Beijing [CHR96]. beijs [EdH15c]. beim [ES13, ES05a]. Being [Ano29d, FC87, Hec12, Eve39, Eve13, Gra94a, Gra96, Mac08]. Beispiel [Fla15]. Beitrag [Isa85, Kie94, Lan21, Pau55, WE24, WE87]. Beiträge [Ein22d, Ein05h, KRW05, SW12, Tre79a, Ein14d, vEPF09, vEPF22]. Beiträgen [Mei05]. Beitragsvermeidung [JDB^{+}97]. Bekampfung [Urb22]. Bekenstein [Bus12b]. Békésy [Kov01]. Belated [CRS97]. Belgian [Kra92]. Belgium [Ber84a, tH20]. belief [Ano81e, BNY81, Ben01, CF80, CH80, GO81, Moo81a, Moo81c, Moo81b, Pai88, Rus81]. beliefs [RL92, Moo81a]. Belies [Lau45]. Believe [Ein30q, Lan14, Dir80d, Leh14a, Ano35]. Believed [Ano45b, Pit16]. believer [Sto96]. believing [Hai10, Ryc14]. Bell [Nau17, Ano77, Ano78, CW12, CM89, Das00, EPW13, GHSZ90, Hel82a, Hel82b, How89, Lau95, Whi16]. bells [BSPKN03]. below [Mac85]. Bemerkung [Abr12a, Abr12b, Ein11c, Ein12a, Ein12g, Ein24b, Ein05d, Ein05-28, vI11b, vB19, Ein09a, Ein11d, Ein11a, Ein12b, Ein14e, Ein14f, Ein18c, Ein18d, Ein19b, Ein19c, Ein20d, Ein22c, Ein22b, Ein22d, Ein23b, Ein23a, Ein25b, Ein05g, Ein05e, Ein05h, Ein05f, Ein05i, vI11a]. Bemerkungen [Ein07f, EE22, Ein22m, Ein53c, Fra37, vI11b, Urb22, Ein07a, EL09, Ein05k, Ein05j, EL05a]. Bench [HRR03]. bend [FSE31]. bending [Jak78]. Benefits [Sch54]. Benjamin [Gor22, Sta06]. Beobachtbarkeit [ST94b]. beobachtet [JDB^{+}97]. beobachtet/Laserlicht [JDB^{+}97]. Beobachtungen [Ein14h, Hof18, vE10]. Bequest [Ano97, Bru03b, Ano98]. bereits [Bre08]. Bergia [Sie97]. Bergson [Eng17b, Her17, Tug98, Wil16a, Can05, Can15b, Can16a, Can16b, Čap71a, Čap71b, Fau79, Fol17, Gen98, Gun69, M.25, Vid17]. Bericht [vE10]. Berichte [KRW05]. Berichtigung [Ein07b, EL08a, Ein11e, Ein15c, EL05b, Ein05l, Ein05m]. Berkeley [Ano81d, Bed01a, HS89a, Kuh54, Sei99a, Tre88]. Berlijn [Ano20b]. Berlin [Ano79g, Ano99c, Bad15, B^{+}79, BR00, Ber07a, Ein31a, Eis00, For80a, GRRS99b, Hen86a, Hof06, Hol05b, Kev05, Kil98, Lac08, NHP^{+}79, Pye80, Pye97, Pye06, Ren05e, Row00, Sta00, Tre66, Wal99, Ano20b, Ano29a, Ano29b, Ano29d, Ano69a, Ano79g, B^{+}79, BSKS04a, BSKS04b, BSR^{+}06a, BSR^{+}06b, BRS^{+}09a, BRS^{+}09b, BIRS12a, BIRS12b, BIR^{+}15a, BIR^{+}15b, BIJ^{+}18, BRI^{+}21a, BRI^{+}21b, CGR^{+}97, Erb10, Goe05, GC20, Gru05c, HRP90, Hof06, Hof13, Hor79b, JSI^{+}02a, JSI^{+}02b, Kat04, KT79, Kle93, Kno97, KKS96, Lam79, LMOS15, Lev03, Lev04a, NHP^{+}79, RCD98, Sch95b, SKJI98a, SKJI98c, SKJI98b, Wei12s, Wei12t, Wei12u, van12]. Berlin/Heidelberg/New [Eis00]. Berlin/Hubert [Ano79g]. Berlin/Hubert-Schardin-Medaille [Ano79g]. Berliner [Kle93, Tre66, Ein32e]. Berliners [Ein32e]. Berlins [LMOS15]. Berman [EO79]. Bermerkungen [Ein11f]. Bern [Flü74, MK56, Flü74, Hen05, Sch95b, Med76]. Bernard [Tee02]. Berne [Lil86, Bus13a, CC96]. Bernese [HG05a]. Bernhard [Fan91, Goo85, Sei84]. Bernstein [Bad96]. Bertrand [Bro72, Kap61, Mar73, Ano16b, Ein44e]. Bertsch [Wat00]. berücksichtigt [CSW97a]. berühmter [Ste89]. beschränkter [Ein20h]. Beschreibung [Hen92a, Kie15]. Beschreibungsmittel [Kra21]. Beschreibungsobjekt [Kra21]. besonderem [EM34]. Besprechung [Ein30d]. Besso [Ein94b, Kle73, Ano05k, Ber89, Ein94b, Fin05, Jan07, KH05, KF74, Med94, Spe72, Tho05, Wes05, Whi05]. Best [Pit17, Dur05b, Fey00, Hes15b, Pur63]. best-kept [Pur63]. bestaan [EdH15c]. Bestand [Egg02]. bestätigt [Kon15]. bestemte [Jak13]. besten [Hes15b]. Bestimmt [Ein27b]. Bestimmung [Ein05a, Ein06e, Ein06c, Ein07i, Ein11e, EM23, Ein05m, Ein05-32, Ein05-36, Ein05r]. Bestrahlung [HS39c]. Besuch [Fre27a]. beta [BRVF84]. beta-decay [BRVF84]. Bethe [BL09, Sch14a, Ano12d, Bet97, Bur68, MB66, Sch14a, Sch00d]. betrachtet [Kie15]. Betrachtungen [Ein11h, Ein17e, Ein05t, HBD71, Kra20b, Mos21, Nic17, Nic19, Cas21]. betrayal [Mil05a, Mil05b]. betrayed [Gro49, Pas10]. betreffend [Ein22g]. betreffenden [Ein05d, Ein05-37]. betreffendes [Ein21s, Ein26d]. better [LOvT12]. Between [Bar99, Ber94, Bro23, Ein05-39, Fau79, Göd49b, Jan95, Kie82, Lan07a, Lan15, Mas02, Mur11, PA21, Pop66, RMI21b, RSS^{+}07a, Sch07b, Stu18, Van22, VW86, Acu14, Ano07g, Bal85, Bar49, B^{+}01, BBL08, BM09, BN88, Bla02, BB71a, BB71b, BB05, Bro72, CS79b, D’A97, Dör05, Dre87, Ein02b, Ein11a, Ein11g, EdS32, Ein05i, Ein05q, FF97, Gan97, Gio20, Hei75b, Hen91, Her66b, HW12, Hub29, Hug54, Ill89c, Inf99, Lam79, LHZ^{+}23, Müh11, OW54, Pul06, Pye82b, RMI21a, Ros79c, Sca14, Sch02a, Sch05b, Sch19, Sch35b, Sch36, Sen21, SE90, SE45, Tym95, Val06, Wal81, Wei12o, Wil01, Wil06, Wil14, WJ21, dM86, dWtH20, Hof08c]. Bewahrer [Jun08]. bewegte [EL08a, EL08b, EL09, EL05b, EL05a, EL05c]. bewegter [Ein05e, Ein20q, Ein05-41]. Bewegung [Ein05c, EH10a, Ein27b, Ein05-35, EH05b, Ein05-42, vS21a, vS21b, Tes69, Ein06b, Ein07f, Ein08b].

Bewegungen [Ein22o]. Bewegungsgesetze [EG27a, EG27b]. Bewegungslehre [Rei21]. Beweis [Büc21, EG23]. Beyond [AC09, Ano04g, Bag04, Bar05g, Cal17, CF11, Cas09, Dat15, KT95, Lau88, Law07, Mad95a, Pap15, Rod19, SL90, SSM21, Smo19b, SMHL23, BW04, Dar06a, Gro12, Grü08, HS06, Hei71b, Hei71a, HB01, Hsu00, HB05, KMN^{+}97, LD08, Mas74, Mat07, Mof08, Sol16, Vec17, Dro20, Wal02b]. Bez [È03]. Beziehung [Ein11a, Ein11g, Ein27a, Ein05i, Ein05q]. Beziehungen [Hei75b, Her66b, Lan21, Sch88b]. Bianchi [Ein50a]. bias [Van07a]. biased [Fri12]. Bible [Ano97b, Com14]. Bibliographical [Ano60]. bibliographies [Egg02, Kli62b]. Bibliography [EKS^{+}08, Hor79b, KSB^{+}09, Shi49, CB84]. bibliometric [GW10, HMW24, TA17, ZXLX14]. Bicák [Fis79]. bicentennial [Woh76]. bicicletta [Kle17a]. bicycle [Kle17a]. Bid [Res50, Jea08]. Biennial [BC71]. Bietet [Ein23c]. Bifurkation [MGG^{+}95]. Big [Ano39g, DM96, FH10, Hür22, Jun14, Kra96b, Ber84a, Lev07a, Mos13a, Mos13b, Sha79b, Che00, Duc11, Fre27a, Gru05b, JMH^{+}05, Mus15b, Ost15, SN93, Wal77a, Wal92a, Bar05g, Cle09, Haw88, HBN00, Kra92, Ler91, Mat07, Mit02, Pee14, Sch96c, Sch99b, Seg11, Sin04b, Sta95c, Str99, Wil13a]. Big-Bang [Ber84a, Ost15]. bigger [Dob06]. Biggest [OM18, Blu22, Bod17, Wei13e]. bigotry [Nav21]. bilayer [EM04]. Bilddokumentation [Ber69b]. Bilddokumenten [Her01b, Ber74d, Her76b, Her84]. Bilder [SKD^{+}87, KDK87]. Bildern [Hah79, Hah88]. Bildung [JP52, Kra88]. bilim [Tip12]. Binaries [Nun22]. binary [Ao04]. binding [McP74]. Biochemie [BS92, NS88]. biochemistry [Nac79, NS88]. Bioenergetics [All02]. Biografie [BHJ^{+}00, Enz05, Ris06]. Biografie/Kuhn/Försterling [BHJ^{+}00]. biographer [Ano00e]. Biographical [Bar05f, BMW89, Ein84a, Rei30, Rei31, Abr88, Abr01b, BW^{+}87, Bro20b, FMR90, Gha82, Hen92a, Kat89, Kur02, MB^{+}92, Pri90, SBB16, Tho97, WB^{+}87, Lin56]. Biographie [Hah79, Hen11c, Hen21, Rec09, Rec10, TGC91, Föl93, Her94a, Her04, HC94a, HC96a, HC04, Nef05, Nef06, See54, Sto98]. Biographies [McL99, Rad08, SS98c, Wei12g, Ano64, Ano65b, Ano72, Don70, FM92, Gla00, Lun92, Nob72, Nob64b, Nob66a, Nob66b, Per97, SS98a]. biographischen [Hen92a]. Biography [Ano07p, Ano07q, Ano07r, Ano07s, Bod05a, Bod05b, Enz05, Fey64, Föl93, Gam61b, Gam61a, Kaw34, Key33, Key89a, Man86, Met95, Nef05, Nef06, Wei12h, Wig96c, Bal75b, Bod00, Bod16, Bri05a, Bro97, Bus12a, Cah55, CL05, Cur37, Cur86, Cur01, Enz02, Fan93, Föl97, Föl98, Gav95, Hah79, Hal59, Her94a, Her04, HC96a, HC04, How97a, Key13, Kra90, LS92, LS94, MM00, Nef07, Pea49, Pip97, Rec10, Ris06, Sch05a, See54, See56a, Sin04a, Sto98, Sug89, Val54b, Wal91a, Smo07a, Lig97, Sto17]. biological [BNY81, DSW99]. Biologie [Nic17, Nic19]. Biologists [Dei96]. Biology [Fru82, Jab11, Nic17, Nic18, Nic19, RMI21b, Bal77, MO95, RMI21a, ZDSF11]. biomedical [HH02]. biophysics [NKY15]. biotechnology [Bau95, Bau97]. Bird [Den84, Pei85]. Birds [Dys09, RMI21a]. Birkhäuser [Gly90, Pec04]. Birkhoff [Bir82]. Birth [Anoxxa, Anoxxb, BW01, Del11b, Gli84, Hub13, Kos03, Mer79, Pye10, Sta19, Ano79c, Ano80a, BF02, BD92, Cla61, Ein24k, Far05c, Hel80, Kel07, Kel09, LHDH21, MW99, McM05, Mel09, Sto96, Swa92, Swa95]. Birthday [Ano29g, Ano46a, CS79b, Ein32e, Ein39f, HBD71, KRW05, Lau55b, Mil49a, Opp39, Opp79c, RH02, Ros79c, Som49c, Tre79a, Wig96b, B^{+}79, EVB^{+}72, FPLR59, Har79c, Jac08, MB66, NHP^{+}79, Rec02, SKD^{+}87, Som49b, Spe79, Wie06, SS97, Lin60]. Birthdays [EHM^{+}79, Ano79h, Ano79g, Lam79]. Birthplace [Ano65a]. bis [BS92, Büh98b, EI61b, Ern92, Fis10, Fis12, For80a, Hen11c, HH14, HW78, NS88, NHP^{+}79, Rec09, Rec10, Ren05e, Sal95, Sch17c, vL56c, vL56d, vW02, vWR04, Bod03]. bisher [Ein19b, Ein19c]. bisherige [Jüt21, Pla20a, Pla20b, Pla22a]. Bit [Ano39h]. Bite [EMH96]. Bite-size [EMH96]. Bittere [Hop07]. Bivector [EB44, Ein44a]. Bizarre [Juc04, Abr01a, Abr02, Abr03, Abr04a]. Björkborn [BCS82]. Black [Ano18b, Bar15, BR21, BT12, Ber96d, CB15, Del11b, Eis93, Fle18, Hub13, Ibu79, Imp19, Lev16, Pen79a, PD15, Pye10, SL91, SL05b, Sus08, TPM86, Tho93, Tho94, Wal98, Whe79a, AAA^{+}18, AMSmS96, Ano10d, Ano18c, BCO^{+}11, Ber96b, Bes04, Bro10b, Cha88, Cha98, Cho08, Gen05, Haw88, Mel09, Mil05a, Mil05b, Mus15b, Shi75, Sri99, Str05b, TW00, VWA95, Wal77a, Wal92a, Wal94a, WF98, BPL05, Har95]. blackboard [TV07]. blackboards [VT22]. Blackett [Nye99, Nye04]. Blair [Kar05]. Blandford [Vog21]. Blast [Ano39h, Lau45]. Blick [JDB^{+}97]. Blind [San81a]. Bloc [Sch54, Dol18]. Block [Ano81e, Dol18, Oyi04, BNY81, GO81, Moo81a, Moo81c, Moo81b, Rus81]. Blocks [Noe11, Coo10, Coo15, ’t 97]. Blomeyer [Hei53a]. Blooming [Dan05]. Blower [Ste95b]. Blowing [Div78, Ano14c]. blows [Hoy94]. Blue [Jun14, LU13, Aka15, CSvB99, Mer85a]. blues [Lev16]. Blunder [BFB82, OM18, Wil13a, Ber90b, Blu22, Gol95b, Gol97, Gol98a, Gol98b, Nav21, Wei13e]. blunders [Liv13]. boardroom [HS06]. Bobby [Bra12]. Bodanis [Sto17]. Boden [Inf31]. Bodies [Ein11h, Ein05q, Ein05-41, Ein05t, Eis93, Wil77, Ein05e, EL08a, EL08b, EL08c, EL09, Ein11g, Ein18j, Ein20q, Ein56i, Ein94e, EL05c, EL05d, Eis07, Lor21, Nor14, Rya06, EL05b, Ryn05, EL05a]. Body [ER36, Fok49, Amo10, Ein05b, Ein05v, Fra37, Gut98, vI10b, vI11a, Pra06, Rob38, vS21a, vS21b]. Bogoliubov [OKSD05]. Bohemia [Gor20, Sta21, Wis21]. Bohm [Pat93, Vig94, Vig95, ZDM08]. Bohmsche [GRE^{+}01]. Bohr [Adl03a, Ano23c, Ano39a, Ano93a, ARK^{+}99, Duf46, Eis00, Fer07a, For09, GRE^{+}01, HEB^{+}80, Hel98, Hen86a, HvMW79, MGG^{+}95, OSWR86, Par06, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Sud11, Tur06a, vMHW85, vM93, vM96, vM99b, vM01, vM05c, Aas90, Aas02, Alb92, Ano23c, Ano93a, Bal85, Bel93, BF94, Boh85a, Bok08, Bol04, Bro71, Bub73, Bun79b, CW12, CS79b, CG14, Dan89, Dör05, Erl72b, FG21, FF94, Fol96, FK85, Fre86, HK69, Hel98, Hen84, HvMW79, Hol05c, Hoo91, Jam74a, Kai94, Kel58, Kle70c, Kle10, Kon15, Kra79, Kum10, Kum11, Lan06, Lau13, Lin07, Lin08, Mar88, MR82d, MR82e, Meh87, Mur90, Mur94, New09, Nik12, OBxx, Pai91, Pat85, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Per85, Per21, Rös95, RRK^{+}64, Ros79c, Rot87, Sac81, Sac88, SM08]. Bohr [VW86, Whi96, Whi04a, Whi06, vCK17, vMHW85]. Bohr-Festschrift [Ano93a]. bold [FFF09]. Boldest [Lau55a]. Bolles [Kar05]. Bologna [Per10, Seg91, Tug98]. bolt [Mer85a]. Boltzmann [Bad09, Faz15, Sch00b, Yos95]. Boltzmann. [Gio20]. Bolyai [Mar15]. Bomb [Ama02, Ano45a, Ano45d, Ano94a, Ano95a, Ano00c, Asa95, Bac50, Bag10, Ber76b, Ber88, Ber93, BC95, Ber03, Bet50a, Bet50b, Bet63a, Boy85, Boy94, Bro15, BM77, Buc93, Cir07, Cir08, Com46b, Cou79a, CR08, Dix00, Duf46, Dys76, Ein45a, Ein49b, FS95, Fie07, Fre08, Fus88, Gol95a, Gon96d, Ham97, HG92, Hei95, Her00, Her03, HM90, Hoh13, Hol81a, Joh46, Jun14, Kae45, Kap20, Kel04, Ken54, Klo97, Lan08, Lan15, Lau45, Lau55a, Lef95, LX88, Lit09, Luc11, Mal09, Mal12, MD67, MM58, Mor79a, Mor57, Nef11, Ove96, Pal95, Ree06, Ree15d, Res50, Rho86, Rho88a, Rho88b, RT90, RE90, Rho95, RA00, Rho12, Rid50, Ros79b, San95, Sch03c, Sig78, Smi58, Str13b, Swi45, VB96, Wal90b, Wel08, WHF^{+}96, Wol67, Wol62, Wol09]. Bomb [You95, You09, vB02, Abr98, Abr09, Alp65, Alp85, Alp94, ATW^{+}95, Amr59, AP01, Aue95, Bak68, Bak76b, Bak76a, Ber83, Bla48, Bla49a, BHM91, Bro97, Bun82, Bun88, Bun90, Cas09, Cla61, Com56a, De 04, De 05, DMPS81, Dee08, Del09a, Dre93, Eli50, Els95, Erm89, Eva53, Far13a, Far13b, Fei61, Fei66, FGMvH14, FV13, For15, FP40c, FP40b, FP40d, Fus90, GS50, GF65, GF67, Gon96a, Gon00, Goo03a, Gos94, Gos99, Gos01, GA79, Gow88, Gow93, Gro67a, Gro67b, Gro68, Gro00, Hal65, Hen00a, Her80, Her82a, Her88, Her02, Hir98, HM97, Hol94, Hug04, Irv67a, Irv83, Jak13, Jen02, Joh17, Jon85, Jon01, Kel06a, Kel07, Kel09, Kha12, Kno90, KSG63, KMN^{+}97, Kur06, Kur86, LS92, LS94, Lau46, Lau72, Lil63, LS93, Log96, Lor08]. bomb [Mac85, Mai99, Mal08, MW46a, MW07, McK03a, Mel00, Mos68, Mos87a, Mos87b, MM83, Nor02b, Nor03a, Pal98, Pos11, Pou93, Pow93, Pow94b, Pow00, Pur63, RS09, Ree15c, Ree15a, Ree17a, Rho88c, Rho93, Ric06a, Rol47, Rot08, Row85, RS07f, Sal95, Sch61, Sch15a, Sch00d, Sei95, SR92, She75, Ska94, Smy45b, Smy76a, Smy78, Smy89, Sti47, Sti76, Tak95, Tau12, Tay95, Tay07, Thi46, Van03, VWA95, Van71, Vig83a, Vig83b, Wal92b, Wal95, WG17, Wil11a, vCH18, vS14, BW91, For01, Ham97, Koj08, Wal17, LRD^{+}91, Duf46, Fre08, Her21, Jon01]. bomba [RA00]. Bombardment [Fer39, Fer70, Fri39b, MF39b]. Bombe [Sal95, Ano05-40, BW91, Buc93, vS14]. bombed [Ful84b, Ful84a]. bomben [Jak13]. Bombing [Gol96]. Bombings [Ano08b, Ber95, Man5 a, Man5 b, Ano81c]. Bombs [Can00, Wal16, CM02b, Hab02, Jay90, Ndi07, Rus90, Seg95, Szi45, Wal97, Wal04a]. Bombshell [AK97]. bomby [Smy46]. Bond [Ein40d]. Boni [Ano60]. Bonn [BR00]. Book [Ach13, Alf09, And81, Ano19b, Ano29b, Ano39d, Ano50a, Ano60, Ano84a, Ano84c, Ano84b, Ano99b, Ano99c, Ano04b, AA75, Arm38, Bad72, Bad83b, Bad96, Bad15, Bal72, Bal75a, Bed01a, Ber07a, Ber06b, Ber03, Ber00, Bey06, Bey09a, Bey09b, Bey12, Bla76a, Bla76b, Bok15, Bot13, Boy88, Bri18, Bri82, Bro15, Bro66, Bro72, Bro39, Bro00, Bru03c, Bru06, Bus13a, C.51a, Cah90b, Cah90a, Cao06, Cas05b, Cli09a, Coh51, Coh54, Con83, Coo04, Cus82, Del09b, Del11b, Dix00, Dro20, Ebu16, Eds82, Ein30d, Far79, Fel11, Fer07a, For71a, For01, For05, For70b, For70a, For80a, For09, Fre08, Fri73, Gea94, Ger15, Gil12, Gil19, Gin82, Gli84, Gly90, Gol84a, Gol84b, Gon05, Goo85, Goo02, Gor02, Gor07a, GRC18, Gre19a, Gre19b, Hah73, Ham11a, Ham97]. Book [Han80a, Har07, Har91, Har96a, Har96b, Har02b, Har79d, Hea13, Hea64, Hei74a, Hei79a, Hei79b, Hen11a, Hen82b, Hen86a, Hen88a, Hen88b, Hen00b, Hen11c, Hen16, Hen21, Her17, Her21, Hes04, Hie68, Hie84, Hoc93, Hoe18, Hoh13, Hol05b, Hou65, Hus15, Ihd74, Ivr99, Jam70, Jan73, Jeh51, Joh95, Jon01, Jos95a, Juc04, Kae38, Kai05a, Kar05, Kat52, Kau14, Ker89b, Kev05, Kil84b, Kil95, Kil98, Kle70b, Kle70a, Kle73, Kle84, Kle14a, Klo63, Kni12a, Kni12b, Koj08, Kra02, Kri08, Kri99, Kuh54, L.89, Lac08, Lan08, Lan15, Leg15, Len39, Lin60, Lit09, Loe88a, Loe88b, Maa13a, Maa13b, Mal65, Mar73, Mar10, Mar11b, Mas01, Mau09, Mcc87, Med76, Mel82, Mic68, MCW47, Mil11a, Moy81, Myr21b, Myr21a, Myr21c, N.38, Nad79, Nav15, Nic20, Noe93]. Book [Noe11, Nye88, Oha79, Ove96, Par06, Pec04, Pep04b, Pep04c, Per10, Plo06, Pri99, Pul11, Pye80, Pye86, Pye89, Pye93, Pye95, Pye97, Pye06, Pye10, PJMS11, Qui40, Ram79b, Ram79c, Ram79a, Ram81, Ram90, Rec86, Red84b, Ree15d, Ree17d, Ree17c, Rei86, Ren21, RS11b, Row00, Row06a, Row06b, Row08, Row17, Rus17, San07, Sar35, Sau16, Sch87a, Sch50a, Sch79c, Sch09a, Sch07c, Sch92c, Sch03c, Sci80, See65b, See65a, Seg91, Sei79, Sei84, Ser88, Set10b, Set10a, Set11, Sho21, Sie97, Sie07, Smi05a, Smi51, Smi89b, Smi89a, Smi02, Sta05a, Sta99b, Sta00, Sta20, Sta21, Ste95b, Ste22, Ste05d, Ste90, Sto17, Str04, Str90, Swa02, Tha77, The12, Tho11b, Top89, Tou05, Tra00, Tre88, Tug98, Tur05, Tur06a, Van22, VB96, Vis94, Vis95, Vog20, Vog21, Wal94b]. Book [WM08, Wal02b, War95, War96a, War96b, War03a, Wat00, Weg08, Wei94, Wei03, Wid07, Wil16a, Wis21, Woo06, Wre66, Yan87, Yea08, You95, dP94, vB01, vB02, Aer99a, AMM99, Aer99b, ABW99, AVvdV99, Ano04a, Bar00a, Bar00b, Bar05d, Bar05e, BAM99, Bus12b, Cor15a, CSvB99, Dan00, Dea13, Her03, HBR99, Kel06b, LPPT75, PP03, Rec02, She15, She16b, Wei15c, dGK11, Ein24c, Ein24i, Hub13]. Books [Ano60, Ano05-31, Bar05f, Bru03c, For80a, Gam42a, Hea13, Hen82b, Hol95a, Kap61, Kar05, L.89, Nah01, Ree17d, Sha54, Tur05, Woo06, AFG90, Hah81, Lew79, Ram79a, WK12]. Books/Henry [Bru03c]. boost [DC12]. boosted [Oha19]. booster [HMW24]. boosts [Kin95]. border [Sch05c]. Bored [Bro21]. Borges [B^{+}99]. boring [Bai07a]. Boris [Kie15, Kie15, Sta06]. Born [ABdB^{+}53, Bey06, Bre89, Bro72, Cas05b, DMPS81, Del09b, For70a, Kle70b, Kle70a, Kov01, Lom05, Mar73, Sch05c, Som05a, Som05b, Sta05d, And81, Ano05f, Ano05h, Ano05i, Ano05j, Bor20, BB69, BB71a, BB71b, BB05, Bre73, Bro72, Car89, CS79b, Ein91d, Far79, For70a, Gre05b, Han80a, Hei79a, Hei79b, NJ05, Pei71, Ros79c, Sch05c, Sch79c, Som05a, Wei12w, Kle70b, Kle70a]. Bose [Ein25a, JDB^{+}97, AK02, BSSP^{+}08, Bel23, Big05, BRH^{+}09, BDR^{+}07, BEC99, BLI^{+}02, BR99, Che00, Col95, CSD^{+}07, CW98, CW02, DGPS99, DSZ12, Das12, DDD^{+}06, DHY10, DCC^{+}01, DCTW02, Ein25a, EM04, EG98, EGW^{+}08, Faz15, Fet09, FS01, FMM96, Fri09, FKZ91, GSPH00, GRT08, GR44, GLG09, GMHB02, GRJ03, HES^{+}12, HHHR01, Hau08, Hea11, Hul01, IAS^{+}98, Jak08, Joh13, JDB^{+}97, Kap40, KRK^{+}06, KF03, Ket99, Ket02, KSVW10, Leg01, Leg03, Lev98, Lev01a, Lev01b, LLSS07, LJGS11, Lon38b, Lon38a, Meh75b, MBD39, Mon19, MO06, NBDD^{+}12, Osh06, OKSD05, PBM^{+}12, PSM^{+}13, RDF^{+}08, SHL^{+}06, SLLS09, SP59, SDCZ01, SISK^{+}98, Sto02, Tom22, Tom99, Wal06a, WNS^{+}08, WHL^{+}05, WH99, Wil11b, YCKG11, ZPSK10, Zol02]. boson [Eng14, Kra14b]. bosons [FKZ91]. Boston [Gly90, Hen82b]. Both [Vid17]. Bottom [Gon05, Lau05]. bound [Pai86]. boundaries [Gro56]. Boundary [LB49, Har13a, Wic95]. boundless [Bar05d, Bar05e]. Bounds [BF16]. Bouquet [Ber18]. Box [Wac06, Har13a, Pat01a, Tre75a, Jet77, Jet07]. Boxes [Hel15]. Boy [Sha14, CM02b]. Boyer [All02]. Boyhood [Tal32a]. Bragg [Jen01]. Brain [Ano13c, Bre08, DSMH85, Fal09, Gal99, Hin98, KaS15, Lau55b, Lep18, Mos16a, Sal99, Sei99b, Vea01, Wad81a, WKH99b, Abr01a, Abr02, Abr03, Abr04a, Ano00a, Cec86, Col00b, Cra00a, Dob06, Goo00, KWGH97, Nes60, Pat00, Pat01a, Pat01b, Pea06, Sti13c, Van96, WKH99c, WKH99a, Ano00b, Pep04c, Juc04]. brain-algorithmic [Van96]. brainchild [Par00]. Brains [Wal16]. Bramcote [Dix00]. Branch [Rob49a]. brand [Bro88]. Brandeis [Sch03d]. Brandmüller [MGG^{+}95]. brane [Ran05]. brane-worlds [Ran05]. Braun [Bri82, KS81]. Braunschweig [Bäu83]. brave [Fre15a, FFF09]. Brazil [GPSR01, Pat99]. Brazilian [KV11]. Bread [Bor71]. break [Ein47c, Tho15]. Breakdown [Mel00]. Breaking [Blu04, Sto02, Wal07, Adl82, Adl83, Nam09, Nam10, SHL^{+}06, Wei12i]. Breakthrough [Ano18b, Lin15, Spa58, Ven94a]. Breakthroughs [Cre10, GR23, Ano14c, Lig05, Mey07, RWL23, Rob10]. breathtaking [Sch04c]. Brechungsexponenten [Ein18j]. Breit [Sch87a]. Brenni [Pye95]. Brésil [Pat99]. Breslau [Ein48g]. Brian [Sta99b]. bridge [Cra97, Sch19, Chu60, Chu98b, Chu01]. bridge-building [Sch19]. Bridging [CCG93]. Bridgman [C.51b]. Brief [Axe50, Ben11, Fol15, Ill89c, McG19, SE79, Ano05n, Bro79b, Dur05a, Ein21b, Enz02, Haw88, Haw92, Haw98, JN14, JDB^{+}97, Tod14]. brief/Max [JDB^{+}97]. brief/student [JDB^{+}97]. Briefe [ES68, Ern92, For70b, Grü90, KRW05, Klo63, Kra25, Prz63, Som05a, Som05b, Win69, IK03, Kil65]. briefer [HM05a, Sch99b]. Briefes [Büc21]. Briefing [Ano99f, Ano00f, RWY97, Sti98]. Briefwechsel [ARK^{+}99, BB69, EB82, Eis00, GRE^{+}01, HEB^{+}80, Her66a, HvMW79, Her63b, LS05, MGG^{+}95, OSWR86, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Win69, vMHW85, vM93, vM96, vM99b, vM01, vM05c, EF33, For70a, Kle70b, Kle70a]. Briefwechsel. [ES68, For70b]. bright [Ano10d, BSSP^{+}08, MRS17, Sow01, See56b]. Brighter [Jun58, Jun82]. brightest [BT15, Bra12]. Brilliant [Liv13, Bai07a, Fey95, Mag09]. Bring [Ein50d]. Bringing [Kai94]. Brings [Ber87b]. Brink [She15, BCD93]. Britain [Arn00, AP01, Bay94, Cla61, Far13b, Gow64a, Gow64b, Gow88, Jon01, Lee06, McM14, Moo81b, Moo10, Sto12, Wel93, Jon01]. Britannica [Ano01a]. British [The12, EG80b, Gol70a, Lau12, Sim22, SR04, Sza92a, Sza92b, Wis82]. Broadcast [Ein29a, B^{+}06, Swi45, Whi67]. Broadcasting [Ein30i]. Broch [Ber94]. Brochure [Ano79g]. Brod [MGG^{+}95]. Brodmann [Kan92b]. Broglie [Cas75, Fes56, For80a, Hum51, L.58, Vig94, Vig95, Ano48b, Ano55c, BH79, Elb13, FS98, Geo53, Hou16, RF69, Smo19a, Vig93]. brokers [Fie15]. Brosche [HEB^{+}80]. Broschüre [Ano79g]. Brose [Bel26]. brother [Rob01, Rob14]. Brotherhood [Her02, Ber03, Her03, Sch03c]. Brought [Ein49d, Ein70b, Ano02a, Swi91]. Brouwer [vN37]. Brouwerschen [vN37]. Brown [Bru68, PBM^{+}12]. Brownian [Ein22o, LG97, Woo28, Bru68, Cha49b, Dim93, Ein06b, Ein07f, Ein08b, Ein56c, Ein05-42, Haw02a, Ian93, Kub86, Lan08, Lav85, Mai90, Per10, Per05b, Ren05a, SP12, SMHL23, Tes69, vS06]. brownien [Lan08, LG97]. Brownsche [Ein07f]. Brownschen [Ein08b, Ein22o, Ein05-42, Tes69, vS06, Ein06b]. Bruce [Fel11, Joh11b, Nuv11, Sho21, Ham11a, Mil11a, Pin00a, Set11]. Bruchstücke [HS39a]. Bruno [Rei41, Clo15]. Brunswick [L.89, Seg91]. Brush [Ano05q]. Brussels [LdB12b]. Bruxelles [LdB12a, LdB12b]. bubble [Cha11]. Buchbesprechung [Krö53, Wil36]. Buchbesprechungen [MGG^{+}95, OSWR86, Win69]. Bücher [ARK^{+}99, GRE^{+}01, HEB^{+}80]. Büchern [Hah81]. Buchrezension [Ein24c]. Buchtipp [BHJ^{+}00]. Buchwald [Bad15, Lac08, Mar11c, Pye95, Pye06]. bucket [BP95, Jan12]. Buddhist [Bha92]. budget [Dum04]. Buenos [Nir25]. buggy [KS12]. Build [GB89, Her21, Hol81a, SR92]. Builders [MD67]. Building [For15, Gol95a, Sch54, Hol96a, Oyi04, Sch19, ’t 97, Ree15d]. Builds [LX88]. buildup [TEM^{+}89]. Bulk [OKSD05, Eck70, Eck73, GLW15, OBxx, Rabxx, vN96]. Bulletin [GR63, GR65, Ano87b, LLH^{+}80]. bumpy [Jan05]. Bund [Ein40d]. Bunde [MGG^{+}95]. Bundesrepublik [Sch05c]. Bundle [Bri18, RdO07, RdO16]. Bunge [Mah01]. Bunin [Ano33c]. Bunnies [DRB05]. burdens [Hol86, Hol98b]. Bureau [CLE27]. Bureaucratic [Sig78]. Bureaucrats [Tor82]. buried [Ano17a]. Burlington [Wac06]. Burmese [Joh11a]. Burov [Jos95a]. bursting [HS39b]. buses [Ano09a]. Bush [Zac97, Zac99a]. Business [Jun14, Koc01, Lun74]. businessman [Tur06b]. Bust [Lei00]. Butsurigaku [EI40a]. büyük [Gol97]. Bye [Mit02].

C [Adl03a, Ano87b, Ber84b, Cah90b, Duf81, For80a, Hen82a, Hen82b, Hol03a, JP52, Kle82a, LG97, Pye80, Pye89, Ram79a, Ste22, You95, Ryc92, dP94]. C. [Del13, Erm89, Gie05, Hor79b, JP55, Rus43, SR99, Sin02, Sin04a, Ven88]. Calaprice [GRC18]. Calcul [Per39]. Calculated [Ano19b, Lam21]. Calculating [Ste16]. calculation [BP79, EH10b, Ein15b, Sla79, Per39]. calculational [AMS04]. Calculations [Lid67, Tur67]. Calculus [EF05, EF13]. Calendar [RR17, Lin17, Qui05]. calf [Nik06]. Calibrating [Lev07a]. California [Bed01a, Cli09b, Kuh54, Ano41, Car04a, MA31]. call [Kle75b]. Callahan [TH79]. Called [Cor12, Cam18, Ein26c, See25]. calling [Ano05-32]. Calls [Ano29b]. Calmthout [Hoe18]. Caltech [Hof97a]. Cambridge [Bar05f, Bru06, Cli09a, Gly90, Har85, Hoc93, Hus15, Jan73, Kai05a, Nav15, Par06, Pri99, Rec86, RK94, Sci80, Tur06a, Wei94, vD15, Ano16b, Cat04, JL14b, JL14a, MMMM96, Oli72, Ste22, War92, War93, War03b]. Came [Del11a, Goo90a, Bod03, Eĭn85a, Fre81b, Log96, Mos68, Sch15a]. camera [Ano05-35]. Cameron [Sho21]. Camp [Mer01]. Campaign [Cau91, Goe93, Spo02]. Campbell [Cam18, Cre83, War92, War93]. Campo [Mas02]. Can [Ano35, Bru76, Bus13b, CF12, Dar21, EPR35, Ein50d, FH10, Lev01b, Mas74, Ros79a, Sta99b, BT15, Bra14b, Die88, Dob06, Ein20g, EM31b, Gio13a, Gro62, Gro83, HMW24, Hen13, Kie15, Lin15, Mau96, Ein18j, Flü39]. Canada [Ano05-40, Inf78b]. Canaday [For01]. Canadian [Cha23]. canal [Ein26b, Van07a]. Canales [Eng17b, Wil16a]. Canaries [War99]. Cancer [Cre15, Kae39, Mil08, Mil13b, Sem08, Kut09, Leo09]. Candid [HH00, HH02, HH03, HH04a, HH05, HH06]. Candidates [Bar98b]. Cannon [Ano39a]. Cannonball [Ano39g]. Canon [Nor00]. Canonical [BB49]. can’t [Gio16]. Canton [For05]. Canty [BHJ^{+}00]. capability [Bay94]. Cape [Ber06b]. Capillaritätserscheinungen [Ein05u, Ein01]. capillarity [Ein01]. Capillarization [Ein05u]. Capitaine [Hau48]. Capital [Tom10, Wal16, Bor04, HS06, Kar10]. Captain [Hau48]. Captives [Klo97]. Caputh [Lut05]. car [GB11, Nac24]. Carbon [Hoy54]. card [Dac11]. Cardinal [Her00, vB02]. Cards [Ghi05]. care [CF09, FL88]. career [CT15, CMT18, LHDH21, She16b]. careers [HF22, VM05a]. Carl [GRE^{+}01, Kie82, Kie94, Sei84, van12, Bjö01, Eck70, Eck73, Goo85, HH14, Kie94, Sta79]. Carlo [Ben16b]. Carnap [BC71, SW94, How96a]. Carne [Ren21]. Carnotite [Bla13]. Carolina [BCEP94]. Caroline [Lem09]. carpenter [Jak13]. Carrel [Ano12]. carrés [Pia31]. Cartan [Ano84a, Deb29, For80a, Deb79, Kil80, Sch19, VT99]. cartel [Du 52]. Carter [Bon94, Smo07a, Ald93, Bur94, Cas94a, Din93, Gea94, Kan94a, Mil93]. cartes [Lam07]. Cartoons [L.89, Har89]. Case [Ash11, BM93, G.33, Gin15, Gol68, Han79a, Jan02, Ken54, Koh95, LU13, Lyn33, Maa07, Mar05, Nug99, SS16, Bru97b, Bru98, CSC23, DMPS81, Dim93, EGR11, Gol83a, HMW24, Hel82a, Hol78b, Hoy77, Hoy79, LR14, Mai90, Nam09, Nam10, Nat05, RWL23, Rob14, Sch10b, Tur03, Žáč12, tH20]. cases [Cec01]. caso [CM02a]. Cassidy [Cah90b, MGG^{+}95, Pye89, Ste22, Bad96]. Cassini [BIT03]. Cassirer [Bre15, Bia16a, Bia16b, Bre15, Ryc99, Sta16]. Cast [Ano29a]. Castellana [Del13]. Casting [Hon04]. castnaja [CH98]. Cat [Ano15e, Ano15d, Sau16, Bru97b, Bru98, Gri84b, Gri84a, Hal15, Win13a, Win13b, Bub79, Hol88e]. Catalan [Gli84]. Catalana [Gli84]. catalog [Bad11, War79, For80a]. Catalogue [Eur69, Bad12, Pye80]. catalyst [Hof79b]. Catastrophe [Wal05a]. Categoricity [How96a, Toa23]. cathedral [Cat04]. Cathode [Bri82, KS81]. Cathode-Ray [Bri82, KS81]. Cats [DRB05]. caught [Ano10c, B^{+}01]. Causal [Coh80, MC80, SSM21, Byr80b, Mea21, Sta06, Vig94, Vig95]. Causality [BM93, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, Coh80, For71b, For80b, Pau94d, Ray95, KKC06]. causation [Bha92]. cause [Nad02, S^{+}10]. Causes [Ein26c, Pep04c, Goo02, Mil01]. Cavazzen [Bad12]. Cavendish [War92, War93, Wil88]. cavities [GLG09]. Cavity [BDR^{+}07, CSD^{+}07, PSM^{+}13]. CCD [Boy10, Smi10]. Ceccarelli [Seg91]. Cech [Hom04]. Celebrate [Ano84c, Fer82b, Gol84b, Woo80b]. Celebrating [CGT11, Coc16, Lam79, Ken05c]. Celebration [Ano29g, Ano79g, Ano81d, Ano84c, Jab11, KRW05, Opp39, Opp79c, BS81, Kil87, Kil89, Man80, Muk05, Wad07, EHM^{+}79]. celebrations [Rod05]. Celebrity [Hal18, Sta05d]. Celestial [Pao84]. cell [KWGH97, Sch44, Sch45, Sch62a, Sch67, Sch92a]. cellulaires [Chr22]. cellular [Chr22]. cenergiëiìa [Smy46]. censored [Wel06]. Censoring [Dee08]. Census [Bar05f, Nye88, CHU87, Cra02a, Tre88]. centenaire [Fre79d]. Centenari [Gli84]. centenario [BF02]. Centenary [BCEP94, For80a, Gli84, Han80b, MG02, Pye80, Sci80, Wei12h, BF02, Blu05, Eri82, Fre79b, Fre79c, Fre79d, FK85, HI79, Hen82b, Jan06, Ken05c, Kil87, Kil89, Kun76, Lew79, Nat79, Ram79a, Wil05b]. Centennial [Ano79c, Ano82c, Ano84c, Ano84b, ABIM15, BW04, For05, Gei15, Gol84b, Kur79, Lin01, Ram79b, Wil17, Adl03b, BS81, Büh15, CH05, Fei01, HE82, Kle81, Lar01, Man80, OONM17, SW94, Woo80b, vM00, Ano79f, Fer82b, Gil84a, Lar84a, Wil84]. Center [For80a, Hen82b, Kli62a, KWM57, Lew79, Ram79a, Sch95b, Ein06a, Ein05n, PA06, Rab70, Lew79]. centième [Ano80a, Sto96]. Central [Tre79b, Hoc83, Rol89]. centrali [Cal75]. centrality [Fie16]. Centre [Isr73, Meh73b, AAA^{+}18]. centrically0symmetrical [EG23]. Centrifuge [Kem12]. centuries [Der15, EI38a, Kao10, Har79d]. Century [Bar05f, Ben11, Ber00, Bro72, Bru06, CC96, Fer14c, Fru82, Hie82, LL82, McC68, Nau17, Neu06, Nor02a, Nov18, Ree20, Ren08, Rup09, Sch04a, Uvn82, Ano15a, Ano01d, Ano05c, Bod17, Bra09, BPP95, BS79, Cas05c, Cea97, Cea98, Cro16, D’A01c, Dav03, ECEP11, Fer14a, FJTS05, Gab95, Gin04, GS05, GS06, Har06b, Her94a, Her04, Hol96b, How14, Kra99, Ler03, LSW12, Lig05, Lon13, Lui11, Meh73b, Moo81c, MT98, MNOW00, Myr21b, Pan04, Ron07, San07, Sch96a, Smi05b, Söd10, Spe05, Ste07, Whi16, WSG00, Wor08, Yos95, vM05b, vM09, vM12, Bar98b, Bre97, Cat98, CM86, CM97, D’A00, Dar04a, Din66, EMH96, FK90, FW60, GHS08, GHS18, Gin01, Has04, Hol00a, Kai02, Kra88, Nye04, Put09a, Pye82a, Rog06, TS01, TB02, Zac97]. Century [Zac99a, Har85, LY95, Bro00]. Cerebral [Kan92b, FLN13]. ceremony [Spe79]. CERN [Acz10, Kra14b, PA21]. CERN’s [PA21]. Certain [Bey06, Cas05b, Del09b, Dir79c, Gre05b]. certainly [KRZ^{+}05]. certificates [HR02, RW02]. Certification [ZXL^{+}19]. certifications [Her76b, Her84, Her01b]. ces [W.25]. Cesarina [CM02a]. Ceská [Brd79]. ceskému [Ein23h]. cesta [Bič79b]. CFCs [Ano97e]. chaîne [Per39]. Chadwick [Ano51a, Bro97, Gow93]. Chain [FP40c, The46, And73, FN13, Hah48, Per39]. Chair [ABdB^{+}53, Bas92, Rob01, AH93, Kle75b]. Chairman [Ein48e]. chaleurs [Ein12c]. Challenge [Ano07t, Ber06b, Sau04a, Sau04b, Sta09a, CT12, Dre07, GZ97, GZ06, Smo19a, Stu14, dB36]. Challenges [DWL98]. chamber [Led14]. champ [Ein10e, Ein30c, Eöt00]. Champignons [Ing05]. Champions [Boy88, Sti13a, Lan06]. champs [EdD29c]. Chance [Gis14, Nic34, Ray95, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, Bun82, Rus81]. Chandra [Wal91a]. Chandrasekhar [BS83b, Sri99, Wal91a]. Chandrasekhara [SR04]. Change [Lil63, Aer99b, Ano79g, Bre08, CÖT15, EE05, Ell90, Kie94, YCL20]. Changed [Can16b, Eng17b, GP19, GM06, Her17, Kem12, Wac06, Wil16a, Ber11b, Can15b, Cla61, Con02, Gui95, Har06b, HEWW15, Liv13, Mah03, aM07, Sta98a, Sta05b, Tho12, Wil11a]. changer [Ree15c]. Changes [Wal88b]. Changing [Ano20a, Duc11, Fis92, LRS12, Müh11, OV86]. Channel [Wac06, EM23]. Chaos [Cap01, MGG^{+}95, McH07, Sto05]. Chaos/Plaschko/Brod [MGG^{+}95]. Chaotische [MGG^{+}95]. Chaplin [Led14, Sol87]. Chapter [Lau50, HM11]. chapters [Cha49a, Hoy94, Ano23a]. Character [FLH85, Fey94, EM34, FL88, FF95, FL06, Har72, Pol89]. characteristic [LRH22]. Characteristics [Ano21d, The46, Ein26b, RWL23]. characterization [CSS^{+}21]. Charakter [Pol89]. Chardin [Ano67]. Charge [Mar08, Mil13, JDB^{+}97]. charged [TN03]. charges [Ein08a]. charging [SU06]. charisma [Bon79]. Charles [Ano50a, Bri82, Hav80, Rav97, Hol03a, Hor00, Nag50]. Charlie [Sol87]. Charter [RFP^{+}45]. Chases [Wei12m]. chastnaìa [HC98]. Chauffør [Pat01a]. Che [Lan67c, Fer24]. Checking [Hen20, Ano10e]. Checklist [Ano60]. Chemical [Boe10a, Hab82, Joh06, Par99, Vis94, Cha49a, Cha05c, DSW99, HS44, Pop99, Sla39, Sla70]. Chemico [Abi91]. Chemico-Thermal [Abi91]. Chemie [BS92, Bre05, MGG^{+}95, NS88, Ros86]. Chemie/Dictionary [MGG^{+}95]. Chemiker [BR00, Joh95, Sto98, Vis95]. chemins [Eis02]. chemische [HS44]. Chemist [BR00, Smi05a, Sto98, Str05a, Ano97e]. Chemistry [Ano82b, Ano99d, BZ10, BHJ^{+}00, Bre05, Coh54, CF82, Cra84a, Don70, Duv64, FF93, Fru82, Gre03b, Hie82, Ihd74, Jab13, Kor94, Mal97, NM12, Nob64a, Nor15, Sch15a, Sch90, Sin07, Tra00, Tre88, Vis94, Aas88, Ano72, Ano97a, Bag92, BSPKN03, CHU87, Cra02a, Far53, Far63, FM92, GW10, Hab02, HF22, Jam93, Käl22, Ler03, Nac79, NS88, Nob64b, Nob66a, Nob66b, Rog06, Tho17, Tho88a, TA17, Ano45b, Ano87b, AO78, Bar05f, CGT11, Jar08a, MGG^{+}95, Opp79a, Gre19a, Gre19b, Nye88]. Chemistry/Wiser [BHJ^{+}00]. Chemists [Sch21, Bad11, Du 52, FB24a, HH00, HH03]. Chen [LY95]. Chernin [Jos95a]. Chicago [Mar10, PJMS11, Smi58, vD09]. Chief [Abb05, Kuh54, Ren05c, RS11b, Sha54, Col68, Dra53, Dra01, Ren05b]. childhood [Tou14]. Children [Cal02, Ema85, Mas95, Dur02, ECS02, Sow01, Sta89c]. Chile [Ano45b]. Chili [Ren21]. Chimie [Bad11, Bro83, Tat64]. Chimney [Phe21]. China [Cao06, Bin06, Cao06, Cao14, Fan06, Hal65, Hu04, Hu05, LX88, Li07, Sch54]. Chinese [Cao04]. chip [BRH^{+}09, CSD^{+}07, HHHR01]. chips [HES^{+}12]. Choice [Sch93c]. Choices [Her00, vB02, BBM78, Bun88, Bun90, Wol93]. Choosing [Lan15, Sca14, Sch61]. choses [Nor21]. Christ [Mit87, PB72]. Christa [Pye80, Ser88]. Christian [Ein34m, Mit87]. Christianity [Bau09]. Christoph [Hus15, Pye06, RSPS49, The12, vD15]. Christopher [Dow80]. Christsein [Bau09]. Chronic [Kan94b, Sri94]. chronicle [Sch10a]. Chronicled [Ano05b]. Chronicles [Man68b, Ano05-33, Sch04b]. Chronik [Man68b]. Chronique [Sch10a]. chronogeometrical [Fok65]. chronogeometry [Fok55]. Chronological [Sch49e]. Chronology [KSB^{+}09]. Chronos [Pes13]. church [MM83, Zan93]. Churchill [Far13a, Far13b]. Ci [Arg09]. CiE [BBL08]. ciencia [Ben01, Fer05a]. Ciencias [Ano25c]. Ciències [Gli84]. científico [B^{+}99]. Cinematic [Waz10]. Cinquantenaire [MK56]. circle [Ein47c]. Circles [Eis93]. Circulation [Eng05b, Leg15]. Citation [LR14, BOS14, FW17, Har13c, LHZ^{+}23, YCL20, FN13]. Citations [EKS^{+}08, LLSH23, San15, ZXLX14]. cited [RN16]. Citizen [Pow94a, Fey98, Wol93]. citizens [Ski04]. citizenship [Ein41f]. City [JJ81, Con05, Olw97, Olw04, Ano21d, Kie13, Ros02a]. Civilization [Ein33a, Boh46a, Hab02]. civilized [Kle65a]. Civita [Lam30, CT24, Goo75, Wei12o]. Claim [Kri19a]. Clarendon [Red84b]. Clarification [Pir07, Kox06]. Clarity [How92, Kae38]. Clark [Bal72]. Clarke [Sch90]. claro [Paz10]. Class [Esh09, Lau55b, BB13, Cam09b, GdS15]. Classic [FL06, Sta99c, GR17, HT10, Hug99, Ren97b]. Classical [Boh62, Eza79b, Hei49b, McC11, Ren07a, RSS^{+}07a, RSS^{+}07b, RS07b, RS07a, Roh60, Row05, Sch05b, Sch07b, WF49, Bäc19, Bäc21, Bes04, BRS03, FP16, GG05a, Har13a, Inf31, Lan07a, Mar97, Vog21]. Classicists [Ker89a]. Classics [Ker89a]. classiques [W.25]. classmates [RK86]. Clausius [Gar70]. Cleaning [Pow15]. clear [Paz10]. clearance [Pol02]. clears [Bre08]. Cleavage [Ano41]. Clerk [Bes04, Gol83b, Har02c, Har98, Hen86b, Kni16, Mah03, Tho31, Tho11a]. Clifford [Bri18, RdO07, RdO16]. Climate [Bre08]. Climbing [MM00]. Clock [Coc66a, Coc66b, Coc66c, Dor68, NS96]. Clocks [Gal00, Gal03, Mar04c, Smo07a, Ste05d, Wei12q, Yer03, Yer04, Gal04a, Gio14, Sta05a]. clos [Bla93]. Close [Fie15, SR01, Wei12e]. Closed [Goe49, WM08, Wid07, Bla93, Elz06]. closely [vS21a, vS21b]. Closing [Poo90]. cloth [Alf09, Bed01a, Bru03c, Bru06, For01, Noe11, Par06, Pri99, vB01]. Cloud [Asa95, PNT87, Wei88, Bes93, Gor09, LML88, Mil91, Win93, Win99]. Cloudcroft [Cle96]. Clouds [Cam18, Mil13b, Ne’06]. Club [Bad96, Ber96c, Bra14b]. Clue [Lau55b, PW17]. Clusters [Ein21e, Zwi37]. cm [CM67, Lid67, Tur67]. CNAM [Bad11]. Co [Ein20h, Ein49d, Ein70b, Woo28, Wyd15, BH01, Fie15, RRK^{+}64]. co-authorship [Fie15]. co-discoverer [BH01]. Co-Operative [Ein70b, Ein49d]. co-workers [RRK^{+}64]. coauthor [LHZ^{+}23]. coauthorship [Fie16]. Code [Del11b, Hub13, Kea12, Pye10, Mel09, Wes15]. codiscoverer [CSW97b]. Coefficient [Lid67, SMHL23, Tur66, Tur67]. Coefficients [CM67, All90, CT24, Eva95]. Coetzee [Hal12]. Cogniot [Ano79b]. Cognition [Kor79]. Cognitive [Cer00]. Cohen [GSW95]. Coherence [Nee79, DCTW02]. Coherent [AKBAE21, BRH^{+}09, Tow65]. Cohn [CD07]. Coincidence [Cam18, Gio13a]. Coincidences [How99a]. Coincidental [Kuz79]. Cold [Cli09a, Ree17c, Sal19, Sha99, Sen21, Tau12, Ada17, Bru03b, CR08, Del09a, Dre07, Ger92, Ger97, Ger02, Ger07, Hec16, Her80, Her82a, Her88, Her89b, Hof10, Hol96a, Joh07, LP05, Leo09, RB04, Uda94, Uda98, Wel99, Wel00]. Colette [Hen82b]. Colin [Tha77]. Coll [Mat86a, Bro83]. Collaboration [Krs04, Pye76, CÖT15, Ill89b, KKKM05, TA17]. collaborative [Ein15c, LOvT12]. Collaborator [Kae38]. Collana [Tug98]. Collapse [GMHB02, NN75b, FTZ^{+}15, GSPH00, See25]. collapsing [DCC^{+}01]. Colleague [Pin00a]. Colleagues [Ano35, EM49, Hof08c, Kle10, Kox93, RKW05, Wei12e]. Collected [Bad15, Bal91, Ber07a, BDR93, BRI^{+}21a, BRI^{+}21b, Cah90b, Cah90a, EKS^{+}08, Fey05, Gol87, Har91, Har96a, Hen88a, Hen88b, Ker89b, Key89a, Kil95, Kle59, Lac08, Loe88a, Loe88b, Mar11b, Mar11c, Mcc87, Pye89, Pye95, Pye97, Pye06, Ram90, Row00, Sch82a, Sch03b, Smi89b, Smi89a, Smo07a, Ste90, Str90, Tau63, Wal99, War96b, Yan87, Ano99b, Ano99c, B^{+}06, Bel87, Bel04, Dic79a, DD95, FS87, Har96b, Hol05b, Kev05, Kil98, Sie97, Sie07, Sta00, War95, War96a, Wig96m, vH62, Bai89, Bal96]. Collection [Ano90b, Hum51, LEMW52, BSW22, Eur69, How50, LEM13, LEM^{+}23, LME23, Mic68, Ren05d, Sea72, Sti13c, Wad07, Bin06]. collections [Lew79, For80a, Ram79a]. Collective [Gin08, HEB^{+}80, Cha17]. Collector [Kov01]. College [Ano21j, Ano22a, Ano21d, Ein21o, SM41]. Collider [Hal09, Acz10, Lin09]. Collocation [Sch17b]. Colloid [Ker89a]. Colloidal [Haw02a, Pus91, WCL^{+}00]. colloque [Sto96]. Colloquium [OKSD05, RDB^{+}09, Ano67, UM86, UM92]. colombe [Mar91]. Colonel [Sar35]. colonial [Hom05, Sin07]. color [Bro10a, Bro11, Bro16]. Colorado [Bla13]. colossal [Liv13]. Colossally [Bro21]. Columbia [Ano39a, Ano39g, Lau51]. com [Ren21]. come [Bru76, DS01, Dor86, Wei18a]. Comedy [Oha19, Dür62, Dür80, Dür06]. Comes [Smo07b, DGG^{+}21, Ehl05a, Ein59f, Smo06]. Comet [DM96]. comets [Van98]. comics [Fol17]. Coming [Fer03, Hol81b, Wal82]. comme [Kle74, Smi15a]. Commemorate [Gol99]. commémoratif [Sto96]. Commemorating [EH05a]. commemoration [Ano01d, Buc80, Hen82b, JDB^{+}97, MFHR06, Tho31, Tho11a]. commemoration/WE [JDB^{+}97]. commemoration/WE-Heraeus-program [JDB^{+}97]. Commemorative [Duf81, Hen82a, Kle82a, Ein21h, Hen82b, Sto96, B^{+}79]. Comment [Ano57a, Ano79f, Ein34a, Ein53b, Ein58, Ein89c, Ein09a, Ein10, Ein15a, Eis37, Her03, Jos80, Kre72, Sha75, Smi15a, Abr12a, Abr12b, Mer12, Oha12a, Oha12b, Ein05g, Ein05e, Ein05h, Ein05f, Ein05i, Ein05k, Sar35]. Commentaries [Col00b, BB05, Bro72, New60, Yan13b]. Commentary [Har79c, JRN^{+}07a, JRN^{+}07b, Kle70b, Kle70a, Spi96, Bet97, BH01, Ern92, Hug99, She16a]. Comments [Ein86d, vI11b, You98, Bjø20a, Bjø20c, Bur90, Löw98, TH79, Ein14e, Ein19b, EL05a]. Commission [Ano93c, Hen13, For82a, For84a, For86, Hac94, HA62, HD69, HH89, HDHA89, Bad83b]. commits [Ben99]. Committee [Bjö01, FR49a, Ein48e, Käl22, Ein49c, Sig78]. Committees [Sig78, Ano82b]. Common [Ein18n, Ein20o, Ein21c, Ein21r, Ein22a, Ein41b, Ein42, FG21, Hni98, Ng11, SG82, W.25, Bon64, Den24, Mos21]. commonly [Ano22a]. communicatie [dGK11]. Communicating [Leg15]. Communication [Bar80a, Can16a, Ein20f, Hil07a, Hil07b, Hil07c, Kni16, Sau05, Sau07b, BMS12, dGK11]. Communications [Szi71]. Communicative [Nug99]. Communism [Sha92]. communist [Mik65]. communities [CZPR17]. Community [Cha74, GSW95, Hab69, Hol81b, Hun07, Bey77, Mal00, Olw97, Rol47, SR04]. Como [See61]. Companion [Hus15, JL14b, JL14a, vD15, Haw92, You11]. companions [Mei05]. Company [Hea64, Lan15, Wyd15]. Comparative [Bar98a, Gli87, Ren93, Sch74, Tor82]. Compared [Lig04]. Comparing [Khr08, War93]. Comparison [Edd24, Jan95, O’L64, dWtH20, Ler73, RN16]. Compartmentalization [Gol95a]. Compass [Gal04b, LS05]. Compatibility [Ein30f]. compatible [EM31b]. Compendium [GHW09]. Competes [Bla88]. Competition [Bjö01, Pin00a, Spe79]. compilation [Kel06b, Bäc19, Bäc21]. Compiled [Ano60, For80a]. complement [Ano97b]. complémentarité [Pat85]. Complementarity [Amo10, Bel93, Jam74b, Jam74c, Rös95, Pat85, Meh87]. Complementary [MST^{+}07b, MST^{+}07a]. Complete [Ano35, Ros79a, EPR35, Hot15, Kie15, Mor00, Mor04, Pen05, Pla59, Ren05d, See25]. completely [Ein02b, Ein27b]. Completeness [Hoo91, Toa23, ST94b]. Completes [Ano31b]. completion [MR00, Pau25]. Complex [Cao04, Gol82, Ree14a, CSS^{+}21, Pau25, RA11, Van04]. complexity [BDH04, HBR99]. Component [Ree14a, NBDD^{+}12]. Components [Ein18k]. Composite [How90b, KK05]. Comprehend [Ano19b]. Comprehensible [Wet80]. comprehensive [Sch77a, Sch77b, Sch77c]. comprehensively [Ang20]. comprendre [Cun22b]. Compton [Duf46, DR89, Ein24d, Stu75, Woo25]. Computability [BBL08]. computation [Hey99]. computational [Day12]. computer [BS10, Ben16b, Pin00a, Pul11]. Computers [BHJ^{+}00, Faz15, Hey99]. Computers/Canty [BHJ^{+}00]. Computing [Agu90, Sla13]. con [Ein94b]. Conant [Ber93, Con17, Her82b, Her89b, Her93]. conceitos [EI39b]. Conceivable [Nor00]. conceived [Gib91a]. Concentration [Hug54, OW54].

Concept [AP65, Din65, Foc55, Hir69, How96a, Kes65a, Kes65b, Noe11, Pia30, Ren93, Tay21, dWtH20, vKL49, Alf01, Coo10, Coo15, Ehl05b, Lon37, Mas99, Ray71]. ConceptDraw [GRE^{+}01]. concepte [EI57]. Conception [Ano25a, Dir79c, How97b, Mar49, Nor49, Ein31g, Ein11b, Lon37, Meh73b, Rei34]. Conceptions [Ano20a, Ein50e, Ein52b, Kra07a]. conceptos [EI65a]. Concepts [Bac49, Bel23, Boh62, Bro39, Čap76, EI38b, EI38c, EI66b, Ell89b, Hon04, Kae38, Kan79, Kil84b, PFM17, Red84b, Roh87, Sla69a, BN97, Byr80b, EI38e, EI39b, EI42, EI47, EI48a, Ein52b, Ein53c, EI57, EI59, EI61d, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, EI98b, EI08b, EI11, EI13, EI15, GHW09, Leg01, Leg03, Lon84, Lon03, RDB^{+}09]. Conceptual [Ber68a, Gra71, Jam89, Jan73, EE59, EE90, Jam66, Kra25, AS91]. Concern [Ein51b]. concernant [Ein14j]. Concerning [Ein22g, Ein36g, Ein49d, Ein70b, Ein86c, Hoo70, Kuh54, Mie07c, Nag82, Sha54, Bau20, Col68, Dra53, Dra01, Eĭn85b, Ein21s, Ein40a, Ein40g, Ein05-37, Sal74, dB85, HS39c]. Concerns [Bro23]. concert [Man80]. concertante [Ein34m]. concetti [EI65b, EI13]. Concezione [Mas99]. conciencia [Fer05a]. Concise [Lor10a, Lor20, Lor10b, Lor10c]. Concluding [Wig96d]. conclusion [Log96]. Conclusions [D^{+}95, Ein07h, Ein07g]. Condensate [Col95, CW98, Joh13, Lev98, BDR^{+}07, DCTW02, EGW^{+}08, GSPH00, GMHB02, GRJ03, IAS^{+}98, JDB^{+}97, LLSS07, NBDD^{+}12, OKSD05, RDF^{+}08, SHL^{+}06, SLLS09, WH99, Wil11b, YCKG11, ZPSK10]. Condensates [Hea11, KF03, Lev01a, Lev01b, BSSP^{+}08, Big05, BRH^{+}09, BR99, Che00, CSD^{+}07, DCC^{+}01, EG98, Fet09, FS01, Fri09, GLG09, Hau08, Jak08, LJGS11, MO06, Osh06, SDCZ01, SISK^{+}98, Tom99, WNS^{+}08, WHL^{+}05]. Condensation [BEC99, Ket99, Lon38b, AK02, BLI^{+}02, CW02, DGPS99, DSZ12, DDD^{+}06, DHY10, EM04, FKZ91, GRT08, HHHR01, Hul01, KU38, KRK^{+}06, Ket02, KSVW10, Leg01, Leg03, PBM^{+}12, PSM^{+}13, Sto02, Tom22, TH79, Zol02]. Condensed [ARK^{+}99, Lau50]. condenser [TN03]. Conditions [Ein22e, Got00, Kel58, Mon04, Pei39, Per39]. conducted [Bjø19]. conducteurs [Ein10e]. conducting [Shi01]. conductivity [S^{+}57]. conductor [EE05]. Cones [PS04]. conféences [Ein55i, Ein64b]. Conference [Ano23a, AS91, BCEP94, Cur82, Ein53l, EK92, GRRS99b, HS89b, JP08, JP89, Sch83b, BB55, Cho16c, Dür82, Gly90, Mag95, Sto96, CHR96, Ein31b, BV09]. conferences [Ein76b, Meh75c, Ein76b]. conferencia [Ano25b, Dür82]. Confession [Ein20e]. Confidence [Rob75]. confinement [JDB^{+}97]. confini [BF16]. Confirm [Ano06h, Cho06, Lau45]. Confirmation [Ano23b, DT23]. Confirmed [Ken19, Nic20, Hec11]. Confirming [Kir80]. Confirms [Kae13a]. Conflict [CS79b, Ros79c, B^{+}01, BO76, BEM12, Feu78, Hof08c, Wal81]. conflit [Feu78]. Conformal [Sch49f, GLW15]. Confrontation [Wil01, Wil06, Wil14, Alp65, Alp85, Alp94, ’t 00, ’t 02]. Confronting [Joh07]. Confronts [Cre83, New09]. Confusion [Kox06, Ang05, Kra21]. congenital [KWM57]. Congo [Wil16b]. Congress [Bar93, Ein48g, Str11a, Dow80]. conjectures [San81b]. connected [Lei08]. connecting [JP55]. Connection [Bar49, Car07a, Gin08, Pin00a, Gan97]. connexion [Fox15]. Conoscere [BF02, BB04a]. Conquest [Sha99]. Conrad [MGG^{+}95]. Conscience [Ano04c, Arr07, EM88, Fer05a, Say85]. Consciousness [RK11, KST99]. Consensus [Can16a]. Consequences [Aga06, Aga07, Bac89a, Ein10c, Ein10d, dSR16a, dSR16b, dSR17, Bac89b, Bla48, Bla49a, Bla49b, CM89, Jun05b, Mun36, Ryc92, SS79, Sel78, Wol87, YGKK05, Ein10c, Ein10d, Ein01, Ein05u]. Consequent [LC07]. Conservation [Ein18e, Lew26, Ein05n, PKLT17, Pit16, Wei13c]. consider [TTB20]. Consideration [Ber94, CSW97a, Gas05]. Considerations [Cas21, Ein17e, Ein40a, Ein40g, Ein05t, Lor07a, Mos21, Ein16d, Ein53d, Ein86e, Eng05a, Khr08, Kra20b, Fer24]. Considerazioni [Fer24, Gas05]. Considered [Ros79a, EPR35, Kie15]. Consistent [KS65, Sla72, Nor72]. Consisting [Ano29b, Ein39d]. conspiracy [AK97, Bow87a, Bow87b]. conspirators [Du 52]. Constance [Bad12]. Constant [Ano06b, Ber07b, Cob40, Col00a, Ear01, GHK^{+}06, GHK^{+}07, GM00, Hol99, Lah20, Lem49b, Mil13, Mos13a, Mos13b, Sch20, VT99, Wil93a, All90, Cur92, DGM04, FM10, Gol95b, Gol98a, GAHS96, JN14, Kra14d, LHF^{+}98, OONM18, Ray90, Spe05, TLW^{+}10, Wei89, Wei13e, ZdSS91, Abb88]. Constants [BF86, Bar02a, Bar02b, Gré19c]. Constituent [RS07c]. constitution [Ein09b, Ein09c, Ein19e]. Constrained [Gou10]. Constraints [Hon04, CPDH^{+}87]. Constructing [Hen86a, Hol79b, Spo02, Viz79, Ner84]. Construction [D’A01a, FP40c, Gin08, MTVB88, SWU09, GP11, Hen97, Hün18, Smi02]. Constructions [Set03]. constructive [Gio20, Hag08, TTHJ68]. constructivism [Sie10]. construite [W.25]. Consumer [Jun14]. Contact [Mil16, BM09]. contacts [SR04]. contain [Fer24]. containing [Chr22]. contemporâneo [Inf57a, Inf60, Inf61a]. contemporaine [dB49c]. contemporanea [Seg76, Seg83]. contemporanee [Cas14]. contemporaneo [Inf63, Inf68, Inf74, Inf92]. Contemporaries [Ein89a, Sig92, Hen03, Hen90, Sta95b]. Contemporary [Bro20a, CW74, FF94, Gra71, Jan73, Mal09, Ush49, Zho85, Bel96a, Cas14, FCKT11, Hug99, Inf57a, Inf60, Inf61a, Inf63, Inf68, Inf74, Inf92, Lan59, Mil94, Sch17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, Seg76, Sha98, dB49c]. contenant [Chr22]. contengono [Fer24]. content [BPS87, Ein05b, Ein05v, Flü39]. contents [Müh11]. Contest [Dac20b, Edg14, Ano21g, Ano21m, Law21, Phe21]. Contested [Ber93, Kut09]. Context [CF82, Fri89, Ren05e, Ren07d, Sta82b, BCR93, Bel96b, EBRC93, Kai94, PL17, Ree15b, SR04, Sin07, Swe79, EBRC93, BCR93, GR17]. contexts [Kra08, RCD98]. Continued [OS39]. Continues [Wad81a]. continuierlichen [Ein23a]. Continuing [Ber06b, Sac88, Tym95]. continuità [Ber81]. Continuity [Sch05b, Sch07b, BN88, Ber81, Duc11]. Continuous [Car07a, Hel09a, Ein23a, Hag14]. continus [Pia31]. Contours [Inf69]. Contra [Sta98b]. Contract [Car04a]. Contraction [Dor68, OS39, Pop66, Rac49, War91]. Contradictory [Lig97]. contrasting [Gra82]. Contrasts [Ano15l, Man86]. contribuicäoes [GdS15]. Contribute [Ano90a]. Contribution [Ein20j, Eza79a, Lee06, LL82, Pau94b, Per23b, WE87, Dim93, Elb13, FK90, Isa85, Kie94, Lan21, Pau55]. Contributions [Eza79b, KRW05, Pau49, Sta94, Tre79a, BNSY98, GdS15, MFHR06, Mei05, Rog06, Sin04a, SW12, Ein14d, Ein90, vEPF09, vEPF22, Ein22d, Ein05h]. Control [Wel08, Ano51b, Dre93, Dre07, FW17, HGS87, Lan15, Ano45a]. Controlling [Har13a, GGW94]. Controverse [Com95]. Controversial [Cho05a, Bri05a, Eck01, Nor12, dB85]. controversies [BNSY98, Sac88]. Controversy [Ano06b, DR89, Hen16, Kif95, Kra96a, McC77, Nav15, Waz09, Waz10, BL96, BHM91, Fel00, Hac94, Hol91b, Ker89c, LS08, Ren97a, Waz14, Wic95, Com95]. Conundrum [Mos21, Sch13b]. convenient [SH01]. Conventionalism [How84]. conventionality [Jan08]. Convergence [Ort95]. Conversation [Bru09, GB03, Whe79a, But23, Wol67]. Conversations [Coh79, Mos70, Mos72, Ray95, Sha63, Sha73a, Sha75, BC01, Bri95, HH00, HH02, HH03, HH04a, HH05, HH06, Hei71b, Hei71a, Hei75a, Hei83b, Hor09, Pal98, Tip10]. Conversion [BHJ^{+}00, Nus14]. conviction [Stu10]. convictions [Alt06b]. cooling [JDB^{+}97, Phi98]. Cooperate [Hec16, Ree17c]. Cooperation [Can05, Sal19]. Coopersmith [Noe11]. Coordinates [Nor89]. Copenhagen [Dör05, Fra00, HB63, McH07, Pat95, Seg07b]. Copernican [Dra53, Dra01, Foc53b, Har72, RR14, Tam14, Kuh54, Sha54]. copernicana^{≫} [Tam14]. Copernicus [Gam42b, Ein53k, Gam42a, Hen91, HB01, Inf55d, Inf73, Jon44, Rei42, Rei57, Rei80]. copier [Sch12d]. Copley [Ano26b]. Corbusier [Loa98]. cordon [Ano33b]. Core [Cas79, RA01, Ree15a]. Corn [Nel77]. Cornelius [Bal75a, Bla76a, Bla76b, Bro66, BCEP94, Hou65, Jam70, Mal65, See65b, See65a, Wre66, Bri66]. Cornell [Lev01a]. corona [Ein23a]. corporation [Bla02]. corps [Ein56i, Ein94e]. corpses [Mar17]. corpuscles [dB49d, dB55a]. Corpuscules [dB49a, dB49d, dB55a]. Correct [New09]. Correct. [Ano35]. Corrected [Kel58]. Correction [Bjø20a, EL08a, Ill20, Lid67, OM22, Ein07b, Ein15c, EL05b, Ein05l, Ein05m]. Corrections [Coc66a, Coc66b, Coc66c, ES46a, Ein11e]. Correlation [KS65, Sla72]. Correlations [Cos87a, Ryf14, dB87, PBM^{+}12, dB65, dB85]. Correspondance [EB72, Bal92, KF74, Spe72, Kle73, Mar73]. Correspondances [EB^{+}89]. Correspondence [ABC^{+}06, Bad15, Ber07a, Ber94, BB69, Bro72, BSKS04a, BSKS04b, BSR^{+}06a, BSR^{+}06b, BRS^{+}09a, BRS^{+}09b, BIRS12a, BIRS12b, BIR^{+}15a, BIR^{+}15b, BIJ^{+}18, BRI^{+}21a, D’A01a, EF33, Ein94b, Eis00, Ern92, Har96b, Her66a, HvMW79, Hol05b, How84, KSB^{+}09, Lac08, LS93, Pye95, Row00, SKJI98a, SKJI98c, SKJI98b, Sta99a, SE90, SE45, Ton97, Van22, War96a, War96b, Wei12o, vM93, vM96, vM99b, vM01, vM05c, Bal96, BB71a, BB71b, BB05, Dur02, Hag08, Han79b, Hen86c, HRP90, Her63b, Kev05, KKS93, Kot93, OB05, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Pei71, Prz11, San11, TH79, WS78, vMHW85, EB82]. Correspondencia [Ein94b]. Corridors [Nye99]. Corrigendum [Ano12a]. corso [Pal99]. Cortex [Kan92b, AH96, FLN13]. cortical [AH96]. cosa [Lan67c]. cose [Cap01]. Cosmic [BOPS99, Bri66, Bro66, DR89, Ein31d, Fra08, Lau55a, Löw78, MGG^{+}95, Mos21, ST98, See65a, Smo07c, Sus06, TVL^{+}09a, Val49, Wre66, Bre08, Don97, Gut97, Hou65, Jae06, KT95, Lan65, Lan67a, Lon13, Mal65, MO13, Mer10, OM14, OM22, Per12, See65b, Spo16, Ste05c, Van85, Alf09, Nun22, Bus12a]. cosmical [Kra13]. Cosmo [Kra82a]. Cosmo-Physics [Kra82a]. cosmogonie [Chr22]. Cosmogony [Gam49, Chr22]. cosmologia [Tam14]. Cosmological [ARK^{+}99, BTM84, BBB^{+}09, Ber07b, BF86, Car02, DR16, Ein31o, Ein45e, Ein86e, Göd49a, Kas21b, Lah20, Lem49a, Lem49b, Sch20, TVL^{+}09b, BT86, BT88, Cur92, Ein19e, Ein22d, Ein51g, Ein05h, Gol95b, Gol98a, Har13b, Kra14d, OONM18, Ray90, Rin69, Rin77, Rin01, Rin06, Web99, Wei89, Wei13e, ZdSS91, Abb88, Ein17e]. cosmologie [Kle74]. cosmologique [Ein51g]. cosmologist [Hoy94]. cosmologists [Fur87, LB90]. Cosmology [BF86, BBBM90, BL14, Boe10a, Boe10b, Cha01, CB15, Din66, FH10, Gal07, GK05, Hal91, Har81, Har00, HBN00, Kra82a, Kra96a, Kra05, Kra07b, Kra18, Kra15b, Lev07a, McC68, Pee93, PK06, Ram81, RBB03, Rei72, RL16, Rin11b, Rin11a, RN68, Rüg88, Sch05b, Sch07b, Sci79, Sol16, Wei08, Arp98, BDH04, CF11, CC84, Che10, Cho08, DB09, DE04, Ein09c, Ell89a, Far05c, Fer07b, Fie88, GH07, Har72, Hea15, HH96, Isr73, Jon17, Kea18, Ker89c, Kle74, Kra07a, Kra08, Ber84a, Lon13, Nor65, Nor90, Nor94, OONM15, Ost15, Pau93, PR11, RP09, Ric01, Rin09, Ryd17, Seg11, Sin61, Sin70a, Sin70b, Tam14, Tol34, Tol87, VT22, Vog21, Wei72, ZdSS91, Sme14, Myr21c, Myr21b]. Cosmos [Gre04b, Gre05a, Gre11, Jam30, Kar05, Kir02, Lev07a, Ste16, vH97, BTM84, Bar06, Dan00, EN97, Gib91b, Gra94b, Kak04, Kak05, Kra07a, Lem09, OMNM14, OONM15, PA06, Sta92a, Wei04, Fau79, Mus15a]. Cosserats [Sch19]. cost [Har13b]. Costa [Rol89]. costly [Car07b]. Could [Ano19b, Ano39h, CW12, Opp79a, Lip00, Lam21]. couldn’t [Jag01]. Couliano [Col93, Zal91]. council [BL11]. councils [MW99]. count [Sla13]. Countdown [SK82a, SK82b, Wal05a, Kee11, Kur86, Wal05b, Wal05c, Wal06c, Wal06d]. counter [Irv67b, TH79]. counter-measures [Irv67b]. counter-revolution [TH79]. Counting [Gar08, Kör96]. countries [GRS16]. country [Gha82, Kle16, LHDH21]. Counts [Gol99]. couple [Bad09]. coupled [LJGS11]. couples [LOvT12, PSAA96]. Coupling [Han22, CSD^{+}07]. Courage [Ein46e]. Courant [Som05a, Som05b, Ein10e, Som05a]. Cours [Pia31]. Course [Ano23a, Ein10b, GO08, Sch85, Sch09b, Con02, Dob06, Ein08b, HEWW15, LT15, Lin17, Pal99, Rei41]. courses [Ein26c]. court [EE82]. Covariance [Cer00, Foc55, Ryc99, Sta80b, EG78b, Nor89, Nor93, EG14]. Covariance/Invariance [Cer00]. Covariant [Nor04, Nor07d, Ein18d, Nor92]. Cover [Ano94a, Ben79, Woo06]. Cover-Up [Ano94a]. coverage [Cor15a]. covering [Dea13]. Coveted [Eds82, Wad81b]. Cowper [Woo28]. Cowperthwaite [Jan73]. CP [Kob09, Mas09]. CPT [Cos87a]. Cracking [Mel09, Del11b, Hub13, Pye10]. crackpots [vD10b]. crash [LT15, Lin17]. Crashes [DM96]. Crashkurs [LT15, Lin17]. Craters [Szi71]. Crawford [Bar05f, Goo85, Hoc93, Nye88, Rec86, Rei86, RK94, Sei84, Tre88, Wei94, Mat86b, Wal94b]. Crazy [Arr06]. Create [Kur12, Sau16, Hal15]. Created [RE90, Abi00a, Acz09, Ein30a, Ein82b, Ein85d, Rol47]. Creates [Ano41]. créateur [HD75]. Creating [Das12, Gar93, Roc03, ZDSF11]. Creation [DM96, Hal91, Hel09a, Rho88a, RT90, Wal16, Acz10, BS83a, BS93, Buc94, CM86, CM97, Ein05-37, Gle10, Hen84, Kra07b, Kra08, Sme14, Tőz15]. Creative [Gin15, Kre58, PSAA96, Rot87, Gol83a, Loa98, Rob10]. Creativity [ADS83, For05, Lin01, Rad08, Bod94, Gar93, JW11, Lar01]. Creator [Fri73, Hei74a, HD75, HD72, HD73, vL55c]. creators [Kel07, Kel09, Str11b]. Credited [Ano39b]. Credo [Ein41c]. creencia [Ben01]. crews [Pos11]. criador [Gar42]. Crichton [BHJ^{+}00]. crime [Ben99, Bor78a]. crimes [Smi92, Wes04]. Crisis [GB88, Smo14, Wal04b, Wil13a, Cas76, Dav74, MM66, Yos95]. cristallisation [HLZJ91]. Cristo [PB72]. Criterion [D’A01a, Erl72a, GHS21]. Critias [Joh04]. Critic [Ber89, Sem08]. critica [Tio09]. Critical [Ber76b, Ein05-29, HHMW93, Log96, Pei39, Sta09a, Tha77, Tre79b, Wen49, Ein10a, How76, Liu92, Ree15a, Tio09, Ein18i, Sch21, Urb22]. Criticism [Cun22a, Ein53b, Foc53a, Hah81, Hu04, Mor16]. Criticisms [ECEP11]. critics [Str68a]. Critique [Ano05o, Est06c, Fin81, Fra37, Grü73a, Kuh80, Leh14b]. Cronica [Pye86]. Cros [Sar35]. cross [Fie15, Gio13b, Nam09, Nam10]. cross-disciplinary [Fie15]. cross-purposes [Gio13b]. crossers [Sch05c]. crosses [Boz14, Len17]. crosshairs [Gru05d]. Crossing [Wat66, Wat70, Wat88]. Crossroads [Row01, BV09, Din72, ET07, Par09]. Crowfoot [KKJ99]. croyant [Sto96]. Cruces [Cle96]. Crucial [Hol69a, Hol69b, CL19, Hol88a]. crusade [HGS87]. Cryptography [MNT^{+}00]. crystallinity [GLG09]. cuantelor [EI57]. cuánticas [W^{+}87, WdC02]. cuantos [EI65a]. cuatro [B^{+}99]. Cubism [LA88, VM99a]. Cuestiones [W^{+}87, WdC02]. culinary [Söd10]. Cult [Bad83b, For82a, For84a]. Cultivation [Mer79]. cults [LD08]. Cultural [Ano83, Ano84b, Bad83a, B^{+}79, Ein52d, Eng05b, Fer83, Gil84a, Gol84b, Hol98a, Kil84a, Lar84a, Nag50, Sch84b, Wil84, HE82, Kil84b, Man55, Owe12]. Culture [Boy85, Boy94, Ebu16, Elz95, For71b, For80b, GHS08, GHS18, Joh05a, Kan84, Koh95, The12, Ash95, Cas14, FCKT11, GB06, GB08, Lau12, LSW12, Rab70, SE79, Thr17, Van02, Leg15]. Cultures [For05, Lin01, Abr09, Cas14, Gra82, Hal04a, Lar01]. Cumulative [EKS^{+}08, KSB^{+}09, Kli62b]. Cunningham [War92, War93]. cura [Per10, Tug98]. Curia [Cur60]. Curie [QP01, Swa02, Ano11a, Cal17, CGT11, CCD^{+}52, Cur37, Cur38b, Cur38c, Cur38a, Cur38d, Cur38e, Cur38f, Cur86, Cur01, Ein59c, Gir86, Gol05, Hor00, Mar95a, Pre05, Qui95, QM96, Qui96, Qui98, Red10, Tho88b]. Curies [Bri05a, Kae48]. Curious [Ash11, Eis06, FLH85, FL88, FF95, FL06]. Current [Kra23a, Kra23b, Ein09d, Ein10e, Ein29j, Ein30n, EK53, Pla20a, Pla20b, Pla22a, Rit09b, RE09, San11, Sie91, Wit62, Ein12c]. Currents [EdH16a, EdH16b, Ein15c, Ein15f, EdH15a, EdH15b, EdH15c]. curriculum [Pla20a]. Curry [Ray92]. Curtis [Kee11]. Curvature [Cal76, Cob40, Ein22c, Ein23f, LC07, Pep04b, Ein27a, Fra11, FH02, Sch19, See25, Sil30]. curve [McP74]. Curved [BLR15, EW88, EW00, Wal94a]. curvy [Ano04a]. cycles [HCD^{+}78]. Cygnus [Shi75]. Cylindrical [WW57, Grø82]. Czech [Bič79b, Brd79, ČK79, Ein23h, EI58, Gut75, Hav77, Hav80, Hom04, Inf58, Inf59, Kol72, Kol74, Kot93, Nac24, Nes60, Pát79a, Těš12, T– D07, Žáč11, Žáč12, Záv25]. Czechoslovakia [JP89]. czycie [HC95].

D [Aga07, B.50, Bad72, Bir82, Bus12b, Cli09a, Del13, Ein09a, Ein91b, Ein05g, Ham97, Hen82b, Jos95a, KP55, RS06, Rei50, Sie97, Vog21, Wis21, Woo28, Pla20a, Ano13c, GB15]. D-quantized [GB15]. D. [Ein09a, Ein05g, LSF94, Wig96j]. d.c. [LLSS07]. dêbat [Kum11]. Dachboden [Som05a, Som05b]. dadi [Pot11]. dag [Hau47, Hau50]. DAGA [Ano79g]. dagegen [CSW97a]. dagli [Bat03]. Dahlem [Aue95, Kan96a, Aue95, Kan96a]. dai [EI65b, EI13, Seg83]. Daitch [Hes04]. dal [Bat03]. d’Albert [Buc80, Gli84, Col57, Elb13, MRS17, dB49c, dB55a]. Dale [O’S01, Tan11]. Dall’ordine [Cap01]. damages [Abb98a]. Dan [Ano60]. d’analyse [Pia31, Pia31]. Dance [Ano05p, Zei10, Ano05-44]. Dancing [BR05, Boh05]. Dane [Jak13]. Danger [Bun88, Bun90, Ein21d, Ein34d, BCD93, DG03, Sch07a]. Dangerous [FC07, Ber07c, Ber09, Hof10]. Dangers [Hec16, Ree17c]. Danian [Cao06]. Daniel [Bad83b, For05, GRC18, Pye06, Nic20]. Daniela [Jun05a, Row06a, Row06b]. Danilo [Bes93]. Danish [EI39a, EI61a, Hau49, Hau63, Hau64, Jak13, MW46b, NN01, Pat01a, RRK^{+}64, Str97a]. dans [Chr22, Dem15, Edd21, Ein10e, Ein10c, Ein10d, LRL15, LRL17, Mar51, Nor21, Pat85, W.25, dB55a]. danske [NN01]. Danskeren [Jak13]. Dardo [Bar05f, Bru06]. darf [How90b, Mau96]. dargestellt [Bor20, Bor22a, Her01b, Her66b, Thi46]. Daring [Ano19b, GP19]. Dario [PLM^{+}0 ]. Dark [Cli09b, Dys16, Eis93, Kir02, Lah20, Law07, Mat10, Moo17, Rho95, RL16, See56b, Smi96, Van98, BSSP^{+}08, Bla10, Dvo12, Gat09, Ne’06, OONM18, Sar08, Buc08]. darkness [Mag09]. Darrigol [Har02b, War03a, Abi04]. Darstellung [Ano22a, Flü74, For80a, EM34]. Darwin [Ano81e, Hor00, Kni12a, Kni12b, Nag50, OB05, BNY81, Bau09, CF80, CH80, GH01, GO81, Hai10, Koc01, Liv13, Mar11d, Moo81a, Moo81c, Moo81b, Rus81, SNA^{+}06]. Darwinism [GPSR01, RA01]. Data [Ano13b, BBB^{+}09, Fra81, LHZA20, Per21, Ao04, BOS14, FW17, Gun11, Har13c].

DataVault [LHZA20]. date [Nor72, Ano79g]. Dated [Jan07]. Daughter [Bed01a, Ivr99, Pal00, Pye90, Gef14, Zac99b]. David [Bad96, BS92, Hea13, Jun05a, Lan08, OBxx, Plo06, Pye89, Pye06, Smo07a, Som05a, Som05b, Ste22, Sto17, Wue05, You95, Cor04a, Fer07a, For09, Koj08, Pat93, Rob01, Row06d, Som05a, Tsi08, Wue05]. Davidson [Fes56]. Davis [Fri73]. Dawn [Boy85, Boy94, Lau46, Lau72, Del09a, Gia03, Gor09, Ham08, Pet72, Rab63, Roc03]. dawning [O’R97, Shi01]. dawns [Hau47, Hau50, Sin01a]. Day [Bar09, Kur86, Lau51, Ōe95a, Pac72, Pac81, Rho88a, RT90, Wyd84, Ano10a, Els95, Far05c, Ful84b, Ful84a, Hau47, Hau50, HHW99, Sza84, Tho85]. Days [Col09, GM06, Wac06, FH11, Gor07b, Leg04, Oli72]. Dayton [Tho17]. DC [Hes04, Kar05]. Dead [Ano68a, Bel98, Fri50, Ano05-36, Ano05-37, Hof10]. Deadly [Mul99, Bai07a]. Dean [TH79]. Dear [ECS02, Cal02, Hoo77]. Dearest [IK03]. Death [Ein53k, Mil13b, Wei12h, Aĭn77, Ano67, Hal55, MNOW00, Nov92, Sti05, Inf55b, Sti05]. Deathbed [Swa02]. deaths [FG20]. Debate [BC24, Can16b, Div78, Eng17b, FG21, Fol96, Hel98, Her17, Jam74a, LLH^{+}80, Mos13a, Mos13b, Smi82, Sud11, Wil16a, Bok08, Bor04, Bun79b, Can15b, Car89, Dör05, FP16, Gio16, Kai94, Kum10, Kum11, Lan06, San11, Sta07d, Wel93]. debates [Cat98, D’A97]. debating [Row06c]. Debatte [Hel98]. Debever [Ano84a, CE79a, For80a, CE79a, Kil80]. Debrock [Ded85]. Debroglie [Pat93]. Debt [Ein38b, Zah78, Sti05, vdB95]. débutants [Sch81b]. Debye [Kan96a, Som05a, Som05b, Alt06b, Kan96a, Som05a]. Decade [Bal75a, Bru03b, Lan74a, Lan74b, Sch07a, Bla76a, Bla76b]. Decades [RMI21b, Mil91, Nor93, RMI21a, Wic95]. Decay [Ein52d, The46, BRVF84]. deceit [Tau86, Tau88]. December [Bad15, Ber07a, BCEP94, BSR^{+}06a, BSR^{+}06b, BRS^{+}09a, BRS^{+}09b, Fer39, Fer70, HY86, Lac08, Wig64b, Wig64a, Wig67a, Ano98, Kra88, Wig64c]. decide [Sch61]. decided [Jak13]. Deciphered [Rus17, Lev15]. Decision [Asa95, Ber93, CRS97, Dys76, Fre08, GB89, Ham97, Hol81a, Lef95, Lit09, Mal09, Mal12, MM58, Mor57, Oli99, Ove96, Ros79b, Sig78, Smi58, VB96, Wal90b, You95, ATW^{+}95, Amr59, Bak68, Bak76b, Bak76a, Bat62, But54, GF65, GF67, Jay90, Käl22, Mad95b, Mad04, Mal08, Sch61, SR01, Sti47, Sti76]. Decisions [KUW82, Ber83, Ray92]. deckchair [LRL15, LRL17]. Declaration [Ein36c]. declare [Mon58]. Declares [Bro21]. Decline [Kit00]. decoding [Bal01a]. Decomposition [Rac49]. Deconstructing [Hug04]. Deconstructionists [Lan00]. Decorative [Pye13]. découverte [Pop07]. découvertes [Hum51, Seg84]. decrease [LLSH23, MDT^{+}03]. Decry [Ano45a]. dédaigné [Sch05f]. Dedicated [Ano93c, CS79b, Kur79, Ros79c, FPLR59, Lin60, RR82]. deduced [LCL^{+}90]. Deduction [Dor71, Ein19f, Ein19g, Ein28d, Zah87, Ein13a, Mey85, Ein28d, Mey92b, Ein13a]. Deduktion [Ein19g]. déeduction [Mey25]. Deep [Gre11, Sch04c, Ree00]. Deeper [Bla92, Krü00]. Deepest [HF05, Bar02b]. Deerfield [Ros86]. defeated [De 12]. Defeats [Kri99]. Defect [HEB^{+}80]. Defects [MGG^{+}95]. Defects/Weiss [MGG^{+}95]. defence [Mit87].

Defenders [Wil13a]. Defending [GW06, Rei06, Rei11b]. Defense [Ano45a, Ein21v, Gol70a, Inf54, Sla13]. Defiant [Kar05, Bol04, Bra14b]. défient [CT12]. Defining [ARK^{+}99, Hal06]. Definite [Dir79b]. definition [Com95, Com95]. Deflection [Cro20a, Gin22, O’L64, Sau21, DED20, Lam21, Wil88, vK60]. Deflexion [Kia46]. deformable [Ein07a, Ein05k]. déformables [Pia31]. deformierbarer [Ein07a, Ein05k]. dégagée [W.25]. Degeneracy [GR44, Lon38a]. degenerate [Khr08]. degli [Fer24]. Dehmelt [HFN06]. d’Einstein [Ano23a, Lam07, W.25, Ano48b, Ber22, Car22a, Car23, Chr22, Com95, Cun22b, Kle74, Lucxx, Sto96, Wei15c, dB65, dB49a, dB49d, Pia31, Ano23a]. Del [Mas99, Mas02, Bal94, Cli09b, De 12, EI65a, KT15, Lau15, Paz10, Pye86, W^{+}87, WdC02]. dela [Inf79a]. Delano [Ein39a, Ein97]. Delay [Res50, MPBL09, PS22]. delayed [Mad89]. delays [Dic79a, Wal81]. Delbrück [Seg11]. Deledda [Bal75b]. delfinek [Szi05]. Delivered [Wig96a, Ano22a, Ano23a, Ein21l, EA22, Ein22p, EA23, Fer39, Fer70, Lan65, Lan67a, Lor31, Sch28]. della [Arg09, BF02, CM02a, Ein21p, Ein22k, EI48b, EI65b, EI13, Inf63, Inf68, Inf74, Inf92, KT15, Kop23b, Ros80, Tug98]. dell’epistemologia [Arg09]. Dell’Etere [Mas99, Ein22k]. Delta [For14]. Deluded [Mil08]. Delusion [Ano28, Red26]. Demanded [Ano45a, Ein05-34]. Demise [Nor07d, Nor92]. democracia [M08]. Democracy [Lan15, Bus49, Bus68, Bus85, Sca14, Sch02a]. Democrat [GC96b, Jay05, GC96a]. Democratic [Sch05a]. demon [Gol83b]. demonstrate [Coo83]. Demonstration [Ein41a, TEM^{+}89]. demrer [Hau47, Hau50]. demselben [Ein07h, Ein07g]. Demut [Ayn34]. demystified [Wol03]. denature [Ban12]. denial [BL96, LM95]. denied [Har96c]. denigrating [Gwy10]. Denis [Sta99b]. Denken [Ank15, Bre05, BT15, Pfe77, Smi15b]. Denkens [JDB^{+}97]. Denkgewohnheiten [Bäu83, Mel79b]. Dennis [Ham97, Smo07a, Joh05b, TA97, vB01]. denounce [Bie03]. Dense [Hen92a]. density [AH96, Duc11, EdS32, Koh99, SLLS09]. Department [Bro15]. Departure [Hei49b]. depend [Ein05b]. dependent [BRH^{+}09, Ein05v, Fie16]. depicted [Her66b]. Depiction [Rad08]. deployment [Bro92]. Deposing [Rai14]. deposition [HS44]. depth [Gie05]. d’équivalence [Kle74]. Derbes [Hea13]. Derivation [Cha72, Ein46c, Flo98, Kre72, Les74, Lie78, Mat40, Vig95, Ein35b, Ein00b, GS80, Lud13, Mer11, Oha12a, Oha12b, TV07, Wei13d, Yil75, Žáč12]. derivative [Ein17c]. derive [Bal11c]. derived [Che13]. derrotado [De 12]. desarrollo [EI65a]. Descartes [Amo10, Le 52, Vig93]. descobertas [Seg87]. describe [Lin15]. Describes [Bro23, Cam18, Rin03]. Description [Ano35, Hor79b, Ros79a, EPR35, FLN13, Hen92a, Kie15, Kra21, SZ88, vH97]. descriptive [Ein16d]. desde [EI39b, EI65a]. desenvolvimento [EI39b]. Design [HM10, Pin00a, Sus06]. designed [Dan97b]. desirable [RSU93]. deska [Pát79a]. despised [Sch05f]. Dessauer [Goe03]. dessus [Rol15, Rol71, Rol19]. Destroyed [Rus17, Lev15, She75, She87, She03]. destruction [Van04, Wal97, Wal04a]. desvoltarea [EI57]. Deswaan [Sch95a, Sch96b]. Detailed [Dra14a]. Details [KS95, AMS04]. detect [Spo16]. Detectably [Ken13]. Detecting [Dys16]. Detection [AAA^{+}18, Cli09b, Ryf14, Col17]. detective [Nat05]. Determination [Ano26a, Bla13, Coh40, Ein14j, Ein86c, Ein05m, Ein05r, GHK^{+}06, GHK^{+}07, LHF^{+}98, QPSD13, DED20, Eĭn86a, Ein05a, Ein06e, Ein06c, Ein07i, Ein11e, EM23, Ein05-32, Gow88, JN14, SEM03, Sil30, TLW^{+}10, vK60, Ein14j]. determine [Ein27b]. determined [Ein18j]. Determining [Ein87b, Ein05-36, Eĭn85b]. determinism [Ano55c, Ian93]. déterminisme [Ano55c]. deterministic [SZ88]. Detonation [Tra16, GHK^{+}96]. detractors [GTP22]. detto [Lan67c]. Detuned [Bre05]. Deuterium [CR12]. Deuterons [SMKW46]. Deutsch [GRE^{+}01, Die24, Ein33i]. Deutsch-Englisch [GRE^{+}01]. Deutsch-Englisch/ConceptDraw [GRE^{+}01]. Deutsche [Ano79g, Ein31a, Hei53a, SM41, vS14]. Deutschen [WvLuAE18, Ein21d, Gru97, Kle75b, Nic17, Nic19]. Deutscher [Joh95, Sto98, Vis95, Sta36]. Deutschland [BS92, Gru97, Kra88, MGG^{+}95, NS88, Som05a, Som05b, Hei46, WW75]. deutschsprachiger [Hah81]. Deutung [Bor55b, Bor63b, Ein16c, HB63, Kra20a, Sch21]. Develop [Lau55a, WHF^{+}96, Jak13]. Developed [RF69, Ang20]. Development [Bai89, Ber79a, Čap76, Cha74, DiS06, Ein22h, EI39a, EI62c, Hen86a, How96a, Jam89, Kra96a, Lee06, LL82, Meh73b, Pau94a, Pla20c, Pla22b, Pla10, Ram81, Ren93, Sch90, AE06, Bad95, Bay94, Ber74b, Ber74c, Bro92, Dir81a, Ein09b, Ein09c, Ein21b, Ein21o, Ein31g, EI38d, EI39b, EI57, EI59, EI61a, EI62a, EI65a, EI65b, EI66a, EI98b, Ein11b, EI13, FR49b, Fer08, Fey66, Gon96a, Hei34, Hei65a, HRP90, How14, Isa85, Jam66, JDB^{+}97, Kal92, LR14, Meh75c, Müh11, Pau58, Pet21, Pet23, Pla20a, Pla20b, Pla22a, Sei95, Sla69a, Smy45b, Smy45c, Smy45d, Smy45e, Smy76a, Smy78, Smy89, War91]. Developments [Ano62, BW04, CW32a, CHR90, Dir79d, Hie82, BB89, JRG02]. deviation [Ein36e, vS21a, vS21b]. Deviations [Fey05, Ken13]. Device [Ano50b]. devient [Bra17]. Devil [Cor03, Du 52]. dévoilée [Wei15c]. Devoted [Lau55a]. devotion [RCRC97]. Dexer [Duf46]. d’explosions [EE82]. Dezember [HY86, Kra88, Wig64a]. Df150.00/$26.50 [Hen82b]. d’Hiroshima [Ano05-40]. d’histoire [Der15]. d’hypothèses [W.25]. diagram [How96b]. diagrams [Kai05b]. Dial [Mul99]. Dialectic [Bac49]. Dialectical [Ein52f]. Dialectics [ABM79, Ein24g]. Dialektik [Ein24g]. Dialog [APWB95, MGG^{+}95, Ein18f]. Dialogue [Ein30m, EM32c, Ein77a, Kle70c, Sha54, Bel99, Boh85a, Gie05, Hen84, Tym95, Dra53, Dra01, Ein18f, Eĭn64a, Col68, Kuh54]. dialogues [MB21]. Diamond [JDB^{+}97]. Diana [Bad15, Lac08, Mar11c, Pye06]. Diaries [Ein22, Ros18]. diary [Ano04d, Dup06, HW55, Sch04b, Tőz15]. Dice [Ano15e, Gor12, Sau16, Tho83, Gor02, Gor17, Hal15, MB11, Pot11]. dichten [Hen92a]. Dick [Sch09a]. Dictatorship [Ein30l, Ein70c]. Dictionary [Ano07q, DG99, MGG^{+}95, MMMM96, BW^{+}87, FMR90, Kan96c, Kat89, MB^{+}92, Pri90, Tho97, WB^{+}87, Ano07p, Ano07r, Ano07s]. Did [Ano90a, Bal11b, Bal11c, Dor78, Hon96, Ita82, Jan07, Lan14, Mul02, Nor16, Oha09, Pal95, Pei83, Pro74, Rös95, Tom22, TP90, WS89, Wei12i, Wei12j, Wei13e, Wei14a, Wil17, You98, Zah73b, Bre05, Cho16c, CSW97b, Ein19c, Lan67c, Leh14a, MH09, NHG89, SBB16, Wei18a, Zah73a, Zah76]. Didactics [Wet80]. Didn’t [PP05, Jar08a, Jar08b, MHM90]. died [Ano10a]. Dielectric [BCO^{+}11]. Diels [Vai89]. Dienstakten [EGO97]. Dies [Ano55d, Ano68b, Bod13, Fri68, Gla00, Ano00e]. dieser [Ein15g, Sch10a]. Dietrich [Joh95, Smi05a, Vis95]. Difference [Ein05-39, Pop66, Ein02b, Sen21]. Differences [Gol68]. Different [Eöt91, Eöt08, HBN00, MSS17, Dob06, Hal04a, Lau05, von90]. Differential [Dra14b, EF05, Sch17b, EF13]. Differentialkalküls [EF05, EF13]. differentiation [EM34]. Differentiationscharakter [EM34]. difficult [GW10]. Difficulties [Hei93, Hen86a, PFM17, MR82d, MR82e, Uff06]. Difficulty [Sil36, Wal21, Ben99]. Diffusion [ARK^{+}99, Mel63, Pik63, SMHL23, Bor02, Hom05]. Digest [MKBF30a, MKBF30b]. Digesting [Par11]. Digital [Ano01b, Ano15b]. digitization [MFO14, Ros12]. diktare [Sch50b]. dilation [Ein09c]. Dilemma [Ano12c, CP15, Ein79b, Whi06, Gré19c, MH12, Pol05, Pou93, RMR96, Whi96, Dow80]. dilemmas [Hei86, Hei00]. Dilettanten [LT15, Lin17]. dilettantes [LT15, Lin17]. dílo [Inf58]. Dilute [BEC99, CW02]. Dimension [vBRvBR77, Cur82, Hak07, Lie08, Thr17, Hea13]. Dimensional [Hal07, Wal07, WCL^{+}00, EBB41, Tsu99]. Dimensionality [GS22]. Dimensions [Bod94, Ein05r, Gre99a, Gre99b, Gre00, Kas21a, Tho83, YN11, Ein05a, Ein06c, Ein11e, Fv85, Gri10, HY15, Le 03, Lor21, Ran05, Van85, YN10, Ein05m]. d’incomplétude [Pat85]. Dinge [Ein52e]. Dingle [McC77]. dintorni [Bri05d]. Dio [Pot11, Ano25b]. Diodes [LU13]. Diploma [Tes99]. diplomacy [Alp65, Alp66, Alp85, Alp94]. Dirac [Ano33c, Ano51a, Bri18, Wig96a, Wil36, Ano34b, BOS96, Bal85, Bok08, DS83a, DD95, EM33b, EM33a, Far09, Faz15, FH10, Gru00, Han22, Kap40, Kra82a, Kra90, KW87, LW34, RA11, RdO07, RdO16, Tau49, Wie29, Wig96d, Wig96k, Zic00]. Dirac-Gleichung [EM33b]. Diracschen [EM33a]. Direct [GSPH00, WCL^{+}00, WF49]. Directed [Ano68a, Wac06, Kle75a]. Direction [Zeh07]. directions [Edd21]. Dirichlet [AKD23]. dirt [Pas10]. Disability [Hal12]. disagreeing [Gio13b]. Disagreement [Räz16, Dic79a]. Disappear [HG05b]. disappearance [Mag09, You91]. Disarmament [Ein31b, FGMvH14]. Disarray [Ren07a, CGR^{+}97]. Disasters [Mah14, Ano07k, Mah15]. Disastrous [Ano45a]. Discarded [JR07]. discards [Ano19a]. Disciple [Ren93]. disciplinarité [Pye75]. disciplinary [Fie15]. discipline [Goe17, Pye75, War89]. Discontents [Gin08, Kle65a, MS09]. Discontinuity [BDH^{+}00]. Discours [Pau47b].

Discourse [Kie94]. Discover [Nor16, Bal11b, RR17, Tho15]. Discovered [RMI21b, Rus17, Ban06, BB00, Ein22r, Lev15, RMI21a, Sch04b, Lin15]. Discoverer [Ano68b, BH01, SN93, vL55c]. Discoverers [Fos13]. Discoveries [Ano00, Bro23, Bar06, Bra09, BB13, Cal17, Cam09b, Die05, HL08a, HL12a, Kle81, Lig05, McG93, McG98, Par04, PS22, RF05, Seg76, Seg80, Seg84, Seg87, Seg07a, Wal02a, Wat00]. Discovering [NB09, HMW24, PA06]. Discovers [Ano13b]. Discovery [Ano39a, Ano39f, Ano82d, Ano05b, BHT86, BC82, CR12, DR89, Fos13, FW67, Gol72, Gra64, Hen86a, Kae13a, Kni13, Max93b, Mei62, Mei63a, Mei63b, Noe11, O’S01, OR21, Pee14, Seg89, Sin02, Spa64, Sta82b, Top07b, Wei12v, Zah80, AKD23, Ano01c, CHW91, Cha80, Cha49a, Cha17, Coo10, Coo15, Das91, DTF97, Dys16, FW85, GA71, GA81, Gro05, HKS10, Her01a, Jun05a, Kae48, Kra88, Lau59, MR82a, Meh98, OP87, Oyi04, Pop07, RS96, Rob12, Row06d, Shi01, Smo07c, Spo16, Wei18b, Wue05]. Discrepant [VT99]. Discrete [Hag14]. discurso [Ein31m, Ein31n]. Discusses [RMI21b, RMI21a]. Discussing [Mil12b, Fri11]. Discussion [Ano20a, BPS87, Boh49, Nag82, Roy19, Sch35b, VW86, Wei12k, Bus49, Bus68, Bus85, DB86, GHK^{+}96, GP11, ST94a, W.25]. Discussions [Hei69a, Hei72, Meh87, BF78, LdB12a, LdB12b, RS80b, RS13]. disease [KWM57]. diseases [RF05]. Disembarked [Ano82d]. disentanglement [Hah42]. Disintegration [MF39a, CW32b]. Disk [CLD^{+}03, Pag21, Sta80a, Sta89a, Wei14b, Wei15b]. Diskurswandel [Kie94]. Diskussionen [Tre66]. disloyal [Gro56]. disorder [Tsu99]. dispatches [Wel06]. Dispelling [PW17]. dispergierenden [Ein22n]. dispersion [Kai05b]. Dispersive [Ein22n]. displaced [Ano03b]. Displacement [Cha23, Cro20b, DR20, Mou20, Sie91]. disposés [Ein10e]. Dispute [CRS97, Hol78a, Mur94, Beh05a, Nor93]. Disputes [Ken54]. disruptive [LLH22]. Dissertation [Bro99, vM00]. Dissidents [Fre15b]. Dissociated [Ein05-39, Ein02b, Ein20l]. dissociierten [Ein02b]. dissociirten [Ein05-39]. dissoziierten [Ein20l]. Distance [CHS97, Fra14, Hub29, Ano15j, CPDH^{+}87, Ein19c, Mus15b, Web99]. distances [Hen91, Pap15]. Distant [IR85, Ein28e, Ein30a, Ein30r, Sal74, Sch12e]. distinction [Gio20, Mon04]. distinguished [Cha83]. distort [Diz05]. Distribution [Ano03, DR08, Pla00b, Pla01]. distributions [MS82]. District [HTS83]. distrust [Abi03]. Disturbing [Dys79, Hon96]. Disunity [PA21]. dit [Mon12]. Divergent [Hoo91]. diversions [Mer16]. Divided [How98, Isa09, Clo15]. Division [Fri39b, SG82]. divulgativa [Ein60c]. dlëiìa [Smy46]. DM [BS92, Hei53a, Hen86a, Kle70b, Kle70a, Ros86]. Dmitrii [Ste07]. Do [CÖT15, Ein38c, Ano15f, EI59, EI62a, EI62c, EI66a, EI98b, FL88, Gin04, Inf55d, Joh14, Mui94, NA92, SBB16, WS88, WS08, Ein19j]. dobrodruzstvi [EI58]. d’observations [Ein14j]. Dobzhansky [Adl03a, Cec01]. Doctor [Ada07, KS12, Jan77]. doctors [GKP12b, Bur45]. Doctrine [ST06, vBT96]. Documentary [BRS^{+}09a, Car04a, Est06a, For80a, Ham97, Mos16a, Rad08, CHW91, Flü74, GA71, Her66b, PLM^{+}0 , See54, See56a, SFW91]. Documentation [Cra90, OP87]. Documenti [Tug98, Tio09, Ver01]. Documents [Bla92, Grü90, Hor79b, SKD^{+}87, Ber74d, CSW97b, DC12, Dör05, Hah79, Tio09, Ver01, RD05]. Does [CF12, Fre31, Kap40, CF09, Fra37, Gor17, Mas09, MB11, Pot11, Ein05b, Ein23c, Ein27b]. doesn’t [Coo83]. dog [Orz12]. Dogs [Bro21]. doi.org [Ree16a]. Doing [Bus13b, Das12]. dokumentarisch [Her66b]. dokumentarische [Flü74, See54]. Dokumente [For80a, KDK87, Grü90, LS05, RD05, SKD^{+}87]. Dokumenten [Hah79]. Dolci [Bes93]. Dollar [Bro79a]. dolphins [Szi92, Szi05]. domains [SISK^{+}98]. domestic [Bus12c]. dominant [Ler91]. Domnele [Kot93]. Don [Gly90, Pye95, War96b]. Dongen [Del12, Ken11]. donkey [Ost15]. Don’t [Pin15]. doom [Sca14, Lan15]. Doomed [Eck90, Hec16, Ree17c]. Doomsday [Smi07b]. Doors [WM08, Wid07, Elz06]. doped [GB15]. dopis [Kot93]. Doppelspalt [Kon15]. Doppler [Edd26, Lod25]. doprinos [Mag95]. Dordrecht [Hen82b, Wal02b]. Dormant [Ein05-35]. Dorothy [But23, KKJ99]. dossier [Ano15a]. dossiers [Gru05c]. Dostoevski [Kuz79]. Dostoyevsky [Kuz72]. Double [Ano10e, RS07a, RB81, Alf01, Kon15, Lie78]. Double-checking [Ano10e]. Doublet [Tur66]. Doubling [Gin22]. Doubly [Cho05a]. Doubt [Con35, Ken19, Nic20]. Doubtful [Kae39]. Doubts [D’A01b]. Dougall [Lam30]. Douglas [Bok15, Ger15, Kle14a]. d’ouvrages [Bal91, Bal96, Bro83, Col68, Don70, Hum51, Lam07, Mat86b, Mat86a, P.50, Ram82, Ram83, Spe62, Spe67, Spe68, Tat64, Wal99]. dove [Mar91]. Down [Par91, Ano05-34, Lau05, Sch04c, Swi91]. Down-to-earth [Par91]. Downfall [Fra99, Gia09]. Downwind [War99, Bal86]. Downwinder [Mil13b]. Downwinders [Ano94a, Ell10, Hum13, Mil08, OO01, PNT87]. Dr [Ano54a, Ein32e, Hei53a, Hei65b, Urb22, Ano20c, Ano10d, Goo04a]. Dr. [Ano12, Ano29a, Ano39e, Ano46b, Ano48a, Ano55d, Bar48b, Bar50b, Bar50c, Bar52b, Bar57b, Bar80b, Bar82, Bar05a, EM49, Gla00, Goo04b, HL10, Ken54, KSG63, Lau55b, MA31, Smi07b, Swa02, VFIS48]. Draft [Voi86]. drafts [Dee08]. Drag [Ari98]. Drago [Cho16b]. dragon [Hac87]. Drake [Col68, Kuh54, Sha54]. Drama [Bod17, Bru77, Hie84, Ken54, Rob58, Val54a, Whi83a]. dramatic [Cas17, Nel14]. Drawing [Kai05b]. Drawings [Bru09]. drawn [Ein07h, Ein07g]. Draws [Bar18c]. dream [Amo10, Gun11, Kir05, Par86, Sal05, Smi07b]. Dreams [Lig94, Haw11, Tau86, Tau88]. Drehwaage [vE10]. Drei [Ano69c, Des72, Erb10, Lor14]. Dreier [Sch05c]. dreißiger [MH71]. Dresden [JP08]. Drie [EI38a]. Drift [Ano25d, Swe13]. Driver [Pat01a]. drivers [Ein10e]. Driving [Ano00a, Cra00a, Goo00, Pat00, Pat01b, Pea06, Vea01, Ano00b]. drogi [Inf33]. Drohen [Bre08]. Dron [Jos95a]. Drop [Ano94b, Ham97, Kae31, Lef95, Sig78, Stu94, BP79, GF65, GF67, Pos11]. dropped [Jak13, Jen02, Tak95]. Drops [Fra81]. Drouin [Ebu16]. Drude [Ren97a]. Drugs [Har07, Pic05]. Dry [Hea13]. Dual [Eks99]. dualism [Bac89b, Sta06, dB55a]. dualisme [dB55a]. Dualismus [Bac89b]. Dualité [dB49a, dB49d]. Duality [Bac89a, Kle64, NM12, Bai92, Bai01, Van07a, dB49d]. Ducommun [Dur12]. Dudley [Ano87b]. due [vS21a, vS21b]. Duel [Eds82, Wad81b]. Duerbeck [Sch09a]. duets [Fre06]. Duhem [Bur90, How90a]. Dukas [Ber79b, For80a, Moy81, Ram79c, Wig79, Hei74a]. dumbed [Ano05-34]. dummies [Cal05c]. dumps [Ano96d]. d’un [De 13]. d’une [Edd21, Ein14c, Lar12, W.25]. dunkle [See56b]. duplicate [Wal81]. DuPont [Ndi07]. dur [Ein34m]. Durata [Tug98]. Duration [Ber22, Ber65]. Durch [Gás54, Ein26b, HS39a, HS39b, Hen90, Ing05, JDB^{+}97, Jun05a, Pye93, Row06a, Row06b, vS21a, vS21b, Tre66, Urb22, Wue05, Yos95]. Durée [Ber22, Car22a, Car23]. During [Uvn82, Bru03b, GC96a, GC96b, HM11, HHMW93, Ill06, Mel00, Rol47, WK12]. Dürr [GRE^{+}01]. Düsünmenin [Bal08]. Dutch [Alp11, Ano20b, Ano20c, Ano15f, Cur38b, Ded85, EdH15c, Ein34g, EI38a, Gui09, Her11, HC94b, HC96b, Ris06, dGK11, vC05a, vC05b, vE95]. Dutton [Woo28]. Dwarfs [Ren06, Sri99, Str05b]. dynamic [Hen97, Ren97b, Smi02]. Dynamical [Cha49b, Cle82, Wis82, Bro05, Gio14]. Dynamics [Buc13, Car07a, Cha49b, Dir49, DCC^{+}01, MO06, Poi07, RWL23, Wei12x, Bäc19, Bäc21, Göt92, JW11, OB05, San15, Wes80]. Dynamik [Bäc19, Bäc21, MGG^{+}95]. dynamischen [Ren97b]. dynamite [Fin90a, Pau42, Pau47a, SS29]. Dyslexia [Kan92b]. Dyson [BHJ^{+}00, Ber18, Sch13a, Sch94]. dzieło [Inf56a].

e-book [Myr21b, Myr21c, She15, Sho21]. e-hû [Ayn34]. E-P-R [Mer85a]. E. [Ein22b, Grü73b, Sch19, Wig96j, Klo63]. E.d.B.B. [Vig95]. Each [Ano39g, Pin15, Rös95]. earliest [EI59, EI62a, EI98b]. Early [Abb02, Bal91, BH87, BN97, Ber88, Ber93, Boy85, Boy94, Bro39, Cus82, Dir82, EI38b, EI38c, EI66b, Fer81, Gin82, Gol68, Gol87, Hel84, Hen88a, Hen88b, Hol68a, Hom06, How90b, Jam74b, Kae38, Kai81, Ker89b, Kle65b, Kle80, Lan48, LL82, Loe88a, Loe88b, McM51, Mcc87, Mei12, Mel82, Nor07d, Pal83a, Pal83b, Pye76, Pye89, Ram83, Ree87, RS03a, Smi89b, Smi89a, SCS87, Sta94, Whi83b, Yan87, Zia94, Ahe03, Bro79b, Byr80b, CMT18, Cho08, EI42, EI47, EI61d, EI71, EI94, EI08b, EI11, Gea90, Got00, Käl22, Ker87, Kle82b, Leg04, Luk02, Mil81, Mil98, Nav91, Nor92, PR11, Ren08, Sau07a, Sch05c, She16b, Smi10, Tan11, ZDSF11, Sal19, Stu93]. Earth [Ano17c, Gut98, HEB^{+}80, HS39c, MFHR06, Par91, Tym95, Ano39h, Duc11, Ein27c, Eöt91, Eöt08, Hen91, WTB09, von90]. Earthbound [Law00]. earthmoving [Kir05]. easier [Ban12]. easily [Ang20]. East [Con05, Ros18, Ein19d, Sut97]. Easy [Ein23h, Fey95, Fey97, Fey11, Par00, Ano20a]. eau [Hau48]. eboluzioa [EI00]. eccentric [Diz05]. eccentricity [Reg87]. Eccles [Amo10]. Eckart [Eck70, Eck73]. Eclipse [Ano19b, Ano19c, Ano24a, C.23, Cam18, Cha23, EG80b, Edd19, Ken19, Ken22, Mat40, Moy79, Nic20, Rus37, Sim22, Sta03b, Bau20, DED20, GTP22, Pan02, Sou19, Spo02, Sta05d, Ren21]. Eclipses [DvdRW37, EG80b, RS03a]. Eclipsing [Ben05]. École [Pia31]. Economic [Zel90, Kat89, Lin99, McC01a, Pet08, Tor03, vN45]. Economics [Tre79b, vN37, Ano97a, Hor09, Kar10, Ng11, Put09a, Tol23, VM05a, ZZY19, BOS14, CMT18, Tom10]. economist [Put09b]. economists [BS86, BH04, BH09, FG20, Ng11, Sch87a]. Economy [Eng05b, Wil77]. écrits [Bor02, EM^{+}91]. Ed [Hie68, Ree17c, For05, Han80b, Hen82b, Hen88b, Kau14, Kil84b, Ros80, Smi51, Vis94]. Eddington [Adl03a, De 24, Ren21, Ano09f, Bro23, Cha83, Coh31, Edd79, Ein25c, Hol30, Ken22, Kil94, MPM^{+}08, McC91, MCS09, NA92, See25, Sou19, Sta86, Sta03b, Sta07c, Sta07d, Tay21]. Eddingtons [Ein25c]. Edgar [Jer02]. edge [Bra12, Fie15, Gle10, Hol18]. Edinburgh [Hen86a, ABdB^{+}53]. Edison [Ein29a, Tho88b]. Edited [Cah04, Dix00, Duf46, Eis00, Hea64, Lin60, Van22, Woo28, Ano84a, Bus12b, For80a, Hea13, Kat52, Kil65, Mic68, Oha79, Pye06, Ram79b, Ram79c, Ram81, Ram90, She15, She16b, Smo07a, Wig79, vD15]. Edith [Chu60, Chu98b, Chu01, WB01]. Editing [Sau04a, Sau04b, Bin06, Ern92, Sta87]. Edition [Ano15b, BRS^{+}09a, Ein21p, Ern92, She16b, Ein23h, Ein53n, Hei53a, vB02, Bor50, Ram79b, Col68]. Editor [Bar05f, GT38, Moo96a, Cah90b, Ein47f, Tou05, DiC15, Don64, Hub13, LRD^{+}91, Wal91b]. Editorial [BGLR17, DiC15, Ehl09, Lum11, RS07d, Rom91, Wal88b, Kni16]. Editors [Ano85, Cah90b, Mar11c, Kan92a, Nev92]. editorship [Whi67]. editrice [Tug98]. Edizioni [Per10]. Edmund [Kar05]. Edoardo [Bat03]. Eds [Gly90, Pec04, Bad15, Goo85, Hus15, Koj08, Sci80, War96b, Lan08]. Eduard [Coh54, Duv64, Jos95a, RK86]. educated [Rei41]. Education [Ano05q, Ein20m, Ein21f, Ein34b, Ein36g, Hun15, Kle75a, Rei07, KST99, Kra88, Tan11, Gon05]. Educational [Ano31f, Hen82b, Pin00a, CT15, Mai99]. Edward [Ber03, Ham97, Her03, Sch03c, You95, BO76, BH89b, Goo04a, Goo04b, Her02, HM05b, Wig96b]. Edwin [SN93]. éeconomie [Dur12]. eeuwen [EI38a]. Effect [Cle47, Cra90, DR89, Fri17, Mat40, MBD39, Val49, Adl82, Adl83, Bau20, BCO^{+}11, Coo83, Ein12d, FN13, Har03a, LLSH23, Sto99, Stu75, Wie06, Wof08, Woo25, Sin02]. Effective [Mar08]. Effects [Fre85, KS65, Mal12, AAA^{+}09, Ano81c, Buc94, EE82, FW17, Gla57, Gla62, GD77, GHK^{+}96, KRZ^{+}05, LLSS07, MDT^{+}03, MHY95]. Effekte [KRZ^{+}05]. Effets [EE82]. Effort [Gon96c]. Efforts [Ano45a]. Egypt [Wit05]. ehemaliger [RK86]. Ehlers [Sch00c]. Ehren [Ano69a]. Ehrendoktoren [GKP12b]. Ehrenfest [Ein05k, Som05a, Som05b, Ano05l, Ein07a, Ein11c, Ein34i, Ein05b, Ein05k, Kle59, Kle70d, Kle81, Kle10, Som05a, vD14]. Ehrenfestschen [Ein11c]. Ehrenhaft [Hol78a, San11]. Ehrlich [Hün18, Hün18, Wit82]. Eidgenössischen [EGO97]. eight [Nor93]. eighteen [BH04, ABH^{+}55, ABH^{+}87]. Eighth [Rho88a, RT90]. Eighty [Ein55a, See25]. einatomigem [Ein11g, Ein05q]. einatomigen [Ein24h, Ein25h]. Einblicke [Mos21, Mos16b, Mos22]. Eindeutigkeit [How92, How96a]. Eindrücke [Ein21g]. Einege [Ein16d]. einem [Des72, Dir60b, EH10a, Ein26d, EM31b, EH05b, Fis02, Rei41, Som05a, Som05b, Spe68]. Einfach [Vaa17]. Einfache [Ein21e]. Einfaches [Ein16e]. Einfluss [Ein27e, JP52, Ein11i, Inf53a, Inf56b, Inf57b, Ein27c]. Einflußder [Ein05-31]. Einführun [Sch20, Sch22a]. Einführung [Sch17, Sch19, Thi46, vB12, vB14, vB18, vB20, Ein10b]. einheitliche [Ein28c, Ein30a, Ein25d, Ein29b, Ein29c, EM31a, EM32a, Ein38a]. Einheitlichen [Ein29k, Ein30o, Ein30f, EM30]. Einige [vI11b, ES13, ES05a]. Einiges [Hah42]. Einleitende [Ein53c]. einmal [Kle85]. Einshtein [HC98, Lan67a]. Einshteina [Kol40]. Einshteine [Inf56c]. Einstein [Adl03a, AMM99, AVvdV99, Aga07, Ank15, Ano20a, Ano21i, Ano22a, Ano28, Ano39d, Ano50a, Ano69c, Ano79d, Ano79h, Ano79f, Ano79g, Ano81e, Ano84a, Ano84c, Ano84b, Ano93c, Ano99b, Ano99c, Ano00b, Ano05o, Ano06i, Arm38, ARK^{+}99, B.50, Bac89b, Bac05, Bad15, Bal75a, Bal85, Bal91, Bal96, Bal99, Bar94b, Bau09, BT15, Bäu83, B^{+}79, Bel26, Ber07a, Ber96b, Bes04, BHJ^{+}00, Bla50, Bla76a, Bla76b, BGLR17, BSW22, Bod03, Bod17, Bor55a, BB69, Bre15, Bre05, Bri18, Bri66, Bro66, BAM99, Bro72, Bro83, Bro39, Buc80, BW91, Buc93, Büh98a, Bus12b, C.51a, Cah90b, Cah90a, CR07, Cao06, Cea98, Chi08, Coh51, Cus82, Del13, Del11b, Der15, Dos17, Dys09, Ebu16, Egg02, Ehl05a, Eis00, Els24, ER13, Eng17b, Erb10, Fer07a, Fer81, Fla15, Flü74]. Einstein [Föl93, For70b, For70a, For80a, For09, FF97, Fre85, Fre19, Fre27a, Fre27b, FTG08, Fri73, Gil19, Gin82, Gio13a, GKP12a, Gli84, Gly90, Gol84b, Gol87, Gol04, Gol97, Gon05, Gor02, GRE^{+}01, GRC18, Grü90, HEB^{+}80, Har07, Har91, Har96a, Har96b, Har95, Hea13, Hei74a, Hei49a, Hei53a, Hel98, Hel84, Hen82b, Hen86a, Hen88a, Hen88b, Hen86c, Hen92a, Her66a, Her78, HvMW79, Her33, Her63a, Her66b, Her67, Her75, Her17, Hes15b, Hie68, HC94a, HC96a, HC04, HY15, Hof08c, HE82, Hol05b, Hor00, Hou65, Hub13, Hun01, Inf53a, Inf56b, Inf57b, Inf69, Inf99, Ing05, Jan77, Jeh51, JW06, JDB^{+}97, Jor69, Jos80, Juc04, Jun05b, Jun05a, Jun08, Kae38, Kai81, Kan96a, Kar05, Ken25, Ker89b, Kev05, Kie82, Kie15, Kil65, Kil84b, Kil95, Kil98]. Einstein [KT79, Kle70b, Kle70a, Kle73, Kle75a, Kle75b, Klo63, Kni12a, Kni12b, Kol74, Kon15, Kra06, Kra20b, Kra20c, Kra25, KRZ^{+}05, Kra14a, Kre72, Krö53, Kru94, Kuh54, L.58, Lac08, Lam79, Lam30, Lau62, Leg03, Leg15, Len39, Lig97, Lil86, Loe88a, Loe88b, LS05, Maa13a, Maa13b, Mal65, Mar73, MGG^{+}95, Mas01, Mat86a, Mcc87, Mcc55, Med76, Mel79a, Mel79b, Mel82, Mic68, Mil11a, Mil12b, Moy81, Müh11, Mur25, N.38, Nag50, Nah01, Nic20, Noe93, OSWR86, Oha79, Oha12a, Oha12b, Ost15, Pal83a, Pal83b, Pap15, Par06, Pau58, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Pep04b, Pie24b, Pir07, Plo06, PB72, Prz63, Prz86, Pul11, Pye80, Pye89, Pye95, Pye97, Pye06, Pye10, Qui40, Ram79b, Ram79c, Ram79a, Ram90, Ram83, Red84b, Rei41, Rei50, RS92, Ren97b, RCD98]. Einstein [Ren05c, Ren05b, Ren05e, Ren05g, Ren21, RS11b, Rob38, Rob48, RR89, Ros05, RS13, Row00, Row06a, Row06b, RK86, RSPS49, Rus17, Sar35, Sch10a, Sch49a, Sch51a, Sch55, Sch79a, Sch79b, Sch83a, Sch50a, Sch81a, Sch88b, SH99, Sch92c, Sch05e, Sch12f, Sci80, See65b, See65a, See52, See53, See54, See55, See60, See64, See80, See86, See91, Seg91, Ser88, Sha75, Sha54, She16a, Sie97, Sim06, Smi51, Smi89b, Smi89a, Smi02, Som05a, Som05b, Som49b, Som49a, Som55b, Spe67, Spe72, Spi96, Sta99b, Sta00, Sta20, Sta95c, Ste05a, Ste05b, Ste22, Ste90, Sti05, Str90, Sud11, Swa02, Tay21, Tob94, Tou05, T– D82b, T– D92, Tre66, Tre71, Tre79a, Tre21, Tug98, Tur06a, TH79, Vaa17, Vir92, Vog20, Wac06, WM08, Wal99, Wal02b, War95]. Einstein [War96a, War96b, Waz10, Whe80, Whi83b, Whi43b, Whi55b, Wig79, Wil16a, Win69, WE24, WE87, Wis79, Wol87, Woo28, Wre66, Wue05, Yan87, You98, YGKK05, Zau05, Zim84, de 15, dP17, vD15, vB05, vL55b, vL55c, vM05b, vM09, vM12, vMHW85, vM93, vM96, vM99b, vM01, vM05c, BBD^{+}11, DC10, Hen85, Abb98a, Abb98b, Abb02, Abb05, AKBAE21, Abi00a, Abi00b, Abi03, Abr12a, Abr12b, Abr01a, Abr02, Abr03, Abr04a, Acz99, Acz00, Acz01, Acz03, Adl82, Adl83, Aer99a, Aer99b, ABW99, Aga06, Agu90, AS79, Alb92, Ald93, All05, All90, Alp11, AH60, Alt13, A^{+}09, AC09, Ame20, Ami87, Amo10, AH96, And15, And17, Ang05, AK02, Ank15, Ano15a, Ano11b, Ano17, Ano19a, Ano19b, Ano19d, Ano20b, Ano20c, Ano20d]. Einstein [Ano20e, Ano20f, Ano21b, Ano21c, Ano21g, Ano21d, Ano21e, Ano21f, Ano21h, Ano21j, Ano21i, Ano21m, Ano21a, Ano22c, Ano22a, Ano22e, Ano23b, Ano23e, Ano23d, Ano24a, Ano25b, Ano25d, Ano25c, Ano25a, Ano26a, Ano26b, Ano27, Ano29a, Ano29g, Ano29b, Ano29e, Ano29d, Ano29c, Ano30a, Ano30b, Ano31a, Ano31b, Ano31e, Ano32, Ano33a, Ano33b, Ano33d, Ano33e, Ano33f, Ano33g, Ano33h, Ano34a, Ano35, Ano39e, Ano39f, Ano46a, Ano46b, Ano47b, Ano48a, Ano48b, Ano48c, Ano49b, Ano50c, Ano54a, Ano54b, Ano54c, Ano55a, Ano55b, Ano55c, Ano55d, Ano57b, Ano59, Ano66, Ano67, Ano68a, Ano69c, Ano69b, Ano70, Ano75, Ano79b, Ano79a, Ano79c, Ano79e, Ano79g, Ano80a, Ano80b, Ano81a, Ano81b, Ano82c, Ano83, Ano84c, Ano84b, Ano87a]. Einstein [Ano90a, Ano90b, Ano91, Ano95b, Ano95c, Ano96c, Ano96b, Ano97b, Ano97d, Ano97c, Ano00a, Ano00c, Ano00e, Ano02a, Ano03a, Ano03c, Ano03b, Ano04a, Ano04c, Ano04f, Ano04g, Ano05c, Ano05e, Ano05g, Ano05f, Ano05h, Ano05k, Ano05i, Ano05j, Ano05l, Ano05d, Ano05n, Ano05p, Ano05-31, Ano05q, Ano05-33, Ano05-32, Ano05-36, Ano05-37, Ano05-34, Ano05-30, Ano05-35, Ano05-39, Ano05-38, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05r, Ano05-41, Ano05-43, Ano05-45, Ano05-44, Ano06c, Ano06e, Ano06f, Ano06h, Ano06g, Ano06j, Ano07b, Ano08a, Ano08c, Ano09a, Ano09d, Ano09e, Ano09f, Ano09g, Ano10a, Ano10b, Ano10c, Ano10e, Ano11b, Ano11a, Ano11c, Ano12a, Ano13b, Ano13c, Ano13a, Ano14a, Ano15b, Ano15e]. Einstein [Ano15c, Ano15d, Ano15f, Ano15g, Ano15j, Ano15l, Ano16a, Ano16d, Ano17c, Ano18a, Ano18c, Apf81, Are05, Ari03, Ari05, Ari14, Ari98, ABM79, AP65, Arr07, Ash11, Ash12, Asi83a, Asp69, Aud06, BF16, BC24, Bac89b, Bac89a, Bac05, Bäc19, Bäc21, BH79, Bad83a, BOS96, Bai92, Bai01, Bai89, Bai07a, Bak97, BV02, Bal05, BR05, Bal92, Bal93, Bal94, Bal95a, Bal95b, Bal99, Bal01a, Bal08, Bal11a, Bal01b, Bal11b, Bal11c, Ban51, Ban06, Ban11, Ban12, Egg95, Bar07, Bar80a, BS81, B^{+}06, Bar48a, Bar48b, Bar50a, Bar50b, Bar50c, Bar51a, Bar51b, Bar52a, Bar52b, Bar54, Bar55, Bar56, Bar57a, Bar57b, Bar58, Bar62, Bar80b, Bar82, Bar98a, Bar05a, Bar18c, Bar18b, Bar05c, Bar94b, BNY81, Bar96, Bar00c, Bar05g, Bar15, Bas65]. Einstein [Bat05, Bau09, Bau20, BT15, BS10, Baw06, B^{+}79, BSSP^{+}08, BBL08, Beh05b, Bel64, Bel96a, Bel98, Bel93, BCR93, BT12, Bel96b, Bel23, BM93, BM09, Ben16a, Ben99, Ben01, Ben05, Ben14, BL11, BC01, BN88, BN97, Ber74a, Ber92a, Ber82b, Ber22, Ber65, Ber74b, Ber74c, Ber82c, Ber94, Ber11a, Ber81, Ber90a, Ber73, Ber76c, BF78, Ber79b, Ber96a, Ber96b, Ber96d, Ber02, Ber06c, Ber90b, BGA^{+}13, Bha92, Bha86, Bia16a, Bič79b, Bič79a, Bič79c, Bič89a, Bič89b, Bič00, BL14, Bie87, Bie91, Big05, Bin06, B^{+}89, Bje02, Bla88, BRS^{+}12, Blu22, Blu89, Bod03, Bod16, Bod13, Boh05, BRH^{+}09, Boh49, BR55, Boh85a, Bok08, Bok15, Bol04, BPL05, BBCPL09, Bon64, Bon94, Bon79, Bon81, Bor66, Bor20, Bor22a, Bor22b, BB24, Bor49a, Bor50, Bor55a, Bor56]. Einstein [Bor62, Bor63c, Bor64, Bor65b, Bor65a, BI67, BB69, BI69, Bor69a, BB71a, BB71b, BB05, vBT82, vBT94, Bou99, Boz14, Bra15, Bra17, Bra22, Bra02, Bra29, BK05, Bra24, Brd79, Bre79, Bre15, Bre73, BDR^{+}07, Bre89, Bre05, Bre08, Bri96, Bri97, BBZ98, Bri05b, Bri05c, Bri49, Bri05d, Bro78a, Bro79a, Bro06, Bro07, BAM99, Bro73a, Bro10a, Bro11, Bro16, Bro19, Bro20b, Bro80, Bro21, Bro23, Bro79b, Bro88, Bro99, Bro22, Bru97a, Bru98, Bru03a, Bru09, Bru82, Bru77, Buc80, Büc21, Buc13, Buc20, BRI^{+}21a, BRI^{+}21b, BW91, BW92, Buc93, Büh98a, Büh15, Bun79b, Bur90, BEC99, Bur94, Bur97, Bus03, BLI^{+}02, But07, BRS03, BR99, Byr80b, Byr80a, Byr81, C.23, C.50, C.51b, Cah85, Cah55, Cal96, Cal00, Cal02, Cal03a, Cal04, CL05, Cal05a, Cal10, Cal13]. Einstein [CKS15, Cal79, Cal80, Cal82, Cal05b, Cal05c, Can05, Can15a, Can15b, Can16a, Can16b, Cao07, Čap71a, CF11, Cap05, Car17, Car82a, Car82b, CT24, CLD^{+}03, CC84, Car22b, Car24, CB06, CM67, CH98, Car89, Car07b, Cas75, Cas74, Cas94a, Cas94b, Cas95, Cas04, Cas05a, CL19, Cas21, Cas23, Cas03a, CGR^{+}97, Cau91, CW12, Cea95, Cea97, Cec86, ČK79, Cer00, Cer11, Cha11, CC67, Cha72, Cha78, Cha01, Cha23, CF80, Cha79, Cha05b, CD07, Che00, Che13, Cho03, Cho06, Cho16a, Cho10a, CB09, Chr20, Chr22, Cla68, Cla71, Cla73, Cla84, Cla78, Cle17, Cle08, Cle06, Cob07, Cob40, Coc16, Coh31, Coh55, Coh79, CS79b, CG90, CC96, Col09, CH80, Col57, Col93, Col95, Col00a, Col68, CSD^{+}07, Col00b, Com95, Com23, Com14, Coo83]. Einstein [Coo10, Coo15, CL07, CSvB99, CW98, CW02, Cor15b, CRS97, Cor04b, Cos80, Cos87a, Cot89, Cou52, Cou79a, Cou79b, Cra97, Cra00a, Cra79, CS60, CG14, Cre81, Cre83, Cre06, Cro16, Cro05, Cro20a, Cro20b, Cro07, CPDH^{+}87, Cug06, Cul80, Cul81, Cun19a, Cun19b, Cun19c, Cun20a, Cun20b, Cun20c, Cun22a, Cun22b, Čvo06, Cwi87, Cze05, D’A50, D’A01a, D’A01b, D’I92, Dac11, Dac20b, Dac20a, Dad78, DL15, DS01, DGPS99, DDDR06, Dam06, DC08, DC12, Dan89, Dan97a, Dan97b, Dan05, Dar06a, Dar00, Dar19, Dar21, DSZ12, DR08, Das12, Dat06, Dav74, Dav81, Dav95, De 24, Deb29, Deb79, Ded85, Del75, Del11a, Del12, DDD^{+}06, Den84, DHY10, Der15, Des79, Des72, DiC15, DiS92, DiS06, DSMH85, Dic81, Dic79a, Dic79b]. Einstein [Dic80, Die05, Die88, Din59, Din93, Dir79a, Dir79d, Dir79c, Dir79b, Dir79e, Dir80c, Dir80b, Dir80a, Dir80d, Dir81a, Dir79f, Diz05, Dob06, Dol18, Don97, DCC^{+}01, DCTW02, DD11, Dor71, Dor78, Dos17, Dot82, Dow80, Dra14a, Dub05, Duf81, Duk49, DH79, DH81, DHR13, Dum04, Dup06, Dur02, Dur05b, Dur05a, Dur08, Dür79, Dür82, Dvo12, DM15, DGG^{+}21, DT23, Dys15, ESA15, EG78a, EG78b, EE99, EH05a, Edd24, Edd26, Edd29a, Edd30a, Edd79, Edd95, Edw00, Egd14, EE05, Egg02, Ehl79, Ehl05a, Ehl05b, Ehr77, Ein21g, Ein21f, Ein24f, Ein30g, Ein31m, Ein31n, Ein36c, Ein39a, Ein45a, Ein45f, EM49, Ein54c, ER55, Ein55f, ENN60a, Ein60a, ENN60b, Ein63a, ENN63, Ein63b, ENN68, EHM^{+}79, EM84, Ein85c, vM90, Ein91g, EBRC93, Ein95a, EMH96, Ein96a]. Einstein [Ein96e, Ein97, EC00, ECS02, EK03, Ein03a, Ein05a, Ein05-45, Ein05b, Ein06a, EKS^{+}08, Ein09b, Ein09c, Ein22, Eis22, Eis23, Eis93, Eis02, Eis06, Eis07, EM04, Eis11, Elb13, EO79, Ell05, Els24, Elz95, Elz06, Ema85, Eng02, Eng05a, ER13, Erb10, Eri82, Erl72b, Erl72a, Esp34, Esp55, EG98, ECS19, EGW^{+}08, Eva95, ECEP11, Eza79a, Eza79b, Ezr82, Ezz87, Fab79, Fal09, FLN13, Fan06, Fan03, Far05b, Far05a, Far03b, FG21, Far05c, Fau79, Faz15, Fel79, Fel82a, FW07, FS98, Fer05a, Fer05b, FS09, Fer14b, Fer82c, Fer82b, Fer83, Fes56, Fet09, Feu71a, Feu71b, Feu74, Feu78, Feu82, Fey74, Fey80, Fey84, Fey97, Fin76, Fin81, Fin86a, Fin86b, Fin90b, Fin93, Fin97, Fin05, FH10, Fis79, Fis11, Fla15, Fle18, Fle33, Flo98, Flo99]. Einstein [Flo05, Flü74, Foe11, Foe23, Fok55, Fok65, Fol17, FS01, Fol96, Föl93, Föl97, Föl98, For14, For80a, For63, Fos05, Fos08, Fow03, FK04, Fox15, Fox18, FMM96, FJTS05, Fra20, Fra37, Fra47, Fra48, Fra49a, Fra49c, Fra49b, Fra49d, Fra55, Fra53, Fra65, Fra88, Fra91, Fra02, Fra15, FF97, Fra11, Fra14, Fre06, Fre15a, Fre79b, Fre79c, Fre79d, Fre81b, Fre85, Fre31, Fre18, Fre16, Fre19, FB19, FB20, Fre20, FB22, FB24b, Fre27a, FSE31, Fre11, Fri64, FTG08, FD85, Fri09, Fri79a, FP16, FH02, Fri11, dFV04, FKZ91, G.20, G.33, Gab95, Gal99, Gal94, Gal07, Gal82, Gal00, Gal03, GB03, Gal04b, Gal04a, GHS08, GHS18, Gal79, Gam42a, Gam88, GO08, Gan97, Gar39, Gar40, Gar41, Gar42, Gar77, Gar72, Gar93]. Einstein [GS80, Gas05, Gat09, GP19, Gea90, Gef02, Gef14, GdS15, Gei15, Gen23c, Gen23a, Gen23b, Gen98, GB11, Ger79, GSPH00, Get06, Gha17, Gha14, GB15, GS22, Gho17, GT80, Gia80, GRT08, Gia70, Gia71, Gib91a, Gib97, Gib91b, Gil80, Gil84a, Gil84b, Gil05, GW06, Gim12, Gim15, Gin82, Gin08, Gin15, Gin22, Gio13a, Gio13b, Gio14, Gio16, Gio20, Gio22, GS05, Giu05, GS06, Gla00, Gla07, GW93, GKP12a, GKP12b, Gli84, Gli88, GH01, GR06, Göd49a, Goe70, Goe93, GC96a, GC96b, Goe03, Goe05, Goe15, GC20, GR44, Gol67, Gol68, Gol69, Gol70a, Gol70b, Gol83a, Gol84b, Gol15, GCO^{+}06, GE97, GMW80, Gol95b, Gol97, Gol98a, Gol98b, GB06, GB08, Goo00, GO81, Goo90a, GM06, Goo75, GLG09, Gor02, Gor20]. Einstein [Gor12, Gor17, Got00, Got01, Got02, Got80, GLW15, Gra99, Gra82, Gra01, GRC18, Gre03a, Gre04a, Gre96, Gre15a, GMHB02, GRJ03, Gri79a, Gri79b, Gri81, Gri93, GG05b, GG15, Gri15, Grø82, GH07, GN11, Gru00, Gru05a, Gru57, Grü73a, Grü73b, Gru97, Gru05c, Gru05d, Grü90, Gui09, Gui18, Gun15, Gup57, GP11, Gut94, GR15, Gut15, GR17, GR20b, GR20a, Gut75, GD90, Gwy02, Gwy10, HH10, Had87, Hag08, Häg46, Hag05, Hai10, Hak07, Hal80, Hal55, Hal07, Hal15, Hal18, Ham11a, Ham11b, Ham85, Ham73, Ham86, HD14, HES^{+}12, Han79b, Han79c, HHHR01, HH16, Han22, Har72, Har97, HS10, Har89, HS06, HJ16, Har03b, Har79a, Har79b, Har79c, Has04, Has07b, Has09, Hau08, HY86, Hav88, Hav77, Hav80, Haw05, HI79, Haw07]. Einstein [Hay91, Hay02, HR13, Hea11, Hec11, Hei10, Hei83a, Hei49a, Hei53a, Hei55, Hei74c, Hei75a, Hei83b, Hei89, Hel80, Hel98, Hel15, Hel82a, Hel82b, Hen11b, Hen82a, Hen20, Hen86c, Hen87, Hen90, Hen92b, Hen94, Hen97, Hen00b, HG05a, Hen09, Hen16, Her66a, Her78, Her79a, HvMW79, Her94a, Her04, Her33, Her63a, Her63b, Her66b, Her67, Her74, Her75, Her76c, HW78, Her79b, Her06, Her11, Hes15a, Hes15b, HTS^{+}88, HW96, HBR99, HC93, HC94a, HC94d, HC94b, HC94c, HC95, HC96a, HC96b, HC96c, HC98, HC04, HEWW15, Hil03, Hin98, Hin49, Hir97, Hni98, Hoa03, Hoe94, HD75, Hof97a, Hof08a, Hof49a, HD72, HD73, Hof79b, Hof79a, Hof82a, Hof82b, Hof97b, Hof03, Hof06, Hof08c, Hof13, HV88, Hol05a, Hol30, Hol18, Hol68a, Hol68b]. Einstein [Hol69a, Hol69b, Hol70, Hol73, Hol79a, Hol79b, Hol80, Hol81b, Hol82, Hol88a, Hol88b, Hol88d, Hol89b, Hol91b, Hol92, Hol95a, Hol95b, Hol96b, Hol98a, Hol00a, Hol00b, HB01, Hol03b, Hol05c, Hol07, Hom05, HW07a, HW07b, Hom04, Hon96, Hon04, HG05b, HG06a, HG06b, Hoo70, Hoo91, Hoo77, Hor79b, Hor79a, Hot15, Hou16, How21, How84, How85, HS89b, How90a, How90b, How92, How93a, How93b, How94b, How94a, How97a, How97b, How98, How99a, HS00, How04, How05a, How05b, How14, H^{+}17, Hsu00, HF05, Hu04, Hu05, HL08b, HL12b, Hug99, Hug54, Hug93, Hul01, Hun10, Hun01, Hun15, Hus15, Igl07, IK03, Ill79, Ill89a, Ill89c, Ill89b, Ill06, Ill12, Ill19, Ill20, Imp19, Inf31, Inf47, Inf50a, Inf50b, Inf50e, Inf50c, Inf50d, Inf51b, Inf51a, Inf53a, Inf53b, Inf54, Inf55a, Inf55b, Inf55d, Inf56a]. Einstein [Inf56b, Inf56c, Inf56d, Inf57b, Inf57a, Inf58, Inf59, Inf60, Inf61b, Inf61a, Inf61c, Inf63, Inf68, Inf69, Inf73, Inf74, Inf79a, Inf79b, Inf83, Inf84, Inf92, Inf99, Inf15, Ing05, IAS^{+}98, Ire89, Iro04, Isa07, Isa09, Isa15, Isa85, Ish23, Ish77, IRW31, Isr98, IR85, Ita82, Ita99, Jac10, Jae06, Jak82, Jak08, Jam30, Jam74a, Jam82, JAvWE95b, JAvWE95a, Jam99, Jam02, Jan02, Jan05, Jan06, JNR^{+}07b, JRN^{+}07b, JR07, Jan07, Jan12, JL14b, JL14a, Jan14b, JR15, JRT^{+}91, Jay05, Jea08, Jer02, JT05, Jin96, Joe19, Jof55, Joh19, Joh05a, Joh13, JMH^{+}05, Jon44, Jon89, Jon06, JDB^{+}97, Jor69, Jos79, JM86, K.12, KaS15, Kae38, Kae13a, Kae13b, KK75, Kai94, Kai02, KT95, Kak04, Kak05, KT15, Kal92, Kan84, Kan87, Kan92a, Kan94a, Kan79, Kan96a]. Einstein [Kan89, Kan92b, Kan94b, Kan95b, Kan96c, Kan96b, Kap80, Kap81, KK89, Kap91b, Kap40, Kar05, Kas20, Kas21b, Kas21a, KRK^{+}06, Kat92, Kat04, Kau85, Kaw34, Kee05, Ken03, Ken05a, Ken12, Ken05b, Ken07, Ken14, Ken19, Ken05c, Ken13, Ken22, Ker89a, Ker89c, KK83, KF03, Ket99, Ket02, Key89b, Khr08, Kia46, Kie82, Kie94, Kie15, KWGH97, Kil70, Kil80, Kil84a, KN19, KST99, Kin95, Kin11, KH05, Kir80, KSVW10, Kle63, Kle64, Kle65a, Kle65b, Kle67, Kle70c, Kle74, Kle80, Kle82a, Kle82b, Kle05, Kle10, Kle16, Kle17a, Kle75b, Kle85, Kle93, Kle17b, Kle14b, Kli62a, Kli62b, Klo85, KS12, Kni16, Kni13, Koc01, Koe95, Koj08, Kol77, Kol07, Kol56, Kol58, Kol40, Kol72, Kon15, Kop21, Kor79, Kos98, Kos00, Kot93, KF74, Kov20, Kox93, Kox95, Kox06]. Einstein [Kox14, KS80, Kra20a, Kra20b, Kra21, Kra25, KRZ^{+}05, Kra14d, Kra14c, Kra92, Kra15b, Kre58, KWM57, KN15, Kro81, Krs91, Krs04, Krü00, Kru94, KK09, Kuh80, Kum10, Kum11, Kur79, Kuz72, Kuz79, Kuz87a, Kuz87b, Kva79, LP11, Lac11, LRL15, LRL17, Lah20, Lai20, Lak85, Lam79, Lan55, Lan59, Lan65, Lan67b, Lan67a, Lan67c, Lan67d, Lan70, Lan72, Lan74a, Lan74b, Lan74c, Lan06, Lan14, Lan02, LCL^{+}90, Lan07b, Lan08, LA88, Lar84a, Lar84b, LD08, Lat88, Lau62, Lau95, Lau98, Lau15, Lau50, Lau55a, Lau55b, Lau13, Law21, Law00, Law07, Laz15, Le 52, Led14, Lee93, Leg01, Leh14a, Leh19, LRS12, Leh14b, Len17, Len49, LLH^{+}80, Lep18, Ler73, Les74, Lev03, Lev04a, Lev15, Lev04b, LC29, Lev98, Lev01a, Lev01b, Lév80, LLSS07, Lew79, Li07]. Einstein [Lid67, Lie45, Lie08, Lie78, Lig94, Lig04, LJGS11, Lin15, Lin07, Lin08, LT15, Lin17, LS08, Lin04b, Lip00, Lis16, Lit91, Liu92, Liv13, Lod20a, Lod20b, Lod25, Loe88b, LS05, Lom05, Lon38b, Lon38a, LFKW06, Lor17, Lor20, Lor09, Lor10b, Lor10a, Lor10c, Lov86, Low01, Lud13, Luk02, LHZA20, Lut05, Lyn33, M.25, Maa07, MNT^{+}00, Mac13, Mac08, MDT^{+}03, Mac86, MSS17, MSS22, Mad89, Mad95a, Mag95, MSK15, MS17, Mal79, Mal91, ME14a, Man86, Man80, Man55, Mar15, Mar40, Mar95a, Mar01, Mar49, MH12, MW44, Mar51, MPBL09, Mar79, Mar97, Mar17, Mar06, Mar08, Mar88, Mar84, MPM^{+}08, MO13, MFHR06, Mar04a, Mar04b, Mar04c, Mar04d, Mar05, Mar07, Mar09, Mar11c, Mar11d, Mas74, Mas05a, Mas02, Mat40, Mau90, Max93a, Max93c]. Einstein [Max93b, McC11, McC84, McC15, McC70, McC57, McC68, McG19, McH07, MKBF30a, MKBF30b, McM14, McP95, McP97, Mcc78, Mcv48, Mea73, Med94, Meh73a, Meh75a, MR82d, MR82e, Meh87, Mel78, Mel79a, Mel79b, Mel63, Mel09, ME14b, MFO14, Men26, Mer10, MP74, MS82, Mer05, Mer11, MD15, Mes83, Met95, Mey92a, Mic78, Mic62, Mic63, MBD39, Mie07c, Mil10a, Mil12a, Mil13a, Mil15, Mil75, Mil81, Mil82, Mil92, Mil93, Mil98, Mil99, Mil01, Mil11b, Mil16, MA31, Mil49a, Min84, Min89, Mir02, Mir04, Mis85, Mit87, Mof08, Mof10, Mol74, Mol63, Mon19, Mon04, Moo81a, Moo81c, Moo81b, Mor96, Mor00, Mor04, Mor16, MRS17, Mor79b, MO06, Mos16a, Mos18, Mos21, MB21, Mos22, Mos70, Mos72, Mos16b, MCS09, Mou20, MNOW00, Müh11]. Einstein [Mui94, Muk05, Mul02, Mul05, MM66, Mun36, Mur94, Mur11, MG02, Mus04, Nat79, Nac24, Nag71, Nat05, NN60a, NN60b, NN75a, NN81, Nat81, Nau16, NS96, Nav91, Nav98, Nav15, Nav21, Nav94, Nav82, Nee79, Ne’06, Nef05, Nef06, Nef07, Nef11, NHP^{+}79, Ner84, Ner86, Nes60, Neu06, NMA^{+}81, Nev92, New56, New60, NJ05, New09, Nic34, NP88, NK84, Nik06, Nik12, Nir25, Nis79, NG07, Noe11, Nol05, Nor21, Nor22b, Nor22a, Nor72, Nor01, Nor49, Nor84, Nor85, Nor89, Nor92, Nor00, Nor04, Nor06, Nor07d, Nor07e, Nor14, Nor16, Nov92, NBDD^{+}12, Nug99, Nug00, Nun22, NA92, Nus14, OM14, OMNM14, OONM15, OONM17, OM18, OOM21, OM22, OV86, Obe16, Oga79, Oha08, Oha09, Oha19, OK81, OB05]. Einstein [OW54, Opp39, Opp56, Opp65, Opp66, Opp79b, Opp79c, Opp79d, Ort95, Osh06, Ost15, Ove00, Ove01, Ove05, Ove13, Owe12, Oxl03, OKSD05, P.30, P.90, Pai79, Pai80, Pai82, Pai88, Pai94, Pal99, Pal83b, Pan04, Pan05, Pao84, Pap15, Par05, PS16, Par86, Par00, Par03, Par04, Par11, Pat00, Pat01a, Pat01b, Pát79a, Pat79b, Pat85, Pat86, Pat93, Pat95, Pau49, Pau55, Pau58, Pau61b, PHvMW01a, PHvMW93, PvM93, Pau94a, Pau94b, Pau94c, PHvMW01b, Pea90, Pea49, Pea06, Pei55, Pei71, Pei83, Pen03, Pep05, Pep04a, Per85, Per05a, PS10, PS79, PBM^{+}12, Per10, Pfl99, Phe21, Phi74, Phi22, PP05, PP07, Phi97, Pia29a, Pia29b, Pic21, Pic93, Pic05, Pie23, Pie24a, Pik63, Pin00b, Pip83, Pip97, Pir05, Pis15, PS03, PS04, Pit16, PP95, PSM^{+}13, PFM17]. Einstein [Pol89, Pol83, Pol91, Poo90, Poo21, Pop98, Pop03, Pop05, Pop57, Pop66, PB72, Pot11, PL17, PB79, Pow02, Pow15, Pra06, Pri99, Pri12a, PW17, Pri10, PP03, PPC05, PP06, Pro74, Prz67, Prz11, Pul06, Put25, Pye73, Pye76, Pye77, Pye85, Pye90, Pye06, PJMS11, Pye13, Qui05, RMI21a, RMI21b, RF55, RF79, Rad08, Raf17, Rai75, Rai46, Rai14, Ram22, RL92, RR82, RA11, Rav97, Ray90, Ray95, Ray71, Räz16, RP09, Rec07, RRR07, Red84a, Red26, Ree87, Reg87, Rei11a, Rei21, Rei42, Rei57, Rei80, Rei06, Rei11b, Rei96a, Rei34, RTS^{+}65, RTS^{+}69, Rei01, RDB^{+}09, RTS^{+}xx, Rei30, Rei31, Rei07, Rei51, Ren93, RS96, Ren97a, Ren97b, RSS97, RCD98, RS99, RS03b, RS03a, Ren05c, Ren05b, RD05, Ren05d, Ren05a, Ren05e, Ren05f, Ren05g, Ren06, RS07a]. Einstein [Ren07d, Ren13, RR14, uJR15, Rey25, Ric83, Rig05b, Rig05a, Rin09, Rin11b, Rin11a, Rin03, Ris06, Rit09b, RDF^{+}08, RB81, Rob75, Rob05, Rob15a, Rob15b, Rod92, Rod05, RdO07, RdO16, RS07e, Rol10, Ron07, Roo39, Ros82a, RMR96, Ros12, Ros79c, Ros98, Ros02b, Ros05, RW07, Ros11, Ros18, RS49, RS80a, RS80b, RS13, Ros82c, Ros80, Rot79, Rot95, Rot87, RMHD11, Row01, Row06c, Row06d, RS07f, Row08, Row12, Row17, Row05, RK86, Rus25, Rus81, Rus85a, Rus91, RR17, Rya06, Ryc99, Ryc14, Ryf14, RR06, Sac81, Sac85, Sac88, SHL^{+}06, Sal05, Sal99, Sal74, San07, San77, San81b, San11, Sar96, SE79, Sau04a, Sau04b, Sau05, Sau06, Sau07b, Sau07a, Sau07c, Sau08, Sau14, Sau16, Sau19, SS21a, SS21b, Sau21, Say85, Sch74, Sch27]. Einstein [Sch00b, Sch02a, Sch05a, Sch02b, Sch10a, SS79, Sch49b, Sch49c, Sch49d, Sch49a, Sch51a, Sch51b, Sch55, Sch70, Sch79a, Sch83a, Sch84a, Sch12c, Sch71, Sch04a, Sch93a, Sch00c, Sch87b, Sch17c, Sch19, Sch95a, Sch96b, Sch22b, SNA^{+}06, Sch84b, Sch88a, Sch88b, ST94a, ST94b, ST98, ST06, Sch13b, SS15, Sch93b, SH95, Sch95b, Sch14b, Sch04b, Sch77a, Sch77b, Sch77c, Sch77d, SM79, Sch79d, Sch81b, Sch92b, SM92, Sch92c, SM99, Sch82b, SS04, Sch05f, Sch05e, Sch12f, Sch03d, Sch06, Sch08, Sch86, Sch02c, Sci61, Sci80, Sco15, Sea79, ST23, See25, See53, See54, See55, See56a, See56b, See60, See61, See64, See80, See86, See91, SU06, SWU09, Sei94, Sei03b, Sei04, Sei99b, Sel78, SZ88, Sel88, Sen21, Sev99, Sha92, Sha14]. Einstein [Sha87, Sha63, Sha73a, Sha79a, Sha73b, Sha79b, She16a, She16b, She88, Shi49, SLLS09, Sic09, Sid76, Sig92, Sil23, Sil36, SSM21, Sim80, SR11, Sim05, Sim06, Sin01a, Sin01b, SP59, Sin10a, Sin10b, Sin12, Sin00a, Sin05, Sla22, Slo22, Slo11, Sme14, SW87, Smi96, Smi03, Smi04, Smi05b, Smi07a, Smi15a, Smi15b, Smi21, Smo07a, Smo19a, Smo19b, SSS^{+}98, Sol16, Sol87, Sol56, Sol59, Som05a, Som49b, Som49a, Som49c, Som55a, Som55b, SDCZ01, Sou19, Sow01, Spa17, Spe79, Spe95, Spe62, SMHL23, Spo16, Sre29, Sri94, Sta80a, Sta80b, Sta82a, Sta82b, Sta86, Sta87, Sta89a, Sta93a, Sta93b, Sta94, Sta98a, Sta98b, Sta99a, Sta02, Sta05b, Sta15, Sta95b, Sta08a, Sta19, Sta95c, Sta06, Sta09b, SE90, Sta93c, Ste05a, Ste05b, Ste05c, Ste95a, SISK^{+}98, Ste22]. Einstein [SKM^{+}03, Ste82, Ste86, Ste99, Ste01, Ste98, Sti05, Sti13c, Sto96, Sto05, Sto13, Sto02, Str97a, Str99, Str94, Str49, Str82, Str68a, Str97b, Str17, Sug89, Sul64, Sul13, Swa30, Swe79, Swi91, Syn65, SE45, TE31b, Tal82, Tal32a, Tan92, Tan66, Tau79, Tay49, Tee02, Tel78, Tes69, Tes99, Tew19, Thi03, Thi21, Tho00, Tho83, Tho05, Tho93, Tho94, Tho15, Tho88b, Thr17, Tip10, Toa23, Tom22, Tom99, Ton97, TV00, TV07, Top07b, Tör81, TC94, Tou14, Tőz15, T– D69, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D95, T– D07, T– D12, TGC91, Tre66, Tre71, Tre75a, Tre79a, TP90, Tsi08, Tul79, Tur66, Tur67, TH79, Uni00, Uff06, Uhl80, Ung98, Ung07, Uns64, Uns81, Urb22, Ush49, Vaa17, Vai89]. Einstein [Val54a, Val54b, Val55, Val64, Van07a, Van07b, Van07c, Van74, Van96, Van11, Van79, VT99, VFIS48, VFIS63, Vea01, Vec17, Vid17, Vie29, Vig93, Vig94, Vig95, VT22, VR06, Vit15, Viz79, Voi86, VW86, Vrá57, Vuc01, Wad81a, Wad05, Wad07, Wal79, Wal88a, Wal21, WS89, Wal90a, Wal61, Wal72, Wal02a, Wal81, Wal82, Wal83, Wal88b, Wan95, Wan00, WL12, War84, War79, War92, War93, War03a, Wat97b, Wat98, Wat99, Waz09, Waz14, WW57, Web93, Wei12a, Wei04, WNS^{+}08, Wei05a, Wei06a, Wei03, Wei12e, Wei12b, Wei12f, Wei12d, Wei12c, Wei12g, Wei12s, Wei12t, Wei12u, Wei12i, Wei12m, Wei12q, Wei12n, Wei12l, Wei12r, Wei12o, Wei12p, Wei12w, Wei13c, Wei13d, Wei13e, Wei13f, Wei13g, Wei14a, Wei14b, Wei15a, Wei15b, Wei18a, Wei15c, vW79]. Einstein [WP04, Wen49, Wes80, Wes05, Wes15, Wes14, Wet80, Wey20, Whe68a, Whe68b, Whe70, Whe79a, Whe79c, Whe91, Whi96, Whi06, Whi05, Whi83a, Whi93, Whi94, WG94, WG95, Whi55a, Whi67, Whi43a, Whi01, Wid07, Wie92, Wie06, Wie29, Wig49, Wig79, Wig80, WW86, Wig96g, Wig96e, Wig96j, Wig96l, WHL^{+}05, Wil13a, Wil72, Wil86, Wil90a, Wil90b, Wil93c, Wil00, Wil17, Wil84, WH99, Wil77, Wil79, Wil11b, WE87, Wis21, WKH99b, WKH99c, WKH99a, Wit05, Wof08, Wol95, WB00, Wol03, Wol00, Wol87, Won05, Woo80a, Woo80b, WJ21, Wue05, Wyd15, Wyk76, XMY^{+}09, Yag85, Yam04, Yan80, Yan06, Yan13a, YC18, Yer03, Yer04, Yil75, Yin15, YCKG11, Yos95, You11, You79, YotY99, You04, YGKK05, You05, Yun15, Žáč12, Zac99b, Zah73a, Zah73b, Zah76, Zah77, Zah78, Zah80, Zah87]. Einstein [Zah89, Zal91, Zan93, Zau05, Záv25, Zee89, Zee90, Zee01, Zee13, Zei99, Zei10, ZDM08, ZXL^{+}19, Zho85, Zia94, Zic00, Zim84, ZPSK10, Zol02, Zuo96, Zwe80, dB65, dB85, dB87, dB36, dB49b, dB49d, dB49c, dB55a, dB55b, dBAS^{+}66, dB72, dBAS^{+}79a, dBAS^{+}79b, de 03, dGK11, dH13, dM86, de 15, dP17, dSR16a, dSR16b, dSR17, dWtH20, dSL10, vC05a, vC05b, vCK17, van06, vD02a, vD02b, vD04, vD09, vD10b, vD10a, van12, vD15, vD17, vD21, vdB95, vB01, vB05, vBT96, vB19, vB73, vF79, vG30, vK60, vL23, vL55a, vL55b, vL55c, vL56b, vL56c, vL56d, vL57, vL59, vM98, vM05b, vM09, vM12, vMHW85, Ano60, Bus13a, Fel11, For71a, For84b, Gam42b, Gea94, Gol84a, Har02b]. Einstein [Hen11a, Inf50b, Jam70, Joh11b, Kle84, Lac08, Mar11c, Nuv11, Pei83, Pol83, Pye80, Pye89, Ram81, Sch78, Set11, Tho11b, Top89, Woo06, Bal72, Bey09a, Bey09b, Bey12, Bot13, Boy88, Con83, Dys15, Est06a, Est06b, For80a, Ger15, Gly90, Goo02, Gor07a, Han80b, Hie84, Hol91a, Ivr99, Kap61, Kar05, Ken11, Kle14a, L.89, Lan00, Lan08, Mar10, Mar11b, Pec04, Pep04c, Sch07c, Set10b, Set10a, Sie07, Smo07a, Sta05a, Sta21, Ste05d, Sto17, Weg08, Yea08]. Einstein-Äquivalenz [Gio13a]. Einstein-Debatte [Hel98]. Einstein-Gruppe [Tre71]. Einstein-Killer [Kle93]. Einstein-Kondensat [JDB^{+}97]. Einstein-locality [Hel82b]. Einstein-Mädchen [Sin12]. Einstein-Marić [Chi08, Ste22, ECS19, Krs91, Mag95, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D07, T– D12, TP90]. Einstein-shift [Fri64]. Einstein-Symposion [Ano79g]. Einstein-Symposium [Tre66]. Einstein-Töter [Kle93]. Einstein-Turm [Hen92a, Vir92]. Einstein-Vorlesung [Ano20e]. Einstein. [For80a, Pye80, Pye80]. Einstein/Arnold [Win69]. Einstein/Mileva [RS92]. Einsteina [Bič79a, Gol98a, Hav77, HC95, Inf55d, Kot93, Těš12, T– D07]. Einsteinfilm [Ano22b]. Einsteinian [D’A97, Dor68, Gor22, GR23, Kop23b, Lor77, Mas99, MC80, Tam14]. einsteiniana [Tam14, Kop23b]. Einsteiniano [Mas99]. Einsteinie [Inf56d]. einsteinienne [Cun22a]. Einstein’ın [Tip12]. Einsteinium [Sch21, CSS^{+}21]. Einsteinjahres [Ren05e]. Einsteinova [Bič79b, Žáč12]. Einsteinové [T– D07]. Einsteinové-Marićové [T– D07]. Einsteinovi [ČK79, Gut75, Pát79a]. einsteinovská [Brd79]. Einsteinovu [Hav80]. Einstein” [Som05a]. Einsteins [Ano69c, Bor20, Bor22a, Ehl05b, Ein16n, vM90, Ein05-38, Gre15a, Gru97, HY86, Hei55, Her63b, IK03, Ill89c, Kie94, Mel78, MD15, Müh11, Pol89, Ren06, SE79, ST94b, Sch09a, Spe68, T– D82a, T– D83, T– D85, T– D88, T– D93, Whe68b, Wie06, vB19, vL23, vL49a, Fey11, Beh05b, Bod17, Bor64, Bor69a, Büh15, Ein21g, Fra20, Fra49d, Gru97, Gui09, Häg46, HK94, Hen90, Hof97b, Hof06, Isa85, Kal35, Kra20a, Mar95b, Pau55, Pye82a, RS96, Ren05e, uJR15, RMHD11, Ste05c, Str97b, Tre79a, Waz09, Wey20, Whe68b, Wie06, vC05a, vG30, Pye93, Pfa12, Sch09a]. Einsteinsche [Ano20a, Ano22a, Rei21]. Einstein’schen [Cas21, Ano21k, Fra37, Fre16, Fre20, Inf31, Kop21, Kra21, Lor77, Tre66, Urb22, Voi86, Cun22b]. Einsteinsturm [Egg95]. Einsteinturm [Gra99]. Einsteintuv [Nes60]. Einsteinu [Záv25]. Einsteinuv [Nac24]. Einwände [Ein18f]. Einwendungen [Kra20b]. EIR [MTVB88]. Eisenhower [HH89].

Eisenstaedt [Pye95, Pye97]. Eismassen [Bre08]. Ejnstejn [CH98]. Ejnstejna [Whe70]. Ejnstejno [T– D12]. Ekspong [Cah04]. Elaboration [Gol76, Sta98c]. Elaine [Lan15]. elämä [QP01]. Elastic [Ein05q, Ian93, Ein11a, Ein11g, Ein05i, Grø82]. elastischen [Ein11a, Ein11g, Ein05i, Ein05q]. Elbe [Boz14, Len17, Boz14, Len17]. Elect [Joh11a]. Elected [Ano82b]. Electric [Mil13, Buc94, Rob01, TN03]. Electrical [Ano23a, Ein05-39, Ein02b, Ein08a, S^{+}57]. Electricity [Ano26a, Bro23, Ein10b, Ein25d, Ein27f, Ein28c, Ein29c, Wey21, EB38, EBB41, Ano26a, EM31a, EM32a]. Electro [Ano29b, Ano29d]. Electro-Gravitational [Ano29d]. Electro-Magnetism [Ano29b]. Electrodynamic [Pra06, SE28, Dar96]. Electrodynamics [Ano54b, Ano07g, Bor07, Dar00, Hof49b, Nor04, War03a, WF49, Wil77, Ein05e, Ein12d, Ein16c, Ein20q, Ein56i, Ein94e, Fey66, FP16, HG06b, Ken05a, Nor14, Nug00, Rya06, Ryn05, Ein05-41, War03a, Har02b]. Electrodynamik [Aga07]. électrodynamique [Ein56i, Ein94e]. electrographite [BJ41]. Electromagnetic [Ano21d, Ano07d, Cle82, D’A75, EL05b, Hir69, Abi03, EL08a, EL08b, EL08c, EL09, Ein10f, Ein31f, EL05a, EL05c, EL05d, FMM96, Gio16, Hen86b, OP87, PL17, Sie91, Wei12c]. électromagnétique [Ein10f]. Electromagnetism [SF47a, SF47b, SF69, IR85, Lan67d]. Electromotive [Mil16]. Electron [Ano26a, Ano39g, BW01, Ein25e, Ein05-36, GHK^{+}06, GHK^{+}07, McC70, Mil24, Poi07, Roh60, Wei12x, And33, Cun15, DTF97, EM04, TEM^{+}89, Tsu99]. Electronic [Koh99, Sla53, Sla54, Sla63, Ben16b, Sla57, War91]. Electronic-Structure [Sla63]. electronics [Alf01, Sla50, Sla69b]. Electrons [Gol69, OSWR86, vD04, Dir65, Ein06e, Ein07a, Ein05k, Kro01, Lan21, Pau25, Arg09, Hir69]. electrostatic [BP79, Ein08a, HH10]. Electrotechnology [Pye82a]. Elegant [Gre99a, Gre99b, Gre00, B^{+}06]. elektrische [Ein02b, Ein05-39, S^{+}57]. Elektrizität [Ein10b, Ein25d, Ein27f, Ein28c, Ein29c, EM31a, EM32a]. Elektrizitätsmengen [Ein08a]. elektrodynamic [Ein20q]. Elektrodynamik [Aga06, Ein05e, Ein05-41, Ein16c]. elektrodynamischen [Ein12d]. Elektrographit [BJ41]. elektromagnetischen [EL08a, EL08b, EL08c, EL09, EL05b, EL05a, EL05c, EL05d]. Elektron [Ein25e]. Elektronen [Ein07a, Ein05k]. Elektronengruppen [Pau25]. Elektronentheorie [Lan21]. Elektrons [Ein05-36, Ein06e]. elektrostatische [Ein08a]. Elektrostatischer [HH10]. Element [Ein21c, Nod34, SMKW46, Ber07c, Ber09, Bla10, Rig02]. élémentaire [Ein56d, Ano23a, Spe67]. Elemental [Lau12, Bor22b, The12]. Elementar [Bor22a]. Elementare [Ein08b, Ein11h, Ein16f, Ein53d, Ein05t]. elementaren [Ein26d, Thi46]. elementarer [Ang13]. Elementargesetzes [Ein05d, Ein12a]. Elementarprozess [Ein21s]. Elementarquanta [Ein05-32, Ein07i]. Elementarteilchen [Ein19j, Hei93]. Elementary [Bha49, Bor22a, Bor49b, Ein26d, Ein35b, Ein46c, Ein00b, HEB^{+}80, Hei93, Lar19, Mil13, NW49, W.25, Ang13, Bor22b, Ein07i, Ein19j, Ein21s, Ein05d, Ein05-32, Les74, Thi46, Vel03, Ein08b, Ein11h, Ein16f, Ein53d, Ein05t]. Elemente [Hah42, HS44]. elementi [Fer24]. Elements [Ano39a, ABM79, BPL05, BBCPL09, Hac94, Hoy54, McM51, Qui38, SL90, Tur40b, CW32b, Fer24, Hab09, Hah42, HS44, Hoy46, Pen79b, Sea58]. elephant [Joh05b]. elettroni [Arg09]. eleven [Nacxx]. Elfenbeinernen [Jos95b]. Elie [Deb29, Deb79, For80a, Kil80, Dur12, Ano84a]. Eliot [Gar93]. Elisabeth [Bar05f, Goo85, Mat86b, Nye88, Rei86, RK94, Sei84, Wal94b, Bar05f, Hoc93, Tre88, Wei94]. Elite [Har79d, MM58, Nad79, Sei79, Zuc77, Zuc96, Zuc18, Cra01, Lar12, Lar12]. elitism [Sch02a]. elitist [Sch05a]. Elizabeth [Rec86]. Elkana [Gol84b, Kil84b, Pye80, Ano83, Bad83a, Fer83, Gil84a, Kil84a, Lar84a, Sch84b, Wil84]. Ellery [Ano05e]. Elmélette [Ein22a]. Elmo [Mur11]. eloadás [Wig65]. Elsasser [And81]. Elsbach [Ein24a]. Elsevier [Ach13, Hea64]. Eltern [Mau96]. Elves [Har07, Pic05]. Elzinga [WM08, Wid07]. Embattled [Sal19]. Embodied [Mor96]. Embracing [Lau55a, Hon04]. Emergence [All02, Cus82, Fer81, Gin82, GRJ03, Hea11, Hel84, Kai81, Mel82, Pal83a, Pal83b, Ram83, Rei82, Ren07a, Whi83b, Gio20, Gro05, Mil81, Mil98, Pla20a, Pla22a, Pla20b]. Emergency [Ein49c]. Emergent [GLG09]. Emerges [RMI21b, Smi96, RMI21a]. Emigration [Möl84]. Emil [Enk18, Ein22f, Enk18, Van07a]. Eminent [EBU^{+}52, Rup09]. Emission [Ein26d, Fox65, Ein16m, Ein21s, Ein16m, Mar04d]. Emissions [Ein26d]. emitted [Ein26b]. emittierten [Ein26b]. Emitting [LU13]. emotionality [ACT23]. Emotions [Sha14]. Emperor [Ōe95a]. Empfelhlungen [Stu87]. Empire [Mil05a, Sto12, Fra99, Mil05b, Pan02, Sta98d]. Empires [Gal03, Smo07a, Sta05a, Ste05d, Gal04a]. Empirical [Dor68, Nor72, Pfl99, Sch21, Acu14, Bra37, DiS92, Gio13a, Tee02, vD21]. Empiricism [Hol89b, How94b, How98, Max93c, Max93b, GR96, Ryc92]. empiricist [D’A97]. Empiricists [Rös95, Gio13b]. empirische [Bra37]. empiristische [Sch21]. Empiristischen [Gio13a]. employees [Mei05]. empty [Grø82]. Enchantment [RB05]. enciklopedia [Was92]. encode [Bar02b]. Encounter [Cas74, Giu05, Kov20, Bro79b, Rin03, Woo80a]. Encounters [Hei83b, Hei89, RK11, HKS10, Hei71b, Hei71a, Hei75a]. Encyclopaedia [GRC18, Was92, GRC18, Ano01a]. Encyclopedia [CKS15, CF98, Kur02, Bar05f]. End [Bey06, Cas05b, Del09b, Ham97, Has09, Kem12, MCW47, Cat98, ECEP11, FK90, Fei61, Fei66, FW07, Fra99, Gor09, Gre05b, Has07a, Hol91b, Hol96b, Hol00a, Sei03a, Wei95]. endeavors [Das00]. Endgame [Bra12]. Endless [Zac97, Zac99a, Bar05d, Bar05e, Sin01a]. Endrezze [Rai14]. endrik [BSW22]. Ends [Her82b, Kae13a]. enduring [CB99, Lac11]. enemies [Row06c]. enemy [Has05]. energia [Pye86]. Energie [Bra37, Ein05n, Ein05-34, Ein06a, Ein07d, Ein10f]. Energieinhalt [Ein05b, Ein05v, Flü39]. Energiekomponenten [Ein18k]. Energien [LMOS15]. Energies [LMOS15, PE49]. Energiesatz [Ein18e]. Energieverteilung [Pla00b, Pla01]. Energy [AGMT12, Ano39c, Ano39a, Ano39g, Ano41, Bad83b, Bai33b, Bet39, Bet68b, Bet68a, BHJ^{+}00, BO76, Bor07, BR11, Cau91, Cli09b, Com46a, Coo10, Coo15, Ein18k, Ein20i, Ein45d, Ein46c, For82a, For84a, For86, FL49, Gol82, Gow64b, Hac94, Har79a, Hei47b, Hei47a, Hen11a, HA62, HD69, HH89, HDHA89, Hoy77, Hoy79, Kir02, Kle14b, Lah20, Law07, Leg15, ME14a, Mei62, Mei63a, Mei63b, MBD39, Pla00b, Pla01, RL16, Roh60, Scl89, Smy45a, Smy48, Stu93, Tre79b, Tur46, Zoe09b, vHJK39a, vL49b, Bai33a, BN97, BBM78, BP79, Bla48, Bla49a, Bla49b, Bla10, Bra37, Cra07, Ein05b, Ein06a, Ein07d, Ein10f, Ein35b, Ein00b, Ein05n, Ein05-34, Ein05v, Far03b, Fer61, Flü39, Gat09, Gow64a, Hei46, Hol94, HM05b, Lau59, Les74, LW71, Mat10, McP74]. energy [OONM18, Oku05, Pit16, PP03, Sar08, Sch65, Smy45b, Smy45c, Smy45d, Smy45e, Smy46, Smy76a, Smy78, Smy89, Spa58, TV07, Wei12j, Wei13c, WW75, Yam09, Zoe09a, Bro15, Buc08, Ein18e, Noe11]. energy-mass [TV07]. energy-momentum [Wei13c]. Enforced [Ano07l]. Engadine [Swa02]. Engagement [Cor15b]. Engelhardt [Ham97]. Engels [Ein24g, Ein24g]. Engineer [Maa07, Ren05c, Abb05, Ren05b, Zac97, Zac99a]. Engineering [Kri01, MNT^{+}00, Pfl99, Ano07k, FFF09, SLLS09]. Engineers [LL82, Tor82]. England [Bes93, Leg15]. Englisch/ConceptDraw [GRE^{+}01]. Englischen [BS92]. English [Ano84a, Cah90b, Ebu16, For05, Har91, Hen88b, Pye89, War96b, BH87, BH89a, BH93, Bro23, BSKS04b, BSR^{+}06b, BRS^{+}09b, BIRS12b, BIR^{+}15b, CE79a, Jak78, Jam30, JSI^{+}02b, Kno97, SKJI98b]. Enigma [Alp11, Dic59, Dos17, RK11, Ber04, Dar14, dSL10, Dos17]. enigmas [Le 03]. Enlisting [Lan07b]. Enola [Gol99, Ham97, Har96c, Koh95, LE96, TW77, TM90, Tib05]. Enough [Smi97, MT98]. enquiries [GBE95]. enquiry [BEM12]. Enrico [All57, Ano47a, Ano51b, Ano82d, Ano01c, Ban03, BB04a, BB04b, Bru07, Cal75, CM02a, Coo99, CDS01, Fer54, Fer82a, Fer87, Peg39, Ras68, Seg70, Ver01, Wig96f]. ensayo [Cam88]. ensemble [Led14]. ensembles [Fin90b, GD90, Pia31]. ensino [GdS15]. Entangled [AKBAE21]. Entanglement [Acz01, Acz03, CHS97, Fra14, TWD03, EGW^{+}08, MPBL09, SDCZ01, Wei05c, Win13a, Win13b, Gil08, Gil09]. entdecken [RR17]. Entdecker [vL55c]. entdeckte [Ein22r]. Entdeckung [Jun05a, Kra88, RS96, Row06a, Row06b, Tre66, Wue05]. Entering [Hol81a]. Enterprise [Pye82a, UM92]. Entfesselung [Ost15, Sch17c]. Entgegennahme [Wil36, HSD34a]. Enthusiasm [Bro21]. entrance [Mal00]. entre [Bal85, Bal92]. entrepreneurship [JDB^{+}97]. Entropic [Kra08, Kra07b]. Entropy [PS16, Ste49, Nor06]. entstehenden [HS39c, HS44]. Entstehung [HS39a, Isa85, Jüt21, Pla20a, Pla22a, Tre66, Pla20b]. entwickelt [Ren05e, Ang20]. Entwicklung [Ein09b, Ein09c, Hei34, Hei65a, HRP90, JDB^{+}97, Jüt21, Pau58, Pet21, Pet23, Pla20a, Pla20b, Pla22a, Tre66, Müh11]. Entwirrung [Hah42]. Entwurf [EG13a, EG13b, Voi86, Wei12s, Wei12t, Wei12u]. Environment [For71b, For80b, Bra15]. environmental [Edg14, MHY95]. Environments [Sie10]. environnement [Bra15]. Envy [MJ90]. èOsztöndíj [BST03, BSST04, CST01, CST02]. Eötvös [Ein05f, FSS^{+}86a, Ill89c, NHG89, BFMNZ91, Dic60, Dic61, Ein11d, Ein05f, FSS^{+}86b, Ill89a, Ill89c, Kol59]. Eötvös-Versuch [Ill89c]. épaisseur [dB65]. epic [Hau48, Nel14, Rob12]. Epicuro [Cam88]. Epicurus [Cam88, Kuz87a]. Epidemic [Mil08, Mil13b]. Epigones [CD07]. Epilog [Des72]. Epilogue [EM32c, Ein77a, Des72, Rob14]. Episode [Elz95, Heh98, Liu92]. episodes [Bra09]. Epistemic [HS10, vD17, Kru94, vDP17]. Epistemological [Boh49, Cas21, CW74, CS79b, GRZ99, How98, Ros79c, GP11, Mey85]. Epistemologie [ER13]. épistémologiques [A^{+}09, Bor02, Del13]. Epistemology [Elz95, Mol63, Arg09, ER13, Pau61a, She88, Wie21, Ein24c]. Epistles [Ano96c]. Epoch [Ano39g]. Epopée [Hau48]. EPR [Ano05-30, BF94, Cor15c, Cos80, EPW13, Fin81, GHS21, How90b, Nik12, Sau07c, Sta16, Tho13, Wei13a, Whi04a]. Epsilon [Fra93a, Fra93b]. Epstein [Ein17h]. Equation [Acz99, Acz00, ARK^{+}99, Bod05a, Bod05b, Flo05, Kae13a, Mar40, Mil16, Pla00a, RS99, RS07a, vN37, Bod00, Dat38, Ein07i, Ein19e, ECEP11, Far03b, Flo06, Kir80, Kra06, Pen03, Pow02, Dat38]. Equation/Gaponenko [ARK^{+}99]. Equations [BB49, Bič00, BBCPL09, Bor61, Bri18, Bru09, Buc13, Cha72, Cre10, EG14, Ein15d, Ein16l, EM33b, Eis22, Far03a, FS98, Göd49a, Han22, Hen86a, Inf50c, Inf50d, Inf56e, Jam30, JR07, JR15, Kas21b, Ken13, KS65, Lau55a, Mar08, Nor84, P.30, PS16, PS03, PS04, Pul11, Sau05, Sau07b, Sch17b, Sei94, Sil23, Tau49, Wei12s, Wei12t, Wei12u, Wei13d, Wei14a, Wie29, BOS96, BS10, BPL05, vBT94, Bre89, CB09, Die88, EL08a, EL08b, Ein09a, EL09, Ein16c, Ein18d, Ein18i, Ein27a, Ein30f, EM30, EM31b, EM33a, EIH38a, EIH38b, EI40b, EP43, EK55, Ein05g, EL05a, EL05c, Far02, Grø82, Gui95, Hou16, Inf51a, Jun05a, KS80, Lan21, Lan67d, Mac13, MR82c, RA11, RS96, Rob38, RdO07, RdO16, Row06d, Sau06, Sch00c, SKM^{+}03, Wue05]. equations [Yil75, EI49a, Kre72, Lam30]. Equilibrium [Ein25a, Ein05w, Kle49, SMHL23, Bos24b, DDD^{+}06, Ein02a, Ein07i, EE23, Ein05-32, vN45]. equipartition [BN97]. Equivalence [Bai33b, BR11, Dor68, DR20, Ein16n, Ein46c, Ein05-38, Gio13a, Kle14b, ME14a, Wei12r, Acu14, Bai33a, Bra22, Ein35b, Ein00b, Nau16, Oha77, RKD64, TV07, Wei12j, Ein12f, CS60, Ein12i, Ein12h, Ein13a, Kle74, Nor85, Tre71, WTB09, Ein13a]. Equivalent [Ein05y, Ein05-30]. Era [MGG^{+}95, Ano86, NS88, Nel14, Pow02, Shi01]. Erasing [AZ05]. Erdalkalimetalle [HS39c]. Erdbewegung [Ein27c]. Erde [Eöt91, Ein27c]. Erdgeschichte [Ing05]. Erdmagnetismus [vE96]. eredità [Pot11]. Ereignis [GHK^{+}96]. Ereignisse [Wig64a]. Erfahrung [Ein21h, Ein21a, GKP12a, Stu14, Stu23]. Erforschung [Ein02b, Ein05-39, Fis10, Fis12, MGG^{+}95]. Ergänzung [Ein21t]. Ergebnisse [Kon15]. Erhaltung [Ein06a, Ein05n]. Erhellung [Kie82]. erhielt [CSW97a]. erhöht [MTVB88]. erhöht/Océ [MTVB88]. erhöht/Océ-van [MTVB88]. Eric [SK09]. Erich [Gra99]. Erik [Per09]. Erika [Van22]. erinnert [For80a, HW78]. Erinnerung [Inf69]. Erinnerungen [Bor75, Grü90, Mei05, RK86, Bor56, Bor63c, Bor65b, BI67, BI69, Ein55b, Sch87c, Wig96g]. Erinnerungen-Souvenirs [Ein55b]. Erkenntnis [Lam79, Rei20, EI38f, EI49c, N.38]. Erkenntnislehre [Wie21, Ein24c]. Erkenntnisse [Hah75]. Erkenntnistheoretische [Cas21, Häg46]. erkenntnistheoretischen [Kru94]. Erkenntnistheorie [Els24, Pau61a]. erklärbar [Ein19b, Ein19c]. Erklärung [Ein23a, GRE^{+}01, Ein15e]. Erleben [Ein21c]. Erlebnisbericht [Com56a]. Erlebnisse [Hah75]. Erling [Kau14, Gil12]. ermann [BSW22]. ermitteln [Ein18j]. Ernest [Ber03, Her03, Sch03c, Ano16b, Her02]. erneut [Kon15]. Ernst [FTG08, Her63b, Ren97b, Bla92, BIT01, CS70, Ein16a, FS06, FTG08, Her63b, Ren97b, Sta16]. Eröffnung [SM41]. Eröffnungsansprache [vL55b]. Err [All02]. Errata [Ano94b, Coc66a, Coc66b, FSS^{+}86a, KSB^{+}09]. Erratum [Adl83, CHSH70, GHK^{+}07, Leg03, Ree16a, SPA^{+}68a, ’t 02]. Erreurs [Pia31]. Errichtung [Ein21w]. Erroneous [Lid67, MNT^{+}00]. Error [Hon89, Top07a]. Errors [Est06a, RS03b, Hen13, Oha19, See25]. Erste [Kra23a, Ano21i, Enk18, GHK^{+}96]. Ersten [Ano23c, Enk18]. Erstes [JDB^{+}97]. erweitert [BS92, MTVB88]. erweitert/Geheimer [MTVB88]. Erweiterte [Ein21h]. Erwiderung [Abr12a, Abr12b, Ein12g, Ein05-28, EP06, Rit09b, WvLuAE18]. Erwin [Adl03a, Cam88, Hen92a, Mic68, Prz86, Bat85, BC94, BD92, Cam88, GGW94, Göt92, Gri12, G^{+}11, Han79a, Hei61b, Hen92a, Hen94, KDK87, McL99, MR87b, MR87a, Meh87, Moo94, Moo96b, Prz11, Rüg88, Sch35c, vK65a, vK65b]. erzählt [FS06]. erzählte [Hes15b]. Erzählungen [Len17]. Erzeugung [Ein05d, Ein05-37]. escape [Sim90]. escaped [Bro10a, Bro11, Bro16, Cou13]. escritos [W^{+}87, WdC02]. Esel [Ost15]. Események [Wig65]. Esipov [Hoh13]. Eskimo [Seg95]. ESP [Ehr77, Kol77]. Espace [Edd21, Mat86a, Ein56d, Eis02, W.25]. espace-temps [Eis02]. Especial [Ein21q, Ein29h, Ein85f]. especially [Rol47]. Esperanto [T– D12]. esperienza [Arg09]. Espionage [Sha92, SH01]. Esposizione [Ein60c]. Espouse [WS89]. esprit [Rol71]. Esq. [Gol04]. Essai [Cun22a, W.25]. Essay [Ano12a, Bar05f, Bon94, Coh80, Cun22b, Dac20b, Del11a, Del12, Gre15b, Mar11c, Pal83b, Pes13, Pol83, SE51, She15, vD15, Ein11c, HOO88, Jak78, Rit09b, Sch77a, Sch77b, Sch77c, WK12]. essay-reviews [WK12]. Essays [ARS15, ADS83, CFW76, D’A00, Ein34c, Ein49d, Ein70b, Fus88, GSW95, Her76c, HOO88, JRN^{+}07b, Key33, Key89a, Key13, KS95, Ano01d, Bar99, Bro06, Bro07, CGR^{+}97, D’A01c, Dör05, EN97, Ein09b, FPLR59, Fus90, Hei89, Hol86, Lau97, Lin60, LEM13, Mah01, Nacxx, NJP92, Pai94, Red26, Sch96a, Sch00c, Sch56d, Sch56b, Sch56c, SH99, SW12, Wad07, Wig67b, Ano28, MB66, Pau61a]. Essential [Edd95, PP07, PP06, Rin69, Rin77, Haw07, Mul10, Rig02]. Essentialism [Max93c]. essentials [Fey95, Ein56b]. est [Dem15, Der15]. established [Ein20g, Log96]. Establishing [GC20]. establishment [CLE27, War89]. estate [BH01, Wal81]. Esterson [Ste22]. Estimates [KWGH97]. Estimating [SR11]. estudio [Ben01]. estudios [B^{+}99]. estupor [Cam88]. était [Ano55c, Ano05o, Kle05]. état [EK53, Ein12c]. etc [Ein07b, Rei41]. etc. [Ein05l, Ein07b, Ein05l]. eternal [How94a]. Eternity [Hel09b]. ETH [EGO97]. Ethel [Rob01]. Ether [Ano21d, Ano25d, Ein30b, Ein34l, Ein47d, Ein07b, Gol70a, Hir76, Jan95, Mas99, Acu14, Chr22, Ein95, Kos00, Kox88, Sta93b, Ein20c, Ein22l, Cun22a, Chr22]. Ethereal [Swe13]. Ethical [Ein44b]. Ethics [Neu06, Tal82]. Ethos [Kat04]. étiage [Eis86]. ëtìseleĭ [Smy46]. Ettore [Mag09]. étude [Chr22, Edd21, Eöt00]. Eucken [Ano07p, BL07a]. Euclid [G.20, Mar15, Mlo01, TH79]. euclidea [Ein26a]. Euclidean [Bia16a, Ein26a, GH93, Gra07]. Euclidian [See25]. Eudora [For14]. Eudoxus [Lin04b]. Eugene [BH89b, HH04b, Wig96m, Wol67]. EUR [Pes13]. Europa [JDB^{+}97, Ein25g]. Europa/Quasiteilchen [JDB^{+}97]. Europäer [ENFB14]. europäischen [Kie94]. Europe [Dro20, Ano12b, Ano13a, Ein34d, Ein39b, Gil80, JDB^{+}97, LSW12, Rod19, RN16]. Europe/quasiparticles [JDB^{+}97]. European [JP89, GRS16, Hoc83, Kie94, Pat99, Vig83a, Vig83b]. Europeans [ENFB14]. européennes [Pat99]. evacuation [Gro49]. Evading [Hig14]. Evaluation [Sch74, Čap71b, CSC23, Tho00]. evaluative [Rol47]. Evan [Est06c]. Evanston [Krö53]. Even [Ano23e, Ano35, Ber06b, Lee06, Poo90, Woi06]. Event [ZDM08, GHK^{+}96]. Event-by-Event [ZDM08]. Events [Gol99, Wig64b, Wig64c, Wig67a, FW17, Wig64a, Wig65]. Eventually [Ano35]. éventuelles [Per39]. eventyr [EI39c]. Ever [Ano17c, Fra05, Pra12, Wei13e]. Every [Tra16, Ano04a, Mul10]. Everyday [Wil00, Yam04, Egd14, Fay10, Roe93, Van96, Wol00]. Everything [EMH96, McG19, RS07c, Bar91, Bry09, Der15, Foe11, Gef14, KT15, Mah03, Mus15b, Oyi04, PP07, PP03, PP06, Whe68b, PPC05]. everywhere [EG23]. everywhere-regular [EG23]. évi [Wig65]. Evidence [Bla10, Fox65, Fri39b, GRS16, Gol99, Mar05, Oga79, RMI21b, Shi75, SLLS09, Elz06, HS39a, LLH22, Mon58, RMI21a, Sar08, Tho00, HS39a, Wid07]. evil [Ban12]. Evoliutsiia [EI66a]. evolucão [EI39b, EI08a]. evolució [EI84]. evolución [EI68]. Evolutia [EI57]. Evolution [Ano26a, Ano39d, Arm38, BHJ^{+}00, Bro39, Ein31h, EI38b, EI38c, EI49c, EI50b, EI50a, Ein54e, EI56, EI59, EI61b, EI62a, EI66a, EI66b, EI68, EI78a, Ein82d, EI84, EI87, EI88, EI89a, EI92, EI95, EI98b, EI98a, EI04, EI08a, Jam30, Jam70, Kae38, Len39, Qui40, Sco15, TA17, BS83a, BS93, D’A50, EI38e, EI39b, EI42, EI47, EI48b, EI48a, EI61c, EI61d, EI63, EI65b, EI71, EI78b, EI89b, EI93, EI94, EI08b, EI11, EI13, EI15, Gun69, HBR99, Inf41a, Inf41b, Inf89, Lan70, Ray87, Bau09, EI57, EI00, LM94, EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. evoluzione [EI65b, EI13]. Evolving [Bar05g, TVL^{+}09a]. evrimi [EI94, EI11]. Ewald [Lig97]. Ewing [Bro10b]. Ewolucja [EI59, EI62a, EI98b]. Ex [Ano50c]. Ex-Associate [Ano50c]. Exact [KS80, Mac13, SKM^{+}03, Die88, Lud13]. Exactas [Ano25c]. Examination [Ner86, CJT22, Ein05-33, Ree17a]. Example [Dra14a, Göd49a, Fla15, MNOW00, SU06]. examples [Gra82]. Exceeding [HG06a]. Excellence [Bar05f, Bru03c, Dir80c, Fri01, Kra02, Tur05, Dir79e, Hün18, AA02].

Excellent [Sch12b, Sch12a]. Except [KS95]. exceptional [WKH99c, WKH99a]. Excerpt [Lor07b, Wey07, Ano16a]. Excerpts [LC07, Mie07a, Poi07, TH79]. Excessive [Bro21]. Exchange [KS65, MBD39, Sla72]. Exchange-Correlation [Sla72]. exchanges [ES68]. Excitation [FL49, MBD39]. excitations [OKSD05]. excited [Hul01]. Exciton [DHY10, KRK^{+}06, PSM^{+}13]. Exciton-polariton [DHY10]. excitons [BLI^{+}02, EM04, YCKG11]. exclusion [Mas05b, Pau47b]. exclusively [Ano01a]. Excursion [Rin11b, Rin11a, Bai92, Bai01, Rin09]. excursions [Hol18]. execution [Mee03]. exelixē [EI78a]. exerted [EL08c, EL05d]. Exhibit [For80a, Har96c, Lew79, Ram79a]. Exhibition [For05, Koh95, Lin01, Bad12, Ano12b, B^{+}79, EO79, Lar01, Nat79, Rec02, RD05, Spe79, Bad11, Ein30i, For80a, Pye80, Han80b, Pye80]. exhibitions [For80a]. Exile [Ano33c, Sch93b, Sim94, Möl84]. Exiles [Fra05]. existed [BM09]. Existence [EdH16a, EdH16b, Cha32, Ein41a, EP43, HS39c]. Exner [Han79a]. Exorcising [Bal01b]. expand [TFC93, Jos95a]. expanded [Bar05b, Cal00, EC00, MTVB88, vB02]. expanded/Secret [MTVB88]. expandierende [Bre08]. Expanding [Acz99, Acz00, GRRS99a, GRRS99b, Sch56a, Sch10c, Top07b, BS83a, BS93, Bre08, Dea13, Ell89a, Ell90, Hea15, Mon58, NB09, Nus14, Smi82]. expansion [EdS32, ES45, ES46a, Per12, Sch12e]. Expedition [Sta03b, MCS09]. Expeditionen [HK94]. Expeditions [GP19, Sim22, Pan02, EG80b]. Expelled [MP01, MP12]. Experience [Com56a, Ein21h, Ein21a, Ein21c, Ein21i, Ein34f, GKP12a, Ein34f, Arg09, Ein22l, Fis85, KWM57, Rod22, Stu14, Stu23]. experiences [Ano96a, Roe93, Hah75]. Experiment [BFMNZ91, But68, Can05, CHSH69, CHSH70, Dic60, Ein16e, EE22, Ein24d, Ein26d, EPW13, Gre04c, Hol69a, Hol69b, Ill89a, Ill89c, Ita99, Jan03, Oga79, Ram22, Wei12b, Wil01, Wil06, Wil14, Ano15j, CL19, Ein21s, FSS^{+}86b, Gib91a, Gol70b, HV88, Hol88a, Kal92, Kon15, Lor21, MS82, Nik12, Rus98, Tau86, Tau88, Teu96, Tre75a, Wil81, Wil93b, Dic61, Ein21s, Ein22g, Ein24d, Ein26d, FSS^{+}86a, Sha64a, Sha64b, vD09]. Experimental [Ano25a, Dic64, Din59, Ein15c, EdH16a, EdH16b, FTZ^{+}15, Gal82, Gin62, Hon89, Ket99, Kov01, Mil24, SEM03, BB89, Duc11, GRZ99, Hah42, JRG02, Mel79a, NA92, Ren08, Stu14, Ein15f, EdH15a, EdH15b, EdH15c, EM23]. expérimentale [Ano22e]. experimentally [Ein18j]. experimentell [Ein18j]. experimentelle [Hah42, Mel79a, EM23]. Experimenteller [Ein15c, EdH15b, Ein15f, EdH15a]. Experimenting [Sch03a]. Experiments [Ano25d, Car17, CW12, CW32a, CW32b, CS60, Gal82, Hos15, How89, Jan95, Maa13a, Maa13b, Mil99, MST^{+}07b, Nor72, PA21, Rob75, Sch22b, Swe13, Wel00, Wil11b, ZDM08, BH01, CW02, DGM04, GHW09, Hen92b, Ill12, Mer97, MST^{+}07a, PL17, Rus85a, SZ88, Sha87, Spo02, Tör81, Van07a, Van07b, Wel99, dH13, Uni96]. Experts [Ano45b, Ano45a]. Explain [Ber93, Bha86, Kae31, Pin15]. explainable [Ein19c]. Explained [Sha14, Che13, Fey95, Gri84b, HMW24, How99b, MB21]. Explains [Ano21d, Kae38, Lev07a, Ein10, Fay10, Sen21]. Explanation [Pia29a, Pia29b, SSM21, Van11, Bar91, Car89, Ein23a, Ein60c, Flo99, Teu96, Van96, Ein15e]. Explanations [Hel09c, Mes83]. Explanatorily [Lan14]. explicitly [Poo90]. Exploded [Kae31]. Exploding [DM96, Kir02, DCC^{+}01]. Exploitation [CD23]. Exploiting [HJ16]. Exploration [Pep04c, CD23, EN97, Uda94, Uda98]. Explorations [Chr22, Chr22]. explore [Wei12j]. Explorer [Nun22]. Explores [Har79a]. Exploring [BRS03, DCL23, GK05, TW00, Har13a, Hey99]. Explosion [Ano39b, Mal85, vHJK39b, Gar76, Sza84, YCKG11]. explosions [Bro82, EE82, Fin90a, Sch96c]. explosives [Lun74, TTHJ68, Wil74]. exponents [Ein18j]. exposé [Ano23a, Edd21, Ein56d, Spe67, W.25]. Exposición [Bor22b]. Exposition [Ano21d, Cun22a, Ein20k, Ein21n, Ein31k, Edd21, Ein36f, Ein54g, Ein60d, Ein77b, Ein78, Tay21, W.25]. Expounds [Kae13b, Ano19a]. express [RS09]. expressibility [Per91]. expuso [Ano25c]. exquisitely [Poo90]. Extended [Ein21h]. extending [HJ16]. Extends [Ano29b]. Extension [Gra07, Mel63, Pik63, Boy10, vM00]. Extensive [Čap71a]. extinction [Hoy77, Hoy79]. Extra [Hub29, Hum36]. Extra-Galactic [Hub29, Hum36]. extrait [Ano16a]. extraordinaire [All05]. Extraordinary [Gib19, All05, Mag09, PA06]. Extravagant [Kir02]. Extreme [Mou19]. Exzellenz [Hün18]. eye [Bre08]. eyes [Hom04, Pin00b]. eyewitness [Wel06]. eyewitnesses [Kel07, Kel09]. Eynshtina [Jof55].

F [Boy88, Car08, Dys76, Flo06, Hen92a, Hol91a, Kop64, Lin60, Pfa12, Wal72]. F. [Kra06]. fa [Seg86b]. Faber [Bon94]. Fabre [Tho88b]. Fabric [Gre04b, Gre05a]. face [Ing05, Sta98a, Sta05b]. Faces [BTR08, Bri18, Mir11, Ano97e, RdO07, RdO16]. Fachhochschulen [JDB^{+}97]. Fachhochschulen/Einstein [JDB^{+}97]. Fachleute [Rei41]. Fachwörterbuch [GRE^{+}01]. Facing [DS83b, Ano47a, DSW99]. Facsimile [Ano07b, She16a, Ein96a, Ein05-47, Spe62]. facsimiles [Bey49]. Fact [dWtH20, Hen98]. Factor [WP04, Wil13a]. Factors [Ano31b, Sau21, Gol95b]. Facts [Get06, Pye89, TTB20, Vel03, Kel06b, WS78]. Factual [Kra05]. Facultad [Ano25c]. Faculty [Ano25c]. Fadenkreuz [Gru05d]. Failed [Can05, WHF^{+}96, Cha09, Oli99]. Failings [Oha08]. fails [Ano96b]. Failure [Ber06b, Eck90, Gou47c, Jan14b, Woi06]. Failures [Ama02, Tel78]. Fair [Ano57a]. faites [Ein55i, Ein64b, Ein76b]. Faith [ACU^{+}54, Ein48h, San81a, How50]. Fall [Abr07a, Cam18, Ell10, Eng17a, Joh06, MGG^{+}95, Smo07b, Bra12, Bur45, Cha05c, Nel14, Smo06]. Fall-Out [Ell10]. fallacies [See25]. Fallacy [Ano21d]. Fallout [Fra89, Fra04, Mil08, Mil13b, Tho13, BHM91, Han97, Pre05, Red10, Ron10]. Fallout-Induced [Mil08]. Falls [McK03b]. Fälschung [Wol87]. falsification [Cea97, Cea98, Wol87]. Fame [Row12, Van07c, SR11]. familiar [Bel96a]. Familie [Man68b]. famille [Dur12]. Family [Man68b, Nor03b, Wig82, Bri05a, Fer54, Fer82a, Fer87, HW08, Mee03, Tol24, Tol76]. famosos [W^{+}87, WdC02]. Famous [Bod05a, Bod05b, Cha75, Cou79b, Cun22a, Dur02, ESA15, Mis85, Ste89, Tho88b, Ano06k, Ano18a, Bod00, FG20, HH00, HH02, HH03, HH04a, HH05, HH06, Jer02, Nor12, Sre29, W^{+}87, WdC02, Wor10]. Fancying [Dac20b]. fand [Ren05g]. fans [Ros12]. Fantasies [Klo97]. Fantastic [WD06]. Fantova [Bus03, Ein03a]. FAPP [RS07d]. Far [Ano39a, Ros18, Ein19b, Ein19c]. Faraday [Ber74b, Ber74c, Ner84]. Farben [Bor78a, Du 52]. Farber [Coh54, Duv64]. Farm [Bad96, BC95, Ber96c, Cas17, Fra93a, Fra93b]. Farming [Dac20b]. Fascinating [Aka15]. Fascist [Pol91]. Fashioned [Duf46]. Fassung [Ein21h]. Fast [Wil46, Kao10]. Faster [Mag03, Fei70, Pri12a]. Fat [CM02b]. Fatal [BR21]. Fate [Bjö01, Ste89, Cha88]. Fateful [Nor07e]. Father [Ber96d, Lau55a, Gef14, HH16, Mau96, Seg95, Kel06a]. Faulkner [Ano07q, BL07b, Hal12]. Faust [MGG^{+}95, Seg07b]. Favor [Bla88, Rus25]. favourable [Spo02]. FBI [BLM19, Gru05c, Gwy02]. Fe [Hel11]. Fear [Bla48, Bla49a, Bur45, Wea88, Wea12]. fears [Cla91]. Fearsome [May95]. feasibility [Log96]. feasible [RSU93]. feast [Ros12]. Feature [Ban51]. Features [Ano20a, DCL23, HES^{+}12]. featuring [GR15, GR17]. Februar [Ano69a]. February [Ano69a, Cli09b, Lew79, Est06c, Hei95]. Federal [Sch05c]. Federally [Cha74]. Fedor [Kuz79]. Feed [Ree14a]. Feedback [GM00]. feet [Kee05]. Fehl [Gio13a]. Fehl- [Gio13a]. Fehlinterpretation [Bac89b]. Fehlinterpretationen [Hen90, Pye93]. Feier [Ano79g, Lam79, EHM^{+}79]. Fejnman [Dai67]. Feld [Ein30h, Ein30b]. Feld-Problem [Ein30b]. Felde [EL08c, EL05d, Ein95]. Felder [HBD71]. Feldes [EG23]. Feldgleichungen [EG14, Ein15d, Ein16c, Ein16l, Ein27a, Ein30f, EM30, EM31b, EM33a, RS96, Ein19e]. Feldtheorie [Ein23c, EG23, Ein25d, Ein28c, Ein29b, Ein23j, Ein29k, Ein29j, Ein30o, Ein30f, Ein30a, EM30, Ein38a]. Felix [Tob94]. Fellow [Pin00a]. Fellows [Lin56, Ano09b]. Fellowship [BEM12, Ein53o]. female [Joh11a, Roe93]. fence [JJ81]. fences [Ano96a]. fenomenologica [Arg09]. Ferdinand [HS14, CD07, HS14, KS81, Bri82]. Fermat [Per00]. Fermi [Wig96f, All57, Ano39a, Ano47a, Ano50b, Ano51b, Ano82d, Ano01c, Ban03, Bat03, Bel88a, Bel88b, BF02, BB04a, BB04b, BRVF84, Bru07, Cal75, CM02a, CA77, Coo99, CDS01, Cro04, Faz15, Fer54, Fer82a, Fer87, GRJ03, Kap40, Lau13, Peg39, Ras68, San81b, Seg70, Ver01]. Fermilab [BDH89, HKW08]. Fern [Ein30a, EM31b]. Fern-Parallelismus [Ein30a, EM31b]. Fernparallelismus} [Gio22, Ein28e, Ein30r, Sau06]. Ferrell [Ham97]. ferromagnetic [Ein10e, SHL^{+}06]. ferromagnétiques [Ein10e]. fertilization [Nam09, Nam10]. Fertilizers [Ano05b]. Feshbach [IAS^{+}98]. Festakt [Spe79]. festen [Ein11g, Ein11h, Ein05q, Ein05t]. Festival [Ano09g]. Festschrift [Ano93a, KRW05, ACH^{+}03, RDS03]. Festvortrages [Ein21h]. Feuer [Seg91, HOO88]. fever [TT70]. few [Poo90]. Feyerabend [Hen85, Obe16, Zah78]. Feynman [Dai67, Hor00, Mil12b, Noe11, Ano54a, Ano54b, Ano05m, BR93, Coo10, Coo15, Dys65, FLS65, FLH85, FLS89, FF95, Fey00, FBD05, Fey05, FL06, FLS06, FGL^{+}06, Fri11, Gle92, Hal17, Hey99, Kai02, Kai05b, Kra11, Lei00, Sau08, SNA^{+}06, Sch94, FGL^{+}06]. fiasco [Fin90a]. Fiction [Ebu16, Bel96a, Cas03b, ET87, Fis92, Hal12, Hen98, Mar17]. fiel [BB89]. Field [Ano29b, Ano31b, Ano07c, Ban51, BB49, Bič00, Cha72, Cle82, D’A01b, Edd29a, EG14, Ein15d, Ein16l, Ein23j, Ein25d, Ein28c, Ein29k, Ein29b, Ein29d, Ein29e, Ein29f, Ein30o, Ein30h, Ein30b, Ein30a, Ein34l, Ein38a, Ein47d, Ein53b, Ein60b, Ein05x, Eis22, FS98, Göd49a, Goe04, Hal18, Hir69, JR07, JR15, Kas21a, Kia46, Kre72, Lam30, LC29, Mar08, Mas02, Mer79, Nor84, P.30, Pia29a, Pia29b, RS99, RSS^{+}07a, RS07a, Ros49, Sau06, Sau14, SS21b, Sci61, Sla72, Sta98b, Str49, Tho83, Vec17, Wei12s, Wei12t, Wei12u, Wei13d, Wei14a, Adl82, Adl83, Ber82a, Bos24b, CT24, CSD^{+}07, Coo83, Die88, DED20, EL08c, Ein10e, EH10a, Ein12e, Ein16c, Ein22b, Ein23c, EG23, Ein23i, Ein27a, Ein29j, Ein30f, Ein30c, EM30, EM31b, EM33a]. field [Ein36e, EP43, EK54, EK55, EL05d, EH05b, Ein05-44, Eöt00, Eva95, ECEP11, Gio16, Gio22, GMHB02, Grø82, Hen86b, JRG02, KP55, KS80, LFKW06, PS95, RS96, Sal74, Sau19, Sch00c, SKM^{+}03, Tel95, Wal94a, WHL^{+}05, Yil75, Zee03, Zee10, vK60, Ein19e, EG23, Has07b]. Field-Theory [Edd29a, LC29]. Fields [Ano21d, Ber74b, Ber74c, Bro10a, Bro11, Bro16, Ein12j, Ein18k, Ein25a, EdD29c, Kle49, MSS22, Nor22b, Sch49f, Wei12p, Ano19, Ein95, Ein19j, Ein31f, Ein41a, EB44, Ein44a, ES45, ES46a, HBD71, Jos80, Kro01, Pai80, Ray71, SS15, Tsu99, Wei00, Wei05b, EGR11, Hen13, Rod22]. Fierz [vBT94]. fifth [BB89, Bla10]. Fiftieth [Ano95a, BRVF84]. Fifty [Ber56, BRVF84, Kae48, McC68, Mer55, VB96, Bun88, Bun90, Ber84a, LM95, Mad95b, MO95, Pal98, PCH77]. Fig [Wil36]. Fight [Ein33b, Moo16, Urb22, Hau49, Hau53, Hau63, Hau64, Hau65, Hau83]. figure [Cas14]. figures [Cas14]. File [Ano15a, Gru05c, Jer02, RM97]. Files [BLM19, Wel99, Wel00]. Film [Dow80, Jan77, MCW47, Sch78, Tho88b, dHvdB49, Ano10c, Shi01, Waz10]. Filme [HEB^{+}80]. filmed [Bre08]. Films [Ano22d, SP59]. filosofia [Alt06a, Cas75]. filosofo [GA02, BC94]. Filtern [EM23]. filters [EM23, SEM03]. Fin [Hei82, Thr17]. Fin-De-Siècle [Hei82, Thr17]. Final [Moo16, Sch10b, Uni96, Wei14a, Ano13a, HM11, Lon84, Rus98, vD14]. final-year [Lon84]. Finally [ACT23, Col95, Sch21]. Finches [Kni12a, Kni12b, Mar11d]. Find [Ano35, Ein46e]. Finding [FG21, Lep18, Ben99, Fey00, LD08, Sto12]. findings [FSE31]. Fine [GHK^{+}06, GHK^{+}07, DL15, JN14, GD90]. fine-structure [JN14]. Finis [Ram81]. Finite [Ano21a, Cal76, TVL^{+}09b]. Finlay [Hen94, Bat85, vK65a, vK65b]. Finlay-Freundlich [Bat85, vK65a, vK65b]. Finn [Ano45b]. finnische [Isa85]. Finnish [EI62b, Hot15, Isa85]. Fire [Abb98a, Vol02]. firing [Wil72]. Firm [Kle80]. First [Alf09, Ano29a, Ano33c, Ano18c, Ano18d, Bar98b, Bus12a, Can00, Cas79, Coh40, Fab79, For01, Giu05, Her21, Hil07a, Hil07b, Hil39, Igl07, Joh13, JDB^{+}97, Ken54, Kle70c, Kri19c, Lau50, Mar22, MGG^{+}95, Mil92, Mil99, MG02, PA21, Sau07b, Sch85, Sch09b, Sta07a, Tou05, Van88, Wel06, Abi00b, Ano15a, Ano21i, Ano23c, Ano25b, Ano01d, Ben16b, BL11, Bro10b, Bru03b, Bun88, Bun90, Cas07, Cho16a, Cho16b, CW02, Dee08, Dur02, EI38e, EI48a, EI57, EI63, EI65a, EI78b, EI89b, EI93, EI15, Enk18, Far13b, FV13, Fra08, GHK^{+}96, GMW80, JN14, Jon17, Kle63, LR01, LE84, Mal91, McK03b, MT98, NK84, Pop03, Rho93, Sch77d, SR92, Smi04, Spe05, Spo16, Str11a, Str17, Tou14, Bag10, Bar93, Fri12, LSW12, Sau05, Stu18]. firsthand [Ros02a, Sha87]. Fischer [Bra12, FS06]. Fish [Dac20b]. Fisica [Alt06a, Cas75, Mas02, GdS15, CM02a, Ein26a, EI39b, EI48b, EI65a, EI65b, EI13, KT15, Pal99, Ras68, Seg76, Seg83, EI68, EI84, EI08a]. fisicos [Bor22b, Seg87, W^{+}87, WdC02]. fisikaim [Seg86a]. Fisikaren [EI00]. Físiques [Gli84]. fissile [FGMvH14]. Fission [Ano41, Ano68b, BHT86, BW39a, BW39b, FP40a, FW67, Gra64, Lau55a, Mei39, MF39c, NBDG40, The46, Seg89, Spa64, Stu94, Tur40c, Van88, vHJK39a, vHJK39c, Bad95, BP79, BH01, CHW91, CSW97b, FW85, GA71, GA81, HS39a, Hah42, HS44, Hah62, Kra88, MF39b, Ree15a, Ree17a, Ano94b, CSW97a]. fission-bomb [Ree17a]. fitting [LHF^{+}98]. FitzGerald [War91, Pop66]. Five [Ano29b, Cha05a, Gor07b, Gui95, Hal07, Van04, Wol67, EBB41, Gab95, Har06b, LW71, Sta98a, Sta05b, Tau12]. Five-Dimensional [Hal07, EBB41]. Fix [Wil13a]. Fixed [vN37]. Fixed-Point [vN37]. Fixer [MGG^{+}95]. Fixer/Wenske [MGG^{+}95]. Fixpunktsatzes [vN37]. Fizeau [CL19, Wei12b]. Fizicheskoi [Ein22j]. fizicii [EI57]. Fizigin [EI94, EI11]. fizikah [Jof62]. fizikami [Jof62]. fizike [Dai67, EI62c, Kol40]. fiziki [Ein40g, EI66a]. fizyce [Gol98a]. fizyki [EI59, EI62a, EI98b]. Flächensatz [Ein07a, Ein05k]. flash [VWA95]. Flat [EW88, EW00, PS03]. Flatland [Nov11]. flavor [Kob09]. Fleck [ER13, ER13]. Fled [MP99a]. flight [Ein16f]. Float [Ano29d]. Flores [Kra06]. Florida [Ros86]. Florida/Basel [Ros86]. Flow [GR44, LB49]. flowering [Swa92, Swa95]. Flückiger [Med76]. Fluctuating [Osh06, Yag85]. fluctuation [Abi00b]. Fluctuations [Ein86c, Ein87b, Kle82b, Eĭn85b, Eĭn86a, Ein14j, Ein19b, Ein19c, Kox14, vB19]. Fluffy [DRB05]. Fluffy-Bunnies [DRB05]. Fluges [Ein16f]. Fluid [SE28, Bre89]. Fluids [Ein05-35, Kle49]. Flüssigkeiten [Ein10a, Ein05-35, Ein05-29, Ein05c]. Flüssigkeitsgemischen [Ein10a, Ein05-29]. Flussläufe [Ein26c]. Flux [Aga75, Nac77]. fly [Cat04]. flyer [Rod05]. Fo [PLM^{+}0 ]. foam [WF98]. Fock [Gás54, Inf53b, Sla51]. Fockschen [Gás54]. focus [HL08a, HL12a, TA17]. Focuses [Ken54]. foiled [Wat97a]. Fois [Ano05o, Kle05]. Föld [von90, CSC23]. folding [Žáč12]. Folgen [Bac89b, Jun05b, YGKK05, Wol87]. Folgerungen [Ein01, Ein07h, Ein07g, Ein05u]. Follow [Fis92, Sha73b]. followed [Edd21, Sto96]. Following [Nye99, CÖT15, Chr22, Sie10]. folly [Ber02, Rho07]. Fölsing [Lig97, MGG^{+}95, Bur97, How97a, Met95, Pip97]. fondamental [Ano55c]. fondamentali [Ein55f]. fondamentaux [Ein53c]. fondamenti [Kop23b]. Fondements [Ein33c, Ano16a]. Font [Gei15]. Fontana [Nor94]. Fonti [Tug98]. Food [MNT^{+}00]. Footnote [Mac85]. Footnotes [McC84]. footprints [Fis92]. Footsteps [Bel98, Hof06, Hof13].

For.Mem.R.S. [Mcc55]. Force [Ano39g, Ano07g, Mil16, Sta07b, Ber74b, Ber74c, Die88, Kre58, Nag50]. Forced [Kae31]. Forces [Ein24l, EES49, Ein05-39, Hec16, Lau55a, Lev01b, Ree17c, Bac05, BP^{+}01, Ein02b, EL08c, Ein10e, EL05d, Pai86, Ree00, Sal05, TN03, Vig95, Ano84c]. Ford [Ree15d, Bad83b]. Foreign [Wal92c]. företagare [Lun74]. Foreword [Boh46a, Bro72, Duf46, Ein31e, Ein48c, Ein50f, Ein51c, Ein51d, Ein53e, Ein54a, Ein55c, Ein55d, Ein69a, Ein93b, Ein01a, Ein11a, Kal13, Ein23h, Kuh54, Sha54, She16a, Ein16o]. Forget [Bru09]. forgive [Büh98a]. Forgotten [Mah17, Woo06, You05, Brü00, Car08, Eis07, Fis02, Kot93, OM14, OM22, Sch12d]. Förhistoria [Hil39]. Form [Lau50, Wei14a, Büc21, EK55, Lud13, MTVB88]. Formal [Gin15, How96a, Kle14b, Byr81, Ein14g, Ein16c, Ein27a]. formale [Ein14g, Ein16c, Ein27a]. Formalisations [MSS17]. Formalism [Ano07t, Wei12w]. Formalisms [Dav14]. Forman [Pye80, FCKT11]. Formation [Cas79, Cer00, Hol81b, HS39a, WNS^{+}08]. Formative [Hol80, Gil19, GR17, HS00, Tal32b]. Formed [Lev98, The46]. Formel [Beh05a, Büh15]. former [RK86, Ski04, Wit05]. Forming [ABM79]. Forms [Dir49, Dra14b, GLW15]. formul [È03]. Formula [Cas14, ME14a, Beh05a, Büh15, È03, Cas14]. Formulae [Lau50, Edd24]. Formulated [AN79]. Formulation [Hen86a, MR82b, SW87]. formule [Ano22e]. Forord [Boh46b]. forsaken [GE97]. Forschen [RR17]. Forschens [Ein18a]. Forscher [Kon15, Ein13d, Ein22f]. Forscherleben [Ger69, GH84]. forskare [Sch50b]. Första [Hil39]. Försterling [BHJ^{+}00]. fortalt [RRK^{+}64]. Fortieth [Rot95]. Fortschritt [MTVB88]. Fortschritte [Ein30o, Hei93]. fortune [Abb02]. fortunes [Hei00]. Forty [Tom10, Gen22, Kar10, LE84, Mon58]. forty-year [Gen22]. Forward [Hea13, Mar22, Ove13]. foster [Bra14b]. fotografica [Bat03]. Foucault [ST23]. Found [Ano33h, Ano39a, Ano41, Bar09, Bro21, Eis06, JR07, JR15, Nor84, Wal88a, FSE31, Kra14b, Ren05g, Som05a, Som05b]. Foundation [Bas65, CGR^{+}97, Dic81, Don70, Ein33c, Fre27b, Goo85, Hea64, Hil39, Ihd74, Kat52, Kle80, RS07c, She16a, Vis94, Ein03, Ein21t, Ein53d, GR15, Meh75a, MR82d, MR82e, Ano33f, Hah73, Ihd74, Kri19a]. Foundational [Ree14a, HR17, Sch13b]. Foundations [Ano16a, BPP^{+}11, Bey49, BHJ62, CCG93, Cun20c, D’A01b, Ein16b, Ein16h, Ein05j, Ein05s, Fre15b, Fre16, Fre20, Fri83, GL23, Gra71, Hen86a, Hil07a, Hil07b, Hil07c, Jan73, Mie07a, Nav98, RDS03, Sau05, Sau07b, Wei05b, ACH^{+}03, Bor20, DiS92, EE59, EE90, FB19, FB20, FB22, FB24b, Fre11, GZ97, GZ06, Ker87, Kop23b, Kra25, Lau98, Löw98, MP74, Mil90, Nor93, PCH77, Pul06, Sch75, vM98, EGS77]. Founded [Bro23]. Founder [Hah79]. Founders [ADS83, Kle17b]. Founding [Ein21w, Sch03d]. Fountain [Pye97, Pye06]. Four [Ano23a, Ber88, Duf81, Ein55i, Ein64b, Hen82a, Hoc93, Kle82a, RK94, Sch28, VFIS48, Wal94b, Wal07, Wei94, Bes93, Bro06, Bro07, B^{+}99, Cal17, Cra92, Ein21l, EA22, Ein22p, EA23, Hen82b, Kis82, Le 03, Ne’06, Sch82a, Sch03b, Ein76b].

Four-Dimensional [Wal07]. Fourth [Hea13, vBRvBR77, Sha64c, SPA^{+}68a, SPA^{+}68b, SAI^{+}71, Hee01, Lie08, Thr17, GRRS99b, Sch65]. Fowler [BS83b]. fra [EI39a, EI61a, Jak13, Val06]. från [EI38d]. Fraassen [Max93a, Max93c, Max93b]. Fractals [MGG^{+}95]. Fractional [Lau99, JDB^{+}97, Sto99, XMY^{+}09]. Frage [Ein14k, Ein40c]. Fragelehrbuch [FTG08]. Fragen [Ano69c]. Fragen/Max [Ano69c]. Fragment [Van88, Brü00]. Fragments [Bed01a, Pal00, HS39a, Pal98]. frame [Grø82]. Frames [Dar19, HF05, Gou13]. Framework [Ano07c, BLR15, HJ16]. Franca [PLM^{+}0 ]. francais [Dan67, Ein05-47, Spe62]. Francaise [LLH^{+}80]. francaises [EB^{+}89]. France [Bes93, EK92, Leg15, Ano33c, Sch65]. Francesco [Ram81]. Francis [Sch78]. Francisco [RFP^{+}45]. Franciscus [Rya06]. Franck [Hon89]. Franco [Put09a]. Francoise [Mat86a]. Frank [C.50, For80a, C.50]. Franklin [Ein39a, Ein91b, Ein97, Com46a]. frankly [Byr47, Byr74]. Franz [Ein05h, Ein05h, Han79a]. Fraser [Alf09, Bus12a]. Frau [Ste89]. fraud [Van07a]. Frauen [Chi94, Föl91]. Frauenglass [Ein53f]. fray [Rol15, Rol19, Rol71]. Frayn [Dör05]. Freak [Dac20b]. Fred [Smo07a, Gre05c, Mit05]. Frederick [Fre79a, Scl89]. fredskämpar [Sch50b]. Free [Abr07a, Ano15b, Eng17a, Ham85, Har79d, Bus49, Bus68, Bus85, Ein41a, FKZ91, HFN06, Ray71, Rol71, RSU93, Wel14, Wel76, Wel88]. Freed [Ano39g]. Freedom [Ein36d, Ein40b, NN75b, EMvL79, Gro05, Wil05a]. Freeman [Bar05f, Bru03c, Kra02, Sch13a]. Freeman/Times [Bru03c]. Frees [Ano39b]. freezing [HLZJ91]. Freiheit [EMvL79]. French [Dan67, For80a, Hau48, Hen82b, Hie84, Pye80, All05, Ano15a, Ano48b, Ano55c, Ano57b, Ano80a, Ano16a, Bac00, Bal85, Bal93, Bal95a, Bal11a, Bar51a, BTM84, Ber22, Bie91, Bla93, BB55, Bra15, Bra17, CE79a, Chr22, CT12, Col57, Com95, Dan67, Dat38, De 13, Dem15, Der15, Dol18, Edd21, Ein10e, Ein10c, Ein10d, Ein10f, Ein12c, Ein13a, Ein14c, Ein14j, Ein21i, Ein21k, Ein21j, Ein23g, Ein28d, EdD29c, Ein30c, Ein33c, Ein34a, Ein34f, EI38e, EI48a, Ein50e, Ein51g, Ein52b, EK53, Ein53c, Ein55i, Ein56i, Ein58, EI63, Ein64b, Ein76b, Ein76a, EI78b, Ein89c, EI89b, EBD^{+}89, Ein90, EM^{+}91, ER91, EI93, Ein94e, Ein09a, Ein15a, EI15, Eis86, Eis02, Eis37, Elb13, Eöt00, EE82, Feu78, Fra91, Fre79d, Geo53, Gou48]. French [Gou10, Hal04b, HD75, Kle74, Kle05, Kle16, Kni16, Kol56, Kum11, LRL15, LRL17, Lan08, LdB12b, Lar12, Lev04b, LL82, Lon37, LB39, Mar51, Mar91, MP74, Mil13a, Mon12, MRS17, Nor21, Pat85, Pat99, Pau47b, Per13, Per39, Poi02, Pop07, PB06, Pye75, Rob15a, Rol15, Rol19, Rol71, Sch81b, Sch05f, Seg84, Smi15a, Spa17, Sto96, Wei15c, Whi83a, dB65, dB49d, dB49c, dB55a, dB55b]. Frenkel [ARK^{+}99, Fre96, Jos95a]. Frenzy [Rin02]. frequencies [Hal06]. frequent [dB85]. Freshman [Ber06a, Wig86]. Freud [Fla15, Ano57b, Ein85c, Fla15, Gar93, GdS15, GH01, MS17, Mul05, Pan04, Par05, Rol10]. Freundlich [Hen92a, Bat85, Hen94, Kop64, vK65a, vK65b]. Freundschaft [FF97, Jun05b, Sol59, YGKK05, Zau05]. FRG [Eck90]. Friction [Cha49b, HEB^{+}80]. fridge [Fri09, Sen21]. Frieden [NN75b, Her33, Her63a, Her67, Her75]. Friedensforschung [HH14].

Friedman [Bar05f, Bru03c, Kra02, Put09b, Tur05]. Friedmann [Ein22c, Ein23f, Jos95a, Ein22c, Ein23f, Ein86d, Tam14, TFC93]. Friedrich [ARK^{+}99, Bri82, Ein16n, Ein05-38, For80a, GRE^{+}01, HH14, Ein16n, Ein18g, Ein05-38, Goe03]. Friend [Gui18, Sta20, Ein84b, Hof55]. Friendly [Ron10, Hen92a]. Friends [Ein49c, Kle10, Med94, Kan96b, RTS^{+}65, RTS^{+}69, RTS^{+}xx, RRK^{+}64, Wei12e]. Friendship [Med94, Pop98, Pop05, You04, Zau05, BB05, Che66, FF97, Mil05a, Mil05b, YGKK05, Sol59, Jun05b]. fringes [Osh06]. Frisch [CSW97a, Lin60, Smo07a, CSW97a, DGG^{+}21]. Fritz [CSW97a, Joh06, Pes13, Ros86, Sie10, Vis95, Bal11c, Bou13, BH01, Cha05c, CSW97a, Gav95, Hen00b, Kra81, Sie10, Smi05a, Sto98, Joh95, Str05a]. Fritzsch [RS11b, Mil12b]. Frogs [Dys09]. Fröhliche [Hen11c, Rec09, Rec10]. Front [Ger07, Kae38, Ano47a, Ger92, Ger97, Ger02, Pla20b]. frontal [AH96]. Frontier [Lin09, FH11, HKW08, Zac97, Zac99a]. Frontiers [CR07, EFH89, Hei74b, Ber96a, Gri10]. Frontispiece [Ano31a, Hea64]. FRS [SR04]. frühen [Luk02, Sch05c]. Frühjahrstagungen [JDB^{+}97]. fruitcake [Hot15]. Frustrated [GR06]. frustration [GLG09]. Fuchs [The12, Goo03a, Lau12, Mos87a, Mos87b, MHM90, Wil87]. Fudge [Wil13a, Gol95b]. fuel [HCD^{+}78]. Fukushima [Mah14, Mah15]. Fulfill [Ano84c]. Full [Duf46, Hot15, MW46a, MW07]. Fuller [Ano35]. Fully [Ein05-39]. Fulvio [Hub13, Pye10, Del11b]. fun [LT15]. Function [Car24, Cas23, Cas03a, Gás54, Lan21]. function-theoretical [Lan21]. functionals [Koh99]. Functions [Her49, Koh99]. Fundamental [BBB^{+}09, BW04, Ein12a, Ein55f, Hen86a, KN08, Mel63, Pik63, Pow15, Ray92, Sil36, Wal02b, Ano55c, Bey49, Bor22b, Ein53c, Gré19c, Hag14, Kil94, Leg01, Leg03, MR82c, Oyi04, Sal05, Sch99a]. Fundamentals [Ric01, Ein40a, Ein40g, MTY79]. Fundamentes [Ein21t]. fundamentos [Bor22b]. funding [GRS16]. Fünfzig [MK56]. Funkausstellung [Ein30i]. funktionentheoretischen [Lan21]. Funnyman [Nad02]. Furchtlos [Hen21]. Fürth [Woo28]. Further [CW32a, Hin98, Pir07, SS79, HS39a, FL88]. fusion [HM05b, JDB^{+}97, MTVB88, Rei72]. fusion/First [JDB^{+}97]. fusion/Nobel [MTVB88]. Future [AZ05, BGA^{+}13, Bla46, ETP31, Mah09, Smo14, Cir07, Cir08, Fer08, GC02, GMG^{+}77, HBFH18, Haw02b, Jor69, Kao10, Lau59, Moo81c, Mul10, MO95, Van04]. Fuzzy [WL12]. fysice [Žáč11]. Fysiikan [EI62b]. Fysikens [EI38d]. fysikern [Bal95b]. Fysikken [EI39c]. Fysikkens [EI39a, EI61a]. Fyzika [EI58]. fyziky [Těš12].

G [Ach13, Adl03a, Bad72, Bor78a, Coh54, Gie05, Hen82b, JP52, JP55, Kat52, Ken25, Ber84a, Mur25, Lan67a]. G. [Ano83, Bad83a, Ded85, Fer83, Gil84a, Kil84a, Lar84a, Sch84b, Van79, Whi67, Wil84]. G.m.b.H [Ein20h]. Gabor [Joh05b, TA97]. Gabriela [Ano45b]. Gadget [Zin01]. gain [Cha09]. Gaining [Wes14]. Galactic [Hub29, Hum36, Val49, AAA^{+}18]. Galaxien [Gen05]. galaxies [Kle74, Gen05]. Galaxy [LCL^{+}90, Pau93]. Galilean [CC67, Nor04]. Galilée [MP74, Spa17, Mat86a]. Galilei [Büh98b, Col68, EI38a, EI39a, Kuh54, MGG^{+}95, Sha54, Dra53, Dra01, EI61a]. Galileo [Buc80, Car82a, Col68, Hor00, Kuh54, PB72, Büh98b, Car82a, Car82b, Cro01, DB09, Dra53, Dra01, EI38a, EI39a, EI61a, Eĭn64a, Fow03, Gam88, MP74, Num09, Pal99, PB72, Ren93, RS11b, SNA^{+}06, Sha87, Sha54, Spa17, Tho88b, VR06, Wal61, Wes14, Zan93, Sha54]. Galison [Smo07a, Lam07, Mar04c, Yer03, Yer04]. Gallery [Ber00, Pai00, BTR08]. Gallium [LU13]. Galvanometer [Ein0x, Ein0x]. Gambled [Bar15]. Game [Fin12, Kal13, Kra96b, New09, Bel05, Cha17, Fin86b, Fin97, Pou93, Ree15c, vE95]. game-changer [Ree15c]. Games [Tau63]. Gamow [Kra96b, Kra05, Rei72, Rin09, Rin11b, Rin11a, Seg11, Sta99c, Wei13e]. Gandhi [Rei41, Ein44c, Gar93]. Ganot [Leg15]. ganz [Ros05]. Ganze [Hei69a, Hei72]. Gap [CCG93]. Gaponenko [ARK^{+}99]. Gaposchkin [PGH96]. Gar [Ham97, Ove96, VB96, You95, VB96]. Garcia [Fel82b]. garde [Kie94]. garden [Kan95a]. Garwin [Ree17d, Shu17]. Gary [Ach13, Pye89]. Gas [Kem12, MBD39, Bur45, Coo83, CW02, GRJ03, PS10, Tsu99]. Gasen [Ein20l]. Gaseous [Col95]. Gases [Ano05b, BEC99, Ein24h, Ein25h, Ein25i, Por81, AK02, DGPS99, Ein20l, Leg01, Leg03, Ste07, Ein24h, Ein25h, Ein25i, Gar70]. Gastvortrag [HY86]. Gate [Smi86]. gates [De 12]. Gathering [Jab13]. Gauge [Ano84c, O’R97]. Gave [HM90, Ano25b]. Gavroglu [ARS15]. Gay [Ham97, Gol99, Har96c, Koh95, LE96, TW77, TM90, Tib05]. GDR [Hor79b]. Geb [Kai81]. Gebiete [Ein16d]. Gebildeten [Rei41]. gebrochenzahliger [JDB^{+}97]. Gebunden [BS92, Bäu83]. Geburt [Ein24k]. Geburtstag [Ano79h, B^{+}79, For80a, HBD71, Jac08, Rec02, Spe79, Tre79a, WvLuAE18, Wie06, Ein32e, KRW05, RH02, SS97, Som49b, vL49a]. Geburtstage [Ano79g, EHM^{+}79, Lam79]. Geburtstages [NHP^{+}79, SKD^{+}87]. Gedächtnisausstellung [B^{+}79, For80a]. Gedächtnisrede [Ein16g]. Gedanken [Ein14l, Vaa17, Mil99]. Gedankenexperiment [Kon15]. Gedankenwelt [Mos21, Mos16b, Mos22]. gedanklichen [Kra25]. Gedenken [JDB^{+}97]. Gedenken/WE [JDB^{+}97]. Gedenken/WE-Heraeus-Programm [JDB^{+}97]. Gedenkfeier [Ano69a]. Gedentag [Ein24k]. Gefahr [Ein21d]. gefordert [Ein05c]. geforderte [Ein07d, Ein05-35, Ein05-34]. gefunden [Som05a, Som05b]. gegebenen [Ein18i]. gegen [Ein18f, FTG08, IRW31, Kra20b, Kra20c, de 15, dP17]. Gegenbewegung [HS14]. gegenwärtige [Ein32a, Kra23a, Kra23b]. gegenwärtigen [Ein09d, Ein13g, Ein13f, Ein14o, Ein29j, Ein30n, Rit09b, RE09, Sch17, Sch19, Sch20, Sch22a, SB06b]. Gegner [Waz09]. gegründete [Ein30a]. gegründeten [EG14]. geh [Jüt21]. Gehalten [Ein20b, Ein22p, Müh11, Pla20a, Ein21h, Jüt21, Lor14, Pla20b, WvLuAE18]. gehaltenen [HSD34a, Wil36]. Geheimdiensten [Gru05d]. geheimen [HC94a, HC96a, HC04]. Geheimer [MTVB88]. Geheimnis [Bod03]. Geheimnisvolle [Rei41]. Gehrcke [Ein18c]. Gehrckes [Ein18c]. Geige [Ehl05a]. Geist [vEA02, G^{+}11]. Geistes [Ano26a]. geistigen [Pet21, Pet23]. Geleitwort [Ein24e, Ein11a]. gemacht [Flü39]. gemeinsam [Ein15c].

Gemeinsame [Ein21c]. gemeinverständlich [Bor20, Thi46, Ein17g, Ein18n, Ein20o, Ein54i]. Gemeinverstandliche [Mos21]. genährten [vB19]. Genauigkeit [Bra37]. Gender [LHDH21]. gendered [Ren08]. General [Ano23a, Ano62, Ano06d, Ano16a, ABIM15, Bia16a, Bič00, BL14, BRS^{+}12, BLR15, Car04b, Cre05, D’A01a, D’A01b, Dat15, Ein14g, Ein16b, Ein16h, Ein16j, Ein17g, Ein18n, Ein19e, Ein20m, Ein20k, Ein20o, Ein21n, Ein21p, Ein21r, Ein21q, Ein21j, Ein22a, Ein23b, Ein23h, EG27a, EG27b, Ein33c, Ein47c, EI49a, Ein54i, Ein76a, Ein82f, Ein88b, Ein05o, Ein13, Ein15b, EK92, Fer14b, Fer14c, For80a, Gei15, Ger78, GS22, Gly90, GRRS99a, Goe17, Gra18, HS89b, How92, vI11b, Inf56e, JP89, Jun05a, Leh19, LC07, LEM^{+}23, LEMW52, Lud13, Max93b, MTY79, Mil49b, Moy79, NM12, Nor93, Oha79, Per00, Pia29a, Pia29b, PK06, Pri82, RBB03, RS03b, Rob38, Row08, Ryc99, Sch17, Sch83b, Sch85, Sha64c, SPA^{+}68b, She16a, She16b, Sig92, Sta80b, Str13a]. General [SO72, Tay21, Tre21, Van11, Van79, Wal84, Web61, Web04, Wei14a, Wil14, dB49b, vL23, Ano18c, BIT03, Bič79b, BB89, Bus12a, Cho08, Cre81, Cur92, Der15, Ein04, Ein18d, EK55, Ein60d, Ein85f, Ein05-40, Fle19, Foc53b, GTP22, Gin62, Gou13, GG15, GR15, HS44, Hot15, Jan05, Jan12, JRG02, Kra21, Kra25, Leh14a, Lin15, Mat91, Meh98, Mil12b, Pen03, Rin69, Rin77, Rin01, Rin06, Rin03, Sch91, Sho21, Smy45c, Smy45d, Smy45e, Sta95b, Sud11, TW00, Unr99, Voi86, Wei12v, Wei13c, WS^{+}01, Whi67, ZCPB11a, ZCPB11b, vBT96, vN45, Ano16d, Bas65, Ber82a, Ber07b, vBT82, Cha78, Cha01, CB09, CB15, CHR90, D’A97, Dir75, Dir96, Dor78, Dra14b, EG78a, EG80a, EJN93, Edd20b, Edd21, Edd29b, Edd59, Edd66, Edd87, Ehl09]. General [Ein15j, Ein15e, Ein15g, Ein17e, Ein18e, Ein18m, Ein19i, Ein21t, Ein23k, Ein24f, Ein25e, Ein30n, Ein31o, Ein31j, Ein31k, Ein33e, Ein34j, ER35, Ein36f, ER36, EI49b, EK53, Ein54g, Ein61a, Ein61b, Ein77b, Ein78, Ein86e, Ein95d, Ein96d, Ein00c, Ein01b, Ein05-27, ELGC05, Ein06b, ELGC06, Eis86, Eis02, EK05, FK90, Gib91a, GRRS99b, Gol96, GH07, Gru57, Gwy10, Har03b, HI79, Hel80, Hen90, Hir74, Ill89b, IS49, Isa85, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, Jan14b, Kee11, Ker87, Kox88, KE05, Lan72, Laz15, MTY79, Mil99, NJP92, Nor72, Nor02b, Nor03a, Poc98, Pye73, Pye75, Rei41, RJS07, RSS^{+}07a, RSS^{+}07b, RS07c, Ren07d, uJR15, Ros37, Row06d, Ryc92, Sch05b, Sch09b, Sci80, SPA^{+}68a, SAI^{+}71, Sil22, Sta03a, Ste04, Tau79, Tho83, Van10, Wal79]. General [Wei72, Wey18, Wey23, Wey09, Wil86, Wil90a, Wil90b, Wil93c, Wil01, Wil06, Wue05, vD02b, Sci80]. General-Relativity [Van79]. Generale [Ein21p, Ano16a, Der15, EK53, Ein56d, Ein76a, Eis86, Eis02, Ein33c]. générales [Pia31]. généralisée [Edd21, Ein51g, Pye75, Ein21j]. generalizada [NA92]. Generalization [vN37, EB38, Ein45b, ES46b]. Generalized [Ban51, Cob40, Ein14e, Kur79, PS16, Dar19, DGM04, Ein48d, Ein50a, HS51, EG13a, EG13b, EG14, Ein14m, Ein50l, Ein51g]. Generally [Wet80]. Generals [Kap20]. Generating [Str19]. Generation [Bey12, Del11a, Hen09, Hür22, Mar10, PJMS11, Set10b, Set10a, Bor69b, Ein05-43, Hee01, Sta08a]. generational [Feu78]. Generations [Bro00, Seg91, Feu74, Feu82, Kra99, Feu78]. generazione [Seg91]. genèse [MP74]. Genesis [Dar06b, HK69, Her71, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, Mil75, Pla20c, RS03b, RJS07, RSS^{+}07a, RSS^{+}07b, Swe79, Wei12v, Dar21, Leh19, Nug88]. Genetic [Kea12]. Geneva [Ein24f, Yor87]. Geniale [Vaa17]. Genie [Fis02, Sch05e, Sch12f, Ste05a, Her11, Moo96a, MP99b, Rob15a, Sch05f]. Genies [Fer14a, See60, Flü74, Gre15a, Hof06]. geniet [Fre15a, Val55]. geniets [Pat01a]. genio [BF16, Bru07]. geniu [Pat01a]. Genius [Ano09e, BF16, Ber00, Bri95, Bur15, Gei15, Gle92, Goo90a, Gut94, Hes04, H^{+}17, How99b, Juc04, Kar05, Kea12, Kle75a, LS92, LS94, Lig97, Lig04, McP97, Oha08, Pai00, Ree17d, Sch13a, SNA^{+}06, See60, Ste05a, Sti13b, Sti13c, Wan00, Abr01a, Abr02, Abr03, Abr04a, Bod17, Bol04, Bru07, Diz05, Fel00, Fer14a, Fis02, Flü74, Fre15a, Gol15, Gol05, Gre15a, Her11, HD02, Hof06, Hof13, Hot15, Ill06, KST99, Lac11, Mag09, McP95, PP07, PP06, Reg87, Rob10, Rob15a, Sch05f, Sch05e, Sch12f, Sch08, Shu17, Smi07a, Tho15, Val55, Wil83, Bey09a, Bey09b]. Geniuses [CD07, Fer14c, Lui11, Seg11]. gennem [Pat01a]. genocidal [LM90]. Genocide [Bor04]. Genome [BHJ^{+}00]. genomics [Seg11]. Gentle [Col13, Her11]. Gentlemen [Ros12]. Genuine [HR13, Row06d]. Geodesics [Sil36]. Geoffrey [Hen16, Nav15]. geographical [Möl84]. geographische [Möl84]. geological [Ing05]. geomagnetism [vE96]. Geometria [Ein26a, Ein22k]. Geometric [Rey25, Kin95]. Geometrical [Jam70, LC07, Fie88, Lan70, Mas74, Sid76]. Geometrie [GKP12a, SE79, Stu14, Stu23, Whe68b, Pia31, Ein21h, Ein21a, Ein21i, Ein34f, Ein28e, Pia31, Cor04a]. géométriques [Pia31]. Geometrization [Gio22, Gio16]. geometrizes [Leh14a]. Geometry [BT12, Bia16a, Bia16b, Car04b, Dra12, Dra14b, Ein22k, Ein26a, Fok49, GKP12a, Gra07, Grü73a, Han22, Kia46, Mar08, Men49, Mlo01, Pie23, Pie24a, Rob49a, vBRvBR77, Stu14, Stu23, Wey07, WJ21, YN11, DiS92, Ein28e, Fri14, GLW15, Inf31, Pul06, RTR87, SE79, SS21a, See25, Ung07, Whe68b, YN10, Ein21h, Ein21a, Ein21i, Ein22l, Ein34f]. geomorphologist [Cro05]. Geons [WF98]. Geophysical [ST06]. geopolitical [Ree15c]. Georg [Jak78, RSPS49, Kov01]. George [B.50, Bar05f, Rei50, Ano05e, Bir82, Hum51, Kra05, Rei72, Seg11, Sta99c, Wei13e, vH62]. Georges [Ber84a, Sto96, Ano79b, Kra18, Lum11]. Georgia [ABC^{+}06]. Georgian [Lem09]. geradlinigen [vS21a, vS21b]. Gerald [For80a, Gol84b, Oha79, Top89, Kil84b, Par06, Tur06a]. Gerlach [EE22]. German [Ano22a, Ein31a, Ein33i, Eis00, Hei53a, Hen00b, Hie68, Kri08, She16a, Smi05a, Smo07a, Str05a, Abr12a, Abr12b, Abr13, ABH^{+}55, ABH^{+}87, AH93, Ama02, Ang13, Ang20, Ank15, Ano11b, Ano20e, Ano21e, Ano21i, Ano23c, Ano69a, Ano69c, Ano79h, Ano79g, Ano81b, Ano93a, Ano93b, Sch12b, Ano17b, Arr06, Ash95, APWB95, Aue95, Bac89b, Bac05, Bäc19, Bäc21, Bad04, Egg95, Bar48a, Bar50a, Bar51b, Bar52a, Bar54, Bar55, Bar56, Bar57a, Bar58, Bar62, Bar94b, Bau09, BT15, Bäu83, B^{+}79, B^{+}01, BR00, Beh05b, Beh05a, Ber69b, Ber74d, Ber96b, Bes04, BDR93, BSW22, Bod03, Bod17, Boh23, B^{+}61, Bor20, Bor22a, BEvL^{+}28, Bor55a, Bor55b, Bor56, BHJ62, Bor63a, Bor63c, Bor63b, Bor64, Bor65b, BI67, BB69, BI69, Bor69a, Bor75, Bos24a, Bos24b, BJ41, Boz14, Bra37, Bre15]. German [Bre05, Bre08, Bri05b, BH01, Büc21, BW91, Buc93, Büh98a, Büh98b, Büh15, Cas17, Cas21, Chi94, Cle17, Com56a, Cot89, CSW97a, CSW98, Cra00b, Cur38c, Cur38d, Cur38e, Des72, Die24, Die11, Dir60b, Dir60a, Dir87, Dos17, Dür62, Dür80, Egg02, Ehl05a, Ehl05b, Ein95, Ein01, Ein02a, Ein02b, Ein03, Ein04, Ein05e, Ein05d, Ein05c, Ein05a, Ein05b, Ein06e, Ein06c, Ein06d, Ein06a, Ein06b, Ein06f, Ein07b, Ein07a, Ein07i, Ein07c, Ein07e, Ein07h, Ein07g, Ein07d, Ein07f, EL08a, EL08b, Ein08b, EL08c, Ein08a, Ein09a, EL09, Ein09b, Ein09c, Ein09d, Ein0x, Ein10b, EH10a, EH10b, Ein10a, Ein11d, Ein11a, Ein11e, Ein11f, Ein11g, Ein11h, Ein11c, Ein11i, Ein11b, Ein12a, Ein12b, Ein12d, Ein12e, Ein12f, Ein12g, Ein12i, Ein12h, Ein12j, ES13]. German [EG13a, EG13b, Ein13b, Ein13g, Ein13f, Ein13c, Ein13d, EF13, Ein13e, ENFB14, Ein14a, Ein14b, Ein14e, Ein14f, Ein14d, Ein14g, Ein14i, Ein14o, EG14, Ein14h, Ein14k, Ein14l, Ein14m, Ein14n, Ein14p, Ein14q, Ein15b, Ein15j, Ein15c, Ein15f, EdH15a, Ein15e, Ein15d, Ein15g, Ein15h, EdH15b, Ein15i, Ein15a, Ein16d, Ein16e, Ein16f, Ein16n, Ein16b, Ein16h, Ein16g, Ein16i, Ein16j, Ein16k, Ein16c, Ein16l, Ein16p, Ein16m, Ein16o, Ein17b, Ein17c, Ein17d, Ein17e, Ein17f, Ein17h, Ein17i, Ein17g, Ein18c, Ein18d, Ein18f, Ein18e, Ein18g, Ein18o, Ein18h, Ein18i, Ein18j, Ein18a, Ein18k, Ein18m, Ein18n, Ein18b, Ein19b, Ein19c, Ein19e, Ein19g, Ein19h, Ein19i, Ein19j, Ein19k, Ein19d, Ein20c, Ein20b, Ein20d, Ein20q, Ein20g, Ein20h, Ein20o]. German [Ein20l, Ein20n, Ein21v, Ein21d, Ein21g, Ein21e, Ein21w, Ein21s, Ein21h, Ein21a, Ein21c, Ein21t, Ein21u, Ein22c, Ein22b, Ein22d, Ein22g, Ein22f, EE22, Ein22m, Ein22n, Ein22o, Ein22p, Ein22r, Ein22q, Ein23j, Ein23k, Ein23c, Ein23b, EG23, Ein23a, EM23, Ein23d, Ein23f, EE23, Ein24b, Ein24j, Ein24c, Ein24e, Ein24k, Ein24d, Ein24g, Ein24h, Ein24i, Ein24l, Ein25a, Ein25b, Ein25e, Ein25d, Ein25c, Ein25h, Ein25i, Ein26b, Ein26c, Ein26d, EG27a, EG27b, Ein27b, Ein27c, Ein27a, Ein27f, Ein27e, Ein28c, Ein28e, Ein29k, Ein29b, Ein29c, Ein29j, Ein30d, Ein30o, Ein30n, Ein30f, Ein30p, Ein30h, Ein30b, Ein30i, Ein30a, Ein30k, Ein30e, Ein30r, EM30, EM31a, Ein31o, Ein31a, EM31b, Ein32e, EM32a, Ein32a, EM32b, Ein33i, EM33b]. German [EM33a, EF33, EM34, Ein34h, Ein34i, Ein36b, Ein38a, EI38f, Ein39c, Ein40c, Ein40d, Ein41d, Ein46b, Ein48a, EI49c, EI50a, Ein51e, Ein52a, Ein52e, Ein53c, Ein53d, Ein53j, Ein54i, Ein55b, Ein56a, EI56, Ein56b, EI61b, ES68, EFA72, EHM^{+}79, Ein79g, EMvL79, Ein80, EB82, EI87, EI88, Ein88b, EI89a, vM90, EI92, EI95, EI98a, Ein99, EI04, Ein05c, Ein05d, Ein05-40, Ein05g, EL05b, Ein05e, Ein05h, Ein05f, Ein05l, Ein05i, Ein05m, Ein05k, Ein05j, EL05a, Ein05q, ES05a, Ein05-41, Ein05t, EL05c, Ein05-31, Ein05-37, Ein05u, Ein05-38, Ein05o, Ein05-32, EL05d, Ein05w, Ein05x, Ein05-36, Ein05-33, Ein05-35, ES05b, Ein05r, EF05, Ein05y, Ein05-27, Ein05n, Ein05p, Ein05-28, Ein05-34, EH05b, EH05c, Ein05-42, Ein05-30, Ein05s, Ein05v]. German [Ein05-43, Ein05-29, Ein05-44, Ein05-39, EP06, Ein11a, Els24, ER13, Enk18, EM88, EGO97, Enz05, Eöt91, Erb10, Erc55, Ern92, Fer14a, Fis02, FS06, Fis10, Fis12, Fla15, Flü74, Flü39, Foc53b, Foc54, Föl91, Föl93, For71b, For80b, Fra20, Fra37, Fra49b, Fra49d, Fra52, FF97, Fre85, Fre16, Fre19, Fre20, Fre27a, Fre27b, FTG08, Gás54, Gen05, Ger69, GH84, Gio13a, GKP12a, GKP12b, GHK^{+}96, Gou47c, Gou47a, Gre15a, Gri15, Gru97, Gru05d, Grü90, Gut15, HH10, Häg46, Hag05, HS39a, HS39c, HS39b, Hah42, HS44, Hah48, Hah62, Hah68, Hah75, Hah79, Hah81, Hah88, Hal55, HS14, Han18, HY86, Hau54, Hei86, Hei00, HG92, Hei34, HSD34a, Hei46, Hei49a, HW49, Hei53a, Hei55, HB63, Hei65a, Hei65b, Hei69a, HBD71, Hei72]. German [Hei75a, Hei75b, Hei77, Hei93, HR02, Hel98, Hen86c, Hen87, Hen90, HRP90, Hen92a, Hen07, Her66a, Her76b, Her77, Her78, HvMW79, Her84, Her94a, Her01b, Her04, Her33, Her63a, Her63b, Her66b, Her67, Her75, Her76c, HW78, Hes15b, HC94a, HC96a, HC04, Hof18, Hof97b, Hof03, HW04, Hof06, Hof08b, Hof08c, HW12, Hor79b, Hün18, Hun01, vI10a, vI10b, vI11a, vI11b, IK03, Ill89c, Inf31, Inf53a, Inf55c, Inf56b, Inf57b, Inf69, Inf99, Ing05, Irv67a, Irv83, Isa85, IRW31, Jac08, Jae06, JAvWE95b, JAvWE95a, Joe19, Jof67, JDB^{+}97, Jor69, Jos95b, Jun05b, Jun05a, Jun08, Kan96a, Kan02, Kie82, Kie94, Kie15, Kle75b, Kle85, Kle93, Kle17b, KRW05, Klo97, Kon15, Kop21, Kra81, Kra88, Kra20a, Kra20b, Kra21, Kra23a, Kra23b, Kra25, KRZ^{+}05, Kru94]. German [Lam79, Lan21, Lau62, LMOS15, Len17, LT15, Lin17, LRD^{+}91, LS05, LEM13, Lor14, LME23, LEM90, Lor77, Luk02, MSK15, Man68b, Mar17, Mau96, Mei63b, Mei05, Mel78, Mel79a, Mel79b, MH71, MD15, Met95, MTVB88, Möl84, Mos21, Mos22, Müh11, MM66, Nac79, NS88, NN75b, Nef05, Nef06, NHP^{+}79, Nic17, Nic19, Nod34, Nor22a, Ost15, Pau21, Pau25, Pau55, Pau58, Pau61a, Pau61b, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Pet21, Pet23, Pfe77, Pla00b, Pla00a, Pla01, Pla20a, Pla20b, Pla22a, Pol89, Pow93, Pow94b, Pow00, Prz63, RW02, Rec02, RKW05, Rec10, Rei20, Rei21, Rei24, Rei27, Rei28a, Rei28b, RK77, Rei41, RH02, RS96, Ren97b, RCD98, Ren05c, RD05, Ren05e, Ren05g, Ren06, uJR15, Rho88a, RT90, RKS08, Rit09a, Rit09b]. German [RE09, Ros05, RS13, RMHD11, RK86, RSPS49, RR17, Sal95, SE79, Sch14a, Sch12a, Sch10a, Sch49a, Sch51a, Sch55, Sch79a, Sch83a, Sch96a, Sch05c, Sch15, Sch17, Sch19, Sch20, Sch21, Sch22a, SB06b, SS97, Sch17c, Sch88b, ST94b, Sch34, Sch35a, Sch35d, SKD^{+}87, Sch87c, Sch05e, Sch12f, See52, See53, See54, See55, See56b, See60, See64, See80, See86, See91, Sic09, Sie10, Sig92, Sim06, Sin12, S^{+}57, Smi15b, vS21a, vS21b, Sol59, Som05a, Som05b, Som49b, Som49a, Som55b, Spe79, SW12, Sta95b, Sta98d, Sta95c, Sta36, SM41, Ste05a, Ste05b, Ste05c, Ste89, Ste99, Ste01, Sti05, Sto98, Str05b, Str97b, Stu14, Stu23, Tes69, Thi46, T– D82a, T– D83, T– D85, T– D88, T– D93, Tre71, Tre79a, TF99, Urb22]. German [Vaa17, Van07c, Vir92, VR06, Vit15, Voi86, Wal89, Wal92b, Wal95, Wal03, Wal05a, Wal17, Waz09, Wey18, Wey23, Wey20, Whe68b, Wie06, Wie21, Wig64a, Wig96g, WW75, WE87, Wol87, Wue05, Wyd15, Yos95, YGKK05, Zau05, Zim84, de 15, dP17, dWtH20, vEA02, vBL96b, vB05, vB12, vB14, vB18, vB20, vB19, vE96, vEPF09, vE10, vEPF22, vG30, vL09, vL11, vL13, vL19, vL21, vL23, vL52, vL55b, vL55c, vL56a, vL65, vM05b, vM09, vM12, vMHW85, vN37, vS14, vS06, vW02, vWR04]. German/French [Ein53c]. Germans [The12, Lau12]. Germany [CR07, Ein48g, GRRS99b, JP08, Lin60, Som05b, Abb98a, Ano03c, Bla02, Bow87a, Bro92, Ein22e, For73, Goe93, Goe17, Gou48, Gra99, Gru97, Hei46, Hei47b, Hei47a, Hen00a, Irv67a, Irv67b, Irv83, JM17, Kra88, Log96, MP99a, NS88, Pye82b, Ric06a, Row06c, Set03, Sim90, Som05a, Sta98c, Ste82, Ste86, WW75, WHF^{+}96]. gern [KRZ^{+}05]. Gerontological [Hir97]. Gershenfeld [ARK^{+}99]. Gertrude [Fol17]. Gerüchte [Bre05]. gesagt [Enz05]. Gesammelte [BDR93]. Gesamtsitzung [Tre66]. Geschichte [AH93, Sch12b, Aue95, BT15, Bre08, Büh15, Inf55c, Kie94, LMOS15, MGG^{+}95, Rho88a, RT90, Sal95, Sch12a, Sch09a, Sta95c, Tes69, Thi46, Wie06, Zau05, vBL96a, vBL96b]. Geschichten [Vit15]. gesehen [KRZ^{+}05]. Gesellschaft [B^{+}01, Ein20h, Ein31a, Hei75b, Sch88a, Sch88b, WvLuAE18, vBL96a, vBL96b, Hof08b]. Gesetz [Bos24a, Ein11d, EMvL79, Ein05f, Pla01, vEPF09]. Gesetze [vEPF22, Str97b]. Gesetzes [Pla00b, Ein26c]. gesetzt [EM31b]. Gesicht [Ing05]. Gesichtspunkt [Ein05-37, Ein05d]. Gespräch [Bre15, Hei65b]. Gesprache [Hei75a, BW91, Buc93, Hei69a, Hei72, Fan03]. Gestalt [Ash95, Mil75]. Gestaltung [Ein27e]. gestation [vE95]. Gestirne [Gut15]. Gests [Lan00]. Get [Ano30b, Ano33c, Ano13c, Sch21, Wil17, Wil69, DC12, Gio16, Jar08a, Jar08b, Rod05]. Gets [Ano33c, Ano34b, Ano45b, Wal83, Ano08c, Dan97b, Gun15, Lee93, Rod04]. Getting [Hul01]. Gewissen [EM88]. gewöhnliche [EM34]. gezogene [Ein07g]. gezogenen [Ein07h]. Ghetto [Ein44d]. ghost [DB86]. GI [SK82a, SK82b]. Gianfranco [Seg91]. Giant [Has04, Ano17a, Grü08]. Giants [MD67, Ren06, Haw02c]. Gibbs [Nav98]. Gibson [Ano33f]. Gibt [Ein12d]. Gift [Lau55b, MP99a, MP01, MP12, Thi03, Wig82]. Gifted [Ste89]. Gilgamesh [Col93, Zal91]. Gimbel [Dys15, Row17]. ging [Bre08]. gioca [Pot11]. gioia [Bal94]. Giora [You98]. Giordano [Rei41, Rei41]. Giorgio [Sha54]. Girl [Wil13b, Sin10b, Sin12]. Girls [Cla97, Moo16, Moo17, Kie13, Swa02]. Gitlin [Ber92a]. Give [EBU^{+}52, Sei04]. given [Ano13a, Ein18i, Ein55i, Ein64b, Pap15, Pau47b, Ein12d]. gives [Car07b]. Glamour [Esh09]. Glanz [Gre15a]. Glänzendsten [Gut15]. Glas [Ein14f]. Glasgow [Ano33f]. Glass [WCL^{+}00, Ein14f, HLZJ91, Ros98, Bot13]. Glasses [Ein52e]. Gleichgewicht [Ein07i, Ein05-32]. Gleichung [EM33b]. Gleichungen [EM33a]. Gleichungssystem [vN37]. Glick [Boy88, Hol91a]. Glimpse [WM08, Wid07, Elz06, Ren97a]. Glimpses [DH79, For80a, Moy81, DH81, DHR13, Roe93, Wig79]. Global [Hec12, MHY95, Ree15c, Ren21, Ash02, S^{+}10, tH20]. Globular [Ein21e]. gloria [FG20]. glories [Wit05]. Glow [Mul99, Nor21]. Gluon [Raf17]. Go [Fra05, Ano03a, Fei70, Hoa03]. Goal [Ein39e, Jam30]. Gockel [Lau62]. god [DC08, Acz99, Acz00, Bar18c, Bar18b, BC01, BW91, Buc93, Cla78, Coh31, Com14, Ein30m, Fus88, Fus90, Gal79, Ghi05, GE97, Gor12, Gor17, Gru05a, Hai10, McG19, Mon58, MB11, Mos68, Pin00b, Pot11, Pow02, Rei41, Sta92a, Tip10, Yin15, vE95, Mac08]. Gödel [Woo06, You98, BPP^{+}11, BBL08, DD11, FW07, Fer05b, Hol18, Hon96, Hon04, Jun05b, Per85, Rin09, Rin11b, Rin11a, Wan95, You91, YotY99, You04, YGKK05, You05]. Goes [Ano31a, Ano33c, Dac20a, Sei03b, Ano95c, Mof08, Num09, Vol02]. Goethe [Rei41, Ein17d]. Goettingen [Pye73, Som05a]. Goetzberger [MGG^{+}95]. Gofman [Sem08]. Going [Pla32, Pla33, Pla77, Sta09a, Pla59, Pla81]. Golden [Ehl09, Lum11, Wey09, Rob58, ES68, Sei99a]. goldenen [ES68, For70b, Win69]. Goldschmidt [CD07]. Goldstone [Hig14]. Gomez [Cob07, Cob07]. Gone [Räz16, Bie91]. Good [Can16a, Kox93, SG82, Sch21, BR93]. Goodman [Wac06]. Google [Har13c]. Gordin [Sta21, Wis21]. Gordon [Bus12a, Alf09]. Gösta [Cah04, Stu10]. Got [Del11a, Kra15b, Reg87]. Gott [BW91, Buc93, Nah01, Rei41]. Göttingen [Row06a, Row06b, BR00, How93a, Sch88a, Sch88b, Som05b]. Göttinger [GRE^{+}01, Som05a, Som05b]. Goudsmit [Hoe18, LRD^{+}91, vC16, vCH18]. governance [AE06]. Government [Cau91, VFIS63, BBD^{+}11, Sno61, Smy45b, Smy45c, Smy45d, Smy45e, Smy76a, Smy78, Smy89]. governmental [Tur03]. gozo [Bal99]. GPS [Coc16]. gradual [Rob10]. graduations [HR02, RW02]. Graetzer [Bad72]. Graham [Gar93]. Grams [Ano18d]. Grand [HM10, Oyi04, Kum11, She75]. grandchildren [Jac10]. grandes [Edd21]. grandeurs [Ein14j]. Grandfather [Inf15, Goo03a]. Grant [Ano33c, Wal83]. granted [Cho10a]. Graph [WD96]. Graphene [Nov11, Gei11]. Graphic [MSK15, FV13]. graphs [Fie16]. Graves [Jan73]. gravest [DG03]. Graveyard [Sne82]. gravifique [De 24]. gravitacii [Ano79a]. gravitacní [Záv25]. Gravität [vEPF09, vEPF22]. gravitating [Ein39d]. Gravitation [Abr12a, Abr12b, Abr07b, Abr07d, Abr07c, Ame20, Ano21f, Ano23a, Ano29b, Ano07a, Ano07d, Ano07c, Ano07h, Ano07l, Ano07t, Ano07f, ABIM15, Bič89a, Bič89b, BGLR17, Bro23, CW95, Cun19a, Cun19b, Cun19c, Cun20a, Cun20c, Cun22a, Dic59, Dic81, Dir79d, EJN93, Edd19, Ein12g, Ein13b, Ein14e, EG14, Ein15d, Ein16l, Ein18k, Ein19a, Ein25d, Ein27f, Ein28c, Ein29c, EM31a, EM32a, Ein05-28, Ein07c, Eis93, Eöt91, Eöt08, Göd49a, Goe70, Gup57, Hak99, Kas20, Kur79, LPPT75, Lor07a, Meh73a, MTY79, Mie07c, Mil49b, MTW73, Nor07c, Nor07b, Nor72, Nor07d, O’L64, Oha76, OR94, Phi22, Pir07, RS96, Ren07a, RS07b, Ren07c, RTR87, Sch83b, Sch22b, SM07, Tay21, Wal07, Wei72, Wei12p, Wey21, Wit62, WJ21, Zen07, ZdSS91, dS19, vE96, Ein05e, Ein05-38]. gravitation [Ano22a, Ano79a, Bäc19, Bäc21, Ber68b, Ber69a, Ber87a, Ber92b, Bir82, BB89, Che10, Dir79e, Dir79f, Duc11, DE04, Edd20b, Edd21, Edd29b, Edd59, Edd66, Edd87, Ein11i, Ein12b, Ein13f, EF13, Ein14c, Ein18d, Ein18i, Ein27a, EBB41, Ein45b, ES45, ES46a, ES46b, Ein48d, Ein50a, Ein51g, EK53, Fle19, Fok59, Fok64a, Fok64b, FB19, FB20, FB22, FB24b, Fre11, FH02, Goe17, HI87, Hel80, HS51, Hoe94, JRG02, Mad95a, Mas74, Nor92, Ray71, Row06d, Rus91, Sil22, Str94, Vel00, Ano22a, Bro20a, Cun22a, Dir80c, Edd20a, Edd06, EG13a, EG13b, Ein13e, Ein14c, Ein14i, Ein14l, Ein14m, Ein14q, Ein15d, Ein16l, Ein20g, Ein21e, Ein27a, Ein50l, EK53, Fre16, Fre20, Lor17, Mal91, RSS^{+}07a, RSS^{+}07b, Rit09a, Sch19, Sch20]. Gravitation [SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, dSR16a, dSR16b, dSR17, Ein16n, Pep04b]. Gravitation* [Pir05]. Gravitational [Ano29d, Ano93c, Ano06b, Ano07g, Cho16a, Dic60, Dic81, Die88, Dys16, Ein18o, Ein31f, Ein05x, Eng17a, Fra11, GS22, JSSZ87, Ken13, Kia46, Kle49, Mie07b, Mos13a, Mos13b, Nor07a, OS39, RSS97, RS99, RS03a, RS07a, RB81, Ros49, Sau21, Tur88, Wal21, WW57, Web61, Web79, Web04, Wei12p, WTF81, Wil93a, Ao04, AAA^{+}18, AHH05, Bar18a, Blu22, Buc20, CF11, CC84, Cha80, Cho16b, Cho16c, Col04, Col17, DGM04, DED20, EG80a, Ein12d, Ein12e, Ein19j, Ein22b, Ein36e, ER37, EIH38a, EIH38b, EI40b, Ein41a, Ein05-44, Jak78, Jun05a, Ken07, Ken14, LHF^{+}98, Nac73, Nac77, Poc98, RS96, Rob38, RKD64, Spo16, Tho18, TLW^{+}10, Wei18b, Wil81, Wil93b, Wue05, Záv25, vK60, Dir60b, Dir60a, Dir87, Tre71]. gravitational-wave [Blu22]. Gravitation“ [Ein12b, Ein16n]. Gravitationsfeld [Ein22b]. Gravitationsfelde [Ein18k]. Gravitationsfelder [Ein19j]. Gravitationsfeldes [Ein05x, Ein05-44, Ein12e, Ein12j]. Gravitationsgesetzes [Ein21e]. Gravitationsgleichungen [Jun05a, Row06a, Row06b, Tre66, Wue05, Ein18d, Ein18i]. Gravitationsproblems [Ein13f]. Gravitationstheorie [Ano22a, Bäc19, Bäc21, EF13, Ein20g, EF05, Tre66, Abr13, Ein13b, Ein13e, Ein14i, EG14, Ein14l, Ein14m, Fre16, Fre20, Sch19, Sch20, Sch22a, Tre71]. Gravitationswellen [Dir60b, Dir60a, Dir87, Ein18o]. Gravitationswirkung [Ein12d]. Gravity [Abr13, BR21, Car07b, Cho03, Cho08, Col04, Col17, Dar06a, DeW79, Ein12j, Ein05e, Ein05-31, EF05, Eis06, FS09, GS22, Har03b, HF05, JRN^{+}07a, Mir11, Mos18, Par79, Pep04a, Pet10, PW14, Sei04, Sta07b, Sul64, Tho83, Vec17, Wat99, Zee13, vL23, AAA^{+}09, Adl82, Adl83, AC09, AS91, BP95, vBT94, CF11, Ein06a, Ein05-28, Ein09c, GG05a, GH93, Jan12, Leh14a, Lin15, Mof08, Nor72, Nug88, Pap15, Pen03, PKLT17, Wal77a, Wal92a, Whe90, Zee89, Zee90, Zee01, de 15, dP17, vE96, vEPF09, vEPF22, Ano11a, Gib91a, Ein05-38]. Gray [Hea13]. Grazia [Bal75b]. Great [Ach13, AFG90, Ano55d, Ber00, CC96, Cre10, Cro01, Far03a, Fri79a, Goo90b, Ham97, HT10, Sau16, Sha87, Sha54, Sin61, Sin70a, Smi82, Sud11, Tib13, Amr59, Bak68, Bak76b, Bak76a, Bre97, Büh98b, Edd21, Far02, Gab95, Gam88, Hal15, Haw02c, How50, Kae48, Kum10, Kum11, KW87, Lau98, Lig05, Lin04a, LY95, Liv13, MW44, Mar51, Moo02, NS88, Pic08, Sch12d, Sub03, Tay49, Wei95, Bes04, vW02, vWR04]. Greatest [Acz01, Ano21d, Ano29b, Ano39a, EMH96, Hec16, Hür22, Ree17c, Sto17, Bod16, BB13, Gol95b, Gol97, Gol98a, Gol98b, Ler73, Mar17, McC01a, Ray90, Sam10, Wal02a, Wes04]. Greatness [Ein51e, Ein80, Rig05a]. Greece [Pap15]. Greek [EI78a, Fur87, Vol02]. Green [Bor71, ABW99, MTVB88]. Greenglass [Rob01]. Greening [Jun14]. Greenspan [Bey06, Cas05b]. Gregg [Ber03, Sch03c, vB02]. Grenzen [HK94]. Grenzgänger [Sch05c]. GRG [Mer79]. GRG-Organization [Mer79]. Grieder [You95]. Griffin [Dow80]. Griminger [Pye89]. Grinten [MTVB88]. Grinten-Preis [MTVB88]. Gröbner [Ano93b]. Gromyko [Bro79b]. Großbritannien [JDB^{+}97]. Grosse [ARK^{+}99, Ein51e, Fra37, Ein80, NS88, BS92, Fre27a, Bes04, Erm89, vW02, vWR04]. Großforschung [MGG^{+}95]. Grossmann [BB89, JRG02, Pit16]. größte [Mar17]. Ground [Ano94a, BMD^{+}11, FG21, How97b, Tur66, Ano09c, Gal93, Inf31, KB60, KB83, KZS99, NBDD^{+}12, SISK^{+}98, Ken54]. ground-state [SISK^{+}98]. groundbreaking [Cor15a]. Grounds [Ein12f, Cea95, Cea97, Cea98, Kir05]. Group [KK05, Tre71, Cat04, RRK^{+}64, Spo16, ZZY19]. Groups [MSS22, FW17, Pau25]. Groves [Erm89, Gol95a, Gol96, Nor02b, Nor03a]. Growing [Ros82b]. Growth [Bro39, EI38b, EI38c, Hun07, Kae38, EI42, EI61d, EI71, GSPH00, Goe17, San15, Dow80]. GRT [Mil92]. Grund [Thi46]. Grundbegriffe [Ein53c]. Grundgedanke [Sch34, Sch35a]. Grundgedanken [Ein15g, Ein23d]. Grundgleichungen [EL08a, EL08b, Ein09a, EL09, Ein05g, EL05b, EL05a, EL05c]. Grundkurs [MGG^{+}95]. Grundlage [Ein14g, Ein05o, GRE^{+}01, Ein16h]. Grundlagen [Cun22b, Ein03, Ein11f, Ein13e, Ein14l, Ein16b, Ein53d, Ein05j, Ein05s, Fre16, Fre20, JDB^{+}97, Kra25, Bor20, Bor22a, Cor04a]. Grundläggning [Hil39]. grundlegenden [vM90]. Grundzüge [Ano21k, Ein56b, Kop21]. Grünes [MTVB88]. Gruppe [Tre71]. Grützmacher [Ano61]. guan [HL10]. Guaranteed [Ein18b]. guard [Ano33g]. guarded [Ano79e]. Guarding [Ano33b]. guerre [ER91]. guest [HY86]. Guglielmo [GMOP12]. Guichelaar [Sta20]. guidance [Mac93]. Guide [Bai07b, Dro20, Gra18, McG19, Rod19, Bag92, Bar05d, Bar05e, Bre79, Cal03b, Cal05b, Doh06, Hei05, Hel11, Kel12, Ken12, MHY95, Mor00, Mor04, Pen05, Ree16a, Ree16b, Van02, Wol93]. guided [Gun11, Hai10, HG07, HH96, Swa92, Swa95]. Gullstrand [MH12]. Gültigkeitsgrenze [Ein07i, Ein05-32]. Gunn [Ahe03]. Gunnar [Isa85, Isa85]. gunpowder [Lor08]. Guppy [Ano96d]. Gustaf [Goo85, Sei84]. Gustav [Bor07]. Gutachten [HR02, RW02]. Gutachten- [HR02, RW02]. Gutfreund [She16a]. Gütigkeitsgrenze [Ein22g]. guys [Jun05a, Row06d, Wue05]. Gymnasium [RK86]. Gyrations [Ano06h, Cho06]. gyro [LS05]. Gyromagnetic [Gal82]. Gyros [Ill19, Ill20].

H [Adl03a, Ano60, Bar05f, BSW22, Bru03c, Coh54, Duf46, Duf81, Ham97, Hen82a, Hor79b, Kat52, Klo63, Kra02, Mic68, Ree15d, Ano00c, AP01, Ber83, BHM91, Bun82, DMPS81, Eli50, For15, Gon96a, Gon96d, HM90, HM97, Jon01, Ken54, Kno90, Mor79a, Mos68, Sch61, Wol09, Jon01]. H-Bomb [Ano00c, Gon96d, HM90, Ken54, Mor79a, Wol09, AP01, Ber83, BHM91, Bun82, DMPS81, Gon96a, HM97, Jon01, Kno90, Mos68, Sch61, Jon01]. H.

[Bad72, Bel26, Ber79b, BW^{+}87, Dre87, Ehl09, Ein13c, Ein16i, Ein17d, Ein18h, Ein28a, Fer82b, MB^{+}92, Ner86, Prz11, Tho97, WB^{+}87]. hû [Ayn34]. ha-ôlam [Ayn34]. Haalvelmo [Zel90]. haar [Cur38b]. Haarlem [Lor14]. Haas [Ein15c, HEB^{+}80, Ein15c]. habe [Hof08b]. Haber [Joh95, Joh06, Smi05a, Sto98, Str05a, Vis95, Cha05c, Sto98]. Habilitationen [HR02, RW02]. Habilitations [HR02, RW02]. habits [Bäu83, Mel79b]. hacedor [Gar40]. Hadron [Acz10, Lin09]. Haftung [Ein20h]. Hahn [Ano69a, Ano79h, Ano79g, B^{+}79, Ber69b, Ber74d, CSW97a, Ern92, For80a, Ger69, GRE^{+}01, Lam79, Lin60, MGG^{+}95, Mei05, RSPS49, Tru96, Wal03, vS14, Ano39b, Ano45b, Ano68b, Ano69a, Ano79g, Bad83c, Bad83d, Bau74, B^{+}79, Ber69b, Ber70, Ber74d, CSW97a, EHM^{+}79, Ern92, FPLR59, Ger69, GH84, Hah66, Hah67, Hah79, Hah81, Hah88, Her79a, Lam79, She83a, She83b, Wal03, Wal06b, vS14]. Hahn-Meitner-Instituts [MGG^{+}95]. Hahn-Meitner-Instituts/Fölsing [MGG^{+}95]. Haifa [RMR96]. Haigerloch [Fec90, Mac85]. Hail [Bro78a]. Hailed [Ano39a, Ano39g, Lau55b]. hajimerareta [EI40a]. halala [Nov92]. Halbleiterbauelemente [OSWR86]. Halbleiterbauelemente/Stitch [OSWR86]. Hale [Ano05e]. Half [Bar98b, Clo15, Joh07, McC68, Dav03, Jon17]. half-century [Dav03]. Half-life [Clo15]. Half-Lives [Joh07]. Half-Truths [Joh07]. Hall [BC95, Ein20p, EO79, Ano33b, Ano13a, Ber96c, Cas17, Fra93a, Fra93b, Sto99, Bad96]. Halle [HH14]. Halley [Van85]. Halpern [Sau16]. Halving [Ano41]. Hamburg [Boz14, Len17, Boz14, Len17]. Hamilton [Ein16j]. Hamiltonian [Ein25c, Ein29b]. Hamiltonsches [Ein25c, Ein29b, Ein16j]. Hamsun [HM11]. Hanbury [PBM^{+}12]. hand [BS83a, BS93, Hof10]. Handbook [BF14, Car15, OSWR86, Car12]. Handelns [JDB^{+}97]. Handling [Mar05]. hands [Sic09]. hands-on [Sic09]. Hanford [D’A93, FH11, Ger92, Ger97, Ger02, Ger07, Rol47, San95, Sne82, Wil11a]. Hanging [Hol88e]. hangja [Szi05]. Hanle [Pye80]. Hanne [Ein03a]. Hanoch [She16a]. Hans [Ano12d, Bur68, Gam42b, Pye80, SE79, Bet97, BL09, Ein91g, GW06, HFN06, MB66, Rei06, Rei11b, SE79]. Hans-Jurgen [Pye80]. Happen [CF12]. Happened [Wol16, Ler91, Pee14, Sch10b]. Happiness [Ein31l]. Happy [HG05a, Mey07, Rec10]. Harald [Mil12b, RS11b]. Harcourt [Sto17]. Hard [RA01, Sha73b, Woo06]. Hardback [Sta21, Coo04, Hen11c, Hes04, Hus15, Myr21b, Nav15, She15, Ste22, Vog20, War03a]. Hardcover [Plo06, Bäu83, Gor02, Nau17, Pes13, Ree17d, Ree17c, Rus17, WM08, Wal02b]. Hardly [Hof08b]. Hardy [Adl03a]. Hargittai [Bar05f]. Harmonious [KN08]. harmony [Lei08, Pye82b, Tym95]. Harness [Cau91]. Harnessing [Ein45d]. Harold [For71a, Kle75a]. Harpoons [Ano00c]. Harriet [Har79d, Nad79, Sei79]. Harris [Est06c, L.89]. Harry [Gol84b, Ham97]. Hart [Fri73]. Hart-Davis [Fri73]. Harteck [Erm89]. Hartree [Gás54, Gás54, Sla51]. Harvard [Par06, Tur06a, Her93]. Harvest [D’A93, Bra09]. Harzer [Ein14b]. Harzers [Ein14b, Ein14f, Ein14f]. hasard [Gis12]. Hasenöhrl [Bal11c, BR11, Bou13]. hasta [EI65a]. Hate [Ein38c]. Hated [Bar15]. hatte [HS14, KRZ^{+}05]. Haukelid [Hau48]. Haunted [Pic93]. Haunting [Tra16]. Hauptsatzes [Ein02a, Ein05w]. Haus [Grü90]. Havas [Pye89]. Havlin [MGG^{+}95]. havoc [Mil01, Pep04c, Goo02]. Hawking [Sci80, Ano10d, Bar15, BCO^{+}11, Cro01, Dan00, Haw92, Pic08, Sus08]. HB [PJMS11]. hbk [The12]. Head [Hos15, Ano47a]. head-and-shoulders [Ano47a]. Heal [Sta03b]. Health [Cla97, MHY95]. Heard [Ree17d, Fri12, Shu17]. hearing [Pol02]. hears [Sut97]. heart [KWM57, Mil05c, Mil06]. hearts [Ein17]. Heat [Ein05q, Ein05w, Ein05-35, MBD39, Ein02a, Ein04, Ein05c, Ein07e, Ein11g, Ein12c, Ein13g, Ein14o, Ein05-40, Ein05p, GBE95, Kox06]. Heathcote [Coh54]. Heats [Ein75, Kle65b, Kox95]. heaven [Cas03b, Tym95]. heavenly [vS21a, vS21b]. Heavens [Ano19b, Kae13b]. Heavily [GB15]. Heavily-doped [GB15]. Heavy [Fri39b, Tur40a, Ber06a, Hau48, Hau49, Hau53, Hau54, Hau63, Hau64, Hau65, Hau83, JDB^{+}97, Wig86]. hebräischen [Ein21w]. Hebrew [Ayn34, Ein21w, Kat04]. Hecker [Ree17c]. hectic [ABC^{+}06]. He’d [Ein54c]. Hedi [Ano05f]. Hedwig [Bro72, For70a, Kle70b, Kle70a, BB69, BB05]. heelalmodel [Gui09]. Hegel [CW84, Pao84, Zah87]. Heidelberg [Eis00, Hen86a, Hen11c]. Heike [RA13]. Heilbron [Nye88, Tre88]. Heinemann [Hen82b]. Heinrich [Ano05g, Buc94]. Heinz [Kap61]. Heisenberg [Adl03a, Ano51a, ARK^{+}99, Bes04, BDR93, B^{+}61, Bro72, BW04, Cas75, Eis00, Fer07a, Fis02, For09, GRE^{+}01, HEB^{+}80, Hen11c, Her76b, HvMW79, Her84, Her01b, Kle17b, KW93, KRW05, LW34, Mar73, MGG^{+}95, Mil12b, OSWR86, Par06, PHvMW01a, PHvMW93, PvM93, PHvMW01b, RW02, Rec02, RKW05, Rec09, Rec10, Sch05c, Tur06a, Wal92c, Wil36, vMHW85, vM93, vM96, vM99b, vM01, vM05c, vS14, AMS04, Alt06a, Ano33c, Arg09, Bes04, Bla86, Boh62, B^{+}61, Bre97, Bro71, BW04, Cam09a, Cap01, Car10, Cas76, Cas79, CB84, Cas92, Cas09, Cre10, CG14, Dör05, D^{+}92, Eck90, Eck01, Erm89, Fis02, Fri11, GA02, Gou47a, Han79a, Hei35, Hei65b, HBD71, HR02, Her76a, Her76b, HvMW79, Her84, Her01b, Hol00b, Hol05c, Kan02, Kle17b]. Heisenberg [KRW05, Lin07, Lin08, LRD^{+}91, LS93, Mac77, Meh87, Mon04, PLS02, PHvMW01a, PHvMW93, PvM93, PHvMW01b, Pfe77, Pow93, Pow94b, Pow00, Raz11, RW02, Rec02, RKW05, Rec10, RH02, Sch05c, Ser77, Tel76, vMHW85, vS14, vW02, vWR04]. Heitjans [ARK^{+}99]. Held [Ein20b, Meh73b, Pla20a, Ano23a, BPS87, BCS82, BB89, Cle06, Ein21h, Ein76b, Isr73, JRG02, LdB12b, Lor14, Pap15, Pla20b]. Helen [For80a, Hei74a, Moy81, Ram79c, Wig79, IK03]. helenchen [IK03]. Helene [IK03, IK03, Pop98, Pop05]. Heleni [Pop98, Pop05]. Helge [Bro00]. Helium [GR44, Lon38a, MBD39, SP59, Hul01]. Hell [Gia09]. Helle [See56b]. Hello [Mit02]. Helmholtz [Ein17d, Bia16b, Ein17d, Pul06]. Helmut [Hen86a, Hen11c]. helped [Jak13, Kie13]. Hendrik [Prz86, Kox88]. Henri [Spe67, Tho88b, Adl03a, Fau79, Fol17, Gen98, Gol67, Tug98, Wei12h].

Henrik [OBxx]. Henry [Bru03c, Fre08, Hes04, Kar05, Lit09, Sne82, Mal08, O’S01, Tan11, Wil88]. Hentschel [Kri08, Pye93, Pye95, Row00, Smi02, Sig92, Sta95b]. Her [Swa02, But23, CSW97b, Cur38b, Cur38c]. Heraclitus [Dan00]. Heraeus [JDB^{+}97]. Herausgegeben [Eis00, Klo63, Krö53, Win69]. Herbert [Dys76]. Here [Ano45b, Ano47b, Ano82c, Bro21, Cho16b, Ros12, Bre97, Coo83, Pai94, vC05b]. Heretical [Ano07e]. Heritage [B^{+}79, Vis94, Kra15a, Owe12]. Herken [vB02, Ber03, Sch03c]. Hermann [Del13, HEB^{+}80, LS05, Pes13, Sie10, A^{+}09, Ber94, Cor97, For70b, Pes13, Pye77, Sie10, Wei12w, WS^{+}01, Win69]. Herneck [For80a]. Hero [Cao07, Fer05b, Kay46, Kay68, Rob48]. Heroes [Ein44d, Ari03, Ari05, MS17]. Herrgott [BHJ^{+}00]. Herrn [Ein07a, Ein15c, Rit09b]. Herschbach [Ano87b]. Herschel [Lem09]. Herta [For80a, HW78]. Hertz [Ein11f, Buc94, D’A75, Ein05j, Hon89, How93a, OP87, Wof08]. Hertzschen [Ein11f, Ein05j]. Herzallerliebstes [IK03]. Hess [Ano51a, Ein20d, Ein20d, Van22]. Heterogeneous [Kri01, Ein20d]. heterogener [Ein20d]. heterostructure [Alf01]. Heuristic [Cer00, Ein05d, Ein05-37, Zah89]. Heuristics [RS99]. heuristischen [Ein05d, Ein05-37]. heute [Whe68b, Sal95]. Hevesy [vH62]. Hidden [Bha86, CHSH69, CHSH70, Far09, Gre99a, Gre99b, Gre00, Gre11, Hol05a, Lei08, Mac85, Pat93, YN11, Bel96b, Far13a, Fv85, Gri10, MFO14, Pan04, Ran05, Tho15, YN10]. Hidden-Variable [Hol05a, Bel96b]. Hidden-Variables [Pat93]. hier [vC05b]. Higgs [Kra14b, Mal79]. High [AGMT12, Dav83, Har79a, Kal13, Sei04, CW32a, CW32b, Jam06, KB60, KB83, Pal99, Str19, Wil74]. High-Energy [AGMT12]. high-intensity [Str19]. High-Precision [Sei04]. Higher [Gon05, Rei07, MS82]. higher-spin [MS82]. Highest [RA13, Kea18]. Highfield [Bon94, Gea94, Smo07a, Ald93, Bur94, Cas94a, Din93, Kan94a, Mil93]. Highlights [Hen05, Nac79]. highly [RN16]. Hike [Swa02]. Hilbert [Jam70, Som05a, Som05b, CRS97, Cor04a, Cor04b, Cor07, Dvo12, EG78a, Jun05a, Lan70, Lev04b, Meh73a, RS07c, Row01, Row06d, Sau05, Sau07b, Som05a, Sta15, Wei14a, Wue05, Row06a, Row06b]. Hildegard [Hei53a]. Hill [AS91, HS89b]. Hillel [UM92]. Hilmar [Sch09a]. Him [Bro21, Nor14]. Himself [Ano97c, Kae38]. Hintertreppe [Fis10, Fis12]. Hirnforschung [Bre08]. Hirochima [Hir98]. Hiroshima [Aue95, Cli09a, Her46a, Her85a, Her89a, Her94b, Sal95, VB96, Alp65, Alp66, Alp85, Alp94, Ano81c, Ano08b, Aue95, BMD^{+}11, BL96, Bur83, Gol96, Gra05, HW55, Her46b, Her85b, Her93, Hog96, Jay90, Jen02, Kif95, Kur86, Lau51, LM95, Man5 a, Man5 b, Mad95b, Mad04, Mad07, Mal85, New89, Ōe81, Ōe95b, Ōe96, OJCP^{+}47, Pac72, Pac81, Pre05, Say85, She87, She03, Sod95, Tak95, TM77, TM90, New65, Ves82, Wal05a, WY05, Wal05b, Wal05c, Wal06c, Wal06d, Wil96, Wyd84, Yor87, Fre08]. Hiroshima/Nagasaki [Sod95]. Hirschmann [Pye06]. Hirzel [Wil36]. Histoire [Bro83, Jac12, Lev04b]. Historian [Gla00, HSW81]. Historians [Gea02, Kel07, Kel09]. Historic [Cha49a, Jos09]. Historical [AH93, Ano23a, Ano83, Ano84b, Bad83a, Cra02b, Fer83, Gil84a, Gol84b, Gol99, GL23, GR23, Hen86a, How97b, Jam74d, Kil84a, Kra96a, Lar84a, OOM21, Pea08, SS16, Sch17a, Sch84b, ST06, Wal04b, Wil84, Bin06, Cha78, CF80, CC74, CC98, Dör05, DGG^{+}21, ER13, GG05a, Gon96a, GR17, Han18, Hen92a, Hof18, Hof97b, HE82, JN14, Kil84b, Liu92, Mac93, Mil90, Sin04a, Wal77b, Wal92d, Wal06e, Wal15, Wil89, Bar05f, Ano04b]. historie [Kot93]. historien [NN01]. Histories [BW01, Nye19, Pei97, Sta98c, LML88]. Historiographical [BLR15, Bur90, Wal90b]. historische [ER13]. Historischen [AH93, Hof18, Hof97b]. History [And17, Sch12b, ARS15, Bag10, Ben11, Ber93, Bič89b, BMW89, BM77, BF14, Byr80a, Cao07, Car04a, Cas86, Cho10a, D’A00, EG80a, EJN93, Eis06, Esh09, Fol15, For05, Fow03, Gam42a, Gam42b, GSW95, GR07, Gly90, GRRS99b, Goe04, Gol82, Grü62, Han80b, Hen86a, HRR03, HOO88, HS89b, Isa15, JM17, Kap20, Koh95, Kox95, Kra82a, KV11, Kur79, Lau55a, LMOS15, LE96, Mac85, Mah09, Mah14, Mao07, Mar04c, Mar05, Mas95, Mau09, McM51, Mil79, Nav21, Pal83b, PJMS11, RS03a, Rho88a, RT90, San95, Sch12a, Sod95, Sta95a, Swe13, Swi45, Tes69, Tre88, UM86, UM92, Vis94, Wac06, Wal17, WM08, War03a, WP85, Wei88, You09, Zau05, vD15, Abr88, Abr01b, AH93, Amr59, Ano98, Ano05n, Aue95, Bag11, Bat03, BT15, Ber07c, Ber09]. history [BL96, Bre08, BB13, Bro00, Bru68, Bry09, Buc20, Büh15, Cal86, CHW91, Cas17, Cha49a, Cir07, Cir08, D’A01c, Dee08, Der15, Dur05a, EG78a, EK92, Ell89a, ET07, Far13a, Fel00, FV13, For15, FFF09, Gib96, Gil19, GHK^{+}96, Gon00, GA71, GHW09, GR15, GR17, Hab02, Han88, Har96c, Haw02a, Haw88, Haw92, Haw98, HM05a, HTS83, HS89a, Hei05, HA62, HD69, HDHA89, HEWW15, HHMW93, Hog96, Hol95b, Hol96b, Hol00a, How96b, Hug04, Inf55c, Ing05, Irv67a, Irv83, Jac11, Jun58, Jun82, Kie94, Kot93, Kra88, Kra99, Kra07a, KK05, LM94, LP05, Lev04b, LSW09, Lin04b, Liu92, Lon13, Mad07, Mah15, Mai99, Mar17, Mar00, Meh98, Myr21b, Nor12, Nor65, Nor90, Nor94, Olw04, Per23b, Pow93, Pow94b, Pow00, RR82, RB04, RS09, Ree14c]. history [RG07, Rob50, SSG13, Sch99b, She16a, Shi75, Smi10, Söd10, SBF00, Tay49, Thi46, Tio09, Tod14, Van07b, Wea88, Wel93, Wes04, Wie06, tH20, vBL96b, Ham97, Mas99, Ree15d]. Hitherto [Rol10]. Hitler [Tru96, Ber06a, Ber96c, Bey77, Bro92, Bro02, Cor03, Cra00b, Dei96, Hen98, Hof55, Kac96, MP99a, MP01, MP12, Say85, vC16, vCH18, Hoe18, Bad96]. hjalp [Jak13]. hjerne [Pat01a]. hoc [Ray92]. Hochschule [EGO97]. Hoddeson [Hes04]. Hodgkin [But23, KKJ99]. Hoe [Ano15f]. Hoffman [For80a, Fri73, Wig79, Mal79]. Hoffmann [Hei74a, Rob38, Ber79b, CC67, Moy81, Ram79c]. Hold [Ano45a]. holdings [Egg02]. Hole [Ano18b, Bar15, Del11b, Fle18, How93a, Hub13, PD15, Pye10, Sus08, Wei15b, Whe79a, AAA^{+}18, Ano18c, Joh19, Lev16, Mel09, Shi75, Wal94a, Wei13g, BPL05]. Holes [BR21, BT12, Ber96d, CB15, Eis93, Har95, Imp19, Pen79a, AMSmS96, Ano10d, BCO^{+}11, Ber96b, Bes04, Bro10b, Cha88, Cha98, Cho08, Gen05, Haw88, Jan05, Mil05a, Mil05b, Mus15b, Sri99, SL91, SL05b, Str05b, TW00, TPM86, Tho93, Tho94, Wal77a, Wal92a, Wal98, WF98]. Holism [How89, Ash95]. Holistic [Tre79b, Hen87]. Holmes [Bad15, Lac08, Bru97a, Bru97b, Bru98]. Holocaust [Bla02, LM90]. holography [GLW15]. Holstebro [Jak13]. Holt [Bru03c]. Holton [Gol84b, Kil84b, Par06, Tur06a, Ano83, Bad83a, Fer83, Gil84a, Kil84a, Lar84a, Sch84b, Top89, Wil84]. Holy [Ano97b]. Home [CC96, Ger07, Hoy94, Wac06, Whe92, Ger92, Ger97, Ger02, Ano13b, Kni13]. homme [Rob15a]. homo [Ano15f]. homodyne [FTZ^{+}15]. homogenen [Ein10a, Ein05-29]. Homogeneous [Ein05-29, Lev07b, Ein10a, SS15]. Hon [Eve39, Eve13, You98]. honeybee [DGG^{+}21]. Honor [Ano29a, Ano63b, ARS15, GSW95, HOO88, KS95, RA13, RDS03, Ano69a, ACH^{+}03, Kea18, Kra15a, MB66, NJP92, SH99]. honorary [GKP12b]. Honored [Ano45b, Pin00a]. Honoring [Dad78]. honors [Ano97a]. Honour [CHR90, Ber84a, RTR87, RMR96, Sch00c, SO72]. Hoover [Ber02, Jer02]. Hope [Ein34d, RFP^{+}45, Lil63, MT98, NKY15, Smi65, Smi70]. Hopkins [Ham11a, Mil11a]. Horizons [BPP^{+}11]. horloges [Lam07]. hormones [Hen03]. horse [HB05, KS12]. Horst [Bäu83, Bäu83]. hose [Ahe03]. Hostile [For71b, For80b]. Hot [Hoy54, Mir02, Kos20, Pee14, Sen21, Kra96b]. Houches [HLZJ91]. Houghton [Nah01, Sto17]. hour [Con17]. House [Duf46, Rus17, Chu60, Chu98b, Chu01, Grü90, Hei53a, Irv67b, Lut05, RS11a, CC96, GW83, Ski04]. housing [Mar00]. Houtermans [ARK^{+}99]. Howard [Gly90, Hen82b, Kle82a, Pye95, War96b, Bur90]. Howson [Tha77]. Hoyle [Gre05c, Mad83, Mit05, Sul64]. Hoyo [M08]. Hrn [Abr12a, Abr12b, Ein18i, Ein05k, vB19]. https [Ree16a]. Hu [Cao06]. Hubble [FM10, HH96, SN93]. Hubert [Ano79g]. Hubert-Schardin-Medaille [Ano79g]. Hudson [Bru97b, Bru98]. Huemul [Pye86, Mar96a]. Huge [MKBF30a, MKBF30b, Bre08]. Huggett [Pri99]. Hugh [Hea13]. Human [Bel23, Ber79b, Boh58, Boh10, Com46a, Duf46, DH79, For80a, Kur12, Moy81, Ōe95a, Oha08, OB05, Ram79c, Ray95, Wal16, BNY81, BF78, DC08, DH81, DHR13, HB01, Kai94, PSC^{+}97, Rab63, Rob58, Tip10, Val55, Wig79]. Humanism [Mar88]. Humanist [Nee79, Sch81a]. humanitaire [Dur12]. Humanitarian [Fri79a]. Humanities [For80a, Hen82b, Lew79, Ram79a, Lew79, vDP17]. Humanity [Neu06, Rot96, Bre05, Her33, Her63a, Her67, Her75]. humankind [ABW99, Jay90]. Humean [Mor16]. Humorous [Ano21d]. Hundert [Ren05c, Ren05b, IRW31]. hundertjährigen [Ein24k]. Hundred [Bar05f, BLR15, BFMNZ91, Nic34, Rob15b, Erb10, Gla06, GMW80, GMOP12, HI87, Hel80, Hen03, Hot15, IRW31, OONM18, Ren05c, Ren05b, Rob05]. Hundredth [Ein24k, Ano80a, NHP^{+}79, Sto96]. Hungarian [Ein22a, Nov92, Szi05, TB02, Tőz15, Wig65, von90]. Hunger [Bra14a, Ano14b]. Hunt [Bow87b, Ham11a, Mil11a, Rus17, Set11, Ber06a, Gat09, Lev15, Sam10, vCH18, Fel11, Joh11b, Nuv11]. Hunter [Nad02]. hunting [Rio87]. hur [RE90]. Huszadik [TB02]. Huxley [Coh31]. hvor [Jak13]. hybrid [Leh19]. Hydrodynamic [EES49]. Hydrodynamics [BBCPL09, Tau63]. Hydrogen [Bac50, Bet50a, Bet50b, Bet63a, Bro15, Dys76, Ken54, Lev98, Res50, Rho95, Rid50, Rig02, Ros79b, Ein20n, Goo03a, Hoy46, WG17]. Hydrogen-Bomb [Dys76]. Hyperbolic [Ung98, Ung07]. Hyperspace [Mlo01]. Hypersurfaces [Cob40]. Hypothèse [Poi02]. Hypothesis [Bos24a, Cra84a, DR20, Ein16n, Ein05-38, Poi02, Cas05a, War91, Ano29d, Mar04d]. Hypothetism [Kos98].

I-stuff [HS06]. I. [Bor78a, Col93, Rabxx, Zal91]. Iberian [GPSR01]. IBM [Bla02]. ice [Ano96d, Bre08]. iced [Dan97b]. ich [Com56a, Ein21u, Ein22r, Hof08b]. icon [Bar05c, Bri05c, Ros11]. iconic [Diz05]. iconoclast [Ros11]. icons [WSG00]. Ida [Hab09]. Idaho [McK03b]. ideōn [EI78a]. Idea [Bar15, Dir79b, Ein31h, Ein54e, Ein82d, Kae31, Ano10d, Bro78b, DS01, Ein15g, Ein23d, EI38d, FSE31, JMH^{+}05, Sch35a, Sch99a, Str99]. Ideal [Ein24h, Ein25h, Ein25i, MBD39, PS10]. idealen [Ein24h, Ein25h, Ein25i]. Idealist [Pye13]. Idealistic [Göd49b, Gri81, Kae31, You91]. Idealization [Sha98]. Ideals [Hoo91, Wig79]. Ideas [Ano22a, Ano25a, Bič89b, Bro39, D’A00, EI38b, EI38c, Ein54b, Ein73, Ein82c, Ein94c, Ein95c, Hof97b, Jam70, Kae38, RF69, Sac81, Thi21, WS89, Wit82, BR05, Bar00b, Ben14, BK05, Bre97, D’A01c, EI38e, EI39b, EI42, EI48a, EI57, EI59, EI61d, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI98b, EI13, EI15, Gra82, HT10, Hen82b, JP55, Lan70, Moo02, MRS17, Sin61, Sin70a, SA85, SA87, SA95, Spi05, Nor00]. idee [EI65b, EI13, Tug98]. Ideen [Hof97b]. idées [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. idei [EI66a]. ideias [EI39b]. ideilor [EI57]. identical [Fer24]. identici [Fer24]. Identification [CZPR17]. Identifying [Kre58]. identities [BEM12, Ein50a]. Identity [Gil05, van12, Hal04a, vD21]. ideological [Müh11]. Ideology [For73, Rol10, Vuc01, Wei03]. idéutveckling [EI38d]. idiosyncracies [Cah04]. idiot [Mor00, Mor04]. idiots [Spi96, Van96]. IEEE [Ano09b, Pin00a]. If [Com46b, Eic94, Ein54c, Joh14]. Ig [Ano96d, Mir94, Nad02, Ren91, Sty94]. Ignited [Bey06, Cas05b, Del09b, Gre05b]. Ignites [Mos13a, Mos13b]. Igniting [HR17]. Ignored [Van88]. ihr [CSW97a, Cur38c, Ein27e]. ihre [Bor20, Bor22a, Hah48, Wol87, Yos95, vBL96a, vBL96b]. ihrer [Ein02b, Ein05-39, Tre66, Wie06, vBL96a, vBL96b]. II [ARK^{+}99, Eis00, For80a, Ham97, Har91, OSWR86, Ree17c, Ser88, Sha75, Wig96m, Wol87, vMHW85, vM99b, Abr04c, Ano82b, Bar18a, BB49, Bet50b, Cas17, CW32b, Con02, Ein22l, EI40b, Ein44a, ES46b, Ein53n, Ein79c, Fei66, Feu71b, Goe49, GR44, Gol96, Gor07b, Gru05b, HH02, HS51, Kes65b, Kie13, MH09, Mel00, MBD39, RB05, Rol47, Rus90, Sch52a, Sch77b, Sha73a, Sta06, Ven94b, War93, Wei12t, Wei95, Whi43b, Wil16b, Wol87, Yan13b, Yer04, Zah73b, vMHW85]. III [BDR93, GKP12b, PvM93, You95, vM93, vM01, Bac50, GKP12b, HH03, Law21, Sch77c, Tho18, Ven94c, Wei12u]. ik [Ano15f]. ikani [EI40a]. Ilha [NA92]. Ilich [Fre96]. iliskinlige [EI94]. ilk [EI94, EI11]. ill [Tur06a, Vis95]. Illinois [Krö53]. Illness [Kan94b, Sri94]. illuminated [Bon94, Van98]. illuminations [MRS17]. Illus [Kuh54, Boy88, Hie68, Jam70, Sci80, Van22]. Illusion [Sus06, Haw07, Moo10, RW07, Zuc82, Zuc83]. Illustrated [Bai07b, Abr88, Abr01b, Gib96]. illustration [FW17]. illustrations [Hea13, Wal77b, Wal92d, Wal06e, Wal15]. Illy [Bad15, Lac08, Maa13a, Maa13b, Pye97, Pye06, Row00, Sie97]. Ilya [Hen82b, Kle82a]. im [Bad04, Bau09, B^{+}79, B^{+}01, Bos24b, Bre15, Chi94, Cun22a, Ein95, EL08c, Ein19j, Ein22p, Ein25a, Ein27b, EL05d, ER13, Enk18, Foc53b, Föl91, Fre27a, Hei69a, Hei72, Hei75b, Her78, Hof08c, JDB^{+}97, Kie94, Kon15, Kra88, Luk02, Pau25, Pes13, Pla00b, Pla01, RKW05, RS96, Ren05e, SE79, Bad12, Sie10, Ste89, Gru05d, Kra81, Ros86, T– D82a, T– D83, T– D85, T– D88, T– D93, Sha92]. image [Ber74d, Gwy10, OV86, Sti13c]. Images [War79, van12, Diz05, FF95, Fis92, Roc03, SBF00, Ver01, Wea88, For80a, Pye80]. Imaginary [Mar04c, Wei06b, Whe79a]. Imagination [Hol82, BF78, Fre18, GP11, Hol78b, Hol98c, Sin01a, Wei12a]. Imaging [CLD^{+}03, WCL^{+}00, Man13, WHL^{+}05]. imagining [Ari03, Ari05, Dan00]. Imbens [Mea21]. immagine [Lau15]. immagini [Ver01]. Immanuel [FTG08, FTG08]. Immigration [Ein19d]. Immortal [MKBF30a, MKBF30b]. Impact [AP00, Bro71, De 10, Ein22h, Jac87, Kan87, Pai88, GS05, GS06, GRS16, Gol84c, Hen82b, Inf79a, LW71, PLS02, Tam14, Yan80]. Impacting [AZ05]. impatto [Tam14]. impensable [Gis12]. Imperative [Ein44b]. imperfect [Gle10]. Imperial [Fra99, HL08b, HL12b]. implausible [Shi75]. Implications [How89, NM12, CT15, Din49, Jay90, Mey85, Sch00c]. Implied [MSS22]. Importance [Wei12p, vW79]. Important [Hol89b, Lee06, Wei12a]. Impossibility [Wei12n]. Impossible [Wal88a]. impressäoes [Sou19]. Impressionen [Pau61b]. Impressions [Ein21g, Ein23e, Ken22, Pau94c, Sou19, Pau61b]. Impressões [Ken22]. Improbable [Nad02]. Improved [QPSD13]. Improvement [MNT^{+}00, Pla00a]. Imre [CFW76]. in-between [Tym95]. Inaugural [EP06, Ein14a]. inaugurate [Van04]. incandescent [Ein29a]. Incarnation [Kat04]. Include [Res50]. Including [Ano07h, KS65, Rol10, SS98c, Ano64, Ano65b, Ano72, Bro20b, Don70, FM92, Lun92, Nob72, Nob64b, Nob66a, Nob66b, Per97, SS98a, See25]. Income [Dac20a]. Incompleteness [BC24, HS10, Hel15, Jam74c, Pat85]. Incomprehensible [Col13]. inconnu [Ein90]. incorrigible [Bje02]. Increase [Rob75]. increased [MTVB88]. increased/Océ [MTVB88]. increasing [Lui11]. indefinitely [Bra14b]. indéformables [Pia31]. indeterminación [Bla86]. Indeterminacy [Han79a]. Indeterminism [Web93]. Index [Ano60, Ano07i, Ano07j, De 10, EKS^{+}08, KSB^{+}09, Bar05f, BCO^{+}11, LLH22, Van22]. India [Abr09, Sin07]. Indian [Abr98]. indicators [BSPKN03]. Indifference [Man86]. Indignation [VB96]. indigo [Aer99a]. indispensable [Nor02b, Nor03a]. Indistinguishable [Bel23]. individual [ES45, ES46a]. Individuation [How97b]. Induced [Ken13, Mil08, BCO^{+}11, vHJK39a]. Induction [Ein19f, Max93a, Max93c, Max93b, Ein12d, PL17, SU06, Ein19g]. Induktion [Ein19g]. Induktionswirkung [Ein12d]. Industrial [Cla97, Van04]. Industrialization [DWL98, Kra82b]. Industry [Tor82, Lun74, Tol76]. inédit [Sto96]. inédito [Ein31m, Ein31n]. Inequalities [YC18, GS80, GHSZ90]. inert [Coo83]. Inertia [CW95, Sci53, vL49b, Ein05b, Ein20n, Ein05n, Ein05-34, Ein05v, Fok65, Ren97b, vEPF09, vEPF22]. Inertial [Ein11h, Nor07a, Ein06a, Ein07d, EL08c, Ein10e, HM05b, JDB^{+}97, RKD64, Vig95, Wei12j]. inevitable [Ber83]. inexplicable [Ein19b]. infamous [Wic95]. infectious [RF05]. Infeld [Cas75, Kae38, Mal79, Rob38, Sch50a, Ano39d, Ano50a, Ano68a, Arm38, Bič79c, Bre89, Bro39, C.51a, CC67, Len39, Lom05, N.38, Qui40, Sta99a, Nag50]. infini [Bla93]. Infinite [TVL^{+}09b, Bar05d, Bar05e, Bru03a]. Infinitesimal [Wey07]. infinity [Bla93, ’t 00, ’t 02]. Inflationary [GK05, Gut97]. influência [Inf57a, Inf60, Inf61a]. Influence [Bir25, C.51a, Cro16, Cun22b, Edd19, Ein27e, Ein31h, Ein54e, Ein82d, Ein87a, Ein05-31, Hof49a, Nag50, Oha79, Pye80, Ros80, Sch50a, Tau79, Ush49, AS79, Ano50a, Dir80b, Ein11i, Ein31g, ES45, ES46a, Ein11b, GB06, GB08, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58, Inf60, Inf61b, Inf61a, Inf61c, Inf63, Inf68, Inf74, Inf83, Inf92, JP55, Kan96b, Nac73, Nac77, Par11, Pat99, Pfe77, Sch79d, Van07b, Ein27c]. influenced [Ank15]. Influences [Cah85, Hol68a, Kie82]. influencia [Inf61c, Inf83]. Influential [Ree17d, Rog10, Pra06, Shu17]. Influenza [Ros80, Inf63, Inf68, Inf74, Inf92]. Information [Ano15a, Fal09, Kni16, WL12, Whi06, Bau95, Bau97, Baw06, Hau08, Row06d, San77, SBB16]. Infrared [MTVB88]. Infrarot [MTVB88]. Infrarot-Observatoriums [MTVB88]. Infrarot-Observatoriums/Neu [MTVB88]. Infrastructure [LHZA20]. infrequent [RN16]. ingen [MW46b]. Ingenieur [Ren05c, Ren05b, Ren05e]. ingenious [HH16, Vaa17]. ingenuity [Top07a]. initial [EI65b, EI13]. initiation [Ber84a]. iniziali [EI65b, EI13]. Injustice [CSW97b, CSW96]. inleiding [Ano15f]. inne [Gol98a]. Inner [Hay02, Ree06, Bar99, Cze05, Gol05, YN10, Pep04c]. Innere [MGG^{+}95]. innermost [Gie05]. Innocence [Stu18, FW07]. Innovation [Ano05b, Ell90, Hab82, LU13, Ros82b, Sie91, Waz10, CK^{+}04, Kee04, Lar08, Mau09]. Innovative [Pin00a]. inquiries [Mil90]. Inquiry [Gib19, Ham85]. ins [Bre08, Erc55, Ing05, MGG^{+}95]. insan [Tip12]. inscriptions [Fei01]. Inselhalle [Bad12]. insider [De 92, RB04]. insiders [Fre81a]. Insight [Ano18b, Hof08a, Sto05, Sti13c]. Insights [HF05, Mos21, RS03b, Hah75, Kel06a]. inspiration [Wes14]. Inspired [Lau55a, Ben79, Tou14]. instability [Edd30a]. Institut [AH93, Kan96a, MTVB88, Ren05e]. Institute [AH93, Ano34a, Ano34b, CR07, GC20, Hoh13, Kan96a, Kat04, MTVB88, MA31, vBL96a, vBL96b, Ano45b, Ano68a, Fle33, GR06, HL08b, HL12b, Kha12, LB02, MTVB88, Reg87, Sha73b]. Institute [?]StuttgartMeyer-Ter-Vehn:1988:ESP. Institution [Rec86, Rei86, Sch28, Cra84b, Mat86b, Koh95]. institutional [Hen92a]. institutionalisation [Pye75]. institutionalization [Pye75]. institutionellen [Hen92a]. Institutions [Ihd74, SBB16, ZZ23, vBL96b]. Instituts/Fölsing [MGG^{+}95]. instruction [FTG08]. instructions [Coo83]. instrumentalism [Luk02]. Instrumentalismus [Luk02]. instruments [SBF00]. insulators [GRT08]. Insuperable [Uff06]. Int [Flo06, Kra06]. integrals [FH65, FHS10, JP08]. Integration [BOS96, Ein16l, Sau08]. integrity [Hal04a, Whi16, Nau17]. intellect [Tho06]. Intellectual [Can05, For71b, For80b, JM86, Kar10, Tom10, Wel08, Gan97, G^{+}11, HS06, LM94, Ris06, Ser88]. Intellectuals [CP15, Ein20j, Kan87, RRR07, Bes93, Ein59b, Ein48g]. intellectuele [Ris06]. Intellettuali [RRR07]. Intelligence [Ano31f, Goo03a, Gru05d, Ric06a]. Intelligent [Sus06]. intelligenten [GRE^{+}01]. Intelligibility [Hoo91]. Intelligible [Ano20a]. intense [Tsu99]. intensity [Str19]. Intention [Ein36c]. Inter [FL49]. Inter-Molecular [FL49]. interacting [RDF^{+}08]. Interaction [HEB^{+}80, BRVF84, vBT94, Tsu99]. Interactions [HHJ^{+}06, Sch06, BSSP^{+}08, GSPH00, Pei71, Vel00]. Interactive [Tra16, Sic09, Sta06]. interacts [Phi97]. interaktiv [Sic09]. intercontinental [CGG^{+}13]. interdisciplinarity [LRH22]. Interdisciplinary [BM09, BAM99, RWL23]. Interest [Bar18c, L.89, Mil12b, She16b, Sho21, Sud11]. Interesting [McC68]. Interests [Ort95]. Interface [Ker89a]. interference [Ein26b, TEM^{+}89]. Interferenzeigenschaften [Ein26b]. interferometric [ZCPB11a, ZCPB11b]. Interim [Wal83, Sig78]. interludes [Des72]. Intermezzi [Des72]. intermolecular [Ein02b]. Internalizing [Nacxx]. International [Bar93, BCEP94, CHR96, Dot82, Ein20j, Ein48k, Ein55j, EK92, GRRS99b, JP08, Meh73b, Mer79, Sch83b, SG82, War89, Ano67, BIK83, CR07, CGG^{+}13, Cat04, Du 52, LP05, Lun74, RMR96, Sin07, Spo16, TA17, Ano05d, CLE27]. Internationalism [Eli92, For73, Hoc93, Wal94b, Cra92, Ene12, RK94, Wei94, Wid95]. internationell [Lun74]. Internet [BHJ^{+}00, JDB^{+}97, JDB^{+}97]. Internet-Adressen [JDB^{+}97]. Internet-Adressen/Blick [JDB^{+}97]. Interoffice [Nol05]. Interparticle [WF49]. Interplanetary [Ein20f]. Interplay [Elz95, Fru82, Tsu99]. Interpretation [Ano94b, BBB^{+}09, Bor53, Cus82, Ein29c, Ein53d, Fer81, Gin82, Gio13a, HB63, Hel84, Hol05a, Hug89, Inf31, Jam74b, Jam74c, JP55, Kai81, Kra20a, McH07, Mel82, Pal83a, Pal83b, PFM17, Ram83, Sch21, Sci61, Sig92, Stu94, Whi83b, BB55, Bor55b, Bor63b, Cam09a, CSW97b, Ein16c, Ein23h, Flo06, Kop23b, Kra06, Mil81, Mil98, Nav21, Pat95, Sla22, Van07b, Vig94, Vig95, XMY^{+}09]. Interpretationen [Hen90, Pye93]. Interpretations [Fin93, Flo05, Jam74d, Kra06, Sta95b, Hen90]. interpretazione [Kop23b]. Interpreter [Ano29d]. interpretive [Tel95]. Interrogating [OM18]. Interscience [Hou65, Lin60, Mal65, See65b]. intersection [Del75, Gim12]. Intersections [Sch06, LSW12]. intertexture [Hen97, Smi02]. intervence [Kot93]. intervention [Kot93]. Interview [Ano05m, Coh55, Ein20f, Ein35f, RS06, Vie29, Wig96j, Yer03, Yer04, Hei65b, Joe19]. Interviews [BH89b, Erm89]. Interwar [Ash11]. intimate [Far79, MW44, Mar51]. Intimations [KK83]. intimité [Mar51]. Intra [FL49]. Intra- [FL49]. intraoral [Ano05-35]. intrarenal [MNOW00]. intriguing [San81b]. Introducing [BE02, D’I92, SM99, Sta89c]. Introduction [Alc08, Ano22a, Ano79b, Ano15g, Ber42, Ber76a, Ber07b, Bro72, Bro20a, Bus03, Car04b, CB15, Col13, COS05, Dav14, Dor71, Duf46, Ein24e, Ein32b, Ein34e, Ein50g, Ein51f, Ein55e, Fan03, Gen98, Gri95c, Gri05, GN11, Hak99, HY15, Hol82, JNR^{+}07b, JNR^{+}07a, JL14a, Lar19, Nar10, PK06, Ren05f, RS07b, Rin82, Rin91, Ryd17, SP12, SF33, Sla39, Sla70, Vai89, vB12, vB14, vB18, vB20, AE01, Ano15f, Ben14, BGLR17, Bor22b, Che10, Cho08, Clo04, Clo09, Fok65, Fra11, Har03b, Her21, Hie68, Hol98b, Hol98c, Ken03, McC11, Mil94, Moo81a, Mul10, PS95, Sch17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, Sch93c, Spe62, Sta08b, Ste04, SFW91, TW92, TW00, Tel95, Thi46, VM05a, WS^{+}01, Wit62]. Introductions [Gru07]. Introductory [Rec02, Bro20b]. intuition [Pul06]. Intuitionistic [WL12]. Intuitive [Lar19, Ben14]. invalidation [San81b]. Invariance [Cer00, Cos87a, Wig64b, Wig64c, HZ01, Sla22, Wig64a, Wig65, Wig67a]. invarianciaelvek [Wig65]. Invariant [How99a, PV49, Lie78]. Invarianzprinzipien [Wig64a]. invasion [Gia09, Ska94]. Inventar [BH01]. invented [Ker89c, Sch12d]. Inventing [Hun07, Mac93]. Invention [Hug93, Mon19, OR21, Ber05, Ren05a, Smi10]. inventions [Ill12, Sch12d, Maa13a, Maa13b]. Inventor [Bri82, Con70, Fok55, Fre13, KS81, Tal32b]. Inventors [Ano56, Hug93, Aas88]. Inventory [BH01, dP94]. Invents [Den84]. inverse [AHH05, Mal79]. inverse-square [AHH05]. Investigation [Hol89b, CT15, Ein95, EH10a, EM31b, EH05b, Ein22o]. Investigations [Ein56c, Els24, Nor04, Woo28, vE96]. Invisibility [Sim22]. invisible [Pan04, Pea90, Tur06b]. Invited [CR07]. involutions [SS21a]. Involvement [Nee79]. Inward [Pai86]. Inwiefern [Ein20g]. Ion [HEB^{+}80, ZPSK10]. Ionen [JDB^{+}97]. ions [CW32a, CW32b, JDB^{+}97]. Iran [Ric06a]. Iraq [Bie03, BEM12]. Irene [Bro72]. irgendwelche [HS14]. Iris [Bad15, Lac08]. ironical [Ryc92]. irradiation [HS39a, HS39c]. irreale [Vir92]. Irreconcilable [Ein48l]. Irregular [Ein14j, Ein14h, Ein86c, Ein87b, Eĭn85b, Eĭn86a, vB19]. irrégulières [Ein14j]. irreverent [Bra14b]. irreversible [Bat62]. Irrtum [Bod17, vM05b, vM09, vM12]. Irrwege [Mei63b]. Isaac [Bes04, Rig00, Aĭn77, Dry14, Ein27d, Ein28b]. Isaacson [Gor07a, Sch07c, Smo07a, Weg08]. Isaak [Hoh13]. ISBN [Ach13, Ber07a, Bon94, Cah04, Coo04, Del11b, Del11a, Del12, Dix00, Eng17b, For05, Gil12, Gor02, Gor07a, GRC18, Har02b, Hea13, Hen86a, Hen88b, Hen00b, Hen11c, Hes04, Hoc93, Hoh13, Hus15, Kar05, Kil84b, Kle14a, Koj08, Kra02, Kri08, Mar10, Mas01, Mau09, Mil11a, Myr21b, Myr21a, Myr21c, Nah01, Nau17, Nav15, Pec04, Per10, Pes13, Rec86, Red84b, Ree15d, Ree17d, Ree17c, Ren21, Row06a, Row06b, Rus17, Sch07c, Seg91, She15, Sho21, Smi02, Sta99b, Sta21, Ste22, Sto17, The12, Tho11b, Tre88, Tur05, Tur06a, Van22, Vog20, Vog21, WM08, War96b, War03a, Weg08, Woo06]. ISBN-13 [Del11b, Del11a, Mar10, Tho11b, Weg08]. Ishiwara [HY86]. Ishiwaras [HY86]. Isidor [Rig00]. Island [Lig18, Pap15, For82b, Wal04b]. Islands [Mah17]. isolated [HFN06]. Isolation [Ano26a]. Isotope [Ano39h]. Isotopes [NBDG40, HS39a]. Isotopic [Mah17]. Isotropic [Lev07b, vKL49]. Israel [Sci80, Ber82c, Nav82, Rol10, Ros11, UM86]. Issue [AP00, Ano95a, Ano04g, Ano05-43, BCR93, BGLR17, COS05, EBRC93]. Issues [Ber76b, W^{+}87, WdC02, Ano01b]. István [Bar05f]. Italian [Aba09, Alt06a, Ano82d, Ano01c, Arg09, BF16, Bal94, Bat03, Bel88a, Bel88b, Ber81, BF02, BES86, Bri05d, Bru07, Cal75, Cap01, Car82a, Car82b, CM02a, Cas75, Cas14, CDS01, Ein21p, Ein22k, EI48b, Ein50k, Ein55f, Ein60c, EI65b, EI13, Fer24, Gas05, GA02, Gen98, HC94c, Inf63, Inf68, Inf74, Inf92, KT15, Kle17a, Kol58, Kop23b, Lan67c, Lau15, LS08, Mal00, MFS^{+}6x, Mas99, Mas02, Pal99, PB72, Pot11, Rec07, RRR07, Ros80, Seg76, Tam14, Tio09, Val06, Ver01]. italiana [Ras68, Tio09]. italiani [CDS01, GM04, MFS^{+}6x]. italiano [Ano82d, LS08, Per10]. Italy [Bes93, Meh73b, Goo75, Ree87]. Itinerari [Per10, LS08]. itineraries [LS08]. Itself [Foe23, BB00, Diz05]. IUPS [Ano93c]. IV [ARK^{+}99, Eis00, vM96, vM99b, vM01, vM05c, Ano21g, GRRS99b, HH04a]. IV/Teil [ARK^{+}99]. Ives [TH79]. Ivor [RTR87]. ivory [Jos95b]. ix [Cah04, Hen00b, Rec86, Sha98]. Izloshenie [Ein21r].

J [Bal96, Cah90a, DTF97, Ein12a, Ein15c, Ein24c, Ein05d, Fel11, Fre08, Gol87, Ham11a, Har96a, Har96b, Hen82b, Hen88a, Her03, Hie68, Joh11b, Ker89b, Kil95, KS95, Loe88a, Mic68, Mil11a, Myr21b, Myr21a, Myr21c, Nah01, Nuv11, Nye88, Pye95, Pye97, Pye06, Ram82, Row00, Set11, Sie97, Smi89a, vS21a, vS21b, Str90, Tre88, TH79, Wal99, War95, War96a, War96b, Whi67, Yan87, You95, dP94]. J. [Ber84b, Ber82d, BieWSK67, BS05, Cas05c, DTF97, Ein12a, Ein24c, EVB^{+}72, Ein05d, Els95, Jer02, Jon89, Kan92a, Kel06a, McM05, Nav94, Nev92, OBxx, PC06, Pie24b, Pol02, Sch92b, SO72, TH79]. ja [Hot15]. jacht [Hoe18, vC16]. Jack [Pin00a]. Jacob [Bus12b]. Jacobi [Ein17c]. Jagdish [Hen86a]. Jahr [Ano93a, Enk18, SE79]. Jahrbuch [Mil92]. Jahre [AH93, Ano23c, Ano61, For80a, Gre15a, Gru97, Hei49a, HW78, MH71, MK56, RS96, RSPS49, vL55b, Som49a]. Jahre/Georg [RSPS49].

Jahre/Nobelpreise [Ano61]. Jahre/Otto [RSPS49]. Jahren [Ern92, Hei53a, Ein52a, For80a, Waz09]. Jahrestages [Tre66]. Jahrhundert [Fer14a, Her94a, Her04]. Jahrhunderte [Erb10]. Jahrhundertgenies [Bod17]. Jahrhunderts [Kra88, vM05b, vM09, vM12]. Jahrhundertwende [Yos95]. Jahrhundertwerk [Büh15]. jährigen [Wie06]. jail [Num09]. Jak [RA00]. jako [EI58]. James [Bes04, Bro97, Fel11, Ham11a, Hen11a, Joh11b, Leg15, Mil11a, Nuv11, Set11, VB96, Vis94, BM09, Con17, Gol83b, Gow93, Ham11b, Har02c, Har98, Hen11b, Hen86b, Her82b, Her89b, Her93, Hun10, Kni16, Mah03, Mil11b, Sch12d, Tho31, Tho11a, Tra00, Ebu16]. Jammer [Gor02, Mas01]. Jan [Sta20]. Jane [AA75]. Jänner [Ano21i]. Janssen [Hus15, Pye06, Row00, Sie97, vD15]. Jantzen [Hea13]. January [Ano23a, Bad11, Bad15, Hol05b, Kev05, Rig00, Ano21i, BSKS04a, BSKS04b, BRS^{+}09a, BRS^{+}09b, BIRS12a, BIRS12b]. Japan [But54, Fei61, Fre08, Gia09, Gru05b, Has05, Hen00a, Jay90, Kan84, Low01, Mal08, Mal12, Nis79, OK81, Ska94, VWA95, Wal97, Wal04a, Wel06, Yam09, Lit09]. Japanese [Fra99, Asa95, Bar98b, Dor86, EI40a, Gol98b, Gro49, HW55, Kan87, Oga79]. Jar [Bur15]. je [Ein34a, Ein58, Ein89c, Ein09a, Ein15a, Eis37, Sar35]. Jean [Bro83, Pye95, Pye97, Nye72, Tho88b]. Jeanne [Chi08]. Jeans [Coh31, Eva95, Hol30, Sta07d]. Jed [Pye95]. jedné [Kot93]. Jedno [Pop98, Pop05]. Jefferson [Hol86]. Jeffery [Ken25, Mur25]. Jeffrey [Ebu16]. jego [Inf56a]. jeho [Inf58, Kol72]. Jena [Sch83b]. Jenciklopediĭ [GUB92]. Jennifer [Bus13a, Noe11]. Jeremiah [TH79]. Jeremy [Bad96]. Jeroen [Ken11, Del12]. Jerome [Smo07a]. Jerusalem [Ano84c, EO79, Ano34a, Ein21w, Gil84a, HE82, Kat04, Lar84a, Wil84]. Jesenská [Chi08]. Jessica [Mar91]. Jesus [Bau09, Bau09]. JET [JDB^{+}97, JDB^{+}97]. JET/JET [JDB^{+}97, JDB^{+}97]. Jets [Ano18b]. jetzt [JDB^{+}97]. jeune [Bra15]. Jew [Dys15, Ein41c, GE97, Sto98, Smi05a, Str05a]. jewel [Ven94b]. Jewish [CD07, Ein21m, Ein32c, Ein35a, Ein40c, Ein46b, Ein53o, EO79, Fis92, Gim12, Nac79, NS88, Rei07, SM41, Tal82]. Jews [Ein38c, Gil05, Isa09]. jiang [HL10, LR02]. Jim [Ano84a, CE79a, CE79a]. Jimena [Eng17b, Wil16a]. Jirí [Fis79]. jmenováni [Hav77]. Joan [For80a, Pye80]. Job [Ano87a]. Jody [Wil13b]. Joel [Ree17d]. Joh [Jüt21]. Johann [Ein34m, EVB^{+}72, Jak78, Wil88]. Johanna [Bus03, Ein03a, Hom04]. Johannes [Her66a, Bau51, Her66a]. Johannesburg [Ano22a]. Johanny [Hom04]. John [Bal91, Cah90b, Gly90, Har91, Hes04, Jan73, Lam30, Leg15, Pye89, Pye95, Ram90, She16a, Smi89b, Sta99b, Wac06, Ano87b, ACH^{+}03, Cou79b, Flo08, For01, GW96, Hal17, Hen88b, HD02, Key89a, Mar11c, Pou93, RDS03, Sch90, Sem08, Tau63, Wal72, Whi16, Wig96c, Wig96h, WG96, vN96, Nau17]. Johns [Ham11a, Mil11a]. joie [Bal93, Bal95a, Bal11a]. Joined [Hec16, Ree17c]. Joint [MS82]. Joke [Nad02, Mad89]. jokes [Hes15b]. Joking [FLH85]. Joliot [CCD^{+}52]. Joliot-Curie [CCD^{+}52]. Jonathan [Ber06b]. Jordan [Sch05c, Ein25b, San81b, Sch05c, Sch13b]. Jordans [Ein25b]. Josef [Zau05]. Joseph [Con83, Hes04, Kar05, TH79, Ein54h, Pin00a, War91]. Josephson [LLSS07]. Jost [Hen21]. Journal [Ein55j, Rom91, HW55, MSYW12]. journalist [Bro10b]. journalists [Ill06, Smi07a]. journals [Bey49, Bjø20a, Bjø20b, Bjø20c, SK94, ZXLX14]. Journey [FR13, Lei00, Mah17, TS01, Ven88, Wei06b, Gen22, Mee03, Whe90, YJ95]. Journeys [Col93, Zal91, Aka15, KZS99, WD06]. joy [Bal93, Bal94, Bal95a, Bal99, Bal08, Bal11a, Ehl05a]. Joyce [Ebu16, BM09]. József [Bad15, Pye97, Pye06, Row00, Lac08, Maa13a, Maa13b]. Jr [Hen82b, Kle82a]. Jubilee [CHR90, MK56]. Jubilees [Sei99a]. Jude [Sto98, Vis95, Joh95]. Judges [Ano20d, Dac20a]. Judgment [Nob95]. Judicious [Ham85]. Jüdische [BS92, NS88, SM41]. jüdischen [Ein40c, Ein46b]. Judy [Ste95b]. Juergen [War96b]. Jugend [Hen11c, Rec09, Rec10]. juillet [HLZJ91]. Juilliard [Man80]. Jules [Ano84a, CE79a, CE79a]. Juli [B^{+}79, Ein31a, GHK^{+}96, MK56]. Julian [MM00]. July [B^{+}79, BL09, CR07, Ein31a, GRRS99b, GHK^{+}96, Ham97, JRG02, Rig00, Bad12, Sch83b, AMS04, Ano00g, MB66, Ree06, SS65, Sza84, Wei95]. July/August [Ham97, Wei95]. Jumps [Sch52b, Sch52a]. Jun [HY86]. June [B^{+}79, Pla20b, BIJ^{+}18, BRI^{+}21a]. Jung [APWB95, MGG^{+}95, APWB95, Gie05, JP55, Lin04a, Mil10b, vE95]. Jung-Dialog [APWB95]. Jungnickel [Ser88]. Juni [B^{+}79, HH14, Jüt21, Pla20b]. Jurgen [Har96a, Pye80, Pye95, Bal96, Cah90b, Noe93, Pye89, She16a, Smo07a, Sch00c]. Jury [Yea08, Ben99, Cre06]. Just [Bor66, EMH96, Kox93, Pye89, Ree00, Rös95, Sta07b, Wal77b, Wal92d, Wal06e, Wal15, Bon81, DL15, Lah20, NHG89]. Justice [Bal86, Moo16]. Justification [Ein05y, Sta82b].

K. [Sig92, Sta95b]. Kōen [Ish77]. Kōen-Roku [Ish77]. Kaempfe [LS05]. Kafka [Kun76]. kaheli [Hot15]. Kahler [Ber94]. Kaila [Sla22]. Kaiser [Kan96a, vBL96a, vBL96b, GC20, HSW09, Kan96a]. Kaiser-Wilhelm- [vBL96a, vBL96b]. Kaiser-Wilhelm-/Max-Planck-Gesellschaft [vBL96a, vBL96b]. Kaiser-Wilhelm-Institut [Kan96a]. Kaiserslautern [B^{+}79]. Kaist [DWL98]. Kaku [Kar05]. Kalender [Lin17, RR17]. Kaluza [EG23, Ein27f, EB38, vD02a]. Kaluzas [Ein27f]. kam [Bod03]. Kamerlingh [RA13]. Kampen [Hau49, Hau53, Hau63, Hau64, Hau65, Hau83]. Kampf [Fer14a]. Kampfe [Kra20c]. Kanalstrahlen [Ein26b]. Kanalweite [EM23]. Kandel [SK09]. Kann [Kie15, How90b, Mau96, Flü39, Gio13a]. Kansas [KK05, KS12]. Kant [FTG08, Rei41, Ano23d, Els24, FTG08, How94b, Mon04, Ryc14, Wan95]. Kantian [D’A97]. Kapitza [Bad85]. Kärger [ARK^{+}99]. Karl [Bla86, Ein16g, Eis00, Erm89, Ein16g, RSPS49, Sch05b, Sch07b]. Karlfeldt [Per09]. Karlovy [Hav80]. Karlskoga [BCS82]. karne [Seg86a]. Karolinska [LB02]. Kart [Kai81, Wil36]. kastes [Jak13]. Katastrophe [Wal05a]. Kathlerine [Duf46]. Kathryn [You95]. Kaum [Hof08b]. Kauntodaun [WY05]. kavramlardan [EI94, EI11]. Kayser [Rob48]. Keenly [Jam30]. Keep [Ano45a]. keepers [Van03]. Keeping [Bal05, Koe95, Gla07, JDB^{+}97]. kefi [Ayn34]. kehitys [EI62b]. Keine [Föl90]. Kelvin [Ein24k, YCKG11]. Kelvins [Ein24k]. Kennefick [GRC18, Nic20, Pye06]. Kenneth [Hie84, Ree15d]. Kepler [Bre15, JP52, Top89, Bau51, Bre15, Ein30p, EVB^{+}72, Fie88, Hal04b, Hol73, Hol88d, JP55, Wil89]. kept [NG07, Pur63]. Kerle [Jun05a, Row06a, Row06b, Wue05]. kernel [Gie05]. Kernenergie [WW75, Mei63b]. Kernfusion [MTVB88]. Kernfusion/Nobelpreis [MTVB88]. kernphysikalischen [Sta98d]. Kernspaltung [BH01, CSW97a, Kra88, BH01]. Kernwaffendetonation [GHK^{+}96]. Kerr [Mar08]. Keswani [Din65]. Kettenreaktion [Hah48]. Ketterle [Lev01a]. Key [Ano14a, Cle17, Ken54, Lau55b, Arn00, Bre15, Kil94, ZZY19, LB02]. keyfi [Bal08]. Keyhole [Gut94]. Keynes [Gal94, Key89a]. kg [Jin96]. Khalatnikov [Hoh13]. Khariton [KS93]. Kholodinin [Hen82b]. Kids [RR17]. Kiel [LS05, SW12]. Kieler [LS05]. Kilby [Pin00a]. Kiley [Wal72]. Killer [Kle93]. Kim [Vog20]. kind [How94a]. Kinematics [Mar09, Wes80]. Kinetic [Ein05-35, Ein05c, vS06, Ein02a, Ein05w, Gar70]. Kinetische [Ein02a, Ein05w]. kinetischen [vS06]. king [Pau42, Pau47a, Ano21j, Ano05b, Ein21o]. kinlige [EI11]. Kinnon [Hen82b]. Kip [Har95, Vog21]. Kipling [Ano07q, BL07b]. Kirche [MGG^{+}95]. Kirchhoff [Sch03a]. Kirschner [OSWR86]. Kirsten [Hor79b, Pye80]. kiss [Col17]. kisses [Seg95]. Kitchen [Sul13]. kittens [Gri95b, Gri95a]. Klassenkameraden [RK86]. Klassiker [Bes04]. klassischen [Bäc19, Bäc21, Inf31, Ren97b]. Klaus [Pye95, Row00, Smi02, The12, Goo03a, Kri08, Lau12, Mos87a, Mos87b, Pye93, Wil87]. Klees [Ing05, Ing05]. Klein [Bal96, Har96a, Har96b, Hie68, Kil95, Mic68, Pye95, Pye97, Sie97, Wal99, War95, War96a, War96b, Hal07, KS95, Tob94, vD02a]. kleiner [Ein08a]. kleinsten [Fis10, Fis12]. Klimawandel [Bre08]. klügsten [BT15]. Kluwer [Wal02b]. Knee [Sha92]. Kniga [GUB92]. knives [Bel05]. Knobloch [MGG^{+}95]. Knot [JR07]. Know [Jan07, PP05, Ber08, Jag01, Mor79a, MHM90, Mui94, Mul02, Mul10, PP07, PP06]. Knowing [Wig96l, Wig80]. Knowledge [AKD23, Boh58, Boh10, ETP31, Ein44e, Ein59f, Len49, LHDH21, Oga79, Pai88, RR65, Wal16, Wei12a, Ano81e, Cas03b, Cha09, EI58, Els24, Fre18, Göt92, GP11, Lam79, Ric06b]. Known [Ano54b, CW12, Kel06b, Sch19]. knows [NHG89]. Knut [HM11, Hau48]. Koby [Hen16, Nav15]. Koch [HEB^{+}80]. Kochen [Gar72]. Koinzidenz [Gio13a]. Kollegen [Hof08c, RKW05]. Kollegiums [SM41]. kom [RE90]. kommentiert [Win69]. Kommentierung [BH01, Ern92]. kommt [Ehl05a]. Komödie [Dür62, Dür80]. Kompatibilität [Ein30f]. Kompatible [EM31b]. Komplexstruktur [Pau25]. Komptonsche [Ein24d]. Kondensat [JDB^{+}97]. Konfliktlösungen [BHJ^{+}00]. königl [Pla20a]. Königlich [Jüt21, Pla20b]. können [BT15, EM31b]. Konrad [Lut05]. Konstanz [SW94]. Konstitution [Ein09b, Ein09c, Ein19e]. Konstruktion [Hün18]. Konstruktivismus [Pes13, Sie10]. Kontext [Ren05e]. Kontexte [RCD98].

Kontroverse [Waz09]. Kontur [Inf69]. Kopenhagener [HB63]. Kopernika [Inf55d]. Kopernikanische [Foc53b]. Kopernikus [Uns64, vL56c, vL56d]. Köpfen [BT15]. Kopff [Ano21k]. Korea [Ric06a]. Korespondence [Kot93]. Kormos [Bad15, Lac08, Mar11c, Pye06]. Koronaspectrums [Ein23a]. Körpen [Ein11g]. Körper [EL08c, EL09, Ein18j, Ein05-41, EL05c, EL05d, vI10b, vI11a, EL05b, EL05a, Ein05e, EL08b, Ein20q]. Körper“ [EL08a]. Körpern [Ein05q, Ein05t, Ein11h]. Körpers [Ein05b, Ein05v, Fra37]. Korrespondenz [Hen86c, HRP90]. kosmische [Bre08, Jae06, Ste05c]. kosmologiczna [Gol98a]. Kosmologie [Sch09a, Ost15]. Kosmologische [Ein17e]. Kosmologischen [Ein22d, Ein31o, Ein19e, Ein05h]. Kosmos [Sch09a]. kosmosa [Lan67a]. Kostas [ARS15]. Kottler [Ein05-38]. Kottlers [Ein16n, Ein05-38, Ein16n]. Kovarienten [Ein18d]. Kovarienzeigenschaften [EG14]. Kowarski [Erm89]. Kox [Bal96, Har96a, Har96b, Kil95, Pye95, Pye97, Pye06, Row00, Sie97, Wal99, War95, War96a, War96b]. Krafft [Ros86]. Kräfte [EL05d, Bac05, EL08c]. Kragh [Bro00]. Krajewski [Flo06]. Kramers [Dre87]. Kraus [GRE^{+}01, Urb22]. kreativitet [Lar01]. kreds [RRK^{+}64]. Kreiselkompass [LS05]. Kremlin [Bad85]. Krieg [EF33, EFA72, Ein15h, Ein16k]. Krieges [Nic17, Nic19]. Krise [Yos95]. Krisis [MM66]. Kritik [Fra37, Hah81]. Kritiken [Sch10a]. kritische [Cun22b, Urb22]. kritischen [Ein10a, Ein05-29]. Kritisches [Ein18i]. Kritizistische [Sch21]. Kron [Hof49b]. Kronprinzenpalais [Ren05e]. Krümmung [Ein22c, Ein23f]. Krümmungstensors [Ein27a]. Krupp [Man64, Man68a, Man68b, Man70, Man03]. kuantumlara [EI94, EI11]. Kugelförmigen [Ein21e]. Kugeltheorie [Wei12y]. Kuhn [BHJ^{+}00, Obe16]. Különbözo [von90]. Különleges [Ein22a]. Kultur [SE79]. Kulturbesitz [B^{+}79]. Kumar [Sud11]. künftige [HBFH18]. Kungl. [Hil39]. Kunst [Hof03]. Künste [Kie82]. künstlerischen [Ein21c]. Kunz [HEB^{+}80]. Kurian [Bar05f]. Kurt [BPP^{+}11, You91]. Kurylo [Bri82]. kurze [Vit15]. kurzgefaßt [JDB^{+}97]. Kurzsichtiger [Ein52e]. Kuwaki [Tan92]. kvantemekanik [EI61a, EI39a]. kvantne [EI62c]. kvantov [EI66a]. kvantteorierna [EI38d]. kvanttiteoriaan [EI62b]. kvarkim [Seg86a]. kwantów [EI59, EI62a, EI98b]. kwanty [Inf33]. Kyi [Joh11a]. Kyoto [HY86, Abi00a, HY86, Ita99].

L [Bad72, Bel26, Fre08, Lit09, Mal08, Nav94, Nye88, Pul11, Ram82, Rob38, SO72, Tre88, GUB92]. L. [Duf81, Hen82a]. lî [Ayn34]. lab [Mac85]. Labor [SMW^{+}09, Lau50, Olw97, Olw04]. Laboratories [Ano77, Ano78, Sei99a]. Laboratory [Ano81d, Swa02, Bru76, Bus12c, Duc11, HSW81, HS89a, Spe05, Ahe03, Car04a, HS89a]. Labour [SG82]. Labs [Das00]. Labyrinth [Hal17, How93a]. Lack [Bro21]. ladder [Web99]. Ladies [Ros12]. Ladung [JDB^{+}97]. Lady [Opf78]. Laennec [Tho88b]. Lagerkvist [Coo04, Kau14, Ano07r, BL07d]. Lagrangian [Lan49]. laid [Ein10e]. Lajos [JS87]. Lakatos [CFW76]. Lake [Bad12]. L’Allemagne [Gou48]. Lambda [Ear01]. Laminar [LB49]. Lämmel [Wei15c]. Lampa [Kle75b, Kle75b]. Lanczos [Bla76a, Bla76b, Bro66, BCEP94, Hou65, Jam70, Mal65, See65b, Wre66, Bal75a, Bri66, Rin09, Rin11b, Rin11a, See65a, Sta94]. land [Pas10, Elz95, RB05]. Landau [Hoh13, Kha12]. landing [Dac11]. Landmark [ZXLX14, Pra06]. Landscape [Sus06, ZZ23]. landscapes [Goi91]. Langenscheidts [GRE^{+}01]. Langevin [FF97, Ano79b, Ein79a, FF97, LG97, Wei12q]. langsame [HS39b]. Language [Bel23, Ein41b, Ein42, Thi21, Ari03, Ari05, Egd14, FS01, Gor22, HBO82, Jak82, MSYW12, Rec10, SW94, Ano22a]. languages [BM09]. Lapin [Mar96b]. lapse [ACT23]. Large [Ber89, Her46b, Her85b, Car02, HE08, Myr21a, Pap15, PSC^{+}97, Acz10, Lin09]. large-scale [Myr21a]. larger [Whe79c]. Largest [GK05]. Larmor [War91]. Lars [She15]. Larson [Erm89]. Larsson [For05, Mau09]. Laser [Mar22, OSWR86, Phi98, Raf17, WTB09, JDB^{+}97, Ket02]. Laserkühlung [JDB^{+}97]. Laserlicht [JDB^{+}97]. lasers [HM05b]. Lassen [Ein18j]. lässt [Ein20g]. Last [Ano11a, Cal03a, Day12, Ham97, Lau50, Lei00, Lev98, RE90, MNOW00, Sal05, ST94b, Sch04b, Wei95]. Late [DWL98, Pye82a, Mal00, Pye82b, Sow01]. lateinisch [Die24]. Latent [AKD23]. Later [C.51b, Cou52, Ein40e, Ein05-45, Inf50b, Jam74c, Ein50o, Ein50p, Ein52a, Ein56h, Ein70a, Ein71, Ein89d, Ein93c, GTP22, Hei53a, Mad95b, RR89, RMR96, Sac85, San07, VB96]. Latest [Ano98, Bar00b]. Latin [Die24]. Latitude [DR89]. lattice [Abr04c]. lattices [MO06]. Laucht [The12]. lauds [Ano29a]. Laue [Ano79h, Ano79g, B^{+}79, For80a, Hen21, Kan96a, Kra25, Lam79, Lin60, RS13, Ano79g, B^{+}79, Ein15b, EHM^{+}79, Ein05c, FPLR59, Her79a, Kan96a, Kra25, Lam79, RS80b, RS13, ST94a]. Laue-Furchtlos [Hen21]. Laue/Einstein [Ano79g]. Laue/Einstein-Symposion [Ano79g]. Laues [Ein15b, Ein05c]. laughter [Sta92a]. Launched [Joh06, Cha05c]. launching [Gib91a]. Laura [Pye95]. Laureate [CGG^{+}13, Car15, Hei65b, Joh06, SS98c, Bad12, Smi05a, Str05a, Zel90, Ano64, Ano65b, Ano72, Bor78b, Car12, Cha05c, Don70, Fei01, FM92, Gha82, HF22, KKJ99, Lun92, Nel77, Nob64b, Nob66a, Nob66b, Per97, Put09a, Rog06, SS98a, SS97, Sin04a, SR04, TA97, Tan11, Ano77, Ano12b, BR00, Bor75, HFN06, Kov01, Sto98, Tis52, And81, Far79, Han80a, Hei79a, Sch79c, Ste95b, Hei79b]. Laureaten [Hof18]. Laureates [Ano07p, Ano07q, Ano07r, Ano07s, Bar05f, BL07a, BL07d, BL07c, Bru06, Cam09b, Dar04a, HBFH18, Har79d, Nad79, Opf78, Sch87a, Sei79, Tra00, Vis94, Zuc77, Zuc96, Zuc18, Abr88, Abr01b, ABH^{+}55, ABH^{+}87, Ano01a, Ano07n, Ano07o, Bad04, Bie03, BS86, BH04, BH09, BL07b, But23, CÖT15, CT15, CJT22, CZPR17, CD23, DCL23, ET87, Fie15, FMR90, Hal04a, Hen13, Hof18, Hoi01, Hor09, Jam93, KKKM05, Kat89, Kos20, McC01a, Pri90, SBB16, Sin75, SSH13, The95, Tho88a, VM05a, Was92, Wir08, Nob72, Ang78, Ano78, Ano82a, BZ10, Bal79, Bun79a, CK^{+}04, EGR11, Fos13, Her66a, Kee04, Nacxx, Ng11, NB79, Orr78, Pau79, Sta79, ZZY19, vS78]. Laureaty [GUB92, Was92, Dai67]. Laurence [Ano60]. Lauritsen [Hol03a]. Laves [Lin60]. Law [Ame20, Ano29d, Ano07l, Bor07, Bos24a, Ein02a, Ein07a, Ein12a, Ein12f, Ein21e, Ein05f, Ein05w, Ein05-30, Fey94, Lau55a, Pla00b, Pla01, RFP^{+}45, AHH05, EMvL79, Eva95, FMM96, Kin95, Rus91, Sch03a, Ung07, Woi06, vEPF09, vEPF22, AH60, Ein11d, Ein12i, Ein12h, Ein13a, Ein18e, Ein26c, EG27a, EG27b, Wal02b, Ein05y, Ein05d]. lawn [Gef14]. Lawrence [Ano81d, Ber03, Her03, Jen01, Sch03c, HSW81, HS89a, Her02, Sei99a]. Laws [Ano29b, Ber06b, FvN78, Gre11, MC80, Wig64b, Wig64c, Koc01, Pen05, PKLT17, Pic08, Pit16, Ray92, Str97b, Wig64a, Wig65, Wig67a]. Lawson [Tay21, Tre21]. Lax [RS06]. Layer [LB49]. Laylin [Tra00, Vis94]. Layman [Ein60c, Gam42a, Gam42b, Pai94]. Laymen [Bro23, Gam42a]. layperson [Cal05b]. lbert [BSW22]. LD [Tho00]. leader [Mul10]. Leaders [Lau51]. leadership [dGK11]. leading [Bro06, Bro07, Cro01, Ein14l, Log96, MDT^{+}03]. League [Ein24f, Can05, Ein53l]. leap [Fis10, Fis12]. Leaps [Dys16]. learn [BT15, Pri10, BF02]. learned [Gio13a, KSG63]. Learning [Bru09, Ein36d, Pri10]. Leave [Ano50c, Ano20b]. Leaves [AH60]. Leben [ARK^{+}99, Bor75, Cle17, Cur38c, Cur38d, Cur38e, For80a, Fra49b, Hah88, Hen11c, Her33, Her63a, Her67, Her75, HC94a, HC96a, HC04, Inf99, Kra81, MGG^{+}95, Rec09, Rec10, Ren05e, Ros86, See60, T– D82a, T– D83, T– D85, T– D88, T– D93, Bri05b, Hah68, Inf69]. Lebensbild [WE24, WE87]. Lebenserinnerungen [BR00]. Lebensfreude [Ehl05a]. Lebenslauf [Pla20a]. Lebensweges [RD05]. Lebenswerk [Jor69]. Lebenszeit [JDB^{+}97]. Leccia [Mar73]. Lechalas [MC80]. Lecons [MP74, Pia31]. Lecture [Aka15, Ano20e, Ano21e, Ano32, Bar18a, Boh65b, Cun22b, Ein20p, Eng14, GO08, Hei95, Hig14, Jam30, JW06, Nam10, Tho18, Wei18b, Wig64a, ’t 02, Ano11b, Ano21i, Ano25b, Bet68b, Boh85b, Dir60b, Ein14a, Ein21o, Ein08b, Fer39, Fer70, Fre18, HY86, HSD34a, Kol40, ST94b, Tow65, Wig64b, Wig64c, Wig67a, Abr04c, Alf01, Boy10, Chu98a, CT98, Com46a, CW02, Dav03, Dys09, Ein10b, Fer08, Gei11, Gen22, Gia03, Gin04, Gla06, Gro05, Grü08, Hal06, Hän06, Har13a, Hee01, Hol86, Ita99, Kao10, Ket02, Kob09, Koh99, Kos03, Kro01, Lau99, Leg04, Mac01, Mas09, Mat07, Mil24, Mou19, Nam09, Nov11, Pau47b, Per12, Phi98, Pla20a, Pla20b, Pol05, Pop99, Rav97, Rie12, Sch12e, Shi01, Smi10, Smo07c, Sto99, Str19, Tsu99, Vel00, Wig65, Wil05a]. Lecture [Win13a, ’t 00, BB55]. Lectures [Ano21h, Ano21j, Ano22a, Ano23a, Bri14, Cah04, CCG93, Dir01, Don70, Duf81, FGL^{+}06, Hen82a, Her76c, Her21, Ihd74, Ish23, Kle82a, Lor31, MP74, Nor15, SS98c, SS98b, She15, SM41, Wey18, Ano64, Ano65b, Ano72, Ano07n, Ano07m, Bal77, B^{+}06, Ein17d, Ein21l, EA22, Ein22p, EA23, Ein55i, Ein64b, FLS65, FLS89, FLS06, FM92, FB24a, GR17, Hen82b, HM05b, Kah69, Lan65, Lan67a, Lor14, Lun92, Nob72, Nob64a, Nob64b, Nob66a, Nob66b, Pau61a, Per97, Pet08, SS98a, Sch28, Sch82a, Sch03b, SR92, Sim06, Tor03, Ein99, Wal92c, UKKP04, Bus12b, Ein06d, Gil19]. led [Nor14]. Lee [Ano87b, Gor22, KP55, SH01]. Leffler [Stu10]. Left [Kan92b, BS83a, BS93, Inf78b, CP15]. Legacies [Joh07, She03]. Legacy [Ano05-43, Ano15h, BB04b, GHS08, GHS18, Ger07, Jun05b, Kel04, Wit82, Woo06, You05, Ano97d, Bat03, BB04a, Bro06, Bro07, Cle17, CB99, DB09, Ger92, Ger97, Ger02, Hof10, Lac11, Obe16, Pot11, Ree15b, San81b, Sch86, Sch02c, Sri99, Tho93, Tho94, Wad05, Wad07, YGKK05, Har95]. legalize [Gio14]. Legend [OM18, Ste05a]. Legende [Ste05a]. Legends [WSG00]. Legitimating [Sem08]. Lehce [Ein23h]. Lehmann [HEB^{+}80]. Lehner [Hus15, Pye06, vD15]. Lehrbuch [HEB^{+}80]. Lehre [Stu14, Stu23, vL23, MM66, Rei24, Rei27, Rei28a, Rei28b, RK77]. Lehrpfad [Ing05]. Lehrsätze [Büc21]. Lehrstuhl [AH93]. Lehrstuhle [Kle75b]. Leibniz [Erb10, Noe11, Coo10, Coo15, EP06, Erb10, Gio13a, LMOS15]. Leibniz-Äquivalenz [Gio13a]. Leibnizschen [LMOS15]. Leibniztag [EP06]. Leichen [Mar17]. leichtfasslich [Ang20]. Leiden [Ein20b, Ein20b, Lor31, Uhl80, van06, van12]. leiderschap [dGK11]. Leinen [Hei53a]. Leipzig [Jüt21, KW93, Wil36, RKW05]. Leipziger [RKW05]. Leistung [Ber69b]. leitende [Ein14l]. Leitungsphänomene [S^{+}57]. lektsii [Lan67a]. Lemaître [Ber84a, Kra18, Lum11, Sto96, Far05c]. LeMay [Kee11]. Lemmerich [Hen21]. L’Empire [Lam07]. Lenard [Sta36, Gun15, HEWW15, Sta36]. Length [Dor68, Hag08, Hag14, vB19]. Lenihan [Ram82]. Lens [Ein36e, Cha80]. Lens-like [Ein36e]. Lenses [RB81, Tur88]. Lensing [Ben05, RSS97, RS03a]. Lenz [Boz14]. Leo [Bes93, You95, BST03, BSST04, CST01, CST02, Ein19h, EE72, FS87, Fra05, Gra94a, Gra96, HGS87, LS92, LS94, SE45, WS78, Wig96i]. Léon [CS79a]. Leopold [Ano39d, Ano50a, Arm38, Bro39, C.51a, Kae38, Len39, N.38, Nag50, Qui40, Ano68a, Bič79c, Sch50a]. Leopoldina [HBFH18, HH14, HBFH18]. Leopoldina-Symposium [HH14]. Leopoldo [dMdA05]. L’eredità [Bat03]. lernen [BT15, Gio13a]. Leroy [Ano84a, CE79a, CE79a]. Leserbriefe [TF99]. Leslie [Nor02b, Nor03a, Erm89]. Lesnické [Pát79a]. Less [Ano19b, She16a]. Lessons [Ano20a]. Lestingi [Sch78]. Let [Cze05, Ano04a]. Lethal [D’A93, Ful84b, Ful84a]. L’éther [Ein21k]. Lett [GHK^{+}07]. Letter [Ano48a, Ano95c, Bar18c, Bar18b, Bel98, Ben79, Ein84b, Don64, Ein24f, Ein33d, Ein39a, Ein48e, Ein49c, Ein50n, Ein53f, Ein63a, Ein79e, Ein85e, Ein85c, Ein95a, Est06c, GT38, GM06, Han92, Han97, HM97, Hoo77, Ill89a, Inf50c, Kno90, LA88, Lau55b, PSM^{+}13, RMI21b, Ree05a, Ree11a, Ree16c, Roo39, VFIS48, Wal91b, Wig96j, Zan93, Ano57b, Büc21, Dem15, DGG^{+}21, Ein47c, Ill89c, Jea08, KS12, Kot93, RMI21a, SE79, WW86, ES68, Wac06]. letteratura [Tio09]. letterature [Cas14]. Letters [Ber94, BHM91, Bre73, Bro72, Cal02, CE79b, CE79a, DFB22, Ein91d, Ein97, Fey05, For80a, Gen98, Grü90, KK75, Kan92a, Kil80, KRW05, LRD^{+}91, MNT^{+}00, MHM90, Nev92, Pop98, Pop05, Prz67, Prz86, Prz11, Ren97a, RFP^{+}45, Rol10, Sch93a, Sch92b, SW95, Smi96, Som05b, Ano84a, Bau51, BB71a, BB71b, BB05, Deb29, Deb79, Ein31c, ES68, ECS02, Eve39, Eve13, Fre96, Har02c, IK03, Kra25, Pop03, RS92, Sal74, San77, Smi04, Tou05, TF99, Prz63, Som05a, Noe93, Hie68, Mic68, Sch92c]. Lettre [Ano57b, Lucxx]. Lettres [CE79a, ES56, Ein05-47, Sol56, Spe62]. lettura [Arg09]. lety [Těš12]. letzter [ST94b]. leurs [Pia31, Seg84]. Level [Sho21, Vog20, Vog21, Bak97, Eis86, Eöt00, Bus12b, Bus12a, Mil12b, Myr21b, Myr21c, She16b, Sud11]. leven [Cur38b, HC94b, HC96b, vC05a]. Levenson [Rus17]. Levi [Lam30, CT24, Goo75, Wei12o]. Levi-Civita [Lam30, CT24, Goo75, Wei12o]. L’Évidence [Ano23a]. Levinas [Gra01]. L’evoluzione [EI48b]. Lew [Erm89]. Lewin [Kri99]. Lewis [For80a, HOO88, Ram79a, Seg91]. L’expérience [Ein21i]. LHC [AGMT12]. L’homme [BTM84]. LI [HLZJ91, Seg86b]. Liangying [Bod13]. Liberalism [B.50, Rei50, Ano47c]. Liberate [Ano39g]. Liberated [Ano39a, vHJK39c, Bra14b, vHJK39a]. Liberation [vHJK39b, Ost15]. Libet [Sta06]. Library [Ano01b, B^{+}79, EO79, Hei53a, Hie68, Jon89, Krö53, New56, New60, Pin00b, Ano81a, Egg02]. libre [Rol71]. licei [Pal99]. Licht [MTVB88, vC05a]. Lichtabsorption [Ein05-43, Ein06f]. Lichte [Cun22a, Foc53b]. Lichtemission [Ein21s]. Lichtenberg [RSPS49]. Lichterzeugung [Ein06f, Ein05-43]. Lichtes [Ein05d, Ein14f, Ein05-31, Ein05-37, Ein11i, Ein26b]. Lichtfortpflanzung [Ein22n]. Lichtgeschwindigkeit [Ein27c, Ein12e, Ein05x]. Lichtquanten [Wie06, vL55b, vL55c]. Lichtquantenhypothese [Bos24a]. Lichtquantum [Bor55a]. Lichtstrahls [vS21a, vS21b]. Liebe [Ein03a]. Lieber [Hea13]. Lies [Smo19b, Tur03]. Lieserl [Zac99b, Ivr99]. liest [Ein52e]. Life [And81, Asi83a, Bal72, Ber87b, Bey06, Bor75, Bri97, Bri05b, Bri82, Cas05b, D’A01b, Del09b, Ehl79, Ein32c, Ein91g, Far79, Far09, Fer82c, For84b, Fra49b, Fri79a, Gil84b, Gil12, Gle92, Gol84a, Gor07a, Hah68, Han80a, Hay91, Hei79a, Hei79b, Hei83a, Hes04, Hil03, Imp19, Kae13b, Kle84, Kra11, Kri99, Lak85, Lar84b, Lau55a, Lov86, MGG^{+}95, Mer01, Mil13a, Mil15, Moo96b, Nav94, Pei83, Pip83, Pol83, Rec10, Red84a, Ree17d, Ren05e, Roe93, Sch79c, Sch07c, See60, Smo07a, Sta99b, Str04, TVL^{+}09a, Tou05, War84, Weg08, Wil00, Win93, Win99, Ano04d, Bau51, Bet03, BO76, Bor78b, Bri96, BS97, But23, Cas92, Cla71, Cla73, Cla84, Cle17, Clo15, Cro01, Cur38b, Cur38c, Cur38d, Cur38e, De 04, De 05, Doh06, Ehl05a, Ein59a, Eve39, Eve13, Fer54]. life [Fer82a, Fer87, Fra47, Fra48, Fra53, Fra65, Fra88, Fra91, Fra02, Fre86, Fre96, Ger69, GH84, Gir86, Gle10, GGW94, GB06, GB08, GO81, Gre05b, G^{+}11, Hah70, Hah88, HM11, Her33, Her63a, Her67, Her75, HD02, Hoy94, Inf99, Isa07, JDB^{+}97, Jor69, Kan96b, Kra81, KS81, Liv13, Mag09, Mah03, Mil10a, Mil12a, Mit05, Mof10, Moo89, Moo94, MO95, Nacxx, Nor10, Pai82, PC06, Par11, Pop03, Qui95, Qui96, Red84b, Rei34, RD05, RRK^{+}64, Sch44, Sch45, Sch56d, Sch56b, Sch56c, Sch62a, Sch67, Sch92a, Shu17, Sim96, Smi04, Tal82, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D07, T– D12, Tym95, Ven88, Wei13e, WF98, Whi93, Whi94, WG94, WG95, WE87, Zew02, vC05a, vF79]. Life-Long [D’A01b]. lifespan [CJT22]. Lifetime [Gin01]. Light [Bos24a, Boy85, Boy94, Bus13a, CR12, Coh40, Cro20a, Dor44, Edd19, Ein22n, Ein05-31, Ein05x, Eis93, Eks99, Fel11, GB89, Gin22, Gra01, Ham11a, Ham11b, Hen11a, Hen11b, HG06a, Jan95, Joh11b, Kas20, Kel14, Kia46, LU13, Leg15, Mal09, MTVB88, Mil11a, Mil11b, Mil24, Nuv11, Pet10, Ree06, Rya06, Sau21, Sch95a, Set11, Ste98, Wei12m, Wei13b, Wig96j, AAA^{+}09, Abi03, Aka15, Ano06f, Ant15, Bai92, Bai01, Bor55a, Car13, DGG^{+}21, DED20, Ein05d, Ein06f, Ein11i, Ein12e, Ein14f, Ein21s, Ein26b, Ein27c, Ein29a, Ein36e, Ein05-37, Ein05-43, Ezr82, Fei70, Fey06, Foc53b, FSE31, Gla06, HS51, Hun10, Jak78, JDB^{+}97, Les74, Lou83, Mag03, Mou19, Nac73, Nac77, Pep05, Pla59, Pri12a, Sch96b, Sie91, Sin04a, vS21a, vS21b, SSS^{+}98, Stu14, Ven88, Wie06]. light [WW86, Wil88, vC05a, vK60, vL55b, vL55c]. light-deflection [vK60]. Light-Quant [Mil24]. Light-Quantum [Bos24a]. light-velocity [Abi03]. Lighton [Flo08]. Lights [Ano19b, Kae13b]. lignes [Edd21]. LIGO [Ao04, Bar18a, Tho18, Wei18b]. Lihotzky [Chi08]. Like [DR08, Ein52e, Gei15, Ste98, Ano30b, BT15, Ein36e, Jan14b, KRZ^{+}05, MDT^{+}03, Smi15a, Smi15b, Tho15, Tör81, Gio20]. Lilienthal [Res50]. Liljestrand [Kat52]. Lillian [Hea13, Hes04]. limelight [Smi21]. Limerick [Ano99e]. limit [AAA^{+}09, Ein05-32, Mas74]. Limitations [Ray71]. limité [Kle74]. limited [Kle74, Rus85a, Wei12a, Bad95]. Limits [Ein22g, BRS03, Cas03b, Ein07i, Hey99, Roe93]. Lindau [Hei93, Bad12, Hei75b, Hei93]. Lindauer [Hei75b]. Lindemann [Fox15]. Lindley [Fer07a, For09, LSF94]. Lindner [MGG^{+}95]. Line [Gon05, How98, vS21a, vS21b, Wil72]. Linear [BB49]. Linearized [FS98]. Linen [Hei53a]. Linenthal [Ham97]. Lines [Chr20, DR20, Lai20, Lod20a, Lod20b, Mlo01, Sil36, Nac24]. lingering [Ryc92]. Linguerri [Per10]. linguistic [Gor22]. Link [RMI21b, RMI21a]. Links [Ano29d, BIT03]. L’interprétation [BB55]. Liquid [Ano94b, Ein05-29, GR44, Lon38a, MBD39, Stu94, BP79, Ein10a]. Liquid-Drop [Ano94b, Stu94]. Liquides [HLZJ91]. Liquids [Ein05-29, HLZJ91, Ein05c, Ein10a]. Lisbon [Ren21]. Lise [Ano38a, Ano69a, Ano79h, Ano79g, B^{+}79, Büh98b, Chi08, CSW97a, For80a, Lam79, Lin60, TF99, Ano69a, Ano79g, B^{+}79, Büh98b, CSW97a, CSW97b, CSW98, EHM^{+}79, Ern92, FPLR59, Fri68, Her79a, Kac96, Kri99, Lam79, McL99, Mei64a, Mei64b, Mei05, Sim90, Sim94, Sim96, Wat97a, Kri99, TF99]. List [Ano82a, Ein48b, Est06a, KSB^{+}09, Sch49e, WK12]. listen [Won05]. listening [Bar00c]. listens [But23]. lists [BSPKN03]. Literary [Ano07p, Ano07q, Ano07r, Ano07s, FK68, WK12]. Literature [All97, Can00, Eng08, For01, FA93a, FA93b, Fre69, Hen11a, Jon89, Leg15, Wis82, AE06, AKD23, AE01, Ano07n, Ano07o, BL07b, Car12, Cas14, ET87, Gyl78, HKS10, HG07, HT10, New56, New60, Pri90, Sin75, TA97, Thi46, Thr17, Tio09, Whi01, Wir08, Ano45b, Ano07m, Ano07p, Ano07q, Ano07r, Ano07s, BL07a, BL07d, BL07c, Car15, Dan67, Hal12]. Literaturnachweise [Egg02]. littérature [Dan67]. Little [Fri79b, Maa07, IK03, Kel06b, Mad89, Mar91, SU06, SWU09, CM02b]. little-known [Kel06b]. Liu [HL10]. liv [RRK^{+}64]. livable [HGS87]. live [Mos68]. lived [Kle16, Pai94]. Lives [Ald93, Ber03, Bon94, BS86, BH04, BH09, Bur94, Cas94a, Cas94b, Con70, Din93, Gea94, Her03, Joh07, Kan94a, Mil93, Moo16, Rup09, Sch03c, Smo07a, Unr99, Wat00, Ano97d, Ano04e, Bre97, CH98, De 13, Gab95, Gar93, Her02, HC93, HC94a, HC94d, HC94b, HC94c, HC95, HC96a, HC96b, HC96c, HC98, HC04, LB90, McG93, McG98, Sch87a]. Living [Gar08, Krö53, MKBF30a, MKBF30b, Row85, War99, Hal97, Mar00, Sch44, Sch45, Sch62a, Sch67, Sch92a, Tho88a, Inf69]. livre [Ano05o, Fre79d]. Ljubavni [Sre29]. Lloyd [Coh54]. lobbying [Har96c]. Local [CHSH69, CHSH70, SZ88, MS82, Sch13b, Sel88, tH20]. localisation [Ein10f]. localité [Gis12]. Locality [But07, CW12, How85, Sta09b, GS80, Hel82b, SS79, Sel78]. localization [Ein10f, RDF^{+}08]. Localized [NW49]. Löcher [Bes04, Ber96b, Gen05, Str05b]. Lodestone [MKBF30a, MKBF30b]. L’Oeuvre [dB49a]. loft [Som05a, Som05b]. Logic [Bub79, Sch74, BBL08, Pop07, Has07b, SW94]. Logical [CW74, Fra49a, Hol89b, How94b, Wil77, GR96, Gio13a, Gio13b, Ryc92, Sch75, Str97b, Rös95]. logical-empirical [Gio13a]. logicians [D’A97]. Logics [LU13]. logique [Pop07]. Logisch [Gio13a]. Logisch-Empiristischen [Gio13a]. logische [Str97b]. loi [Ein13a]. London [Ber06b, Bla76a, Bla76b, Bro72, Cur82, Fri73, Gil12, Hea64, Hen82b, Lan15, Par06, Sch28, She15, Ano21j, Gav95]. Lone [Ano45a]. loneliest [EF69]. Lonergan [Tee02]. Long [Ano11a, D’A01b, Kae13a, Lau50, Lev98, Ano10d, Duc11, Man13, Pin00a]. Long-Awaited [Lau50]. long-predicted [Ano10d]. long-term [Duc11]. longevity [CJT22]. longitude [Ein19b]. Longitudinal [Ein05-36, Ein06e, Har13c]. longitudinalen [Ein06e, Ein05-36]. Look [Bla92, GP19, Ghi05, Mah09, Sch21, AMS04, BNSY98]. Looking [Ein48f, JDB^{+}97, Ing05, Kle74, Ros98]. looking-glass [Ros98]. Looks [Ano87a, Mer85b, Mei64a, Mei64b]. Loop [Gef02]. Lord [Ein24k, Eve39, Eve13, For84b, Kle84, Pai82, Pol83, Red84b, Smo07a, Asi83a, Fer82c, Gil84b, Gol84a, Hei83a, Lak85, Lar84b, Lov86, Pei83, Pip83, Red84a, War84, Ein24k, Jar08a, Tod14]. Lorentz [BSW22, Ein13c, Ein16i, Kil65, Klo63, Prz63, Tre71, Acu14, Ano22e, Ano04e, Blu04, Bor61, Bra22, Bro80, Cos87a, Ein28a, Gia70, Gia71, Gol69, Hir69, HZ01, Ill89b, Jan95, Jan02, Kan95b, Kox88, Kox93, McC70, Ner86, Nor92, Nor07d, Pop66, Pro74, Prz63, Prz67, Prz86, Prz11, Sch74, Ste95a, Tre71, War91, Zah73a, Zah73b, Zah76, Zah78, vL19, vL21, vL52, vL56a, vL65, Hie68, Mic68]. Lorentz-and-CPT [Cos87a]. Lorentz-covariant [Nor92]. Lorentz-Gruppe [Tre71]. Lorentzian [Dor68]. Lorentztransformation [vL19, vL21, vL52, vL56a, vL65]. Lorna [Jon01]. löscht [Bre08]. Loses [Wat97b]. Losing [Kea18, Sto12]. Loss [Ano55d, Ano97e]. Lost [EG78b, Eis06, Kea12, Mar09, Pac72, Pac81, Bre08, Kot93]. Lösung [Ein18i, Ein23c, MGG^{+}95]. Lösung/Meÿenn [MGG^{+}95]. Lösungen [Ein02b, Ein05-39, EM30]. Lösungssystem [Ein18d]. Louis [Ano48b, Fes56, For80a, Hum51, L.58, Ano55c, Cor15b, Elb13, Geo53, Pat93, Tho88b, L.58].

Louise [Bes93]. lourde [Hau48]. Louvain [Ber84a, Sto96, Sto96]. Louvain-la-Neuve [Ber84a, Sto96, Sto96]. Love [Kea12, Krs04, Noe93, Ren97a, Sch92c, Smo07a, vB01, Ano96d, CA77, CG14, Fos05, KSG63, Ove00, Ove01, Red10, RS92, Kan92a, Nev92, Sch92b]. Loved [Bar15]. lover [Sch92c]. Loving [Pri12a, Gla07]. Low [Dum04, Löw78, Eis86, LRH22]. Low-budget [Dum04]. low-water [Eis86]. Lowest [RA13]. loyal [Gro56]. Loyalties [Ber03, Her03, Sch03c, Her02]. Loyd [Hen82b, Kle82a]. Ltd. [Ein20h]. luck [Wal02a]. Lucretius [Die24, Die11, Vai89]. Lüders [Khr08]. Ludwig [ER13, ER13]. lueur [Nor21]. Lugones [dMdA05]. Lukrez [Die24, Die11]. luminaries [Bri95]. lumineuses [Ein10f]. luminous [Ein10f]. Luminy [EK92]. Lunar [Szi71, WTB09, Rod04, vB19]. Lunch [Jan77]. l’Université [Ein55i, Ein64b]. lus [Mat86a]. lxix [Ber07a]. lxvi [Hen88b].

M [And81, Ano88a, Del13, DD95, Ein06d, Ein12g, Ein15b, Ein35f, Ein05-28, For05, Gor02, Har96b, Hen82b, Hoh13, Kat52, Kil95, Klo63, Nye99, Ram82, Row00, Sie97, Smi02, Sta21, Vog20, WvLuAE18, War95, War96a, Wig96j, Wil36, dP94, Wil36]. M. [Ano48b, Ein12g, Ein15b, Ein28d, Ein05c, Ein05-28, SR99, LdB12a, LdB12b]. M.v [Ein05c]. Mêlée [Rol15, Rol19, Rol71]. MA [For05, Gly90, Ste22, WM08]. Maänderbildung [Ein26c]. Mac [Sch99b]. MacGibbon [Fri73]. Mach [Her66b, Wol87, Ano23d, BP95, Bar07, Bla88, Bla92, BIT01, Bra24, CS70, Ein16a, Fey80, Fey84, Fra49a, Gas05, Goe70, Her63b, Her66b, Hol68b, Hol70, Hol88b, Hol89b, Hol92, HW07b, Hon04, Ita82, Mey92a, Ren97b, Ren07d, Rin09, Rin11b, Rin11a, Ros80, ST06, Ste98, Tan92, Wie92, Wol87, Zah77, dWtH20, vBT96, Hen85]. Machian [Ber11a, Hoe94]. Machine [Maa07, SU06, SWU09]. Machines [SE28, Pia31]. Macmillan [Bro72, The12]. macroeconomist [Put09a]. Mad [Pri12b]. Madame [Ano11a, CGT11, Cur37, Cur38b, Cur38c, Cur38a, Cur38d, Cur38e, Cur86, Cur01, Ein59c]. Mädchen [Sin12]. Maddox [VB96]. Made [Ano17c, Ein30o, HG05b, Joh13, Jos95a, Lev01b, RFP^{+}45, Wil11a, Bal75b, BB13, Cli87, DED20, Flü39, Haw11, Jan12, Nic87, Par00, Sch94, TFC93, Wei13e, vE10]. Madison [For01]. madness [Bra12]. Madrid [GR06]. maestro [Far05a]. Maggio [Wac06]. Magic [Cal03b, Eis93, Hun15]. magical [AMS04]. Magill [WK12]. Magnetic [Lev01b, Ein95, Ein10e, Eöt00, GRT08, Man13, San11, Tsu99, WHL^{+}05]. magnetic-field [WHL^{+}05]. magnétique [Ein10e, Eöt00]. magnetishen [Ein95]. Magnetism [Ano29b, Ein10b]. Magnetismus [Ein10b]. Magnetohydrodynamics [PB79]. magnetoresistance [Grü08]. Magnitude [Cos87b, LLSH23]. magnons [DDD^{+}06, NBDD^{+}12]. Magnus [Sch17c, Sch17c]. Magritte [AMM99, AVvdV99, BAM99, Aer99a, Aer99b, ABW99, BAM99, CSvB99, HBR99]. Mai [Ein20b, Ein22p, Ren05e]. main [Gas05, Hen88b, VM05a]. Maine [Lig18]. Maiocchi [Bur90]. maj [Mag95]. major [CT15, Ken12]. Majorana [Mag09, RRR07]. Make [Lau55a, Sus08, Abb02, Ein59f]. Maker [Ano20d, Gar41, Gar39, Gar40, Gar42]. Makers [AG02, Hol89a, Ano99a, CM86, CM97]. Making [Ano39g, Bru09, Gla07, Her93, Kel04, McK03a, Res50, Rho86, Rho88b, Rho95, RA00, Rho12, Rod22, Str13b, The12, Wol09, Yor87, Zol02, Abr98, Bel99, Gos94, Gos99, Gos01, Gro67a, Gro67b, Gro68, Gro00, Käl22, Lau12, Mai99, McK84, Nacxx, Pal98, Rho88c, Rho07, Rol47, Spo16, WS88, WS08]. Maley [You95]. Mall [Bro78a]. Malloy [Fre08, Lit09]. Man [Ano38b, Ano79e, Ano04c, Bel98, Ber85, Bro73b, Buc93, Fac82, Con17, Coo04, Cot89, Ein51b, Far09, FC07, Hea64, Jos95a, Kat52, Kra11, Mic62, Mic63, Oha79, Pac72, Pac81, Peg39, Smi05b, Tau79, Tel78, Ano15f, Ban12, Ber62, Ber05, Boy10, Bri05c, Bro06, Bro07, Ein93c, Gib91b, Gio13a, Gre96, HS14, Hei86, Hei00, Her94a, Her04, HEWW15, Hol03a, Inf63, Inf68, Inf74, Inf92, Inf99, LS92, LS94, Lau98, Mah03, MW44, Mar51, MP99b, Mos87a, Mos87b, Sch62b, Ode72a, Nor02b, Nor03a, Ode72b, PP03, RTS^{+}65, RTS^{+}69, RTS^{+}xx, Rob15a, Rou64, Rou65, Sch57, S^{+}62, Sta87, Sta92a, Str83, Stu10, TFC93, Wal06a, Whi67, Zee89, Zee90, BTM84, Bre05, BW91, CM02b, Fau79, ZDSF11]. managed [Gin04]. Management [Bus13b, HCD^{+}78]. Managerial [LL82]. Manchester [Ano21j, Dys80]. Mandarin [HL10]. Mandl [RS03a]. Manhattan [Jos09, Sho21, Ano07k, Ben16b, CB99, Dei95, Del09a, FS95, Gon00, Gon12, Gos94, Gos99, Gos01, Gro67a, Gro67b, Gro68, Gro00, Gro62, Gro83, Hac87, Hal97, HTS83, HHW99, Jon85, Kel04, Kel06a, Kel07, Kel09, Kel12, Kem12, aM07, Nie89, Nor02b, Nor03a, Pal98, Ree05a, Ree09, Ree11b, Ree11a, Ree11c, Ree14a, Ree14c, Ree14b, Ree14d, Ree15c, Ree15b, Ree16c, Ree20, Sch15a, SFW91, Str13b, Sza92a, Sza92b, Tho17, Wes04]. mania [Oha19]. Manifest [Dra14a]. Manifesto [Rot95, Ano06i, ENFB14, ER55, HY15]. Manifold [LC07, WL12]. Manipulating [CT98]. Manipulation [For73, BRH^{+}09, Chu98a]. Manjit [Sud11]. mankind [Pet21, Pet23, Wel14]. Mann [Inf99, Ber94, Car17, Sch93a]. Manne [Kac96]. mannen [Str83]. Männer [Sch05c, Ste89]. människan [Val55]. Människor [Lar01]. Manoscritto [Mas99, Mas02]. Mantell [Kle75a]. manufacture [Du 52]. Manufactures [Pye13]. Manuscript [EK03, SS21b, She16a, Abb02, Blu22, Ein96a, Ezz87, GR15, Mas99, Mas02, Sau07c]. Manuscripts [Dry14, For80a, Sau04a, Sau04b, Lew79, Ram79a, Ric83]. Many [Bri18, SDCZ01, Tho83, Vil06, Ein39d, RdO07, RdO16]. Many-particle [SDCZ01]. Mao [PB72, Hir97, PB72]. Map [Cun20b, Tra16]. Mapping [ZZ23]. Maps [Gal03, Smo07a, Ste05d, Yer03, Yer04, Gal04a, Sta05a]. Marc [Bar05f, Bru03c, Kra02, Tur05, Nad02]. Marcel [BB89, JRG02]. March [Ano16c, Bad11, Bad15, BS81, B^{+}79, Bre79, BL09, BIRS12a, BIRS12b, EO79, Flo08, Lew79, Nat79, NHP^{+}79, RMR96, Sch28, Whe80, Wis79, Ndi07, Lam79, Wei12t, Wei12u]. Märchen [Jos95b]. Marco [Cho16b]. Marconi [GMOP12]. Marcus [FTG08, FTG08]. Marcuse [PB72, PB72]. Margarete [Chi08]. Maria [Cur60]. Marian [Ein17f]. Marić [Bar98a, Chi08, Hil03, Mil10a, Mil12a, Mil13a, Mil15, Noe93, Sch92c, Ste22, Tou05, T– D82b, Edw00, Alt13, ECS19, Kan92a, Krs91, Mag95, Nev92, Pop03, RS92, Sch92b, Smi04, Sre29, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D07, T– D12, TP90, Wal91b, Wei12i, Est06b]. Marić-Einstein [T– D82b, Alt13]. Marićové [T– D07]. Marie [CGT11, Hor00, QP01, Cur38f, Gir86, Gol05, HM11, Pre05, Qui95, QM96, Qui96, Qui98, Red10, Swa02, Tho88b]. Marina [Cli09b]. Mario [Pye86, Ano97e, Mah01]. Mariscotti [Pye86]. Marketing [Gon05]. Markets [DWL98]. marks [JDB^{+}97]. marks/inertial [JDB^{+}97]. Marriage [Goo90a, Gab95]. marslakó [Har06a]. Martians [Har06b]. Martijn [Hoe18]. Martin [ARK^{+}99, Bal96, Fes56, Har96a, Hie68, Mic68, Pye95, Pye97, Wal99, Ano61, Die95, KS95, War96b]. Martínez [Mar07]. Marvels [Gis14]. Marx [MS17]. Mary [Bed01a]. März [B^{+}79, Lam79, NHP^{+}79]. más [W^{+}87, WdC02]. Masaryk [Kot93]. Masarykem [Kot93]. masks [Bur45]. Mass [Bai33b, BR11, Ein46c, FS09, Hei10, Jan73, Kle14b, Log96, ME14a, Mar08, Nor07a, Nun22, Oku05, Par06, Pri99, Stu93, Sul64, Tur06a, Wei12y, Bai33a, Bra37, Ein06e, Ein35b, Ein41a, Ein00b, Ein05n, FKZ91, LCL^{+}90, Nie89, Ree15a, RKD64, TV07, Van04, Wei12j, Wey20]. Mass-Energy [Stu93, Bra37]. Massachusetts [AS91, HS89b, Kai81]. Masse [Ein06e, Ein05-36, Bra37]. Masse-Energie-Verhältnisses [Bra37]. Massenpunkte [Ein22b]. Massensuggestion [Wey20]. Masses [Ano04f, Ein05-36, Bai33c, Bre08, Ein22b, Ein39d, Zwi37]. Massive [Sam10, AAA^{+}18, Gen05]. Master [Cha05c, Lau50, Joh06]. Mastering [FS01]. Mastermind [Wei12g]. masterpiece [GR15]. Masters [Duf46, War03b, Kai05a]. masterwork [GG15]. Mastery [Ser88, JM86, Wei12c]. Matched [FW17]. Matematica [Mas02, Arg09]. Matemátiques [Gli84]. materia [BES86, Pot11]. Material [BPS87, Mar04c, FGMvH14]. Materialien [KDK87, SKD^{+}87]. Materialism [ABM79, Ein52f]. Materials [HES^{+}12, LU13, Nov11, Pin00a, Ree14a, Hav80, Hee01, Ren08, SKD^{+}87, von90]. Materiály [Hav80]. Materie [Ein19e, MGG^{+}95, Wey18, Bos24b, Ein18h, Ein25a, G^{+}11, JDB^{+}97, WS^{+}01, vEA02]. Materie/Brandmüller [MGG^{+}95]. materiellen [Ein19j]. materja [Inf33]. Mathematica [BHJ^{+}00]. Mathematical [Ano23a, Ano28, ARK^{+}99, Edd23, Edd30b, Kai05a, Mas02, Nor00, Per91, RS99, vD02b, Bla93, Cha98, Edd21, FGPP06, KP55, Kop23a, Lin04b, War03b]. Mathematically [GN11]. mathematician [Ula76]. Mathematicians [Acz01, For71b, For80b, Seg03, Acz03]. Mathematics [Ano07f, BPP^{+}11, Col13, CCG93, DD11, GSW95, Hel14, Hol89a, Jon89, Kat04, New60, Pit17, RSS^{+}07b, Row01, Ser88, Ste16, TP90, Zah80, Ano93b, Arg09, Ari03, Ari05, Bar96, Ber84b, FF91, Gui95, HHJ^{+}06, Hun01, Jag01, Loa98, Mor95a, New56, Pye82b, RR82, War92, War93, Kat04]. Mathematik [Ano93b, Hun01]. mathématique [Bla93, Edd21]. mathématiques [Pia31]. Matica [Mag95, Mag95]. Matilda [Fri17]. Matilde [Ren21]. matriculation [Hun01]. Matrix [BHJ62, Cos87a, Hen86a, Mac77, Lud13, MR82b].

Matrizenmechanik [BHJ62]. Matter [Ano29b, Ano07d, ARK^{+}99, Ber82d, Cli09b, Ein25a, FS98, Fre79a, Kor94, Mie07a, Sil36, SM07, Wey18, Wis79, Ben14, Bos24b, BES86, Dvo12, Ein19e, Ein19j, Fey06, Gat09, GMHB02, Inf33, Inf34, JDB^{+}97, Koh99, Pai86, Phi97, Pol02, Pot11, PP03, Sac81, Sch67, Van98, War91, Wei13f, Wey22a, Wey22b, Wey50, WH07, vEA02, Ano26a, Sch92a, Ein18h]. matter-wave [Koh99]. Matter/Friedrich [ARK^{+}99]. Matters [Bus13b, Gre15b, McG19, Raf17, Hag08, Lin17, Rei72]. Matthew [LLSH23]. Maturitätsprüfung [Hun01]. Maurice [KW87, Spe62, ES56, Ein05-47, Sol56]. Mauro [Bar05f, Bru06]. Maverick [Sch13a]. Max [Adl03a, And81, Ano61, Ano69c, Ano79h, Ano79g, Ano05f, ABdB^{+}53, B^{+}79, Bey06, Bre08, Bro72, Cas05b, Del09b, EP06, Far79, Fis10, Fis12, For70a, For80a, Gut15, Han80a, Hei79a, Her77, Hof08b, Hof08c, Jac08, JDB^{+}97, Jüt21, Kan96a, Kle70b, Kle70a, Kra25, Lam79, Lin60, Mar73, Med76, MTVB88, Mic68, Prz86, Ren05e, Sch05c, Sch79c, vBL96b, Ano69c, Ano79g, Ano05h, Ano05i, Ano05j, B^{+}79, BB71a, BB71b, BB05, Bra14a, Bre08, Ano14b, CS79b, Ein06d, Ein13d, Ein54d, EHM^{+}79, Ein53g, Ein91d, Ein91f, EP06, Fis10, Fis12, FPLR59, Gol76, Gor22, Gre05b, Gut15, Hei79b, Hei86, Hei00, Hen21, Her77, Her79a, Hof08b, Hof08c, Jac08, JDB^{+}97, Kan96a, Kra00, Kra25, Lam79, Mas01, Mil75, Nag82, NJ05, Prz11, Rav97, Ren07c, Ros79c, Sch05c]. Max [Seg11, Wei12w]. Max-Planck-Gesellschaft [vBL96a, vBL96b]. Max-Planck-Institut [Ren05e]. Max-Planck-Institute [MTVB88]. Maxim [Rob48]. Maxwell [Bes04, Bri18, Ber84b, Bes04, BE69, Ein16c, Ein31g, Ein31h, Ein54e, Ein82d, EM84, Ein11b, Faz15, FMM96, Gar70, GBE95, Gol83b, Hal80, Hal04b, Har02c, Har98, Hen86b, Hou16, Kni16, Mah03, Por81, RA11, RdO07, RdO16, Sie91, Tho31, Tho11a, Wei12c, Wis82]. Maxwellian [Lan21]. Maxwellians [Hun91, OP87]. Maxwellsche [Lan21]. Maxwellschen [Ein16c]. May [Ano23a, Ano30b, AS91, Ber07a, BIR^{+}15a, BIR^{+}15b, BIJ^{+}18, BRI^{+}21a, Edd19, Ein20b, Ein21l, EA22, Ein22p, EA23, Gly90, HS89b, Lac08, Mag95, SW94, Sch54, Sti13c, Ano19c, Bau20, BSR^{+}06a, BSR^{+}06b, DED20, Mil24, Roe93]. May-December [Lac08]. Maya [How97b]. Maybe [Rai14]. McCormmach [Ser88]. McGrayne [Wat00]. McGuinness [Con83]. MD [Ano13c]. Mé [Inf59, Ano06c, Büh98a, Cle06]. Mean [Fre31, Duc11, EdS32]. meandering [Ein26c]. Meaning [Ano23a, Ano82d, APWB95, Bey09a, Bey09b, Bor50, Duf46, Ein21l, EA22, Ein22p, EA23, Ein35c, Ein45c, Ein46d, Ein50j, Ein50h, Ein50i, Ein53h, Ein53i, Ein55g, Ein56e, Ein56f, Ein56g, Ein65, Ein67a, EA67, Ein69b, Ein74, Ein81, Ein84c, Ein88a, Ein94d, Ein96b, Ein03b, Ein05-48, EG05, Hol99, Mor96, Rai46, Sch15, W.25, Ano01c, Bag92, Bag04, Ein53n, Fey98, Gef14, Gil19, GHS21, GR15, GR17, Jay90, MW46a, MW07, Ner84, Oku05, RR89, Sch08, She16a, You11]. meanings [Owe12]. Means [Tho83, VW86, Her82b, Kra21, Mus15b, Per23b]. Measure [De 10, Asi83b, Bag04, Nor65, Nor90]. Measurement [Ano26a, Bel75, Car17, Cha11, GM00, Leg05, Mos13a, Mos13b, Bel90, DCL23, Ein08a, Gré19c, WZ83]. Measurements [VT99, FTZ^{+}15, Khr08, Poo90]. Measures [Lau13, Bus12c, Fie16, Irv67b]. Measuring [Per12, Van85, Web99, YCL20, Fri64, GAHS96, Hal06]. mecânica [GdS15]. mécanique [BD92, BB55, EBD^{+}89, LB39, Pia31, CT80, CTDL73]. mécaniques [Pia31]. Meccanica [Ros80]. Mechanical [CHS97, Ein11f, EI64, Ein05j, Kae31, Ros79a, Ser77, EPR35, Kie15, Nav91, TN03]. Mechanics [Ano07g, Ano07e, Bor53, BHJ62, Cro07, Dav14, Ein27e, ETP31, EI39a, Ein48a, Ein87a, Eza79a, Eza79b, Fre15b, GGK02, Ghi05, GL23, HSD34b, Hel98, Hel15, Hen86a, Hie68, Hol05a, How90b, Hug89, Jam74d, Jam89, Kel85, Kub86, Mac77, Mic68, Min07, Nav98, Nor07a, Pao84, Pat93, Pau47b, Prz63, Prz67, RSS^{+}07a, Rüg88, SF47c, SF83, Slo11, Sus08, VW86, vdW67, Bel87, Bel04, BD92, Bit96, BB55, Bor55b, Bor63b, Bun79b, Cam09a, CTDL77, Dir67, Dir81b, Dir01, Dir03, Dir09, EE59, EE90, Ein19b, Ein19c, Ein27b, Ein53d, EI61a, EBD^{+}89, EDK05, ET07, FH65, FHS10, FCKT11, Fri11, GBE95, Gea90, GdS15, Gor17, GZ97, GZ06, Gri95c, Gri05, Han79b, Hei34, Hei65a, Hen84, Inf31, Jam66, KKC06, Lau98, Leh14b, LB39, Mar97, McC11]. mechanics [McH07, Meh75a, MR82a, MR82c, MR82b, MR87b, MR87a, MR00, Mor16, Nau16, Nug00, Pat95, PCH77, Prz86, Prz11, Raz11, Ren97a, Sch28, Sch82a, Sch99a, Sch03b, Sel88, Sty00, Van07b, Vig94, Vig95, Wil79, dM86, vdW68, vdW07, Mil90, Sch35a, Mil12b, Ros80]. Mechanik [Ein19b, Ein19c, Ein27e, GRE^{+}01, Ein48a, Ein53d, Inf31]. Mechanische [Ein11f]. mechanischen [Ein05j]. Mechanism [BW39b, FW67, Eng14, Mal79]. Mechanistic [Hir76]. Mechoulam [Bus12b]. méconnu [Dur12]. med [Pat01a]. Medaille [Ano61, Ano79g, BEvL^{+}28]. Medal [Ano26b, Ano29a, Com46a, Ano79g, BEvL^{+}28, Rod22]. medalists [Hen13, EGR11]. medarbejdere [RRK^{+}64]. Media [Can15a, Car07a, Ein22n]. Medical [Kli62a, KWM57, Wel00, Ano81c, Kan96b, Mey07, Wel99]. Medicine [Ano82b, BC82, Coh54, Cra02b, HBFH18, Rei82, Aas88, Ano97a, BSPKN03, FMR90, HF22, Kut09, Ler03, Nob03, Sou66, Ste53, Tho88a, Tho12, Bar05f, Bjö01, Enk18, Fie15, LB02, Ano04b]. Medicus [Sie10, Pes13]. Medien [Ein22n]. medieval [Gra94b]. medium [Grø82]. Medizin [HBFH18]. Mediziner [HEB^{+}80]. Mediziner/Hermann [HEB^{+}80]. Medizinnobelpreis [Enk18]. Meer [Bre08]. meet [Lan06]. Meeting [Ano39a, Ano93c, BC71, JRG02, LLH^{+}80, Bad12, Wig96a, Ano33g, BB89, CR07, Dör05, Hof08b, LdB12b, Lin04a, BPS87, Van07c]. Meetings [Hei75b, Jof62, Jof67, JDB^{+}97]. Meets [Row01, Aer99a, AMM99, Aer99b, ABW99, AVvdV99, Bou99, BAM99, CSvB99, HBR99, Ill06, McH07, Smi07a, YotY99]. megascience [HKW08]. megatons [CG01, GC02]. Megawatts [CG01, GC02]. Mehra [Hen86a]. Meilensteine [RKS08]. Mein [See53, See55, See64, See80, See86, See91, Bor75, Ein34g, Ein34h, Ein35d, Ein53j, ES05b, Esp55, Hah68]. Meine [Ein15h, Ein16k, Ein20h, Ein40c]. meinen [Ein52a, For80a, Hei53a]. meiner [Ehl05a, Ein07b, Ein11a, Ein11e, Ein12f, Ein15c, Ein23b, Ein05l, Ein05i, Ein05m, Ein05y, vI11b]. Meinung [Ein15h, Ein16k, Ein24g]. meiste [Ehl05a]. Meitner [Ano69a, Ano79g, B^{+}79, Chi08, CSW97a, CSW98, Ern92, Kac96, Kri99, Lam79, Lin60, MGG^{+}95, TF99, Ano38a, Ano69a, Ano79h, Ano79g, B^{+}79, Brü00, Büh98b, CSW97a, CSW97b, CSW98, EHM^{+}79, Ern92, For80a, Fri17, FPLR59, Fri68, Her79a, Kri99, Lam79, McL99, Mei64a, Mei64b, Sim90, Sim94, Sim96, Tru96, Wat97a]. Melcher [Bäu83]. Melia [Del11b, Hub13, Pye10]. Meltdown [For86, For82b]. Meltdowns [Mah14, Mah15]. Members [Ano82b, Sch88a, Sch88b]. membrane [TPM86]. Memo [Jan07, Nol05]. Memoir [Bed01b, ČA99, Wig96e, HB05, Seg95, Yin15, Dix00]. mémoires [De 13]. Memoirs [BR00, Con70, Els78, Lin56, LEMW52, RK86, TS01, Erm89, LEM^{+}23, Ric93, SSSS95, And81]. Memorabilia [Cla68, Sha79b]. Memorable [Rom91]. Memorandum [FP40b, Ber11b]. Memorial [Bus12b, Ano69a, Ano80a, FW60, Rei72, Ano80a, Ein16g, Pát79a, VM05a, Wig96a]. Memoriam [Ein54d, Ein53g, JS87, RF79, See56b]. memorie [Ein55f]. Memories [AGMT12, Ein55f, Inf56c, Inf56d, Inf59, Mei05, Whe79b, BR93, De 13, Kol56, Bor56, Bor63c, Bor65b, BI67, BI69, Jof55, Sch87c, Wig96g]. Memorium [Ein22i]. Memory [BC71, CFW76, Gin08, Her79a, Hog96, Inf69, Lin99, RR82, SK09, See56b]. Memphis [BS81]. Men [Ano19b, Ano41, Bro21, Cha75, Cli87, Cou79b, Goo90b, Lau59, Med94, Mos68, Ste89, Tor82, Bus49, Bus68, Bus85, Dor10, Ein17, Sch05c, Sch94, Smi07b, Van03]. Mendeleev [Ste07]. Mendelssohn [Gra99]. Mensch [Bre05, Cot89]. Menschen [BW91, Buc93]. Menschheit [Bre05, Pet21, Pet23]. Menschlichkeit [Her33, Her63a, Her67, Her75]. menswording [vE95]. Mental [Kri08, Hen07]. Mentality [Ein47b, Ein47e, Kri08, Hen07, LM90]. mentee [CZPR17]. Mention [Van88]. mentioned [Gio14]. mentor [CZPR17, Rav97]. mentoring [CZPR17]. Mercer [FW07, Whe79b]. Mercury [Van11, Ein15e, Lan02, Ros82c, Van10, Wei14b]. Merely [Ber01]. merged [MTVB88]. merit [BSPKN03]. Merkur [Ein15e]. Merleauponty [Nav94]. Mermin [Plo06, Oha12a, Oha12b]. merveilles [Gis12]. Mesa [BS97]. Meson [Sch49f, Yuk49]. Mesotron [Kap40]. Message [Ano48c, Ein48g, Ein50k, Ein53l, Ein53k, Ein59b, Lau97]. Messung [Ein08a]. Metadebates [CSvB99]. metafora [Cas14]. Metalle [Ein22m]. Metallen [Ein02b, Ein05-39]. metallic [Hee01]. metallurgy [Hab02]. Metals [Ein05-39, SE28, Ein02b, Ein22m, Ein05-49, HS39c, Mac01]. Metamorphosis [Tre75b]. metaphor [Cas14, Whi01]. Metaphysical [How89, Col57]. metaphysics [Bar94b, Moo81b]. Metaphysik [Bar94b]. metaphysique [Col57]. métaphysiques [W.25]. Meteorology [Tau63, CLE27]. Method [Ach13, AO78, CW32a, Cor07, Ein14j, Ein14h, Ein33f, Ein34k, Ein86c, Ein87b, Ein05-36, Ein05-39, Fri89, Gha17, How76, Mol63, Ner86, Sla51, Tau49, Tib13, Ben16b, Eĭn85b, Eĭn86a, Ein02b, Ein06e, Ein08a, LHF^{+}98, Tee02, TLW^{+}10, Tha77]. Methode [Ein02b, Ein06e, Ein08a, Ein14h, Ein05-36, Ein05-39, Ein14j]. Methodological [Mor79b, CT15, Duc11, Mon04]. methodological-experimental [Duc11]. Methodology [Wei12f, vD04]. Methods [PFM17, QPSD13, Sch17b, Yuk49, HMW24, Smy45c, Smy45d, Smy45e]. Métiers [Bad11]. metodo [CM02a]. Metric [PKLT17, Grø82, Ein30a]. metrics [Ein30r, NP88]. Metrik [Ein30a, Ein30r]. Mexico [GHK^{+}96, Ree17c, GHK^{+}96, Lau45, LSW09, Mai99, Mar11a, Roe93, SSG13, Swa92, Swa95]. Mey [RSPS49]. Mey/Albert [RSPS49]. Meyenn [Eis00, GRE^{+}01, HEB^{+}80, ARK^{+}99, MGG^{+}95]. Meyerson [Ein28d, Zah80, Zah87]. MG1131 [HTS^{+}88]. MG1654 [LCL^{+}90]. MGIX [JRG02]. Mgr. [Sto96]. Mi [Seg86a]. Mi-karne [Seg86a]. mian [HL10]. Michael [Con83, Dix00, JW06, Jos95a, Dör05, Wis21]. Michel [Pye06, Row00, vD15, Hus15]. Michele [Ivr99, Kle73, Spe72, Ein94b, Ano05k, Fin05, KH05, Tho05, Wes05, Whi05, Ein94b]. Michelle [Ano05m]. Michelson [Gre04c, Hol69b, Hol88a, Ita99, Lor21, MST^{+}07a, MST^{+}07b, Oga79, Rus98, Sha64a, Sha64b, Sta82b, Swe13, vD09]. Michigan [Lan65, Lan67a]. Michiganskom [Lan67a]. Michio [Kar05]. Microbes [Coo04]. Microbiology [Lag03, Coo04]. microcavity [KSVW10]. microchip [Ber05]. microelectronic [HHHR01]. microfiche [Hen88b]. microphysical [Hel82a]. Microphysics [BW01, Fes56]. MicroReviews [Hub13]. Microscope [BR85]. Microscopic [WHL^{+}05]. Microwave [Löw78, Sla42, Sla50, Sla69b, BRH^{+}09, Smo07c]. Microwaves [Ano14a]. Mie [Bor07, Cor04b, SM07]. Mifflin [Nah01, Sto17]. Might [Ken13]. Mighty [Ser88]. Migration [FL49, Hol84]. Mikhailovich [Kuz79]. Milan [Bra15, Tou05]. milanaise [Bra15]. Mile [For82b, Wal04b]. Milena [Chi08, Hen16, Nav15]. Milestones [BC82, Gon96b, Gon96c, Gon96d, Gon96e, RBS^{+}08, Sea72, RKS08]. Mileva [Est06b, IK03, Krs91, Noe93, RS92, Sch92c, Ste22, Tou05, T– D82a, T– D83, T– D85, T– D88, T– D93, Alt13, Bar98a, Chi08, Edw00, Ein95a, ECS19, Goo90a, Hil03, IK03, Kan92a, Krs04, Mag95, Mil10a, Mil12a, Mil13a, Mil15, Nev92, Pop98, Pop03, Pop05, Sch92b, Smi04, Sre29, T– D82a, T– D82b, T– D83, T– D85, T– D88, T– D92, T– D93, T– D07, T– D12, TGC91, TP90, Wal91b, Wei12i]. Mileve [Mag95, Pop98, Pop05, Sre29]. Milevy [T– D07]. milieux [Pia31]. Militant [Ein31i]. Military [Bla49b, Ein47b, Ein47e, Smy45a, Smy48, Ano13c, Bla48, Bla49a, Ein48e, Jac11, Kut09, Lor08, MSW88, Roe93, Smy45b, Smy45c, Smy45d, Smy45e, Smy46, Smy76a, Smy78, Smy89, Van04]. miljöer [Lar01]. Milky [Fer03, Pau93]. Mill [Mog02]. millennium [Ano99a, Ano97b]. Miller [Kai81, Mel82, Pal83b, Cus82, Fer81, Hel84, Pal83a, Ram83, Swe13, Whi83b, Zah78]. Millikan [DR89, Fra81, Hol78a, Hol99]. Million [Bar18b, Bro79a, Gar62, Jon17, TT70]. millionaire [EF69]. Mills [BT12, HF05]. Milne [Grü73b, Smi21]. Milton [Put09b]. Mimicking [Ano10d]. Min [Cur38f]. Mind [Ano14c, EMH96, Hür22, Inf50b, Joh06, Kur12, Sch67, Amo10, Ano06j, BC01, Cha05c, MS17, Opp55, Opp63, PSC^{+}97, WD06, Sch92a]. Mind-blowing [Ano14c]. minded [Rad08]. Minds [CC96, Sau16, BT15, GM04, Gar93, Hal15, Lin04a, McC01a, Pic08]. mindset [HBO82]. ming [Seg86b]. Minimal [Sil36]. minimum [Gin04, Ree15a]. Minkovskis [Sch27]. Minkowski [BSW22, Ano07t, BSW22, Cor97, Jan03, PS04, Pye77, Sch27, Sla22, Wei12w]. Minorities [Ano46b]. Minority [FR49a]. mins [Wac06]. Minute [Lau50, Par11]. minutes [For82b, Kee11]. mir [Büh98a]. mirabilis [GG05b, Jan06, Mas05a, RR06]. miracle [Ano05-39]. miraculous [Nor06, San07, Sta98a, Sta05b]. mirage [Mat10]. Miralles [Pfa12]. Miramare [Meh73b]. Miranda [You95]. Mirimanoff [Ein05g, Ein09a, Ein05g]. Mirror [van12, HW96]. mis [Gio13a]. Miscamble [Cli09a]. Mises [Ano93b]. misgivings [Ein10]. Misinterpretation [Bac89a, Sig92, Bac89b]. misinterpretations [Hen90, Sta95b]. misrepresentation [Hni98]. Miss [Ano38a]. Missed [PW17, Bun82, Kau14, Lag12]. Missile [Edg14, Mac93, Sla13]. Missing [FS09, Hel14, QN08, Sau07b, Tur40a, Hab09, Sam10, Sau05, Van98]. Mission [Ein25f, Dei95, Gou48, Oli99, TM77, TM90, Wil16b, Egl97, Pas69, Pas70]. Missionary [Ren13]. Mistake [Sto17, Blu22, Bod16, Bod17, Ray90, vM05b, vM09, vM12]. Mistaken [van12, Log96]. Mistakes [GCO^{+}06, Oha08, Wei05a, Liv13, Ng11, Wei06a]. mistayyer [Ayn34]. místicos [W^{+}87, WdC02]. Mistral [Ano45b]. misu [Gol98b]. Misunderstanding [Ein33g, SHW^{+}49]. Misunderstood [Rös95, WW75]. Mitarbeitern [Mei05]. Mitentdecker [BH01]. Mitführung [Ein14f]. Mitgliedern [Sch88b]. Mittag [Stu10]. Mittag-Leffler [Stu10]. Mitte [BHJ^{+}00]. mittels [HS39c]. Mixed [AKBAE21, Big05, Led14]. mixing [EG98, Kob09]. Mixtures [Ein05-29, Ein10a]. Mk [Jüt21]. Mladjenovic [ARK^{+}99]. MM [JRG02]. Mme [Swa02]. Mobility [Mel63, Pik63]. Mod [Leg03, Nam10, ’t 02]. mode [Gra94a, Gra96]. Model [Ano94b, Cas79, Hol79b, SMHL23, Stu86, Stu94, BP79, Gui09, KP55, OM14, OMNM14, OONM15, OONM17, OOM21, OM22, vN45, HBRD97]. Modeling [ARK^{+}99, Cha17, FW17]. Modeling/Origin [ARK^{+}99]. Models [Mac77, PS03, TVL^{+}09b, Yuk49, Cor15c, Nug88, Pop99, Shi75]. Modern [Ano24b, Anoxxa, Anoxxb, ADS83, APWB95, BTM84, BF86, BBBM90, Bir23b, BW04, Bus49, Bus68, Bus85, C.50, Čap71b, Ein10c, Ein10d, EI39a, Far03a, Fey80, Foc53a, GHS08, GHS18, Gib19, HSD34b, Hei49b, Hie84, Hol98a, Kle17b, Kra18, Leg15, LL82, Nye19, Pau94d, PB79, Sin70b, Sla55, TWD03, Tha77, Vis94, Wei05b, Wol95, Zah77, Bag04, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, Bra09, Cal03b, Coo99, DB09, Die05, Ein20g, EI61a, ES68, Els24, ER13, Far02, Far05c, Fer07b, Foc54, GdS15, GZ97, GZ06, GH07, Hag14, Hal12, Hei61a, H^{+}61, Hen82b, How76, Inf34, Joh11a, Kol40, Lac11, LB90, Liu92, MW99, Mar17, MB66, McC11, McC01a, McM05, Mey07, Myr21b, Ndi07, Nor65, Nor90, Nor89, Nug88, Par04, Pla59, Rio87, SS97]. modern [Seg80, Seg87, Seg07a, Sin61, Sin70a, Uns00, Vog21, Whi83a, WB00, ZdSS91, Ano23a, Ano81e, Cob07, HSD34a, Seg84]. Moderna [GdS15]. moderne [APWB95, BTM84, Ein10c, Ein10d, Ein20g, EI39a, EI61a, ER13, Foc54, HSD34a, MGG^{+}95, Uns00, Wil36, Bre15]. modernen [ES68, Els24, For70b, Kle17b, Mar17, SS97, Win69]. modernes [Seg84]. modernim [Seg86a]. modernism [VM99a]. Modernist [Ebu16, Whi01]. Modernity [GB03, Bre15]. Modernization [Rei07]. modernos [Seg87]. Modification [LC29]. Modifications [Hen86a, Pap15, MR82b]. modiga [Fre15a]. Modigliani [Put09a]. Moet [vC05b]. Möglichkeit [Ein07i, Ein07c, Ein05-32, Ein05-33]. Möglichkeiten [Ein23c, Ein28c]. moi [Jof62, Inf56c]. moindres [Pia31]. Mój [Ein35e]. Moje [Inf56d]. moleculaire [EdH15c]. Molecular [CLD^{+}03, Cas07, ES13, EdH16a, EdH16b, Ein05m, ES05a, Ein05t, Ein05-35, Ein05r, Ein05-39, FL49, Her49, Nye72, Bal77, Ein04, Ein05c, Ein05a, Ein06c, Ein11e, Ein15c, Ein15f, EdH15a, EdH15b, EdH15c, Ein16e, Ein05-40, GRJ03, Haw02a, Per10, Per05b, Sie91, Sla79, Uff06, ZDSF11]. Molecular-Kinetic [Ein05-35, Ein05c]. Molecularkräfte [Ein05-39]. Molecule [Ein05q, DCTW02, Ein20n]. Molecules [Sla63, Ein11g, GBE95, Kon15, Sla53, Tow65, Zol02]. Molekül [Ein05q, Ein11g]. molekular [ES13]. Molekularbewegung [Ein11h, Ein05t, vS06]. molekularen [Ein04, Ein05-40, ES05a]. molekularkinetischen [Ein05c, Ein05-35]. Molekularkräfte [Ein02b]. Molekularströme [Ein15c, EdH15b, Ein15f, EdH15a, Ein16e]. Moleküldimensionen [Ein05r, Ein05m, Ein05a, Ein06c, Ein11e]. Molekülen [Kon15]. Molina [Ano97e]. moment [CG14, Ein20n]. Momentous [Wat00, McG93, McG98]. moments [Bag11, Büh98b]. Momentum [Bor07, Wei13c]. Mommsen [Fla15]. monads [Car22c]. Monarchy [Lan15, Sca14]. Monatomic [Ein24h, Ein25h, Ein05q, Ein11g, PS10]. Monde [Hum51, Bla93, Ein34a, Ein58, Ein89c, Ein09a, Ein15a, Eis37, Sar35]. mondes [Chr22]. mondiale [Mar91]. Mondknoten [vB19]. Mondlänge [Ein19b, Ein19c, vB19]. mondo [Ano82d, Arg09, Inf63, Inf68, Inf74, Inf92, Lau15]. mondo-della-vita [Arg09]. monetary [Put09b]. Money [BHJ^{+}00, Wol16, MTVB88]. Monique [Ano60]. monograph [Myr21b, Vog20]. Monomorphism [How96a]. monopole [Lom05]. Monopoly [CP15, Sch54, Gor09]. Monsignor [Sto96]. Monsters [Ami87, Imp19]. monstrous [Nik06]. Montale [AMB^{+}77]. Monte [Ben16b]. Montgolfiers [Eis93]. Month [Cho10a, Meh98]. months [EG78a]. Montias [Spe67]. Monument [Vir92]. Monumental [Spo16]. Monuments [Far05b]. moon [DvdRW37, Hen91, Van70, Ein19b, Ein19c, Gut98, Mer85b, WTB09, vB19]. moons [RA11]. Moonwalking [Foe11]. mor [Cur38f]. Moral [Ano69b, Cro16, Ein50m, Ein52c, Ein50e, Ein52b, Ein79f, Sch02a, Sch00d, Wal77b, Wal92d, Wal06e, Wal15]. morales [Ein50e, Ein52b]. moralist [Sch05a]. morgen [Ano15f]. Moritz [ER13, Ein05a, ER13, Pul06, SW12]. Morley [Sha64a, Sha64b, Swe13, vD09, Gre04c, Ita99, Lor21, MST^{+}07a, MST^{+}07b, Oga79, Rus98]. Morocco [Ada17]. Morozov [Hoh13]. Morris [You95]. Morrison [Hal12]. morsels [Par11]. Mortal [Wes04]. Mose [Jon89, New56, New60]. Moses [Pye06]. Moss [BHJ^{+}00]. Most [Ano17c, Ban51, Bod05a, Bod05b, BR93, Col13, ESA15, Eds82, Ein46a, FC07, Gut15, Kha20, Ree17d, Rog10, Sch54, Ano51b, Ber07c, Ber09, Bla02, Bod00, Bri05a, Cha83, EM33a, Eng02, Fey95, Ful84b, Ful84a, Gla00, Haw11, Jer02, Lui11, Nor12, PP07, PP06, Rod22, Shu17, Wad81b, W^{+}87, WdC02, Wor10, Ehl05a]. Mostly [Mes83]. Mother [Sha14, Cur38f]. Motion [Bru68, Cha49b, Ein06b, Ein08b, ES13, EI49a, Ein05t, Ein05-42, Föp07, Fri07, Get06, IS49, Inf50c, Inf50d, Inf56e, Kub86, Lan08, Lav85, LG97, Rei06, SP12, Sil23, SMM^{+}60, SBSW60, Van11, Wei14b, Ano05p, Dim93, Ein07f, Ein15e, Ein27b, EIH38a, EIH38b, EI40b, EI49b, Ein05n, EH05b, GW06, Haw02a, Ian93, Inf51a, Jan12, Leh19, Mai90, Rei21, Rei11b, Ren05a, Rin03, Rob38, vS21a, vS21b, vS06, Ein22o, EG27a, EG27b, W.25]. Motions [Cle47, Ein11h, Poo21, Vig94]. Motivation [Nor01]. motivations [Cha87]. Motive [Ein18a]. Motors [Gwy10]. Mountain [Hun15, MM00]. Mountains [Mah14, Mah15]. Mourns [Ano55d]. mouth [Gla07]. mouvement [Ein56i, Ein94e, Lan08, LG97]. Moved [Sta05c]. Movement [Ash11, Ein05-35, Str68b, SE28, Tes69, Wal07, Woo28, Ein05c, EH10a, Ein27c, Ein56c, Per10, Per05b, Smi65, Smi70]. movements [Lin15, Yam09]. Movie [Wac06, KMN^{+}97]. Moving [EL05b, Ein05-41, Fok49, Ryn05, Sta05c, Wil77, Ano19a, Ein05e, EL08a, EL08b, EL09, Ein20q, Ein56i, Ein94e, EL05a, EL05c, Eis07, HS06, Lor21, Nor14, Pra06, Rya06, TN03]. mozek [Nes60]. MP [Ree05a, Ree11a, Ree16c]. MP-1 [Ree05a]. MP-2 [Ree11a]. MP-3 [Ree16c]. Mr [Ano00b, Mil12b, RS11b, Din65, Ein15c]. Mr. [Abr12a, Abr12b, Ano00a, CB06, Cra00a, Ein07a, Ein18i, Ein05k, FLH85, Fri11, Goo00, LdB12b, Pat00, Pat01b, Pea06, Rit09b, Sta99c, Vea01, vB19]. MR1503389 [Rob38]. MR2201925 [Kra06]. MR2253243 [Flo06]. MRI [Man13]. Mrs [Bru98]. Mrs. [Bru97b, DL15, RMR^{+}50]. Ms. [Wal90a]. Much [Ano41, Mad89]. Mulino [Seg91]. Multi [Kni13]. Multi-beam [Kni13]. multimode [GLG09]. Multinomial [DR08]. Multipartite [HR13]. Multiple [Fra05]. multiplication [Pei39]. Multiplier [Lan49, HH10]. Multiverse [Gre12, Gri10]. München [SM41]. mundane [Ano04d]. mundi [FG20]. Mundo [Esp34, Esp55, Inf57a, Inf60, Inf61a, Inf61c, Inf83, Phi74, See61, W^{+}87, WdC02, de 03]. Munich [Kle70b, Kle70a, Pye82a, SM41, TEW02]. Murder [Ber92a, Mah17]. Muse [Can00, For01, FD85]. Musée [Bad11]. Museum [B^{+}79, For05, Han80b, Ano13c, Fec90]. Mushroom [Asa95, PNT87, Bes93]. Mushrooming [Wei88]. mushrooms [Ing05]. Music [Ein80, Rod92, Dub05, EE05, Ehl05a, Fos05, Gla07, Sch10a]. Musical [Sch82b]. Musik [Ein51e, Ein80, Sch10a, Ehl05a]. Musiklexicon [Sch02b]. Musings [Cal03a]. Muslim [Alf09, Bus12a, Fra08, Gha82]. Must [Ein35f, Far03a, RFP^{+}45, Far02, Gio14]. Mutation [Cre15]. Mutter [Föl90]. My [Bet03, Bor78b, Bro06, Bro07, Con70, Cou52, Cug06, Ein12f, Ein23e, Ein40e, Ein48h, Ein05-45, Ein05m, Fre06, Hah70, How50, Inf15, Jof62, Rie12, See53, See55, See64, See80, See86, See91, Sus08, Wil13b, Bai07a, Bor69b, C.51b, Ehl05a, Ein07b, Ein11a, Ein11e, Ein15c, Ein21v, Ein23b, Ein50o, Ein50p, Ein52a, Ein56h, Ein70a, Ein71, Ein89d, Ein93c, Ein05l, Ein05i, Ein05y, Fer54, Fer82a, Fer87, Fis92, Hei53a, Hof55, vI11a, vI11b, Inf50b, Kel06b, Mof10, Sta87, Ayn34, Bor75, Cur38f, Ein15h, Ein16k, Ein20h, Ein34a, Ein34g, Ein34h, Ein35e, Ein40c, Ein53j, ES05b, Esp55, Hah68, Inf56c, Inf56d, Inf59, And81, Far79, Han80a, Hei79a, Hei79b, Sch79c]. Mynttinen [Pye06]. Myrdal [Fel82b]. Myriad [Mar11c]. mystére [Nor21]. mystères [Spa17]. Mysteries [Ghi05, Bru97a, Bru97b, Bru98, DB86, Gri95b, Gri95a, Mar11a, Spa17, Vel03]. Mysterious [Ano16d, Ano18b, Jea30a, Jea30b, Jea32, Jea38, Jea43, Jea48, Jea76]. Mystery [Abb88, Acz01, FR13, Gre12, QN08, Bes04, Bod03, GO81, Nor21, Rei41]. Mystic [Ano29c, Far09, RS07e, Sta07c]. mystical [W^{+}87, WdC02]. Myth [Bay94, Elk82, Goo90a, Ven12, ATW^{+}95, Dat06, FD85, Gre96, TH79, Wal95, Ham97, Ove96, VB96, You95, TH79]. Mythical [Wei13f]. Myths [Mar11a, Wal92b, Dor10, Kra07a, Mad07, Mar11d, Num09, Uda94, Uda98, Kni12a, Kni12b].

N [Ein14f, For80a, Hen82b, Kat52, Mur25, Plo06, Ree17d, SR99, Tay21, Gol97, Mah17, Nor03b]. N-rays [Mah17]. N-Stein [Nor03b]. N.J [Ano60]. N.Y [Bar05f]. N.Y.U. [Ano29d]. nablëiìudeniem [Smy46]. Nabo [NN01]. Nach [Dir60b, Aue95, Bod03, Ein16m, Ein19b, Ein19c, EG23, FTG08, HH10, Hei53a, Her77, Jac08, Ste05b]. Nachlasses [BH01]. Nachmansohn [BS92]. Nachruf [Ein34i]. Nacht [vS14]. Nachtrag [Fis02, Ein12f, Ein05y]. Nachträgliche [Ein14k]. Nachweis [Ein15c, Ein15f, EdH15a, EdH15b, Ein16e, HS39a, HS39c, HS39a]. Nachwirken [Sch09a]. Nachwirkung [Fre85, Ian93]. Nagasaki [Ano81c, Ano08b, Gol96, Jay90, Man5 a, Man5 b, Mal85, Oli99, Sod95, Wel06, YJ95]. Nähe [Ein10a, Ein05-29, vS21a, vS21b]. naheliegende [Ein21t]. Näherungsweise [Ein16l]. Nährpflicht [Ein18b]. naissance [Ano80a, BD92, Sto96]. naixenca [Gli84]. najdawniejszych [EI59, EI62a, EI98b]. Najwieksza [Gol98a]. Naked [Mat91]. Name [Wil13b, Cug06]. Named [Ano33c, Ano77, Ano78, Ano11b]. Nancy [Cas05b, Bey06]. Nanocrystals [ARK^{+}99]. Nanocrystals/Gershenfeld [ARK^{+}99]. Nanoscale [MDT^{+}03]. Nanotechnologies [BZ10]. nanotechnology [HES^{+}12]. naplója [Tőz15]. naplók [Tőz15]. narození [Bič79a]. narrative [Com56b, VM99a]. nas [Inf79a]. Nascita [Seg91, BF02]. Nath [Das12]. Nathan [Kap61, Kie15, Blu22, CHR90, Kie15, RMR96, Sta06]. Nation [Ano33c, Ein21d, Ein21d]. National [Asa95, BIK83, EO79, Gol68, GR63, GR65, Han80b, HF22, Sei99a, Wal89, BSPKN03, DMPS81, B^{+}06, Ein18l, HSW09, Hen96, HW04, RW94, Wal92c]. Nationalism [Cra92, Eli92, Hoc93, Koj08, Lan08, Sta19, RS07f, RK94, Wei94, Wal94b]. Nationalismus [Stu87]. Nationen [Ein40d]. Nations [Ein24f, Ein40d, Ein47c, Can05]. Native [Eic94]. NATO [Sto12]. Natu [Ein26d]. Natur [Die24, Die11, Ein24g, Ein86c]. natura [Die24]. Natural [ABdB^{+}53, BH17, GSW95, Kor79, Pye90, Ein21t, EM33a, Föl91, FTG08, Har98, Joh04, Koc01, LM94, LT15, Lin17, Som55b, Sta36]. Naturalist [Nag50]. Naturalness [vD02b, Cur92]. Nature [ARK^{+}99, Bir25, BH17, Edd28, Ein22j, Ein26d, Eks99, Gol76, Gor12, JW06, Kat05, ME14a, Mor96, Ser88, Sho21, Sud11, TE31a, Wig64b, Wig64c, Aer99a, ABW99, dCENdCA64, Bar02a, Bar02b, Bru03b, Die24, Die11, Ein09b, Ein09c, Gho17, JP55, JM86, KK75, Kum10, Kum11, Meh73b, Nor13, Obe16, Pat95, Wig64a, Wig65, Wig67a, Tay21, Nor00, Ein24g]. Naturerklärung [JP52]. Naturforscher [Sch49a, Sch51a, Sch55, Sch79a, Sch83a, Som55b, Sta36]. Naturforschers [Nic17, Nic19]. Naturgesetze [Wig64a]. natürlichsten [EM33a]. Naturwissenschaft [AH93, Rei41, RKS08, AH93, Ein39c, Ein41d, FTG08]. Naturwissenschaften [LT15, Lin17]. naturwissenschaftlerinnen [Föl91]. naturwissenschaftlicher [JP52, Müh11]. nauce [Inf56a]. nauci [Mag95]. naucnika [Sre29]. Nauheim [Van07c]. nauki [Inf33]. Naval [Ahe03]. Navigator [Ano82d]. navigatore [Ano82d]. Navy [Ano21a]. Nazi [Ama02, Bla02, Bow87a, Bow87b, Bro92, HM11, HEWW15, Irv67a, Irv83, LM90, Mac85, MP99a, MP01, MP12, Ric06a, Wal95, WHF^{+}96]. Nazis [Cou13, Seg03]. Nazism [Rei07]. NBC [Jan77]. Near [Ano29c, Jam30, WCL^{+}00, AAA^{+}18, Ano18c, Boz14, Ein10a, Ein05-29, FSE31, Len17]. nearly [Bry09]. Nears [Ano39f, Gib91a]. Nearsighted [Ein52e]. neben [Inf99]. Nebensächliches [Ank15]. nebst [WvLuAE18]. Nebulae [Hub29, Hum36, Zwi37]. necessary [Baw06]. Necessità [Mas02]. Necessity [Ein08a, Dro15]. Necrological [Kap91b]. nécrologique [dB55b]. Nederland [Ris06]. Need [Ano19b, Ein18l, Lev01b, Mas02, Ber08, Dor78, PP07, PP03, PP06]. Needed [Bet88]. needs [Mul10]. Neffe [Smo07a]. negative [Sta16]. Neglected [How92, Hof08b]. Negro [Ein47f]. Negroes [Ein32d]. neighbor [Rol89, NN01]. neighborhood [Pes13]. Neither [Vid17]. nel [Ano82d, Arg09, BF02, Mas99]. Nell [Ano60]. nella [Seg76, Seg83]. nelle [Cas14]. Nelson [Jar08a]. Neo [Pye13, D’A97]. Neo-Idealist [Pye13]. neo-Kantian [D’A97]. neocortex [KWGH97]. neokantianism [Hen87]. Neokantianismus [Hen87]. neptunium [Nod34]. Nernst [Som05a, Som05b, Kox06, Som05a]. neroutta [Hot15]. n’est [Bie91]. Net [Ano04f, Hea64]. Netherlands [Wal02b, Ris06]. network [CZPR17, HBFH18]. network-based [CZPR17]. Networks [Mau09, FW17, Hof18, Lar08]. Netzwerk [HBFH18]. Netzwerken [Hof18].

Neu [MTVB88]. Neubauer [Hop07]. Neuberg [Bjö01]. Neubesetzung [Kle75b]. Neue [Ein05a, Ein23a, Abr13, Ein06c, Ein08a, Ein11e, Ein16c, Ein05m, Ein05r, Rec02, RCD98, Sta95c, Ein28c]. neuen [Bäc19, Bäc21, Ein07i, Ein07c, Ein05-32, Ein05-33, Inf31, Sch21]. Neuer [Ein40d]. Neues [Jun05a, Row06a, Row06b, Flü74, Mos21, Wue05]. Neufassung [Dür80]. Neumann [Wig96h, GW96, Khr08, Pou93, Tau63, Wig96c, WG96, vN96]. neuronal [AH96]. Neuroscience [Bur15, Sti13b]. neuroscientist [SK09]. neutral [Chu98a, Phi98]. Neutrality [LSW12, Ene12, Fri12, Wid95]. neutrino [Bro78b, Kos03, Rei96b]. neutrinos [CT12, Dav03, CT12]. Neutron [Ano38b, Bro71, Fer39, Fer70, Fri39b, Kae39, Lau13, Stu93, Ao04, Bro97, Cha32, HS39a, HS39c, MF39b, Pei39, Str05b]. Neutronen [BJ41, HS39c, HS39b]. Neutronenbestrahlung [HS39a]. Neutronensterne [Str05b]. Neutrons [Gol72, MF39a, vHJK39a, vHJK39b, vHJK39c, BJ41, HS39b]. Neuve [Ber84a, Sto96, Sto96]. Nevada [FG00]. Never [Bru09, Ree17d, Ank15, Gio13b, Kno90, Ler91, Shu17]. New-Found [Ano33h]. Newcomb [Bra29]. Newest [Bro23]. Newly [Sch04b, Ros12]. Newman [Jon89, Mar08, NJP92]. Newmark [Rob48]. News [Ano98, Ano29f, Ano31c, Ano31d, Ano52, Ano04f, Ant15, Axe50, Axe52, Fow03, HES^{+}12, Jun05a, McC01b, MTVB88, Nad02, Wue05]. Newson [Sne82]. Newspapers [Ein52e]. Newstein [Sta07b]. Newton [Ano20a, Bes04, Büh98a, EI39a, EI61b, Mat86a, Mil12b, Pep04b, Pia31, Whi43a, Whi43b, Aĭn77, Ano19d, Ano27, Ano06b, Bai92, Bai01, BP95, Bru03a, Büh98a, Büh98b, D’A50, Dav09, DiS92, DiS06, Dry14, Ein19c, Ein27d, Ein27e, Ein28b, EI39a, EI61b, EI61a, EI62b, EI62c, Ein87a, Eis93, Eis07, FH02, Fri11, G.20, Gib97, GAHS96, GM00, Koc01, Lam21, Ler73, Lig04, McC11, Nik06, Pao84, Phi97, Qui05, SNA^{+}06, Sil23, Spe05, Str68a, VT99, Vec17, Vig93, Wil89, Yin15, dB36]. Newtona [EI62c]. Newtonian [CC67, Ein19b, Ein21e, Eis22, Fle19, GS22, LHF^{+}98, O’L64, PW14]. Newtonians [Bar15]. Newtonista [EI62b]. Newtons [Büh98b, Ein27e]. Newtonschen [Ein19b, Ein19c, Ein21e]. Next [Jam30, Smo07b, Ano11c, Che00, Inf99, MO95, Smo06]. nian [HL10, LR02]. Nicholas [Eds82]. nicht [CSW97a, Ein19b, Ein19c, How90b, Mau96, Voi86, How90b]. nicht-relativistischen [Voi86]. Nichtexistenz [EG23]. nichtjüdische [BS92, NS88]. Nichtlineare [MGG^{+}95]. Nick [Pri99]. Nickel [Hoy54]. Niels [Adl03a, Ano23c, Coh54, Duf46, Aas90, Aas02, Ano23c, Dör05, FF94, FK85, Fre86, Kle10, Kra79, Meh87, Mur90, OBxx, Pai91, Per21, RRK^{+}64, SM08, VW86]. niemals [Ank15]. niewiadome [Gol98a]. Nigh [Top07b]. night [Ema85, Van70, vS14, Swa02]. Nightmare [Ein17b]. NIMBY [Wel93]. Nina [Fre08]. Nine [CCD^{+}52, Whi55a, Van03]. Nineteenth [Uvn82, D’A01c, Har85, D’A00, FK90, Har79d, Pye82a]. Nineteenth-Century [FK90, Pye82a]. Ning [LY95]. Ninth [JRG02, Sch83b]. Nishida [Has07b]. Nishiwaki [NKY15]. Nitride [LU13]. niveau [Eöt00]. NJ [Ano84a, Boy88, Cah90b, L.89, Myr21b, Myr21c, Plo06, She15, Sta21]. njegova [Inf79a]. njihov [Inf79a]. No [Ano21d, BSPKN03, Ein14b, Ein50d, Enz02, Goo90a, Ken19, Kle80, KB60, KB83, KS95, TT70, Ein26a, Far03b, Flo06, Gol98b, Inf57a, Inf60, Inf61a, Kra06, Lee06, NHG89, Ano19b, Jan14b, Nic20]. No-bells [BSPKN03]. Nobel [Aba09, Adl03b, AA02, Alf09, And81, Ano78, Ano97a, Ano04b, Ano12d, Bad11, Bar05f, BCS82, BS83b, Bey06, Bjø20a, Bjø20c, Bri82, Bru03c, Bus12a, Cas05b, Coo04, Cur82, Dai67, Dan67, De 13, Del09b, Don70, Dor86, Eds82, Enz05, Far79, Fey64, For05, Goo85, Gre19a, Gre19b, Hah73, HBFH18, Han80a, Har79d, HMW24, Hea64, Hei79a, Hei79b, HSD34a, Hel83, Hes04, Hoc93, HL10, Ihd74, Jar08a, Jen01, Joh06, Jüt21, Kat52, Kau14, Kra02, Kri01, MFS^{+}6x, Mat86b, Mau09, Mei54, Mir94, Myr21b, Myr21a, Myr21c, Nad79, NN01, Nye88, Opp79a, P.50, Pau47b, Pla20a, Put09b, Ram82, Rec86, Rec10, Rei86, Ren91, RK94, Sch87a, S^{+}10, Sch79c, SS97, Bad12, Sei79, Sin07, Smi05a, Ste95b, Str04, Str05a, Sty94, Tat64, Tis52, Tre88, TF99]. Nobel [Tur05, UKKP04, Van22, Vis94, Wal94b, WM08, Wei94, Wid07, Wig64a, Wig65, Zel90, dP94, Aas88, Aas02, Aba09, Abr88, Abr01b, Abr04c, AE06, ABH^{+}55, ABH^{+}87, Adl12, Aka15, Alf01, AKD23, All02, AE01, AGMT12, Ang78, Ano97, Ano98, Ano00, Ano03, Ano11a, Ano12, Ano33c, Ano34b, Ano36, Ano45b, Ano45c, Ano56, Ano64, Ano65b, Ano65c, Ano68b, Ano72, Ano77, Ano82b, Ano82a, Ano85, Ano87b, Ano93a, Ano97e, Ano99f, Ano00d, Ano00f, Ano01d, Ano01a, Ano02b, Ano05a, Ano05b, Ano06a, Ano07n, Ano07m, Ano07p, Ano07q, Ano07r, Ano07s, Ano07o, Ano08d, Ano09b, Sch12b, Ano12b, Ano12c, Ano16b, Ano17a, Ano19, ACT23, AMB^{+}77, AO78, BZ10, Bad04, BTR08, Bad11, Bai06, Bal79, Bal75b, Bal77, Bar18a, Bar98b, Bar05f, BNSY98]. Nobel [B^{+}01, BR00, Ber62, Bet63b, Bet67, Bet68b, Bie03, Bjö01, BOS14, Bjø19, Bjø20b, Boh65b, Boh85b, Bor04, BB55, Bor75, Bor78b, Boy10, BS86, BH04, BH09, Bri95, BW^{+}87, Bri14, BL07a, BL07b, BL07d, BL07c, Bru06, Bun79a, Bus13b, But23, Cah04, Cah05, Cam09b, Cam00, Cao04, Cao14, CGG^{+}13, Car12, Car15, CÖT15, CT15, CMT18, CJT22, CZPR17, Cha05c, Cha17, CD23, CSC23, CGT11, Chu98a, CT98, CCD^{+}52, CW02, Cra84a, Cra84b, CHU87, Cra90, Cra92, CSW96, CSW97a, CSW97b, CSW98, Cra00b, Cra01, Cra02b, Cra02a, CK^{+}04, Dan67, Dar04a, Das00, Das91, Dav03, Day12, De 13, DCL23, DS83a, Doh06, DWL98, Dor10, DFB22, Dur12, Dys65, EGR11, ET87, Eks99, Elz95, Elz06, Ene12, Eng14, Enk18, Erc55, Esh09, EF69, FN38, Fan91, Fan93, Far53, Far63]. Nobel [FW17, Fei01, Fel82b, Fel00, Fer38, Fer39, Fer70, Fer08, Fie15, Fie16, Föl91, Fos13, FMR90, FN13, FM92, Fra08, Fre13, FK68, Fri81, Fri89, Fri01, Fri12, GM04, Gei11, Gen22, Gha82, Gia03, GNS99, Gil12, GW10, Gin04, Gla06, Goo03b, GMOP12, Gra49, GW83, Gre05b, Gro05, Grü08, Gun11, Gyl78, H.89, Hac98, Hal59, Hal06, Hal12, Hal04a, HM11, Hän06, HS14, Han18, Har02a, Har13a, Har13c, dVNHdVH53, Hea64, dVNHdVH71, Hee01, Hef10, Heh98, HF22, Hei32, HSD34b, Hei65b, Hel14, Hen13, HFN06, Hen03, Her66a, Hig14, HD02, Hof18, Hoi01, Hor09, HR17, Hün18, IP76, Jab11, Jab13, Jac12, Jam93, Jar08b, Joh11a, Joh05b, Joh14, JDB^{+}97, KKJ99, KKKM05, Käl22, Kao10, KM04, Kar10, Kat89, Kea18, Kee04, Ket02, Kir13]. Nobel [KJY24, Kob09, Koh99, Kor94, Kos03, Kos20, Kov01, Kra17, KWB^{+}02, Kri19a, Kro01, Kro80, Kuc11, KUW82, Kur02, KS81, LW34, Lag03, Lag12, Lar12, Lar01, Lar08, Lau99, Leg04, Ler03, Lev01a, LR01, LRH22, LHDH21, LLH22, LLSH23, Lin99, Lin01, LR14, LB02, LHZ^{+}23, Low01, Löw78, Lud57, Lun92, Lun74, MSYW12, MNT^{+}00, Mac01, Mag89, MFS^{+}6x, MH12, Mar91, Mas09, Mat07, MJ12, McC01a, McC48, McC01c, McC01b, McG93, MB94, McG98, MB^{+}92, McL99, Mea21, M08, MTVB88, Mil24, Mir11, Mon12, MS09, Mor95a, Mot79, Mou19, Nacxx, Nad02, Nag82, Nam09, Nam10, Nel77, Ng11, NN01, Sch62b, Nob64a, Nob64b, Nob66a, Nob66b, Ode72a, NB79, Nor15, Nor12, Nor02a, Nor10, Nor13, Nov11, Nye19, O’S01, Ode72b]. Nobel [Opf78, Orr78, OR21, PA21, Pau42, Pau45, Pau47a, Pau47b, Pau79, Paz10, Peg39, Per09, Per12, Per97, Pet08, Phi98, Pin00a, PLM^{+}0 , Pla20a, Pla20b, PS22, Pol05, Pop99, Pow94a, Pra12, Pri90, Put09a, RA13, RWL23, RWY97, Rie12, RF05, RN16, Rod22, Rog06, Rol89, SB82, SS98a, SS98c, SS98b, San15, Sch14a, Sch12a, Sch17a, SBB16, SS86, SS91, SS96, Sch12e, Bad12, Sch87c, S^{+}62, SK09, She15, SO02, Shi01, SAY^{+}82, SR99, SR01, Sin04a, SR04, Sin75, Smi92, Smi10, Smo07c, Söd10, SS29, Sou66, Sta93a, Sta89b, Sta79, Stê60, SSH13, SL93, Ste53, Ste07, Sti13a, Sti98, Sto98, Sto99, Str83, Str19, Sut97, TA97, Tan11, Tau86, Tau88, The95, Tho88a, Tho97, Tho12, Tho18, Tio09, Tol23, Tol24]. Nobel [Tol76, Tom10, TA17, Tor03, Tow65, Tsu99, TTB20, Van04, VM05a, Vel00, Vla79, Wad81b, WD96, WB^{+}87, Was92, WL80, Wei18b, Wes01, Wid95, Wig64b, Wig64c, Wig65, Wig67a, Wig82, Wil05a, Wil74, Wil13b, Wil69, Win13a, Wir08, Wol67, Wol16, Wor08, Wor10, Wyd15, YCL20, Zew02, ZZY19, ZZ23, Zuc67, Zuc77, Zuc96, Zuc18, ’t 00, ’t 02, vS78, Coh54, Ihd74, Sei84, Cah04, Coh54, Don70, Duv64, Hum51, Kat52, She15, Tra00, Wat00]. Nobel-eloadás [Wig65]. Nobel-Frauen [Föl91]. Nobel-Prize-winner [Bal75b]. Nobel-Vortrag [Jüt21, Pla20a, Wig64a]. Nobel-winning [RWL23]. Nobelevskoj [Was92, Dai67, GUB92]. Nobelist [Ano50b, Mar22, Wat97a]. Nobelists [BSPKN03, ZXLX14]. Nobelmen [Ano63a]. Nobelpreis [Ano51a, Sch12b, CSW97a, Hen11c, Hün18, MTVB88, Rec09, Rec10, Sch12a, Ano93a, CSW98, JDB^{+}97, Sch14a, Sch87c, HS14]. Nobelpreise [Ano61, Stu87, Hop07, Mei54]. Nobelpreises [HSD34a, Wil36]. Nobelpreisforschung [Han18, Hof18]. Nobelpreis” [TF99]. Nobelpreistrager [Joh95, ABH^{+}55, ABH^{+}87, B^{+}01, BR00, Erc55, HBFH18, Hei65b, Bad12, Sto98, Wyd15, Bad04, Her66a, Vis95]. Nobelpreisträgers [Enz05, SS97, Bor75, Hen21]. Nobelprisen [Sch50b]. Nobelpriser [NN01]. Nobels [Ano96d]. nobelska [Lun74]. Nobelstiftung [Mei54]. Nobelvortrag [Pla20b]. Nobelwoman [Ano63a]. Noble [Ano05b, Das00, Mir11, Ste07, But68, Jan95, Teu96]. Nobody [Ano19b, Mer85b]. Noch [Kle85]. Nochmals [Ein12b, Ein05e, Abr12a]. Nodal [Her49]. Noddack [Hab09]. nodes [vB19]. noirs [Sta92b]. nombre [Der15]. Nomination [DFB22]. Nominators [Bar05f, Nye88, Tre88, CHU87, Cra02a, Sin07]. Nominees [Bar05f, Nye88, Tre88, CHU87, Cra02a, GW10, Sin07]. Non [BB49, Bia16a, Boh62, Bra22, Fre08, Gra07, Hof49b, Kur79, Kva79, MBD39, Bus12b, Com95, Ein41a, EP43, Gis12, Kha12, Kle74, MDT^{+}03, Myr21b, Myr21c, NS88, NBDD^{+}12, PS10, Pot11, RDF^{+}08, See25, She16b, Voi86, Wol95, WB00, Ein26a, Hoh13]. Non-Classical [Boh62]. non-Einstein-like [MDT^{+}03]. Non-equivalence [Bra22]. Non-Euclidean [Bia16a, Ein26a]. non-Euclidian [See25]. non-existence [Ein41a, EP43]. Non-Ideal [MBD39]. non-interacting [RDF^{+}08]. non-Jewish [NS88]. Non-Linear [BB49]. non-localité [Gis12]. Non-Relativistic [Kva79, Voi86]. Non-Riemannian [Hof49b]. non-scientists [Wol95, WB00]. non-secret [Kha12, Hoh13]. non-séparabilité [Com95]. non-separability [Com95]. non-specialists [Bus12b, Myr21b, Myr21c, She16b]. non-statistical [PS10]. Non-Technical [Kur79]. non-uniform [NBDD^{+}12]. Non-Use [Fre08]. non-vanishing [Ein41a]. nondegenerate [Khr08]. None [Duf46, MW46a, MW46b, MW07]. Nonequilibrium [Kub86, SMHL23]. Nonexistence [Tur40b, EG23]. nonlinear [LHF^{+}98, McH07]. nonlocal [FTZ^{+}15]. Nonlocality [BC24, CW12, Gis14, Har97, dM86]. nonparallel [Kin95]. nonproliferation [FGMvH14]. Nonscientists [Fow03]. Nonseparability [Web93]. Nonseparable [Lau95, Kel58]. Nonsingularity [PS03]. nor [Vid17]. Norden [Kap61]. Nordström [Isa85, EF13, EF05, Isa85, Nor92, Nor07d, Wei12l]. Nordströmsche [EF13, EF05]. Norfolk [Ano79e]. Normal [Pla00b, Pla01]. Normalspektrum [Pla00b, Pla01]. Norman [Pes13]. norms [Käl22]. Norrby [Gil12, Kau14]. North [Adl03a, AS91, BCEP94, Gly90, HS89b, Ric06a]. Norton [Kar05, Lan15, Pye95, Sie97]. Norway [Ano63b, FN38, HM11, Hau54]. Norwegian [EI39c, Ene12, Hau47, Hau50, Hau53, Hau65, Hau83]. Nostrify [Wei14a]. not-so-easy [Fey97, Fey11]. Notable [Kha20]. Notas [Sch84a]. notation [HJ16]. Note [Din65, Hol88e, Kas20, Lau55b, Mac86, Nam10, NK84, PE49, Phi22, Pop66, Ree15a, Rey25, Rob38, BP79, Bro20b, Dim93, Ehl09, Ein07a, Ein18d, Ein23a, Ein84a, Ein05k, GG05a, Gan97, Hei35, vI11b, Lum11, Moo96a, NP88, PS22, Sch17a, Sin75, Ein18k]. Notebook [JNR^{+}07b, JRN^{+}07b, Ree05b, RS96, Ano07b, JRN^{+}07a, JNR^{+}07a, JR07, RS03b]. Notebooks [HRR03]. Noted [Whi55a]. Notes [AG13, Ano52, Ano79f, Axe52, Col13, Ein10b, Ein18c, Ein33e, Ein34j, Ein49a, JRN^{+}07a, Mor96, Pow94a, Ram79b, RS11b, Bar05f, EE05, Ein79d, Ein91c, Gen98, New60, Ōe81, Ōe95b, Ōe96, Sch79b, Woo28, Bla50, For80a, Han18]. Nothing [Clo09, Kra92, Mor95b, Bar00a, Bar00b, Fri12, Gef14]. Notice [dB55b, EdH15b, Ein23f]. Notices [vK65b]. Notion [Car07a, LC07]. Notiz [Ein07a, Ein18c, Ein18d, Ein23a, Ein05k, EdH15b, Ein18k, Ein23f].

Notizbuch [RS96]. Nottingham [Dix00, Ano30a, Ein30g]. Noulens [Lit91]. nouvelle [Chr22, Lon37, W.25]. nova [GdS15]. Novel [RMHD11, Ben16a, Gri15, Kan97, Mac01, MSK15, Mul05, Nes14, Sre29, Wat66, Wat70, Wat88, MSK15]. novelist [HM11]. November [LdB12b, Sto96, Tre66, Cho10a, Lin15, Räz16, Wei12s, Wei12t, Wei18a]. novembre [LdB12a, LdB12b, Sto96]. Novi [Mar95b]. novosadana [Mar95b]. Nowe [Inf33]. Noyce [Ber05]. Noz [Vog20]. Nr [Ein14b]. NSF [Wal83]. ntoon [BSW22]. nu [ER91]. nucléaires [EE82]. Nuclear [Ano68b, Ano94b, Ano01b, ARK^{+}99, BHT86, Ber93, Ber08, Bet88, BW39b, BEI^{+}55a, Bro82, Cal75, Can00, D’A93, Del09a, Div78, Dix00, D^{+}95, Dre07, Eck90, Fin90a, For01, FC07, Fre08, Fre81a, FP40c, Fri68, Gar08, Ger92, Ger97, Ger02, Ger07, Gib96, Goi91, Gol82, GMG^{+}77, GB88, Gou47a, Gra64, Hec12, Hei53b, Hei53c, Hei69b, Hen00a, Hol81a, Hoy54, Kap20, Kem12, Kit00, Kri19b, Lau55a, Lee06, Mah09, Mah14, Mah17, MHY95, Mal85, MGG^{+}95, May95, MF39a, Mei62, Mei63a, Mei63b, Moo10, Mor95c, NBDG40, Ree14a, Ree15b, Ree16c, Rin02, Ros72, Sea72, Seg89, Sem08, She83a, She83b, Stu94, Stu18, The12, Tra16, Tre79b, Tur06b, Tur40c, Wal04b, Wea88, Wee87, Wel08, Whi55a, Wig96m, Wol93, Yam09, Zuc82, Zuc83, vHJK39a, vHJK39b, vHJK39c]. nuclear [Aas90, Acz09, ABH^{+}55, ABH^{+}87, Ahe03, AG13, Ama02, And73, Ano96d, Ano00g, AA75, Arn00, Bad95, Ban03, Bau95, Bau97, BP79, BRVF84, Bet91, Bey49, BH01, Bro92, Bru03b, BP^{+}01, BCD93, Bus12c, CHW91, CG01, Cir07, Cir08, Cra07, CSW97b, DS83b, DG03, EE82, Far13b, FGMvH14, Fel79, Fel82a, Fra89, Fra04, Gal93, GC02, Gil62, Gla57, Gla62, GD77, GHK^{+}96, Gor07b, Gow88, GA71, GA81, Hac94, Ham08, Han88, HGS87, HCD^{+}78, Hen98, Her89b, HY15, HBO82, HW08, Hom06, Hoy77, Hoy79, Hug04, Hym12, Irv67a, Irv83, Kae48, Kee11, Kir05, Kra88, Lau12, Lec86, LM90, Mac93, Mag09, Mah15, McK03b, MTVB88, Mil91, MP99b, NKY15, New88, New89, Nic87, RC13, Ree05a, RS09, Ree11a, Ree15c, Rho07, Ric93, Ric06a, RSU93, Sch07a]. nuclear [Sch96c, Sta98d, SH01, Sto12, Sza84, TTHJ68, Tur03, Wal89, Wea12, Wel93, Wil75, WW75, ČA99, Her93]. Nuclear-Physics [Fri68]. nuclear-weapon-free [RSU93]. nucleari [Cal75]. Nuclei [CR12, Fri39b, Goe49, The46, Tur40a, Bai33c, Flü39]. Nucleon [GRE^{+}01]. Nucleon/Sube [GRE^{+}01]. Nucleus [FR13, Mei39, MF39c, Stu86, HS39b, HW49]. nuclides [Ree17a]. nuestro [Inf61c, Inf83]. Nukespeak [HBO82]. Null [PS04, Ryf14]. Null-Result [Ryf14]. Nullpunkt [ES05a, ES13]. Numata [Hir98]. Number [Bia16b, Hay91, vHJK39c, KWGH97]. Numbers [Ano15h, Bar02b, Der15, Fie16, Ree00]. Numeric [HJ16]. Numerical [AMSmS96, BS10, BBCPL09, CR07, Alc08, BPL05, EFH89, FGPP06, Pul11]. Nuobei’er [LR02, HL10]. Nuova [Mas99]. nuove [KT15]. nuovo [Ano82d]. nur [CSW97a, Ein27b, Ein52e]. Nutshell [Zee13, Haw01, Zee03, Zee10]. nutzbar [Flü39]. ny [Gar41]. nyertesei [BST03, BSST04, CST01, CST02]. Nylon [Ndi07]. Nymphenburger [Kle70b, Kle70a].

O [Lin60, Pye06, Urb22, Žáč11, EI39b, Gar42, GdS15, Inf56d, Jof62, Ken22, Kol40, See61, Smy46, Sou19, de 03, Ein22j]. O.M. [Eve39, Eve13]. o/and [Ken22]. Oak [JJ81, Olw97, Olw04, Ove87, Rob50]. Oath [Pol91]. ob [Inf56c]. obbediente [Bru07]. obecná [Ein23h]. obecné [Bič79b]. obedient [Bru07]. obeisance [Dav81]. Oberflächen [Ano17b]. Obey [Kap40]. Obituary [Ano55d, Ano00e, Bro10b, Kop64, Tel76, vK65b, Ein34i, dB55b]. Object [CLD^{+}03, Kra21]. Objection [Hoo70, Jam74c]. objections [Ein18f, Pat95, Kra20b]. Objective [McC57, Bel75]. Objectivity [Hoo91, Ash95, Dor95]. Objects [Sta05c]. Objekten [GRE^{+}01]. Obligation [Ein46f, Ein50m, Ein52c, Ein79f]. oblivion [Bre08, Yor70]. obosnovaniiu [Ein40g]. obra [Inf57a, Inf60, Inf61a, Inf61c, Inf83]. obraz [Ein35e]. Obscure [Ano79b]. Observability [AP00, ST94b]. observables [Sau07c]. Observation [IAS^{+}98, Rob49b, GSPH00, LB39, She88]. Observational [Cre83, Dic81]. Observations [Ano19b, BBB^{+}09, Ben05, Ber01, Ein11h, Ein14j, Ein14h, Ein86c, Ein87b, Hir97, Mat40, Bau20, DED20, Eĭn85b, Eĭn86a, Ein22m, Nor72, Per23b, Sch12e, Yag85, vE10, Hof18]. Observatories [JSSZ87]. Observatoriums/Neu [MTVB88]. Observatory [MTVB88, Cul80, Dic80, Fab79, GT80, Gia80]. Observatory/News [MTVB88]. Observed [Col95, JDB^{+}97]. observed/Laser [JDB^{+}97]. Observers [Kov20, Sta05c]. Obsession [Mil10b, Mil05a, Mil05b]. Obsessive [Gol05]. Obshchedostypnoe [Ein21r]. obstacles [Bor02]. Obtaining [CW32a]. occasion [Aĭn77, Ano80a, BRVF84, FPLR59, Hal55, Lin60, MB66, NHP^{+}79]. occasioni [Seg91]. occidental [Tym95]. occupation [HM11]. Occurrence [BHJ^{+}00]. Occurring [Hoy54, Hah42]. Océ [MTVB88]. Océ-van [MTVB88]. oceans [Ham08]. ocima [Hom04]. October [Ano81d, Ano93c, Ano13a, BPS87, Cur82, KK05, LdB12b, Ber84a, Shi49, Wei12s]. octobre [LdB12a, LdB12b]. Od [Inf55d, EI59, EI62a, EI62c, EI98b]. odd [Bad09]. Odds [Kri99, Lau13]. Odvození [Žáč12]. Odyssey [Dry14, Juc04, Sch13a, Abr01a, Abr02, Abr03, Abr04a, Yor87]. œuvre [Ano48b, dB49d, dB49c, dB55a]. off [Rod05]. offenbar [HS14]. Offene [Kra25]. offenen [Büc21]. Öffentliche [GHK^{+}96, Waz09]. öffentliches [Ber69b]. offer [Ein23c]. Office [Lil86, Tre88, Baw06, Reg87, Sch93b, Wei12e]. official [Bad11, Bad12, Smy45b, Smy76a, Smy78, Smy89]. Officials [Res50]. offsets [Kro01]. ofiëtìsialsnyi [Smy46]. Oh [Mil05c, Mil06, CM67, Lid67, Tur66, Tur67]. Ohanian [Mer12]. Ohio [Pin00a, Tho17]. Ohm [Ser88, JM86]. ohne [Ein20g, Ein52e, Mar17]. Ohnesorge [Sta98d]. Oil [Fra81, Tol76]. oincidence [Ryc92]. oint [Ryc92]. ökonomisches [vN37]. Old [Fin76, Jun08, Lor07b, Mar95b, Sha92, Ber06c, Ein60b, Hei49a, Lau98, RCD98, Sti05, Zee89, Zee90]. older [EI38d]. oldest [Gut98]. Oldie [Ehl09, Lum11, Wey09]. Olga [Pye89]. Oliveira [BHJ^{+}00]. Olivier [Har02b, Abi04, War03a]. Olson [For05, HEB^{+}80]. olvassa [Tőz15]. olycklige [Fre13]. Om [Jak13, Hau49, Hau53, Hau63, Hau64, Hau65, Hau83, NN01]. ombro [T– D12].

Omega [Bar02a, Bar02b, Sei03a]. omission [Hen13]. omitted [HM11]. Once [Dam06, Tre75b, Kle85]. oncle [Sta92b]. onde [BES86]. Ondes [dB49a, dB49d, dB55a]. One [Ano29d, Ano35, Bar05f, Ber87b, BLR15, BFMNZ91, CC96, Čvo06, Gib19, Gla06, Kae31, Kra23a, Lau55a, MW46a, MW07, Max93a, Meh98, Ne’06, Nic34, OONM18, Ren05c, Ren05b, RFP^{+}45, Sta09a, Wyd84, Ano80a, Ber06c, Cas03b, De 13, Gio14, Gla00, Gla07, GMOP12, Hau47, Hau50, Hel80, Hen03, Hot15, Jar08b, Kot93, Mau96, Mee03, MP99b, Nacxx, NHG89, San07, Vil06, Ein48i, IRW31, MW46b, Duf46]. ones [Sub03]. ongoing [Bus12c, Sch13b]. Onkel [Sta95c]. Online [Ein03d, Ano03a, DC12, Hoa03, CF98]. Only [Ano29b, Ein46e, Ein52e, Hes04, Ano19a, Ano39g, CSW97a, Diz05, Ein27b, Hen88b, HD02, Wei18a, Ano29d]. Onnes [RA13]. onruststoker [vC05b]. onset [XMY^{+}09]. op [Hoe18, vC16]. Opalescence [Ein05-29, Tes69, Ein10a]. Opaleszenz [Ein10a, Ein05-29, Tes69]. Open [Ein47c, Ein49c, Ein50n, VFIS48, Ano81a, Büc21, DSZ12, NBDD^{+}12, Opp55, Opp63, VT22, YCKG11, Kra25]. Opening [GO08, Ein08b, MFO14, SM41, vL55b]. Opens [Ano29d]. Oper [Des72]. opera [Des72, OV86]. Operate [Car04a]. Operation [Ber06a, Fra93a, Fra93b, Gia09, Wig86]. Operational [Bas65, Bri49]. Operations [WL12]. Operative [Ein70b, Ein49d]. Ophir [Pye80]. Opinion [Ein15h, Ein16k, FR49b, Stu14, Stu23, Ein24g]. Opinions [Ein54b, Ein73, Ein82c, Ein94c, Ein95c, Ein31d, EI38d, Hah81]. Oppenheimer [Ber03, Bey09a, Bey09b, Dys76, Her03, Sch03c, ACU^{+}54, Ano54c, Ber82d, Ber04, BieWSK67, BS05, CH05, Cas05c, Cas07, Che66, Con05, Els95, Her02, HHMW93, Inf78a, Kai02, Kel06a, Ken54, McM05, MS15, Opp79d, OBxx, PC06, Pol02, Rou64, Rou65, Sch00d, Sch06, Sch08, SW95, Tho06, Val06, Wes14, Yor89, Tho11b]. Opponent [Gui18, Sta20]. Opponents [Hen16, Nav15, Waz09, Waz14, vD10b]. Opportunism [How98]. Opportunity [Yan06, Yan13a]. Opposition [Nye99]. Optical [ARK^{+}99, CR12, Hau08, Lev01b, CSD^{+}07, Hal06, KSVW10, MO06, Str19]. Optics [Per00, Eis07, Mas74, Ryn05]. Optik [Ein06d]. Optoelektronische [OSWR86]. Oral [Mas95, San95, Wei88]. orange [Aer99b]. Orbit [Lan02, AAA^{+}18, LJGS11, Mes83, vB19]. orbit-coupled [LJGS11]. orbitals [Sla79]. Orbiting [RA11, WTF81]. orbs [Gra94b]. Orchester [Ein34m]. Order [Cap01, Mer38, NN75b, Hol79a, Lan65, Lan67a, Bro66, Wre66, Hou65, Mal65, See65b, Bri66, See65a]. Ordered [MSS22]. Ordinary [Gol87, McP95, McP97, Seg11, EM34]. Ores [Bla13]. organic [Mac01]. Organisation [Hil39]. Organization [Hil39, Mer79]. organizational [HF22]. Organized [Hu04, Ano67]. oriental [Tym95, Tym95]. oriental/occidental [Tym95]. oriented [Max93c, Max93b]. origen [Pye86]. Origin [Ano94b, ARK^{+}99, Bir25, Cor07, Ein34j, Fer08, Kes65a, Kes65b, Pla22b, Pla10, RSS97, Sch00b, Sci53, Stu94, TVL^{+}09a, Val49, Abi03, BS83a, BS93, Bar94a, Gol84c, Gow88, Ler91, Mas05b, Pen79b, Ren97a, Sau08, Sin04b, Ano29b, Din65]. Original [Ano84a, BDR93, CE79a, LEMW52, Bro19, Bro20b, EM20, EI66a, GR15, LEM^{+}23, Thi46]. Originalarbeiten [BDR93]. Originalliteratur [Thi46]. originally [Bey49]. originaux [CE79a]. Origins [Abi91, Bey12, CR08, Del11a, Ein33e, FG00, GR96, Hen09, Hir69, Hir76, How94b, Kat05, LP05, LB90, Mar09, Mar10, Meh87, PJMS11, Set10b, Set10a, Top89, Bar00b, Byr81, Dar96, Gut97, HY15, Hol73, Hol88c, Hol88d, Seg11, She87, Sta08a, Van98]. Osborne [BE69]. oscillating [SLLS09]. Oscillations [BSSP^{+}08, FH10]. oscilloscope [KS81, Bri82]. Osers [Lig97]. Osgood [AS91, HS89b]. Oslo [Bor04]. Ossietzky [Sta79]. Osten [Ein19d]. Osterling [Kat52]. Österreichische [Ano93b]. Ostwald [DFB22]. ot [EI66a, Rol19, Har06a]. otchet [Smy46]. Other [Ano31b, Ber18, Bir23a, Bro21, Edw00, Fus88, Gis14, Gru07, Gup57, Ham97, Kea12, Kni12a, Kni12b, L.89, Pin15, Rös95, Ros12, Sta98b, Sty94, Ano28, Ano15d, Ben16a, Bri95, Bru97a, Bru97b, Cha05b, Coh79, Ein31d, FL88, Fus90, Gol95b, Gol98a, Hei89, Hol86, Hol95b, Hol96b, Hol00a, Lev16, LE96, MGG^{+}95, Mar96b, Mar11d, Mer16, Num09, PGH84, PGH96, Pia29a, Pia29b, Red26, Ree17a, Rei72, Sch96c, Sch56d, Sch56b, Sch56c, Sin10a, Smo07a, Szi92, Vil06, Whe79c, Whe79b, vD10b]. Otherness [Gra01]. Otherworldly [Col93, Zal91]. Otnositel’nosti [Ein21r, EI66a]. Otowi [Chu60, Chu98b, Chu01, Wat66, Wat70, Wat88]. Otto [Ano45b, Ano69a, Ano79h, Ano79g, B^{+}79, CSW97a, Ern92, For80a, GRE^{+}01, Hah81, Kap61, Lam79, Lin60, Mei05, RSPS49, Ano68b, Ano69a, Ano79g, Bad83c, Bad83d, Bau74, B^{+}79, BR00, Ber69b, Ber70, Ber74d, CSW97a, EHM^{+}79, Ern92, FPLR59, Ger69, GH84, Hah66, Hah67, Hah79, Hah81, Hah88, Her79a, Lam79, She83a, She83b, Wal03, Wal06b]. Our [Bai75, Can16b, Ein25f, Ein38b, Ein39e, Ein46a, Ein91g, Eng17b, Hol82, Kar05, Kea12, Lev07a, RFP^{+}45, Roh87, Sch50a, See60, TVL^{+}09a, Wil16a, Ank15, Ano50a, AA75, Bar06, C.51a, Can15b, Cas95, Cas04, Cle09, Diz05, Egd14, EL09, Ein09b, Ein09c, EdH15b, ES46a, EL05a, Fay10, Ger69, GH84, GB06, GB08, Her17, Inf49, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf58, Inf61b, Inf61c, Inf79a, Inf83, Kak04, Kak05, Kan95a, Lem09, Liv13, Moo02, Myr21b, Nag50, Ove87, PA06, Rol89, Sch12d, Uda94, Uda98, Wei04, Wil75, Won05]. outer [Bar99, Lev16]. Outline [Ein14e, Tay21, Edd20b, Edd29b, Edd59, Edd66, Edd87, Ein21b, EG13a, EG13b]. outlines [Car22c]. outlook [Ash12]. outputs [GRS16]. Outrageous [Har95, Tho93, Tho94]. outsider [Das12, De 92]. Outskirts [Bri05d]. Outstanding [Lau55a, Whi27, Ano97a, WK12]. ouvertes [A^{+}09, Del13]. ouvrage [Wei15c]. ove [Seg86a]. ove-ad [Seg86a]. ove-tagliyotehe [Seg86a]. Overbye [Smo07a, vB01]. overtook [Far13b]. overview [Ano97a, Das00, JN14]. Own [Ano15b, Ein05-45, Sch54, Ein93c]. Oxford [Alf09, Bar05f, Ber07a, Ber00, Goo85, Har02b, Kri08, Lan08, Nau17, Noe11, Red84b, War03a, Ano31e, BF14, Cal03b, Cru79, Fox15, Fox18, Hei05, Ken05c, VT22]. Oxford/New [Bar05f]. oxygen [Kon15]. Ozark [Mah14, Mah15]. ozone [Ano97e].

P [Ano54a, Bes93, BH89b, Col93, Ded85, Duf81, Ein11f, Ein14b, Ein14f, Ein25b, Ein05j, Fey64, FF95, Fey00, Fey05, For80a, HH04b, Hen82a, Hen82b, Kle82a, Krö53, Lin60, Mer85a, Myr21b, Myr21a, Myr21c, Pye80, Ryc92, Wig67b, Wig96j, Wil36, Wol67, Woo28, Zal91, dP94, Ano50a, Sto17]. P. [Ald93, Bur94, Cas94a, Din93, DD95, Ein11f, Ein14b, Ein14f, Ein25b, Ein05j, Kan94a, Mil93, Nye99, Wei94]. P.O [Wac06]. Paarläufe [FS06]. Pablo [Fol17]. Pace [Ano20d]. Pace-Maker [Ano20d]. Pacific [Fei61, Has07a, Ken54]. Pacifism [Ein22h, Ein31i]. pacifisme [Dur12]. Pacifist [GC96b, Lau55a, Sal19, GC96a, Jay05, Rot79]. pact [Cor03]. pad [Gib91a]. Page [Sau07b, Sau05]. Pageant [Ano60]. Pages [Ano29b, Alf09, Bed01a, Bru03c, Bru06, Eis00, For01, GRC18, Ham11a, Noe11, Par06, Pri99, vB01, vB02, Bäu83, Woo28]. Paid [Moo16]. Pain [Ano21d]. Painting [Fol17]. pair [FS06]. Pais [Ber00, BHJ^{+}00, Kle84, Red84b, Smo07a, Ano00e, Asi83a, Fer82c, For84b, Gil84b, Gla00, Gol84a, Hei83a, Lak85, Lar84b, Lov86, Pei83, Pip83, Pol83, Red84a, War84]. paix [Dur12, ER91, Mar91]. Pajarito [SG11]. Pakistan [Abr09, FC07]. palace [Con02, Con05]. pale [HB05]. Palestine [Ros18, Bor04, Ein21m, Ein23e, EK44]. Palevsky [Bed01a]. Palgrave [Jon01, The12]. Palle [Jun05b, Woo06]. Pametni [Pát79a]. Pamphlet [Par05]. Pamyati [Jof55]. Pan [Ein25g]. Pandora [Van03]. Panel [Law07, Ano51b, GHK^{+}96]. Paneth [Lin60]. Paolo [Pye95, Tug98]. Papagei [Hes15b]. Paper [AP65, BRI^{+}21a, Cah90b, CF82, Dra14a, Ein12f, Igl07, Kox95, McC68, Mil92, MG02, RSS97, Van11, Wei13b, Whi04a, AMS04, Cas07, Ein07b, EL08a, Ein11a, Ein14f, Ein15b, Ein16n, Ein20d, Ein22b, Ein25b, ES46a, Ein05c, Ein05-38, Ein05y, For01, Ham11a, HMW24, Isr98, Ken05a, Kle63, LG97, NK84, Pal83b, Pra06, Pri99, Ree15d, Rei96a, Rig05b, Rob38, Rya06, Sch77d, Yag85, vB02, ZDSF11]. Paperback [Sta99c, Ste90, Coo04, Dix00, Fis02, For05, Gil12, GRC18, Hea13, Hen88b, Mil11a, Myr21a, She15, Ste22]. Paperclip [Bow87a, Bow87b]. Papers [Ano75, Ano15b, Bad83b, Bad15, Bai89, Bal91, Bal96, Ber07a, BF86, BRI^{+}21b, Cah90b, Cah90a, CS79a, Dry14, EKS^{+}08, Gol87, Har91, Har96a, Hav88, Hen88a, Hen88b, Jos95b, Ker89b, Kil95, Lac08, Loe88a, Loe88b, MSYW12, Mar11b, Mar11c, Mcc87, Nat81, Pye89, Pye95, Pye97, Pye06, Ram90, Row00, Smi89b, Smi89a, Smo07a, Ste90, Str90, Wal81, Wal82, Wal83, Wal88b, Wal99, War96b, Yan87, dP94, vN96, Ano81a, Ano99b, Ano99c, ABdB^{+}53, Bel87, Bel04, BF78, CR07, Cha05a, Dre07, Eur69, Eck70, Eck73, EM20, FS87, For82a, For84a, For86, FCKT11, Gea90, Har02c, Har96b, HMW24, Haw11, Hoa03, Hol05b, Ken12, Kev05, Kil98, Kle59, LFKW06, OBxx, Rabxx, Ren05d, Ren05f, RN16, Rom91, Ros98, SMvdVS79, Sch82a, Sch03b, Sie97, Sie07, Sta87, Sta98a]. papers [Sta05b, Sta00, SO72, TH79, War95, War96a, Wei12i, Wig96m, Yan13b, ZXLX14, vH62, BDR93]. para [GdS15, de 03]. Paradigm [dB87, TPM86, Wil05a]. paradigms [LM94]. Paradise [Hol03b]. Paradox [BC24, Bha86, Bro22, Car22b, Com23, Erl72b, Gen23c, Gen23a, Gen23b, Hol88e, RR89, Roh87, Sch71, dB87, All05, Ano10b, Bel64, Bru97a, Com95, Cos80, CPDH^{+}87, Ein11c, Kau85, Mer85a, RDB^{+}09, Sau07c, Sel88, SW87, Van74, Wil05a, dB65, dM86, Ano05-30, Nag71, Sac85, Web93, Zuo96]. paradoxe [All05, Com95, dB65]. Paradoxes [McH07, Van98]. Paradoxon [Ein11c]. Parallel [Gre11, Har07, Mlo01, Pic05, Gri10, Žáč12]. Parallelism [For80a, Kil80, LC07, Ano84a, CE79b, CE79a, Deb29, Deb79, Ein28e, Ein30a, Ein30r, EM31b, IR85, Sal74]. parallélisme [CE79a]. Parallelismus [Ein30a, EM31b]. parallelních [Žáč12]. Parallels [Kuz79, McH07]. parcourus [Ein10e]. parent [Mau96]. Parikia [Pap15]. Paris [Bad11, LG97, Bie87, Bie91]. park [Kra15a, Con02]. Parkes [Kni13]. parla [LS08, Per10]. parle [Ano48b]. Parmenide [Pis15]. Paros [Pap15]. Parrot [Mir04, Hes15b]. Part [BDR93, Ein41e, Kra23a, Kra23b, LL82, Max93a, Max93b, MR87b, MR87a, Rai14, Smi89a, vM96, vM99b, vM01, vM05c, Ano16b, Cla61, Eis00, HS44, Hei69a, Hei72, Nacxx, Ano07p, Ano07q, Ano07r, Ano07s, BL07a, BL07b, BL07d, BL07c, Feu71a, Feu71b, Gol96, Max93c, Sch52b, Sch52a, Sch77a, Sch77b, Sch77c, War92, War93, Wig96m, Yer03, Yer04]. Partial [Sch17b]. Participant [Sta98c, Yor70]. Participants [Pin00a]. participation [BEM12]. Particle [ARK^{+}99, Clo04, ER35, Hen86a, Kle64, Lar19, Raf17, Ros49, SMHL23, Ano51b, Bac89b, Bai92, Bai01, BDH89, FTZ^{+}15, GG05a, HFN06, HBRD97, Kan95a, Nam09, Nam10, NM12, Rei96b, Ric06b, Sam10, Sch96c, Sch04c, SDCZ01, Van07a, Vel03, Vig94, vD02a]. Particles [Bac89a, Bha49, Bor49b, Fei70, Hei93, IS49, Pan94, Chu98a, Ein05c, Ein19j, EI49b, Fis10, Fis12, Gou10, Gou13, Hal09, Hei89, Jos80, Pai80, Ran05, SS97]. particularism [KJY24]. particularly [SE28]. particules [Gou10]. partisan [Ano55c]. Partisans [Sal19]. partly [Ein20l]. Partnership [Jen01, Tau12]. parts [Hen86a, Won05]. Party [Cha23]. Pasadena [Ano31b]. Pasado [Paz10]. Pascual [Sch05c, Sch05c]. Passage [Pei85, Wei13a]. passages [Ran05]. passenger [Ano18a]. passes [Mer10, vS21a, vS21b]. passing [FSE31]. Passion [CHS97, COS05, Hän06, Jon08]. Passion-at-a-Distance [CHS97]. passions [Hol95b, Hol96b, Hol00a, Par03]. passive [RKD64]. Past [AZ05, Ber85, ETP31, Mar22, Paz10, Kao10, LE96, Lui11, RWL23, Ham97]. Pasteur [Tho88b]. patch [Nel77]. Patent [Ano50b, Cho10a, Lil86, PB79, TA97, Sch93b, Wei12e]. Patente [Mel78]. Patenting [Wel08, Tur06b]. Patents [Maa13a, Maa13b, Mel78, Ill12]. Paterson [Ano60]. Path [JP08, KN08, PS79, Sch95b, Wil13b, Bič79b, FH65, FHS10, Lan72, Rei96a, Rie12, Rob10, Sau08, SH99, Wei12i]. Pathfinder [Got80]. Pathway [Wei12r, RD05, Wei15a]. Pathways [RS07a, Inf33]. Patricius [Rya06]. Patrick [Dow80]. patriotism [Gro56]. pattern [CD23, TEM^{+}89]. patterns [CÖT15, HF22, OB05]. Paul [Bla50, Bon94, Coh51, Ein07a, Ein34i, Ein05k, Gea94, Hea13, Jeh51, MTVB88, Nag50, OSWR86, Pye80, Ram79b, Smi51, Smo07a, Wig96a, All02, Ano79b, Ano05l, Bok08, Ein34i, Ein79a, Ein05b, Erm89, Far09, FCKT11, FF97, Hün18, Kle59, Kle70d, Kle81, Kle10, Kle93, KW87, LG97, MTVB88, Sau16, Wig96k, Wit82, vD14]. Paul-Scherrer-Institut [MTVB88]. Pauli [APWB95, ARK^{+}99, Ein22q, Ein79e, Ein85e, Eis00, EM88, EGO97, Enz05, GRE^{+}01, HEB^{+}80, HvMW79, MGG^{+}95, OSWR86, Uns00, vEA02, vMHW85, vM93, vM96, vM99b, vM01, vM05c, Ano45b, Ano45c, Ano93b, APWB95, vBT94, DPS97, Eur69, Ein22q, EM88, EvM94, EGO97, Enz02, Enz05, FW07, FW60, Gie05, Hen81, Hen84, JP55, Lau88, Lau97, Lin04a, Mas05b, Mil10b, PHvMW93, Ser77, Uns00, vE95, vEA02, vM00, vMHW85]. Pauli-Jung-Dialog [MGG^{+}95]. Pauling [Ano63b, Don64, Par99]. Pavia [San77]. Pay [Gri93, Gia09]. Payne [PGH96]. pays [Kle16]. paz [M08]. PB [Mar10, PJMS11]. PBR [Wei13g]. Peace [Abr88, Abr04b, Abr05, AL09, Ang78, Ano16b, Ash11, Bal79, Bor04, Bor71, Bun79a, Ein34b, Ein35f, Ein45g, Ein47a, Ein53o, FN38, Fel82b, FA97, Hab82, HA99, Hef10, Hol96a, HL10, Joh11a, Koj08, Lan08, NN75a, NB79, Nor12, Orr78, Pau79, Pea58, RFP^{+}45, Sal19, Sta79, Sti13a, The95, Wil13b, vS78, Ano12b, Bes93, BK05, Ein84b, ENN60a, Ein60a, ENN60b, ENN63, ENN68, ER91, Ene12, Her33, Her63a, Her67, Her75, HH89, Hir98, Mar91, NN60a, NN60b, NN81, New88, Pau42, Pau47a, RS07f, SS29, Yor87, Abr01b, Kap61]. Peace-building [Hol96a]. Peaceful [Sal19, Fin90a, Rol89]. Peacock [ARK^{+}99]. pedagogical [BP79]. Pedagogy [DFB22]. pedestal [Dor10]. Pedigree [DRB05]. Peebles [Myr21b, Myr21a, Myr21c]. Peek [How97b, Jak08]. peer [Beh05b, Ben99, Beh05b]. Peer-Review-System [Beh05b]. peerless [Put09a]. Peierls [The12, Lau12]. Pending [Res50]. Pendragon [Per10]. penetration [Wes04]. pensamiento [EI65a]. pensar [Bal99]. pensée [Bal93, Bal95a, Bal11a]. penser [Smi15a]. penseur [Geo53, L.58]. pensiero [Bal94, Ros80]. Pentagon [Wei06b]. People [Lib79, Ano06k, BB13, FL88, Hei89, Jam06, aM07, Pas10, Pea49, Rob50, Bro21]. perceived [YCL20]. Perception [Lee06, Yan13a, Hol79a, Yan06]. Perceptions [Asa95]. perceptual [Per91]. Perché [Aba09]. Percy [C.51b]. Peredacha [Sla47]. Peregrinations [Wad81a]. Perfect [Fer14c, Fer14b, Fer14a]. perfection [Dir79f]. Perfectly [Fey05]. perfekte [Fer14a]. Performance [Gil05, Ber69b, CMT18, Van96]. Pergamon [Goo85]. Perihelbewegung [Ein15e]. perihelical [Ein15e]. Perihelion [Van10, Van11, Wei14b, Ros82c]. Peril [Whi55a, NKY15, Smi65, Smi70]. perils [Kea18]. period [Bra15, Gru05c, Tal32b, vB19]. Periode [vB19, Bra15]. periodic [Ein19b, Ein19c, Kau14, Lag12]. periodical [Zia94]. periodische [Ein19b, Ein19c]. Peripatetic [Hen05]. Peripheral [Ano07f, Das12]. Periphery [Sch95b, Hei69a, Hei72]. permutations [Bus12c]. perpetual [Rin03]. perplexed [Bre79]. Perrett [Ken25, Mur25]. Perrin [Bru68, Nye72]. persistent [Haw07, RW07]. Person [Ein52e, Cho16b]. personable [Ros05]. Personaggi [Seg76, Seg83]. Personal [Bed01b, Ber74a, Ber01, Com56a, Nee79, Ree15d, Tal32a, Bet91, Com56b, For15, GBE95, Hah81, Jun58, Jun82, Kai94, Nic87, Rin03, Ros98, Sch82b, Uda94, Uda98]. Personalities [Ano21m, Gin01, Seg76]. Personality [How94a, Ber69b]. persönlich [Ros05]. persönlicher [Com56a, Hah81]. Persönlichkeit [Ber69b]. Persons [Ano29d]. Perspective [Ang78, ABIM15, Bal79, Bun79a, For14, Gru97, Jam74d, Kil84a, Nan78, NB79, Orr78, Pau79, Pea08, Sta79, TWD03, Wal04b, vS78, Bet91, Bro05, CGG^{+}13, Fri81, Gib91b, Kil84b, Kun76, Nye72, OONM15]. Perspectives [Ano83, Ano01d, Bad83a, Bar98a, BL14, Bla92, BH17, Fer83, GT80, Gil84a, Kel04, Lar84a, MB66, Sch84b, Wil84, Cha78, Ein22k, ET07, FCKT11, GRZ99, HE82, HL10, Hug04, JP08, KN19, Vog20, Gol84b, Ano84b]. Perspektiven [Tre66]. pertinence [G^{+}11]. perturbation [BCO^{+}11]. perturbations [PKLT17]. pervonachalnykh [EI66a]. pesanteur [Eöt00]. Peter [Ber06b, Kan96a, Lam07, Pye89, Rei86, Smo07a, Alt06b, Kan96a, Lan00, RS06, Yer03, Yer04]. petite [Mar91]. Pg [Cas94b, Ein86c, Ano05j]. Phalén [Fra37]. Phaléns [Fra37]. Phänomen [Hag05]. Phantom [Mil08]. Pharmazeuten [HEB^{+}80]. Phase [Bar49, Hol89b, Kle70c, Vig94]. Phenomena [Ano29d, DR08, Dor71, Ein05u, Jam30, Cam18, Fri39a]. Phenomenalism [BIT01]. Phenomenological [Gio14, Arg09]. Phenomenology [Tym95, Sta07d]. Phenomenon [Ano39b, Lon38a, Ano10d, Ein01, Hag05, Man55, Mus15b, Spo16]. Philadelphia [Hea13]. philanthropy [Aas90]. Philip [Bot13]. Philipp [C.50, For80a, Gun15, HEWW15, Sta36]. Philos [Flo06, Kra06]. Philosoph [Sch49a, Sch51a, Sch55, Sch79a, Sch83a, Som55b]. Philosopher [Ano15i, Bas92, Can16b, CS70, Dys15, Eng17b, Fol96, GA02, Hen81, Her17, How05b, Sch49b, Sch49c, Sch49d, Sch51b, Sch70, Smi51, Wil16a, Zei99, Cam09a, Can15b, Cha09, Ein93c, How05a, Sch49a, Sch51a, Sch55, Sch79a, Sch83a, Som55b, Krö53, Coh51, Jeh51]. Philosopher-Scientist [Sch49b, Sch49c, Sch49d, Sch51b, Zei99, Krö53, Coh51, Jeh51]. Philosophers [Krö53, Acz03]. Philosophic [Bac49, Kor79]. Philosophical [Ano23a, ABM79, BH17, CM89, DiS06, Din49, Gri79a, Gri79b, Gru57, Grü73b, Grü73c, Hei53a, Hie68, HW07b, How92, Kra20b, ME14a, Rei49, Sch15, SMvdVS79, Zuo96, HBD71, Kat92, Lan37, Lau88, Mil90, OOM21, Pol89, Urb22, vM98, Qui05]. Philosophie [Bar94b, ER13, HRP90, HH14, Rei27, Rei28a, Rei28b, BD92, Fra49d, Hen86c, Her77, RK77, LLH^{+}80]. Philosophies [Ano23d, Mat86a]. Philosophische [Kra20b, HBD71, Sch15]. philosophischem [Pol89]. philosophischen [Spe68, Urb22]. Philosophy [Alt06a, ABdB^{+}53, BIT01, BC71, BE07, Coh80, CHR96, CCG93, CF98, Cun22b, Ded85, Ein48j, FF94, Fra49c, GSW95, Göd49b, Gol76, Gri81, Grü62, Grü73a, Hin49, HOO88, Hor79a, How84, How04, Jam74d, KV11, Mur90, Nav94, Pye80, Rei27, Rei28a, Rei28b, RF57, RK77, Ryc05, SS16, Skl85, UM86, UM92, Ush49, vW79, Wen49, Zho85, vD15, AS79, Ano05d, Bag04, Bar99, Bar94b, Bel87, Bel04, BD92, Bit96, Car22c, Car13, Cas75, Cha09, Coh49, Ein50b, Ein05z, ER13, EvM94, ET07, Fra49d, Fri83, GHW09, Har98, HHJ^{+}06, Hen86c, HRP90, Her77, How14, Joh04, Mik65, Mor16, MM66, Mun36, Pai91, Pal83b, Pla59, Rei58, RS07e, Ron07, Rya06, Sch79d, Sta16, Wan95, You91, d’E06, C.50, Boh62]. Philosphers [Acz01]. photo [SSH13]. Photochemical [Ein12f, Ein12h, Ein13a, Ein05y, Ein05-30, Ein12i]. photochemischen [Ein12f, Ein12i, Ein12h, Ein05y, Ein05-30]. photochimique [Ein13a]. Photoeffekte [Wie06]. Photoelectric [Mil16, Wof08, Wie06]. Photoelectrically [Fri64]. photoemission [Gha14]. Photograph [Sol87, Kur06]. photographic [Bat03, Sug89]. Photographies [Bad11, Spe62]. Photographs [FS95, Man5 b, Bad12, FLN13, LML88, YJ95]. Photon [AP65, Wat97b, Kel14]. Photons [Dor71, Lew26, O’L64, CT98, Har13a, KSVW10, Wil11b, Zei10]. photos [Jos09]. Photovoltaik [MGG^{+}95]. Photovoltaik/Lindner [MGG^{+}95]. Phys [Ein86c, GHK^{+}07, Her03, Leg03, Nam10, ’t 02]. Physica [Ree16a, EI38a]. Physical [Ano12, Ano23a, Ano35, Ano84c, BHJ^{+}00, Bro05, Čap71a, Cea95, Cea97, Cea98, Cos87b, D’A01b, Edd28, EWJ^{+}25, Ein31h, Ein54e, Ein82d, Fey94, Foc53a, Fri39b, GL23, Got01, Got02, HCD^{+}78, Hel98, Hie82, HW04, HW12, Hoo91, How97b, Hun15, Jam30, JP89, Mas02, Mur25, Pee93, Roh87, Ros79a, Tha77, VW86, Wal16, Zah80, Zeh07, Ano81c, Bac05, Bor20, Bor22a, Cha49a, Com95, Ein30c, Ein31g, EPR35, Ein11b, Gin04, Gré19c, Hei50, HBD71, How76, Kie15, MP74, Mil90, Myr21c, Pai86, Pit16, Sch00c, Sch44, Sch45, Sch62a, Sch67, Sch92a, See25, Whi22, Whi04b, Woi06, ZDSF11, vG30, vH97, Anoxxc, Ein13e, Ein14c, Ein14l, Ein22j, Ein31a, Hof08b, Ken05b]. physician [HW55]. physicians [Ano85]. physicien [Geo53, L.58]. physiciens [CT12, Seg84]. Physicist [And81, Ano25a, Ano68a, Ano15i, Ber87b, Bey06, Bro15, Can16b, Cao06, Cas05b, CS70, Cot89, Egg02, Ein18g, Ein19h, Eng17b, Gla00, Hal18, Her17, JM17, Kov01, Mcc78, Nee79, Nye99, Pei85, Row05, Wil16a, dP94, And15, Bal95b, Cam09a, Can15b, Car08, Cat98, Clo15, Els78, Fan06, For63, Fra37, Geo53, Gre05b, Hag05, HH16, Hu05, Isa85, Kra13, KW87, Li07, LY95, Meh73b, Mof08, Ree16a, Ree16b, Ric93, Seg70, Sti13c, Yor87, Yor95, C.51b, Jam30]. Physicists [Ano33c, Ano39a, Ber88, Dür62, Dür80, For71b, For73, For80b, Hür22, Jof62, Kri08, McL99, Sal19, TEW02, vW02, vWR04, Ber00, Bes04, Bre97, Cho10b, Cli65, Cli87, CT12, Cro01, Dür06, Gam88, HH04a, Har06b, Har85, Hen07, Hoc83, HW12, Hol84, Jof67, Kan95a, Poo90, Bad12, Seg80, Seg84, Seg87, Seg07a, Sno81, Wig82, W^{+}87, WdC02, vS14, Ber00]. Physics [Acz01, AGMT12, Ano21d, Ano23a, Ano24b, Ano29d, Ano34a, Ano34b, Ano39b, Ano45c, Ano69b, Ano82b, Anoxxa, Anoxxb, Ano07h, Ano10e, Ano12d, Ano19, Asp69, Asp80, ARK^{+}99, Bar05f, Bet67, BS83b, Bet88, BGA^{+}13, Bič79c, Bir23b, Boh49, Boh58, Boh10, Bor66, BBSR69, Bor69b, Bri14, Bro17, Bro20a, Bro39, BF14, Cah04, Can00, Cao07, Čap71b, Cap05, CD23, Cho10a, Coh54, CW74, COS05, Cor04a, CCG93, CF82, Cra84a, CSW97b, Cun22b, D’A00, Dar04a, DiS06, Ein10c, Ein10d, Ein14a, Ein19f, Ein19g, Ein26a, Ein27e, Ein30h, Ein30b, Ein31a, Ein33f, Ein34k, Ein34l, Ein36a, EI38b, EI38c, EI39c, EI39a, Ein47d, Ein48j, EI49c, EI50b, EI50a, Ein50b, EI56, EI57, EI61b, EI62b, EI65a, EI66b, EI68, EI78a, EI84, EI87, Ein87a, EI88]. Physics [EI89a, EI92, EI95, EI98a, Ein03c, EI04, Ein05z, EI08a, Eks93, Eks97, Esh09, Fel11, Fer38, Fey64, For01, Fos13, Fow03, Fri81, Fri68, Gal82, Gam61a, Gam85, GSW95, Gil08, Gil09, Gin15, Gin01, Gor02, Got80, Ham11a, Ham11b, HMW24, Hei82, Hei32, HSD34a, Hei53b, Hei53c, Hei65b, Hei69b, Hei71b, Hei71a, Hei95, Hei49b, Hen11a, Hen11b, Hen82b, Hen86a, Hen96, Hes04, Hil07a, Hil07b, Hil07c, Hol81b, HB01, Ihd74, Jac87, JP89, Jof62, Joh11b, JM17, Kae38, Kai02, Kai05a, Kan96a, Kat05, Kil84b, Kle17b, Kra82a, Kri01, Lar19, Leg15, Lin60, Lor07b, Löw78, Lun92, Mag89, Mas01, MJ12, MC80, Meh73b, Mil11a, Mil11b, Mil24, Mul10, Mur90, Mus15a, Myr21b, Myr21a, Myr21c, Nau17, New09, Nob67, Nuv11, Nye88]. Physics [Nye19, Pau45, Pau94a, Pau94d, Peg39, Per21, Por81, Pow15, Pye80, RMI21b, RW64, Red84b, Ree09, RA13, Ren93, RDS03, Ren07a, RSS^{+}07a, RSS^{+}07b, RS07b, RS07c, RS07a, Rob49a, RS07e, RK11, RS49, Row01, Ryc05, Sac10, SS98c, SS98b, Sau07b, Sch17a, Sch21, ST98, Sch15b, Ser88, Set11, She83a, She83b, She15, Smo07b, Smo14, SM41, Stu18, TEW02, Tre88, Wal92c, Wal17, Wal02b, Wei12g, vW79, dWtH20, vCK17, Aas88, Aas90, AS79, AG13, Alf01, Ano65b, Ano65c, Ano81b, Ano81e, Ano93b, Ano97a, Ano05c, Sch12b, ACH^{+}03, Bad09, Bag92, Bag04, BBL08, BL11, Ber96a, BRVF84, Bey49, Bey77, Bjø20b, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, B^{+}61, Bor22b, BB05, Bra09, BSPKN03, BDH89, BPP95, Bro00, Büh98b, BS79, BW04, BE07, BRS03]. physics [Cal17, Cas09, Cat98, Cea97, Cea98, Che13, Cli87, Clo04, Con62, Coo99, CHU87, Cra02a, CM86, CM97, D’A01c, DB86, Dea13, Dir80b, Dir81a, Dor95, Dre93, EE05, EI38a, EI38d, EI38e, EI39b, Ein40a, Ein40g, EI42, EI47, EI48b, EI48a, EI59, EI61c, EI61d, EI61a, EI62a, EI62c, EI63, EI65b, EI66a, ES68, EI71, EI78b, EI89b, EBD^{+}89, EI93, EI94, EI98b, EI00, EI08b, EI11, EI13, EI15, Eng02, EM88, EvM94, ET07, ECEP11, FF91, FLS65, FLS89, Fey95, FLS06, FGL^{+}06, FW60, FJTS05, Fri83, FPLR59, Gam61b, Gam66a, Gam66b, Gam72, GdS15, Gie05, Gim12, GW10, GS05, GS06, GC20, Gol70b, Gol95b, Gol98a, GRZ99, GHW09, Gri84b, Gri84a, Gun69, Hag14, Hal15, Har85, Haw02c, Haw11, dVNHdVH53, dVNHdVH71, Hei05]. physics [HF22, Hei61a, H^{+}61, Hei69a, Hei72, HHJ^{+}06, Hen86c, Her78, Her79a, HBRD97, HD02, HKW08, Hom05, Hom06, Hun10, Jam82, Jam99, Jam02, Joh14, Jor69, JM86, Kae48, Kai05b, KT15, Käl22, Kel14, Kle82b, Kol40, Kox14, Kra99, Lan59, Lau05, LRS12, Lei08, Ler03, Liu92, Lon84, Lon03, Lor31, Löw98, MSYW12, MW99, Mar95a, MB66, Meh75c, Mel79a, Mel09, Mil82, Mil94, Mof10, Mou19, Nac79, NS88, Nacxx, Nam09, Nam10, Nye04, Pai91, Pal99, Pal83b, Pau58, Pau61a, Pla59, Pye75, Pye82b, RMI21a, RR82, RC13, Rec02, Ree11b, Ree14d, Ree15c, Rei05, Ric06b, Rio87, Sac88, SE51, SS98a, Sch12a, Sch17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, SS97, Sch13b, Sch04c, Seg76, Seg07b, Set03, Sha87, Sha98, Sho21, Sim94]. physics [Sim96, Skl85, SF33, Sla39, Sla55, Sla69a, Sla70, Sla13, Smo06, Sop88, Sta98a, Sta05b, Sta16, Sta98d, Sto96, Str11b, Stu75, TW66, TW92, Thi46, Top07a, Tur06b, TH79, Uff06, Uns00, Vel03, Ven94b, Wan95, War92, War93, War03b, WP85, WL80, WF98, Whi16, WH07, Wic95, Wie21, Wil81, Wil93b, Wol95, WB00, Yan80, Žáč11, ZZY19, ZXLX14, Zic00, d’E06, dB49c, vD21, vM98, Adl03b, Alt06a, Ano56, Ano64, Ano05d, Bag10, Bar93, Bar05f, Ber06b, Bru06, Cas75, Dai67, Del11b, DCL23, Dys65, Ein36b, EI38f, EI58, FvN78, GC20, HW49, Jen01, JDB^{+}97, KM04, Kle75b, LW34, MFHR06, OR21, Pau47b, Pin00a, Sau05, SR01, Sin07, UKKP04, Wad07, Wol16, Wig96m, Ano39d, Arm38, Hub13, Kri99, Len39, Pye10, Qui40, Ram82, Sau16]. physics [Vog21, Boh62, Est06c, Fes56, FGL^{+}06, For71a, Rom91]. physics-astrophysics [RR82]. Physics/Grosse [ARK^{+}99]. Physics/Peacock [ARK^{+}99]. Physik [Ano26a, Ano51a, Ano93b, Sch12b, B^{+}01, B^{+}61, Büh98b, Ein31a, EI61b, GRE^{+}01, HEB^{+}80, HW49, Hen86c, HH14, Kan96a, MGG^{+}95, Pau61a, Rec02, Sch12a, Sch17, Sch19, Sch20, Sch22a, SS97, SM41, Wie21, Ein19g, Ein27e, Ein30h, Ein30b, Ein36b, EI38f, EI49c, EI50a, EI56, ES68, EI87, EI88, EI89a, EI92, EI95, EI98a, EI04, EM88, Hei65b, Her78, JDB^{+}97, Jor69, Kle17b, Mel79a, Pau58, Sch21, SB06b, Uns00, Win69, For70b, Hen21, AP65, Cor04a, N.38]. Physik-Lehrbuch [HEB^{+}80]. Physik-Nobelpreis [JDB^{+}97]. Physik/Vilenkin/Shellard [MGG^{+}95]. physikē [EI78a]. Physikalische [Hof08b, HBD71, Hel98, vG30, Ein13e, Ein14l, Ein31a]. Physikalischen [WvLuAE18, Bac05, Bor20, Bor22a, Kie15, Kle75b]. Physiker [Bes04, Dür62, Dür80, Fra37, Hag05, Isa85, Sch81a, vS14, vW02, vWR04, Cot89, Egg02, Ein18g, Ein19h]. Physikern [Jof67]. Physiologie [HBFH18]. Physiology [Ano82b, Ano93c, Coh54, HBFH18, Kov01, Nob03, Uvn82, FMR90, HF22, Sou66, Ste53, Tho88a, Fie15]. physiology/medicine [HF22]. physique [Bro83, Com95, Ein10c, Ein10d, Ein30c, EI38e, EI48a, EI63, EI78b, EI89b, EBD^{+}89, EI93, EI15, Kum11, Pau47b, Pye75, Sto96, Tat64, dB49c]. physiques [Ano23a, Ein14c, MP74]. PI [Sym98]. Picasso [Goo02, Pep04c, Fol17, Gar93, Mar96b, Mil01, Mul02, Pye13, Sch00a]. Pickering [Hen86a]. pictorial [Ber69b, Cah55]. picture [Boz14, Lau15, Rei34, SMM^{+}60, SBSW60]. Pictures [Har79b, Hah79, Hah88, Her76b, Her84, Her01b, SKD^{+}87, vF79]. Picturing [Diz05, FS95]. Pieces [Mar11c, Fey95, Fey97, Fey11]. Pierre [Kle73, Mar73, Red10, Tat64]. Pieter [dP94]. Pillars [Cor04b]. pilot [UKKP04]. ping [HL10]. Pioneer [Fri68, Kra18, Ren07c, RCRC97, Wil74]. Pioneering [Sch13a, SS97]. Pioneers [Lag03, WL80, Nac79, NS88, Wil75, Coo04, AA75, Ram82]. Pioniere [BS92, NS88]. Pionierleistungen [SS97]. Pions [BDH89]. pisma [Pop98, Pop05]. Pitagora [Tug98]. PKS1830 [JRT^{+}91]. PKS1830-211 [JRT^{+}91]. Place [Gal00, Ano10d, PA06]. Places [OR21, Hei89]. plage [LRL15, LRL17]. plagiarist [Bje02]. plagiarizer [Mar17]. Plagiator [Mar17]. Plague [Bur45]. Planck [Adl03a, Ano61, Bre08, Ein06d, Ein07b, Ein13d, Ein05d, Fis10, Fis12, Gut15, Hen86a, Her77, Hie68, Hof08b, Hof08c, JDB^{+}97, Jüt21, Kil65, Klo63, MTVB88, Mic68, Prz63, Prz86, Ren05e, RS13, Som05a, Som05b, Yos95, vBL96a, vBL96b, Ano21d, Ano29a, Ano69c, Bad09, Bas92, Bla88, BEvL^{+}28, Bos24a, Bra14a, Bra14b, Bre08, BRS03, Cas75, Ano14b, CG14, Ein06d, Ein07e, Ein12a, Ein13d, Ein54d, Ein53g, Ein91f, Ein05p, EP06, Eva95, Far03b, Fis10, Fis12, FMM96, Gea02, Gol76, Gut15, Hei86, Hei00, Her77, Hof08b, Hof08c, Hol99, Jac08, JDB^{+}97, Kir80, Kra00, Kuh80, Kuh84, Lev04b, MR82d, MR82e, Mun36, Nag82, Prz67, Prz11, RS80b, RS13, Sch00b, Sch03a, Som05a, WvLuAE18, Yos95, Ein05l]. Planck-Medaille [Ano61, BEvL^{+}28]. Plancks [Ano69c, WvLuAE18, Bos24a]. Plancksche [Ein07b, Ein07e, Ein05l, Ein05p]. Planckschen [Ein12a, Ein05d]. Plane [Ros37]. Planet [Rus17, Lev15]. Planetary [Cle47, Poo21]. Planets [Gra94b, Lin15, Van98]. planning [Has09]. Plans [Ban11, Poc98]. Plant [Sch54, Hau54]. plaque [Pát79a]. Plaschko [MGG^{+}95]. Plasma [Raf17]. plastic [Pat01a]. plasticboks [Pat01a]. Plateau [SG11]. Plates [Cha23]. Plato [Joh04]. platonic [Cas03b, Her77]. Platonische [Her77]. Platonism [How98]. play [Ein19j, Gor17, MB11, Mos68, Pot11, Sta93c, Zee89, Zee90, Zee01]. played [How94a]. Playhouse [Tho17]. Playing [Gor12, Man80, Gor02]. plays [Ema85, Mar96b]. Plea [Ein48k, Whi55a]. Pleads [Ano46b]. Pleasant [Mal12, Lin17]. Please [CB06]. pleasure [Fey00]. Pleasures [Kör96]. Plight [Ein21d]. Plowshare [Kir05]. Plumber [Ein54c]. Plus [New65, Bie91, LW71]. Pluto [RA11]. Plutonium [Ber07c, Ber09, Wel99, Wel00, Ban12, SK94]. Pocta [ČK79]. pod [Smy46]. Podiumsdiskussion [GHK^{+}96]. Podolsky [AKBAE21, Bel64, Bha86, Com95, Cos87a, CPDH^{+}87, DR08, Erl72b, Erl72a, Har97, HR13, Hni98, HV88, Kau85, Kie15, Krü00, MPBL09, MS82, Nag71, RDB^{+}09, Rus85a, Ryf14, Sch71, SZ88, Sel88, SW87, Sta06, Tör81, Van74, Web93, YC18, ZDM08, ZXL^{+}19, Zuo96, dB85, dB87, dM86, Alb92, Hoo70, Khr08, Kie15, Mol74, Per05a, RMR96, Sta06]. Poems [Bus03, Ein03a, Coh79]. Poet [Ano33c]. Poetic [Cor15b, Sta05c, Yin15]. Poetry [dMdA05, CGG^{+}13, ET87, Gui95]. pogladów [EI59, EI62a, EI98b]. pohlednice [Brd79]. Poincaré [Adl03a, DDDR06, Gia70, Smo07a, Sta05a, Ste05d, Gal03, GB03, Gal04a, Gia71, Gin08, Gol67, Gol70b, GLW15, HZ01, Kat05, Lev04b, Mar04c, PB06, Sin00b, Wei12h, Wei12j, Wei12x, Yer03, Yer04, Bor02, Lam07]. Point [BRS^{+}12, Bri49, Con35, EF05, Eis22, How99a, Mil24, vN37, CG01, Ein19e, Ein22b, Ein05-37, Gio13a, Lum11, RR06, Stu75]. point-coincidence [Gio13a]. Pointer [How99a]. points [BM09]. Poison [Bur45, Ham08, Eic94]. poisoned [Pas10]. pojeć [EI59, EI62a, EI98b]. poke [Jak08]. Poland [Ano50c, Bič79c]. Polanyi [Ano87b]. polariton [DHY10]. polaritons [KRK^{+}06, PSM^{+}13]. Polarized [OSWR86, Ros37]. Pole [Ano68a]. polemiche [LS08, Per10]. police [Ano33b, Ano33g, Gru05d]. policies [CHW91, Nic87, Put09b]. Policy [Ano45a, Eme48, Eck90, EBU^{+}52, Mog02, Gil62, Her82b, Pan94, Sch65]. Polish [Ein35e, Ano48c, EI59, EI62a, EI98b, Gol98a, HC95, Inf33, Inf55d, Inf56a, Inf56d, RA00]. Political [Ano06g, Goe03, Gol82, Wil96, Ash12, Bla48, Bla49a, Bla49b, FF97, Gra82, Gru05d, Lan08, RS09, Sch05a, Sch05c, Sim05, SR04, EM^{+}91]. politicians [Hym12]. Politicized [Ree87]. Politics [Aro96, Bar05f, Bru03c, Cro16, Elz95, Fri01, Gru97, Hab69, Koj08, Kra02, Lan08, Sig78, Sim22, Sod95, TS01, Tur05, vW79, Wol62, AS79, Bey77, BB05, CGR^{+}97, Far13a, Fel79, Fel82a, Gim12, GMG^{+}77, Gro49, Gru05c, Hof08c, Inf78b, LSW12, LW71, Nye04, RS07f, Sch14b, Sch79d, SH01, Str68b, TB02, Tur03, Wir08, Pye80, Smo07a, Tsi08]. Politik [Hof08c, Gru97]. politikában [TB02]. politique [Dur12]. politiques [EM^{+}91]. politische [Sch05c]. politischer [FF97, Gru05d]. polity [Pai91]. Polizei [Gru05d]. Poll [Ano69c]. Poll/Three [Ano69c]. Polonium [Tho17]. poltergeist [Rei96b]. polyacetylene [Shi01]. polymath [Kil87, Kil89]. polymer [PSM^{+}13]. polymeric [Hee01]. polymers [Hee01, Mac01, Shi01]. Polytechnic [Wei12c]. polytechnics [JDB^{+}97]. polytechnics/Einstein [JDB^{+}97]. polytechnique [Pia31]. pomyłka [Gol98a]. ponderomotive [EL08c, EL05d, Ein10e]. ponderomotorischen [EL08c, EL05d]. pondéromotrices [Ein10e]. poniatii [EI66a]. Pontecorvo [Clo15, Tur03]. Pontifical [Ano80b, Dir80a]. Pontificia [Buc80]. Pontoppidan [Ano07r, BL07d]. Pool [Sch54]. Poor [Han79c]. Popovic [Tou05]. Popper [Gar72, vM98]. Popular [Ano22a, Cun22a, Ein17g, Ein20m, Ein20k, Ein21n, Ein21p, Ein31k, Elz95, GH01, Tay21, Wet80, Ein36f, Ein54g, Ein60d, Ein77b, Ein78, Van02]. Popularization [Cor15b]. Population [Bar05f, Hoc93, Nye88, Tre88, CHU87, Cra92, Cra01, Cra02a, Wei94, RK94, Wal94b]. pore [SEM03]. pork [Mer16]. portals [VT22]. Portfolio [Sin01b, Sin01a]. Portrait [Ber00, Ein59c, Inf47, Pai00, Rei30, Rei31, Rob48, SS97, Bad12, Sch78, Ano47a, Ber04, Ein59e, Föl91, KKJ99, Kay46, Kay68, McM14, Bad04]. Portraits [For80a, LMOS15, Bad04, Lew79, Ram79a]. portraitures [AA02]. Porträt [Föl91]. Porträts [LMOS15]. Portrayals [Tho88b]. portrayed [Ein93c]. Portugal [Ren21, MCS09]. Portuguese [EI39b, EI08a, Gar42, GdS15, Inf57a, Inf60, Inf61a, MCS09, See61, Seg87, de 03]. porzadek [Lan67b]. pose [TV00]. Posing [TV00]. Position [Ein40c, Ano79g, Ein21o, Fra49d, Pet21, Pet23, SE79, Szi45]. Positioning [Ash02, GR07, RG07]. positive [And33, CW32a, CW32b]. Positivism [Fra49a, Hol89b]. Positivism/Empiricism [Hol89b]. positrons [Dir65]. Possessing [Abr01a, Abr02, Abr03, Abr04a, Juc04]. Possibilità [Mas02]. Possibilities [Ein28c, Got01, Got02, Kae39, BC01, Ein23c, Mon04]. Possibility [Ber68a, Fri64, Mas02, Ein07i, Ein07c, Ein05-32, Ein05-33]. Possible [Ano35, Din59, McH07, Vig94, dH13, BBD^{+}11, HTS^{+}88, Per39, Poo90, San81b]. possibly [Jag01]. Post [Ano34b, CC67, Des79, Hec16, JDB^{+}97, Ree17c, BEM12, Dre07, Hol96a, PW14, Wel06]. post-atomic [Wel06]. Post-Cold [Ree17c, Hec16, Dre07, Hol96a]. post-conflict [BEM12]. Post-Einstein [Des79]. Post-Galilean [CC67]. Post-Newtonian [CC67, PW14]. postcard [Brd79]. postcolonial [Abr98]. postgraduate [Myr21b, Myr21c, She16b]. Posthumous [Gre19a, Gre19b]. Postmaster [Sta98d]. Postscript [RSS97]. Postulate [Din59, HG06a, Min07, Rob49b, Abi03, Dro15, Les74, Cor97, Wyk76]. postulates [Khr08]. Postulational [Bha49]. Postwar [CSW97b, Jun14, Kai05b, Szi45, Yam09]. posuv [Nac24]. Potential [Bar49, Ein05-39, Mie07b, BLI^{+}02, Dat38, Ein02b, Gás54, HH10, HV88, Ree17a]. Potentialdifferenz [Ein02b, Ein05-39]. Potentialfunktion [Gás54]. Potentiality [CHS97]. Potentialmultiplikator [HH10]. Potentials [Gás54, BRH^{+}09]. potentiel [Dat38]. Potere [RRR07, Val06]. pots [Jan05]. Potsdam [Egg95, BR00, CR07, Cli09a, Fre27a, Alp65, Alp66, Alp85, Alp94, BR00, Fre27a, Gra99]. Pour [Cun22b, Ein14j, Kle74, Sch81b]. pouvoir [ER91]. Power [Ano16c, Cra07, Fel11, Gin15, Gou47a, Ham11a, Ham11b, Hen11a, Hen11b, Joh11b, Kae39, Lau55b, Leg15, Mah09, Mil11a, Mil11b, Nuv11, Nye99, RRR07, Res50, RK39, Set11, Alp65, Alp85, Alp94, Bau95, Bau97, Bet91, ER91, GC02, Gow88, Gui95, Hun10, SSS^{+}98, Tau86, Tau88, Val06, Wal89, Wei05c, dP17]. powerful [Ano51b, Bla02]. Powerless [Lan14]. powstała [RA00]. pozapomenutá [Kot93]. poznani [EI58]. Pp [Bro72, Cah04, Coo04, Dix00, For05, Gil12, Goo85, Har79d, Hea64, Hen82b, Hen88b, Hen11c, Hes04, Hoc93, Hus15, Kil84b, Mil11a, Rec86, Red84b, Ste22, Tre88, Van22, WM08, War96b, War03a, Wig79, Ach13, Ano60, Bar05f, Ber07a, Boy88, Bro66, Cah90b, Duf46, Dys15, Ein14b, Fri73, Gly90, Gor02, Gor07a, Har02b, Hea13, Hei53a, Hen86a, Hen00b, Hie68, Hoh13, Jam70, Jan73, Kae38, Kar05, Kle70b, Kle70a, Kle14a, Koj08, Kra02, Kri08, Kuh54, Lin60, Mar10, Mas01, Mau09, Myr21b, Myr21a, Myr21c, Nah01, Nau17, Per10, Pes13, Plo06, PJMS11, Ree15d, Ree17d, Ree17c, Ren21, Rus17, Sch07c, Sci80, She15, Sho21, Smi02, Sta99b, Sta21, The12, Tur05, Tur06a, Vis94, Vis95, Vog20, Vog21, Wal02b, Weg08, Woo06, Wre66, Hou65, Mal65, See65b]. pp. [Ano84a, Lan08, Lan15, Tug98]. Practical [Fre79a, Maa13a, Maa13b, Sta07c, Ill12, Ric06b, dGK11]. Practice [CD07, Goe03, Sta09a, Käl22]. practices [BMS12, Ren08, Tur06b]. Prag [Kle75b, K.12]. Pragmatic [Mar88]. Prague [JP89, Bič89a, Bič89b, BL14, Brd79, ČK79, Hav77, Hav80, Ill79, Kle75b, Kol72, Kol74, Pát79a]. Praha [Bič79a, Brd79, ČK79, Fis79]. Praise [Cou79b]. praktijkboek [dGK11]. pravitelstva [Smy46]. Praze [Hav77, Kol72, Kol74]. Prazi [Pát79a]. prazské [Těš12]. Prazskemu [Hav80]. prazským [Těš12]. Pre [MC80, Rol10, She16b]. Pre-Einsteinian [MC80]. Pre-state [Rol10]. pre-university [She16b]. Preceded [Ken54]. preceding [Ein12b, Rob38]. precise [Poo90]. Precision [Jon17, Sei04, Gré19c, Hän06, Lin15]. preconditions [FTG08]. précurseurs [Rol71, Cun22a]. precursors [Rol71]. pred [Těš12]. Predecessors [EG80b]. predict [GW10, Lin15, Tom22]. Predicted [BR99, Opp79a, Ano10d, Spo16]. Predicting [AO78]. Prediction [DT23]. predictions [Häg46, Rus85a]. Predispositions [Gal82]. predmluvou [Ein23h]. Predvidenie [Whe70]. preestablished [Pye82b]. Preface [Ein53m, Ein55h, Ren07b, Her88, Kap61]. pregiudizio [CM02a]. Prehistory [Gol72, Hil39, How90b, Cle09]. Preis [Ano61, CSW97a, Jüt21, MTVB88, Wil36]. Preisgeld [MTVB88]. Prejudice [Nor07e]. prejudices [Bäu83, Mel79b]. préliminaires [Ein53c]. Preliminary [Ein53c, SPA^{+}68a, SPA^{+}68b, FLN13, Har13c]. Prelude [Bai75]. Premi [MFS^{+}6x]. premières [Ano15a]. premiers [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. premii [Dai67, GUB92, Was92]. premio [AMB^{+}77, Tio09, Paz10]. Preparation [Rus85a, Rus90]. Preparation-limited [Rus85a]. Prepare [Ano29b]. Preparing [WH99]. Presence [Ein25a, Bos24b]. Present [Acz10, Bič89b, Ein32a, Ein07c, Pla20c, CHW91, Ein13g, Ein13f, Ein14o, Ein21o, Paz10]. Presentation [Ano26b, Lam30, P.30, SS98c, Ano64, Ano65b, Ano72, Don70, FM92, Lun92, Nob72, Nob64b, Nob66a, Nob66b, Per97, SS98a, Woh76]. Presentations [Bri66]. presente [Paz10]. Presented [Lau50, Ano25c, ABdB^{+}53, New60, Thi46]. Presents [Ano31b, Pin00a, Ano13a]. Preservation [Van07c, Ein06a, Ein28e]. Preserver [Jun08]. preserving [Bus49, Bus68, Bus85]. Presidency [Nav82]. President [Lau55a, Ein39a, Ein97, Joh11a, Res50, Lucxx]. President-Elect [Joh11a]. Presidential [Her00, vB02]. Presidents [Kap20, Mul10, Ski04]. Press [Alf09, Ano84a, Bar05f, Bed01a, Ber07a, Ber00, Bla76a, Bla76b, Boy88, Bru06, Cah90b, Cli09a, Dix00, Dys15, Eng17b, For01, Gil12, Goo85, Gor02, GRC18, Hal18, Ham11a, Har02b, Har79d, Hen86a, Hen88b, Hen00b, Hes04, Hoc93, Hus15, Jam70, Jan73, Kar05, Kil84b, Kle14a, Koj08, Kri08, Kuh54, L.89, Lan08, Mar10, Mas01, Mil11a, Myr21b, Myr21a, Myr21c, Nav15, Noe11, Par06, Plo06, Pri99, PJMS11, Rec86, Red84b, Ree17c, RK94, Sci80, Smi02, Sta21, Ste22, Tur06a, Van22, Vog21, War96b, War03a, Wei94, vB02, Cho16c, Wal81, Zia94]. pressure [Cha11]. Prestige [Eng05b, Fel00]. Presuppositions [Hol78a, RS49]. Pretty [Sta07b]. Preußischen [Sim06, Ein21h]. Preussischer [B^{+}79]. prevalently [Hen13]. Prevent [FH10, SHW^{+}49]. prevention [JDB^{+}97]. previous [Ein05e]. Previously [RMI21a]. Price [Nau17, Pes13, Ree15d, Ree17d, Ree17c, Row12, Rus17, CSW97a, GRC18]. Prices [How21]. Pricey [Sei04]. Priest [Kal13, Kra92]. Prigogine [Hen82b, Duf81, Hen82a, Kle82a]. prijateljstvo [Pop98, Pop05]. Primacy [Eck90]. primary [CH80, Hen96]. Primas [MGG^{+}95]. prime [Gas05]. primele [EI57]. Primer [Her21, Ser43, Fer07b, Hea15, Pri82, SR92, Ree16a, Ree16b, Ree17b]. primera [Ano25b]. primeros [EI65a]. primitive [EI39b]. primitivos [EI39b]. Princeton [Ano23a, Ano84a, Ber07a, Boy88, Cah90b, Eng17b, Gil19, Gor02, GRC18, Hen88b, Hen00b, Kil84b, Kle14a, Koj08, Lan08, Mas01, Myr21b, Myr21a, Myr21c, Plo06, Sta21, Vog21, War96b, Wig79, Ano45b, Ano81a, Dav81, Ein21l, EA22, Ein22p, EA23, Ein55i, Ein64b, Ein76b, GR17, Myr21c, Ste05b, Wal81]. princip [Záv25]. Principal [Sch49e, EL08a, EL08b, Ein09a, EL09]. Principe [Cun22a, Cun22b, BTM84, Ein10c, Ein10d, Kle74, MCS09, NA92]. Principes [Pia31]. Principia [Dav09]. Principio [Gas05, Bla86]. Principle [Ang13, Ang20, Ano23a, Bar07, BTM84, BSW22, Dra14a, Ein10c, Ein10d, Ein13c, Ein14n, Ein16j, EWJ^{+}25, Hel15, vI10a, Lan14, LEM13, Lor14, LME23, LEM90, MSS17, MSS22, Mie07b, Mur25, Per00, SSM21, Wei12r, vB12, vB14, vB18, vB20, vL09, vL11, vL13, vL19, vL21, vL52, vL56a, vL65, BP95, BT86, BT88, Ein06a, Ein05n, Gio20, JP55, Mas05b, Nau16, Oha77, Žáč11, Záv25, Bla86, Car22c, Car13, CS60, Ein07c, Ein07h, Ein07g, Ein07d, EM20, Ein25c, Ein29b, Ein05-33, Ein05-34, Gas05, Goe70, Gre04c, Haw02d, Hof97b, Hum12, vI10b, vI11a, vI11b, Kle74, LEM^{+}23, LEMW52, Nor07c, Nor07b, Nor85, Pau47b, PCH77, Sch15, Tre71, Whi22, Whi04b, WTB09]. Principles [BHJ^{+}00, Ein14a, Ein96c, Pee93, Set03, Wig64b, Wig64c, Dir67, Dir81b, Dir03, Dir09, EG78b, Hei50, JN14, Wei72, Wig64a, Wig67a, Ein14m, Ein18a, Ein18m, Ein05-27, Meh87, Myr21c]. principu [Žáč11]. Pringle [Pye97, Pye06]. Print [Ebu16]. Prinzip [Ein06a, Ein16j, Ein05n, JP52, Ein25c, Ein29b]. Prinzipelles [Ein18m]. Prinzipielles [Ein14m, Ein05-27]. Prior [Nor04]. Priori [RR65]. Priorities [Van88]. Priority [CRS97]. prípadé [Žáč12]. Prirodie [Ein22j]. Prism [UM86]. Prisoner [Pou93, Van71]. prisoners [Wel06]. Privacy [BBD^{+}11]. privadas [HC96c]. privat [For80a, HW78, Ros05]. Private [Ald93, Bon94, Bur94, Cas94a, Cas94b, Din93, Gea94, Kan94a, Koj08, Lan08, Mil93, Smo07a, BW91, BW92, Buc93, CH98, HC93, HC94a, HC94d, HC94b, HC94c, HC95, HC96a, HC96b, HC96c, HC98, HC04, Ros05, RS07f, Ski04]. privately [HW78]. Prix [Hah73, Ihd74, P.50, Tat64, De 13, Jac12, Lar12, Pau47b, Dur12]. Prize [Abr01b, Adl03b, AA02, Ano45c, Ano56, Ano63b, Ano78, Ano07m, Ano07p, Ano07q, Ano07r, Ano07s, Ano12b, Ano12d, Bar05f, Bet67, BS83b, Bjö01, BOS14, Bjø20a, Bjø20c, BL07a, BL07d, BL07c, Bru03c, Cah05, Car15, CMT18, CGT11, Coh54, Dai67, Dan67, Dys65, Enk18, Enz05, Fer38, Fey64, For05, Gre19a, Gre19b, Hal12, Hei32, HSD34a, Hel14, Her66a, Jar08a, Jen01, JDB^{+}97, KM04, Kra02, Lau55b, LB02, Mil24, OR21, Pau45, Pau47b, Pin00a, Ram82, Rec10, RA13, SR01, Tom10, Tur05, WM08, Wat00, Wid07, Wol16, dP94, Ano65c, Ano01d, CÖT15, Joh14, Lar01, Lin99, LR14, MSYW12, Mag89, Mea21, MTVB88, Myr21b, Myr21a, Myr21c, Nor12, SBB16, Wad81b, Wor10, Abr88, AE06, Adl12, AKD23, All02, AE01, Ano11a, Ano12, Ano33c]. Prize [Ano34b, Ano36, Ano45b, Ano87b, Ano93a, Ano07o, Sch12b, Ano12c, Ano16b, Ano19, ACT23, BTR08, Bal75b, B^{+}01, Bjø19, Bjø20b, Boh85b, Bor04, BW^{+}87, BL07b, Bus13b, Cam00, Cao04, Cao14, Car12, CSC23, Coo04, CCD^{+}52, Cra90, CSW97a, CSW97b, CSW98, Das00, Das91, De 13, DS83a, Doh06, Don64, DFB22, Elz06, Ene12, Erc55, Esh09, FN38, Far53, Far63, FW17, Fei01, Fel82b, Fel00, FN13, FK68, Fri81, Fri01, GW10, GMOP12, GW83, Gun11, Gyl78, Hac98, HS14, Han18, Har13c, HMW24, dVNHdVH53, dVNHdVH71, Hef10, Hen03, Hof18, HL10, HR17, Hün18, Jar08b, Joh11a, Joh05b, Kar10, Kea18, KJY24, Kor94, Kra17, Kri01, Kro80, KUW82, KS81, Lag03, Lag12, Lev01a, LR01, LRH22, LHDH21, LLH22, LLSH23, Lin01, Low01, Lud57, MSYW12, MNT^{+}00]. Prize [MFS^{+}6x, Mat07, McC48, McG93, MB94, McG98, MB^{+}92, MTVB88, MS09, Mor95a, Nag82, NN01, Nor12, O’S01, Opf78, PA21, Paz10, Peg39, Per09, PS22, Pra12, Rei86, RF05, Rod22, SB82, San15, Sch14a, Sch12a, SS86, SS91, SS96, Sch87c, SK09, SO02, SR99, Sin07, Sou66, Sta93a, SL93, Ste53, Sti13a, Sty94, The95, Tho88a, Tho97, Tio09, TA17, TTB20, Van04, Wal62, WB^{+}87, Was92, WL80, Wil13b, Wir08, Wol67, Wor08, Wor10, Wyd15, Zew02, ZZ23, Eds82, Kau14, Str04, TF99]. Prize-producing [ZZ23]. Prize-winner [Wol67]. Prize-Winners [CCD^{+}52, Van04]. prize-winning [SBB16, Das00, LHDH21, LLH22]. Prizes [Ano97a, Ano04b, Bar05f, BNSY98, Cra84a, Cra02b, DWL98, Eng05b, Gil12, Gra49, Hea64, Hes04, LW34, Mat86b, Nye88, Rec86, Rei86, Tre88, Aas02, Ano09b, Ano17a, Cra84b, CHU87, Cra00b, Cra02a, Day12, Dor10, GW10, Goo03b, Käl22, Löw78, Mot79, Sch62b, Ode72a, Ode72b, S^{+}62, Ste07, Wid95, Aas88, Ano00, Ano03, Ano45b, Ano99f, Ano00d, Ano00f, Ano01d, Ano02b, Ano05a, Ano06a, Ano08d, Bar05f, CF82, Dor86, Fri89, Fri12, GNS99, H.89, Har02a, HD02, IP76, Lar12, Ler03, McC01c, McC01b, Mir94, Nor02a, Nor10, Nor13, Ren91, RWY97, RN16, Sta89b, Stê60, Sti98, Tho12, Vla79, Wes01, Zuc67, Kat52]. Prizewinner [Bri82, Opp79a]. Prizewinners [AO78, Duv64, Wil69]. pro [Rol89, Žáč12]. pro-Yankee [Rol89]. Probability [Ein05c, EH05c, MNT^{+}00, Sch36, CMT18, EH10b, Ein15b, Gor17, Sch35b]. probably [Bre97]. Probe [Ano11a, Gib91a, Law00, Law07, Sei04, Car07b, Pep04a, Rei11a]. Problem [Ano07t, Ber07b, Can16a, Dor78, Ein30b, Ein30e, Ein31o, Ein34l, ER35, ER36, Ein45e, Ein46a, Ein47d, Ein54f, Ein75, Ein79g, Ein07c, Ein07d, FS09, Hir76, Krü00, Leg05, LPPT75, Max93a, Sch20, Sta05c, Sto05, Van10, Viz79, AVvdV99, Dat38, Ein09d, Ein12c, Ein13g, Ein13f, Ein14o, Ein19e, EIH38a, EI40b, Ein51g, Ein17, FGL^{+}06, GG05a, Gut98, Leh19, Mal79, Rit09b, RE09, Rob38, Wei89, ZdSS91, dM86, Cre83, Cun22a, Ein14p, Ein30e, Ein79g]. problem-solving [FGL^{+}06]. Probleme [Beh05b, Ein23d, Dat38, Ein12c, Ein51g, Ein56d]. Problems [B^{+}89, Boh49, Ein13g, Ein14o, Ein19e, Ein30b, Gin01, Grü73c, Hel09b, KK89, Kor79, Min89, Whi27, Zuo96, Ano81e, AS91, Beh05b, Ein23c, Ein23d, EIH38b, Nor14, Set03, BNY81]. Procedures [Bjö01]. Proceedings [BC71, HS89b, Isr73, JP89, AS91, BS81, BPS87, BBM78, BB89, Cli09b, EK92, JP08, JRG02, Ber84a, SW94, Sto96, Ano81d, Ano93c, BCEP94, Mag95, Sch83b, Sim06, MK56]. Process [Ano39g, Bor04, Ein26d, MNT^{+}00, Ros82b, Bin06, CSW97b, Ein21s, LM94]. Processes [Ein14h, SP12, Bru68, Cha17]. processing [Hau08]. Prochitannykh [Lan67a]. prodigy [Bra12]. produce [Lui11]. Produced [Eöt91, Eöt08, Fer39, Fer70, Dow80, Jan77, Kle75a, Sch78, Wac06]. Produces [Ano39g]. producing [ZZ23]. Product [Lau50]. Production [Bet39, BHJ^{+}00, HEB^{+}80, Leg15, Tow65, Bet68b, Bet68a, Ein05d, Ein06f, MHY95]. Production/Olson [HEB^{+}80]. productivity [KKKM05]. Products [Mei39, MF39c, MF39b]. Proefondervindelijk [EdH15c]. Prof [Ano45c, Urb22, vC05b]. Prof. [Ano20c, Ano21j, Ano21i, Ano34b, Ano49b, Cam18, Ein45f, Ein96e, Kop64, Mcc55, Urb22]. profane [Wei15c]. profanes [Wei15c]. profesorem [Hav77]. professé [Pia31]. Professional [Cas79, Tor82, Cas07]. professionals [Rei41]. Professor [Ano26b, Ano29d, Ano54b, ARK^{+}99, Büc21, Cal02, Hei65b, Rad08, Don97, Hav77, vC05b, Ano20c, Ano21h, Ano27, Ano29a, Ano30a, Ano31e, Ano32, Ano33a, Ano33d, Ano33e, Ano33f, Ben99, Büc21, Bur68, Ein30g, ECS02, Fle33, Roo39, SK94, Swa30]. Professors [Ano39a, LW34]. Professzori [BST03, BSST04, CST01, CST02]. Profile [Ano97e, Mic62, Mic63, CC74, CC98, Sim05]. profiles [Ber91]. Profilováni [Těš12]. profit [Ano06g]. Profound [Bet91, Ano04d, Sti13c]. Profs. [Ano33c]. Program [Gon96d, Hol80, Ree14a, Sau14, Ama02, Bal86, Ber11a, JDB^{+}97]. Programm [JDB^{+}97]. Programme [Pro74, Zah73a, Zah73b, Zah76, BEM12]. Programmes [Sch74, Elb13, Gon96a]. programs [Elb13]. Progress [Ano20d, Dac20a, Ein30o, Ein48j, Ein50b, Ein05z, MNOW00]. Progressive [DMPS81]. Project [Bai89, MNT^{+}00, BBD^{+}11, MFO14, Ree05b, Ros12, Ano07k, Ano15b, Ben16b, Cas17, CB99, Dei95, Del09a, FS95, Gon00, Gon12, Gos94, Gos99, Gos01, Gou47a, Gro67a, Gro67b, Gro68, Gro00, Gro62, Gro83, Hac87, Hal97, HTS83, HHW99, Jos09, Kan02, Kel04, Kel06a, Kel07, Kel09, Kel12, Kir05, aM07, Nie89, Nor02b, Nor03a, Pal98, Ree05a, Ree09, Ree11b, Ree11a, Ree11c, Ree14a, Ree14c, Ree14b, Ree14d, Ree15c, Ree15b, Ree16c, Ree20, Rol47, Rus90, Sch15a, SFW91, Str13b, Sza92a, Sza92b, Tho17, Wal83, Wes04, Sho21]. Projection [Hoo91]. projective [SS21a]. proliferation [Hym12, RS09]. Prolog [Des72]. Prologue [EM32c, Ein77a, Des72]. Prometheus [Bar05f, Ree17d, BS05]. Promise [RSS^{+}07b, RMHD11, Rob58]. promote [Spo02]. Promoting [Bra14b]. Promotionen [HR02, RW02]. Prompt [Wal97, Wal04a]. prononcé [Pau47b]. pronouncements [Kru94]. Proof [EdH16a, EdH16b, Foe23, Büc21, Ein07c, Ein15c, Ein15f, EdH15a, EdH15b, EdH15c, Ein16e, FTZ^{+}15, Str17, EG23]. Proofs [Sau07b, FSE31, Sau05, Sch03a]. Propaganda [Wal92c]. Propagation [Ein22n, Ein05-31, Ken13, Sta98c, Ein11i, Ein20l]. proper [ZCPB11a, ZCPB11b]. Properties [ARK^{+}99, EG14, FP40b, FP40d, PE49, EK54, Ano26a]. Property [Wel08]. Prophet [Bac05, EE72, Pow02]. Proportion [Ano84c, Fer82b, Gol84b, Woo80b]. Proportionalität [vEPF09, vEPF22]. proportionality [vEPF09, vEPF22]. propos [Ber22, Car22a, Car23, Ein28d, ER91]. Proposal [AP65, Sul64]. Proposed [CHSH69, CHSH70, LC29]. proposes [Kon15]. Proposition [Ano31b]. Propositions [EL05b, Büc21]. proposito [Tug98]. propulsion [Ahe03]. Prose [Sta05c, ET87]. Prosperity [Pea58, Bra14b]. Prospettive [Ein22k]. Prostranstva [Ein22j]. Protactinium [BW39a]. Protest [Tis52]. Protests [Dic79b]. Protons [Wil46, vD04, CW32b]. prove [Oha09]. Proved [Acz01, Lau55a, Acz03]. Proves [Ano33h, Ano11a]. provide [Sti13c]. Provided [Ano35]. provides [Ano04d]. Proving [Ano91, Ano17c, GP19, Ken54, Kir05]. provocation [PB72]. provocazione [PB72]. proximity [Fie15]. prozesses [Ein26d]. Prüfung [Ein07c, Ein05-33, Ein19i]. Prüfungsprotokolle [HR02, RW02]. Prussian [B^{+}79, Sim06, Ano33e, Ein21h, Ein33d]. Prywatne [HC95]. Przibram [Hie68, Klo63, Kil65, Mic68]. PS [AGMT12]. PSA [BC71]. PSR [Cal86]. Psyche [JP52, JP55]. Psychedelics [Har07, Pic05]. Psyches [Asa95]. Psychoanalysis [Rol10]. Psychoanalytic [Eri82]. Psychological [DR16]. Psychologist [Mil75]. Psychology [SAY^{+}82, Ash95, Gie05]. Psychopathology [Hay91]. Ptolemaic [Kuh54, Sha54, Dra53, Dra01, Foc53b]. Ptolemäische [Foc53b]. Ptolemäus [Uns64]. Ptolemy [Hen91]. Pu [Ree17a]. Public [BEI^{+}55a, Duf46, Hen16, Koj08, Lan08, Nav15, Waz09, Waz10, Ber69b, Car10, Fre18, GHK^{+}96, MW46a, MW07, Ng11, Owe12, Rol47, RS07f, Waz14]. Publication [Ano79b, LLH^{+}80, Dic79a, KKKM05, MSYW12, Pra06, Wal81, Ano93a]. Publications [Ano21k, Bel26, Bri66, EWJ^{+}25, For05, Lam30, Mau09, Mur25, Pie24b, Rai46, Swa30, WM08, Woo28, For80a, HMW24, HR17, Ken05c, ZZY19]. Publications/USA [WM08]. Publicity [Cau91]. publics [Abr09]. publié [Ano80a]. Publiés [CE79a]. publish [Bjø20a, Bjø20b, Bjø20c, NHG89]. Published [Ano60, Fra81, Igl07, Ano80a, Ein15c, Hea13, Jak78, Kli62b, VM05a, WK12, CE79a]. Publisher [Nam10]. Publishers [Ano19b, Hou65, Lin60, Mal65, See65b]. Publishes [Ano15b, Lau50]. Publishing [Hea64, Hei53a, Ano06e, Bin06]. puertas [De 12]. puisobení [Hav80]. pull [von90]. Pulsar [Ano06h, Cho06, Kni13, Sta03a, WTF81]. Pulsars [Ano13b, Kni13]. pulses [Str19]. pump [Cha11]. pumping [DDD^{+}06]. punishment [Bor78a]. Punkt [Gio13a]. Punkt-Koinzidenz-Arguments [Gio13a]. pupils [RKW05]. purchasers [Hen88b]. Pure [Kat04, Van07c, Mil05c, Mil06]. Purely [Wey07]. purpose [Hai10]. purpose-guided [Hai10]. Purposes [Smy45a, Smy48, Gio13b, Smy45b, Smy45c, Smy45d, Smy45e, Smy46, Smy76a, Smy78, Smy89]. Pursglove [Dix00]. Pursuer [Sch95a, Sch96b]. Pursuing [Ham11b, Hen11b, Hun10, Mil11a, Mil11b, Ham11a, Hen11a, Leg15, Set11, Fel11, Joh11b, Nuv11]. Pursuit [Cao14, Mil10b]. Pursuits [Dys16, MJ12, Duc11]. pusobením [Těš12]. Put [Kae31]. Putting [Dor10, Gib91a, Wil86, Wil90b, Wil93c]. Puzzle [HS05, Kra92, RL16, Pou93]. Puzzling [Mos13a, Mos13b]. Pythagoras [Jam70, Cre10, Der15, Lan70, Ung98]. Pythagore [Der15]. Pythagorean [Mao07, Ung98].

Q [Hof08a, Rud17]. Q&A [Lui11]. QCD [Gro05]. QED [GHK^{+}07, BDR^{+}07, Fey06, GHK^{+}06, Sch94, Ven94b]. Quads [Mar00]. Quaker [Sta03b]. quale [Ber81]. qualified [Ano81a]. qualities [Per91]. Quality [IP76, YCL20]. Quand [Bra17]. Quandary [Kae31]. Quant [Mil24]. Quanta [Cun22b, Dra14a, EI38b, EI38c, EI66b, EBD^{+}89, Per21, Ram81, vCK17, Che13, Ein07i, EI38e, EI39b, EI42, EI47, EI48a, EI61d, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, Ein05-32, EI08b, EI11, EI13, EI15, Gla06, Inf33, Inf34, Kle63, LdB12a, LdB12b, Les74, Mil94, Nor06, Stu14, Wie06, vL55b, vL55c, EBD^{+}89, HBD71, Kae38, Bro39]. Quanten [LT15, Lin17, ST94b, Ein48a, Ein53d, HBD71]. Quanten- [ST94b]. Quanten-Dilettanten [LT15, Lin17]. Quanten-Mechanik [Ein48a, Ein53d]. Quantenmechanik [GRE^{+}01, Hel98, Bor55b, Bor63b, Hei34, Hei65a]. Quantenmechanik/Thomas/Weise [GRE^{+}01]. quantenmechanische [Kie15]. Quantenphysik [GRE^{+}01]. Quantenproblems [Ein23c]. Quantenrätsel [Bes04]. quantensatz [Ein17h]. Quantensprung [Fis10, Fis12]. Quantenstrahlung [Ein25b]. Quantentheoretische [EE22]. Quantentheorie [Ein16p, Ein17i, EE23, Hei93, Jüt21, Klo63, Pla20a, Pla20b, Ein14d, Ein16m, Ein24h, Ein25h, Ein25i, EI61b, HB63, Pau55, Pla22a]. Quantenwelt [Kon15, Arr06]. quanti [EI65b, EI13]. quantique [Ano15a, BD92, BB55, CTDL73, CT80, Com95, EBD^{+}89, Kum11, LB39, dB65]. quantiques [Gis12]. quantistica [BES86]. Quantitative [Möl84]. quantités [Ein10f]. Quantities [Ein14j, Ein87b, Eĭn85b, Eĭn86a, Ein10f]. Quantization [BB49, Ber81, Fer24, Lau99]. quantized [GB15, NBDD^{+}12]. Quantizing [Sto05, Tho83]. quantizzazione [Ber81, Fer24]. Quants [Col12]. Quantum [Aer99a, Aga93, AZ05, Ano35, Ano15d, Ano15j, Arr06, BZ10, Bel99, Bel23, Ber91, Bes04, Bey06, Boh62, Boh89a, Bok15, Bor53, vBT82, Bos24a, BES86, Bub79, BDH^{+}00, Cas05b, Cho03, CTDL77, Coh80, CHS97, CHR90, CF12, DR08, Dav14, DeW79, Del09b, Ein14d, Ein25b, ETP31, EI39a, Ein51a, EI62b, Ein67b, ET07, Eza79b, FH65, FHS10, Fin81, Fin93, FH10, For71b, For80b, Fre15b, GGK02, Gam85, Gar72, Gea02, Ger15, Ghi05, Gil08, Gil09, Gis14, GL23, Hal91, Hal17, HS10, HSD34b, HB63, Hei93, Hel98, Hel15, Hen86a, Her71, Hol05a, HW07a, Hoo70, How90b, Hug89, Hür22, Jam74d, Jam89, Jon08, Kar05, Ken13, Kle65b, Kle80, Kle14a, Kra99, Kra11, Kum10, Lan74c, Leg05, Lin09, Mer85b, Mil12b, Mus15a, Ove05, Par79, Pat93, Pau47b, Pau49]. Quantum [PV49, Pau94b, Pea08, Pla20b, Pla20c, Pla22b, Pla10, Pop99, PW17, Ros79a, RK11, Sau16, Sch52b, Sch52a, Sch53, Sch90, Sel88, Sla60, Sla63, Smo19b, Sop88, Sta06, Sus08, TWD03, Ven94a, Ven94b, Ven94c, VW86, Wal94a, Wat97b, Wei13b, WZ83, Whi06, Zee03, Zee10, ZCPB11a, ZCPB11b, vCK17, vdW67, AAA^{+}09, Abi00b, Adl82, Adl83, Ano15a, AS91, Bag92, Bag04, Bag11, BP95, BDH04, Bel75, Bel87, Bel04, BN97, BD92, Bit96, Bol04, Bor55a, BB55, Bor55b, Bor63b, vBT94, Bun79b, Cam09a, Cas76, Cas05a, Cas09, Cli65, Cli87, Com95, Con62, CG14, CPDH^{+}87, CM89, Dar14, DB86, Dir67, Dir81b, Dir01, Dir03, Dir09, Ein16m, Ein17h, Ein23c, EE23, EPR35, EI38a, EI38d, Ein53d, EI57, EI59, EI61b, EI61a, EI62a, EI62c, EBD^{+}89]. quantum [Ein90, EI98b, Ein10, EDK05, Far03b, Fay10, Fey66, FBD05, Fin86b, Fin97, Fis10, Fis12, FCKT11, Fri11, Gam66a, Gam66b, Gam72, GH93, Gie05, GC20, Gor17, Got00, GRZ99, GHW09, Gre05b, GZ97, GZ06, Gri84b, Gri84a, Gri95b, Gri95a, Gri12, Gri95c, Gri05, Hal15, Har13a, Haw11, Hei34, Hei50, Hei65a, Hen84, HV88, Jam66, Jam82, Jos80, KKC06, Kel58, Kie15, Kon15, Lan07a, Lau98, Leh14b, Lin15, LT15, Lin17, LB39, Lou83, Lum11, McH07, MR82a, MR82c, MR82d, MR82e, MR00, Mil90, MB11, Mor16, Nau16, Nav91, Oyi04, Pai79, Pai80, Pat95, Pau55, PS10, PS95, Pla20a, Pla22a, PCH77, Raz11, Rud17, ST94b, Sch13b, Sla69a, Sto13, Sto99, Str11b, Sty00, Tel95, Van07b, Vig94, Vig95, Wei00, Wei05b, Wei13g, WF98]. quantum [Whi96, Wic95, Wil79, Wol95, WB00, Zei10, dB65, dM86, vdW68, vdW07, Ano15a, BV09, Ein16p, Ein17i, EE22, Ein24h, Ein25h, Ein25i, Ein48a, Kum11, W^{+}87, WdC02, Bro00, Sud11]. Quantum-Induced [Ken13]. Quantum-Mechanical [Ros79a]. Quantum-Mechanics [Pat93]. Quantum-Theoretical [Gar72]. quantumtheorie [EI38a]. Quark [Raf17, Mer16, Rio87, San81b, Seg83]. Quarks [Hen86a, BDH89, Seg80, Seg84, Seg87, Seg07a]. quarrel [Gio22]. quartet [Man80]. Quasar [RB81]. Quasi [CLD^{+}03, DDD^{+}06]. quasi-equilibrium [DDD^{+}06]. Quasi-Stellar [CLD^{+}03]. Quasielectric [Kro01]. Quasimodo [Ano07s, BL07c]. quasiparticles [JDB^{+}97]. Quasiperiodic [Ein14h]. quasiperiodisch [Ein14h]. Quasiteilchen [JDB^{+}97]. Quatre [Ein55i, Ein64b, Ein76b]. quenched [SHL^{+}06]. Quest [Ano12a, Ber82a, Bok15, Del12, Fle18, Ger15, Gre99a, Gre99b, Gre00, Har07, Inf41a, Inf41b, Inf42, Inf80, Inf89, Inf06, Kem12, Kle14a, Pic05, vD02b, Ash95, Bar91, Com56b, Dar19, Ell05, GE97, Gri10, Gut97, Hag14, Jan14b, KT95, Ken07, Mil05a, Mil05b, PR11, Sto13, Tau12, Wal89]. Question [Ein40c, FTG08, Pal95, Pye13, Spe95, Ein10f, Ein14k, TV00]. questioners [Cli65]. Questioning [TV00]. Questions [DB09, Del13, Lor07b, Pal00, A^{+}09, And73, Ano69c, Mat91, Mea21, Wol00, Bed01a]. questions/Max [Ano69c]. qu’habitait [Kle16]. Qui [Dem15, Ein10e]. Quick [Fri50]. Quicksand [Fri89]. Quiet [Ano46a, JDB^{+}97]. quiet/Spring [JDB^{+}97]. Químiques [Gli84]. Quirky [Top07a]. quiz [Ano05-38]. Quo [EDK05]. quotable [Cal96, Cal00, Cal05a, Cal10, Cal13, EC00, Hor00]. Quotations [Cle84, EMH96, Ein21o, EM84]. quotes [Kra14c].

R [Ano54a, Bal72, Ein20d, Erm89, Fel82b, Har96b, Hea13, Jon89, Kan92a, Kat52, Kil95, LG97, Lin60, Nev92, Nor02b, Nor03a, Pye80, Pye06, Sch09a, Sch92b, Sie97, War95, War96a, Woo28, Mer85a, Sch17b]. R. [Ald93, Bur94, Cas94a, Din93, Ein35f, Hol99, Kan94a, Kil80, Mil93]. Rôle [Mer79]. Rabi [Rig00, Rabxx]. Rabindranath [Sin01a]. Race [Duf46, Eds82, Gon96e, Hol81a, Lau51, Smo07a, Yea08, Yor70, Cat04, Cre06, Far13b, Hen00a, HY15, Hof10, JT05, McM05, Rho07, She87, Wad81b]. Racing [Has05, Nor02b, Nor03a]. racism [JT05, Smo07a]. Radar [SAI^{+}71, Leo09]. Radial [Hub29, Hum36]. radiance [Nel14]. radiant [Mil05c, Mil06]. Radiation [Aga93, Cre15, Ein07b, Ein25a, Ein26d, Ein67b, Ein05c, EH05c, HEB^{+}80, Lau45, Löw78, Mal12, Sem08, Val49, Web79, Abi00b, BN88, Bos24b, Ein07e, Ein09b, Ein09c, Ein09d, EH10a, EH10b, Ein15b, Ein05l, Ein05p, EH05b, FMM96, FP16, Hac87, Hac94, Iro04, NKY15, Nat05, OP87, Rit09b, RE09, Sch03a, Smo07c, Tow65, Ein16p, Ein16m, Ein17i, Ein25b, LdB12b, Ein06d]. radiative [EE23]. Radical [Jay05, Gle10, Her01a, Sim05]. Radio [Ano16d, BIT03, HTS^{+}88, JRT^{+}91, B^{+}06]. Radioactivators [Ano36]. Radioactive [Fre79a, FP40b, FP40d, Joh07, Mil13b, Red10, SMKW46, De 92, Ham08]. Radioactivité [A^{+}84]. Radioactivity [Fer39, Fer70, Fri39a, RCRC97]. Radiochemistry [Adl12]. radioisotope [vH62]. Radiological [De 92, Wil46, Bro92, HY15]. Radiometer [BE69]. Radiometerkräfte [Ein24l]. Radiometric [Ein24l]. Radiothor [Hah62]. radiothorium [Hah62]. radiovola [Sla47]. Radium [Cla97, Kae48, Moo16, Moo17, Mul99, Pri12b, Ren08, Tod14, Swa02]. radius [See25, Sil30]. Raffaella [Per10]. Raffiniert [BHJ^{+}00]. rage [Has09]. raggi [Seg83]. raid [Wig86]. Railroad [Ano12]. Railways [Mar03]. rain [Ibu79, VWA95]. raios [Seg87]. raise [Mat91]. Raises [Con35]. raising [Win13a, Win13b]. raisons [Bla93]. Raleigh [BCEP94]. Raman [SR04, SR99, Sin02, Sin04a, Ven88]. Rame [PLM^{+}0 ]. Ranch [WA03, CC74, CC98]. Random [Gei11, Mus15a, Rus17, Web73, Bru68, Haw05]. randomly [Yag85]. Randomness [Sau16, Hal15, Hel82a]. Range [Edg14, Lau45, Edg14]. Ranging [WTB09]. ranked [BSPKN03]. Ranotgen [Seg86a]. Raphael [Bus12b]. Rapporto [RRR07]. rapports [LdB12a, LdB12b]. Rapprochement [Wal72, Kil70]. rare [Wei12c]. Rate [GT38, dHvdB49]. rates [Owe12]. Rath [GRE^{+}01]. Rathenau [Ein22i]. rather [Jan12, Ros05]. ratio [Ein06e, Lie78]. rational [Edd21, FTG08]. Rationale [Ein05-30]. rationalist [Ryc14]. Rationalität [ER13]. Rationality [Nug99, ER13]. rationnel [Edd21]. Raum [Ein18h, Ein30b, Ein30e, Foc54, Inf99, Rei24, Rei27, Rei28a, Rei28b, RK77, Ein30h, Sch17, Sch19, Sch20, Sch22a, SB06b, Wey18, WS^{+}01, Cun22a]. Raum- [Ein30b]. Raum-Zeit [Ein30e]. Räume [Ein30r, EM31b, Stu14, Stu23]. Raumes [Ein23f, Ein22c]. Raumstruktur [Tre71]. Raumzeit [Bar94b, Gre15a, Ren97b, HK94]. Raumzeit-Philosophie [Bar94b]. Raw [ER91]. Ray [Bri82, DR89, Gia80, Per00, Cul80, Gia03, Har79b, KS81, vS21a, vS21b, Van07a, Cur92]. Rayleigh [Eva95]. Raymond [Swi45]. rayonnement [LdB12a, LdB12b]. rayons [Seg84]. Rays [Kae39, Kae13b, Kox95, Ein18j, Ein26b, FSE31, Mah17, Seg80, Seg84, Seg87, Seg07a]. razón [Yun15]. razrabotke [Smy46]. razvitie [EI66a]. Razvoj [EI62c]. RBRVS [Ein95b]. Re [Čap71b, LLH^{+}80, Whi04a, GTP22]. Re-Assessment [Whi04a]. Re-evaluation [Čap71b]. Re-Publication [LLH^{+}80]. re-told [GTP22]. Reach [Ano15h, SS65]. Reaction [BHT86, FP40c, Goe93, MF39a, Wil96, And73, Hah48]. Reactionaries [Van07c]. Reactions [GT38, Hoy54, FN13, MCS09, Tör81]. Reactor [Cal75, Sch54, Ban03]. Read [Ano05-41, Ein52e, Lip00]. Readable [Dur03, Ano26a]. Reader [Pia29a, Pia29b, Ein06a, ET87, GS59, Haw92, HT10, Rin03, Woh76]. readers [Mil12b, Pri12a, TF99]. readership [Bus12a, She16b, Sho21, Sud11]. Reading [Ein48b, Kai81]. readings [Hug99, Tay49, WH07]. Ready [Fra05]. Real [Ano45a, Ste22, vCK17, Arg09, Bri05c, Ein17, ECS19, Goo04a, Goo04b, Jun05a, Smi07b, Wue05]. Realigns [Sul64]. Realisation [Nor00]. realisiert [Kon15]. Realism [AP00, Bes93, CHR96, Dan89, Fin86a, Fol96, Gar72, How84, Lau95, Max93a, Max93c, Max93b, Pat79b, Sch00b, Wen49, Fin86b, Fin97, Leh14b, Mah01, MS82, Nol88, Pol89, Sel88, Smo19a, Luk02]. Realismus [Luk02, Pol89]. Realist [How93b]. realistic [Stu14, Stu23]. realistische [Stu14, Stu23]. Realität [Kie15, RS13, ST94b, Ein36b, FF97]. réalité [Com95, Kum11]. realities [WD06]. Reality [AP00, Ano35, Ein31h, Ein36a, Ein36b, Ein48a, Ein54e, Ein82d, Ein03c, Erl72a, GHS21, Gre04b, Gre05a, Gre11, Hal17, Hel98, Hol68b, Hol70, HBN00, Isa15, Mar49, Mit02, Mor96, New09, Ray95, Roh87, Ros79a, RS80b, RS13, TE31a, VW86, WW57, Aer99a, BDH04, Bar96, Bay94, Com95, Dor95, Ein31g, EPR35, EI39c, Ein11b, FF97, Gho17, Göt92, Gri84a, Gri95b, Gri95a, Gri10, Haw02a, Hol88b, Kie15, Kum10, Kum11, MH71, Moo10, Nye72, Pen05, Per10, Per05b, ST94b, Ven94c, Vig94, Zuc82, Zuc83, vG30, Lee06, Mer85b, Sud11]. Realization [Ber11a, Rei20, EI38f]. realized [Kon15]. realizing [Amo10]. Really [Ano15l, Ein21g, HM90, How93b, Lan14, Wis79, Baw06, Bry09, Hef10, Lan67c, Sch10b]. realm [Kae48, Lin15]. realtà [Arg09]. Reanalysis [FSS^{+}86b, FSS^{+}86a]. Reappraisal [Dys76, HW07a]. reappraisals [Has07a]. Reappraising [CH05, Iro04]. Reason [FC87, Gha17, MA31, Mor96, MT98, Ezr82]. Reasonable [Fey05, Hol03a]. Reasoned [Ach13, Tib13]. Reasoning [Har97, Mon19, Lon84, Lon03]. reasons [Bla93, Ein12i, Ein12h]. Reassessing [FP16]. Rebecca [Nye88, Tre88]. rebel [HD75, HD72, HD73, Fri73, Hei74a]. rebelle [HD75]. rebellion [Hol96b, Hol00a]. Rebellious [Wei12d]. rebooted [Fer14b]. Reborn [Gil08, Gil09, Smo14]. Rebuilding [Fre15b]. Rebuttal [Inf53b]. recalled [Ber00]. Recalling [Arn00]. Recalls [Lau51]. Receive [Ano54b, CSW97b]. received [CSW97a]. Receives [Ano29a, Pin00a, KS12]. Receiving [Ein48i, Pau47b]. recency [LHDH21]. Recensioni [Del13]. Recently [CSW97b, RMI21b, RMI21a]. Reception [Bie87, Cao06, Gli87, GPSR01, Gra82, Hen86a, Man86, Ree87, ACT23, CÖT15, Cre81, Fan06, Hoc83, Hu05, Li07, MR82c, Pat99, Pye73, Pye75, Spo02, Zia94, Pat99, Pau47b, Pye75]. Receptions [GH01]. Rechenberg [Hen86a, Hen11c]. recipients [Ler03]. Reciprocal [Kie82]. Reciprocity [Bor49b]. récit [Cun22a]. reclaiming [BEM12]. Recognition [CJT22, Nye19, VW86, CMT18]. Recognized [Pin00a]. recognizing [Ano19a]. Recoil [Tre75b]. Recollections [And81, Far79, Grü90, Han80a, Hei79a, Hei79b, Mor95c, Pei85, Sch79c, SW95, Tal32a, Bor78b, Oli72, PGH84, PGH96, WS78, Bor75, Mar95b]. Recommendations [Ano15k, KJY24]. Reconciliation [Ein20j, Nug00]. reconfirmed [Kon15]. Reconsidered [Ber95, Cho05b]. Reconsidering [Jan02, BV09]. Reconstructing [RS03b]. Reconstruction [NM12, Joh05b]. Record [Ish23, Jea08, Ber69b, JDB^{+}97]. Recordings [Bad96, Ber96c]. records [Anoxxc]. Recovery [Boy88, Hol91a, Gli88]. Recurrences [BN88]. Red [Dic60, Gor09, Jun14, Sch54, AMM99, Arp98, EG80a]. Rede [Ein20b, Ein30i]. Reden [Jos95b]. redheaded [vC05b]. Redirecting [Aas90]. Rediscovering [Obe16]. rediscovery [Pen03]. Redshift [O’L64, AAA^{+}18]. redshifts [Arp98]. Reduces [Ano29d]. Reducing [BCD93]. Reduction [GS22, Hea11, Fle19]. Reed [Sho21]. Reexamination [Mol74]. Reexamined [Nag71]. reference [Ber65, Myr21c, Van02]. references [LRH22, TTB20, UKKP04]. refined [Bon81]. Reflected [Eks99]. Reflection [Sea79, BAM99]. Reflections [Ber18, Bub79, C.51b, Gin01, Gor12, Har79c, Mar79, Nic17, Nic19, Rya06, AE06, CH05, CM89, Dav09, Eri82, GR20b, Hof18, Inf78b, OOM21, Rab63, Ree15b, Wig67b]. Reform [Cla97]. refraction [Ein18j]. refractive [BCO^{+}11]. Reframing [Ash11]. Refrigerating [SE28]. Refrigeration [ES30, Cha11]. Refrigerator [Cho10a, And17, Dan97b]. Refrigerators [Dan97a]. refugee [Hoc83]. Refugees [Rei07]. Refuses [Ear01, Kif95]. refutation [Ler91]. Regal [Erc55]. Regard [Swa30, LFKW06]. regardant [Kle74]. Regarded [Ano45a, Del11a]. Regarding [Ein14j, Les74]. regia [Erc55]. Regime [MP01, MP12]. regress [Ken14]. regular [EG23, EP43, Lom05]. regulären [EG23]. regularities [KM04]. Regularization [PV49]. Reich [Bey77, HW12]. Reichenbach [SW94, Gam42b, GW06, Gio16, Gio22, HRP90, Rei06, Rei11b, SSM21]. Reichs [Ein20b]. Reichs-Universität [Ein20b]. Reichspostministers [Sta98d]. Reidel [Hen82b]. Reign [Hal21, Ryc05]. Reihe [Bäu83]. reimagines [Mus15b]. Reimagining [BEM12]. Reinhold [Hen91]. Reinterpretation [Čap71b, CL19]. Reinvented [Isa15, Sta93b]. Reinventing [Mof08, Lau05]. Reinvention [BLR15]. Reise [BHJ^{+}00]. Reissner [Ein14k, Ein14k]. Reject [Ita82, Wei18a]. Rejecting [Cam09b]. Rejection [Ser77, Gor17, Mal91]. Related [Ree16c, AA02, Ano19a, Eĭn86a, Ree05a, Ree11a]. relatif [Per39]. Relating [Ein87b, Ano45d]. Relation [Bro23, CMT18, Ein27f, Gra07, Her66a, Hub29, Hug54, MBD39, OW54, Sil23, Sul64, WJ21, EdS32, Fey84, GB15, Hen85, Hom05, Pau25, Sch13b, XMY^{+}09, dM86]. Relations [Mer79, Gla07, Sch88b, Sch35b, Sch36]. Relationship [Ein05q, Ein05-36, Fau79, Göd49b, Mel63, Pik63, RRR07, Sch87b, Sch88a, Sta94, Bra37, Ein11a, Ein11g, Ein27a, Ein05i, Els24, Hei75b, Hen86c, Kol72, LHZ^{+}23, Müh11, Sid76]. Relationships [The46, CZPR17, Her66b, Lan21]. Relatitivätstheorie [Cun22a]. relativ [Fra52]. Relative [Föp07, Fri07, Get06, Hug54, OW54, Abb02, Apf81, Ein27c, EI38d, Fra52, Jan12]. Relatively [Cha88, Bel96a, Egd14, Par00]. Relatividad [Ein21q, Bor22b, EI65a, Ein85f, Rus43]. relatividade [EI39b, NA92]. relativism [Rot82, Nol88]. Relativist [Fri79a]. relativiste [Kle74, Mey25, Mey92b, Ein28d]. Relativistic [Ben11, BBCPL09, Coc66a, Coc66b, Coc66c, Dir49, Ein22k, Ein28d, Gam49, Hak99, Kat05, Kov20, Kra05, Kva79, Lat88, MW11, Nor07a, O’L64, PE49, PV49, PS03, RDS03, Ren07c, Sch05b, Sch07b, Sch53, Zah87, ACH^{+}03, BB89, EP43, Ein45b, ES46b, EK54, EK55, Eis07, Fri83, FKZ91, Hea15, JRG02, Ker89c, Kle74, KRZ^{+}05, Liu92, Mey85, PR11, PW14, RP09, Rei24, She16a, SW87, Tam14, Voi86, Wal98, Wal91b, Wes80, ZCPB11a, ZCPB11b, Rus37, Sme14]. relativistica [Tam14]. Relativistiche [Ein22k]. relativistischen [Rei24, Tre66, Voi86]. relativistischer [KRZ^{+}05]. Relatività [Ein21p, Ein60c, Bri05d, EI65b, EI13, Inf63, Inf68, Inf74, Inf92, Kop23b]. Relativitaätstheorie [Bor64, Bor69a]. Relativitás [Ein22a]. Relativität [Abr12a, Ein12b, Ein20g, Ein05e, Abr12b, Ein12g, Ein05-28]. relativitatii [EI57]. Relativitäts [Ein20b, Fra20, Kra20a, Lan21, Sch19, Sch20, Sch22a, Wei15c, Ein79g]. Relativitäts- [Sch19, Sch20, Sch22a]. Relativitäts-Problem [Ein79g]. Relativitäts-Theorie [Ein20b, Fra20, Kra20a]. Relativitätsprinzip [Ang13, Ang20, BSW22, Ein07h, Ein07g, Ein07d, Ein05-34, Hof97b, vI11a, LEM13, LME23, vB12, vB14, vB18, vB20, vL19, vL21, vL52, vL56a, vL65, Lor14, Ein13c, Ein14n, vI10a, vI10b, vI11b, LEM90, vL09, vL11, vL13]. Relativitätsprinzips [Ein05-33, Ein07c, Sch15]. Relativitätsproblem [Ein14p]. Relativitätstheorie [Bes04, Bor20, Bor22a, Cas21, Cun22a, Cun22b, EG13a, EG13b, Ein14g, EG14, Ein14m, Ein15g, Ein17g, Ein18n, Ein19e, Ein19k, Ein20o, Ein22r, Ein23b, EG27a, EG27b, Ein54i, vM90, Ein05o, Ein05-27, Föl90, FF97, Häg46, Hen90, Hen92a, Isa85, Jun05a, Kra23a, Kra23b, Mau96, MK56, MD15, Mos21, Pet21, Pet23, Pye93, Ren05g, Row06a, Row06b, Sch09a, Urb22, Waz09, Wey20, Wie21, Wol87, Wue05, vG30, vL19, vL21, vL23, vL52, vL55b, vL55c, vL56a, vL65, Ein11b, Ein15j, Ein15e, Ein15i, Ein16d, Ein16b, Ein16h, Ein16j, Ein17e, Ein18f, Ein18e, Ein18m, Ein19i, Ein20c, Ein21t, Ein22q, Ein23k, Ein23d, Ein24c, Ein25e, Ein30n, Ein31o, Ein32a, Ein56b, Ein99, Els24, Fer14a, Foc53b, Fra37, Fre19, Inf55c, Kop21, Kra20b, Kra20c]. Relativitätstheorie [Kra21, Kra25, Pau21, Rei20, uJR15, Sch17, ST94b, Voi86, Wey18, Wey23, vL56b, Ein88b, Ano21k, Ano20a, Spe68]. Relativitätstheorien [Ehl05b]. Relativité [Cun22a, Ein21k, Ein21j, Ein33c, Gou10, Ano23a, Ano16a, Bor02, Cun22b, Der15, Edd21, Ein10c, Ein10d, Ein23g, EI38e, EI48a, Ein55i, Ein56d, EI63, Ein64b, Ein76b, Ein76a, EI78b, EI89b, EI93, EI15, Eis86, Eis02, LRL15, LRL17, Lev04b, MK56, Pat99, Pye75, Spe67, W.25, PB06, Mat86a]. relativiteit [vC05a]. relativiteitstheorie [EI38a]. relativiteten [Str97a]. relativitets [EI38d, Sch27, Gru07]. relativitets- [EI38d]. relativitets-teorye [Sch27, Gru07]. Relativitetsteori [EI39a, EI61a]. Relativity [Acz99, Acz00, Ano17, Ano22a, Ano23e, Ano23a, Ano29b, Ano31b, Ano33f, Ano62, Ano70, Ano07e, Ano15d, Ano15k, Ano16a, ABIM15, Bac49, Bai07b, Bar80a, Ber42, Ber56, Bes04, Bey12, Bia16a, Bič00, BL14, Bir23b, Bir25, BRS^{+}12, BLR15, BBCPL09, Boy88, Bri05d, Bro20a, Bro39, C.50, CR07, Can15a, Car04b, Cas21, Cer00, Che10, Cle47, Col13, Cor15b, Cun19a, Cun19b, Cun19c, Cun20a, Cun22b, Cus82, D’A01a, D’A01b, Dat15, Del11b, Del11a, Dic64, Din49, DE04, Edd20b, Edd23, Edd29b, Edd59, Edd66, Edd87, Ehl09, Ein11b, EG13a, EG13b, Ein14e, Ein14m, Ein16j, Ein19k, Ein21k, Ein23b, Ein23d, Ein24c, EWJ^{+}25, ER36, EI38b, EI38c, EI39a, EI49a, EI49b, Ein60c, EI66b, Ein82f, Ein96f, Ein05o, Ein07b, Ein07d, Ein09c, Ein15b]. Relativity [EK92, Fer14c, Foc55, For14, Fra20, Fra37, Fre68, GGK13, Gam42b, Gei15, Gha17, GS22, Gia71, Gin82, GRRS99a, GRRS99b, Gol69, Goo75, Gra18, Häg46, Hal21, Has07b, Hea13, Hen09, Hen16, Hol91a, HS89b, How92, Hub13, Inf56e, Ita82, Jan73, Jan14a, Jun05a, Kae38, Kai81, Ken19, Kra20a, Kra23a, Kra23b, LRL15, Lat88, Lau50, Lau55a, LPPT75, LEM13, Lor10a, MSS17, MSS22, Mar09, Mar10, McG19, Mel82, Men49, Mie07b, Mik65, Mil79, Mil82, Mil49b, Min07, Mos21, Moy79, Mur25, Nag50, Nar10, Nav15, Oha79, Pal83b, Per00, Pie24b, PK06, PFM17, Poo90, Pul11, Pye10, PJMS11, Raf17, Ram83, RBB03, Ree87, RS03b, Rob75, Rob15b, Row06d, Roy19, Rus17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, Sch83b, Sch85]. Relativity [Sci80, Set10b, Set10a, Sha68, Sha14, Sha64c, SPA^{+}68b, Smi02, Spe95, Sta19, Str97a, Str13a, Sul13, Tal32b, Tay21, Tre21, Van11, Van79, Wal84, Waz09, Web61, Web04, Wei12k, Wei12w, Wei12x, Wei13b, Wei14a, Wei15c, Wen49, Wey23, Wey09, Wie92, Wil00, Wil14, WL12, Wol87, dH13, Acu14, Alc08, Ano06d, Ano07u, Ano09c, Ano18c, BS10, Ben14, BIT03, Bha92, Bič79b, BB89, BPL05, Bor65a, Bro05, CL19, Cas23, Cas03a, Cea95, Cea97, Cea98, Cha76, Che13, Cho08, Cre05, Cre81, Cre06, Cur92, Dar19, Dar21, Der15, Dro15, Ein23h, EI38a, EI38e, EI39b, EI48a, EI57, EI59, EI61a, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, Ein82b, Ein85d, EI89b, EI93, EI94, Ein95b, EI98b, Ein07f, Ein07g, EI08b, EI11, EI13]. relativity [EI15, Ell89b, EFH89, Fer14b, Fey11, Fle19, Foc53b, FGPP06, Ger78, Gib91b, Gil19, GTP22, Gin62, Goe17, Gol84c, Gol15, Gor22, Gou13, GG15, GR15, GR17, HH16, Her01a, Hot15, Inf63, Inf68, Inf74, Inf92, Jag01, Jan05, Jan12, JRG02, Ker87, Kop21, Kop23b, Kop23a, Kra20b, Kra21, Kra25, LRL17, LD08, Leh14a, Leh19, Les74, Lev15, Lie08, Lin15, Lor20, Lor09, Lor10b, Lor10c, Mar03, Mat91, Mau96, Meh98, MK56, MD15, Nor93, Oku05, Orz12, Par91, Pen03, Per91, Pet21, Pet23, Pri82, Pye85, Rai75, Rin60, Rin66, Rin69, Rin77, Rin82, Rin91, Sar96, Sch88c, Sch91, Sch93c, Sch75, Sch77a, Sch77b, Sch77c, Sch77d, See25, Sha88, She16a, Sta95a, Sta95b, Sta89c, Swe79, SO72, TW92, TW00, Unr99, Urb22, Voi86]. relativity [War89, Wei12v, Wei13c, Whi01, Wol95, WB00, You11, Žáč11, Záv25, dB36, tH20, vC05a, vBT96, vL23, vL55b, vL55c, Abi91, Abi00a, Abr12a, Abr12b, Ang13, Ang20, AMSmS96, Ano20a, Ano21d, Ano21f, Ano21l, Ano22d, Ano22a, Ano23a, Ano24b, Ano25a, Ano28, Ano30b, Ano05-42, Ano07c, Ano07l, Ano15h, Ano16d, Apf81, App11, Ash02, Bai92, Bai01, Ban06, Bas65, BE02, Ber76a, Ber79a, Ber82a, Ber07b, Bir23a, BSW22, Boh65a, Boh79, Boh89b, Boh96, Boh06, Bon64, Bor20, Bor22a, Bor22b, BB24, Bor50, Bor62, Bor64, Bor69a, vBT82, Boy88, Bro19, Bro20b, Bru99, Car02, Car22c, Car24, Car13, Cas86, Cha88, Cha75, Cha78, Cha01, CSM06, Cho05b, CB09, CB15, Cob07, CHR90, Cor97, Cre83, Cul81, Cun15, Cun22a, Cun22b, Cwi87]. Relativity [D’A97, D’I92, Dar96, Dar06b, Del11a, Den24, Dir75, Dir82, Dir96, Dor78, Dra12, Dra14b, Dur03, EG78a, EG80a, EG80b, EJN93, Edd20a, Edd21, Edd30b, Edd06, Ehl05b, Eĭn85a, Eĭn86b, Ein07c, Ein07h, Ein07g, Ein07d, Ein10c, Ein10d, Ein12g, Ein14g, EG14, Ein14n, Ein14p, Ein15j, Ein15e, Ein15g, Ein15i, Ein16d, Ein16b, Ein16h, Ein17e, Ein17g, Ein18f, Ein18e, Ein18m, Ein18n, Ein19e, Ein19i, Ein20c, Ein20b, Ein20g, EM20, Ein20k, Ein20o, Ein21b, Ein21l, Ein21t, Ein21o, Ein21n, Ein21p, Ein21r, Ein21q, Ein21j, Ein22a, EA22, Ein22l, Ein22p, Ein22r, Ein23k, EA23, Ein23g, Ein25e, EWJ^{+}25, EG27a, EG27b, Ein30n, Ein31o, Ein31j, Ein31k, Ein32a, Ein33c, Ein33e, Ein34j, Ein35c, ER35, Ein36f, EI42, Ein45c, Ein46d, EI47]. Relativity [Ein50j, Ein50h, Ein50i, Ein53h, Ein53i, Ein53n, Ein54f, Ein54g, Ein54i, Ein55g, Ein55i, Ein56b, Ein56e, Ein56f, Ein56g, Ein60d, EI61d, Ein61a, Ein61b, Ein64b, Ein65, Ein67a, EA67, Ein69b, Ein74, Ein76b, Ein76a, Ein77b, Ein78, Ein79g, Ein81, Ein83, Ein84c, Ein86e, Ein88a, Ein88b, vM90, Ein94d, Ein95d, Ein96a, Ein96b, Ein96d, Ein99, Ein00c, Ein01b, EK03, Ein03b, Ein05-33, Ein05-48, EG05, Ein05-27, Ein05-28, Ein05-34, ELGC05, Ein06b, ELGC06, Ein08b, Ein13, Eis86, Eis02, EK05, Eis06, Els24, FK90, Fer14a, Fer81, Feu71a, Feu71b, Fey97, For80a, Fra50, Fra51, FF97, Fre19, FB24a, Gam42a, GO08, Gar62, Gar76, Gia70, Gib91a, Gin59, Giu05, Gli87, Gli88, Göd49b, Goe93, GRRS99b, Gol67, Gol68, Gol70a, Gol70b, Gol76]. Relativity [Gol83a, Gou10, Gra71, Gre04c, GG05b, GH07, Gru57, Grü62, Gru07, Hag08, Har03b, HI79, Haw02d, Hel80, Hel84, Hen90, Hen92b, Hen92a, Hen94, Hen97, Hir74, Hir76, Hof97b, Hof99, Hol68a, Hol88c, HG05b, How96a, Hsu00, HZ01, Hu04, Hum12, vI10a, vI10b, vI11a, vI11b, Ill89b, IS49, Inf55c, Isa85, Isr73, Jac87, Jan08, Jan95, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, Jan14b, KKC06, Kat92, Ken25, Ker87, Kes65a, Kes65b, KK83, Kox88, KE05, Lan21, Lan37, Lan55, Lan72, LA88, Lau55a, Laz15, Lea07, Lem49a, LLH^{+}80, Lev04b, Lie45, Lor14, LEM^{+}23, LME23, LEMW52, LEM90, Lud13, Mac86, Mal00, Mar97, Max93b, McC84, Mel09, Men26, Mer55, MTY79, Mer05, Mey85, Mil75, Mil81, Mil98, Mil99, MST^{+}07a, MST^{+}07b, NS05, NJP92, Nor07c, Nor07b]. Relativity [Nor72, Nor14, Pag21, Pai88, Pal83a, Pal83b, Par00, Pat03, Pat99, Pau21, Pau81, Pau94e, Per23b, Poc98, PB06, Pye73, Pye75, Pye77, Pye82b, PJMS11, Rai46, Ray87, Rei20, Rei49, RR65, RR69, Ren05g, RJS07, RSS^{+}07a, RSS^{+}07b, RS07c, Ren07d, uJR15, Rin01, Rin06, Rob38, Rob39a, Rob39b, Rob49b, RN68, Rob05, Rod04, RR89, Ros37, Rot82, Row06c, Row12, vBRvBR77, Rus31, Rus43, Rus58, Rus69, Rus85b, Rus97, Rus09, Ryc92, Ryc05, Sch27, Sch05b, Sch15, Sch17, ST94b, Sch09b, SS04, Sci80, Sei03b, SPA^{+}68a, SAI^{+}71, Sig92, Sil22, Sil24, Sin00a, Sin00b, Sin05, Sta94, Sta05c, Sta03a, Sta98c, Sta08a, Sta08b, Ste04, Ste98, Str99, Tau79, Tay21, Thi21, Tho83, Tol34, Tol87, Van07c, Van10, VM99a, W.25, Wal79, War92, War93, Waz14]. Relativity [Wei72, Wei15a, Wet80, Wey18, Wey20, Whi22, Whi04b, Whi83b, Whi27, Wie21, Wil86, Wil90a, Wil90b, Wil93c, Wil01, Wil05b, Wil06, Wol03, Woo03, Wue05, You04, Zia94, dMdA05, vD02b, van12, vB12, vB14, vB18, vB20, vG30, vL09, vL11, vL13, vL19, vL21, vL52, vL56a, vL65, She16a, Bel26, Ein13c, Ein22q, Gly90, Mur25, Nic20, Plo06, You11, Din65, Ein12b, Ein05e, Ram81, Row08, She16b, Vog21, Yea08]. Relativity’s [Ano21d]. relativnosti [Záv25]. Relativtätstheorie [Ein22p]. Relativtheorie [Wei12y]. Relaxation [WCL^{+}00, YCKG11]. Released [Ano39c, Gei15, CSW97b]. Relevance [Grü62]. Relevant [Zuo96]. relief [Pep05]. Religion [Ein30j, Ein39c, Ein40f, Ein41d, Ein41e, Ein48l, Gor02, JAvWE95b, JAvWE95a, Mas01, Mor79b, Müh11, Rei41, Rup09, Ben01, Ein31d, Ein82e, Ein91e, Gim12, Hol30, Jae06, Jam99, Jam02, Kal35, Mit87, Mon58, Num09, Par09, Sta07c, Ben01, Ein30k, Ein39c, Ein41d, Jae06, JAvWE95b, JAvWE95a]. religiös [Müh11]. religiös-weltanschaulichen [Müh11]. Religiosität [Ste05c]. religiosity [Ste05c]. religious [Kra08, Müh11, Pow02]. religious-ideological [Müh11]. relive [VWA95]. Relocating [ARS15]. Reluctant [Ber96d, Bra14a, Ano14b, Kri19b, Kra00]. Remain [Ein47f, Pal95, Lin15]. Remains [Law00, Mar11c, FG20, Lin15]. Remark [Ein11c, Ein24b, Göd49b, Ein12a, Ein12g, Ein05d, Ein05-28, Ein05-49, vB19, Ein07a, Ein09a, EL09, Ein11d, Ein11a, Ein12b, Ein14f, Ein18c, Ein18d, Ein19c, Ein20d, Ein22b, Ein22d, Ein23a, Ein25b, vI11a]. Remarkable [Kae38, Rin11b, Rin11a, Rus25, Bra12, Cal17, Cam18, Die05, Jam06, Par04, Rin09]. Remarks [D’A97, EE22, Ein44e, Ein49d, Ein70b, FL49, Fra37, Gol72, Mie07c, Sau08, Wig96d, Ein07f, Ein18i, ES46a, Ein53c, Man80, Urb22, Ein11f, Ein22c, Ein23b, Ein05j]. Remarques [Ein53c]. Remember [Fri79b, Neu06, Rot96]. remembered [Cro04, KK05]. Remembering [Kel04, Wig96k, Foe11]. remembers [HW78]. Remembrance [Sha79a]. Remembrances [Ber74a, Ein55b]. Remind [Duf46]. reminiscence [Kis80]. Reminiscences [BHB80, Ber82b, Ein91g, Ger79, Hof82a, Jof62, KW87, PP95, Ros82a, Ros72, Rot95, Str82, AA75, Far79, Nie89, Sch82b, Wil75, BFB82]. Reminiscenses [RS80a, Uhl80]. remote [EM31b]. ren [Seg86b]. Renaissance [BGLR17]. Rene [Tho88b]. Renewal [Rec07]. Renn [Bal96, Cah90b, Har96a, Kan92a, Kil95, Nev92, Noe93, Pye89, Pye95, Sch92b, Sie97, War95, War96a, War96b, She16a]. renormalizable [KP55]. renovation [Gra99]. replace [GE97]. replica [SWU09]. Replies [Ein47e, Ein63b]. Replik [Ein14b]. Reply [Abi04, Ein14b, Ein48m, Ein53f, KH05, Mer12, Whi05, Zah78, Ano69c, EP06, Flo06, Kra06, Mun36, Ray92, Abr12b, Ein05d]. Report [Ano39a, Ano41, Com56a, Duf46, Edd20a, Edd06, Ein21o, FR49a, HCD^{+}78, Smy76b, Buc08, MW46a, MW07, Rol89, Smy45b, Smy76a, Smy78, Smy89, Uni96, Kae45, vE10]. reported [Bey49, Bro10b]. Reporter [Her46b, Her85b]. Reports [Hay91, HSD34b, KRW05, LdB12b]. Representation [RS99, Ano22a, EBB41, Flü74, EM34]. Repression [Wal06b, Wal03]. Reprinted [Ein86c, Ein08b]. reprints [Eur69]. reproduites [Ein05-47, Spe62]. Republic [Hom04, Sch05a, Sch65, Sch05c]. Republication [Wey09]. repulsive [Die88]. Required [Ein05-35, Ein07d]. rerum [Die24]. Research [Ahe03, Ano15d, Bus13b, Cra90, Eds82, Ein96c, For80a, Hei47b, Hei47a, Hen82b, HRR03, IP76, Lew79, Ram79a, Ree16c, RN16, RR17, Sch74, BN88, Bjø19, Bjø20a, Bjø20b, Bjø20c, Bre08, BEM12, CD23, DC10, Das00, DCL23, Duc11, Ein18a, Fey84, GRS16, GZ97, GZ06, Han18, Hen85, Hof18, Hom06, Irv67a, Irv67b, Irv83, Kut09, LRH22, MH09, PS22, Ree05a, Ree11a, Rei05, RWL23, Ren08, Sch13b, Wad81b, Wit62, vH62]. Researcher [Bus13b, Ein13d, Ein22f, Ger69, GH84, Myr21b, Myr21c, She16b, Som55b, Vog20]. Researchers [Ein08a, Bra14b, Kon15, Spo16]. Researches [D’A75, See25, Cha49a]. Residents [Mar95b]. Residual [Kox95]. Resistance [Jun14, Ell90, Bau95, Bau97]. Resister [Ein53l]. resisting [Cam09b]. résolu [Ano55c]. résolution [Dat38]. resolved [Van98]. resonance [Man13, Par99]. resonances [IAS^{+}98]. resonator [EH05b]. Resonators [EH10a, EH05b]. Resource [Ree05a, Ree11a, Ree16c, Ein95b]. resource-based [Ein95b]. Resources [And17, BHJ^{+}00]. respect [Ein20g]. respected [Ber00]. Response [Alb92, BF94, Bjø20c, Ein47f, Fin90b, Gol70a, Whi04a, Abr12a, Ano69c, Cur92, Flo06, Gol68, Mil82, Rit09b, Bjø20a, Ein05-28]. responsibilities [Ein07e]. Responsibility [Ein50c, Tre79b, Wal06b, Zim82, Bad83c, Bad83d, Lam79, Sch00d, Wal03, Wee87, Ein50n]. Rest [Hei10, Ros49, Sch07d, Ein05c, HFN06, Jag01, Sta93c]. resting [EL05d]. Restreinte [Ein21j, Ein56d, Ein76a, Gou10, Spe67]. Result [Ryf14, SAI^{+}71]. resulting [HS39c, HS44]. Results [Ano19b, Ano19c, MA31, SPA^{+}68a, SPA^{+}68b, Str49, Fab79, GTP22, Kon15, NHG89]. Résumé [Bau20]. resurfaces [Dac11]. Resurrected [Bel98, Ber90b, DC10]. Retention [Gru57]. Rethinking [Lan06, Pfe77]. retirement [ABdB^{+}53]. Retort [Ein12g]. Retracted [Kox95]. Retraction [Kri19a]. retrait [Pat85]. Retreat [Sch93b, Abb98b]. Retrocausal [Cor15c]. Retrospect [BBBM90]. retrospective [Row08]. Return [Tib05, SKD^{+}87]. returns [Abb98b]. reunion [KK05, LdB12a, LdB12b]. Rev [GHK^{+}07, Leg03, Nam10, ’t 02]. reveal [CSW97b, Wil11b]. Revealed [Kae45, Kur12, Egd14]. Revealing [BOPS99]. reveals [Ano06d, RMI21a]. Revelations [WD96]. reversal [Bel75]. Review [Ach13, Alf09, Ano39d, Ano50a, Ano60, Ano84a, Ano84c, Ano84b, Ano99b, Ano99c, Ano04b, Ano12a, AA75, Arm38, B.50, Bad72, Bad83b, Bad96, Bal72, Bal75a, Bar05f, Bed01a, Beh05b, Ber07a, Ber06b, Ber03, Ber00, Bey06, Bey09a, Bey09b, Bey12, Bla50, Bla76a, Bla76b, Bok15, Bor50, Bot13, Boy88, Bri18, Bri82, Bro66, Bro72, Bro39, Bro00, Bru03c, Bru06, Bus13a, C.51a, Cah90b, Cah90a, Cao06, Cas86, Cas05b, Cli09a, Coh51, Con83, Cou52, Cus82, Del09b, Del11b, Del11a, Del12, Dro20, Duf46, Dys76, Ebu16, Eds82, Ein24a, Ein24i, Ein30d, Far79, Fel11, Fer07a, Fes56, Fis79, For71a, For01, For05, For70b, For70a, For09, Fri73, Gam42b, Gea94, Ger15, Gil12, Gil19, Gin82, Gli84, Gly90, Gol84a, Gon05, Goo85, Goo02, Gor02, Gor07a, GRC18]. Review [Gre19a, Gre19b, Hah73, Ham11a, Han80a, Har07, Har96a, Har96b, Har02b, Har79d, Hea13, Hea64, Hei74a, Hei79a, Hei79b, Hen82b, Hen88a, Hen88b, Hen00b, Hen11c, Hen16, Hen21, Her17, Her21, Hes04, Hie68, Hie84, Hoc93, Hoe18, Hoh13, Hol05b, Hot15, Hou65, Hub13, Ivr99, Jam70, Jan73, Jan77, Jeh51, Joh95, Jon01, Jos95a, Juc04, Kae38, Kai05a, Kat52, Kau14, Ken25, Ken11, Ker89b, Kev05, Kil65, Kil84b, Kil95, Kil98, Kle70b, Kle70a, Kle73, Kle84, Kle14a, Klo63, Kni12a, Kni12b, Koj08, Kra02, Kri08, Kri99, Kuh54, L.58, L.89, Lac08, Lan08, Lan15, Len39, Lin56, Lin60, Lit09, Loe88a, Loe88b, Maa13a, Maa13b, Mal65, Mar73, Mar10, Mar11b, Mar11c, Mas01, Mau09, Mcc87, Med76, Mel82, Mic68, Mil11a, Moy81, Myr21b, Myr21a]. Review [Myr21c, N.38, Nic20, Noe93, Noe11, Nye88, Oha79, Ove96, Par06, Pec04, Pep04b, Per10, Pes13, Plo06, Pol83, Pri99, Pul11, Pye86, Pye93, Pye97, Pye10, Qui40, Ram79b, Ram79c, Ram79a, Ram81, Ram90, Rec86, Red84b, Ree15d, Ree17d, Ree17c, Rei50, Rei51, Ren21, Rob48, Row00, Row06a, Row06b, Row08, Row17, Rus17, Sar35, Sau16, Sch87a, Sch50a, Sch79c, ST06, Sch09a, Sch07c, Sch78, Sch92c, Sch03c, Sci80, See65b, See65a, Seg91, Sei79, Sei84, Ser88, Set10b, Set10a, Set11, She15, Sho21, Sie97, Sie07, Smi05a, Smi51, Smi89b, Smi89a, Smi02, Sta05a, Sta99b, Sta00, Sta20, Sta21, Ste95b, Ste05d, Ste90, Sto17, Str04, Str90, Tha77, The12, Tho11b, Top89, Tou05, Tra00, Tre88, Tre21, Tug98, Tur05, Tur06a, Van22, Vis94, Vis95, Vog20, Vog21]. Review [Wac06, Wal94b, WM08, Wal02b, War95, War96a, War96b, Wat00, Weg08, Wei94, Wei03, Wid07, Wig79, Wil16a, Wis21, Woo06, Wre66, Yan87, Yea08, You95, dP94, vD15, vB01, vB02, Cah04, Coo04, Dix00, Duv64, Ein24c, For84b, Han80b, Har95, Her03, Hol91a, Joh06, Kli62b, MFHR06, OONM15, OONM17, Pal83b, Pap15, Put09b, San07, She16a, Str05a, War03a, Anoxxc, Ein06d, Ein13c, Ein16i, Ein18h, Ein22q, GKP12a, Jun05a, Ken05b, ABC^{+}06]. Reviewed [Ano82c]. Reviews [And81, Ano21k, Ano15k, Bad15, Bel26, Bon94, C.50, C.51b, Coh54, Dow80, EWJ^{+}25, For80a, Fre08, Gol84b, Ham97, Har91, Hen11a, Hen86a, Hus15, Ihd74, Kar05, Lam30, Leg15, Mur25, Nad79, Nav15, Nuv11, Pep04c, Pie24b, Pye80, Pye89, Pye95, Pye06, Rei86, Ste22, Swa02, Tay21, VB96, Hah81, Sch10a, SW12, WK12, Bäu83, Jun05b, Rai46, Woo28]. revised [Col68, Sha54, vB02, Dür80, Lau50]. Revises [Sul64]. Revision [Lor77, Lor77]. revisionism [Mad07]. Revisited [Aga06, Aga07, But07, Van10, Heh98, MH12, Nau16, Nik12, OM14, OM22, SL93, Teu96, Ein05e, Ein12b]. Revisiting [ACH^{+}03, BDH^{+}00, Gor17, Kuh84, LLH22, PJMS11, Ree17b, RDS03]. revival [GMHB02]. revivals [HMW24]. Revolution [Aga93, Ano19d, Bor71, Gal94, GR23, Jan02, Kae48, Moy79, Nug99, Pea08, RR14, Smo19b, van12, App11, Bel99, Bol04, Bro80, BRS03, CM86, CM97, Dav95, Dre87, Far03b, Gol84c, Gre05b, Gri12, Isr98, Kle05, Lor08, Nug00, Ren06, Rob58, Spo02, Sta08a, Tam14, TH79, Ven94a, Ven94b, Ven94c, Wol95, WB00, Zah89, Ano05o, Kle05, Del11a, Hen09, PJMS11, Ren06, Bey06, Cas05b, Kar05, Mar10, Set10b, Bey12, Del09b, Set10a]. Revolutionär [Jun08]. revolutionaries [Far03b]. Revolutionary [Bra14a, Ano14b, Dra14a, Jun08, And15, Kra00, Rig05b]. Revolutionized [Hal17, Lem09]. Revolutions [BHJ^{+}00, Ne’06, Coo99]. Revolutions/Pais [BHJ^{+}00]. Reworking [HRR03]. Rey [Cli09b]. Reynolds [BE69]. Rezension [Ein06d, Ein24i, GKP12a, BS92, Eis00, Hen09, Hop07, Jun05a, RK94]. Rezensionen [Hah81, Jüt21, SW12, Bäu83, Eng17b, Jun05b, Kai81, Ros86]. Rheology [Sim80]. Rhetoric [Bel93, Cec01]. rhetorical [Hal04b]. rhétoriques [Hal04b]. Rhodes [Her21]. rhymes [Mer16]. Ria [Del13]. Rica [Rol89]. Ricci [Agu90, CT24, Mal91, Sid76]. Rice [Pie24b]. Richard [Ano93b, Dow80, Hen09, Her21, Hor00, Mar10, Ree17d, Set10b, Ber94, BR93, CD07, Del11a, Fey64, FF95, Fey00, Fey05, Gle92, Hal17, Kra11, Lei00, PJMS11, Shu17]. Richardson [Hen82b]. Richart [Nah01]. Richer [C.50, Fra50, Fra51]. riches [Rei96a]. Richthofen [SR11]. rid [Gio16]. Riddle [Lan74c, Ren07a, Ber68b, Ber69a, Ber87a, Ber92b]. Ridenour [Ein47e]. Rider [HB05]. Ridge [JJ81, Olw97, Olw04, Ove87, Rob50]. ridges [Rot92, Rot97]. Riemann [Ein30a, Ein30r, Ein27a, Ein28e, Ein30a, Ein30r, Kov20, Pie23, Pie24a, Sch02b, See25]. Riemann-geometrie [Ein28e]. Riemann-Metrik [Ein30a, Ein30r]. Riemannian [Ein28e, EM31b, Hof49b, Lan49, LC07]. Riemannschen [Ein27a, EM31b]. Riesen [Ren06]. riesige [Bre08]. Right [Ano33h, Ano91, Ano11a, Ano17c, GP19, Mei62, Mei63a, Mei63b, Mus04, Sch15a, Wil86, Wil90a, Wil90b, Wil93c, Wil17, Yun15, Buc20, Bun79b, Hes15a, Mui94, Sta93a]. Righting [War03a]. rights [Blu89]. Rigid [Fok49, vI10b, vI11a]. Rigidly [Sta80a, Sta89a]. Rigorous [GN11]. Rim [War99]. rims [Rot92, Rot97]. Ring [CLD^{+}03, Wal88a, HTS^{+}88, JRT^{+}91, LCL^{+}90, SLLS09]. Ringen [ER13, Flü74]. Rings [Ano33g, Wat98, Pea90]. Rinnovamento [Rec07]. ripple [FN13]. ripples [Cho16a, Dys16]. Rise [BC82, EI64, Hen86a, Hol89b, Joh06, Kai05a, Mac77, She83a, She83b, Smo07b, Uvn82, Zah77, Aas90, Bra12, Cha05c, Col12, Flü74, HBRD97, JDB^{+}97, Ker89c, MR82d, MR82e, MR87b, MR87a, Nel14, Smo06, War03b, Wea12]. rise/Inertial [JDB^{+}97]. Rising [RB05]. Risks [Sem08]. risky [Dys16]. Rita [Pye97]. Ritter [Ano84a, CE79a, CE79a]. Ritz [FP16, Mar04d]. rivalry [Acz09]. River [Nov98, Ein26c]. Riverhead [Ivr99]. rivoluzione [Tam14]. Rjukan [Wig86]. Road [Ano07a, Ano15k, Bar05f, Ber98, Kra96b, Sch05b, Bet91, BCD93, Eis11, Erc55, Goo03b, GR15, Har02a, Jan05, Man13, Nic87, Pen05, Per09, uJR15, RS11a, She16a, Uff06]. Roads [Hor09, Mei62, Mei63a, Mei63b, Wei15c]. Robert [Ano84a, Bal96, Bar05f, Ber82d, Ber03, BieWSK67, BS05, Bot13, Cah90b, CE79a, Cas05c, Els95, For80a, GRC18, Ham97, Har96a, Hea13, Her03, Kel06a, Koj08, Kra02, Lan08, Mar11c, McM05, Noe93, Pol02, Pye89, Pye95, Pye97, Pye06, Ram82, Row00, Sch03c, Smo07a, Tay21, Tre21, Tsi08, VB96, Wal99, War96b, Ber05, Bru03c, But23, Car08, CE79a, Con05, GSW95, Her02, Her21, Lan00, Rou64, Rou65, SW95, Tur05, VB96]. Robertson [PS03]. Robinson [But23, RTR87]. Robles [Fel82b]. Rock [Zoe09b, Zoe09a]. Rocked [Ano39h]. Rockefellers [Tol76]. rocky [Boz14]. Rodger [Smo07a]. rods [Gio14]. Roemer [Coh40]. Roger [Gea94, Sch87a, Smo07a, Vog21, Bon94]. Roku [Ish77]. rola [Inf56a]. Role [Bič00, Eck90, How94a, Ken54, Kri19b, Leg15, May95, MM58, Nad02, Ros82b, Tre79b, Wal16, Zah80, vD09, Arn00, Bus49, Bus68, Bus85, Das12, Ein19j, Eva95, Gol70b, Khr08, Lan59, Mac01, She88, Sme14, Sto12, Wal91b, War91]. Roll [Rod92]. Rolle [Ein19j]. Rollin [dHvdB49]. Rolling [Tho83]. roman [Kum11, Sre29, Gri15, RMHD11]. romance [Ove00, Ove01, Pri12a, Smo07a, Sre29, vB01]. Romancing [DR89]. Romanes [Cun22b]. Romanian [EI57]. Romantic [Rad08]. Rome [JRG02]. Romon [Cob07]. Röntgen [MGG^{+}95, Mar95a]. Röntgenstrahlen [Ein18j]. rooie [vC05b]. Room [Joh13, PSM^{+}13, DDD^{+}06]. Room-temperature [PSM^{+}13]. rooms [Mar00]. Roosevelt [Cli09a, Ein91b, Dem15, Ein39a, Ein97, Lucxx, RMR^{+}50]. Roots [Fin81, GGK13, GR23, Hol98a, Kat04, Sti13b, Sti13c, TVL^{+}09a, Cea97, Cea98, Feu71a, Feu71b, Hof97b, Hof99, Mar03]. Rosbaud [Lin60]. rose [Sza84, Tho85]. Rosen [Kie15, Nag71, Web93, ZDM08, Zuo96, AKBAE21, Alb92, Bel64, Bha86, Blu22, CC84, Com95, CHR90, Cos87a, CPDH^{+}87, DR08, Erl72b, Erl72a, Har97, HR13, Hni98, HV88, Hoo70, Kau85, Khr08, Krü00, MPBL09, MS82, Mol74, Per05a, RDB^{+}09, RMR96, Rus85a, Ryf14, Sch71, SZ88, Sel88, SW87, Sta06, Tör81, Van74, WW57, YC18, ZXL^{+}19, dB85, dB87, dM86]. Rosenberg [RM97, Rob01, Rob14, Sch10b]. Rosenfeld [CS79a]. Rosenkranz [Bad15, Lac08]. Rosmorduc [Bro83]. Ross [Ahe03]. Rossi [Del13]. Rostock [SW12, GKP12b, SW12]. Rostocker [GKP12b]. Roswitha [BS92]. Rotating [Fet09, MST^{+}07b, Pag21, Sta80a, Sta89a, Wei14b, Wei15b, You98, BR99, Grø82, Hon96, MST^{+}07a]. Rotation [Ano17, Ano17c, HEB^{+}80, Val49, CT24, Ker87]. Rotation/Haas [HEB^{+}80]. Rotter [Fre08]. Rough [EES49, Dee08]. round [Ano33b]. Rouse [You95]. route [Her79a]. Routes [Ne’06]. Routledge [CF98]. Rovira [Pfa12]. Row [Ree17c, Ano81b, DC10]. Rowe [Koj08, Lan08, Pye06, Smo07a, Tsi08]. Rowland [HG06b]. Royal [Ano26b, Lin56, Pla20a, Ano13a, B^{+}06, BPS87, Hil39, Pla20b, Sch28]. rozwój [EI59, EI62a, EI98b]. Rt. [Eve39, Eve13]. Rubbia [Bel88a, Bel88b]. Rubiner [BSS^{+}17]. Rudder [Foe23]. Rudiments [Ano22a]. Rudolf [BC71, Rob48, The12, How96a, Lau12, Wei15c]. Ruetsche [Pye95]. ruhende [EL08c, EL05d]. ruhenden [Ein05c, Ein05-35]. ruhu [Tip12]. Ruin [TM77, TM90, McM05]. Rule [Bro21]. Rules [Ano70, Tho15]. Rummel [Van22]. Rumors [Bre05]. Rundfrage [Ano69c]. Rundfrage/Drei [Ano69c]. Runs [Bro15, FS06]. Rupp [Van07a, Van07b]. Ruptured [CG90]. rural [Abb98b]. rush [Hom06]. Russ [Ano60]. Russell [Bro72, Mar73, Ser88, Ano06i, Ano16b, Ein44e, FW07, HY15, Kap61, Rot95]. Russia [Ano96d, BV02, Pec04]. Russian [Hoh13, Jof62, Ree17c, Ano33c, Ano79a, BP^{+}01, CH98, Dai67, È03, Ein21r, Ein22j, Ein40g, EI66a, Hec16, HC98, Inf56c, Jof55, Kol40, Lan67a, Rol19, Smy46, Tol76, Was92, Whe70]. Russian/French [Rol19]. Rutgers [L.89]. Ruth [Kri99]. Rutherford [Eve39, Eve13, Jar08a, dCENdCA64, Ano16b, Bad85, Bou99, BS79, Cam99, Cam00, Eve39, Eve13, Jar08b, Kri19a, Kri19b, Kri19c, Oli72, Stu86, Tod14, Tre75b, Wil83]. rutschen [Bre08]. rychlostí [Žáč12].

S [Aga07, AA75, Duf81, Ebu16, Ein14b, Eng17b, GSW95, Ham97, Har95, Hei53a, Hen82a, Hen82b, Hen16, Hie84, Kle82a, Nav15, Nye99, Ree17c, Sch78, Sci80, Sie97, Sta07c, Tay21, Tho11b, Vog20, Vog21, Wil36, Jof62, Kot93, NHG89, Sal19, Těš12]. S. [Ano57b, Klo63, Sri99, Wal91a]. S2 [AAA^{+}18]. sa [Ano80a, Fra91, Mag95, Mil10a, Mil12a, Sto96]. Saale [HH14]. Sabbath [Bar93, BL11]. sabotage [Hau54]. SAC [FR49a]. Sachen [LT15, Lin17]. Sad [Mar95b]. Safe [Sus08]. Safeguards [Ano12]. Safety [Sem08, Gré19c, Hac87, Hac94]. Saga [Rin02, Rob50, Tol76]. Sagamore [Mau09, WM08]. saggio [Cap01]. sah [Ein21g]. Saha [SR99]. Said [Ano20f, Enz05, Res50, Ano19b, Mau90, Mon12, Pra12]. Sailing [Ano31a, dGK11]. Sakharov [Mar88]. Salam [Alf09, Bus12a, Fra08, Gha82]. sale [Had87]. Salts [Ein05-39, Ein02b]. Salutary [Cro16]. Salvation [Sch93b]. Salze [Ein02b, Ein05-39]. Sam [Dac20a, Hoe18, vC16, vCH18]. Sammlung [BSW22, LEM13, LME23]. samtiden [Str83]. Samuel [Seg91]. San [Joh11a, RFP^{+}45]. Sand [Kao10]. Sandra [Per10]. Sands [Edg14, Mog02]. Sanho [You95]. Santa [Hel11]. Santillana [Sha54]. Sanzsaenz [Nav94]. Sapienza [JRG02]. Sarasohn [Ste95b]. Sat [Sha92]. sata [Hot15]. Satellite [Ano11a, Stu86]. Satellites [ESA15, Gin59]. satisfied [Pit16]. Satyendra [Das12, Meh75b]. Satz [EH10b, Ein15b, Ein05c, EH05c]. Satzes [Ein07i, Ein05-32]. Sauer [Bad15, Lac08, Pye95, Sie97]. Sauerbruch [Fla15, HS14, Fla15, HS14]. Sauerstoff [Kon15]. Sauerstoff-Molekülen [Kon15]. savant [Sto96, Van96]. Savants [Hum51, Spi96, Van96]. save [Cra07]. saves [Ano15d]. savetovanja [Mag95]. Savić [IK03, Pop98, Pop05]. Saw [Ein21g]. Say [Ano39h, Ano41, Lan67c, Nor12, Wei13e]. saying [ET20]. Says [Ano29d, Bro21, Ein54c, Law07]. sbarcato [Ano82d]. Scaffold [JR15]. Scalar [Nor07d, Eng14, Nor92]. scale [Car02, HE08, Myr21a]. Scaled [Ren07c]. Scales [GK05]. Scan [Nad02, McC01b]. scandaleuse [Sch10a]. Scandalous [Sch10a]. Scanning [BR85]. scare [Cir07, Cir08]. Scarry [Lan15]. Scattering [Bar49, Sin04a]. scenario [Mar15]. scenes [LHZ^{+}23]. Schalit [Gru07]. Schallausbreitung [Ein20l]. Schardin [Ano79g, Ano79g]. Scharl [Zau05, Ein54h, Zau05]. Schatten [Kra81, Ros86, Ste89, T– D82a, T– D83, T– D85, T– D88, T– D93, Chi94]. schauen [Ing05]. Scheidweg [Bre05]. scheinbar [Ank15, Ein14h]. Scherrer [MTVB88]. Schicksal [ARK^{+}99, Ste89]. Schicksal/Kärger [ARK^{+}99]. schienen [Ein19b, Ein19c]. Schilpp [For80a, Krö53, Bla50, Coh51, Jeh51, Ram79b, Smi51]. schlägt [Kon15]. Schlick [ER13, Hen86c, SW12, Ein05a, ER13, Hen86c, How84, Pul06, SW12]. Schlüssel [Bre15]. Schmid [BS92]. Schoenberg [Sch10a, Sch87b]. scholar [Sto96, Har13c]. Scholars [Wal81, Ano81a]. Scholarship [Mal09, Räz16]. Schönberg [Sch10a, Ton97]. Schönbergs [Sch10a]. School [Ano26a, Pap15, WA03, CC74, CC98, Isr73, RK86]. schooldays [Tan11]. schools [Pal99]. Schopenhauer [Ein59d, How97b]. Schöpfer [vL55c]. Schöpfungstages [Rho88a, RT90]. Schou [Kat52]. Schreurer [Ded85]. Schritte [Rec02]. Schrödinger [Adl03a, Ano51a, ARK^{+}99, BC94, Cam88, Hei61b, Hie68, Hou16, KDK87, Kil65, Klo63, LW34, Prz86, Sau16, Wil36, Ano15e, Bal92, Bas92, BD92, Bit96, Bub79, Cam88, Cec01, Dat38, DS83a, Ein18d, Ein18k, Ein27b, GGW94, Göt92, Gri84b, Gri84a, Gri95b, Gri95a, Gri12, G^{+}11, Hal15, Han79a, Han79b, Hol88e, Kil87, Kil89, McL99, MR87b, MR87a, Meh87, Moo89, Moo94, Moo96b, Prz63, Prz67, Prz11, RF69, RA11, Rüg88, Sch35c, SKD^{+}87, Sci61, Win13b, dB65]. Schrödingers [Ein18d, Ein18k, Ein27b, Win13a]. Schroedinger [Mic68, Ano33c]. Schück [Kat52]. Schucking [SH99]. Schuld [Sti05]. Schüler [RKW05]. Schülerwettbewerb [Spe79]. Schulman [Wal99]. Schulmann [Cah90b, GRC18, Har96a, Koj08, Lan08, Mar11c, Pye89, Pye95, Pye97, Pye06, Row00, Sie97, Smo07a, Tsi08, War96b, Har96b, Kan92a, Kil95, Nev92, Noe93, Sch92b, War95, War96a, Bal96]. Schultern [Ren06]. Schumacher [HEB^{+}80]. Schuster [Gor07a, Kae38, Sch07c, Weg08]. Schütte [Chi08]. Schütte-Lihotzky [Chi08]. Schütz [Ein17a]. Schwankungen [Ein19b, Ein19c, vB19]. Schwartz [Con83]. Schwarze [Gen05, Str05b]. Schwarzen [Ber96b, Bes04]. Schwarzenbach [Sch05e, Sch12f]. Schwarzschild [Ein16g, Sch05b, Sch07b, Wei14b]. Schweber [Tho11b]. Schwed [Pla20a]. Schwedischen [Jüt21, Pla20b]. Schweitzer [Ein55a]. Schweiz [Sch05e, Sch12f, See52]. Schwelle [Kra88]. schwerelos [Vaa17]. schweren [JDB^{+}97]. Schwerkraft [Ein11i, Ein05-31, de 15, dP17, vL23]. Schwerpunktsbewegung [Ein06a, Ein05n]. Schwierigkeiten [Hei93]. Schwinger [Dai67, Dys65, MM00, Sch94]. Sci [Ein86c, Flo06, Kra06, LG97]. Sciascia [RRR07]. Science [Abb98b, AMM99, Aer99b, Aga75, AH93, AG02, Ano19b, Ano19d, Ano38b, Ano47b, Ano67, Ano81e, Ano99f, Ano04b, Ano09g, Sch12b, Ano12c, ARS15, ADS83, Aro96, Arr07, Asi83a, APWB95, BDH04, BNY81, Bar05f, BMS12, BF78, Ber01, Ber18, Ber00, Bey06, BIT01, Boh46a, Bru03b, BC71, Byr80a, C.50, Can16a, Cao04, Cao14, Cas05b, Cha17, Coh80, CHR96, CCG93, CF82, Cra02b, Cre10, Ded85, Del09b, Din72, Dir79c, Dir79b, Ebu16, Ein21d, Ein22h, Ein30j, Ein30k, Ein30l, Ein30m, Ein31l, Ein33a, Ein34c, Ein40b, Ein40f, Ein41b, Ein41e, Ein42, Ein48b, Ein48l, Ein50m, Ein50c, Ein59f, Ein70c, Ein82e, Ein91e, Ein08a, Eli92, Elz95, Esh09, Far03a, Fer82c, Fey80, For05, For84b, Fra49c, Fri89, Fri01, GGK02, GHS08, GGK13, GHS18, Gam42a, GR07].

Science [Gib19, Gil84b, Gla00, Gle92, GR06, Gol84a, Goo85, GO81, Gra94a, Gun15, Hab69, Has04, Hei83a, Hei83c, Her00, Hes04, Hoc93, HRR03, Hol98a, How84, How94a, How04, How05b, Jon08, Kae31, Kae39, Kae13b, Käl22, Ker89a, Kha20, Kle84, KV11, Kra11, KUW82, Lak85, Lar84b, Lee93, Leg15, LMOS15, Lov86, Mah17, MGG^{+}95, Mar11d, Mat86b, Mau09, MB94, MSW88, Mer38, Moo81a, Moo81c, Moo81b, Mor79b, Moy79, Nad02, Nor49, Pai00, Par06, Par09, Pau94e, Pei83, PP05, Pip83, Pla32, Pla33, Pla77, Poi02, Pol83, Pow94a, Rab70, Ram82, Rec07, Rec86, Rec10, Red84a, Rei86, Ren13, RW94, RWY97, RKS08, RBS^{+}08, Ros82b, Rot87, Rup09, Rus81, Sch12a, Sch57, Seg91, Smo07a, Smo07b, Sno61, Sti98, TS01, TB02, Tha77, Tre79b, Tre88, Tur06a]. Science [UM86, UM92, Van07c, Van02, Vol02, Wal16, Wal94b, WM08, War84, Wat00, Web73, Wei88, Wid95, Zah77, Zho85, Zim82, vD15, vB02, Aas90, Abr98, Ano06g, Ano15f, Arp98, Bad83c, Bad83d, Bar99, B^{+}01, Ben01, Ber84b, Bet91, BSPKN03, Bri05a, BB13, Bru97a, Bru97b, Bru76, Buc20, Bus49, Bus68, Bus85, Cal03b, Car10, Car13, Cas92, CGR^{+}97, Cec01, Cha09, Cha87, CF80, CSC23, Coh49, CH80, Con02, Coo83, Cor15a, CSvB99, Cor03, Cra84b, Cra92, Cur82, Das12, Day12, Die05, Diz05, Doh06, EN97, Ein59d, Ein09b, Ene12, ER13, Erb10, Far02, Far13a, FW07, Fer05a, Feu74, Feu82, FS06, Foe11, Fri83, Fri12, FFF09, Gim12, Gli88, Goo03b, GS59, Gou47c, Gra96, Gre05b, GM80, Gru05b, Gru05c, HKS10, Hal04b, HES^{+}12, Han18, HH00]. science [HH02, Har02a, HH03, HH04a, HH05, HH06, Har89, Har81, Har00, Har03a, HT10, Hei86, Hei00, Hei75b, Hei77, Hes15b, HD02, Hof08c, Hol30, Hol86, Hol91b, Hol96b, Hol98b, Hol00a, Hol05c, How05a, How14, How76, Inf33, Inf34, Inf78b, Jak82, Kea18, Kel06b, KJY24, LM94, LSW12, LW71, Lig05, Lin07, Lin08, LR14, Mag95, Mar17, McG93, McG98, Mit05, Mot79, Nac79, NS88, Nol88, Num09, Obe16, Pai82, Par04, Pic08, Pla59, Pla81, Rab63, RR82, RCRC97, Red84b, Ree14c, Rei41, RG07, Ren21, Rod22, RK94, Sam10, SE79, Sch96a, SNA^{+}06, Smo06, Som55b, Spo02, Sta07c, SSH13, Stu14, Stu23, Tay49, Tip10, Tur03, Val06, Ven88, Wal95, WL80, Wei94, Whi93, Whi94, WG94, WG95, Whi22, Whi04b, Wig82, AH93]. Science [Ano97, Ano23a, Ano97a, Bar05f, BNSY98, BCS82, Bru03c, EWJ^{+}25, Ein39c, Ein41d, Ein50k, EMvL79, EBRC93, FR49a, GNS99, Har06b, Hil39, Kae31, Kra88, Leg15, Möl84, Moo81b, Poi02, Tur05, Zel90, Boy88, Fer07a, Hen11a, Hol91a, Kra02, L.89, Mur25, BCR93, For09, Kni12a, Kni12b, RSS97, Sei84, Ste95b]. Science/Goetzberger/Voß/Knobloch [MGG^{+}95]. Sciences [BH17, CW84, Cra90, GSW95, GN86, Kie82, Kra17, Pla20a, Sim06, Tha77, Vis94, vDP17, LOvT12, BNY81, FTG08, How76, Kat89, Lin99, LT15, Lin17, MFHR06, Nor10, Pet08, PSAA96, Tor03, Ano25c, Ein33d, Fri12, Gil12, Pla20b, RS11b]. scientiarum [Buc80]. Scientific [Ach13, AP00, Ano00, Ano22a, ADS83, Bäu83, BHJ^{+}00, Cah04, Cha74, CD07, Coo04, Cun22b, Din49, Dro20, Eck90, Ein21c, Ein29g, Ein48j, Ein50e, Eis00, FS87, For73, Gal94, GSW95, Gil12, Gin08, GH01, GW96, Har69, Hay91, Hol79b, Hol80, Jan02, Kor79, Krs04, LS93, MH09, Mah01, Max93a, Max93c, Max93b, McC77, Mil10b, Mol63, MM58, MG02, Ne’06, Ner86, Pye76, Ram81, Ree15d, Rei82, Rod19, RS13, Sau04a, Sau04b, Sch00b, Sch88b, She15, Smo07a, Sta05c, Tib13, Top89, UM92, Van07a, Viz79, Wal16, Wei06b, Wet80, WG96, Zuc77, Zuc96, Zuc18, Zwe80, dB49b, vB01, vM93, vM96, vM99b, vM01, vM05c, Acz09, Ano69c, Ano97b, Bey49, BDR93, Bra15, BB13, Bro80, Buc94, But23, BRS03, Cas03b, Chr22, Cor15a]. scientific [Cra01, D’A50, Eur69, Ein50b, Ein05z, ER13, Enz02, Fla15, FTG08, Gav95, GRS16, Gol84c, Gra99, GP11, GR20b, Hah62, Hah66, Hah67, Har02c, HG07, HMW24, Haw07, Haw11, Hei55, Hei74c, Hen82b, HRP90, Her01a, Hol73, Hol78b, Hol88d, Hol98c, Hün18, JW11, JP55, Käl22, Kle59, Kra90, Lau15, Mag03, Mal00, Mas05b, MM00, Mer16, Müh11, Ner84, Nye72, Ove00, Ove01, Pal99, Pat99, PHvMW93, Pop07, PL17, Rec10, RWL23, RS80b, SW94, Sch56d, Sch56b, Sch56c, Sin04a, SR04, Tay49, TTB20, Vog20, Wal02a, Wee87, WS^{+}01, Wey20, Wid95, Wig67b, Wig96m, Wol00, ZZ23, vMHW85, ABdB^{+}53, Ein52b, ER13, Har79d, HvMW79, PHvMW01a, PvM93, PHvMW01b, Ano05b, MKBF30a, MKBF30b, Nad79, Sei79]. scientifica [Lau15]. scientifiche [Seg91]. scientifici [Pal99]. scientifique [Bra15, Chr22, Pop07]. scientifiques [Ein50e, Ein52b, Pat99]. Scientist [Ano29a, Ano29b, Ano35, Ano50c, Ano55d, Ano77, Bro21, Cam99, Cra79, De 10, Del09b, DSMH85, Eds82, Ein46f, Ein52c, Ein79b, GA02, Kae45, Leg15, Nic17, Nic19, Ree17d, Sch49b, Sch49c, Sch49d, Sch51b, Sch70, Sul64, Wil96, Bal72, B^{+}99, Cha09, Con17, Diz05, Eck01, Ein79f, Ein93c, Ein07e, Fey98, Fra08, GE97, Hah70, HEWW15, Hor00, Inf41a, Inf41b, Inf63, Inf68, Inf74, Inf89, Inf92, Jer02, Lam79, Mar79, Myr21b, Myr21c, Par03, Say85, Sch02a, Sch49a, Sch51a, Sch55, Sch79a, Sch83a, Sch00d, Sev99, She16b, Shu17, Smi65, Smi70, Sre29, Sta36, Wad81b, Zei99, Alf09, Krö53, Vog20, Bus12a, Coh51, Jeh51, Smi51]. Scientists [Ach13, Acz01, ACU^{+}54, Ano31b, Ano39h, Ano41, Ano45a, Ano78, Ant15, Bad95, Bar05f, Ber01, Ber18, Bey77, BMW89, Bow87b, BEI^{+}55a, BEI^{+}55b, BLM19, CD07, ČA99, DG99, Duf46, Ein48m, EBU^{+}52, Fie07, GB89, Gol95a, GR63, GR65, Hec16, Jam30, Kae31, Klo97, LL82, MP99a, MP01, MP12, Mer79, MMMM96, Ree17c, Rog10, Str68b, Tho88b, Tib13, Van07c, VFIS48, Wea76, Whi55a, Wol62, Acz03, Bri95, Cas14, Cat04, Cor03, Cou13, Cra00b, Das91, Dix00, Dre07, ET20, Erm89, Föl91, Gil62, Goo03b, HH02, Har02a, HH05, HH06, How50, Hym12, Jun58, Jun82, Kur02, Ler73, Liv13, Log96, Lui11, MH09, Man80, Mon58, Rob12, Sch12d, Str11b, Sza92a, Sza92b, Tay49, Top07a, Wol95, WB00, Bar05f, GR63, GR65]. Scientometric [KKJ99, UKKP04]. scienza [Val06]. scienzato [GA02]. Scienze [Rec07]. scienziati [Cas14]. scienziato [BC94, Inf63, Inf68, Inf74, Inf92]. Scope [Sho21, Vog20, Vog21, Bus12b, Bus12a, Mil12b, Myr21b, Myr21c, She16b, Sud11]. scoperta [Ano82d, Ano01c]. scoperte [Seg76, Seg83]. scrapbook [Ros02b]. Scribe [Jon89, New56, New60]. Scribner [Ano50a]. Scripta [Ree16a]. SDI [Her00, vB02]. Se [Ein23h]. sea [Bre08]. Seaborg [SK94]. Sean [Fre08, Lit09]. Search [Hal18, Hol68b, Hol70, Ivr99, Kae13a, Nun22, RS99, RS07a, Smo19b, Sta80b, TVL^{+}09a, Wan00, Ada17, Bod03, Col04, Gri84b, Gri84a, Gri95b, Gri95a, Gri10, Hal04a, Hal09, Har13b, Her77, Hol88b, Jac08, Kil94, Pan04, Par86, Sau19, SK09, Sei03a, Vil06, Vol02, Woi06, Zac99b, ’t 97]. searcher [MB21]. Searching [Lig18]. Sears [Nag50]. season [Ano14c]. seated [Ano51b]. Seawater [Ano18d]. Sećanje [Mar95b]. secchi [CM02a]. Sechzig [For70b, Win69]. sechzigstem [WvLuAE18]. Second [Din59, Hil07c, JM17, Kra79, OJCP^{+}47, Pin00a, Ber81, Blu22, Col68, CM86, CM97, Dro15, Jar08a, Ein02a, Ein05w, Har03a, Stu18, Wyk76]. second-biggest [Blu22]. seconda [Ber81]. secondary [Per91]. Seconds [Cle17]. Secrecy [Cra90, Abr98]. Secret [Alt13, Ano45a, Bad83b, Bad96, Bag10, BM77, Dee08, FS95, Gru05b, HM90, Kae45, Kae13b, Kap20, Kur12, Mah17, Mar96a, Mor79a, Pos11, Wel00, Ada17, AK97, Amr59, Ber96c, Cob07, Con02, Con05, CM02b, DMPS81, Dei95, Dup06, Eva53, For82a, For84a, For86, Ful84b, Ful84a, Gal93, Gon12, Gou48, Han88, Jac11, Jer02, Kha12, Kno90, Mel00, Pow93, Pow94b, Pow00, Pur63, Ree05b, Tho17, Wea76, WG17, Wel99, MTVB88, Hoh13]. Secretary [Ein24f]. Secretary-General [Ein24f]. Secreto [Mar96a, Pye86]. Secrets [Ber06c, Bow87a, Bru09, KaS15, Kni12a, Kni12b, Mos16a, PP05, Ano06d, Bar02b, HM97, Lau98, Mar11d, Tur03]. secure [Sar08]. Security [BBD^{+}11, Dot82, Ein50d, Ken54, May95, BIK83, Bus12c, DMPS81, Pol02, Tur03, Vig83a, Vig83b]. SEDs [Bed01b]. See [Ano45a, Ein52e, Fle18, Rei51, See61, Ein34n, Ein40h, EH41, Ein49e, Ein58, Ein89c, Ein00d, Ein06c, Ein09a, Ein15a, Eis37, KRZ^{+}05, Ano35]. Seeing [Arp98, Ree06]. Seek [Res50]. seekers [Whe79c]. Seeking [Dur08]. Seeks [Ano29b, Ano50c, Kae13b, Pin00a]. seem [Ein19c]. Seemed [Ano19b]. Seen [Ano29b, Mic78, Sta99a, Ano05d, Fri12, Gar93, Leg04, RKW05, Roe93]. Sees [Ber85, Jam30]. segrete [HC94c]. seige [Joh11a]. Sein [For80a, Fra49b, Inf56b, Jor69, Bri05b, Enk18, EGO97, Her94a, Her04, Her76c, How90b, Inf53a, Inf57b, Isa85, Mau96, WE24, WE87]. seine [Ano21e, Ano21i, APWB95, Bre08, EH10b, Ein15b, Ein05c, EH05c, For80a, Fra49b, Hof97b, MGG^{+}95, Mos21, Mos22, Mos16b, Sta95c]. seinem [Ein05b, Ein05v, Hes15b, KRW05]. seinen [Kru94, Sch88b]. seiner [Fla15, Fre19, Müh11, RKW05, vS21a, vS21b]. seines [NHP^{+}79]. seit [Hah81]. Seiten [BS92, Eis00, Jüt21, Ros86, Bäu83, RK94]. Seiya [Dar04b]. Seizing [Ber93]. Sekunden [Cle17]. Selbstbiographie [Hah62]. selbstvergessene [Fis02]. Selbstzeugnissen [Ber74d, Her76b, Her84, Her01b]. Selected [Bet97, Bič00, CS79a, For80a, RR82, Wig79, Yan13b, Dre07, FCKT11, Gre04a, Gun11, HH96, Lew79, Mah01, Ram79a, Sch00c, WB01, WS78, Wir08, Bor63a, Egg02]. selection [Hah81, Sch49a, Sch51a, Sch55, Sch79a, Sch83a]. Selections [Cra90, Haw02d, SB82]. Selective [GT38, Har69]. Selety [Ein22d, Ein05h]. Seletyschen [Ein22d, Ein05h]. Self [Ein59e, KS65, PE49, Roh60, Sla72, Bal75b, vBT94, Her76b, Her84, Her01b].

Self-Consistent [KS65, Sla72]. Self-Energies [PE49]. Self-Energy [Roh60]. self-interaction [vBT94]. self-made [Bal75b]. Self-portrait [Ein59e]. sell [Ano96b]. Sellars [Lan00]. Sells [Bar18b]. semantic [DCL23]. Semantics [How96a]. Semester [Ein10b]. Semi [EM33a, EM34, Kun76, EM32b, EM33b]. Semi-centenary [Kun76]. Semi-Vektoren [EM33a, EM34, EM32b, EM33b]. semicentennial [Ein29a]. Semiclassical [Kel85, Par79]. Semiconducting [Hee01]. Semiconductor [ARK^{+}99]. Seminar [DDDR06]. Semivectors [EM33b, EM33a, vD04, EM34, EM32b]. Senate [Ein48e]. senci [T– D69, T– D95]. Send [Kao10]. sens [Pat85]. Sensation [Kra81, Ros86]. Sensationalized [Hal18]. Sense [Fie07, W.25, Bon64, Den24, Ein27b, Gio16, Lee06]. Sensitive [Ano17c]. Sensors [Ano17c]. sent [Rob01]. sentiment [AKD23, YCL20]. senza [BF16]. séparabilité [Com95]. Separability [How85, How89, Com95]. Separable [Bha86]. separate [Car89]. Separated [NBDG40, Sch35b, Sch36]. Separation [Hel15]. Sept [De 13, Ano05o, Kle05]. September [Ano93c, BF02, Ein79e, Ein85e, HW55, JP08, Meh73b, Pap15, Ren05e, Bad12, Sne82, Ano00g]. Serb [Sre29]. Serber [Her21]. Serbian [Mil10a, Mil12a, Inf79a, Mag95, Mar95b, Pop98, Pop05, Sre29, T– D69, T– D95]. serendipity [Mey07]. Sergay [Dix00]. Sergei [Hoh13]. Series [Bäu83, BDR93, Hay91, Myr21c, Ano06k, B^{+}06, BOS14, KWMS05, Whi67, Yag85]. Serna [Cob07]. Service [EGO97, Roe93]. ses [Ano23a, Cun22a, Edd21, Ein10c, Ein10d]. Sesión [Ein29h]. sesquicentennial [Ber84b]. Session [Ano80b, Dir80a, HLZJ91]. sessions [SSH13]. Set [ESA15, Abb02, Dub05, EM31b, Wel14, Wel76, Wel88, Ein05k]. sets [JDB^{+}97]. settembre [BF02]. setting [Hom05]. setzt [JDB^{+}97]. Seven [RMI21b, Sch87a, SA85, SA87, SA95, Spi05, BS86, BB13, Kle05, Pal98, RMI21a, Wic95, De 13]. seven-times [Kle05]. seventh [Sch07a, Pap15, Rav97]. Seventieth [Ein32e, Mil49a, Som49c]. Seventy [Wig49]. Several [Ber94, Con70, Kah69]. Sex [Bro21, PP05, Pic05, Har07]. Sexl [Pye80]. sezione [Tug98]. shōtai [Gol98b]. shackles [Ost15]. Shackling [Moo96a]. shade [Chi94]. Shadow [Ein49b, Fle18, Hol07, Ken19, Nic20, Sim22, Tou05, BL96, Col04, Dre93, Gru05b, Kra81, LML88, Pop03, Sch00d, Smi04, T– D69, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D95, T– D07, T– D12, WB01]. Shadowitz [Sha92]. Shadows [Bur83, Goo90b, Ste89, Chi08, LS92, LS94]. shakes [Mai99]. Shakespeare [Gon05]. Shaky [Fin12, Fin86b, Fin97]. Shall [Ein46e]. Shannon [Kov20, Per05a]. shape [Dob06, Ree00, Sch07a, YN10]. Shaped [Gib19, Zoe09b, Die05, McC01a, Moo02, Par04, Zoe09a]. shapes [PS22]. Shaping [Cec01, Ein27e, Hol82, Har13b]. Shapiro [Pul11]. Share [Ano33c, Lev01a, Rob50]. Sharon [Wat00]. sharpened [Bel05]. sheets [Sau19]. Shellard [MGG^{+}95]. Shelley [Smo07a]. Shells [Goe49]. Sherlock [Bru97a, Bru97b, Bru98]. Shest [Lan67a]. Shield [Ano46b, HD69, HAD90]. Shift [Chr20, Dic60, Edd26, Lai20, Lod20a, Lod20b, Lod25, EG80a, Fri64, Nac24, Sne82, Tre75a]. Shifting [Nye19]. Shifts [Bar49]. Shimony [CHS97]. Shining [Gut15, Moo17]. Shirakawa [Low01]. shite [Sch27]. shock [Han92]. Shockwave [Wal05b, Wal05c, Wal06c, Wal06d]. shoeboxes [Mar00]. Shook [Col09, Gam85, Cha05a, Gam66a, Gam66b, Gam72, Haw11, SA85, SA87, SA95, Spi05]. Short [Ano33c, Bar05d, Bar05e, Bry09, Clo04, Clo09, ET87, Fey00, Haw02a, Sta08b, Tay49, Vit15, EE82]. Short-Story [Ano33c]. Should [Law07, vC05b, CF09, Jak13, Joh14, Won05]. Shoulders [Ren06, Ano47a, Haw02c]. show [Boz14, Dob06, FSE31]. Showed [Bro21]. showing [See25]. Shows [Tra16, Lui11]. Shurkin [Ree17d]. shut [Gla07]. Shvinger [Dai67]. si [EI57]. Sic [FG20]. sich [Ein18j, Ein20g, For80a, HW78, JDB^{+}97]. sicher [KRZ^{+}05]. Sicht [Gru97]. Side [Ber79b, DH79, For80a, Moy81, Smi96, Cha05b, DH81, DHR13, RS11a, Sin10a, Wig79, Ram79c]. Sidelights [Ein22l, Ein83, Ken25, Mur25, EWJ^{+}25]. sides [Top07a]. Sidney [L.89]. siebzigsten [Ein32e]. Siècle [Hei82, Ano15a, Thr17]. siècles [Der15]. Siegbahn [Kac96, Kat52, Tru96]. Siege [Fre19]. Siegfried [Boz14, Ree17c, Boz14]. sieht [Ein52e]. Sieroka [Pes13]. Sigmund [Ano57b, Ein85c, Rol10]. Signal [CW12]. Signaling [Kov20]. signalling [HV88]. Signals [Ano16d]. signatures [BR99]. Signed [Whi55a]. Signers [Whi55a]. significado [Ein85f]. Significance [Byr80a, Dic64, Fin81, Fre79a, Has07b, Kan92b, KUW82, Rei49, Wen49, Bel96b, Duc11, Ein41f, Hah48, Wel93]. significato [Ano82d, Ano01c]. sigue [Yun15]. silence [Gol70b]. Silent [RB05]. Silicon [Ber05]. Silko [Rai14]. Silvan [Tho11b]. Sime [Kri99]. similar [Ein12d]. similarity [DCL23]. Simili [Per10]. Similiarity [vKL49]. Simões [Ren21]. Simon [Gor07a, Kae38, Sch07c, Weg08, Bra29, dHvdB49]. Simple [Ano22a, Ein21e, Thi21, BBD^{+}11, Cun22a, Egd14, NP88, Sta07d, Tho15, Wil83, Ein16e]. Simplest [Nor00]. Simplicity [Elk82, Kae38, Nor00, She88]. Simplification [Sla51]. Simplified [L.89, Lam30, P.30, Har89, Sar96, Tal32b]. simplifies [CT24]. Simplifying [vCK17]. Simply [Kor94, Wol03, Vaa17]. Simulation [ZDM08]. Simulations [BBCPL09, BPL05]. simultaneità [Tug98]. Simultanéité [Car22a, Ber22, Car23]. Simultaneity [Jan08, Mar04c, Ber22, Ber65]. SIN [MTVB88, Gio14, MTVB88]. Since [Ano09g, Fre08, Giu05, Sci80, EVB^{+}72, GMOP12, Meh75c, Rot92, Rot97, Tra16, Hah81]. Sinfonia [Ein34m]. Singapore [Gil12, Ree15d, She15, Vog20, Mas02]. Single [Hür22, Cha11, FTZ^{+}15, TEM^{+}89, ZPSK10]. Singular [Laz15, Med94, Bal72]. Singularities [EE99, Ein41a, Mat91]. Singularity [FH10]. Sinne [Ein27b]. Sir [Adl03a, Bro97, JW06, SR04]. sista [RE90]. sistemi [Fer24]. Sister [PA21, Rob01]. Site [Lau45, Owe12, FG00, Ger92, Ger97, Ger02, Ger07, Sza84]. sites [Kel12]. Sitter [Ein18i, Gui09, Sta20, Ein18i, Gui09, Gui18, KK75, Ker89c, OONM15, OOM21, RP09]. situation [Rol47]. Sitzung [Hof08b]. Sitzungsberichte [Sim06]. Six [Fey95, Fey97, Fey11, Kas21a, GRS16, Lan65, Lan67a, Ree00]. Sixtieth [Ein39f, Opp39, Opp79c]. Sixty [FR13, Gar08, Mil90]. Sixty-two [Mil90]. Size [dMdA05, EMH96, Fra37, PR11, SEM03, Sil30, Swi91]. Sizes [Ein86c]. skaparen [Gar41]. skeptic [Ein10]. Skeptical [Kol77]. skeptics [Ano97e]. sketch [Bro19, Bro20b, Ein56a, Gha82, Sch35c, vH97]. Sketches [SSH13, Sch92a]. skhvatki [Rol19]. Skis [Hau54, Hau74]. Skizze [Ein56a]. skládání [Žáč12]. Skłodowska [CGT11, Cur38f]. skulle [Jak13]. Skurrile [Arr06]. Sky [Har79b, Kee05]. Slant [EPW13]. slash [RSPS49]. Slater [Sch90]. slavnog [Sre29]. Sleep [Ano55d]. slept [Bre97]. Slice [Jan03, Ano04d]. slip [Bre08]. slit [Kal92, Kon15]. Slovenian [EI62c]. Slow [Gol72, HS39b]. slunecních [Nac24]. Small [HS05, Jon89, MNT^{+}00, Mac85, SMM^{+}60, SBSW60, Ein08a, Fay10, New56, New60, PSC^{+}97]. Small-Project [MNT^{+}00]. smaller [GRS16]. Smallest [GK05, Fis10, Fis12, Hal09]. smart [NG07]. smasher [Ano51b]. Smichov [Gut75, Gut75]. Smith [Pye89]. Smithsonian [BL96, Koh95, Nob95]. Smoke [Lis16]. Smoluchowski [Ein17f, Tes69, Ein17f, Tes69]. Smuggler [FC07]. Smyth [Kae45]. Snake [Wei13f]. Sneaking [Ghi05]. Sneered [Sei03b]. Snow [Duf81, Hen82a, Hen82b, Kle82a]. snub [TA17]. Snyder [Rav97]. so-called [Ein26c, See25]. Sobral [de 03]. Sobre [Ein85f, Ben01, Cam88, PPC05]. Social [Con70, Ein46f, Ein50n, Ein50c, Hin49, HOO88, Lan00, Mer38, Por81, Tre79b, Wil96, Zim82, Aer99b, Ano81c, Bad83c, Bad83d, EE05, Ein50e, Ein52b, Feu71a, Feu71b, Gro56, Olw04, Owe12, SR04, Tal82]. sociales [Ein50e, Ein52b]. Socialism [Ein48o, Ein89b, Ein92, Ein93a, Ein94a, Ein09d, Ein98, Ein00a, Ein02, Ein04, HSW09, Hen96, HW04, RW94, Wal89, Wal92c]. Societat [Gli84]. Société [LLH^{+}80]. Society [Ano26b, B^{+}06, BPS87, Ein31a, Ein50k, GGK02, GGK13, Gib19, Goo85, JP89, Kra82b, Sei84, Vis94, B^{+}01, BCS82, Ein59f, Hei75b, Hen82b, Ano22a, Ein18b, Ein50n, HCD^{+}78, HSW09, HW04, Hof08b, HW12, Pin00a, Sch88b, vBL96b, Lin56]. Sociological [Hen86a]. sociologie [Lar12]. Sociology [UM86, UM92, Zuc67, Lar12, Mac93]. Socken [Ank15]. socks [Ank15]. Socrate [PB72]. Socrates [PB72]. Socratic [EM32c, Ein77a]. Soddy [Cru79, Fre79a, Scl89, Tre79b]. Sofie [Van22]. softcover [Vog21]. Software [ARK^{+}99, GRE^{+}01, MNT^{+}00, Pfl99]. sogenannten [Ein26c]. sohbetler [Tip12]. Sohlman [Kat52]. Sohn [Lin60]. Solar [Cam18, Ein45d, Eis22, Kas21a, Nor72, Nun22, Ree06, Sim22, Bau20, Dav03, Nac24, Pan02, Spo02]. soldier [Roe93]. Soldner [Eis23, Jak78, Sau21, ST23, Wil88]. Solid [Ein05q, Ein05t, Ein11g]. solides [Pia31]. Solids [Ein05-35, HEB^{+}80, Sla63, Sla54, Sla57]. soliton [SLLS09]. soliton/vortex [SLLS09]. solitons [BSSP^{+}08]. solitude [Ein59a]. Solovine [Spe62, ES56, Ein05-47, Sol56]. Solution [Ein18d, Jam30, Mar08, Dat38, Ein18i, Ein23c, Hou16, Rus98]. Solutions [Bič00, Ein05-39, Göd49a, Bre89, Die88, Ein02b, EM30, EP43, Grø82, KS80, Lom05, Mac13, RA11, SKM^{+}03]. Solvay [BV09, Bar93, Meh75c, BL11, CW12, LdB12a, LdB12b, MW99, Str11a]. Solve [Ano12c, Max93a, Dor78]. solved [Bru97a, Bru97b, Dor78]. solves [Bru98]. Solving [Pul11, Sei94, Tau49, BS10, FGL^{+}06, Gri95b, Gri95a]. som [EI39c, Lun74]. sombra [T– D92]. Some [Aga06, Ano21m, Bin06, Bub79, Cha75, Dav09, Ein36g, Fer82b, Hec16, Hof82b, vI11b, Kuz79, L.89, Leg01, Leg03, Lid67, Löw98, Mor96, Nie89, PP95, Ree17c, Rol10, Sch22b, Str49, Uhl80, Woo80b, Aga07, Ano26a, CW02, Jam06, Kle65a, LR14, Ryc92, San77, San15, Wal02a, dB85, ES13, Ein16d, ES05a, Gol84b, Ano84c]. Something [Hah42]. sommaire [Pia31]. Sommerfeld [Ein17h, For70b, Hen86a, Kle85, Som05a, Som05b, Win69, Ein17h, Ein05-46, Kel58, Kle85, MR82d, MR82e, Som05a]. Son [Lau55b, Spe95, Fra91, Mee03, Ros02a, Seg95]. songs [Lev16]. Sonnenenergie [MGG^{+}95]. sonnets [Joh11a]. sono [Arg09]. Sons [Ano50a, RB05]. Soobrazheniia [Ein40g]. sora [CM02a]. Sörbom [Goo85, Sei84]. Sorry [SL05a]. Sought [Lau55a, Lau55b]. Soul [Alt13, Bra14a, Ano14b, Fer07a, For09, Lin07, Lin08, Seg07b]. soumises [Ein14j]. Sound [Ein20l]. Sounds [Ste98, Bar00c, Bel98]. Source [JNR^{+}07b, Kae39, Har13c, HTS^{+}88, JRT^{+}91]. Sources [Ein20i, How96b, vdW67, vdW68, vdW07, CH80, Han79b, Hen96, Cli09b]. sous [LdB12a, LdB12b]. Sousa [Ren21]. South [Wac06, Abr09, Ein22]. Southwestern [Mil08]. souvenirs [Kol56, Ein55b]. Soviet [Dix00, Wei03, Alp65, Alp85, Alp94, ČA99, Dix00, Ein48m, Gon96a, Gon96d, HG92, Hol81a, Hol94, Mar79, Mik65, MM66, Res50, Rho93, Sch54, SSSS95, VFIS48, Vuc01, WG17, Wes04]. sovremennoi [Kol40]. sowjetische [HG92]. Sowjetphilosophie [MM66]. Space [Ano21d, Ano22a, Ano15c, Ano16d, Bia16b, BLR15, Bro20a, Cal76, Čap76, Car07a, CSM06, DiS06, Dor78, Dys15, Edd20b, Edd21, Edd29b, Edd59, Edd66, Edd87, Ein18h, Ein19a, Ein22j, Ein23f, Ein29i, Ein30e, Ein34l, Ein47d, Ein54f, Ein01c, Foc55, Fok49, Fri14, Gib91a, Goo02, Gre04b, Gre05a, Gru57, Grü73c, How97b, Hug99, Jam30, Jan03, JW06, Kar05, Kas21a, Ken03, Lan70, Mil94, Pau94d, Pep04c, Pia30, Rei27, Rei28a, Rei28b, RF57, RK77, Rei06, Row17, Rüg88, SB20, Sch63, Sch05d, SB06a, Tay21, Wal77a, Wal92a, Wei12w, Wey22a, Wey22b, Wey50, WL12, dS19, Ano19a, Ano04a, Bar99, Bar00c, Bro05, Car02, Cle96, EGS77, Ein30r, EM31b, ES45, ES46a, Eis02, EW88, EW00, Fey66, Fey97, Fey11, Foc54, Fok59, Fok64a, Fok64b]. space [Fok65, Fri83, GW06, Gim15, Gre15a, HE08, HFN06, Inf99, Kak04, Kak05, Kin95, Le 03, Lev16, Mau90, Mil01, Mus15b, Nor89, PR86, Pot11, PP03, Pul06, Rei24, Rei58, Rei11b, Ren97b, Sal05, Sau06, Sch86, Sch02c, See25, Sti13c, Stu14, Stu23, Tre71, TN03, Vig94, Wei04, YN10, Ano25a, Ein30h, Ein30b, Pri99, Sch17, Sch19, Sch20, Sch22a, SB06b, Wey18, Ein22c, Jam70, dMdA05]. space-time [Sau06]. spacecraft [BIT03]. Spaces [Cob40, Lan49, Syn65, Hal97]. Spacetime [Car04b, How99a, Pep04b, PS04, Rei05, TW66, TW92, Cho16a, DiS92, Dor95, Dys16, Ein09c, Fv85, FH02, Haw02b, LSS16, Oha76, OR94, Sil30, Skl85, Wal94a, Whe90, Bar94b]. Spacetime-philosophy [Bar94b]. Spain [Boy88, Hol91a, Ros18, Gli88, GPSR01]. Spaltung [Hah42, HS44, EM33a]. Spanish [Ano25b, Ano25c, Bal99, Ben01, Bla86, Bor22b, B^{+}99, Cam88, De 12, Dür82, Ein21q, Ein26a, Ein31m, Ein31n, EI65a, EI68, EI84, Ein94b, Esp34, Esp55, Fer05a, Gar40, GPSR01, HC96c, Inf61c, Inf83, Mar96a, Moo96b, Nav94, Nir25, Paz10, Phi74, PPC05, Rus43, Sch96b, T– D92, W^{+}87, WdC02, Yun15]. Spannungsfeld [B^{+}01]. sparked [Sam10]. sparking [HR17]. späten [Ein52a, For80a, Hei53a]. spatial [Pul06]. Spatially [NBDD^{+}12]. spazio [Pot11]. speak [Fre81a]. Speakable [Bel87, Bel04]. Speaking [Byr47, Byr74, Cha88]. Speaks [Inf50b, Ano48b, LS08]. Special [Ano95a, Ano99f, Ano00f, Ano04g, Ano05-43, Ano09c, Bai07b, Ber94, Bri05d, Cho05a, Cus82, Din49, Ein20k, Ein21n, Ein21j, Ein54i, Ein76a, Ein82f, Ein99, Ein13, Ein15b, Fer81, Fre68, GGK13, Gin82, Gou13, Häg46, Hel84, Jan14a, Kai81, MSS22, Mel82, Mil79, MW11, Pal83a, Pal83b, PS03, Ram83, RWY97, Rin60, Rin66, Sch27, Sha68, Sha88, She16b, Sti98, SSSS95, Tau49, Tay21, Tre21, Wei12k, Wei12w, Whi83b, Acu14, Ano01a, BCR93, Blu05, BGLR17, CL19, COS05, Dar21, Dro15, Ein23h, EM34, Ein60d, EBRC93, Ein95b, Ein07f, Kra21, Kra25, Les74, Rin69, Rin77, Rin82, Rin91, Rin01, Rin06, Sch93c, Sch75, Sta95b, TW92, Voi86, Wes80, Abi91, Ano21d, Ano07c, Ano07l, Bai92, Bai01, Ban06, Boh65a, Boh79, Boh89b, Boh96, Boh06, Car02]. Special [Cas86, Cho05b, Cul81, Dor78, Dra12, Eĭn86b, Ein17g, Ein18n, Ein20o, Ein21p, Ein21r, Ein21q, Ein22a, Ein23h, Ein31j, Ein31k, Ein36f, Ein54g, Ein61a, Ein61b, Ein77b, Ein78, Ein88b, Ein95d, Ein96a, Ein96d, Ein00c, Ein01b, EK03, ELGC05, Ein06b, ELGC06, Giu05, Gol68, Gol70a, Gol76, Gol83a, Grü62, Hag08, Hen90, Hir74, Hol88c, Jac87, Jan95, KKC06, Lan72, LEM^{+}23, LEMW52, Lud13, Max93b, McC84, Mil75, Mil81, Mil82, Mil98, NS05, Nor14, Pal83b, Poo90, Pye77, Rob49b, SS04, Sig92, Sin00a, Sin00b, Sin05, Ste04, Ste98, Wei15a, Woo03, dH13]. Special-Relativistic [MW11, Wes80]. Speciale [Ein21p, Bri05d]. specialist [Myr21b, Myr21c, She16b, Vog20]. specialists [Bus12b, Myr21b, Myr21c, She16b]. speciálni [Ein23h]. Species [Nag50]. Specific [De 10, Ein75, Ein05q, Kle65b, Kox95, MBD39, Ein07e, Ein11g, Ein12c, Ein13g, Ein14o, Ein05p]. Specification [LC07]. spécifiques [Ein12c]. Specimen [MKBF30a, MKBF30b]. Specker [Gar72]. spectra [Cas07, Khr08, Pau25]. Spectral [Chr20, Cro20b, Lod20a, Lod20b, Mou20, Nac24]. spectrometry [Nie89]. spectroscopic [CSS^{+}21]. Spectrum [DR20, Lai20, Pla00b, Pla00a, Pla01, Ein23a]. speculation [Bac00, Mag03, Ric06b, Bac00]. Speculations [Bro23, MO95]. Speech [Ein16g, Ein20b, Ano13a, Ein31m, Ein31n, EP06, Ein30i]. Speeches [Ano81d, Jos95b, SS98c, Ano64, Ano65b, Ano72, ACT23, Buc93, Don70, Ein31c, FM92, Lun92, Nob72, Nob64b, Nob66a, Nob66b, Per97, SS98a, Sea72]. Speed [Lau13, AAA^{+}09, Ano02a, Ein27c, Ken07, Mag03, Ein12e, Ein05x]. speeds [Žáč12]. Spektralgleichung [Pla00a]. spektrálních [Nac24]. Spektren [Pau25]. spel [vE95]. Spell [Bub73]. Spencer [Sch87a]. Spengler [Hol91b]. spetsyele [Gru07, Sch27]. Speziali [Kle73]. Spezialinventar [For80a]. Spezielle [Ein99, Ein17g, Ein18n, Ein20o, Ein54i, Ein88b]. speziellen [Cun22a, Häg46, Hen90, Kra21, Kra25, Pye93, Voi86]. spezifischen [Ein07e, Ein11g, Ein13g, Ein14o, Ein05q, Ein05p]. sphere [Car10]. Spheres [Hug54, OW54]. spherical [Edd30a, Ein39d]. spherically [Lom05]. Spiegel [Hei75b, RKW05]. Spielen [Ein19j]. Spies [Lan94, Wil16b]. Spin [LJGS11, SISK^{+}98, DPS97, Grü08, MS82, Sau07c]. spin-statistics [DPS97]. Spinor [EM33a, SHL^{+}06]. Spinor-Gleichungen [EM33a]. Spinoren [EM32b]. Spinors [PR86, EM32b]. Spinoza [Rei41, Rob48, Gal79, GN86, Kay46, Kay68, Kuz87b, Pat86]. Spins [Big05]. spintronics [Fer08]. spirals [Ste07]. Spirit [Gib19, Ham85, Koe95, Ban12, Rol71, Tip10, vEA02]. spirits [Dav83]. Spiritual [Rob48, Kay46, Kay68, Mac08, Pet21, Pet23]. splendor [Gre15a]. Spline [Sch17b]. Split [Ano39g, Ano41, Cat04]. Splits [Kae39]. Splitting [EM33a]. Spoils [Bow87a]. Spoke [Ōe95a]. spokesman [Hei86]. sponsor [Ano87b]. Sponsored [Cha74]. Spontaneous [FP40a, Nam09, Nam10, SHL^{+}06, WNS^{+}08]. Spookiest [Acz01, Acz03]. Spookiness [Wat97b]. Spooks [PW17, Bal01b]. Spooky [Ano09d, Fra14, Mus15b, Ano15j, Kon15]. spot [Cho16b]. spotted [Cho16a]. Spouse [WS89]. Sprache [Hen11c, Rec09, Rec10]. sprängämnesindustrin [Lun74]. Spring [JDB^{+}97, Lan65, Lan67a]. Springer [Eis00, Hen86a, Hen11c, Hoh13, Klo63, Sho21]. Springer-Verlag [Hen86a, Klo63]. Springs [ADS83]. spukhaften [Kon15]. Spuren [Hof06]. spurned [Sch05e, Sch12f]. spy [AK97, Clo15, Hel11, Rob01, SH01, Wil87]. Spying [Ric06a]. spymaster [SSSS95]. square [AHH05]. Squeezing [EGW^{+}08]. srozumitelný [Ein23h]. Srpkinje [Sre29]. srpska [Mag95]. SRT [Mil92]. SS [Kai81]. SShA [Smy46]. St. [Wal61]. stē [EI78a]. Staatsbibliothek [B^{+}79]. Stability [GB88, Roh60, BIK83, Pit16]. Stachel [Cah90b, Gly90, Har91, Hen88b, Mar11c, Pye89, She16a, ACH^{+}03, Bal91, Cah90a, Gol87, Hen88a, Ker89b, Loe88a, Ram90, RDS03, Smi89b, Smi89a, Str90, Yan87]. Stadtmuseum [Bad12]. stage [CL07, Kai94, LHDH21]. stages [Hom06]. Stähl [Kat52]. staircase [Fis10, Fis12]. stała [Gol98a]. Staley [Mar10, Hen09, Set10b]. Stalin [Gor09, Has05, Hol94]. stalked [HEWW15]. Stallion [Smi86]. stamps [Fei01]. Stand [Ein29j, Ein30n, Ein32a, Kra23a, Kra23b, Rit09b]. Standard [Boe10b, Rig05a, HBRD97]. Stande [Ein09d, Ein13g, Ein13f, Ein14o, RE09]. Standing [Shi17, WS88, WS08]. Standpoint [Nor07b, Wen49, Wie21]. Standpunkt [EF13, EF05, Wie21]. Standpunkte [Ein19e]. Stands [Koj08, Lan08, Ano09c, RS07f]. Stanford [Smi02, vB02]. Stanley [Ham97, McC91]. Star [Ein21e, Kae13b, AAA^{+}18, Ein36e, FSE31, Lem09, Shi75, Sta07d, Cor12, New89, Vol02]. Starburst [CLD^{+}03]. starih [Mar95b]. Stark [Ein12a, Ein05d, Her66a, Ein12a, Ein05d, Her66a]. starlight [Lam21]. Starling [Hen03]. starre [vI10b, vI11a]. Stars [Bet39, DM96, Gut15, Hoy54, Kir02, Lig18, RS03a, Ao04, Bet68b, Bet68a, ES45, Gra94b, Lin15, Mil05a, Mil05b, Str05b, Wal98, Ano19b, ES46a]. Start [Law07]. started [EI40a]. Starting [Mar15]. startling [Gri84b, Ler91]. starts [GB11]. State [BOPS99, BS81, Ein20b, Ein32a, Ein05-29, Ein07c, Kra23a, Kra23b, McC68, Pla20c, Tur66, Abr98, Abr09, BRH^{+}09, Ein95, Ein09d, Ein10a, Ein12c, Ein13g, Ein13f, Ein14o, Ein30n, EK53, Fer07b, NBDD^{+}12, OMNM14, Rit09b, RE09, Rol10, SISK^{+}98, B^{+}79, Pin00a]. state-dependent [BRH^{+}09]. Statement [Eme48, Ein39f, Ein41f, Lor10a, Swa30, Lor20, Lor10b, Lor10c]. Statements [Ano45d, OJCP^{+}47, dB85]. States [AKBAE21, Fre08, HS10, Har79d, HA62, HD69, HDHA89, NW49, Smy45b, Smy45c, Smy45d, Smy45e, Smy76a, Smy78, Smy89, WH99, Ano00g, De 92, Far13b, Gib96, Hol84, JDB^{+}97, Moo10, RN16, Szi45, WK12, Zuc77, Zuc96, Zuc18, Nad79, Sei79]. States/UK [JDB^{+}97]. Statesmanship [Ein44c]. Static [Ein22b, HBN00, Ros49, Wei12p, DS01, EM30, Ein05-44, Grø82, Nus14, OONM17]. Statics [Ein05x, Ein12e]. Statik [Ein12e, Ein05x]. Stationary [Bre89, Ein39d, EP43]. Statische [Ein22b, EM30]. statischen [Ein12j, Ein05-44]. statistica [Ber81, BES86]. Statistical [Bor49a, Byr80b, Ein12j, Ein14j, Ein14h, Ein16i, Ein86c, Ein87b, EBD^{+}89, Eza79a, Eza79b, Kub86, Mon19, Nav98, Por81, Sch89, Sla72, Abi00b, BB55, Bor55b, Bor63b, EE59, EE90, Eĭn85b, Eĭn86a, GBE95, Gea90, Kle82b, Kox14, Meh75a, Nav91, Pau93, PS10, Ren97a, Uff06, EH10a, EH05b]. statistical-mechanical [Nav91]. Statistics [Bel23, DR08, Kap40, Mon19, SP59, Bad09, Ber81, BES86, Das12, DPS97, Ein27b, FMM96, Iro04, PS10, Wal06a]. Statistik [Ein27b]. statistique [BB55, EBD^{+}89]. statistiques [Ein14j, Ein16i]. statistische [Bor55b, Bor63b, EH10a, EH05b]. statistischen [Ein14h]. Statue [Bro78a, Dic79b]. Stature [McL99]. Status [Nor72, VW86, Bla10, Buc08, Ein29j, KKKM05, YCL20]. Stauffer [BHJ^{+}00]. staunen [RR17]. Stavis [SE90]. stay [BL14, San77]. Stayley [Del11a]. steady [OMNM14]. steady-state [OMNM14]. Steering [AKBAE21, HR13, YC18, ZXL^{+}19]. steht [Whe68b]. Stein [Duf81, Hen82a, Hen82b, Kle82a, Fol17, Nor03b]. Stellar [CLD^{+}03, Cha49b]. Stellung [Ein40c, Fra49d, Pet21, Pet23, SE79]. Stellungnahmen [Hah81]. Stellungswechsel [Ano79g]. stému [Bič79a]. Step [Ano97c, Ano06j, Jam30, Mil92, Coo83]. step-by-step [Coo83]. Stephen [Haw92, Sus08]. Steps [Mil99, Rec02]. Stern [EE22, Hen00b, EE22]. Sternhaufen [Ein21e]. Sternstunden [Büh98b]. Steven [Dys15, Row17]. Stevenson [Coh54]. Stewart [Nau17, Whi16]. Stiftung [Lor14, Fre27b, Lor14]. stigma [KST99]. Still [Cao07, Yun15, Ano97e, Ano06e, Gio16, Kae48, Pra06]. Stillborn [Tre75b]. Stillhalten [JDB^{+}97]. Stillhalten/Frühjahrstagungen [JDB^{+}97]. Stillman [Col68, Kuh54, Sha54]. Stimson [Fre08, Lit09, Ber93, Mal08]. stimulated [Nat05]. stínu [T– D07]. Stirred [Lau55a]. Stitch [OSWR86]. Stochastic [But07, Hel82b, SP12, Vig94, Vig95]. Stockholm [Bar05f, Fer39, Fer70, HS14, HSD34a, Jüt21, Pla20a, Pla20b, Wil36, Goo03b, HS14, Har02a, Man13, Pla20b, WD06]. Stockpile [Kit00]. Stoddard [B.50, Rei50]. Stokes [XMY^{+}09]. stole [Mos87a, Mos87b]. stolen [Mel00]. Stoltzenberg [Smi05a, Vis95, Joh95]. Stone [Bok15, Čvo06, Kle14a, Cha09, Ger15]. stop [Bun82, KSG63]. Stopped [Ano47b]. Stopping [Phe21]. Storage [LHZA20]. Storia [Tug98, Bat03, Tio09, Val06]. stories [Ano96a, Len17, Mar11a, Roc03, Szi92, Vit15]. storonye [Rol19]. Story [Acz01, Ano26a, Ano33c, Ari14, Bus13a, Dan97b, Fos13, FC07, Gam85, Jon08, Kae45, Laz15, McP97, MP01, MP12, Mlo01, Moo17, Ree20, Sta07b, Ste22, Wei14b, Acz03, Acz10, AK97, Bag11, Che66, Chu60, Chu98b, Chu01, Cla61, CM02b, Dar06a, Don97, ECS19, Fer61, Gam66a, Gam66b, Gam72, GTP22, Gol83b, Gro67a, Gro67b, Gro68, Gro00, Gro62, Gro83, Gun15, HTS83, Hof10, Jon01, Kea18, Kie13, KWB^{+}02, Lau46, Lau72, LB02, Mag03, Man13, McK03b, McP95, Mos68, NN01, Ove87, Pas10, Pur63, Ree15c, RTS^{+}65, RTS^{+}69, RTS^{+}xx, Rin03, Rio87, Rob01, Rob14, Rob50, Sal95, SK94, Sta95c, Sza84, Val06, Wel14, Wor10, SS65]. STR [Grü73b]. Strahlung [Ein07b, Ein07e, Ein05l, Ein05p, Ein09b, Ein09c, Ein16p, Ein17i]. Strahlungs [Ein26d, Ein16m]. Strahlungs-Emission [Ein16m]. Strahlungs-Emissions-prozesses [Ein26d]. Strahlungsfeld [Bos24b, Ein25a, EH05b, EH10a]. Strahlungsgleichgewichts [EE23]. Strahlungsproblems [Rit09b, Ein09d, RE09]. Strahlungstheorie [EH05c, Ein05c, EH10b]. Strahlungstheorie“ [Ein15b]. straight [vS21a, vS21b]. straight-line [vS21a, vS21b]. Strange [Ban51, Dry14, Bru97b, Bru98, Fey06, Liu92, Sty00]. Strangelove [Goo04a, Goo04b, KSG63, Smi07b]. Strangeness [Ano84c, Fer82b, Gol84b, Woo80b]. Strangest [Far09, Ove05]. Strassmann [BH01, CSW97a, BH01, CSW97a, Kra81, Ros86]. strategic [Bla02, BP^{+}01]. strategies [Bes93, Par99]. Strategy [RS07a, Ros79b, Pit16]. Stravinsky [Gar93]. stream [Ein16e]. street [Pát79a, Con02, FW07, Whe79b]. Streit [Beh05a]. strenge [EM30]. Strengthening [Hel82a]. stresses [Dar19]. strict [EM30]. String [Ber06b, Smo07b, Sus06, YN11, vD21, Smo06, Woi06, YN10]. Strings [MGG^{+}95, Ran05, Sch91, SS04]. stroenie [Lan67a]. Strong [Ano21a, CSD^{+}07, Pei71, Sch88c, JRT^{+}91]. stroomen [EdH15c]. Structural [CSS^{+}21, WCL^{+}00, XMY^{+}09]. Structure [Ano26a, Boh23, Boh65b, Čap76, GHK^{+}06, GHK^{+}07, GRE^{+}01, How99a, Hug89, Inf49, Rüg88, Sla63, Ano23c, Bai33c, Boh85b, Bro80, Bro05, Buc08, BRS03, Car02, Ein19j, HE08, JN14, KP55, Koh99, LCL^{+}90, Myr21a, Nug88, Pau25, SW94, Sla53, Sla54, Sla57, Sla60, Tre71]. Structures [LL82, Aer99a, GB15, Hal04b]. Struggle [Fer07a, For09, Hol99, ER13, Fer14a, Flü74, Hoe94, How94a, Lin07, Lin08, Seg07b]. Struggles [HW07a, Wat00, EG78b, McG93, McG98, Wei12p]. Struggling [BM93]. Stuart [Pul11]. Stubborn [Wei12o]. stubbornly [Haw07]. Stud [Flo06, Kra06]. Student [Gra18, Her06, Pin00a, JDB^{+}97, Ken12, Spe79]. Studenten [Ein46b]. Students [Ano21d, Ein32b, Ein46b, Ein08a, Bag92, Ng11]. Studied [Ano50c]. Studien [vBL96a, vBL96b]. Studierender [JDB^{+}97]. Studies [Ano04b, Bar05f, Coh49, CW74, Coh80, CHS97, Cra02b, EJN93, Gly90, HS51, Hoc93, Ket99, KV11, Row08, SS16, UM86, UM92, Wal94b, Wei94, BV02, Bey49, B^{+}99, CH05, CF80, Cra92, D’A97, EGR11, Har85, Hol78b, Pec04, RR82, RWL23, RK94, Wil89, vBL96b, EK92]. Study [Ano34b, Ano15l, BM09, Ein05-39, Gol68, Hon89, Lau55b, Man86, Nug99, Ren93, Spa64, Arg09, Ben01, Chr22, Edd21, Ein02b, Fis10, Fis12, Har13c, HMW24, HH96, Joh04, LLSH23, Liu92, LR14, Lui11, MW44, Mar51, OP87, See25, TA17, UKKP04, Zah89, Eöt00, Fle33, Reg87, Sha73b]. Studying [Fri89, PL17]. stuff [BR93, HS06, Haw11]. stunt [OONM18]. stupor [Cam88]. Stuttgart [BS92, Hei53a, AH93, MTVB88]. Stuttgarter [MTVB88]. Style [Cas79, Mon19, Cas07]. sua [Inf57a, Inf60, Inf61a, Inf63, Inf68, Inf74, Inf92, Seg91]. suas [Seg87]. Sub [Nun22, Goe17, YCKG11]. sub-discipline [Goe17]. sub-Kelvin [YCKG11]. Sub-Solar [Nun22]. subatomic [Fri39a, Lin15]. subdued [Fei61]. Sube [GRE^{+}01]. Subelectrons [Hol78a]. Subject [Ein14j, Eĭn85b]. Subjected [Ein87b, Eĭn86a]. Submitted [Ein86c]. Subsistence [Ein18b]. Substance [Ano29d, Car24, Cas23, Cas03a, Bar94b]. substance-metaphysics [Bar94b]. Substances [Eöt91, Eöt08, Mau90]. substantial [Ein19j]. Substanz [Bar94b]. Substanz-Metaphysik [Bar94b]. Substanzen [Eöt91]. Subtle [Asi83a, Bat05, Fer82c, For84b, Gil84b, Gol84a, Hei83a, Kle84, Lak85, Lar84b, Lov86, Noe11, Pai82, Pip83, Red84b, Red84a, War84, Coo10, Coo15, Smo07a, Pei83, Pol83]. succeeded [Cha09]. Success [Tre75b, CT15, Jan14b, Koc01]. Successes [Tel78]. Successful [Ken54, Kri19c, Ano18c, CZPR17]. such [Ank15]. Suche [Bod03, Her77, Jac08]. Sudden [Rob10]. sued [Gwy10]. Suffering [Gor12]. Suggestion [Tör81, Wey20, Ein26d]. suggestions [Kah69]. suhteellisuusteoriasta [Hot15]. Suisse [Kol56, Sch05f]. suite [Chr22]. suivi [Edd21, Sto96]. sul [Inf63, Inf68, Inf74, Inf92, Ros80]. sulla [Fer24, Tam14, Ein21p]. sum [Cla91, Won05]. Summary [D^{+}95]. Summer [Pap15, Isr73, Lut05]. Summerfield [Pye89]. Summit [Ren07c]. Sun [BHJ^{+}00, Rho95, DED20, DvdRW37, FSE31, Hen91, HHW99, Sza84, Tho85, vK60, Gut98, Pan02, RB05]. Sundberg [Wig96j]. Sündermann [HEB^{+}80]. sunken [Mar00]. suns [Jun58, Jun82]. Super [Cao07, FP40c, FP40b, FP40d, FR49b]. Super-bomb [FP40c, FP40b, FP40d]. Super-Hero [Cao07]. Superbomb [GB89, Yor89, Dys76]. Supercomputers [Sei94]. Superconductivity [Fos13, Ein22m, Ein05-49, Gin04, Lon37, Sau07a]. superconductors [Abr04c]. Superfluid [Leg04, DSZ12, Sto02]. superfluidity [Gin04]. Superfluids [EG98, Lon54, Lon61]. superfluous [OONM18]. Supergravity [Des79, FvN78]. superhuman [Dor10]. Superluminal [Wei12n, Vig94]. supermassive [Ano18c]. supernatural [Mac08]. Supernovae [Bet88, Per12, Sch12e]. Superposition [Win13a, Win13b]. Supersede [Pro74, Zah73b, Zah73a, Zah76]. Supersonic [LB49]. Superspace [AN79]. superstars [WSG00]. Superstrings [Gre99a, Gre99b, Gre00]. Supersymmetry [AN79, Wei05b]. superweapon [Smi07b]. Supplement [BH87, BH89a, BH93, BSKS04b, BSR^{+}06b, BRS^{+}09b, BIRS12b, BIR^{+}15b, BRI^{+}21b, Duk49, Ein53n, JSI^{+}02b, Kno97, SKJI98b, Ste90, Tho97, FGL^{+}06, Fis02, Ein12f, MB^{+}92]. Supplementary [Ano23a, Ber07a, BSR^{+}06a, BSR^{+}06b, Lac08]. supplies [Hau54]. Support [Ano69b, LHZ^{+}23, VFIS63]. Supporting [Hir97]. Supposedly [Hen13, Kot93]. supraconductibilité [Lon37]. Supraleitung [Ein22m]. Supreme [Cam99]. Supremely [McC68]. Surely [FLH85]. Surface [GR44, Ein05k]. Surfaces [Her49, OSWR86, Eöt00, Ano17b]. Surfaces/Paul [OSWR86]. surmounting [Yos95]. Surpassing [Ano23e]. Surprise [Kir13, You98, Hon96]. surprises [Nor13]. Surprising [McG19, Hon96]. surrender [But54, Has05]. Surrounding [CLD^{+}03, ES45, ES46a]. Surveillance [BLM19]. Survey [Har69, Kni13, Por81, dB49b, CF11, HI79, PCH77, Sci80]. Survival [Eme48, Bun88, Bun90]. survive [GS50]. survivor [Hir98]. sus [Ano25c, Bor22b]. Sushi [Har07, Pic05]. Suspended [Ein05-35, Ein05c]. suspendierten [Ein05c, Ein05-35]. Suspensionen [vS06]. Suspensions [Hug54, Mar40, OW54, Haw02a, Pus91, vS06]. Susskind [Bri82]. sustainability [S^{+}10]. sustained [And73]. Sutherland [CPDH^{+}87, Hom05]. Suu [Joh11a]. Suzuko [Hir98]. sv [Inf58]. Svante [Kra13]. Svenska [Hil39]. svet [Inf79a]. sviluppo [EI65b, EI13, Seg91]. Svizzera [Kol58]. Swabian [Sto13, Bok15, Ger15, Kle14a]. Swallows [Wei13f]. swan [ZZY19]. Swann [Hag08]. Swartz [Ano88a, Ano88b]. Swastika [Luc11]. Sweden [BCS82, Sim94, Wid95]. Swedish [Cur38f, Hil39, Pla20a, Pla20b, Bal95b, CF82, CSW97b, EI38d, Fre13, Fre15a, Fri12, Gar41, Hil39, Lun74, RE90, Str83, Tis52, Val55]. Sweet [Cha74]. Swenson [Duf81, Hen82a, Hen82b, Kle82a]. świata [Ein35e]. swing [TLW^{+}10, Swi45]. Swiss [Ano99b, Bal96, Cah90b, Cah90a, Har91, Har96a, Har96b, Kil95, Pye95, Ram90, Sie97, Str90, War95, War96a, War96b, BH89a, BH93, KKS93, KKRS93, KKRS95, Med94, SCRS89]. Switzerland [Sho21, CC96, Kol56, Kol58, Sch05f, Sch05e, Sch12f, See52]. symbolic [NP88, Sie10]. Symbolischer [Pes13, Sie10]. symmetric [Lom05]. Symmetrie [Bac05]. Symmetries [GM80, Pit16, Wig67b]. symmetrischen [EG23]. Symmetry [HG05b, HG06b, Sch15b, Adl82, Adl83, Bac05, Blu04, Ein39d, Fey97, Fey11, Hsu00, Nam09, Nam10, SHL^{+}06]. symmetry-breaking [Adl82, Adl83]. symphony [Bar00c]. Symposion [Ano79g]. Symposium [Ano79g, Ano81d, Ano84c, BW04, Cli09b, Fer82b, Gil84a, Gol84b, HH14, Lar84a, Meh73b, NHP^{+}79, PJMS11, Wil84, Ber84b, BCS82, HE82, Ber84a, RMR96, Woo80b, Tre66, Wig96d, BDH89]. Symptoms [Ein52d]. Synchronicity [JP55]. Synchronisation [Dor68]. Synchronization [Coc66a, Coc66b, Coc66c]. Synchronizität [JP52]. synchronous [UKKP04]. Synchrotron [HEB^{+}80]. syndrome [Jam06, Sow01, Wel93]. Synge [Flo08, SO72]. Synthase [All02]. synthèse [Ano67]. Synthesis [Hoy54, AVvdV99, Ano67, Hoy46]. Synthetic [EdD29c, Mac01]. synthétique [EdD29c]. synthetischen [Cun22a]. System [Foc53b, Kra17, Mos21, Nor72, Ree06, vN37, Ein05h, Beh05b, Cha11, Ein18d, Ein27b, Ein39d, Kle74, NP88, SE79, Ash02, Beh05b, Ein22d, Ein05h]. Systematic [GS80, Pie24b, See25, EM31b]. Systematische [EM31b]. Systeme [MGG^{+}95, Kle74, Ein20d]. Systeme/Bunde/Havlin [MGG^{+}95]. systèmes [Chr22]. System“ [Ein22d]. Systems [Col68, Ein27b, How90b, How97b, Kuh54, NW49, Sha54, Ano19a, Chr22, Dra53, Dra01, Ein20d, EM04, Fer24, Kel58, Sch35b, Sch36, RS11b, Sha54]. századi [TB02]. Szilard [Bes93, Wig96i, Cha11, Cho10a, Dan97a, Dan97b, EE72, FS87, Fra05, Gra94a, Gra96, HGS87, LS92, LS94, PB79, SE45, SMM^{+}60, SBSW60, WS78, BST03, BSST04, CST01, CST02].

T [Ano87b, Ham97, SK94, Sie97, Inf58]. T. [Ber92a, KP55]. tōn [EI78a]. Tabelle [Fis02, HH14]. table [Fis02, Kau14, Lag12]. tables [Ste07]. Taboo [Fre08]. Tackles [Nad02]. Tactical [Sch93b, Sig78]. tagliyotehe [Seg86a]. Tagore [Gho17, Ray95, Sin01a, Sin01b, TE31b]. Tagungen [Hei75b]. Tail [Wei13f, Hac87]. Tailings [Mog02]. taistelu [Hot15]. Tait [ABdB^{+}53]. Taken [Cha23, CSW97a]. Takes [Bro79a]. Taking [Ano31b, Neu06]. Tale [CSW97b, OO01, CSW96, Jos95b, Red10, Tou14]. talent [Rod22]. Tales [Bro60, BS97, Kea12, Lan48, Lau45, Top07a]. talk [Bri95, Sow01, TT70]. Talking [Gio13b, Bre15, Yor87]. talks [BW91, Pap15, TE31b, Whi67, Ano21d]. tall [Ano96a]. Talmey [Gor22, Rav97]. tamped [Ree15a]. Tangent [Ein0x]. Tangentenbussole [Ein0x]. Tangled [Ber03, Her03, Sch03c, Her02]. tänkaren [Bal95b]. Tannenwald [Fre08]. Tanrısı [Tip12]. Taroni [Tug98]. Taschenbuchausgabe [Fis02]. Task [Ein55j]. tasks [SSSS95]. Tauber [For80a, Oha79, Van79]. taught [CG14]. Täydellinen [Hot15]. Taylor [Smo07a]. täysi [Hot15]. Tea [AH60]. Teach [Ano50c, Fow03, Orz12]. teacher [Fey95, Myr21b, Myr21c, She16b, Sho21]. Teachers [Ein08a, Pin00a]. teaches [Ill89b]. Teaching [Bru09, GdS15, Kro01, MM66, Par99]. team [Spo16]. tear [Gle10]. Technical [Bar80a, Ein20m, Ein31a, EGO97, Hei47b, Hei47a, Kur79, Hei46, HHMW93]. Technically [Cha74, Flü39]. technician [And17]. Technik [AH93, RKS08]. technique [GAHS96]. Techniques [HEB^{+}80, Wet80]. technisch [Flü39]. technische [Ein31a]. Technischen [EGO97, Hei46]. Technological [MGG^{+}95, Wel08, ZZ23]. Technologies [Can15a]. Technology [Fel11, Goo85, Ham11a, Ham11b, Ham85, Hen11a, Hen11b, Joh11b, Leg15, Mil11a, Mil11b, Nuv11, RW94, RKS08, RBS^{+}08, Scl89, Sei84, Set11, Tre88, Aer99b, AH93, Alf01, Bau95, Bau97, Ber84b, BCS82, Hun10, LW71, MSW88, Mul10, Wol93, Han80b, MA31]. Ted [NJP92]. Teil [ARK^{+}99, BDR93, Eis00, For80a, HS44, Hei69a, Hei72, OSWR86, vM96, vM99b, vM01, vM05c, Kra23a, Kra23b]. Teilchen [Bac89b, Ein05-35, Fis10, Fis12, SS97, Ein05c]. Teilhard [Ano67]. teilweise [Ein20l]. Telecommunications [Can15a]. Telemark [Hau74]. teleparallel [Sau06]. Teleparallelism [VT99]. Teleportation [Gis14, Zei10, Gis12]. Telescope [Fre27b, Ano17a, Fre27a, Gat09, Kin11, KN15, Poc98, Ban11, Nun22]. Teleskop [Fre27a]. Tell [Ano39a, Fos13, Mas09]. Teller [Dys76, BO76, BH89b, Goo04a, Goo04b, Her02, HM05b, Wig96b, Yor89, Ber03, Her03, Sch03c]. Telling [LSW09, Mar01]. Tells [Ano21d, FS06]. tem [vL49a]. Temperature [Joh13, Löw78, DDD^{+}06, PSM^{+}13]. Temperatures [RA13, YCKG11]. tempo [Pot11]. temporal [Dor95, KWGH97, dB65]. temporelle [dB65]. temps [Bie91, Edd21, Eis02, Fra91, Lam07, Spa17, W.25]. Ten [Ano29b, Fos13, GM06, Jon08, Ano23c, Hor09, Wac06]. tendencies [KM04]. teniendo [Yun15]. Tennessee [JJ81, Olw97, Olw04]. tensions [Lin15]. Tensor [Wei12s, Wei12t, Wei12u, Ein27a, HJ16, Mal91, Wei18a]. tensorial [Mal91]. Tensors [Agu90, Rac49, EG78b, Lud13, Sid76]. tenth [Ano67]. tenue [LdB12a, LdB12b]. teoreticheskoi [Ein40g]. teoretické [Těš12]. Teoria [Ein21p, Ein21q, Mas02, EI57, Hot15, Inf63, Inf68, Inf74, Inf92, KT15, Kop23b, NA92, Tug98, Bor22b, Ein85f]. Teoriaa [Ein21r]. teorías [Ano25c]. teorie [Seg91, Záv25]. teorii [Bič79b, EI59, EI62a, EI66a, EI98b]. teorija [Ano79a]. teorije [EI62c]. teorye [Sch27, Gru07]. teremtést [Tőz15]. Term [Ste49, Duc11, EE82]. terme [EE82]. Termessos [Row06a, Row06b]. természettörvények [Wig65]. Terminänderung [Ano79g]. Terms [WF49, RN16]. ternary [How96b]. Terrain [Ber93]. Terrestrial [Coc66a, Coc66b, Coc66c, Sch22b]. Terrible [Duf46, Res50]. terror [Bro02, DSW99]. Test [Ano22a, Ano16d, Bad95, CHSH69, CHSH70, CS60, Din59, Div78, ESA15, Ein19i, FG00, Gib91a, IS49, Ken54, Moy79, PD15, Ree06, Sei04, Sha64c, SPA^{+}68b, SAI^{+}71, Wat97b, Wil86, Wil90b, Wil93c, Yea08, Ano04a, Ano18c, BIT03, Cam88, Cap01, Car07b, Cre06, Dum04, EG80a, Har13c, HR02, Mer10, Pep04a, Per23b, RW02, Rod04, SPA^{+}68a, Vig94, dH13]. tested [Mas74]. Testimonial [Ein45f, Ein96e]. testimony [Hir98]. Testing [AHH05, Ano94a, Ano14a, Chr20, D^{+}95, Duc11, Hen92b, Hen94, Jon06, Lod20b, Mar06, Ove05, SK82a, SK82b, Sta03a, Bal86, Hac94, Mil91, Poo90, SWU09]. Tests [Lau45, MST^{+}07b, Rus25, WTB09, Ano00g, Ano07u, Jon01, MST^{+}07a]. Teubner [OSWR86]. Texas [Hen82b, Lew79]. Text [Fri89, HY86, Jam30, Ano84a, GR17, Sto96, Lau50]. textbook [FTG08, Sho21, Vog21]. Textbooks [Leg15, NM12]. texte [Sto96]. Texten [Hah88]. Textes [CE79a]. Texts [Ano79b, Sta05c, AKD23, CE79a, Hah88, HH96]. Texture [Ano09e, Gre04b, Gre05a]. Teylers [Lor14, Lor14]. Th [EG23]. Thad [You95]. Thales [Ano26a, Bro83]. Thank [Fus88, Fus90]. Thanking [Lau55b]. Thanks [Has09]. Theatrical [Lan00]. theft [Wes04]. Their [AP00, Ber93, Bič00, Čap76, EG80b, EBU^{+}52, Ein05-39, Fos13, HHW99, KKKM05, Klo97, Krs04, Moo16, Sch87b, Smo07c, Sta99a, Wat00, Ano25a, Bes93, Cal17, CÖT15, Ein02b, FG20, FPLR59, GTP22, HM11, Ill89b, Kan96b, Lin60, Lui11, MSYW12, McG93, McG98, Mon58, Par99, PL17, RCRC97, SBB16, Sch00c, Seg80, Seg84, Seg87, Seg07a, Tau12, Van07b, VM05a, Wie06, YCL20, ZXLX14, vBL96b]. them [Ano19a, BB13, Pic08]. Thematic [Hol73, Hol88d, Top89]. Theme [How92, TA17]. themes [TA17]. themselves [KK05]. Theodor [B^{+}79, B^{+}79]. Theodor-Zink-Museum [B^{+}79]. Theodosius [Adl03a]. Theological [Ric06b, Moo81c]. Theology [Mas01, Jam99, Jam02, Gor02]. Theorem [Ein05c, EH05c, KN08, Mao07, Ung98, BN97, CM89, DPS97, EH10b, Ein17c, Ein05-32, GHSZ90, Hig14, Lie78, Oyi04, Wei13g, Žáč12, Kox06, vN37]. Theorems [Ein17c, Har97, Hel82b]. theoremu [Žáč12]. Theoretical [ARK^{+}99, Bro23, D’A00, Dic64, Ein14a, Ein20h, EE22, Ein27e, Ein33f, Ein34k, Ein87a, Ein05-49, FW60, Gal82, Gar72, Jac87, JM17, Kan79, Kra18, Lon84, Lon03, Meh73b, Ser88, SM41, Wei12g, BB89, CSW97b, D’A01c, Ein40a, Ein40g, Flo99, For63, Hen92a, Hom05, JRG02, JM86, Lan21, Lor31, Set03, SF33, Van07a, War89, Yan80, Ein07f, Ein15a, Ein22m, Häg46]. theoretical-historical [Hen92a]. theoretician [Ken14]. Theoretische [MGG^{+}95, SM41, Ein07f, Ein15a, Ein22m]. theoretischen [Ein27e]. Theorie [Ano21e, Ano21i, Ano23c, Bäc19, Bäc21, Ein02a, Ein02b, Ein03, Ein04, Ein05c, Ein06d, Ein06b, Ein06f, Ein07b, Ein07e, Ein08b, Ein10b, Ein12j, EG13a, EG13b, Ein14q, Ein15g, Ein16f, Ein20b, Ein20d, Ein22o, Ein23h, Ein24l, Ein25b, Ein25c, Ein27f, Ein30r, EM31a, EM32a, Ein05-40, Ein05l, Ein05w, Ein05-35, Ein05p, Ein05-42, Ein05s, Ein05-43, Ein05-44, Ein05-39, Fer14a, Foc54, Fra20, Kie94, Kra20a, Pla00b, vS06, Ano23a, Bac00, Cun22a, Ein21k, Ein21j, Ein23g, Ein30c, Ein33c, Ano15a, Ano16a, Ber22, Car22a, Car23, Cun22b, Edd21, Ein10f, Ein14c, EdD29c, EK53, Ein55i, Ein56d, Ein64b, Ein76b, Ein76a, Lan08, LdB12a, LdB12b, LG97, Lev04b, LB39, MK56, Pat99, Pia31, Spe67, W.25, Wei15c, Ano17b, Ein10a, Ein22n, Ein05-29, Sta95c, vB12]. Theorie [vB14, vB18, vB20, Ein51g, Ano22a, Wei15c]. Theoriebildung [Müh11]. Theorien [JP52]. theoriengeschichtlichen [Hen92a]. Theorienholismus [Hen87]. Theories [Abr07d, Ano21f, Ano22c, Ano84c, Ano90a, Bar91, Ben05, BB49, Bor49a, Bri49, CHSH69, CHSH70, D’A01b, Dor68, Ein16i, Foc53a, Fox65, Goe04, Gup57, How99a, Kae31, Kra96a, Lan14, Nor07d, Pia29a, Pia29b, Pic21, Poo21, SM07, Vec17, Wal02b, Ein16i, Abi00b, Ano25c, Ber82a, BMS12, BB89, CF11, CT24, Chr22, Col57, DGM04, EI38d, EI38e, EI48a, Ein60b, EI63, EI78b, EI89b, EI93, EI15, Eng02, Eng05a, Flo99, Fri83, Gib91b, Gio20, JRG02, JP55, Kai05b, Kat92, Koc01, LSS16, Lom05, MP74, Mus15b, Ner84, Nor92, Par11, PKLT17, PP03, Put09b, Rou64, Rou65, SW94, Sar96, See25, Sin61, Sin70a, Str68a, Chr22, Cun22b, EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, MP74, W.25, EGS77]. Theorist [SM08, vD17, Kra13, Leg04, Log96]. Theory [Abr13, Abr07b, Abr07c, Acz01, Aga93, Ano17, Ano19b, Ano20a, Ano21d, Ano22a, Ano23b, Ano23e, Ano23a, Ano25d, Ano25a, Ano28, Ano29b, Ano29e, Ano29d, Ano31b, Ano35, Ano39f, Ano06h, Ano07a, Ano07d, Ano07c, Ano07e, Ano07h, Ano07f, Ano14a, Ano15d, Ano16a, Ano16d, Bac00, Bai92, Bai01, Ban51, Ban06, Bas65, Bel26, Ber06b, Ber42, Bet88, Bha49, Bha86, BGLR17, Boh62, Boh65a, Boh79, Boh89b, Boh96, Boh06, Bor20, Bor22a, Bor22b, BB24, Bor49b, Bor62, Bor64, Bor69a, Bro23, BEC99, Cao06, Cas21, Cer00, CC67, Cha01, Cho06, Cle82, Cob07, Coh80, CHR90, Cor15b, Cor04b, Cro20b, Cun19a, Cun19b, Cun19c, Cun20a, Cun20b, Cun20c, Cun22b, Cus82, DGPS99, Dic81, Din49, Dir65, Dir75, Dir79d, Dir96, Dra14a, ESA15, Edd20a, Edd21, Edd23]. Theory [Edd29a, Edd30b, Edd06, Eĭn86b, Ein06b, Ein07b, Ein08b, Ein11b, Ein14e, Ein14d, EG14, Ein15j, Ein15e, Ein15g, Ein16p, Ein17e, Ein17i, Ein17g, Ein18e, Ein18m, Ein18n, Ein19i, Ein19k, Ein20d, Ein20o, Ein21k, Ein21t, Ein21p, Ein21r, Ein21j, Ein22g, Ein22a, Ein22q, Ein23j, Ein23k, Ein23b, EG23, Ein23d, Ein23i, Ein24h, Ein24l, Ein25b, Ein25e, Ein25c, Ein25h, Ein25i, EG27a, EG27b, Ein27f, Ein28c, Ein29d, Ein29e, Ein29f, EdD29c, Ein30n, EM31a, Ein31o, EM32a, Ein33e, Ein34j, ER35, Ein38a, EI39a, Ein44e, Ein50l, Ein51a, EK53, EI62b, Ein67b, Ein76a, Ein82f, Ein86e, Ein88b, Ein96a, Ein96f, EK03, Ein05c, Ein05o, Ein05w, Ein05-35, EF05, Ein05-27, EH05c, Ein05-42, Ein05s, Ein05-39, Ein07b, Ein13, Eis06, Eng17a, FS09]. Theory [Fer14c, Fer81, Feu71b, Fin81, Fin93, Foc55, Foe23, Fol15, For71b, For80a, For80b, Fra20, Fra37, Gam42a, Gam42b, Gam85, GO08, Gar70, Gei15, Gia70, Gia71, Gin82, Goe70, Goe93, Gol68, Gol69, Gol70a, Gol76, Gra07, Gre99a, Gre99b, Gre00, GN11, Gru57, Grü62, Hag08, Häg46, Hal18, Has07b, Hea13, HSD34a, HB63, Hei93, Hen86a, Hen90, Hen16, Her71, Hol79b, Hol88c, HW07a, Hoo70, Inf50e, Inf50c, Inf50d, Inf51b, Inf56e, Isa85, Ita82, Jan95, Kae13b, Kai81, Kai05a, Kal13, Kas20, Ken19, Kle65b, Kle80, Kol07, Kox88, Kra79, Kra20a, Kra23a, Kra23b, Kur79, Lai20, Lan08, Lau50, Lau55a, LG97, Len49, LC29, Lie45, LEM^{+}23, LEMW52, Lor10a, MNT^{+}00, Mas02, McC70, McG19, MC80, Meh73a, Mel82, Men49, Mer55, Mer85b, Mie07a, Mie07c]. Theory [Mil75, Mil79, Mil81, Mil98, Mos18, Mos21, Mou20, NS05, Nav15, Nic20, Nor07b, Nor72, Nor14, Oha79, Ove05, Pal83a, Pal83b, Par79, Pau49, PV49, Pau81, Pau94b, Pau94e, Phi22, Pla20b, Pla20c, Pla22b, Pla10, PFM17, Put25, Pye73, Pye77, Ram83, Rei20, Rei27, Rei28a, Rei28b, RK77, RSS^{+}07a, RS07c, Ren07c, uJR15, Rey25, Ric06b, Rob49b, Ros37, Roy19, Rus25, Sau14, Sau16, SS21b, Sch74, Sch27, Sch22b, Sch53, Sci61, Sha14, Sig92, Sil36, Sin00a, Sin00b, Sin05, Sla63, Smi02, Smo07b, Sta98b, Ste98, Sul64, Sus06, Tal32b, Tau79, Tay21, Thi21, Tho83, Tre75b, Van11, Van79, Viz79, W.25, Wal21, Wat98, Waz09, Wei72, Wei12l, Wei12p, Wei12x, Wei15a, Wen49, Whi06, Whi83b, Wie21, Wie29, Wil81, Wil93b, Wis82, Wol87, Woo28]. Theory [WJ21, YN11, Yuk49, vB12, vB14, vB18, vB20, vKL49, vL19, vL21, vL52, vL56a, vL65, Abi00b, Abi03, Acu14, Acz03, Adl82, Ano15a, Ano19a, Ano21e, Ano21i, Ano23c, Ano79a, Ano17b, BV09, Bäc19, Bäc21, Bag92, Bag04, BDH04, Bel75, Bel96b, Ber22, Ber65, BRVF84, Bes04, Bha92, Bič79b, Bir82, Boh89a, Bor65a, vBT82, Bro10a, Bro11, Bro16, Car22c, Cas76, Cas23, Cas03a, Cea95, Cea97, Cea98, Cha76, Cha98, Cli65, Cli87, Cre81, Cun15, Cur92, CM89, Dar96, Dar19, Dir79e, Dir80d, Dir79f, Edd79, Ein02a, Ein02b, Ein03, Ein04, Ein05c, Ein06d, Ein06f, Ein07e, EH10b, Ein10f, EF13, Ein15b, Ein16f, Ein16m, Ein17h, Ein23c, EE23, Ein29j, Ein30c, EI38a, EB38, Ein45b, ES46b, Ein48d, Ein50a, Ein51g, EK54]. theory [Ein56c, EI57, EI59, Ein60d, EI61b, EI61a, EI62a, EI62c, EI66a, Ein82b, Ein85d, Ein90, Ein95b, EI98b, Ein05-40, Ein05l, Ein05p, Ein05-43, Ein07f, Ein07g, Ein09c, Ein10, Ein15b, Els24, Fan06, Fay10, Fer14b, Fer14a, Fey84, FBD05, Fey06, Fin86b, Fin97, Flo99, Foc53b, Foc54, Fok59, Fok64a, Fok64b, Fok65, FMM96, Fra11, FB19, FB20, FB22, FB24b, Fre11, Gam66a, Gam66b, Gam72, Gin62, Gio22, Gol70b, Gol15, Got00, GG15, GH07, Gut97, Hal15, Hei50, Hen86b, Hen85, Hen87, HS51, Hoe94, Hot15, Hu05, Inf51a, Inf63, Inf68, Inf74, Inf92, Jon06, Jos80, KU38, KT95, KT15, KP55, Kie94, Kil94, Kop21, Kop23b, Kop23a, Kra20b, Kra21, Kra25, Kru94, Lan59, LdB12b, LSS16, Leh19, Ler91, Li07, Lie08, Lin15, LB39, LFKW06]. theory [Lor20, Lor09, Lor10b, Lor10c, Lou83, Lud13, Lum11, Mad95a, Mar06, Mas74, Mau96, MR82d, MR82e, Meh98, MK56, MD15, Mik65, Müh11, Nav91, Nug88, O’R97, OMNM14, Oyi04, Pai79, Pai80, Pap15, Par99, Par86, Pau55, PS10, PS95, Pet21, Pet23, Pla20a, Pla22a, Poo90, Pou93, Ray71, Rei21, Rei24, Rin03, Sal74, Sau06, Sau19, Sch03a, ST94b, Sch57, See25, Sie91, Sil22, Sla60, Sla22, Smo06, Sta95b, Sta95c, Str94, Tel95, Tre71, Urb22, VM99a, Voi86, Wal77a, Wal92a, Wal94a, War91, War92, War03b, Wei00, Wei05b, Wei12c, Wei15c, WZ83, Woi06, Yag85, YN10, Záv25, Zee03, Zee10, dB36, vD21, vL23, vL55c, vS06, Abi00a, Apf81, Ber76a, Ber79a, Bro19, Bro20a, Bro20b, Car24, Cha75, Cre83, Cun22a, Cun22b, Cwi87]. Theory [Dar06b, Dir80c, Edd20b, Edd29b, Edd59, Edd66, Edd87, Ehl05b, Ehl09, Eĭn85a, Ein10a, Ein12j, EG13a, EG13b, Ein13b, Ein13e, Ein14g, Ein14i, Ein14l, Ein14m, Ein14q, Ein15i, Ein16d, Ein16b, Ein16h, Ein16j, Ein18f, Ein19e, Ein20c, Ein20b, Ein20g, Ein20k, Ein21b, Ein21n, Ein21q, Ein22n, Ein22o, Ein22r, Ein23g, Ein24c, Ein25d, Ein29k, Ein29b, Ein30o, Ein30f, Ein30a, Ein30r, EM30, Ein31j, Ein31k, Ein32a, Ein33c, Ein36f, ER36, EI49b, Ein53b, Ein54g, Ein55i, Ein56b, Ein61a, Ein61b, Ein64b, Ein76b, Ein77b, Ein78, vM90, Ein95d, Ein96d, Ein99, Ein00c, Ein01b, ELGC05, Ein05-29, Ein05-44, Ein06b, ELGC06, Ein08b, Els24, Fer14a, Feu71a, Fre16, Fre19, Fre20, FB24a, Gin59, Göd49b, Gol67, Gra71, GG05b, Gru07, Hen92b]. Theory [Hen92a, Hen94, Hen97, Hir69, Hir76, Hol68a, Inf55c, Lan21, Lan55, LLH^{+}80, Lev04b, Lor17, Mal91, Mar97, Men26, MTY79, Mey85, Mil82, Pai88, Pat03, Pat99, Pau21, Per23b, Poo90, Rei49, RR65, RR69, Ren05g, RR89, Row06d, Sau07a, Sch17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, Sil24, Sta98c, Str49, Tau63, Thi21, Tre21, Waz14, Wet80, Wey23, Wey09, Wey20, dH13, dSR16a, dSR16b, dSR17, vG30, Adl83, Ano22a, Ein14c, Jan73, Pie24b, She16b, Ein05c, Wei15c, Ein15b]. Therapy [Rol10]. There [Ein32c, Mer85b, Sch52b, Sch52a, Ein12d, HS14, Roe93]. Thermal [Abi91, Ein25a, Ein05t, vHJK39a, Abi00b, BJ41, Ein06d, Bos24b]. thermische [Ein11h, Ein05t]. thermischer [BJ41]. Thermodynamic [BT12, Ein05-39, Ein02b, Ein07i, Ein05-32, Ein12i, Ein12h, Ein13a, Ein05-30, Ein12f, Ein05y]. Thermodynamical [Kle49]. Thermodynamics [Ein05s, Kle67, Abi00b, Ein03, Gio20, Iro04, Kra08, Liu92, Nug00, Rus81, Sch89, Tol34, Tol87, Wal94a, Ein02a, Ein11f, Ein05w, Ein16i, Ein05j]. Thermodynamik [Ein05w, Ein05s, Ein05j, Ein02a, Ein03, Ein11f]. thermodynamique [Ein13a, Ein16i]. Thermodynamische [Ein12f, Ein05y, Ein02b, Ein05-39, Ein12i, Ein12h, Ein05-30]. thermodynamischen [Ein07i, Ein05-32]. Thermomolecular [Ein11h]. Thermonuclear [GT38, Gon96b, Gon96c, Gon96d, Gon96e, Sca14, Bay94, Kah60, Kah69, KJ07, Lan15]. ThermoPoetics [Leg15, Hen11a]. These [Ove87, Ano25a, BM09]. thesis [FBD05, FCKT11]. thickness [AH96, dB65]. Thief [Phe21]. Thing [Col13, Che00, Sta07d]. Things [Bir23a, Bro21, Ein52e, PP05, Sch07d, Fey00, Nor21, Pra12, Sch04c]. think [Bre05, FL88, Smi15a, Tho15]. thinker [Bal95b, Geo53, Ham86]. thinkers [Bro06, Bro07, Wes01]. Thinking [Rot87, Ank15, RMI21a, Ryc92, Sch05a, BT15, Pfe77, Smi15b]. Thinks [Bro21]. Third [Ano41, Hol03b, Ren07d, Ano97b, Bey77, HW12]. thirteen [Bey49, NN01]. Thirties [Kra82a, MH71, Stu93]. Thirty [Gam66a, Gam66b, Gam72, Gam85, Wig80, Wig96l, For82b]. Thomas [Bar05f, Boy88, Hol91a, Pul11, Rus17, Wal61, Ber94, Ein29a, Sch93a, Tho88b]. Thompson [Bes93]. Thomson [DTF97]. Thorium [Ano41, HS39a, Mei39, MF39b]. Thorndike [Bey06, Cas05b]. Thorne [Vog21, Har95]. Thorstein [Gan97]. Those [HG05a, Cho16b, Ein52e]. Though [Ano35]. Thought [Boy85, Boy94, Car17, Cun22b, Dir79a, Har69, Hei49b, Hos15, Kle67, Kur12, MNT^{+}00, Mos21, Rai14, Ram81, Top89, Bal93, Bal94, Bal95a, Bal99, Bal08, Bal11a, Byr80b, D’A50, EI65a, Gio14, GB06, GB08, Har13b, Her01a, Hol18, Hol73, Hol88d, Jag01, Ken07, Kon15, Lau88, McC01a, Mel79b, Moo89, Mos22, Nik12, PL17, Ros80, Tal82, Tay49, Bäu83]. Thoughts [EGR11, Ein36g, Koj08, Lan08, McG19, Tel78, Bat05, Ein14l, Fey98, Mul05, RS07f, Vaa17]. thousand [Jun58, Jun82]. Threat [Duf46, DS83b, DSW99, LM90, Sal95]. threaten [Bre08]. Three [And73, Ano33c, Ano69c, Ano79b, BieWSK67, DM96, HI87, HSD34b, MSS17, Max93b, Med94, San77, WCL^{+}00, Ano09b, BH09, Des72, Erb10, FB24a, Gut98, Kah69, Lor14, Ne’06, TA17, EI38a, For82b, Sch05c, Wal04b, ZDSF11]. three-body [Gut98]. Three-Dimensional [WCL^{+}00]. Three-Man [ZDSF11]. threshold [Kra88, PBM^{+}12]. throughout [MNOW00]. throw [Ano04a]. Throwing [Gef02]. Thumb [Kea12]. Thursday [Spo16]. Thus [Ano39a, Res50]. Tibbetts [Ach13]. Tidal [HEB^{+}80]. tide [Ian93]. til [EI39a, EI61a, Jak13, NN01]. Till [Räz16, EI38d, RE90]. Tilman [Bad15, Lac08, Pye95]. Tilting [P.90]. Timaeus [Joh04]. Time [AZ05, Ano25a, Ano09g, Ano16c, Bai75, Ber68a, Ber85, BOS14, BH17, Bro20a, Can16b, Čap76, Car17, Car07a, CSM06, Coh80, Cos87b, DR16, Dor95, DD11, Ein18h, Ein19a, Ein46a, Ein01c, Fau79, Fie16, Foc55, Fok49, Fok65, Fol15, Gal00, Gal03, Goo85, Got01, Got02, Gre04b, Gre05a, Grü73c, Hal17, Har95, Hay91, Jan03, Joh13, M.25, MC80, Mir02, Nov98, Ove13, Pau94d, Pep04c, Plo06, Rei27, Rei28a, Rei28b, RF57, RK77, Rei06, Rog10, Rüg88, Rya06, See60, Sei84, Smi97, Smo14, Str04, Tay21, Ven12, Wan95, Wei13a, Wei12w, Wey18, Woo06, You98, You05, Zeh07, dS19, Ano19a, Ano04a, ACT23, AA75, Bal05, Bar00c, Bel75, BCS82, Bie91, BLR15, Bon81, B^{+}61, Bro05, Can15b, Car02, Cha83, Cho16a, Cle96, DS01]. time [Dav95, DiS06, EGS77, Edd20b, Edd21, Edd29b, Edd59, Edd66, Edd87, Ehl05b, Ein29i, Ein30e, Ein09c, Eis02, Enz02, Far05a, Fey66, Fey97, Fey11, Foc54, Fok59, Fok64a, Fok64b, Fri83, Fri14, Gal04a, Ger69, GH84, GW06, Goo02, Gre15a, Haw88, Haw98, HM05a, HE08, Hon96, Inf99, Kak04, Kak05, Ken03, Kin95, Le 03, Mau90, Mer05, MD15, Mil01, Mil94, Mus15b, Nor89, PR86, Pot11, PP03, Rei24, Rei58, Rei11b, Ren97b, RR89, Sal05, Sau06, Sch17, Sch19, Sch20, SB20, Sch22a, Sch63, Sch05d, SB06b, SB06a, Sch99b, Sch86, Sch02c, SD93, Spa17, Sti13c, Tho93, Tho94, TLW^{+}10, Wal77a, Wal92a, Wei04, Wey22a, Wey22b, Wey50, WL12, Wil75, You91, Zew02, ZCPB11a, ZCPB11b, Ano22a, Eng17b, Kar05, Rus81, Smo07a, Haw92, Wil16a, Her17, Nah01]. Time [Ste05d, Sta05a]. Time-dependent [Fie16]. time-of-swing [TLW^{+}10]. timeless [Bar05d, Bar05e]. Timeline [BHJ^{+}00]. timely [Rol89]. Times [Ano15c, Bru03c, Fra49b, Fri79a, Imp19, MGG^{+}95, Mos18, See56b, Shi17, Tur05, BO76, BB05, Bre97, Cla71, Cla73, Cla84, Cro01, EW88, EW00, Fra47, Fra48, Fra53, Fra65, Fra88, Fra91, Fra02, Gim15, Jon06, Kle05, Mar06, Pai91, Sac81, Str83, Unr99, Cor15a, Dea13, RFP^{+}45, Bal72, Dys15, Row17]. Timing [Sta03a, CJT22]. Timor [Sut97]. Timothy [Dix00, Pin00a]. tips [FGL^{+}06]. titles [WK12]. Toast [Sch35d]. Tobé [Swa02]. Tod [Sti05]. Today [Ano98, How21, RMR^{+}50, Ano15f, Sal95, Whe68b, Est06c]. Todes [Hal55]. Todestag [Sti05]. Todo [PPC05]. Together [Ein49d, Ein70b, Ein91g, Shi17, HM11, Sch82a, Sch03b]. Tokyo [Bar05f, Ano41, Lau45]. Told [CW12, Sha92, FSE31, GTP22, Gro62, Gro83, Hes15b, KK05, RRK^{+}64, Wol00]. Toll [D’A93, Lau45]. Tolls [Kuc11]. Tolstoi [Rei41]. Tolstoy [Rei41]. Tom [Ham97]. Tomonaga [Dai67, Dys65, Sch94]. Tomorrow [SAY^{+}82, Ano15f, Rob58]. Tompkins [Sta99c]. tømrerlærlingen [Jak13]. Too [Hür22, Dob06, Mor95b]. Took [Ano29b, Lau45, Ano97a]. Toolkit [LHZA20]. Tools [BHJ^{+}00, Har13b]. Top [Dee08, Hoh13, Kha12, Kha20, CM02b, Jac11]. Topological [BT12, MGG^{+}95, WH99]. Torch [Ser88]. torno [Bla86]. Toronto [Ano23a, Van22]. Torsion [GM00, Sch19, vE10]. torus [DSZ12]. tot [Ano15f, EI38a]. Total [Ano19c, Edd19, Sim22, DED20, Ein41a]. totalisierenden [Zim84]. totalizing [Zim84]. Töter [Kle93]. Tough [Pal95]. tour [HG07, Swa92, Swa95]. Tourist [Hun15]. tout [Der15]. Tower [Fre27b, Abb98a, Fre27a, Jos95b, Smi02, Egg95, Fre27a, Hen92a, Hen97, Vir92]. Town [Mac85, Mas95]. Townsend [Hon89]. Toy [Bro21, Rin03, Zee89, Zee90]. tr [Seg91]. Tra [Mas02, RRR07]. Trace [Alt13, Kae38]. Traces [Jam30]. Tracing [ZZY19]. Track [Fey05]. Trade [Hec12]. Trademark [Rho93]. trading [Das12, Tur06b]. Tradition [Hei83c, Dre87, You91, Hei77]. traditions [Pat99, War93]. traduction [CE79a]. traduites [Ein05-47, Spe62]. Trafficking [Ren08]. Tragedy [Bro17, Fre08, Jon08, Lit09, Mul99, BS05, Fra89, Fra04, Mal08]. Trägheit [Ein05b, Ein06a, Ein07d, Ein05n, Ein05-34, Ein05v, Ren97b, vEPF09, vEPF22]. Trägheitsfusion [JDB^{+}97]. Tragheitsmoment [Ein20n]. tragic [Tho06, T– D82a, T– D83, T– D85, T– D88, T– D92, T– D93, T– D07, T– D12, Uda94, Uda98]. trágica [T– D92]. tragický [T– D07]. tragische [T– D82a, T– D83, T– D85, T– D88, T– D93]. trail [Bai92, Bai01, Ing05]. trailblazer [Joh11a]. train [Str97b]. trained [Pos11]. training [Ein48e, JDB^{+}97]. Trains [For14]. Trakl [Chi08]. Trans [Bro72, Man86, Smi02]. Trans-Atlantic [Man86]. Transaction [Seg91]. Transactions [Qui05]. transat [LRL15, LRL17]. Transcendence [Pic05, Har07]. Transcription [Ano07b, Gen98]. transcripts [Fra93a, Fra93b]. Transfer [dHvdB49]. Transformation [Bor61, CC67, vL19, vL21, vL52, vL56a, vL65, Ein05d, Ein05-37]. Transformations [MSS22, PB06, Sch49f, Bus12c]. Transformed [Kar05, Bar06, Kak04, Kak05, Wei04]. transience [RW07]. Transistor [Ano56]. transit [FG20]. Transition [Lid67, Tur67, WCL^{+}00, YCKG11, ABW99, Ell90, HLZJ91, dB65]. Transitions [CM67, Tur66]. Translated [Dix00, Hie68, Kuh54, Nav15, Woo28, For80a, Hoh13, Lig97, Mic68, Sha54, Smo07a, Van22]. Translating [Slo22, Bin06]. Translation [AP65, BH87, BH89a, BH93, BSKS04b, BSR^{+}06b, BRS^{+}09b, BIRS12b, BIR^{+}15b, BRI^{+}21b, Ein07a, Ein05k, Hav88, Hei53a, JSI^{+}02b, Kno97, SKJI98b, Vai89, Ano84a, Cah90b, CE79a, For05, Gen98, Hen88b, Jak78, OM14, OM22, War96b, Pye89, Har91]. translational [Sch19]. Translator [Bod13, Nav94, Cah90b, Mcc87, Ste90]. translators [Par99]. Transmission [Nis79, Sla42, Ein14f]. Transmutation [Kri19a, Kri19b, Per39]. Transnational [HY15, Lar12]. transnationale [Lar12]. Tranströmer [CGG^{+}13]. Transurane [MH71]. Transuranic [Tur40b, MH71]. Transuranium [McM51, Qui38, Sea58]. transversalen [Ein06e, Ein05-36]. Transverse [Ein05-36, Ein06e]. Trap [Lev01b]. trapped [DGPS99, Fet09, ZPSK10]. trapping [Phi98]. traps [BLI^{+}02, JDB^{+}97]. traps/in [JDB^{+}97]. Traum [HK94, Ein17b]. Trautman [Bič79c]. Travel [Ano12, Ber68a, Ein22, Fol15, Got01, Got02, Ken22, Ove13, Smi97, TB02, You98, Hon96, Ill06, Nah01, Ros18, Sou19, YotY99]. traveled [She16a]. Traveling [Ken07, Ost15, Dro20]. Travelling [Rod19]. Travels [Le 03, HW08, Smi07a]. traversed [Ein10e]. Tread [Sei03b]. treason [Gro56]. treasure [Ano17a]. Treasures [Ros12, MFO14]. treasury [FF91, GW83, McC48, Ree11c]. treated [Ill06, Smi07a]. Treatise [EL05b, Coo83, Joh11a]. Treatment [Kan79, Pie24b, Sta05c, Ang13, Bor22a, MNOW00]. Treatments [O’L64, Gio14]. Treaty [Bad95]. Treder [Pye80, Hor79b]. Tree [You95].

Trefftz [Ein22b, Ein22b]. Trend [Lau55b]. Trends [FPLR59, JP89, Nye19, BMS12, JP08, Rei05, Lin60]. Trespassing [Gef14]. tretten [NN01]. treu [Hen21]. Trial [Ste95b]. Triangle [BOPS99]. Tribute [Ano49b, Ein54h, Ein91f, HFN06, LSF94, Tho88a, ČK79]. Tributes [Ein89a, BieWSK67]. Trickery [Ove05]. tricks [Kro01]. Trieste [Meh73b]. Trinity [Bai75, Els95, FV13, Kis80, KMN^{+}97, Mer97, Mer01, Nic87, Ree06, SS65, Sza84]. Trinkspruch [Sch35d]. Trip [Hea13, Hos15, Ove13, Ano00a, Ano00b, Cle96, Cra00a, Goo00, Lie08, Pat00, Pat01b, Pea06, Swa92, Swa95, Vea01, WD06]. Tripadvisor [Owe12]. triple [Shi75]. triumph [BS05, Das91]. Triumphs [Ano19b, Kri99]. trivialities [Ank15]. trois [Spe62]. Tropical [BC82]. Tropp [Jos95a]. Trouble [Cas05b, Smo07b, Dic80, Mar95a, Smo06]. Troubled [Mos18, Rei11a, Shi17, Mag09, San11]. troublemaker [vC05b]. trous [Sta92b]. Trouton [But68, Jan95, Jan03, Teu96]. True [Dry14, FC07, Hes04, HD02, MP01, MP12, Shu17, Sil23, Wat98, Abr03, Cas03b, Har13b, Jon01, Lev04b, Mar11a, Mor95b, Rio87, SE79, Top07a, Ree17d]. trug [Ank15]. truly [Ryc14]. Truman [Cli09a, Ham97, Gor09, Has05, Lef95, Wal97, Wal04a]. trust [Dar21]. Truth [BPP^{+}11, Cha87, Ein29g, Kni12a, Kni12b, KS95, MNT^{+}00, RS13, Arg09, Cra07, Fra50, Fra51, Her33, Her63a, Her67, Her75, How94a, Mar01, Mar11d, RS80b, Wal95, Wal02a, C.50, Fra52]. Truths [Joh07]. Trygve [Zel90]. Trying [Cau91]. Tsukerman [Dix00]. Tube [Wei12b]. Tudományban [TB02]. Tullio [Lam30]. Tumbling [Buc13]. tumult [Pat85]. tumulte [Pat85]. Tumultszenen [Ano20e]. Tunable [MPBL09]. tunge [Hau49, Hau63, Hau64]. tungtvannet [Hau53, Hau65, Hau83]. Tunneling [BR85]. Turbulence [vKL49]. Turbulent [LB49]. Turing [Sha14]. Turkey [Rei07]. Turki [Bjø20a, Bjø20c]. Turkish [Bal08, EI94, EI11, Gol97]. Turm [Fre27a, Hen92a, Jos95b, Vir92, Fre27a]. Turm-Teleskop [Fre27a]. Turmteleskop [Fre27b]. Turn [Fru82, Hie82, BS79, Kat92, Wal81, Yos95]. turned [Ahe03, Ano10d, CL19]. Turner [TH79]. Turning [BRS^{+}12, CG01, RR06, Stu75]. Turns [Foe23, JDB^{+}97]. tutto [KT15]. Tuva [Lei00]. Tuvia [Gru07]. Tuxedo [Con02]. Twain [LD08]. Twelve [Vit15, AA75, Wil75]. Twentieth [Bar05f, Ber00, BPP95, Bru06, Cro16, D’A00, Har79d, LL82, Nau17, Rup09, Bro00, D’A01c, Gab95, GS05, GS06, Har06b, Hol96b, How14, Kra99, LSW12, Lig05, LY95, Moo81c, Whi16, Bar98b, Bro72, CM86, CM97, Dar04a, Din66, EMH96, FW60, Hol00a, Nye04, Put09a, TS01, TB02]. Twentieth-Century [Bar05f, Ber00, Bru06, Nau17, Rup09, Cro16, Gab95, How14, LSW12, Lig05, CM86, CM97, Dar04a, Din66, Put09a, TS01]. Twenty [Wol67, BH09, Bro06, Bro07, LW71, MT98, See25, Spe05]. twenty-first [MT98, Spe05]. Twenty-Five [Wol67, LW71]. twenty-four [Bro06, Bro07]. twenty-three [BH09]. Twice [Cho05a, Sza84, Tho85]. Twilight [RSS^{+}07a, RSS^{+}07b, RS07b]. Twin [Ano07u, RR89, Ano10b, Sac85]. Twins [Wei12q, Jan12]. Twiss [PBM^{+}12]. Two [Ano35, Bes04, But23, CHS97, Cor04b, Eds82, ER36, Hes04, Hon89, Jun05a, Kra96a, Kra23b, Kuh54, Lau51, Max93c, McL99, QPSD13, Sal74, Sau16, Sau21, Sha54, Sim22, Sta07a, CSC23, Col68, Des72, Dra53, Dra01, Dür62, Dür80, Dür06, Ein22b, Ema85, Gra82, Hal15, Hen86a, Her66a, HD02, Kal92, Ler73, Lin04a, Ne’06, NBDD^{+}12, PLM^{+}0 , Rob38, TV07, Tsu99, Wad81b, vH62, EG78a, Ein17d, EM30, Gut15, RS11b, Row06d, Wue05, Mil90]. two-blackboard [TV07]. Two-Body [ER36]. two-component [NBDD^{+}12]. two-dimensional [Tsu99]. two-fold [CSC23]. two-slit [Kal92]. tycoon [Con02]. Tyndall [Leg15]. Type [Göd49a, MF39a, EM33a, Joh14, Sta87, Abr04c]. Type-II [Abr04c]. Types [Hon89, Flo99, Hah42, HS44]. Typewritten [Lau50]. Typologisierung [Fla15]. typology [Fla15]. Typus [EM33a]. Tyranny [M.25].

U [GRE^{+}01, Kle14a, Pye80, ARK^{+}99, HEB^{+}80, MGG^{+}95, OSWR86, PHvMW93, Wil36, Ree17a, T– D69, T– D95]. u.a [HvMW79, PvM93, vMHW85, vM93, vM96, vM99b, vM01, vM05c]. u.a. [Eis00]. U.K. [Hoc83]. U.P [Nau17]. U.S. [Ano45a, BBD^{+}11, Lau45, Lau55a, Res50, SK82a, SK82b]. U.S.A. [Hoc83]. U.S.S.R. [Hoc83]. ua. [PHvMW01a, PHvMW01b]. UA1 [PA21]. UA2 [PA21]. überall [EG23]. überarbeitet [BS92]. Überlegungen [Ein16d, Hof18, Ein53d]. überquert [Boz14, Len17]. Übersetzung [Hei53a]. Überwindung [Yos95]. Uchū [Gol98b]. Uchū-no-shōtai [Gol98b]. UCLA [Cli09b, Ein32b]. Udvejen [Ein46g]. Udvikling [EI39a, EI61a]. Ueber [Ein05-39, Kra20a]. Ugly [Gun15]. UK [Gor02, Har02b, JDB^{+}97, Sta99b]. Ulam [Han92]. Ulf [Coo04, For05, Kau14, Mau09]. ulici [Pát79a]. Ullrich [Tre88, Nye88]. Ulm [Spe79, Ste05b, Spe79, Ste05b]. ulstrakorotkikh [Sla47]. Ultimate [Gre99a, Gre99b, Gre00, Hel09c, Mor96, Bar91, BDH04, Cal10, Cal13, Gri10, Tau86, Tau88, Vol02, ’t 97]. ultra [Chr22]. ultra-ether [Chr22]. ultrashort [Str19]. Umdenken [Pfe77]. Umgebungen [Sie10, Pes13]. Umkreis [Hei69a, Hei72]. Umlaufs [vB19]. Umsturz [Her78]. unable [Lin15]. unbounded [Kae48]. uncensored [Jac11]. Uncertain [Ano45a, BB05]. uncertainties [Gol98a]. Uncertainty [Bes04, Cas92, Kae31, Lin07, Lin08, Meh87, Cas09, CG14, LSF94, Lin06, Lin13, Mil90, Bla86, Fer07a, PCH77, For09]. uncited [Hen13]. uncitedness [EGR11]. Uncle [Dac20a, SL05a, SL91, Sta95c, SL05b]. Unconventional [Tel78]. Uncovered [Hür22]. Uncovering [MNT^{+}00]. unde [Ein19g]. Undergraduate [Pin00a, Myr21b, Myr21c, She16b, Sho21]. undergraduates [Lon84, Mil12b, Vog21]. Underlying [Mor79b]. Understand [Nic34, Pin15, Ein23h, Fra37, Jag01, MD15, PP03]. understandable [Vaa17]. understandably [Thi46]. Understanding [AMS04, Can16b, Cas86, DiS06, Die95, Ein18n, Ein20o, Ein21r, Ein22a, Ein48k, Eng17b, FG21, Fri89, Gib91b, Gol84c, Her17, Kar05, Krü00, Lin06, Lin13, Mer11, Oha12a, Oha12b, Plo06, Sar96, VB96, Wil16a, Bar06, Can15b, Diz05, Kak04, Kak05, Lem09, Liv13, Mer05, Mor00, Mor04, Myr21b, Sch17, Sch19, Sch20, Sch22a, Wei04, Whe68b]. understood [Ano22a, Fri12, Kan95a]. Underworld [Wei06b]. Undulationstheorie [Ein22g]. Unending [Ano45a]. UNESCO [Ano67]. Unexpected [Hol88e, Ano10d, Bri05c, LRH22]. Unexpectedly [GM06, Wac06]. unexplained [Mar11a]. Unfashionable [Dys16]. Unfinished [Smo19b, Bar00c, Dav95, Ren06]. Ungar [Wal02b]. ungewöhnlichen [Jun05b, YGKK05]. unhappy [Fre13]. Unholy [MM83]. Unification [Ano84c, Des79, FvN78, Hal07, Ne’06, VT99, vD02b, vD04, vD10a, Cat98, DM15, Ell05, Gio22, Sal05, Del12, Ken11]. Unificato [Mas02]. Unified [Ano31b, Ano07h, Cor04b, D’A01b, Ein28c, Goe04, Hal18, Jam30, Kol07, Lam30, Mas02, P.30, Sau14, Sau16, SS21b, Sci61, Asp80, CT24, Hal15, Oyi04, Par86, Sau06, Sau19, Ein25d, Ein29k, Ein29b, Ein29c, Ein30o, EM31a, EM32a, Ein38a, Ein53b, Str49]. uniform [NBDD^{+}12, Ein30f, Ein30a, EM30]. Uniformity [Foc55]. Uniformly [Wei15b]. Unify [Ber06b]. Unintelligible [Bro23]. Union [Hol94]. Unique [Jen01]. Unit [Ano81e, BNY81, GO81, Moo81a, Moo81c, Moo81b, Rus81]. unitaire [Ein30c]. Unitary [Pia29a, Pia29b, Ber82a, Ein30c]. Unite [Ano29b, Jam30]. United [Fre08, Har79d, Nad79, Sei79, Ano00g, De 92, Ein47c, Far13b, Gib96, HA62, HD69, HDHA89, Hol84, JDB^{+}97, Moo10, RN16, Smy45b, Smy45c, Smy45d, Smy45e, Smy76a, Smy78, Smy89, Szi45, WK12, Zuc77, Zuc96, Zuc18]. Unity [Ano12a, Del12, Ein35a, Mor79b, PA21, Ber82a, Sch00a, Sch86, Sch02c, Woi06]. Univ [Ano84a, Kuh54]. Univers [Nor21, Bar51a, Bla93, Chr22, Dol18, Kle74]. univers-bloc [Dol18]. Universal [Bar05f, Cao14, Ano15f, Duc11, Ein48e, Gás54, Tal82]. universalis [Ano15f]. Universalism [KJY24]. Universe [Acz99, Acz00, Ano22a, Ano49a, BOPS99, Bar51a, Bar05g, Bar09, BR21, Büc21, Cal76, Cli09b, Fac82, CF12, Cun22a, DM96, GP19, Got01, Got02, Gre99a, Gre99b, Gre00, GK05, Hal91, Hel09b, Hel09c, HBN00, Inf49, Jam30, Jea30a, Jea30b, Jea32, Jea38, Jea43, Jea48, Jea76, Jos95a, Kae31, Kir02, Kir13, Kra96a, Kra92, Lev07a, PP05, Ren05c, Sch12e, Sch07d, Smo14, Ste16, Tau79, TVL^{+}09a, TVL^{+}09b, Top07b, Whe92, YN11, You98, Abb05, Ano10c, Asi83b, B^{+}06, Bar48b, Bar50b, Bar50c, Bar52b, Bar57b, Bar80b, Bar82, Bar05a, BS83a, BPS87, BS93, Bar94a, Bar95, Bar00b, BB00, Bar02b, Bar05b, Bar06, Bla93, Bru03a, Cal79, Cal80, Cal82, Cal03b, Cal05b, Cho08, Chr22, Cle09, Coh31, DS01, Dan00, De 12, Dea13, Dol18, Egd14, EdS32, EK05, Ell89a]. universe [Ell90, Fer07b, Gat09, Gle10, Gre05c, Gui09, Gut97, Hai10, Har81, Har00, Har13b, Haw01, Hea15, Hol79a, Hon96, Ing05, Isa07, KT95, Kan95a, Ker89c, Kil94, Kle74, KE05, Kra07a, LD08, Lau05, Ler91, Lev15, Liv13, Mon58, Nor65, Nor90, NB09, Nus14, OM14, OONM17, OM22, Par86, Pen05, PR11, Pla59, Pri12a, PA06, Ree00, Ren05b, Row08, Sac10, Sei03a, Sen21, SN93, Sil30, Sin04b, Smi82, SA85, SA87, SA95, Spi05, SBF00, TFC93, Van85, Web99, YN10, YotY99, Zee89, Zee90, Zee01, vH97, Ano20a, Ano21a, Bal01a, Bar48a, Bar50a, Bar51b, Bar52a, Bar54, Bar55, Bar56, Bar57a, Bar58, Bar62, Bau09, Bre05, Bre08, Buc13, Cha88, Dys79, Gol95b, Gol98a, KK75, Lau55a, Nor21, Nor22a, Ran05, Rie12, Gor07a, Myr21a, Oha79]. Universe [Rus17, Sch07c, Smo07a, Weg08, Myr21b, Nah01]. Universelle [Gás54]. Universes [Gre11, Har07, Lev07b, Pic05, Gar39, Gar40, Gar42, Pan04, Sch56a, Sch10c, Vil06]. Universidad [Ano25b]. Universität [AH93, Ein20b, Ein21w, Ein22p, HY86, Kle75b, SM41]. Université [Ein76b]. universitete [Lan67a]. universities [Bra14b]. University [AH93, Alf09, Ano22a, Ano23a, Ano31f, Ano81a, ABdB^{+}53, BS81, Bar05f, Bed01a, Ber07a, Ber00, BB89, Boy88, Bru06, Cah90b, Cli09a, Dys15, Ein20b, Ein21w, Ein21l, EA22, Ein22p, EA23, Ein25f, EO79, Eng17b, EGO97, For01, Gor02, GRC18, Ham11a, Har02b, Hen82b, Hen86a, Hen88b, Hen00b, Hoc93, Hus15, JRG02, Kat04, Kil84b, Kle75b, Koj08, Kri08, L.89, Lan08, Lew79, Mar10, Mas01, Mil11a, Myr21b, Myr21a, Myr21c, Nav15, Noe11, Par06, Plo06, PJMS11, Rec86, RK94, SW94, Sci80, Smi02, Sta21, SM41, Tur06a, Van22, Vog21, War96b, War03a, Wei94, Wig79, vB02, She16b, Ano21j, Ano25b, Ano30a, Ano41, Car04a, Ein10b, Ein30g, Ein55i, Ein64b, Ein76b, HY86, Hav80, Lan65, Lan67a, Lau51, Lor31, Sch03d]. universo [De 12]. universos [Gar40, Gar42]. Universum [Bau09, Bre05, Bre08, Ing05, Bar48a, Bar50a, Bar51b, Bar52a, Bar54, Bar55, Bar56, Bar57a, Bar58, Bar62]. Universums [Ren05c, Ren05b, Ren05e]. univerzite [Těš12]. Univerzity [Hav80]. unjust [Wal77b, Wal92d, Wal06e, Wal15]. Unknown [SS65, Sto05, ABC^{+}06, AK97, Ein48n, Foc53c, RMI21a]. Unleashing [Sch17c]. Unlikely [Acz01, Acz03]. Unmaking [FGMvH14]. unmasked [Fos08]. Unmechanischer [Ser77]. Unmeltable [Ano96d]. Unobservability [AP00]. Unobservable [AP00]. unpack [Owe12]. Unpredictability [CW12]. Unpublished [Ber94, Ein08b, Fra81, GO08, Rol10, Bel96b, Ein31m, Ein31n, FLN13, Sto96]. Unraveling [Ghi05]. Unravelling [FR13]. Unreal [Vir92]. unrealized [Kir05]. unreasonable [Hol03a]. unregelmässig [Ein14h]. unregelmäßigen [vB19]. Unresolved [McC77]. Unschärfe [Bes04]. Unscientific [Foc53a]. Unseeable [Fle18]. unser [Ank15]. unsere [Inf53a, Inf56b, Inf57b]. unserer [B^{+}61, EL09, Ein09b, Ein09c, EdH15b, EL05a, Ger69, GH84, See60]. Unsöld [Kro81, NMA^{+}81, Uns81]. unsolved [Mar11a]. unspeakable [Bel87, Bel04, Lau97]. Untermeyer [Cor15b]. unternehmerischen [JDB^{+}97]. Unterredung [Joe19]. Unterricht [FTG08]. Untersuchung [Ein95, EH10a, EM31b, EH05b, Cun22b, Ein22o]. Untersuchungen [Els24, Sch09a, vE96]. Unterwegs [Ost15]. until [Sal95, Wei12s, Wei12t]. unto [Cas05b]. untold [Cla61, Gro67a, Gro67b, Gro68, Gro00, Hof10, Kie13, McK03b, Rob01, Rob14]. Untrained [GN11]. Untying [JR07]. Unusual [HTS^{+}88, Cah05, Jun05b, Rol47, YGKK05]. unusually [JRT^{+}91]. unveil [Pea90]. unveiled [Ban11, Ros12, SBF00, Wei15c]. Unveils [Ste16]. unverstandene [WW75]. unvollendete [Ren06]. unwanted [SSSS95]. unyielding [Bol04]. uomo [Inf63, Inf68, Inf74, Inf92]. Update [Ano18c, Wal90b, Ant15]. Updated [Mil13b]. upheaval [Her78]. upholds [Rei11a]. upon [Dam06, GRRS99b]. uppfinnaren [Fre13]. upright [Hei86, Hei00]. Uproar [Ein20p, Ano20e]. Uran [HS39a]. Urania [Ano21e, Ano21e]. Uranium [Acz09, Ano39a, Ano39g, Ano39h, Ano41, Bad96, Bla13, FP40a, FP40c, Git63, Hec12, Kae39, Kan02, Lib79, MF39a, MF39c, Mog02, NBDG40, Rin02, RK39, SMKW46, SL90, TT70, Zoe09a, Zoe09b, vHJK39a, vHJK39b, vHJK39c, Ada17, Ber96c, Eic94, HS39a, HS39c, HS39b, Hah42, HS44, Hah48, Hah62, MF39b, Per39]. Urankernes [HS39b]. Urans [HS39c, Hah42, HS44, Hah48]. Uranspaltung [GRE^{+}01, Hah62, HS39a]. Urge [Whi55a]. Urged [Res50]. Urgent [Ein46a]. Urknall [Sta95c, Ost15]. Urknall-Kosmologie [Ost15]. Ursache [Ein26c]. USA [Bes93, JDB^{+}97, Plo06, Sta99b, WM08, Woo06, Pat01a, Sto12]. USA/Großbritannien [JDB^{+}97]. Use [Ber93, BHJ^{+}00, Byr80a, Ein14h, Fre08, Ham85, Ham97, Lit09, Mal12, MM58, Mor57, Ove96, Pal95, Sig78, Smi58, VB96, Wal90b, Wil46, You95, Alp65, Alp85, Alp94, ATW^{+}95, Byr81, Hei46, Jay90, Lie78, Loa98, Mal08, Rus90, Sti47, Sti76, Wal97, Wal04a]. Use/Moss [BHJ^{+}00]. Used [Com46b, Kle74]. Uses [Ano39g, Lau59, TTHJ68, War92]. Using [Ano17c, Fow03, GS22, GM00, QPSD13, WL12, BIT03, CZPR17, Fla15, FTZ^{+}15, HMW24, Kia46, Kin95, LR14, McK03a, Owe12, Per12, Smy45c, Smy45d, Smy45e, Tör81]. Usual [Gol87]. Utah [Bla13, Ful84b, Ful84a, Mil08, Mil13b]. utazás [TB02]. uticaj [Inf79a]. utilisé [Kle74]. utility [Sta16]. utilization [Smo07c]. utopia [Bes93]. utter [Wal97, Wal04a]. utveckling [EI38d]. üzerine [Tip12].

V [Coh54, Ein11c, LSF94, SR99, Sin02, Sin04a, Ven88, Ano79h, Ein17d, Ein05c, Gar72, Hav77, Kol72, Kol74, Lan67a, Pát79a, Žáč12, Bro02, HH05, Rol19]. V-1 [Bro02]. V. [Ein11c, Inf53b]. vacation [HW08]. vacuums [Bar00b]. vadis [EDK05]. Vaihinger [SE79, SE79]. valeurs [Ein14j]. Valiant [Bok15, Ger15, Kle14a, Sto13]. validation [Mas05b]. Validity [Ein22g, Joh19, Ein05-32, Ein07i]. Valley [Ber05]. valore [Kop23b]. valuable [Lui11, Ryc14]. Value [Eng05b, GHK^{+}06, GHK^{+}07, GCO^{+}06, Kop23b, Wei06a]. Valued [Ano45b]. Values [Cao14, Ein14j, Ein86c, Ein87b, Eĭn85b, Eĭn86a]. Vanadium [Bla13]. vand [Hau49, Hau63, Hau64]. vandaag [Ano15f]. Vandervert [Spi96]. vanishing [Ein41a]. Vanity [You79]. Vannevar [Zac97, Zac99a]. vapnet [RE90]. Variable [CHSH69, CHSH70, Hei10, Hol05a, Bel96b, Nac77]. Variables [Bha86, Pat93]. Variation [Wei12y, AAA^{+}09, Eöt00]. Variational [Vec17]. Varicaks [Ein11c]. världsbild [Gar41]. Vast [Ano39g]. Vater [Mau96]. Vatican [Klo85]. Vault [Ree11c]. vazby [Těš12]. VCH [Vis95]. VDPG [Ano61]. Ve [T– D07, EI94, EI11, Tip12, Žáč11]. Veblen [Gan97]. Vector [Ano23a]. Vectors [Wal07, EM34]. Vehler [Bre05, Bre05]. Veil [How97b]. vejo [See61]. Vektoren [EM33a, EM34, EM32b, EM33b]. Velocities [HG06a, Hum36, Put25, Wei12n]. Velocity [Coh40, Dor44, HG06a, Hub29, Wal02b, Wei12y, Abi03, Car02, CW32a, CW32b, HS51, Lie78, Ung07]. velt [Sch27]. Velthuys [dP94]. Velthuys-Bechthold [dP94]. vendetta [Cra00b]. Vengeance [Hen98]. Veniamin [Dix00]. Venkata [SR04]. venner [RRK^{+}64]. venture [Dys16]. venues [Bad12]. Venus [MJ90]. verallgemeinerte [EG14]. verallgemeinerten [EG13a, EG13b, Ein14m]. Verallgemeinerung [vN37]. veramente [Lan67c]. Verantwortung [GRE^{+}01, Lam79, Wal03]. Verbesserung [Pla00a]. verborgen [HC94b, HC96b]. Verden [MW46b, Jak13]. Verdensbillede [EI39a, EI61a]. verdict [Sch10b]. Verdrängung [Wal03]. Verein [Ein18b]. verfilmt [Bre08]. Vergessen [Bre08]. vergnüglicher [LT15, Lin17]. Verhalten [Ein11g, Ein05i, Ein05q, HS39c, Ein11a]. Verhältnis [Els24, Hen86c, Her66a]. Verhältnisses [Ein06e, Ein05-36, Bra37]. Verhandlungen [MK56]. véridique [Lev04b]. verificão [NA92]. Vérification [Ano22e]. verificationism [Bro88]. verifications [Gin62]. verified [GTP22, Spo16]. verifies [Ano15j]. Verifying [Cha75, Cha76]. verket [Str83]. Verlag [Hen86a, Jüt21, Klo63, Pes13, Ros86, Wil36]. Verlagsanstalt [Hei53a]. Verlagsgesellschaft [BS92]. verlaten [Ano20b]. verlaufender [Ein14h]. verliehen [Ano79g]. verliehen/Broschüre [Ano79g]. verloren [Bre08]. Vermächtnis [Jun05b, YGKK05, Cle17]. Vermont [Wil13b]. vernunftwissenschaftlichen [FTG08]. veröffentlichten [Ein15c]. Verrier [Ros82c]. versäumt [Hof08b]. verschiedene [Eöt91]. verschmähte [Sch05e, Sch12f]. Version [Ano48c, GRE^{+}01, Ein21h, Hen85, KS93, WS78]. Versions [Jam74b, Jam74c]. Versprechen [RMHD11, Pfa12]. Verständigung [Ein33i]. verständlich [Vaa17]. verstandlichen [Ano22a]. Verständnis [Sch17, Sch19, Sch20, Sch22a]. verstehen [MD15]. versteht [Fra37]. Verstimmtes [Bre05]. Versuch [Ano22a, Cun22a, Ill89c, Zim84]. Versuche [BH01]. Versus [Duf46, Hel09a, Kar05, Ken05b, Max93a, Max93c, Max93b, Sch74, Zah87, Ano19d, Bol04, Fol96, Gré19c, Her82b, Hon89, Jan02, Mar97, Phi97, Ric06b, Rob49b, Sac88, Sel88, tH20, vM98]. Vertrieb [ARK^{+}99]. verwandelt [JDB^{+}97]. Verwandlung [Ein05d, Ein05-37]. Verwechselungen [Kra21]. Verwertung [Ein14h]. Very [Bai07b, Col13, Hoy54, Mal12, Ne’06, Blu05, Cho08, Clo04, Clo09, Jam06, Sta08b, W.25]. verzeih [Büh98a]. vesnoi [Lan67a]. vessel [How94a]. Vetenskapsakademien [Hil39]. veterans [KK05, LML88, Pal98]. Vexation [Par06, Tur06a, Hol05c]. vi [Bar05f, Bro66, Hen86a, Hou65, Mal65, See65b, Wre66, HH06, Tau63]. via [DSZ12, GLW15, Erc55]. Viagem [Ken22, Sou19]. Viagem/Travel [Ken22]. Vicious [Sta19, Ein47c]. Vicissitudes [Rol10]. Vicki [Hes04]. Victim [Ein33g]. Victims [SK82a, SK82b]. Victor [Jos95a]. Victorian [Hen11a, Leg15, Her01a, Moo81b, Pan02, Zim82]. victories [Fre19]. Victory [Ham97, Hol05c, VB96, Mad95b, Mad04, Wei95, Par06, Tur06a]. vida [Moo96b, T– D92]. vidas [HC96c]. Video [Wac06]. vie [Fra91, KT15, TGC91]. Vienna [BIT01, MR87b, MR87a, Ren08, SW12]. Vier [Ein22p]. vies [De 13]. View [Ayn34, Bia16a, Bri49, Ein32c, Ein34a, Ein34g, Ein34h, Ein35e, EI39a, Ein53j, EI64, ES05b, Eis22, Gut94, HS10, Her76c, Kan84, Mil75, Mil24, Rei82, See53, See55, See64, See80, See86, See91, SAY^{+}82, Ano03c, Bai07a, Ber74b, Ber74c, Boy10, Ein19e, Ein05-37, Fey66, Flü74, Gar41, Göt92, HSW81, Inf53b, Lon84, Lon03, Luk02, Lum11, MO13, Nor89, PA06, Ric83, RR89, Sac85, Stu14, Stu23, Whe79c, WE87, Yor70]. Viewed [Ano29c, Wen49]. Vieweg [Bäu83, Lin60]. viewpoint [Ein05d, EF13]. viewpoints [Bet91]. Views [Ano29f, Ano31c, Ano31d, ABM79, EBU^{+}52, Goe03, Gri79a, Gri79b, Jam30, Mos22, Sch87b, AVvdV99, Ant15, De 92, Ein09b, Ein09c, GP11, HBD71, Mon58, Moo81c]. vii [Hen86a, Tre88]. viii [Bar05f, Hen86a, Woo06, Woo28]. Viking [vB01]. világformáló [Har06a]. Vilenkin [MGG^{+}95]. Vimana [Bro02]. Vinci [Ano15f]. Vindicates [Wat99, Dum04]. vindicating [Hol30]. vindication [McH07]. Vintage [Ami04]. violation [Kob09, Mas09]. violence [Ash12]. violet [HBR99]. Violin [Gil05, Sch91, EE05, Ehl05a]. Violine [Ein34m]. Violinist [Kea12]. Violoncello [Ein34m]. Virginia [Bad15, Lac08]. virke [RRK^{+}64]. virkelighetens [EI39c]. Virtual [Sch54]. Virtues [vDP17, vD17]. Virtuous [vD17]. virus [Irv67b]. Viruses [Ano05b]. Visa [EBU^{+}52]. Viscosität [Ein20d]. Viscosity [Hug54, Mar40, OW54, Ein20d, MDT^{+}03]. Visibility [Ree06, ZCPB11a, ZCPB11b]. Vision [Ano39h, Ano84c, Kar05, Leg15, Nau17, Pep04c, Per21, Die05, Gle10, Kak04, Kak05, Par04, Wei04, Whe68a, Whe68b, Whe70, Whi16]. Visionär [Ste05a]. Visionary [Ste05a, Sev99]. Visions [Ano29d, Tom99, BK05, HBO82, HL10]. Visit [MA31, Ort95, Fre27a]. Visiting [CC96]. visits [Uhl80]. vistas [Ber84b]. Visual [WD96, Wes01]. Visualisierung [KRZ^{+}05]. Visualization [MW11, KRZ^{+}05]. visualize [Sti13c]. Visualizing [KF03]. vita [Arg09, Qui98]. vitae [Pla20a]. vital [Lee06]. vite [HC94c]. vitreuse [HLZJ91]. Vivekananda [Rei41, Rei41]. Vivre [Mil13a]. Vizgin [Pec04]. vjroii [Bič79a]. Vladimir [Pec04]. vliv [Inf58]. Voci [MFS^{+}6x]. voennykh [Smy46]. Voice [Ōe95a, Szi92, Szi05]. Voices [Ano47c, B.50, Cho16c, Rei50, Hal04a, Kao10, MFS^{+}6x, Ove87, RB05]. voids [Bar00b]. vois [Ein34a, Ein58, Ein89c, Ein09a, Ein15a, Eis37, Sar35]. Vol [Cah90b, Cah90a, Cas94b, Ein86c, Gol87, Hen88a, Ker89b, Kil95, Lin56, Loe88a, Loe88b, Mcc87, OSWR86, Ram90, Ree17c, Sch09a, Str90, Wig96m, Yan87, Ano99b, Ano99c, Ano05j, Bal91, Bal96, Gly90, Har96b, Hol05b, Kev05, Kil98, Sie97, Sta00, Wal99, War95, War96a]. Volgarizzazione [Ein21p]. vollständig [Ein02b, Ein27b, Ein05-39, Kie15]. Vollständigkeit [ST94b]. vols [Hen11c]. Volt [Ano39g]. Volts [Ano39b, Ano39g]. Volume [Bad15, Ber07a, BRI^{+}21b, CHR90, Ein49d, Ein70b, Eis00, Gre19a, Gre19b, Har91, Har96a, Hen86a, Hen88b, HA62, HD69, Hug54, Key89a, Lac08, Lau50, OW54, PvM93, Pye89, Pye95, Pye97, Pye06, Sta09a, Ste90, UM92, War96b, vMHW85, vM93, vM96, vM99b, vM01, vM05c, FW60, Fre79b, Fre79c, Fre79d, FK85, Hen82b, RR82, Rei72, RTR87, Row08, Tho31, Tho11a, Vog21, Coh80, CHS97, HvMW79, JNR^{+}07b, JRN^{+}07b, Pye80, Pye89, RSS^{+}07a, RSS^{+}07b, Tau63, UM86, vL19, vL21, vL23, vL52, vL56a, vL65, Hen88b, For80a].

Volumes [Ano07i, Ano07j, EKS^{+}08, JNR^{+}07a, KSB^{+}09, vH62]. vom [Ano69c, B^{+}79, Ein07i, Ein07d, EF13, Ein19e, EM33a, Ein05-32, EF05, Ein05-34, HH14, Jos95b, Ren05e, SE79, Stu14, Stu23, Tre66, Wie21, Hah62]. Vonzása [von90]. voor [Ano15f, EdH15c, Gui09]. voprosu [Kol40]. vor [ABH^{+}55, ABH^{+}87, Jüt21, Pla20a, Pla20b]. vorangehender [Ein12b, Ein05e]. Vorausseteungen [Häg46]. Vorbedingungen [FTG08]. vorbeigeht [vS21a, vS21b]. Vorgänge [Ein14h]. Vorgeschichte [GHK^{+}96, Kra88]. Vorlage [Tre66]. Vorlesung [ST94b, Ano20e]. Vorlesungen [Ein22p, Lor14, Wey18, Ein06d, Ein99]. Vorschlag [Ein26d]. Vorsstellungen [JP52]. vortex [Abr04c, SLLS09]. vortices [NBDD^{+}12, Sie91, Tom99, WNS^{+}08]. Vortrag [Ano21e, Ano21i, Jüt21, Pla20a, Wig64a, Ano11b, Dir60b]. Vorträge [Ein17d, Pau61a, Sim06, SM41, Tre66, Wil36, HSD34a, Her76c]. Vorurteile [Bäu83, Mel79b]. Vorwort [Ein16o]. vospominaniia [Jof62]. Vospominaniya [Inf56c]. Voß [MGG^{+}95]. vote [Lee93]. vous [Mon12]. Voyage [Zew02, Str04]. vriendelijk [Her11]. vs [Gio13a, Gio14, Gio22, Gwy02, Wie92]. Vstrechi [Jof62]. VT [Wac06]. Vucinich [Wei03]. Vulcan [Rus17, Lev15]. vulgarisation [Cun22a]. Vulnerable [Ano45a]. vuotta [Hot15]. vyd [Ein23h]. výklad [Ein23h]. vystupleniiu [Kol40]. vzpomínky [Inf59]. vztah [Kol72].

W [Bal72, Bar05f, BW^{+}87, Ein15c, Ein20d, Ein23a, Ein24b, For80a, HW78, JP52, Kar05, Ken25, Kra02, Lan15, MB^{+}92, Mur25, Pul11, Ree15d, Sch87a, Sch09a, Sci80, Sne82, Tay21, Tho97, Tre21, WB^{+}87, Wil36, Gol98a, Inf56a]. W. [Ein20d, Ein22q, Ein23a, Ein24b, Ein79e, Ein85e, Flo06, HW78, JP55]. wa [EI40a]. Wachsmann [Lut05]. Waddington [Adl03a]. Wade [Eds82]. Waged [Ein35f]. wahre [SE79]. Wahrheit [Her33, Her63a, Her67, Her75, Fra52, RS13]. Wahrscheinlichkeitsrechnung [EH10b, Ein15b, Ein05c, EH05c]. Wainstock [Ham97]. Waiting [Mil08, Wal81]. wake [Whi01]. Wald [PS16]. Walk [Web73, Gei11, Haw05, MP99b]. walked [Hol18]. Walker [Est06c, PS03]. Walks [Str04, Zew02]. Wall [EES49, Con02]. Wallace [Cre83]. Wallach [BR00, BR00]. Walter [And81, Gor07a, Sch07c, Smo07a, Weg08, Ein47f]. Walther [Ein22i, Kox06]. Wann [Bre05]. want [Joh14]. wanted [Hef10, Jag01]. wants [Ano20b]. War [Ada17, Ano20d, Ano45a, Bag10, BHJ^{+}00, BEI^{+}55a, BEI^{+}55b, Cas17, Con02, Ein15h, Ein16k, Ein33b, EF33, Ein33h, Ein45g, Ein47a, Ein53l, EFA72, Ein79c, Ein85c, Fei66, For73, GC96a, GC96b, Gor07b, GB88, Gro12, Gru05b, Hab82, Ham97, Kap20, Kea12, Kie13, Koj08, Lan08, MH09, Mel00, Mor95c, New88, Nic17, Nic18, Nic19, Par05, Ree17c, RFP^{+}45, RB05, Rol47, Rus90, RE63, Scl89, Sta03b, Sta19, Sus08, Wei95, Whi55a, Wil16b, Zoe09b, Amb97, Ano12b, Bla48, Bla49a, Cor03, ER91, Eva53, Far13a, Fei61, Gal93, Has07a, HH89, Jer02, Kah60, Kah69, KJ07, Mar17, Nye04, Pow93, Pow94b, Pow00, RS07f, Vol02, Wel06, Zoe09a, vB05, Bru03b, CR08, Del09a, Dre07, Fri12, Ger92, Ger97, Ger02, Ger07, Har03a, Hec16, Her80, Her82a, Her88]. War [Her89b, Hof10, Hol96a, How21, Joh07, LP05, Leo09, LSW12, Nye99, RB04, Sal19, Uda94, Uda98, Wel99, Wel00, Cli09a]. Warburg [Ein22f, Bre15, Ein22f]. waren [MH71]. warfare [Cha05c, MM83, Joh06]. Warm [Ano05-44, Mog02]. Wärme [Ein05c, Ein11g, Ein05-40, Ein05q, Ein05-35, Ein05p, Ein04, Ein07e, Ein13g, Ein14o]. Warmegleichgewicht [Ein25a, Bos24b]. Wärmegleichgewichtes [Ein02a, Ein05w]. Wärmestrahlung [Ein06d]. warn [ABH^{+}55, ABH^{+}87]. warnen [ABH^{+}55, ABH^{+}87]. Warner [Chu60, Chu98b, Chu01, WB01]. Warning [Lau55a, Bur45, Whi55a]. Warnow [For80a, Pye80]. Warped [Ran05]. Warps [Har95, Tho93, Tho94]. warrior [Con17, Rot79]. warriors [Tau12]. Wars [Aro96, Ham97, Koh95, Acz09, Du 52, Edg14, LE96, Wal77b, Wal92d, Wal06e, Wal15, Cor12, New89, Stu18, Vol02]. Warsaw [Ein44d, Ano68a]. wartime [CSW96, Dei95, WS88, WS08]. Warum [CSW97a, Ank15, EF33, EFA72]. Warwick [Kai05a]. Was [Ano33h, BT15, Bre08, Ein21g, Eis06, Gil08, Gil09, Gio13a, Goo90a, How93b, Moo16, Mus04, New09, Nor07e, Rai14, RF69, Sha92, Whi55a, Wil86, Wil90a, Wil90b, Wil93c, Ano55c, Ano15l, Bra02, Bru99, Buc20, Bun79b, CC74, CC98, Dob06, EI40a, Fer05b, HM11, HS14, Hes15a, Hof55, How90b, Jen02, KAS96, Kno90, Kov01, Lee93, Mad83, Mar17, Mau96, Mel00, Mui94, Nor85, RE90, Spo16, Sta93a, Wei13e, vB05, Ein19k, Fra37, Hes15b, KRZ^{+}05, MH71]. Washington [Hes04, Kar05, Vis94, Rol47, SS21b, SS01]. Wasn’t [Kri19c, Ner86]. Wasserstoffmolekuls [Ein20n]. Wasserwellen [Ein16f]. Waste [Ein39b, Ano96d, De 92, Ham08, HCD^{+}78]. wastelands [MHY95]. Water [Wei12b, Ber06a, Ein16f, Eis86, Hau48, Hau49, Hau53, Hau54, Hau63, Hau64, Hau65, Hau83, Wig86, XMY^{+}09]. watershed [RC13]. Watt [Fel11, Ham11a, Hen11a, Joh11b, Leg15, Mil11a, Nuv11, Set11, Ham11b, Hen11b, Hun10, Mil11b, Sch12d]. Watts [SS01]. Wave [Ein22g, Her49, Hie68, JSSZ87, Ken13, Kle64, Mic68, NM12, Prz63, Prz67, Rüg88, Bac89b, Bai92, Bai01, Blu22, Cho16c, Ein27b, GMHB02, Han79b, Hou16, Koh99, MR87b, MR87a, OP87, Prz86, Prz11, Sch28, Sch35a, Sch82a, Sch99a, Sch03b, Van07a]. Wave-particle [NM12, Bai92, Bai01, Van07a]. wavefront [Joh05b]. wavefunction [FTZ^{+}15]. Waves [Ano70, Bac89a, D’A75, Dys16, Ein18o, FS98, WW57, Web61, Web04, Ao04, Bar18a, BES86, Buc94, Buc20, CC84, Cho16a, Cho16b, Col04, Col17, Dir60b, Dir60a, Dir87, Ein16f, ER37, Grü08, Han92, Ken07, Ken14, Ket02, Ros37, Spo16, Tho18, Vig94, WTF81, Wei18b, dB49d, dB55a, SS97]. Way [Ano41, BRS^{+}12, Duf46, Erc55, JR07, Mal12, Ren07d, Wil69, Ein46h, Ein07h, KAS96, Mac08, Str94, Fer03, Pau93]. Wayne [Cou79b]. ways [Eis02, KT15, Tho15]. Wazeck [Hen16, Nav15]. WE-Heraeus-program [JDB^{+}97]. WE-Heraeus-Programm [JDB^{+}97]. weak [BRVF84, Sch88c, Vel00]. weak-interaction [BRVF84]. weaknes [DC08]. Weapon [Res50, RE90, Cla61, GHK^{+}96, Hen98, Her80, Her82a, Her88, RSU93, Vol02]. weaponizability [Ree17a]. weaponry [HW08]. Weapons [Fre08, Gar08, Kem12, Lee06, Mad95b, Mad04, May95, Nye99, Ree14a, Wei06b, Wel08, ABH^{+}55, ABH^{+}87, Bad95, Ber08, Bet91, Bro92, Bro02, Cir07, Cir08, De 92, DS83b, DSW99, DG03, Dre07, Fel79, Fel82a, GC02, Gib96, Gil62, Gla57, Gla62, GD77, GMG^{+}77, Gru05b, Hac94, Han88, Her89b, MHY95, Moo10, Ree15b, Sto12, Van04, Wee87, Yor87, VB96]. wear [Bur45]. web [Owe12]. Weber [Ram82]. Wechselseitige [Kie82]. Wechselverhältnis [Müh11]. Wedding [For14]. Week [Kae39]. weekly [How50]. Weg [uJR15]. Wege [Erc55, Mei63b, Wei15c]. Weggefährten [Mei05]. Weide [Pin00a]. Weigh [Wil13a]. Weighing [Ano33h]. weight [Eöt00, Fok65]. weightless [Vaa17]. Weil [Mau96]. Weimar [For71b, For73, For80b, FCKT11, Sch05a]. Weinberg [MGG^{+}95, Wig96j]. Weinheim [Ros86, Vis95]. Weinheim/Deerfield [Ros86]. Weintraub [Ham97]. Weird [Str11b]. Weise [GRE^{+}01]. Weiss [Bes93, MGG^{+}95]. Weiße [Str05b]. Weisskopf [HEB^{+}80]. Weiterbildung [JDB^{+}97]. weiterer [HS39a]. Weizsäcker [GRE^{+}01, HH14, vS14, Sch14a, vS14]. welche [Ein19b, Ein19c, EM31b]. welchem [vS21a, vS21b]. Welcome [Ano63b, KZS99]. welfare [Rol47]. Well [Top07b, Gin04, Ham08]. Well-Nigh [Top07b]. Welle [Bac89b]. Wellen [SS97]. Wellenmechanik [Ein27b, Prz63, Sch34, Sch35a, Kil65, Klo63]. Welt [Bre08, Bod03, Inf53a, Inf56b, Inf57b]. Weltall [Büc21, Nor22a]. weltanschaulichen [Müh11]. Weltanschauung [Spe68]. Weltansicht [Stu14, Stu23]. Weltbild [Flü74, Her78, Her76c, Luk02, Ein34g, Ein34h, Ein35d, Ein53j, ES05b, Esp55, See53, See55, See64, See80, See86, See91]. Weltgeometrie [Inf31]. Weltkörpers [vS21a, vS21b]. Weltordnung [NN75b]. Weltrekordmarken [JDB^{+}97]. Weltrekordmarken/Trägheitsfusion [JDB^{+}97]. Weltsystem [Mos21]. Weltuntergang [NN75b]. Weltweise [Her94a, Her04]. Welty [For14]. Wen [SH01]. Wendover [Pos11]. Wenske [MGG^{+}95]. wer [vB05, Ein52e]. werden [EM31b, Flü39, Kie15, CSW97a]. We’re [SL05a, Ano19b, Ein54c, Ber83, Bey49, Gio13b, MH71, Nic87, Pit16, Wir08]. Werk [Hah88, Inf53a, Inf56b, Inf57b, Kie94, MGG^{+}95, Pfe77, Ren05e, See60, Cur38b, Hei55]. Werk/Weinberg [MGG^{+}95]. Werke [BDR93]. Werken [Sch10a]. Werner [Adl03a, Bes04, Bro72, GRE^{+}01, Hen11c, Mar73, MGG^{+}95, Pfe77, Rec02, Sch05c, vW02, vWR04, Alt06a, Arg09, Bes04, Bla86, BDR93, B^{+}61, BW04, Cap01, Cas76, CB84, Cas92, Dör05, D^{+}92, Eck01, Erm89, Fis02, GA02, Hei35, HBD71, HR02, Her76a, Her76b, Her84, Her01b, Hol00b, Kan02, KW93, KRW05, Meh87, PLS02, Pfe77, RW02, Rec02, RKW05, Rec09, Rec10, RH02, Sch05c, Tel76, Wal92c, vW02, vWR04]. Wertheimer [Mil75]. Wesen [Ein09b, Ein09c]. wesentliche [Ein19j]. Wesley [Kai81]. West [Ein48g, HF98, Rin02, FH11, Mog02, War99]. Wester [BB89]. Western [AFG90, Kra82b, Zho85]. Westerners [Sch17c]. Westgren [Kat52]. wetenschappen [Ano15f]. Weyl [Del13, Ein18h, Pes13, Sie10, Som05a, Som05b, A^{+}09, De 24, Ehl09, Pes13, Sie10, Som05a, WS^{+}01].

Weyland [Kle93, Kle93]. Whale [Ano00c]. Wheeler [Hal17]. Wheland [Par99]. Where [Ano19b, Car89, Dac20a, Dor86, Mad83, Mas95, Pla32, Pla33, Pla77, Pla81, Sei03b, Hoy94, Jak13, Mar84, Pla59]. which [Bjø19, Ein10e, Ein19b, Ein19c, EM31b, How94a, Lui11, Wei13f]. while [Roe93]. Whistle [Ste95b]. Whistle-Blower [Ste95b]. Whitacher [Dow80]. Whitaker [Nau17]. White [Jun14, VWA95, BAM99, Joh05b, Sri99, Edg14, Ein47f, Hie84, Ski04, Str05b]. Whitehead [Adl03a, Edd24, Mur25, Tay21]. Whitrow [Whi67]. Whittaker [Nav21]. Whittlesey [Duf46]. Who [Ano45b, Ano15l, Ber85, Bey06, Bra02, Cas05b, Duf46, Ein52e, HM90, Joh06, Jos95a, McC84, MP99a, Opf78, RF69, Reg87, SS86, SS91, SS96, Sha14, SO02, TP90, Ano79e, BB13, Bun79b, Cam18, Cha05c, Cli87, Cou13, Das91, Del09b, Dem15, Du 52, Gre05b, Har06b, HEWW15, Jak13, Kie13, Lee93, Mah03, Mos68, Mos87a, Mos87b, Rob50, Sch94, Sow01, TFC93, vB05, Dem15]. Whole [Ano35, Hei69a, Hei72]. Whom [Kuc11]. Whorf [Gor22]. whose [DS01]. Wichita [KK05]. wicked [Bai07a]. Widely [Ano39a]. wider [Bäu83, Mel79b]. Widerstreit [Hof08c]. width [EM23]. Wie [Ank15, Whe68b, BT15, Bre08, Ein52e, Smi15b, Ein21u, Ein22r, Ren05g]. Wiederkehr [NHP^{+}79, SKD^{+}87]. Wieman [Lev01a]. Wien [Klo63, SW12, HG06a, Ill89a, Ill89c, Pla00a]. Wienschen [Pla00a]. Wiesner [SMM^{+}60, SBSW60]. Wife [Ano90a, Ein95a, Est06b, Get06, Hil03, Kni12a, Kni12b, Mar05, Mar07, SH95, Ste22, ECS19, Gab95, Mar04a, Mar04b, Mar11d, Pop03, Smi04, Tou05, Est06a]. Wigner [BH89b, HH04b, Wig67b, Wig96m, Wig96j, Wol67]. wijsheid [vE95]. wil [Ano20b]. Wild [LMOS15, Mah17]. wilde [LMOS15]. Wiley [Bro66, Sta99b, Wre66]. Wilhelm [GC20, HSW09, Kan96a, MGG^{+}95, vBL96a, vBL96b, DFB22, HG06a, Rei86]. Wilhelm/Max [vBL96b]. Wilhelmian [Pye82b]. Will [Ano54b, Cao07, Ein39b, Pal95, Dor86, Kra15a]. Willem [Gui18, Sta20]. William [Sch87a, Ano88a, Ano88b, Cre83, Ein53f, Hom05, Jen01, Lem09]. Williams [C.51b, You95, Wil13b]. Williamson [You95]. Willy [Ill89c, Ill89a, Ill89c]. Wilson [AA75, Bot13, Cli09a, BW^{+}87, Cec01, Lan00, MB^{+}92, Tho97, WB^{+}87]. Win [All02, Car15, Eds82, Ano45b, Bus13b, Car12, Kie13, Wad81b]. Wind [Div78, Hoy94]. Winding [DSZ12, Wil13b]. Window [Mlo01]. Windows [GRE^{+}01]. Winner [Ano34b, Ano68b, Hes04, dP94, Ano65c, Ano05-38, Bal75b, Enz05, HD02, HL10, Joh11a, KS81, Myr21b, Myr21a, Myr21c, SK09, Wol67, Wyd15]. Winners [Bet63b, Coh54, CCD^{+}52, Ram82, Sti13a, Sty94, AA02, AKD23, Ano16b, BTR08, B^{+}01, BW^{+}87, Erc55, Far53, Far63, FK68, GW10, GW83, Hac98, Har13c, dVNHdVH53, dVNHdVH71, HR17, LR14, Lud57, MSYW12, Mag89, MFS^{+}6x, MB^{+}92, NN01, Pra12, San15, SS86, SS91, SS96, SO02, Sou66, Ste53, Tho88a, Tho97, Van04, WB^{+}87, WL80, Wor08]. Winning [DWL98, O’S01, BNSY98, CMT18, Das00, Doh06, Her80, Her82a, Her88, LRH22, LHDH21, LLH22, RWL23, RF05, SBB16]. Wins [Wat97b]. Winstead [You95]. Winter [Ein10b]. Winternitz [Ein24c]. Wintersemester [Ein10b]. wir [BT15]. wireless [GMOP12]. Wirken [EGO97, Hen11c, Kra81, Rec09, Rec10, Ros86, Sch09a, Ber69b, Cur38c, Cur38d, Cur38e]. wirklich [Ein21g]. wirkliche [Jun05a, Row06a, Row06b, Wue05]. Wirklichkeit [MH71, Ein48a, Hel98, vG30]. Wirkung [Fre85, Pfe77, Spe68]. Wirkungen [GHK^{+}96]. Wisconsin [For01]. Wisdom [Ein52f, Hor09, Hor00, Pra12, vE95]. Wise [Ano19b, Ano05-45, Her94a, Her04]. Wiser [BHJ^{+}00]. wisest [Pra12]. wish [BB13]. Wishing [Cle08]. Wiss [Pla20a]. Wissenschaft [BS92, Bäu83, B^{+}01, Ein21d, ER13, Erb10, Hei75b, Hen11c, Hof08c, MGG^{+}95, NS88, Rec09, Rec10, Sch12b, APWB95, Ein30k, EMvL79, FS06, Hei77, Hes15b, Kra88, Möl84, Rei41, Sch12a]. Wissenschaft/Cassidy [MGG^{+}95]. Wissenschaften [Jüt21, Kie82, Pla20b, Ren05e, Sch88b, Sim06, Sch88a]. Wissenschaftler [Lam79]. Wissenschaftlers [GRE^{+}01]. Wissenschaftlers/Dürr [GRE^{+}01]. Wissenschaftliche [BS92, BDR93, RS13, Ano69c, Ber69b, BHJ^{+}00, ER13, Fla15, Hah62, Hen11c, Rec09, Rec10, Wey20, ER13]. wissenschaftlichen [Ein21c, HRP90, Spe68]. Wissenschaftlicher [ARK^{+}99, Eis00, GRE^{+}01, HEB^{+}80, HvMW79, MGG^{+}95, OSWR86, vMHW85, vM93, vM96, vM99b, vM01, vM05c, Hün18, PHvMW93, PHvMW01a, PvM93, PHvMW01b]. wissenschaftliches [Hei55]. Wissenschaftsakademie [LMOS15]. Wissenschaftsgeschichte [Mar17]. wissenschaftshistorischen [Han18]. Wissenschaftskrimi [Mar17]. Wissenschaftssystem [SE79]. Witche [Bar93, BL11]. withdrawal [Pat85]. within [Bar88, Bar90, Roe93, Sil36]. Without [Ano21d, BF16, Hel15, Woo06, You05, Asp69, Ein20g, Ein52e, Far05c, For71a, FKZ91, GHSZ90, HG06a, Kia46, Lud13, Mac08, Mar17, Mil49b, Sch61, Wes80, È03]. witness [Hir98, SSSS95, ZCPB11a, ZCPB11b]. witnesses [Fre81a]. Witze [Hes15b]. wives [Nes14]. Wizards [Kap91a]. Woit [Ber06b]. Wolfgang [Ano45b, Ano45c, Ano93b, ARK^{+}99, Eis00, Enz05, HEB^{+}80, MGG^{+}95, Sch09a, Eis00, EM88, EvM94, EGO97, Enz02, FW60, Gie05, HvMW79, Lau88, Lau97, PHvMW93, Uns00, vEA02, vM00, vMHW85, vM93, vM96, vM99b, vM01, vM05c]. Wolfsröte [Bre05]. Woman [Hol07, Ste89, TP90, Bal75b, Sre29, Wat66, Wat70, Wat88]. Women [Blu89, Bro21, Chi08, Cla97, MB94, Moo17, Opf78, Sti13a, Wat00, Cal17, Das91, Fis92, Föl91, FFF09, Gab95, HHW99, Kie13, McG93, McG98, RCRC97, VWA95, WS88, WS08, Chi94]. Won [Ein45g, New09, Opf78, Cat04, Das91, MSYW12]. Wonder [Fie07, WW75]. wonderful [Mor95b]. Woolf [Gol84b, Fer82b]. Wootters [Kal92]. Word [Day12, Ein05-45, Ano15f, Tou14]. worden [vC05b]. Words [Ano15b, Ano05-45, Ano05-44, Dur02, Ein93c, Hor00, Kel09]. wore [Ank15]. Work [Ano29b, Ano54b, BB04b, C.51a, Ehl79, Ein05m, GW96, Hol68a, Kva79, Lau51, McC57, Nag50, Nav94, Nis79, Rec10, Ree17d, Ren05e, RFP^{+}45, Rüg88, Sch05b, Sch07b, Sch50a, See60, Ste98, WG96, Zwe80, dB49b, Ano50a, Ano97a, BNSY98, Ber62, BN97, BB04a, Ber69b, Bro06, Bro07, Büh15, Bur45, Coo83, Cur38b, Cur38c, Cur38d, Cur38e, Ein09a, EL09, Ein11e, Ein11f, Ein15c, EdH15b, Ein18k, Ein22c, Ein22d, Ein23b, Ein23f, Ein86d, vM90, Ein05g, Ein05h, Ein05l, Ein05i, EL05a, Fre86, Fre96, Gab95, Hah88, Hav80, Hei46, Hei55, Hei74c, vI11a, vI11b, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58, Inf60, Inf61b, Inf61a, Inf61c, Inf79a, Inf83, Jor69, Kie94, Kle82b, Kra81, Leh19, Lui11, MB21, Nye72, Olw04, Pfe77, Sch05c]. work [SBB16, Shu17, Sin04a, Sta93c, Str83, Tho17, WS^{+}01, YCL20, Zee89, Zee90, Zee01, dB49d, dB49c, dB55a]. Worked [Ano68a, Phe21]. Worker [Mul99, Ano05-39]. workers [Olw97, RRK^{+}64]. Working [CF82, San95, Sau19]. Works [Ano21d, Ano82c, Ein05j, Hor79b, Rai14, Sch49e, Tau63, Ano48b, Bet97, BDR93, Bor63a, Dic79a, DD95, EGO97, FS87, Fey00, Haw02c, Haw07, KV11, LME23, PLS02, Pla59, RRK^{+}64, Sch10a, VM05a, Wig96m]. Workshop [EP06]. World [AVvdV99, Ano19b, Ano45a, Ano55d, Ano06k, Ano09g, Arr06, Ayn34, Bey06, Bod05a, Bod05b, Bri66, Bro66, Bru09, Cah04, Čap71a, Cas05b, Cau91, Col09, Col93, Col68, Coo04, Ded85, Del09b, Duf46, Edd28, Eds82, Ein34a, Ein34b, Ein34g, Ein34h, Ein35e, EI39a, Ein48i, Ein53j, ES05b, FS95, FC07, Fre08, Fri12, Gil12, GPSR01, GR63, GR65, Ham97, Har03a, Her76c, HA62, Hür22, Jon89, Kri19c, Kuh54, LSW12, Mah17, Mos21, NN75b, New60, Ray95, Ree15d, Rei51, RS11b, Roh87, See65a, See53, See55, See61, See64, See80, See86, See91, Sha54, She15, Sta99c, Stu18, Sus08, VFIS63, Vog20, Wre66, You05, Zal91, Zoe09b, AFG90, ABW99, Ano51b, Arg09, Ari03, Ari05, Bar88, Bar90, Ber74b, Ber74c, BMS12, Ber07c, Ber09, Ber11b, Bla93]. world [Bod00, Bod03, Bre08, Bro06, Bro07, Cas95, Cas04, Cha05a, Chr22, Cla61, Cra07, Dra53, Dra01, Edd30a, Ein34n, Ein40h, EH41, Ein49e, Ein58, Ein59f, Ein89c, Ein00d, Ein06c, Ein09a, Ein15a, Eis37, Esp34, Esp55, Fay10, Fer57, FF91, Flü74, Gar41, Gol05, GMG^{+}77, Göt92, GMOP12, Gre05b, Gri84b, Gui95, Hal09, Haw11, HGS87, Her78, HW08, Hol03a, Hol96a, Hou65, Inf31, Inf34, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58, Inf60, Inf61b, Inf61a, Inf61c, Inf63, Inf68, Inf74, Inf79a, Inf83, Inf92, Jer02, JDB^{+}97, Kon15, Lan65, Lan67a, Lau15, LW71, Luk02, Lum11, Mal65, Mar91, MW46a, MW46b, MW07, McC01a, aM07, Moo02, MT98, Mos68, Mos22, Mul10, New56, Nor12, Pai86, Phi74, PP07, PP06, Rei34, Roc03]. world [Rod22, RSU93, Rot08, See65b, She75, She87, She03, Ste99, Ste01, Stu14, Stu23, Sty00, SMM^{+}60, SBSW60, Wad81b, Wei13f, Wel14, Wel76, Wel88, Wig82, W^{+}87, WdC02, Wil11a, Wor10, Zoe09a, de 03, Ano82d, Cas17, Con02, Ein48g, Ein79c, Fei66, For73, GC96a, GC96b, Gor07b, Gru05b, Hen00b, Kie13, MH09, MFHR06, Mel00, Nag50, RB05, Rol47, Rus90, Sch27, See25, Sta19, Wei95, Wil16b, Ano50a, C.51a, Sch50a, Woo06]. world. [Bur45]. worldly [Her94a, Her04]. Worlds [GRRS99a, GRRS99b, Gri10, Lei08, LB90, Ran05, Vil06]. Worldview [Hir76, EI61a, MSS22, Esp55]. worldviews [LRS12]. Worldwide [Gol82].

Worries [How90b]. Worry [Ano19b]. worrying [KSG63]. worse [LOvT12]. Wörterbuch [MGG^{+}95]. Would [Ano45a, Car15, Ano04a, Car12, KRZ^{+}05, Mau90]. Would-be [Car15, Car12]. Wounds [Sta03b]. Wranglers [Har85]. Wreck [Ano39h]. Wrinkles [SD93]. Write [Ano33c, Gei15, Nye19, Bur45]. Writer [Bar80a, Gla00]. Writing [Min84, Pit17, Rei11b, Wei12i]. Writings [Ano60, Ano99b, Ano99c, Bad15, BH89a, BH93, BIRS12a, BIRS12b, BIR^{+}15a, BIR^{+}15b, BIJ^{+}18, BRI^{+}21a, Cah90b, Cah90a, Duk49, EvM94, Har91, Har96a, JSI^{+}02b, Key89a, Kil95, KKRS93, KKRS95, Kno97, Mar07, Pye95, Pye97, Pye06, Ram90, Rei06, Sie97, SCRS89, Str90, War95, Bal96, BL96, BSW22, Bro19, Bro20b, CB84, EM^{+}91, GW06, Gre04a, JSI^{+}02a, Kli62b, KKS96, Shi49, Sta00, Wal99, WB01, W^{+}87, WdC02]. Written [Kea12, RMI21b, Ano01a, Kat52, Kle75a, MSYW12, RMI21a, ZXLX14]. Wrong [Ano91, Ber06b, Hol05a, Kra15b, Mei62, Mei63a, Mei63b, Mil12b, RS11b, Cha17, Fri11, Woi06]. Wrote [McC84, Mof10]. wspomnienia [Inf56d]. Wszechsświata [Gol98a]. wszechświata [Lan67b]. Wu [Seg86b]. Wuensch [Row06a, Row06b, Jun05a]. Wunder [WW75]. Wunderkind [Wei12d]. wurde [Ren05e, Ein21u]. Würdigung [Zim84]. Wurzeln [Hof97b]. WWII [San95]. wzgledności [EI59, EI62a, EI98b].

X [Coo04, Gia80, Hen00b, Red84b, Bar05f, Duf46, Jam70, Kae38, Mar10, Mil11a, Per10, PJMS11, Cul80, Ein18j, Gia03, Har79b, Kae39, Kar05, Key89a, Seg80, Seg83, Seg84, Seg87, Seg07a, Shi75]. X-1 [Shi75]. X-Ray [Gia80, Cul80, Gia03, Har79b]. X-Rays [Kae39, Ein18j, Seg80, Seg84, Seg87, Seg07a]. xi [Bro72, Bru06, Hen86a, Hes04, Jan73, Pri99, Sta21, War03a]. xian [Seg86b]. Xiaobo [HL10]. xii [Gly90, Hea64, Hoc93, She15, WM08, Wei94]. xiii [Alf09]. xiv [Bed01a, Boy88, Hen82b, Noe11, Par06, The12, Tur06a, Vis95]. xix [Har02b]. xlii [Wal02b]. xlvii [Hen86a]. Xu [Bod13]. xv [Bar05f, Bru03c, Fri73, Har79d, Plo06, Tur05, vB01, vB02]. xvi [Cah90b, Gil12, Hus15, Red84b]. XVII [Ros86, Ano84a, Bar05f, Goo85, War96b]. xviii [For01, Hea13, Sci80, Vis94]. xx [Hen82b, Hie68, Ron07]. xxi [Hen11c, Gin04]. xxii [Gor07a]. xxix [Koj08]. XXVIII [Kai81]. xxxii [Kil84b]. xxxiv [Dix00]. xxxvi [Cah90b].

Y. [Ano83, Bad83a, Fer83, Gil84a, Kil84a, Lar84a, Sch84b, Wil84]. Ya [Jos95a]. Yakov [Fre96]. Yale [Dys15]. Yamahata [YJ95]. Yang [BT12, LY95, HF05]. yanılgısı [Gol97]. Yankee [Rol89]. Yaqui [Rai14]. Yasushi [NKY15]. Year [Ano93a, Ano05d, FR13, MFHR06, Sha79b, Ano05-41, BBM78, Coh79, Eds82, Gen22, Lon84, Mao07, Pan05, PCH77, RC13, San07, SR99, Sta98a, Sta05b, Wad81b, Wad07]. Years [AH93, AGMT12, Ano29b, Ano05b, ARK^{+}99, Bad15, Bar05f, BH87, BH89a, BH93, Ber07a, Ber56, Ber88, BLR15, BFMNZ91, BSKS04a, BSKS04b, BSR^{+}06a, BSR^{+}06b, BRS^{+}09a, BRS^{+}09b, BIRS12a, BIRS12b, BIR^{+}15a, BIR^{+}15b, BIJ^{+}18, BRI^{+}21a, BRI^{+}21b, Cah90b, Cah90a, CDS01, Dir82, Ein40e, Ein05-45, Gam85, Gar08, Giu05, Gol87, Gru97, Har91, Har96a, Hen88a, Hen88b, HG05a, Hol80, HF05, JSI^{+}02b, Ker89b, Kil95, KKRS93, KKRS95, Kno97, Kov01, Lac08, Lau50, Loe88a, Loe88b, McC68, Mcc87, Mer55, Oha79, PLS02, Pye89, Pye95, Pye97, Pye06, Ram90, Rob39a, Rob39b, Rob15b, Row00, SKJI98a, SKJI98c, SKJI98b, Sch21, Sha92, Smi89b, Smi89a, SCS87, SCRS89, Str90, Tau79, VB96, War96b, Wig96l, Wol67, Yan87, Ano23c, Ano99b, Ano99c, Ano01a, Ano07m, Bal91, Bal96]. years [BRVF84, BL14, BGG^{+}07, Bun88, Bun90, CGR^{+}97, Con62, Cor15a, CW02, Dac11, Dea13, Ein50o, Ein50p, Ein52a, Ein56h, Ein70a, Ein71, Ein89d, Ein93c, Erb10, Ern92, Gam66a, Gam66b, Gam72, Gib91a, Gil19, Gla06, GGW94, GMW80, GMOP12, Gre15a, GR17, Har96b, HI87, HSW81, Hei49a, Hei53a, Hel80, Hen03, HW78, HHMW93, Hoi01, Hol05b, Hot15, HS00, HZ01, JSI^{+}02a, Jon17, Kae48, Kar10, Kev05, Kil98, KKS93, KKS96, Ber84a, LR01, LW71, LM95, Lin15, LE84, Mad95b, Mil90, MO95, OONM18, Pal98, Paz10, Poo90, Pra06, Qui05, RWL23, Rob05, RMR96, RSPS49, Sch04b, Sie97, Som49a, Sop88, Spo16, Sta00, Sza92a, Sza92b, Tom10, Wal99, War95, War96a, Wig80, WA03, vD14, vL55b, C.51b, Cou52, Inf50b]. Yeats [Ano07s, BL07c]. Yehuda [Gol84b, Kil84b, Pye80]. Yellow [Pas10, AVvdV99]. Yellowcake [Ano18d]. Yesterday [How21, Ano25b, Far05c]. Yiddish [Gru07, Sch27]. Yields [Ano18d, Mal85, The46]. yleisestä [Hot15]. York [Ano50a, Bar05f, Bla76a, Bla76b, Bru03c, Cli09a, Cor15a, Dea13, Duf46, Eis00, Goo85, Gor07a, Har02b, Har79d, Hea64, Hei53a, Hen86a, Hie68, Hus15, Jam70, Kae38, Kar05, Kra02, Lan15, Lin60, Nah01, Nau17, Ree17d, Rus17, Sch07c, Sci80, Sho21, Sta99b, Tur05, Woo06, Woo28, vB01, Ano39h, Dys76]. Yosuke [YJ95]. Young [Ari14, Ein54c, Fin76, Sch90, Bra15, Ham86, Pea49, Pye85, Roe93]. You’re [FLH85]. Yourgrau [Woo06, Jun05b]. yours [Dob06]. Youth [Loe88b, Tal32a, Bra14b, Rec10]. You’ve [Ree17d, Shu17]. Yuan [Ano87b]. Yuri [Pec04]. Yurii [Hoh13].

z [Hav80, Hom04, FK04, Sta02]. Zachary [Pin00a]. Zackheim [Ivr99]. Zahar [Fey74, Fey80]. Zangger [Ano05g, Med94]. zarubezhnykh [Jof62]. Zbornik [Mag95]. Zealand [B^{+}06, Pri12b]. Zeeman [dP94]. Ze’ev [Bad15, Lac08]. zehn [Ano23c]. Zeichnen [GRE^{+}01]. zeilen [dGK11]. Zeilinger [Fis10, Fis12]. Zeit [BHJ^{+}00, Cun22a, Ein18h, Ein30e, MD15, Sch17, Sch19, Sch20, Sch22a, SB06b, See56b, Wey18, WS^{+}01, B^{+}61, Foc54, For80a, Fra49b, Ger69, GH84, Inf99, Rei24, Rei27, Rei28a, Rei28b, RK77, See60]. Zeit-Lehre [Rei24, Rei27, Rei28a, Rei28b, RK77]. Zeit/Dyson [BHJ^{+}00]. Zeitalter [ES68, For70b, Sal95, Win69]. Zeitbegriff [Ehl05b]. Zeitgenossen [Hen90, Pye93]. Zeitungen [Ein52e]. Zen [Pri10]. Zeno [Pri99, Hug99]. Zentrale [Cle17]. zentrisch [EG23]. Zero [Ano94a, BMD^{+}11, ES13, ES05a, Sha99, Gal93, KZS99, Lau46, Lau72, SK82a, SK82b]. Zerplatzen [HS39b]. Zewail [Str04]. zhizn [HC98]. zhu [HL10]. Zijn [Hoe18]. Zinaida [Dix00]. Zink [B^{+}79]. Zionism [Koj08, Lan08, Ash12, Ein31c, Ein38b, Hof79a, RS07f, Sch87b]. Zionist [Ein21u, Ein21u, Kat04, Ros11]. Zivot [Mil10a, Mil12a, T– D07]. zizn [CH98]. zone [Ano08c, Das12]. ztracený [Kot93]. Zuckerman [Har79d, Nad79, Sei79]. Zug [Str97b]. Zukunft [Jor69]. Zum [Fre19, Hen86c, Jac08, Wie06, Ano79g, B^{+}79, Cle17, Ein16e, Ein22d, EE22, Ein05h, Fis10, Fis12, For80a, FTG08, HBD71, vI11b, JDB^{+}97, Kra88, MTVB88, Rec02, RH02, SS97, Sic09, Som49b, Spe79, Tre79a, vEPF09, Ano79h, Ein09d, Ein11c, Ein13g, Ein13f, Ein14o, Ein14p, Ein17h, Ein24k, Ein31o, Ein79g, Her63b, Kra20c, Rit09b, RE09, Sti05]. Zur [Aga06, Aga07, Ein23b, Lam79, Ost15, Sch17, Spe68, Tes69, BH01, Ein02b, Ein06e, EL08a, Ein08a, Ein14d, Ein14h, Ein14m, Ein17e, Ein18m, Ein20d, Ein22m, Ein27c, Ein30i, Ein40c, Ein53d, EI61b, EL05b, Ein05-36, Ein05-27, Ein05-39, Els24, Fis02, Fra49d, GRE^{+}01, Hah62, Han18, Hei46, Hen86c, HRP90, Hof18, Isa85, Jun05a, Kie94, Kil65, Klo63, Lan21, Mei63b, Pau55, Prz63, Rec02, Ren97b, RD05, Ren05e, uJR15, Row06a, Row06b, Sch19, Sch20, Sch22a, Sch09a, SKD^{+}87, SM41, Urb22, Voi86, Wei15c, Wue05, vEPF22, BHJ62, Cas21, Ein04, Ein05e, Ein06b, Ein06f, Ein12j, Ein14q, Ein15j, Ein16p, Ein17i, Ein20q, Ein21v, Ein21w, Ein23j, Ein23k, EE23, Ein24l, Ein29k, Ein30r, Ein33i, Ein05-40, Ein05-41, Ein05-42, Ein05-43, Ein05-44, Fra20, Pla00b, Wey23]. Zur [vS06]. Zürcher [RK86]. Zurek [Kal92]. Zurich [Ein10b, Pes13, Ein10b, EGO97, Jos79, Vit15, JNR^{+}07b, JRN^{+}07b, Kan89, MR87b, MR87a, RS96, RK86, Vit15, Wei12c, Wyd15, EGO97, Ano07b, JRN^{+}07a, JNR^{+}07a, JR07, RS03b]. Züricher [RS96]. zusammengelegt [MTVB88]. zusammengelegt/Stuttgarter [MTVB88]. Zusammenhang [Pau25]. Zusammenhänge [JP52, JP52]. Zusammenhanges [Ein27f]. Zusammenstellung [Bäc19, Bäc21]. Zustandes [Ein05-29, Ein10a]. Zuwanderung [Ein19d]. zvlástni [Ein23h]. Zwang [Ser77]. Zwei [Bes04, Ein17d, Jun05a, Row06a, Row06b, Wue05, Des72, Dür62, Dür80, EM30, Gut15]. zweien [Ein22b]. Zweipersonenspiele [BHJ^{+}00]. Zweipersonenspiele/Adobe [BHJ^{+}00]. Zweite [Kra23b]. zweiten [Ein02a, Ein05w]. Zwerge [Str05b]. Zwergen [Ren06]. zwischen [Ein02b, Ein11a, Ein11g, Ein05i, Ein05q, Ein05-39, FF97, Hei75b, Her66b, Inf99, Lam79, LS05, Müh11, Hof08c]. Zwölf [Vit15].

References

Amaldi:1984:RAA

[A^{+}84] Edoardo Amaldi et al. La Radioactivité artificielle à 50 ans: 1934-1984. Éditions de physique, Les Ulis, France, 1984. ISBN 2-902731-71-X. x + 164 pp. LCCN QC795.55.I5 R33 1984. 250F. Preface by André Guinier.

Alunni:2009:AEH

[A^{+}09] Charles Alunni et al. Albert Einstein et Hermann Weyl (1955– 2005): questions épistémologiques ouvertes, volume 1 of Collection Pensée des Sciences. Barbieri Selvaggi, Manduria, Italia, 2009. ISBN 88-6187-023-6. 249 pp. LCCN Q175.32.K45 A43 2009.

Argo:1975:BRJ

[AA75] Mary Argo and Harold Argo. Book review: Jane S. Wilson, All in our time: the reminiscences of twelve nuclear pioneers. Physics Today, 31(5):74–76, May 1975. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v29/i5/p74_s1.

Afulezi:2002:AAR

[AA02] Uju Nkwocha Afulezi and Ugochukwu Uju Afulezi, editors. African and Africa-related Nobel Prize winners: portraitures in excellence. University Press of America, Lanham, MD, USA, 2002. ISBN 0-7618-2101-5 (hardcover), 0-7618-2102-3 (paperback). xvi + 357 pp. LCCN AS911.N9 A38 2002.

Abdo:2009:LVS

[AAA^{+}09] A. A. Abdo, M. Ackermann, M. Ajello, K. Asano, W. B. Atwood, M. Axelsson, L. Baldini, J. Ballet, G. Barbiellini, M. G. Baring, D. Bastieri, K. Bechtol, R. Bellazzini, B. Berenji, P. N. Bhat, E. Bissaldi, E. D. Bloom, E. Bonamente, J. Bonnell, A. W. Borgland, A. Bouvier, J. Bregeon, A. Brez, M. S. Briggs, M. Brigida, P. Bruel, J. M. Burgess, T. H. Burnett, G. A. Caliandro, R. A. Cameron, P. A. Caraveo, J. M. Casandjian, C. Cecchi, Ö. Çelik, V. Chaplin, E. Charles, C. C. Cheung, J. Chiang, S. Ciprini, R. Claus, J. Cohen-Tanugi, L. R. Cominsky, V. Connaughton, J. Conrad, S. Cutini, C. D. Dermer, A. de Angelis, F. de Palma, S. W. Digel, B. L. Dingus, E. do Couto e Silva, P. S. Drell, R. Dubois, D. Dumora, C. Farnier, C. Favuzzi, S. J. Fegan, J. Finke, G. Fishman, W. B. Focke, L. Foschini, Y. Fukazawa, S. Funk, P. Fusco, F. Gargano, D. Gasparrini, N. Gehrels, S. Germani, L. Gibby, B. Giebels, N. Giglietto, F. Giordano, T. Glanzman, G. Godfrey, J. Granot, J. Greiner, I. A. Grenier, M.-H. Grondin, J. E. Grove, D. Grupe, L. Guillemot, S. Guiriec, Y. Hanabata, A. K. Harding, M. Hayashida, E. Hays, E. A. Hoversten, R. E. Hughes, G. Jóhannesson, A. S. Johnson, R. P. Johnson, W. N. Johnson, T. Kamae, H. Katagiri, J. Kataoka, N. Kawai, M. Kerr, R. M. Kippen, J. Knödlseder, D. Kocevski, C. Kouveliotou, F. Kuehn, M. Kuss, J. Lande, L. Latronico, M. Lemoine-Goumard, F. Longo, F. Loparco, B. Lott, M. N. Lovellette, P. Lubrano, G. M. Madejski, A. Makeev, M. N. Mazziotta, S. McBreen, J. E.

McEnery, S. McGlynn, P. Mészáros, C. Meurer, P. F. Michelson, W. Mitthumsiri, T. Mizuno, A. A. Moiseev, C. Monte, M. E. Monzani, E. Moretti, A. Morselli, I. V. Moskalenko, S. Murgia, T. Nakamori, P. L. Nolan, J. P. Norris, E. Nuss, M. Ohno, T. Ohsugi, N. Omodei, E. Orlando, J. F. Ormes, M. Ozaki, W. S. Paciesas, D. Paneque, J. H. Panetta, D. Parent, V. Pelassa, M. Pepe, M. Pesce-Rollins, V. Petrosian, F. Piron, T. A. Porter, R. Preece, S. Rainò, E. Ramirez-Ruiz, R. Rando, M. Razzano, S. Razzaque, A. Reimer, O. Reimer, T. Reposeur, S. Ritz, L. S. Rochester, A. Y. Rodriguez, M. Roth, F. Ryde, H. F.-W. Sadrozinski, D. Sanchez, A. Sander, P. M. Saz Parkinson, J. D. Scargle, T. L. Schalk, C. Sgrò, E. J. Siskind, D. A. Smith, P. D. Smith, G. Spandre, P. Spinelli, M. Stamatikos, F. W. Stecker, M. S. Strickman, D. J. Suson, H. Tajima, H. Takahashi, T. Takahashi, T. Tanaka, J. B. Thayer, J. G. Thayer, D. J. Thompson, L. Tibaldo, K. Toma, D. F. Torres, G. Tosti, E. Troja, Y. Uchiyama, T. Uehara, T. L. Usher, A. J. van der Horst, V. Vasileiou, N. Vilchez, V. Vitale, A. von Kienlin, A. P. Waite, P. Wang, C. Wilson-Hodge, B. L. Winer, K. S. Wood, X. F. Wu, R. Yamazaki, T. Ylinen, and M. Ziegler. A limit on the variation of the speed of light arising from quantum gravity effects. Nature, 462(7271):331–334, November 19, 2009. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v462/n7271/full/nature08574.html.

Abuter:2018:DGR

[AAA^{+}18] R. Abuter, A. Amorim, N. Anugu, M. Bauböck, M. Benisty, J. P. Berger, N. Blind, H. Bonnet, W. Brandner, A. Buron, C. Collin, F. Chapron, Y. Clénet, V. Coudé du Foresto, P. T. de Zeeuw, C. Deen, F. Delplancke-Ströbele, R. Dembet, J. Dexter, G. Duvert, A. Eckart, F. Eisenhauer, G. Finger, N. M. Förster Schreiber, P. Fédou, P. Garcia, R. Garcia Lopez, F. Gao, E. Gendron, R. Genzel, S. Gillessen, P. Gordo, M. Habibi, X. Haubois, M. Haug, F. Haußmann, Th. Henning, S. Hippler, M. Horrobin, Z. Hubert, N. Hubin, A. Jimenez Rosales, L. Jochum, L. Jocou, A. Kaufer, S. Kellner, S. Kendrew, P. Kervella, Y. Kok, M. Kulas, S. Lacour, V. Lapeyrère, B. Lazareff, J.-B. Le Bouquin, P. Léna, M. Lippa, R. Lenzen, A. Mérand, E. Müler, U. Neumann, T. Ott, L. Palanca, T. Paumard, L. Pasquini, K. Perraut, G. Perrin, O. Pfuhl, P. M. Plewa, S. Rabien, A. Ramírez, J. Ramos, C. Rau, G. Rodríguez-Coira, R.-R. Rohloff, G. Rousset, J. Sanchez-Bermudez, S. Scheithauer, M. Schöller, N. Schuler, J. Spyromilio, O. Straub, C. Straubmeier, E. Sturm, L. J. Tacconi, K. R. W. Tristram, F. Vincent, S. von Fellenberg, I. Wank, I. Waisberg, F. Widmann, E. Wieprecht, M. Wiest, E. Wiezorrek, J. Woillez, S. Yazici, D. Ziegler, and G. Zins. Detection of the gravitational redshift in the orbit of the star S2 near the Galactic centre massive black hole. Astronomy and Astrophysics, 615:L15, July 2018. CODEN AAEJAF. ISSN 0004-6361 (print), 1432-

0746 (electronic). URL https://www.aanda.org/articles/
aa/abs/2018/07/aa33718-18/aa33718-18.html.

Aaseng:1988:INP

[Aas88] Nathan Aaseng. The inventors: Nobel Prizes in chemistry, physics, and medicine. Nobel Prize winners. Lerner Publications, Minneapolis, MN, USA, 1988. ISBN 0-8225-0651-3. 79 pp. LCCN T48 .A27 1988.

Aaserud:1990:RSN

[Aas90] Finn Aaserud. Redirecting science: Niels Bohr, philanthropy, and the rise of nuclear physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35366-1, 0-521-53067-9 (paperback). xiii + 356 pp. LCCN

QC789.2.D4 A27 1990. URL http://www.loc.gov/catdir/
description/cam024/89048317.html;   http://www.loc.
gov/catdir/toc/cam028/89048317.html.

Aaserud:2002:NBN

[Aas02] Finn Aaserud. Niels Bohr and the Nobel prizes. In Crawford [Cra02b], pages 39–64. ISBN 4-946443-69-X. ISSN 0282-1036. LCCN AS911.N9 .H57 2002.

Abate:2009:PN

[Aba09] Marco Abate. Perché Nobel?. (Italian) [Why Nobel?]. I blu. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 88-470-0811-5, 88-470-0810-7. 143 pp. LCCN AS911.N9 P47 2009eb; AS911 .N9 INTERNET.

Abbott:1988:MCC

[Abb88] Larry Abbott. The mystery of the Cosmological Constant. Scientific American, 258(5):106–113, May 1988. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v258/n5/pdf/scientificamerican0588-106.pdf.

Abbott:1998:FDG

[Abb98a] Alison Abbott. Fire damages Germany’s Einstein tower. Nature, 391(6664):217, January 15, 1998. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v391/n6664/full/391217b0.
html.

Abbott:1998:SRE

[Abb98b] Alison Abbott. Science returns to Einstein’s rural retreat. Nature, 394(6694):608, August 13, 1998. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v394/n6694/
full/394608a0.html.

Abbott:2002:EEM

[Abb02] Alison Abbott. Early Einstein manuscript set to make a relative fortune. Nature, 419(6903):103, September 12, 2002. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v419/n6903/
full/419103b.html.

Abbott:2005:AEC

[Abb05] A. Abbott. Albert Einstein: Chief engineer of the universe. Nature, 435(7041):426–427, May 26, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Aiken:2006:HUC

[ABC^{+}06] Conrad Aiken, Pearl S. Buck, Leo Connellan, Harry Crews, James Dickey, Albert Einstein, T. S. Eliot, Jesse Hill Ford, John Gardner, George Garrett, Richard Howard, Katherine Anne Porter, Lewis P. Simpson, Allen Tate, Lionel Trilling, Evelyn Waugh, E. B. White, Edmund Wilson, and Stephen Corey. “Into the hectic unknown”: Correspondence from the archives of The Georgia Review, 1947–76. Georgia Review, 60(3–4):606–698, Fall /Winter 2006. ISSN 0016-8386.

Appleton:1953:SPP

[ABdB^{+}53] Sir Edward Appleton, David Bohm, Louis de Broglie, Richard Courant, Albert Einstein, Pascual Jordan, Theodore V. Kármán, S. S. Penner, Alfred Landé, Erwin Schrödinger, and Hermann Weyl, editors. Scientific papers presented to Max Born on his retirement from the Tait Chair of Natural Philosophy in the University of Edinburgh. Oliver and Boyd, Edinburgh, UK; London, UK, 1953. vi + 94 pp. LCCN QC71 .S35.

Adler:1955:ANW

[ABH^{+}55] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. Keesing’s Archiv der Gegenwart, 25(??):5251H–??, ???? 1955. The original typescript, entitled “Mainauer Kundgebung” [Mainau rally], signed by authors, is reproduced in [ABH^{+}87].

Adler:1987:ANW

[ABH^{+}87] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. In ????, editor, 30 Jahre Göttinger Erklärung. Nachdenken über die Rolle des Wissenschaftlers in der Gesellschaft. (German) [30 years Göttingen declaration. Reflection on the role of the scientist in the company], page 42. ????, Munich, West Germany, 1987. Reproduction of original signed typescript of [ABH^{+}55].

Abiko:1991:CTO

[Abi91] Seiya Abiko. On the chemico-thermal origins of Special Relativity. Historical Studies in the Physical and Biological Sciences, 22(1):1–24, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.jstor.org/

stable/27757671.

Abiko:2000:EKA

[Abi00a] Seiya Abiko. Einstein’s Kyoto address: ‘how I created the Theory of Relativity’. Historical Studies in the Physical and Biological Sciences, 31(1):1–35, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Abiko:2000:ETF

[Abi00b] Seiya Abiko. Einstein’s theories of the fluctuation and the thermal radiation: the first quantum theory through statistical thermodynamics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2): 130–147, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Abiko:2003:EDE

[Abi03] Seiya Abiko. On Einstein’s distrust of the electromagnetic theory: The origin of the light-velocity postulate. Historical Studies in the Physical and Biological Sciences, 33(2):193–215, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See comments [Dar04b] and reply [Abi04].

Abiko:2004:ROD

[Abi04] Seiya Abiko. Reply to Olivier Darrigol. Historical Studies in the Physical and Biological Sciences, 35(1):157–160, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Abi03, Dar04b].

Ashtekar:2015:GRG

[ABIM15] Abhay Ashtekar, B. (Beverly) Berger, James A. Isenberg, and M. A. H. MacCallum, editors. General Relativity and Gravitation: a Centennial Perspective. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-03731-X (hardcover). xxi + 674 pp. LCCN QC173.6 .G452 2015.

URL http://www.cambridge.org/us/academic/subjects/
physics/cosmology-relativity-and-gravitation/general-
relativity-and-gravitation-centennial-perspective.

Aronov:1979:EMD

[ABM79] R. A. Aronov, B. M. Bolotovskii, and N. V. Mitskevich. Elements of materialism and dialectics in the forming of Albert Einstein philosophical views. Voprosy Filosofii, ??(11):56–66, ???? 1979. ISSN 0042-8744.

Abraham:1912:NRG

[Abr12a] Max Abraham. Nochmals Relativität und Gravitation: Erwiderung auf eine Bemerkung des Hrn. A. Einstein. (German) [Relativity and gravitation again: response to a comment by Mr. A. Einstein]. Annalen der Physik (1900), 344 (39)(12):444–448, ???? 1912. ISSN 1521-3889. See reply [Ein12b, Ein12g, Abr12b].

Abraham:1912:RGE

[Abr12b] Max Abraham. Relativität und Gravitation. Erwiderung auf eine Bemerkung des Hrn. A. Einstein. (German) [Relativity and gravitation. Reply to a comment by Mr. A. Einstein]. Annalen der Physik (1900), 343 (38)(10):1056–1058, ???? 1912.

ISSN 1521-3889. URL https://onlinelibrary.wiley.com/
doi/10.1002/andp.19123431014. See reply [Ein12b, Ein12g,

Abr12a].

Abraham:1913:NGG

[Abr13] Max Abraham. Eine neue Gravitationstheorie. (German) [A new theory of gravity]. Archiv der Mathematik under Physik, 20(??):193–209, ???? 1913.

Abrams:1988:NPP

[Abr88] Irwin Abrams. The Nobel Peace Prize and the laureates: an illustrated biographical history, 1901–1987. G. K. Hall, Boston, MA, USA, 1988. ISBN 0-8161-8609-X. xv + 269 pp. LCCN JX1962.A2 A25 1988.

Abraham:1998:MIA

[Abr98] Itty Abraham. The making of the Indian atomic bomb: science, secrecy, and the postcolonial state. Postcolonial encounters. Zed Books, London, UK, 1998. ISBN 1-85649-629-5 (hardcover), 1-85649-630-9 (paperback). ix + 180 pp. LCCN QC773 .A27 1998. UK£39.95, UK£13.95. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/bios/hol058/98027616.html;   http://www.loc.
gov/catdir/description/hol056/98027616.html;   http:
//www.loc.gov/catdir/toc/hol052/98027616.html.

Abraham:2001:PGB

[Abr01a] Carolyn Abraham. Possessing genius: the bizarre odyssey of Einstein’s brain. St. Martin’s Press, New York, NY, USA, 2001. ISBN 0-312-28117-X. xi + 388 + 8 pp. LCCN RB17.H365 A274 2002. URL http://www.loc.

gov/catdir/toc/fy033/2001048604.html;   https://www.
math.utah.edu/pub/tex/bib/einstein.bib.

Abrams:2001:NPP

[Abr01b] Irwin Abrams. The Nobel Peace Prize and the laureates: an illustrated biographical history, 1901–2001. Science History Publications/USA, Nantucket, MA, MA, centennial edition, 2001. ISBN 0-88135-388-4. xvii + 350 pp. LCCN JZ5540 .A27 2001.

Abraham:2002:PGB

[Abr02] Carolyn Abraham. Possessing genius: the bizarre odyssey of Einstein’s brain. Penguin Canada, Toronto, ON, Canada, 2002. ISBN 0-14-029368-X. xiii + 388 + 8 pp. LCCN ???? US$24.00.

Abraham:2003:PGT

[Abr03] Carolyn Abraham. Possessing genius: the true account of the bizarre odyssey of Einstein’s brain. St. Martin’s Press, New York, NY, USA, 2003. ISBN 0-312-30304-1. xi + 388 pp. LCCN RB17.H365 A274 2003.

Abraham:2004:PGB

[Abr04a] Carolyn Abraham. Possessing genius: the bizarre odyssey of Einstein’s brain. Icon, Cambridge, UK, 2004. ISBN 1-84046-549-2 (cased), 1-84046-625-1 (paperback). 388 + 8 pp. LCCN ????

Abrams:2004:P

[Abr04b] Irwin Abrams, editor. Peace 1996–2000. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-238-001-9, 981-238-002-7 (paperback). xii + 165 pp. LCCN JZ5537 .P43 2004.

Abrikosov:2004:NLT

[Abr04c] A. A. Abrikosov. Nobel Lecture: Type-II superconductors and the vortex lattice. Reviews of Modern Physics, 76(3):975–979, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.975;   http://rmp.aps.org/
abstract/RMP/v76/i3/p975_1.

Abrams:2005:P

[Abr05] Irwin Abrams, editor. Peace 1996–2000. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-238-001-9, 981-238-002-7 (paperback). xii + 165 pp. LCCN JX1953 .P37 2005eb.

Abraham:2007:FF

[Abr07a] Max Abraham. The free fall. In Renn et al. [RSS^{+}07a], pages 341–345. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.1007/978-1-4020-4000-

9_24. Originally published in Rendiconti del R. Instituto Lombardo di scienze e lettere. German translation published as “Der freie Fall”, Physikalische Zeitschrift 13, 310–311 (1912).

Abraham:2007:NTG

[Abr07b] Max Abraham. A new theory of gravitation. In Renn et al. [RSS^{+}07a], pages 347–362. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_25. Lecture presented on 19 October 1912 to the Società italiana per il progresso delle scienze. German translation published as “Eine neue Gravitationstheorie”, Archiv der Mathematik und Physik, Third Series 20, 193–209 (1913).

Abraham:2007:TG

[Abr07c] Max Abraham. On the theory of gravitation. In Renn et al. [RSS^{+}07a], pages 331–339. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_23. Originally published in Rendiconti della R. Accademia dei Lincei. German translation published as “Zur Theorie der Gravitation”, Physikalische Zeitschrift 13, 1–4, 176 (1912).

Abraham:2007:RTG

[Abr07d] Max Abraham. Recent theories of gravitation. In Renn et al. [RSS^{+}07a], pages 363–410. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_26. Originally published as “Neuere Gravitationstheorien”, Jahrbuch der Relativität und Elektronik 11, 470–520 (1915).

Abraham:2009:SAC

[Abr09] Itty Abraham, editor. South Asian cultures of the bomb: atomic publics and the state in India and Pakistan. Indiana University Press, Bloomington, IN, USA, 2009. ISBN 0-253-35253-3 (cloth), 0-253-22032-7 (paperback). 222 pp. LCCN

UA840 .S645 2009. URL http://www.loc.gov/catdir/toc/
ecip0824/2008031607.html.

Aerts:1999:WTH

[ABW99] Diederik Aerts, Jan Broekaert, and Willy Weyns, editors. A world in transition: humankind and nature: the green book of “Einstein meets Magritte”, volume 5 of Einstein meets Magritte. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5761-2 (hardcover). LCCN GE350 .W674 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//lcweb.loc.gov/catdir/toc/99029564.html. See other

volumes in this series [Aer99a, AMM99, Aer99b, AVvdV99, BAM99, CSvB99, HBR99].

Amelino-Camelia:2009:BEG

[AC09] Giovanni Amelino-Camelia. Beyond Einstein’s gravity. Physics World, 22(7):48–49, July 2009. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/22/07/phwv22i07a45.
pdf.

Ashtekar:2003:RFR

[ACH^{+}03] Abhay Ashtekar, Robert S. Cohen, Don Howard, Jürgen Renn, Sahotra Sarkar, and Abner Shimony, editors. Revisiting the foundations of relativistic physics: Festschrift in honor of John Stachel, volume 234 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1284-5. xxii + 649 pp. LCCN

Q174 .B67 vol. 234. URL http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-d.html;   http://www.
loc.gov/catdir/enhancements/fy0813/2003046757-t.html.

Achinstein:2013:BRH

[Ach13] Peter Achinstein. Book review: How the Great Scientists Reasoned: The Scientific Method in Action, Gary G. Tibbetts, Elsevier, 2013. $74.95 (176 pp.). ISBN 978-0-12-398498-2. Physics Today, 66(11):50, November 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Aranzales:2023:FHT

[ACT23] Iván Aranzales, Ho Fai Chan, and Benno Torgler. Finally! How time lapse in Nobel Prize reception affects emotionality in the Nobel Prize banquet speeches. Scientometrics, 128(7):

4089–4115, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04739-w.

Allison:1954:SAF

[ACU^{+}54] Samuel K. Allison, Edward Condon, Harold C. Urey, Farrington Daniels, F. W. Loomis, Edward Shils, Linus Pauling, S. A. Goudsmit, Julian Schwinger, Albert Einstein, Hugh C. Wolfe, Cyril S. Smith, Victor F. Weisskopf, and Hans Bethe. Scientists affirm faith in Oppenheimer. Bulletin of the Atomic Scientists, 10(5):188–190, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Acuna:2014:EEB

[Acu14] Pablo Acuña. On the empirical equivalence between special relativity and Lorentz’s ether theory. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):283–302, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000124.

Aczel:1999:GEE

[Acz99] Amir D. Aczel. God’s Equation: Einstein, Relativity, and the Expanding Universe. Four Walls Eight Windows, New York, NY, USA, 1999. ISBN 1-56858-139-4. xvii + 236 pp. LCCN QB981 .A35 1999.

Aczel:2000:GEE

[Acz00] Amir D. Aczel. God’s Equation: Einstein, Relativity, and the Expanding Universe. Delta, New York, NY, USA, 2000. ISBN 0-385-33485-0. xvii + 236 pp. LCCN ????

Aczel:2001:EGM

[Acz01] Amir D. Aczel. Entanglement: The Greatest Mystery in Physics: The Unlikely Story of How Scientists, Mathematicians, and Philosphers Proved Einstein’s Spookiest Theory. Four Walls Eight Windows, New York, NY, USA, 2001. ISBN 1-56858-232-3. xviii + 284 pp. LCCN QC174.12 .A29 2001. URL http://

www.loc.gov/catdir/enhancements/fy0831/2002069338-d.
html.

Aczel:2003:EUS

[Acz03] Amir D. Aczel. Entanglement: the unlikely story of how scientists, mathematicians, and philosophers proved Einstein’s spookiest theory. Plume, New York, NY, USA, 2003. ISBN 0-452-28457-0 (paperback). xviii + 284 pp. LCCN QC174.12 .A29 2003.

Aczel:2009:UWS

[Acz09] Amir D. Aczel. Uranium wars: the scientific rivalry that created the nuclear age. Palgrave Macmillan, New York, NY, USA, 2009. ISBN 0-230-61374-8. 248 + 8 pp. LCCN QC773.A1 A28 2009.

Aczel:2010:PCS

[Acz10] Amir D. Aczel. Present at the creation: the story of CERN and the Large Hadron Collider. Harmony Books, New York, NY, USA, 2010. ISBN 0-307-59167-0. xvi + 271 + 8 pp. LCCN QC787.P73 A29 2010.

Adams:2007:DA

[Ada07] John Adams. Doctor atomic. Opera in two acts, published by Hendon Music (New York), April 9, 2007. Libretto by Peter Sellars. Based on the 16 July 1945 test of the first atomic bomb at Trinity, NM, USA.

Adamson:2017:SSU

[Ada17] Matthew Adamson. The secret search for uranium in Cold War Morocco. Physics Today, 70(6):54–60, June 2017. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Adler:1982:EGS

[Adl82] Stephen L. Adler. Einstein gravity as a symmetry-breaking effect in quantum field theory. Reviews of Modern Physics, 54 (3):729–766, July 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.54.729;   http://rmp.
aps.org/abstract/RMP/v54/i3/p729_1. See erratum [Adl83].

Adler:1983:EEG

[Adl83] Stephen L. Adler. Erratum: Einstein gravity as a symmetrybreaking effect in quantum field theory. Reviews of Modern Physics, 55(3):837, July 1983. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.55.837;   http:/
/rmp.aps.org/abstract/RMP/v55/i3/p837_1. See [Adl82].

Adler:2003:HPM

[Adl03a] Mortimer J. Adler, editor. Henri Poincaré, Max Planck, Alfred North Whitehead, Albert Einstein, Sir Arthur Eddington, Niels Bohr, G. H. Hardy, Werner Heisenberg, Erwin Schrödinger, Theodosius Dobzhansky, C. H. Waddington, volume 56 of Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, 2003. ISBN 0-85229-531-6. xvi + 749 pp. LCCN AC1 .G72 1990.

Adloff:2003:CNP

[Adl03b] Jean-Pierre Adloff. The centennial of the 1903 Nobel Prize for Physics. Radiochimica Acta, 91(12):681–688, December 2003. CODEN RAACAP. ISSN 0033-8230 (print), 2193-3405

(electronic). URL http://www.degruyter.com/view/j/ract.
2003.91.issue-12-2003/ract.91.12.681.23428/ract.91.
12.681.23428.xml.

Adloff:2012:NPA

[Adl12] Jean-Pierre Adloff. Nobel Prize awards in radiochemistry. Radiochimica Acta, 100(8–9):509–521, August 2012. CODEN RAACAP. ISSN 0033-8230 (print), 2193-3405 (electronic). URL

http://www.degruyter.com/view/j/ract.2012.100.issue-
8-9/ract.2012.1953/ract.2012.1953.xml.

Aris:1983:SSC

[ADS83] Rutherford Aris, H. Ted (Howard Ted) Davis, and Roger H. Stuewer, editors. Springs of Scientific Creativity: Essays on Founders of Modern Science. University of Minnesota Press, Minneapolis, MN, USA, 1983. ISBN 0-8166-1087-8. viii + 342 pp. LCCN QC7 .S77 1983.

Allen:2001:NPL

[AE01] Sture Allén and Kjell Espmark. The Nobel Prize in literature: an introduction. Swedish Academy, Stockholm, Sweden, 2001. ISBN 91-1-301042-5. 6 + 58 pp. LCCN ???? English translation by Erik T. Frykman.

Adeoti:2006:ANP

[AE06] Gbemisola Adeoti and Mabel I. E. Evwierhoma, editors. After the Nobel Prize: reflections on African literature, governance, and development. Association of Nigerian Authors, Lagos, Nigeria, 2006. ISBN 978-039-185-1. LCCN PL8010 .A3548 2006.

Aerts:1999:QSN

[Aer99a] Diederik Aerts, editor. Quantum structures and the nature of reality: the indigo book of “Einstein meets Magritte”, volume 7 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 0-7923-5763-9. LCCN QC174.17.M35 Q36 1999. See other volumes in this series [AMM99, Aer99b, ABW99, AVvdV99, BAM99, CSvB99, HBR99].

Aerts:1999:STS

[Aer99b] Diederik Aerts, editor. Science, technology, and social change: the orange book of “Einstein meets Magritte”, volume 3 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 0-7923-5759-0 (hardcover). LCCN HM831 .S37 1999. See other volumes in this series [Aer99a, AMM99, ABW99, AVvdV99, BAM99, CSvB99, HBR99].

Adler:1990:GBW

[AFG90] Mortimer Jerome Adler, Clifton Fadiman, and Philip W. Goetz, editors. Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, second edition, 1990. ISBN 0-85229-531-6 (set). ???? pp. LCCN AC1 .G72 1990.

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002. URL http:

//www.gbv.de/dms/goettingen/360423353.pdf;   http://
www.loc.gov/catdir/enhancements/fy0611/2001048396-d.
html.

Al-Ghazi:2013:NNP

[AG13] Muthana Al-Ghazi. Notes on nuclear physics in 1932. Physics Today, 66(11):11, November 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [RC13].

Agassi:1975:SFb

[Aga75] Joseph Agassi. Science in Flux, volume 28 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346.

xxvi + 559 pp. LCCN ???? URL https://link.springer.
com/book/10.1007/978-94-010-1810-4.

Agassi:1993:RTQ

[Aga93] Joseph Agassi. Radiation Theory and the Quantum Revolution. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1993. ISBN 0-8176-2905-X, 3-7643-2905-X. 12 + 178 pp. LCCN ????

Agashe:2006:EEL

[Aga06] S. D. Agashe. Einstein’s “Zur Elektrodynamik … ” (1905) revisited, with some consequences. Foundations of Physics, 36 (7):955–1011, July 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-006-9055-y. See addendum

[Aga07].

Agashe:2007:AEE

[Aga07] S. D. Agashe. Addendum to “Einstein’s ‘Zur Electrodynamik … ’ (1905) Revisited, with some Consequences”^{(1)} by S. D. Agashe. Foundations of Physics, 37(2):306–309, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-006-9103-7. See [Aga06].

Alvarez-Gaum:2012:PLY

[AGMT12] Luis Alvarez-Gaum, Michelangelo Mangano, and Emmanuel Tsesmelis. From the PS to the LHC — 50 Years of Nobel Memories in High-Energy Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-30843-0, 3-642-30844-9 (e-book). ???? pp. LCCN ????

Aguirregabiria:1990:CRE

[Agu90] J. Aguirregabiria. Computing the Ricci and Einstein tensors. Mathematica Journal, 1(2):51–54, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Alpher:1960:TLB

[AH60] R. A. Alpher and R. Herman. Tea leaves, Baer’s Law, and Albert Einstein. American Journal of Physics, 28(8):748, August 1960. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/28/748/

Albrecht:1993:NTG

[AH93] Helmuth Albrecht and Armin Hermann, editors. Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. (German) [Science and technology in history: 25 Years of the Chair of History of Science and Technology at the Historical Institute of the University of Stuttgart]. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. 401 pp. LCCN Q183.4.G293 U556 1993.

Anderson:1996:ACT

[AH96] Britt Anderson and Thomas Harvey. Alterations in cortical thickness and neuronal density in the frontal cortex of Albert Einstein. Neuroscience Letters, 210(3):161–164, June 7, 1996. CODEN NELED5. ISSN 0304-3940 (print), 1872-7972 (electronic).

Ahern:2003:WHT

[Ahe03] Joseph-James Ahern. ’we had the hose turned on us!’: Ross Gunn and the Naval Research Laboratory’s early research into nuclear propulsion, 1939–1946. Historical Studies in the Physical and Biological Sciences, 33(2):217–236, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Adelberger:2005:TGI

[AHH05] Eric Adelberger, Blayne Heckel, and C. D. Hoyle. Testing the gravitational inverse-square law. Physics World, 18(4):41–45, April 3, 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://www.tiptop.iop.org/
full/pwa-pdf/18/4/phwv18i4a38.pdf. See [vE10] for the orig-

inal torsion balance research.

Ainscain:1977:INO

[Aĭn77] Albert Aĭnščaĭn. Isaac Newton (1643–1727) (on the occasion of the 250th anniversary of his death). Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 20(53)(4):303–305, 1977. ISSN 0015-3265. Translated from the Russian.

Albright:1997:BSS

[AK97] Joseph Albright and Marcia Kunstel. Bombshell: the secret story of America’s unknown atomic spy conspiracy.

Times Books, New York, NY, USA, 1997. ISBN 0-8129-2861-X. xv + 399 pp. LCCN QC773.3.U5 A45

1997. URL http://www.loc.gov/catdir/enhancements/
fy0601/97203804-b.html;   http://www.loc.gov/catdir/
enhancements/fy0601/97203804-d.html;   http://www.loc.
gov/catdir/enhancements/fy0601/97203804-s.html.

Anglin:2002:BEC

[AK02] James R. Anglin and Wolfgang Ketterle. Bose–Einstein condensation of atomic gases. Nature, 416(6877):211–218, March 14, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v416/n6877/full/416211a.html.

Akasaki:2015:FJB

[Aka15] Isamu Akasaki. Fascinating journeys into blue light (Nobel lecture). Annalen der Physik (Berlin), 527(5-6):311–326, 2015. CODEN ANPYA2. ISSN 1521-3889.

Abdel-Khalek:2021:EPR

[AKBAE21] Sayed Abdel-Khalek, Kamal Berrada, Mariam Algarni, and Hichem Eleuch. Einstein–Podolsky–Rosen steering for mixed entangled coherent states. Entropy, 23(11):1442, ???? 2021. CODEN ENTRFG. ISSN 1099-4300. URL https://dblp.org/

db/journals/entropy/entropy23.html#Abdel-KhalekBAE21.

Alkan:2023:KDT

[AKD23] Bilal Baris Alkan, Leyla Karakus, and Bekir Direkci. Knowledge discovery from the texts of Nobel Prize winners in literature: sentiment analysis and Latent Dirichlet Allocation. Scientometrics, 128(9):5311–5334, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://

link.springer.com/article/10.1007/s11192-023-04783-6.

Abrams:2009:P

[AL09] Irwin Abrams and Scott London, editors. Peace 2001–2005. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2009. ISBN 981-279-432-8 (hardcover), 981-279-433-6 (paperback). xvi + 154 pp. LCCN JZ5537 .P43 2009.

Albert:1992:BRE

[Alb92] David Z. Albert. Bohr’s response to Einstein, Podolsky, and Rosen. In The Scientific Enterprise: The Bar-Hillel Colloquium:

Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 17, pages 269–272. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. LCCN Q174 .S354

1992. URL https://link.springer.com/chapter/10.1007/
978-94-011-2688-5_17.

Alcubierre:2008:INR

[Alc08] Miguel Alcubierre. Introduction to 3 + 1 numerical relativity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-920567-1 (hardcover). xiv + 444 pp. LCCN QC173.55 .A43 2008.

Aldridge:1993:PLA

[Ald93] S. Aldridge. The private lives of Albert Einstein — R. Highfield, P. Carter. New Statesman Society, 6(268):37–38, September 3, 1993. CODEN NESSEF. ISSN 0954-2361.

Alferov:2001:NLD

[Alf01] Zhores I. Alferov. Nobel Lecture: The double heterostructure concept and its applications in physics, electronics, and technology. Reviews of Modern Physics, 73(3):767–782, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.767;   http://rmp.aps.org/
abstract/RMP/v73/i3/p767_1.

Alford:2009:BRG

[Alf09] Mark Alford. Book review: Gordon Fraser, Cosmic Anger: Abdus Salam — The First Muslim Nobel Scientist. Oxford: Oxford University Press, 2008, xiii + 305 pages. $49.95 (cloth). Physics in Perspective (PIP), 11(3):347–348, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Allison:1957:EF

[All57] Samuel King Allison. Enrico Fermi, 1901–1954, volume XXX– 6th memoir of National Academy of Sciences. Biographical memoirs. National Academy of Sciences, New York, NY, USA, 1957. 125–155 pp. LCCN Q141 .N2 vol. 30, 6th memoir.

Allen:1990:ECC

[All90] Les Allen. Are Einstein A coefficients constant? Physics World, 3(1):19, January 1990. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/3/1/phwv3i1a18.
pdf.

Allen:1997:L

[All97] Sture Allén, editor. Literature, 1991–1995. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ix + 98 pp. LCCN ????

Allchin:2002:EWN

[All02] Douglas Allchin. To err and win a Nobel Prize: Paul Boyer, ATP synthase and the emergence of bioenergetics. Journal of the History of Biology, 35(1):149–172, March 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1014583721788;   http://link.springer.com/content/
pdf/10.1023/A:1014583721788.pdf.

Allais:2005:AEE

[All05] Maurice Allais. Albert Einstein, un extraordinaire paradoxe. (French) [Albert Einstein, an extraordinary paradox], volume IV of Contributions de Maurice Allais à la physique expérimentale et théorique. C. Juglar, Paris, France, 2005. ISBN 2-908735-20-2. 89 pp. LCCN QC173.55 .A45 2005.

Alperovitz:1965:ADH

[Alp65] Gar Alperovitz. Atomic diplomacy: Hiroshima and Potsdam; the use of the atomic bomb and the American confrontation with Soviet power: Hiroshima and Potsdam. Simon and Schuster, New York, NY, USA, 1965. 317 pp. LCCN E813 .A75.

Alperovitz:1966:ADH

[Alp66] Gar Alperovitz. Atomic diplomacy:: Hiroshima and Potsdam. Secker and Warburg, London, UK, 1966. ???? pp. LCCN E813 .A75 1966.

Alperovitz:1985:ADH

[Alp85] Gar Alperovitz. Atomic diplomacy: Hiroshima and Potsdam: the use of the atomic bomb and the American confrontation with Soviet power. Penguin, London, UK and New York, NY, USA, expanded and updated edition, 1985. ISBN 0-14-008337-5 (paperback). xi + 427 pp. LCCN E813 .A75 1985.

Alperovitz:1994:ADH

[Alp94] Gar Alperovitz. Atomic diplomacy: Hiroshima and Potsdam: the use of the atomic bomb and the American confrontation with Soviet power. Pluto Press, London, UK, second expanded edition, 1994. ISBN 0-7453-0948-8 (clothbound). xii + 402 pp. LCCN E183 .A75 1994.

Alpert:2011:EED

[Alp11] Mark Alpert. Het Einstein Enigma. (Dutch) [The Einstein Enigma]. Poema Pocket, Amsterdam, The Netherlands, third edition, 2011. ISBN 90-210-1079-8. 351 pp. LCCN ???? Dutch translation by Pon Ruiter of Final Theory.

Altavilla:2006:FFW

[Alt06a] Costanza Altavilla. Fisica e filosofia in Werner Heisenberg. (Italian) [Physics and Philosophy of Werner Heisenberg], volume 4 of Transazioni. Guida, Napoli, Italia, 2006. ISBN 88-6042-391-0. 368 pp. LCCN QC16.H35 L48 2006. 23.00

EUR. URL http://www.loc.gov/catdir/toc/casalini05/
07041675.pdf.

Altschuler:2006:CPD

[Alt06b] Glenn C. Altschuler. The convictions of Peter Debye. Dædalus, 135(4):96–103, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.jstor.org/

stable/20028076;   http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.4.96.

Alter:2013:STS

[Alt13] Svetlana Alter. Secret Trace of the Soul of Mileva Marić-Einstein. Dorrance Publishing, Pittsburgh, PA, USA, 2013. ???? pp.

Melnick:2007:TCW

[aM07] aj Melnick. They changed the world: people of the Manhattan Project. Sunstone Press, Santa Fe, NM, USA, 2007. ISBN 0-86534-530-9 (softcover). ???? pp. LCCN Q141

.M45 2007. URL http://www.loc.gov/catdir/toc/ecip071/
2006030482.html;   http://www.southwestpeoplepix.com/
;   http://www.sunstonepress.com/cgi-bin/bookview.cgi?
_recordnum=413.

Amacher:2002:NBF

[Ama02] John Amacher. The Nazi bomb: Failures of the German nuclear program. Report, University of California, Santa Barbara, Santa Barbara, CA, USA, 2002. URL http:

//www.history.ucsb.edu/projects/holocaust/Research/
Proseminar/johnamacher.htm.

Arpino:1977:MPN

[AMB^{+}77] Giovanni Arpino, Eugenio Montale, Arrigo Bongiorno, et al. Montale premio Nobel, volume 1 of Studi e testimonianze. M. Boni, Bologna, Italy, 1977. xxv + 77 pp. LCCN ????

Ambrose:1997:AW

[Amb97] Stephen E. Ambrose. Americans at war. University Press of Mississippi, Jackson, MS, USA, 1997. ISBN 1-57806-026-5 (hardcover). xiii + 201 pp. LCCN E181 .A34 1997.

Ames:1920:ELG

[Ame20] J. S. Ames. Einstein’s law of gravitation. Science, 51(1315): 253–261, March 12, 1920. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/51/1315/253.full.pdf.

Amis:1987:EM

[Ami87] Martin Amis. Einstein’s Monsters. Harmony Books, New York, NY, USA, 1987. ISBN 0-517-56520-X. 149 pp. LCCN PR6051.M5 E36 1987.

Amis:2004:VA

[Ami04] Martin Amis. Vintage Amis. Vintage readers. Vintage Books, New York, NY, USA, 2004. ISBN 1-4000-3399-3 (paperback). 197 pp. LCCN PR6051.M5 A6 2004. URL http://

www.loc.gov/catdir/description/random044/2003057552.
html;   http://www.loc.gov/catdir/samples/random042/
2003057552.html.

Aerts:1999:SAR

[AMM99] Diederik Aerts, Ernest Mathijs, and Bert Mosselmans, editors. Science and art: the red book of “Einstein meets Magritte”, volume 2 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 0-7923-5758-2 (vol. 2, Kluwer), 0-7923-5765-5 (set, Kluwer), 90-5487-227-6 (vol. 2, VUB), 90-5487-234-9 (set, VUB). LCCN N72.S3 S24 1999. See other volumes in this series [Aer99a, Aer99b, ABW99, AVvdV99, BAM99, CSvB99, HBR99].

Amoroso:2010:CMB

[Amo10] Richard L. Amoroso, editor. Complementarity of mind and body: realizing the dream of Descartes, Einstein and Eccles. World philosophy series. Nova Science Publishers, New York, NY, USA, 2010. ISBN 1-61668-203-5 (hardcover). xix + 362 pp. LCCN BD418.3 .C624 2010.

Amrine:1959:GDS

[Amr59] Michael Amrine. The great decision: the secret history of the atomic bomb. Putnam, New York, NY, USA, 1959. 251 pp. LCCN D767.2 A5.

Aitchison:2004:UHM

[AMS04] Ian J. R. Aitchison, David A. MacManus, and Thomas M. Snyder. Understanding Heisenberg’s “magical” paper of July 1925: a new look at the calculational details. American Journal of Physics, 72(11):1370–1379, November 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v72/i11/p1370_s1.

Anninos:1996:NRB

[AMSmS96] Peter Anninos, Joan Massó, Edward Seidel, and Wai mo Suen. Numerical Relativity and black holes. Physics World, 9(7):43– 48, July 1996. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/9/7/phwv9i7a26.pdf;   http://www.
nobelprize.org/nobel_prizes/physics/laureates/1993/.

Arnowitt:1979:SFS

[AN79] R. Arnowitt and Pran Nath. Supersymmetry formulated in superspace. In Perlmutter and Scott [PS79], pages 103–125. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178 .O72 1979. URL http://link.

springer.com/chapter/10.1007/978-1-4684-3596-2_5/.

Anderson:1933:PE

[And33] Carl D. Anderson. The positive electron. Physical Review, 43(6): 491–494, March 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/ abstract/PR/v43/i6/p491_1. This paper reports the discovery of the positron, for which Anderson received the 1936 Nobel Prize in Physics.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/April 1973. ISSN 0041-9508. URL http://library.

ucsd.edu/dc/object/bb3960604v.

Anderson:1981:BRB

[And81] David L. Anderson. Book reviews: My Life: Recollections of a Nobel Laureate, by Max Born and Memoirs of a Physicist in the Atomic Age, by Walter M. Elsasser. American Journal of Physics, 49(1):94–95, January 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v49/i1/p94_s1.

Anderson:2015:AER

[And15] Jennifer Joline Anderson. Albert Einstein: revolutionary physicist. Great minds of science. Abdo Publishing Company, Minneapolis, MN, USA, 2015. ISBN 1-62403-379-2, 1-62968-479-1 (e-book). 48 pp. LCCN QC16.E5 A72 2015eb.

Anderson:2017:AER

[And17] Mark Anderson. Albert Einstein — refrigerator technician? [resources history]. IEEE Spectrum, 54(11):19, November 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Angersbach:1913:REB

[Ang13] Adam Angersbach. Das Relativitätsprinzip in elementarer Behandlung. (German) [The Principle of Relativity in an elementary treatment]. A. Cramer, Weilburg, Germany, 1913. ???? pp. LCCN ????

Angersbach:1920:RLE

[Ang20] Adam Angersbach. Das Relativitätsprinzip leichtfasslich entwickelt. (German) [The Principle of Relativity easily comprehensively developed], volume 39 of Mathematisch-physikalische Bibliothek. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1920. 57 pp. LCCN ????

Angell:1978:PNP

[Ang78] Norman Angell. Perspective: Nobel Peace Laureates: 1933. Bulletin of the Atomic Scientists, 34(10):1, December 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 12 June 1935, Oslo, Norway.

Angelidis:2005:EC

[Ang05] Thomas D. Angelidis. Einstein confusion. Physics World, 18(4):20–21, April 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/4/phwv18i4a29.
pdf.

Ankowitsch:2015:WEN

[Ank15] Christian Ankowitsch. Warum Einstein niemals Socken trug: Wie scheinbar Nebensächliches unser Denken beeinflusst. (German) [Why Einstein never wore socks: how such apparent trivialities influenced our thinking]. Rowohlt, Reinbek bei Hamburg, Germany, second edition, 2015. ISBN 3-87134-793-0. 224 pp. LCCN ????

Anonymous:1897:NBSb

[Ano97] Anonymous. The Nobel Bequest to Science. Scientific American, 77(2):18, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v77/n2/
pdf/scientificamerican07101897-18.pdf.

Anonymous:1898:LNN

[Ano98] Anonymous. The latest news of the Nobel Bequest. Scientific American, 79(19):290, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v79/n19/
pdf/scientificamerican11051898-290c.pdf.

Anonymous:1900:NPS

[Ano00] Anonymous. The Nobel Prizes for scientific discoveries. Scientific American, 83(15):227, October 13, 1900. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v83/n15/pdf/scientificamerican10131900-227a.
pdf.

Anonymous:1903:DNP

[Ano03] Anonymous. Distribution of the Nobel Prizes. Scientific American, 89(26):479, December 26, 1903. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v89/n26/
pdf/scientificamerican12261903-479e.pdf.

Anonymous:1911:ANP

[Ano11a] Anonymous. Award of the Nobel Prize to Madame Curie. Scientific American, 105(21):467, November 18, 1911. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v105/n21/pdf/scientificamerican11181911-467.
pdf.

Anonymous:1911:VEG

[Ano11b] Anonymous. Vortrag Einstein. (German) [Einstein lecture]. Bohemia, ??(141):10, May 23, 1911.

Anonymous:1912:PSRb

[Ano12] Anonymous. Physical safeguards in railroad travel, the Nobel Prize awarded to Dr. Carrel, and more. Scientific American, 107(17):342–343, October 26, 1912. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v107/n17/
pdf/scientificamerican10261912-342.pdf.

Anonymous:1917:RRE

[Ano17] Anonymous. On the relativity of rotation in Einstein’s theory. Scientific American, 84:102, August 18, 1917. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v84/n2172supp/pdf/scientificamerican08181917-
102csupp.pdf.

Anonymous:1919:EEH

[Ano19a] Anonymous. Einstein expounds his new theory: it discards absolute time and space, recognizing them only as related to moving systems. New York Times, ??(??):19, December 3, 1919. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/100464356/13EDE115846CF57707/8.

Anonymous:1919:LAA

[Ano19b] Anonymous. Lights all askew in the heavens: Men of science more or less agog over results of eclipse observations. Einstein theory triumphs. Stars not where they seemed or were calculated to be, but nobody need worry. A book for 12 wise men. No more in all the world could comprehend it, said Einstein when his daring publishers accepted it. New York Times, November 9, 1919. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/

docview/100454918. According to [Pai94, pages 146–147], this is the first mention of Albert Einstein in the New York Times. However, now that electronic archives are available (Pais had to work from printed indexes), we know that to be incorrect; see entry [Cam18]. There were further Einstein reports in that newspaper on 11 and 16 November 1919 that have yet to be located.

Anonymous:1919:RTE

[Ano19c] Anonymous. Results of the total eclipse of May 29 [1919]. London Times, November 7, 1919.

Anonymous:1919:RSE

[Ano19d] Anonymous. Revolution in science: Einstein versus Newton. London Times, November 8, 1919.

Anonymous:1920:ELE

[Ano20a] Anonymous. (1) Easy Lessons in Einstein: A Discussion of the More Intelligible Features of the Theory of Relativity; (2) From Newton to Einstein: Changing Conceptions of the Universe; (3) Die Einsteinsche Relativitätstheorie. Nature, 106 (2667):466, December 9, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v106/n2667/pdf/106466b0.pdf.

Anonymous:1920:EWB

[Ano20b] Anonymous. Einstein wil Berlijn verlaten. (Dutch) [Einstein wants to leave Berlin]. Algemeen Handelsblad, ??(??):??, August 27, 1920. Evening Edition.

Anonymous:1920:PDE

[Ano20c] Anonymous. Prof. dr. Einstein. (Dutch) [Professor Dr. Einstein]. Nieuwe Rotterdamsche Courant, ??(??):??, August 27, 1920. Evening Edition.

Anonymous:1920:PEJ

[Ano20d] Anonymous. The progress of the Einstein judges, are we to become the war pace-maker?, and more. Scientific American, 123(26):626–627, December 25, 1920. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v123/n26/
pdf/scientificamerican12251920-626.pdf.

Anonymous:1920:TEV

[Ano20e] Anonymous. Tumultszenen bei einer Einstein-Vorlesung. (German) [Uproar in Einstein’s lecture]. 8-Uhr-Abendblatt, ??(??): ??, February 13, 1920.

Anonymous:1920:WBS

[Ano20f] Anonymous. What has been said about Einstein. Scientific American, 123(9):200–201, August 28, 1920. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v123/n9/pdf/scientificamerican08281920-200.
pdf.

Anonymous:1921:SNE

[Ano21a] Anonymous. “A Strong Navy”, Einstein’s finite Universe, and more. Scientific American, 124(11):202–203, March 12, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v124/n11/pdf/scientificamerican03121921-202.
pdf.

Anonymous:1921:AE

[Ano21b] Anonymous. An alternative to Einstein. Scientific American, 124(24):468, June 11, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v124/n24/
pdf/scientificamerican06111921-468.pdf.

Anonymous:1921:EA

[Ano21c] Anonymous. The Einstein Award. Scientific American, 3: 99, February 1, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v3/n2msupp/pdf/scientificamerican02011921-
99asupp.pdf.

Anonymous:1921:EEF

[Ano21d] Anonymous. Einstein explains fallacy of ether: Space without electromagnetic fields has no characteristics, he tells students. Humorous at City College. Talks about “the greatest pain of physics” in exposition of his Relativity theory. “Special Relativity’s” action. How Planck’s theory works. New York Times, ??(??):12, April 21, 1921. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/98436746/.

Anonymous:1921:EST

[Ano21e] Anonymous. Einstein über seine Theorie: Vortrag in der ‘Urania’. (German) [Einstein on his theory: Lecture in the ‘Urania’]. Prager Tagblatt, ??(6):3, January 8, 1921.

Anonymous:1921:ETR

[Ano21f] Anonymous. “Einstein’s theories of Relativity and gravitation”. Scientific American, 124(21):409, May 21, 1921. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v124/n21/pdf/scientificamerican05211921-409a.
pdf.

Anonymous:1921:ECI

[Ano21g] Anonymous. From the Einstein Contest — IV. Scientific American, 124(12):226, March 19, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v124/n12/
pdf/scientificamerican03191921-226.pdf.

Anonymous:1921:LPE

[Ano21h] Anonymous. Lectures by Professor Einstein. Science, 53(1373): 382, April 22, 1921. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/53/1373/382.1.full.pdf.

Anonymous:1921:PES

[Ano21i] Anonymous. Prof. Einstein über seine Theorie: Der erste Vortrag: 7. Jänner. (German) [Prof. Einstein on his theory: The first lecture: January 7th]. Deutsche Zeitung Bohemia, 94(7): 2–3, January 9, 1921.

Anonymous:1921:PEL

[Ano21j] Anonymous. Prof. Einstein’s lectures at King’s College, London, and the University of Manchester. Nature, 107(2694):504, June 16, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v107/n2694/pdf/107504a0.pdf.

Anonymous:1921:RPRbd

[Ano21k] Anonymous. Recent publications: Reviews: Grundzüge der Einsteinschen Relativitätstheorie, by August Kopff. American Mathematical Monthly, 28(11/12):456–457, November/ December 1921. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:1921:R

[Ano21l] Anonymous. Relativity. Scientific American, 124(6):106– 107, February 5, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v124/n6/pdf/scientificamerican02051921-
106.pdf.

Anonymous:1921:SEC

[Ano21m] Anonymous. Some Einstein Contest personalities. Scientific American, 124(6):107, February 5, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v124/n6/
pdf/scientificamerican02051921-107.pdf.

Anonymous:1922:RUP

[Ano22a] Anonymous. Relativity and the Universe: a Popular Introduction into Einstein’s Theory of Space and Time; The Ideas of Einstein’s Theory: A Theory of Relativity in Simple Language; An Introduction to the Theory of Relativity; Relativity and Gravitation; The Rudiments of Relativity: Lectures delivered under the Auspices of the University College, Johannesburg Scientific Society; Die Einsteinsche Gravitationstheorie: Versuch einer allgemein verstandlichen Darstellung der Theorie (German) [Einstein’s theory of gravitation: Test of a commonly understood representation of the theory]. Nature, 109(2739):544–545, April 29, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v109/n2739/pdf/109544a0.pdf.

Anonymous:1922:E

[Ano22b] Anonymous. Der Einsteinfilm. Die Umschau, 26(16):247–249, April 16, 1922.

Anonymous:1922:ET

[Ano22c] Anonymous. Einstein’s theories. Nature, 110(2759):398, October 16, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/
journal/v110/n2759/pdf/110398a0.pdf.

Anonymous:1922:RF

[Ano22d] Anonymous. Relativity in the films. Scientific American, 127(2):92, August 1922. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v127/n2/
pdf/scientificamerican0822-92.pdf.

Anonymous:1922:VEF

[Ano22e] Anonymous. Vérification expérimentale de la formule de Lorentz–Einstein. Nature, 109(2735):406–407, April 1, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/
v109/n2735/pdf/109406a0.pdf.

Anonymous:1923:MTR

[Ano23a] Anonymous. The Mathematical Theory of Relativity; The Principle of Relativity with Applications to Physical Science; The Meaning of Relativity: Four Lectures delivered at Princeton University, May 1921; Modern Electrical Theory: Supplementary Chapters; La Théorie de la relativité d’Einstein et ses bases physiques: exposé élémentaire; The General Principle of Relativity in its Philosophical and Historical Aspect; The Theory of General Relativity and Gravitation: Based on a Course of Lectures delivered at the Conference on Recent Advances in Physics held at the University of Toronto, in January 1921; The Mathematical Theory of Relativity; Vector Analysis and the Theory of Relativity; L’Évidence de la théorie d’Einstein. Nature, 111 (2795):697–699, May 26, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v111/n2795/pdf/111697a0.pdf.

Anonymous:1923:CET

[Ano23b] Anonymous. Confirmation of the Einstein theory. Science, 58(1512):517, December 21, 1923. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/58/1512/517.1.full.pdf.

Anonymous:1923:EZJ

[Ano23c] Anonymous. Die Ersten zehn Jahre der Theorie von Niels Bohr über den Bau der Atome. (German) [The first ten years of Niels Bohr’s theory of the structure of atoms]. Die Naturwissenschaften, 11(27):536–624, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Anonymous:1923:EPK

[Ano23d] Anonymous. Einstein and the philosophies of Kant and Mach. Nature, 112(2807):253, August 18, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v112/n2807/
pdf/112253a0.pdf.

Anonymous:1923:EAT

[Ano23e] Anonymous. Einstein to announce theory ‘surpassing even relativity’. New York Times, ??(??):1, March 22, 1923. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
103150807/.

Anonymous:1924:ERE

[Ano24a] Anonymous. Einstein and the recent eclipse. Scientific American, 130(2):95, February 1924. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v130/n2/
pdf/scientificamerican0224-95c.pdf.

Anonymous:1924:RMP

[Ano24b] Anonymous. Relativity and modern physics. Scientific American, 130(3):170, March 1924. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v130/n3/
pdf/scientificamerican0324-170a.pdf.

Anonymous:1925:ETR

[Ano25a] Anonymous. (1) Einstein’s Theory of Relativity; (2) Space and Time: an Experimental Physicist’s Conception of these Ideas and of their Alteration; (3) The Theory of Relativity. Nature, 115 (2890):408–410, March 21, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v115/n2890/pdf/115408a0.pdf.

Anonymous:1925:EDA

[Ano25b] Anonymous. Einstein dio ayer su primera conferencia en la Universidad. (Spanish) [Einstein yesterday gave his first lecture at the University]. La Nación [Buenos Aires??], ??(??):4, March 28, 1925.

Anonymous:1925:EES

[Ano25c] Anonymous. Einstein expuso sus teorías en la Facultad de Ciencias Exactas. (Spanish) [Einstein presented his theories at the Faculty of Sciences]. La Nación [Buenos Aires??], ??(??):??, March 29, 1925.

Anonymous:1925:EDE

[Ano25d] Anonymous. Ether drift experiments and the Einstein theory. Science, 61(1586):x, May 22, 1925. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/61/1586/x.1.full.pdf.

Anonymous:1926:ESM

[Ano26a] Anonymous. (1) Electricity and the Structure of Matter; (2) The Story of Electricity from Thales to Einstein; (3) Readable School Electricity; (4) The Electron: its Isolation and Measurement and the Determination of some of its Properties; (5) Die Evolution des Geistes der Physik 1873–1923. Nature, 117(2931): 8–9, January 2, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v117/n2931/pdf/117008a0.pdf.

Anonymous:1926:PCM

[Ano26b] Anonymous. Presentation of the Copley Medal of the Royal Society to Professor Einstein. Science, 63(1618):11–12, January 1, 1926. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/63/
1618/11.full.pdf.

Anonymous:1927:PEN

[Ano27] Anonymous. Professor Einstein on Newton. Science, 65(1684): 347–348, April 8, 1927. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/65/1684/347.full.pdf.

Anonymous:1928:MTR

[Ano28] Anonymous. (1) The Mathematical Theory of Relativity; (2) The Einstein Delusion and other Essays. Nature, 121(3060): 979–980, June 23, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v121/n3060/pdf/121979a0.pdf.

Anonymous:1929:DPL

[Ano29a] Anonymous. Dr. Planck lauds Einstein: Professor receives first medal cast in Berlin in honor of scientist. New York Times, ??(??):15, June 30, 1929. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/104972983/. Re-

port of 29 June 1930 of first award of Planck Medal, to Albert Einstein.

Anonymous:1929:EER

[Ano29b] Anonymous. Einstein extends relativity theory. New work seeks to “unite laws of field of gravitation and electro-magnetism.” He calls it his greatest “book,” consisting or only five pages, took Berlin scientist ten years to prepare. Origin of matter theory seen. New York Times, ??(??):1, 2, January 12, 1929. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/105133961/.

Anonymous:1929:EVN

[Ano29c] Anonymous. Einstein is viewed as near the mystic. New York Times, ??(??):??, February 4, 1929. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1929:ERA

[Ano29d] Anonymous. Einstein reduces all physics to 1 law. The new electro-gravitational theory links all phenomena, says Berlin interpreter. Only one substance also. Hypothesis opens visions of persons being able to float in air, says N.Y.U. professor.

New York Times, ??(??):1, 5, January 25, 1929. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
105048956/.

Anonymous:1929:ENT

[Ano29e] Anonymous. Einstein’s new theory. Science, 69(1779):xlvi, February 1, 1929. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/69/1779/xlvi.1.full.pdf.

Anonymous:1929:NV

[Ano29f] Anonymous. News and views. Nature, 123(3092):174–179, February 2, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/123174a0.

Anonymous:1929:EBC

[Ano29g] Anonymous. THE Einstein birthday celebration. Science, 69(1791):447–448, April 26, 1929. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/69/1791/447.2.full.pdf.

Anonymous:1930:PEA

[Ano30a] Anonymous. Professor Einstein’s address at the University of Nottingham. Science, 71(1850):608–610, June 13, 1930. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/71/
1850/608.full.pdf.

Anonymous:1930:WMG

[Ano30b] Anonymous. We may not “get” Relativity, but we like Einstein. Literary Digest, ??(??):29–30, December 27, 1930. URL http:

//www.unz.org/Pub/LiteraryDigest-1930dec27-00029.

Anonymous:1931:BFE

[Ano31a] Anonymous. Back of frontispiece: Einstein goes sailing. Scientific American, 144(3):151, March 1931. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v144/n3/
pdf/scientificamerican0331-151.pdf.

Anonymous:1931:ECN

[Ano31b] Anonymous. Einstein completes unified field theory: Presents to Pasadena scientists proposition taking in relativity and all other factors. New York Times, ??(??):17, January 23, 1931. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/99498000/.

Anonymous:1931:NVa

[Ano31c] Anonymous. News and views. Nature, 127(3212):788–794, May 23, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
127764b0.

Anonymous:1931:NVb

[Ano31d] Anonymous. News and views. Nature, 127(3213):826–832, May 30, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v127/n3213/pdf/127826b0.pdf.

Anonymous:1931:PEO

[Ano31e] Anonymous. Professor Einstein at Oxford. The Spectator, ?? (5370):845, May 30, 1931.

Anonymous:1931:UEI

[Ano31f] Anonymous. University and educational intelligence. Nature, 127(3211):764–765, May 16, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/articles/127764b0.

Anonymous:1932:LPE

[Ano32] Anonymous. Lecture by Professor Einstein. Cambridge Review, 53(1310):382, May 13, 1932.

Anonymous:1933:APE

[Ano33a] Anonymous. Address by Professor Einstein. Science, 77(1994): 274–275, March 17, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/77/1994/274.full.pdf.

Anonymous:1933:GEC

[Ano33b] Anonymous. Guarding Einstein: cordon of police round the Albert Hall. Daily Mail, ??(??):14, October 4, 1933.

Anonymous:1933:NPG

[Ano33c] Anonymous. Nobel Prize goes to Bunin, Russian; short-story write and poet, exile in France, is first of his nation to get award; three physicists named; Profs. Dirac and Schroedinger share for 1933 — Heisenberg gets the 1932 grant. New York Times, ?? (??):23, November 10, 1933. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/100827281.

Anonymous:1933:PE

[Ano33d] Anonymous. Professor Einstein. Science, 77(2002):444, May 12, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/77/
2002/444.1.full.pdf.

Anonymous:1933:PEP

[Ano33e] Anonymous. Professor Einstein and the Prussian Academy. Science, 77(1997):346, April 7, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/77/1997/346.2.full.pdf.

Anonymous:1933:RPE

[Ano33f] Anonymous. Relativity. Professor Einstein at Glasgow. Gibson Foundation. The Scotsman, ??(??):10, June 21, 1933.

Anonymous:1933:RPG

[Ano33g] Anonymous. Rings of police guard Einstein meeting. Daily Herald, ??(??):1–2, October 4, 1933.

Anonymous:1933:WNF

[Ano33h] Anonymous. Weighing new-found atoms proves Einstein was right. Science News (Washington, DC), 24(638):10, July 1, 1933. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-

0930 (electronic). URL http://www.sciencenews.org/view/
url/id/302853/goto/Weighing_New-Found_Atoms_Proves_
Einstein_Was_Right;   https://www.jstor.org/stable/
3908936.

Anonymous:1934:EIP

[Ano34a] Anonymous. The Einstein Institute of Physics in Jerusalem. Science, 79(2045):224–225, March 9, 1934. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/79/2045/224.2.full.pdf.

Anonymous:1934:PDA

[Ano34b] Anonymous. Prof. Dirac appointed: Nobel Prize winner in physics gets advance study institute post. New York Times, ??(??):19, March 16, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1935:EAQ

[Ano35] Anonymous. Einstein attacks quantum theory: Scientist and two colleagues find it is not ‘complete’, even though ‘correct.’ See fuller one possible. Believe a whole description of ‘the physical reality’ can be provided eventually. New York Times, ??(??): 11, May 4, 1935. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/

docview/101522305. This news story is based on prepublication access (by 11 days) to the famous EPR article [EPR35]. See [Bag11, Chapter 16] for a discussion.

Anonymous:1936:NPR

[Ano36] Anonymous. Nobel Prize radioactivators. Scientific American, 154(2):83, February 1936. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v154/n2/
pdf/scientificamerican0236-83.pdf.

Anonymous:1938:MLM

[Ano38a] Anonymous. Miss Lise Meitner. Nature, 142(3602):856–866, November 12, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v142/n3602/pdf/142865d0.pdf.

Anonymous:1938:SNM

[Ano38b] Anonymous. Science: Neutron man. Time, ??(??):??, November 21, 1938. CODEN TYMEA9. ISSN 0040-

781X. URL http://www.time.com/time/magazine/article/

0,9171,760318,00.html. News of the award of the 1938 Nobel Prize for Physics to Enrico Fermi.

Anonymous:1939:EFU

[Ano39a] Anonymous. 6 elements found in uranium atom: Physicists bare discovery of greatest amount of energy liberated thus far: Report widely hailed: Professors Bohr and Fermi, at Columbia meeting, tell of atomic ‘cannon ball’. New York Times, ??

(??):17, February 25, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/102952660/.

Anonymous:1939:AEF

[Ano39b] Anonymous. Atom explosion frees 200,000,000 volts; new physics phenomenon credited to Hahn. New York Times, ?? (??):2, January 29, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/102763891. From

the article: “American scientists heard today of a new phenomenon in physics — explosion of atoms with a discharge of 200,000,000 volts of energy. … Dr. Enrico Fermi of the University of Rome told yesterday that this had been accomplished by Dr. G. [sic] Hahn of Berlin. … Scientists at the meeting said the discovery was comparable in significance to the original discovery of radioactivity thirty years ago.”.

Anonymous:1939:AER

[Ano39c] Anonymous. Atomic energy released. The Science News-Letter, 35(6):??, February 11, 1939. CODEN SNLEAI. ISSN 0096-

4018. URL http://www.sciencenews.org/view/feature/
id/199036/title/Atomic_Energy_Released. According to

[BHT86, page 212, column 1], this news story “appears to be the first connection between fission and the possibility of explosives made in print.” Flügge’s scientific article on that topic in Naturwissenschaften [Flü39] did appear until June 1939.

Anonymous:1939:BRE

[Ano39d] Anonymous. Book review: The Evolution of Physics by Albert Einstein and Leopold Infeld. Revue de métaphysique et de morale, 46(1):173, 1939. ISSN 0035-1571 (print), 2102-5177

(electronic). URL http://www.jstor.org/stable/40898826.

Anonymous:1939:DAE

[Ano39e] Anonymous. Dr. Albert Einstein. Science, 89(2307):242, March 17, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/89/
2307/242.1.full.pdf.

Anonymous:1939:END

[Ano39f] Anonymous. Einstein nears discovery of new theory. The Christian Science Monitor, ??(??):??, March 14, 1939. ISSN 0882-7729 (print), 1540-4617 (electronic).

Anonymous:1939:VEF

[Ano39g] Anonymous. Vast energy freed by uranium atom: Split, it produces 2 ‘cannonballs’, each of 100,000,000 electron volts: Hailed as epoch making; new process, announced at Columbia, uses only 1–30 volt to liberate big force. New York Times, ??(??):18, January 31, 1939. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/102759255/.

Anonymous:1939:VER

[Ano39h] Anonymous. Vision Earth rocked by isotope blast: Scientists say bit of uranium could wreck New York. New York Times, ?? (??), April 30, 1939. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/102790674/fulltextPDF.

Anonymous:1941:TWS

[Ano41] Anonymous. Third way to split atom is found by halving uranium and thorium: Scientists at University of California say cleavage creates much energy — Tokyo men also report uranium fission. New York Times, ??(??):17, March 3, 1941. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/105534449.

Anonymous:1945:EDL

[Ano45a] Anonymous. 400 experts decry lone atom policy; see ‘unending’ war: Los Alamos scientists assert efforts to keep bomb secret would be disastrous. Hold U.S. is vulnerable. Control by world authority is demanded — real defense is regarded as ‘uncertain’. New York Times, ??(??):1, 4, October 14, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107252091/.

Anonymous:1945:EAW

[Ano45b] Anonymous. Experts on atoms win Nobel Prizes: [Wolfgang] Pauli of Princeton Institute and [Otto] Hahn, who is believed to be here, are honored. Chemistry award to Finn. Gabriela Mistral of Chile gets Literature Prize, valued at about $28,950. New York Times, ??(??):19, November 16, 1945. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/107323828/pageviewPDF/13F38F649E969AA05FF.

Anonymous:1945:NPP

[Ano45c] Anonymous. Nobel Prize for Physics for 1945: Prof. Wolfgang Pauli. Nature, 156(3970):657–658, December 1, 1945. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v156/n3970/
pdf/156657c0.pdf.

Anonymous:1945:SRA

[Ano45d] Anonymous. Statements relating to the Atomic Bomb. His Majesty’s Stationery Office, London, UK, 1945. 23 + 1 pp. LCCN UF767 .G7 1945. URL http://www.osti.

gov/includes/opennet/includes/MED_scans/Book%20I%20-
General%20-%20Vol.%204-Chapter%208-Press%20Releases-
Part%202.pdf.

Anonymous:1946:EBQ

[Ano46a] Anonymous. Einstein’s 67th birthday quiet. New York Times, ??(??):12, March 15, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/107430896.

Anonymous:1946:SMD

[Ano46b] Anonymous. Shield minorities, Dr. Einstein pleads. New York Times, ??(??):26, March 28, 1946. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/hnpnewyorktimes/docview/
107367493.

Anonymous:1947:EFH

[Ano47a] Anonymous. Enrico Fermi, head-and-shoulders portrait, fac-

ing front, 1947. URL http://hdl.loc.gov/loc.pnp/cph.
3c22125.

Anonymous:1947:SES

[Ano47b] Anonymous. Science: Einstein stopped here. Time, ??(??): ??, February 10, 1947. CODEN TYMEA9. ISSN 0040-781X (print), 2169-1665 (electronic). URL http://content.time.

com/time/magazine/article/0,9171,854585,00.html.

Anonymous:1947:VL

[Ano47c] Anonymous. Voices of liberalism. Beacon Press, Boston, MA, USA, 1947. ISSN 0199-9540. ???? pp. LCCN BR1615 .V6. Two volumes.

Anonymous:1948:DEL

[Ano48a] Anonymous. Dr. Einstein’s letter. Bulletin of the Atomic Scientists, 4(11):332, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ein48g, Ano48c].

Anonymous:1948:MLB

[Ano48b] Anonymous. M. Louis de Broglie parle de l’œuvre d’Einstein. (French) [M. Louis de Broglie speaks of the works of Einstein]. Le Monde [Paris], ??(??):??, December 20, 1948. ISSN 0395-

2037. URL http://www.lemonde.fr/archives/article/
1949/12/20/m-louis-de-broglie-parle-de-l-ouvre-d-einstein_
1931475_1819218.html.

Anonymous:1948:PVE

[Ano48c] Anonymous. Polish version of Einstein’s message. Bulletin of the Atomic Scientists, 4(10):320, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The article begins: “This is a translation of the message attributed to Dr. Einstein in Poland. It is completely dissimilar to the actual message [Ein48g]. See explanation [Ano48a].

Anonymous:1949:AU

[Ano49a] Anonymous. The age of the universe. Reviews of Modern Physics, 21(3):374–378, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.374;
http://rmp.aps.org/abstract/RMP/v21/i3/p374_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Anonymous:1949:TPE

[Ano49b] Anonymous. Tribute to Prof. A. Einstein. Nature, 164(4177): 863, November 19, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v164/n4177/pdf/164863a0.pdf.

Anonymous:1950:BRI

[Ano50a] Anonymous. Book review: Infeld, Leopold. Albert Einstein: His work and its influence on our world. New York: Charles Scribner’s Sons, 1950. 132 p. $2.00. Science Education, 34(5): 333, 1950. CODEN SEDUAV. ISSN 1098-237X.

Anonymous:1950:NFA

[Ano50b] Anonymous. Nobelist Fermi awarded patent on atomic device. The Science News-Letter, 58(20):311, November 11, 1950. CO-DEN SNLEAI. ISSN 0096-4018. URL http://www.jstor.org/

stable/3927761.

Anonymous:1950:SLS

[Ano50c] Anonymous. Scientist’s leave studied: Einstein ex-associate seeks to teach again in Poland. New York Times, ??(??):10, March 17, 1950. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/111497409.

Anonymous:1951:NPH

[Ano51a]  Anonymous.  Der Nobelpreis für Physik: Heisenberg–

Schrödinger 1932 Dirac 1933/Chadwick 1935/1936 Hess– Anderson. Physikalische Blätter, 7(9):404–409, September 1951. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19510070904/abstract.

Anonymous:1951:EFS

[Ano51b] Anonymous. Enrico Fermi seated at control panel of a particle accelerator, the “world’s most powerful atom smasher”, 1951.

URL http://hdl.loc.gov/loc.pnp/cph.3c20917.

Anonymous:1952:NNc

[Ano52] Anonymous. News and notes. Bulletin of the Atomic Scientists, 8(9):309–312, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:EAD

[Ano54a] Anonymous. Einstein Award: Dr. R. P. Feynman. Nature, 173 (4403):524, March 20, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v173/n4403/pdf/173524a0.pdf.

Anonymous:1954:EAP

[Ano54b] Anonymous. Einstein Award to professor, 35: Feynman is known for his electrodynamics work — will receive $15,000. New York Times, ??(??):87, March 14, 1954. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
112897379/.

Anonymous:1954:EBO

[Ano54c] Anonymous. Einstein backs Oppenheimer. New York Times, ??(??):20–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/113110447.

Anonymous:1955:AEa

[Ano55a] Anonymous. Albert Einstein. Cechoslovackij fiziceskij zurnal, 5(4):435, December 1955. ISSN 0009-0700 (print), 1804-8536

(electronic). URL http://link.springer.com/article/10.
1007/BF01687206.

Anonymous:1955:AEb

[Ano55b] Anonymous. Albert Einstein. Die Naturwissenschaften, 42(15): 425, August 1, 1955. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF00599812.

Anonymous:1955:LBE

[Ano55c] Anonymous. Louis de Broglie: Einstein était le partisan résolu du déterminisme fondamental. (French) [Louis de Broglie: Einstein was the partisan of fundamental determinism]. Le Monde [Paris], ??(??):??, April 20, 1955.

ISSN 0395-2037. URL http://www.lemonde.fr/archives/
article/1955/04/20/louis-de-broglie-einstein-etait-
le-partisan-resolu-du-determinisme-fondamental_1959819_
1819218.html.

Anonymous:1955:ODA

[Ano55d] Anonymous. Obituary: Dr. Albert Einstein dies in sleep at 76; world mourns loss of great scientist. New York Times, April 19, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL http://www.nytimes.com/learning/
general/onthisday/bday/0314.html.

Anonymous:1956:NPP

[Ano56] Anonymous. Nobel Prize in Physics awarded to transistor inventors. The Bell System Technical Journal, 35(6):i–iv, November 1956. CODEN BSTJAN. ISSN 0005-8580. URL

http://bstj.bell-labs.com/BSTJ/images/Vol35/bstj35-
6-i.pdf;   http://www.alcatel-lucent.com/bstj/vol35-
1956/articles/bstj35-6-i.pdf.

Anonymous:1957:FC

[Ano57a] Anonymous. Fair comment? Bulletin of the Atomic Scientists, 13(7):250, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from an editorial What Goes on? in the Chicago Daily Tribune, 16 August 1957, objecting to the presence of a ‘Red’ Chinese scientist at the Pugwash Conference, and charging Bertrand Russell and Albert Einstein with having “a certain fondness for communism”.

Anonymous:1957:LFA

[Ano57b] Anonymous. Lettre de S. Freud à Albert Einstein. (French) [A letter of Sigmund Freud to Albert Einstein]. Revue française de psychanalyse, 21(6):757–768, November/December 1957. ISSN 0035-2942.

Anonymous:1959:AE

[Ano59] Anonymous. Albert Einstein. Zeitschrift für Physik, 154(4): i–??, ???? 1959. CODEN ZEPYAA. ISSN 0044-3328. URL

http://link.springer.com/article/10.1007/BF01328959.

Anonymous:1960:BRB

[Ano60] Anonymous. Book review: A Bibliographical Checklist and Index to the Published Writings of Albert Einstein. Compiled by Nell Boni, Monique Russ, and Dan H. Laurence. Pageant Books, Paterson, N.J., 1960. 84 pp. $6. Science, 132(3436): 1307, November 4, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/132/3436/1307.1.full.pdf.

Anonymous:1961:MGJ

[Ano61] Anonymous. Martin Grützmacher 60 Jahre/Nobelpreise 1961/ Max Planck-Medaille 1961/Preis VDPG 1961. Physikalische Blätter, 17(11):525–526, 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1962:RDG

[Ano62] Anonymous, editor. Recent Developments in General Relativity. Pergamon Press, London, UK, 1962. 472 pp. LCCN QC 6 .R295. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Anonymous:1963:ANN

[Ano63a] Anonymous. Awards: Nobelmen & Nobelwoman. Time, ??(??): ??, November 15, 1963. CODEN TYMEA9. ISSN 0040-781X. News of the award of the 1963 Nobel Prize in Physics to Eugene Wigner, Maria Goeppert Mayer and J. Hans D. Jensen.

Anonymous:1963:PPW

[Ano63b] Anonymous. The Pauling prize: A welcome honor from Norway. Bulletin of the Atomic Scientists, 19(10):18, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Don64].

Anonymous:1964:NLI

[Ano64] Anonymous, editor. Nobel lectures, including presentation speeches and laureates’ biographies. Physics, 1942–1962. Elsevier, Amsterdam, The Netherlands, 1964. xiii + 619 pp. LCCN ????

Anonymous:1965:BAA

[Ano65a] Anonymous. Birthplace of the atomic age. Los Alamos Chamber of Commerce, Los Alamos, NM, USA, 1965. ISBN ???? ???? pp. LCCN ????

Anonymous:1965:NLI

[Ano65b]  Anonymous, editor.  Nobel lectures: including presen-

tation speeches and laureates’ biographies: physics 1922– 1941. Elsevier, Amsterdam, The Netherlands, 1965. x + 456 pp. LCCN QC71 .P455 1965. URL http:/

/nobelprize.org/nobel_prizes/physics/laureates/1922/
bohr-lecture.html;   http://nobelprize.org/nobel_prizes/
physics/laureates/1922/bohr-lecture.pdf.

Anonymous:1965:NPP

[Ano65c] Anonymous. Nobel physics prize winner, October 21, 1965. 1 photographic print from a photograph taken in 1963.

Anonymous:1966:AE

[Ano66] Anonymous. Albert Einstein. I. Journal of the American Optometric Association, 37(9):893–??, September 1966. ISSN 0003-0244.

Anonymous:1967:SSS

[Ano67] Anonymous. Science et synthèse. Science and synthesis; an international colloquium organized by UNESCO on the tenth anniversary of the death of Albert Einstein and Teilhard de Chardin. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 0-387-05344-1. xvi + 206 pp. LCCN Q175 .S418. French contributions translated by Barbara M. Crook.

Anonymous:1968:LIP

[Ano68a] Anonymous. Leopold Infeld, physicist, dead: Pole worked with Einstein – directed Warsaw Institute. New York Times, ?? (??):51, January 17, 1968. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/118493262/.

Anonymous:1968:OHN

[Ano68b] Anonymous. Otto Hahn, Nobel winner, dies; discoverer of nuclear fission. New York Times, ??(??):1, 20, July 29, 1968. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/118497858/fulltextPDF. From the article: “Their

historic report [about the discovery of nuclear fission], which appeared two weeks later on Jan. 6, 1939, was a masterpiece of fence sitting. They presented their extraordinary results but refused to draw any conclusions from them, lest they offend the physicists. As chemists, they said in effect, they could report only their experimental observations. They stopped short of saying what the results meant — that the uranium nucleus had divided. ‘As nuclear chemists,’ the Hahn–Strassmann report read, ‘we cannot bring ourselves to take this step so contradictory to all the experience of nuclear physics,’ adding: ‘lt is possible that a number of rare accidents may have fooled us into making erroneous observations.’ From the end of the article: “He was fond of couching his feelings in humorous terms, such as this rhyme [limerick] he liked to recite: To smash the simple atom All mankind was intent Now any day The atom may Return the compliment.”.

Anonymous:1969:AGE

[Ano69a] Anonymous. Akademische Gedenkfeier zu Ehren von Otto Hahn und Lise Meitner am 21. Februar 1969 in Berlin. (German) [Academic memorial in honor of Otto Hahn and Lise Meitner on 21 February 1969 in Berlin]. Maz-Planck-Gesellschaft [zur Förderg. d. Wissenschaften], München, West Germany, 1969. ???? pp. LCCN QC15 .A38.

Anonymous:1969:PMS

[Ano69b] Anonymous. Physics: Moral support for Einstein. Nature, 222(5199):1117–1118, June 21, 1969. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v222/n5199/pdf/
2221117b0.pdf.

Anonymous:1969:PEB

[Ano69c] Anonymous. Planck und Einstein beantworten eine wissenschaftliche Rundfrage/Drei Fragen/Max Plancks Antwort vom 14.12.1930/Albert Einsteins Antwort vom 31.1.1930. (German) [Planck and Einstein to answer a scientific poll/three questions/Max Planck’s response of 12.14.1930/Albert Einstein’s reply of 01.31.1930]. Physikalische Blätter, 25(12):558, 1969. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1970:RER

[Ano70] Anonymous. Relativity: Einstein rules the waves. Nature, 226 (5252):1203, June 27, 1970. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v226/n5252/pdf/2261203b0.pdf.

Anonymous:1972:NLC

[Ano72] Anonymous. Nobel lectures — chemistry: including presentation speeches and laureates’ biographies, 1963–1970, volume 1963. Elsevier, Amsterdam, The Netherlands, 1972. ISBN 0-444-40987-4. x + 359 pp. LCCN ???? Published for the Nobel Foundation.

Anonymous:1975:PAE

[Ano75] Anonymous. Papers of Albert Einstein. Historia Mathematica, 2(1):73, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900452.

Anonymous:1977:BLS

[Ano77] Anonymous. Bell Laboratories scientist named Nobel Laureate. The Bell System Technical Journal, 56(9):i–iii, November 1977. CODEN BSTJAN. ISSN 0005-8580. URL

http://bstj.bell-labs.com/BSTJ/images/Vol56/bstj56-
9-i.pdf;   http://www.alcatel-lucent.com/bstj/vol56-
1977/articles/bstj56-9-i.pdf.

Anonymous:1978:BLS

[Ano78] Anonymous. Bell Laboratories scientists named 1978 Nobel Prize Laureates. The Bell System Technical Journal, 57(9):i–ii, November 1978. CODEN BSTJAN. ISSN 0005-8580. URL

http://bstj.bell-labs.com/BSTJ/images/Vol57/bstj57-
9-i.pdf;   http://www.alcatel-lucent.com/bstj/vol57-
1978/articles/bstj57-9-i.pdf.

Anonymous:1979:AET

[Ano79a] Anonymous. Al’bert Einstein i teorija gravitacii. (Russian) [Albert Einstein and gravitation theory]. Izdatel’stvo Mir, Moscow, USSR, 1979. 592 pp.

Anonymous:1979:AEP

[Ano79b] Anonymous. Albert Einstein, Paul Langevin and Georges Cogniot: Publication of three obscure texts — introduction. Pensée, ??(205):73, ???? 1979. ISSN 0031-4773.

Anonymous:1979:CAE

[Ano79c] Anonymous. Centennial of Albert Einstein’s birth. Soviet Physics Journal, 22(3):336–337, March 1979. ISSN 0038-5697 (print), 1573-9228 (electronic). URL http://link.springer.

com/article/10.1007/BF00889884.

Anonymous:1979:EE

[Ano79d] Anonymous. Einstein: E = mc^{2}. Centre Georges Pompidou, Bibliothèque publique d’information, Paris, France, 1979. ISBN 2-85850-016-9. 40 pp. LCCN QC16.E5 E44 1979.

Anonymous:1979:MWG

[Ano79e] Anonymous. Man who guarded Einstein in Norfolk. Eastern Daily Press, ??(??):??, March 22, 1979.

Anonymous:1979:NCA

[Ano79f] Anonymous. Notes and comment: Albert Einstein Centennial. The New Yorker, 55(3):28–29, ???? 1979. ISSN 0028-792X.

Anonymous:1979:TDA

[Ano79g] Anonymous. Terminänderung!/Deutsche Arbeitsgemeinschaft Akustik (DAGA)/Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner und Max von Laue/Einstein-Symposion Berlin/Hubert-Schardin-Medaille verliehen/Broschüre zum Stellungswechsel. (German) [Date change! /German Association for Acoustics (DAGA) /celebration of the 100th birthdays of Albert Einstein, Otto Hahn, Lise Meitner and Max von Laue /Einstein Symposium Berlin /Hubert Schardin medal awarded /Brochure on the change of position]. Physik Journal, 35(1):43–46, January 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1979:GAE

[Ano79h] Anonymous. Zum 100. Geburtstag von Albert Einstein, Max v. Laue, Otto Hahn und Lise Meitner. (German) [On 100th birthdays …]. Physik Journal, 35(5):189–191, May 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1980:AEM

[Ano80a] Anonymous. Albert Einstein, 1879–1955: mémorial Albert Einstein publié à l’occasion du centième anniversaire de sa naissance. (French) [Albert Einstein, 1879–1955: Albert Einstein memorial published on the occasion of the one hundredth anniversary of his birth]. L’Académie, Bruxelles, Belgique, 1980. ISBN 2-8031-0020-7. 138 + xxviii pp. LCCN QC16.E5 A646. 300F.

Anonymous:1980:ESP

[Ano80b] Anonymous. Einstein session of the Pontifical Academy. Science, 207(4436):1159–1161, March 14, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/207/4436/1159.full.pdf.

Anonymous:1981:AEP

[Ano81a] Anonymous. Albert Einstein papers are now open to qualified scholars at the Princeton University Library. Applied Optics, 20 (14):2521, July 15, 1981. CODEN APOPAI. ISSN 0003-6935.

Anonymous:1981:GPR

[Ano81b] Anonymous. German physics in row about Einstein. Nature, 290 (5807):535, April 16, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v290/n5807/pdf/290535b0.pdf.

Anonymous:1981:HNP

[Ano81c] Anonymous. Hiroshima and Nagasaki, the physical, medical, and social effects of the atomic bombings, volume CN 5088 of Harper Colophon Books. Basic Books, New York, NY, USA, 1981. ISBN 0-465-02985-X, 0-465-02987-6 (paperback). xlv + 706 pp. LCCN D767.25.H6 H6713 1981. URL http://www.

gbv.de/dms/bowker/toc/9780465029853.pdf.

Anonymous:1981:PCA

[Ano81d] Anonymous, editor. Proceedings of the Celebration of the 50th Anniversary of the Lawrence Berkeley Laboratory: Symposium and Banquet Speeches October 1981. Lawrence Berkeley Laboratory, National Technical Information Service U.S. Department of Commerce, Springfield, VA 22161, USA, 1981. LCCN ????

URL http://escholarship.org/uc/item/9772r6gq. Report

number LBL-13613.

Anonymous:1981:SBD

[Ano81e] Anonymous. Science and belief: from Darwin to Einstein. Block 4: Unit 6–9, Modern physics and problems of knowledge. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11003-7. 153 pp. LCCN QC6 .M513 1981.

Anonymous:1982:ALN

[Ano82a] Anonymous. Appendix I — list of Nobel Laureates, 1901–1930. In Bernhard et al. [BCS82], pages 404–405. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500311.

Anonymous:1982:AIM

[Ano82b] Anonymous. Appendix II — members of the Nobel Committees for physics, chemistry and physiology or medicine elected 1900–1930. In Bernhard et al. [BCS82], pages 406–409. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500323.

Anonymous:1982:ECW

[Ano82c] Anonymous. Einstein centennial works reviewed here. Science, 216(4548):847, May 21, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/216/4548/847.full.pdf.

Anonymous:1982:EFS

[Ano82d] Anonymous. Enrico Fermi: significato di una scoperta: il navigatore italiano sbarcato nel nuovo mondo. (Italian) [Enrico Fermi: The Meaning of a Discovery. A Disembarked Italian Navigator in the New World]. Forum italiano dell’energia nucleare, Roma, Italy, 1982. 121 pp. LCCN QC16.F46 E57 1982.

Anonymous:1983:AEH

[Ano83] Anonymous. Albert Einstein — historical and cultural perspectives — G. Holton, Y. Elkana. Science, Technology and Human Values, ??(43):66, ???? 1983. ISSN 0162-2439 (print),1552-8251 (electronic).

Anonymous:1984:BRBa

[Ano84a] Anonymous. Book review: Elie Cartan — Albert Einstein, letters on absolute parallelism 1929–1932: Original text, English translation by Jules Leroy and Jim Ritter; edited by Robert Debever. Princeton, NJ (Princeton Univ. Press). 1979. xvii + 233 pp. Historia Mathematica, 11(1):94–96, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086084900119.

Anonymous:1984:BREb

[Ano84b] Anonymous. Book review: The 1979 Einstein centennial: Albert Einstein: Historical and Cultural Perspectives. Journal for the History of Astronomy, 15(3):222, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:BREa

[Ano84c] Anonymous. Book review: The 1979 Einstein centennial: Some Strangeness in the Proportion: a Centennial Celebration to Celebrate the Achievements of Albert Einstein, to Fulfill a Vision: Jerusalem Einstein Centennial Symposium on Gauge Theories and Unification of Physical Forces. Journal for the History of Astronomy, 15(3):222, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1985:ENP

[Ano85] Anonymous. From the editors: Nobel for physicians. Bulletin of the Atomic Scientists, 41(11):2, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cal86].

Anonymous:1986:ABE

[Ano86] Anonymous. Los Alamos, beginning of an era, 1943–1945. Los Alamos National Laboratory, Los Alamos, NM, USA, July 1986. 59 pp. LCCN QC773.3.U5 L67 1986.

Anonymous:1987:AEL

[Ano87a] Anonymous. Albert Einstein looks for a job. The Scientist (Philadelphia, PA), 1(19):22–23, August 10, 1987. ISSN 0890-3670 (print), 1945-5127 (electronic). URL http://

www.the-scientist.com/?articles.view/articleNo/8871/.

Reprinted from the Collected Papers of Albert Einstein, Vol 1, 1987.

Anonymous:1987:NPB

[Ano87b] Anonymous. Nobel Prize for Bulletin sponsor: [John C. Polanyi, Yuan T. Lee, and Dudley Herschbach, in Chemistry]. Bulletin of the Atomic Scientists, 43(1):59–60, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WMS

[Ano88a] Anonymous. William M. Swartz, 1912–1987. Bulletin of the Atomic Scientists, 44(1), January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WS

[Ano88b] Anonymous. William Swartz, 1912–1987. Bulletin of the Atomic Scientists, 44(5):0, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:DEW

[Ano90a] Anonymous. Did Einstein’s wife contribute to his theories? New York Times, ??(??):C5, March 27, 1990. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
108569061/.

Anonymous:1990:ECA

[Ano90b] Anonymous. Einstein collection available. Science, 248(4956): 680, May 11, 1990. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/248/4956/680.3.full.pdf.

Anonymous:1991:PER

[Ano91] Anonymous. Proving Einstein right (or wrong). Science, 251(4996):870, February 22, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/251/4996/870.3.full.pdf.

Anonymous:1993:JBF

[Ano93a] Anonymous. Das Jahr 1945 Die Bohr-Festschrift und der Nobelpreis (German) [The year 1945: The Bohr Anniversary Publication and the Nobel Prize]. In von Meyenn [vM93], pages 253– 336. ISBN 3-540-54911-0. ISSN 0172-6315. LCCN QC16.P37

A34. URL http://www.springer.com/physics/book/978-3-
540-54911-6.

Anonymous:1993:OMP

[Ano93b] Anonymous. Österreichische Mathematik und Physik: Wolfgang Gröbner, Richard von Mises, Wolfgang Pauli. (German) [Austrian mathematics and physics: …]. Die Zentralbibliothek, Wien, Austria, 1993. ISBN 3-900490-03-1. 71 pp. LCCN QA27.A9 O87 1993.

Anonymous:1993:PAM

[Ano93c] Anonymous. Proceedings of the 14th Annual Meeting of the IUPS Commission on Gravitational Physiology, September 29– October 2, 1992. Dedicated to Albert Einstein 1897–1955. The Physiologist, 36(1S):S1–S172, February 1993. ISSN 0031-9376.

Anonymous:1994:CGZ

[Ano94a] Anonymous. Cover-up at ground zero: Atomic bomb testing and the “downwinders”. ABC News Production, New York, NY, USA 1994. 44 minute video.

Anonymous:1994:EOL

[Ano94b] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2(2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Anonymous:1995:ABF

[Ano95a] Anonymous. Atomic bomb fiftieth anniversary (special issue). The Journal of American History, 82(3):??, ???? 1995. ISSN 0021-8723 (print), 1945-2314 (electronic).

Anonymous:1995:E

[Ano95b] Anonymous. Einstein. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 3(1):121, December 1995. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02913707.

Anonymous:1995:LEG

[Ano95c] Anonymous. Letter from Einstein goes to auction. Nature, 374 (6523):581, April 13, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v374/n6523/pdf/374581c0.pdf.

Anonymous:1996:BTF

[Ano96a] Anonymous. Behind tall fences: stories and experiences about Los Alamos at its beginning. Los Alamos Historical Society, Los Alamos, NM, USA, 1996. ISBN 0-941232-91-3. xiv + 210 pp. LCCN QC773.3.U5 B43 1996.

Anonymous:1996:EFS

[Ano96b] Anonymous. Einstein fails to sell. Nature, 380(6571):193, March 21, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v380/n6571/pdf/380193b0.pdf.

Anonymous:1996:EE

[Ano96c] Anonymous. Einstein’s epistles. Science, 274(5293):1617–0, December 6, 1996. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/274/5293/1617.3.full.

Anonymous:1996:RDN

[Ano96d] Anonymous. Russia dumps its nuclear waste. Guppy love. Unmeltable ice. the Ig Nobels for 1996. Scientific American, 275 (6):20–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1296issue/1296currentissue.html.

Anonymous:1997:NPS

[Ano97a] Anonymous. The 1996 Nobel Prizes for Science: An overview of the outstanding work that took the honors in physics, chemistry, medicine and economics. Scientific American, 276(1): 14–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Anonymous:1997:BAE

[Ano97b] Anonymous. The Bible according to Einstein: a scientific complement to the Holy Bible for the Third Millennium. Jupiter Scientific Publishers, New York, NY, USA, 1997. ISBN 0-9655176-8-3. 634 pp. LCCN QC21.2 .B493 1997.

Anonymous:1997:ESA

[Ano97c] Anonymous. Einstein — a step ahead of himself. Science, 275(5297):163, January 10, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/275/5297/163.2.full.

Anonymous:1997:ELL

[Ano97d] Anonymous. Einstein’s legacy lives on. Physics World, 10(12):5, December 1997. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/10/12/phwv10i12a3.pdf.

Anonymous:1997:PNC

[Ano97e] Anonymous. Profile: Nobel chemist Mario Molina still faces skeptics over CFCs and ozone loss. Scientific American, 277(5): 40–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Anonymous:1998:THD

[Ano98] Anonymous, editor. Today in history. December 2. Library of Congress, Washington, DC, USA, 1998. LCCN QC173. URL

http://memory.loc.gov/ammem/today/dec02.html.

Anonymous:1999:MM

[Ano99a] Anonymous. 1000 makers of the millennium. DK Pub., New York, NY, USA, 1999. ISBN 0-7894-4709-6, 0-7513-5664-6. 256 pp. LCCN CT107 .A16 1999.

Anonymous:1999:BRCa

[Ano99b] Anonymous. Book review: The collected papers of Albert Einstein, vol 4, The Swiss years: Writings, 1912–1914. History of Science (UK), 37(117):369, September 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Anonymous:1999:BRCb

[Ano99c] Anonymous. Book review: The collected papers of Albert Einstein, vol 6, the Berlin years: Writings, 1914–1917. History of Science (UK), 37(116):245, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Anonymous:1999:C

[Ano99d] Anonymous. Chemistry, 1942–1962. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3407-4. xi + 710 pp. LCCN QD39 .C483 1999.

Anonymous:1999:L

[Ano99e] Anonymous. Ein limerick. Physics in Perspective (PIP), 1(2): 214, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:1999:SBN

[Ano99f] Anonymous. Special briefing: The 1998 Nobel Prizes in science. Scientific American, 280(1):16–19, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-16.pdf.

Anonymous:19xx:AEL

[Anoxxa] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.

lanl.gov/history/road/birthofmodernphysics.shtml. Sec-

tion The Einstein Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:19xx:BMP

[Anoxxb] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.

lanl.gov/history/road/birthofmodernphysics.shtml. Sec-

tion The Einstein Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:19xx:PRR

[Anoxxc] Anonymous. Physical Review records 1940–1947. 19xx. 3001 pp. LCCN MMC-3069.

Anonymous:2000:DMAa

[Ano00a] Anonymous. Driving Mr. Albert: A trip across America with Einstein’s brain. New York Times Book Review, ??(??):22, August 6, 2000. ISSN 0028-7806.

Anonymous:2000:DMAb

[Ano00b] Anonymous. Driving Mr. Albert: A trip across America with Einstein’s brain. Nature Medicine, 6(10):1090, October 2000. CODEN NAMEFI. ISSN 1078-8956 (print), 1546-170X (elec-

tronic). URL http://www.nature.com/nm/journal/v6/n10/
full/nm1000_1090.html.

Anonymous:2000:EHB

[Ano00c] Anonymous. Einstein, the H-bomb and whale harpoons. Scientific American, 282(4):14, April 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v282/n4/
pdf/scientificamerican0400-14.pdf.

Anonymous:2000:NP

[Ano00d] Anonymous. The Nobel Prizes for 2000. Scientific American, 283(6):36, December 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v283/n6/
pdf/scientificamerican1200-36b.pdf.

Anonymous:2000:OEB

[Ano00e] Anonymous. Obituary: Einstein’s biographer Pais dies. Physics World, 13(9):6, September 2000. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
13/9/phwv13i9a8.pdf.

Anonymous:2000:SBN

[Ano00f] Anonymous. Special briefing: The Nobel Prizes for 1999. Scientific American, 282(1):14–17, January 2000. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v282/n1/pdf/scientificamerican0100-14.pdf.

Anonymous:2000:USN

[Ano00g] Anonymous. United States nuclear tests: July 1945 through September 1992. Office of Scientific and Technical Information, Oak Ridge, TN, USA, 2000. xviii + 162 pp. LCCN U264.3

.U547 2000. URL http://www.doe.gov.bridge.

Anonymous:2001:YNL

[Ano01a] Anonymous. 100 years with Nobel laureates: 100 special articles written by Nobel laureates exclusively for Encyclopædia Britannica. Encyclopædia Britannica (India), New Delhi, India, international edition, 2001. ISBN 0-85229-794-7. xiii + 1098 pp. LCCN ????

Anonymous:2001:ADL

[Ano01b] Anonymous. Alsos Digital Library for Nuclear Issues. Washington and Lee University, Lexington, VA, USA, 2001. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://alsos.wlu.edu.

Anonymous:2001:EFS

[Ano01c] Anonymous. Enrico Fermi: significato di una scoperta. (Italian) [Enrico Fermi: the meaning of a discovery]. ENEA, Roma, Italy, revised and expanded edition, 2001. ISBN 88-8286-022-1. 184 pp. LCCN QC16.F46 E57 2001.

Anonymous:2001:PPE

[Ano01d] Anonymous. Perspectives on the prize: essays in commemoration of the first century of the Nobel Prizes. Minerva [Washington, DC], 39(4):363–464, 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Anonymous:2002:EBS

[Ano02a] Anonymous. Einstein brought up to speed. Nature, 419(6909): 783–784, October 24, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v419/n6909/full/419783a.html.

Anonymous:2002:NP

[Ano02b] Anonymous. The Nobel Prizes for 2002. Scientific American, 287(6):37, December 2002. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v287/n6/
pdf/scientificamerican1202-37a.pdf.

Anonymous:2003:EAG

[Ano03a] Anonymous. Einstein archives go online. Physics World, 16(6): 13, June 2003. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/16/6/phwv16i6a18.pdf.

Anonymous:2003:EYD

[Ano03b] Anonymous. Einstein not yet displaced. Nature, 422(6932): 563–564, April 10, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v422/n6932/full/422563a.html.

Anonymous:2003:EVG

[Ano03c] Anonymous. Einstein’s view of Germany. Nature, 422(6934): 811–812, April 24, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v422/n6934/full/422811a.html.

Anonymous:2004:AWA

[Ano04a] Anonymous. Albert would approve — Let’s throw every test in the book at Einstein’s curvy space–time. New Scientist, 184 (2470):3, October 23, 2004. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Anonymous:2004:BRBg

[Ano04b] Anonymous. Book review: Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. Isis, 95(2):344, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/
426281.

Anonymous:2004:EMC

[Ano04c] Anonymous. Einstein = man of conscience. Scientific American, 291(3):10, September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v291/n3/
pdf/scientificamerican0904-10.pdf.

Anonymous:2004:PMD

[Ano04d] Anonymous. From the profound to the mundane — diary provides a slice of life. Princeton Weekly Bulletin, 93(25):??, April

26, 2004. URL http://pr.princeton.edu/pwb/04/0426/6a.

shtml. From Conversations With Einstein by Johanna Fantova; summaries and translation by Alice Calaprice.

Anonymous:2004:LL

[Ano04e] Anonymous. Lorentz lives on. Physics World, 17(3):15, March 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/17/3/phwv17i3a20.pdf.

Anonymous:2004:NNE

[Ano04f] Anonymous. Net news: Einstein for the masses. Science, 306 (5703):1871, December 10, 2004. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/306/5703/1871.3.full.pdf.

Anonymous:2004:SIB

[Ano04g] Anonymous. Special issue: Beyond Einstein. Scientific American, 291(3), September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:2005:NP

[Ano05a] Anonymous. The 2005 Nobel Prizes. Scientific American, 293(6):34, December 2005. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v293/n6/
full/scientificamerican1205-34a.html;   http://www.
nature.com/scientificamerican/journal/v293/n6/pdf/scientificamerican1205-
34a.pdf.

Anonymous:2005:YAId

[Ano05b] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Nobel viruses

• noble gases • King of fertilizers. Scientific American, 292(4):18, April 2005. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v292/n4/full/
scientificamerican0405-18.html;   http://www.nature.
com/scientificamerican/journal/v292/n4/pdf/scientificamerican0405-
18.pdf.

Anonymous:2005:AEC

[Ano05c] Anonymous. Albert Einstein and 20th century physics. Chemie in unserer Zeit, 39(4):298, August 2005. ISSN 0009-2851.

Anonymous:2005:AES

[Ano05d] Anonymous. Albert Einstein, seen from philosophy in the International Year of Physics. Pensamiento, 61(231):355–361, September/December 2005. ISSN 0031-4749.

Anonymous:2005:AEG

[Ano05e] Anonymous. Albert Einstein to George Ellery Hale. Physics Today, 58(9):14, September 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEHb

[Ano05f] Anonymous. Albert Einstein to Hedi and Max Born. Physics Today, 58(11):14, November 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEHa

[Ano05g] Anonymous. Albert Einstein to Heinrich Zangger. Physics Today, 58(10):18, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEMa

[Ano05h] Anonymous. Albert Einstein to Max Born. Physics Today, 58 (5):16, May 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEMc

[Ano05i] Anonymous. Albert Einstein to Max Born. Physics Today, 58(8):16, August 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEMd

[Ano05j] Anonymous. Albert Einstein to Max Born (vol 58, pg 16, 2005). Physics Today, 58(9):17, September 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEMb

[Ano05k] Anonymous. Albert Einstein to Michele Besso. Physics Today, 58(7):14, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:AEP

[Ano05l] Anonymous. Albert Einstein to Paul Ehrenfest. Physics Today, 58(4):88, April 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:2005:BFI

[Ano05m] Anonymous. Basic Feynman: An interview with Michelle Feynman. Web site, 2005. Comments about the preparation of [Fey05].

Anonymous:2005:BHA

[Ano05n] Anonymous. A brief history of Albert Einstein. Physics World, 18(1):14, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a22.
pdf.

Anonymous:2005:CLB

[Ano05o] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (electronic). URL http:/

/www.larecherche.fr/idees/livres/il-etait-sept-fois-
revolution-albert-einstein-autres-01-05-2005-73924.

Anonymous:2005:DEM

[Ano05p] Anonymous. Dance: Einstein in motion. Nature, 435(7043): 741, June 8, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v435/n7043/full/435741a.html.

Anonymous:2005:EBY

[Ano05q] Anonymous. Education: Brush up your Einstein. Science, 310(5750):949, November 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/310/5750/949.1.full;   https:
//www.einstein-online.info/en.

Anonymous:2005:Em

[Ano05r] Anonymous. Einstein 100 not out. Physics World, 18(12):48, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/12/phwv18i12a41.pdf.

Anonymous:2005:Ea

[Ano05s] Anonymous. Einstein 2005. Physics World, 18(1):15, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/1/phwv18i1a23.pdf.

Anonymous:2005:Eb

[Ano05t] Anonymous. Einstein 2005. Physics World, 18(2):16, February 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/2/phwv18i2a22.pdf.

Anonymous:2005:Ec

[Ano05u] Anonymous. Einstein 2005. Physics World, 18(3):14, March 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/3/phwv18i3a18.pdf.

Anonymous:2005:Ed

[Ano05v] Anonymous. Einstein 2005. Physics World, 18(4):16, April 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/4/phwv18i4a23.pdf.

Anonymous:2005:Ee

[Ano05w] Anonymous. Einstein 2005. Physics World, 18(5):14, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/18/5/phwv18i5a23.pdf.

Anonymous:2005:Ef

[Ano05x] Anonymous. Einstein 2005. Physics World, 18(6):14, June 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/18/6/phwv18i6a22.pdf.

Anonymous:2005:Eg

[Ano05y] Anonymous. Einstein 2005. Physics World, 18(7):14, July 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/7/phwv18i7a21.pdf;   http://
physicsworldarchive.iop.org/full/pwa-pdf/18/8/phwv18i8a20.
pdf.

Anonymous:2005:Ei

[Ano05z] Anonymous. Einstein 2005. Physics World, 18(9):3, September 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/9/phwv18i9a1.pdf.

Anonymous:2005:Ej

[Ano05-27] Anonymous. Einstein 2005. Physics World, 18(10):3, October 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/10/phwv18i10a1.pdf.

Anonymous:2005:Ek

[Ano05-28] Anonymous. Einstein 2005. Physics World, 18(11):3, November 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/11/phwv18i11a1.pdf.

Anonymous:2005:El

[Ano05-29] Anonymous. Einstein 2005. Physics World, 18(12):3, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/12/phwv18i12a1.pdf.

Anonymous:2005:EEP

[Ano05-30] Anonymous. Einstein and the EPR Paradox. APS News

Online, November 2005. URL http://en.wikipedia.org/
wiki/EPR_paradox;   http://www.aps.org/apsnews/1105/
110516.cfm.

Anonymous:2005:EBM

[Ano05-31] Anonymous. Einstein books: More on Einstein. Nature, 435 (7041):427, May 26, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v435/n7041/full/435427a.html.

Anonymous:2005:ECb

[Ano05-32] Anonymous. Einstein calling. Physics World, 18(2):15, February 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/2/phwv18i2a21.pdf.

Anonymous:2005:ECa

[Ano05-33] Anonymous. The Einstein chronicles. Nature, 433(7023):195– 196, January 19, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v433/n7023/full/433195a.html.

Anonymous:2005:EDc

[Ano05-34] Anonymous. Einstein dumbed down. Physics World, 18(10):43, October 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/10/phwv18i10a39.pdf.

Anonymous:2005:EIC

[Ano05-35] Anonymous. Einstein intraoral camera. British Dental Journal, 199(12):791, December 24, 2005. CODEN BDJOHQ. ISSN ????

URL http://www.nature.com/bdj/journal/v199/n12/full/
4813099a.html.

Anonymous:2005:EDa

[Ano05-36] Anonymous. Einstein is dead. Nature, 433(7023):179, January 19, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v433/n7023/full/433179a.html.

Anonymous:2005:EDb

[Ano05-37] Anonymous. Einstein is dead. Nature Digest, 2(2):2–3, February 2005. CODEN ???? ISSN ???? URL http://www.nature.com/

ndigest/journal/v2/n2/pdf/ndigest.2005.050202.pdf.

Anonymous:2005:EQW

[Ano05-38] Anonymous. Einstein quiz winner. Physics World, 18(4):21, April 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/4/phwv18i4a31.pdf.

Anonymous:2005:EMW

[Ano05-39] Anonymous. Einstein the miracle worker. Physics World, 18(7): 43, July 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/7/phwv18i7a39.pdf.

Anonymous:2005:CBD

[Ano05-40] Anonymous. Le Canada et la bombe d’Hiroshima. Web site.,

August 1, 2005. URL http://nonviolence.ca/index.php/
le-canada-et-la-bombe-dhiroshima/.

Anonymous:2005:RAE

[Ano05-41] Anonymous. Read about Einstein year by year. Physics World, 18(2):43, February 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/2/phwv18i2a43.
pdf.

Anonymous:2005:RB

[Ano05-42] Anonymous. Relativity for beginners. Physics World, 18(5): 43, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/5/phwv18i5a43.pdf.

Anonymous:2005:SIE

[Ano05-43] Anonymous. Special issue: Einstein’s legacy. Science, 307 (5711):797–996, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:2005:WWEb

[Ano05-44] Anonymous. Warm words for Einstein dance. Physics World, 18 (6):10, June 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/7/phwv18i7a17.pdf.

Anonymous:2005:WWEa

[Ano05-45] Anonymous. Wise words from Einstein. Physics World, 18(5): 43, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/5/phwv18i5a42.pdf.

Anonymous:2006:NP

[Ano06a] Anonymous. The 2006 Nobel Prizes. Scientific American, 295(6):40, December 2006. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v295/n6/
full/scientificamerican1206-40d.html;   http://www.
nature.com/scientificamerican/journal/v295/n6/pdf/scientificamerican1206-
40d.pdf.

Anonymous:2006:CNG

[Ano06b] Anonymous. The controversy over Newton’s gravitational constant. Wed document., September 1, 2006. URL http:

//web.archive.org/web/20060901061147/;   http://www.
npl.washington.edu/eotwash/gconst.html.

Anonymous:2006:EM

[Ano06c] Anonymous. Einstein and me. Physics World, 19(12):42, December 2006. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/19/12/phwv19i12a43.pdf.

Anonymous:2006:GRR

[Ano06d] Anonymous. General relativity reveals its secrets. Physics World, 19(5):5, May 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/19/5/phwv19i5a7.
pdf.

Anonymous:2006:LES

[Ano06e] Anonymous. … but Einstein is still publishing. Physics World, 19(1):3, January 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://arXiv.org/abs/

physics/0510251;   http://physicsworldarchive.iop.org/
full/pwa-pdf/19/1/phwv19i1a3.pdf.

Anonymous:2006:NLE

[Ano06f] Anonymous. New light on Einstein’s affairs. Physics World, 19(8):3, August 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/19/8/phwv19i8a2.
pdf.

Anonymous:2006:PPE

[Ano06g] Anonymous. Political profit from Einstein’s science. Physics World, 19(7):3, July 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/19/7/phwv19i7a3.
pdf.

Anonymous:2006:PGC

[Ano06h] Anonymous. Pulsars’ gyrations confirm Einstein’s theory. Science, 313(5793):1556, September 15, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/cgi/content/abstract/1132305;
http://www.sciencemag.org/content/313/5793/1556.2.summary.

Anonymous:2006:REM

[Ano06i] Anonymous. The Russell–Einstein Manifesto. The SPS Observer, 23(2):??, Summer 2006. URL http://www.

spsobserver.org/2006/observer_ethics.pdf. The Mani-

festo is signed by Max Born, Percy W. Bridgman, Albert Einstein, Leopold Infeld, Frederic Joliot-Curie, Herman J. Muller, Linus Pauling, Cecil F. Powell, Joseph Rotblat, Bertrand Russell, and Hideki Yukawa.

Anonymous:2006:SIE

[Ano06j] Anonymous. Step inside Einstein’s mind. Physics World, 19(3): 41, March 2006. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/19/3/phwv19i3a44.pdf.

Anonymous:2006:WFP

[Ano06k] Anonymous. World famous people series. Unik Smart Educ-Tech, Kajang, Selangor, 2006. LCCN ???? 10 computer optical discs.

Anonymous:2007:ARN

[Ano07a] Anonymous. An astronomical road to a new theory of gravitation. In Renn et al. [RSS^{+}07a], pages 153–154. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:EZN

[Ano07b] Anonymous. Einstein’s Zurich Notebook: Transcription and facsimile. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 313–487. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v.

1; QC173.6 .G469 2007. URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-9_5.

Anonymous:2007:FTG

[Ano07c] Anonymous. A field theory of gravitation in the framework of Special Relativity. In Renn et al. [RSS^{+}07a], pages 411–412. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:ETM

[Ano07d]  Anonymous.  From an electromagnetic theory of mat-

ter to a new theory of gravitation. In Renn et al. [RSS^{+}07b], pages 621–622. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL http://

d-nb.info/985765933/34;   http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html;
 https://link.springer.com/book/10.1007/978-1-4020-
4000-9.

Anonymous:2007:HMN

[Ano07e] Anonymous. From heretical mechanics to a new theory of relativity. In Renn et al. [RSS^{+}07a], pages 567–568. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:PMN

[Ano07f] Anonymous. From peripheral mathematics to a new theory of gravitation. In Renn et al. [RSS^{+}07b], pages 1039–1040.

ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL http://d-nb.info/985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/2007440325-d.html;
http://www.loc.gov/catdir/enhancements/fy0824/2007440325-
t.html;   https://link.springer.com/book/10.1007/978-
1-4020-4000-9.

Anonymous:2007:GFB

[Ano07g] Anonymous. The gravitational force between mechanics and electrodynamics. In Renn et al. [RSS^{+}07a], pages 19–20. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:IGU

[Ano07h] Anonymous. Including gravitation in a unified theory of physics. In Renn et al. [RSS^{+}07b], pages 757–758. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL http://

d-nb.info/985765933/34;   http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html;
 https://link.springer.com/book/10.1007/978-1-4020-
4000-9.

Anonymous:2007:IVa

[Ano07i] Anonymous. Index: Volumes 1 and 2. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 927–938. ISBN 1-4020-3942-5. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

Anonymous:2007:IVb

[Ano07j] Anonymous. Index: Volumes 3 and 4. In Renn et al. [RSS^{+}07b], pages 1131–1152. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL http://

d-nb.info/985765933/34;   http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html;
 https://link.springer.com/book/10.1007/978-1-4020-
4000-9.

Anonymous:2007:MPM

[Ano07k] Anonymous. The Manhattan Project: More engineering disasters. The History Channel, New York, NY, USA, 2007. ISBN 0-7670-9767-X. LCCN T20 .M83 2007 V.3.

Anonymous:2007:NLG

[Ano07l] Anonymous. A new law of gravitation enforced by Special Relativity. In Renn et al. [RSS^{+}07a], pages 191–192. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:NLY

[Ano07m] Anonymous. Nobel lectures: 20 years of the Nobel Prize for Literature lectures. Icon, Cambridge, UK, 2007. ISBN 1-84046-834-3 (hardcover). xviii + 325 pp. LCCN PN6122 .N63 2007.

Anonymous:2007:NLL

[Ano07n] Anonymous. Nobel lectures: from the literature laureates, 1986 to 2006. New Press, New York, NY, USA, 2007. ISBN 1-59558-201-0 (hardcover), 1-59558-409-9 (paperback). ix + 293 pp. LCCN PN771 .N67 2007.

Anonymous:2007:NPLe

[Ano07o] Anonymous. Nobel Prize laureates in literature. Gale/Cengage Learning, Detroit, MI, USA, 2007. ISBN 1-4144-2863-4. LCCN ????

Anonymous:2007:NPLa

[Ano07p] Anonymous. Nobel Prize Laureates in Literature — Part 1. Dictionary of Literary Biography. Agnon–Eucken, volume 329. Gale/Cengage Learning, Detroit, MI, USA, 2007. ISBN 1-4144-2864-2. LCCN ????

Anonymous:2007:NPLb

[Ano07q] Anonymous. Nobel Prize Laureates in Literature — Part 2 Dictionary of Literary Biography. Faulkner–Kipling, volume 330. Gale/Cengage Learning, Detroit, MI, USA, 2007. ISBN 1-4144-2865-0. LCCN ????

Anonymous:2007:NPLc

[Ano07r] Anonymous. Nobel Prize Laureates in Literature — Part 3. Dictionary of Literary Biography. Lagerkvist–Pontoppidan, volume 331. Gale/Cengage Learning, Detroit, MI, USA, 2007. ISBN 1-4144-2866-9. LCCN ????

Anonymous:2007:NPLd

[Ano07s] Anonymous. Nobel Prize Laureates in Literature — Part 4 Dictionary of Literary Biography. Quasimodo–Yeats, volume 332. Gale/Cengage Learning, Detroit, MI, USA, 2007. ISBN 1-4144-2867-7. LCCN ????

Anonymous:2007:PGC

[Ano07t] Anonymous. The problem of gravitation as a challenge for the Minkowski formalism. In Renn et al. [RSS^{+}07a], pages 303–304. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:TTR

[Ano07u] Anonymous. Twin tests of relativity. Physics World, 20 (8):5, August 2007. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://arXiv.org/0706.

2031v1;   http://physicsworldarchive.iop.org/full/pwa-
pdf/20/8/phwv20i8a8.pdf.

Anonymous:2008:AE

[Ano08a] Anonymous. Albert Einstein. Jeden Dzień z… = A Day with. … Wydawnictwo Ibis, Poznań, Poland, 2008. ISBN 83-61482-41-5. 25 pp. LCCN ????

Anonymous:2008:ABH

[Ano08b] Anonymous. The atomic bombings of Hiroshima and Nagasaki. Web site of Lillian Goldman Law Library, 127 Wall Street, New Haven, CT 06511, USA., 2008. URL http://avalon.law.

yale.edu/subject_menus/mpmenu.asp.

Anonymous:2008:EGZ

[Ano08c] Anonymous. Einstein gets in the zone. Physics World, 21(5): 3, May 2008. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/21/05/phwv21i05a1.pdf.

Anonymous:2008:NP

[Ano08d] Anonymous. Nobel Prizes. Scientific American, 299(6): 42, December 2008. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v299/n6/full/
scientificamerican1208-42d.html;   http://www.nature.

com/scientificamerican/journal/v299/n6/pdf/scientificamerican1208-
42d.pdf.

Anonymous:2009:EB

[Ano09a] Anonymous. Einstein on the buses. Physics World, 22(2):3, February 2009. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/22/02/phwv22i02a2.pdf.

Anonymous:2009:NPT

[Ano09b] Anonymous. Nobel prizes for three IEEE fellows. IEEE Spectrum, 46(11):15, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:SRS

[Ano09c] Anonymous. Special relativity stands its ground. Physics World, 22(12):4, December 2009. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/22/12/phwv22i12a5.

pdf. See [AAA^{+}09] for the original research.

Anonymous:2009:SE

[Ano09d] Anonymous. Spooky Einstein. Physics World, 22(3):3, March 2009. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/22/03/phwv22i03a4.pdf.

Anonymous:2009:TEG

[Ano09e] Anonymous. The texture of Einstein’s genius. Science, 324 (5926):445, April 24, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/324/5926/445.3.full.

Anonymous:2009:WEM

[Ano09f] Anonymous. When Eddington met Einstein. Physics World, 22(1):37, January 2009. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/22/01/phwv22i01a45.

pdf. Brief review of BBC film Einstein and Eddington.

Anonymous:2009:WSF

[Ano09g] Anonymous. World science festival: Time since Einstein. New York Blog, June 15, 2009. CODEN ???? ISSN ????

Anonymous:2010:DED

[Ano10a] Anonymous. The day Einstein died. Physics World, 23(6): 15, June 2010. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/23/06/phwv23i06a25.pdf.

Anonymous:2010:ETP

[Ano10b] Anonymous. Einstein’s twin paradox in action. Physics World, 23(11):4, November 2010. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/23/11/phwv23i11a5.
pdf.

Anonymous:2010:EUC

[Ano10c] Anonymous. Einstein’s universe caught on film. Physics World, 23(8):9, August 2010. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/23/08/phwv23i08a18.
pdf.

Anonymous:2010:MBH

[Ano10d] Anonymous. Mimicking black holes: Dr Hawking’s bright idea: a long-predicted phenomenon has turned up in an unexpected place. The Economist, ??(??):??, September 30, 2010. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic). URL

http://www.economist.com/node/17144843.

Anonymous:2010:PDC

[Ano10e] Anonymous. Physics: Double-checking Einstein. Nature, 465 (7297):401, May 26, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v465/n7297/full/465401b.html.

Anonymous:2011:LLG

[Ano11a] Anonymous. At long last, Gravity Probe B satellite proves Einstein right. Science, 332(6030):649, May 6, 2011. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/332/

6030/649.full. From the introduction: “Fifty years after it was conceived, a $760 million NASA spacecraft has confirmed general relativity, Einstein’s theory of gravity, albeit less precisely than hoped.”.

Anonymous:2011:ANA

[Ano11b] Anonymous. ATV-4 named Albert Einstein. ESA Bulletin — European Space Agency, ??(147):53, August 2011. ISSN 0376-4265.

Anonymous:2011:NE

[Ano11c] Anonymous. The next Einstein? Physics World, 24(5):3, May 2011. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/24/05/phwv24i05a2.pdf.

Anonymous:2012:CER

[Ano12a] Anonymous. Corrigendum: Essay review: Einstein’s quest for unity. Annals of Science, 69(4):593, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Del12, Sau14, JL14b].

Anonymous:2012:EWP

[Ano12b] Anonymous. Europe from war to peace: The Nobel Peace Prize Laureate exhibition. Nobel Peace Center, Oslo, Norway, 2012. 71 pp. LCCN ????

Anonymous:2012:SASb

[Ano12c] Anonymous. Science agenda: Solve the Nobel Prize dilemma. Scientific American, 307(4):12, October 2012. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-12.html;   http://www.
nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
12.pdf.

Anonymous:2012:NPP

[Ano12d] Anonymous. The Nobel Prize in Physics 1967: Hans Bethe.

Web site., 2012. URL http://www.nobelprize.org/nobel_
prizes/physics/laureates/1967/.

Anonymous:2013:OAE

[Ano13a] Anonymous. 3 October 1933 — Albert Einstein presents his final speech given in Europe, at the Royal Albert Hall. Web

site, October 2013. URL https://www.royalalberthall.
com/about-the-hall/news/2013/october/3-october-1933-
albert-einstein-speaks-at-the-hall/.

Anonymous:2013:EDN

[Ano13b]  Anonymous.  Einstein@Home discovers 24 new pulsars

in archival data. Web news story., August 29, 2013.

URL http://www.scientificcomputing.com/news/2013/08/
einsteinhome-discovers-24-new-pulsars-archival-data.

See [Kni13] for the published research.

Anonymous:2013:GIE

[Ano13c] Anonymous. Get inside Einstein’s brain: Einstein’s brain in 3D at MD military museum. Web site, April 5, 2013. URL http:

//www.nbcphiladelphia.com/news/local/Einstein-Brain-
201580961.html.

Anonymous:2014:AMK

[Ano14a] Anonymous. Ancient microwaves key to testing Einstein theory. Scientific Computing, May 14, 2014. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2014/05/
ancient-microwaves-key-testing-einstein-theory.

Club:2014:MPR

[Ano14b] Anonymous. Max Planck: A reluctant revolutionary with a hunger of the soul. In Promoting the Planck Club: how defiant youth, irreverent researchers and liberated universities can foster prosperity indefinitely [Bra14b], pages 38–49. ISBN 1-118-54635-0, 1-306-47279-2, 1-118-54642-3, 1-118-54638-5 (ePub), 1-118-54636-9 (Adobe PDF). LCCN Q141. URL http://

proquest.tech.safaribooksonline.de/9781118546383.

Anonymous:2014:MBB

[Ano14c] Anonymous. Mind-blowing breakthroughs — season 1. Astrolab Motion / GFX, ???, ??, USA, 2014. 25 video files (ca. 325 min.)

pp. LCCN ???? URL http://d1n8i81n8cnkc4.cloudfront.
net/asl/mbb/001/dv1/asl\_mbb001dv1\_thumbnail.jpeg.

Anonyme:2015:IQS

[Ano15a] Anonyme. Information quantique : un siècle après Einstein, les premières applications de la théorie quantique : dossier. (French) [Quantum information: a century after Einstein, the first applications of quantum theory: File]. La Recherche, ??(501–502):26–77, July/August 2015. CODEN RC-CHBV. ISSN 0029-5671 (print), 1625-9955 (electronic). URL

http://www.sophiaboutique.fr/la-recherche-numerique-
mensuel-501.html.

Anonymous:2015:AEH

[Ano15b] Anonymous. Albert Einstein, in his own words: Einstein Papers Project publishes free digital edition. Scientific Computing, ??(??):??, March 23, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2015/03/
albert-einstein-his-own-words-einstein-papers-project-
publishes-free-digital-edition.

Anonymous:2015:EHS

[Ano15c] Anonymous. Einstein: His space and times. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v313/n3/
full/scientificamerican0915-94b.html;   http://www.
nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
94b.pdf.

Anonymous:2015:ESQ

[Ano15d] Anonymous. Einstein saves the quantum cat: Relativity theory applicable in other research areas. Scientific Computing, ??(??):??, June 6, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic). URL http://www.

scientificcomputing.com/news/2015/06/einstein-saves-
quantum-cat-relativity-theory-applicable-other-research-
areas.

Anonymous:2015:EDS

[Ano15e] Anonymous. Einstein’s dice and Schrödinger’s cat. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v313/n3/
full/scientificamerican0915-94c.html;   http://www.
nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
94c.pdf.

Anonymous:2015:HWI

[Ano15f] Anonymous. Hoe word ik Einstein of Da Vinci?: een inleiding tot wetenschappen vandaag voor de homo universalis van morgen. (Dutch) [How do I become Einstein or Da Vinci ?: an introduction to science today for the universal man of tomorrow].

LannooCampus, Leuven, Belgium, 2015. ISBN 94-014-2632-5 (gebonden). 581 pp. LCCN ????

Anonymous:2015:IE

[Ano15g] Anonymous. Introduction: Einstein. Scientific American, 313(3):32–34, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v313/n3/full/
scientificamerican0915-32.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
32.pdf.

Anonymous:2015:LNR

[Ano15h] Anonymous. A legacy in numbers: Relativity’s reach. Scientific American, 313(3):56–59, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-56.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-56.pdf.

Anonymous:2015:PPb

[Ano15i] Anonymous. The physicist and the philosopher. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v313/n3/
full/scientificamerican0915-94d.html;   http://www.
nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
94d.pdf.

Anonymous:2015:QEV

[Ano15j] Anonymous. Quantum experiment verifies Einstein’s ‘spooky action at a distance’. R&D Magazine, ??(??):??, March 24,

2015. URL http://www.rdmag.com/news/2015/03/quantum-
experiment-verifies-einsteins-spooky-action-distance.

Anonymous:2015:RRR

[Ano15k] Anonymous. Reviews and recommendations: The road to relativity. Scientific American, 313(3):94, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-94a.html;

 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-94a.pdf.

Anonymous:2015:SCW

[Ano15l] Anonymous. A study in contrasts: Who was Einstein, really? Scientific American, 313(3):80–81, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-80.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-80.pdf.

Anonymous:2016:AEF

[Ano16a] Anonymous. Albert Einstein: Les fondements de la théorie de la relativité générale [extrait]. (French) [Albert Einstein: The foundations of the theory of general relativity [excerpt]]. Bull. d’Informatique Approfondie et Appl., 103(??):73–74, ????

2016. CODEN ????  ISSN ????  URL http://biaa.

eu/-upload/articleno1032.pdf;   https://dblp.org/db/
journals/biaa/biaa103.html#X16.

Anonymous:2016:CNP

[Ano16b] Anonymous. Cambridge’s 92 Nobel Peace Prize winners — part 1: from Ernest Rutherford to Bertrand Russell. Cambridge News, ??(??):??, January 11, 2016. URL http://

www.cambridge-news.co.uk/Cambridge-s-92-Nobel-Peace-
Prize-winners-1-Ernest/story-28495103-detail/story.
html.

Anonymous:2016:MTA

[Ano16c] Anonymous. The march of time: Atomic power! Periscope Film LLC archive, April 17, 2016. URL https://www.youtube.com/

watch?v=iTJJ7tkD5L4.

Anonymous:2016:MRS

[Ano16d] Anonymous. Mysterious radio signals from space test Einstein’s General Relativity theory. Scientific Computing, ??(??):??, January 8, 2016. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic). URL http://

www.scientificcomputing.com/news/2016/01/mysterious-
radio-signals-space-test-einsteins-general-relativity-
theory.

Anonymous:2017:NPG

[Ano17a] Anonymous. Nobel prizes, giant telescope and buried treasure. Nature, 550(7674):12–13, October 5, 2017. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:2017:OT

[Ano17b] Anonymous. Oberflächen der Theorie. (German) [Surfaces of the theory], volume 1–2 of Bildwelten des Wissens: Kunsthistorisches Jahrbuch für Bildkritik. De Gruyter (A), Berlin, Germany; Boston, MA, USA, 2017. ISBN 3-11-055146-2, 3-11-055585-9. ISSN 1611-2529. LCCN ???? URL http://www.

degruyter.com/doc/cover/9783110551464.jpg.

Anonymous:2017:PER

[Ano17c] Anonymous. Proving Einstein right using the most sensitive earth rotation sensors ever made. R&D Magazine, ??

(??):??, May 10, 2017. URL https://www.rdmag.com/news/
2017/05/proving-einstein-right-using-most-sensitive-
earth-rotation-sensors-ever-made.

Anonymous:2018:AEF

[Ano18a] Anonymous. Albert Einstein, a famous passenger. Red Star

Line Web site, 2018. URL https://www.redstarline.be/en/
story/albert-einstein-famous-passenger.

Anonymous:2018:BHB

[Ano18b] Anonymous. Black hole breakthrough: New insight into mysterious jets. R&D Magazine, ??(??):??, January 11,

2018. URL https://www.rdmag.com/news/2018/01/black-
hole-breakthrough-new-insight-mysterious-jets.

Anonymous:2018:FST

[Ano18c] Anonymous. First successful test of Einstein’s general relativity near supermassive black hole (update). Web site., July

26, 2018. URL https://phys.org/news/2018-07-gravity-
relativity-galactic-centre-massive.html.

Anonymous:2018:SYF

[Ano18d] Anonymous. Seawater yields first grams of yellowcake. R&D

Magazine, June 14, 2018. URL https://www.rdmag.com/news/
2018/06/seawater-yields-first-grams-yellowcake.

Anonymous:2019:NPP

[Ano19] Anonymous. The Nobel Prize in physics fields. Nobel Media Web site., March 26, 2019. URL https://www.nobelprize.

org/prizes/uncategorized/the-nobel-prize-in-physics-
fields.

Anthes:2015:NSU

[Ant15] Gary Anthes. News: Scientists update views of light. Communications of the ACM, 58(10):15–17, October 2015. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.acm.org/magazines/2015/10/
192377/fulltext.

Ashton:1978:MPN

[AO78] S. V. Ashton and C. Oppenheim. A method of predicting Nobel Prizewinners in Chemistry. Social Studies of Science, 8 (3):341–348, August 1, 1978. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277800800306.

Abbott:2004:ALD

[Ao04] B. Abbott and 373 others. Analysis of LIGO data for gravitational waves from binary neutron stars. Physical Review D (Particles and Fields), 69(12):122001, June 2004. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.69.122001.

Arons:1965:EPP

[AP65] A. B. Arons and M. B. Peppard. Einstein’s proposal of the photon concept — a translation of the Annalen der Physik paper of 1905. American Journal of Physics, 33(5):367–374, May 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/33/5/10.1119/1.1971542. English trans-

lation, with commentary in abstract, of Einstein’s 1905 paper [Ein05d] for which he won the 1921 Nobel Prize in Physics.

Agazzi:2000:RUO

[AP00] Evandro Agazzi and Massimo Pauri. The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism, volume 215 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346.

vii + 377 pp. LCCN ???? URL https://link.springer.com/
book/10.1007/978-94-015-9391-5.

Arnold:2001:BHB

[AP01] Lorna Arnold and Katherine Pyne. Britain and the Hbomb. St. Martin’s Press, New York, NY, USA, 2001. ISBN 0-312-23518-6 (hardcover). xiv + 273 + 12 pp. LCCN UG1282.A8 A76 2001. URL http://www.loc.gov/

catdir/enhancements/fy0610/00036898-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0610/00036898-d.
html;   http://www.loc.gov/catdir/enhancements/fy0610/
00036898-t.html.

Apfel:1981:ARE

[Apf81] Necia H. Apfel. It’s all relative: Einstein’s Theory of Relativity. Lothrop, Lee and Shepard Books, New York, NY, USA, 1981. ISBN 0-688-41981-X, 0-688-51981-4 (hardcover). 141 pp. LCCN QC173.575 .A63. Diagrams by Yukio Kondō.

Appell:2011:RNR

[App11] David Appell. Relativity’s new revolution. Physics World, 24(10):36–41, October 2011. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/24/10/phwv24i10a36.
pdf.

Atmanspacher:1995:PJD

[APWB95] Harald Atmanspacher, Hans Primas, and E. (Eva) Wertenschlag-Birkhäuser, editors. Der Pauli–Jung-Dialog und seine Bedeutung für die moderne Wissenschaft. (German) [The Pauli–Jung Dialog and its Meaning for Modern Science]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58518-4. viii + 365 pp. LCCN Q175.3 .P38

1995. URL http://www.zentralblattmath.org/zmath/en/
search/?an=0820.01012.

ArenasGomez:2005:AE

[Are05] Albino Arenas Gómez. Albert Einstein, volume 42 of Grandes biografías. EDIMAT, Arganda del Rey (Madrid), Spain, 2005. ISBN 84-9764-583-9. 190 pp. LCCN QC16.E5 G66 2005.

Argentieri:2009:CSE

[Arg09] Niccolò Argentieri. Ci sono elettroni nel mondo-della-vita?: esperienza, matematica, realtà: una lettura fenomenologica dell’epistemologia di Werner Heisenberg. (Italian) [Electrons in the real world: experience, mathematics, truth: a phenomenological study of the epistemology of Werner Heisenberg], volume 7 of Tascabili Bonanno. Filosofia. Bonanno, Acireale, Italia, 2009. ISBN 88-7796-549-5. 126 pp. LCCN QC16.H35

A76 2009. 10.00 EUR. URL http://www.loc.gov/catdir/
toc/casalini08/09777733.pdf.

Ariza:1998:ED

[Ari98] Luis Miguel Ariza. Einstein’s drag. Scientific American, 279(1):24–25, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v279/n1/
pdf/scientificamerican0798-24.pdf.

Arianrhod:2003:EHI

[Ari03] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. University of Queensland Press, St Lucia, Queensland, Australia, 2003. ISBN 0-7022-3408-7. viii + 323 pp. LCCN QC19.6 .A75 2003. US$24.00.

Arianrhod:2005:EHI

[Ari05] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-518370-3 (hardcover), 0-19-530890-5 (paperback). xiii + 323 pp. LCCN QC19.6 .A75 2005. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip053/2004026055.html.

Arianrhod:2014:YES

[Ari14] Robyn Arianrhod. Young Einstein and the Story of E = mc^{2}. Hampress, Amazon Kindle Singles, ????, 2014. ISBN ???? ???? pp. LCCN ????

Atmanspacher:1999:BSM

[ARK^{+}99] Harald Atmanspacher, Helmut Rechenberg, Horst Kant, Rolf Hempelmann, Reinhold Blümel, Gerhard Börner, Houjun Mo, Philippe Blanchard, Marius Grundmann, Dietrich Stauffer, and Christian Uebing. Bücher und Software: Meÿenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Band IV/Teil II: 1953–1954/Mladjenovic: The Defining Years in Nuclear Physics. 1932–1960s/Frenkel: Professor Friedrich Houtermans Arbeiten, Leben, Schicksal/Kärger and Heitjans: Diffusion in Condensed Matter/Friedrich: Theoretical Atomic Physics/Peacock: Cosmological Physics/Grosse u Martin: Particle Physics and the Schrödinger Equation/ Gaponenko: Optical Properties of Semiconductor Nanocrystals/Gershenfeld: The Nature of Mathematical Modeling/ Origin 6.0: Vertrieb. Physikalische Blätter, 55(9):79–83, September 1999. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19990550919/abstract.

Armstrong:1938:BRE

[Arm38] E. F. Armstrong. Book review: The Evolution of Physics by Albert Einstein, Leopold Infeld. Journal of the Royal Society of Arts, 86(4477):1045–1046, September 9, 1938. CODEN JR-

SAA4. ISSN 0035-9114. URL http://www.jstor.org/stable/
41361408.

Arnold:2000:RBK

[Arn00] Lorna Arnold. Recalling Britain’s key nuclear role. Physics World, 13(2):17–18, February 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http:/

/stacks.iop.org/2058-7058/13/i=2/a=18.

Aronowitz:1996:PSW

[Aro96] Stanley Aronowitz. The politics of the science wars. Social Text, 62(46/47):177–197, Spring/Summer 1996. CODEN ???? ISSN 0164-2472 (print), 1527-1951 (electronic). URL http://www.

jstor.org/stable/466853.

Arp:1998:SRR

[Arp98] Halton C. Arp. Seeing red: redshifts, cosmology and academic science. Apeiron, Montréal, Canada, 1998. ISBN 0-9683689-0-5. iv + 306 + 8 pp. LCCN QB857 .A764 1998.

ArroyoCamejo:2006:SQG

[Arr06] Silvia Arroyo Camejo. Skurrile Quantenwelt. (German) [Crazy Quantum World]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-29720-0. xiii + 246 pp. LCCN QC174.12 .A77 2006. EUR

29.95, SFR 51.00. URL http://www.springer.com/physics/
quantum+physics/book/978-3-540-29720-8.

Arrizabalaga:2007:AES

[Arr07] Jon Arrizabalaga. Albert Einstein: Science and conscience. Dynamis, 27(??):418–421, ???? 2007. ISSN 0211-9536.

Arabatzis:2015:RHS

[ARS15] Theodore Arabatzis, Jürgen Renn, and Ana Simões, editors. Relocating the History of Science: Essays in Honor of Kostas Gavroglu, volume 312 of Boston Studies in the Philosophy and History of Science. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). vii + 383 pp.

LCCN Q125 .R45 2015. URL http://www.loc.gov/catdir/
enhancements/fy1516/2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/fy1516/2015936523-t.html.

Aichelburg:1979:AEH

[AS79] Peter C. Aichelburg and Roman Ulrich Sexl, editors. Albert Einstein: his influence on physics, philosophy and politics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1979. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). xv + 220 pp. LCCN

QC16.E5 A66. DM48.00. URL http://link.springer.com/
book/10.1007/978-3-322-91080-6/.

Ashtekar:1991:CPQ

[AS91] Abhay Ashtekar and John J. Stachel, editors. Conceptual problems of quantum gravity: based on the proceedings of the 1988 Osgood Hill Conference, North Andover, Massachusetts, 15–19 May 1988, volume 2 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1991. ISBN 0-8176-3443-6, 3-7643-3443-6. LCCN QC178 .C63 1991.

Asada:1995:MCN

[Asa95] Sadao Asada. The mushroom cloud and national psyches: Japanese and American perceptions of the A-bomb decision, 1945–1995. Journal of American–East Asian Relations, 4(2): 95–116, ???? 1995.

Ash:1995:GPGk

[Ash95] Mitchell G. Ash. Gestalt psychology in German culture, 1890– 1967: holism and the quest for objectivity. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47540-6 (hardcover). xii + 513 pp. LCCN BF203 .A84 1995. URL http://catdir.loc.

gov/catdir/description/cam026/94036273.html;   http:
//catdir.loc.gov/catdir/toc/cam026/94036273.html;
http://digitool.hbz-nrw.de:1801/webclient/DeliveryManager?
pid=3213996.

Ashby:2002:RGP

[Ash02] Neil Ashby. Relativity and the Global Positioning System. Physics Today, 55(5):41–47, May 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Ashkenazi:2011:RIP

[Ash11] Ofer Ashkenazi. Reframing the interwar peace movement: The curious case of Albert Einstein. Journal of Contemporary History, 46(4):741–766, October 2011. ISSN 0022-0094 (print), 1461-7250 (electronic).

Ashkenazi:2012:ZVA

[Ash12] Ofer Ashkenazi. Zionism and violence in Albert Einstein’s political outlook. Journal of Jewish Studies, 63(2):331–355, Fall 2012. ISSN 0022-2097 (print), 2056-6689 (electronic).

Asimov:1983:SLS

[Asi83a] I. Asimov. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. American Jewish History, 72(4):531–534, ???? 1983. ISSN 0164-0178.

Asimov:1983:MU

[Asi83b] Isaac Asimov. The measure of the universe. Harper & Row, New York, NY, USA, 1983. ISBN 0-06-015129-3. 339 pp. LCCN QC39 .A76 1983. US$15.95.

Aspden:1969:PE

[Asp69] Harold Aspden. Physics without Einstein. Sabberton Publications, Southampton, UK, 1969. ISBN 0-85056-001-2. xiii + 224 pp. LCCN QC75 .A838p 1969. See critical review [For71a].

Aspden:1980:PUH

[Asp80] Harold Aspden. Physics unified. Sabberton Publications, Southampton, UK, 1980. ISBN 0-85056-009-8 (hardcover), 0-85056-010-1 (paperback). xiii + 206 pp. LCCN QC173.7 .A85.

Alperovitz:1995:DUA

[ATW^{+}95] Gar Alperovitz, Sanho Tree, Edward Rouse Winstead, Kathryn C. Morris, David J. Williams, Leo C. Maley III, Thad Williamson, and Miranda Grieder. The decision to use the atomic bomb and the architecture of an American myth. Alfred A. Knopf, New York, NY, USA, 1995. ISBN 0-679-44331-2. xiv + 847 pp. LCCN D769.2 .A5 1995. US$32.50; CDN$45.50.

Audeguy:2006:AE

[Aud06] S. Audeguy. Albert Einstein. Nouvelle Revue Française, ?? (577):131, April 2006. ISSN 0029-4802.

Auer:1995:DNH

[Aue95] Peter Auer. Von Dahlem nach Hiroshima: die Geschichte der Atombombe. (German) [From Dahlem to Hiroshima: the history of the atom bomb]. Aufbau-Verlag, Berlin, Germany, 1995. ISBN 3-351-02429-0. 336 + 14 pp. LCCN QC773 .A84 1995.

Aerts:1999:WVP

[AVvdV99] Diederik Aerts, Hubert Van Belle, and Jan van der Veken, editors. World views and the problem of synthesis: the yellow book of “Einstein meets Magritte”, volume 4 of Einstein meets Magritte. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5760-4 (hardcover). LCCN B20 .W68 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//lcweb.loc.gov/catdir/toc/99029566.html. See other vol-

umes in this series [Aer99a, AMM99, Aer99b, ABW99, BAM99, CSvB99, HBR99].

Axel:1950:NB

[Axe50] Peter Axel. News in brief. Bulletin of the Atomic Scientists, 6 (12):380–382, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNa

[Axe52] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(1):31–32, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aynshtayn:1934:DOO

[Ayn34] Albert Aynshtayn. Děmut ôlami: ha-ôlam kěfi e-hû mis.t.ayyer lî. (Hebrew) [My World View]. Št.ībel, Tel Aviv, Israel, 1934. 224 pp. LCCN ???? Hebrew translation by Š. Et.înger of [Ein34h].

Aharonov:2005:TQE

[AZ05] Yakir Aharonov and M. Suhail Zubairy. Time and the quantum: Erasing the past and impacting the future. Science, 307(5711): 875–879, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

B:1950:RVL

[B.50] J. L. B. Review: Voices of Liberalism: 2, by Albert Einstein; George D. Stoddard. The Journal of Philosophy, 47(3): 79–80, February 1950. ISSN 0022-362x (print), 1939-8549 (elec-

tronic). URL 2-5;   http://links.jstor.org/sici?sici=
0022-362X(19500202)47:3<79:VOL2>2.0.CO.

Bopp:1961:WHP

[B^{+}61] F. (Fritz) Bopp et al., editors. Werner Heisenberg und die Physik unserer Zeit. (German) [Werner Heisenberg and the physics of his time]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961. 310 pp. LCCN QC3 .B59.

Beck:1979:GGA

[B^{+}79] F. (Friedrich) Beck et al., editors. Gedächtnisausstellung zum 100 Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner: in der Staatsbibliothek Preussischer Kulturbesitz, Berlin, vom 1. März–12. April 1979: im Theodor-Zink-Museum, Kaiserslautern, vom 21. Juni–31. Juli 1979. (German) [Commemorative exhibition for the 100th birthday of Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner at the State Library of Prussian Cultural Heritage, Berlin, from 1 March to 12th April 1979: the Theodor Zink Museum, Kaiserslautern, 21 June–31. July 1979]. Physik Kongress-Ausstellungs- und Verwaltungs GmbH, Bad Honnef, West Germany, 1979. 151 pp. LCCN QC16.E5 G43 1979.

Birshtein:1989:PAE

[B^{+}89] Barbara K. Birshtein et al. Problems at Albert Einstein. Nature, 339(6219):10, May 4, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v339/n6219/pdf/339010a0.pdf.

Bunge:1999:BCC

[B^{+}99] Mario Bunge et al. Borges científico: cuatro estudios. (Spanish) [Borges scientist: four studies]. Colección Fin del milenio. Ediciones Biblioteca Nacional, Buenos Aires, Argentina, 1999. ISBN 987-503-197-6. 79 pp. LCCN PQ7797.B635 Z63446 1999.

Beckers:2001:ANP

[B^{+}01] Tatjana Beckers et al. Aachener Nobelpreisträger?: Physik im Spannungsfeld von Wissenschaft und Gesellschaft um 1900. (German) [Nobel Prize winners from Aachen?: caught in the conflict between science and society around 1900]. Deutsches Museum, Bonn, Germany, 2001. ISBN 3-89796-047-8 (hardcover). 175 pp. LCCN QC9.G3 A23 2001.

Barnes:2006:EUA

[B^{+}06] T. H. (Thomas Heinrich) Barnes et al. The elegant universe of Albert Einstein: the collected lectures of the Royal Society of New Zealand E = mc^{2} series, broadcast on National Radio. Awa science. Awa Press, Wellington, New Zealand, 2006. ISBN 0-9582629-2-6 (paperback). 170 pp. LCCN QC16.E5

E67 2006. URL http://www.loc.gov/catdir/enhancements/

fy0705/2006389375-d.html. With an introduction by Rebecca Priestley.

Bachelier:1900:TSF

[Bac00] Louis Bachelier. Théorie de la spéculation. (French) [theory of speculation]. Annales Scientifiques de l’École Normale Supérieure, 3(17):21–86, 1900. CODEN ASENAH. ISSN 0012-9593 (print), 1873-2151 (electronic). URL http:

//archive.numdam.org/article/ASENS_1900_3_17__21_0.
pdf;   http://en.wikipedia.org/wiki/Louis_Bachelier.

Backlund:1919:ZTK

[Bäc19] A. V. Bäcklund. Zusammenstellung einer Theorie der klassischen Dynamik und der neuen Gravitationstheorie von Einstein. (German) [Compilation of a theory of classical dynamics and the new theory of gravitation of Einstein], volume 14, 11 of Arkiv för matematik, astronomi och fysik. Almqvist and Wiksell, Stockholm, Sweden, 1919. 64 pp. LCCN ????

Backlund:1921:ZTK

[Bäc21] A. V. Bäcklund. Zusammenstellung einer Theorie der klassischen Dynamik und der neuen Gravitationstheorie von Einstein.

(German) [Compilation of a theory of classical dynamics and the new theory of gravitation of Einstein], volume 15, 17 of Arkiv för matematik, astronomi och fysik. Almqvist and Wiksell, Stockholm, Sweden, 1921. 12 pp. LCCN ????

Bachelard:1949:PDC

[Bac49] Gaston Bachelard. The philosophic dialectic of the concepts of relativity. In Schilpp [Sch49c], chapter 22, pages 563–580. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Bacher:1950:HBI

[Bac50] Robert F. Bacher. The hydrogen bomb: III. Scientific American, 182(5):11–15, May 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v182/n5/
pdf/scientificamerican0550-11.pdf. See also Parts I

[Rid50] and II [Bet50b].

Bach:1989:MCA

[Bac89a] A. Bach. A misinterpretation with consequences — Albert Einstein and the duality of waves and particles. Archive for History of Exact Sciences, 40(2):173–206, ???? 1989. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Bach:1989:FFA

[Bac89b] Alexander Bach. Eine Fehlinterpretation mit Folgen: Albert Einstein und der Welle–Teilchen Dualismus. (German) [A misinterpretation with consequences: Albert Einstein and the wave–particle dualism]. Archive for History of Exact Sciences, 40(2):173–206, June 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=2&spage=173.

Bachert:2005:AEP

[Bac05] Peter Bachert. Albert Einstein — Prophet der Symmetrie aller physikalischen Krafte. (German) [Albert Einstein — prophet of symmetry of all physical forces]. Zeitschrift für medizinische Physik, 15(2):69, ???? 2005. ISSN 0939-3889.

Badash:1972:BRH

[Bad72] Lawrence Badash. Book review: H. G. Graetzer and D. L. Anderson. Physics Today, 25(3):65–??, March 1972. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v25/i3/p65_s1.

Badash:1983:AEH

[Bad83a] L. Badash. Albert Einstein — historical and cultural perspectives — G. Holton, Y. Elkana. Annals of Science, 40(5):505–506, ???? 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Badash:1983:BRBb

[Bad83b] Lawrence Badash. Book review: The Cult of the Atom: The Secret Papers of the Atomic Energy Commission by Daniel Ford. Isis, 74(4):622–623, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232269.

Badash:1983:OHSa

[Bad83c] Lawrence Badash. Otto Hahn, science, and social responsibility. In Shea [She83a], pages 167–180. ISBN 94-009-7133-8 (e-book), 94-009-7135-4. ISSN 1566-659X. LCCN D1-DX301.

Badash:1983:OHSb

[Bad83d] Lawrence Badash. Otto Hahn, science, and social responsibility. In Shea [She83b], pages 167–180. ISBN 90-277-1584-X. LCCN

QC773 .O87 1983. URL http://link.springer.com/book/
10.1007/978-94-009-7133-2.

Badash:1985:KRK

[Bad85] Lawrence Badash. Kapitza, Rutherford, and the Kremlin. Yale University Press, New Haven, CT, USA, 1985. ISBN 0-300-01465-1. xi + 129 pp. LCCN QC16.K25 B3 1985.

Badash:1995:SDN

[Bad95] Lawrence Badash. Scientists and the development of nuclear weapons: from fission to the Limited Test Ban Treaty, 1939– 1963. The control of nature. Humanities Press, Atlantic Highlands, NJ, USA, 1995. ISBN 0-391-03873-7, 0-391-03874-5 (paperback). ix + 129 pp. LCCN U264 .B34 1995.

Badash:1996:BRB

[Bad96] Lawrence Badash. Book review: Hitler’s Uranium Club: The Secret Recordings at Farm Hall by Jeremy Bernstein; David Cassidy. Isis, 87(3):569, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236043.

Badge:2004:NIP

[Bad04] Peter Badge. Nobelpreisträger im Portrait. (German) [Nobel laureates in portraits]. ars vivendi, Nürnberg, Germany, 2004. ISBN 3-89716-519-8. 318 pp. LCCN Q141 .B33 2004.

Badino:2009:OCB

[Bad09] Massimiliano Badino. The odd couple: Boltzmann, Planck and the application of statistics to physics (1900–1913). Annalen der Physik 8 (Berlin, Germany), 18(2–3):81–101, ???? 2009. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See also [Gea02].

Badge:2011:NC

[Bad11] Peter Badge. Nobel chemists: [official catalog of the Exhibition “Nobel de Chimie — Photographies de Peter Badge” at the Musée des Arts et Métiers le CNAM, Paris, 25 January–27 March 2011]. Walter de Gruyter, Berlin, Germany, 2011. ISBN 3-11-025482-4. ???? pp. LCCN AS911.N9 .B33 2011.

Schmidt:2012:PBN

[Bad12] Peter Badge. Nobel physicists: [official catalogue of the Exhibition “Nobel Physicists — Photographs by Peter Badge”, at the 62nd Lindau Nobel Laureate Meeting at Lake Constance 2012, with venues at Inselhalle, and the Exhibition “Nobelpreisträger im Portrait” at the Stadtmuseum Cavazzen, Lindau, 1 July– 30 September 2012]. Stiftung Lindauer Nobelpreisträgertreffen am Bodensee, Lindau am Bodensee, Germany, 2012. ISBN 3-939201-02-2. 105 (est.) pp. LCCN ???? EUR 29.90 (DE). Foreword by Bettina Bernadotte. Preface by Angela Merkel. Introduction by Brian Schmidt. Essay by Nikolaus Turner.

Badino:2015:BRA

[Bad15] Massimiliano Badino. Book reviews: Albert Einstein; Diana Kormos Buchwald; Ze’ev Rosenkranz; Tilman Sauer; József Illy; Virginia Iris Holmes, eds. The Collected Papers of Albert Einstein. Volume 12: The Berlin Years: Correspondence, January December 1921. Albert Einstein; Diana Kormos Buchwald; József Illy; Ze’ev Rosenkranz; Tilman Sauer, eds. The Collected Papers of Albert Einstein. Volume 13: The Berlin Years: Writings and Correspondence, January 1922 March 1923. Isis, 106 (1):209–211, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681876.

Baggott:1992:MQT

[Bag92] J. E. Baggott. The meaning of quantum theory: a guide for students of chemistry and physics. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-855575-X, 0-19-855576-8 (hardcover). xi + 230 pp. LCCN QC174.12 .B34

1992. URL http://www.loc.gov/catdir/enhancements/
fy0640/91034937-d.html;   http://www.loc.gov/catdir/
enhancements/fy0640/91034937-t.html.

Baggott:2004:BMM

[Bag04] J. E. Baggott. Beyond measure: modern physics, philosophy, and the meaning of quantum theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852927-9 (hardcover), 0-19-852536-2 (paperback). xvi + 379 pp. LCCN

QC174.12 .B33 2004. URL http://catdir.loc.gov/catdir/
enhancements/fy0615/2003066207-d.html.

Baggott:2010:FWP

[Bag10] Jim Baggott. The First War of Physics: The Secret History of the Atomic Bomb, 1939–1949. Pegasus Books, New York, NY, USA, 2010. ISBN 1-60598-084-6. xiv + 576 pp.

LCCN QC773 .B24 2010. URL http://pegasusbooks.us/pdf/
Spg10_Catalog.pdf.

Baggott:2011:QSH

[Bag11] J. E. Baggott. The quantum story: a history in 40 moments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-956684-4 (hardcover). xix + 469 + 16 pp. LCCN QC173.98 .B34 2011.

Bainbridge:1933:EMEb

[Bai33a] K. T. Bainbridge. The equivalence of mass and energy. Journal of The Franklin Institute, 216(2):255–256, August 1933. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Bainbridge:1933:EMEa

[Bai33b] Kenneth T. Bainbridge. The equivalence of mass and energy. Physical Review (2), 44(2):123, July 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bainbridge:1933:MAS

[Bai33c] Kenneth T. Bainbridge. The masses of atoms and the structure of atomic nuclei. Journal of The Franklin Institute, 215(5):509– 534, May 1933. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Bainbridge:1975:AOTa

[Bai75] Kenneth T. Bainbridge. “All in our time”: Prelude to Trinity. Bulletin of the Atomic Scientists, 31(4):42–46, April 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bailey:1989:CPA

[Bai89] Herbert Smith Bailey, Jr. On the Collected Papers of Albert Einstein: The development of the project. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 133(3):347–359, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://www.jstor.org/stable/987104.

Baierlein:1992:NET

[Bai92] Ralph Baierlein. Newton to Einstein: the trail of light: an excursion to the wave-particle duality and the Special Theory of Relativity. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41171-8 (hardcover), 0-521-42323-6 (paperback). xvi + 329 pp. LCCN QC476.W38 B35 1992.

Baierlein:2001:NET

[Bai01] Ralph Baierlein. Newton to Einstein: the trail of light: an excursion to the wave-particle duality and the Special Theory of Relativity. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-42323-6 (paperback). xvi + 329 pp. LCCN QC476.W38

B35 2001. URL http://www.cambridge.org/uk/catalogue/
catalogue.asp?isbn=0521423236.

Bailey:2006:AN

[Bai06] Ellen Bailey. Alfred Nobel. Great Neck Publishing, Toledo, OH, USA, 2006. ISBN 1-4298-0704-0. LCCN TP268.5.N7; TP268.5.N7 INTERNET.

Bairstow:2007:MVA

[Bai07a] Jeffrey Bairstow. In my view — Albert Einstein: wicked brilliant or deadly boring? Laser Focus World, 43(8):144, August 2007. CODEN LFWOE8. ISSN 1043-8092.

Bais:2007:VSR

[Bai07b] Sander Bais. Very Special Relativity: an Illustrated Guide. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02611-X (hardcover), 0-674-01967-9 (paperback). 120 pp. LCCN

QC173.65 .B35 2007. URL http://www.loc.gov/catdir/toc/
ecip0712/2007009144.html.

Baker:1968:ABG

[Bak68] Paul R. Baker, editor. The atomic bomb: the great decision. American problem studies. Holt, Rinehart, and Winston, New York, NY, USA, 1968. 122 pp. LCCN D842 .B34.

Bakercomp:1976:ABG

[Bak76a] Paul R. Baker, editor. The atomic bomb: the great decision. American problem studies. Dryden Press, Hinsdale, IL, USA, second edition, 1976. ISBN 0-03-089873-0. viii + 193 pp. LCCN D842 .B34 1976.

Baker:1976:ABG

[Bak76b] Paul R. Baker, editor. The atomic bomb: the great decision. American problem studies. Dryden Press, Hinsdale, IL, USA, second revised edition, 1976. ISBN 0-03-089873-0. viii + 193 pp. LCCN D842 .B34 1976.

Baker:1997:AEL

[Bak97] Candice Baker. Albert Einstein, level 2. SmartReader. Recorded Books, Prince Frederick, MD, USA, 1997. ISBN 0-7887-3307-9 (sound), 0-7887-9439-6 (paperback). LCCN QC16.E5 B35 1997. 2 sound cassettes.

Balazs:1972:BRS

[Bal72] Nandor Balazs. Book review: A singular scientist: Einstein: The Life and Times, by R. W. Clark. Physics Today, 25(11): 51–53, November 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3071093.

Balazs:1975:BRE

[Bal75a] Nandor L. Balazs. Book review: The Einstein Decade (1905– 1915), by Cornelius Lanczos. Physics Today, 28(8):70, August 1975. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Balducci:1975:SMW

[Bal75b] Carolyn Balducci. A self-made woman: biography of Nobel-Prize-winner Grazia Deledda. Houghton Mifflin, Boston, MA, USA, 1975. ISBN 0-395-21914-0. 200 pp. LCCN PQ4811.E6 Z58.

Baltimore:1977:NLM

[Bal77] David Baltimore. Nobel lectures in molecular biology, 1933– 1975. Elsevier North-Holland, New York, NY, USA, 1977. ISBN 0-444-00236-7. ix + 534 pp. LCCN QH506 .N58.

Balch:1979:PNP

[Bal79] Emily Greene Balch. Perspective: Nobel Peace Laureates: 1946. Bulletin of the Atomic Scientists, 35(4):1, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 7 April 1948, Oslo, Norway.

Balibar:1985:BEE

[Bal85] Françoise Balibar. Bohr entre Einstein et Dirac. (French) [Bohr between Einstein and Dirac]. Revue d’Histoire des Sciences, 38(3–4):293–307, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/
article/rhs_0151-4105_1985_num_38_3_4008.

Ball:1986:JDA

[Bal86] Howard Ball. Justice downwind: America’s atomic testing program in the 1950’s. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503672-7. xviii + 280 pp. LCCN U264 .B35 1986. US$21.95. URL http://www.loc.

gov/catdir/enhancements/fy0636/85008777-d.html.

Balibar:1991:AOB

[Bal91] Françoise Balibar. Analyses d’ouvrages: The Collected Papers of Albert Einstein, vol. I: Early years, 1879–1902 par John Stachel. Revue d’Histoire des Sciences, 44(2):255–257, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633154.

Balibar:1992:CEE

[Bal92] F. Balibar. La correspondance entre Einstein et Schrödinger. In Bitbol and Darrigol [BD92], pages 307–312. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Balibar:1993:EJP

[Bal93] Françoise Balibar. Einstein: la joie de la pensée. (French) [Einstein: the joy of thought], volume 193 of Découvertes Gallimard. Sciences. Gallimard, Paris, France, 1993. ISBN 2-07-053220-8 (paperback). 144 pp. LCCN QC16.E5 .B35 1993.

Balibar:1994:EGD

[Bal94] Françoise Balibar. Einstein: la gioia del pensiero. (Italian) [Einstein: the joy of thought], volume 49 of Universale Electa/Gallimard. Electa/Gallimard, Torino, Italia, 1994. ISBN 88-445-0057-4. 176 pp. LCCN ???? Italian translation of [Bal93].

Balibar:1995:EJP

[Bal95a] Françoise Balibar. Einstein: la joie de la pensée. (French) [Einstein: the joy of thought], volume 193. Sciences of Découvertes Gallimard. Gallimard, Paris, France, 1995. ISBN 2-07-053220-8. 144 pp. LCCN ????

Balibar:1995:ETO

[Bal95b] Françoise Balibar. Einstein: tänkaren och fysikern. (Swedish). [Einstein: the thinker and physicist]. En värld av vetande. Bergh, Stockholm, Sweden, 1995. ISBN 91-502-1204-4s. 144 pp. LCCN ????

Balibar:1996:AOB

[Bal96] Françoise Balibar. Analyses d’ouvrages: Collected Papers of Albert Einstein (The), vol. 3: The Swiss years: writings, 1909– 1911. vol. 5. The Swiss years: correspondence, 1902–1914 par Martin J. Klein; A. J. Kox; Jürgen Renn; Robert Schulmann. Revue d’Histoire des Sciences, 49(4):583–584, October 1996. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23633772.

Balibar:1999:EGP

[Bal99] Françoise Balibar. Einstein: el gozo de pensar. (Spanish). [Einstein: the joy of thought]. Ediciones B, Barcelona, España, 1999. ISBN 84-406-9192-0. 144 pp. LCCN ????

Balibar:2001:EDU

[Bal01a] Françoise Balibar. Einstein: decoding the Universe. Discoveries. Harry N. Abrams, New York, NY, USA, 2001. ISBN 0-8109-2980-5 (paperback). 143 pp. LCCN QC16.E5 B3513

2001. URL http://catalog.hathitrust.org/api/volumes/
oclc/45263897.html.

Ball:2001:EES

[Bal01b] Philip Ball. Exorcising Einstein’s spooks. Nature News, November 29, 2001. CODEN ???? ISSN ???? URL http://www.

nature.com/news/2001/011129/full/news011129-15.html.

Balazs:2005:KTE

[Bal05] F. Balazs. Keeping time with Einstein (Albert Einstein). Smithsonian, 36(5):12, August 2005. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Balibar:2008:EDK

[Bal08] Françoise Balibar. Einstein: Düşünmenin keyfi. (Turkish) [Einstein: the joy of thought], volume 2020; 15. Yapı Kredi Yayınları, Istanbul, Turkey, 2008. ISBN 975-08-0775-8. 144 pp. LCCN ???? Turkish translation by Aykut Derman of [Bal93].

Balibar:2011:EJP

[Bal11a] Françoise Balibar. Einstein: la joie de la pensée. (French) [Einstein: the joy of thought], volume 193 of Découvertes Gallimard. Sciences. Gallimard, Paris, France, 2011. ISBN 2-07-053220-8, 2-07-034784-2. 144 pp. LCCN QC16.E5 .B35 2011.

Ball:2011:DED

[Bal11b] Philip Ball. Did Einstein discover E = mc^{2}. Web article., Au-

gust 23, 2011. URL http://physicsworld.com/cws/article/

news/2011/aug/23/did-einstein-discover-e-equals-mc-
squared.

Ball:2011:DFH

[Bal11c] Philip Ball. Did Fritz Hasenöhrl derive E = mc^{2} before Einstein? Physics World, 24(10):13, October 2011. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
24/10/phwv24i10a24.pdf.

Broekaert:1999:EMM

[BAM99] Jan Broekaert, Diederik Aerts, and Ernest Mathijs, editors. Einstein meets Magritte: an interdisciplinary reflection: the white book of “Einstein meets Magritte”, volume 1 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 90-5487-226-8 (vol. 1, VUB), 0-7923-5757-4 (vol. 1, Kluwer). LCCN Q174 .E38 1999. See other volumes in this series [Aer99a, AMM99, Aer99b, ABW99, AVvdV99, CSvB99, HBR99].

Bandyopadhyay:1951:SFE

[Ban51] Gaganbihari Bandyopadhyay. A strange feature of Einstein’s most recent generalized field theory. Nature, 167(4251):648– 649, April 21, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v167/n4251/pdf/167648a0.pdf.

Bankston:2003:EFN

[Ban03] John Bankston, editor. Enrico Fermi and the nuclear reactor. Unlocking the secrets of science. Mitchell Lane Publishers, Bear, DE, USA, 2003. ISBN 1-58415-184-6. 48 pp. LCCN QC16.F46 B36 2003.

Banerji:2006:HED

[Ban06] S. Banerji. How Einstein discovered the Special Theory of Relativity. Resonance, 11(2):27–42, February 2006. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X

(electronic). URL http://www.springerlink.com/content/
x775u6603k8n4880/.

Banks:2011:PET

[Ban11] Michael Banks. Plans for Einstein Telescope unveiled. Physics World, 24(6):13, June 2011. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/24/06/phwv24i06a24.
pdf.

Banner:2012:IED

[Ban12] William Banner. “It is easier to denature plutonium than to denature the evil spirit of man” — Albert Einstein. Clinical Toxicology, 50(7):537–538, August 2012. ISSN 1556-3650.

Barnett:1948:EUG

[Bar48a] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe]. Fischer, Frankfurt, West Germany, 1948. 159 pp. LCCN ????

Barnett:1948:UDE

[Bar48b] Lincoln Kinnear Barnett. The universe and Dr. Einstein. W. Sloane Associates, New York, 1948. 127 pp. LCCN QC6 .B33. With a foreword by Albert Einstein.

Bargmann:1949:CBP

[Bar49] V. Bargmann. On the connection between phase shifts and scattering potential. Reviews of Modern Physics, 21(3):488– 493, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.488;   http://rmp.aps.
org/abstract/RMP/v21/i3/p488_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Barnett:1950:EUG

[Bar50a] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe]. Bermann-Fischer, Amsterdam, The Netherlands, 1950. 147 pp. LCCN ???? With a foreword by Albert Einstein.

Barnett:1950:UDEa

[Bar50b] Lincoln Kinnear Barnett. The universe and Dr. Einstein. Gollancz, London, UK, 1950. 112 pp. LCCN QC6 .B33 1950a.

Barnett:1950:UDEb

[Bar50c] Lincoln Kinnear Barnett. The universe and Dr. Einstein. Sloane, New York, NY, USA, 1950. 127 pp. LCCN QC6 .B33 1950. With a foreword by Albert Einstein.

Barnett:1951:EUF

[Bar51a] Lincoln Kinnear Barnett. Einstein et l’univers. (French) [Einstein and the Universe]. Gallimard, Paris, France, 1951. 222 pp. LCCN 4QC 5009. Préface d’Albert Einstein. (The Universe and Dr Einstein). Traduit de l’américain par Julien Néquaud.

Barnett:1951:EUG

[Bar51b] Lincoln Kinnear Barnett. Einstein und das Universum. (German) [Einstein and the Universe]. S. Fischer, Frankfurt am Main, West Germany, 1951. 160 pp. LCCN ???? With a foreword by Albert Einstein.

Barnett:1952:EUG

[Bar52a] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe]. Fischer Bücherei, Frankfurt am Main, West Germany, 1952. 158 pp. LCCN QC16.E5 .B2.

Barnett:1952:UDE

[Bar52b] Lincoln Kinnear Barnett. The universe and Dr. Einstein, volume M71 of A Mentor book. New American Library, New York, NY, USA, revised edition, 1952. 140 pp. LCCN QC6 .B33 1952. With a foreword by Albert Einstein.

Barnett:1954:EUG

[Bar54] Lincoln Kinnear Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer Bücherei, Frankfurt am Main, West Germany, 1954. 158 pp. LCCN ???? With a foreword by Albert Einstein.

Barnett:1955:EUG

[Bar55] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer, Frankfurt am Main, West Germany, 1955. 158 pp. LCCN QC16.E5 .B2. With a foreword by Albert Einstein.

Barnett:1956:EUG

[Bar56] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer Bücherei, Frankfurt am Main, West Germany, 1956. 158 pp. LCCN QC16.E5 .B2. With a foreword by Albert Einstein.

Barnett:1957:EUG

[Bar57a] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer Bücherei, Frankfurt am Main, West Germany, 1957. 158 pp. LCCN QC16.E5 .B2. With a foreword by Albert Einstein.

Barnett:1957:UDE

[Bar57b] Lincoln Kinnear Barnett. The universe and Dr. Einstein. W. Sloane Associates, New York, NY, USA, second revised edition, 1957. 127 pp. LCCN QC6 .B33 1957. With a foreword by Albert Einstein.

Barnett:1958:EUG

[Bar58] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer Bücherei, Frankfurt am Main, West Germany, 1958. 158 pp. LCCN QC16.E5 .B2. With a foreword by Albert Einstein.

Barnett:1962:EUG

[Bar62] Lincoln Barnett. Einstein und das Universum. (German) [Einstein and the Universe], volume 21 of Fischer-Bücherei. Fischer Bücherei, Frankfurt am Main, West Germany, 1962. 158 pp. LCCN QC16.E5 .B2. With a foreword by Albert Einstein.

Baresich:1980:RCA

[Bar80a] M. J. Baresich. The relativity of communication — Albert Einstein as technical writer. Journal of Technical Writing and Communication, 10(2):125–132, ???? 1980. ISSN 0047-2816.

Barnett:1980:UDE

[Bar80b] Lincoln Kinnear Barnett. The universe and Dr. Einstein. American Reprint, Mattituck, NY, USA, revised edition, 1980. ISBN 0-89190-069-1. 140 pp. LCCN QC6 .B33 1980z.

Barnett:1982:UDE

[Bar82] Lincoln Kinnear Barnett. The universe and Dr. Einstein. Time reading program special edition. Time-Life Books, Alexandria, VA, USA, 1982. ISBN 0-8094-3736-8 (deluxe), 0-8094-3737-6 (paperback). xix + 121 pp. LCCN QC173.57 .B37 1982. US$12.95; US$7.95.

Barrow:1988:WWW

[Bar88] John D. Barrow. The world within the world. Clarendon Press, Oxford, UK, 1988. ISBN 0-19-851979-6. xiv + 398 pp. LCCN QC24.5 .B37 1988. US$24.95.

Barrow:1990:WWW

[Bar90] John D. Barrow. The world within the world. Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 1990. ISBN 0-19-286108-5. 398 pp. LCCN Q175.

Barrow:1991:TEQ

[Bar91] John D. Barrow. Theories of everything: the quest for ultimate explanation. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-853928-2. xi + 223 pp. LCCN Q175 .B225 1991.

US$22.95. URL http://www.zentralblatt-math.org/zmath/
en/search/?an=0867.00009.

Barkan:1993:WSF

[Bar93] Diana Kormos Barkan. The witches’ sabbath: The First International Solvay Congress in Physics. Science in Context, 6 (1):59–82, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Barrow:1994:OU

[Bar94a] John D. Barrow. The origin of the universe. Science masters series. Basic Books, New York, NY, USA, 1994. ISBN 0-465-05354-8. xv + 150 pp. LCCN QB981 .B2798 1994. US$20.00, CAN$26.75.

Bartels:1994:EAR

[Bar94b] Andreas Bartels. Von Einstein zu Aristoteles. Raumzeit-Philosophie und Substanz-Metaphysik. (German) [From Einstein to Aristotle. Spacetime-philosophy and substance-metaphysics]. Philosophia Naturalis, 31(2):293–308, 1994. CODEN ???? ISSN 0031-8027.

Barrow:1995:AU

[Bar95] John D. Barrow. The artful universe. Clarendon Press, Oxford, UK, 1995. ISBN 0-19-853996-7. x + 274 pp. LCCN BH301.N3

B37 1995. URL http://www.zentralblattmath.org/zmath/
en/search/?an=0867.00008.

Bartnik:1996:EMR

[Bar96] Robert Bartnik. Einstein, mathematics, and reality. University of New England, Armidale, NSW, Australia, 1996. ISBN 1-86389-296-6. 16 pp. LCCN QC20.5 .B37 1996. An inaugural public lecture delivered at the University of New England, Armidale, New South Wales on 26 July 1995.

Barnett:1998:CAP

[Bar98a] Carol Covington Barnett. A Comparative Analysis of Perspectives of Mileva Marić Einstein. Ph.D. thesis, Florida State University, Tallahassee, FL, USA, Spring 1998. vii +

225 pp. URL https://search.proquest.com/pqdtglobal/
docview/304423009.

Bartholomew:1998:JNC

[Bar98b] James R. Bartholomew. Japanese Nobel candidates in the first half of the Twentieth Century. Osiris (Series 2), 13(??):238– 284, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/
301885.

Barrow:1999:BIS

[Bar99] John D. Barrow. Between inner space and outer space: essays on science, art, and philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-850254-0. xii + 274 pp. LCCN Q162 .B364 1999.

Barrow:2000:BN

[Bar00a] John D. Barrow. The book of nothing. Jonathan Cape, London, UK, 2000. ISBN 0-224-05962-9. 380 pp. LCCN QA141

.B35 2000.  URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.loc.gov/catdir/bios/
random054/2002421946.html;   http://www.loc.gov/catdir/
description/random041/2002421946.html.

Barrow:2000:BNV

[Bar00b] John D. Barrow. The book of nothing: vacuums, voids, and the latest ideas about the origins of the universe. Pantheon Books, New York, NY, USA, 2000. ISBN 0-375-42099-1. xv + 361 pp. LCCN QA141 .B36 2000. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/random052/00058894.html;

http://www.loc.gov/catdir/description/random042/00058894.
html;   http://www.loc.gov/catdir/samples/random041/
00058894.html.

Bartusiak:2000:EUS

[Bar00c] Marcia Bartusiak. Einstein’s unfinished symphony: listening to the sounds of space–time. Joseph Henry Press, Washington, DC, USA, 2000. ISBN 0-309-06987-4. xii + 249 pp. LCCN QC173.59.S65 B39 2000.

Barrow:2002:CNAa

[Bar02a] John D. Barrow. The constants of nature: from Alpha to Omega. Jonathan Cape, London, UK, 2002. ISBN 0-224-06135-6. xv + 352 pp. LCCN QC39 .B38 2002. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/bios/random055/2003464349.html;   http://www.
loc.gov/catdir/description/random0410/2003464349.html;
 http://www.loc.gov/catdir/samples/random051/2003464349.
html.

Barrow:2002:CNAb

[Bar02b] John D. Barrow. The constants of nature: from Alpha to Omega—the numbers that encode the deepest secrets of the universe. Pantheon Books, New York, NY, USA, 2002. ISBN 0-375-42221-8. xv + 352 pp. LCCN QC39 .B37 2002. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/random051/2002075975.html;
 http://www.loc.gov/catdir/description/random042/2002075975.
html;   http://www.loc.gov/catdir/samples/random051/
2002075975.html;   http://www.loc.gov/catdir/toc/fy035/
2002075975.html.

Barnett:2005:UDE

[Bar05a] Lincoln Kinnear Barnett. The universe and Dr. Einstein. Dover, New York, NY, USA, 2005. ISBN 0-486-44519-4 (paperback). 128 pp. LCCN QC6 .B33 2005. URL http://www.loc.gov/

catdir/enhancements/fy0625/2005053025-d.html.

Barrow:2005:AUE

[Bar05b] John D. Barrow. The artful universe expanded. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-280569-X. xii + 321 pp. LCCN BH301.N3

B37 2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/

gutenberg/;   http://www.loc.gov/catdir/toc/ecip055/
2004030354.html.

Barrow:2005:EI

[Bar05c] John D. Barrow. Einstein as icon. Nature, 433(7023):218– 219, January 19, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v433/n7023/full/433218a.html.

Barrow:2005:IBSa

[Bar05d] John D. Barrow. The infinite book: a short guide to the boundless, timeless and endless. Jonathan Cape, London, UK, 2005. ISBN 0-224-06917-9. xvi + 328 pp. LCCN BD411 .B37 2005.

Barrow:2005:IBSb

[Bar05e] John D. Barrow. The infinite book: a short guide to the boundless, timeless, and endless. Pantheon Books, New York, NY, USA, 2005. ISBN 0-375-42227-7. xvi + 328 pp. LCCN QA9 .B282 2005.

Bartholomew:2005:ERO

[Bar05f] James R. Bartholomew. Essay review: One hundred years of the Nobel Science Prizes: Elisabeth Crawford (Editor). Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. viii + 161 pp., Tokyo: Universal Academy Press, 2002. Elisabeth Crawford. The Nobel Population, 1901–1950: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry. vi + 420 pp., Tokyo: Universal Academy Press, 2002. Mauro Dardo. Nobel Laureates and Twentieth-Century Physics. x + 515 pp., Cambridge: Cambridge University Press, 2004. Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. xv + 400 pp., notes, index. New York: W. H. Freeman, 2001. István Hargittai. The Road to Stockholm: Nobel Prizes, Science, and Scientists. xvii + 342 pp., Oxford/New York: Oxford University Press, 2002. George Thomas Kurian. The Nobel Scientists: A Biographical Encyclopedia. 675 pp., Amherst, N.Y.: Prometheus Books, 2002. Isis, 96(4):625–632, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498605;   http:/
/www.jstor.org/stable/3652242. See [CHU87, Cra92, Fri01,

Cra02b, Har02a, Kur02, Dar04a].

Bartusiak:2005:BBB

[Bar05g] Marcia Bartusiak. Beyond the Big Bang: Einstein’s evolving universe. National Geographic Magazine, ??(??):111–121,

May 2005. URL http://www7.nationalgeographic.com/ngm/
0505/feature6/.

Bartusiak:2006:AUD

[Bar06] Marcia Bartusiak, editor. Archives of the universe: 100 discoveries that transformed our understanding of the cosmos. Vintage Books, New York, NY, USA, 2006. ISBN 0-375-71368-9. xvii + 695 pp. LCCN QB15 .A75 2006. US$18.95.

Barbour:2007:EMP

[Bar07] Julian B. Barbour. Einstein and Mach’s principle. In Renn et al. [RSS^{+}07a], pages 569–604. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_32.

Bartusiak:2009:DWF

[Bar09] Marcia Bartusiak. The Day We Found the Universe. Pantheon Books, New York, NY, USA, 2009. ISBN 0-375-42429-6. xviii + 337 pp. LCCN QB15 .B37 2009.

Bartusiak:2015:BHH

[Bar15] Marcia Bartusiak. Black Hole: How an Idea Abandoned by Newtonians, Hated by Einstein, and Gambled on by Hawking Became Loved. Yale University Press, New Haven, CT, USA, 2015. ISBN 0-300-21085-X (hardcover). xii + 237 pp. LCCN QB843.B55 B37 2015.

Barish:2018:NLL

[Bar18a] Barry C. Barish. Nobel lecture: LIGO and gravitational waves II. Reviews of Modern Physics, 90(4):040502–??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL https://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.040502.

Barron:2018:EGLb

[Bar18b] James Barron. Einstein’s 1954 ‘God letter’ sells for $2.9 million. New York Times, 168(58168):A25, December 6, 2018. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Barron:2018:EGLa

[Bar18c] James Barron. Einstein’s ‘God letter’ draws auction interest. New York Times, 168(58165):A25, December 3, 2018. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Basri:1965:OFE

[Bas65] Saul A. Basri. Operational foundation of Einstein’s General Theory of Relativity. Reviews of Modern Physics, 37(2):288– 315, April 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.37.288;   http://rmp.aps.
org/abstract/RMP/v37/i2/p288_1.

Bass:1992:SPP

[Bas92] Ludvik Bass. Schrödinger: a philosopher in Planck’s chair. The British Journal for the Philosophy of Science, 43(1):111– 127, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/43/1/111.full.pdf+html.

Batchelder:1962:ID

[Bat62] Robert C. Batchelder. The irreversible decision, 1939–1950. Houghton Mifflin, Boston, MA, USA, 1962. 306 pp. LCCN BR115.A85 B34.

Batten:1985:EFF

[Bat85] Alan H. Batten. Erwin Finlay-Freundlich, (1885–1964). Journal of the British Astronomical Association, 96(1):33–35, December 1985. CODEN JBAAA6. ISSN 0007-0297. URL http:

//esoads.eso.org/full/1985JBAA...96...33B.

Battimelli:2003:LFS

[Bat03] Giovanni Battimelli. L’eredità di Fermi: storia fotografica dal 1927 al 1959 dagli archivi di Edoardo Amaldi. (Italian) [The legacy of Fermi: photographic history from 1927 to 1959 from the archives of Edoardo Amaldi]. Editori riuniti, Roma, Italia, 2003. ISBN 88-359-5428-2. 220 pp. LCCN Q127.I8 B24 2003.

Batten:2005:SET

[Bat05] Alan H. Batten. Subtle are Einstein’s thoughts. Physics World, 18(9):16–17, September 2005. CODEN PHWOEW.

ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/9/phwv18i9a21.
pdf.

Bauer:1920:ROC

[Bau20] Louis A. Bauer. Résumé of observations concerning the solar eclipse of May 29, 1919, and the Einstein effect. Science, 51(1317):301–311, March 26, 1920. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/51/1317/301.full.pdf.

Baumgardt:1951:JKH

[Bau51] Carola Baumgardt. Johannes Kepler: his life and letters. Philosophical Library, New York, NY, USA, 1951. 209 pp. LCCN QB36.K4 A3 1951. Introduction by Albert Einstein.

Baumer:1974:OH

[Bau74] Franz Baumer. Otto Hahn, volume 78 of Köpfe des XX. Jahrhunderts. Colloquium-Verlag, Berlin, West Germany, 1974. ISBN 3-7678-0367-4. 92 pp. LCCN QD22.H2 B36.

Baumer:1983:RHM

[Bäu83] Änne Bäumer. Rezensionen: Horst Melcher: Albert Einstein wider Vorurteile und Denkgewohnheiten. (Reihe Wissenschaft) Braunschweig: Vieweg 1979, gebunden, 107 Seiten. (German) [Reviews: Horst Melcher: Albert Einstein against prejudices and habits of thought. (Scientific Series) Braunschweig: Vieweg, 1979, hardcover, 107 pages]. Berichte zur Wissenschaftsgeschichte, 6(1–4):251–252, ???? 1983. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Bauer:1995:RNT

[Bau95] Martin W. Bauer, editor. Resistance to new technology: nuclear power, information technology, and biotechnology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-45518-9 (hardcover). LCCN T174.5 .R48 1995. URL http://www.loc.

gov/catdir/description/cam026/94026745.html;   http:
//www.loc.gov/catdir/toc/cam021/94026745.html.

Bauer:1997:RNT

[Bau97] Martin Bauer, editor. Resistance to new technology: nuclear power, information technology, and biotechnology. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-59948-2 (paperback), 0-521-45518-9 (hardcover). LCCN T174.5 .R48 1997.

Baudler:2009:DEJ

[Bau09] Georg Baudler. Darwin, Einstein — und Jesus: Christsein im Universum der Evolution. (German) [Darwin, Einstein — and Jesus: Christianity in the Universe of Evolution]. Patmos, Düsseldorf, Germany, 2009. ISBN 3-491-72535-6. 256 pp. LCCN

????  URL http://d-nb.info/991116216/04;   http:/
/digitool.hbz-nrw.de:1801/webclient/DeliveryManager?
pid=2844334\%26custom\%5Fatt\%5F2=simple\%5Fviewer.

Bawden:2006:EOI

[Baw06] David Bawden. Einstein in the office: is information really necessary? Journal of Documentation, 62(3):??, ???? 2006. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-

7379 (electronic). URL https://dblp.org/db/journals/
jd/jd62.html#Bawden06d;   https://www.wikidata.org/
entity/Q58631434.

Baylis:1994:DBT

[Bay94] John Baylis. The development of Britain’s thermonuclear capability 1954–61: Myth or reality? Contemporary Record: the journal of the Institute of Contemporary British History, 8(1): 159–174, June 1994. ISSN 0950-9224.

Born:1924:ETR

[BB24] Max Born and Henry L. (Henry Leopold) Brose. Einstein’s Theory of Relativity. Methuen and Co. Ltd., London, UK, third edition, 1924. xi + 293 pp. LCCN QC6 .B66; 530 B73; QC6 .B645rE 1924.

Bergmann:1949:NLF

[BB49] Peter G. Bergmann and Johanna H. M. Brunings. Nonlinear field theories II. Canonical equations and quantization. Reviews of Modern Physics, 21(3):480–487, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.480;   http://rmp.aps.org/

abstract/RMP/v21/i3/p480_1. Special issue in honor of Albert Einstein on his 70th birthday.

Born:1955:LSM

[BB55] Max Born and E. Bauer. L’interprétation statistique de la mécanique quantique — (conférence Nobel, 1954). (French) [The statistical interpretation of quantum mechanics (Nobel lecture, 1954)]. Journal de Physique et le Radium, 16(10):737–743, ???? 1955. CODEN JPRAAJ. ISSN 0368-3842.

Born:1969:AEH

[BB69] Max Born and Hedwig Born, editors. Albert Einstein — Hedwig und Max Born. Briefwechsel 1916–55. (German) [Albert Einstein — Hedwig and Max Born. Correspondence 1916–55]. Nymphenburger Verlagshandlung, Munich, Germany, 1969. 329 pp. Preface by Bertrand Russell and foreword by Werner Heisenberg.

Born:1971:BELa

[BB71a] Max Born and Hedwig Born, editors. The Born–Einstein letters: correspondence between Albert Einstein and Max and Hedwig Born from 1916–1955. Macmillan Publishing Company, New York, NY, USA, 1971. ISBN 0-333-11267-9. xi + 240 + 5 pp. LCCN QC16.E5 A4513 1971b. Preface by Bertrand Russell and foreword by Werner Heisenberg. Translated by Irene Born.

Born:1971:BELb

[BB71b] Max Born and Hedwig Born, editors. The Born–Einstein letters; correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1971. ISBN 0-8027-0326-7. x + 240 pp. LCCN QC16.E5 A4513.

Blair:1989:FMG

[BB89] David G. Blair and Michael J. Buckingham, editors. The fifth Marcel Grossmann meeting on recent developments in theoretical and experimental general relativity, gravitation and relativistic fiel theories: proceedings of the meeting held at the University of Wester Australia: 8–13 August 1988. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0784-6 (hardcover), 9971-5-0785-4 (paperback). LCCN QC173.6 .M37 1988.

Barrow:2000:UDI

[BB00] John D. Barrow and John D. Barrow, World within the world. The universe that discovered itself. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-286200-6 (paperback). xvi + 448 pp. LCCN QC24.5 .B368 2000.

Bernardini:2004:CFE

[BB04a] C. (Carlo) Bernardini and Luisa Bonolis, editors. Conoscere Fermi: Enrico Fermi: his work and legacy. Springer, Società Italiana di Fisica, Bologna, Italia, 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag, Berlin, Heidelberg, New York). xii + 410 pp. LCCN

QC774.F4 C6613 2004. URL http://www.loc.gov/catdir/
enhancements/fy0818/2004108212-d.html;   http://www.
loc.gov/catdir/enhancements/fy0818/2004108212-t.html.

Bernardini:2004:EFH

[BB04b] C. (Carlo) Bernardini and Luisa Bonolis, editors. Enrico Fermi: His Work and Legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag, Berlin, Heidelberg, New York). xii + 410 pp. LCCN

QC774.F4 C6613 2004. URL http://www.loc.gov/catdir/
enhancements/fy0818/2004108212-d.html;   http://www.
loc.gov/catdir/enhancements/fy0818/2004108212-t.html.

Born:2005:BEL

[BB05] Max Born and Hedwig Born, editors. The Born–Einstein letters: friendship, politics, and physics in uncertain times: correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Macmillan Publishing Company, New York, NY, USA, 2005. ISBN 1-4039-4496-2. xxxii + 235 pp. LCCN

QC16.E5 A4 2005. URL http://www.loc.gov/catdir/bios/
hol059/2004061027.html;   http://www.loc.gov/catdir/
description/hol054/2004061027.html;   http://www.loc.
gov/catdir/enhancements/fy0618/2004061027-t.html. Note

on the new edition by Gustav Born. New preface by Diana Buchwald and Kip S. Thorne. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. Translated by Irene Born.

Brody:2013:SCY

[BB13] David Eliot Brody and Arnold R. Brody. The science class you wish you had: the seven greatest scientific discoveries in history and the people who made them. Perigee/Penguin Group, New York, NY, USA, second edition, 2013. ISBN 0-399-16032-9 (paperback). xxi + 392 pp. LCCN ???? Foreword by Martin Rodbell.

Bartelmann:2009:FCO

[BBB^{+}09] Matthias Bartelmann, Charles L. Bennett, Carlo Burigana, Cesare Chiosi, Mauro D’Onofrio, Alan Dressler, Isabella Gioia, Günther Hasinger, Juan Francisco Macias-Perez, Piero Madau, Paola Marziani, John Mather, Francesca Matteucci, Keith Olive, John Peacock, Wolfgang Reich, Pierre-Marie Robitaille, Michael Rowan-Robinson, Gart Steigman, Matthias Steinmetz, Jack W. Sulentic, Massimo Turatto, and Simon D. M. White. Fundamental cosmological observations and data interpretation. In D’Onofrio and Burigana [DB09], chapter 2, pages 7–201. ISBN 3-642-00791-0 (print), 3-642-00792-9 (e-book). LCCN QB981 .Q47 2009eb.

Bertotti:1990:MCR

[BBBM90] B. Bertotti, R. Balbinot, S. Bergia, and A. Messina, editors. Modern Cosmology in Retrospect. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37213-5. xx + 426 pp. LCCN QB981 .M774 1990. URL http://www.loc.

gov/catdir/description/cam024/90041803.html;   http:
//www.loc.gov/catdir/toc/cam028/90041803.html.

Bona:2009:ENR

[BBCPL09] Carles Bona, C. (Carles) Bona-Casas, and Carlos Palenzuela-Luque. Elements of Numerical Relativity and Relativistic Hydrodynamics: from Einstein’s Equations to Astrophysical Simulations, volume 783 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 3-642-01163-2, 3-642-01164-0 (ebook). ISSN 0075-8450 (print), 1616-6361 (electronic). xiv + 214 pp. LCCN QC6 .B663 2009.

Bellovin:2011:PSS

[BBD^{+}11] S. M. Bellovin, S. O. Bradner, W. Diffie, S. Landau, and J. Rexford. Privacy and security: As simple as possible—but not more so: An assessment of the U.S. government’s EINSTEIN project. Communications of the ACM, 54(8):30–33, August 2011. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beckmann:2008:GEC

[BBL08] Arnold Beckmann, Edwin Beggs, and Benedikt Löwe. From Gödel to Einstein: Computability between logic and physics at CiE 2006. Theoretical Computer Science, 394(3):141–143, April 8, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergman:1978:AEC

[BBM78] Elihu Bergman, Hans Albrecht Bethe, and Robert Eugene Marshak, editors. American energy choices before the year 2000: [proceedings]. Lexington Books, Lexington, MA, USA, 1978. ISBN 0-669-02398-1. viii + 150 pp. LCCN TJ163.25.U6 C63 1978.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989.

US$11.95. URL http://www.loc.gov/catdir/description/
dover031/89012033.html.

Brian:1998:E

[BBZ98] Denis Brian, Jarosław Bielas, and Maria Zborowska. Einstein. Biografie. Wydawnictwo “Amber”, Warszawa, Poland, 1998. ISBN 83-7169-449-0, 83-7169-685-X. 303 + 1 + 8 pp. LCCN ????

Buck:1971:PMR

[BC71] Roger C. Buck and Robert S. Cohen, editors. PSA 1970: In Memory of Rudolf Carnap. Proceedings of the 1970 Biennial Meeting Philosophy of Science Association, volume 8 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1971. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-

0346. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Bruce-Chwatt:1982:RTM

[BC82] Leonard Jan Bruce-Chwatt. The rise of tropical medicine: Milestones of discovery and application. In Bernhard et al. [BCS82], pages 167–185. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.sciencedirect.

com/science/article/pii/B9780080279398500165.

Bertotti:1994:ESS

[BC94] B. Bertotti and Umberto Curi, editors. Erwin Schrödinger: scienziato e filosofo, volume 3 of Percorsi della scienza, storia, testi, problemi. Il poligrafo, Padova, Italy, 1994. ISBN 88-7115-034-1. LCCN QC16.S265 E77 1994. L30000. Papers presented at the congress held in Venice, 1987, now revised.

Bernstein:1995:BAF

[BC95] Jeremy Bernstein and David Cassidy. Bomb apologetics: Farm Hall, August 1945. Physics Today, 48(8):32–36, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v48/i8/p32_s1.

Berenson:2001:PMC

[BC01] Bernard G. Berenson and Robert R. Carkhuff. The possibilities mind: conversations with God, Einstein and others. Possibilities Pub., Amherst, MA, USA, 2001. ISBN 0-87425-630-5, 1-4237-1639-6 (e-book). xxiv + 164 pp. LCCN BC199.P7 B47 2001.

Bacciagaluppi:2024:EPD

[BC24] Guido Bacciagaluppi and Elise Crull. The Einstein Paradox: the Debate on Nonlocality and Incompleteness in 1935. Cambridge University Press, Cambridge, UK, 2024. ISBN 1-107-01445-X.

???? pp. LCCN ???? URL https://www.cambridge.org/us/
universitypress/subjects/physics/history-philosophy-
and-foundations-physics/einstein-paradox-debate-nonlocality-
and-incompleteness-1935.

Bundy:1993:RND

[BCD93] McGeorge Bundy, William J. Crowe, and Sidney D. (Sidney David) Drell. Reducing nuclear danger: the road away from the brink. Council on Foreign Relations Press, New York, NY, USA, 1993. ISBN 0-87609-149-4. ix + 107 pp. LCCN UA23 .B7864 1993.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Belgiorno:2011:DBH

[BCO^{+}11] F. Belgiorno, S. L. Cacciatori, G. Ortenzi, L. Rizzi, V. Gorini, and D. Faccio. Dielectric black holes induced by a refractive index perturbation and the Hawking effect. Physical Review D (Particles and Fields), 83(2):024015, January 2011. CO-DEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic),

1538-4500 (CD-ROM). URL http://link.aps.org/doi/10.
1103/PhysRevD.83.024015.

Beller:1993:ECS

[BCR93] Mara Beller, R. S. (Robert Sonné) Cohen, and Jürgen Renn, editors. Einstein in context: a special issue of Science in Context. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44834-4. 368 pp. LCCN QC16.E5 E512 1993. URL http:/

/www.loc.gov/catdir/description/cam031/94134891.html.

Bernhard:1982:STS

[BCS82] Carl Gustaf Bernhard, Elisabeth T. Crawford, and Per Sörbom, editors. Science, technology, and society in the time of Alfred Nobel: Nobel symposium 52, held at Björkborn, Karlskoga, Sweden, 17–22 August 1981. Published for the Nobel Foundation by Pergamon Press, Oxford, UK, 1982. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.).

Bitbol:1992:ESP

[BD92] Michel Bitbol and Olivier Darrigol, editors. Erwin Schrödinger: philosophy and the birth of quantum mechanics = philosophie et naissance de la mécanique quantique. Editions Frontières, Gifsur-Yvette, France, 1992. ISBN 2-86332-116-1. xiii + 460 pp. LCCN QC16.S265 E78 1992. 263.75F.

Brown:1989:PQP

[BDH89] Laurie M. Brown, Max Dresden, and Lillian Hoddeson, editors. Pions to quarks: particle physics in the 1950s: based on a Fermilab symposium. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-30984-0. LCCN M90.E01541; QC793.

Buttner:2000:RQD

[BDH^{+}00] Jochen Büttner, Olivier Darrigol, Dieter Hoffmann, Jürgen Renn, and Matthias Schemmel. Revisiting the Quantum Discontinuity, volume 150 of Preprint (Max-Planck-Institut für Wissenschaftsgeschichte). Max-Planck Institut für Wissenschaftsgeschichte, Berlin, Germany, 2000. ISSN 0948-9444. 68 pp.

LCCN Q124.6 .P74 no.150. URL http://www.mpiwg-berlin.
mpg.de/Preprints/P150.PDF.

Barrow:2004:SUR

[BDH04] John D. Barrow, P. C. W. Davies, and Charles L. Harper, editors. Science and ultimate reality: quantum theory, cosmology, and complexity. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83113-X. xx + 721 pp. LCCN QC174.12

.S4 2004. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
cam051/2003055903.html;   http://www.loc.gov/catdir/
toc/cam051/2003055903.html.

Blum:1993:WHG

[BDR93] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung A. Teil III. Wissenschaftliche Originalarbeiten. (German) [Werner Heisenberg: Collected works. Series A. Part III. Original scientific papers]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-13848-X. x + 700 pp. Annotated by Rolf Hagedorn, Heinz Koppe, H.-P. Dürr and H. Rechenberg.

Brennecke:2007:CQB

[BDR^{+}07] Ferdinand Brennecke, Tobias Donner, Stephan Ritter, Thomas Bourdel, Michael Köhl, Tilman Esslinger, et al. Cavity QED with a Bose–Einstein condensate. Nature, 450(7167):268–271, November 8, 2007. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v450/n7167/full/nature06120.html.

Brush:1969:MOR

[BE69] S. G. Brush and C. W. F. Everitt. Maxwell, Osborne Reynolds, and the radiometer. Historical Studies in the Physical Sciences, 1(??):105–125, ???? 1969. CODEN HSPSAS. ISSN 0073-2672.

URL http://www.jstor.org/stable/27757296.

Bassett:2002:IR

[BE02] Bruce Bassett and Ralph Edney. Introducing Relativity. Icon, Cambridge, UK, 2002. ISBN 1-84046-372-4. 176 pp. LCCN QC 173.55 .B37 2002.

Butterfield:2007:PP

[BE07] Jeremy Butterfield and John Earman, editors. Philosophy of physics. Handbook of the philosophy of science. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-51560-7 (set), 0-444-53001-0 (part A), 0-444-53002-9 (part B). ???? pp.

LCCN QC6 .P44 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006050854-d.html.

Burnett:1999:TBE

[BEC99] Keith Burnett, Mark Edwards, and Charles W. Clark. The theory of Bose–Einstein condensation of dilute gases. Physics Today, December 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.882899.

Bederson:2001:BRM

[Bed01a] Benjamin Bederson. Book review: Mary Palevsky, Atomic Fragments: A Daughter’s Questions. Berkeley and Los Angeles: University of California Press, 2000, xiv + 289 pages. $24.95 (cloth). Physics in Perspective (PIP), 3(2):253–254, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2001:SAP

[Bed01b] Benjamin Bederson. SEDs at Los Alamos: a personal memoir. Physics in Perspective (PIP), 3(1):52–75, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050056.

Behr:2005:SFG

[Beh05a] André Behr. Der Streit um die Formel. (German) [the dispute about the formula]. Neue Zürcher Zeitung, ??(??):??, July 6,

2005. URL https://www.nzz.ch/articleCVPKL-1.148578.

Behr:2005:EPP

[Beh05b] André Behr. Einsteins Probleme mit dem Peer-Review-System. (German) [Einstein’s problems with the peer review system].

Neue Zürcher Zeitung, ??(??):??, December 28, 2005. URL

https://www.nzz.ch/articled7uu6-1.176461.

Bridgman:1955:NWA

[BEI^{+}55a] P. W. Bridgman, Albert Einstein, L. Infeld, H. J. Muller, C. F. Powell, J. Rotblat, Bertrand Russell, and Hideki Yukawa. Nuclear war: An appeal to scientists and the public. The Lancet, 266(6881):131–132, July 16, 1955.

Bridgman:1955:SAA

[BEI^{+}55b] Percy W. Bridgman, Albert Einstein, Leopold Infeld, Hermann Joseph Muller, Cecil F. Powell, Bertrand Russell, and Hideki Yukawa. Scientists appeal for abolition of war. Bulletin of the Atomic Scientists, 11(7):236–237, September 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint in [RE63].

Bell:1926:RPR

[Bel26] E. T. Bell. Recent publications: Reviews: Einstein’s Theory of Relativity, by H. L. Brose. American Mathematical Monthly, 33 (3):145–147, March 1926. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bell:1964:EPR

[Bel64]  John S. Bell.  On the Einstein–Podolsky–Rosen para-

dox. Physics, 1(3):195–200, 1964. ISSN ???? URL

http://en.wikipedia.org/wiki/EPR_paradox;   http://
plato.stanford.edu/entries/bell-theorem/;   http://
scienceworld.wolfram.com/physics/Einstein-Podolsky-
RosenParadox.html;   http://www.ncsu.edu/felder-public/
kenny/papers/bell.html;   http://www.physics.princeton.
edu/~mcdonald/examples/QM/bell_physics_1_195_64.pdf.

Reprinted in [Bel87, pp. 14–21].

Belinfante:1975:MTR

[Bel75] F. J. (Frederik Jozef) Belinfante. Measurement and time reversal in objective quantum theory, volume 75 of International series in natural philosophy. Pergamon, New York, NY, USA, 1975. ISBN 0-08-018152-X. xix + 142 pp. LCCN QC174.12 .B44.

Bell:1987:SUQ

[Bel87] J. S. Bell. Speakable and unspeakable in quantum mechanics: collected papers on quantum philosophy. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33495-0. xii + 212 pp. LCCN QC173.97 .B45 1987.

Belloni:1988:FR

[Bel88a] Lanfranco Belloni. Da Fermi a Rubbia. (Italian) [From Fermi to Rubbia]. Rizzoli, Milano, Italia, 1988. ISBN 88-17-53095-6. 217 pp. LCCN QC9.I8 B44 1988. L22000.

Belloni:1988:FRI

[Bel88b] Lanfranco Belloni. Da Fermi a Rubbia. (Italian) [From Fermi to Rubbia]. Rizzoli, Milano, Italia, 1988. ISBN 88-17-53095-6. 217 pp. LCCN QC9.I8 B44 1988. L22000.

Bell:1990:AM

[Bel90] John Bell. Against ‘measurement’. Physics World, 3(8):33–41, August 1990. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Beller:1993:EBR

[Bel93] Mara Beller. Einstein and Bohr’s rhetoric of complementarity. Science in Context, 6(1):241–255, Spring 1993. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bell:1996:HRF

[Bel96a] E. S. Bell. He’s relatively familiar: Albert Einstein in contemporary American fiction. Journal of American Culture, 19(2): 119–125, Summer 1996. ISSN 0191-1813.

Belousek:1996:EUH

[Bel96b] Darrin W. Belousek. Einstein’s 1927 unpublished hiddenvariable theory: Its background, context and significance. Studies in History and Philosophy of Modern Physics, 27(4):437– 461, December 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219896000159.

Bell:1998:SRD

[Bel98] M. Bell. ‘Sounds of the resurrected dead man’s footsteps #10.2. letter to Albert Einstein’. Poetry, 172(3):145, June 1998. ISSN 0032-2032.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999. URL http://www.loc.

gov/catdir/description/uchi052/99035499.htm;   http:
//www.loc.gov/catdir/toc/uchi052/99035499.htm.

Bell:2004:SUQ

[Bel04] John S. Bell. Speakable and unspeakable in quantum mechanics: collected papers on quantum philosophy. Cambridge University Press, Cambridge, UK, second edition, 2004. ISBN 0-521-52338-9 (paperback), 0-521-81862-1 (hardcover). xxxix + 248 pp. LCCN QC173.97 .B45 2004. URL http://www.loc.gov/

catdir/description/cam051/2004557644.html;   http://
www.loc.gov/catdir/toc/cam051/2004557644.html.

Belton:2005:GSK

[Bel05] Neil Belton. A game with sharpened knives. Weidenfeld and Nicolson, London, UK, 2005. ISBN 0-297-64359-2 (cased), 0-297-84814-3 (trade paperback). 328 pp. LCCN

PR6102.E45 G36 2005. URL http://www.loc.gov/catdir/
enhancements/fy0628/2005432799-b.html;   http://www.
loc.gov/catdir/enhancements/fy0628/2005432799-d.html.

Beltran:2023:QBE

[Bel23] Lester Beltran. Quantum Bose–Einstein statistics for indistinguishable concepts in human language. Foundations of Science, 28(1):43–55, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09794-1.

Brunskell-Evans:2012:RRR

[BEM12] Heather Brunskell-Evans and Michele Moore, editors. Reimagining research for reclaiming the academy in Iraq: identities and participation in post-conflict enquiry: the Iraq Research Fellowship Programme, volume 15 of Studies in inclusive education. Sense Publishers, Rotterdam, The Netherlands, 2012. ISBN 94-6091-897-2 (e-book). xiii + 97 pp. LCCN LA1468 .B78 2012.

Benson:1979:LIC

[Ben79] Ruth Benson. Letter: an inspired cover. Bulletin of the Atomic Scientists, 35(3):86, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benedikt:1999:PAE

[Ben99] M. Benedikt. ‘Professor Albert Einstein allegedly commits a crime; &, of the difficulty of finding a jury of his peers’. Paris Review, ??(151):37–42, Summer 1999. ISSN 0031-2037.

Benitez:2001:AER

[Ben01] Hermes H. Benítez. Albert Einstein y la religión: un estudio sobre ciencia y creencia. (Spanish) [Albert Einstein and religion: a study on science and belief]. RiL Ed., Providencia, Santiago de Chile, Chile, 2001. ISBN 956-284-190-1, 956-284-191-x. 131 pp. LCCN ???? EUR 15.71.

Bennett:2005:AOL

[Ben05] Charles L. Bennett. Astrophysical observations: Lensing and eclipsing Einstein’s theories. Science, 307(5711):879–884, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/307/5711/879.full.pdf. Special Issue: Einstein’s

Legacy.

Benedetto:2011:BHR

[Ben11] E. Benedetto. Brief history of a relativistic century. International Journal of Theoretical Physics, 50(8):2493–2513, ???? 2011. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575

(electronic). URL http://www.springerlink.com/content/
5162r573h13mq311/.

Bennett:2014:WRI

[Ben14] Jeffrey O. Bennett. What is relativity?: an intuitive introduction to Einstein’s ideas, and why they matter. Columbia University Press, New York, NY, USA, 2014. ISBN 0-231-16726-1 (hardcover), 0-231-53703-4 (e-book). viii + 192 pp. LCCN QC173.57 .B46 2014.

Benedict:2016:OEN

[Ben16a] Marie Benedict. The other Einstein: a novel. Sourcebooks Landmark, Naperville, Illinois, 2016. ISBN 1-4926-3726-2, 1-4926-3727-0. U? pp. LCCN PS3620.E75 O84 2016. URL

http://banq.lib.overdrive.com/ContentDetails.htm?id=
C8059040-E457-418E-911A-8FDD880FC3F8.

Benov:2016:MPF

[Ben16b] Dobriyan M. Benov. The Manhattan Project, the first electronic computer and the Monte Carlo method. Monte Carlo Methods and Applications, 22(1):73–??, March 2016. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.
2016.22.issue-1/mcma-2016-0102/mcma-2016-0102.xml.

Bergson:1922:DSP

[Ber22] Henri Bergson. Durée et simultanéité à propos de la théorie d’Einstein. (French) [Duration and simultaneity: on Einstein’s theory]. Alcan, Paris, France, 1922. viii + 245 pp. LCCN ????

Bergmann:1942:ITR

[Ber42] Peter Gabriel Bergmann. Introduction to the Theory of Relativity. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1942. xvi + 287 pp. LCCN ????

Bergmann:1956:FYR

[Ber56] Peter Gabriel Bergmann. Fifty years of relativity. Science, 123(3195):487–494, March 23, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/123/3195/487.

Bergengren:1962:ANM

[Ber62] Erik Bergengren. Alfred Nobel, the man and his work. T. Nelson and Sons, London, UK, 1962. xviii + 222 pp. LCCN TP268.5.N7 B4 1962. With a supplement on the Nobel institutions and the Nobel Prizes by Nils K. Stêahle. Translated by Alan Blair.

Bergson:1965:DSR

[Ber65] Henri Bergson. Duration and simultaneity, with reference to Einstein’s theory, volume 199 of The library of liberal arts. Bobbs-Merrill, Indianapolis, IN, USA and New York, NY, USA, 1965. xlvi + 190 pp. LCCN BD638 .B413. Translated by Leon Jacobson. With an introduction by Herbert Dingle.

Berger:1968:CPT

[Ber68a] George Berger. The conceptual possibility of time travel. The British Journal for the Philosophy of Science, 19(2):152–155, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/19/2/152.full.pdf+html.

Bergmann:1968:RG

[Ber68b] Peter Gabriel Bergmann. The riddle of gravitation. Scribner, New York, NY, USA, 1968. xvi + 270 pp. LCCN QC6 .B454.

Bergmann:1969:RG

[Ber69a] Peter Gabriel Bergmann. The riddle of gravitation. J. Murray, London, UK, 1969. ISBN 0-7195-1919-5. xvi + 271 pp. LCCN QC6 .B454 1969.

Berninger:1969:OHB

[Ber69b]  Ernst Berninger.  Otto Hahn, eine Bilddokumentation:

Persönlichkeit, wissenschaftliche Leistung, öffentliches Wirken. (German) [Otto Hahn, a pictorial record: personality, academic performance, public work]. H. Moos, München, West Germany, 1969. 105 pp. LCCN QD22.H2 B4.

Berninger:1970:OH

[Ber70] Ernst Berninger. Otto Hahn. Inter Nationes, Bonn–Bad Godesberg, West Germany, 1970. 79 pp. LCCN QD22.H2 B38.

Bernstein:1973:E

[Ber73] Jeremy Bernstein. Einstein. Viking, New York, NY, USA, 1973. ISBN 0-670-29077-7, 0-670-01959-3 (paperback). xii + 242 pp. LCCN QC16.E5 B45. Most of the material appeared originally in the New Yorker.

Bergman:1974:PRA

[Ber74a] H. Bergman. Personal remembrances of Albert Einstein. In Logical and Epistemological Studies in Contemporary Physics [CW74], chapter 15, pages 388–394. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56. URL https://link.springer.

com/chapter/10.1007/978-94-010-2656-7_15.

Berkson:1974:FFDa

[Ber74b] William Berkson. Fields of force; the development of a world view from Faraday to Einstein. Routledge and K. Paul, London, UK, 1974. ISBN 0-7100-7626-6. xiii + 370 pp. LCCN QC173.7 .B47 1974b.

Berkson:1974:FFDb

[Ber74c] William Berkson. Fields of force; the development of a world view from Faraday to Einstein. Wiley, New York, NY, USA, 1974. ISBN 0-470-07029-3. xiii + 370 pp. LCCN QC173.7 .B47 1974.

Berninger:1974:OHS

[Ber74d] Ernst Berninger. Otto Hahn in Selbstzeugnissen und Bilddokumenten. (German) [Otto Hahn in autobiographies and image documents]. Rowohlt, Reinbek, West Germany, 1974. 155 pp. LCCN QD22.H2 B42.

Bergmann:1976:ITR

[Ber76a] Peter Gabriel Bergmann. Introduction to the Theory of Relativity. Dover, New York, NY, USA, 1976. ISBN 0-486-63282-2. xii + 307 pp. LCCN QC173.55 .B47 1976. URL http://www.loc.

gov/catdir/description/dover031/75032903.html.

Bernstein:1976:ABC

[Ber76b] Barton J. Bernstein, editor. The Atomic Bomb: The Critical Issues. Critical issues in American history series. Little, Brown and Co., Boston, MA, USA, 1976. xix + 169 pp. LCCN E183 .A85.

Bernstein:1976:E

[Ber76c] Jeremy Bernstein. Einstein. Penguin, London, UK and New York, NY, USA, 1976. ISBN 0-14-004317-9. xii + 242 pp. LCCN QC16.E5 B45 1976. Most of the material appeared originally in the New Yorker. Reprint of the 1973 Viking Press edition.

Bergmann:1979:DTR

[Ber79a] Peter G. Bergmann. The development of the Theory of Relativity. In Aichelburg and Sexl [AS79], pages 1–16. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_1/.

Bernstein:1979:AEH

[Ber79b] J. Bernstein. Albert Einstein — human side — H. Dukas and B. Hoffmann. New York Times Book Review, 84(10):3, ???? 1979. ISSN 0028-7806.

Bernardini:1981:SES

[Ber81] Paolo Bernardini. La statistica di Einstein e la seconda quantizzazione: quale continuità?. (Italian) [Einstein’s statistics and second quantization: what continuity?]. Physis: Rivista Internazionale di Storia della Scienza, 23(3):337–374, 1981. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Bergmann:1982:QUG

[Ber82a] Peter G. Bergmann. Quest for unity: General Relativity and unitary field theories. In Holton and Elkana [HE82], pages 27– 38. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Bergmann:1982:RAE

[Ber82b] Peter G. Bergmann. Reminiscences [of Albert Einstein]. In Holton and Elkana [HE82], pages 397–400. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Berlin:1982:EI

[Ber82c] Isaiah Berlin. Einstein and Israel. In Holton and Elkana [HE82], pages 281–292. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Bernstein:1982:MJR

[Ber82d] Barton J. Bernstein. In the matter of J. Robert Oppenheimer. Historical Studies in the Physical Sciences, 12(2):195–252, ???? 1982. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757496.

Bernstein:1983:HBD

[Ber83] Barton J. Bernstein. The H-bomb decisions: were they inevitable? In Brodie et al. [BIK83], pages 330–336. ISBN 0-89946-172-7. LCCN UA10.5 .N28 1983.

Lemaitre:1984:BBG

[Ber84a] André Berger, editor. The Big bang and Georges Lemaître: proceedings of a symposium in honour of G. Lemaître fifty years after his initiation of Big-Bang cosmology, Louvain-la-Neuve, Belgium, 10–13 October 1983. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1848-2. LCCN QB980 .B54 1984.

Berger:1984:JCM

[Ber84b] Melvyn S. (Melvyn Stuart) Berger, editor. J. C. Maxwell, the sesquicentennial symposium: new vistas in mathematics, science, and technology. North-Holland, Amsterdam, The Netherlands, 1984. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Bernstein:1985:MWS

[Ber85] Jeremy Bernstein. The man who sees past time. Johns Hopkins Magazine, ??(??):22–33, ???? 1985. ISSN 0021-7255.

Bergmann:1987:RG

[Ber87a] Peter Gabriel Bergmann. The riddle of gravitation. Scribner, New York, NY, USA, revised and updated edition, 1987. ISBN 0-684-18460-5. xix + 233 pp. LCCN QC173.55 .B49 1987. US$18.95.

Bernstein:1987:LIB

[Ber87b] Jeremy Bernstein. The Life It Brings: One Physicist’s Beginnings. Ticknor and Fields, New York, NY, USA, 1987. ISBN 0-89919-470-2. xv + 171 + 8 pp. LCCN QC16.B458 A3 1987.

Bernstein:1988:FPB

[Ber88] Barton J. Bernstein. Four physicists and the bomb: The early years, 1945–1950. Historical Studies in the Physical and Biological Sciences, 18(2):231–263, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://

www.jstor.org/stable/27757603.

Bernstein:1989:CLB

[Ber89] Jeremy Bernstein. A critic at large: Besso. The New Yorker, 65 (??):86–??, February 27, 1989. ISSN 0028-792X. URL http://

www.newyorker.com/archive/1989/02/27/1989_02_27_086_
TNY_CARDS_000351993.

Bernhof:1990:AAE

[Ber90a] R. Bernhof. ‘Akte Albert Einstein’. Sinn und Form, 42(1): 127–128, January/February 1990. ISSN 0037-5756.

Bertschinger:1990:EBR

[Ber90b] Edmund Bertschinger. Einstein’s blunder resurrected. Nature, 348(6303):675–676, December 27, 1990. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v348/n6303/pdf/
348675a0.pdf.

Bernstein:1991:QP

[Ber91] Jeremy Bernstein. Quantum profiles. Princeton University Press, Princeton, NJ, USA, 1991. ISBN 0-691-08725-3. viii + 178 pp. LCCN QC174.12 .B464 1991. URL http://www.

loc.gov/catdir/description/prin031/90040843.html.

Bergman:1992:MAE

[Ber92a] A. Bergman. The ‘murder of Albert Einstein’ — T. Gitlin. New York Times Book Review, ??(??):36, October 18, 1992. ISSN 0028-7806.

Bergmann:1992:RG

[Ber92b] Peter Gabriel Bergmann. The riddle of gravitation. Dover, New York, NY, USA, revised and updated edition, 1992. ISBN 0-486-27378-4, 0-486-26778-4 (paperback). xxii + 233 pp. LCCN

QC173.55 .B49 1992. URL http://www.loc.gov/catdir/
description/dover031/92022218.html.

Bernstein:1993:SCT

[Ber93] Barton J. Bernstein. Seizing the contested terrain of early nuclear history: Stimson, Conant, and their Allies explain the decision to use the atomic bomb. Diplomatic History, 17 (1):35–72, January 1993. ISSN 0145-2096 (print), 1467-7709

(electronic). URL http://onlinelibrary.wiley.com/doi/
10.1111/j.1467-7709.1993.tb00158.x.

Berlin:1994:UCB

[Ber94] J. B. Berlin. Unpublished correspondence between Antoinette Von Kahler and Hermann Broch with special consideration of several unpublished letters of Richard Beerhofmann, Albert Einstein and Thomas Mann. Modern Austrian Literature, 27(2): 39–76, ???? 1994. ISSN 0026-7503.

Bernstein:1995:ABR

[Ber95] Barton J. Bernstein. The atomic bombings reconsidered. Foreign Affairs, 74(1):135–152, ???? 1995. CODEN FRNAA3. ISSN 0015-7120.

Bernstein:1996:AEF

[Ber96a] Jeremy Bernstein. Albert Einstein and the frontiers of physics. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-509275-9, 0-19-512029-9. 189 pp. LCCN QC16.E5 B44 1996.

Bernstein:1996:AES

[Ber96b] Jeremy Bernstein. Albert Einstein und die Schwarzen Löcher. (German) [Albert Einstein and black holes]. Spektrum der Wissenschaft (German translation of Scientific American), August 1, 1996. CODEN SPEKDI. ISSN 0170-2971. URL http://www.

spektrum.de/alias/dachzeile/albert-einstein-und-die-

schwarzen-loecher/823187. German translation of [Ber96d].

Bernstein:1996:HUC

[Ber96c] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Bernstein:1996:RFB

[Ber96d] Jeremy Bernstein. The reluctant father of black holes [Einstein]. Scientific American, 274(6):80–85, June 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

https://www.math.utah.edu/pub/tex/bib/einstein.bib;
 http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-80.pdf;   http://www.
sciam.com/0696issue/0696currentissue.html.

Bernstein:1998:R

[Ber98] Jeremy Bernstein. The road to E = mc^{2}. Times Higher Education Supplement, ??(??):??, September 25, 1998. ISSN 1364-2995.

Berry:2000:GPR

[Ber00] Michael Berry. Great physicists recalled and respected: [book review:] The Genius of Science: A Portrait Gallery of Twentieth-Century Physicists, Abraham Pais, 2000, Oxford University Press, 364pp, £25.00/$35.00hb. Physics World, 13 (6):56, June 2000. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://stacks.iop.org/2058-
7058/13/i=6/a=32.

Bernstein:2001:MPO

[Ber01] Jeremy Bernstein. The Merely Personal: Observations on Science and Scientists. Ivan R. Dee, Chicago, IL, USA, 2001. ISBN 1-56663-344-3. 245 pp. LCCN Q171 .B5375 2001.

Bernstein:2002:HFE

[Ber02] Jeremy Bernstein. Hoover’s folly over Einstein. Physics World, 15(10):43, October 2002. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/15/10/phwv15i10a39.
pdf.

Bernstein:2003:BRB

[Ber03] Jeremy Bernstein. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. American Journal of Physics, 71(4):411–415, April 2003. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/
v71/i4/p411_s1;   http://link.aip.org/link/?AJP/71/

411/1. See comment [Her03].

Bernstein:2004:OPE

[Ber04] Jeremy Bernstein. Oppenheimer: portrait of an enigma. Ivan R. Dee, Chicago, IL, USA, 2004. ISBN 1-56663-569-1. xi + 223 pp. LCCN QC16.O62 B43 2004.

Berlin:2005:MBM

[Ber05] Leslie Berlin. The man behind the microchip: Robert Noyce and the invention of Silicon Valley. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-516343-5. xi + 402 pp. LCCN TK7807.N69 B47 2005.

Berglyd:2006:OFH

[Ber06a] Jostein Berglyd. Operation Freshman: the hunt for Hitler’s heavy water. Leandoer and Ekholm Förlag, Stockholm, Sweden, 2006. ISBN 91-975895-9-4 (hardcover). 202 pp. LCCN D794.5 .B4713 2006. Translation by Tim Dinan from the Norwegian original.

Bergman:2006:BRE

[Ber06b] Aaron Bergman. Book review: Not Even Wrong: The Failure of String Theory and the Continuing Challenge to Unify the Laws of Physics, Peter Woit, Jonathan Cape, London 2006. World Wide Web document, August 18, 2006. URL http://zippy.

ph.utexas.edu/~abergman/Review.pdf.

Bernstein:2006:SOO

[Ber06c] Jeremy Bernstein. Secrets of the old one: Einstein, 1905. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-26005-6 (hardcover), 0-387-25900-7 (e-book). xvi + 216 pp. LCCN QC173.6 .B47 2006.

Bergia:2007:BRA

[Ber07a] Silvio Bergia. Book review: Albert Einstein, The Collected Papers of Albert Einstein, Volume 10: the Berlin Years: Correspondence, May–December 1920, and Supplementary Correspondence, 1909–1920. Princeton and Oxford: Princeton University Press, 2006. lxix + 686 pp., ISBN 978-0-691-12825-2. Nuncius, 22(2):404–405, ???? 2007. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00932.

Berman:2007:IGR

[Ber07b] Marcelo Samuel Berman. Introduction to General Relativity and the Cosmological Constant Problem. Nova Science Publishers, New York, NY, USA, 2007. ISBN 1-59454-717-3. xiv + 213 pp. LCCN QC173.6 .B468 2007. URL http://www.loc.gov/

catdir/toc/ecip066/2006000430.html.

Bernstein:2007:PHW

[Ber07c] Jeremy Bernstein. Plutonium: a history of the world’s most dangerous element. Joseph Henry Press, Washington, DC, USA, 2007. ISBN 0-309-10296-0 (hardcover), 1-280-84457-4, 0-309-10773-3 (e-book). x + 194 + 8 pp. LCCN QD181.P9

B47 2007. URL http://www.loc.gov/catdir/toc/ecip075/
2006038466.html.

Bernstein:2008:NWW

[Ber08] Jeremy Bernstein. Nuclear weapons: what you need to know. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88408-X. xi + 299 pp. LCCN U264 .B453 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0743/2007016625-
b.html;   http://www.loc.gov/catdir/enhancements/fy0743/
2007016625-d.html;   http://www.loc.gov/catdir/toc/
ecip0716/2007016625.html.

Bernstein:2009:PHW

[Ber09] Jeremy Bernstein. Plutonium: a history of the world’s most dangerous element. Cornell University Press, Ithaca, NY, USA, 2009. ISBN 0-8014-7517-1. x + 194 + 8 pp. LCCN

???? URL http://catdir.loc.gov/catdir/toc/ecip075/
2006038466.html.

Berman:2011:REM

[Ber11a] Marcelo Samuel Berman. Realization of Einstein’s Machian program. Nova Science Publishers, Hauppauge, NY, USA, 2011. ISBN 1-61942-164-X (hardcover). xiv + 164 pp. LCCN QB991.C65 B47 2011.

Bernstein:2011:MCW

[Ber11b] Jeremy Bernstein. A memorandum that changed the world. American Journal of Physics, 79(5):440–446, May 2011. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bernstein:2018:BDO

[Ber18] Jeremy Bernstein. A Bouquet of Dyson and Other Reflections on Science and Scientists. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2018. ISBN 981-323-192-0 (hardcover), 981-323-828-3 (paperback). xi + 175 pp. LCCN QC16.D95 A4 2018.

Bose:1986:SQO

[BES86] S. N. Bose, A. Einstein, and E. Schrödinger. La statistica quantistica e le onde di materia. (Italian) [Quantum statistics and matter waves], volume 3 of Saggi di Scienze e Filosofia Naturale [Essays on Natural Philosophy and Natural Science]. Bibliopolis, Naples, Italy, 1986. ISBN 88-7088-128-8. 99 pp. Translated from the German by Paolo Bernardini.

Bess:1993:RUM

[Bes93] Michael Bess. Realism, utopia, and the mushroom cloud: four activist intellectuals and their strategies for peace, 1945–1989: Louise Weiss (France), Leo Szilard (USA), E. P. Thompson (England), Danilo Dolci (Italy). University of Chicago Press, Chicago, IL, USA and London, UK, 1993. ISBN 0-226-04420-3 (hardcover), 0-226-04421-1 (paperback). xxviii + 322 pp. LCCN JX1962.A2

B47 1993. URL http://www.loc.gov/catdir/description/
uchi052/93009707.html;   http://www.loc.gov/catdir/
enhancements/fy0608/93009707-t.html;   http://www.loc.
gov/catdir/enhancements/fy0609/93009707-b.html.

Beste:2004:GPQ

[Bes04] Dieter Beste, editor. Grosse Physiker: Quantenrätsel — Werner Heisenberg und die Unschärfe; Relativitätstheorie — Albert Einstein und die Schwarzen Löcher; Zwei Klassiker —- Isaac Newton und James Clerk Maxwell. (German) [Great physicists: Quantum mystery — Werner Heisenberg and Uncertainty; Relativity theory — Albert Einstein and black holes; Two classical authors — Isaac Newton and James Clerk Maxwell]. Spektrum der Wissenschaft Dossier. Spektrum-der-Wiss.-Verl.-Ges., Heidelberg, Germany, 2004. ISBN 3-936278-75-X. 82 pp. LCCN ????

Bethe:1939:EPSa

[Bet39] Hans A. Bethe. Energy production in stars. Physical Review, 55(5):434–456, March 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.55.434;   http://
prola.aps.org/abstract/PR/v55/i5/p434_1. See further

work [Hoy46, Hoy54]. According to [BL09, page 41], this paper won Bethe the A. Cressy Morrison Prize of the New York Academy of Sciences, and almost three decades later, the 1967 Nobel Prize in Physics.

Bethe:1950:HB

[Bet50a] Hans A. Bethe. The hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):99–104, 125, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bet63a].

Bethe:1950:HBI

[Bet50b] Hans A. Bethe. The hydrogen bomb: II. Scientific American, 182(4):18–23, April 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v182/n4/
pdf/scientificamerican0450-18.pdf. See also Parts I

[Rid50] and III [Bac50]. See [Hol03a] for a comment about the delay in publishing this paper, due to confiscation by the Atomic Energy Commission of the printing plates.

Bethe:1963:AHB

[Bet63a] Hans Bethe. The American hydrogen bomb. In Grodzins and Rabinowitch [GR63], pages 144–155. LCCN D842 .B78. Reprint of [Bet50a].

Bethe:1963:NAW

[Bet63b] Hans A. Bethe. Nobel Award winners announced. Science, 142 (3594):937–938, November 15, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/142/3594/937.full.pdf.

Bethe:1967:NPP

[Bet67] Hans Bethe. The Nobel Prize in Physics 1967. Nobelprize.org, 1967. The Nobel Prize in Physics 1967 was awarded to Hans Bethe “for his contributions to the theory of nuclear reactions, especially his discoveries concerning the energy production in stars”.

Bethe:1968:EPSb

[Bet68a] Hans A. Bethe. Energy production in stars. Physics Today, 21 (9):36–44, September 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v21/i9/p36/s1.

Bethe:1968:EPSa

[Bet68b] Hans A. (Hans Albrecht) Bethe. Energy production in stars. Nobel lecture. Norstedt & Söner, Stockholm, Sweden, December 11, 1968. 16 pp. LCCN QB464 .B46.

Bethe:1988:NPN

[Bet88] Hans A. Bethe. Nuclear physics needed for the theory of supernovae. Annual Review of Nuclear and Particle Science, 38:1–29, December 1988. CODEN ARPSDF. ISSN 0163-8998 (print), 1545-4134 (electronic). Reprinted in Autobiographies by Nobel Laureates in Physics Vol. 1 (2009): 1–29.

Bethe:1991:RAP

[Bet91] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1997:SWH

[Bet97] Hans Albrecht Bethe. Selected works of Hans A. Bethe: with commentary, volume 18 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2876-7. viii + 605 pp. LCCN QC780 .B452 1996. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0916.01027.

Bethe:2003:MLA

[Bet03] Hans A. Bethe. My life in astrophysics. Annual Review of Astronomy and Astrophysics, 41:1–14, 2003. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-4282 (electronic).

URL http://adsabs.harvard.edu/abs/2003ARA%26A..41...

.1B. Reprinted in Autobiographies by Nobel Laureates in Physics Vol. 1 (2009): 1–14.

Born:1928:PMG

[BEvL^{+}28] M. Born, A. Einstein, M. v. Laue, E. Schrödinger, and A. Sommerfeld. Planck-Medaille. (German) [Planck Medal]. Die Naturwissenschaften, 16(20):368, May 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/gg356437362203j2/fulltext.
pdf.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39c] and later translation [Gra64].

Beyerchen:1977:SUH

[Bey77] Alan D. Beyerchen. Scientists under Hitler: politics and the physics community in the Third Reich. Yale University Press, New Haven, CT, USA, 1977. ISBN 0-300-01830-4. xii + 287 + 5 pp. LCCN QC9.G3 B48.

Beyler:2006:BRN

[Bey06] Richard H. Beyler. Book review: Nancy Thorndike Greenspan: The End of the Certain World: The Life and Science of Max Born: The Nobel Physicist Who Ignited the Quantum Revolution. Isis, 97(3):569–570, September 2006. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/10.1086/509982;   http://www.
jstor.org/stable/pdfplus/10.1086/509982.pdf.

Beyler:2009:BRB

[Bey09a] Richard H. Beyler. Book review: Einstein and Oppenheimer: The Meaning of Genius. Technology and Culture, 50(3):722– 723, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270547.

Beyler:2009:BRE

[Bey09b] Richard H. Beyler. Book review: Einstein and Oppenheimer: The Meaning of Genius. Technology and Culture, 50(3):722–723, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL http://muse.jhu.edu/journals/
technology_and_culture/v050/50.3.beyler.html.

Beyler:2012:BRE

[Bey12] Richard H. Beyler. Book review: Einstein’s Generation: The Origins of the Relativity Revolution. Technology and Culture, 53(2):492–493, 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

article/476816.

Bernstein:1978:SHI

[BF78] Jeremy Bernstein and Gerald Feinberg. Science and the human imagination: Albert Einstein: papers and discussions, volume 5 of The Leverton lecture series. Fairleigh Dickinson University Press, Rutherford, NJ, USA, 1978. ISBN 0-8386-2223-2. 94 pp. LCCN QC16.E5 B46. US$4.50.

Bernstein:1986:CCP

[BF86] Jeremy Bernstein and Gerald Feinberg, editors. Cosmological Constants: Papers in Modern Cosmology. Columbia University Press, New York, NY, USA, 1986. ISBN 0-231-06376-8. ix + 328 pp. LCCN QB985 .C66 1986.

Beller:1994:BRE

[BF94] M. Beller and A. Fine. Bohr’s response to EPR. In Faye and Folse [FF94], pages 1–32. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL http://www.loc.gov/

catdir/enhancements/fy0823/93024825-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0823/93024825-t.
html;   https://link.springer.com/book/10.1007/978-
94-015-8106-6.

Bernardini:2002:CFN

[BF02] Carlo Bernardini and Enrico Honoree Fermi. Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001]. Edizioni scientifiche SIF, Bologna, Italy, second edition, 2002. ISBN 88-7438-000-3. viii + 383 pp. LCCN QC774.F4. 10.33 EUR. URL http:

//www.gbv.de/dms/casalini/02/02220733.pdf;   http://
www.sif.it/fermiindex.html; http://www.sif.it/libri/
conoscere_fermi;   http://www.zentralblatt-math.org/
zmath/en/search/?an=1073.01512. English translation in

[BB04b].

Buchwald:2014:OHH

[BF14] Jed Z. Buchwald and Robert Fox, editors. The Oxford Handbook of the History of Physics. Oxford Handbooks in Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-969625-X. 960 pp. LCCN Q124.6-127.2.

Baccalario:2016:EGS

[BF16] Pierdomenico Baccalario and Giuseppe Ferrario. Einstein: genio senza confini. (Italian) [Einstein: Genius Without Bounds]. Grandissimi. Edizioni El, San Dorligo della Valle, Italy, 2016. ISBN 88-477-3334-0. 76 pp. LCCN ????

Baker:1982:AAB

[BFB82] C. L. Baker, Richard P. Feynman, and Bernice Brode. Anecdotes: The Adventures of a Blunder 9; Reminiscences of Los Alamos. Annals of the History of Computing, 4(1):60–64, January/March 1982. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1982/pdf/a1060.
pdf;   http://www.computer.org/annals/an1982/a1060abs.
htm.

Bod:1991:OHY

[BFMNZ91] L. Bod, E. Fischbach, G. Marx, and Maria Náray-Ziegler. One hundred years of the Eötvös experiment. Acta Physica Hungarica, 69(3–4):335–355, ???? 1991. URL http://www.kfki.hu/

eotvos/onehund.html.

Bethe:2013:EPF

[BGA^{+}13] Hans Bethe, Marvin L. Goldberger, Stephen L. Adler, Steven Weinberg, Freeman Dyson, and C. N. Yang. Einstein and the physics of the future. In Yang [Yan13b], pages 21–34. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). LCCN QC21.3.

URL http://www.worldscientific.com/doi/abs/10.1142/
9789814449021_0006.

Breinschmid:2007:Y

[BGG^{+}07] Georg Breinschmid, Mario Gonzi, Herwig Gradischnig, Martin Koller, Anna Lauvergnac, Adrian Mears, Mauro Negri, Joris Roelofs, Mathias Rüegg, Andy Scherrer, Harry Sokal, Tobias Weidinger, et al. 3 : 30 years. Universal Music Austria, ????, 2007. LCCN ???? Three CDs.

Blum:2017:EIS

[BGLR17] Alexander Blum, Domenico Giulini, Roberto Lalli, and Jürgen Renn. Editorial introduction to the special issue “The Renaissance of Einstein’s Theory of Gravitation”. European Physical Journal H, 42(2):95–105, June 2017. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1140/epjh/e2017-80023-3.pdf.

Bacry:1979:EB

[BH79] Henri Bacry and Banesh Hoffmann. Einstein and de Broglie. Physics Today, 32(8):11–13, August 1, 1979. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://physicstoday.scitation.org/doi/abs/10.1063/1.
2995657.

Beck:1987:EYE

[BH87] Anna Beck and Peter Havas, editors. The Early Years: 1879– 1902. English Translation Supplement, volume 1 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08475-0 (paper), 0-691-08463-7 (microfiche). xxii + 196 pp. LCCN QC16.E5 A25 x, 1987.

Beck:1989:SYW

[BH89a] Anna Beck and Peter Havas, editors. The Swiss Years: Writings, 1900–1909. English Translation Supplement, volume 2 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08549-8. xiv + 399 pp. LCCN QC16 .E5.A2 1987 v.2.

Brink:1989:IET

[BH89b] Jean R. Brink and Roland Haden. Interviews with Edward Teller and Eugene P. Wigner. Annals of the History of Computing, 11(3):177–178, July/September 1989. CO-DEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1989/pdf/a3177.pdf;   http:/
/www.computer.org/annals/an1989/a3177abs.htm.

Beck:1993:SYW

[BH93] Anna Beck and Don Howard, editors. The Swiss Years: Writings, 1909–1911. English Translation Supplement, volume 3 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-10250-3. vii + 437 pp. LCCN QC16.E5 A2 1987. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://einsteinpapers.press.princeton.edu/vol3-doc;
 http://einsteinpapers.press.princeton.edu/vol3-trans;
 http://press.princeton.edu/titles/5259.html;   http:
//press.princeton.edu/titles/5276.html;   http://www.
loc.gov/catdir/description/prin031/87160800.html;
http://www.loc.gov/catdir/toc/prin031/87160800.html.

English translation by Anna Beck in consultation with Don Howard, With a preface by Beck and Howard.

Brommer:2001:FSM

[BH01] Peter Brommer and Günter Herrmann. Fritz Strassmann (1902– 1980), Mitentdecker der Kernspaltung: Inventar des Nachlasses und Kommentierung der Versuche zur Kernspaltung. (German) [Fritz Strassmann (1902–1980), co-discoverer of fission: Inventory of the estate and commentary the experiments on nuclear fission], volume 95 of Veröffentlichungen der Landesarchivverwaltung Rheinland-Pfalz, 0556-834X. Landesarchivverwaltung Rheinland-Pfalz, Koblenz, Germany, 2001. ISBN 3-931014-57-6 (paperback). 277 pp. LCCN QC16.S815 B76 2001.

Breit:2004:LL

[BH04] William Breit and Barry T. Hirsch. Lives of the laureates: eighteen Nobel economists. MIT Press, Cambridge, MA, USA, fourth edition, 2004. ISBN 0-262-02562-0. xiv + 351 pp. LCCN HB76 .L58 2004.

Breit:2009:LLT

[BH09] William Breit and Barry T. Hirsch, editors. Lives of the laureates: twenty-three Nobel economists. MIT Press, Cambridge, MA, USA, fifth edition, 2009. ISBN 0-262-01276-6 (hardcover). xiv + 459 pp. LCCN HB76 .L58 2009.

Bouton:2017:TNN

[BH17] Christophe Bouton and Philippe Huneman. Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences, volume 326 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346. xiii + 403 pp. LCCN QC173.59.S65 T56 2017. URL https://

link.springer.com/book/10.1007/978-3-319-53725-2.

Bhabha:1949:PBT

[Bha49] H. J. Bhabha. On the postulational basis of the theory of elementary particles. Reviews of Modern Physics, 21(3):451– 462, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.451;   http://rmp.aps.
org/abstract/RMP/v21/i3/p451_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Bhave:1986:SHV

[Bha86] S. V. Bhave. Separable hidden variables theory to explain Einstein–Podolsky–Rosen paradox. The British Journal for the Philosophy of Science, 37(4):467–475, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Bharucha:1992:BTC

[Bha92] Filita P. Bharucha. Buddhist theory of causation and Einstein’s theory of relativity, volume 111 of Bibliotheca Indo-Buddhica series. Sri Satguru Publications, Delhi, India, 1992. ISBN 81-7030-331-1. xiii + 198 pp. LCCN BQ4245 .B53 1992. Rs225.00.

Badash:1980:RA

[BHB80] Lawrence Badash, Joseph Oakland Hirschfelder, and Herbert P. Broida, editors. Reminiscences of Los Alamos, 1943–1945, volume 5 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). xxi + 188 pp. LCCN QC791.96 .R44.

Born:1962:BMG

[BHJ62] Max Born, Werner Heisenberg, and Pascual Jordan, editors. Zur Begründung der Matrizenmechanik. (German) [On the Foundations of Matrix Mechanics], volume 2 of Dokumente der Naturwissenschaft. Abteilung Physik. E. Battenberg, Stuttgart, Germany, 1962. 135 pp. LCCN QC174.3 .Z8.

Birkl:2000:CTR

[BHJ^{+}00] Gerhard Birkl, Stephan Hartmann, Manfred Jacobi, Hans Christoph Wolf, Erhard Keppler, Armin Bunde, Jens Eisert, and Max Rauner. Crichton: Timeline. Eine Reise in die Mitte der Zeit/ Dyson: The Sun, The Genome, and The Internet. Tools of Scientific Revolutions/Pais: Raffiniert ist der Herrgott Albert Einstein. Eine wissenschaftliche Biografie/Kuhn/Försterling: Principles of Physical Chemistry/Wiser: Energy Resources. Occurrence, Production, Conversion, Use/Moss de Oliveira, de Oliveira und Stauffer: Evolution, Money, War, and Computers/Canty: Konfliktlösungen mit Mathematica. Zweipersonenspiele/Adobe Acrobat 4.0/Buchtipp. Physikalische Blätter, 56 (10):69–73, October 2000. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.20000561016/abstract.

Bowes:1991:LMF

[BHM91] Stanley A. Bowes, Daniel Hirsch, and William G. Mathews. Letters: More fallout from H-bomb controversy. Bulletin of the Atomic Scientists, 47(9):45–46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90] and comment [Han92].

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/987181.

Born:1967:EEG

[BI67] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1967. ?? pp. LCCN ????

Born:1969:EEG

[BI69] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1969. 74 pp. LCCN ????

Biagioli:2016:NEG

[Bia16a] Francesca Biagioli. Non-Euclidean geometry and Einstein’s general relativity: Cassirer’s view in 1921. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 189–228. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064

(electronic). URL https://link.springer.com/chapter/10.
1007/978-3-319-31779-3_7.

Biagioli:2016:SNG

[Bia16b] Francesca Biagioli. Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2016. ISBN 3-319-31777-6 (hardcover), 3-319-31779-2 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xx + 239 pp. LCCN QA685 .B49 2016.

Bicak:1979:EPK

[Bič79a] Jiří Bičák. Einstein a Praha; K stému vjroii narození Alberta Einsteina. Jednota československých matematiků a fyziků, Prague, Czechoslovakia, 1979. 63 pp.

Bicak:1979:ECK

[Bič79b] Jiří Bičák. Einsteinova cesta k obecné teorii relativity. (Czech) [Einstein’s path to the general theory of relativity ]. Československý Časopis pro Fyziku, A29(??):222–243, ???? 1979.

Bicak:1979:TAE

[Bič79c] Jiří Bičák. With Trautman on Albert Einstein, Leopold Infeld and on physics in Poland. Ceskoslovensky Časopis pro Fyziku Sekce A, 29(3):255–259, ???? 1979. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Bicak:1989:EPA

[Bič89a] Jiří Bičák. Einstein’s Prague articles on gravitation. In Blair and Buckingham [BB89], pages 1325–1333. ISBN 9971-5-0784-6 (hardcover), 9971-5-0785-4 (paperback). LCCN QC173.6 .M37 1988.

Bicak:1989:EPI

[Bič89b] Jiří Bičák. Einstein’s Prague ideas on gravitation: The history and the present. In Janta and Pantoflíček [JP89], pages 65–75. LCCN QC1 .E857 1984.

Bicak:2000:SSE

[Bič00] Jiří Bičák. Selected solutions of Einstein’s field equations: Their role in general relativity and astrophysics. In Einstein’s field equations and their physical implications: selected essays in honour of Jürgen Ehlers [Sch00c], pages 1– 126. ISBN 3-540-67073-4, 3-540-46580-4 (e-book). ISSN 0075-8450. LCCN QC173.6 .E39 2000. URL http:

//rave.ohiolink.edu/ebooks/ebc/3540465804;   http://
www.springerlink.com/openurl.asp?genre=book&isbn=978-
3-540-67073-5;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=1616-6361&volume=540.

Biezunski:1987:ERP

[Bie87] Michel Biezunski. Einstein’s reception in Paris in 1922. In Glick [Gli87], chapter 5, pages 169–188. ISBN 90-277-2498-9. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.com/chapter/10.1007/
978-94-009-3875-5_5.

Biezunski:1991:EPT

[Bie91] Michel Biezunski. Einstein à Paris: le temps n’est plus …. (French) [Einstein in Paris: the time has gone …]. Histoires de science. Presses universitaires de Vincennes, Saint-Denis, France, 1991. ISBN 2-903981-74-4. ISSN 1140-3373. 222 pp. LCCN QC16.E5 B54 1991.

Biever:2003:NLD

[Bie03] Celeste Biever. Nobel laureates denounce a US attack on Iraq. New Scientist, ??(??):??, January 29, 2003. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL https://

www.newscientist.com/article/dn3324-nobel-laureates-
denounce-a-us-attack-on-iraq/.

Bethe:1967:TTJ

[BieWSK67] Hans A. Bethe, Henry DeWolfe [i. e. Wolf] Smyth, and George Kennan. Three tributes to J. Robert Oppenheimer. Institute for Advanced Study, Princeton, NJ, USA, 1967. 29 pp. LCCN QC16.O62 B45 1967.

Bigelow:2005:BEC

[Big05] Nicholas Bigelow. Bose–Einstein condensates: Spins mixed up. Nature Physics, 1(2):89–90, November 2005. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (elec-

tronic). URL http://www.nature.com/nphys/journal/v1/
n2/full/nphys163.html.

Buchwald:2018:BYW

[BIJ^{+}18] Diana Kormos Buchwald, József Illy, Jennifer Nollar James, Ann M. Hentschel, Mary Jane Teague, Andreas Aebi, and Klaus Hentschel, editors. The Berlin Years: Writings and Correspondence, June 1925–May 1927, volume 15 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17881-X (hardcover), 0-691-17882-8 (paperback). cxii + 1075 pp. LCCN QC16.E5 A2 1987.

Brodie:1983:NSI

[BIK83] Bernard Brodie, Michael D. Intriligator, and Roman Kolkowicz, editors. National security and international stability. Oelgeschlager, Gunn and Hain, Cambridge, MA, USA, 1983. ISBN 0-89946-172-7. vi + 441 pp. LCCN UA10.5 .N28 1983.

Bing:2006:SHA

[Bin06] Liu Bing. Some historical analysis of the translating, editing, and publishing process of the Collection of Albert Einstein in China. Synthesis Philosophica, 21(2):285–298, ???? 2006. ISSN 0352-7875.

Bird:1923:ROT

[Bir23a] J. Malcolm Bird. Relativity — and other things. Scientific American, 128(1):58, January 1923. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v128/n1/
pdf/scientificamerican0123-58.pdf.

Birkhoff:1923:RMP

[Bir23b] George David Birkhoff. Relativity and Modern Physics. Harvard University Press, Cambridge, MA, USA, 1923. xi + 283 pp. LCCN QC6 .B55. With the cooperation of Rudolph Ernest Langer.

Birkhoff:1925:ONI

[Bir25] George David Birkhoff. The Origin, Nature, and Influence of Relativity. Macmillan Publishing Company, New York, NY, USA, 1925. ix + 185 pp. LCCN QC6 .B53.

Birkhoff:1982:GDB

[Bir82] Garrett Birkhoff. George D. Birkhoff’s theory of gravitation. In Selected studies: physics-astrophysics, mathematics, history of science, pages 57–77. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86161-0. See [?].

Buchwald:2015:BYWa

[BIR^{+}15a] Diana Kormos Buchwald, József Illy, Ze’ev Rosenkranz, Tilman Sauer, and Osik Moses, editors. The Berlin Years: Writings and Correspondence, April 1923–May 1925, volume 14 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16410-X (hardcover), 0-691-16422-3 (paperback). ci + 1103 pp. LCCN QC16.E5.

URL http://press.princeton.edu/titles/10366.html;
https://einsteinpapers.press.princeton.edu/vol14-doc/

.

Buchwald:2015:BYWb

[BIR^{+}15b] Diana Kormos Buchwald, József Illy, Ze’ev Rosenkranz, Tilman Sauer, and Osik Moses, editors. The Berlin Years: Writings and Correspondence, April 1923–May 1925. English Translation Supplement, volume 14 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16410-X (hardcover), 0-691-16422-3 (paperback). ci + 1103 pp. LCCN QC16.E5. URL http:

//press.princeton.edu/titles/10367.html;   https://
einsteinpapers.press.princeton.edu/vol14-trans.

Buchwald:2012:BYWa

[BIRS12a] Diana Kormos Buchwald, József Illy, Ze’ev Rosenkranz, and Tilman Sauer, editors. The Berlin Years: Writings and Correspondence, January 1922–March 1923, volume 13 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15674-3 (paperback). xxx + 414 pp. LCCN QC16.E5.

URL http://press.princeton.edu/titles/9873.html;
https://einsteinpapers.press.princeton.edu/vol13-doc.

English translation of selected texts.

Buchwald:2012:BYWb

[BIRS12b] Diana Kormos Buchwald, József Illy, Ze’ev Rosenkranz, and Tilman Sauer, editors. The Berlin Years: Writings and Correspondence, January 1922–March 1923. English Translation Supplement, volume 13 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15674-3 (paperback). xxx + 414 pp. LCCN

QC16.E5. URL http://press.princeton.edu/titles/9874.
html;   https://einsteinpapers.press.princeton.edu/

vol13-trans. English translation of selected texts.

Bitbol:1996:SPQ

[Bit96] Michel Bitbol. Schrödinger’s philosophy of quantum mechanics, volume 188 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4266-6. xi + 285 pp. LCCN Q174 .B67 vol. 188; QC173.98. URL http:

//www.loc.gov/catdir/enhancements/fy0813/96042123-d.
html;   http://www.loc.gov/catdir/enhancements/fy0813/
96042123-t.html.

Blackmore:2001:EMV

[BIT01] J. Blackmore, R. Itagaki, and S. Tanaka, editors. Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science, volume 218 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-015-9690-9.

Bertotti:2003:TGR

[BIT03] B. Bertotti, L. Iess, and P. Tortora. A test of general relativity using radio links with the Cassini spacecraft. Nature, 425(6956): 374–376, September 25, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v425/n6956/full/nature01997.html.

Bothe:1941:ATN

[BJ41] W. Bothe and P. Jensen. Die Absorption thermischer Neutronen in Elektrographit. (German) [The absorption of thermal neutrons in electrographite]. Report G-71, ????, ????, January 20, 1941. Captured German report cited in [Wea76, page 29] that contained incorrect results on the neutron capture cross-section of carbon. Those results caused the German Uranium Project scientists to switch from graphite to the hard-to-get deuterium (in heavy water obtained from the Rjukan Plant in Norway) as a neutron moderator, and likely, substantially delayed their progress in achieving a working nuclear reactor as a precursor to an atomic bomb.

Bjerknes:2002:AEI

[Bje02] Christopher Jon Bjerknes. Albert Einstein: the incorrigible plagiarist. XTX, Downers Grove, IL, USA, 2002. ISBN 0-9719629-8-7 (paperback). 408 pp. LCCN QC16.E5 B57 2002.

Bjork:2001:INC

[Bjö01] Ragnar Björk. Inside the Nobel Committee on Medicine: Prize competition procedures 1901–1950 and the fate of Carl Neuberg. Minerva [Washington, DC], 39(4):393–408, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012767418228.

Bjork:2019:AWN

[Bjø19] R. Bjørk. The age at which Nobel Prize research is conducted. Scientometrics, 119(2):931–939, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/s11192-019-03065-4.

Bjork:2020:CRC

[Bjø20a] R. Bjørk. Correction to: Response to the comments of Turki et al. on “The journals that publish Nobel Prize research”. Scientometrics, 124(1):795, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/s11192-
020-03565-8;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03565-8.pdf. See [Bjø20a, Bjø20c,

TTB20].

Bjork:2020:JPP

[Bjø20b] R. Bjørk. The journals in physics that publish Nobel Prize research. Scientometrics, 122(2):817–823, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s11192-019-03312-8. See comments [TTB20] and response [Bjø20c, Bjø20a].

Bjork:2020:RCT

[Bjø20c] R. Bjørk. Response to the comments of Turki et al. on “The journals that publish Nobel Prize research”. Scientometrics, 124(1):791–793, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03459-9. See

[Bjø20b, TTB20, Bjø20a].

Braun:2005:EPN

[BK05] Reiner Braun and David Krieger, editors. Einstein — peace now!: visions and ideas. History of knowledge. Wiley-VCH, Weinheim, Germany, 2005. ISBN 3-527-40604-2 (hardcover). vii + 305 pp. LCCN JZ5538 .E37 2005. URL http://www.loc.

gov/catdir/enhancements/fy0740/2007435276-b.html;
http://www.loc.gov/catdir/enhancements/fy0740/2007435276-
d.html;   http://www.loc.gov/catdir/enhancements/fy0740/
2007435276-t.html.

Bird:1996:HSW

[BL96] Kai Bird and Lawrence Lifschultz, editors. Hiroshima’s shadow: writings on the denial of history and the Smithsonian controversy. Pamphleteer’s Press, Stony Creek, CT, USA, 1996. ISBN 0-9630587-3-8 (hardcover), 0-9630587-4-6 (paperback). lxxvii + 584 pp. LCCN 9609 BOOK NOT YET IN LC.

Bruccoli:2007:NPLa

[BL07a] Matthew Joseph Bruccoli and Richard Layman. Nobel Prize Laureates in Literature. Part 1. Agnon–Eucken, volume 329 of Dictionary of Literary Biography. Thomson Gale, Detroit, MI, USA, 2007. ISBN 0-7876-8147-4. xxvi + 563 + 4 pp. LCCN ????

Bruccoli:2007:NPLb

[BL07b] Matthew Joseph Bruccoli and Richard Layman. Nobel Prize laureates in literature. Part 2. Faulkner–Kipling, volume 330 of Dictionary of Literary Biography. Thomson Gale, Detroit, MI, USA, 2007. ISBN 0-7876-8148-2. xvii + 1 + 507 + 3 pp. LCCN ????

Bruccoli:2007:NPLd

[BL07c] Matthew Joseph Bruccoli and Richard Layman. Nobel Prize Laureates in Literature. Part 4. Quasimodo–Yeats, volume 332 of Dictionary of Literary Biography. Thomson Gale, Detroit, MI, USA, 2007. ISBN 0-7876-8150-4. xvii + 1 + 642 + 12 pp. LCCN ????

Bruccoli:2007:NPLc

[BL07d] Matthew Joseph Red Bruccoli and Richard. Red Layman. Nobel Prize Laureates in Literature. Part 3. Lagerkvist–Pontoppidan, volume 331 of Dictionary of Literary Biography. Thomson Gale, Detroit, MI, USA, 2007. ISBN 0-7876-8149-0. xvii + 1 + 622 + 16 pp. LCCN ????

Brown:2009:HAB

[BL09] Gerald E. Brown and Sabine Lee. Hans Albrecht Bethe: July 2, 1906–March 6, 2005. Biographical memoirs — National Academy of Sciences of the United States of America, 91:31–56, 2009. CODEN BMNSAC. ISBN 0-309-14560-0. ISSN 0077-2933.

URL http://www.nap.edu/catalog.php?record_id=12776.

Berends:2011:EWS

[BL11] Frits Berends and Franklin Lambert. Einstein’s witches’ sabbath: the first Solvay council on physics. Europhysics News, 42(5):15–17, September/October 2011. CODEN EU-PNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

URL http://www.europhysicsnews.org/articles/epn/abs/
2011/05/epn2011425p15/epn2011425p15.html.

Bicak:2014:GRC

[BL14] Jiří Bičák and Tomáš Ledvinka, editors. General Relativity, Cosmology and Astrophysics: Perspectives 100 years after Einstein’s stay in Prague, volume 177 of Fundamental Theories of Physics. Springer International Publishing, Cham, Switzerland, 2014. ISBN 3-319-06349-9. xv + 535 + 117 + 80 pp. LCCN ????

Blair:1913:DUV

[Bla13] Andrew A. Blair. The determination of uranium and vanadium in the carnotite ores of Colorado and Utah. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 52(209):201–205, April 1913. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983866.

Blakeslee:1946:AF

[Bla46] Howard W. Blakeslee. The atomic future. Chicago Sun, ??(??): ??, May 3, 1946. 28-page supplement.

Blackett:1948:FWB

[Bla48] Baron P. M. S. (Patrick Maynard Stuart) Blackett. Fear, war, and the bomb: military and political consequences of atomic energy. McGraw-Hill, New York, NY, USA, 1948. 244 pp. LCCN ????

Blackett:1949:FWB

[Bla49a] Baron P. M. S. (Patrick Maynard Stuart) Blackett. Fear, war, and the bomb: military and political consequences of atomic energy. Whittlesey House, New York, NY, USA, 1949. vii + 244 pp. LCCN UF767 .B58 1949.

Blackett:1949:MPC

[Bla49b] Baron P. M. S. (Patrick Maynard Stuart) Blackett. Military and political consequences of atomic energy. Turnstile Press, London, UK, 1949. viii + 222 pp. LCCN UF767 .B58 1949a.

Black:1950:RAN

[Bla50] Max Black. Review: Autobiographical Notes, by Albert Einstein and Paul Arthur Schilpp. Journal of Symbolic Logic, 15 (2):157, June 1950. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL 2-I; http://links.jstor.

org/sici?sici=0022-4812(195006)15:2<157:AN>2.0.CO.

Blackmore:1976:BRB

[Bla76a] John T. Blackmore. Book review: The Einstein decade: By Cornelius Lanczos. New York and London (Academic Press), 1974. $12.50. Historia Mathematica, 3(2):236, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900598.

Blackmore:1976:BRE

[Bla76b] John T. Blackmore. Book review: The Einstein decade: By Cornelius Lanczos. New York and London (Academic Press), 1974. $12.50. Historia Mathematica, 3(2):236, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900598.

BlancoCendon:1986:TAP

[Bla86] Fernando Blanco Cendon. En torno al principio de indeterminación de Werner Karl Heisenberg. (Spanish) [On the Uncertainty Principle of Werner Karl Heisenberg], volume 9 of Publicaciones de los Institutos Pontificios de Filosofía y Teología. Serie III, Cuadernos de filosofía. Universidad pontificia de Santo Tomás de Manila, Instituto de filosofía de Madrid, Madrid, Spain, 1986. ISBN 84-600-4226-X. 133 pp. LCCN QC174.17.H4 B56 1986.

Blackmore:1988:MCP

[Bla88] John Blackmore. Mach competes with Planck for Einstein’s favor. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):45–89, 1988. CODEN HISCDU. ISSN 0285-4821. URL https://ci.

nii.ac.jp/naid/110006435088;   https://ci.nii.ac.jp/
naid/110009837382.

Blackmore:1992:EMD

[Bla92] John Blackmore. Ernst Mach — A Deeper Look: Documents and New Perspectives, volume 143 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. xix + 463 pp. LCCN C5.53; QC6.9. URL https://link.springer.

com/book/10.1007/978-94-011-2771-4.

Blay:1993:RID

[Bla93] Michel Blay. Les raisons de l’infini: du monde clos à l’univers mathématique. (French) [The reasons for infinity: from a closed world to a mathematical universe]. nrf essais. Gallimard, Paris, France, 1993. ISBN 2-07-072992-3. 258 pp. LCCN BD411 .B55

1993. URL http://catalog.hathitrust.org/api/volumes/
oclc/29673358.html.

Black:2002:IHS

[Bla02] Edwin Black. IBM and the Holocaust: the strategic alliance between Nazi Germany and America’s most powerful corporation. Three Rivers Press, New York, NY, USA, 2002. ISBN 0-609-80899-0 (paperback). 551 pp. LCCN HD9696.2.U64 I253 2002.

Blanchard:2010:EFE

[Bla10] Alain Blanchard. Evidence for the fifth element: Astrophysical status of dark energy. Astronomy and Astrophysics Review, 18(4):595–645, ???? 2010. CODEN AASREB. ISSN 0935-4956 (print), 1432-0754 (electronic). URL http://www.

springerlink.com/content/q372r215448x6x33/.

Butov:2002:TBE

[BLI^{+}02] L. V. Butov, C. W. Lai, A. L. Ivanov, A. C. Gossard, and D. S. Chemla. Towards Bose–Einstein condensation of excitons in potential traps. Nature, 417(6884):47–52, May 2, 2002. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v417/n6884/
full/417047a.html.

Brown:2019:SUS

[BLM19] J. Patrick Brown, Beryl Lipton, and Michael Morisy. Scientists under Surveillance: the FBI Files. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-53688-9 (paperback), 0-262-35302-4 (ebook). xviii + 413 pp. LCCN Q125 .S4366 2019. URL http:/

/mitpress.mit.edu/9780262536882. Foreword by Steven Af-

tergood and introduction by Walter V. Robinson.

Blum:2015:RGR

[BLR15] Alexander Blum, Roberto Lalli, and Jürgen Renn. The reinvention of general relativity: A historiographical framework for assessing one hundred years of curved space–time. Isis, 106 (3):598–620, 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/683425.

Blumenfeld:1989:WRE

[Blu89] Olga O. Blumenfeld. Women’s rights at ‘Einstein’. Nature, 340 (6231):259, July 27, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v340/n6231/pdf/340259a0.pdf.

Bluhm:2004:BLS

[Blu04] Robert Bluhm. Breaking Lorentz symmetry. Physics World, 17(3):41–46, March 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/17/3/phwv17i3a33.
pdf.

Bluhm:2005:VSC

[Blu05] Robert Bluhm. A very special centenary. Physics World, 18(1): 43, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/1/phwv18i1a31.pdf.

Blum:2022:ESB

[Blu22] Alexander S. Blum. Einstein’s second-biggest blunder: the mistake in the 1936 gravitational-wave manuscript of Albert Einstein and Nathan Rosen. Archive for History of Exact Sciences, 76(6):623–632, November 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00295-6.

Brown:1977:SHA

[BM77] Anthony Cave Brown and Charles Brown MacDonald, editors. The Secret History of the Atomic Bomb. Dial Press, New York, NY, USA, 1977. ISBN 0-8037-8164-4. xxi + 582 + 4 pp. LCCN QC773.3.U5 S44. US$15.00.

Ben-Menahem:1993:SCE

[BM93] Yemima Ben-Menahem. Struggling with causality: Einstein’s case. Science in Context, 6(1):291–310, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-Merre:2009:WPC

[BM09] David Ben-Merre. “What points of contact existed between these languages?”: James Joyce, Albert Einstein, and interdisciplinary study. James Joyce Quarterly, 47(1):25–49, Fall 2009. CODEN ???? ISSN 0021-4183 (print), 1938-6036 (electronic).

URL https://muse.jhu.edu/article/411793.

Barnett:2011:HGZa

[BMD^{+}11] Erin Barnett, Phil Mariani, John W. Dower, Adam Harrison Levy, and David Monteyne, editors. Hiroshima: Ground Zero 1945. International Center of Photography, New York, NY, USA, 2011. ISBN 3-86930-334-4. 247 pp. LCCN TR820.5 .B365 2011.

Bernard:2012:SCW

[BMS12] Schiele Bernard, Claessens Michel, and Shi Shunke, editors. Science communication in the world: practices and theories and trends. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4278-9, 94-007-4279-7 (e-book). LCCN 18 f.

Boorse:1989:ASB

[BMW89] Henry Abraham Boorse, Lloyd Motz, and Jefferson Hane Weaver. The Atomic Scientists: a Biographical History. Wiley, New York, NY, USA, 1989. ISBN 0-471-50455-6. vii + 472 pp. LCCN QC773 1989.

Bergia:1988:RCE

[BN88] S. Bergia and L. Navarro. Recurrences and continuity in Einstein’s research on radiation between 1905 and 1916. Archive for History of Exact Sciences, 38(1):79–99, March 1988. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=38&issue=1&spage=79.

Bergia:1997:EQC

[BN97] Silvio Bergia and Luis Navarro. Early quantum concepts and the theorem of equipartition of energy in Einstein’s work (1905– 1925). Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 34(1–2):183–223, 1997. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Beardsley:1998:NPS

[BNSY98] Tim Beardsley, Sasha Nemecek, Gary Stix, and Philip Yam. The 1997 Nobel Prizes for Science: a look at the contributions and controversies of the winning work. Scientific American, 278(1): 14–16, 18 (Intl. ed. 8–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0198issue/0198currentissue.html.

Bartholomew:1981:SBD

[BNY81] Michael Bartholomew, Bernard Norton, and Robert M. Young. Science and belief: from Darwin to Einstein. Block 6: Unit 12– 14, Problems in the biological and human sciences. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11005-3. 110 pp. LCCN QH361 .B36 1981.

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37 B58 1976. US$12.95. URL http://alsos.wlu.edu/

information.aspx?id=367.

Boden:1994:DC

[Bod94] Margaret A. Boden, editor. Dimensions of creativity. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-02368-7. 242 pp. LCCN BF408 .D56 1994.

Bodanis:2000:BWM

[Bod00] David Bodanis. E = mc^{2}: a biography of the world’s most famous equation. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2000. ISBN 0-8027-1352-1 (hardcover). ix + 337 pp. LCCN QC73.8.C6 B63 2000.

Bodanis:2003:BEK

[Bod03] David Bodanis. Bis Einstein kam: die abenteuerliche Suche nach dem Geheimnis der Welt. (German) [Einstein came: the adventurous search for the mystery of the world], volume 15399 of Fischer. Fischer-Taschenbuch-Verl., Frankfurt am Main, Germany, 2003. ISBN 3-596-15399-9. 351 pp. LCCN ????

Bodanis:2005:BWMa

[Bod05a] David Bodanis. E = mc^{2}: a Biography of the World’s Most Famous Equation. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2005. ISBN 0-8027-1352-1 (hardcover), 0-8027-1463-3. xiii + 337 pp. LCCN QC73.8 .C6 B63 2000; QC73.8.C6 B63 2000. With a new introduction by Simon Singh.

Bodanis:2005:BWMb

[Bod05b] David Bodanis. E = mc^{2}: a Biography of the World’s Most Famous Equation. Berkley Books, New York, NY, USA, 2005. ISBN 0-425-18164-2. ix + 337 pp. LCCN QC73.8.C6 B63 2001.

Bodeen:2013:ETX

[Bod13] Christopher Bodeen. Einstein translator Xu Liangying dies. Scientific Computing [online], ??(??):??, February 5, 2013. URL

http://www.scientificcomputing.com/news-DS-Einstein-
Translator-Xu-Liangying-Dies-020413.aspx. Obituary for

92-year-old Xu, who began translating Einstein’s collected works to Mandarin Chinese in 1962.

Bodanis:2016:EGM

[Bod16] David Bodanis. Einstein’s greatest mistake: a biography. Houghton Mifflin Harcourt, Boston, MA, USA, 2016. ISBN 0-544-80856-8 (hardcover), 0-544-80858-4 (e-book). xiv + 280 pp. LCCN QC16.E5 B66 2016.

Bodanis:2017:EID

[Bod17] David Bodanis. Einsteins Irrtum Das Drama eines Jahrhundertgenies. (German) [Einstein’s biggest mistake: The drama of a century’s genius]. DVA, München, Germany, 2017. ISBN 3-421-04754-5. 350 pp. LCCN ????

Boeyens:2010:CC

[Boe10a] J. C. A. (Jan C. A.) Boeyens. Chemical Cosmology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3828-0. xvii + 419 pp. LCCN QD453.3 .B64 2010.

Boeyens:2010:SC

[Boe10b] Jan C. A. Boeyens. Standard cosmology. In Chemical Cosmology [Boe10a], chapter 6, pages 183–226. ISBN 90-481-3828-0. LCCN QD453.3 .B64 2010.

Bohr:1923:BAG

[Boh23] Niels Bohr. Über den Bau der Atome. (German) [On the structure of atoms]. Die Naturwissenschaften, 11(27):606–624, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Lecture held in Stockholm, 11 December 1922, on the occasion of the receipt of the Nobel Prize for physics for the year 1922. Translated to German by W. Pauli, Jr.

Bohr:1946:FSC

[Boh46a] Niels Bohr. Foreword: Science and civilization. In Masters and Way [MW46a], page ?? LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Bohr:1946:F

[Boh46b] Niels Bohr. Forord. In Masters and Way [MW46b], page ?? LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Bohr:1949:DEE

[Boh49] Niels Bohr. Discussion with Einstein on epistemological problems in atomic physics. In Schilpp [Sch49b], chapter 7, pages 199–241. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. URL http://www.

marxists.org/reference/subject/philosophy/works/dk/

bohr.htm. Reprinted 1951, 1969, and 1982.

Bohm:1957:CCMa

[Boh57a] David Bohm. Causality and chance in modern physics. Routledge and Paul, London, UK, 1957. 170 pp. LCCN QC6 .B597 1957.

Bohm:1957:CCMb

[Boh57b] David Bohm. Causality and chance in modern physics. Van Nostrand, Princeton, NJ, USA, 1957. 170 pp. LCCN QC6 .B597 1957a.

Bohr:1958:APHb

[Boh58] Niels Bohr. Atomic Physics and Human Knowledge. Wiley, New York, NY, USA, 1958. viii + 101 pp. LCCN QC6 .B598. URL

http://www.questia.com/PM.qst?a=o&d=21083222.

Bohm:1961:CCM

[Boh61] David Bohm. Causality and chance in modern physics. Harper torchbooks. The science library. Harper, New York, NY, USA, 1961. 170 pp. LCCN QC6.4.C3 B63x 1961.

Bohm:1962:CNC

[Boh62] David Bohm. Classical and non-classical concepts in the quantum theory: an answer to Heisenberg’s Physics and Philosophy. The British Journal for the Philosophy of Science, XII(48):265– 280, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/XII/48/265.full.pdf+html.

Bohm:1965:STR

[Boh65a] David Bohm. The Special Theory of Relativity. W. A. Benjamin, New York, NY, USA, 1965. xiv + 256 pp. LCCN QC6 .B59719.

Bohr:1965:NLS

[Boh65b] Niels Bohr. Nobel lecture: The structure of the atom. In Anonymous, editor, Nobel lectures: including presentation speeches and laureates’ biographies: physics 1922–1941, pages 7–43. Elsevier, Amsterdam, The Netherlands, 1965. LCCN QC71

.P455 1965. URL http://nobelprize.org/nobel_prizes/
physics/laureates/1922/bohr-lecture.html;   http://
nobelprize.org/nobel_prizes/physics/laureates/1922/

bohr-lecture.pdf. Foreword by Arne Tiselius.

Bohm:1971:CCM

[Boh71] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, 1971. 170 pp. LCCN QC6 .B6755 1971.

Bohm:1979:STR

[Boh79] David Bohm. The Special Theory of Relativity. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1979. ISBN 0-8053-1001-0. xiv + 236 pp. LCCN QC6 .B59719.

Bohm:1984:CCMa

[Boh84a] David Bohm. Causality and chance in modern physics. Routledge and Kegan Paul, London, UK, 1984. ISBN 0-7102-0031-5 (paperback). xv + 170 pp. LCCN QC6.4.C3 B63x 1984.

Bohm:1984:CCMb

[Boh84b] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, third edition, 1984. ISBN 0-8122-1002-6. xi + 170 pp. LCCN QC6 .B6755 1984.

Bohr:1985:BED

[Boh85a] Niels Bohr. The Bohr–Einstein dialogue. In French and Kennedy [FK85], pages 121–140. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:NPL

[Boh85b] Niels Bohr. Nobel Prize lecture: the structure of the atom. In French and Kennedy [FK85], pages 91–97. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohm:1989:QT

[Boh89a] David Bohm. Quantum theory. Dover, New York, NY, USA, 1989. ISBN 0-486-65969-0. ix + 646 pp. LCCN QC174.12 .B632 1989. US$10.95. URL http://www.loc.gov/

catdir/description/dover031/89031187.html;   http://
www.loc.gov/catdir/toc/dover031/89031187.html.

Bohm:1989:STR

[Boh89b] David Bohm. The Special Theory of Relativity. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-09448-7. xiv + 236 pp. LCCN QC173.65 .B64 1989.

Bohm:1996:STR

[Boh96] David Bohm. The Special Theory of Relativity. Routledge, London, UK, 1996. ISBN 0-415-14808-1, 0-415-14809-X (paperback). xiv + 236 pp. LCCN QC6 .B59719.

Bohm:1999:CCM

[Boh99] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, 1999. ISBN 0-585-12650-X (e-book), 0-8122-1002-6. xi + 170 pp. LCCN QC6.4.C3 B64 1999eb.

Bohannon:2005:DE

[Boh05] John Bohannon. Dancing Einstein. Science, 309(5731):59, July 1, 2005. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/309/5731/59.1.full.pdf.

Bohm:2006:STR

[Boh06] David Bohm. The Special Theory of Relativity. Routledge classics. Routledge, London, UK, 2006. ISBN 0-415-40425-8. xxi + 282 pp. LCCN QC173.65 .B64 2006.

URL http://www.loc.gov/catdir/enhancements/fy0668/
2006017434-d.html;   http://www.loc.gov/catdir/toc/

ecip0614/2006017434.html. With a foreword by Basil Hiley and a new preface by John Barrow.

Bohr:2010:APH

[Boh10] Niels Bohr. Atomic Physics and Human Knowledge. Dover books on physics. Dover, New York, NY, USA, 2010. ISBN 0-486-47928-5. viii + 101 pp. LCCN QC6 .B598

2010. URL http://catdir.loc.gov/catdir/enhancements/
fy1010/2010022935-d.html.

Bokulich:2008:PDE

[Bok08] Alisa Bokulich. Paul Dirac and the Einstein–Bohr debate. Perspectives on Science, 16(1):103–114, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://muse.jhu.edu/journals/perspectives_on_science/
v016/16.1bokulich.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.1bokulich.pdf;   http:
//www.mitpressjournals.org/doi/abs/10.1162/posc.2008.
16.1.103.

Bokulich:2015:BRD

[Bok15] Alisa Bokulich. Book review: A. Douglas Stone, Einstein and the Quantum: The Quest of the Valiant Swabian. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):177–179, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/678188.

Bolles:2004:EDB

[Bol04] Edmund Blair Bolles. Einstein defiant (Bohr unyielding): genius versus genius in the quantum revolution. Joseph Henry Press, Washington, DC, USA, 2004. ISBN 0-309-08998-0 (hardcover), 0-309-09617-0 (paperback). viii + 348 pp.

LCCN QC173.98 .B65 2004. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip0410/2003023735.html.

Bondi:1964:RCS

[Bon64] Hermann Bondi. Relativity and common sense: a new approach to Einstein. Dover, New York, NY, USA, 1964. ISBN 0-486-24021-5. xi + 177 pp. LCCN ????

Bonnor:1979:EC

[Bon79] W. B. Bonnor. Einstein’s charisma. Nature, 278(5707): 823, April 26, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v278/n5707/pdf/278823a0.pdf.

Bonnor:1981:ERJ

[Bon81] W. B. Bonnor. Einstein refined — just in time. Nature, 291 (5810):96, May 7, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v291/n5810/pdf/291096a0.pdf.

Bondi:1994:ERE

[Bon94] H. Bondi. Essay reviews: Einstein illuminated [Roger Highfield and Paul Carter, The Private Lives of Albert Einstein. Faber & Faber, 1993. £15.99. ISBN 0-571-16744-6]. Notes and Records of the Royal Society of London, 48(2):313–316, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532171.

Bahcall:1999:CTR

[BOPS99] Neta A. Bahcall, Jeremiah P. Ostriker, Saul Perlmutter, and Paul J. Steinhardt. The cosmic triangle: Revealing the state of the universe. Science, 284(5419):1481–1488, May 28, 1999.

URL http://www.sciencemag.org/content/284/5419/1481.

Born:1920:REI

[Bor20] Max Born. Die Relativitätstheorie Einsteins und ihre physikalischen grundlagen gemeinverständlich dargestellt von Max Born. (German) [Einstein’s Theory of Relativity and its physical foundations]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1920. x + 242 pp. LCCN QC6 .B65.

Born:1922:REI

[Bor22a] Max Born. Die Relativitätstheorie Einsteins und ihre physikalischen grundlagen. Elementar dargestellt. (German) [Einstein’s Theory of Relativity and its physical basis. Elementary treatment]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1922. xi + 267 pp. LCCN QC6 .B65.

Born:1922:TRE

[Bor22b] Max Born. La teoría de la relatividad de Einstein y sus fundamentos fisicos: Exposición elemental. (Spanish) [Einstein’s Theory of Relativity and its fundamental physics: elementary introduction]. Calpe, Madrid, Spain, 1922. 384 pp. LCCN ????

Born:1949:EST

[Bor49a] Max Born. Einstein’s statistical theories. In Schilpp [Sch49b], chapter 5, pages 161–177. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Born:1949:RTE

[Bor49b] Max Born. Reciprocity theory of elementary particles. Reviews of Modern Physics, 21(3):463–473, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.463;   http://rmp.aps.org/abstract/RMP/v21/i3/p463_

1. Special issue in honor of Albert Einstein on his 70th birthday.

Born:1950:ERR

[Bor50] Max Born. Einstein and Relativity: Review of Einstein: The Meaning of Relativity, 4th edition. Nature, 166(4227): 751, November 4, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v166/n4227/pdf/166751a0.pdf.

Born:1953:IQM

[Bor53] Max Born. The interpretation of quantum mechanics. The British Journal for the Philosophy of Science, IV(14):95–106, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/IV/14/95.full.pdf+html.

Born:1955:AEL

[Bor55a] Max Born. Albert Einstein und das Lichtquantum. (German) [Albert Einstein and the quantum of light]. Die Naturwissenschaften, 42(15):425–431, August 1, 1955. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1955:SDQ

[Bor55b] Max Born. Die statistische Deutung der Quantenmechanik. (German) [The statistical interpretation of quantum mechanics]. In Les Prix Nobel en 1954 Stockholm, page ?? ????, ????, 1955.

Born:1956:EAE

[Bor56] Max Born. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. Mathematische und Naturwissenschaftliche Unterricht, 9(3):97–105, ???? 1956. CODEN MNWUAL. ISSN 0025-5866.

Born:1961:MAL

[Bor61] Max Born. More about Lorentz transformation equations. The British Journal for the Philosophy of Science, XII(46):150–151, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/XII/46/150.full.pdf+html.

Born:1962:ETR

[Bor62] Max Born. Einstein’s Theory of Relativity. Dover, New York, NY, USA, 1962. vii + 376 pp. LCCN QC6 .B66 1962. With the collaboration of Günther Leibfried and Walter Biem.

Born:1963:AAGb

[Bor63a] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:SDQ

[Bor63b] Max Born. Die statistische Deutung der Quantenmechanik. (German) [The statistical interpretation of quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 68, pages 430–441. LCCN ???? Reprinted from Nobelvortrag, gehalten am 11. Dezember 1954, Les Prix Nobel en 1954, Stockholm 1955, pp. 79–90.

Born:1963:EAE

[Bor63c] Max Born. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 90, pages 660–668. LCCN ???? Reprinted from Der Mathematische und Naturwissenschaftliche Unterricht IX 97–105 (1956).

Born:1964:REG

[Bor64] Max Born. Die Relativitaätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1964. xi + 328 pp. LCCN QC6 .B65 1964. With the collaboration of Walter Biem.

Born:1965:ETR

[Bor65a] Max Born. Einstein’s theory of relativity. Dover, New York, NY, USA, revised edition, 1965. ISBN 0-486-60769-0. vii + 376 pp. LCCN QC6 .B653 1965. URL http://www.loc.gov/

catdir/description/dover032/65001214.html;   http:/
/www.loc.gov/catdir/enhancements/fy1318/65001214-t.

html. Prepared with the collaboration of Günther Leibfried and Walter Biem.

Born:1965:EEG

[Bor65b] Max Born. Erinnerungen an Einstein. (German) [Memories of Einstein]. Physikalische Blätter, 21(7):297–306, July 1965. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19650210701/abstract.

Bork:1966:PJB

[Bor66] Alfred M. Bork. Physics just before Einstein. Science, 152 (3722):597–603, April 29, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/152/3722/597.full.pdf.

Born:1969:REG

[Bor69a] Max Born. Die Relativitaätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 1969. xii + 328 pp. LCCN QC6 .B65 1969. With the collaboration of Walter Biem.

Born:1969:PMG

[Bor69b] Max Born. Physics in my generation. Heidelberg science library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1969. vii + 172 pp. LCCN QC71 .B66 1969.

Borlaug:1971:GRB

[Bor71] Norman E. Borlaug. The green revolution: For bread and peace. Bulletin of the Atomic Scientists, 27(6):6–9, 42–48, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint of the author’s 1971 Nobel Peace Prize address.

Born:1975:MLE

[Bor75] Max Born. Mein Leben. Die Erinnerungen des Nobelpreisträgers. (German) [My Life. Recollections of a Nobel Laureate]. Nymphenburger Verlagshandlung GmbH, München, West Germany, 1975. 400 pp. Translation to German by Helmut Degner with the collaboration of Ilse Krewinkel qnd Werner Rau.

Borkin:1978:CPG

[Bor78a] Joseph Borkin. The crime and punishment of I. G. Farben. Free Press, New York, NY, USA, 1978. ISBN 0-02-904630-0. 250 + 4 pp. LCCN HD2769.C53 G293.

Born:1978:MLR

[Bor78b] Max Born. My life: recollections of a Nobel laureate. Scribner, New York, NY, USA, 1978. ISBN 0-684-15662-8. xi + 308 pp. LCCN QC16.B643 A3213 1978. US$17.50.

Borella:2002:EEP

[Bor02] Vincent Borella. Les écrits épistémologiques de Poincaré, obstacles à la diffusion de la relativité? Revue d’Histoire des Sciences, 55(1):45–81, January 2002. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633659.

Borgir:2004:GDO

[Bor04] Arne Borgir. Genocide: debate in Oslo: the capital of the Nobel Peace Prize and the Oslo Process of Palestine. Connector Press, Oslo, Norway, 2004. ISBN 82-92555-01-3. 148 pp. LCCN ????

Born:2007:MEL

[Bor07] Max Born. The momentum–energy law in the electrodynamics of Gustav Mie. In Renn et al. [RSS^{+}07b], pages 745–756. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_39. Originally published as “Der Impuls–Energie-Satz in der Elektrodynamik von Gustav Mie”, Nachrichten von der Königlichen Gesellschaft der Wissenschaften zu Göttingen 1, 23–36 (1914).

Bose:1924:PGL

[Bos24a] Satyendra Nath Bose. Plancks Gesetz und Lichtquantenhypothese. (German) [Planck’s law and the light-quantum hypothesis]. Zeitschrift für Physik, 26(1):178–181, ???? 1924. CODEN

ZEPYAA. ISSN 0044-3328. URL http://link.springer.com/
article/10.1007/BF01327326.

Bose:1924:WIS

[Bos24b] Satyendra Nath Bose. Wärmegleichgewicht im Strahlungsfeld bei Anwesenheit von Materie. (German) [Thermal equilibrium in the radiation field in the presence of matter]. Zeitschrift für Physik, 27(1):384–393, December 1, 1924. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/
10.1007/BF01328037;   http://www.encyclopedia.com/topic/
S.N._Bose.aspx.

Bagrov:1996:IED

[BOS96] V. G. Bagrov, V. V. Obukhov, and A. G. Sakhapov. Integration of the Einstein–Dirac equations. Journal of Mathematical Physics, 37(11):5599–5610, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjork:2014:TSC

[BOS14] Samuel Bjork, Avner Offer, and Gabriel Söderberg. Time series citation data: the Nobel Prize in Economics. Scientometrics, 98 (1):185–196, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-013-0989-5.

Bottoms:2013:BRE

[Bot13] Stephen Bottoms. Book review: Einstein on the Beach, by Robert Wilson and Philip Glass. Theatre Journal, 65(1):99–101, March 2013. CODEN ???? ISSN 0192-2882 (print), 1086-332X (electronic).

Bouffard:1999:RME

[Bou99] Karen Bouffard. Rutherford meets Einstein. The Physics Teacher, 37(2):125, 1999. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Boughn:2013:FHM

[Bou13] Stephen Boughn. Fritz Hasenöhrl and E = mc^{2}. European Physical Journal H, 38(2):261–278, March 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-30061-5.

Bower:1987:PCB

[Bow87a] Tom Bower. The Paperclip Conspiracy: The Battle for the Spoils and Secrets of Nazi Germany. M. Joseph, London, UK, 1987. ISBN 0-7181-2744-7. xiv + 336 + 12 pp. LCCN D810.S2 B69x 1987b.

Bower:1987:PCH

[Bow87b] Tom Bower. The Paperclip Conspiracy: The Hunt for the Nazi Scientists. Little, Brown and Co., Boston, MA, USA, 1987. ISBN 0-316-10399-3. x + 309 pp. LCCN D810.S2 B69 1987. US$17.95.

Boyer:1985:BEL

[Boy85] Paul S. Boyer. By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age. Pantheon Books, New York, NY, USA, 1985. ISBN 0-394-52878-6, 0-394-74767-4. xx + 440 pp. LCCN E169.12 .B684 1985. US$22.50.

Boyd:1988:CRB

[Boy88] Carolyn P. Boyd. Champions of Relativity: Book review: Einstein in Spain: Relativity and the Recovery of Science, Thomas F. Glick, Princeton University Press, Princeton, NJ, 1988, xiv + 391 pp., illus. $42. Science, 240(4856):1209–1212, May 27, 1988. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/240/4856/1209.full.pdf. See [Gli88].

Boyer:1994:BEL

[Boy94] Paul S. Boyer. By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age. University of North Carolina Press, Chapel Hill, NC, USA, 1994. ISBN 0-8078-4480-2 (paperback). xxii + 440 pp. LCCN E169.12 .B684

1994. URL http://www.loc.gov/catdir/enhancements/
fy0828/94004241-b.html;   http://www.loc.gov/catdir/
enhancements/fy0828/94004241-d.html.

Boyle:2010:NLC

[Boy10] Willard S. Boyle. Nobel Lecture: CCD — an extension of man’s view. Reviews of Modern Physics, 82(3):2305– 2306, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2305;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2305_1.

Bozzi:2014:ERP

[Boz14] Paola Bozzi. Einstein’s rocky picture show. Einstein überquert die Elbe bei Hamburg di Siegfried Lenz. (German) [Einstein crosses the Elbe near Hamburg, by Siegfried Lenz]. In Castellari [Cas14], pages 255–266. ISBN 88-6705-207-1. LCCN ????

Bernstein:1979:CEE

[BP79] Jeremy Bernstein and Franklin Pollock. The calculation of the electrostatic energy in the liquid drop model of nuclear fission — a pedagogical note. Physica A, 96(1–2):136–140, April 1979. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378437179902012.

Barbour:1995:MPN

[BP95] Julian B. Barbour and Herbert Pfister, editors. Mach’s principle: from Newton’s bucket to quantum gravity, volume 6 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3823-7, 3-7643-3823-7 (paperback). vii + 536 pp. LCCN QC137 .M33 1995.

URL http://www.gbv.de/dms/bowker/toc/9780817638238.
pdf;   http://www.zentralblattmath.org/zmath/en/search/
?an=0846.01008.

Bukharin:2001:RSN

[BP^{+}01] Oleg Bukharin, P. L. (Pavel Leonardovich) Podvig, et al., editors. Russian strategic nuclear forces. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-16202-4 (hardcover). xxi + 692 pp. LCCN UA776.R37.R87; UA776.R37.R87 2001. An Englishlanguage book by The Center for Arms Control, Energy and Environmental Studies at the Moscow Institute of Physics and Technology.

Bona:2005:ENR

[BPL05] Carles Bona and Carlos Palenzuela-Luque. Elements of numerical relativity: from Einstein’s equations to Black Hole simulations, volume 673 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-25779-9 (hardcover). ISSN 0075-8450 (print), 1616-6361 (electronic). xi + 151 pp. LCCN QC6 .B615 2005.

Brown:1995:TCP

[BPP95] Laurie M. Brown, Abraham Pais, and Sir A. Brian Pippard, editors. Twentieth century physics. Institute of Physics Publishers, Bristol, UK, 1995. ISBN 1-56396-314-0 (US: set), 1-56396-047-8 (US: vol. 1), 1-56396-048-6 (US: vol. 2), 1-56396-049-4 (US: vol. 3), 0-7503-0310-7 (UK: set), 0-7503-0353-0 (UK: vol. 1), 0-7503-0354-9 (UK: vol. 2), 0-7503-0355-7 (UK: vol. 3). xviii + 2059 + 175 (3 volumes) pp. LCCN QC7 .T84 1995. URL http://www.

loc.gov/catdir/enhancements/fy0668/95041186-d.html.

Baaz:2011:KGF

[BPP^{+}11] Matthias Baaz, Christos H. Papadimitriou, Hilary W. Putnam, Dana S. Scott, and Charles L. Harper, Jr., editors. Kurt Gödel and the Foundations of Mathematics: Horizons of Truth. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-76144-1 (print), 0-511-97423-X (electronic). xxiv + 515 pp. LCCN QA9.65 .K87 2011.

Barrow:1987:MCU

[BPS87] John D. Barrow, P. J. E. (Phillip James Edwin) Peebles, and D. W. (Dennis William) Sciama. Material content of the universe: proceedings of a Royal Society Discussion Meeting held on 23 and 24 October 1985. Royal Society, London, UK, 1987. ISBN 0-85403-303-3. v + 179 pp. LCCN QB790 .R68 1985.

Bohr:1955:AE

[BR55] Niels Bohr and I. I. Rabi. Albert Einstein: 1879–1955. Scientific American, 192(6):31–32, June 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v192/n6/
pdf/scientificamerican0655-31.pdf.

Binnig:1985:STM

[BR85] Gerd Binnig and Heinrich Rohrer. The scanning tunneling microscope. Scientific American, 253(2):50–56, August 1985. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v253/n2/pdf/scientificamerican0885-50.pdf;
http://www.nobelprize.org/nobel_prizes/physics/laureates/
1986/.

Brown:1993:MGS

[BR93] Laurie M. Brown and John S. Rigden, editors. Most of the good stuff: memories of Richard Feynman. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 0-88318-870-8. 181 + 16 pp. LCCN QC16.F49 A3 1993.

Butts:1999:PSR

[BR99] D. A. Butts and D. S. Rokhsar. Predicted signatures of rotating Bose–Einstein condensates. Nature, 397(6717):327– 329, January 28, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v397/n6717/full/397327a0.html.

Beer:2000:OWC

[BR00] Günther Beer and Horst Remane. Otto Wallach 1847–1931: Chemiker und Nobelpreisträger: Lebenserinnerungen: Potsdam, Berlin, Bonn, Göttingen. (German) [Otto Wallach 1847–1931: Chemist and Nobel Laureate: Memoirs: Potsdam, Berlin, Bonn, Göttingen], volume 12 of Studien und Quellen zur Geschichte der Chemie. Verlag für Wissenschafts- und Regionalgeschichte Dr. Michael Engel, Berlin, Germany, 2000. ISBN 3-929134-34-9 (paperback). ISSN 0941-8059. 270 pp. LCCN QD22.W35 O88 2000.

Baldwin:2005:DIE

[BR05] Mark Baldwin and Ray Rivers. Dancing to the ideas of Einstein. Physics World, 18(5):16–17, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/5/phwv18i5a25.pdf.

Boughn:2011:HEM

[BR11] Stephen Boughn and Tony Rothman. Hasenöhrl and the equivalence of mass and energy. arxiv.org, August 10, 2011. URL

http://arxiv.org/abs/1108.2250.

Begelman:2021:GFA

[BR21] Mitchell C. Begelman and Martin J. Rees. Gravity’s Fatal Attraction: Black Holes in the Universe. Cambridge University Press, Cambridge, UK, third edition, 2021. ISBN 1-108-81905-2 (paperback), 1-108-87112-7 (e-pub). ???? pp. LCCN QB843.B55 B44 2020.

Brauner:1924:EM

[Bra24] Bohuslav Brauner. Einstein and Mach. Nature, 113(2852): 927, June 28, 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v113/n2852/pdf/113927b0.pdf.

Brasch:1929:EAS

[Bra29] Frederick E. Brasch. Einstein’s appreciation of Simon Newcomb. Science, 69(1783):248–249, March 1, 1929. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/69/1783/248.3.full.pdf.

Braunbek:1937:EGM

[Bra37] Werner Braunbek. Die empirische Genauigkeit des Masse-Energie-Verhältnisses. (German) [the empirical accuracy of the mass-energy relationship]. Zeitschrift für Physik, 107(1–2):1– 11, January 1937. CODEN ZEPYAA. ISSN 0044-3328. URL

http://link.springer.com/article/10.1007/BF01330221.

Brallier:2002:WWA

[Bra02] Jess M. Brallier. Who was Albert Einstein? Grosset and Dunlap, New York, NY, USA, 2002. ISBN 0-448-42659-5 (GB), 0-448-42496-7 (paperback). 105 pp. LCCN QC16.E5 B7339 2002. Illustrated by Robert Andrew Parker.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brady:2012:EBF

[Bra12] Frank Brady. Endgame: Bobby Fischer’s remarkable rise and fall — from America’s brightest prodigy to the edge of madness. Crown Publishers, New York, NY, USA, 2012. ISBN 0-307-46391-5 (paperback), 0-307-46392-3 (e-book). xv + 411 + 8 pp. LCCN GV1439.F5 B68 2012.

Braben:2014:MPR

[Bra14a] D. W. Braben. Max Planck: A reluctant revolutionary with a hunger of the soul. In Promoting the Planck Club: how defiant youth, irreverent researchers and liberated universities can foster prosperity indefinitely [Bra14b], pages 38–49. ISBN 1-118-54635-0, 1-306-47279-2, 1-118-54642-3, 1-118-54638-5 (ePub), 1-118-54636-9 (Adobe PDF). LCCN Q141. URL http://

proquest.tech.safaribooksonline.de/9781118546383.

Braben:2014:PPC

[Bra14b] D. W. Braben. Promoting the Planck Club: how defiant youth, irreverent researchers and liberated universities can foster prosperity indefinitely. Wiley, New York, NY, USA, 2014. ISBN 1-118-54635-0, 1-306-47279-2, 1-118-54642-3, 1-118-54638-5 (ePub), 1-118-54636-9 (Adobe PDF). LCCN

Q141. URL http://proquest.tech.safaribooksonline.de/
9781118546383.

Bracco:2015:ESD

[Bra15] Christian Bracco. L’environnement scientifique du jeune Albert Einstein : La période milanaise (1899–1901). (French) [the scientific environment of the young Albert Einstein: the Milan period (1899–1901)]. Revue d’Histoire des Sciences, 68(1): 109–144, January 2015. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.cairn.info/

article.php?ID_ARTICLE=RHS_681_0109.

Bracco:2017:QAD

[Bra17] Christian Bracco. Quand Albert devient Einstein: 1895–1901. (French) [When Albert becomes Einstein: 1895–1901]. CNRS, Paris, France, 2017. ISBN 2-271-09189-6 (paperback). 230 pp. LCCN ????

Bradley:2022:NEE

[Bra22] Clara Bradley. The non-equivalence of Einstein and Lorentz. The British Journal for the Philosophy of Science, 72(4):1039– 1059, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Brdicka:1979:EPC

[Brd79] Miroslav Brdička. Einstein a Praha: Česká einsteinovská pohlednice. (Czech) [Einstein and Prague: Czech Einstein postcard]. Československý časopis fyziky, A29(??):269–275, ???? 1979.

Bredig:1973:BEL

[Bre73] Max A. Bredig. The Born–Einstein letters. Science, 180(4091): 1118, June 15, 1973. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/180/4091/1118.3.full.pdf.

Brecher:1979:AEM

[Bre79] Kenneth Brecher. Albert Einstein: 14 March, 1879–18 April, 1955. A guide for the perplexed. Nature, 278(5701):215– 218, March 15, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v278/n5701/pdf/278215a0.pdf.

BretonB:1989:SAF

[Bre89] Nora Bretón B. Stationary axisymmetric fluid solutions of the Einstein–Born–Infeld equations. Journal of Mathematical Physics, 30(11):2607–2610, November 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/
i11/p2607_s1.

Brennan:1997:HPS

[Bre97] Richard P. Brennan. Heisenberg probably slept here: the lives, times, and ideas of the great physicists of the 20th Century. Wiley, New York, NY, USA, 1997. ISBN 0-471-15709-0 (cloth). xi + 274 pp. LCCN QC15 .B74 1997; 97.E02533.

Breuer:2005:WBM

[Bre05] Reinhard Breuer. Wann begann der Mensch zu denken?; Angriff auf die Wolfsröte; Verstimmtes Universum; Einstein in der Chemie; Gerüchte und Vehler; Menschheit am Scheidweg. (German) [When did Man begin to think?; Attack on the Wolfsröte; Detuned Universe; Einstein in Chemistry; Rumors and Vehler; humanity on Scheidweg], volume Dezember 12/2005 of Spektrum der Wissenschaft. Spektrum der Wissenschaft Verlagsgesellschaft, Heidelberg, Germany, 2005. 122 pp. LCCN ????

Breuer:2008:KVE

[Bre08] Reinhard Breuer. Das kosmische Vergessen: das expandierende Universum löscht seine Geschichte. Was ging bereits verloren?; Hirnforschung: wie das Auge die Welt verfilmt; Klimawandel: Drohen riesige Eismassen ins Meer zu rutschen?; Geschichte: Max Planck und Albert Einstein. (German) [The cosmic oblivion: the expanding Universe clears its history. What has already been lost?; Brain research: how the eye filmed the world; Climate change: threaten huge masses of ice to slip into the sea?; History: Max Planck and Albert Einstein], volume Mai 2008 of Spektrum der Wissenschaft. Spektrum der Wissenschaft Verlagsgesellschaft, Heidelberg, Germany, 2008. 106 pp. LCCN ????

Bredekamp:2015:WCE

[Bre15] Horst Bredekamp. Warburg, Cassirer und Einstein im Gespräch: Kepler als Schlüssel der Moderne. (German) [Warburg, Cassirer and Einstein talking: Kepler as a key of modernity], volume 88 of Kleine kulturwissenschaftliche Bibliothek. Wagenbach, Berlin, Germany, 2015. ISBN 3-8031-5188-0 (hardcover). 144 pp. LCCN ????

Bohi:2009:CMB

[BRH^{+}09] Pascal Böhi, Max F. Riedel, Johannes Hoffrogge, Jakob Reichel, Theodor W. Hänsch, Philipp Treutlein, et al. Coherent manipulation of Bose–Einstein condensates with state-dependent microwave potentials on an atom chip. Nature Physics, 5(8):592– 597, July 5, 2009. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v5/n8/full/nphys1329.html.

Bridgman:1949:ETO

[Bri49] P. W. Bridgman. Einstein’s theories and the operational point of view. In Schilpp [Sch49c], chapter 12, pages 333–354. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Brill:1966:RPP

[Bri66] Dieter Brill. Recent publications and presentations: Albert Einstein and the Cosmic World Order, by Cornelius Lanczos. American Mathematical Monthly, 73(10):1145, December 1966. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brittain:1982:BRB

[Bri82] James E. Brittain. Book review: Ferdinand Braun: A Life of the Nobel Prizewinner and Inventor of the Cathode-Ray Oscilloscope by Friedrich Kurylo; Charles Susskind. Isis, 73(3):482– 483, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
231512.

Brian:1995:GTC

[Bri95] Denis Brian. Genius talk: conversations with Nobel scientists and other luminaries. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-45089-5. xii + 423 pp. LCCN fa1941.

Brian:1996:EL

[Bri96] Denis Brian. Einstein: a life. Wiley, New York, NY, USA, 1996. ISBN 0-471-11459-6 (hardcover), 0-471-19362-3 (paperback). xiv + 509 pp. LCCN QC16.E5 B737 1996. URL http://

www.gbv.de/dms/bowker/toc/9780471114598.pdf;   http:
//www.gbv.de/dms/ilmenau/toc/183614305.PDF;   http:
//www.loc.gov/catdir/enhancements/fy0706/95012075-b.
html;   http://www.loc.gov/catdir/enhancements/fy0706/
95012075-d.html;   http://www.loc.gov/catdir/toc/onix04/
95012075.html;   http://www.zentralblatt-math.org/zmath/
en/search/?an=0849.01023.

Brian:1997:EL

[Bri97] Denis Brian. Einstein: a Life. Wiley, New York, NY, USA, 1997. ISBN 0-471-11459-6 (hardcover), 0-471-19362-3 (paperback). xiv + 509 pp. LCCN QC16.E5 B737 1996.

Brian:2005:CBM

[Bri05a] Denis Brian. The Curies: a biography of the most controversial family in science. Wiley, New York, NY, USA, 2005. ISBN 0-471-27391-0 (hardcover). ix + 438 pp. LCCN QD22.C79 B75

2005. URL http://opac.nebis.ch/cgi-bin/showAbstract.
pl?u20=0471273910.

Brian:2005:ESL

[Bri05b] Denis Brian. Einstein: sein Leben. (German) [Einstein: His Life]. Wiley-VCH, Weinheim, Germany, 2005. ISBN 3-527-40562-3. 727 pp. LCCN ????

Brian:2005:UER

[Bri05c] Denis Brian. The unexpected Einstein: the real man behind the icon. Wiley, New York, NY, USA, 2005. ISBN 0-471-71840-8 (clothbound). xi + 260 pp. LCCN

QC16.E5 B7375 2005. URL http://www.loc.gov/catdir/
enhancements/fy0617/2004025914-b.html;   http://www.
loc.gov/catdir/enhancements/fy0617/2004025914-d.html;
 http://www.loc.gov/catdir/toc/ecip053/2004025914.html.

Brissoni:2005:AER

[Bri05d] Armando Brissoni. Albert Einstein: relativita speciale e dintorni 1889–1905. (Italian) [Albert Einstein: Special Relativity and Outskirts]. Gangemi Editore, Roma, Italy, 2005. ISBN 88-492-0480-9. 191 pp. LCCN QC16.E5 B7376 2004.

Brink:2014:NLP

[Bri14] Lars Brink, editor. Nobel Lectures, Physics, 2006–2010. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4612-67-7 (hardcover), 981-4612-68-5 (paperback). xii + 371 pp. LCCN QC71 .N63 2014.

Bridges:2018:BRM

[Bri18] Thomas J. Bridges. Book review: The Many Faces of Maxwell, Dirac and Einstein Equations: A Clifford Bundle Approach. Contemporary Physics, 59(3):323, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Buchwald:2021:CPAa

[BRI^{+}21a] Diana Kormos Buchwald, Ze’ev Rosenkranz, József Illy, Daniel Kennefick, Anne J. Kox, Dennis Lehmkuhl, Tilman Sauer, and Jennifer Nollar James, editors. The Collected Paper of Albert Einstein: The Berlin Years: Writings and Correspondence, June 1927–May 1929. Princeton University Press, Princeton, NJ, USA, 2021. ISBN 0-691-21681-9 (hardcover). civ + 1021 + 4 pp. LCCN QC16.E5 A2 1987. URL http://einsteinpapers.

press.princeton.edu/.

Buchwald:2021:CPAb

[BRI^{+}21b] Diana Kormos Buchwald, Ze’ev Rosenkranz, József Illy, Daniel Kennefick, Anne J. Kox, Dennis Lehmkuhl, Tilman Sauer, Jennifer Nollar James, William D. Brewer, and Steven Rendall, editors. Collected Papers of Albert Einstein, Volume 16 (Translation Supplement): The Berlin Years, volume 16. Princeton University Press, Princeton, NJ, USA, 2021. ISBN 0-691-21682-7 (paperback). xxxiv + 491 pp. LCCN QC16.E5 A2 1987.

Brose:1919:TRS

[Bro19] Henry Herman L. A. Brose. The Theory of Relativity, sketch based on Einstein’s original writings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1919. 300 pp. LCCN 1986.

Brose:1920:STC

[Bro20a] Henry L. (Henry Leopold) Brose, editor. Space and Time in Contemporary Physics: An Introduction to the Theory of Relativity and Gravitation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1920. 32 pp. LCCN QC6 .B75. English translation of [Sch17]. Reprinted in [SMvdVS79, pp. 207–269].

Brose:1920:TRI

[Bro20b] Henry L. (Henry Leopold) Brose. The Theory of Relativity: an introductory sketch based on Einstein’s original writings including a biographical note. Basil Blackwell, Oxford, UK, revised edition, 1920. 32 pp. LCCN NCP918; 1986; 108.VAR.RYLE; QC6 .B75.

Brown:1921:EDW

[Bro21] Cyril Brown. Einstein declares women rule here; scientist says he found American men the toy dogs of the other sex. People colossally bored; showed excessive enthusiasm over him for lack of other things, he thinks. New York Times, ??(??):9, July 8, 1921. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/98298316/.

Browne:1922:EP

[Bro22] H. C. Browne. Einstein’s paradox. Nature, 110(2768):668– 669, November 18, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v110/n2768/pdf/110668d0.pdf.

Brown:1923:EDH

[Bro23] Cyril Brown. Einstein describes his newest theory. It concerns the relation between electricity and gravitation. Unintelligible to laymen. Founded on theoretical speculations and discoveries of the English astronomer Eddington. New York Times, ??(??): 18, March 27, 1923. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/100235767/.

Brown:1939:BRE

[Bro39] G. Burniston Brown. Book review: The Evolution of Physics: The Growth of Ideas from the Early Concepts to Relativity and Quanta, by Albert Einstein and Leopold Infeld. Philosophy, 14 (54):242, April 1939. ISSN 0031-8191 (print), 1469-817X (elec-

tronic). URL http://www.jstor.org/stable/3746034.

Brode:1960:TA

[Bro60] Bernice Brode. Tales of Los Alamos, 1943–1945. University of California, Berkeley, Berkeley, CA, USA, 1960. 33 pp. LCCN ???? Reprinted from LASL Community News, June 2–September 22, 1960.

Broadbent:1966:BRA

[Bro66] T. A. A. Broadbent. Book review: Albert Einstein and the Cosmic World Order, by Cornelius Lanczos. vi + 139 pp. 1956 (Wiley). Mathematical Gazette, 50(371):60, February 1966. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL http://www.jstor.org/stable/3614838.

Bromberg:1971:INB

[Bro71] Joan Bromberg. The impact of the neutron: Bohr and Heisenberg. Historical Studies in the Physical Sciences, 3(??):307–341, ???? 1971. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757321.

Bromberg:1972:BRT

[Bro72] Joan Bromberg. Book review: Twentieth Century The Born– Einstein Letters. Correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Trans. by Irene Born. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. London: Macmillan, 1971. Pp. xi + 240. £3.85. British Journal for the History of Science, 6(2):222–223, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4025309.

Bronowski:1973:AE

[Bro73a] Jacob Bronowski. Albert Einstein. In The Ascent of Man [Bro73b], pages 244–257. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:AM

[Bro73b] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Broad:1978:SMH

[Bro78a] William J. Broad. Statue on the Mall to hail Einstein’s 100th. Science, 202(4371):951, December 1, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/202/4371/951.full.pdf.

Brown:1978:IN

[Bro78b] Laurie M. Brown. The idea of the neutrino. Physics Today, 31(9):23–28, September 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

nobelprize.org/nobel_prizes/physics/laureates/1935/;
 http://www.physicstoday.org/resource/1/phtoad/v31/

i9/p23_s1. Page 27 contains an English translation of Pauli’s famous letter proposing the existence of an as-yet-undiscovered particle, which he called the ‘neutron’. Fermi later renamed it ‘neutrino’ because of its small (and possibly zero) mass. The real neutron was first discovered by James Chadwick in 1932, a result for which he received the 1935 Nobel Prize in Physics.

Broad:1979:ATM

[Bro79a] William J. Broad. Academy takes a ‘million dollar bath’ with Einstein. Science, 206(4420):798–799, November 16, 1979.

CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/206/
4420/798.full.pdf.

Brown:1979:EBE

[Bro79b] Harrison Brown. An early brief encounter [Albert Einstein and Andrei Gromyko]. Bulletin of the Atomic Scientists, 35(3):17– 19, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Brouwer:1980:ELS

[Bro80] W. Brouwer. Einstein and Lorentz: The structure of a scientific revolution. American Journal of Physics, 48(6):425– 431, 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/48/6/10.1119/1.11997.

Broyles:1982:NE

[Bro82] Arthur A. Broyles. Nuclear explosions. American Journal of Physics, 50(7):586–594, July 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Brouzeng:1983:AOBb

[Bro83] Paul Brouzeng. Analyses d’ouvrages: De Thalès à Einstein — Histoire de la physique et de la Chimie, (coll.^{≪}Axes^{≫}) par Jean Rosmorduc. Revue d’Histoire des Sciences, 36(3–4):368, juillet– décembre 1983. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23632228.

Brown:1987:EBV

[Bro88] James Robert Brown. Einstein’s brand of verificationism. International Studies in the Philosophy of Science, 2(1):33–54, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/02698598708573301.

Brooks:1992:HNW

[Bro92] Geoffrey Brooks. Hitler’s nuclear weapons: the development and attempted deployment of radiological armaments by Nazi Germany. L. Cooper, London, UK, 1992. ISBN 0-85052-344-3. 212 pp. LCCN UA710 .B8442 1992.

Brown:1997:NBB

[Bro97] Andrew Brown. The neutron and the bomb: a biography of Sir James Chadwick. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-853992-4. xiv + 384 + 16 pp. LCCN QC774.C43 B76

1997. URL http://www.loc.gov/catdir/enhancements/
fy0640/97200691-d.html;   http://www.loc.gov/catdir/
enhancements/fy0640/97200691-t.html.

Brown:1999:EAD

[Bro99] Harold I. Brown. Einstein: “all but the dissertation”. Science, 284(5418):1273, May 21, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/284/5418/1273.2.full.

Brown:2000:BRBb

[Bro00] Neil Brown. Book review: Quantum Generations. A history of physics in the twentieth century, by Helge Kragh. Physics Education, 35(2):145, 2000. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://iopscience.iop.

org/0031-9120/35/2/703.

Brooks:2002:HTW

[Bro02] Geoffrey Brooks. Hitler’s terror weapons: from V-1 to Vimana. Leo Cooper, Barnsley, South Yorkshire, UK, 2002. ISBN 0-85052-896-8. 214 pp. LCCN UF505.G3 B76 2002. URL http:

//www.loc.gov/catdir/toc/fy036/2003428026.html.

Brown:2005:PRS

[Bro05] Harvey R. Brown. Physical relativity: space–time structure from a dynamical perspective. Clarendon Press, Oxford, UK, 2005. ISBN 0-19-927583-1. xiii + 225 pp. LCCN

QC173.65 .B76 2005. URL http://www.loc.gov/catdir/
enhancements/fy0639/2005023506-d.html;   http://www.
loc.gov/catdir/toc/ecip0517/2005023506.html.

Brockman:2006:MEE

[Bro06] John Brockman, editor. My Einstein: essays by twentyfour of the world’s leading thinkers on the man, his work, and his legacy. Pantheon Books, New York, NY, USA, 2006. ISBN 0-375-42345-1. xvi + 261 pp. LCCN

QC16.E5 M9 2006. URL http://www.loc.gov/catdir/
enhancements/fy0643/2005048286-b.html;   http://www.
loc.gov/catdir/enhancements/fy0643/2005048286-d.html;
 http://www.loc.gov/catdir/enhancements/fy0643/2005048286-
s.html;   http://www.loc.gov/catdir/enhancements/fy0661/
2005048286-t.html.

Brockman:2007:MEE

[Bro07] John Brockman, editor. My Einstein: essays by twenty-four of the world’s leading thinkers on the man, his work, and his legacy. Vintage Books, New York, NY, USA, 2007. ISBN 1-4000-7950-0. 288 (est.) pp. LCCN QC16.E5 A5 2006.

Brooks:2010:FCT

[Bro10a] Rodney Allen Brooks. Fields of color: the theory that escaped Einstein. R. A. Brooks, Wanaka, New Zealand, 2010. ISBN 0-473-17976-8 (paperback). v + 178 pp. LCCN QC173.98 .B76 2010.

Brown:2010:OAE

[Bro10b] Emma Brown. Obituary: Ann E. Ewing, journalist first reported black holes. Boston Globe, August 3, 2010. ISSN

0743-1791.  URL http://www.boston.com/bostonglobe/
obituaries/articles/2010/08/03/ann_e_ewing_journalist_
first_reported_black_holes/.

Brooks:2011:FCT

[Bro11] Rodney Allen Brooks. Fields of color: the theory that escaped Einstein. R. A. Brooks, Wanaka, New Zealand, second edition, 2011. ISBN 0-473-17976-8. vi + 159 pp. LCCN QC173.98 .B76 2011.

Broad:2015:HBP

[Bro15] William J. Broad. Hydrogen bomb physicist’s book runs afoul of Energy Department. New York Times, ??(??):??, March 23, 2015. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL http://www.nytimes.com/2015/03/
24/science/hydrogen-bomb-physicists-book-runs-afoul-
of-energy-department.html.

Brooks:2016:FCT

[Bro16] Rodney Allen Brooks. Fields of color: the theory that escaped Einstein. Rodney A. Brooks, Wanaka, New Zealand, third edition, 2016. ISBN 0-473-17976-8 (paperback). v + 149 pp. LCCN ????

Brooks:2017:PT

[Bro17] Rodney A. Brooks. A physics tragedy. arXiv.org, ??(??):1–9,

October 21, 2017. URL https://arxiv.org/abs/1710.10291.

Buttner:2003:ELC

[BRS03] Jochen Büttner, Jürgen Renn, and Matthias Schemmel. Exploring the limits of classical physics: Planck, Einstein, and the structure of a scientific revolution. Studies in History and Philosophy of Modern Physics, 34(1):37–59, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980200031X.

Buchwald:2009:BYCa

[BRS^{+}09a] Diana Kormos Buchwald, Ze’ev Rosenkranz, Tilman Sauer, József Illy, and Virginia Iris Holmes, editors. The Berlin Years: Correspondence, January–December 1921 (Documentary Edition), volume 12 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14190-8. LCCN ???? URL

http://einsteinpapers.press.princeton.edu/vol12-doc;
 http://einsteinpapers.press.princeton.edu/vol12-trans;
 http://press.princeton.edu/titles/8954.html.

Buchwald:2009:BYCb

[BRS^{+}09b] Diana Kormos Buchwald, Ze’ev Rosenkranz, Tilman Sauer, József Illy, and Virginia Iris Holmes, editors. The Berlin Years: Correspondence, January–December 1921. English Translation Supplement, volume 12 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14190-8. LCCN ????

URL http://einsteinpapers.press.princeton.edu/vol12-
trans;   http://press.princeton.edu/titles/8966.html.

Blum:2012:TPE

[BRS^{+}12] Alexander S. Blum, Jürgen Renn, Donald C. Salisbury, Matthias Schemmel, and Kurt Sundermeyer. 1912: A turning point on Einstein’s way to general relativity. Annalen der Physik (1900), 524(1):A11–A13, January 2012. ISSN 1521-3889.

Brush:1968:HRP

[Bru68] Stephen G. Brush. A history of random processes. I. Brownian motion from Brown to Perrin. Archive for History of Exact Sciences, 5(1):1–36, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=1&spage=1.

Brush:1976:CSC

[Bru76] Stephen G. Brush. Can science come out of the laboratory now? Bulletin of the Atomic Scientists, 32(4):40–43, April 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [?, ?, ?].

Brustein:1977:DAE

[Bru77] Robert Brustein. Drama in the age of Einstein. New York Times, ??(??):D1, August 7, 1977. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
123261664/.

Brush:1982:BE

[Bru82] Stephen G. Brush. Before Einstein. Nature, 299(5886): 845, October 28, 1982. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v299/n5886/pdf/299845b0.pdf.

Bruce:1997:EPO

[Bru97a] Colin Bruce. The Einstein paradox and other science mysteries solved by Sherlock Holmes. Perseus Books, Reading, MA, USA, 1997. ISBN 0-7382-0023-9 (paperback). x + 254 pp. LCCN PS3552.R7917 S77 1997b. US$12.00, CAN$17.50.

Bruce:1997:SCM

[Bru97b] Colin Bruce. The strange case of Mrs. Hudson’s cat: and other science mysteries solved by Sherlock Holmes. Helix books. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-46139-0. x + 254 pp. LCCN PS3552.R7917 S77 1997.

Bruce:1998:SCM

[Bru98] Colin Bruce. The strange case of Mrs Hudson’s cat: or Sherlock Holmes solves the Einstein mysteries. A Vintage original. Vintage Books, New York, NY, USA, 1998. ISBN 0-09-926769-1. viii + 254 pp. LCCN fa2466.

Brush:1999:WWR

[Bru99] Stephen G. Brush. Why was Relativity accepted? Physics in Perspective (PIP), 1(2):184–214, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050015.

Bruckner:2000:MFF

[Brü00] Reinhold Brückner. Meitner and the forgotten fragment. Physics World, 13(5):20, May 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http:/

/stacks.iop.org/2058-7058/13/i=5/a=23.

Bruce:2003:IUE

[Bru03a] Barry Bruce. The infinite universe of Einstein and Newton. Brown Walker Press, Parkland, FL, USA, 2003. ISBN 1-58112-410-4. ii + 108 pp. LCCN QB981 .B787 2003.

Bruno:2003:BNB

[Bru03b] Laura A. Bruno. The bequest of the nuclear battlefield: Science, nature, and the atom during the first decade of the Cold War. Historical Studies in the Physical and Biological Sciences, 33 (2):237–260, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Brush:2003:BRR

[Bru03c] Stephen G. Brush. Book review: Robert Marc Friedman, The Politics of Excellence: Behind the Nobel Prize in Science. New York: A. H. Freeman/Times Books/Henry Holt, 2001, xv + 379 pages. $30.00 (cloth). Physics in Perspective (PIP), 5(2):235– 238, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brush:2006:BRM

[Bru06] Stephen G. Brush. Book review: Mauro Dardo, Nobel Laureates and Twentieth-Century Physics. Cambridge: Cambridge University Press, 2004, xi + 533 pages. $85.00 (cloth). Physics in Perspective (PIP), 8(1):105–106, March 2006. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bruzzaniti:2007:EFG

[Bru07] Giuseppe Bruzzaniti. Enrico Fermi: il genio obbediente. (Italian) [Enrico Fermi: the obedient genius], volume 882 of Saggi. G. Einaudi, Torino, Italy, 2007. ISBN 88-06-16682-4 (paperback). xiii + 386 pp. LCCN QC774.F4 B78

2007. 24.50 EUR. URL http://matematica.unibocconi.
it/libri/enrico-fermi-il-genio-obbediente;   http://
www.loc.gov/catdir/toc/casalini06/06541291.pdf.

Brunner:2009:NFM

[Bru09] A. Brunner. Never forget this in making your drawings and equations! A conversation with Albert Einstein on learning, teaching and the secrets of the world. Gesundheitswesen, 71(3): 175–179, March 2009. ISSN 0941-3790.

Bertin:1984:FYW

[BRVF84] Antonio Bertin, R. A. Ricci, A. (Antonio) Vitale, and Enrico Fermi, editors. Fifty years of weak-interaction physics: on the occasion of the fiftieth anniversary of Fermi’s theory on nuclear beta-decay. Italian Physical Society, Bologna, Italy, 1984. 31 + xl + 797 + 3 pp. LCCN QC794.8.W4 F54 1984.

Bryson:2009:RSH

[Bry09] Bill Bryson. A really short history of nearly everything. Delacorte Press, New York, NY, USA, 2009. ISBN 0-385-73810-2. 169 pp. LCCN Q163 .B875 2008.

Bunge:1979:RPT

[BS79] Mario Bunge and William R. Shea, editors. Rutherford and physics at the turn of the century. Dawson, Kent, UK, 1979. ISBN 0-88202-184-2 (New York), 0-7129-0918-4 (Cannon House Folkestone). 184 pp. LCCN QC7.5 .R87 1979.

Barker:1981:AEP

[BS81] Peter Barker and Cecil G. Shugart, editors. After Einstein: proceedings of the Einstein centennial celebration at Memphis State University, 14–16 March, 1979. Memphis State University Press, Memphis, TN, USA, 1981. ISBN 0-87870-095-1. viii + 244 pp. LCCN QC16.E5 A63. US$24.95.

Barrow:1983:LHC

[BS83a] John D. Barrow and Joseph Silk. The left hand of creation: the origin and evolution of the expanding universe. Basic Books, New York, NY, USA, 1983. ISBN 0-465-03895-6. xi + 256 pp. LCCN QB991.E94 B37 1983. US$17.95.

Bethe:1983:NPP

[BS83b] Hans A. Bethe and E. B. Salpeter. The 1983 Nobel Prize in Physics — Chandrasekhar and Fowler. Science, 222(4626):881– 883, November 25, 1983. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/222/4626/881.extract.

Breit:1986:LLS

[BS86] William Breit and Roger W. Spencer, editors. Lives of the laureates: seven Nobel economists. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-02255-9. xii + 135 pp. LCCN HB76 .L58 1986.

Baumer-Schleinkofer:1992:RDN

[BS92] Änne Bäumer-Schleinkofer. Rezension: David Nachmansohn: Die große Ära der Wissenschaft in Deutschland 1900 bis 1933. Jüdische und nichtjüdische Pioniere in der Atomphysik, Chemie und Biochemie. Aus dem Englischen überarbeitet und erweitert von Roswitha Schmid. Stuttgart: Wissenschaftliche Verlagsgesellschaft 1988. 398 Seiten. Gebunden, DM 88. Berichte zur Wissenschaftsgeschichte, 15(2):97–98, ???? 1992. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Barrow:1993:LHC

[BS93] John D. Barrow and Joseph Silk. The left hand of creation: the origin and evolution of the expanding universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-508675-9, 0-19-508676-7 (paperback). xxv + 262 pp. LCCN QB991.E94 B37 1993.

Brode:1997:TAL

[BS97] Bernice Brode and Barbara Storms. Tales of Los Alamos: life on the Mesa, 1943–1945. Los Alamos Historical Society, Los Alamos, NM, USA, 1997. ISBN 0-941232-17-4 c(paperback). v + 157 pp. LCCN QC773.3.U5 B76 1997.

Bird:2005:APT

[BS05] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-375-72626-8, 0-375-41202-6. xiii + 721 + 32 pp. LCCN QC16.O62 B57 2005. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://site.ebrary.com/lib/ucsc/Doc?id=10078784.

Baumgarte:2010:NRS

[BS10] Thomas W. Baumgarte and Stuart L. (Stuart Louis) Shapiro. Numerical relativity: solving Einstein’s equations on the computer. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51407-X. xviii + 698 pp. LCCN QC173.6 .B38 2010.

Buchwald:2004:BYCa

[BSKS04a] Diana Kormos Buchwald, Robert Schulmann, József Illy’ Daniel J. Kennefick, and Tilman Sauer. The Berlin Years: Correspondence, January 1919–April 1920, volume 9 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-12088-9. lxxiii + 697 + 8 pp. LCCN QC 16 .E5A214 1987. URL http://

einsteinpapers.press.princeton.edu/vol9-doc;   http:
//einsteinpapers.press.princeton.edu/vol9-trans;   http:
//press.princeton.edu/titles/7910.html;   http://press.
princeton.edu/titles/7911.html.

Buchwald:2004:BYCb

[BSKS04b] Diana Kormos Buchwald, Robert Schulmann, József Illy’ Daniel J. Kennefick, and Tilman Sauer. The Berlin Years: Correspondence, January 1919–April 1920. English Translation Supplement, volume 9 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-12124-9. xxi + 358 pp.

URL http://press.princeton.edu/titles/7911.html;
https://einsteinpapers.press.princeton.edu/vol9-trans.

Braun:2003:NBA

[BSPKN03] Tibor Braun, Zsuzsa Szabadi-Peresztegi, and Éva Kovács-Németh. No-bells for ambiguous lists of ranked Nobelists as science indicators of national merit in physics, chemistry and medicine, 1901–2001. Scientometrics, 56(1):3–28, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021998006078.

Buchwald:2006:BYCa

[BSR^{+}06a] Diana Kormos Buchwald, Tilman Sauer, Ze’ev Rosenkranz, Jószef Illy, and Virginia Iris Holmes, editors. The Berlin Years: Correspondence, May–December 1920, and Supplementary Correspondence, 1909–1920, volume 10 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12825-1. lxix + 683 pp. LCCN QC16.E5 EIN; QC16.E5 A2

1987. URL http://einsteinpapers.press.princeton.edu/
vol10-doc;   http://einsteinpapers.press.princeton.
edu/vol10-trans;   http://press.princeton.edu/titles/
8331.html.

Buchwald:2006:BYCb

[BSR^{+}06b] Diana Kormos Buchwald, Tilman Sauer, Ze’ev Rosenkranz, Jószef Illy, and Virginia Iris Holmes, editors. The Berlin Years: Correspondence, May–December 1920, and Supplementary Correspondence, 1909–1920. English Translation Supplement, volume 10 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12825-1. lxix + 683 pp. LCCN QC16.E5 EIN; QC16.E5 A2

1987. URL http://einsteinpapers.press.princeton.edu/
vol10-doc;   http://einsteinpapers.press.princeton.
edu/vol10-trans;   http://press.princeton.edu/titles/
8331.html.

Ball:2017:BBE

[BSS^{+}17] Hugo Ball, Joerg K. Sommermeyer, Orlando Syrg, Carl Einstein, and Ludwig Rubiner. Balleinrubin : Ball, Einstein, Rubiner. Books on Demand, Norderstedt, Germany, 2017. ISBN 3-7448-1673-7. 148 pp. LCCN

???? URL http://deposit.d-nb.de/cgi-bin/dokserv?id=
e7ccf0dd4afb47a9a585cf16d3b6f5b9%26prov=M%26dok_var=
1%26dok_ext=htm.

Becker:2008:OID

[BSSP^{+}08] Christoph Becker, Simon Stellmer, Parvis Soltan-Panahi, Sören Dörscher, Mathis Baumert, Eva-Maria Richter, Jochen Kronjäger, Kai Bongs, Klaus Sengstock, et al. Oscillations and interactions of dark and dark–bright solitons in Bose–Einstein condensates. Nature Physics, 4(6):496–501, May 4, 2008. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic).

URL http://www.nature.com/nphys/journal/v4/n6/full/
nphys962.html.

Batyi:2004:SLP

[BSST04] Emese Bátyi, Leo Szilard, Orsolya Szîucs, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2004. Oktatási Minisztérium, Budapest, Hungarian, 2004. 60 pp. LCCN MLCS 2008/45435 (Q).

Batyi:2003:SLP

[BST03] Emese Bátyi, Leo Szilard, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2003. MF Plusz, Budapest, Hungary, 2003. 64 pp. LCCN MLCS 2008/45436 (Q).

Blumenthal:1922:RSA

[BSW22] Otto Blumenthal, Arnold Sommerfeld, and Hermann Weyl. Das Relativitätsprinzip: Eine Sammlung von Abhandlungen [von] H[endrik] A[ntoon] Lorentz, A[lbert] Einstein, H[ermann] Minkowski. (German) [The Principle of Relativity: A collection of writings of H[endrik] A[ntoon] Lorentz, A[lbert] Einstein, H[ermann] Minkowski], volume 2 of Fortschritte der mathematischen Wissenschaften in Monographien …. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fourth edition, 1922. iv + 159 pp. LCCN ????

Barrow:1986:ACP

[BT86] John D. Barrow and Frank J. Tipler. The anthropic cosmological principle. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-851949-4. xx + 706 pp. LCCN BD511 .B34 1986. US$29.95.

Barrow:1988:ACP

[BT88] John D. Barrow and Frank J. Tipler. The anthropic cosmological principle. Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-282147-4. xx + 706 pp. LCCN BD511 .B34 1988. US$15.95.

Bellucci:2012:TGT

[BT12] Stefano Bellucci and Bhupendra Nath Tiwari. Thermodynamic geometry and topological Einstein–Yang–Mills black holes. Entropy, 14(6):1045–1078, ???? 2012. CODEN ENTRFG. ISSN

1099-4300. URL https://dblp.org/db/journals/entropy/
entropy14.html#BellucciT12;   https://www.wikidata.
org/entity/Q62558493.

Bauerlein:2015:DWE

[BT15] Theresa Bäuerlein and Shai Tubali. Denken wie Einstein: Was wir von den klügsten Köpfen der Geschichte lernen können. (German) [Thinking like Einstein: What we can learn from the brightest minds in history]. Lagato Verlag, Leipzig, Germany, 2015. ISBN 3-942748-69-X. LCCN ???? Audio CDs.

Barrow:1984:LCP

[BTM84] John D. Barrow, Frank J. Tipler, and Marie-Odile Monchicourt. L’homme et le cosmos: le principe anthropique en astrophysique moderne. (French) [Man and the cosmos: The Anthropic Cosmological Principle and Modern Astrophysics]. Editions Imago, Paris, France, 1984. ISBN 2-902702-19-1. 113 pp. LCCN BD512 .B37 1984. FF68.00.

Badge:2008:NFG

[BTR08] Peter Badge, Nikolaus Turner, and Chris Richmond, editors. Nobel Faces: a gallery of Nobel Prize winners. Wiley-VCH, Weinheim, Germany, 2008. ISBN 3-527-40678-6. vii + 621 pp. LCCN AS911.N9 B34 2008.

Bub:1973:USB

[Bub73] Jeffrey Bub. Under the spell of Bohr. The British Journal for the Philosophy of Science, 24(1):78–90, March 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
24/1/78.full.pdf+html.

Bub:1979:SRQ

[Bub79]  Jeffrey Bub.  Some reflections on quantum logic and

Schrödinger’s Cat. The British Journal for the Philosophy of Science, 30(1):27–39, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/30/1/27.full.pdf+html.

Buchler:1921:LWB

[Büc21] Robert Büchler. Lehrsätze über das Weltall mit Beweis in form eines offenen Briefes an Professor Einstein. (German) [Propositions about the Universe with proof in the form of an open letter to Professor Einstein]. ????, Aachen, Germany, 1921.

Bucciarelli:1980:EGC

[Buc80] B. Bucciarelli, editor. Einstein, Galileo: commemoration d’Albert Einstein, 1979, Pontificia academia scientiarum, Collana Scienza e fede. Pontificia Accademia delle scienze, Libreria editrice vaticana, Vatican City, 1980. URL http://books.

google.com/books?id=pnpkYgEACAAJ.

Bucky:1993:PAE

[Buc93] Peter A. Bucky. Der private Albert Einstein: Gespräche über Gott, die Menschen und die Bombe. (German) [The private Albert Einstein: speeches on God, Man, and the Bomb], volume 26055 of ECON–Sachbuch. ECON-Taschenbuch-Verlag, Düsseldorf, Germany, 1993. ISBN 3-612-26055-3. 303 pp. LCCN ???? Translation to German by Kurt Simon of [BW92].

Buchwald:1994:CSE

[Buc94] Jed Z. Buchwald. The creation of scientific effects: Heinrich Hertz and electric waves. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-07887-6, 0-226-07888-4 (paperback). xiv + 482 pp. LCCN

QC661 .B85 1994. URL http://www.loc.gov/catdir/
description/uchi052/93041783.html;   http://www.loc.
gov/catdir/toc/uchi051/93041783.html.

Buchert:2008:DES

[Buc08] Thomas Buchert. Dark Energy from structure: a status report. General Relativity and Gravitation, 40(2–3):467–527, ???? 2008. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532

(electronic). URL http://www.springerlink.com/content/
d3531254w7276862/.

Buchner:2013:ADE

[Buc13] Johannes Buchner. Ancient Dynamics of the Einstein Equations and the Tumbling Universe. Freie Universität Berlin, Berlin, Germany, 2013. vii + 90 pp. LCCN ???? URL http://d-nb.

info/1045859281/34.

Buchwald:2020:EWR

[Buc20] Jed Z. Buchwald, editor. Einstein was right: the science and history of gravitational waves. Princeton University Press, Princeton, NJ, USA, 2020. ISBN 0-691-19454-8. xi + 235 pp. LCCN QC16.E5 .E367 2020; QC179 .E53 2020.

Buhrke:1998:LNV

[Büh98a] Thomas Bührke. “Newton, verzeih’ mir!” Albert Einstein (1879–1955). (German) [“Newton, forgive me!” Albert Einstein (1879–1955)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 84–105. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Buhrke:2015:EJG

[Büh15] Thomas Bührke. Einsteins Jahrhundertwerk: die Geschichte einer Formel. (German) [Einstein’s centennial work: the history of a formula], volume 26052 of dtv premium. DTV premium, München, Germany, 2015. ISBN 3-423-26052-1 (paperback).

280 pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/
12148/cb44276101w.

Bunche:1979:PNP

[Bun79a] Ralph J. Bunche. Perspective: Nobel Peace Laureates: 1950. Bulletin of the Atomic Scientists, 35(1):1, January 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 11 December 1950, Oslo, Norway.

Bunge:1979:EBD

[Bun79b] Mario Bunge. The Einstein–Bohr debate over quantum mechanics: who was right about what? In H. Nelkowski, A. Hermann, H. Poser, R. Schrader, and R. Seiler, editors, Einstein Symposion, Berlin (1979), volume 100 of Lecture Notes in Physics, pages 204–219. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. URL https://ui.

adsabs.harvard.edu/#abs/1979LNP...100..204B.

Bundy:1982:MCS

[Bun82] McGeorge Bundy. The missed chance to stop the H-bomb. New York Review of Books, 29(8):13–21, May 13, 1982. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:

//www.nybooks.com/articles/archives/1982/may/13/the-
missed-chance-to-stop-the-h-bomb/.

Bundy:1988:DSC

[Bun88] McGeorge Bundy. Danger and survival: choices about the bomb in the first fifty years. Random House, New York, NY, USA, 1988. ISBN 0-394-52278-8. xiii + 735 pp. LCCN UA23 .B786 1988.

Bundy:1990:DSC

[Bun90] McGeorge Bundy. Danger and survival: choices about the bomb in the first fifty years. Vintage Books, New York, NY, USA, 1990. ISBN 0-679-72568-7 (paperback). xiii + 735 pp. LCCN UA23 .B7862 1990.

Burchett:1945:AP

[Bur45] Wilfred G. Burchett. The atomic plague: ‘I write this as a warning to the world.’ Doctors fall as they work. poison gas fear: all wear masks. Daily Express (London, UK), ??(??):??, September 2, 1945. URL http:/

/www.factsandopinions.com/dispatches/science/people/
wilfred-burchett-a-journalists-warning-to-the-world/.

Burhop:1968:PHB

[Bur68] E. H. S. Burhop. Professor Hans Bethe. Physics Education, 3 (1):16, 1968. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-

6552 (electronic). URL http://stacks.iop.org/0031-9120/
3/i=1/a=304.

Burchett:1983:SH

[Bur83] Wilfred G. Burchett. Shadows of Hiroshima. Verso, London, UK, 1983. ISBN 0-86091-783-5 (paperback), 0-86091-080-6 (hardcover). 123 pp. LCCN D767.25.H6 B86 1983.

Burian:1990:MDH

[Bur90] Richard M. Burian. Maiocchi on Duhem, Howard on Duhem and Einstein: Historiographical comments. Synthese, 83(3):401–408, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/
10.1007/BF00413424.

Burton:1994:PLA

[Bur94] H. D. Burton. The private lives of Albert Einstein — R. Highfield, P. Carter. Library Journal, 119(10):120, June 1994. CO-DEN LIBJA7. ISSN 0363-0277.

Burton:1997:AEF

[Bur97] H. Burton. Albert Einstein — A. Fölsing. Library Journal, 122 (3):159, February 15, 1997. CODEN LIBJA7. ISSN 0363-0277.

Burrell:2015:NGJ

[Bur15] Brian D. Burrell. Neuroscience: Genius in a jar. Scientific American, 313(3):82–87, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-82.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-82.pdf.

Bush:1949:MAF

[Bus49] Vannevar Bush. Modern arms and free men: a discussion of the role of science in preserving democracy. Simon and Schuster, New York, NY, USA, 1949. 273 pp. LCCN U102 .B985.

Bush:1968:MAF

[Bus68] Vannevar Bush. Modern arms and free men; a discussion of the role of science in preserving democracy. MIT Press, Cambridge, MA, USA, 1968. xiv + 273 pp. LCCN U102 .B985 1968.

Bush:1985:MAF

[Bus85] Vannevar Bush. Modern arms and free men: a discussion of the role of science in preserving democracy. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1985. ISBN 0-313-24985-7. 273 pp. LCCN U102 .B985 1985.

Bush:2003:EPJ

[Bus03] Alfred Bush. Einstein’s poems to Johanna Fantova: An introduction. Princeton University Library Chronicle, 65(1):79–82, Autumn 2003. ISSN 0032-8456. URL https://www.jstor.

org/stable/10.25290/prinunivlibrchro.65.1.0079.

Bussey:2012:CAA

[Bus12a] P. J. Bussey. Cosmic Anger: Abdus Salam — The First Muslim Nobel Scientist, by Gordon Fraser, Scope: biography. Level: general readership. Contemporary Physics, 53(5):442–443, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:AEM

[Bus12b] Peter J. Bussey. Albert Einstein Memorial Lectures, by Jacob D. Bekenstein and Raphael Mechoulam. Scope: edited book. Level: non-specialists. Contemporary Physics, 53(6):511–512, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussolini:2012:ALD

[Bus12c] Jeffrey Bussolini. Los Alamos as laboratory for domestic security measures: nuclear age battlefield transformations and the ongoing permutations of security. In Grondin [Gro12], pages 77–106. ISBN 0-415-52368-0. LCCN U21.5 .W37 2012.

Bush:2013:BRB

[Bus13a] Elizabeth Bush. Book review: On a Beam of Light: A Story of Albert Einstein by Jennifer Berne. Bulletin of the Center for Children’s Books, 66(10):454, 2013. CODEN ???? ISSN 0008-9036 (print), 1558-6766 (electronic).

Busi:2013:DRM

[Bus13b] Marco Busi. Doing Research That Matters: Can a Management Researcher win the Nobel Prize? Emerald Group Publishing Limited, Bradford, UK, 2013. ISBN 0-85724-707-7. 218

(est.) pp. LCCN ???? URL http://gbv.eblib.com/patron/
FullRecord.aspx?p=1250100.

Butow:1954:JDS

[But54] Robert J. C. (Robert Joseph Charles) Butow. Japan’s decision to surrender, volume 24 of The Hoover Library on War, Revolution, and Peace. Publication. Stanford University Press, Stanford, CA, USA, 1954. xi + 259 pp. LCCN D821.J3 B8 1954. Foreword by Edwin O. Reischauer.

Butler:1968:TNE

[But68] J. W. Butler. On the Trouton–Noble experiment. American Journal of Physics, 36(11):936–, November 1968. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/
v36/i11/p936_s1.

Butterfield:2007:SEL

[But07] Jeremy Butterfield. Stochastic Einstein locality revisited. The British Journal for the Philosophy of Science, 58(4):805–867, December 2007. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/58/4/805.full.pdf+html.

Butler:2023:TNL

[But23] Stella V. F. Butler. Two Nobel laureates in conversation: Robert Robinson listens to Dorothy Hodgkin’s account of her life scientific. Notes and Records of the Royal Society of London, 77(3):537–556, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Balashov:2002:ESR

[BV02] Yuri Balashov and Vladimir Pavlovich Vizgin, editors. Einstein studies in Russia, volume 10 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002. ISBN 0-8176-4263-3, 3-7643-4263-3. xii + 314 pp. LCCN QC16.E5 E518 2002. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/fy031/2002018301.html.

Bacciagaluppi:2009:QTC

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96

.B33 2009. URL http://www.loc.gov/catdir/toc/ecip0817/
2008019585.html.

Bohr:1939:FP

[BW39a] Niels Bohr and John A. Wheeler. The fission of protactinium. Physical Review, 56(10):1065–1066, November 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v56/i10/p1065_2.

Bohr:1939:MNF

[BW39b] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p426_1.

Brieger:1987:NPW

[BW^{+}87] Gert H. Brieger, Tyler Wasson, et al., editors. Nobel Prize winners: an H. W. Wilson biographical dictionary. H. W. Wilson, New York, NY, USA, 1987. ISBN 0-8242-0756-4. xxxiv + 1165 pp. LCCN AS911.N9 N59 1987.

Bucky:1991:PAE

[BW91] Peter A. Bucky and Allen G. Weakland. Der private Albert Einstein: Gespräche über Gott, die Menschen und die Bombe. (German) [The private Albert Einstein: talks on God, Man, and the Bomb]. ECON-Verl, Düsseldorf, Germany, 1991. ISBN 3-430-11589-2. 303 pp. LCCN ???? Translation to German by Kurt Simon of [BW92].

Bucky:1992:PAE

[BW92] Peter A. Bucky and Allen G. Weakland. The private Albert Einstein. Andrews and McMeel, Kansas City, MO, USA, 1992. ISBN 0-8362-7997-2. xii + 171 + 16 pp. LCCN QC16.E5 B83 1992. US$18.95, CAN$26.95.

Buchwald:2001:HEB

[BW01] Jed Z. Buchwald and Andrew Warwick, editors. Histories of the Electron: the Birth of Microphysics. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02494-2 (hardcover). xi + 514 pp. LCCN QC793.5.E62 H57 2001.

URL http://www.e-streams.com/es0509/es0509_2104.htm;
 http://www.loc.gov/catdir/toc/fy033/00046580.html.

Buschhorn:2004:FPH

[BW04] Gerd W. Buschhorn and Julius Wess, editors. Fundamental physics — Heisenberg and beyond: Werner Heisenberg Centennial Symposium “Developments in Modern Physics”. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20201-3. LCCN QC173.96 .W47

2004. URL http://www.springer.com/978-3-540-20201-1.

Byrnes:1947:SF

[Byr47] James F. (James Francis) Byrnes. Speaking frankly. Harper, New York, NY, USA, 1947. xii + 324 pp. LCCN D815 .B9.

Byrnes:1974:SF

[Byr74] James F. (James Francis) Byrnes. Speaking frankly. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1974. ISBN 0-8371-7480-5. xii + 324 pp. LCCN D815 .B9 1974.

Byrne:1980:SEU

[Byr80a] Patrick H. Byrne. The significance of Einstein’s use of the history of science. Dialectica: International Review of Philosophy of Knowledge, 34(4):263–276, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Byrne:1980:SCC

[Byr80b] Patrick H. Byrne. Statistical and causal concepts in Einstein’s early thought. Annals of Science, 37(2):215–228, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Byrne:1981:OEU

[Byr81] Patrick H. Byrne. The origins of Einstein’s use of formal asymmetries. Annals of Science, 38(2):191–206, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Babkin:2010:NLN

[BZ10] V. A. (Vladimir Aleksandrovich) Babkin and G. E. (Gennadiĭ Efremovich) Zaikov. Nobel Laureates and Nanotechnologies of Applied Quantum Chemistry. Nova Science Publishers, New York, NY, USA, 2010. ISBN 1-61668-849-1. LCCN ????

C:1923:ERE

[C.23] A. C. D. C. Einstein and the recent eclipse. Nature, 111 (2790):541, April 21, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v111/n2790/pdf/111541a0.pdf.

C:1950:RMS

[C.50] H. T. C. Reviews: Modern Science and its Philosophy, by Philipp Frank and Relativity: A Richer Truth by Philipp Frank and Albert Einstein. The Journal of Philosophy, 47(23):666– 671, November 1950. ISSN 0022-362x (print), 1939-8549 (elec-

tronic). URL 2-H;   http://links.jstor.org/sici?sici=
0022-362X(19501109)47:23<666:MSAIP>2.0.CO.

C:1951:BRA

[C.51a] C. W. C. Book review: Albert Einstein. His Work and its Influence on our World by Leopold Infeld. The Journal of Philosophy, 48(7):222, 1951. ISSN 0022-362X (print), 1939-8549

(electronic). URL http://www.jstor.org/stable/2020458.

C:1951:RML

[C.51b] H. T. C. Reviews: Out of my Later Years, by Albert Einstein, and Reflections of a Physicist by Percy Williams Bridgman. The Journal of Philosophy, 48(13):427, June 1951. ISSN 0022-362x (print), 1939-8549 (electronic). URL 2-Z; http:/

/links.jstor.org/sici?sici=0022-362X(19510621)48:13<
427:OOMLY>2.0.CO.

Chimbidis:1977:FLS

[CA77] James Chimbidis and Jay Andre, editors. To Fermi–with love. Argonne National Laboratory, Lemont, IL, USA, 1977. Sound recording (33 1/3 rpm).

Cukerman:1999:ASS

[ČA99] Veniamin Aronovich Čukerman and Zinaida Matveevna Azarch, editors. Arzamas-16: Soviet Scientists in the Nuclear Age: a Memoir, volume 3 of Russian memoirs series. Bramcote Press, Nottingham, UK, 1999. ISBN 1-900405-04-0. xxxiv + 182 pp. LCCN QC773.3.R9 T7813 1999. URL http://www.gbv.de/

dms/bowker/toc/9781900405041.pdf.

Cahn:1955:EPB

[Cah55] William Cahn. Einstein, a pictorial biography. Citadel Press, New York, NY, USA, 1955. 126 pp. LCCN QC16.E5 C3.

Cahan:1985:IE

[Cah85] David Cahan. Influences on Einstein. Science, 229(4715):751– 752, August 23, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/229/4715/751.2.full.pdf.

Cahan:1990:BRC

[Cah90a] D. Cahan. Book review: The Collected Papers of Albert Einstein, Vol 2, the Swiss Years — Writings, 1900–1909 — J. Stachel. Science, 248(4957):878–879, May 18, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Cahan:1990:BRB

[Cah90b] David Cahan. Book review: The Collected Papers of Albert Einstein. Vol. 2, The Swiss Years: Writings, 1900– 1909. John Stachel, editor. David C. Cassidy, Jürgen Renn, and Robert Schulmann, associate editors. Princeton University Press, Princeton, NJ, 1989. xxxvi, 656 pp. $85. English translation (Anna Beck, translator), xvi, 399 pp. Paper, $25. Science, 248(4957):878–879, May 18, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/248/4957/878.full.pdf.

Cahn:2004:NIR

[Cah04] Robert W. Cahn. Nobel idiosyncracies: a review of Nobel Lectures: Physics (1996–2000). Edited by Gösta Ekspong. (World Scientific, 2002). Pp. ix + 485. ISBN 981-238-004-3. Contemporary Physics, 45(1):71–73, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cahn:2005:UNP

[Cah05] Robert W. Cahn. An unusual Nobel prize. Notes and Records of the Royal Society of London, 59(2):145–153, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041485.

Calo:1975:EFC

[Cal75] Vincenzo Calò. Enrico Fermi e le centrali nucleari. (Italian) [Enrico Fermi and the Nuclear Reactor]. Mezzina, Molfetta, Italia, 1975. 64 pp. LCCN TK1078 .C34. L2500.

Callahan:1976:CSF

[Cal76] J. J. Callahan. The curvature of space in a finite universe. Scientific American, 235(2):90–100, August 1976. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Calder:1979:EU

[Cal79] Nigel Calder. Einstein’s universe. Viking, New York, NY, USA, 1979. ISBN 0-670-29076-9. 154 + 5 pp. LCCN QC173.55 .C34.

Calder:1980:EU

[Cal80] Nigel Calder. Einstein’s universe. Penguin, London, UK and New York, NY, USA, 1980. ISBN 0-14-005499-5. 254 + 8 pp. LCCN QC173.55 .C34 1980.

Calder:1982:EU

[Cal82] Nigel Calder. Einstein’s universe. Greenwich House, New York, NY, USA, 1982. ISBN 0-517-38570-8. vii + 154 + 12 pp. LCCN QC173.55 .C34 1982. URL http://www.loc.gov/

catdir/description/random0415/82009218.html.

Caldicott:1986:PH

[Cal86] Helen Caldicott. PSR history. Bulletin of the Atomic Scientists, 42(3):57, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano85].

Calaprice:1996:QE

[Cal96] Alice Calaprice. The quotable Einstein. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02696-3 (hardcover). xxxiv + 269 pp. LCCN QC16.E5 A25

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
prin031/96003543.html;   http://www.loc.gov/catdir/
toc/prin031/96003543.html.

Calaprice:2000:EQE

[Cal00] Alice Calaprice. The expanded quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-07021-0 (hardcover). xliii + 407 pp. LCCN QC16.E5

A25 2000.  URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.loc.gov/catdir/bios/
prin051/00026873.html;   http://www.loc.gov/catdir/
description/prin022/00026873.html;   http://www.loc.
gov/catdir/samples/prin031/00026873.html;   http://
www.loc.gov/catdir/toc/prin031/00026873.html.

Calaprice:2002:DPE

[Cal02] Alice Calaprice. Dear Professor Einstein: Albert Einstein’s Letters to and from Children. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-61592-276-8. ???? pp. LCCN ????

Calaprice:2003:ELM

[Cal03a] Alice Calaprice. Einstein’s last musings. Princeton University Library Chronicle, 65(1):65–78, Autumn 2003. ISSN 0032-

8456. URL https://library.princeton.edu/libraries/
firestone/rbsc/aids/libchron/pulc_v_65_n_1_contents.
pdf;   https://www.jstor.org/stable/10.25290/prinunivlibrchro.
65.1.0065.

Calder:2003:MUO

[Cal03b] Nigel Calder. Magic universe: the Oxford guide to modern science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-850792-5. xii + 756 pp. LCCN Q125 .C286 2003. URL http://catdir.loc.

gov/catdir/enhancements/fy0620/2004297022-d.html;
http://catdir.loc.gov/catdir/enhancements/fy0620/2004297022-
t.html;   http://catdir.loc.gov/catdir/enhancements/
fy0722/2004297022-b.html.

Calaprice:2004:EA

[Cal04] Alice Calaprice. The Einstein almanac. The Johns Hopkins University Press, Baltimore, MD, USA, 2004. ISBN 0-8018-8021-1. xvii + 176 pp. LCCN QC16.E5 C35 2004. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/jhu051/2004009048.html;
http://www.loc.gov/catdir/description/jhu051/2004009048.
html;   http://www.loc.gov/catdir/toc/ecip0417/2004009048.
html.

Calaprice:2005:NQE

[Cal05a] Alice Calaprice, editor. The new quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12074-9 (hardcover), 0-691-12075-7 (paperback). xxxvii + 407 pp.

LCCN QC16.E5 A25 2005. URL http://press.princeton.

edu/titles/7921.html. Foreword by Freeman Dyson.

Calder:2005:EUL

[Cal05b] Nigel Calder. Einstein’s universe: the layperson’s guide. Penguin, London, UK and New York, NY, USA, 2005. ISBN 0-14-102056-3 (paperback). x + 190 pp. LCCN QC173.55 .C34 2005.

Calle:2005:ED

[Cal05c] Carlos I. Calle. Einstein for dummies. For dummies. Wiley, New York, NY, USA, 2005. ISBN 0-7645-8348-4 (paperback). xx + 364 pp. LCCN QC16.E5 C355 2005.

Calaprice:2010:UQE

[Cal10] Alice Calaprice, editor. The ultimate quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13817-6. xxviii + 578 pp. LCCN QC16.E5 A25 2010. URL

http://press.princeton.edu/titles/9268.html.

Calaprice:2013:UQE

[Cal13] Alice Calaprice, editor. The ultimate quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-16014-7. xxviii + 578 pp. LCCN QC16.E5 A25 2013. URL

http://press.princeton.edu/titles/10176.html.

Calvin:2017:BCF

[Cal17] Scott Calvin. Beyond Curie: four women in physics and their remarkable discoveries, 1903 to 1963. IOP concise physics. Morgan and Claypool Publishers, San Rafael, CA, 94903, USA, 2017. ISBN 1-68174-644-1 (print), 1-68174-645-X (ebook), 1-68174-647-6 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC774.A2 C353 2017. URL http://

iopscience.iop.org/book/978-1-6817-4645-6.

Campbell:1918:CFS

[Cam18] W. W. Campbell. Clouds fall away for solar eclipse: Coincidence called remarkable by Prof. Campbell, who describes phenomena. New York Times, ??(??):7, June 10, 1918. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/100213850/13EDE115846CF57707.

Campa:1988:EEE

[Cam88] Riccardo Campa. El estupor de Epicuro: ensayo sobre Erwin Schrödinger. (Spanish) [The stupor of Epicurus: test of Erwin Schrödinger], volume 7 of Alianza ensayo. Alianza, Buenos Aires, Argentina, 1988. ISBN 950-40-0036-3. 205 pp. LCCN QC16.S265 C35 1988.

Campbell:1999:RSS

[Cam99] John Campbell. Rutherford: Scientist Supreme. AAS Publications, Christchurch, New Zealand, 1999. ISBN 0-473-05700-X (hardcover). xvi + 516 + 32 + 16 pp. LCCN QC16.R8 C35 1999.

Campbell:2000:RNP

[Cam00] John Campbell. Rutherford and the Nobel Prize. Physics World, 13(11):21, November 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http:/

/stacks.iop.org/2058-7058/13/i=11/a=22.

Camilleri:2009:HIQ

[Cam09a] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp. LCCN QC16.H395 .C34 2009. URL http:

//www.loc.gov/catdir/toc/ecip0825/2008034113.html.

Campanario:2009:RRN

[Cam09b] Juan Miguel Campanario. Rejecting and resisting Nobel class discoveries: accounts by Nobel laureates. Scientometrics, 81(2): 549–565, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-008-2141-5.

Canaday:2000:NML

[Can00] John Canaday. The Nuclear Muse: Literature, Physics, and the First Atomic Bombs. University of Wisconsin Press, Madison, WI, USA, 2000. ISBN 0-299-16850-6 (hardcover), 0-299-16854-9 (paperback). xviii + 310 pp. LCCN QC791.96 .C36 2000. URL

http://www.e-streams.com/es0409/es0409_1493.htm.

Canales:2005:EBE

[Can05] Jimena Canales. Einstein, Bergson, and the experiment that failed: Intellectual cooperation at the League of Nations. MLN, 120(5):1168–1191, 2005. CODEN ???? ISSN 0026-7910 (print), 1080-6598 (electronic). URL https://muse.jhu.edu/

article/193244.

Canales:2015:MRE

[Can15a] Jimena Canales. The media of relativity: Einstein and telecommunications technologies. Technology and Culture, 56(3):610– 645, 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

Canales:2015:PPE

[Can15b] Jimena Canales. The physicist and the philosopher: Einstein and Bergson and the debate that changed our understanding of time. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16534-3 (hardcover), 1-4008-6577-8 (e-book). viii + 479 pp. LCCN BD638 .C326 2015.

Canales:2016:EBP

[Can16a] Jimena Canales. Einstein’s Bergson problem: Communication, consensus and good science. In Cosmological and Psychological Time [DR16], chapter 4, pages 53–72. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346.

LCCN BD638 C67 2016. URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-6_4.

Canales:2016:PPE

[Can16b] Jimena Canales. Physicist and the Philosopher: Einstein, Bergson, and the Debate That Changed Our Understanding of Time. Princeton University Press, Princeton, NJ, USA, 2016. ISBN 0-691-16534-3, 0-691-17317-6, 1-4008-9722-X. viii + 479

pp. LCCN ???? URL http://public.eblib.com/choice/
publicfullrecord.aspx?p=1936462.

Cao:2004:CSN

[Cao04] Cong Cao. Chinese science and the ‘Nobel Prize complex’. Minerva [Washington, DC], 42(2):151–172, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000030020.28625.7e.

Cao:2006:BRD

[Cao06] Tian Yu Cao. Book review: Danian Hu: China and Albert Einstein: The Reception of the Physicist and His Theory in China, 1917–1979. Isis, 97(3):571–573, September 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509984.

Cao:2007:WES

[Cao07] Tian Yu Cao. Will Einstein still be the super-hero of physics history in 2050? In Renn and Gavroglou [RG07], chapter 4, pages 27–32. ISBN 1-4020-5420-3, 1-4020-5419-X. ISSN 0068-0346.

LCCN Q125 .P77 2007eb. URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_4.

Cao:2014:UVS

[Cao14] Cong Cao. The universal values of science and China’s Nobel Prize pursuit. Minerva [Washington, DC], 52(2):141–160, June 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9249-y.

Capek:1971:BEP

[Čap71a] Milič Čapek. Bergson and Einstein. The physical world as extensive becoming. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71b], chapter 33, pages 238–256. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.
1007/978-94-010-3096-0_33.

Capek:1971:BMP

[Čap71b] Milič Čapek. Bergson and Modern Physics: A Reinterpretation and Re-evaluation, volume 7 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1971. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346. xvi + 418 pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-010-3096-0.

Capek:1976:CST

[Čap76] Milič Čapek. The Concepts of Space and Time: Their Structure and Their Development, volume 22 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346. lvii +

570 pp. LCCN BD632. URL https://link.springer.com/
book/10.1007/978-94-010-1727-5.

Cappelletti:2001:DAC

[Cap01] Valentina Cappelletti. Dall’ordine alle cose: saggio su Werner Heisenberg. (Italian) [From Order to Chaos: test on Werner Heisenberg], volume 16, 471 of Lo spoglio dell’occidente; ; Di fronte e attraverso. Jaca Book, Milano, Italia, 2001. ISBN 88-16-40571-6. 234 pp. LCCN QC174.12 .C348 2001. 16.53 EUR.

Capria:2005:PBA

[Cap05] Marco Mamone Capria, editor. Physics before and after Einstein. IOS Press, Amsterdam, The Netherlands, 2005. ISBN 1-58603-462-6 (hardcover), 1-60129-087-X, 1-4294-0238-5 (ebook). vii + 324 pp. LCCN QC21.3 .P3 200.

Carus:1913:PRL

[Car13] Paul Carus. The Principle of Relativity in the light of the philosophy of science. Open Court Pub. Co., Chicago, IL, USA, 1913. 105 pp. LCCN B945.C22 P92. With an appendix containing a letter from James Bradley (1693?–1762) on the motion of the fixed stars, 1727.

Carr:1922:DSP

[Car22a] H. Wildon Carr. Durée et simultanéité: À propos de la théorie d’Einstein. Nature, 110(2763):503–505, October 14, 1922. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v110/n2763/
pdf/110503a0.pdf.

Carr:1922:EP

[Car22b] H. Wildon Carr. Einstein’s paradox. Nature, 110(2768):669, November 18, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v110/n2768/pdf/110669a0.pdf.

Carr:1922:TMO

[Car22c] Herbert Wildon Carr. A theory of monads: outlines of the philosophy of the Principle of Relativity. Macmillan Publishing Company, New York, NY, USA, 1922. viii + 351 pp. LCCN DB221 .C28.

Carr:1923:DSP

[Car23] H. Wildon Carr. Durée et simultanéité: á propos de la théorie d’Einstein. Nature, 112(2812):426–428, September 22, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v112/n2812/pdf/112426a0.pdf.

Carr:1924:SFE

[Car24] H. Wildon Carr. Substance and function and Einstein’s Theory of Relativity. Nature, 114(2858):187–188, August 9, 1924. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v114/n2858/
pdf/114187a0.pdf.

Carazza:1982:GE

[Car82a] Bruno Carazza. Da Galileo a Einstein. (Italian) [From Galileo to Einstein]. Edizioni Zara, Parma, Italia, 1982. ???? pp. LCCN QC7 .C237 1982.

Carazza:1982:GEI

[Car82b] Bruno Carazza. Da Galileo a Einstein. (Italian) [From Galileo to Einstein]. Zara, Parma, Italy, 1982. ???? pp. LCCN QC7 .C237 1982.

Cartwright:1989:BED

[Car89] Nancy Cartwright. The Born–Einstein debate: Where application and explanation separate. Synthese, 81(3):271–282, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/
10.1007/BF00869317.

Carmeli:2002:CSR

[Car02] Moshe Carmeli. Cosmological Special Relativity: the large scale structure of space, time and velocity. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2002. ISBN 981-02-4936-5. xvi + 207 pp. LCCN

QB981 .C363 2002. URL http://www.loc.gov/catdir/toc/
fy031/2002025893.html.

Carr:2004:UCC

[Car04a] Alan B. Carr. The University of California contract to operate the Los Alamos Laboratory, 1942–1947: A documentary history. Technical Report LA-UR-04-789, Los Alamos National Laboratory, Los Alamos, NM, USA, 2004. ???? pp.

Carroll:2004:SGI

[Car04b] Sean M. Carroll. Spacetime and Geometry: an Introduction to General Relativity. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-8053-8732-3. xiv + 513 pp. LCCN QC173.6 .C377 2004.

Cartan:2007:DCM

[Car07a] Elie Cartan. The dynamics of continuous media and the notion of an affine connection on space–time. In Renn et al.

[RSS^{+}07b], pages 1107–1129. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL http://

d-nb.info/985765933/34;   http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html;
 https://link.springer.com/book/10.1007/978-1-4020-

4000-9. Originally published as “Sur les variétés à connexion affine et la théorie de la relativité généralisée”, Annales Scientifiques de l’École Normal Supérieure, 325–412 (1923).

Cartwright:2007:GPG

[Car07b] Jon Cartwright. Gravity probe gives Einstein costly test. Physics World, 20(5):8, May 2007. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/20/5/phwv20i5a12.
pdf.

Carr:2008:FPR

[Car08] Alan B. Carr. The forgotten physicist: Robert F. Bacher, 1905– 2004, volume Monograph 6 of The Los Alamos story. Los Alamos Historical Society, Los Alamos, N.M., 2008. ISBN 0-941232-36-0 (paperback). viii + 75 pp. LCCN QC16.B23

C37 2008. URL http://www.loc.gov/catdir/toc/fy0805/
2008000961.html.

Carson:2010:HAA

[Car10] Cathryn Carson. Heisenberg in the atomic age: science and the public sphere. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 2010. ISBN 0-521-82170-3 (hardcover). xvi + 541 + 8 pp. LCCN QC16.H45 C37 2010.

Carter:2012:HWN

[Car12] David Carter. How to win the Nobel Prize in literature: a handbook for the would-be laureate. Hesperus Press, London, UK, 2012. ISBN 1-84391-374-7, 1-78094-040-8 (e-book). LCCN PN171.P75 C378 2012. URL http:

//www.pensu.eblib.com/EBLWeb/patron/?target=patron\%
26extendedid=P_1093665_0.

Carter:2015:HWN

[Car15] David Carter. How to Win the Nobel Prize in Literature: a Handbook for the Would-be Laureate. Hesperus Press Ltd, New York, NY, USA, 2015. ISBN 1-84391-374-7, 1-78094-040-8 (ebook). 215 pp. LCCN PN171.P75. URL http://gbv.eblib.

com/patron/FullRecord.aspx?p=1093665.

Caracheo:2017:MTM

[Car17] Armando Caracheo. The measurement of time: Mann and Einstein’s thought experiments. Configurations: a Journal of Literature, Science, and Technology, 25(1):29–55, 2017. CODEN ???? ISSN 1063-1801 (print), 1080-6520 (electronic). URL

https://muse.jhu.edu/article/646281.

Cassirer:1921:ERE

[Cas21] Ernst Cassirer. Zur Einstein’schen Relativitätstheorie : Erkenntnistheoretische Betrachtungen. (German) [Einstein’s Relativity Theory: Epistemological Considerations]. Bruno Cassier, Berlin, Germany, 1921. 134 pp. LCCN ????

Cassirer:1923:SFE

[Cas23] Ernst Cassirer. Substance and function, and Einstein’s theory of relativity. The Open Court Publishing Company, Chicago, IL, USA, 1923. 2 + ii–xii + 465 pp. LCCN BD221 .C285. Authorized translation by William Curtis Swabey and Marie Collins Swabey.

Cassel:1974:EE

[Cas74] Hans M. Cassel. Encounter with Einstein. Physics Today, 27 (7):15, July 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Casertano:1975:FFE

[Cas75] Giovanni Casertano. Fisica e filosofia: Einstein, Infeld, de Broglie, Heisenberg, Planck. (Italian) [Physics and philosophy: Einstein, Infeld, de Broglie, Heisenberg, Planck]. Edizione Il Tripode, Napoli, Italia, 1975. 151 pp. LCCN QC6 .F55.

Cassidy:1976:WHC

[Cas76] David C. Cassidy. Werner Heisenberg and the crisis in quantum theory, 1920–1925. Thesis (Ph.D.), Purdue University, West Lafayette, IN, USA, 1976. xiv + 520 pp.

Cassidy:1979:HFC

[Cas79] David C. Cassidy. Heisenberg’s first core model of the atom: The formation of a professional style. Historical Studies in the Physical Sciences, 10(??):187–224, ???? 1979. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757390.

Cassidy:1986:RUH

[Cas86] David Cassidy. Review: Understanding the history of Special Relativity. Historical Studies in the Physical and Biological Sciences, 16(1):177–188, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757561.

Cassidy:1992:ULS

[Cas92] David C. Cassidy. Uncertainty: the life and science of Werner Heisenberg. W. H. Freeman and Company, New York, NY, USA, 1992. ISBN 0-7167-2243-7. xii + 669 + 16 pp. LCCN QC16.H35 C37 1992; QC 16 .H35C37 1992. US$29.95.

Cassidy:1994:PLAa

[Cas94a] D. Cassidy. The private lives of Albert Einstein — R. Highfield, P. Carter. Science, 263(5149):997–998, February 18, 1994. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Cassidy:1994:PLAb

[Cas94b] D. Cassidy. The private lives of Albert Einstein (vol 263, pg 997, 1994). Science, 263(5151):1303, March 4, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Cassidy:1995:EOW

[Cas95] David C. Cassidy. Einstein and our world. The control of nature. Humanities Press, Atlantic Highlands, NJ, USA, 1995. ISBN 0-391-03876-1 (hardcover); 0-391-03875-3 (paperback). viii + 100 pp. LCCN QC16.E5 C37 1995.

Cassirer:2003:SFE

[Cas03a] Ernst Cassirer. Substance and function and Einstein’s theory of relativity. Dover phoenix editions. Dover, New York, NY, USA, 2003. ISBN 0-486-49547-7. xii + 465 pp. LCCN QC173.55 .C3813 2003. URL http://www.loc.gov/

catdir/enhancements/fy0615/2003067492-d.html. Autho-

rized translation by William Curtis Swabey and Marie Collins Swabey.

Casti:2003:OTP

[Cas03b] J. L. Casti. The one true platonic heaven: a scientific fiction on the limits of knowledge. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08547-0 (hardcover), 0-309-09510-7 (paperback). xviii + 160 pp. LCCN Q175 .C4339 2003. URL

http://www.loc.gov/catdir/toc/fy037/2003002279.html;
 http://www.nap.edu/catalog.php?record_id=10533.

Cassidy:2004:EOW

[Cas04] David C. Cassidy. Einstein and our world. Control of nature. Humanity Books, Amherst, NY, USA, second edition, 2004. ISBN 1-59102-256-8 (paperback). 162 pp. LCCN QC16.E5

C37 2004. URL http://www.loc.gov/catdir/toc/ecip0416/
2004007648.html.

Cassidy:2005:EQH

[Cas05a] D. C. Cassidy. Einstein and the quantum hypothesis. Annalen der Physik (Berlin), 14(1):15–22, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200410125.

Cassidy:2005:BRB

[Cas05b] David C. Cassidy. Book review: Born unto trouble: The End of the Certain World: The Life and Science of Max Born: The Nobel Physicist Who Ignited the Quantum Revolution, by Nancy Thorndike Greenspan. American Scientist, 93(4):372–374, July/ August 2005. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
pdfplus/27858620.pdf.

Cassidy:2005:JRO

[Cas05c] David C. Cassidy. J. Robert Oppenheimer and the American century. Pi Press, New York, NY, USA, 2005. ISBN 0-13-147996-2. xviii + 462 + 16 pp. LCCN QC16.O62 C37 2005. US$27.95.

Cassidy:2007:OFP

[Cas07] David C. Cassidy. Oppenheimer’s first paper: Molecular band spectra and a professional style. Historical Studies in the Physical and Biological Sciences, 37(2):247–270, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Castellari:2014:FMF

[Cas14] Marco Castellari, editor. Formula e metafora: figure di scienziati nelle letterature e culture contemporanee. (Italian) [Formula and metaphor: figures of scientists in the literature and contemporary cultures], volume 8 of Di/segni. Ledizioni, Milano, Italy, 2014. ISBN 88-6705-207-1. LCCN ????

Cassidy:2017:FHG

[Cas17] David C. In Cassidy. Farm Hall and the German Atomic Project of World War II: a dramatic history. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-59577-6 (print), 3-319-59578-4 (e-book). xiv + 125 + 32 pp. LCCN PS3553.A87 F37 2017.

Cat:1998:PDU

[Cat98] Jordi Cat. The physicists’ debates on unification in physics at the end of the 20th Century. Historical Studies in the Physical and Biological Sciences, 28(2):253–300, ???? 1998. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cathcart:2004:FCH

[Cat04] Brian Cathcart. The fly in the cathedral: how a group of Cambridge scientists won the international race to split the atom. Farrar, Strauss, and Giroux, New York, NY, USA, 2004. ISBN 0-374-15716-2 (hardcover). xii + 308 + 4 pp. LCCN Q141 .C2515 2004.

Caudill:1991:THA

[Cau91] S. Caudill. Trying to harness atomic energy, 1946–1951 — Albert Einstein publicity campaign for world government. Journalism Quarterly, 68(1–2):253–262, Spring/Summer 1991. CO-DEN JOQUAX. ISSN 0196-3031.

Cassidy:1984:WHB

[CB84] David C. Cassidy and Martha Baker. Werner Heisenberg: a bibliography of his writings, volume 9 of Berkeley papers in history of science, 0145-0379. Office for History of Science and Technology, University of California, Berkeley, Berkeley, CA, USA, 1984. ISBN 0-918102-10-3 (paperback). vi + 153 pp. LCCN Z8395.513 .C38 1984; QC16.H395.

Crease:1999:MPE

[CB99] Robert P. Crease and Stony Brook. The Manhattan Project: an enduring legacy. Physics World, 12(12):59–63, December 1999. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058

(electronic). URL http://stacks.iop.org/2058-7058/12/i=
12/a=21.

Carriere:2006:PME

[CB06] Jean-Claude Carrière and John Brownjohn. Please, Mr. Einstein. Harcourt, Inc., Orlando, FL, USA, 2006. ISBN 0-15-101422-1. 186 pp. LCCN PQ2663.A78 E3613 2006.

Choquet-Bruhat:2009:GRE

[CB09] Yvonne Choquet-Bruhat. General Relativity and the Einstein equations. Oxford mathematical monographs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-923072-2 (hardcover), 0-19-155226-7 (e-book). xxiv + 785 pp. LCCN QC173.6 .C474 2009.

Choquet-Bruhat:2015:IGR

[CB15] Yvonne Choquet-Bruhat. Introduction to General Relativity, Black Holes and Cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-966645-8 (hardcover), 0-19-966646-6 (paperback), 0-19-164452-8 (e-book). xx + 279 pp. LCCN QC173.6 .C46 2015. With a foreword by Thibault Damour.

Chandrasekhar:1967:PGT

[CC67] S. Chandrasekhar and G. Contopoulos. On a post-Galilean transformation appropriate to the post-Newtonian theory of Einstein, Infeld and Hoffmann. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 298 (1453):123–141, April 4, 1967. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic). URL http://www.jstor.

org/stable/2415974.

Church:1974:WAW

[CC74] Fermor S. Church and Peggy Pond Church, editors. When Los Alamos was a ranch school: historical profile. Los Alamos Historical Society, Los Alamos, NM, USA, 1974. 60 pp. LCCN LD7501.L7298 C48.

Carmeli:1984:ERG

[CC84] M. Carmeli and Ch. Charach. The Einstein–Rosen gravitational waves and cosmology. Foundations of Physics, 14(10):963–986, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01889248.

Cohen:1996:EHB

[CC96] Paul Cohen and Brenda Cohen. The Einstein House in Berne, Switzerland: Visiting the home of one of the century’s great minds. Journal of College Science Teaching, 25(6):440–441, May 1996. CODEN JSCTBN. ISSN 0047-231X (print), 1943-4898

(electronic). URL https://www.jstor.org/stable/42990968.

Church:1998:WAW

[CC98] Fermor S. Church and Peggy Pond Church, editors. When Los Alamos was a ranch school: historical profile. Los Alamos Historical Society, Los Alamos, NM, USA, second edition, 1998. ISBN 0-941232-18-2 (paperback). viii + 67 pp. LCCN LD7501.L7298 C48 1998.

Cori:1952:NNP

[CCD^{+}52] Carl F. Cori, Gerti T. Cori, C. J. Davisson, Joseph Erlanger, Philip S. Hench, H. J. Muller, William P. Murphy, Harold C. Urey, and George H. Whipple. Nine Nobel Prize-winners address Joliot-Curie. Bulletin of the Atomic Scientists, 8(5):131, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Corsi:1993:BGP

[CCG93] Giovanna Corsi, Maria Luisa Dalla Chiara, and Gian Carlo Ghirardi. Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science, volume 140 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346. xi + 329 pp. LCCN ???? URL https://link.

springer.com/book/10.1007/978-94-011-2496-6.

Charpa:2007:JSG

[CD07] Ulrich Charpa and Ute Deichmann. Jewish scientists as geniuses and epigones: Scientific practice and attitudes towards Albert Einstein, Ferdinand Cohn, Richard Goldschmidt. Studia Rosenthaliana, 40(??):75–108, ???? 2007. ISSN 0039-3347.

Chen:2023:EEA

[CD23] Yifan Chen and Jingda Ding. Exploitation and exploration: an analysis of the research pattern of Nobel laureates in physics. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000536.

Cordella:2001:EFA

[CDS01] Francesco Cordella, Alberto De Gregorio, and Fabio Sebastiani. Enrico Fermi: gli anni italiani. (Italian) [Enrico Fermi: The Italian Years]. Nuova biblioteca di cultura. Editori riuniti, Roma, Italy, 2001. ISBN 88-359-5097-X. 332 pp. LCCN QC16.F46 C67 2001.

Cartan:1979:LPA

[CE79a] Élie Cartan and Albert Einstein. Lettres sur le parallélisme absolu, 1929–1932. Textes originaux, traduction anglaise par Jules Leroy et Jim Ritter. Publiés par Robert Debever. (French) [Letters on absolute parallelism, 1929–1932. Original texts and English translation by Jules Leroy and Jim Ritter. Published by Robert Debever]. Princeton University Press and L’Académie Royale de Belgique, Princeton, NJ, USA and Bruxelles, Belgique, 1979. xvii + 235 pp.

Cartan:1979:LAP

[CE79b] Élie Joseph Cartan and Albert Einstein. Letters on absolute parallelism, 1929–1932. Princeton University Press, Princeton, NJ, USA, 1979. ISBN 0-691-08229-4. xvii + 235 + 2 + (loose erratum) pp. LCCN QC15 .C27 1979. Original text with English translation by Jules Leroy and Jim Ritter. Edited by Robert Debever.

Ceapa:1995:PGE

[Cea95] A. C. V. Ceapa. Physical grounds of Einstein’s theory of relativity. A. Ceapa, Bucharest, Romania, 1995. ISBN 973-97340-1-4. 96 pp. LCCN QC173.55 .C43 1995.

Ceapa:1997:PGE

[Cea97] A. C. V. Ceapa. Physical grounds of Einstein’s theory of relativity: roots of the falsification of 20th century physics. A. C. V. Ceapa, Bucharest, Romania, second edition, 1997. ISBN 973-9318-06-1. xii + 128 pp. LCCN MLCM 99/02567 (Q).

Ceapa:1998:PGE

[Cea98] A. C. V. Ceapa. Physical grounds of Einstein’s theory of relativity: roots of the falsification of 20th century physics. A. C. V. Ceapa, Bucharest, Romania, third edition, 1998. ISBN 973-9318-06-1. xviii + 137 pp. LCCN QC173.55 .C43 1998.

Cecil:1986:EB

[Cec86] Richard Cecil. Einstein’s brain. University of Utah Press poetry series. University of Utah Press, Salt Lake City, UT, USA, 1986. ISBN 0-87480-255-5 (paperback). 78 pp. LCCN PS3553.E32 E36 1986. US$7.95.

Ceccarelli:2001:SSR

[Cec01] Leah Ceccarelli. Shaping science with rhetoric: the cases of Dobzhansky, Schrödinger, and Wilson. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-09906-7 (cloth), 0-226-09907-5 (paperback). xi + 204 pp. LCCN QH303.6 .C433 2001. URL http://www.loc.

gov/catdir/bios/uchi051/00012179.html;   http://www.
loc.gov/catdir/description/uchi051/00012179.html;
http://www.loc.gov/catdir/toc/uchi051/00012179.html.

Cerroni:2000:CIC

[Cer00] Andrea Cerroni. Covariance/invariance: a cognitive heuristic in Einstein’s relativity theory formation. Foundations of Science, 5(2):209–224, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011304431615.

CervantesCastro:2011:AEH

[Cer11] Jorge Cervantes Castro. Albert Einstein and his abdominal aortic aneurysm. Gaceta Medica de Mexico, 147(1):74–76, January/ February 2011. ISSN 0016-3813.

Chant:1980:DEH

[CF80] Colin Chant and John Fauvel, editors. Darwin to Einstein: historical studies on science and belief. Longman, London, UK, 1980. ISBN 0-582-49157-6 (paperback), 0-582-49156-8. x + 335 pp. LCCN Q174.8 .D37.

Crawford:1982:PPC

[CF82] Elisabeth Crawford and Robert Marc Friedman. The Prizes in physics and chemistry in the context of Swedish science: a working paper. In Bernhard et al. [BCS82], pages 311–331. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500268.

Craig:1998:REP

[CF98] Edward Craig and Luciano Floridi, editors. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN

B51. URL http://www.rep.routledge.com/.

Cox:2009:WDW

[CF09] Brian Cox and J. R. (Jeffrey Robert) Forshaw. Why does E = mc^{2}: (and why should we care?). Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 2009. ISBN 1-4416-2111-3 (e-book), 0-7867-5216-5 (e-book), 0-306-81758-6. xiii + 249 pp. LCCN QC173.6 .C68 2009.

Capozziello:2011:BEG

[CF11] Salvatore Capozziello and Valerio Faraoni. Beyond Einstein gravity: a survey of gravitational theories for cosmology and astrophysics, volume 170 of Fundamental theories of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-0164-0. ISSN 0168-1222. xix + 428 pp. LCCN QB335 .C37 2011; QC178 .C176 2011.

Cox:2012:QUW

[CF12] Brian Cox and Jeff Forshaw. The Quantum Universe: (And Why Anything That Can Happen, Does). Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 2012. ISBN 0-306-81964-3, 0-306-82060-9 (e-book). 255 (est.) pp. LCCN QC174.12

.C68 2012. URL http://www.economist.com/node/21536548.

Cohen:1976:EMI

[CFW76] R. S. Cohen, P. K. Feyerabend, and M. W. Wartofsky. Essays in Memory of Imre Lakatos, volume 39 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346. xii +

768 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-010-1451-9.

Cohen:1990:RAA

[CG90] J. R. Cohen and L. M. Graver. The ruptured abdominal aorticaneurysm of Albert Einstein. Surgery Gynecology & Obstetrics, 170(5):455–458, May 1990. ISSN 0039-6087.

Charpak:2001:MMT

[CG01] Georges Charpak and Richard L. Garwin. Megawatts and megatons: a turning point in the nuclear age? Alfred A. Knopf, New York, NY, USA, 2001. ISBN 0-375-40394-9. xvi + 412 pp. LCCN QC792 .C4713 2001. URL http://www.loc.

gov/catdir/bios/random053/2001029863.html;   http://
www.loc.gov/catdir/description/random047/2001029863.
html;   http://www.loc.gov/catdir/samples/random043/
2001029863.html;   http://www.loc.gov/catdir/toc/fy037/
2001029863.html.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Carlsson:2013:TII

[CGG^{+}13] Kristian Carlsson, Ángela García, Azita Ghahreman, Anisur Rahman, Tomas Tranströmer, et al. Tranströmer international: an intercontinental perspective on the poetry of Nobel Laureate Tomas Tranströmer. Dracopis Press, Berkeley, CA, USA, 2013. ISBN 91-87341-01-8. ???? pp. LCCN ???? Translated by Nimao Ahmed Bulaleh from the original Swedish.

Castagnetti:1997:FDE

[CGR^{+}97] Giuseppe Castagnetti, Hubert Goenner, Jürgen Renn, Tilman Sauer, and Britta Scheideler. Foundation in disarray: essays on Einstein’s science and politics in the Berlin years. Preprint 63, Max-Planck-Institut für Wissenschaftsgeschichte, Berlin, Ger-

many, May 1997. iv + 80 pp. URL http://www.mpiwg-berlin.
mpg.de/en/forschung/Preprints/P63.PDF.

Chiu:2011:CAM

[CGT11] M.-H. (Mei-Hung) Chiu, Penny J. Gilmer, and David F. Treagust. Celebrating the 100th anniversary of Madame Marie Skłodowska Curie’s Nobel Prize in Chemistry. Sense Publishers, Rotterdam, The Netherlands, 2011. ISBN 94-6091-717-8, 94-6091-718-6, 94-6091-719-4 (e-book). LCCN QD22.C8 C45 2011.

Coley:1980:DEP

[CH80] Noel George Coley and Vance M. D. Hall, editors. Darwin to Einstein: primary sources on science and belief. Longman in association with the Open University Press, Harlow, Essex, UK, 1980. ISBN 0-582-49158-4, 0-582-49159-2 (paperback). ix + 358 pp. LCCN BD215 .D37.

Carter:1998:ECZ

[CH98] Paul Carter and Roger Highfield. Ejnstejn: castnaja zizn’. (Russian) [The private lives of Albert Einstein]. Zacharov, Moskva, Russia, 1998. ISBN 5-8159-0009-5. 367 + 16 pp. LCCN ???? Russian translation of [HC93].

Carson:2005:ROC

[CH05] Cathryn Carson and David A. Hollinger, editors. Reappraising Oppenheimer: centennial studies and reflections, volume 21 of Berkeley papers in history of science. Office for History of Science and Technology, Berkeley, CA, USA, 2005. ISBN 0-9672617-3-2. xii + 413 pp. LCCN QC16.O62 R47 2005.

Chant:1923:EDP

[Cha23] C. A. Chant. Einstein displacement on the plates taken by the Canadian party at the Australian eclipse. Science, 57(1477): 469, April 20, 1923. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/57/1477/469.2.full.pdf.

Chadwick:1932:EN

[Cha32] James Chadwick. The existence of a neutron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):692–708, June 1, 1932. ISSN 0950-1207. URL http://www.jstor.org/ stable/95816. This paper reports the discovery of the neutron, for which Chadwick received the Nobel Prize in Physics in 1935.

Chalmers:1949:HRCa

[Cha49a] T. W. (Thomas Wightman) Chalmers. Historic researches: chapters in the history of physical and chemical discovery. Morgan Brothers Publishers Ltd., London, UK, 1949. 288 pp. LCCN QC7 .C47.

Chandrasekhar:1949:BMD

[Cha49b] S. Chandrasekhar. Brownian motion, dynamical friction, and stellar dynamics. Reviews of Modern Physics, 21(3):383–388, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.383;   http://rmp.aps.org/

abstract/RMP/v21/i3/p383_1. Special issue in honor of Albert Einstein on his 70th birthday.

Chandrasekhar:1972:DEF

[Cha72] S. Chandrasekhar. On the “derivation” of Einstein’s field equations. American Journal of Physics, 40(2):224–234, 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/40/2/10.1119/1.1986496. See comment

[Kre72].

Chambers:1974:TSA

[Cha74] Marjorie Bell Chambers. Technically Sweet Los Alamos: The Development of a Federally Sponsored Scientific Community. Ph.D. dissertation, Department of History, University of New Mexico, Albuquerque, NM, USA, May 1974. x + 391 pp. URL

http://search.proquest.com/docview/302670321.

Chandrasekhar:1975:SFM

[Cha75] S. Chandrasekhar. Of some famous men: Verifying the Theory of Relativity. Bulletin of the Atomic Scientists, 31(6):17–22, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chandrasekhar:1976:VTR

[Cha76] S. Chandrasekhar, F.R.S. Verifying the theory of relativity. Notes and Records of the Royal Society of London, 30(2):249– 260, January 1976. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
531756.

Chandrasekhar:1978:EGR

[Cha78] S. Chandrasekhar. Einstein and General Relativity: historical perspectives. Report LASL 78-91, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1978. v + 16 pp. Seventh Oppenheimer memorial lecture, August 17, 1978, Los Alamos New Mexico.

Chapman:1979:AE

[Cha79] P. Chapman. Albert Einstein. Electronics and Power, 25(3): 169–171, ???? 1979. CODEN ELPWAQ. ISSN 0013-5127.

Chaffee:1980:DGL

[Cha80] F. H. Chaffee, Jr. The discovery of a gravitational lens. Scientific American, 243(5):60–68, November 1980. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chandrasekhar:1983:EMD

[Cha83] S. (Subrahmanyan) Chandrasekhar. Eddington, the most distinguished astrophysicist of his time. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25746-8. 64 pp. LCCN QB36.E33 C48 1983. URL http://www.loc.gov/

catdir/enhancements/fy0914/84123568-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0914/84123568-t.
html.

Chandrasekhar:1987:TBA

[Cha87] S. (Subrahmanyan) Chandrasekhar. Truth and beauty: aesthetics and motivations in science. University of Chicago Press, Chicago, IL, USA and London, UK, 1987. ISBN 0-226-10086-3, 0-226-10087-1 (paperback), 0-226-16277-X (e-book), 1-306-15807-9 (e-book). x + 170 pp. LCCN Q175 .C453 1987.

Chaisson:1988:RSR

[Cha88] Eric Chaisson. Relatively speaking: Relativity, black holes, and the fate of the Universe. W. W. Norton & Co., New York, NY, USA, 1988. ISBN 0-393-02536-5. 254 pp. LCCN QC173.55 .C46

Chandrasekhar:1998:MTB

[Cha98] S. (Subrahmanyan) Chandrasekhar. The mathematical theory of black holes. Oxford classic texts in the physical sciences. Clarendon Press, Oxford, UK, 1998. ISBN 0-19-850370-9. xxi + 646 pp. LCCN QB843.B55 C48 1998. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0912.53053.

Chandrasekhar:2001:EGT

[Cha01] Subrahmanyan Chandrasekhar. Einstein’s General Theory of Relativity and cosmology. In Hoiberg [Hoi01], pages 198–231. ISBN 81-88237-00-0. LCCN AS911.N9 A15 2001.

Chalmers:2005:FPS

[Cha05a] Matthew Chalmers. Five papers that shook the world. Physics World, 18(1):16–17, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworld.com/cws/article/print/2005/jan/05/five-
papers-that-shook-the-world;   http://physicsworldarchive.
iop.org/full/pwa-pdf/18/1/phwv18i1a24.pdf.

Chapman:2005:OSA

[Cha05b] Tim Chapman. The other side of Albert Einstein. Physics World, 18(1):52, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a35.
pdf.

Charles:2005:MMR

[Cha05c] Daniel Charles. Master mind: the rise and fall of Fritz Haber, the Nobel laureate who launched the age of chemical warfare. Ecco, New York, NY, USA, 2005. ISBN 0-06-056272-2. xvii + 313 pp. LCCN QD22.H15 C48 2005. URL http://www.loc.

gov/catdir/enhancements/fy0910/2004057532-b.html;
http://www.loc.gov/catdir/enhancements/fy0910/2004057532-
d.html.

Chalmers:2009:SAP

[Cha09] A. F. (Alan Francis) Chalmers. The scientist’s atom and the philosopher’s stone: how science succeeded and philosophy failed to gain knowledge of atoms, volume 279 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2361-5. ISSN 0068-0346. xii + 287 pp. LCCN QD461 .C375 2009.

Chan:2011:MAB

[Cha11] Keng Wai Chan. Measurement and analysis of bubble pump and Einstein–Szilard single pressure absorption refrigeration system. Thesis (D.Phil.), Mathematical, Physical & Life Sciences Division, Oxford University, Oxford, UK, 2011. ix + 237 pp.

URL http://ora.ox.ac.uk/objects/ora:5962;   http:/
/ora.ox.ac.uk/objects/uuid:f10706c5-6952-4c50-9aef-
065f1627a19b;   http://search.proquest.com/docview/
1414910237.

Chavalarias:2017:WWS

[Cha17] David Chavalarias. What’s wrong with science?: Modeling the collective discovery processes with the Nobel game. Scientometrics, 110(1):481–503, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11192-
016-2109-9.

Chevalier:1966:OSF

[Che66] Haakon Chevalier. Oppenheimer: the story of a friendship. Andre Deutsch, London, UK, 1966. xvi + 219 pp. LCCN QC16.O62 C5 1966.

Chen:2000:BEC

[Che00] Josette Chen. Bose–Einstein condensates: The next big thing. Nature, 407(6804):575, October 5, 2000. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v407/n6804/
full/407575b0.html.

Cheng:2010:RGC

[Che10] Ta-Pei Cheng. Relativity, gravitation, and cosmology: a basic introduction. Oxford master series in particle physics, astrophysics, and cosmology; Oxford master series in physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2010. ISBN 0-19-957363-8 (hardcover), 0-19-957364-6 (paperback). xiii + 435 pp. LCCN QC173.6 .C4724 2010.

Cheng:2013:EPA

[Che13] Ta-Pei Cheng. Einstein’s physics: atoms, quanta, and relativity derived, explained, and appraised. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966991-0. xx + 350 pp. LCCN QC16.E5 C455 2013. URL http://www.loc.

gov/catdir/enhancements/fy1403/2012554428-b.html;
http://www.loc.gov/catdir/enhancements/fy1403/2012554428-
d.html;   http://www.loc.gov/catdir/enhancements/fy1403/
2012554428-t.html.

Chiu:1994:FIS

[Chi94] Charles S. Chiu. Frauen im Schatten. (German) [Women in the shade]. J and V, Wien, Austria, 1994. ISBN 3-224-17669-5. 228 pp. LCCN CT3310 .C45 1994.

Chiu:2008:WSM

[Chi08] Ch. S. Chiu. Women in the shadows: Mileva Einstein-Marić, Margarete Jeanne Trakl, Lise Meitner, Milena Jesenská, and Margarete Schütte-Lihotzky, volume 40 of Austrian culture. Peter Lang, New York, NY, USA, 2008. ISBN 0-8204-8856-9 (hardcover). viii + 203 pp. LCCN CT3310 .C4513 2008. URL http:

//www.loc.gov/catdir/toc/ecip0819/2008023764.html. In-

troduction and translation by Edith Borchardt.

Cho:2003:EQG

[Cho03] Adrian Cho. Einstein 1, quantum gravity 0. Science, 301(5637): 1169–1171, August 29, 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/301/5637/1169.1.full.pdf.

Cho:2005:DST

[Cho05a] Adrian Cho. Doubly special, twice as controversial. Science, 307 (5711):867, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Cho:2005:SRR

[Cho05b] Adrian Cho. Special Relativity reconsidered. Science, 307 (5711):866–868, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Cho:2006:PGC

[Cho06] Adrian Cho. Pulsars’ gyrations confirm Einstein’s theory. Science, 313(5793):1556–1557, September 15, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/313/5793/1556.
2.full.pdf.

Chow:2008:GBH

[Cho08] Tai L. Chow. Gravity, black holes, and the very early universe: an introduction to general relativity and cosmology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-73629-8, 0-387-73631-X (e-book). xv + 280 pp. LCCN QC178 .C29 2008. URL http://www.loc.

gov/catdir/toc/fy0802/2007936678.html.

Chodos:2010:MPH

[Cho10a] Alan Chodos. This month in physics history: November 11, 1930: Patent granted for Einstein–Szilard refrigerator. APS News, ??(??):??, November 2010. URL http://www.aps.

org/publications/apsnews/201011/physicshistory.cfm;
 http://www.aps.org/publications/capitolhillquarterly/
201012/physicshistory.cfm.

Choudhuri:2010:AP

[Cho10b] Arnab Rai Choudhuri. Astrophysics for physicists. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-81553-3 (hardcover). xviii + 471 pp. LCCN QB461 .C535 2010.

URL http://www.cambridge.org/gb/academic/subjects/
astronomy/astrophysics/astrophysics-physicists.

Cho:2016:GWE

[Cho16a] Adrian Cho. Gravitational waves, Einstein’s ripples in spacetime, spotted for first time. Science News (Washington, DC), ??(??):??, February 11, 2016. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://

www.sciencemag.org/news/2016/02/gravitational-waves-
einstein-s-ripples-spacetime-spotted-first-time.

Cho:2016:HFP

[Cho16b] Adrian Cho. Here’s the first person [Marco Drago] to spot those gravitational waves. Science News (Washington, DC), ??(??):??, February 11, 2016. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://

www.sciencemag.org/news/2016/02/here-s-first-person-
spot-those-gravitational-waves.

Cho:2016:WDI

[Cho16c] Adrian Cho. ‘We did it!’: Voices from the gravitational wave press conference. Science News (Washington, DC), ??(??):??, February 11, 2016. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http:

//www.sciencemag.org/news/2016/02/we-did-it-voices-
gravitational-wave-press-conference.

Chree:1920:TES

[Chr20] Charles Chree. Testing Einstein’s shift of spectral lines. Nature, 106(2663):343, November 11, 1920. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v106/n2663/pdf/
106343b0.pdf.

Christesco:1922:EUE

[Chr22] Stéfan Christesco. Explorations dans l’ultra — éther de l’univers et les anomalies des théories d’Einstein: suite à la nouvelle étude de cosmogonie scientifique, contenant un atlas de cosmogonie scientifique; systèmes cellulaires des mondes. (French) [Explorations in the ultra-ether of the universe and the anomalies of Einstein theories: following the new study of scientific cosmogony, containing an atlas of scientific cosmogony; worlds’ cellular systems], volume 8 of La Science universelle de l’énergie. Alcan, Paris, France, 1922. 440 pp.

Cooperstock:1990:DGR

[CHR90] F. (Fred) Cooperstock, L. P. (Lawrence Paul) Horwitz, and Joe Rosen, editors. Developments in General Relativity, Astrophysics and Quantum Theory: a Jubilee Volume in Honour of Nathan Rosen, volume 9 of Annals of the Israel Physical Society. Institute of Physics Publishers, Bristol, UK, 1990. ISBN 0-7503-0053-1. viii + 377 pp. LCCN QB460 .D48

1990. URL http://www.zentralblatt-math.org/zmath/en/
search/?an=0734.53005.

Cohen:1996:RAR

[CHR96] Robert S. Cohen, Risto Hilpinen, and Qiu Renzong, editors. Realism and Anti-Realism in the Philosophy of Science: Beijing International Conference, 1992, volume 169 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 94-015-8638-1. ISSN 0068-0346. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-015-8638-2.

Cohen:1997:PEP

[CHS97] Robert S. Cohen, Michael Horne, and John Stachel, editors. Potentiality, Entanglement and Passion-at-a-Distance: Quantum Mechanical Studies for Abner Shimony. Volume Two, volume 194 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4809-X, 94-017-2732-5 (ebook). ISSN 0068-0346. LCCN QC173.96-174.52. URL https:

//link.springer.com/book/10.1007/978-94-017-2732-7.

Clauser:1969:PET

[CHSH69] John F. Clauser, Michael A. Horne, Abner Shimony, and Richard A. Holt. Proposed experiment to test local hidden variable theories. Physical Review Letters, 23(15):880–884, ???? 1969. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://prl.aps.org/abstract/

PRL/v23/i15/p880_1. See erratum [CHSH70].

Clauser:1970:EPE

[CHSH70] John F. Clauser, Michael A. Horne, Abner Shimony, and Richard A. Holt. Erratum: “Proposed Experiment to Test Local Hidden Variable Theories”. Physical Review Letters, 24(10): 549, ???? 1970. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://prl.aps.org/

abstract/PRL/v23/i15/p880_1. See [CHSH69].

Church:1960:HOB

[Chu60] Peggy Pond Church. The house at Otowi Bridge: the story of Edith Warner and Los Alamos. University of New Mexico Press, Albuquerque, NM, USA, 1960. ix + 149 pp.

Crawford:1987:NPC

[CHU87] Elisabeth T. Crawford, J. L. Heilbron, and Rebecca Ullrich. The Nobel population 1901–1937: a census of the nominators and nominees for the prizes in physics and chemistry, volume 11; 4 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1987. ISBN 0-918102-15-4 (paperback). 337 pp. LCCN QC28 .C74 1987.

Chu:1998:NLM

[Chu98a] Steven Chu. Nobel Lecture: The manipulation of neutral particles. Reviews of Modern Physics, 70(3):685–706, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.70.685;   http://rmp.aps.org/
abstract/RMP/v70/i3/p685_1.

Church:1998:HOB

[Chu98b] Peggy Pond Church. The house at Otowi Bridge: the story of Edith Warner and Los Alamos. University of New Mexico Press, Albuquerque, NM, USA, 1998. ISBN 0-585-18785-1 (electronic bk.), 0-8263-0014-6 (cloth), 0-8263-0281-5 (paperback). ix + 149 pp. LCCN F804.L6 W3 1998eb.

Church:2001:HOB

[Chu01] Peggy Pond Church. The house at Otowi Bridge: the story of Edith Warner and Los Alamos. University of New Mexico Press, Albuquerque, NM, USA, 2001. ISBN 0-8263-0281-5. 149 pp. LCCN ???? Drawings by Connie Fox Boyd.

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Cirincione:2007:BSH

[Cir07] Joseph Cirincione. Bomb scare: the history and future of nuclear weapons. Columbia University Press, New York, NY, USA, 2007. ISBN 0-231-13510-6 (cloth), 0-231-50940-5 (electronic). xiv + 206 pp. LCCN U264 .C57 2007. URL http://www.loc.

gov/catdir/toc/ecip0620/2006029174.html.

Cirincione:2008:BSH

[Cir08] Joseph Cirincione. Bomb scare: the history and future of nuclear weapons. Columbia University Press, New York, NY, USA, 2008. ISBN 0-231-13511-4. xiv + 232 pp. LCCN M08.E12019.

Chan:2022:RLE

[CJT22] Ho Fai Chan, Franklin G. Mixon Jr, and Benno Torgler. Recognition and longevity: an examination of award timing and lifespan in Nobel laureates. Scientometrics, 127(6):3629–3659, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-022-04379-6.

Cernohorsky:1979:PEP

[ČK79] Martin Černohorský and Jiří Komrska. Pocta Einsteinovi — Praha 1979. (Czech) [Tribute to Einstein — Prague 1979]. Československý Časopis pro Fyziku, A29(??):309–311, ???? 1979.

Cribb:2004:ANL

[CK^{+}04] Julian Cribb, John Keeney, et al., editors. Australia’s Nobel Laureates: adventures in innovation 1915–1996. ETN Communications Pty Ltd, Roseville, NSW, Australia, 2004. ISBN 0-9580207-4-4. 117 pp. LCCN ????

Calaprice:2015:EE

[CKS15] Alice Calaprice, Daniel Kennefick, and Robert J. Schulmann. An Einstein Encyclopedia. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-14174-6 (hardcover), 1-4008-7336-3 (e-book). xxiii + 347 pp. LCCN QC16.E5 C343 2015.

Calaprice:2005:AEB

[CL05] Alice Calaprice and Trevor Lipscombe. Albert Einstein: a biography. Greenwood biographies, 1540-4900. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2005. ISBN 0-313-33080-8. xxiii + 161 pp. LCCN QC16.E5 C34 2005. URL http:

//www.loc.gov/catdir/toc/ecip054/2004028175.html.

Corie-Lok:2007:ES

[CL07] Corie-Lok. Einstein on stage. Boston Blog, April 23, 2007.

CODEN ???? ISSN ???? URL http://blogs.nature.com/
boston/2007/04/23/einstein-on-stage.

Cassini:2019:ERF

[CL19] Alejandro Cassini and Marcelo Leonardo Levinas. Einstein’s reinterpretation of the Fizeau experiment: How it turned out to be crucial for special relativity. Studies in History and Philosophy of Modern Physics, 65(??):55–72, February 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816302131.

Clark:1961:BBU

[Cla61] Ronald W. Clark. The birth of the bomb: the untold story of Britain’s part in the weapon that changed the world. Phoenix, London, UK, 1961. 209 pp. LCCN ????

Clark:1968:EM

[Cla68] Ronald W. Clark. Einstein memorabilia. Science, 162(3855): 751, November 15, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/162/3855/751.full.pdf.

Clark:1971:ELT

[Cla71] Ronald William Clark. Einstein: the life and times. World Publishing Company, New York, NY, USA, 1971. xv + 718 pp. LCCN QC16.E5 C5.

Clark:1973:ELT

[Cla73] Ronald William Clark. Einstein: the life and times. Hodder and Stoughton, London, UK, 1973. ISBN 0-340-10960-2. 670 + 16 pp. LCCN QC16.E5 C5 1973.

Clarke:1978:GE

[Cla78] Arthur C. Clarke. God and Einstein. Science and Public Policy, 5(5):399–400, October 1978. ISSN 0302-3427 (print),

1471-5430 (electronic). URL http://spp.oxfordjournals.
org/content/5/5/399.full.pdf+html.

Clark:1984:ELT

[Cla84] Ronald (Ronald William) Clark. Einstein: the life and times. Avon Books, New York, NY, USA, 1984. ISBN 0-380-01159-X. 878 + 16 pp. LCCN QC16.E5 C5 1995. US$5.95.

Clancy:1991:SAF

[Cla91] Tom Clancy. The sum of all fears. Putnam, New York, NY, USA, 1991. ISBN 0-399-13615-0, 0-399-13631-2 (limited edition). 798 pp. LCCN PS3553.L245 S8 1991.

Clark:1997:RGW

[Cla97] Claudia Clark. Radium Girls, Women and Industrial Health Reform: 1910–1935. University of North Carolina Press, Chapel Hill, NC, USA, 1997. ISBN 0-8078-2331-7 (hardcover), 0-8078-4640-6 (paperback), 0-8078-6081-6 (e-book). xii + 289 pp. LCCN HD6067.2.U6 C55 1997.

Carilli:2003:MER

[CLD^{+}03] C. L. Carilli, G. F. Lewis, S. G. Djorgovski, A. Mahabal, P. Cox, F. Bertoldi, and A. Omont. A molecular Einstein ring: Imaging a starburst disk surrounding a quasi-stellar object. Science, 300(5620):773–775, May 2, 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/300/5620/773.full.pdf.

Curie:1927:EIB

[CLE27] M. Curie, H. A. Lorentz, and A. Einstein. The establishment of an International Bureau of Meteorology. Science, 65(??):415– 417, ???? 1927. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Clemence:1947:REP

[Cle47] G. M. Clemence. The relativity effect in planetary motions. Reviews of Modern Physics, 19(4):361–364, October 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.19.361;   http://rmp.aps.org/
abstract/RMP/v19/i4/p361_1.

ClerkMaxwell:1982:DTE

[Cle82] James Clerk Maxwell. A Dynamical Theory of the Electromagnetic Field. Scottish Academic Press, Edinburgh, UK, 1982. ISBN 0-7073-0324-9 (paperback). xiii + 103 pp. LCCN

QC670 .M38 1982. URL https://hdl.handle.net/2027/mdp.
39015017148431;   https://hdl.handle.net/2027/mdp.39076002001753.

With an appreciation by Albert Einstein. Edited and introduced by Thomas F. Torrance.

ClerkMaxwell:1984:QMa

[Cle84] James Clerk Maxwell. Quotations from Maxwell. In Berger [Ber84b], pages 1–9. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Clemons:1996:TTS

[Cle96] Russell E. Clemons. A trip through space and time: Las Cruces to Cloudcroft, volume 15 of Scenic trips to the geologic past. New Mexico Bureau of Mines and Mineral Resources, Socorro, NM, USA, 1996. 194 + 14 pp. LCCN QE143 .A3 no. 15.

Clewell:2006:AEH

[Cle06] David Clewell. ‘Albert Einstein held me in his arms’. Georgia Review, 60(2):292–293, Summer 2006. ISSN 0016-8386.

Clery:2008:WAE

[Cle08] Daniel Clery. Wishing for an African Einstein. Science, 320(5876):604–605, May 2, 2008. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/320/5876/604.full.pdf.

Clegg:2009:BBB

[Cle09] Brian Clegg. Before the Big Bang: the prehistory of our universe. St. Martin’s Press, New York, NY, USA, 2009. ISBN 0-312-38547-1. 306 pp. LCCN QB981 .C627 2009. URL http://

www.loc.gov/catdir/enhancements/fy0908/2008046035-b.
html;   http://www.loc.gov/catdir/enhancements/fy0908/
2008046035-d.html.

Clegg:2017:ESZ

[Cle17] Brian Clegg. Einstein in 30 Sekunden: 50 Zentrale Aspekte zum Leben und Vermächtnis. (German) [Einstein in 30 Seconds: 50 Key aspects of life And legacy]. Librero, Kerkdriel, Germany, 2017. ISBN 90-8998-879-3. 160 pp. LCCN ????

Cline:1965:QPQ

[Cli65] Barbara Lovett Cline. The questioners: physicists and the quantum theory. Crowell, New York, 1965. vii + 274 pp. LCCN QC15

.C4. URL http://catalog.hathitrust.org/api/volumes/
oclc/372589.html.

Cline:1987:MWM

[Cli87] Barbara Lovett Cline. Men who made a new physics: physicists and the quantum theory. University of Chicago Press, Chicago, IL, USA and London, UK, 1987. ISBN 0-226-11027-3 (paperback). xii + 274 pp. LCCN QC15

.C4 1987. US$11.95. URL http://www.loc.gov/catdir/
description/uchi051/87010786.htm;   http://www.loc.
gov/catdir/enhancements/fy0608/87010786-t.htm;   http:
//www.loc.gov/catdir/enhancements/fy0609/87010786-b.
htm.

Clifford:2009:BRB

[Cli09a] J. Garry Clifford. Book review: From Roosevelt to Truman: Potsdam, Hiroshima, and the Cold War. By Wilson D. Miscamble. New York: Cambridge University Press. 2007. American Studies, 50(1/2):226–227, Spring/Summer 2009. ISSN 0026-3079 (print), 2153-6856 (electronic). URL

http://muse.jhu.edu/journals/american_studies/v050/
50.1.clifford.html.

Cline:2009:SDD

[Cli09b] D. (David) Cline, editor. Sources and Detection of Dark Matter and Dark Energy in the Universe: proceedings of the 8th UCLA Symposium, Marina del Rey, California, 20–22 February 2008, volume 1166 of AIP conference proceedings, 0094-243X. American Institute of Physics, Woodbury, NY, USA, 2009. ISBN 0-7354-0703-7. LCCN QB791.3 .I579 2008. URL http://www.

loc.gov/catdir/enhancements/fy1006/2009907622-d.html.

Close:2004:PPV

[Clo04] Frank E. Close. Particle physics: a very short introduction, volume 109 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-280434-0 (paperback). 148 pp. LCCN

QC778 .C56 2004. URL http://www.loc.gov/catdir/
enhancements/fy0636/2004049295-d.html;   http://www.
loc.gov/catdir/enhancements/fy0723/2004049295-b.html.

Close:2009:NVS

[Clo09] Frank E. Close. Nothing: a very short introduction. Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-922586-9 (paperback). 157 pp. LCCN QC6 .C588 2009.

Close:2010:A

[Clo10] Frank E. Close. Antimatter. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-955016-6 (hardcover), 0-19-957887-7 (paperback). x + 166 pp. LCCN QC173.3 .C56 2010.

Close:2015:HLD

[Clo15] Frank E. Close. Half-life: the divided life of Bruno Pontecorvo, physicist or spy. Basic Books, New York, NY, USA, 2015. ISBN 0-465-06998-3 (hardcover), 0-465-04487-5 (e-book), 1-78074-582-6 (e-book). xix + 378 pp. LCCN QC774.P66 C56 2014. US$29.99.

Carrington:1967:ECC

[CM67] Alan Carrington and Terry A. Miller. Einstein A coefficients for the 18 cm transitions of OH. Nature, 214(5092):998–999, June 3, 1967. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v214/n5092/pdf/214998a0.pdf.

Crease:1986:SCM

[CM86] Robert P. Crease and Charles C. Mann. The second creation: makers of the revolution in Twentieth-Century physics. Macmillan Publishing Company, New York, NY, USA, 1986. ISBN 0-02-521440-3. xi + 480 pp. LCCN 509.

Cushing:1989:PCQ

[CM89] James T. Cushing and Ernan McMullin, editors. Philosophical consequences of quantum theory: reflections on Bell’s theorem, volume 2 of Studies in science and the humanities from the Reilly Center for Science, Technology, and Values. University of Notre Dame Press, Notre Dame, IN, USA, 1989. ISBN 0-268-01578-3, 0-268-01579-1 (paperback). xiii + 314 pp. LCCN QC174.12 .P43 1989. US$39.95, US$19.95. URL http://en.

wikipedia.org/wiki/Bell%27s_theorem.

Crease:1997:SCM

[CM97] Robert P. Crease and Charles C. Mann. The second creation: makers of the revolution in Twentieth-Century physics. Quartet, London, UK, new edition, 1997. ISBN 0-7043-8038-2. ix + 484 pp. LCCN 98.E04413.

Cardone:2002:EFS

[CM02a] Fabio Cardone and Roberto Mignani. Enrico Fermi e i secchi della sora Cesarina: metodo, pregiudizio e caso in fisica. (Italian) [????], volume 14 of Collana Arcobaleno. Di Renzo, Roma, Italy, nuova edition, 2002. ISBN 88-8323-050-7. 119 pp. LCCN QC16.F46 C37 2002.

Coster-Mullen:2002:ABT

[CM02b] John Coster-Mullen. Atom bombs: the top secret inside story of Little Boy and Fat Man. [privately published], Waukesha, WI, USA, 2002. 400 (est.) pp. LCCN QC773.3.U5 C678 2002.

Chan:2018:REC

[CMT18] Ho F. Chan, Franklin G. Mixon, Jr., and Benno Torgler. Relation of early career performance and recognition to the probability of winning the Nobel Prize in Economics. Scientometrics, 114(3):1069–1086, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2614-5.

Coburn:1940:GEH

[Cob40] N. Coburn. Generalized Einstein hypersurfaces of spaces of constant curvature. Journal of Mathematics and Physics (MIT), 19(1–4):140–152, April 1940. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/
10.1002/sapm1940191140.

Cobo:2007:EGS

[Cob07] Ancena Lopez Cobo. From Einstein to Gomez de la Serna. The Theory of Relativity and the secret of Modern Art (Albert Einstein, Romon Gomez de la Serna ). Arbor — Ciencia Pensamiento y Cultura, 183(728):911–921, November/December 2007. ISSN 0210-1963.

Cocke:1966:ERCa

[Coc66a] W. J. Cocke. Errata: Relativistic corrections for terrestrial clock synchronization. Physical Review Letters, 16:779, April 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.16.779.2.

Cocke:1966:ERCb

[Coc66b] W. J. Cocke. Errata: Relativistic corrections for terrestrial clock synchronization. Physical Review Letters, 16:1233, June 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.16.1233.

Cocke:1966:RCTa

[Coc66c] W. J. Cocke. Relativistic corrections for terrestrial clock synchronization. Physical Review Letters, 16:662–664, April 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/

PhysRevLett.16.662. This may be the first published work on the need for corrections from General Relativity for the Global Positioning System (GPS); see errata [Coc66a, Coc66b] and comment [Coc16].

Cocke:2016:CEG

[Coc16] W. J. Cocke. Celebrating Einstein [GPS]. Scientific American, 314(1):6, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). The author claims that his classified work in 1966 was the first to apply corrections from General Relativity for the Global Positioning System (GPS), but a sanitized version was published in [Coc66c].

Cohen:1931:GUE

[Coh31] Chapman Cohen. God and the universe, Eddington, Jeans, Huxley, & Einstein. Pioneer Press, London, UK, third edition, 1931. 133 pp. LCCN ???? With a reply by A. S. Eddington.

Cohen:1940:RFD

[Coh40] I. Bernard Cohen. Roemer and the first determination of the velocity of light (1676). Isis, 31(2):327–379, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225757.

Cohen:1949:SPS

[Coh49] Morris Raphael Cohen. Studies in philosophy and science. H. Holt, New York, NY, USA, 1949. 278 pp. LCCN B945.C53 S7.

Cohen:1951:BRBb

[Coh51] I. Bernard Cohen. Book review: Albert Einstein: Philosopher-Scientist by Paul Arthur Schilpp. Isis, 42(1):76–79, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226691.

Cohen:1954:BRBc

[Coh54] I. Bernard Cohen. Book reviews: Nobel Prize Winners in Physics, 1901–1950 by Niels H. de V. Heathcote; Nobel Prize Winners in Medicine and Physiology, 1901–1950 by Lloyd G. Stevenson; Nobel Prize Winners in Chemistry, 1901–1950 by Eduard Farber. Isis, 45(4):407–408, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226800.

Cohen:1955:IE

[Coh55] I. Bernard Cohen. An interview with Einstein. Scientific American, 193(1):68–73, July 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v193/n1/
pdf/scientificamerican0755-68.pdf.

Cohen:1979:CAE

[Coh79] William Howard Cohen. Conversations with Albert Einstein, and other poems for the Einstein year. Caves Books, Taipei, Republic of China, 1979. 56 pp. LCCN PS3553.O4337 C6.

Cohen:1980:TCQ

[Coh80] Robert S. Cohen. Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time, volume 19-1 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346. xx +

308 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-009-8935-1.

Collin:1957:MAE

[Col57] R. Collin. La metaphysique d’Albert Einstein. (French) [Metaphysical theories of Albert Einstein]. Concours medical, 79(37): 3947–3949, September 14, 1957. ISSN 0010-5309.

Colnort:1968:AOBa

[Col68] S. Colnort. Analyses d’ouvrages: Dialogue concerning the two chief World Systems, second revised edition par Galileo Galilei; Stillman Drake; Albert Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):182, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904480.

Collins:1993:WOJ

[Col93] J. J. Collins. Out of this world — otherworldly journeys from Gilgamesh to Albert Einstein — I. P. Couliano. Journal of Religion, 73(1):144, January 1993. ISSN 0022-4189.

Collins:1995:GBE

[Col95] Graham P. Collins. Gaseous Bose–Einstein condensate finally observed. Physics Today, 48(8):17–20, August 1, 1995. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2808119.

Collins:2000:EC

[Col00a] Graham P. Collins. Einstein’s constant. Scientific American, 283(1):28, July 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v283/n1/
pdf/scientificamerican0700-28a.pdf.

Colombo:2000:CBA

[Col00b] J. A. Colombo. Commentaries on the brain of Albert Einstein. Medicina — Buenos Aires, 60(4):530–532, ???? 2000. ISSN 0025-7680.

Collins:2004:GSS

[Col04] H. M. (Harry M.) Collins. Gravity’s shadow: the search for gravitational waves. University of Chicago Press, Chicago, IL, USA and London, UK, 2004. ISBN 0-226-11377-9 (hardcover), 0-226-11378-7 (paperback). xxiii + 870 pp. LCCN QC179 .C65 2004. URL http://www.loc.gov/

catdir/description/uchi052/2003023823.html;   http:
//www.loc.gov/catdir/enhancements/fy0614/2003023823-
b.html;   http://www.loc.gov/catdir/toc/ecip0410/2003023823.
html.

Colbert:2009:AED

[Col09] David Colbert. Albert Einstein: 10 Days That Shook Your World. Paw Prints, ????, 2009. ISBN 1-4395-7513-4. 160 (est.) pp. LCCN ????

Colin:2012:RQM

[Col12] Read Colin. The rise of the Quants. Palgrave Macmillan, Basingstoke, UK, 2012. ISBN 1-137-02614-6. LCCN

???? URL http://www.palgraveconnect.com/doifinder/
10.1057/9781137026149.

Collier:2013:MIT

[Col13] Peter Collier. Most Incomprehensible Thing: Notes Towards a Very Gentle Introduction to the Mathematics of Relativity. Incomprehensible Books, Harlow, UK, 2013. ISBN 0-9573894-4-2. 364 (est.) pp. LCCN ????

Collins:2017:GKD

[Col17] H. M. (Harry M.) Collins. Gravity’s kiss: the detection of gravitational waves. MIT Press, Cambridge, MA, USA, 2017. ISBN 0-262-03618-5 (hardcover). vi + 408 pp. LCCN QC179 .C647 2017.

Combridge:1923:EP

[Com23] J. T. Combridge. An Einstein paradox. Nature, 112(2804): 134, July 28, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v112/n2804/pdf/112134a0.pdf.

Compton:1946:FML

[Com46a] Arthur H. Compton. Franklin Medal Lecture: Atomic energy as a human asset. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 90 (1):70–79, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/3301042.

Compton:1946:IAB

[Com46b] Karl T. Compton. If the atomic bomb had not been used. Atlantic Monthly, ??(??):54–??, December 1946. ISSN 1072-7825 (print), 2151-9463 (electronic). URL http://

www.theatlantic.com/magazine/archive/1946/12/if-the-
atomic-bomb-had-not-been-used/376238/.

Compton:1956:AIP

[Com56a] Arthur H. Compton. Die Atombombe und ich: Ein persönlicher Erlebnisbericht. (German) [The atomic bomb and I: a Personal Experience Report]. Nest, Frankfurt/M., West Germany, 1956. 480 pp. LCCN ???? Translation to German by Erwin Schuhmacher of [Com56b].

Compton:1956:AQP

[Com56b] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Combourieu:1995:CAD

[Com95] Marie-Christine Combourieu. Controverse autour de la définition de la réalité physique. Le paradoxe d’Einstein– Podolsky–Rosen (1935) et la non-séparabilité quantique. (French) [Controversy around the definition of physical reality. The paradox of Einstein–Podolsky–Rosen (1935) and the quantum non-separability]. Dialectica: International Review of Philosophy of Knowledge, 49(1):47–74, March 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Comfort:2014:EGB

[Com14] Ray Comfort. Einstein, God and The Bible. WND Books, Washington, DC, USA, 2014. ISBN 1-936488-17-5 (hardcover), 1-936488-18-3 (e-book). xxvi + 278 pp. LCCN BL473 .C655 2014.

Condon:1935:RPD

[Con35] Edward U. Condon. Raises point of doubt. New York Times, ??(??):11, May 4, 1935. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/101505481. This short article imme-

diately follows, and comments on, the reports of the upcoming EPR paper [Ano35].

Condon:1962:YQP

[Con62] Edward U. Condon. 60 years of quantum physics. Physics Today, 15(10):37–49, October 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//www.physicstoday.org/resource/1/PHTOAD/v15/i10. De-

layed 1951 Presidential address at the 1500th regular meeting of the American Philosophical Society of Washington, 2 December 1962, at the Natural History Museum Auditorium of the Smithsonian Institution, on the 60th anniversary of Planck’s constant, h. Reprinted in [WP85, pages 310–318].

Conant:1970:MSL

[Con70] James Bryant Conant. My Several Lives; Memoirs of a Social Inventor. Harper & Row, New York, NY, USA, 1970. xvi + 701 pp. LCCN CT275.C757 A3.

Conklin:1983:BRE

[Con83] Nan Conklin. Book review: Einstein for Beginners, by Joseph Schwartz and Michael McGuinness. Leonardo (Oxford, England), 16(1):63, January 1983. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL https://muse.jhu.

edu/article/599505/summary.

Conant:2002:TPW

[Con02] Jennet Conant. Tuxedo Park: a Wall Street tycoon and the secret palace of science that changed the course of World War II. Simon and Schuster, New York, NY, USA, 2002. ISBN 0-684-87287-0. xvi + 330 + 16 pp. LCCN

QC16.L647 C66 2002. URL http://www.loc.gov/catdir/
bios/simon052/2002021001.html;   http://www.loc.gov/
catdir/description/simon033/2002021001.html;   http:
//www.loc.gov/catdir/enhancements/fy0645/2002021001-
t.html;   http://www.loc.gov/catdir/samples/simon031/
2002021001.html.

Conant:2005:EPR

[Con05] Jennet Conant. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. Simon and Schuster, New York, NY, USA, 2005. ISBN 0-7432-5007-9. xviii + 425 + 16 pp. LCCN QC773.A1 C66 2005.

Conant:2017:MHJ

[Con17] Jennet Conant. Man of the hour: James B. Conant, warrior scientist. Simon and Schuster, New York, NY, USA, 2017. ISBN 1-4767-3088-1 (hardcover), 1-4767-3091-1 (paperback), 1-4767-3092-X (e-book). x + 587 + 16 pp. LCCN CT275.C757 C66 2017.

Cooke:1983:EDW

[Coo83] Maurice B. Cooke. Einstein doesn’t work here anymore: a treatise on the new science: with step-by-step instructions for apparatus to demonstrate the inert gas field effect. Marcus Books, Toronto, ON, Canada, 1983. ISBN 0-919951-00-7. vi + 133 pp.

Cooper:1999:EFR

[Coo99] Dan Cooper. Enrico Fermi: and the revolutions in modern physics. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-511762-X. 117 pp. LCCN QC16.F46 C66 1999. URL http://www.loc.

gov/catdir/enhancements/fy0603/98034471-d.html.

Cooke:2004:BRM

[Coo04] Anne Cooke. Book review: From microbes to man: Ulf Lagerkvist, Pioneers of microbiology and the Nobel Prize. World Scientific, 2003. Pp. 192, £33.00 (hardback). ISBN 981-238-233-X. £16.00 (paperback). ISBN 981-238-234-8. Notes and Records of the Royal Society of London, 58(2):234–236, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142054.

Coopersmith:2010:ESC

[Coo10] Jennifer Coopersmith. Energy, the subtle concept: the discovery of Feynman’s blocks from Leibniz to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-954650-9 (hardcover). xiv + 400 pp. LCCN QC72 .C66 2010.

Coopersmith:2015:ESC

[Coo15] Jennifer Coopersmith. Energy, the subtle concept: the discovery of Feynman’s blocks from Leibniz to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 2015. ISBN 0-19-871674-5 (paperback). xviii + 422 pp. LCCN QC72 .C66 2015.

Corry:1997:HMP

[Cor97] Leo Corry. Hermann Minkowski and the Postulate of Relativity. Archive for History of Exact Sciences, 51(4):273–314, December 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=4&spage=273.

Cornwell:2003:HSS

[Cor03] John Cornwell. Hitler’s scientists: science, war, and the Devil’s pact. Viking, New York, NY, USA, 2003. ISBN 0-670-03075-9. xvi + 535 pp. LCCN Q127.G3 C67 2003.

Corry:2004:DHA

[Cor04a] Leo Corry. David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2004. ISBN 1-4020-2777-X (hardcover), 1-4020-2778-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 513 pp. LCCN

QA29.H5 C67 2004. URL http://www.loc.gov/catdir/
enhancements/fy0823/2006280837-d.html;   http://www.
loc.gov/catdir/enhancements/fy0823/2006280837-t.html.

Corry:2004:EMT

[Cor04b] Leo Corry. Einstein and Mie: Two pillars of Hilbert’s unified theory. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 287–316. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.com/chapter/10.1007/978-1-
4020-2778-9_7.

Corry:2007:OHA

[Cor07] Leo Corry. The origin of Hilbert’s axiomatic method. In Renn et al. [RSS^{+}07b], pages 759–855. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL https://link.

springer.com/chapter/10.1007/978-1-4020-4000-9_40.

Correll:2012:TCI

[Cor12] John T. Correll. They called it Star Wars. Air Force magazine, 95(6):66–70, June 2012. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.airforce-

magazine.com/MagazineArchive/Pages/2012/June%202012/
0612starwars.aspx.

Corcoran:2015:NYT

[Cor15a] David Corcoran, editor. The New York Times book of science: more than 150 years of groundbreaking scientific coverage. Sterling Publishing, New York, NY, 2015. ISBN 1-4027-9321-9. xx + 540 pp. LCCN Q225 .N49 2015. Foreword by Brian Greene.

Cornell:2015:LUP

[Cor15b] Elizabeth Cornell. Louis Untermeyer’s poetic engagement of the popularization of Einstein’s relativity theory. South Central Review, 32(2):48–66, 2015. CODEN ???? ISSN 0743-6831 (print), 1549-3377 (electronic).

Corry:2015:RME

[Cor15c] Richard Corry. Retrocausal models for EPR. Studies in History and Philosophy of Modern Physics, 49(??):1–9, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981400121X.

CostaDeBeauregard:1980:EEP

[Cos80] Olivier Costa De Beauregard. The 1927 Einstein and 1935 EPR paradox. Physis: Rivista Internazionale di Storia della Scienza, 22(2):211–242, 1980. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

CostadeBeauregard:1987:MLC

[Cos87a] Olivier Costa de Beauregard. S-matrix, Lorentz-and-CPT invariance, and Einstein–Podolsky–Rosen correlations. In Time, The Physical Magnitude [Cos87b], chapter 19, pages 265–284. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.
1007/978-94-009-3811-3_19.

CostaDeBeauregard:1987:TPM

[Cos87b] Olivier Costa De Beauregard. Time, The Physical Magnitude, volume 99 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346. xxiii + 340 pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-009-3811-3.

Coontz:2005:ISI

[COS05] Robert Coontz, Ian Osborne, and Phil Szuromi. Introduction to special issue: a passion for physics. Science, 307(5711):865, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Cotti:1989:EMP

[Cot89] Piero Cotti. Einstein als Mensch und Physiker. (German) [Einstein as man and physicist]. Techinfo [Winterthur], 1(4):10–14, ???? 1989.

Chan:2015:DNL

[CÖT15] Ho Fai Chan, Ali Sina Önder, and Benno Torgler. Do Nobel laureates change their patterns of collaboration following prize reception? Scientometrics, 105(3):2215–2235, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11192-015-1738-8.

Coulson:1952:RML

[Cou52] C. A. Coulson. Review: Out of My Later Years, by Albert Einstein. The British Journal for the Philosophy of Science, III(9):92–93, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/III/9/92.full.pdf+html;
http://links.jstor.org/sici?sici=0007-0882%28195205%
293%3A9%3C92%3AOOMLY%3E2.0.CO%3B2-8.

Cousins:1979:AEB

[Cou79a] N. Cousins. Albert Einstein and the bomb. Saturday Review, 6 (5):10, ???? 1979. ISSN 0361-1655.

Cousins:1979:PFM

[Cou79b] N. Cousins. In praise of famous men: Albert Einstein and John Wayne. Saturday Review, 6(17):8, ???? 1979. ISSN 0361-1655.

Coughlan:2013:SWE

[Cou13] Sean Coughlan. The scientists who escaped the Nazis. BBC News Web story., July 17, 2013. URL http://www.bbc.com/ news/business-23261289. Interview with 92-year-old Gustav Born, son of Max and Hedi Born.

Callaghan:2015:ILA

[CP15] John Callaghan and Mark Phythian. Intellectuals of the Left and the atomic dilemma in the age of the US atomic monopoly, 1945–1949. Contemporary British History, 29(4):441–463, January 2015. ISSN 1361-9462 (print), 1743-7997 (electronic).

Cufaro-Petroni:1987:EPR

[CPDH^{+}87] N. Cufaro-Petroni, C. Dewdney, P. R. Holland, A. Kyprianidis, and J. P. Vigier. Einstein–Podolsky–Rosen constraints on quantum action at a distance: The Sutherland paradox. Foundations of Physics, 17(8):759–773, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00733265.

Campanelli:2007:IPI

[CR07] M. Campanelli and L. Rezzolla. Invited papers from the international meeting on ‘New Frontiers in Numerical Relativity’ (Albert Einstein Institute, Potsdam, Germany, 17–21 July 2006). Classical and Quantum Gravity, 24(12):??, June 21, 2007. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Craig:2008:ABO

[CR08] Campbell Craig and Sergey Radchenko. The Atomic Bomb and the Origins of the Cold War. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-11028-6. xxv + 201 pp. LCCN

D843 .C67 2008. URL http://www.ebrary.com.

Clark:2012:LAN

[CR12] Charles W. Clark and Joseph Reader. Light, atoms and nuclei: The optical discovery of deuterium. Optics and Photonics News, 23(5):36–41, May 2012. CODEN OPPHEL. ISSN 1047-6938 (print), 1541-3721 (electronic). URL https://www.osa-opn.

org/home/articles/volume_23/issue_5/features/light,_
atoms_and_nuclei_the_optical_discovery_of/. See also

[RC13].

Cranberg:1979:EAS

[Cra79] Lawrence Cranberg. Einstein: Amateur scientist. Physics Today, 32(12):9, 11, December 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Crawford:1984:AAH

[Cra84a] Elisabeth Crawford. Arrhenius, the atomic hypothesis, and the 1908 Nobel prizes in physics and chemistry. Isis, 75(3):503– 522, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
232940.

Crawford:1984:BNI

[Cra84b] Elisabeth T. Crawford. The beginnings of the Nobel institution: the science prizes, 1901–1915. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26584-3. ix + 281 + 6 pp. LCCN QC28 .C73 1984. URL http://www.loc.

gov/catdir/description/cam022/84005844.html;   http:
//www.loc.gov/catdir/toc/cam026/84005844.html.

Crawford:1990:SNP

[Cra90] Elisabeth Crawford. The secrecy of Nobel Prize selections in the sciences and its effect on documentation and research. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 134(4):408–419, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986896.

Crawford:1992:NIS

[Cra92] Elisabeth T. Crawford. Nationalism and internationalism in science, 1880–1939: four studies of the Nobel population. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover). xii + 157 pp. LCCN

Q126.9 .C73 1992. URL http://www.loc.gov/catdir/
description/cam024/91033702.html;   http://www.loc.
gov/catdir/samples/cam031/91033702.html;   http://www.
loc.gov/catdir/toc/cam028/91033702.html.

Cramer:1997:EB

[Cra97] John Cramer. Einstein’s bridge. Avon Books, New York, NY, USA, 1997. ISBN 0-380-97510-6 (hardcover), 0-380-79279-6 (softcover). 354 pp. LCCN PS3553.R2674 E39 1997.

Cramer:2000:DMA

[Cra00a] M. D. Cramer. Driving Mr. Albert: A trip across America with Einstein’s brain. Library Journal, 125(12):125, July 2000. CODEN LIBJA7. ISSN 0363-0277.

Crawford:2000:GSH

[Cra00b] Elisabeth Crawford. German scientists and Hitler’s vendetta against the Nobel prizes. Historical Studies in the Physical and Biological Sciences, 31(1):37–53, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Crawford:2001:NPA

[Cra01] Elisabeth Crawford. Nobel population 1901–50: anatomy of a scientific elite. Physics World, 14(11):31–35, November 2001. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://stacks.iop.org/2058-7058/14/i=11/
a=29.

Crawford:2002:NPC

[Cra02a] Elisabeth Crawford. The Nobel population 1901–1950: a census of the nominators and nominees for the prizes in physics and chemistry, volume 30 of Uppsala studies in the history of science. Universal Academy Press, Tokyo, Japan, 2002. ISBN 4-946443-70-3. 420 pp. LCCN fa3475.

Crawford:2002:HSN

[Cra02b] Elisabeth T. Crawford, editor. Historical Studies in the Nobel Archives: the Prizes in Science and Medicine, volume 31 of Uppsala studies in the history of science. Universal Academy Press, Tokyo, Japan, 2002. ISBN 4-946443-69-X. ISSN 0282-1036. vii + 161 pp. LCCN AS911.N9 .H57 2002.

Cravens:2007:PSW

[Cra07] Gwyneth Cravens. Power to save the world: the truth about nuclear energy. Alfred A. Knopf, New York, NY, USA, 2007. ISBN 0-307-26656-7. xv + 439 pp. LCCN TK9146

.C65 2007. URL http://www.loc.gov/catdir/enhancements/
fy0804/2007017611-b.ht;   http://www.loc.gov/catdir/
enhancements/fy0804/2007017611-d.ht;   http://www.loc.
gov/catdir/enhancements/fy0804/2007017611-s.ht;   http:
//www.loc.gov/catdir/toc/ecip0716/2007017611.html.

Crelinsten:1981:REG

[Cre81] Jeffrey Michael Crelinsten. The reception of Einstein’s general theory of relativity among American astronomers, 1910– 1930. Ph.D. thesis, Université de Montréal, Montréal, Québec, Canada, 1981. xxvi + 533 pp. URL http://search.proquest.

com/docview/303272287.

Crelinsten:1983:WWC

[Cre83] Jeffrey Crelinsten. William Wallace Campbell and the “Einstein Problem”: An observational astronomer confronts the Theory of Relativity. Historical Studies in the Physical Sciences, 14(1):

1–91, ???? 1983. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757525.

Crease:2005:GR

[Cre05] Robert P. Crease. General relativity. Physics World, 18 (12):16–17, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/12/phwv18i12a24.
pdf.

Crelinsten:2006:EJR

[Cre06] Jeffrey Crelinsten. Einstein’s jury: the race to test relativity. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12310-1 (hardcover), 1-4008-4967-5 (e-book). xxix + 397 pp. LCCN QC173.585 .C74 2006. URL http://www.loc.

gov/catdir/enhancements/fy0654/2005032681-d.html;
http://www.loc.gov/catdir/enhancements/fy0734/2005032681-
b.html;   http://www.loc.gov/catdir/toc/ecip063/2005032681.
html.

Crease:2010:GEB

[Cre10] Robert P. Crease. The Great Equations: Breakthroughs in Science from Pythagoras to Heisenberg. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33793-6. 315 pp. LCCN Q175 .C884 2010.

Creager:2015:RCM

[Cre15] Angela N. H. Creager. Radiation, cancer, and mutation in the Atomic Age. Historical Studies in the Natural Sciences, 45(1): 14–48, February 2015. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/
10.1525/hsns.2015.45.1.14.

Crommelin:1920:EDL

[Cro20a] A. C. D. Crommelin. The Einstein deflection of light. Nature, 105(2627):23–24, March 4, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v105/n2627/pdf/105023a0.pdf.

Crommelin:1920:ETS

[Cro20b] Andrew C. D. Crommelin. The Einstein theory and spectral displacement. Nature, 104(2621):532, January 22, 1920. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v104/n2621/
pdf/104532d0.pdf.

Cropper:2001:GPL

[Cro01] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-

5. xii + 500 pp. LCCN QC15 .C76 2001. URL http://www.loc.

gov/catdir/enhancements/fy0637/2001021611-d.html;
http://www.loc.gov/catdir/enhancements/fy0724/2001021611-
b.html.

Cronin:2004:FR

[Cro04] James W. Cronin, editor. Fermi remembered. University of Chicago Press, Chicago, IL, USA and London, UK, 2004. ISBN 0-226-12111-9 (cloth). xi + 287 pp. LCCN QC16.F46 F49 2004. URL http://www.loc.gov/

catdir/description/uchi052/2003020524.html;   http:
//www.loc.gov/catdir/enhancements/fy0614/2003020524-
b.html;   http://www.loc.gov/catdir/toc/ecip049/2003020524.
html.

Crom:2005:AEG

[Cro05] W. Crom. Albert Einstein as a geomorphologist. Zeitschrift für Geomorphologie, 49(3):273–274, September 2005. ISSN 0372-8854.

Crowe:2007:MAE

[Cro07] Michael J. Crowe. Mechanics from Aristotle to Einstein. Green Lion Press, Santa Fe, NM, USA, 2007. ISBN 1-888009-32-2 (paperback). xxii + 331 pp. LCCN QA802 .C76 2007. URL http:

//catdir.loc.gov/catdir/toc/fy0803/2007925664.html.

Crockatt:2016:ETC

[Cro16] Richard Crockatt. Einstein and Twentieth-century Politics: ‘a Salutary Moral Influence’. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-878549-6 (hardcover), 0-19-108829-3 (e-book), 0-19-182740-1 (e-book). viii + 236 pp. LCCN QC16.E5 C76 2016.

Corry:1997:BDH

[CRS97] Leo Corry, Jürgen Renn, and John Stachel. Belated decision in the Hilbert–Einstein priority dispute. Science, 278(5341):1270– 1273, November 14, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/278/5341/1270.full.pdf.

Cruickshank:1979:SO

[Cru79] A. D. Cruickshank. Soddy at Oxford. British Journal for the History of Science, 12(3):277–288, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cranshaw:1960:ETE

[CS60] T. E. Cranshaw and J. P. Schiffer. Experiments to test Einstein’s Principle of Equivalence. Nature, 185(4714):653–654, March 5, 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v185/n4714/pdf/185653a0.pdf.

Cohen:1970:EMPb

[CS70] Robert S. Cohen and Raymond J. Seeger. Ernst Mach: Physicist and Philosopher, volume 6 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1970. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346. viii +

298 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-017-1462-4.

Cohen:1979:SPL

[CS79a] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. xxxiv + 929 pp. LCCN Q174

.B67 vol. 21 QC7. URL https://link.springer.com/book/
10.1007/978-94-009-9349-5.

Cohen:1979:ECB

[CS79b] Robert S. Cohen and John J. Stachel. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday). In Cohen and Stachel [CS79a], chapter II.10, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67

vol. 21 QC7. URL http://link.springer.com/chapter/10.
1007/978-94-009-9349-5_35.

Chen:2023:TFE

[CSC23] Lingzhi Chen, Yutao Sun, and Cong Cao. A two-fold evaluation in science: the case of Nobel Prize. Scientometrics, 128 (11):6267–6291, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04830-2.

Colombe:2007:SAF

[CSD^{+}07] Yves Colombe, Tilo Steinmetz, Guilhem Dubois, Felix Linke, David Hunger, Jakob Reichel, et al. Strong atom–field coupling for Bose–Einstein condensates in an optical cavity on a chip. Nature, 450(7167):272–276, November 8, 2007. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v450/n7167/
full/nature06331.html.

Chaturvedi:2006:STR

[CSM06] S. Chaturvedi, R. Simon, and N. Mukunda. Space, time and Relativity. Resonance, 11(7):14–29, July 2006. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://

www.springerlink.com/content/t35676764n02430k/.

Carter:2021:SSC

[CSS^{+}21] Korey P. Carter, Katherine M. Shield, Kurt F. Smith, Zachary R. Jones, Jennifer N. Wacker, Leticia Arnedo-Sanchez, Tracy M. Mattox, Liane M. Moreau, Karah E. Knope, Stosh A. Kozimor, Corwin H. Booth, and Rebecca J. Abergel. Structural and spectroscopic characterization of an einsteinium complex. Nature, 590(7844):85–88, February 3, 2021. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Czinkos:2001:SLP

[CST01] Tímea Czinkos, Leo Szilard, and Jánosne Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2001. Oktatási Minisztérium, Budapest, Hungary, 2001. 44 pp. LCCN MLCS 2008/45433 (Q).

Czinkos:2002:SLP

[CST02] Tímea Czinkos, Leo Szilard, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2002. MF Plusz, Budapest, Hungary, 2002. 44 pp. LCCN MLCS 200845422 (Q).

Cornelis:1999:MSB

[CSvB99] Gustaaf C. Cornelis, Sonja Smets, and Jean Paul van Bendegem, editors. Metadebates on science: the blue book of “Einstein meets Magritte”, volume 6 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 0-7923-5762-0 (vol. 6), 0-7923-5765-5 (set), 90-5487-231-4 (VUB: vol. 6), 90-5487-234-9 (VUB: set). LCCN Q174 .M45 1999. See other volumes in this series [Aer99a, AMM99, Aer99b, ABW99, AVvdV99, BAM99, HBR99].

Crawford:1996:NTW

[CSW96] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of wartime injustice. Nature, 382(6590):393–395, August 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v382/n6590/abs/382393a0.html.

Crawford:1997:KIP

[CSW97a] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Die Kernspaltung und ihr Preis. Warum nur Otto Hahn den Nobelpreis erhielt, Otto Frisch, Lise Meitner und Fritz Straßmann dagegen nicht berücksichtigt werden. (German) [Fission and its price. Why only Otto Hahn received the Nobel Prize: Otto Frisch, Lise Meitner and Fritz Strassmann are not taken into consideration]. Kultur & Technik. Zeitschrift des Deutschen Museums München, 21(2):30–35, ???? 1997. ISSN 0344-5690. URL http://www.deutsches-

museum.de/fileadmin/Content/data/Insel/Information/
KT/heftarchiv/1997/21-2-30.pdf. Translation from English

by Dieter Beisel of [CSW96].

Crawford:1997:NTP

[CSW97b] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of postwar injustice: Recently released Swedish documents reveal why Lise Meitner, codiscoverer of nuclear fission, did not receive the 1946 Physics Prize for her theoretical interpretation of the process. Physics Today, 50(9):26–32, September 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Crawford:1998:LMN

[CSW98] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Lise Meitner und der Nobelpreis. (German) [Lise Meitner and the Nobel Prize]. Physik in unserer Zeit, 29(6):234–241, 1998. CO-DEN PHUZAH. ISSN 0031-9252.

Cohen-Tannoudji:1980:MQ

[CT80] Claude Cohen-Tannoudji. Mécanique quantique. Hermann, Paris, France, 1980. ISBN 2-7056-5733-9 (vol. 1), 2-7056-5767-3 (vol. 2). xvi + xv + 1515 (two volumes) pp. LCCN QC174.12 .C63 1977.

Cohen-Tannoudji:1998:NLM

[CT98] Claude N. Cohen-Tannoudji. Nobel Lecture: Manipulating atoms with photons. Reviews of Modern Physics, 70(3):707– 719, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.70.707;   http://rmp.aps.
org/abstract/RMP/v70/i3/p707_1.

Cohen-Tannoudji:2012:NDP

[CT12] Gilles Cohen-Tannoudji. Les neutrinos défient les physiciens. (French) [Neutrinos challenge physicists]. La Recherche, ??(466): 38, June 19, 2012. CODEN RCCHBV. ISSN 0029-5671 (print),

1625-9955 (electronic). URL http://www.larecherche.fr/
physique-du-xxie-siecle/neutrinos-defient-physiciens-
01-07-2012-91376.

Chan:2015:IEM

[CT15] Ho Fai Chan and Benno Torgler. The implications of educational and methodological background for the career success of Nobel laureates: an investigation of major awards. Scientometrics, 102 (1):847–863, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-014-1367-7.

Cardin:2024:LCS

[CT24] Franco Cardin and Rossana Tazzioli. Levi-Civita simplifies Einstein. The Ricci rotation coefficients and unified field theories. Archive for History of Exact Sciences, 78(1):87–126, January 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00322-0.

Cohen-Tannoudji:1973:MQ

[CTDL73] Claude Cohen-Tannoudji, Bernard Diu, and Franck Laloë. Mécanique quantique, volume 16 of Collection Enseignement des sciences. Hermann, Paris, France, 1973. ISBN 2-7056-5733-9 (v. 1). xv + 1493 (two volumes) pp. LCCN QC174.12 .C63.

Cohen-Tannoudji:1977:QM

[CTDL77] Claude Cohen-Tannoudji, Bernard Diu, and Franck Laloë. Quantum mechanics. Wiley, New York, NY, USA, 1977. ISBN 0-471-16432-1 (vol. 1), 0-471-16433-X (vol. 1: paperback). xv + 1524 (two volumes) pp. LCCN QC174.12 .C6313

1977. URL http://www.loc.gov/catdir/enhancements/
fy0607/76005874-b.html;   http://www.loc.gov/catdir/
enhancements/fy0607/76005874-d.html;   http://www.loc.
gov/catdir/toc/wiley023/76005874.html.

Cugota:2006:MNA

[Cug06] Lluis Cugota. My name is — Albert Einstein. My name is. Barrons Educational Series, Inc., Hauppauge, NY, USA, 2006. ISBN 0-7641-3391-8. 96 pp. LCCN QC16.E5

C82 2006. URL http://www.loc.gov/catdir/enhancements/
fy0630/2005929509-d.html.

Culhane:1980:XRA

[Cul80] J. L. Culhane. X-ray astronomy with the Einstein observatory. Nature, 284(5756):509–510, April 10, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v284/n5756/
pdf/284509a0.pdf.

Cullwick:1981:ESR

[Cul81] E. G. Cullwick. Einstein and Special Relativity. The British Journal for the Philosophy of Science, 32(2):167–176, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
32/2/167.full.pdf+html.

Cunningham:1915:RET

[Cun15] Ebenezer Cunningham. Relativity and the electron theory. Monographs on physics. Longmans, Green and Co., London, UK, 1915. vii + 96 pp. LCCN QC6 .C8.

Cunningham:1919:ERTa

[Cun19a] E. Cunningham. Einstein’s relativity theory of gravitation. Nature, 104(2614):354–356, December 4, 1919. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v104/n2614/
pdf/104354c0.pdf.

Cunningham:1919:ERTb

[Cun19b] E. Cunningham. Einstein’s relativity theory of gravitation. Nature, 104(2615):374–376, December 11, 1919. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v104/n2615/
pdf/104374b0.pdf.

Cunningham:1919:ERTc

[Cun19c] E. Cunningham. Einstein’s relativity theory of gravitation. Nature, 104(2616):394–395, December 18, 1919. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v104/n2616/
pdf/104394b0.pdf.

Cunningham:1920:ERT

[Cun20a] E. Cunningham. Einstein’s relativity theory of gravitation. Scientific American, 1:116–119, February 1, 1920. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v1/
n2msupp/pdf/scientificamerican02011920-116supp.pdf.

Cunningham:1920:ETM

[Cun20b] E. Cunningham. Einstein’s theory and a map analogue. Nature, 104(2618):437, January 1, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v104/n2618/pdf/104437b0.pdf.

Cunningham:1920:FET

[Cun20c] E. Cunningham. The foundations of Einstein’s theory of gravitation. Nature, 105(2638):350–351, May 20, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v105/n2638/
pdf/105350a0.pdf.

Cunningham:1922:CEH

[Cun22a] E. Cunningham. A criticism of Einstein and his Problem; Relativity for All; Einstein and the Universe: A Popular Exposition of the Famous Theory; Le Principe de Relativité et la Théorie de la Gravitation; La Théorie einsteinienne de la Gravitation: Essai de vulgarisation de la théorie L’Éther actuel et ses précurseurs (simple récit); Raum und Zeit im Lichte der speziellen Relativitätstheorie Versuch eines synthetischen Aufbaus der speziellen Relatitivätstheorie. Nature, 109(2746):770– 772, June 17, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v109/n2746/pdf/109770a0.pdf.

Cunningham:1922:PCE

[Cun22b] E. Cunningham. Pour comprendre Einstein: Die Grundlagen der einstein’schen Relativitätstheorie: Eine kritische Untersuchung philosophy and the new physics: An essay on the relativity theory and the theory of quanta le principe de la relativité et les théories d’Einstein le principe de la relativité et la théorie d’Einstein the romanes lecture, 1922 the Theory of Relativity and its influence on scientific thought. Nature, 110(2765): 568–570, October 28, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v110/n2765/pdf/110568a0.pdf.

Curie:1937:MCB

[Cur37] Eve Curie. Madame Curie: a biography. The Da Capo series in science. Doubleday and Co., Garden City, NY, USA, 1937. xvii + 393 pp. LCCN QD22.C8 C85 1937. Translation from French original by Vincent Sheean.

Curie:1938:MCLa

[Cur38a] Eve Curie. Madame Curie. Éditions de la Nouvelle Revue française. Gallimard, Paris, France, 1938. 311 pp. LCCN ???? German translation of the French original.

Curie:1938:MCH

[Cur38b] Eve Curie. Madame Curie: haar leven en werk. (Dutch) [Madame Curie: her life and work]. Éditions de la Nouvelle Revue française. Leopold, Den Haag, The Netherlands, 1938. 405 pp. LCCN ???? Dutch translation by Willy Corsari (pseudonym for Wilhelmina Angela Schmidt) of the French original.

Curie:1938:MCI

[Cur38c] Eve Curie. Madame Curie: ihr Leben und Wirken. (German) [Madame Curie: her life and work]. Knaur, Berlin, Germany, 1938. 467 pp. LCCN ???? German translation of the French original.

Curie:1938:MCLb

[Cur38d] Eve Curie. Madame Curie: Leben und Wirken. (German) [Madame Curie: life and work]. Bermann-Fischer, Wien, Austria, 1938. 458 pp. LCCN ???? German translation of the French original.

Curie:1938:MCLc

[Cur38e] Eve Curie. Madame Curie: Leben und Wirken. (German) [Madame Curie: life and work]. Büchergilde Gutenberg, Zürich, Switzerland, 1938. 408 pp. LCCN ???? German translation of the French original.

Curie:1938:MMM

[Cur38f] Eve Curie. Min mor Marie Skłodowska Curie. (Swedish). [My mother, Marie Skłodowska Curie]. Schildts, Helsingfors, Finland, 1938. 408 pp. LCCN ???? German translation of the French original.

Curie:1960:MC

[Cur60] Ewa Curie. Maria Curia. Panstwowe Wydawnictwo Naukowe, Warszawa, Poland, 1960. ???? pp. LCCN ????

Curtin:1982:ADS

[Cur82] Deane W. Curtin, editor. The aesthetic dimension of science: 1980 Nobel Conference [7–8 October, London, 1980]. Philosophical Library, New York, NY, USA, 1982. ISBN 0-8022-2393-1. LCCN Q174 .N6 1980. US$12.50.

Curie:1986:MCB

[Cur86] Eve Curie. Madame Curie: a biography. The Da Capo series in science. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1986. xvii + 393 pp. LCCN QD22.C8 C85 1986. Translation from French original by Vincent Sheean.

Curry:1992:NCC

[Cur92] Charles Curry. The naturalness of the cosmological constant in the general theory of relativity: a response to Ray. Studies in History and Philosophy of Science Part A, 23(4):657–660, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/003936819290016Y. See [Ray90,

Ray92].

Curie:2001:MCB

[Cur01] Eve Curie. Madame Curie: a biography. The Da Capo series in science. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, second edition, 2001. ISBN 0-306-81038-7 (paperback). xvii + 393 pp. LCCN QD22.C8 C85 2001. Translation from French original by Vincent Sheean.

Cushing:1982:BRB

[Cus82] James T. Cushing. Book review: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911), by Arthur I. Miller. American Journal of Physics, 50(5):476, May 1982. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?
AJP/50/476/1.

Cvoric:2006:OSE

[Čvo06] Mirjana Čvorić. One stone — Einstein. Production TV

Šabac, Serbia, 2006. URL http://www.youtube.com/watch?
v=htfBTfVD75o.

Cockcroft:1932:EHVa

[CW32a] John D. Cockcroft and Ernest T. S. Walton. Experiments with high velocity positive ions: I. further developments in the method of obtaining high velocity positive ions. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):619–630, June 1, 1932. ISSN 0950-1207. URL http://www.jstor.org/ stable/95811. The work in this paper, and its companion [CW32b], won the Nobel Prize in Physics in 1951 for the authors “for their pioneer work on the transmutation of atomic nuclei by artificially accelerated atomic particles”.

Cockcroft:1932:EHVb

[CW32b] John D. Cockcroft and Ernest T. S. Walton. Experiments with high velocity positive ions: II. The disintegration of elements by high velocity protons. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 137(831):229–242, July 1, 1932. ISSN 0950-

1207. URL http://www.jstor.org/stable/95941. The work

in this paper, and its companion [CW32a], won the Nobel Prize in Physics in 1951 for the authors “for their pioneer work on the transmutation of atomic nuclei by artificially accelerated atomic particles”.

Cohen:1974:LES

[CW74] Robert S. Cohen and Marx W. Wartofsky. Logical and Epistemological Studies in Contemporary Physics, volume 13 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 462 pp. LCCN Q174 .B67 vol. 13;

QC5.56. URL https://link.springer.com/book/10.1007/
978-94-010-2656-7.

Cohen:1984:HS

[CW84] Robert S. Cohen and Marx W. Wartofsky. Hegel and the Sciences, volume 64 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346. ix + 374 pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-009-6233-0.

Ciufolini:1995:GI

[CW95] Ignazio Ciufolini and John Archibald Wheeler. Gravitation and Inertia. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03323-4 (hardcover), 0-691-03654-3. xi + 498 pp. LCCN

QC173.59.G44 C58 1995. URL http://www.loc.gov/catdir/
description/prin021/94029874.html;   http://www.loc.
gov/catdir/toc/prin031/94029874.html.

Cornell:1998:BEC

[CW98] Eric A. Cornell and Carl E. Wieman. The Bose–Einstein condensate. Scientific American, 278(3):40–45 (Intl. ed. 26– 31), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
40.pdf;   http://www.sciam.com/1998/0398issue/0398currentissue.
html.

Cornell:2002:NLB

[CW02] E. A. Cornell and C. E. Wieman. Nobel Lecture: Bose–Einstein condensation in a dilute gas, the first 70 years and some recent experiments. Reviews of Modern Physics, 74(3):875–893, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.875;   http://rmp.aps.org/
abstract/RMP/v74/i3/p875_1.

Cavalcanti:2012:BNS

[CW12] Eric G. Cavalcanti and Howard M. Wiseman. Bell nonlocality, signal locality and unpredictability (or what Bohr could have told Einstein at Solvay had He known about Bell experiments). Foundations of Physics, 42(10):1329–1338, October 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
012-9669-1.

Cwiklik:1987:AET

[Cwi87] Robert Cwiklik. Albert Einstein and the Theory of Relativity. Solutions. Barron’s, New York, NY, USA, 1987. ISBN 0-8120-3921-1 (paperback). vi + 184 pp. LCCN QC16.E5 C88 1987.

US$4.95. URL http://www.loc.gov/catdir/enhancements/
fy0630/87019549-d.html.

Czerski:2005:LIE

[Cze05] Helen Czerski. Let the inner Einstein out. Physics World, 18(5): 56, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/5/phwv18i5a50.pdf.

Chariker:2017:ISM

[CZPR17] Julia H. Chariker, Yihang Zhang, John R. Pani, and Eric C. Rouchka. Identification of successful mentoring communities using network-based analysis of mentor–mentee relationships across Nobel laureates. Scientometrics, 111(3):1733–1749, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2364-4.pdf.

Durr:1992:WH

[D^{+}92] H.-P. (Hans-Peter) Dürr et al., editors. Werner Heisenberg. Hanser, München, Germany, 1992. ISBN 3-446-17217-3. 123 pp. LCCN QC16.H35 W47 1992.

Drell:1995:NTS

[D^{+}95] Sidney D. Drell et al. Nuclear testing: Summary and conclusions. Report JSR-95-320, JASON, The MITRE Corporation, McLean, VA, USA, 1995.

DAbro:1950:EST

[D’A50] A. D’Abro. The evolution of scientific thought from Newton to Einstein. Dover, New York, NY, USA, second edition, 1950. ISBN 0-486-20002-7. xx + 481 pp. LCCN QC6 .A3 1950.

DAgostino:1975:HRE

[D’A75] Salvo D’Agostino. Hertz’s researches on electromagnetic waves. Historical Studies in the Physical Sciences, 6(??):261–323, ???? 1975. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757343.

DAntonio:1993:AHH

[D’A93] Michael D’Antonio. Atomic Harvest: Hanford and the Lethal Toll of America’s Nuclear Arsenal. Crown Publishers, New York, NY, USA, 1993. ISBN 0-517-58981-8. xiii + 304 pp. LCCN TD898.12.W2 D36 1993. US$22.50 (US$29.50 Canada).

DAgostino:1997:RRE

[D’A97] Salvo D’Agostino. Remarks on recent Einsteinian studies: General Relativity in the debates of the 1920s between neo-Kantian and empiricist logicians. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 34(3):643–658 (1998), 1997. CO-DEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

DAgostino:2000:HIT

[D’A00] Salvo D’Agostino. A History of the Ideas of Theoretical Physics: Essays on the Nineteenth and Twentieth Century Physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6094-X, 94-010-9034-3. ISSN 0068-0346. xviii + 381 pp. LCCN Q174 .B67 no.

213. URL https://link.springer.com/book/10.1007/978-
94-010-9034-6.

DAgostino:2001:ECC

[D’A01a] Salvo D’Agostino. Einstein’s correspondence criterion and the construction of general relativity. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01c], chapter 10, pages 223–237. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. LCCN QC5.53;

QC6. URL https://link.springer.com/chapter/10.1007/
978-94-010-9034-6_10.

DAgostino:2001:ELL

[D’A01b] Salvo D’Agostino. Einstein’s life-long doubts on the physical foundations of the general relativity and unified field theories. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01c], chapter 11, pages 239–252. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. LCCN QC5.53; QC6. URL https://link.

springer.com/chapter/10.1007/978-94-010-9034-6_11.

DAgostino:2001:HIT

[D’A01c] Salvo D’Agostino. A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0244-0, 94-010-9034-3 (ebook). ISSN 0068-0346. xviii + 381 pp. LCCN QC5.53; QC6.

Dacy:1920:WUS

[Dac20a] G. H. Dacy. Where Uncle Sam’s income goes, the progress of the Einstein judges. Scientific American, 123(26):625, December 25, 1920. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v123/n26/pdf/
scientificamerican12251920-625.pdf.

Dacy:1920:FFF

[Dac20b] George H. Dacy. Fish fancying and freak farming, the $5,000 Einstein essay contest, and more. Scientific American, 123(2):31–33, July 10, 1920. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v123/n2/
pdf/scientificamerican07101920-31.pdf.

Dacey:2011:ELC

[Dac11] James Dacey. Einstein’s landing card resurfaces after 80 years. Web blog., May 12, 2011.

Daddario:1978:HAE

[Dad78] Emilio Q. Daddario. Honoring Albert Einstein. Science, 202 (4373):1149, December 15, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/202/4373/1149.full.pdf.

Daison:1967:TSF

[Dai67] F. Daison. Tomonaga, Shvinger i Fejnman — laureaty nobelevskoj premii po fizike. (Russian) [Tomonaga, Schwinger and Feynman — Nobel Prize in Physics]. Uspekhi Fizicheskikh Nauk, 91(1):71–73, 1967. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL http://ufn.ru/ru/articles/
1967/1/f/.

Damour:2006:OUE

[Dam06] Thibault Damour. Once upon Einstein. A. K. Peters, Ltd., Wellesley, MA, USA, 2006. ISBN 1-56881-289-2. xiii + 185 pp. LCCN QC16.E5 D35813 2006. URL http://www.crcnetbase.

com/ISBN/978-1-4398-7009-9.

Dansel:1967:NFL

[Dan67] Michel Dansel. Les Nobel français de littérature. (French) [The French Nobel Prize in Literature]. Collection Les Grands documentaires illustrés. A. Bonne, Paris, France, 1967. 221 pp. LCCN PQ150.L5 D3.

Daniel:1989:BER

[Dan89] Wojciech Daniel. Bohr, Einstein and realism. Dialectica: International Review of Philosophy of Knowledge, 43(3):249–261, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dannen:1997:ESR

[Dan97a] Gene Dannen. The Einstein–Szilard refrigerators. Scientific American, 276(1):90–95, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v276/n1/
pdf/scientificamerican0197-90.pdf;   http://www.sciam.
com/0197issue/0197currentissue.html.

Dannen:1997:SRD

[Dan97b] Gene Dannen. Story of refrigerator designed by Einstein and Szilard gets iced. Physics Today, 50(12):102, December 1997.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1997PhT....
50l.102D.

Danielson:2000:BCI

[Dan00] Dennis Richard Danielson, editor. The book of the cosmos: imagining the universe from Heraclitus to Hawking. Helix books. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0247-9. xxxiii + 556 pp. LCCN QB981 .B724

2000. URL http://catdir.loc.gov/catdir/enhancements/
fy0831/2001268173-d.html.

Danson:2005:BE

[Dan05] Roy Danson. Blooming Einstein. Physics World, 18(8):22, August 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/8/phwv18i8a26.pdf.

Darrigol:1996:EOR

[Dar96] Olivier Darrigol. The electrodynamic origins of Relativity theory. Historical Studies in the Physical and Biological Sciences, 26(2):241–312, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:2000:EAE

[Dar00] Olivier Darrigol. Electrodynamics from Ampère to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850594-9. xix + 532 pp. LCCN 0.2dar a0165 a4100.

Dardo:2004:NLT

[Dar04a] M. (Mauro) Dardo. Nobel Laureates and Twentieth-Century Physics. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83247-0, 0-521-54008-9 (paperback). xi + 533 pp. LCCN QC7 .D27 2004. URL http://www.loc.gov/

catdir/description/cam041/2004049240.html;   http://
www.loc.gov/catdir/toc/cam041/2004049240.html.

Darrigol:2004:RAS

[Dar04b] Olivier Darrigol. On a recent article by Seiya Abiko. Historical Studies in the Physical and Biological Sciences, 35(1):153–155, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Abi03] and reply [Abi04].

Darling:2006:GAS

[Dar06a] David J. Darling. Gravity’s arc: the story of gravity, from Aristotle to Einstein and beyond. Wiley, New York, NY, USA, 2006. ISBN 0-471-71989-7. ix + 278 pp. LCCN

QC178 .D373 2006. URL http://www.loc.gov/catdir/
enhancements/fy0650/2005030772-d.html;   http://www.
loc.gov/catdir/toc/ecip062/2005030772.html.

Darrigol:2006:GTR

[Dar06b] Olivier Darrigol. The genesis of the Theory of Relativity. In Damour et al. [DDDR06], pages 1–31. ISBN 3-7643-7435-7 (hardcover), 3-7643-7436-5 (e-book). LCCN QC173.5 .P65 2005eb.

Darrigol:2014:QE

[Dar14] Olivier Darrigol. The quantum enigma. In Janssen and Lehner [JL14b], chapter 4, pages 117–142. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Darrigol:2019:FSE

[Dar19] Olivier Darrigol. Frames and stresses in Einstein’s quest for a generalized theory of relativity. Studies in History and Philosophy of Modern Physics, 68(??):126–157, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301217.

Darrigol:2021:CWT

[Dar21] Olivier Darrigol. Can we trust Einstein’s accounts of the genesis of special relativity? Studies in History and Philosophy of Science Part A, 89(??):138–154, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000492.

Dash:1991:TDW

[Das91] Joan Dash. The triumph of discovery: women scientists who won the Nobel Prize. Julian Messner, Englewood Cliffs, NJ, 1991. ISBN 0-671-69332-8 (hardcover), 0-671-69333-6 (paperback). xi + 148 pp. LCCN R692 .D36 1990.

Das:2000:NEO

[Das00] Saswato R. Das. Noble endeavors: an overview of Nobel Prizewinning research at Bell Labs. Bell Labs Technical Journal, 5 (1):95–106, Spring 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dasgupta:2012:CPT

[Das12] Deepanwita Dasgupta. Creating a peripheral trading zone: Satyendra Nath Bose and Bose–Einstein statistics, doing science in the role of an outsider. International Studies in the Philosophy of Science, 26(3):259–287, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/02698595.2012.
731731.

Datzeff:1938:PBP

[Dat38] Assène Datzeff. Sur le problème des barrières de potentiel et la résolution de l’équation de Schrödinger. (French) [On the problem of potential barriers and the solution of Schrödinger’s equation ]. Masson et cie, Paris, France, 1938. 2 + 91 + 1 pp. LCCN QC174.2 .D38.

Dattagupta:2006:MAE

[Dat06] Sushanta Dattagupta. The myth about Einstein. Resonance, 11(1):63–78, January 2006. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.

springerlink.com/content/q7005313620403nw/.

Date:2015:GRB

[Dat15] Ghanashyam Date. General Relativity: Basics and Beyond. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-5271-9 (hardcover), 1-4665-5272-7 (e-book). xi + 255 pp. LCCN QC173.6 .D38 2015.

Davies:1974:EC

[Dav74] P. C. W. Davies. Einstein in crisis? Nature, 249(5457): 510–511, June 7, 1974. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v249/n5457/pdf/249510a0.pdf.

Davies:1981:POE

[Dav81] Paul Davies. Princeton’s obeisance to Einstein. Nature, 291 (5813):362–363, May 28, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v291/n5813/pdf/291362a0.pdf.

Davies:1983:HS

[Dav83] Robertson Davies. High spirits. Penguin, London, UK and New York, NY, USA, 1983. ISBN 0-14-006505-9 (paperback). iv + 198 pp. LCCN PR9199.3.D3 H5 1982.

Davies:1995:ATE

[Dav95] Paul C. W. Davies. About time: Einstein’s unfinished revolution. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-671-79964-9. 316 pp. LCCN QC173.59.S65 D37 1995.

Davis:2003:NLH

[Dav03] Raymond Davis, Jr. Nobel Lecture: a half-century with solar neutrinos. Reviews of Modern Physics, 75(3):985–994, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.985;   http://rmp.aps.org/
abstract/RMP/v75/i3/p985_1.

Davies:2009:SRN

[Dav09] E. B. Davies. Some reflections on Newton’s principia. British Journal for the History of Science, 42(2):211–224, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

David:2014:FQM

[Dav14] François David. The Formalisms of Quantum Mechanics: an Introduction, volume 893 of Lecture Notes in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-10538-8 (paperback), 3-319-10539-6 (e-book). vii + 157 pp. LCCN

QC174.12 .F72 2015. URL http://www.loc.gov/catdir/
enhancements/fy1504/2014954643-d.html;   http://www.
loc.gov/catdir/enhancements/fy1504/2014954643-t.html.

Day:2012:NPC

[Day12] Charles Day. Nobel prizes for computational science [the last word]. Computing in Science and Engineering, 14(6):88, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

deBothezat:1936:BNC

[dB36] Georges de Bothezat. Back to Newton; a challenge to Einstein’s theory of relativity. G. E. Stechert, New York, NY, USA, 1936. vii + 152 pp. LCCN QC6 .B67.

deBroglie:1949:LED

[dB49a] L. de Broglie. L’oeuvre d’Einstein et la dualité des ondes et des corpuscules. Reviews of Modern Physics, 21(3):345–347, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.345;   http://rmp.aps.org/

abstract/RMP/v21/i3/p345_1. Special issue in honor of Albert Einstein on his 70th birthday.

deBroglie:1949:GSS

[dB49b] Louis de Broglie. A general survey of the scientific work of Albert Einstein. In Schilpp [Sch49b], chapter 2, pages 107–127. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Translated from the French manuscript by Forrest W. Williams.

deBroglie:1949:PCO

[dB49c] Louis de Broglie. La physique contemporaine et l’œuvre d’Albert Einstein. (French) [Contemporary physics and the work of Albert Einstein]. Public lecture of 19 December 1949., 1949.

deBroglie:1949:LUE

[dB49d] Louis de Broglie. L’œuvre d’Einstein et la dualité des ondes et des corpuscules. (French) [The work of Einstein and the duality of waves and corpuscles]. Reviews of Modern Physics, 21 (3):345–347, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.345;   http://rmp.
aps.org/abstract/RMP/v21/i3/p345_1. Special issue in honor

of Albert Einstein on his 70th birthday.

deBroglie:1955:DOC

[dB55a] Louis de Broglie. Le dualisme des ondes et des corpuscules dans l’œuvre d’Albert Einstein. (French) [The dualism of waves and corpuscles in the work of Albert Einstein]. Public lecture of 5 December 1955, 1955.

deBroglie:1955:NNA

[dB55b] Louis de Broglie. Notice nécrologique sur Albert Einstein. (French) [Obituary of Albert Einstein]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 240(??): 1741–1745, ???? 1955. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBeauregard:1965:PCE

[dB65] O. Costa de Beauregard. Le · paradoxe · des correlations d’Einstein: et de Schrödinger et l’épaisseur temporelle de la transition quantique. (French) [The · paradox · of the correlations of Einstein: and Schrödinger and the temporal thickness of the quantum transition]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):280–289, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deBroglie:1972:E

[dB72] Louis de Broglie. Einstein. Réalités Hachette, Paris, France, 1972. ISBN ???? ???? pp. LCCN ????

deBeauregard:1985:SFC

[dB85] O. Costa de Beauregard. On some frequent but controversial statements concerning the Einstein–Podolsky–Rosen correlations. Foundations of Physics, 15(8):871–887, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00738320.

Davies:1986:GAD

[DB86] P. C. W. Davies and J. R. (Julian Russell) Brown. The ghost in the atom: a discussion of the mysteries of quantum physics. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30790-2, 0-521-31316-3 (paperback). ix + 157 pp. LCCN

QC174.12 .D365 1986. URL http://www.loc.gov/catdir/
description/cam031/85025478.htm.

deBeauregard:1987:PPE

[dB87] Olivier Costa de Beauregard. Paradox and paradigm: The Einstein–Podolsky–Rosen correlations. In Time, The Physical Magnitude [Cos87b], chapter 18, pages 244–264. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-009-3811-3_18.

DOnofrio:2009:QMC

[DB09] Mauro D’Onofrio and Carlo Burigana, editors. Questions of modern cosmology: Galileo’s legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-00791-0 (print), 3-642-00792-9 (e-book). xxx + 530 pp. LCCN QB981 .Q47 2009eb.

deBroglie:1966:E

[dBAS^{+}66] Louis de Broglie, Louis Armand, Pierre Henri Simon, et al., editors. Einstein. Hachette, Paris, France, 1966. ???? pp.

deBroglie:1979:E

[dBAS^{+}79a] Louis de Broglie, Louis Armand, Pierre Henri Simon, et al., editors. Einstein. Peebles Press, New York, NY, USA, 1979. ISBN 0-85690-070-2. 219 pp. LCCN QC16.E5 E34E 1979; QC16.E5 E3613; QC16.E5 D4. Translation to English of [dBAS^{+}66].

deBroglie:1979:EE

[dBAS^{+}79b] Louis de Broglie, Louis Armand, Pierre Henri Simon, et al., editors. Einstein. Peebles Press, New York, NY, USA, 1979. ISBN 0-85690-070-2. 219 pp. LCCN QC16.E5 E34E 1979; QC16.E5 E3613; QC16.E5 D4.

Daniel-Cressey:2008:EGH

[DC08] Daniel-Cressey. Einstein: ‘god is human weakness’. Nature News Blog, May 14, 2008. CODEN ???? ISSN ????

URL http://blogs.nature.com/news/2008/05/einstein_
god_is_human_weakness_1.html.

Daniel-Cressey:2010:BER

[DC10] Daniel-Cressey. ‘Baby Einstein’ research row resurrected. Nature News Blog, January 13, 2010. CODEN ???? ISSN

???? URL http://blogs.nature.com/news/2010/01/baby_
einstein_research_row_res.html.

Daniel-Cressey:2012:EDG

[DC12] Daniel-Cressey. Einstein’s documents get online boost. Nature News Blog, March 20, 2012. CODEN ???? ISSN ???? URL

http://blogs.nature.com/news/2012/03/einstein%e2%80%
99s-documents-get-online-boost.html.

Donley:2001:DCE

[DCC^{+}01] Elizabeth A. Donley, Neil R. Claussen, Simon L. Cornish, Jacob L. Roberts, Eric A. Cornell, Carl E. Wieman, et al. Dynamics of collapsing and exploding Bose–Einstein condensates. Nature, 412(6844):295–299, July 19, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v412/n6844/
full/412295a0.html.

Andrade:1964:RNA

[dCENdCA64] E. N. da C. (Edward Neville da Costa) Andrade. Rutherford and the nature of the atom, volume S35 of Science study series. Doubleday, Garden City, NY, USA, 1964. xix + 218 pp. LCCN QC16.R8 A5.

Ding:2023:ERF

[DCL23] Jingda Ding, Yifan Chen, and Chao Liu. Exploring the research features of Nobel laureates in Physics based on the semantic similarity measurement. Scientometrics, 128(9):5247– 5275, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04786-3.

Donley:2002:AMC

[DCTW02] Elizabeth A. Donley, Neil R. Claussen, Sarah T. Thompson, and Carl E. Wieman. Atom–molecule coherence in a Bose–Einstein condensate. Nature, 417(6888):529–533, May 30, 2002. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v417/n6888/
full/417529a.html.

Dirac:1995:CWP

[DD95] P. A. M. (Paul Adrien Maurice) Dirac and R. H. (Richard Henry) Dalitz. The collected works of P. A. M. Dirac, 1924– 1948. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-36231-8. xxiii + 1310 pp. LCCN QC21.2

.D57 1995. URL http://www.loc.gov/catdir/description/
cam026/95001010.html;   http://www.loc.gov/catdir/enhancements/
fy1006/95001010-t.html.

Doria:2011:EGM

[DD11] Francisco Antonio Doria and Manuel Doria. Einstein, Gödel, and the mathematics of time. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 20, pages 269–279. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346. LCCN Q175.3 .B73

2011. URL https://link.springer.com/chapter/10.1007/
978-90-481-9422-3_20.

Demokritov:2006:BEC

[DDD^{+}06] S. O. Demokritov, V. E. Demidov, O. Dzyapko, G. A. Melkov, A. A. Serga, B. Hillebrands, A. N. Slavin, et al. Bose–Einstein condensation of quasi-equilibrium magnons at room temperature under pumping. Nature, 443(7110):430–433, September 28, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v443/n7110/full/nature05117.html.

Damour:2006:EPS

[DDDR06] Thibault Damour, Olivier Darrigol, Bertrand Duplantier, and Vincent Rivasseau, editors. Einstein, 1905–2005: Poincaré Seminar 2005, volume 47 of Progress in mathematical physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 3-7643-7435-7 (hardcover), 3-7643-7436-5 (e-book). viii + 293 pp. LCCN QC173.5 .P65 2005eb.

DeDonder:1924:GWE

[De 24] Théophile De Donder. La gravifique de Weyl–Eddington– Einstein. Gauthier-Villars, Paris, France, 1924. 32 + 14 pp. LCCN ???? Extrait des Bulletins de l’Académie royale de Belgique (Classe des Sciences), nos. 5–9, 298–324, 1924.

DeLaBruheze:1992:RWR

[De 92] Adri De La Bruheze. Radiological weapons and radioactive waste in the United States: insiders’ and outsiders’ views, 1941—1955. British Journal for the History of Science, 25(2): 207–227, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4027299.

deFreitasMourao:2003:ESM

[de 03] Ronaldo Rogério de Freitas Mourão. Einstein: de Sobral para o mundo. (Portuguese) [Einstein: from Sobral to the world]. Edições UVA, Sobral, Brazil, 2003. ISBN 85-87906-15-1. 329 pp. LCCN QC173.6 .M68 2003.

DeGroot:2004:BL

[De 04] Gerard J. De Groot. The bomb: a life. Jonathan Cape, London, UK, 2004. ISBN 0-224-06232-8. xiii + 397 + 16 pp. LCCN U264 .D43 2005.

DeGroot:2005:BL

[De 05] Gerard J. De Groot. The bomb: a life. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01724-2. xiii + 397 + 16 pp. LCCN U264 .D43 2005.

DeVisscher:2010:IMS

[De 10] Alex De Visscher. An index to measure a scientist’s specific impact. Journal of the Association for Information Science and Technology, 61(2):319–328, February 2010. ISSN 2330-

1643. URL http://onlinelibrary.wiley.com/doi/10.1002/
asi.21240/abstract.

DeGreiffA:2012:PDU

[De 12] Alexis De Greiff A. A las puertas del universo derrotado. (Spanish) [At the gates of the defeated universe]. Colección obra selecta. Universidad Nacional de Colombia, Vicerrectoría Académica, Bogotá, D.C., Colombia, 2012. ISBN 958-761-141-1 (hardcover), 958-761-142-X (paperback), 958-761-143-8 (ebook). 180 pp. LCCN QC7.5 .D425 2012.

DeDuve:2013:SVM

[De 13] Christian De Duve. Sept vies en une: mémoires d’un prix Nobel. (French) [Seven lives in one: memories of a Nobel Prize]. Sciences. Jacob, Paris, France, 2013. ISBN 2-7381-2843-2 (paperback). 334 pp. LCCN QP26.D44 A3 2013. 25,90 EUR.

dePadova:2015:AGS

[de 15] Thomas de Padova. Allein gegen die Schwerkraft: Einstein 1914–1918. (German) [Alone against gravity: Einstein 1914– 1918]. Hanser, Carl, München, Germany, 2015. ISBN 3-446-44481-5. 280 pp. LCCN ????

Dvoeglazov:2004:RGC

[DE04] Valeri V. Dvoeglazov and Augusto A. Espinoza Garrido, editors. Relativity, gravitation, cosmology. Contemporary fundamental physics. Nova Science Publishers, Hauppauge, NY, USA, 2004. ISBN 1-59033-981-9. x + 195 pp. LCCN QC173.58 .R45 2004.

dEspagnat:2006:PP

[d’E06] Bernard d’Espagnat. On physics and philosophy. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11964-3. ix + 503 pp. LCCN QC6 .E81713 2006. URL http://www.

loc.gov/catdir/enhancements/fy0654/2006926978-d.htm;
 http://www.loc.gov/catdir/enhancements/fy1007/2006926978-
t.htm.

Dean:2013:NYT

[Dea13] Cornelia Dean, editor. The New York Times book of physics and astronomy: more than 100 years of covering the expanding universe. Sterling, New York, NY, USA, 2013. ISBN 1-4027-9320-0 (hardcover). xviii + 557 pp. LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Debever:1929:ECA

[Deb29] Robert Debever. Elie Cartan — Albert Einstein letters on absolute parallelism 1929–1932. Princeton University Press, Princeton, NJ, USA, 1929. ???? pp. LCCN ????

Debever:1979:ECA

[Deb79] Robert Debever, editor. Elie Cartan — Albert Einstein: letters on absolute parallelism, 1929–1932. Princeton Legacy Library. Princeton University Press, Princeton, NJ, USA, 1979. ISBN 0-691-60804-0, 1-4008-6804-1 (e-book). 260 pp. LCCN QC15

.E454 1979. URL http://www.jstor.org/stable/10.2307/j.

ctt13x1bd2. English translation by Jules Leroy and Jim Ritter.

Dyson:1920:DDL

[DED20] F. W. Dyson, A. S. Eddington, and C. Davidson. A determination of the deflection of light of the sun’s gravitational field from observations made at the total eclipse of May 29, 1919. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 220(??):291–333, January 1, 1920. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Dedijn:1985:AEH

[Ded85] H. Dedijn. Albert Einstein — his world, science and philosophy — Dutch — G. Debrock and P. Schreurer. Tijdschrift voor Filosofie, 47(1):152–153, ???? 1985. ISSN 0040-750X.

DeepeKeever:2008:TSC

[Dee08] Beverly Ann Deepe Keever. Top secret: Censoring the first rough drafts of atomic-bomb history. Media History, 14(2): 185–204, July 2008. CODEN ???? ISSN 1368-8804 (print),

1469-9729 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/13688800802176805.

Deitch:1995:MPS

[Dei95] Kenneth M. Deitch. The Manhattan Project: a secret wartime mission. Perspectives on history series. Discovery Enterprises, Ltd., Lowell, MA, USA, 1995. ISBN 1-878668-41-2 (paperback). 64 pp. LCCN QC773.3.U5 D44 1995.

Deichmann:1996:BUH

[Dei96] Ute Deichmann. Biologists under Hitler. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-07404-1. xviii + 468 pp. LCCN QH305.2.G3 D4513 1996.

Delany:1975:EI

[Del75] Samuel R. Delany. The Einstein intersection. The Garland library of science fiction. Garland Publishing, New York, 1975. ISBN 0-8240-1407-3. 155 pp. LCCN PZ4.D338 Ei6.

Delgado:2009:NDA

[Del09a] James P. Delgado. Nuclear dawn: the atomic bomb, from the Manhattan Project to the Cold War. Osprey Publishing, Limited, Oxford, UK, 2009. ISBN 1-84603-396-9 (hardcover). 216 pp. LCCN U264 .D45 2009.

Deltete:2009:BRB

[Del09b] Robert J. Deltete. Book review: The End of the Certain World: The Life and Science of Max Born, the Nobel Scientist who Ignited the Quantum Revolution. Annals of Science, 66(3):433– 436, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Gre05b].

Deltete:2011:ERH

[Del11a] Robert Deltete. Essay review: How Relativity got accepted and how Einstein came to be regarded as its author. Richard Stayley, Einstein’s Generation: The Origins of the Relativity Revolution, ISBN-13 978-0-226-77057-5. Annals of Science, 68(2):261–265, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2011:BRF

[Del11b] Robert J. Deltete. Book review: Fulvio Melia, Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics, ISBN-13 978-0-226-51951-7. Annals of Science, 68(4):581–583, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2012:ERE

[Del12] Robert Deltete. Essay review: Einstein’s quest for unity: Jeroen van Dongen, Einstein’s Unification, ISBN 978-0-521-88346-7. Annals of Science, 69(2):283–288, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See corrigendum [Ano12a].

DellAglio:2013:RCA

[Del13] L. Dell’Aglio. Recensioni: C. Alunni, M. Castellana, D. Ria, A. Rossi, Albert Einstein et Hermann Weyl, 1955–2005. Questions épistémologiques ouvertes. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 49(1–2):??, 2013. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “the letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125– 152, ???? 2015. CODEN RQSCAN. ISSN 0035-2160.

URL https://www.unamur.be/sciences/philosoc/revueqs/
textes-en-ligne/rqs_186_1_demortier.

Denton:1924:RCS

[Den24] Francis Medforth Denton. Relativity and common sense. Cambridge University Press, Cambridge, UK, 1924. xvii + 279 + 1 pp. LCCN QC6 .D4.

Denberg:1984:AEI

[Den84] K. Denberg. ‘1923, Albert Einstein invents birds’. Shenandoah, 35(1):61, ???? 1984. ISSN 0037-3583.

Deruelle:2015:PET

[Der15] Nathalie Deruelle. De Pythagore à Einstein, tout est nombre: la relativité générale, 25 siècles d’histoire. (French) [From Pythagoras to Einstein, everything is numbers: general relativity, 25 centuries of history]. Belin, Paris, 2015. ISBN 2-7011-9501-2. ???? pp. LCCN ????

Dessau:1972:EOD

[Des72] Paul Dessau. Einstein : Oper in drei Akten, Prolog, zwei Intermezzi und einem Epilog. (German) [Einstein: opera in three acts, prologue, two interludes and an epilogue]. Henschelverlag, Berlin, West Germany, 1972. 291 pp. Libretto von Karl Mickel; Klavierauszug von Horst Karl Hessel.

Deser:1979:SPE

[Des79] Stanley Deser. Supergravity: a post-Einstein unification. In Perlmutter and Scott [PS79], pages 39–54. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-

20.85; QC178 .O72 1979. URL http://link.springer.com/
chapter/10.1007/978-1-4684-3596-2_3/.

DeWitt:1979:QG

[DeW79] Bryce DeWitt. Quantum gravity. In Perlmutter and Scott [PS79], pages 127–143. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178

.O72 1979. URL http://link.springer.com/chapter/10.
1007/978-1-4684-3596-2_6/.

DosSantosPereira:2022:WOP

[DFB22] Letícia Dos Santos Pereira, Olival Freire Júnior, and Gisela Boeck. Wilhelm Ostwald’s pedagogy: an analysis of the Nobel Prize nomination letters. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):139–162, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00026980.2022.2052529.

FrutosSanchez:2004:AE

[dFV04] David de Frutos Sánchez and Carmen Albaladejo Vivero. Albert Einstein. Grandes biografías ilustradas. Dastin Export, Madrid, Spain, 2004. ISBN 84-96249-75-1. 302 + 8 pp. LCCN QC15.E5 F78 2004.

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999.

ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp. LCCN Q141 .D52 1994. URL http://www.loc.gov/

catdir/enhancements/fy0637/99488304-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0637/99488304-t.
html;   http://www.netLibrary.com/urlapi.asp?action=
summary&v=1&bookid=12306.

Drell:2003:GDN

[DG03] Sidney D. (Sidney David) Drell and James E. Goodby. The gravest danger: nuclear weapons. Hoover Institution Press, Stanford, CA, USA, 2003. ISBN 0-8179-4472-9. xii + 134 pp. LCCN U264 .D56 2003.

Dyer:2021:EFH

[DGG^{+}21] Adrian G. Dyer, Andrew D. Greentree, Jair E. Garcia, Elinya L. Dyer, Scarlett R. Howard, and Friedrich G. Barth. Einstein, von Frisch and the honeybee: a historical letter comes to light. Journal of Comparative Physiology, A, 207(4):449–456, May 2021. CODEN JCPADN. ISSN 0340-7594 (print), 1432-1351 (electronic).

deGraaf:2011:EKV

[dGK11] Anne de Graaf and Klaas Kunst. Einstein en de kunst van het zeilen: praktijkboek over leiderschap en communicatie (Dutch) [Einstein and the Art of Sailing: practical book on leadership and communication]. SWP, Amsterdam, The Netherlands, ninth edition, 2011. ISBN 90-8850-207-2. 175 pp. LCCN ????

DeSabbata:2004:GCG

[DGM04] Venzo De Sabbata, George T. Gillies, and Vitalii Nikolaevich Mel’nikov, editors. The gravitational constant: generalized gravitational theories and experiments, volume 141 of NATO science series. Series II, Mathematics, physics, and chemistry. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-1955-6, 1-4020-2242-5 (e-book). xxix + 416 pp.

LCCN QB341 .G74 2003. URL http://www.loc.gov/catdir/
enhancements/fy0814/2004042148-d.html;   http://www.
loc.gov/catdir/enhancements/fy0814/2004042148-t.html.

Dalfovo:1999:TBE

[DGPS99] Franco Dalfovo, Stefano Giorgini, Lev P. Pitaevskii, and Sandro Stringari. Theory of Bose–Einstein condensation in trapped gases. Reviews of Modern Physics, 71(3):463–512, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.71.463;   http://rmp.aps.org/
abstract/RMP/v71/i3/p463_1.

Dukas:1979:AEH

[DH79] Helen Dukas and Banesh Hoffmann. Albert Einstein: The Human Side: New Glimpses From His Archives. Princeton University Press, Princeton, NJ, USA, 1979. ISBN 0-691-08231-6. viii + 167 pp. LCCN QC16.E5 A33. URL http://www.loc.

gov/catdir/description/prin051/78070289.html.

Dukas:1981:AEH

[DH81] Helen Dukas and Banesh Hoffmann, editors. Albert Einstein, the human side: new glimpses from his archives. Princeton University Press, Princeton, NJ, USA, 1981. ISBN 0-691-02368-9. viii + 167 pp. LCCN QC16.E5 A33. URL http://press.

princeton.edu/titles/562.html.

deHaan:2013:PET

[dH13] Victor-Otto de Haan. Possible experiments to test Einstein’s Special Relativity Theory. Journal of Computational Methods in Sciences and Engineering, 13(1–2):51–57, ???? 2013. CO-DEN JCMSCR. ISSN 1472-7978 (print), 1875-8983 (electronic).

URL https://dblp.org/db/journals/jcmse/jcmse13.html#
Haan13.

Dukas:2013:AEH

[DHR13] Helen Dukas, Banesh Hoffmann, and Ze’ev Rosenkranz, editors. Albert Einstein, the human side: glimpses from his archives. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-16023-6 (paperback), 1-4008-4812-1 (e-book). LCCN QC16.E5 A33 2013.

deHaas:1949:TRR

[dHvdB49] W. J. de Haas and G. J. van den Berg. On the transfer rate of the Rollin–Simon film. Reviews of Modern Physics, 21(3):524– 526, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.524;   http://rmp.aps.
org/abstract/RMP/v21/i3/p524_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Deng:2010:EPB

[DHY10] Hui Deng, Hartmut Haug, and Yoshihisa Yamamoto. Excitonpolariton Bose–Einstein condensation. Reviews of Modern Physics, 82(2):1489–1537, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.1489;
 http://rmp.aps.org/abstract/RMP/v82/i2/p1489_1.

DInverno:1992:IER

[D’I92] Ray D’Inverno. Introducing Einstein’s Relativity. Clarendon Press, Oxford, UK, 1992. ISBN 0-19-859686-3s (paperback), 0-19-859653-7 (hardcover). xi + 383 pp. LCCN QC173.55 .D56 1992. US$30.00 (paperback); US$75.00 (hard-

cover). URL http://www.loc.gov/catdir/enhancements/
fy0604/91024894-d.html;   http://www.loc.gov/catdir/
enhancements/fy0604/91024894-t.html.

Dicke:1959:GE

[Dic59] R. H. Dicke. Gravitation — an enigma. American Scientist, 47(1):25–40, March 1959. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/27827244.

Dicke:1960:EEG

[Dic60] R. H. Dicke. Eötvös experiment and the gravitational red shift. American Journal of Physics, 28(4):344–347, 1960. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/28/4/10.1119/1.1935801.

Dicke:1961:EE

[Dic61] R. H. Dicke. The Eötvös Experiment. Scientific American, 205(6):84–95, December 1961. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v205/n6/
pdf/scientificamerican1261-84.pdf.

Dicke:1964:TSE

[Dic64] Robert H. (Robert Henry) Dicke, editor. The Theoretical Significance of Experimental Relativity. Documents on modern physics. Gordon and Breach, New York, NY, USA, 1964. xii + 153 pp. LCCN QC6 .D476.

Dickson:1979:EDD

[Dic79a] David Dickson. Einstein: disagreement delays publication of collected works. Nature, 278(5702):294–295, March 22, 1979. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v278/n5702/
pdf/278294a0.pdf.

Dickson:1979:PES

[Dic79b] David Dickson. Protests over Einstein statue. Nature, 277 (5691):4, January 4, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v277/n5691/pdf/277004a0.pdf.

Dickson:1980:EOT

[Dic80] David Dickson. Einstein Observatory in trouble. Nature, 287(5781):379–380, October 2, 1980. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v287/n5781/pdf/
287379b0.pdf.

Dicke:1981:OFE

[Dic81] Robert H. Dicke. The observational foundation of Einstein’s theory of gravitation the gravitational acceleration. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 125(1):16–19, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986182.

DiChristina:2015:EEE

[DiC15] Mariette DiChristina. From the Editor: Editorial and Einstein. Scientific American, 313(3):6, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-6.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-6.pdf.

Diels:1924:RNL

[Die24] Hermann Diels, editor. De rerum natura: lateinisch und deutsch. 2: Lukrez, von der Natur. (Latin and German) [On nature: Latin and German. 2: Lucretius, On nature]. Weidmann’sche Buchhandlung, Berlin, Germany, 1924. 312 pp.

Dietz:1988:NES

[Die88] Werner Dietz. New exact solutions of Einstein’s field equations: Gravitational force can also be repulsive! Foundations of Physics, 18(5):529–547, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732744.

Diedrick:1995:UMA

[Die95] James Diedrick. Understanding Martin Amis. Understanding contemporary British literature. University of South Carolina Press, Columbia, SC, USA, 1995. ISBN 1-57003-058-8. xi + 207 pp. LCCN PR6051.M5 Z62 1995.

Diehl:2005:AEV

[Die05] R. D. Diehl. Albert Einstein’s vision: remarkable discoveries that shaped modern science. Physics in Perspective (PIP), 7 (4):491–492, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Diels:2011:LNG

[Die11] Hermann Diels, editor. Lukrez: Von der Natur. (German) [Lucretius: On nature]. Sammlung Tusculum. Akademie-Verlag, Berlin, Germany, second edition, 2011. ISBN 3-05-005392-5, 3-05-006392-0. 712 pp. LCCN ????

Dimand:1993:CBM

[Dim93] Robert W. Dimand. The case of Brownian motion: a note on Bachelier’s contribution. British Journal for the History of Science, 26(2):233–234, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4027100.

Dingle:1949:SPI

[Din49] Herbert Dingle. Scientific and philosophical implications of the special theory of relativity. In Schilpp [Sch49c], chapter 20, pages 535–554. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Dingle:1959:PET

[Din59] Herbert Dingle. A possible experimental test of Einstein’s second postulate. Nature, 183(4677):1761, June 20, 1959. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v183/n4677/
pdf/1831761a0.pdf.

Dingle:1965:NMK

[Din65] Herbert Dingle. Note on Mr Keswani’s articles, Origin and Concept of Relativity. The British Journal for the Philosophy of Science, XVI(63):242–246, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/XVI/63/242.full.pdf+
html;   http://links.jstor.org/sici?sici=0007-0882%
28196511%2916%3A63%3C242%3ANOMKAO%3E2.0.CO%3B2-0;
http://www.jstor.org/stable/686307. See [Kes65a, Kes65b].

Dingle:1966:TCC

[Din66] Herbert Dingle. Twentieth-Century cosmology. British Journal for the History of Science, 3(2):184–187, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dingle:1972:SC

[Din72] Herbert Dingle. Science at the crossroads. Martin, Brian and O’Keefe Ltd., London, UK, 1972. ISBN 0-85616-060-1. 256 pp. LCCN Q175 .D6349. This book is critical of the Special Theory of Relativity. See refutation in [McC77].

Dinnage:1993:PLA

[Din93] R. Dinnage. The private lives of Albert Einstein — R. Highfield, P. Carter. TLS — The Times Literary Supplement, ??(4733):8, December 17, 1993. ISSN 0307-661X.

Dirac:1949:FRD

[Dir49] P. A. M. Dirac. Forms of relativistic dynamics. Reviews of Modern Physics, 21(3):392–399, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.392;
http://rmp.aps.org/abstract/RMP/v21/i3/p392_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Dirac:1960:GGGb

[Dir60a] P. A. M. Dirac. Gravitationswellen (German) [Gravitational waves]. Naturwissenschaftliche Rundschau, 13(??):165–168, ???? 1960. CODEN NARSAC. ISSN 0028-1050. Unabridged version of Lindau lecture given at the 9th Lindau Meeting, June 29–July 3, 1960.

Dirac:1960:GGGa

[Dir60b] P. A. M. Dirac. Gravitationswellen: Nach einem Vortrag (German) [Gravitational waves: After a lecture]. Physikalische Blätter, 16(7):364–366, July 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/phbl.19600160703/

abstract. Abridged version of Lindau lecture given at the 9th Lindau Meeting, June 29–July 3, 1960.

Dirac:1965:TEP

[Dir65] P. A. M. Dirac. Theory of electrons and positrons. In Anonymous [Ano65b], pages 320–225. LCCN QC71 .P455

1965. URL http://nobelprize.org/nobel_prizes/physics/
laureates/1922/bohr-lecture.html;   http://nobelprize.
org/nobel_prizes/physics/laureates/1922/bohr-lecture.

pdf. Nobel Prize in Physics lecture, December 12, 1933.

Dirac:1967:PQM

[Dir67] P. A. M. (Paul Adrien Maurice) Dirac. The principles of quantum mechanics, volume 27 of International series of monographs on physics (Oxford, England). Clarendon Press, Oxford, UK, fourth revised edition, 1967. ISBN 0-19-852011-5 (paperback). xii + 314 pp. LCCN ????

Dirac:1975:GTR

[Dir75] P. A. M. (Paul Adrien Maurice) Dirac. General Theory of Relativity. Wiley, New York, NY, USA, 1975. ISBN 0-471-21575-9. viii + 69 pp. LCCN QC173.6 .D57.

Dirac:1979:BET

[Dir79a] P. Dirac. Beauty of Einstein’s thought. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 101(6):46–47, ???? 1979. CODEN MEENAH. ISSN 0025-6501 (print), 1943-5649 (electronic).

Dirac:1979:EDI

[Dir79b] P. Dirac. Einstein, definite idea about science. Folia Humanistica, 17(196):199–201, ???? 1979. ISSN 0015-5594.

Dirac:1979:ECC

[Dir79c] P. Dirac. Einstein, or a certain conception of science. Chimia, 33(9):346–347, ???? 1979. ISSN 0009-4293.

Dirac:1979:DET

[Dir79d] P. A. M. Dirac. Developments of Einstein’s theory of gravitation. In Perlmutter and Scott [PS79], pages 1–13. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178 .O72 1979. URL http://adsabs.

harvard.edu/abs/1979opae.conf....1D;   http://link.
springer.com/chapter/10.1007/978-1-4684-3596-2_1/;
http://link.springer.com/content/pdf/bfm:978-1-4684-
3596-2/1.pdf.

Dirac:1979:EET

[Dir79e] P. A. M. Dirac. The excellence of Einstein’s theory of gravitation. Impact of Science on Society [UNESCO], 29(??):11–14, 1979. CODEN ISSOA8. ISSN 0019-2872. Reprinted in [Dir80c].

Dirak:1979:PET

[Dir79f] Pol A. M. Dirak. The perfection of Einstein’s theory of gravitation. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 22(55)(3):216–218, 1979. CODEN FMBMAC. ISSN 0015-3265. Translated from the French by D. Vačov.

Dirac:1980:ESP

[Dir80a] P. A. M. Dirac. Einstein session of the Pontifical Academy. Science, 207(4436):1161–1162, ???? 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Dirac:1980:EIP

[Dir80b] P. A. M. Dirac. Einstein’s influence on physics. In Bucciarelli

[Buc80], pages 19–23. URL http://books.google.com/books?
id=pnpkYgEACAAJ.

Dirac:1980:EET

[Dir80c] P. A. M. Dirac. Excellence of Einstein’s Theory of Gravitation. In Goldsmith et al. [GMW80], pages 41–46. ISBN 0-08-025019-X. LCCN QC16.E5 E52 1980. Reprint of [Dir79e].

Dirac:1980:WWB

[Dir80d] P. A. M. Dirac. Why we believe in the Einstein theory. In Gruber and Millman [GM80], pages 1–11. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

Dirac:1981:EDP

[Dir81a] P. A. M. Dirac. Einstein and the development of physics. In Colette M. Kinnon, A. N. Kholodilin, and J. G. Richardson, editors, The Impact of Modern Scientific Ideas on Society: in commemoration of Einstein: papers presented at the UNESCO symposium on the impact of modern scientific ideas on society, Munich–Ulm, 18–20 September 1978, and the addresses delivered on the occasion of UNESCO’s celebration of the hundredth anniversary of Einstein’s birth, Paris, 9 May 1979, page ?? D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1238-7. LCCN Q175.4 .I54.

Dirac:1981:PQM

[Dir81b] P. A. M. (Paul Adrien Maurice) Dirac. The principles of quantum mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised edition, 1981. ISBN 0-19-852011-5. xii + 314 pp. LCCN QC6 .D55 1981M.

Dirac:1982:EYR

[Dir82] P. A. M. Dirac. The early years of Relativity. In Holton and Elkana [HE82], pages 79–90. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Dirac:1987:GGG

[Dir87] P. A. M. Dirac. Gravitationswellen (German) [Gravitational waves]. In R. Schmid and W. Wessinger, editors, Nobel Prize Winners, pages 232–235. ????, ????, 1987. ISBN ???? LCCN ???? Reprint of [Dir60b, Dir60a].

Dirac:1996:GTR

[Dir96] P. A. M. (Paul Adrien Maurice) Dirac. General Theory of Relativity. Princeton landmarks in mathematics and physics; Princeton paperbacks. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01146-X (paperback). viii + 69 pp. LCCN QC173.6 .D57 1996. URL http:/

/press.princeton.edu/titles/5813.html;   http://www.
loc.gov/catdir/description/prin021/95046196.html;
http://www.loc.gov/catdir/toc/prin031/95046196.html.

Dirac:2001:LQM

[Dir01] P. A. M. (Paul Adrien Maurice) Dirac. Lectures on quantum mechanics. Dover, New York, NY, USA, 2001. ISBN 0-486-41713-1 (paperback). v + 87 pp. LCCN QC174.125

.D55 2001. URL http://www.loc.gov/catdir/description/
dover032/00065608.html.

Dirac:2003:PQM

[Dir03] Paul Adrien Maurice Dirac. The principles of quantum mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised reprinted edition, 2003. ISBN 0-19-852011-5. xii + 314 pp. LCCN QC6

.D55 2003. URL 04; 2;   http://www.gbv.de/dms/bowker/
toc/9780198520115..

Dirac:2009:PQM

[Dir09] Paul Adrien Maurice Dirac. The principles of quantum mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised reprinted edition, 2009. ISBN 0-19-852011-5 (paperback). xii + 314 pp. LCCN QC6 .D55 2009.

DiSalle:1992:ENE

[DiS92] Robert DiSalle. Einstein, Newton and the empirical foundations of spacetime geometry. International Studies in the Philosophy of Science, 6(3):181–189, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.

tandfonline.com//doi/abs/10.1080/02698599208573429.

DiSalle:2006:UST

[DiS06] Robert DiSalle. Understanding Space–time: The Philosophical Development of Physics from Newton to Einstein. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85790-2. xiii + 173 pp. LCCN QC173.59.S65 D57 2006. UK£45.00.

URL http://www.cambridge.org/uk/catalogue/catalogue.
asp?isbn=0521857902.

Divine:1978:BWN

[Div78] Robert A. Divine. Blowing on the Wind: the Nuclear Test Ban Debate, 1954–1960. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1978. ISBN 0-19-502390-0. ix + 393

pp. LCCN E835 .D53. URL http://catdir.loc.gov/catdir/
toc/fy053/77025057.html.

Dixon:2000:BRS

[Dix00] Henry B. F. Dixon. Book review: The Soviet bomb: Veniamin Tsukerman and Zinaida Azarkh, Arzamas-16. Soviet scientists in the Nuclear Age: a memoir. Translated by Timothy Sergay. Edited by Michael Pursglove. Bramcote Press, Nottingham, 1999. Pp. xxxiv + 182, £13.95 (paperback). ISBN 1-900405-04-0. Notes and Records of the Royal Society of London, 54(1): 124–125, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532068.

Dizikes:2005:PEW

[Diz05] Peter Dizikes. Picturing Einstein: Why the iconic images of Albert Einstein as an aging, eccentric genius distort our understanding not only of the scientist, but of science itself. Boston Globe, ??(??):??, March 20, 2005. ISSN 0743-1791. URL https:

//search.proquest.com/pqdtglobal/docview/404949356/.

Dahl:2015:AJF

[DL15] Michael Dahl and F. (Fabio) Leone. Albert is just fine, Mrs. Einstein! Science Biographies. Cantata Learning, Minneapolis, MN, USA, 2015. ISBN 1-63290-083-1 (hardcover), 1-63290-131-5 (e-book), 1-63290-053-X (audio). ???? pp. LCCN ????

deMuynck:1986:RBE

[dM86] Willem M. de Muynck. On the relation between the Einstein– Podolsky–Rosen paradox and the problem of nonlocality in quantum mechanics. Foundations of Physics, 16(10):973–1002, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01892937.

Dauber:1996:TBB

[DM96] Philip M. Dauber and Richard A. Muller. The Three Big Bangs: Comet Crashes, Exploding Stars, and the Creation of the Universe. Helix books. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-15495-1 (paperback), 0-201-40752-3 (hardcover). viii + 207 + 16 pp. LCCN QB991.B54.D38.

Dvoeglazov:2015:EU

[DM15] Valeri V. Dvoeglazov and Alberto Molgado, editors. Einstein and others: unification. Contemporary fundamental physics.

Nova Publishers, New York, NY, USA, 2015. ISBN 1-63463-276-1 (hardcover), 1-63463-278-8 (e-book). x + 227 pp. LCCN

QC174.45 .E56 2015. URL https://www.novapublishers.
com/catalog/product_info.php?products_id=52058.

deMendoza:2005:PRL

[dMdA05] Diego Hurtado de Mendoza and Miguel de Asúa. The poetry of Relativity: Leopoldo Lugones’ The Size of Space. Science in Context, 18(2):309–315, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

DeVolpi:1981:BSH

[DMPS81] A. DeVolpi, G. E. Marsh, T. A. Postol, and G. S. Stanford, editors. Born secret: the H-bomb, the Progressive case and national security. Pergamon policy studies. Pergamon, New York, NY, USA, 1981. ISBN 0-08-025995-2. xiii + 305 pp. LCCN KF228.U5 B67 1981.

Dobson:2006:IEB

[Dob06] Roger Dobson. Inside Einstein’s brain: Of course it was bigger than yours. But now we can show it was a different shape, too. Independent, ??(??):??, June 3, 2006. URL http://

www.independent.co.uk/news/science/inside-einsteins-
brain-624651.html.

Doherty:2006:BGW

[Doh06] P. C. (Peter C.) Doherty. The beginner’s guide to winning the Nobel Prize: a life in science. Columbia University Press, New York, NY, USA, 2006. ISBN 0-231-13896-2. xxi + 294 pp. LCCN ????

Dolbeault:2018:EUB

[Dol18] Joël Dolbeault. Einstein et l’univers-bloc. (French) [Einstein and the block universe]. Revue d’Histoire des Sciences, 71(1):79–109, ???? 2018. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL https://www.cairn.info/article.
php?ID_ARTICLE=RHS_711_0079.

Donnay:1964:LEP

[Don64] J. D. H. Donnay. Letter to the Editor: Pauling Prize. Bulletin of the Atomic Scientists, 20(3):25, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano63b].

Donghi:1970:AOBb

[Don70] Leticia Halperin Donghi. Analyses d’ouvrages: Nobel Foundation, Nobel Lectures, including presentation speeches and laureates’ biographies, Chemistry, 1901–1921. Revue d’Histoire des Sciences et de Leurs Applications, 23(4):374, October 1970. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23904064.

Donkin:1997:CPS

[Don97] Andrew Donkin. The cosmic professor: the story of Albert Einstein. Macdonald Young, Hove, UK, 1997. ISBN 0-7500-2300-7, 0-7500-2304-X (paperback). 48 pp. LCCN ????

Dorsey:1944:VL

[Dor44] N. Ernest Dorsey. The velocity of light. Transactions of the American Philosophical Society, New Series, 34(1):1–110, October 1944. CODEN TAPSAY. ISSN 0065-9746. URL http://

www.jstor.org/stable/1005532. This lengthy review reports

on work on determining the velocity of light in work by Fizeau (1849), Foucault (1862), Cornu (1872–1874), Perrotin and Prim (1908), Newcomb (1880–1882), Michelson (1878–1935), Karolus and Mittelstaedt (1929), Anderson (1937–1941), and Hüttel (1940).

Dorling:1968:LCC

[Dor68] Jon Dorling. Length contraction and clock synchronisation: the empirical equivalence of the Einsteinian and Lorentzian theories. The British Journal for the Philosophy of Science, 19(1): 67–69, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/19/1/67.full.pdf+html.

Dorling:1971:EIP

[Dor71] Jon Dorling. Einstein’s introduction of photons: Argument by analogy or deduction from the phenomena? The British Journal for the Philosophy of Science, 22(1):1–8, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
22/1/1.full.pdf+html.

Dorling:1978:DEN

[Dor78] Jon Dorling. Did Einstein need General Relativity to solve the problem of absolute space? Or had the problem already been solved by Special Relativity? The British Journal for the Philosophy of Science, 29(4):311–323, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/29/4/311.
full.pdf+html.

Dore:1986:WWJ

[Dor86] Ron Dore. Where will the Japanese Nobel Prizes come from? Science and Public Policy, 13(6):347–361, December 1986. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dorato:1995:TRS

[Dor95] Mauro Dorato. Time and reality: spacetime physics and the objectivity of temporal becoming, volume 11 of Collana di studi epistemologici. CLUEB, Bologna, 1995. ISBN 88-8091-172-4. xiii + 235 pp. LCCN QC6.4.R42 D67 1995.

Dorries:2005:MFC

[Dör05] Matthias Dörries, editor. Michael Frayn’s Copenhagen in debate: historical essays and documents on the 1941 meeting between Niels Bohr and Werner Heisenberg, volume 20, 33 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California, Berkeley, CA, USA, 2005. ISBN 0-9672617-2-4. viii + 195 pp. LCCN PR6056.R3 C6636 2005; PR6056.R3 C6436 2005; PR6056.R3 C6765 2005.

Dorling:2010:PMP

[Dor10] Danny Dorling. Putting men on a pedestal: Nobel prizes as superhuman myths? Significance (Oxford, England), 7(3):142– 144, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

DosSantos:2017:EE

[Dos17] José Rodrigues Dos Santos. Das Einstein Enigma. (German) [The Einstein Enigma]. RealMystery, Pohlheim, German, 2017. ISBN 3-946621-00-7. 512 pp. LCCN ????

Doty:1982:EIS

[Dot82] Paul Doty. Einstein and international security. In Holton and Elkana [HE82], pages 347–368. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Dowling:1980:FRBb

[Dow80] John Dowling. Film reviews: An Act of Congress; Growth Dilemma, produced by Richard Whitacher and Christopher Aikenhead; Einstein, produced by Patrick Griffin. Bulletin of the Atomic Scientists, 36(3):55, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

dePater:1994:BRB

[dP94] C. de Pater. Book review: Inventory of the Papers of Pieter Zeeman (1865–1943), Physicist and Nobel Prize Winner: c. 1877–1946 by P. J. M. Velthuys-Bechthold. Isis, 85(3):557– 558, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
235543.

dePadova:2017:AGS

[dP17] Thomas de Padova. Allein gegen die Schwerkraft: Einstein 1914–1918. (German) [Alone against the power of gravity: Einstein 1914–1918]. Piper, München, Germany, 2017. ISBN 3-492-31028-1. 310 pp. LCCN ????

Duck:1997:PSS

[DPS97] Ian Duck, Wolfgang Pauli, and E. C. G. Sudarshan. Pauli and the spin-statistics theorem. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-3114-8. x + 512 pp. LCCN QC793.3.S6 D83

1997. URL http://www.zentralblatt-math.org/zmath/en/
search/?an=0941.81002.

Duffield:1920:DSL

[DR20] W. G. Duffield and M. N. Roy. The displacement of spectrum lines and the equivalence hypothesis. Astron. Soc., LXXX(3): 262–??, 1920.

DeMaria:1989:CRR

[DR89] M. De Maria and A. Russo. Cosmic ray romancing: The discovery of the latitude effect and the Compton–Millikan controversy. Historical Studies in the Physical and Biological Sciences, 19(2):211–266, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.jstor.org/

stable/27757626.

Dasgupta:2008:MDQ

[DR08] Ratan Dasgupta and Sisir Roy. Multinomial distribution, quantum statistics and Einstein–Podolsky–Rosen like phenomena. Foundations of Physics, 38(4):384–394, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
008-9207-3.

Dolev:2016:CPT

[DR16] Yuval Dolev and Michael Roubach. Cosmological and Psychological Time, volume 285 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346. xiv + 218 pp. LCCN BD638 C67 2016. URL https://link.

springer.com/book/10.1007/978-3-319-22590-6.

Drake:1953:GGD

[Dra53] Stillman Drake, editor. Galileo Galilei: Dialogue concerning the two chief world systems, Ptolemaic and Copernican. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1953. xxvii + 496 pp. LCCN QB41 .G1356. Translation by Stillman Drake of Galileo’s original work. Foreword by Albert Einstein.

Drake:2001:GGD

[Dra01] Stillman Drake, editor. Galileo Galilei: Dialogue concerning the two chief world systems, Ptolemaic and Copernican. Modern Library science series. Modern Library, New York, NY, USA, 2001. ISBN 0-375-75766-X (paperback). xxxvii + 586 pp. LCCN QB41 .G129413 2001. URL http://www.loc.

gov/catdir/bios/random053/2001030842.html;   http://
www.loc.gov/catdir/description/random047/2001030842.
html;   http://www.loc.gov/catdir/samples/random043/

2001030842.html. Translation by Stillman Drake of Galileo’s original work. Foreword by Albert Einstein. Introduction by J. L. Heilbron.

Dray:2012:GSR

[Dra12] Tevian Dray. The Geometry of Special Relativity. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2012. ISBN 1-4665-1047-1. xvii + 131 pp. LCCN QC173.65 .D73 2012.

Drago:2014:ERP

[Dra14a] Antonino Drago. Einstein’s 1905 “revolutionary” paper on quanta as a manifest and detailed example of a “principle theory”. Advances in Historical Studies, 3(??):130–154, ???? 2014. ISSN 2327-0438 (print), 2327-0446 (electronic). URL http://

www.oalib.com/references/7981965.

Dray:2014:DFG

[Dra14b] Tevian Dray. Differential Forms and the Geometry of General Relativity. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-1000-5. ???? pp. LCCN QC173.6 .D73 2014.

Dunningham:2005:PCF

[DRB05] Jacob Dunningham, Alexander Rau, and Keith Burnett. From pedigree cats to fluffy-bunnies. Science, 307(5711):872–875, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Dresden:1987:HKB

[Dre87] Max Dresden. H. A. Kramers: between tradition and revolution. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96282-4. xxiv + 563 pp. LCCN QC16.K69 D74 1987. US$57.00.

Drell:1993:SBP

[Dre93] Sidney D. (Sidney David) Drell. In the shadow of the bomb: physics and arms control, volume 6 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 1-56396-058-3. xx + 358 pp. LCCN QC773.3.U5 D74 1993.

Drell:2007:NWS

[Dre07] Sidney D. (Sidney David) Drell. Nuclear weapons, scientists, and the post-Cold War challenge: selected papers on arms control. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-256-896-4, 981-256-897-2 (paperback). vi + 323 pp. LCCN U264 .D567 2007.

Drory:2015:NSP

[Dro15] Alon Drory. The necessity of the second postulate in special relativity. Studies in History and Philosophy of Modern Physics, 51 (??):57–67, August 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219814001038.

Dronsfield:2020:BRT

[Dro20] Alan Dronsfield. Book review: Traveling with the Atom: a Scientific Guide to Europe and Beyond. Web site, March 20, 2020.

URL https://www.chemistryworld.com/review/traveling-
with-the-atom-a-scientific-guide-to-europe-and-beyond/
4011154.article.

Dry:2014:NPS

[Dry14] Sarah Dry. The Newton Papers: the Strange and True Odyssey of Isaac Newton’s Manuscripts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-995104-7 (hardcover). xi + 238 pp. LCCN QC16.N7 D79 2014.

deSitter:1919:STG

[dS19] W. de Sitter. Space, time and gravitation. The Observatory, 505:412–??, 1919. CODEN OBSEAR. ISSN 0029-7704.

Dirac:1983:ANP

[DS83a] Paul Dirac and Erwin Schrödinger. 50th anniversary of the Nobel Prize for Schrödinger and Dirac. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 25(58)(4):277–297, 1983. CODEN FMBMAC. ISSN 0015-3265. Translated from the French and German by D. Vachov.

Drell:1983:FTN

[DS83b] Sidney D. (Sidney David) Drell and Andrei Sakharov. Facing the threat of nuclear weapons. The Jessie and John Danz lectures. University of Washington Press, Seattle, WA, USA, 1983. ISBN 0-295-96082-5, 0-295-96083-3 (paperback). x + 120 pp. LCCN JX1974.7 .D85 1983.

Daigneault:2001:ESU

[DS01] Aubert Daigneault and Arturo Sangalli. Einstein’s static universe: an idea whose time has come back? Notices of the American Mathematical Society, 48(1):9–16, 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

dosSantos:2010:EE

[dSL10] José Rodrigues dos Santos and Carter Lisa. The Einstein enigma. William Morrow, New York, NY, USA, 2010. ISBN 0-06-171924-2 (hardcover). 485 pp. LCCN PQ9319.A685 F6713 2010.

Diamond:1985:BSA

[DSMH85] Marian C. Diamond, Arnold B. Scheibel, Greer M. Murphy, Jr., and Thomas Harvey. On the brain of a scientist — Albert Einstein. Experimental Neurology, 88(1):198–204, April 1985. CODEN EXNEAC. ISSN 0014-4886 (print), 1090-2430 (electronic).

deSitter:1916:ETGa

[dSR16a] W. de Sitter and M. N. Roy. On Einstein’s Theory of Gravitation and its astronomical consequences. Astron. Soc., LXXVI: 699–??, 1916.

deSitter:1916:ETGb

[dSR16b] W. de Sitter and M. N. Roy. On Einstein’s Theory of Gravitation and its astronomical consequences. Astron. Soc., LXXVII: 155–??, 1916.

deSitter:1917:ETG

[dSR17] W. de Sitter and M. N. Roy. On Einstein’s Theory of Gravitation and its astronomical consequences. Astron. Soc., LXXVIII: 3–??, 1917.

Drell:1999:NTF

[DSW99] Sidney D. (Sidney David) Drell, Abraham D. Sofaer, and George D. Wilson, editors. The new terror: facing the threat of biological and chemical weapons. Hoover national security forum series. Hoover Institution Press, Stanford, CA, USA, 1999. ISBN 0-8179-9701-6, 0-8179-9702-4 (paperback). xxx + 512 pp. LCCN UG447.8 .N48 1999.

Das:2012:WST

[DSZ12] Arnab Das, Jacopo Sabbatini, and Wojciech H. Zurek. Winding up superfluid in a torus via Bose–Einstein condensation open. Scientific Reports, 2, April 12, 2012. CODEN

SRCEC3. ISSN ???? URL http://www.nature.com/srep/
2012/120411/srep00352/full/srep00352.html.

Dyson:1923:CEP

[DT23] Frank Dyson and H. H. Turner. The confirmation of the Einstein prediction. Science, 57(1481):571–573, May 18, 1923. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/57/
1481/571.full.pdf.

Davis:1997:JJT

[DTF97] E. A. (Edward Arthur) Davis, J. J. (Joseph John) Thomson, and I. J. (Isobel J.) Falconer. J. J. Thomson and the discovery of the electron. Taylor and Francis, London, UK and Boca Raton, FL, USA, 1997. ISBN 0-7484-0696-4 (hardcover). xxvii + 243 pp. LCCN QC16.T45 D38 1997.

DuBois:1952:DCC

[Du 52] Josiah Ellis Du Bois, Jr. The Devil’s chemists: 24 conspirators of the international Farben cartel who manufacture wars. Beacon Press, Boston, MA, USA, 1952. 374 pp. LCCN HD9654.9.I5 D8. In collaboration with Edward Johnson.

Dubiel:2005:ESM

[Dub05] Stanisław Dubiel. Einstein set to music. Physics World, 18(12):20, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/12/phwv18i12a27.
pdf.

Ducheyne:2011:TUG

[Duc11] Steffen Ducheyne. Testing universal gravitation in the laboratory, or the significance of research on the mean density of the Earth and big G, 1798–1898: changing pursuits and long-term methodological-experimental continuity. Archive for History of Exact Sciences, 65(2):181–227, March 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=2&spage=181.

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??):BR1–BR2, March 17, 1946. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/107455627.

Duffy:1981:AEF

[Duf81] M. C. Duffy. Albert Einstein — four commemorative lectures — L. S. Swenson, C. P. Snow, H. Stein, and I. Prigogine. Annals of Science, 38(3):354–357, ???? 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Dukas:1949:SAE

[Duk49] Helen Dukas. Supplement to addenda to Einstein’s writings. In Schilpp [Sch49c], pages 758a–758b. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Dume:2004:LBT

[Dum04] Belle Dumé. Low-budget test vindicates Einstein. Physics World, 17(11):7, November 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/17/11/phwv17i11a7.
pdf.

Dupre:2006:ESD

[Dup06] Mhairi Dupre. Einstein’s secret diary? Nature, 441(7094):782, June 7, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/naturejobs/
2006/060608/full/nj7094-782c.html.

Durrenmatt:1962:PKZ

[Dür62] Friedrich Dürrenmatt. Die Physiker: Eine Komödie in zwei Akten. (German) [The Physicists: A comedy in two acts]. Verlag Die Arche, Zürich, Switzerland, 1962. 71 pp. LCCN ????

Durrenmatt:1979:AE

[Dür79] Friedrich Dürrenmatt. Albert Einstein. Diogenes, Zürich, Switzerland, 1979. ISBN 3-257-20729-8 (paperback). 65 + 1 pp. LCCN QC16.E5 D78.

Durrenmatt:1980:PKZ

[Dür80] Friedrich Dürrenmatt. Die Physiker: Eine Komödie in zwei Akten. Neufassung 1980. (German) [The Physicists: A comedy in two acts. Revised 1980]. Diogenes, Zürich, Switzerland, 1980. ISBN 3-257-20837-5. 93 pp. LCCN ????

Durrenmatt:1982:AEC

[Dür82] Friedrich Dürrenmatt. Albert Einstein: una conferencia. (Spanish) [Albert Einstein: a conference], volume 75 of Marginales.

Tusquets Editores, Barcelona, Spain, 1982. ISBN 84-7223-075-9. 60 pp. LCCN ????

Durrani:2002:FFW

[Dur02] Matin Durrani. Famous first words: Einstein’s correspondence with children. Physics World, 15(12):41, December 2002. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/15/12/phwv15i12a34.pdf.

Durell:2003:RR

[Dur03] Clement V. (Clement Vavasor) Durell. Readable Relativity. Dover, New York, NY, USA, 2003. ISBN 0-486-43257-2 (paperback). xi + 146 pp. LCCN QC173.57 .D87 2003. URL http:

//www.loc.gov/catdir/enhancements/fy0614/2003055297-
d.html.

Durrani:2005:BHA

[Dur05a] M. Durrani. A brief history of Albert Einstein. Physics World, 18(1):14–15, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Durrani:2005:BE

[Dur05b] Matin Durrani. The best of Einstein. Physics World, 18(12):41, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/12/phwv18i12a34.pdf.

Durrenmatt:2006:PCT

[Dür06] Friedrich Dürrenmatt. The physicists: a comedy in two acts. Grove Press, New York, NY, USA, 2006. ISBN 0-8021-4427-6. 76 pp. LCCN PT2607.U493 P4913 2006. Translated by Joel Agee from the German original, Die Physiker (The Physicists), published by Verlag der Arche, Zürich 1962 and revised version by Diogenes Verlag AG, Zürich 1980.

Durrani:2008:SAE

[Dur08] Matin Durrani. Seeking an African Einstein. Physics World, 21 (7):11, July 2008. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/21/07/phwv21i07a22.pdf.

Durand:2012:EDP

[Dur12] Roger Durand. Élie Ducommun: Prix Nobel de la paix méconnu: famille, politique, éeconomie, humanitaire, pacifisme. Institut National Genevois, Genéve, Suisse, 2012. ISBN 2-940336-04-0. 254 pp. LCCN ???? With the collaboration of Serge Paquier.

Duveen:1964:NPC

[Duv64] Denis Duveen. Nobel Prizewinners in Chemistry 1901–1961 by Eduard Farber (review). Technology and Culture, 5(1):114, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/894942/pdf.

Dyson:1937:ESM

[DvdRW37] Frank Watson Dyson and Richard van der Riet Woolley. Eclipses of the sun and moon. The international series of monographs on physics. Clarendon Press, Oxford, UK, 1937. vii + 1 + 160 pp. LCCN QB541 .D85.

Heathcote:1953:NPW

[dVNHdVH53] Niels H. de V. (Niels Hugh de Vaudrey) Heathcote. Nobel Prize winners in physics, 1901-1950, volume 30 of The Life of science library. H. Schuman, New York, NY, USA, 1953. 473 pp. LCCN QC15 .H4. With a foreword by Herbert Dingle.

Heathcote:1971:NPW

[dVNHdVH71] Niels H. de V. (Niels Hugh de Vaudrey) Heathcote. Nobel Prize winners in physics, 1901–1950. Essay index reprint series. Books for Libraries Press, Freeport, NY, USA, 1971. ISBN 0-8369-2455-X. xvi + 473 pp. LCCN QC15 .H4 1971. With a foreword by Herbert Dingle.

Dvoeglazov:2012:EHD

[Dvo12] Valeri V. Dvoeglazov, editor. Einstein and Hilbert: dark matter. Nova Science Publishers, Hauppauge, NY, USA, 2012. ISBN 1-61324-840-7 (hardcover), 1-61324-887-3 (ebook). xi + 199 pp. LCCN QC178 .E337 2012. URL

https://www.novapublishers.com/catalog/product_info.
php?products_id=25609.

Dong-Won:1998:WMW

[DWL98] Kim Dong-Won and Stuart W. Leslie. Winning markets or winning Nobel prizes? Kaist and the challenges of late industrialization. Osiris (Series 2), 13(??):154–185, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301882.

deWaal:2020:CFG

[dWtH20] Elske de Waal and Sjang L. ten Hagen. The concept of fact in German physics around 1900: A comparison between Mach and Einstein. Physics in Perspective (PIP), 22(2):55–80, June 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-020-00256-y.pdf.

Dyson:1965:TSF

[Dys65] Freeman J. Dyson. Tomonaga, Schwinger, and Feynman awarded Nobel Prize for Physics. Science, 150(3696):588– 589, October 29, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1717147;   http://www.sciencemag.org/content/
150/3696/588.full.pdf.

Dyson:1976:HBD

[Dys76] F. J. Dyson. The hydrogen-bomb decision: a reappraisal: Review of The Advisors: Oppenheimer, Teller and The Superbomb, by Herbert F. York. Science, 193(4254):668–669, August 20, 1976. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/193/4254/668.full.pdf.

Dyson:1979:DU

[Dys79] Freeman J. Dyson. Disturbing the Universe. Harper & Row, New York, NY, USA, 1979. ISBN 0-06-011108-9. x + 283 pp. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1980:MA

[Dys80] Freeman J. Dyson. Manchester and Athens. Talk given to Nobel Conference at Gustavus Adolphus College, St. Peter, Minnesota., October 8, 1980.

Dyson:2009:BFE

[Dys09] Freeman J. Dyson. Birds and frogs: 2008 Einstein Lecture. Notices of the American Mathematical Society, 56(2):212–223, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/
200902/rtx090200212p.pdf.

Dyson:2015:EJP

[Dys15] Freeman Dyson. Einstein as a Jew and a philosopher: Einstein: His Space and Times by Steven Gimbel, Yale University Press, 191 pp., $25.00. New York Review of Books, 62 (8):??, May 7, 2015. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2015/may/07/albert-einstein-jew-and-philosopher/.

Dyson:2016:UPL

[Dys16] Freeman Dyson. Unfashionable pursuits, leaps in the dark, and detecting gravitational waves: How the recent discovery of spacetime ripples began as a risky venture. The Institute Letter,

??(??):1, 3, Spring 2016. URL https://www.ias.edu/ideas/
2016/dyson-gravitational-waves.

Einshtein:2003:BFR

[È03] Al′bert Èĭnshteĭn. Bez formul. (Russian) [Without formula]. Mysl′, Moscow, Russia, 2003. ISBN 5-244-00981-8. 224 pp. Edited by K. A. Kedrov.

Einstein:1922:MRF

[EA22] Albert Einstein and Edwin P. (Edwin Plimpton) Adams. The Meaning of Relativity: four lectures delivered at Princeton University, May, 1921. Methuen and Co. Ltd., London, UK, 1922. v + 123 pp. LCCN QC6.

Einstein:1923:MRF

[EA23] Albert Einstein and Edwin P. (Edwin Plimpton) Adams. The Meaning of Relativity: four lectures delivered at Princeton University, May, 1921. Princeton University Press, Princeton, NJ, USA, 1923. 3 + 123 + 1 pp. LCCN QC173.58; QC6.

Einstein:1967:MRb

[EA67] Albert Einstein and Edwin P. (Edwin Plimpton) Adams. The Meaning of Relativity. Science Paperbacks, London, UK, 1967. xiv + 161 pp. LCCN QC173.55.

Earman:2001:LCR

[Ear01] John Earman. Lambda: The constant that refuses to die. Archive for History of Exact Sciences, 55(3):189–220, January 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=3&spage=189.

Einstein:1938:GKT

[EB38] Albert Einstein and Peter Bergmann. On a generalization of Kaluza’s theory of electricity. Annals of Mathematics (2), 39(3):683–701, July 1938. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL 2-9; http:

//links.jstor.org/sici?sici=0003-486X(193807)2:39:3<
683:OAGOKT>2.0.CO.

Einstein:1944:BF

[EB44] Albert Einstein and Valentine Bargmann. Bivector fields. Annals of Mathematics (2), 45:1–14, January 1944. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL 2-7;   http://links.jstor.org/sici?sici=
0003-486X(194401)2:45:1<1:BF>2.0.CO.

Einstein:1972:C

[EB72] Albert Einstein and Michèle Angelo Besso. Correspondance 1903–1955. Hermann, Paris, France, 1972. lxiii + 559 pp. LCCN QC 16 .E5 A4212. Traduction, notes et introduction de Pierre Speziali, Avec des notes aux passages traitant de la relativité, par Marie-Antoinette Tonnelat, Collection Histoire de la Pensée, 17.

Einstein:1982:BGC

[EB82] Albert Einstein and Max Born. Briefwechsel: 1916–1955. (German) [Correspondence: 1916–1955]. Edition Erbrich, Frankfurt am Main, Germany, 1982. ISBN 3-88682-005-X. 330 pp. LCCN QC16.E5 A25 1982. With contributions by Hedwig Born, with an introduction by Bertrand Russell, and with a foreword by Werner Heisenberg. Reprint of the 1969 edition.

Einstein:1989:CF

[EB^{+}89] Albert Einstein, Michel Biezunski, et al. Correspondances françaises, volume 4 of Sources du savoir; Oeuvres choisies / Albert Einstein. Éditions du Seuil, Paris, France, 1989. ISBN 2-02-010177-7 (Seuil), 2-222-04332-8 (CNRS), 2-02-011390-2 (Collection). 345 + 4 pp. LCCN QC16.E5 A612 B34. Translation to French by Emanuelle Aurenche from German and English.

Einstein:1941:FDR

[EBB41] A. Einstein, V. Bargmann, and P. G. Bergmann. On the fivedimensional representation of gravitation and electricity. In Theodore von Kármán Anniversary Volume, pages 212–225. California Institute of Technology, Pasadena, CA, USA, 1941.

Einstein:1989:QMS

[EBD^{+}89] Albert Einstein, Françoise Balibar, Olivier Darrigol, Bruno Jech, et al. Quanta: mécanique statistique et physique quantique. (French) [Quanta: Statistical mechanics and quantum physics], volume 1 of Sources du savoir; Oeuvres choisies / Albert Einstein. Éditions du Seuil, Paris, France, 1989. ISBN 2-02-010027-4 (Seuil), 2-222-04333-6 (CNRS), 2-02-011390-2 (collection). 270 + 2 pp. LCCN QC16.E5 A612 B34. Translation by Françoise Balibar from German to French.

Einstein:1993:ECS

[EBRC93] Albert Einstein, Mara Beller, Jürgen Renn, and R. S. (Robert Sonné) Cohen, editors. Einstein in context: a special issue of Science in Context, Science in context. [Special issue]. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44834-4 (paperback). LCCN QC16.E5 E512 1993.

Einstein:1952:ESG

[EBU^{+}52] Albert Einstein, Hans A. Bethe, Harold C. Urey, James Franck, Samuel Goudsmit, Cyril S. Smith, Arthur H. Compton, and William P. Murphy. Eminent scientists give their views on American visa policy. Bulletin of the Atomic Scientists, 8(7): 217–220, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Ebury:2016:BRJ

[Ebu16] Katherine Ebury. Book review: James Joyce, Science, and Modernist Print Culture: “The Einstein of English Fiction” by Jeffrey S. Drouin. Modernism/modernity, 23(3):698–699, 2016. CODEN ???? ISSN 1071-6068 (print), 1080-6601 (electronic).

Einstein:2000:EQE

[EC00] Albert Einstein and Alice Calaprice. The expanded quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-07021-0 (cloth). xliii + 407 pp. LCCN

QC16.E5 A25 2000. URL http://www.loc.gov/catdir/bios/
prin051/00026873.html;   http://www.loc.gov/catdir/
description/prin022/00026873.html;   http://www.loc.
gov/catdir/samples/prin031/00026873.html;   http://
www.loc.gov/catdir/toc/prin031/00026873.html.

Evans:2011:CEF

[ECEP11] Myron W. Evans, Stephen J. Crothers, Horst Eckardt, and Kerry Pendergast. Criticisms of the Einstein field equation: the end of 20th century physics. Cambridge International Science Publishers, Cambridge, UK, 2011. ISBN 1-907343-28-8, 1-907343-29-6 (e-book). v + 461 pp. LCCN QC173.6 .E93 2011.

Eckart:1970:CEP

[Eck70] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970),

1970. URL http://hdl.loc.gov/loc.mss/eadmss.ms001014.

Eckart:1973:CEP

[Eck73] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). US Library of Congress archival manuscript material (collec-

tion)., 1973. URL http://hdl.loc.gov/loc.mss/eadmss.
ms001014.

Eckert:1990:PDF

[Eck90] Michael Eckert. Primacy doomed to failure: Heisenberg’s role as scientific adviser for nuclear policy in the FRG. Historical Studies in the Physical and Biological Sciences, 21(1):29–58, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355

(electronic). URL http://www.jstor.org/stable/27757654.

Eckert:2001:WHC

[Eck01] Michael Eckert. Werner Heisenberg: controversial scientist. Physics World, 14(12):35–40, December 2001. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL

http://stacks.iop.org/2058-7058/14/i=12/a=31.

Einstein:2002:DPE

[ECS02] Albert Einstein, Alice Calaprice, and Robert Schulmann. Dear Professor Einstein: Albert Einstein’s letters to and from children. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-59102-015-8. 232 pp. LCCN QC16.E5 A4 2002. URL http:/

/www.loc.gov/catdir/toc/fy035/2002073570.html.

Esterson:2019:EWR

[ECS19] Allen Esterson, David C. Cassidy, and Ruth Lewin Sime. Einstein’s wife: the real story of Mileva Einstein-Marić. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-03961-3 (hardcover), 0-262-53897-0 (paperback). xxi + 313 pp. LCCN QC16.E52 E88 2019.

Eddington:1919:TEM

[Edd19] A. S. Eddington. The total eclipse of 29th May 1919, and the influence of gravitation on light. The Observatory, 505:??, March 1919. CODEN OBSEAR. ISSN 0029-7704.

Eddington:1920:RRT

[Edd20a] Sir Arthur Stanley Eddington. Report on the Relativity Theory of Gravitation. Fleetway Press, Ltd., London, UK, second edition, 1920. xi + 91 pp. LCCN QC178 .E3.

Eddington:1920:STG

[Edd20b] Sir Arthur Stanley Eddington. Space, time and gravitation: an outline of the General Relativity Theory. University Press, Cambridge, UK, 1920. vi + 1 + 218 + 1 + 1 pp. LCCN QC6 .E4.

Eddington:1921:ETG

[Edd21] Sir Arthur Stanley Eddington. Espace, temps et gravitation: la théorie de la relativité généralisée dans ses grandes lignes: exposé rationnel, suivi d’une étude mathématique de la théorie. (French) [Space, time, and gravitation: the General Theory of Relativity in its great directions: rational exposition followed by a mathematical study]. J. Hermann, Paris, France, 1921. xii + 262 + iv + 149 + 1 pp. LCCN QC173.59.S65 E3314 1921.

Eddington:1923:MTR

[Edd23] Sir Arthur Stanley Eddington. The Mathematical Theory of Relativity. Cambridge University Press, Cambridge, UK, 1923. ix + 247 pp. LCCN QC6 E21 1923.

Eddington:1924:CWE

[Edd24] A. S. Eddington. A comparison of Whitehead’s and Einstein’s formulæ. Nature, 113(2832):192, February 9, 1924. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v113/n2832/
pdf/113192a0.pdf.

Eddington:1926:ESD

[Edd26] A. S. Eddington. Einstein shift and Doppler shift. Nature, 117 (2933):86, January 16, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v117/n2933/pdf/117086a0.pdf.

Eddington:1928:NPW

[Edd28] Sir Arthur Stanley Eddington. The Nature of the Physical World, volume 1927 of Gifford lectures. Cambridge University Press, Cambridge, UK, 1928. xix + 361 + 1

pp. LCCN Q175 .E3. URL https://archive.org/details/
natureofphysical00eddi.

Eddington:1929:EFT

[Edd29a] A. S. Eddington. Einstein’s field-theory. Nature, 123(3095):280– 281, February 23, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v123/n3095/pdf/123280a0.pdf.

Eddington:1929:STG

[Edd29b] Sir Arthur Stanley Eddington. Space, time and gravitation: an outline of the General Relativity Theory. Cambridge University Press, Cambridge, UK, 1929. vi + 218 pp. LCCN QC6 .E4 1929.

Einstein:1929:TSC

[EdD29c] Albert Einstein and Th. de Donder. La théorie synthétique des champs. (French) [The synthetic theory of fields]. Revue général de l’électricité, 25(??):35–39, 1929. ISSN 0035-3116.

Eddington:1930:IES

[Edd30a] A. S. Eddington. On the instability of Einstein’s spherical world. Monthly Notices of the Royal Astronomical Society, 90(??):668– 678, May 1930. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic).

Eddington:1930:MTR

[Edd30b] Sir Arthur Stanley Eddington. The Mathematical Theory of Relativity. Cambridge University Press, Cambridge, UK, second edition, 1930. ix + 270 pp. LCCN QC6 E21 1930.

Eddington:1959:STG

[Edd59] Sir Arthur Stanley Eddington. Space, time, and gravitation: an outline of the General Relativity Theory, volume TB510 of Harper torchbooks, The science library. Harper, New York, NY, USA, 1959. 213 pp. LCCN QC6 .E4 1959.

Eddington:1966:STG

[Edd66] Sir Arthur Stanley Eddington. Space, time, and gravitation: an outline of the General Relativity Theory. Cambridge University Press, Cambridge, UK, 1966. vi + 218 pp. LCCN QC6 .E4 1966.

Eddington:1979:EEN

[Edd79] A. S. Eddington. Eddington on Einstein’s new theory. Nature, 278(5701):213–214, March 15, 1979. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v278/n5701/pdf/
278213a0.pdf.

Eddington:1987:STG

[Edd87] Sir Arthur Stanley Eddington. Space, time, and gravitation: an outline of the General Relativity Theory. Cambridge science classics. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33709-7 (paperback). xii + 218 pp. LCCN QC173.59.S65 E33 1987. URL http://www.loc.

gov/catdir/description/cam023/86033348.html;   http:
//www.loc.gov/catdir/toc/cam028/86033348.html. Reprint

of 1920 edition.

Eddington:1995:EE

[Edd95] Allen Boyce Eddington, editor. Essential Einstein. Pomegranate Artbooks, San Francisco, CA, USA, 1995. ISBN 0-87654-472-3 (softcover). ix + 120 pp. LCCN QC16.E5 E392 1995. Introduction by Alan Bisbort.

Eddington:2006:RRT

[Edd06] Sir Arthur Stanley Eddington. Report on the Relativity Theory of Gravitation. Dover Phoenix editions. Dover, New York, NY, USA, 2006. ISBN 0-486-45080-5. xi + 91 pp. LCCN QC178

.E3 2006. URL http://www.loc.gov/catdir/enhancements/
fy0629/2006042282-d.html.

Edgington:2014:RWE

[Edg14] Ryan H. Edgington. Range wars: the environmental contest for White Sands Missile Range. University of Nebraska Press, Lincoln, NE, USA, 2014. ISBN 0-8032-3844-4 (hardcover), 0-8032-5535-7 (paperback), 0-8032-5562-4 (PDF), 0-8032-5563-2 (ePub), 0-8032-5564-0 (mobi). xiii + 268 pp. LCCN GE155.N6 E34 2014.

Einstein:1915:ENAb

[EdH15a] Albert Einstein and Wander Johannes de Haas. Experimenteller Nachweis der Ampèreschen Molekularströme. (German) [Experimental proof of the Ampére molecular currents]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(8):152– 170, April 30, 1915. CODEN VDPEAZ. ISSN 0372-5448.

URL https://hdl.handle.net/2027/nyp.33433090827365?

urlappend=%3Bseq=170. See corrections [Ein15c, EdH15b]. See also Dutch translation in [EdH15c] and English translation in [EdH16b].

Einstein:1915:NUA

[EdH15b] Albert Einstein and Wander Johannes de Haas. Notiz zu unserer Arbeit: “Experimenteller Nachweis der Ampèreschen Molekularströme”. (German) [Notice on our work: “Experimental proof of the Ampére molecular currents”]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(22):420,

November 30, 1915. URL https://hdl.handle.net/2027/
nyp.33433090827365?urlappend=%3Bseq=436. See [EdH15a].

Einstein:1915:PBV

[EdH15c] Albert Einstein and Wander Johannes de Haas. Proefondervindelijk beijs voor het bestaan der moleculaire stroomen van Ampère. (Dutch) [Experimental proof of the Ampére molecular currents]. Verslag Akademi van wetenschappen, Amsterdam, series 4, 23(??):1449–1464, 1915. Dutch translation of [EdH15a]. See also English translation in [EdH16b].

Einstein:1915:EPE

[EdH16a] Albert Einstein and Wander Johannes de Haas. Experimental proof of the existence of Ampére’s molecular currents. Koninklijke Akademie van Wetenschappen te Amsterdam: Proceedings, 18:696–711, 1915–1916.

Einstein:1916:EPE

[EdH16b] Albert Einstein and Wander Johannes de Haas. Experimental proof of the existence of Ampére’s molecular currents. Verslag Akademi van wetenschappen, Amsterdam, series 4, 18(??): 696–711, 1916. English translation of [EdH15a]. See also Dutch translation in [EdH15c].

Elitzur:2005:QVQ

[EDK05] Avshalom C. Elitzur, S. (Shahar) Dolev, and N. (Nancy) Kolenda. Quo vadis quantum mechanics? Frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-22188-3 (hardcover), 3-540-26669-0. xiv + 421 pp. LCCN

QC174.12 .E446 2005. URL http://www.loc.gov/catdir/
enhancements/fy0662/2004116224-d.htm;   http://www.
loc.gov/catdir/enhancements/fy0817/2004116224-t.htm.

Einstein:1932:RBE

[EdS32] Albert Einstein and Willem de Sitter. On the relation between the expansion and the mean density of the universe. Proceedings of the National Academy of Sciences of the United States of America, 18(3):213–214, March 15, 1932. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.pnas.org/cgi/reprint/18/3/213.pdf.

Edsall:1982:BRB

[Eds82] John T. Edsall. Book review: The Nobel Duel: Two Scientists’ 21-year Race to Win the World’s Most Coveted Research Prize by Nicholas Wade. Isis, 73(3):484–485, September 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231514.

Edwards:2000:MMO

[Edw00] Harriet Edwards. Mileva Maric: The other Einstein. Library Journal, 125(18):150–??, November 1, 2000. CODEN

LIBJA7. ISSN 0363-0277. URL http://ezproxy.lib.utah.
edu/docview/19681358.

Einstein:1922:QBE

[EE22] Albert Einstein and Paul Ehrenfest. Quantentheoretische Bemerkungen zum Experiment von Stern und Gerlach. (German) [Quantum theoretical remarks on the experiment of Stern and Gerlach]. Zeitschrift für Physik, 11(??):31–34, ???? 1922. CO-DEN ZEPYAA. ISSN 0044-3328.

Einstein:1923:QSG

[EE23] Albert Einstein and Paul Ehrenfest. Zur Quantentheorie des Strahlungsgleichgewichts. (German) [On the quantum theory of the radiative equilibrium]. Zeitschrift für Physik, 19(??):301– 306, 1923. CODEN ZEPYAA. ISSN 0044-3328.

Ehrenfest:1959:CFS

[EE59] Paul Ehrenfest and Tatiana Ehrenfest. The conceptual foundations of the statistical approach in mechanics. Cornell University Press, Ithaca, NY, USA, 1959. xvi + 114 pp. LCCN QC175 .E353. Translated by Michael J. Moravcsik.

Esterer:1972:PAA

[EE72] Arnulf K. Esterer and Louise A. Esterer. Prophet of the atomic age: Leo Szilard. Julian Messner, New York, NY, USA, 1972. ISBN 0-671-32522-1 (hardcover), 0-671-32523-X (MCE). 189 pp. LCCN QC16.S95 E83.

Errera:1982:STE

[EE82] M. Errera and A. Einstein. Effets à court terme d’explosions nucléaires. (French) [Short term effects of nuclear explosions]. Bulletin et mémoires de l’Académie royale de médecine de Belgique, 137(8–9):560–568, ???? 1982. ISSN 0377-8231.

Ehrenfest:1990:CFS

[EE90] Paul Ehrenfest and Tatiana Ehrenfest. The conceptual foundations of the statistical approach in mechanics. Dover books on physics and chemistry. Dover, New York, NY, USA, 1990. ISBN 0-486-66250-0. xiv + 114 pp. LCCN QC175 .E353 1990. US$4.95.

Earman:1999:ES

[EE99] John Earman and Jean Eisenstaedt. Einstein and singularities. Studies in History and Philosophy of Modern Physics, 30(2): 185–235, June 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219899000052.

Eger:2005:EVC

[EE05] Joseph Eger and Albert Einstein. Einstein’s violin: a conductor’s notes on music, physics, and social change. Jeremy P. Tarcher/Penguin, New York, NY, USA, 2005. ISBN 1-58542-388-2. xiv + 417 pp. LCCN ML3800 .E3 2005.

Einstein:1949:HFR

[EES49] Hans Albert Einstein and El-Sayed Ahmed El-Samni. Hydrodynamic forces on a rough wall. Reviews of Modern Physics, 21 (3):520–524, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.520;   http://rmp.
aps.org/abstract/RMP/v21/i3/p520_1. Special issue in honor

of Albert Einstein on his 70th birthday.

Einstein:1913:NGS

[EF13] Albert Einstein and Adriaan Daniel Fokker. Die Nordströmsche Gravitationstheorie vom Standpunkt des absoluten Differentialkalküls. (German) [Nordström gravitation theory from the viewpoint of absolute differential calculus]. Annalen der Physik (1900) (series 4), 349(10):321–328, 1913. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1933:WKB

[EF33] Albert Einstein and Sigmund Freud. Warum Krieg? Ein Briefwechsel. (German) [Why war? A correspondence]. Technical report, Institut International de Coopération Intellectuele, 1933.

Evlanoff:1969:ANL

[EF69] Michael Evlanoff and Marjorie Fluor. Alfred Nobel, the loneliest millionaire. W. Ritchie Press, Los Angeles, CA, USA, 1969. xv + 336 pp. LCCN TP268.5.N7 E88. Foreword by Simon Ramo. Commentaries by Arnold O. Beckman and Henry T. Mudd.

Einstein:2005:NGS

[EF05] A. Einstein and A. D. Fokker. Die Nordströmsche Gravitationstheorie vom Standpunkt des absoluten Differentialkalküls [AdP 44, 321 (1914)]. (German) [Nordström’s theory of gravity from the point of absolute differential calculus [AdP 44, 321 (1914)]]. Annalen der Physik (Berlin), 14(1):500–508, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/andp.200590042.

Einstein:1972:WKG

[EFA72] Albert Einstein, Sigmund Freud, and Isaac Asimov. Warum Krieg?. (German) [Why War?], volume 28 of Diogenes-Taschenbuch. Diogenes-Verlag, Zürich, Switzerland, 1972. ISBN 3-257-20028-5. 62 pp. LCCN ????

Evans:1989:FNR

[EFH89] Charles R. Evans, Lee S. Finn, and David W. Hobill, editors. Frontiers in numerical relativity. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36666-6. x + 435 pp. LCCN QB461.

Einstein:1913:EVRa

[EG13a] Albert Einstein and Marcel Grossmann. Entwurf einer verallgemeinerten Relativitätstheorie und eine Theorie der Gravitation. (German) [Outline of a Generalized Relativity Theory and a Theory of Gravitation]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1913. 38 pp. LCCN ???? 1. Physikalische Teil von Albert Einstein in Zürich. 2. Mathematische Teil von Marcel Grossmann.

Einstein:1913:EVRb

[EG13b] Albert Einstein and Marcel Grossmann. Entwurf einer verallgemeinerten Relativitätstheorie und eine Theorie der Gravitation. (German) [Outline of a Generalized Relativity Theory and a Theory of Gravitation]. Zeitschrift für Mathematik und Physik, 62(??):225–261, 1913. Reprint of [EG13a].

Einstein:1914:KFV

[EG14] Albert Einstein and Marcel Grossmann. Kovarienzeigenschaften der Feldgleichungen der auf die verallgemeinerte Relativitätstheorie gegründeten Gravitationstheorie. (German) [Covariance properties in the field equations of gravitation theory based on the Generalized Theory of Relativity]. Zeitschrift für Mathematik und Physik, 63(??):215–225, 1914.

Einstein:1923:BNU

[EG23] Albert Einstein and Jakob Grommer. Beweis der Nichtexistenz eines überall regulären zentrisch symmetrischen Feldes nach der Feldtheorie von Th. Kaluza. (German) [Proof of the nonexistence of an everywhere-regular centrically0symmetrical field according to the Field Theory of Th. Kaluza]. Jerusalem Univ. Scripta [Scripta Universitatis atque Bibliothecae Hierosolymitanarum: Mathematica et Physica], 1(7):??, 1923.

Einstein:1927:ARBa

[EG27a] Albert Einstein and Jakob Grommer. Allgemeine Relativitätstheorie und Bewegungsgesetze. (German) [The General Theory of Relativity and the Law of Motion]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):2–13, 1927.

Einstein:1927:ARBb

[EG27b] Albert Einstein and Jakob Grommer. Allgemeine Relativitätstheorie und Bewegungsgesetze. (German) [The General Theory of Relativity and the Law of Motion]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):235–245, 1927.

Earman:1978:EHT

[EG78a] John Earman and Clark Glymour. Einstein and Hilbert: Two months in the history of General Relativity. Archive for History of Exact Sciences, 19(3):291–308, September 1978. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=19&issue=3&spage=291.

Earman:1978:LTE

[EG78b] John Earman and Clark Glymour. Lost in the tensors: Einstein’s struggles with covariance principles 1912–1916. Studies in History and Philosophy of Science Part A, 9(4):251–278, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368178900080.

Earman:1980:GRS

[EG80a] John Earman and Clark Glymour. The gravitational red shift as a test of General Relativity: History and analysis. Studies in History and Philosophy of Science Part A, 11(3):175–214, September 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368180900254.

Earman:1980:REB

[EG80b] John Earman and Clark Glymour. Relativity and eclipses: The British Eclipse Expeditions of 1919 and their predecessors. Historical Studies in the Physical Sciences, 11(1):49–85, ???? 1980. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.

org/stable/27757471.

Esry:1998:BEC

[EG98] Brett D. Esry and Chris H. Greene. Bose–Einstein condensates: Superfluids mixing it up. Nature, 392(6675):434–435, April 2, 1998. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v392/n6675/full/392434a0.html.

Einstein:2005:MRb

[EG05] Albert Einstein and Brian Greene. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, expanded Princeton science library edition, 2005. ISBN 0-691-12027-7. xxiv + 166 pp. LCCN QC6.

URL http://press.princeton.edu/titles/484.html.

Egdall:2014:ERS

[Egd14] Ira Mark Egdall. Einstein relatively simple: our universe revealed in everyday language. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4525-58-8 (hardcover), 981-4525-59-6 (paperback). 300 (est.) pp. LCCN QC16.E5 E24 2013. URL http://www.

worldscientific.com/worldscibooks/10.1142/8915.

Barbara:1995:EPA

[Egg95] Barbara Eggers. Der Einsteinsturm in Potsdam: Architektur und Astrophysik. (German) [The Einstein Tower in Potsdam: architecture and astrophysics]. ARS Nicolai, Berlin, Germany, 1995. ISBN ???? ???? pp. LCCN ????

Eggert:2002:ALB

[Egg02] Marianne Eggert. Ausgewählte Literaturnachweise aus dem Bestand der Akademiebibliothek: Albert Einstein, Physiker. (German) [Selected bibliographies from the holdings of the Academy Library: Albert Einstein, physicist]. Report, Berlin-Brandenburgische Akademie der Wissenschaften Akademiebibliothek, Berlin, Germany, 2002. 43 pp. URL http://

bibliothek.bbaw.de/kataloge/literaturnachweise/einstein/
literatur.pdf.

Egleton:1997:AM

[Egl97] Clive Egleton. The Alsos Mission. Severn House, Sutton, Surrey, UK, 1997. ISBN 0-7278-5201-9. 244 pp.

Enz:1997:WPS

[EGO97] Charles P. (Charles Paul) Enz, Beat Glaus, and Gerhard Oberkofler, editors. Wolfgang Pauli und sein Wirken an der ETH Zürich: aus den Dienstakten der Eidgenössischen Technischen Hochschule. (German) [Wolfgang Pauli and his works at the Technical University of Zürich: from the Service of the Technical University]. VDF Hochschulverlag AG an der ETH Zürich, Zürich, Switzerland, 1997. ISBN 3-7281-2317-X. xi + 463 pp. LCCN QC16.P37 W66 1997.

Egghe:2011:TUN

[EGR11] Leo Egghe, Raf Guns, and Ronald Rousseau. Thoughts on uncitedness: Nobel Laureates and Fields Medalists as case studies. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1637–1644, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Earman:1977:FST

[EGS77] John Earman, Clark N. Glymour, and John J. Stachel, editors. Foundations of space–time theories, volume 8 of Minnesota studies in the philosophy of science. University of Minnesota Press, Minneapolis, MN, USA, 1977. ISBN 0-8166-0807-5. LCCN Q175 .M64 vol. 8; QC173.59.S65.

Esteve:2008:SEB

[EGW^{+}08] J. Estève, C. Gross, A. Weller, S. Giovanazzi, and M. K. Oberthaler. Squeezing and entanglement in a Bose–Einstein condensate. Nature, 455(7217):1216–1219, October 1, 2008. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v455/n7217/
full/nature07332.html.

Einstein:1910:SUB

[EH10a] A. Einstein and L. Hopf. Statistische Untersuchung der Bewegung eines Resonators in einem Strahlungsfeld. (German) [Statistical investigation of the movement of resonators in a radiation field]. Annalen der Physik (1900) (series 4), 338(16): 1105–1115, 1910. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1910:SWS

[EH10b] A. Einstein and L. Hopf. Über einen Satz der Wahrscheinlichkeitsrechnung und seine Anwendung in der Strahlungstheorie. (German) [On a theorem of probability calculation and its application in radiation theory]. Annalen der Physik (1900) (series 4), 338(16):1096–1104, 1910. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1941:WSI

[EH41] Albert Einstein and Allen Harris. The world as I see it. J. Lane, The Bodley Head, London, UK, 1941. 214 pp.

Eckern:2005:CAE

[EH05a] U. Eckern and F. W. Hehl. Commemorating Albert Einstein. Annalen der Physik (1900), 14(1–3):3–4, January/March 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:2005:SUB

[EH05b] A. Einstein and L. Hopf. Statistische Untersuchung der Bewegung eines Resonators in einem Strahlungsfeld [AdP 33, 1105 (1910)]. (German) [Statistical investigation of the motion of a resonator in a radiation field [AdP 33, 1105 (1910)]]. Annalen der Physik (Berlin), 14(1):357–367, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590025.

Einstein:2005:SWS

[EH05c] A. Einstein and L. Hopf. Über einen Satz der Wahrscheinlichkeitsrechnung und seine Anwendung in der Strahlungstheorie [AdP 33, 1096 (1910)]. (German) [On a theorem of probability theory and its application in radiation theory [AdP 33, 1096 (1910)]]. Annalen der Physik (Berlin), 14(1):347–356, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/andp.200590024.

Ehlers:1979:LWA

[Ehl79] J. Ehlers. Life and work of Albert Einstein. Umschau in Wissenschaft und Technik, 79(1):7–10, ???? 1979. ISSN 0041-6347.

Ehlers:2005:MLK

[Ehl05a] Anita Ehlers. ,,Die meiste Lebensfreude kommt aus meiner Geige” — Albert Einstein und die Musik. (German) [“Most of the joy of life comes from my violin” — Albert Einstein and music]. In Steiner [Ste05a], pages 171–189. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http://link.springer.

com/chapter/10.1007/3-540-30595-5_8/.

Ehlers:2005:ZER

[Ehl05b] Jürgen Ehlers. Der Zeitbegriff in Einsteins Relativitätstheorien. (German) [The concept of time in Einstein’s Theory of Relativity]. In Steiner [Ste05a], pages 79–91. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http://link.springer.

com/chapter/10.1007/3-540-30595-5_3/.

Ehlers:2009:GOE

[Ehl09] Jürgen Ehlers. Golden oldie editorial: Editorial note to: H. Weyl, on the General Relativity Theory. General Relativity and Gravitation, 41(7):1655–1660, ???? 2009. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL http:

//www.springerlink.com/content/h3865j5750077338/. See

[Wey23, Wey09].

Einstein:1979:FGA

[EHM^{+}79] Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue, et al. Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue. (German) [Celebration of the 100th Birthdays of Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue]. Paperback der Zeitschrift Naturwissenschaftliche Rundschau. Wissenschaftliche Verlagsgesellschaft, Stuttgart, West Germany, 1979. ISBN 3-8047-0594-4. 99 pp. LCCN QC15 .F44.

Ehrenwald:1977:EE

[Ehr77] Jan Ehrenwald. Einstein on ESP. Science, 198(4316):448, November 4, 1977. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/198/4316/448.4.full.pdf.

Einstein:1938:DEP

[EI38a] Albert Einstein and Leopold Infeld. Drie eeuwen physica: van Galilei tot relativiteitstheorie en quantumtheorie. (Dutch) [Three centuries of physics: from Galileo to relativity theory and quantum theory]. D. B. Centen, Amsterdam, The Netherlands, 1938. 319 pp. LCCN ???? Translation to Dutch of [EI38b].

Einstein:1938:EPGa

[EI38b] Albert Einstein and Leopold Infeld. The Evolution of Physics: the Growth of Ideas from Early Concepts to Relativity and Quanta. Simon and Schuster, New York, NY, USA, 1938. x + 319 + 1 pp. LCCN QC7 .E5.

Einstein:1938:EPGb

[EI38c] Albert Einstein and Leopold Infeld. The Evolution of Physics. The Growth of Ideas from the Early Concepts to Relativity and Quanta. Cambridge library of modern science. Cambridge University Press, Cambridge, UK, 1938. 319 pp. LCCN QC 7 EIN.

Einstein:1938:FUI

[EI38d] Albert Einstein and Leopold Infeld. Fysikens utveckling : idéutveckling från äldre åsikter till relativitets- och kvantteorierna. (Swedish) [The development of physics: idea development from older opinions to the relative and quantum theories]. Natur och Kultur, Stockholm, Sweden, 1938. 253 pp. LCCN ????

Einstein:1938:LIP

[EI38e] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, 1938. vi + 298 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1938:PAE

[EI38f] Albert Einstein and Leopold Infeld. Physik als Abenteuer der Erkenntnis. (German) [Physics as adventures of the realization]. Sijthoff’s Uitgeversmaatschappij, Leiden, The Netherlands, 1938. viii + 222 pp. LCCN ???? URL http://bouquin.

de/ch/88335.php.

Einstein:1939:MVF

[EI39a] Albert Einstein and Leopold Infeld. Det moderne Verdensbillede: Fysikkens Udvikling fra Galilei og Newton til Relativitetsteori og Kvantemekanik. (Danish) [The Modern World View: The Development of Physics from Galileo and Newton to Relativity Theory and Quantum Mechanics]. Schultz, Copenhagen, Denmark, 1939. 287 pp. LCCN ???? Translation to Danish by Niels Arley of [EI38b].

Einstein:1939:EFD

[EI39b] Albert Einstein and Leopold Infeld. A evolução da fisica: o desenvolvimento das ideias desde os primitivos conceitos ate a relatividade e os quanta. (Portuguese) [The evolution of physics: the development of ideas from primitive concepts to relativity and quanta], volume 1 of Biblioteca do espirito moderno. Serie 2a, Ciencia. Companhia Editora Nacional, São Paulo, Brazil, 1939. 344 + 3 pp. LCCN ???? Translation to Japanese of [EI38b].

Einstein:1939:FSV

[EI39c] Albert Einstein and Leopold Infeld. Fysikken som virkelighetens eventyr. (Norwegian) [Physics as reality’s adventure]. Aschehoug, Oslo, Norway, 1939. 310 pp. LCCN ???? Translation to Norwegian by Gunnar Randers of [EI38b].

Einstein:1939:BWI

[EI40a] Albert Einstein and Leopold Infeld. Butsurigaku wa ikani hajimerareta ka. (Japanese) [How was it started?]. Iwanami Shoten, Tokyo, Japan, 1939–1940. ???? (two volumes) pp. LCCN ???? Translation to Japanese of [EI38b].

Einstein:1940:GEP

[EI40b] Albert Einstein and Leopold Infeld. The gravitational equations and the problem of motion. II. Annals of Mathematics (2), 41(2):455–464, April 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL 2-6; http:

//links.jstor.org/sici?sici=0003-486X(194004)2:41:2<

455:TGEATP>2.0.CO. See also part I [EIH38a].

Einstein:1942:EPG

[EI42] Albert Einstein and Leopold Infeld. The evolution of physics: the growth of ideas from early concepts to Relativity and quanta. Simon and Schuster, New York, NY, USA, 1942. x + 319 pp. LCCN QC7 .E5 1942.

Einstein:1947:EPE

[EI47] Albert Einstein and Leopold Infeld. The evolution of physics from early concepts to Relativity and quanta. A Touchstone book. Cambridge University Press, Cambridge, UK, 1947. x + 319 pp. LCCN QC7 .E5 1947.

Einstein:1948:LIP

[EI48a] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, eleventh edition, 1948. 298 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1948:LDF

[EI48b] Albert Einstein and Leopold Infeld. L’evoluzione della fisica (Italian). [The evolution of physics]. G. Einaudi, Torino, Italy, 1948. 310 pp. LCCN QC7 .E515.

Einstein:1949:EMG

[EI49a] Albert Einstein and Leopold Infeld. [Equations of motion in general relativity]. Cited in [Inf78b, page 149] as the authors’ last joint paper, but never finally completed or published., 1949.

Einstein:1949:MPG

[EI49b] Albert Einstein and Leopold Infeld. On the motion of particles in General Relativity Theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):209–241 + 1, ????

1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1949:PAE

[EI49c] Albert Einstein and Leopold Infeld. Physik als Abenteuer de Erkenntnis. (German) [The Evolution of Physics]. A. W. Sijthoff, Leiden, The Netherlands, second edition, 1949. 200 pp. LCCN QC7 .E514.

Einstein:1950:EPG

[EI50a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Zsolnay-Verlag, Wien, Austria, 1950. 352 pp. LCCN ????

Einstein:1950:EP

[EI50b] Albert Einstein and Leopold Infeld. The Evolution of Physics. Scientific Book Club, London, UK, 1950. x + 320 pp. URL

https://archive.org/details/evolutionofphysi033254mbp.

Einstein:1956:EPG

[EI56] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Hamburg, West Germany, 1956. ???? pp. LCCN ????

Einstein:1957:EFD

[EI57] Albert Einstein and Leopold Infeld. Evoluţia fizicii: desvoltarea ideilor de la primele concepte la teoria relativităţii şi teoria cuantelor. (Romanian) [Evolution of Physics: The development of ideas from the first concepts to the theory of relativity and quantum theory]. Editura Tehnică, Bucureşti, Romania, 1957. 240 pp. LCCN ???? Translation to Romanian of [EI38b].

Einstein:1958:FJD

[EI58] Albert Einstein and Leopold Infeld. Fyzika jako dobrodruzstvi poznani. (Czech) [Physics as a knowledge adventure]. Orbis, Praha, Czechoslovakia, 1958. 282 pp. LCCN ???? Translation to Czech by Jan Rey of [EI49c].

Einstein:1959:EFR

[EI59] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogla̧dów od najdawniejszych pojȩć do teorii wzglȩdności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, 1959. 327 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [EI38b].

Einstein:1961:MVF

[EI61a] Albert Einstein and Leopold Infeld. Det moderne verdensbillede, fysikkens udvikling fra Galilei og Newton til relativitetsteori og kvantemekanik. (Danish) [The modern worldview: the development of physics from Galileo and Newton to relativity theory and quantum mechanics], volume 7 of Gyldendals uglebøger. Gyldendal, Copenhagen, Denmark, 1961. 226 pp. LCCN ???? Translation to Danish by Niels Arley of [EI38b].

Einstein:1961:EPN

[EI61b] Albert Einstein and Leopold Infeld. Die Evolution der Physik : von Newton bis zur Quantentheorie. (German) [The Evolution of Physics: from Newton to quantum theory]. Rowohlt, Hamburg, West Germany, 1961. 210 pp. LCCN ????

Einstein:1961:EP

[EI61c] Albert Einstein and Leopold Infeld. The evolution of physics. Cambridge University Press, Cambridge, UK, second edition, 1961. ISBN 0-521-08371-0, 0-521-09687-1. 302 pp. LCCN QC 7 EIN.

Einstein:1961:EPG

[EI61d] Albert Einstein and Leopold Infeld. The evolution of physics: the growth of ideas from early concepts to Relativity and quanta. An Essandess paperback. Simon and Schuster, New York, NY, USA, 1961. xvi + 302 pp. LCCN QC7 .E5 1961.

Einstein:1962:EFR

[EI62a] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogla̧dów od najdawniejszych pojȩć do teorii wzglȩdności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, 1962. 263 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [EI38b].

Einstein:1962:FKN

[EI62b] Albert Einstein and Leopold Infeld. Fysiikan kehitys : Newtonista kvanttiteoriaan. (Finnish) [Physics: From Newton to Quantum Theory], volume 11 of Jokamiehen korkeakoulu. Kustannusosakeyhtio Otava, Helsinki, Finland, 1962. 256 pp. LCCN ???? Translation to Finnish by Heikki Hirvinen of [EI50a].

Einstein:1962:RFO

[EI62c] Albert Einstein and Leopold Infeld. Razvoj fizike od Newtona do kvantne teorije. (Slovenian) [Development of Newton’s physics to quantum theory]. Mladinska Knjiga, Ljubljana, 1962. 223 pp. LCCN ????

Einstein:1963:LIP

[EI63] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta], volume 47. Payot, Paris, France, 1963. ???? pp. LCCN ???? Reprint of [EI38e].

Einstein:1964:RMV

[EI64] Albert Einstein and Leopold Infeld. The rise of the mechanical view. In Rapport and Wright [RW64], pages 13–47.

Einstein:1965:FAD

[EI65a] Albert Einstein and Leopold Infeld. La fisica, aventura del pensamiento: el desarrollo de las ideas desde los primeros conceptos hasta la relatividad y los cuantos. (Spanish) [Physics, adventure of thought: the development of ideas from the first concepts to relativity and quanta]. Editorial Losada, Buenos Aires, Argentina, 1965. 254 pp. LCCN ???? Translation by Rafael Grinfeld of [EI38b].

Einstein:1965:LDF

[EI65b] Albert Einstein and Leopold Infeld. L’evoluzione della fisica: sviluppo delle idee dai concetti iniziali alla relativita e ai quanti. (Italian) [The evolution of physics: development of ideas from initial concepts to relativity and quanta]. Boringhieri, Torino, Italia, 1965. 310 pp. LCCN ????

Einstein:1966:EFR

[EI66a] Albert Einstein and Leopold Infeld. Evoliutsiia fiziki: razvitie idei ot pervonachalnykh poniatii do teorii otnositel’nosti i kvantov. (Russian) [Evolution of physics: The development of ideas from the original concepts to the theory of relativity and quanta]. Molodaya gvardiya, Moscow, USSR, fourth edition, 1966. 266 + 4 pp. LCCN ???? Translation to Russian by S. G. Suvorova of [EI38b].

Einstein:1966:EPE

[EI66b] Albert Einstein and Leopold Infeld. The Evolution of Physics from Early Concepts to Relativity and Quanta. A Touchstone book. Simon and Schuster, New York, NY, USA, 1966. ISBN 0-671-20156-5. xvi + 302 pp. LCCN QC7 .E5 1961.

Einstein:1968:LFS

[EI68] Albert Einstein and Leopold Infeld. L’evolución de la física. (Spanish) [The Evolution of Physics], volume 18 of Biblioteca bàsica de cultura contemporània. Edicions 62, Barcelona, Spain, 1968. 244 pp. LCCN ????

Einstein:1971:EPG

[EI71] Albert Einstein and Leopold Infeld. The evolution of physics: the growth of ideas from early concepts to relativity and quanta. A Touchstone book. Cambridge University Press, Cambridge, UK, second edition, 1971. ISBN 0-521-09687-1, 0-521-08371-0. xvi + 302 pp. LCCN QC7 .E5 1971.

Einstein:1978:ETI

[EI78a] Albert Einstein and Leopold Infeld. Ē exelixē tōn ideōn stē physikē. (Greek) [The Evolution of Physics], volume 1 of Epistēmologikē bibliothēkē. Ekdoseis Dōdōnē, Athēna, Greece, 1978. 298 pp. LCCN ???? Translation to Greek by Eut. Mpitsakē.

Einstein:1978:LIP

[EI78b] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta], volume 47 of Petite bibliothèque Payot. Payot, Paris, France, 1978. ISBN 2-228-30473-5, 2-228-30474-3. 280 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1984:LFS

[EI84] Albert Einstein and Leopold Infeld. L’evolució de la física. (Spanish) [The Evolution of Physics], volume 12 of Clàssics del Pensament Modern. Edicions 62, Barcelona, Spain, 1984. ISBN 84-297-2107-X. 236 pp. LCCN ???? Translation to Spanish by Humbert Padellans of [EI38b]. Edited by David Jou. Preface by Antoni Lloret i Orriols.

Einstein:1987:EPG

[EI87] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 8342 of rororo-Sachbuch. Rowohlt, Reinbek bei Hamburg, Germany, 1987. ISBN 3-499-18342-0. 265 pp. LCCN ???? Translation to German by Werner Preusser.

Einstein:1988:EPG

[EI88] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Reinbek bei Hamburg, Germany, ninth edition, 1988. ISBN 3-499-18342-0. 265 pp. LCCN ????

Einstein:1989:EPG

[EI89a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Reinbek bei Hamburg, Germany, thirteenth edition, 1989. ISBN 3-499-18342-0. 265 pp. LCCN ????

Einstein:1989:LIP

[EI89b] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. France Loisirs, Paris, France, 1989. ISBN 2-7242-4426-5. 445 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b]. Preface by Igor Bogdanov and Grichka Bogdanov.Klein.

Einstein:1992:EPG

[EI92] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Weltbild-Verlag, Augsburg, Germany, 1992. ISBN 3-89350-161-4. 332 pp. LCCN ????

Einstein:1993:LIP

[EI93] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, 1993. ISBN 2-08-081119-3. 280 pp. LCCN ???? Reprint of [EI38e].

Einstein:1994:FEI

[EI94] Albert Einstein and Leopold Infeld. Fizigin evrimi: ilk kavramlardan iliskinlige ve kuantumlara. (Turkish) [The evolution of physics: from early concepts to relativity and quanta]. Onur Yayinlari, Ankara, Turkey, 1994. ISBN 975-351-006-3. 262 pp. LCCN ???? Translation to Turkish by Öner Ünalan of [EI47].

Einstein:1995:EPG

[EI95] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 1995. ISBN 3-499-19921-1. 286 pp. LCCN ???? Translation to German by Werner Preusser. With an introduction by Albrecht Fölsing.

Einstein:1998:EPG

[EI98a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 1998. ISBN 3-499-19921-1. 286 pp. LCCN ???? Translation to German by Werner Preusser. With an introduction by Albrecht Fölsing.

Einstein:1998:EFR

[EI98b] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogla̧dów od najdawniejszych pojȩć do teorii wzglȩdności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Klasycy nauki. Prószyński i S-ka, Warszawa, Poland, fourth edition, 1998. ISBN 83-7180-985-9. 257 pp. LCCN QC7 .E5 1998.

Einstein:2000:FEB

[EI00] Albert Einstein and Leopold Infeld. Fisikaren eboluzioa. (Basque) [Evolution of physics]. Gaiak Argitaldaria, Donostia, Spain, 2000. ISBN 84-89772-44-4. 257 pp. LCCN ???? Translation to Basque by Juan Ignazio Abrisketa of [EI47].

Einstein:2004:EPG

[EI04] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 2004. ISBN 3-499-19921-1. 286 pp. LCCN ????

Einstein:2008:EFP

[EI08a] Albert Einstein and Leopold Infeld. A evolução da física. (Portuguese) [The Evolution of Physics]. Jorge Zahar, Rio de Janeiro, Brazil, 2008. ISBN 85-378-0052-X. 248 pp. LCCN ???? Translation to Portuguese by Giasone Rebuá of [EI47].

Einstein:2008:EPE

[EI08b] Albert Einstein and Leopold Infeld. The evolution of physics: from early concepts to relativity and quanta. Simon and Schuster, New York, NY, USA, 2008. ISBN 1-4165-5945-0 (hardcover). xxx + 302 pp. LCCN QC7 .E5 2008. Reprint of [EI38b]. With a new foreword by Walter Isaacson.

Einstein:2011:FEI

[EI11] Albert Einstein and Leopold Infeld. Fiziğin evrimi: ilk kavramlardan ili kinliğe ve kuantumlara. (Turkish) [The evolution of physics: from early concepts to relativity and quanta]. Evrensel Basım Yayın, İstanbul, Turkey, 2011. ISBN 605-4156-85-3. 262 pp. LCCN ???? Translation to Turkish by Öner Ünalan of [EI47].

Einstein:2013:LDF

[EI13] Albert Einstein and Leopold Infeld. L’evoluzione della fisica: sviluppo delle idee dai concetti iniziali alla relatività e ai quanti. (Italian) [The evolution of physics: development of ideas from initial concepts to relativity and quanta], volume 20 of I grandi pensatori / [Bollati Boringhieri]. Boringhieri, Torino, Italia, second edition, 2013. ISBN 88-339-2258-8. 279 + 2 pp. LCCN ???? Translation to Italian by Adele Graziadei of [EI47]. Preface by Carlo Castagnoli.

Einstein:2015:LIP

[EI15] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Champs. Flammarion, Paris, France, 2015. ISBN 2-08-137310-6 (paperback). ISSN 0151-8089. 344 pp. LCCN ???? URL http:

//catalogue.bnf.fr/ark:/12148/cb44450120n. Translation

to French by Maurice Solovine of [EI38b]. Preface by Étienne Klein.

Eichstaedt:1994:IYP

[Eic94] Peter H. Eichstaedt. If you poison us: uranium and Native Americans. Red Crane Books, Santa Fe, NM, USA, 1994. ISBN 1-878610-40-6. xvi + 263 + 16 pp. LCCN E99.N3 E29 1994.

Einstein:1938:GEPa

[EIH38a] Albert Einstein, Leopold Infeld, and Banesh Hoffmann. The gravitational equations and the problem of motion. Annals of Mathematics (2), 39(1):65–100, January 1938. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-
486X(193801)2:39:1<65:TGEATP>2.0.CO%3B2-P.

Einstein:1938:GEPb

[EIH38b] Albert Einstein, Leopold Infeld, and Banesh Hoffmann. The gravitational equations and the problems of motion. Annals of Mathematics (2), 41(2):455–564, January 1938. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-
486X(193801)2:39:1<65:TGEATP>2.0.CO%3B2-P.

Einstein:1895:UAI

[Ein95] Albert Einstein. Über die Untersuchung des Aetherzustandes im magnetishen Felde. (German). [On the investigation of the state of the ether in magnetic fields]. Dated between 05/01/1895 and

09/25/1895., 1895. URL http://www.alberteinstein.info/
db/ViewDetails.do?DocumentID=34390.

Einstein:1901:FCG

[Ein01] A. Einstein. Folgerungen aus den Capillaritätserscheinungen. (German) [Consequences of the capillarity phenomenon]. Annalen der Physik (1900) (series 4), 309(3):513–523, 1901. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). This is Einstein’s first published paper.

Einstein:1902:KTW

[Ein02a] A. Einstein. Kinetische Theorie des Wärmegleichgewichtes und des zweiten Hauptsatzes der Thermodynamik. (German) [Kinetic theory of heat equilibrium and the Second Law of Thermodynamics]. Annalen der Physik (1900) (series 4), 314(10): 417–433, 1902. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1902:TTP

[Ein02b] A. Einstein. Über die thermodynamische Theorie der Potentialdifferenz zwischen Metallen und vollständig dissociierten Lösungen ihrer Salze und über eine elektrische Methode zur Erforschung der Molekularkräfte. (German) [On the thermodynamic theory of the difference of potential between metals and completely dissociated solutions of their salts and more on an electrical method for the study of intermolecular forces]. Annalen der Physik (1900) (series 4), 313(8):798–814, 1902. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1903:TGT

[Ein03] A. Einstein. Eine Theorie der Grundlagen der Thermodynamik. (German) [A theory of the foundation of thermodynamics]. Annalen der Physik (1900) (series 4), 316(5):170–187, 1903. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1904:AMT

[Ein04] A. Einstein. Zur allgemeinen molekularen Theorie der Wärme. (German) [Towards a general molecular theory of heat]. Annalen der Physik (1900) (series 4), 319(7):354–362, 1904. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1905:NBM

[Ein05a] Albert Einstein. Eine Neue Bestimmung der Moleküldimensionen. (German) [A new determination of molecular dimensions]. Inaugural dissertation, Bern Wyss., Bern, Switzerland, 1905. URL

http://en.wikipedia.org/wiki/Alfred_Kleiner. Published

in [Ein06c].

Einstein:1905:TKS

[Ein05b] Albert Einstein. Ist die Trägheit eines Körpers von seinem Energieinhalt abhängig?. (German) [Does the inertia of a body depend on its energy content?]. Annalen der Physik (1900) (series 4), 323(13):639–641, 1905. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

www.fourmilab.ch/etexts/einstein/E_mc2/www/;   http:
//www.gsjournal.net/Science-Journals/Essays/View/2493;
 http://www.itba.edu.ar/cargrado/fismat/fismod/transf/
htm/einstein_2.htm;   http://www.zbp.univie.ac.at/einstein/

einstein4.pdf. English translation in [LEMW52, 67–71].

Einstein:1905:MTW

[Ein05c] Albert Einstein. Über die von der molekularkinetischen Theorie der Wärme gefordert Bewegung von in ruhenden Flüßigkeiten suspendierten Teilchen. (German) [On the molecular-kinetic theory of the movement by heat of particles suspended in liquids at rest]. Annalen der Physik (1900) (series 4), 322(8):549–560, 1905. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://www.

nobelprize.org/nobel_prizes/physics/laureates/1926/;
 http://www.zbp.univie.ac.at/einstein/einstein2.pdf.

See the historical overview in [Haw02a], the earlier discovery

[Bac00] applied to modeling of stock market fluctuations, and a simpler derivation, building on Einstein’s work [Lan08, LG97]. See also [Nye72] for a book-length treatment of theory and experiment of Brownian motion, particularly the careful, and numerous, experiments of Jean Perrin that demonstrated close agreement of theory with experiment, and provided credible proof of the existence of atoms and molecules. Perrin received the Nobel Prize in Physics in 1926 “for his work on the discontinuous structure of matter, and especially for his discovery of sedimentation equilibrium”.

Einstein:1905:EVL

[Ein05d] Albert Einstein. Über einen die Erzeugung und Verwandlung des Lichtes betreffenden heuristischen Gesichtspunkt. (German) [On the production and transformation of light from a heuristic viewpoint]. Annalen der Physik (1900) (series 4), 322(6):132–148, 1905. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://www.gsjournal.

net/Science-Journals/Essays/View/2490;   http://www.
gsjournal.net/Science-Journals/Essays/View/2491;   http:
//www.zbp.univie.ac.at/einstein/einstein1.pdf. English

translation in [AP65].

Einstein:1905:EBK

[Ein05e] Albert Einstein. Zur Elektrodynamik bewegter Körper. (German) [On the electrodynamics of moving bodies]. Annalen der Physik (1900) (series 4), 322(10):891–921, 1905. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic).  URL http://www.gsjournal.net/Science-
Journals/Essays/View/2492;   http://www.itba.edu.ar/
cargrado/fismat/fismod/transf/htm/einstein_1.htm;
http://www.zbp.univie.ac.at/einstein/einstein3.pdf. En-

glish translations in [LEMW52, 35–65] and [Sch77d].

Einstein:1906:PES

[Ein06a] A. Einstein. Das Prinzip von der Erhaltung der Schwerpunktsbewegung und die Trägheit der Energie. (German) [The principle of the preservation of the center of gravity and inertial energy]. Annalen der Physik (1900) (series 4), 325(8):627–633, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1906:TBB

[Ein06b] A. Einstein. Zur Theorie der Brownschen Bewegung. (German) [On the theory of Brownian motion]. Annalen der Physik (1900) (series 4), 324(2):371–381, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1906:NBM

[Ein06c] Albert Einstein. Eine neue Bestimmung der Moleküldimensionen. (German) [A new determination of molecular dimensions]. Annalen der Physik (1900) (series 4), 324(2):289–306, 1906. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See corrections [Ein11a, Ein11e]. This is a slightly revised version of Einstein’s doctoral dissertation [Ein05a].

Einstein:1906:ORM

[Ein06d] Albert Einstein. Optik: [Rezension von] M[ax] Planck: Vorlesungen über die Theorie der Wärmestrahlung. (German) [Review of Max Planck: Lectures on the theory of thermal radiation]. Beiblätter zu den Annalen der Physik, 30 (15):764–766, ???? 1906. CODEN BEAPAI. ISSN 0375-

846X. URL http://hdl.handle.net/2027/uva.x000269580?
urlappend=%3Bseq=772.

Einstein:1906:MBV

[Ein06e] Albert Einstein. Über eine Methode zur Bestimmung des Verhältnisses der transversalen und longitudinalen Masse des Elektrons. (German) [On a method for the determination of the ratio of the transverse and the longitudinal mass of electrons]. Annalen der Physik (1900) (series 4), 326(13):583–586, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1906:TLL

[Ein06f] Albert Einstein. Zur Theorie der Lichterzeugung und Lichtabsorption. (German) [On the theory of light production and light absorption]. Annalen der Physik (1900) (series 4), 325(6):199– 206, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1907:BNH

[Ein07a] A. Einstein. Bemerkungen zu der Notiz des Herrn Paul Ehrenfest: “Die Translation deformierbarer Elektronen und der Flächensatz”. (German) [Remark on note of Mr. Paul Ehrenfest: “The translation of deformable electrons and the Area Law”]. Annalen der Physik (1900) (series 4), 328(6):206–208, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1907:BMA

[Ein07b] A. Einstein. Berichtigung zu meiner Arbeit: Die Plancksche Theorie der Strahlung etc.. (German) [Correction to my paper: “The Planck Theory of Radiation etc.”]. Annalen der Physik (1900) (series 4), 327(4):800, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Ein07e].

Einstein:1907:MNP

[Ein07c] A. Einstein. Über die Möglichkeit einer neuen Prüfung des Relativitätsprinzips. (German) [On the possibility of a new proof of the Principle of Relativity]. Annalen der Physik (1900) (series 4), 328(6):197–198, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1907:RGT

[Ein07d] A. Einstein. Über die vom Relativitätsprinzip geforderte Trägheit der Energie. (German) [On the inertial energy required by the Relativity Principle]. Annalen der Physik (1900) (series 4), 328(7):371–384, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1907:PTS

[Ein07e] Albert Einstein. Die Plancksche Theorie der Strahlung und die Theorie der spezifischen Wärme. (German) [Planck’s theory of radiation and the theory of specific heat]. Annalen der Physik (1900) (series 4), 327(1):180–190, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See correction [Ein07b].

Einstein:1907:TBB

[Ein07f] Albert Einstein. Theoretische Bemerkungen über die Brownsche Bewegung. (German) [Theoretical remarks on Brownian motion]. Zeitschrift für Elektrochemie und angewandte physikalische Chemie, 13(6):41–42, 1907. ISSN 0005-9021.

Einstein:1907:RDGb

[Ein07g] Albert Einstein. Über das Relativitätsprinzip und die aus demselben gezogene Folgerungen. (German) [On the Relativity Principle and the conclusions drawn from it]. Jahrbuch der Radioaktivität und Elektronik, 5:98–99, 1907.

Einstein:1907:RDGa

[Ein07h] Albert Einstein. Über das Relativitätsprinzip und die aus demselben gezogenen Folgerungen. (German) [On the Relativity Principle and the conclusions drawn from it]. Jahrbuch der Radioaktivität und Elektronik, 4:411–462, 1907. See corrections in [Ein07g]. English translations in [Sch77a, Sch77b, Sch77c, ?, Volume 2]. This is Einstein’s first statement of the Principle of Equivalence.

Einstein:1907:GST

[Ein07i] Albert Einstein. Über die Gültigkeitsgrenze des Satzes vom thermodynamischen Gleichgewicht und über die Möglichkeit einer neuen Bestimmung der Elementarquanta. (German) [Validity limits of the equation of thermodynamic equilibrium and more on the possibility of a new determination of elementary quanta]. Annalen der Physik (1900) (series 4), 327(3):569–572, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1908:NEM

[Ein08a] Albert Einstein. Eine neue elektrostatische Methode zur Messung kleiner Elektrizitätsmengen. (German) [New electrostatic method for the measurement of small electrical charges]. Physikalische Zeitschrift, 9(??):216–217, 1908.

Einstein:1908:ETB

[Ein08b] Albert Einstein. Elementare Theorie der Brownschen Bewegung. (German) [Elementary theory of Brownian motion]. Zeitschrift für Elektrochemie und angewandte physikalische Chemie, 14(17):235–239, 1908. ISSN 0005-9021.

Einstein:1909:BAD

[Ein09a] A. Einstein. Bemerkung zu der Arbeit von D. Mirimanoff: ,,Über die Grundgleichungen. … “. (German) [Remark on the work of D. Mirimanoff: On the principal equations …]. Annalen der Physik (1900) (series 4), 333(4):885–888, 1909. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1909:EUAa

[Ein09b] A. Einstein. Über die Entwicklung unserer Anschauungen über das Wesen und die Konstitution der Strahlung. (German) [On the development of our views about the nature and constitution of radiation]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 11(20):482–500, October 30, 1909. CODEN VD-PEAZ. ISSN 0372-5448. Republished in [Ein09c].

Einstein:1909:EUAb

[Ein09c] A. Einstein. Über die Entwicklung unserer Anschauungen über das Wesen und die Konstitution der Strahlung. (German) [On the development of our views about the nature and constitution of radiation]. Physikalische Zeitschrift, 10(??):817–826, 1909. Republication of [Ein09b].

Einstein:1909:GSS

[Ein09d] Albert Einstein. Zum gegenwärtigen Stande des Strahlungsproblems. (German) [On the current state of the radiation problem]. Physikalische Zeitschrift, 10(??):185–193, 1909.

Einstein:190x:TGG

[Ein0x] Albert Einstein. Tangentenbussole und Galvanometer. (German) [Tangent galvanometer and galvanometer]. Matura Examination (D) Physics, 190x.

Einstein:1910:TOH

[Ein10a] A. Einstein. Theorie der Opaleszenz von homogenen Flüssigkeiten und Flüssigkeitsgemischen in der Nähe des kritischen Zustandes. (German) [Theory of opalescence of homogeneous liquids and liquid mixtures near the critical state]. Annalen der Physik (1900) (series 4), 338(16):1275–1298, 1910. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1910:ETE

[Ein10b] Albert Einstein. Einführung in die Theorie der Elektrizität und des Magnetismus. Zürich, Wintersemester 1910–1911. (German) [Lecture notes for course on electricity and magnetism at the University of Zurich, Winter Semester 1910–1911]. Unpublished, but reproduced in [?, pages 316–396]., 1910.

Einstein:1910:PRSa

[Ein10c] Albert Einstein. Le principe de relativité et ses conséquences dans la physique moderne. (French) [The Principle of Relativity and its consequences in modern physics]. Arch. Sci. phys. nat. (4), 29(??):5–28, 1910.

Einstein:1910:PRSb

[Ein10d] Albert Einstein. Le principe de relativité et ses conséquences dans la physique moderne. (French) [The Principle of Relativity and its consequences in modern physics]. Arch. Sci. phys. nat. (4), 29(??):125–244, 1910.

Einstein:1910:FPQ

[Ein10e] Albert Einstein. Les forces pondéromotrices qui agissent sur les conducteurs ferromagnétiques disposés dans un champ magnétique et parcourus par un courant. (French) [Ponderomotive (inertial??) forces which act on the ferromagnetic drivers laid out in a magnetic field and traversed by a current]. Arch. Sci. phys. nat. (4), 30(??):323–324, 1910.

Einstein:1910:TQL

[Ein10f] Albert Einstein. Sur la théorie des quantités lumineuses et la question de la localisation de l’énergie électromagnétique. (French) [On the theory of luminous quantities and the question of the localization of electromagnetic energy]. Arch. Sci. phys. nat. (4), 29(??):525–528, 1910.

Einstein:1911:BMAa

[Ein11a] A. Einstein. Bemerkung zu meiner Arbeit: Eine Beziehung zwischen dem elastischen Verhalten. (German) [Remark on my paper: “A relationship between the elastic behavior … ”]. Annalen der Physik (1900) (series 4), 339(3):590, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Ein11g].

Einstein:1911:RGT

[Ein11b] A. Einstein. Die Relativitätstheorie. (German) [The theory of relativity]. Naturforschende Gesellschaft in Zürich, Vierteljahrsschrift, 56(??):1–14, 1911.

Einstein:1911:EPB

[Ein11c] A. Einstein. Zum Ehrenfestschen Paradoxon. Bemerkung zu V. Varicaks Aufsatz. (German) [On the Ehrenfest paradox: Remark on V. Varicaks’s essay]. Physikalische Zeitschrift, 12(??): 509–510, 1911.

Einstein:1911:BGE

[Ein11d] Albert Einstein. Bemerkung zu dem Gesetz von Eötvös. (German) [Remark on the Law of Eötvös]. Annalen der Physik (1900) (series 4), 339(1):165–169, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1911:BMAb

[Ein11e] Albert Einstein. Berichtigung zu meiner Arbeit: Eine neue Bestimmung der Moleküldimensionen. (German) [Corrections to my work: a new determination of molecular dimensions]. Annalen der Physik (1900) (series 4), 339(3):591–592, 1911. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Ein06c].

Einstein:1911:BPH

[Ein11f] Albert Einstein. Bermerkungen zu den P. Hertzschen Arbeiten: Mechanische Grundlagen der Thermodynamik. (German) [Remarks on the work of P. Hertz: “Mechanical Bases of Thermodynamics]. Annalen der Physik (1900) (series 4), 339(1):175–176, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1911:BZE

[Ein11g] Albert Einstein. Eine Beziehung zwischen dem elastischen Verhalten und der spezifischen Wärme bei festen Körpen mit einatomigem Molekül. (German) [A relationship between the elastic behavior and the specific heat of solid bodies with monatomic molecules]. Annalen der Physik (1900) (series 4), 339(1):170–174, 590, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See remarks [Ein11a, Ein11e].

Einstein:1911:EBT

[Ein11h] Albert Einstein. Elementare Betrachtungen über die thermische Molekularbewegung in festen Körpern. (German) [Elementary observations on the thermomolecular motions of inertial bodies]. Annalen der Physik (1900) (series 4), 340(9):679–694, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1911:ESA

[Ein11i] Albert Einstein. Über den Einfluß der Schwerkraft auf die Ausbreitung des Lichtes. (German) [On the influence of gravitation on the propagation of light]. Annalen der Physik (1900) (series 4), 340(10):898–908, 1911. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://www.itba.

edu.ar/cargrado/fismat/fismod/transf/htm/einstein_5.

htm. English translation in [?, Volume 3] and [LEMW52, 97– 108].

Einstein:1912:ABJ

[Ein12a] Albert Einstein. Antwort auf eine Bemerkung von J. Stark: Über eine Anwendung des Planckschen Elementargesetzes. (German) [Answer to a remark of J. Stark: “Application of Planck’s fundamental law”]. Annalen der Physik (1900) (series 4), 343(9):888, May 30, 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1912:BAV

[Ein12b] Albert Einstein. Bemerkung zu Abrahams vorangehender Auseinandersetzung: ,,Nochmals Relativität und Gravitation“. (German) [Remark on Abrahams’ preceding argument: Relativity and gravitation revisited]. Annalen der Physik (1900) (series 4), 344(13):704, August 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Abr12b, Abr12a].

Einstein:1912:EAD

[Ein12c] Albert Einstein. État actuel du problème des chaleurs spécifiques. (French) [Current state of the problem of specific heat]. In Langevin and de Broglie [LdB12a], pages 407–435 +

436–450 (discussion). URL https://archive.org/details/
lathoriedurayo00inst.

Einstein:1912:GGE

[Ein12d] Albert Einstein. Gibt es eine Gravitationswirkung, die der elektrodynamischen Induktionswirkung Analog ist?. (German) [Given that there is a gravitational effect on electrodynamics, is the induction effect similar?]. Vierteljahrsschrift für gerichtliche Medizin und öffentliches Sanitätswesen (3), 44(??):37–40, 1912.

Einstein:1912:LSG

[Ein12e] Albert Einstein. Lichtgeschwindigkeit und Statik des Gravitationsfeldes. (German) [Speed of light and statics of the gravitational field]. Annalen der Physik (1900) (series 4), 343(7): 355–369, 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1912:NMA

[Ein12f] Albert Einstein. Nachtrag zu meiner Arbeit: ‘Thermodynamische Begründung des photochemischen Äquivalentgesetzes’. (German) [Supplement to my paper: ‘Thermodynamic Grounds for the Law of Photochemical Equivalence’]. Annalen der Physik (1900) (series 4), 343(9):881–884, ???? 1912.

URL http://onlinelibrary.wiley.com/doi/10.1002/andp.
19123430910/abstract.

Einstein:1912:RGE

[Ein12g] Albert Einstein. Relativität und Gravitation: Erwiderung auf eine Bemerkung von M. Abraham. (German) [Relativity and gravitation: Retort on a remark of M. Abraham]. Annalen der Physik (1900) (series 4), 343(10):1059–1064, 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Abr12b, Abr12a].

Einstein:1912:TBPb

[Ein12h] Albert Einstein. Thermodynamische Begründung des photochemischen Äquivalentgesetzes. (German) [Thermodynamic reasons for the Law of Photochemical Equivalence]. Annalen der Physik (1900) (series 4), 38(??):881–884, 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1912:TBPa

[Ein12i] Albert Einstein. Thermodynamische Begründung des photochemischen Äquivalentgesetzes. (German) [Thermodynamic reasons for the photochemical Law of Equivalence]. Annalen der Physik (1900) (series 4), 342(4):832–838, 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1912:TSG

[Ein12j] Albert Einstein. Zur Theorie des statischen Gravitationsfeldes. (German) [Theory of statistical gravity fields]. Annalen der Physik (1900) (series 4), 343(7):443–458, 1912. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). Also Nachtrag zur Korrektur (Addendum for the correction) in same volume.

Einstein:1913:DTL

[Ein13a] Albert Einstein. Déduction thermodynamique de la loi de l’équivalence photochimique. (French) [Thermodynamic deduction of the Law of Photochemical Equivalence]. Journal de physique (3), 3(??):277–282, 1913.

Einstein:1913:GGT

[Ein13b] Albert Einstein. Gravitationstheorie. (German) [Theory of gravitation]. Schweizerische Naturforschende Gesellschaft. Verhandlungen, 96(2):??, ???? 1913.

Einstein:1913:HLR

[Ein13c] Albert Einstein. H. A. Lorentz: Das Relativitätsprinzip. (German) [Review of H. A. Lorentz: The Principle of Relativity]. Die Naturwissenschaften, 2(??):1018, 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1913:MPF

[Ein13d] Albert Einstein. Max Planck als Forscher. (German) [Max Planck as researcher]. Die Naturwissenschaften, 1(45):1077– 1079, November 7, 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01493259. Reprinted in [?, 561–563].

Einstein:1913:PGG

[Ein13e] Albert Einstein. Physikalische Grundlagen einer Gravitationstheorie. (German) [Physical basis of a Theory of Gravitation]. Naturforschende Gesellschaft in Zürich, Vierteljahrsschrift, 58 (??):284–290, 1913. Lecture given on 9 September 1913 to the Ninety-Sixth Annual Meeting of the Schweizerische Naturgesellschaft (Swiss Society of Natural Scientists) in Frauenfeld, Switzerland.

Einstein:1913:GZG

[Ein13f] Albert Einstein. Zum gegenwärtigen Stande des Gravitationsproblems. (German) [On the present state of the problem of gravitation]. Physikalische Zeitschrift, 14(25):1249–1266, De-

cember 15, 1913. URL https://hdl.handle.net/2027/mdp.
39015021268936?urlappend=%3Bseq=1391.

Einstein:1913:GSP

[Ein13g] Albert Einstein. Zum gegenwärtigen Stande des Problems der spezifischen Wärme. (German) [On the present state of the problem of specific heat]. Abhandlungen der Deutsche Bunsengesellschaft, ??(7):330–364, 1913. German edition of the Proceedings of the First Solvay Conference; see [Ein12c]. Republished in [Ein14o].

Einstein:1914:AGI

[Ein14a] Albert Einstein. Antrittsrede. (German) [Inaugural lecture: Principles of theoretical physics]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):739– 742, 1914. Included in the book Mein Weltbild. English translation in [Ein82c, pp. 220–223].

Einstein:1914:ARP

[Ein14b] Albert Einstein. Antwort auf eine Replik P. Harzers (Nr. 4753, S. 10 und 11). (German) [Answer to a reply of P. Harzer (no. 4753, pp. 10–11)]. Astronomische Nachrichten, 199(3):47–48, 1914. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Einstein:1914:BPT

[Ein14c] Albert Einstein. Base physiques d’une théorie de la gravitation. (French) [Physical basis of a theory of Gravitation]. Arch. Sci. phys. nat. (4), 37(??):5–12, 1914. French translation by E. Guillaume of [Ein13e].

Einstein:1914:BQG

[Ein14d] Albert Einstein. Beiträge zur Quantentheorie. (German) [Contributions to quantum theory]. Ber. dtsch. phys. Ges., 16(??): 820–828, 1914.

Einstein:1914:BGC

[Ein14e] Albert Einstein. Bemerkung. (German) [Comments on “Outline of a Generalized Theory of Relativity and of a Theory of Gravitation”]. Zeitschrift für Mathematik und Physik, 62(??): ??, ???? 1914. CODEN ???? ISSN ????

Einstein:1914:BPH

[Ein14f] Albert Einstein. Bemerkung zu P. Harzers Abhandlung: Über die Mitführung des Lichtes in Glas und die Aberration (A. N. 4748). (German) [Remark on P. Harzers’ paper: “The transmission of light in glass and its aberration” (A. N. 4748)]. Astronomische Nachrichten, 199(1):7– 10, 1914. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-

3994 (electronic). URL http://www.gsjournal.net/Science-
Journals/Essays/View/4232.

Einstein:1914:FGA

[Ein14g] Albert Einstein. Die formale grundlage der allgemeinen relativitätstheorie. (German) [The formal basis of the Theory of General Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):1030–1085, 1914.

Einstein:1914:MSV

[Ein14h] Albert Einstein. Eine Methode zur statistischen Verwertung von Beobachtungen scheinbar unregelmässig quasiperiodisch verlaufender Vorgänge. (German) [A method for the statistical use of observations of apparently irregular, quasiperiodic processes ]. ????, ??(??):603–??, February 28, 1914.

Einstein:1914:GGT

[Ein14i] Albert Einstein. Gravitationstheorie. (German) [Theory of Gravitation]. Verhandlungen der Schweizerische naturforschende Gesellschaft, part 2, ??(??):137–138, 1914. Same lecture as in [Ein13e].

Einstein:1914:MPD

[Ein14j] Albert Einstein. Méthode pour la détermination de valeurs statistiques d’observations concernant des grandeurs soumises à des fluctuations irrégulières. (French) [method for the determination of statistical values of observations regarding quantities subject to irregular fluctuations]. Archives des sciences physiques et naturelles, 37(??):254–256, ???? 1914. CODEN ???? ISSN ???? English translation in [Ein87b].

Einstein:1914:NAF

[Ein14k] Albert Einstein. Nachträgliche Antwort auf eine Frage von Reißner. (German) [Additional answer to a question of Reissner]. Physikalische Zeitschrift, 15(??):108–110, 1914.

Einstein:1914:PGL

[Ein14l] Albert Einstein. Physikalische Grundlagen und leitende Gedanken für eine Gravitationstheorie. (German) [Physical basis and leading thoughts for a Theory of Gravitation]. Verhandlungen der Schweizerische naturforschende Gesellschaft, part 2, 96(??):146, 1914. Title only. Lecture delivered at the 96th session of the Swiss Natural Science Research Society, 19 September, 1913. Same lecture as in [Ein13e].

Einstein:1914:PVR

[Ein14m] Albert Einstein. Prinzipielles zur verallgemeinerten Relativitätstheorie und Gravitationstheorie. (German) [Principles of Generalized Relativity Theory and Theory of Gravitation]. Physikalische Zeitschrift, 15(??):176–180, 1914.

Einstein:1914:RGP

[Ein14n] Albert Einstein. Relativitätsprinzip. (German) [The Principle of Relativity]. Die Vossische Zeitung, pages 33–34, April 26, 1914.

Einstein:1914:GSP

[Ein14o] Albert Einstein. Zum gegenwärtigen Stande des Problems der spezifischen Wärme. (German) [On the present state of the problem of specific heat]. In Arnold Eucken, editor, Die Theorie der Strahlung und der Quanten. (German) [The Theory of Radiation and the Quantum], pages 330–352. Knapp, Halle a.s., Germany, 1914. Republication of [Ein13g].

Einstein:1914:RGT

[Ein14p] Albert Einstein. Zum Relativitätsproblem. (German) [Toward the Problem of Relativity]. Scientia (Bologna), 15(??):337–348, 1914.

Einstein:1914:TGG

[Ein14q] Albert Einstein. Zur Theorie der Gravitation. (German) [Toward a Theory of Gravitation]. Naturforschende Gesellschaft in Zürich, Vierteljahrsschrift, 59(2):4–6, 1914.

Einstein:1915:TAG

[Ein15a] A. Einstein. Theoretische Atomistik. (German) [Theoretical atomism]. In Emil Warburg, editor, Die Kultur der Gegenwart. Ihre Entwicklung und ihre Ziele. Paul Hinneberg, ed. Part 3, Mathematik, Naturwissenschaften, Medizin. Section 3, Anorganischen Naturwissenschaften, volume 1 of Die Physik, pages 251–263. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1915. LCCN ????

Einstein:1915:AAM

[Ein15b] Albert Einstein. Antwort auf eine Abhandlung M. von Laues: ,,Ein Satz der Wahrscheinlichkeitsrechnung und seine Anwendung auf die Strahlungstheorie“. (German) [Answer to a paper of M. von Laue: “A theory of probability calculation and its application to radiation theory”]. Annalen der Physik (1900) (series 4), 352(15):879–885, 1915. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1915:BMG

[Ein15c] Albert Einstein. Berichtigung zu meiner gemeinsam mit Herrn J. W. de Haas veröffentlichten Arbeit ,,Experimenteller Nachweis der Ampèreschen Molekularströme”. (German) [Correction to my collaborative published work with Mr J. W. de Haas “Experimental proof of Ampère’s molecular currents”]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(10):203, May 30, 1915. CODEN VDPEAZ. ISSN 0372-5448. See [EdH15a].

Einstein:1915:FGG

[Ein15d] Albert Einstein. Die Feldgleichungen der Gravitation. (German) [The field equations of gravitation]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):844– 847, 1915.

Einstein:1915:EPM

[Ein15e] Albert Einstein. Erklärung der Perihelbewegung des Merkur aus der allgemeinen Relativitätstheorie. (German) [Explanation of the perihelical motion of Mercury from the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):831–839, 1915.

Einstein:1915:ENAa

[Ein15f] Albert Einstein. Experimenteller Nachweis der Ampèreschen Molekularströme. (German) [Experimental proof of the Ampére molecular currents]. Die Naturwissenschaften, 3(19):237–238, May 7, 1915. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A preliminary note about [EdH15a].

Einstein:1915:GAR

[Ein15g] Albert Einstein. Grundgedanken der allgemeinen Relativitätstheorie und Anwendung dieser Theorie in Astronomie. (German) [Basic idea of the General Theory of Relativity and application of this theory in astronomy]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??): 315, 1915. Abstract only. Published in full in [Ein15j, Ein15e].

Einstein:1915:MMK

[Ein15h] Albert Einstein. Meine Meinung über den Krieg. (German) [My opinion of the war]. Published later in [Ein16k]., October/November 1915.

Einstein:1915:RGT

[Ein15i] Albert Einstein. Relativitätstheorie. (German) [Theory of Relativity]. In Die Kultur der Gegenwart. Ihre Entwicklung und ihre Ziele. Paul Hinneberg, ed. Part 3, Mathematik, Naturwissenschaften, Medizin. Section 3, Anorganischen Naturwissenschaften, pages 703–713. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1915. See Schilpp number 192 for a revised edition.

Einstein:1915:ARG

[Ein15j] Albert Einstein. Zur allgemeinen Relativitätstheorie. (German) [Toward a General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??): 778–786, 799–801, 1915.

Einstein:1916:EM

[Ein16a] A. Einstein. Ernst Mach. Physikalische Zeitschrift, 17(??):101– 104, ???? 1916. CODEN PHZTAO. ISSN 0369-982X.

Einstein:1916:GARa

[Ein16b] Albert Einstein. Die Grundlagen der Allgemeinene Relativitätstheorie. (German) [The foundations of the Theory of General Relativity]. Annalen der Physik (1900) (series 4), 354 (7):769–822, ???? 1916. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1916:NFD

[Ein16c] Albert Einstein. Eine neue formale Deutung der Maxwellschen Feldgleichungen der Elektrodynamik. (German) [A new formal interpretation of the Maxwell field equations of electrodynamics]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):184–187, 1916.

Einstein:1916:EAU

[Ein16d] Albert Einstein. Einege anschauliche Überlegungen aus dem Gebiete der Relativitätstheorie. (German) [Some descriptive considerations from the areas of Relativity Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):423, 1916.

Einstein:1916:EEN

[Ein16e] Albert Einstein. Einfaches Experiment zum Nachweis der Ampèreschen Molekularströme. (German) [Simple experiment for the proof of the Ampère molecular stream]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 18(6–7):173– 177, April 15, 1916. CODEN VDPEAZ. ISSN 0372-5448.

Einstein:1916:ETW

[Ein16f] Albert Einstein. Elementare Theorie der Wasserwellen und des Fluges. (German) [Elementary theory of water waves and of flight]. Die Naturwissenschaften, 4(??):509–510, 1916. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1916:GKS

[Ein16g] Albert Einstein. Gedächtnisrede auf Karl Schwarzschild. (German) [Memorial speech for Karl Schwarzschild (1873–1916)]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):768–770, 1916.

Einstein:1916:GARb

[Ein16h] Albert Einstein. Grundlage der allgemeinen Relativitätstheorie. (German) [The Foundations of the Theory of General Relativity]. Johann Ambrosius Barth, Leipzig, Germany, 1916. 64 pp. LCCN ???? See English translation in [EM20] and French translation in [Ein33c].

Einstein:1916:HLT

[Ein16i] Albert Einstein. H. A. Lorentz: Théories statistiques en thermodynamique. (German) [Review of H. A. Lorentz: Statistical Theories in Thermodynamics]. Die Naturwissenschaften, 4(??): 480–481, 1916. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1916:HPA

[Ein16j] Albert Einstein. Hamiltonsches Prinzip und allgemeine Relativitätstheorie. (German) [Hamilton’s Principle and General Relativity Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):1111–1116, 1916.

Einstein:1916:MMK

[Ein16k] Albert Einstein. Meine Meinung über den Krieg. (German) [My opinion of the war]. In Isidor Landau and Eugen Zabel, editors, Das Land Goethes 1914–1916: ein vaterländisches Gedenkbuch. (German) [The Land of Goethes 1914–1916: A Patriotic Memorial Book], page ???? Deutsche Verlags-Anstalt, Stuttgart and Berlin, Germany, October/November 1916.

Einstein:1916:NIF

[Ein16l] Albert Einstein. Näherungsweise Integration der Feldgleichungen der Gravitation. (German) [Approximate integration of the field equations of gravitation]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):688– 696, 1916.

Einstein:1916:SEA

[Ein16m] Albert Einstein. Strahlungs-Emission und -Absorption nach der Quantentheorie. (German) [Radiation emission and absorption according to quantum theory]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 18(13–14):318–323, July 30, 1916. CODEN VDPEAZ. ISSN 0372-5448.

Einstein:1916:FKA

[Ein16n] Albert Einstein. Über Friedrich Kottlers Abhandlung: ,,Über Einsteins Äquivalenzhypothese und die Gravitation“. (German) [On Friedrich Kottlers’s paper: “On Einstein’s Equivalence Hypothesis and Gravitation”]. Annalen der Physik (1900) (series 4), 356(22):639–642, 1916. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1916:VGF

[Ein16o] Albert Einstein. Vorwort. (German) [Foreword]. In Die Grundlagen der Einsteinschen Gravitationstheorie (German) [The Foundations of Einstein’s Theory of Gravitation] [Fre16], page ?? LCCN QC178 .F67 1916.

Einstein:1916:QSG

[Ein16p] Albert Einstein. Zur Quantentheorie der Strahlung. (German) [On the Quantum Theory of Radiation]. Mitt. phys. Ges. Zürich, 16(??):47–62, 1916. Republished in [Ein17i]. English translation in [Ein67b].

Einstein:1917:AS

[Ein17a] Albert Einstein. Alexander von Schütz. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19(5–6):41, March 30, 1917. CODEN VDPEAZ. ISSN 0372-5448.

Einstein:1917:ATG

[Ein17b] Albert Einstein. Der Angst-Traum. (German) [The nightmare]. Berliner Tageblatt, ??(??):??, December 25, 1917. This essay proposes abolishing high school final examinations.

Einstein:1917:ATJ

[Ein17c] Albert Einstein. Eine Ableitung des Theorems von Jacobi. (German) [A derivative of the theorem of Jacobi]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):606–608, 1917.

Einstein:1917:HVH

[Ein17d] Albert Einstein. H. v. Helmholtz: Zwei Vorträge über Goethe. (German) [H. v. Helmholtz: Two lectures on Goethe]. Die Naturwissenschaften, 5(??):675, 1917. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1917:KBA

[Ein17e] Albert Einstein. Kosmologische Betrachtungen zur allgemeinen Relativitätstheorie. (German) [Cosmological considerations in the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):142– 152, 1917. English translation in [Ein86e, ?].

Einstein:1917:MSG

[Ein17f] Albert Einstein. Marian von Smoluchowski [1872–1917]. (German) [Marian von Smoluchowski [1872–1917]]. Die Naturwissenschaften, 5(??):737–738, 1917. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1917:SAR

[Ein17g] Albert Einstein. Über die spezielle und die allgemeine Relativitätstheorie, gemeinverständlich. (German) [On the Special and General Theory of Relativity: a Popular Account], volume 38. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1917. 70 pp. LCCN ????

Einstein:1917:QSE

[Ein17h] Albert Einstein. Zum quantensatz von Sommerfeld und Epstein. (German) [On the quantum theory of Sommerfeld and Epstein]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19 (9–10):82–92, May 30, 1917. CODEN VDPEAZ. ISSN 0372-5448.

Einstein:1917:QSG

[Ein17i] Albert Einstein. Zur Quantentheorie der Strahlung. (German) [On the Quantum Theory of Radiation]. Physikalische Zeitschrift, 18(??):121–128, 1917. Republication of [Ein16p].

Einstein:1918:MFG

[Ein18a] A. Einstein. Motive des Forschens. (German) [Principles of research]. In Zu Max Plancks sechzigstem Geburtstag. Ansprachen, gehalten am 26. April 1918 in der Deutschen Physikalischen Gesellschaft. (German) [For Max Planck’s sixtieth birthday. Meeting held on 26 April 1918 in the German Physical Society], pages 29–32. C. F. Müller, Karlsruhe, Germany, 1918. LCCN ???? English translation in [Ein82c, pp. 224–227].

Einstein:1918:VAN

[Ein18b] Albert Einstein. An den Verein “Allgemeine Nährpflicht. (German) [To the Society “A Guaranteed Subsistence for All”]. Published in [?, document 16]., December 12, 1918.

Einstein:1918:BGN

[Ein18c] Albert Einstein. Bemerkung zu Gehrckes Notiz: Über den Äther. (German) [Remark on Gehrcke’s notes: “On the Aether”]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21–24):261, December 30, 1918. CODEN VD-PEAZ. ISSN 0372-5448.

Einstein:1918:BSN

[Ein18d] Albert Einstein. Bemerkung zu Schrödingers Notiz: Lösungssystem der allgemein Kovarienten Gravitationsgleichungen. (German) [Remark on Schrödinger’s note: Solution system of general covariant gravitation equations]. Physikalische Zeitschrift, 19(??): 165–166, 1918.

Einstein:1918:EAR

[Ein18e] Albert Einstein. Der Energiesatz in der allgemeinen Relativitätstheorie. (German) [The Law of Conservation of Energy in the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):448– 459, 1918.

Einstein:1918:DEG

[Ein18f] Albert Einstein. Dialog über Einwände gegen die Relativitätstheorie. (German) [Dialogue on objections to the Theory of Relativity]. Die Naturwissenschaften, 6(??):697–702, 1918. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1918:FAP

[Ein18g] Albert Einstein. Friedrich Adler als Physiker. (German) [Friedrich Adler as a physicist]. Die Vossische Zeitung [Morgen Ausgabe], ??(259):2, May 23, 1918.

Einstein:1918:HWR

[Ein18h] Albert Einstein. H. Weyl: Raum, Zeit, Materie. (German) [Review of H. Weyl: Space, Time, Matter]. Die Naturwissenschaften, 6(??):373, 1918. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1918:KHS

[Ein18i] Albert Einstein. Kritisches zu einer von Hrn. de Sitter gegebenen Lösung der Gravitationsgleichungen. (German) [Critical remarks on a solution of the gravitation equations given by Mr. de Sitter]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):270–272, 1918.

Einstein:1918:LSB

[Ein18j] Albert Einstein. Lassen sich Brechungsexponenten der Körper für Röntgenstrahlen experimentell ermitteln. (German) [Can refraction exponents of the bodies for X-rays be determined experimentally?]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(9–12):86–87, June 30, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Einstein:1918:NSA

[Ein18k] Albert Einstein. Notiz zu Schrödingers Arbeit: Energiekomponenten des Gravitationsfelde. (German) [Note on Schrödinger’s work: “Energy Components of Gravitation Fields”]. Physikalische Zeitschrift, 19(??):115–116, 1918.

Einstein:1918:NNA

[Ein18l] Albert Einstein. On the need for a National Assembly. Speech given to the pacifist organization Bund “Neues Vaterland”. Published in [NN60a]., November 13, 1918.

Einstein:1918:PAR

[Ein18m] Albert Einstein. Prinzipelles zur allgemeinen Relativitätstheorie. (German) [Principles of the General Theory of Relativity]. Annalen der Physik (1900) (series 4), 360(4):241–244, 1918. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1918:SAR

[Ein18n] Albert Einstein. Über die spezielle und die allgemeine Relativitätstheorie, gemeinverständlich. (German) [On the Special and General Theory of Relativity in Common Understanding], volume 38. Friedrich Vieweg und Sohn, Braunschweig, Germany, third edition, 1918. 83 pp. LCCN ????

Einstein:1918:GGG

[Ein18o] Albert Einstein. Über Gravitationswellen. (German) [On gravitational waves]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):154–167, 1918.

Einstein:1919:TSG

[Ein19a] A. Einstein. Time, space, and gravitation. The Times (London), ??(??):13–14, November 28, 1919. Reprinted as “What is the Theory of Relativity?” In [Ein54b, 227–232].

Einstein:1919:BPSa

[Ein19b] Albert Einstein. Bemerkung über periodische Schwankungen der Mondlänge, welche bisher nach der Newtonschen Mechanik nicht erklärbar schienen. (German) [Comments on periodic fluctuations of the Moon’s longitude, which so far appeared to be inexplicable in Newtonian mechanics]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):433– 436, April 24, 1919.

Einstein:1919:BPSb

[Ein19c] Albert Einstein. Bemerkung über periodische Schwankungen der Mondlänge, welche bisher nach der Newtonschen Mechanik nicht erklärbar schienen. (German) [Remark on periodic fluctuations of the Moon’s distance, which did not seem explainable so far after Newton’s mechanics]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):711, 1919. See [vB19].

Einstein:1919:ZOG

[Ein19d] Albert Einstein. Die Zuwanderung aus dem Osten. (German) [Immigration from the East]. Berliner Tageblatt, ??(??):??, December 30, 1919. Morning edition.

Einstein:1919:FAR

[Ein19e] Albert Einstein. Feldgleichungen der allgemeinen Relativitätstheorie vom Standpunkte des kosmologischen Problems und des problems der Konstitution der Materie . (German) [Field equation of the Theory of General Relativity from the point of view of the cosmological problem and the problem of the constitution of matter]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, 26(??):463, May 15, 1919. Title only. Essentially the paper [Ein19j].

Einstein:1919:IDP

[Ein19f] Albert Einstein. Induction and deduction in physics. Berliner Tageblatt, ??(??):??, December 25, 1919. Reprinted, with commentary, in [Kos98].

Einstein:1919:IUD

[Ein19g] Albert Einstein. Induktion unde Deduktion der Physik. (German) [Induction and deduction in physics]. Berliner Tageblatt, ??(??):??, December 25, 1919. Section 4.

Einstein:1919:LAP

[Ein19h] Albert Einstein. Leo Arons als Physiker. (German) [Leo Arons as physicist]. Sozialistische Monatshefte (Jarhgang 52, part 2), ??(??):1055–1056, 1919.

Einstein:1919:PAR

[Ein19i] Albert Einstein. Prüfung der allgemeinen Relativitätstheorie. (German) [A test of the General Theory of Relativity]. Die Naturwissenschaften, 7(??):776, 1919. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1919:SGI

[Ein19j] Albert Einstein. Spielen die Gravitationsfelder im Aufbau der materiellen Elementarteilchen eine wesentliche Rolle?. (German) [Do gravitational fields play a substantial role in the structure of the elementary particles of matter?]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, 20(??):349–356, April 10, 1919. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/69YV118A.

Einstein:1919:WRG

[Ein19k] Albert Einstein. Was ist Relativitätstheorie? (German) [What is the theory of relativity?]. ????, ??(??):??, 1919. Reprinted in [See91, pp. 209ff].

Einstein:1920:AA

[Ein20a] Albert Einstein. Assimilation und Antisemitismus. ????, ?? (??):??, April 3, 1920.

Einstein:1920:ART

[Ein20b] Albert Einstein. Äther und die Relativitäts-Theorie: Rede Gehalten am 5. Mai 1920 an der Reichs-Universität zu Leiden. (German) [Aether and the Theory of Relativity. Speech Held on 5 May 1920 at the State University of Leiden]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1920. 15 pp. URL http://www.mountainman.com.au/aether_

0.html. An address delivered on May 5th, 1920, in the University of Leyden. English translation in [Ein22l, Ein83]. French translation in [Ein21k] and Italian translation in [Ein22k].

Einstein:1920:ARG

[Ein20c] Albert Einstein. Äther und Relativitätstheorie. (German) [Ether and The Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1920. ???? pp. Einstein’s lecture at the University of Leiden after his appointment as Visiting Professor.

Einstein:1920:BAW

[Ein20d] Albert Einstein. Bemerkung zur Abhandlung von W. R. Hess: Theorie der Viscosität heterogener Systeme. (German) [Remark on the paper on W. R. Hess: “Theory of the viscosity of heterogeneous systems”]. Kolloidzschr, 27(??):137, 1920.

Einstein:1920:C

[Ein20e] Albert Einstein. A confession. Israelistisches Wochenblatt für die Schweiz, ??(??):??, September 24, 1920. Comments on anti-Semitism.

Einstein:1920:IIC

[Ein20f] Albert Einstein. Interview on interplanetary communication. Daily Mail (London), ??(??):??, January 31, 1920. Quoted in New York Times, February 2, 1920.

Einstein:1920:ILS

[Ein20g] Albert Einstein. Inwiefern lässt sich die moderne Gravitationstheorie ohne die Relativität begrunden. (German) [In what respect can the modern Theory of Gravitation be established without Relativity?]. Die Naturwissenschaften, 8(51):1010–1011, December 17, 1920. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1920:MAA

[Ein20h] Albert Einstein. Meine Antwort über die antirelativitätstheoretische G.m.b.H. (Gesellschaft mit beschränkter Haftung). (German) [My answer to Antirelativity Theoretical Co. Ltd.]. Berliner Tageblatt und Handelszeitung, ??(??):1–2, August 17, 1920.

Einstein:1920:NSE

[Ein20i] Albert Einstein. On new sources of energy. Berliner Tageblatt, ??(??):??, July 25, 1920. Morning edition. Comment on possible use of atomic energy.

Einstein:1920:CII

[Ein20j] Albert Einstein. On the contribution of intellectuals to international reconciliation. In Thoughts on Reconciliation, page ?? Deutscher Gesellig-Wissenschaftlicher Verein von New York, New York, NY, USA, 1920. LCCN D645 .D4 FT MEADE. Invited and compiled by the Social and scientific society of New York, in commemoration of the fiftieth anniversary of its foundation. 1870–1920.

Einstein:1920:RSG

[Ein20k] Albert Einstein. Relativity, the Special and the General Theory: a Popular Exposition. Methuen and Co. Ltd., London, UK, 1920. xiii + 138 pp. LCCN ???? URL http://www.

gsjournal.net/Science-Journals/Essays/View/2494. Au-

thorized translation by Robert W. Lawson from the fifth German edition [Ein18n].

Einstein:1920:STD

[Ein20l] Albert Einstein. Schallausbreitung in teilweise dissoziierten Gasen. (German) [Sound propagation in partly dissociated gases]. S. B. brews. Akad. Wzk., ??(??):380–385, 1920.

Einstein:1920:GAP

[Ein20m] Albert Einstein. To the “General Assembly for Popular Technical Education”. Neue Freie Presse, ??(??):??, July 24, 1920.

Einstein:1920:TWG

[Ein20n] Albert Einstein. Tragheitsmoment des Wasserstoffmolekuls. (German) [Moment of inertia of the hydrogen molecule]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):65, 1920.

Einstein:1920:SAR

[Ein20o] Albert Einstein. Über die spezielle und die allgemeine Relativitätstheorie, gemeinverständlich. (German) [On the Special and General Theory of Relativity in Common Understanding]. Friedrich Vieweg und Sohn, Braunschweig, Germany, tenth edition, 1920. 91 pp. LCCN ????

Einstein:1920:ULH

[Ein20p] Albert Einstein. Uproar in the lecture hall. 8-Uhr Abendblatt, ?? (??):??, February 13, 1920. News story of Einstein lecture being heckled by anti-Semitic “students”.

Einstein:1920:EBK

[Ein20q] Albert Einstein. Zur elektrodynamic bewegter korper. (German) [On the electrodynamics of moving bodies]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1920. 26–50 pp. LCCN ????

Einstein:1921:GEG

[Ein21a] A. Einstein. Geometrie und Erfahrung. (German) [Geometry and experience]. In Erweiterte Fassung des Festvortrages gehalten an der Preussischen Akademie der Wissenschaften zu Belin am 27. Januar 1921, pages 123–130. Julius Springer, Berlin, Germany, 1921. LCCN ???? English translation in [Ein54b, 232–246].

Einstein:1921:BOD

[Ein21b] Albert Einstein. A brief outline of the development of the Theory of Relativity. Nature, 106(2677):782–784, February 17, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v106/n2677/pdf/106782a0.pdf;   http://www.nature.com/
physics/looking-back/einstein/einstein.pdf. Translated

by Robert W. Lawson from the German manuscript. Special issue on Relativity. This is one of only three papers by foreign authors in this issue; the others are [Lor21, Wey21].

Einstein:1921:GKW

[Ein21c] Albert Einstein. Das Gemeinsame am künstlerischen und wissenschaftlichen Erleben. (German) [The common element in scientific and artistic experience]. Menschen. Zeitschrift neuer Kunst, 4(??):19, 1921.

Einstein:1921:DWG

[Ein21d] Albert Einstein. Die Not der deutschen Wissenschaft. Ein Gefahr für die Nation. (German) [The plight of German science: a danger for the nation]. Neue Freie Presse, ??(??):??, December 25, 1921.

Einstein:1921:EAN

[Ein21e] Albert Einstein. Eine Einfache Anwendung des Newtonschen Gravitationsgesetzes auf die Kugelförmigen Sternhaufen. (German) [A simple application of the Newtonian Law of Gravitation to globular star clusters]. Kaiser Wilhelm Gesellschaft zür Forderung der Wissenchaft, Festschrift, ??(??):50–52, 1921. English translation in [?, pp. 420–425].

Einstein:1921:EE

[Ein21f] Albert Einstein. Einstein on education. Nation and Athenaeum, 30(??):378–379, 1921.

Einstein:1921:EAE

[Ein21g] Albert Einstein. Einsteins amerikanische Eindrücke. Was er wirklich sah. (German) [Einstein’s impressions of America: What he really saw]. Die Vossische Zeitung [Morgen Ausgabe], ??(??):??, July 10, 1921. Partial quotation, and English translation, in [Sop88].

Einstein:1921:GEE

[Ein21h] Albert Einstein. Geometrie und Erfahrung: Erweiterte Fassung des Festvortrages gehalten an der preussischen Akademie. (German) [Geometry and Experience: Extended version of the commemorative address held at the Prussian Academy]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1921. 20 pp. LCCN ????

Einstein:1921:GLF

[Ein21i] Albert Einstein. La Géométrie et L’expérience. (French) [Geometry and Experience]. Gauthier-Villars, Paris, France, 1921. 20 pp. LCCN ???? French translation by Maurice Solovine of [Ein21h].

Einstein:1921:TRR

[Ein21j] Albert Einstein. La Théorie de la Relativité Restreinte de Généralisée. (French) [The Special and General Theory of Relativity]. Gauthier, Paris, France, 1921. xii + 120 pp. LCCN ???? French translation by Mlle. J. Rouviere of the tenth German edition [Ein20o].

Einstein:1921:LTR

[Ein21k] Albert Einstein. L’éther et la Théorie de Relativité. (French) [Aether and the Theory of Relativity]. Gauthier, Paris, France, 1921. 15 pp. LCCN ???? French translation by Maurice Solovine of [Ein20b].

Einstein:1921:MRF

[Ein21l] Albert Einstein. The Meaning of Relativity: four lectures delivered at Princeton University, May, 1921. Princeton University Press, Princeton, NJ, USA, 1921. v + 123 pp. LCCN QC6. Reprinted in 1922 and 1923 by the same publisher, and by Methuen (London) in 1922 and 1924.

Einstein:1921:JP

[Ein21m] Albert Einstein. On a Jewish Palestine. Jüdische Rundschau, ??(??):371, July 1, 1921.

Einstein:1921:RSG

[Ein21n] Albert Einstein. Relativity, the Special and the General Theory: a Popular Exposition. Holt, New York, NY, USA, 1921. xiii + 168 pp. LCCN ???? This differs from the English edition [Ein20k] only in pagination. Reprinted in 1931 (Smith, New York) and in 1947 (Hartsdale House, Inc., New York).

Einstein:1921:RLK

[Ein21o] Albert Einstein. Report of lecture at King’s College on the development and present position of Relativity, with quotations. Nation and Athenaeum, 29(??):431–432, 1921.

Einstein:1921:STS

[Ein21p] Albert Einstein. Sulla Teoria Speciale e Generale della Relatività: Volgarizzazione. (Italian) [On the Special and General Theory of Relativity: Popular Edition]. Zanichelli, Bologna, Italy, 1921. xii + 125 pp. LCCN ???? Italian translation G. L. Calisse of [Ein20o].

Einstein:1921:TRE

[Ein21q] Albert Einstein. Teoria de la Relatividad Especial y General. (Spanish) [Theory of Special and General Relativity]. Peláez, Toledo, Spain, 1921. 79 pp. LCCN ???? Translation to Spanish of the 12th edition by F. Lorente.

Einstein:1921:TOO

[Ein21r] Albert Einstein. Teoriaa Otnositel’nosti: Obshchedostypnoe Izloshenie. (Russian) [On the Special and General Theory of Relativity in Common Understanding]. Slowo, Berlin, Germany, 1921. 150 pp. LCCN ???? Russian translation by G. B. Itel’son of [Ein20o].

Einstein:1921:ELB

[Ein21s] Albert Einstein. Über ein den Elementarprozess der Lichtemission betreffendes Experiment. (German) [On an experiment concerning the elementary process of light emission]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 2, ??(??):882–883, 1921.

Einstein:1921:NEF

[Ein21t] Albert Einstein. Über eine naheliegende Ergänzung des Fundamentes der allgemeinen Relativitätstheorie. (German) [A natural addition to the foundation of the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Part 1, ??(??):261–264, 1921. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/EVSSPDP8.

Einstein:1921:WIZ

[Ein21u] Albert Einstein. Wie ich Zionist Wurde. (German) [How I became a Zionist]. Jüdische Rundschau, ??(??):351–352, June 21, 1921.

Einstein:1921:AGM

[Ein21v] Albert Einstein. Zur Abwehr. (German) [In my defense]. Die Naturwissenschaften, 9(??):261–264, 1921. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1921:EHU

[Ein21w] Albert Einstein. Zur Errichtung der hebräischen Universität in Jerusalem. (German) [On the founding of the Hebrew University of Jerusalem]. Jüdische Pressezentrale Zürich, ??(??):??, August 26, 1921.

Einstein:1922:KAR

[Ein22a] Albert Einstein. A Különleges es az Általános Relativitás, Elmélette. (Hungarian) [On the Special and General Theory of Relativity in Common Understanding]. Pantheon irodalmi, Budapest, Hungary, 1922. 94 pp. LCCN ???? Hungarian translation of [Ein20o].

Einstein:1922:BAT

[Ein22b] Albert Einstein. Bemerkung zu der Abhandlung von E. Trefftz: Statische Gravitationsfeld zweien Massenpunkte. (German) [Remark on the paper by E. Trefftz: “Static gravitational field of two point masses”]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):448–449, 1922.

Einstein:1922:BAF

[Ein22c] Albert Einstein. Bemerkung zu der Arbeit von A. Friedmann: Über die Krümmung des Räumes. (German) [Remarks on the work of A. Friedmann: On the Curvature of Space]. Zeitschrift für Physik, 11(??):326, 1922. CODEN ZEPYAA. ISSN 0044-3328. English translation in [Ein86d].

Einstein:1922:BSA

[Ein22d] Albert Einstein. Bemerkung zu der Seletyschen Arbeit: ,,Beiträge zum Kosmologischen System“. (German) [Remark on the Selety work: “Contributions to the cosmological system”]. Annalen der Physik (1900) (series 4), 374(22):436–438, 1922. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1922:CG

[Ein22e] Albert Einstein. Conditions in Germany. New Republic, 32(??): 197, 1922. ISSN 0028-6583.

Einstein:1922:EWF

[Ein22f] Albert Einstein. Emil Warburg als Forscher. (German) [Emil Warburg as a researcher]. Die Naturwissenschaften, 10(??):823– 828, 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1922:EBG

[Ein22g] Albert Einstein. Experiment betreffend die Gütigkeitsgrenze der Undulationstheorie. (German) [Experiment concerning the limits of validity of the wave theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):4, 1922.

Einstein:1922:ISD

[Ein22h] Albert Einstein. Impact of science on the development of pacifism. In Die Friedensbewegung. (German) [The Peace Movement], pages 78–79. Schwetschke, Berlin, Germany, 1922.

Einstein:1922:MWR

[Ein22i] Albert Einstein. In memorium Walther Rathenau. Neue Rundschau, 33 (part 2)(??):815–816, 1922.

Einstein:1922:FPP

[Ein22j] Albert Einstein. O Fizicheskoi Prirodie Prostranstva. (Russian) [Physical Nature of Space]. Slowo, Berlin, Germany, 1922. 52 pp. LCCN ???? Russian translation of [Ein20b, Ein21h].

Einstein:1922:PRD

[Ein22k] Albert Einstein. Prospettive Relativistiche dell’etere e della Geometria. (Italian) [Relativistic perspectives of the Aether and Geometry]. Andare, Milan, Italy, 1922. 54 pp. LCCN ???? Translation to Italian by R. Cantù and Te. Bembo of [Ein20b, Ein21h].

Einstein:1922:SRE

[Ein22l] Albert Einstein. Sidelights on Relativity: I. Ether and Relativity. II. Geometry and experience. Methuen and Co. Ltd., London, UK, 1922. 3–56 pp. LCCN QC6 .E53; Microfilm 30075 QC. Translation of the German Äther und Relativitäts-Theorie and Geometrie und Erfahrung by G. B. (George Barker) Jeffery and W. Perrett.

Einstein:1922:TBS

[Ein22m] Albert Einstein. Theoretische Bemerkungen zur Supraleitung der Metalle. (German) [Theoretical observations on superconductivity of metals]. In Leyden Rijksuniversitet Natuurkundig Laboratorium. Gedenkboek aangeboden aan H. Kamerlingh Onnes, pages 429–435. Ijdo, Leyden, The Netherlands, 1922. English translation in [Ein05-49].

Einstein:1922:TLD

[Ein22n] Albert Einstein. Theorie der Lichtfortpflanzung in dispergierenden Medien. (German) [Theory of light propagation in dispersive media]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ?? (??):18–22, 1922.

Einstein:1922:UTB

[Ein22o] Albert Einstein. Untersuchung über die Theorie der Brownschen Bewegungen. (German) [Investigation on the Theory of Brownian Motion], volume 199 of Ostwalds Klassiker der exakten Wissenschaften. Akademische Verlagsgesellschaft, Leipzig, Germany, 1922. 72 pp. LCCN ???? Edited by R. Fürth.

Einstein:1922:VVR

[Ein22p] Albert Einstein. Vier Vorlesungen über Relativtätstheorie, Gehalten im Mai, 1921, an der Universität Princeton. (German) [The Meaning of Relativity: four lectures delivered at Princeton University, May, 1921]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1922. 70 pp. LCCN ???? German text of [Ein21l].

Einstein:1922:WPR

[Ein22q] Albert Einstein. W. Pauli: Relativitätstheorie. (German) [Review of W. Pauli: The Theory of Relativity]. Die Naturwissenschaften, 10(??):184–185, 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1922:WIR

[Ein22r] Albert Einstein. Wie ich die Relativitätstheorie entdeckte. (German) [How I discovered the Theory of Relativity]. Lecture at the University of Kyoto, Japan. Transcription from notes by Yun Ishiwara and translation into English by Y. A. Ono in Physics Today, August 1932, page 45., December 14, 1922.

Einstein:1923:BNW

[Ein23a] Albert Einstein. Bemerkung zu der Notiz von W. Anderson: Neue Erklärung des continuierlichen Koronaspectrums. (German) [Remark on the note of W. Anderson: “New explanation of the continuous corona spectrum”]. Astronomische Nachrichten, 219(1):19–20, 1923. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Einstein:1923:BMA

[Ein23b] Albert Einstein. Bemerkung zu meiner Arbeit “Zur allgemeinen Relativitätstheorie”. (German) [Remarks on my work “On the General Theory of Relativity”]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, 1923(??):76–77, 1923. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/EVSSPDP8.

Einstein:1923:BFM

[Ein23c] Albert Einstein. Bietet die Feldtheorie Möglichkeiten für die Lösung des Quantenproblems?. (German) [Does field theory offer possibilities for the solution of quantum problems?]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):359–

364, 1923. URL http://einstein-annalen.mpiwg-berlin.
mpg.de/related_texts/sitzungsberichte/MXYDBZW5.

Einstein:1923:GPR

[Ein23d] Albert Einstein. Grundgedanken und Probleme der Relativitätstheorie. (German) [Basic idea and problems of the Theory of Relativity], page 10. Imprimerie royale, Stockholm, Sweden, 1923. Nobel lecture delivered in Göteborg, Sweden on July 11, 1923.

Einstein:1923:MIP

[Ein23e] Albert Einstein. My impressions of Palestine. New Palestine, ??(??):341, 1923.

Einstein:1923:NAF

[Ein23f] Albert Einstein. Notiz zu der Arbeit von A. Friedmann “Über die Krümmung des Raumes”. (German) [Notice on the work of A. Friedmann “On the Curvature of Space”]. Zeitschrift für Physik, 16(1):228, December 1923. URL http:

//www.springerlink.com/content/h3544p7u24671460/. En-

glish translation in [Ein86d].

Einstein:1923:TRF

[Ein23g] Albert Einstein. Théorie de relativité. (French) [Theory of Relativity]. Bull. Soc. philom., Paris, 22(??):97, 98, 101, 107, 111– 112, 1923.

Einstein:1923:TRS

[Ein23h] Albert Einstein. Theorie relativity speciálni i obecná: Lehce srozumitelný výklad. Se zvláštni předmluvou autorovou k českému vyd. (Czech) [Special and General Relativity: Easy to understand interpretation. With a special foreword by the author to the Czech edition ]. Fr. Borový, Prague, Czechoslovakia, 1923. 103 pp.

Einstein:1923:TAF

[Ein23i] Albert Einstein. Theory of the affine field. Nature, 112(2812): 448–449, September 22, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Einstein:1923:AFG

[Ein23j] Albert Einstein. Zur affinen Feldtheorie. (German) [On affine field theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte,

??(??):137–140, 1923.  URL http://einstein-annalen.
mpiwg-berlin.mpg.de/related_texts/sitzungsberichte/
GBEHMB2D.

Einstein:1923:ARG

[Ein23k] Albert Einstein. Zur allgemeinen Relativitätstheorie. (German) [On the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, 1923(??):32–38, 1923. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/B98WXTVX.

Einstein:1924:RE

[Ein24a] A. Einstein. Review of Elsbach 1924. Deutsche Literaturzeitung, 45(??):1688–1689, ???? 1924. See [Els24].

Einstein:1924:ABW

[Ein24b] Albert Einstein. Antwort auf eine Bemerkung von W. Anderson. (German) [Answer to a remark of W. Anderson]. Astronomische Nachrichten, 221(20):329–330, 1924. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Einstein:1924:BJW

[Ein24c]  Albert Einstein.  Buchrezension: J. Winternitz, Rela-

tivitätstheorie und Erkenntnislehre. (German) [Book review: J.

Winternitz, Theory of Relativity and Epistemology]. Deutsche Literaturzeitung für Kritik der internationalen Wissenschaft, 45 (NS 1)(??):20–22, ???? 1924.

Einstein:1924:KEG

[Ein24d] Albert Einstein. Das Komptonsche Experiment. (German) [The Compton experiment]. Berliner Tageblatt, ??(??):??, April 20, 1924.

Einstein:1924:GGI

[Ein24e] Albert Einstein. Geleitwort. (German) [introduction]. In Diels [Die24], page ??

Einstein:1924:LAE

[Ein24f] Albert Einstein. Letter from Albert Einstein to the Secretary-General of the League of Nations in Geneva. Web document,

June 25, 1924. URL http://www.wdl.org/en/item/11595/.

Einstein:1924:MED

[Ein24g] Albert Einstein. Meinung zu Engels ‘Dialektik der Natur’. (German) [Opinion on Engels’ Dialectics of Nature]. ??, ??(??):??, June 30, 1924.

Einstein:1924:QEI

[Ein24h] Albert Einstein. Quantentheorie des einatomigen idealen Gases. (German) [Quantum theory of monatomic ideal gases]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):261– 267, 1924. See part 2 [Ein25h]. In the two papers, Einstein extended Bose’s work on monatomic gases [Bos24a, Bos24b] to predict the Bose–Einstein effect (and likely, develop Bose– Einstein statistics).

Einstein:1924:RGB

[Ein24i] Albert Einstein. [Rezension]. (German) [Book review]. Deutsche Literaturzeitung, 1(??):20–??, ???? 1924.

Einstein:1924:AGA

[Ein24j] Albert Einstein. Über den Äther. (German) [On the aether]. Verh. natur. Ges. Basel, 105 (part 2)(??):85–93, 1924.

Einstein:1924:HGL

[Ein24k] Albert Einstein. Zum hundertjährigen Gedentag von Lord Kelvins Geburt. (German) [On the hundredth anniversary of Lord Kelvin’s birth]. Die Naturwissenschaften, 12(??):601–602, 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1924:TRG

[Ein24l] Albert Einstein. Zur Theorie der Radiometerkräfte. (German) [On the theory of radiometric forces]. Zeitschrift für Physik, 27 (??):1–6, 1924. CODEN ZEPYAA. ISSN 0044-3328.

Einstein:1925:ABW

[Ein25a] Albert Einstein. Addendum to Bose, Warmegleichgewicht im Strahlungsfeld bei Anwesenheit von Materie. (German) [Addendum to Bose, “thermal equilibrium in radiation fields in the presence of matter”]. Zeitschrift für Physik, 27(??):392–393, 1925. CODEN ZEPYAA. ISSN 0044-3328.

Einstein:1925:BPJ

[Ein25b] Albert Einstein. Bemerkung zu P. Jordans Abhandlung: Theorie der Quantenstrahlung. (German) [Remark on P. Jordan’s paper: “Theory of Quantum Radiation”]. Zeitschrift für Physik, 31(??):784–785, 1925. CODEN ZEPYAA. ISSN 0044-3328.

Einstein:1925:ETH

[Ein25c] Albert Einstein. Eddingtons Theorie und Hamiltonsches Prinzip. (German) [Eddington’s theory and the Hamiltonian Principle]. In Sir Arthur Stanley Eddington, editor, Relativitätstheorie in mathematischer Behandlung. (German) [The Theory of Relativity in Mathematical Treatment], pages 366– 371. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1925. Appendix.

Einstein:1925:EFG

[Ein25d] Albert Einstein. Einheitliche Feldtheorie von Gravitation und Elektrizität. (German) [Unified Field Theory of gravitation and electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte,

??(??):414–419, 1925.  URL http://einstein-annalen.
mpiwg-berlin.mpg.de/related_texts/sitzungsberichte/
ZHDAXTXQ.

Einstein:1925:EAR

[Ein25e] Albert Einstein. Elektron und allgemeine Relativitätstheorie. (German) [The electron and the General Theory of Relativity]. Physica, 5(??):330–334, 1925. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Einstein:1925:MOU

[Ein25f] Albert Einstein. Mission of our university. New Palestine, 8 (??):294, 1925.

Einstein:1925:PE

[Ein25g] Albert Einstein. Pan Europa. La Prensa [Buenos Aires??], ?? (??):4, March 24, 1925.

Einstein:1925:QEI

[Ein25h] Albert Einstein. Quantentheorie des einatomigen idealen Gases. (German) [Quantum theory of monatomic ideal gases]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):3–

14, 1925. URL http://www.condmat.uni-oldenburg.de/
TeachingSP/einstein_2.ps. See part 1 [Ein24h].

Einstein:1925:QIG

[Ein25i] Albert Einstein. Quantentheorie des idealen Gases. (German) [Quantum theory of ideal gases]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):18–25, 1925. This paper predicts Bose–Einstein condensation.

Einstein:1926:GNE

[Ein26a] Albert Einstein. Geometria no euclidea y fisica. (Spanish) [Non-Euclidean geometry and physics]. Revista mathematica hispanoamerica (Buenos-Aires), series 2, 1(??):72–76, April 16, 1926.

Einstein:1926:IDK

[Ein26b] Albert Einstein. Über die Interferenzeigenschaften des durch Kanalstrahlen emittierten Lichtes. (German) [On the interference characteristics of light emitted by canal rays]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):334–340, 1926.

Einstein:1926:UMF

[Ein26c] Albert Einstein. Über die Ursache der Maänderbildung der Flussläufe und des sogenannten Baerschen Gesetzes. (German) [On the causes of meandering of river courses and the so-called Baer Law]. Die Naturwissenschaften, 14(??):223–224, 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Read before the Prussian Academy of Sciences on 7 January 1926. English translation in [Ein34n] and [Ein82c, pp. 249– 253].

Einstein:1926:VEN

[Ein26d] Albert Einstein. Vorschlag zu einem die Natu des elementaren Strahlungs-Emissions-prozesses betreffendes Experiment. (German) [Suggestion for an experiment on the nature of the elementary radiation emission process]. Die Naturwissenschaften, 14(??):300–301, 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1927:FBR

[Ein27a] A. Einstein. Über die formale Beziehung des Riemannschen Krümmungstensors zu den Feldgleichungen der Gravitation. (German) [The formal relationship of the Riemann curvature tensor to the field equations of gravitation]. Mathematische Annalen, 97(1):99–103, 1927. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Einstein:1927:BSW

[Ein27b] Albert Einstein. Bestimmt Schrödingers Wellenmechanik die Bewegung eines Systems vollständig oder nur im Sinne der Statistik? (German) [Does Schrödinger’s wave mechanics determine the motion of a system completely, or only in the sense of statistics?]. Read before the Prussian Academy of Sciences, but never published., May 5, 1927. URL http://alberteinstein.

info/vufind1/Record/EAR000034338.

Einstein:1927:EEL

[Ein27c] Albert Einstein. Einfluss der Erdbewegung auf die Lichtgeschwindigkeit relative zur Erde. (German) [Influence of the Earth movement on the speed of light relative to the Earth]. Forsch. Fortschr. dtsch. Wiss., 3(??):36–37, 1927.

Einstein:1927:IN

[Ein27d] Albert Einstein. Isaac Newton. Nature, 119(2995):467, March 26, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). Letter to the Royal Society on the two-hundredth anniversary of Newton’s death. English translation in [Ein82c, pp. 253–261].

Einstein:1927:NMI

[Ein27e] Albert Einstein. Newtons Mechanik und ihr Einfluss auf die Gestaltung der theoretischen Physik. (German) [Newton’s mechanics and its influence on the shaping of theoretical physics]. Die Naturwissenschaften, 15(??):273–276, 1927. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Essay written on the two-hundredth anniversary of Newton’s death. English translation in [Ein34n] and [Ein54b].

Einstein:1927:KTZ

[Ein27f] Albert Einstein. Zu Kaluzas Theorie des Zusammenhanges von Gravitation und Elektrizität. (German) [On Kaluza’s theory on the relation of gravitation and electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):23–30, 1927. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/75ASGAHY.

Einstein:1928:HL

[Ein28a] Albert Einstein. H. A. Lorentz. Mathematisch-naturwissenschaftliche Blätter, 22(??):24–25, 1928. Reprinted as “Address at the Grave of H. A. Lorentz” in [Ein54b].

Einstein:1928:IN

[Ein28b] Albert Einstein. Isaac Newton. In C. G. Abbot, editor, Annual Report of the Board of Regents of The Smithsonian Institution, Showing the Operations, Expenditures, and Condition of the Institution for the Year Ending June 30, 1927, pages 201–208. United States Government Printing Office, Washing-

ton, DC, USA, 1928. URL https://archive.org/stream/
annualreportofbo1927smit/annualreportofbo1927smit_djvu.
txt.

Einstein:1928:NME

[Ein28c] Albert Einstein. Neue Möglichkeiten für eine einheitliche Feldtheorie von Gravitation und Elektrizität. (German) [New possibilities for a unified field theory of gravitation and electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??): 224–227, 1928.

Einstein:1928:PDR

[Ein28d] Albert Einstein. A propos de ‘La déduction relativiste’ de M. E. Meyerson. (French) [on M. E. Meyerson’s article “Relativistic Deduction”]. Revue Philosophique de France, 105(??):161–166, 1928.

Einstein:1928:RGA

[Ein28e]  Albert Einstein.  Riemann-geometrie mit Aufrechterhal-

tung des Begriffes des Fernparallelismus. (German) [Riemannian geometry with preservation of distant parallelism]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):217–

221, 1928. URL http://einstein-annalen.mpiwg-berlin.
mpg.de/related_texts/sitzungsberichte/B8AG2G5E.

Einstein:1929:BST

[Ein29a] Albert Einstein. Broadcast on the semicentennial of Thomas A. Edison’s incandescent light. Reprinted in the New York Times, October 23, 1929, 1929.

Einstein:1929:EFH

[Ein29b] Albert Einstein. Einheitliche Feldtheorie und Hamiltonsches Prinzip. (German) [Unified Field Theory and the Hamiltonian Principle]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte,

??(??):156–159, 1929.  URL http://einstein-annalen.
mpiwg-berlin.mpg.de/related_texts/sitzungsberichte/
EVNERTBY.

Einstein:1929:EIG

[Ein29c] Albert Einstein. Einheitliche Interpretation von Gravitation un Elektrizität. (German) [Unified interpretation of gravitation and electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ?? (??):102, 1929.

Einstein:1929:NFTa

[Ein29d] Albert Einstein. The new field theory. New York Times, ?? (??):??, February 3, 1929. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. Report on [Ein29k].

Einstein:1929:NFTb

[Ein29e] Albert Einstein. The new field theory. The Times, London, ?? (??):??, February 4, 1929.

Einstein:1929:NFTc

[Ein29f] Albert Einstein. The new field theory. The Observatory, 51(??): 82–87, 1929. CODEN OBSEAR. ISSN 0029-7704.

Einstein:1929:STb

[Ein29g] Albert Einstein. On scientific truth. Gelegentliches, ??(??):??, 1929. Answers to questions from a Japanese scholar. English translation in [Ein82c, pp. 261–262].

Einstein:1929:SEA

[Ein29h] Albert Einstein. Sesión especial de la academia, 16 April 1925. An. Soc. cient. argent., 107(??):337–347, 1929.

Einstein:1929:STa

[Ein29i] Albert Einstein. Space–time. In Encyclopaedia Brittanica, volume 21, pages 105–108. ????, ????, 14th edition, 1929. LCCN ????

Einstein:1929:GSF

[Ein29j] Albert Einstein. Über den gegenwärtigen Stand der Feldtheorie. (German) [On the current status of field theory]. In E. Honegger, editor, Festschrift für Professor A. Stodola zum 70. Geburtstag. Überreicht von seinen Freunden und Schülern. (German) [Festschrift for Professor A. Stodola on his 70th birthday. Presented by his friends and students], pages 126–132. Orel Füssli Verlag, Zürich, Switzerland, 1929.

Einstein:1929:EFG

[Ein29k] Albert Einstein. Zur Einheitlichen Feldtheorie. (German) [On the Unified Field Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):2–7, 1929.

Einstein:1930:RMF

[Ein30a] A. Einstein. Auf die Riemann-Metrik und den Fern-Parallelismus gegründete einheitliche Feldtheorie. (German) [On the Riemann Metric and the distant parallelism created by Uniform Field Theory]. Mathematische Annalen, 102(1): 685–697, 1930. CODEN MAANA3. ISSN 0025-5831 (print),

1432-1807 (electronic). URL http://www.lrz-muenchen.de/~
u7f01bf/WWW/einstein1930.intro.html.

Einstein:1930:RAFb

[Ein30b] A. Einstein. Das Raum-, Äther- und Feld-Problem der Physik. (German) [Space, ether, and field problems of physics]. Transactions of the World Power Conference, 2nd, Berlin, 19(??): 1–5, 1930. English translation in [Ein54b, 276–285].

Einstein:1930:TUD

[Ein30c] A. Einstein. Théorie unitaire du champ physique. (French) [The unitary theory of the physical field]. Ann. Inst. H. Poincaré, 1 (1):1–24, 1930. ISSN 0365-320X. URL http://www.numdam.

org/item?id=AIHP_1930__1_1_1_0.

Einstein:1930:BGB

[Ein30d] Albert Einstein. [Besprechung]. (German) [book review]. Die Naturwissenschaften, 18(22):536, May 30, 1930. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01513429.

Einstein:1930:RZP

[Ein30e] Albert Einstein. Das Raum-Zeit Problem. (German) [The space–time problem]. Koralle, 5(??):486–488, 1930.

Einstein:1930:KFE

[Ein30f] Albert Einstein. Die Kompatibilität der Feldgleichungen in der einheitlichen Feldtheorie. (German) [Compatibility of the field equations in Uniform Field Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):18–23, 1930. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/NMS7SWKE.

Einstein:1930:PEA

[Ein30g] Albert Einstein. Professor Einstein’s address at the University of Nottingham. Science, 71(1850):608–611, June 13, 1930. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://science.sciencemag.org/content/71/
1850/608.

Einstein:1930:RAFa

[Ein30h] Albert Einstein. Raum, Äther und Feld in der Physik. (German) [Space, aether and field in physics]. Forum philosophicum, 1(??): 173–180, 1930.

Einstein:1930:RFG

[Ein30i] Albert Einstein. Rede zur Funkausstellung. (German) [Speech at the Broadcasting Exhibition]. Die Naturwissenschaften, 49: 33, 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Transcribed by F. Herneck from the sound recording made in Berlin, August 22, 1930.

Einstein:1930:RS

[Ein30j] Albert Einstein. Religion and science. New York Times, ?? (??):??, November 9, 1930. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. Reprinted in [Ein31d] and [Ein54b, pages 36–40].

Einstein:1930:RWG

[Ein30k] Albert Einstein. Religion und Wissenschaft. (German) [Religion and science]. Berliner Tageblatt, ??(??):??, November 11, 1930.

Einstein:1930:SD

[Ein30l] Albert Einstein. Science and dictatorship. In Otto Forst-Battaglia, editor, Prozess der diktatur. (German) [Dictatorship on its trial], page ?? G. G. Harrap & Co. Ltd., London, UK, 1930. LCCN D107 .F65.

Einstein:1930:SGD

[Ein30m] Albert Einstein. Science and God: a dialogue. Forum and Century, 83(??):373–379, 1930.

Einstein:1930:GSA

[Ein30n] Albert Einstein. Über den gegenwärtigen Stand der allgemeinen Relativitätstheorie. (German) [On the current state of the General Theory of Relativity]. Yale Univ. Lib. Gaz., 6(??):3–6, 1930.

Einstein:1930:FEF

[Ein30o] Albert Einstein. Über die Fortschritte der einheitlichen Feldtheorie. (German) [On progress made by the Unified Field Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):102, 1930.

Einstein:1930:KGA

[Ein30p] Albert Einstein. Über Kepler. (German) [About Kepler]. Frankfurter Zeitung, ??(??):??, November 9, 1930. Written on the three-hundredth anniversary of the death of Johannes Kepler. English translation in [Ein34n] and [Ein82c, pp. 262–266].

Einstein:1930:WB

[Ein30q] Albert Einstein. What I believe. Forum and Century, 84(??): 193–194, 1930. Reprinted in [Ein54b, pages 8–11].

Einstein:1930:TRR

[Ein30r] Albert Einstein. Zur Theorie der Räume mit Riemann-metrik und Fernparallelismus. (German) [Theory of space with Riemann metrics and distant parallelism]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):401–402, 1930. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/K17BZ9SW.

Einstein:1931:PGB

[Ein31a] A. Einstein. Physikalische Gesellschaft zu Berlin und Deutsche Gesellschaft für technische Physik. Berlin, 17. Juli 1931. (German) [Physical Society of Berlin and the German Society for Technical Physics, Berlin 17 July 1931]. Angewandte Chemie, 44(33):685, August 15, 1931. CODEN ANCEAD. ISSN 1521-3757.

Einstein:1931:DC

[Ein31b] Albert Einstein. The 1932 Disarmament Conference. The Nation, 133(??):300, September 1931. ISSN 0027-8378. Reprinted in [Ein54b].

Einstein:1931:AZS

[Ein31c] Albert Einstein. About Zionism: speeches and letters. Macmillan Publishing Company, New York, NY, USA, 1931. 94 (est.) pp. LCCN DS149.E5 1931. Translated and edited with an introduction by Leon Simon.

Einstein:1931:CRO

[Ein31d] Albert Einstein. Cosmic religion: with other opinions and aphorisms. Covici-Friede, New York, NY, USA, 1931. 109 pp. LCCN ????

Einstein:1931:F

[Ein31e] Albert Einstein. Foreword. In Sir Isaac Newton, editor, Opticks, or, A treatise of the reflections, refractions, inflections & colours of light, page ?? McGraw-Hill, New York, NY, USA, 1931. LCCN QC353 .N57. Introduction by Prof. E. T. Whittaker.

Einstein:1931:GEF

[Ein31f] Albert Einstein. Gravitational and electromagnetic fields. Science, 74(1922):438–439, October 30, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Einstein:1931:MID

[Ein31g] Albert Einstein. Maxwell’s influence on the development of the conception of physical reality. In Thomson [Tho31], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Einstein:1931:MIE

[Ein31h] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In J. J. (Joseph John) Thomson, editor, James Clerk Maxwell: A Commemoration Volume, 1831–1931, page ?? Cambridge University Press, Cambridge, UK, 1931. LCCN QC16.M4 J3. English translation in [Ein82c, pp. 266– 270].

Einstein:1931:MP

[Ein31i] Albert Einstein. Militant pacifism. World Tomorrow, 14(??):9, 1931. Reprinted in [Ein54b] as “Active Pacifism”.

Einstein:1931:RSGa

[Ein31j] Albert Einstein. Relativity: the Special and the General Theory. Peter Smith, New York, NY, USA, 1931. ???? pp. Authorized translation by Robert W. Lawson.

Einstein:1931:RSGb

[Ein31k] Albert Einstein. Relativity: the Special and the General Theory, a Popular Exposition. Methuen and Co. Ltd., London, UK, tenth edition, 1931. xiii + 1 + 138 pp. Authorized translation by Robert W. Lawson.

Einstein:1931:SH

[Ein31l] Albert Einstein. Science and happiness. Science (new series), 73(1893):375–381, April 10, 1931. Speech given at the California Institute of Technology, and reprinted in the New York Times, 22 February, 1931, section 9.

Einstein:1931:DIA

[Ein31m] Albert Einstein. Un discurso inédito de Alberto Einstein. (Spanish) [An unpublished speech of Albert Einstein]. La Vida Literaria, III(30):1–??, April 1931.

Einstein:1931:DIE

[Ein31n] Albert Einstein. Un discurso inédito de Einstein. (Spanish) [An unpublished speech of Einstein]. Davar, 61(??):110–113, ???? 1931.

Einstein:1931:KPA

[Ein31o] Albert Einstein. Zum Kosmologischen Problem der allgemeinen Relativitätstheorie. (German) [On the cosmological problem of the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):235–237, 1931.

Einstein:1932:GSR

[Ein32a] Albert Einstein. Die gegenwärtige Stand der Relativitätstheorie. (German) [The present state of Relativity Theory]. Die Quelle, 82(??):440–442, 1932.

Einstein:1932:IAS

[Ein32b] Albert Einstein. Introduction and address to students of UCLA. In Builders of the Universe, page ?? US Library Association, Inc., Westwood Village, Los Angeles, CA, USA, February 1932. LCCN BD493 .B8.

Einstein:1932:TJV

[Ein32c] Albert Einstein. Is there a Jewish view of life? Opinion, 2(??): 7, September 26, 1932. Reprinted in [Ein54b].

Einstein:1932:AN

[Ein32d] Albert Einstein. To American Negroes. Crisis, 39(??):45, 1932.

Einstein:1932:DBS

[Ein32e] Albert Einstein. Zu Dr. Berliners siebzigsten Geburtstag. (German) [On Dr. Berliner’s seventieth birthday]. Die Naturwissenschaften, 20(??):913, 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Reprinted in [Ein54b].

Einstein:1933:CS

[Ein33a] Albert Einstein. Civilization and science. Speech given at Royal Albert Hall, London. Published in [Ein34d]., October 4, 1933.

Einstein:1933:FAW

[Ein33b] Albert Einstein. The fight against war. In Alfred Lief, editor, The Fight against War, page 64. John Day, New York, NY, USA, 1933. LCCN ????

Einstein:1933:FTR

[Ein33c] Albert Einstein. Les Fondements de la Théorie de la Relativité Générale. (French) [The Foundation of the Theory of General Relativity]. Hermann, Paris, France, 1933. 109 pp. LCCN ???? Translated by M. Solovine.

Einstein:1933:LPA

[Ein33d] Albert Einstein. Letter to the Prussian Academy of Sciences. Science (new series), 77(??):444, April 5, 1933.

Einstein:1933:NOG

[Ein33e] Albert Einstein. Notes on the Origins of the General Theory of Relativity. Number 20 in Glasgow University Publications. Jackson, Glasgow, Scotland, 1933. 11 pp. LCCN ???? The George A. Gibson lecture at Glasgow University, June 20, 1933. Reprinted in [Ein54b].

Einstein:1933:MTP

[Ein33f] Albert Einstein. On the Method of Theoretical Physics. Clarendon Press, Oxford, UK, 1933. 15 pp. The Herbert Spencer Lecture, delivered at Oxford, 10 June 1933. Reprinted in [Ein34n]. New translation by Sonja Bargmann in [Ein54b, 270–276].

Einstein:1933:VM

[Ein33g] Albert Einstein. Victim of misunderstanding. Times (London), ??(??):??, September 16, 1933.

Einstein:1933:WW

[Ein33h] Albert Einstein. Why war? Pamphlet from The International Institute of Intellectual Cooperation, League of Nations, Paris, France., 1933.

Einstein:1933:DAV

[Ein33i] Albert Einstein. Zur deutsch–amerikanischen Verständigung. (German) [On German–American agreement]. California Institute of Technology Bulletin, 43(138):4–8 (German), 9–12 (English), 1933. Reported in New York Times, January 24, 1933.

Einstein:1934:CJV

[Ein34a] Albert Einstein. Comment je vois le monde. (French) [My World View]. Ernst Flammarion, Paris, France, 1934. vi + 258 pp. LCCN ???? French translation by Georges Cros of [Ein34h].

Einstein:1934:EWP

[Ein34b] Albert Einstein. Education and world peace. Progressive Education, 11(??):440, 1934.

Einstein:1934:ES

[Ein34c] Albert Einstein. Essays in Science. Philosophical Library, New York, NY, USA, 1934. xi + 114 pp. LCCN ????

Einstein:1934:EDE

[Ein34d] Albert Einstein. Europe’s danger, Europe’s hope. Friends of Europe Publications, 4:??, 1934. Reprint of speech [Ein33a].

Einstein:1934:I

[Ein34e] Albert Einstein. Introduction. In The World in Modern Science, page ?? Gollancz, London, UK, 1934. LCCN ????

Einstein:1934:GEF

[Ein34f] Albert Einstein. La Géométrie et l’Expérience. (French) [Geometry and Experience]. Gauthier-Villars, Paris, France, 1934. 24 pp. French translation by Maurice Solovine.

Einstein:1934:MWD

[Ein34g] Albert Einstein. Mein Weltbild. (Dutch) [My World View]. Van Holkema & Warendorf, Amsterdam, The Netherlands, 1934. 284 pp. LCCN ???? Dutch translation by T. Landré of [Ein34h].

Einstein:1934:MWG

[Ein34h] Albert Einstein. Mein Weltbild. (German) [My World View]. Querido Verlag, Amsterdam, The Netherlands, 1934. 269 pp. LCCN AC35 .E3.

Einstein:1934:NPE

[Ein34i] Albert Einstein. Nachruf für Paul Ehrenfest. (German) [Obituary for Paul Ehrenfest]. In Almanak van het Leidsche Studencorps, page ?? Doesburg, Leiden, The Netherlands, 1934.

Einstein:1934:NOG

[Ein34j] Albert Einstein. Notes on the origin of the General Theory of Relativity. In Mein Weltbild. (German) [My World View] [Ein34h], page ?? LCCN AC35 .E3. English translation in [Ein34c, pp. 78–84] and [Ein82c, pp. 285–290].

Einstein:1934:MTP

[Ein34k] Albert Einstein. On the method of theoretical physics. Philosophy of Science, 1(2):163–169, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184387.

Einstein:1934:PSE

[Ein34l] Albert Einstein. The problem of space, ether, and the field in physics. In Mein Weltbild. (German) [My World View] [Ein34h], page ?? LCCN AC35 .E3. English translation in [Ein82c, pp. 276–285].

Einstein:1934:SCD

[Ein34m] Albert Einstein, editor. Sinfonia concertante A dur, für Violine und Violoncello mit Orchester [von Johann Christian Bach 1735–1782]. Ernst Eulenburg, Leipzig, Germany, 1934. 38 pp. LCCN ????

Einstein:1934:WSI

[Ein34n] Albert Einstein. The world as I see it. Covici-Friede, New York, NY, USA, 1934. ???? pp. LCCN AC35 .E52. Translated by Alan Harris.

Einstein:1935:AJU

[Ein35a] Albert Einstein. Appeal for Jewish unity. New Palestine, 25(9): 1, 1935.

Einstein:1935:EDE

[Ein35b] Albert Einstein. Elementary derivation of the equivalence of mass and energy. Bulletin of the American Mathematical Society, 41(??):223–230, 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://www.ams.org/

bull/2000-37-01/S0273-0979-99-00805-8/S0273-0979-99-

00805-8.pdf. J. W. Gibbs Lecture to the American Association for the Advancement of Science (AAAS), 28 December 1934. Reprinted in [Ein00b].

Einstein:1935:MR

[Ein35c] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, 1935. ???? pp. LCCN ????

Einstein:1935:MW

[Ein35d] Albert Einstein. Mein Weltbild. Wisdom Library, New York, NY, 1935. xiii + 112 pp. LCCN ???? English translation by Alan Harris of [Ein34h].

Einstein:1935:MOS

[Ein35e] Albert Einstein. Mój obraz świata. (Polish) [My World View]. M. Fruchtman, Warszawa, Poland, 1935. 283 + 4 pp. LCCN ???? Polish translation by S. Łukowski of [Ein34h].

Einstein:1935:PMW

[Ein35f] Albert Einstein. Peace must be waged: Interview by R. M. Bartlett. Survey Graphic, 24(??):384, 1935.

Einstein:1936:PR

[Ein36a] A. Einstein. Physics and reality. Journal of The Franklin Institute, 221(3):349–382, March 1936. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic). Reprinted in [Ein82c, 290–323].

Einstein:1936:PRG

[Ein36b] A. Einstein. Physik und Realität. (German) [Physics and reality]. Journal of The Franklin Institute, 221(3):313–347, March 1936. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic). Reprinted in [Ein54b]. English translation in [Ein36a].

Einstein:1936:DIA

[Ein36c] Albert Einstein. Declaration of intention for Albert Einstein. Web document, January 15, 1936. URL http://hdl.loc.gov/

loc.wdl/dna.2745;   http://www.wdl.org/en/item/2745.

Einstein:1936:FL

[Ein36d] Albert Einstein. Freedom of learning. Science (new series), 83 (??):372–373, 1936.

Einstein:1936:LLA

[Ein36e] Albert Einstein. Lens-like action of a star by deviation of light in the gravitational field. Science, 84(??):506–507, 1936. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See postscript [RSS97].

Einstein:1936:RSG

[Ein36f] Albert Einstein. Relativity: the Special and the General Theory: a popular exposition. Methuen and Co. Ltd., London, UK, eleventh edition, 1936. xiii + 138 pp. Authorized translation by Robert W. Lawson.

Einstein:1936:STC

[Ein36g] Albert Einstein. Some thoughts concerning education. School and Society, 44(??):589–592, 1936. Translated by Linda Arronet. Reprinted in [Ein54b].

Einstein:1938:EFG

[Ein38a] Albert Einstein. Einheitliche Feldtheorie. (German) [Unified field theory]. The story of this never published work is documented in [SS21b]. It is Einstein’s last work on a 5dimensional theory. The authors of [EB38] developed an extension of Kaluza’s 5-dimensional theory. In this paper, Einstein alone, without his assistant Peter Bergmann, redeveloped the results of that paper from three axioms, ignoring Kaluza’s work. Einstein contributed the manuscript to the US Library of Congress archives, but before doing so, his secretary Helen Dukas made a 15-page typescript in which Einstein filled in the equations by hand. The typescript was then sent to Peter Bergmann, but neither Einstein nor Dukas retained a copy. Fortunately, Bergmann made copies of the typescript, which have since been added to the Einstein archives, but the original has been lost., July 6, 1938.

Einstein:1938:ODZ

[Ein38b] Albert Einstein. Our debt to Zionism. New Palestine, 28(2): 2–4, April 2, 1938. Address to National Labor Committee for Palestine on 17 April 1938. Reprinted in [Ein54b].

Einstein:1938:WDT

[Ein38c] Albert Einstein. Why do they hate the Jews? Collier’s Weekly, 102(??):9–10, 38, November 26, 1938. Translated by Ruth Norden. Reprinted in [Ein54b].

Einstein:1939:AEL

[Ein39a] Albert Einstein. Albert Einstein’s letter to President Franklin Delano Roosevelt. World-Wide Web document, August 2,

1939. URL http://hypertextbook.com/eworld/einstein.
shtml;   http://www.anl.gov/Science_and_Technology/
History/Anniversary_Frontiers/aetofdr.html;   http:/
/www.lanl.gov/history/road/pdf/Einstein.pdf. Reprinted

in [Ein63a].

Einstein:1939:EWB

[Ein39b] Albert Einstein. Europe will become a barren waste. New Palestine, 29(??):1–2, March 24, 1939.

Einstein:1939:NRG

[Ein39c] Albert Einstein. Naturwissenschaft und Religion. (German) [Science and religion]. Address delivered at the Princeton Theological Seminary. Reprinted in [Ein52a, pages 25–29]., May 19, 1939.

Einstein:1939:SSS

[Ein39d] Albert Einstein. On a stationary system with spherical symmetry consisting of many gravitating masses. Annals of Mathematics (2), 40(4):922–936, October 1939. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL 2-E;   http://links.jstor.org/sici?sici=
0003-486X(193910)2:40:4<922:OASSWS>2.0.CO.

Einstein:1939:OG

[Ein39e] Albert Einstein. Our goal. Mimeographed address at a conference at the Princeton Theological Seminary, 19 May 1939. Reprinted in [Ein54b, pages 41–44]., 1939.

Einstein:1939:SBS

[Ein39f] Albert Einstein. Sixtieth birthday statement. Science (new series), 89(??):242, 1939.

Einstein:1940:CCF

[Ein40a] Albert Einstein. Considerations concerning the fundamentals of theoretical physics. Science, 91(??):487–492, May 24, 1940. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Address for the Eighth American Scientific Congress, Washington, DC, May 24, 1940. Reprinted in [Ein82c, pp. 323–335] as “The Fundaments of Theoretical Physics”.

Einstein:1940:FS

[Ein40b] Albert Einstein. Freedom and science. In Ruth Nanda Anshen, editor, Freedom: Its Meaning, page ?? Harcourt Brace, New York, NY, USA, 1940. LCCN JC585 .A55. Translated by James Gutmann.

Einstein:1940:MSJ

[Ein40c] Albert Einstein. Meine Stellung zur jüdischen Frage. (German) [My position on the Jewish question]. Aufbau, 6(52):9, 1940.

Einstein:1940:NBN

[Ein40d] Albert Einstein. Neuer Bund der Nationen. (German) [New bond among nations]. Aufbau, 6(26):1–2, December 27, 1940.

Einstein:1940:MLY

[Ein40e] Albert Einstein. Out of My Later Years. Watts & Company, London, UK, 1940. ???? pp. LCCN QC16.E5 A3.

Einstein:1940:SR

[Ein40f] Albert Einstein. Science and religion. Nature, 146(3706):605– 607, November 9, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v146/n3706/abs/146605a0.html.

Einstein:1940:SKO

[Ein40g] Albert Einstein. Soobrazheniia k obosnovaniiu teoreticheskoi fiziki (Russian) [considerations concerning the fundamentals of theoretical physics]. Pod znamenem marksizma, ??(12):106–113, 1940. Russian translation of [Ein40a]. See response [Kol40].

Einstein:1940:WSI

[Ein40h] Albert Einstein. The world as I see it. Thinker’s library. Watts, London, UK, 1940. xv + 112 pp. LCCN ???? Translated by Alan Harris.

Einstein:1941:DNE

[Ein41a] A. Einstein. Demonstration of the non-existence of gravitational fields with a non-vanishing total mass free of singularities. Univ. Nac. Tucumán. Revista A., 2:5–15, 1941.

Einstein:1941:CLS

[Ein41b] Albert Einstein. The common language of science. Advancement of Science, 2(5):109–110, ???? 1941. CODEN ADSCAH.

ISSN 0001-866X. URL http://www.youtube.com/watch?v=

td2SReWxogY. Broadcast recording for Science Conference, London, UK, September 28, 1941.

Einstein:1941:CJ

[Ein41c] Albert Einstein. Credo as a Jew. Universal Jewish Encyclopedia, 4(??):32–33, 1941.

Einstein:1941:NRG

[Ein41d] Albert Einstein. Naturwissenschaft und Religion. (German) [Science and religion]. In Anonymous, editor, Science, Philosophy and Religion: a Symposium, page ?? Conference on Science, Philosophy and Religion in Their Relation to the Democratic Way of Life, Inc., New York, NY, USA, 1941. LCCN ???? Reprinted in [Ein52a, pages 25–29].

Einstein:1941:SRP

[Ein41e] Albert Einstein. Science and religion, part 2. In Conference on Science, Philosophy and Religion in Their Democratic Way of Life. ????, New York, NY, USA, 1941.

Einstein:1940:SSA

[Ein41f] Albert Einstein. Statement on the significance of American citizenship. In Robert Spiers Benjamin, editor, I am an American:

By Famous Naturalized Americans, page ?? Alliance Book, New York, NY, USA, 1941. LCCN E169.1 .B45.

Einstein:1942:CLS

[Ein42] Albert Einstein. The common language of science. Advancement of Science (London), 2(5):109, 1942. Radio address to the British Association for the Advancement of Science, 28 September 1941. Reprinted in [Ein82c, pp. 335–337].

Einstein:1944:BFI

[Ein44a] A. Einstein. Bivector fields. II. Annals of Mathematics (2), 45(1):15–23, January 1944. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL 2-I; http:

//links.jstor.org/sici?sici=0003-486X(194401)2:45:1<
15:BFI>2.0.CO.

Einstein:1944:EI

[Ein44b] Albert Einstein. The ethical imperative. Opinion, 14(??):10, 1944.

Einstein:1944:GS

[Ein44c] Albert Einstein. Gandhi’s statesmanship. In Mahatma Gandhi: Essays and Reflections on His Life and Work, page ?? Allen and Unwin, London, UK, 1944. LCCN ????

Einstein:1944:HBW

[Ein44d] Albert Einstein. On the heroes of the Battle of the Warsaw Ghetto. Bulletin of the Society of Polish Jews (New York), ?? (??):??, 1944.

Einstein:1944:RBR

[Ein44e] Albert Einstein. Remarks on Bertrand Russell’s theory of knowledge. In Paul A. Schilpp, editor, The Philosophy of Bertrand Russell, volume 5 of Library of Living Philosophers, pages 277– 291. Open Court, LaSalle, IL, USA, 1944. ISSN 0075-9139. LCCN B1649.R94.

Einstein:1945:EAB

[Ein45a] Albert Einstein. Einstein on the atomic bomb. Atlantic Monthly, 176:43–45, November 1945. Edited by Raymond Swing. Reprinted as “Atomic War or Peace, Part 1”, in [Ein54b, pages 118–123]. See part 2 in [Ein47a].

Einstein:1945:GRT

[Ein45b] Albert Einstein. A generalization of the relativistic theory of gravitation. Annals of Mathematics (2), 46(4):578–584, October 1945. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980

(electronic). URL 2-1;   http://links.jstor.org/sici?
sici=0003-486X(194510)2:46:4<578:AGOTRT>2.0.CO.

Einstein:1945:MR

[Ein45c] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, 1945. 135 pp.

Einstein:1945:HSE

[Ein45d] Albert Einstein. On harnessing solar energy. New York Times, August 14, 1945. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Einstein:1945:CP

[Ein45e] Albert Einstein. On the cosmological problem. American Scholar, 14(??):137–156, 269, 1945.

Einstein:1945:TPE

[Ein45f] Albert Einstein. A testimonial from Prof. Einstein. In Jacques Hadamard, editor, An Essay on the Psychology of Invention in the Mathematical World, pages 142–143. Princeton University Press, Princeton, NJ, USA, 1945. LCCN ???? Reprinted in [Ein96e].

Einstein:1945:WWP

[Ein45g] Albert Einstein. The war is won, but the peace is not. Address to the Fifth Nobel Anniversary Dinner at the Hotel Astor, New York City., December 10, 1945. URL http://www.

americanrhetoric.com/speeches/alberteinsteinpostwarworld.
htm;   http://www.openculture.com/2013/06/listen_as_
albert_einstein_calls_for_peace_and_social_justice_

in_1945.html. Audio recording, with text reprinted in [Ein50o, Ein54b, RS07f].

Einstein:1946:MUP

[Ein46a] Albert Einstein. E = mc^{2}: The most urgent problem of our time. Science Illustrated, 1:16–17, April 1946. Reprinted in [Ein82c, pp. 337–341].

Einstein:1946:JSG

[Ein46b] Albert Einstein. An die jüdischen Studenten. (German) [To the Jewish students]. Aufbau, 12(1):16, 1946.

Einstein:1946:EDE

[Ein46c] Albert Einstein. Elementary derivation of the equivalence of mass and energy. Technion Opinion, 5(??):16–17, 1946.

Einstein:1946:MR

[Ein46d] Albert Einstein. The Meaning of Relativity. Methuen and Co. Ltd., London, UK, third edition, 1946. v + 129 pp. LCCN ????

Einstein:1946:OTS

[Ein46e] Albert Einstein. Only then shall we find courage. The New York Times Magazine, ??(??):??, June 23, 1946. ISSN 0362-1308. Interview by M. Amrine.

Einstein:1946:SOS

[Ein46f] Albert Einstein. Social obligation of the scientist. In Ben Raeburn, editor, Treasury for the Free World, page ?? Books for Libraries Press, Freeport, NY, USA, 1946. LCCN D445 .R2 1972. Introduction by Ernest Hemingway.

Einstein:1946:U

[Ein46g] Albert Einstein. Udvejen. In Masters and Way [MW46b], page ?? LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Einstein:1946:W

[Ein46h] Albert Einstein. The way out. In Masters and Way [MW46a], page ?? LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Einstein:1947:AWP

[Ein47a] Albert Einstein. Atomic war or peace. Atlantic Monthly, 180 (??):29–33, November 1947. Reprinted as “Atomic War or Peace, Part 2” in [Ein54b]. See part 1 in [Ein45a].

Einstein:1947:MM

[Ein47b] Albert Einstein. The military mentality. American Scholar, 16 (??):353–354, 1947. Reprinted in [Ein54b, pages 132–134].

Einstein:1947:OLU

[Ein47c] Albert Einstein. Open letter to the United Nations General Assembly on how to “break to vicious circle”. United Nations World, 1(??):13–14, October 1947.

Einstein:1947:PSE

[Ein47d] Albert Einstein. The problem of space, ether and the field in physics. In Saxe Commins and R. N. Linscott, editors, Man and the Universe, volume 4 of The Great Thinkers, page ?? Random House, New York, NY, USA, 1947. LCCN ????

Einstein:1947:RRM

[Ein47e] Albert Einstein. Replies to Ridenour: The military mentality. Bulletin of the Atomic Scientists, 3(8):223–224, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1947:REW

[Ein47f] Albert Einstein. Response to the editor on Walter White’s article “Why I remain a Negro”. Saturday Review of Literature, 30(??):21, November 1, 1947.

Einstein:1948:QMW

[Ein48a] A. Einstein. Quanten-Mechanik und Wirklichkeit. (German) [Quantum mechanics and reality]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):320–324, ???? 1948. ISSN 1746-8361.

Einstein:1948:ASR

[Ein48b] Albert Einstein. Atomic science reading list. ’48: Magazine of the Year, pages 60–61, January 1948.

Einstein:1948:F

[Ein48c] Albert Einstein. Foreword. In The Hebrew University of Jerusalem, page ?? Goldberg’s Press, Jerusalem, Israel, 1948. LCCN ????

Einstein:1948:GTG

[Ein48d] Albert Einstein. A generalized theory of gravitation. Reviews of Modern Physics, 20(1):35–39, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/abstract/RMP/v20/p35;
http://prola.aps.org/pdf/RMP/v20/i1/p35_1.

Einstein:1948:LUM

[Ein48e] Albert Einstein. Letter on universal military training addressed to the Chairman of the Senate committee. US Congress, Senate, Committee on Armed Services, Hearings on Universal Military Training, 1948. Read 24 March 1948.

Einstein:1948:LA

[Ein48f] Albert Einstein. Looking ahead. Rotarian, ??(??):8–10, June 1948.

Einstein:1948:MWC

[Ein48g] Albert Einstein. A message to the World Congress of Intellectuals [Breslau, West Germany, August 1948]. Bulletin of the Atomic Scientists, 4(10):295, 299, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano48c, Ano48a].

Einstein:1948:MF

[Ein48h] Albert Einstein. My faith. American Weekly, ??(??):3–4, ???? 1948.

Einstein:1948:ROW

[Ein48i] Albert Einstein. On receiving the One World Award. Address given at Carnegie Hall, April 27, 1948. Reprinted in [Ein54b, pages 146–147]., 1948.

Einstein:1948:PPS

[Ein48j] Albert Einstein. Physics, philosophy, and scientific progress. Journal of the International College of Surgeons, 14(??):755– 758, 1948. From tape-recorded address to a convention of the International College of Surgeons.

Einstein:1948:PIU

[Ein48k] Albert Einstein. A plea for international understanding. Bulletin of the Atomic Scientists, 4(1):1, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1948:RSI

[Ein48l] Albert Einstein. Religion and science: Irreconcilable? Christian Register, 127(??):19–20, June 1948. Reprinted in [Ein54b].

Einstein:1948:RSS

[Ein48m] Albert Einstein. A reply to the Soviet scientists. Bulletin of the Atomic Scientists, 4(2):35–37, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [VFIS48]. Reprinted in [Ein63b].

Einstein:1948:U

[Ein48n] Albert Einstein. [unknown]. AAAS Bulletin, ??(??):??, April 14, 1948.

Einstein:1948:WS

[Ein48o] Albert Einstein. Why socialism? Monthly Review: An Independent Socialist Magazine, 1(1):9–15, May 1948.

Einstein:1949:AN

[Ein49a] Albert Einstein. Autobiographical notes. In Schilpp [Sch49b], pages 3–94. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Einstein:1949:SAB

[Ein49b] Albert Einstein. In the shadow of the atomic bomb. Southern Patriot, 7(3):??, 1949.

Einstein:1949:OLF

[Ein49c] Albert Einstein. Open letter to friends of the Emergency Committee. Bulletin of the Atomic Scientists, 5(6–7):206, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1949:RCE

[Ein49d] Albert Einstein. Remarks concerning the essays brought together in this co-operative volume. In Schilpp [Sch49c], pages 665–688. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Einstein:1949:WSI

[Ein49e] Albert Einstein. The world as I see it. Philosophical Library, New York, NY, USA, 1949. ISBN 0-8065-0711-X. xiii + 112 pp. LCCN AC35 .E528. Translated by Alan Harris.

Einstein:1950:BIG

[Ein50a] A. Einstein. The Bianchi identities in the generalized theory of gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2:120–128, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1950:PPS

[Ein50b] A. Einstein. Physics, philosophy and scientific progress. The Journal of the International College of Surgeons, 14(6):755–758, December 1950. ISSN 0096-557X.

Einstein:1950:SRS

[Ein50c] A. Einstein. Social responsibility in science. Science, 112(2921): 760–761, December 22, 1950. CODEN SCIEAS. ISSN 0036-8075.

Einstein:1950:ACB

[Ein50d] Albert Einstein. Arms can bring no security. Bulletin of the Atomic Scientists, 6(3):71, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1950:CSM

[Ein50e] Albert Einstein. Conceptions: scientifiques, morales, et sociales. (French) [Scientific, moral, and social conceptions]. Flammarion, Paris, France, 1950. 316 pp. LCCN QC16.E5 A314. French translation of Out of my later years.

Einstein:1950:F

[Ein50f] Albert Einstein. Foreword. In Frank [Fra50], page ?? LCCN QC6 F642. With a foreword by Albert Einstein.

Einstein:1950:I

[Ein50g] Albert Einstein. Introduction. In Eliot [Eli50], page ?? LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Einstein:1950:MRa

[Ein50h] Albert Einstein. The Meaning of Relativity. Methuen and Co. Ltd., London, UK, fourth edition, 1950. 145 pp. LCCN ????

Einstein:1950:MRb

[Ein50i] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, third edition, 1950. 150 pp. LCCN QC6. Includes the Generalized Theory of Gravitation.

Einstein:1950:MRT

[Ein50j] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, third edition, 1950. iv + 165 pp. Includes the Generalized Theory of Gravitation.

Einstein:1950:MIS

[Ein50k] Albert Einstein. Message to the Italian Society for the Advancement of Science. Impact, page ??, Fall 1950. Sent to the group’s forty-second meeting in Lucca, Italy. English translation in [Ein82c, pp. 356–359].

Einstein:1950:GTG

[Ein50l] Albert Einstein. On the Generalized Theory of Gravitation. Scientific American, 182(4):13–17, April 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v182/n4/
pdf/scientificamerican0450-13.pdf. Reprinted in [Ein82c,

pp. 341–356].

Einstein:1950:MOS

[Ein50m] Albert Einstein. On the moral obligation of science. Impact, 1 (??):104–105, 1950.

Einstein:1950:OLS

[Ein50n] Albert Einstein. An open letter to the Society for Social Responsibility. Science, 112(??):760–761, 1950. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Reprinted as “The State and the Individual Conscience” in [Ein54b, pp. 26– 27].

Einstein:1950:MLYa

[Ein50o] Albert Einstein. Out of my later years. Philosophical Library, New York, NY, USA, 1950. vii + 282 pp. LCCN QC16.E5 A3.

Einstein:1950:MLYb

[Ein50p] Albert Einstein. Out of my later years. Thames and Hudson, London, UK, 1950. viii + 282 pp. LCCN QC16.E5 A3.

Einstein:1951:AQT

[Ein51a] Albert Einstein. The advent of the quantum theory. Science, 113 (2926):82–84, January 26, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://science.

sciencemag.org/content/113/2926/82.

Einstein:1951:CM

[Ein51b] Albert Einstein. Concern for man. Bulletin of the Atomic Scientists, 7(7–8):204, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1951:Fa

[Ein51c] Albert Einstein. Foreword. In Dagobert D. Runes, editor, Spinoza: Dictionary, page ?? Philosophical Library, New York, NY, USA, 1951. LCCN B3951 .S5.

Einstein:1951:Fb

[Ein51d] Albert Einstein. Foreword. In Homer Smith, editor, Man and His Gods, page ?? Little, Brown and Co., Boston, MA, USA, 1951. LCCN ????

Einstein:1951:GMG

[Ein51e] Albert Einstein. Grösse in der Musik. (German) [Greatness in Musik]. Pan-Verlag, Zürich, Switzerland, 1951. 252 pp.

Einstein:1951:I

[Ein51f] Albert Einstein. Introduction. In Carola Baumgardt, editor, Johannes Kepler: Life and Letters, page ?? Philosophical Library, New York, NY, USA, 1951. LCCN ????

Einstein:1951:PCT

[Ein51g] Albert Einstein. Sur le problème cosmologique. Théorie de la gravitation généralisée. (French) [On the cosmological problem. Generalized theory of gravitation]. Gauthier-Villars, Paris, France, 1951. 54 pp. Translated from English par Maurice Solovine.

Einstein:1952:MSJ

[Ein52a] Albert Einstein. Aus meinen späten Jahren. (German) [Out of my later years]. Deutsche Verlagsanstalt, Stuttgart, West Germany, 1952. 278 pp. LCCN ???? German translation by Hildegard Blomeyer of [Ein50o].

Einstein:1952:CSM

[Ein52b] Albert Einstein. Conceptions scientifiques, morales, et sociales. (French) [Scientific, moral, and social concepts]. Flammarion, Paris, France, 1952. 316 pp. LCCN QC16.E5 A314. French translation of [Ein50p].

Einstein:1952:MOS

[Ein52c] Albert Einstein. On the moral obligation of the scientist. Bulletin of the Atomic Scientists, 8(2):34–35, February 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1952:SCD

[Ein52d] Albert Einstein. Symptoms of cultural decay. Bulletin of the Atomic Scientists, 8(7):166–167, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders. Reprinted in [Ein54b, pp. 166–167].

Einstein:1952:WNZ

[Ein52e] Albert Einstein. Wer nur Zeitungen liest, sieht die Dinge wie ein Kurzsichtiger ohne Augengläser. (German) [Those who read only newspapers see things like a nearsighted person without glasses]. Der Jungkaufmann (Zürich), 27(4):73, 1952. Reprinted in English translation in [Ein54b, pp. 64–65].

Einstein:1952:WDM

[Ein52f] Albert Einstein. The wisdom of dialectical materialism. Eightline poem, reprinted in [Lan08], 1952.

Einstein:1953:A

[Ein53a] A. Einstein. Appendice. In Louis de Broglie, physicien et penseur (French) [Louis de Broglie, physicist and thinker] [Geo53], pages 337–342. LCCN QC71 .L88.

Einstein:1953:CCU

[Ein53b] Albert Einstein. A comment on a criticism of Unified Field Theory. Physical Review (2), 89(1):321, January 1, 1953. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/abstract/PR/v89/p321;
 http://prola.aps.org/pdf/PR/v89/i1/p321_1.

Einstein:1953:EBG

[Ein53c] Albert Einstein. Einleitende Bemerkungen über Grundbegriffe. Remarques préliminaires sur les concepts fondamentaux. (German/French) [preliminary remarks on fundamental concepts]. In Louis de Broglie, physicien et penseur (French) [Louis de Broglie, physicist and thinker] [Geo53], pages 4–15. LCCN QC71 .L88.

Einstein:1953:EUI

[Ein53d] Albert Einstein. Elementare überlegungen zur Interpretation der Grundlagen der Quanten-Mechanik. (German) [Elementary considerations for the interpretation of the foundation of quantum mechanics]. In Appleton et al. [ABdB^{+}53], pages 33–40. LCCN QC71 .S35.

Einstein:1953:F

[Ein53e] Albert Einstein. Foreword. In Drake [Dra53], page ?? LCCN QB41 .G1356. Translation by Stillman Drake of Galileo’s original work. Foreword by Albert Einstein.

Einstein:1953:LRW

[Ein53f] Albert Einstein. Letter in reply to William Frauenglass. Bulletin of the Atomic Scientists, 9(6):229–230, July 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Ein54b, pages 33–34].

Einstein:1982:MPM

[Ein53g] Albert Einstein. Max Planck in memoriam. In Ideas and Opinions [Ein82c], pages 78–79. ISBN 0-517-88440-2. LCCN AC35 .E526 1982x.

Einstein:1953:MRa

[Ein53h] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, fourth edition, 1953. 165 pp. LCCN QC6.

Einstein:1953:MRb

[Ein53i] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, fourth edition, 1953. iv + 168 pp.

Einstein:1953:MWG

[Ein53j] Albert Einstein. Mein Weltbild. (German) [My World View]. Europa Verlag, Zürich, Switzerland, 1953. 257 pp. LCCN ???? Edited by Carl Seelig.

Einstein:1953:MAD

[Ein53k] Albert Einstein. Message on the 410th anniversary of the death of Copernicus. On the occasion of the commemoration evening, Columbia University, New York, NY., December 1953. Reprinted in [Ein82c, pp. 359–360].

Einstein:1953:MAC

[Ein53l] Albert Einstein. Message to the 24th Annual Conference of the War Resister’s League. New York, NY, USA, August 10, 1953.

Einstein:1953:P

[Ein53m] Albert Einstein. Preface. In Galileo Galilei, editor, Dialogue concerning the two chief world systems, Ptolemaic & Copernican, page ?? University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1953. LCCN QB41 .G1356. Translated by Stillman Drake, with foreword by Albert Einstein.

Einstein:1953:SAI

[Ein53n] Albert Einstein. Supplement to Appendix II of “The meaning of Relativity, 4th edition”. Princeton University Press, Princeton, NJ, USA, 1953. 8 pp.

Einstein:1953:JPF

[Ein53o] Albert Einstein. To the Jewish peace fellowship. Tidings, 8(1): 3, 1953. Also in volume 9(2) 5–5.

Einstein:1954:F

[Ein54a] Albert Einstein. Foreword. In Max Jammer, editor, Concepts of Space: the history of theories of space in physics, page ?? Harvard University Press, Cambridge, MA, USA, 1954. LCCN QC6 .J3. See also later editions [Ein69a, Ein93b].

Einstein:1954:IO

[Ein54b] Albert Einstein. Ideas and Opinions. Crown Publishers, New York, NY, USA, 1954. vi + 377 pp. LCCN AC35 .E526 1954.

Einstein:1954:IEW

[Ein54c] Albert Einstein. If Einstein were young again, he says he’d become a plumber. New York Times, ??(??):??, November 10, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. Based on letter written to the editor of The Report Magazine on 13 October 1954.

Einstein:1954:MPM

[Ein54d] Albert Einstein. Max Planck in memoriam. In Ideas and Opinions [Ein54b], pages 78–79. LCCN AC35 .E526 1954.

Einstein:1954:MIE

[Ein54e] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In Ideas and Opinions [Ein54b], pages 266–270. LCCN AC35 .E526 1954.

Einstein:1954:RPS

[Ein54f] Albert Einstein. Relativity and the problem of space. In Relativity: the Special and the General Theory: a popular exposition [Ein54g], pages ix + 165. Portion reprinted in [Ein82c, pp. 360– 377].

Einstein:1954:RSG

[Ein54g] Albert Einstein. Relativity: the Special and the General Theory: a popular exposition. Methuen and Co. Ltd., London, UK, fifteenth edition, 1954. ix + 165 pp. Authorized translation by Robert W. Lawson.

Einstein:1954:TJS

[Ein54h] Albert Einstein. Tribute to Joseph Scharl. New York Times, December 9, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Einstein:1954:SAR

[Ein54i] Albert Einstein. Über die spezielle und allgemeine Relativitätstheorie: gemeinverständlich. (German) [On Special and General Relativity]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1954. viii + 104 pp.

Einstein:1955:ASE

[Ein55a] Albert Einstein. Albert Schweitzer at eighty. Christian Century, 72(2):42, 1955.

Einstein:1955:ESG

[Ein55b] Albert Einstein. Erinnerungen-Souvenirs. (German) [Remembrances]. Schweizerische Hochschulzeitung, 28:143–153, 1955. Special volume.

Einstein:1955:Fa

[Ein55c] Albert Einstein. Foreword. In Physics and Microphysics, page ?? Pantheon Books, New York, NY, USA, 1955. LCCN ????

Einstein:1955:Fb

[Ein55d] Albert Einstein. Foreword. In William Esslinger, editor, Politics and Science, page ?? Philosophical Library, New York, NY, USA, 1955. LCCN JA71 .E7.

Einstein:1955:I

[Ein55e] Albert Einstein. Introduction. In Henry I. Wachtel, editor, Security for all and free enterprise; a summary of the social philosophy of Josef Popper-Lynkeus [pseudonym], page ?? Philosophical Library, New York, NY, USA, 1955. LCCN H59.P55 W3.

Einstein:1955:MFA

[Ein55f] Albert Einstein. Le memorie fondamentali di Albert Einstein. (Italian) [The fundamental memories of Albert Einstein]. In Cinquant’anni di Relatività, 1905–1955. (Italian) [Fiftieth Anniversary of Relativity, 1905–1955], pages 477–611. Editrice Universitaria, Firenze, Italia, 1955.

Einstein:1955:MR

[Ein55g] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, fifth edition, 1955. vi + 169 pp.

Einstein:1955:P

[Ein55h] Albert Einstein. Preface. In Jules Salvador Moch, editor, Human folly: to disarm of perish?, page 222. Gollancz, London, UK, 1955. LCCN UF767 .M613. Translated by Edward Hyams.

Einstein:1955:QCT

[Ein55i] Albert Einstein. Quatre conféences sur la théorie de la relativité faites à l’Université de Princeton. (French) [Four lectures on the Theory of Relativity, given at Princeton University]. Gauthier-Villars, Paris, France, 1955. 99 pp. LCCN ???? French translation by Maurice Solovine.

Einstein:1955:WTI

[Ein55j] Albert Einstein. What is the task of an international journal? Common Cause (Florence, Italy), 1(1):3, 1955.

Einstein:1956:ASG

[Ein56a] Albert Einstein. Autobiographische Skizze. (German) [Autobiographical sketch]. In Seelig [See56b], page ?? ISBN 3-322-84225-8. LCCN ???? With contributions and letters from Hedwig Born.

Einstein:1956:GRG

[Ein56b] Albert Einstein. Grundzüge der Relativitätstheorie. (German) [Essentials of the Theory of Relativity]. ????, ????, sixth edition, 1956. ???? pp. LCCN ????

Einstein:1956:ITB

[Ein56c] Albert Einstein. Investigations on the theory of the Brownian movement. Dover, New York, NY, USA, 1956. ISBN 0-486-60304-0. vi + 122 pp. LCCN QC183 .E513 1956. Edited with notes by R. Fürth, Translated by A. D. Cowper.

Einstein:1956:TRR

[Ein56d] Albert Einstein. La théorie de la relativité restreinte et générale: exposé élémentaire: La relativité et le problème de l’espace. Gauthier-Villars, Paris, France, 1956. LCCN ???? Translated from the German by Maurice Solovine.

Einstein:1956:MRa

[Ein56e] Albert Einstein. The Meaning of Relativity. Chapman and Hall, Ltd., London, UK, 1956. ???? pp. LCCN ????

Einstein:1956:MRb

[Ein56f] Albert Einstein. The Meaning of Relativity. Methuen and Co. Ltd., London, UK, sixth edition, 1956. 161 pp. LCCN ????

Einstein:1956:MRc

[Ein56g] Albert Einstein. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, fifth edition, 1956. 166 pp. LCCN QC6.

Einstein:1956:MLY

[Ein56h] Albert Einstein. Out of my later years. Citadel Press, New York, NY, USA, 1956. 283 pp. LCCN QC16.E5 A3 1956.

Einstein:1956:ECM

[Ein56i] Albert Einstein. Sur l’électrodynamique des corps en mouvement. (French) [On the electrodynamics of moving bodies]. Gauthier-Villars, Paris, France, 1956. 59 pp. French translation by Maurice Solovine.

Einstein:1958:CJV

[Ein58] Albert Einstein. Comment je vois le monde. (French) [How I see the world]. Bibliothèque de Philosophie Scientifique. Flammarion, Paris, France, 1958. 218 pp. LCCN ???? Translated from the German, and revised and augmented by Maurice Solovine.

Einstein:1959:LS

[Ein59a] Albert Einstein. The life of solitude. In Gordon and Sorkin [GS59], page ?? LCCN PN6110.S37 G6.

Einstein:1959:MI

[Ein59b] Albert Einstein. A message to intellectuals. In Gordon and Sorkin [GS59], page ?? LCCN PN6110.S37 G6.

Einstein:1959:PMC

[Ein59c] Albert Einstein. Portrait of Madame Curie. In Gordon and Sorkin [GS59], page ?? LCCN PN6110.S37 G6.

Einstein:1959:SSA

[Ein59d] Albert Einstein. [Schopenhauer, science, and art]. Bulletin of the Atomic Scientists, 15(2):85, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1959:SP

[Ein59e] Albert Einstein. Self-portrait. In Gordon and Sorkin [GS59], page ?? LCCN PN6110.S37 G6.

Einstein:1959:WWM

[Ein59f] Albert Einstein. The world we make. knowledge comes: Science and society. In Gordon and Sorkin [GS59], page ?? LCCN PN6110.S37 G6.

Einstein:1960:EPb

[Ein60a] Albert Einstein. Einstein on peace. Simon and Schuster, New York, NY, USA, 1960. xvi + 704 pp.

Einstein:1960:FTO

[Ein60b] Albert Einstein. Field theories, old and new. Readex Microprint, New York, NY, USA, 1960. ???? pp. LCCN QC 173.7 EIN.

Einstein:1960:RED

[Ein60c] Albert Einstein. Relatività: Esposizione divulgativa. (Italian) [Relativity: Layman’s explanation], volume 40. Editore Boringhieri, Torino, Italia, 1960. 192 pp.

Einstein:1960:RSG

[Ein60d] Albert Einstein. Relativity: the special and the general theory: a popular exposition. University paperbacks. Methuen and Co. Ltd., London, UK, 15th edition, 1960. ix + 165 pp. LCCN ???? Authorised translation by Robert W. Lawson.

Einstein:1961:RSGa

[Ein61a] Albert Einstein. Relativity: the Special and the General Theory. Three Rivers Press, New York, NY, USA, 1961. xiii + 188 pp. LCCN ???? Authorized translation by Robert W. Lawson.

Einstein:1961:RSGb

[Ein61b] Albert Einstein. Relativity: the Special and the General Theory. Crown Trade Paperbacks, New York, NY, USA, second edition, 1961. ISBN 0-517-88441-0. xiii + 188 pp. LCCN QC173.55 .E384513 1961. Authorized translation by Robert W. Lawson.

Einstein:1963:EL

[Ein63a] Albert Einstein. The Einstein letter. In Grodzins and Rabinowitch [GR63], pages 11–12. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Einstein:1963:ER

[Ein63b] Albert Einstein. Einstein replies. In Grodzins and Rabinowitch [GR63], pages 130–134. LCCN D842 .B78. Reply to [VFIS63]. Reprint of [Ein48m].

Einstein:1964:GHD

[Eĭn64a] A. Eĭnšteĭn. On Galileo and his Dialogue. Voprosy Istor. Estest. i Tehn. No., 16:29–33, 1964.

Einstein:1964:QCT

[Ein64b] Albert Einstein. Quatre conféences sur la théorie de la relativité faites à l’Université de Princeton. (French) [Four lectures on the Theory of Relativity, given at Princeton University]. Gauthier-Villars, Paris, France, 1964. ii + 98 pp. LCCN ???? French translation by Maurice Solovine.

Einstein:1965:MR

[Ein65] Albert Einstein. The Meaning of Relativity. Oxford Book, Inc., Calcutta, India, 1965. 161 pp. LCCN ????

Einstein:1967:MRa

[Ein67a] Albert Einstein. The Meaning of Relativity. Chapman and Hall, Ltd., London, UK, sixth edition, 1967. xiv + 161 pp. LCCN QC173.55.

Einstein:1967:QTR

[Ein67b] Albert Einstein. On the quantum theory of radiation. In van der Waerden [vdW67], pages 63–77. LCCN QC174.12 S655.

Einstein:1969:F

[Ein69a] Albert Einstein. Foreword. In Max Jammer, editor, Concepts of Space: the history of theories of space in physics, page ?? Harvard University Press, Cambridge, MA, USA, second edition, 1969. LCCN QC6 .J3. See other editions [Ein54a, Ein93b].

Einstein:1969:MR

[Ein69b] Albert Einstein. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, fifth edition, 1969. 166 pp. LCCN QC173.55.

Einstein:1970:MLY

[Ein70a] Albert Einstein. Out of my later years. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1970. ISBN 0-8371-2086-1. 283 pp. LCCN QC16.E5 A3 1970.

Einstein:1970:RCE

[Ein70b] Albert Einstein. Remarks concerning the essays brought together in this co-operative volume. In Schilpp [Sch70], pages 665–688. LCCN ????

Einstein:1970:SD

[Ein70c] Albert Einstein. Science and dictatorship. In Otto Forst de Battaglia, editor, Dictatorship on its trial, by eminent leaders of modern thought, Essay index reprint series, page ?? Books for Libraries Press, Freeport, NY, USA, 1970. ISBN 0-8369-1607-7. LCCN D107 .F65 1970. Translated by Huntley Paterson. With an introduction by Winston S. Churchill.

Einstein:1971:MLY

[Ein71] Albert Einstein. Out of my later years. Bonanza Books, New York, NY, USA, 1971. viii + 282 pp. LCCN QC16.E5 A3 1989.

Einstein:1973:IO

[Ein73] Albert Einstein. Ideas and Opinions. Souvenir Press, London, UK, 1973. ISBN 0-285-64724-5. 378 pp. LCCN AC35 .E48 1973. Based on Mein Weltbild, edited by Carl Seelig, and other sources. New translations and revisions by Sonja Bargmann.

Einstein:1974:MR

[Ein74] Albert Einstein. The Meaning of Relativity. Princeton University Press, Princeton, NJ, USA, fifth edition, 1974. ISBN 0-691-02352-2. 166 pp. LCCN QC6 .E5 1979.

Einstein:1975:PSH

[Ein75] Albert Einstein. 12. the problem of the specific heats. In Mehra [Meh75c], pages 61–72. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Einstein:1976:TRR

[Ein76a] Albert Einstein. La théorie de la relativité restreinte et générale. (French) [The Special and General Theory of Relativity]. Gauthier-Villars, Paris, France, 1976. ISBN 2-04-002566-9. vi + 177 pp. Exposé élémentaire, La relativité et le problème de l’espace, Translated from the German by Maurice Solovine, New printing, Collection “Discours de la Méthode”.

Einstein:1976:QCT

[Ein76b] Albert Einstein. Quatre conférences sur la théorie de la relativité faites à l’Université de Princeton. (French) [Four conferences on the Theory of Relativity held at Princeton University]. Gauthier-Villars, Paris, France, 1976. ISBN 2-04-002618-5. ii + 96 pp. Translated from the German by Maurice Solovine, Reprinting of the 1925 original, Collection “Discours de la Méthode”.

Einstein:1977:WSG

[Ein77a] Albert Einstein. Prologue and epilogue: a Socratic dialogue. In Max Planck, editor, Where is Science Going?, page 221. AMS Press, New York, NY, USA, 1977. ISBN 0-404-14696-1. LCCN Q175 .P57 1977.

Einstein:1977:RSG

[Ein77b] Albert Einstein. Relativity: the Special and the General Theory: a popular exposition, volume 10 of University paperbacks. Methuen and Co. Ltd., London, UK, 1977. ISBN 0-416-67600-6. x + 165 pp. LCCN ???? Authorized translation by Robert W. Lawson. Seventh reprinting.

Einstein:1978:RSG

[Ein78] Albert Einstein. Relativity: the Special and the General Theory: a popular exposition. Wings Books, New York, NY, USA, 1978. ISBN 0-517-02530-2 (paperback), 0-517-02961-8 (hardcover). xiii + 188 pp. LCCN ???? Authorized translation by Robert W. Lawson.

Einstein:1979:LP

[Ein79a] A. Einstein. Paul Langevin. Pensée, ??(205):74–75, ???? 1979.

ISSN 0031-4773.

Einstein:1979:SD

[Ein79b] A. Einstein. Scientists’ dilemma. UNESCO Courier, ??(5):31– 32, ???? 1979. ISSN 0041-5278.

Einstein:1979:AWW

[Ein79c] Albert Einstein. After World War II. Bulletin of the Atomic Scientists, 35(3):35, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Einstein:1979:AN

[Ein79d] Albert Einstein. Autobiographical notes. Open Court Publishing Co., La Salle, IL, USA, centennial edition, 1979. ISBN 0-87548-352-6. v + 89 + 1 pp. Translated from the German and edited by Paul Arthur Schilpp.

Einstein:1979:LWP

[Ein79e] Albert Einstein. Letter to W. Pauli, 19 September 1938, AE 19-175. In Hermann et al. [HvMW79], page ?? ISBN 0-387-08962-4, 3-540-08962-4. LCCN QC16.P37 A34; QC16.P37

W64. URL http://www.springer.com/physics/book/978-3-

540-08962-9. Edited by Armin Hermann, K. von Meyenn and Victor F. Weisskopf, With a foreword by Victor F. Weisskopf, With an introduction by Armin Hermann.

Einstein:1979:MOS

[Ein79f] Albert Einstein. On the moral obligation of the scientist. Bulletin of the Atomic Scientists, 35(3):1, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Einstein:1979:RPG

[Ein79g] Albert Einstein. Zum Relativitäts-Problem. (German) [On the Relativity problem]. Scientia (Milan), 114(1-4):5–28, 1979. CO-DEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic). With a French translation by M. E. Philippi.

Einstein:1980:GMG

[Ein80] Albert Einstein. Größe in der Musik. (German) [Greatness in Music]. Deutscher Taschenbuch Verlag, München, West Germany, 1980. ISBN 3-423-01609-4, 3-7618-1609-X. 173 pp. LCCN ????

Einstein:1981:MR

[Ein81] Albert Einstein. The Meaning of Relativity. Franklin Library, Franklin Center, PA, USA, fifth limited edition, 1981. 208 pp. LCCN ????

Einstein:1982:A

[Ein82a] Albert Einstein. Appreciation. In A Dynamical Theory of the Electromagnetic Field [Cle82], page ?? ISBN 0-7073-0324-9 (paperback). LCCN QC670 .M38 1982. URL https:

//hdl.handle.net/2027/mdp.39015017148431;   https://
hdl.handle.net/2027/mdp.39076002001753. With an appre-

ciation by Albert Einstein. Edited and introduced by Thomas F. Torrance.

Einstein:1982:HCT

[Ein82b] Albert Einstein. How I created the theory of relativity. Physics Today, 35(8):45–47, August 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v35/i8/p45_s1. In-

troduction and translation by Yoshimasa A. Ono of a lecture given by Einstein in Kyoto, Japan on 14 December 1922. It was first published in the Japanese periodical Kaizo in 1923, and more than a half-century later, in a partial English translation in [Oga79]. Reprinted in [WP85, 243–245].

Einstein:1982:IO

[Ein82c] Albert Einstein. Ideas and Opinions. Three Rivers Press, New York, NY, USA, 1982. ISBN 0-517-88440-2. 377 pp. LCCN AC35 .E526 1982x.

Einstein:1982:MIE

[Ein82d] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In Ideas and Opinions [Ein82c], pages 266–270. ISBN 0-517-88440-2. LCCN AC35 .E526 1982x.

Einstein:1982:SR

[Ein82e] Albert Einstein. Science and religion. In Faculty of Lynchburg College [Fac82], page ?? ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). LCCN Q125 .M316 1982.

Einstein:1982:SGT

[Ein82f] Albert Einstein. The special and general theory of relativity. In Faculty of Lynchburg College [Fac82], page ?? ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). LCCN Q125 .M316 1982.

Einstein:1983:SR

[Ein83] Albert Einstein. Sidelights on Relativity. Dover, New York, NY, USA, 1983. ISBN 0-486-24511-X (paperback). 56 pp. LCCN QC173.58 .E5513 1983. US$2.25. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/description/dover032/83005275.html.

Einstein:1984:BN

[Ein84a] Albert Einstein. Biographical note. In Berger [Ber84b], pages ix–x. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

CHP:CHP4760040103

[Ein84b] Albert Einstein. Letter to a friend of peace. Möbius: A Journal for Continuing Education Professionals in Health Sciences, 4(1): 5, 1984. ISSN 0894-1912 (print), 1554-558X (electronic). Check: is the author of this letter the subject of this bibliography?? Is this letter is a reprint of work prior to his death in 1955??

Einstein:1984:MR

[Ein84c] Albert Einstein. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, sixth edition, 1984. ISBN ???? 166 pp. LCCN ????

Einshtein:1985:HTR

[Eĭn85a] A. Eĭnshteĭn. How the Theory of Relativity came about. In Einstein collection, 1980–1981, pages 5–9, 334. Nauka, Moscow, USSR, 1985.

Einshtein:1985:MDS

[Eĭn85b] A. Eĭnshteĭn. A method of determining statistical values of observations concerning quantities subject to irregular fluctuations. Problemy Peredachi Informatsii, 21(4):99–100, 1985. CO-DEN PPDIA5. ISSN 0555-2923. Translated from the French by B. E. Yavelov.

Einstein:1985:WWL

[Ein85c] A. Einstein. Why war — a letter from Albert Einstein to Sigmund Freud. UNESCO Courier, ??(5):5, ???? 1985. ISSN 0041-5278.

Einstein:1985:HCT

[Ein85d] Albert Einstein. How I created the theory of relativity. In Weart and Phillips [WP85], pages 243–245. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Ein82b].

Einstein:1985:LWP

[Ein85e] Albert Einstein. Letter to W. Pauli, 19 September 1938, AE 19-175. In Pauli et al. [PHvMW01b], page ?? ISBN 0-387-08962-4 (vol. 1). LCCN QC16.P37 A34.

Einstein:1985:STE

[Ein85f] Albert Einstein. Sobre la teoría especial y la teoría general de la relatividad: El significado de la relatividad, volume 2 of Obras maestras del pensamiento contemporáneo. Planeta-Agostini, Barcelona, Spain, 1985. ISBN 84-395-0002-5. 237 pp. LCCN ????

Einshtein:1986:MDSa

[Eĭn86a] A. Eĭnshteĭn. A method for the determination of the statistical values of observations related to quantities subjected to irregular fluctuations. In Einstein collection, 1982–1983 (Russian), pages 15–16. Nauka, Moscow, USSR, 1986. Translated from the French by B. E. Yavelov.

Einshtein:1986:STR

[Eĭn86b] A. Eĭnshteĭn. The Special Theory of Relativity. In Einstein collection, 1982–1983 (Russian), pages 7–14. Nauka, Moscow, USSR, 1986. Translated from the Japanese by L. Z. Ponizovskiĭ.

Einstein:1986:MDS

[Ein86c] A. Einstein. Method for the determination of statistical values of observations concerning sizes submitted to irregular fluctuations (reprinted from Arch. Sci. Phys. et Natur, vol 37, pg 254–2568 1914). Izvestiya Akademii Nauk SSSR Fizika Atmosfery i Okeana, 22(1):99–100, January 1986. ISSN 0002-3515.

Einstein:1986:CWF

[Ein86d] Albert Einstein. Comments on the work of A. Friedmann. In Bernstein and Feinberg [BF86], pages 66–67. ISBN 0-231-06376-8. LCCN QB985 .C66 1986. English translation of [Ein22c, Ein23f].

Einstein:1986:CCG

[Ein86e] Albert Einstein. Cosmological considerations on the General Theory of Relativity. In Bernstein and Feinberg [BF86], pages 16–26. ISBN 0-231-06376-8. LCCN QB985 .C66 1986. English translation of [Ein17e].

Einstein:1987:NMI

[Ein87a] A. Einstein. Newton mechanics and its influence on theoretical physics. Revista de Occidente, ??(68):67–76, January 1987. ISSN 0034-8635.

Einstein:1987:MDS

[Ein87b] Albert Einstein. Method for determining the statistical values of observations relating to quantities subjected to irregular fluctuations. IEEE ASSP Magazine, 4(4):6–11, October 1987. ISSN 0740-7467 (print), 1558-1284 (electronic). English translation of [Ein14j].

Einstein:1988:MR

[Ein88a] Albert Einstein. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, fifth edition, 1988. ISBN 0-691-08007-0, 0-691-02352-2 (paperback). 166 pp. LCCN QC6 .E5 1979; QC173.55.

Einstein:1988:SAR

[Ein88b] Albert Einstein. Über die spezielle und die allgemeine Relativitätstheorie. (German) [On the Special and General Theory of Relativity], volume 26 of Facetten der Physik [Facets of Physics]. Friedrich Vieweg und Sohn, Braunschweig, Germany, third edition, 1988. ISBN 3-528-16059-4. x + 112 pp.

Einstein:1989:TC

[Ein89a] A. Einstein. Tributes to contemporaries. Sinn und Form, 41(2): 232–240, March/April 1989. ISSN 0037-5756.

Einstein:1989:WS

[Ein89b] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 40(8):14–21, January 1989. ISSN 0027-0520.

Einstein:1989:CJV

[Ein89c] Albert Einstein. Comment je vois le monde. (French) [How I see the world], volume 183 of Champs. Flammarion, Paris, France, 1989. ISBN 2-08-081183-5. 189 pp. LCCN QC71 .E3614 1989. Translation from German by Régis Hanrion.

Einstein:1989:MLY

[Ein89d] Albert Einstein. Out of my later years. Bonanza Books, New York, NY, USA, 1989. ISBN 0-517-69417-4. 282 pp. LCCN QC16.E5 A3 1989.

Einstein:1990:IFC

[Ein90] Albert Einstein. [inconnu]. (French) [Contributions to quantum theory]. Annales de la Fondation Louis de Broglie, 15 (??):121–129, ???? 1990. ISSN 0182-4295 (print), 2108-6397

(electronic). URL http://aflb.ensmp.fr/AFLB-Web/fldb-

annales-index.htm. TO DO: Find original French title; online references to this paper omit the title, and the journal does not appear to have online archives before 1999??

Einstein:1991:EMC

[Ein91a] Albert Einstein. E = mc^{2}. In Ferris and Fadiman [FF91], pages 56–59. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1991:AEF

[Ein91b] Albert Einstein. Albert Einstein to Franklin D. Roosevelt, August 2, 1939. In Cantelon et al. [CHW91], pages 9–10. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). LCCN UA23 .A597 1991.

Einstein:1991:AN

[Ein91c] Albert Einstein. Autobiographical notes. In Ferris and Fadiman [FF91], pages 577–589. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1991:LMB

[Ein91d] Albert Einstein. From letters to Max Born. In Ferris and Fadiman [FF91], pages 808–809. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1991:SR

[Ein91e] Albert Einstein. Science and religion. In Ferris and Fadiman [FF91], pages 828–836. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1991:TMP

[Ein91f] Albert Einstein. Tribute to Max Planck. In Ferris and Fadiman [FF91], page 590. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1991:HAE

[Ein91g] Elizabeth Roboz Einstein. Hans Albert Einstein: Reminiscences of His Life and Our Life Together. Iowa Institute of Hydraulic Research, University of Iowa, Iowa City, IA, USA, 1991. ISBN 0-87414-083-8. xiii + 112 pp. LCCN TC140.E38 E38 1991. With a foreword by Robert W. Clark.

Einstein:1992:WS

[Ein92] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 44(1):1–8, May 1992. ISSN 0027-0520.

Einstein:1993:WSR

[Ein93a] A. Einstein. Why socialism? Pensée, ??(294–95):113–117, July/ October 1993. ISSN 0031-4773. Reprinted from Monthly Review, Vol 1, May 1949.

Einstein:1993:F

[Ein93b] Albert Einstein. Foreword. In Max Jammer, editor, Concepts of Space: the history of theories of space in physics, page ?? Dover, New York, NY, USA, third edition, 1993. ISBN 0-486-27119-6. LCCN QC6 .J3. See also earlier editions [Ein54a, Ein69a].

Einstein:1993:MLY

[Ein93c] Albert Einstein. Out of my later years: the scientist, philosopher, and man portrayed through his own words. Wings Books, New York, NY, USA, 1993. ISBN 0-517-09380-4. 282 pp. LCCN QC16.E5 A3 1993. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/description/random0414/93024932.html.

Einstein:1994:WS

[Ein94a] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 46(1):46–53, May 1994. ISSN 0027-0520.

Einstein:1994:CCM

[Ein94b] Albert Einstein. Correspondencia con Michèle Besso (1903– 1955). (Spanish) [Correspondence with Michèle Besso (1903– 1955)], volume 36 of Metatemas. Tusquets Editores S. A., Barcelona, Spain, 1994. ISBN 84-7223-771-0. 476 pp. Edited and with an introduction and notes by Pierre Speziali. Translated from the French by Manuel Puigcerver.

Einstein:1994:IO

[Ein94c] Albert Einstein. Ideas and Opinions. Modern Library, New York, NY, USA, 1994. ISBN 0-679-60105-8. xix + 418 pp. LCCN QC71 .E54 1994. With an introduction by Alan Lightman. New translations and revisions by Sonja Bargmann.

Einstein:1994:MR

[Ein94d] Albert Einstein. The Meaning of Relativity. The Stafford Little lectures. Easton Press, Norwalk, CT, USA, collector’s edition, 1994. ISBN ???? xii + 166 pp. LCCN QC6.

Einstein:1994:ECM

[Ein94e] Albert Einstein. Sur l’électrodynamique des corps en mouvement. (French) [On the electrodynamics of moving bodies]. Les Grands Classiques Gauthier-Villars. [Gauthier-Villars Great Classics]. éditions Jacques Gabay, Sceaux, France, 1994. ISBN 2-87647-155-8. ii + 146 pp. Reprint of the 1925 translation, Also including reprints of the French translations of Ether and Relativity (1953), Geometry and experience (1953), Four lectures on the Theory of Relativity (1925), On the cosmological problem (1960), The Unified Field Theory (1960) and Generalized gravitation theory (1950).

Einstein:1995:LEM

[Ein95a] A. Einstein. Letter to Mileva Einstein, 1901: Wife of Albert Einstein. Du — die Zeitschrift der Kultur, ??(10):14–15, October 1995. ISSN 0012-6837.

Einstein:1995:RTS

[Ein95b] A. Einstein. RBRVS 2.0: toward a special theory of resourcebased relativity. Connecticut medicine, 59(1):21–26, January 1995. ISSN 0010-6178.

Einstein:1995:IO

[Ein95c] Albert Einstein. Ideas and Opinions. Crown Trade Paperbacks, New York, NY, USA, third edition, 1995. ISBN 0-517-88440-2. 377 pp. LCCN AC30 .E48 1982. Based on Mein Weltbild, edited by Carl Seelig, and other sources. New translations and revisions by Sonja Bargmann.

Einstein:1995:RSG

[Ein95d] Albert Einstein. Relativity: the Special and the General Theory. Crown Trade Paperbacks, New York, NY, USA, second edition, 1995. ISBN 0-517-88441-0. xiii + 188 pp. LCCN ???? Authorized translation by Robert W. Lawson.

Einstein:1996:EMS

[Ein96a] Albert Einstein. Einstein’s 1912 manuscript on the Special Theory of Relativity: a facsimile. George Braziller, in association with the Jacob E. Safra Philanthropic Foundation and the Israel Museum, New York, NY, USA, 1996. ISBN 0-8076-1417-3, 0-8076-1532-3. 192 pp. LCCN QC173.65 .E3613 1996.

Einstein:1996:MR

[Ein96b] Albert Einstein. The Meaning of Relativity. MJF Books, New York, NY, USA, fifth edition, 1996. ISBN 1-56731-136-9. 166 pp. LCCN QC6.

Einstein:1996:PR

[Ein96c] Albert Einstein. Principles of research. In Schirmacher [Sch96a], page ?? ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996.

Einstein:1996:RSG

[Ein96d] Albert Einstein. Relativity: the Special and the General Theory. Crown Publishers, New York, NY, USA, second edition, 1996. ISBN 0-517-88441-0. xiii + 188 pp. Authorized translation by Robert W. Lawson.

Einstein:1996:MMP

[Ein96e] Albert Einstein. A testimonial from Prof. Einstein. In Jacques Hadamard, editor, The mathematician’s mind: the psychology of invention in the mathematical field. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02931-8 (paperback).

LCCN QA8.4 .H3 1996. URL http://www.loc.gov/catdir/
description/prin021/96027544.html;   http://www.loc.
gov/catdir/toc/prin031/96027544.html.

Einstein:1996:WTR

[Ein96f] Albert Einstein. What is the theory of relativity? In Schirmacher [Sch96a], page ?? ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996.

Einstein:1997:AEL

[Ein97] Albert Einstein. Albert Einstein’s letters to President Franklin Delano Roosevelt. World-Wide Web document, 1997.

URL http://hypertextbook.com/eworld/einstein.shtml;
 http://www.anl.gov/Science_and_Technology/History/
Anniversary_Frontiers/aetofdr.html.

Einstein:1998:WSR

[Ein98] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 50(1):1–7, ???? 1998. ISSN 0027-0520. Reprinted from Monthly Review, vol 1, 1938.

Einstein:1999:VSR

[Ein99] Albert Einstein. Vorlesungen Über Spezielle Relativitätstheorie. (German) [Lectures on the Theory of Special Relativity]. Nuncius, 14(2):650–661, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:

//booksandjournals.brillonline.com/content/10.1163/
182539199x00102.

Einstein:2000:WS

[Ein00a] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 52(1):36–43, May 2000. ISSN 0027-0520.

Einstein:2000:EDE

[Ein00b] Albert Einstein. Elementary derivation of the equivalence of mass and energy. Bull. Amer. Math. Soc., New Ser., 37(1): 39–44, 2000. CODEN BAMOAD. ISSN 0273-0979. URL http:

//www.ams.org/bull/2000-37-01/S0273-0979-99-00805-8/
S0273-0979-99-00805-8.pdf. Reprint of [Ein35b].

Einstein:2000:RSG

[Ein00c] Albert Einstein. Relativity: the Special and the General Theory. Routledge classics. Henry Holt and Company, Incorporated, New York, NY, USA, ???? edition, 2000. ISBN 1-58734-092-

5. xiii + 168 pp. LCCN QC173.55. Authorized translation by Robert W. Lawson.

Einstein:2000:WSI

[Ein00d] Albert Einstein. The world as I see it. Citadel Press, New York, NY, USA, 2000. ISBN 0-8065-0711-X. xiii + 112 pp. LCCN AC35 .E35 2000. US$10.95.

Einstein:2001:F

[Ein01a] Albert Einstein. Foreword. In Drake [Dra01], pages xxiii– xxx. ISBN 0-375-75766-X (paperback). LCCN QB41 .G129413

2001. URL http://www.loc.gov/catdir/bios/random053/
2001030842.html;   http://www.loc.gov/catdir/description/
random047/2001030842.html;   http://www.loc.gov/catdir/
samples/random043/2001030842.html. Translation by Still-

man Drake of Galileo’s original work. Foreword by Albert Einstein. Introduction by J. L. Heilbron.

Einstein:2001:RSG

[Ein01b] Albert Einstein. Relativity: the Special and the General Theory. Routledge classics. Routledge, London, UK, fifteenth edition, 2001. ISBN 0-415-25538-4 (hardcover), 0-415-25384-5 (paperback). x + 166 pp. LCCN QC173.55. URL http:// www.bartleby.com/173/. Authorized translation by Robert W. Lawson.

Einstein:2001:ST

[Ein01c] Albert Einstein. Space–time. In Hoiberg [Hoi01], pages 290–296. ISBN 81-88237-00-0. LCCN AS911.N9 A15 2001.

Einstein:2002:WS

[Ein02] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 54(1):56–62, May 2002. ISSN 0027-0520.

Einstein:2003:LHE

[Ein03a] Albert Einstein. Liebe Hanne! Einstein’s poems to Johanna Fantova. Princeton University Library Chronicle, 65(1):83–88, Autumn 2003. ISSN 0032-8456. URL https://www.jstor.

org/stable/10.25290/prinunivlibrchro.65.1.0083. Trans-

lated by Alfred Engel.

Einstein:2003:MR

[Ein03b] Albert Einstein. The Meaning of Relativity. Routledge, London, UK, sixth edition, 2003. ISBN 0-415-28588-7, 0-203-44953-3 (ebook). 177 pp. LCCN QC173.55.

Einstein:2003:PR

[Ein03c] Albert Einstein. Physics & reality. Dædalus, 132(4):22–25, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.org/stable/20027877.

Project:2003:EAO

[Ein03d] Einstein Papers Project. Einstein Archives online. World-Wide Web document, May 19, 2003. URL http://www.

alberteinstein.info/.

Einstein:2004:WS

[Ein04] A. Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 56(1):44–50, May 2004. ISSN 0027-0520.

Einstein:2005:AEMa

[Ein05a] A. Einstein. Albert Einstein to Moritz Schlick. Physics Today, 58(12):17, December 2005. CODEN PHTOAD. ISSN 0031-9228.

Einstein:2005:AEP

[Ein05b] A. Einstein. Albert Einstein to Paul Ehrenfest. Physics Today, 58(4):88, April 2005. CODEN PHTOAD. ISSN 0031-9228.

Einstein:2005:AAM

[Ein05c] A. Einstein. Antwort auf eine Abhandlung M.v. Laues ,,Ein Satz der Wahrscheinlichkeitsrechnung und seine Anwendung auf die Strahlungstheorie“ [AdP 47, 879 (1915)]. (German) [Answer to a paper M. v. Laue: A Theorem of Probability Theory and Its Application to Radiation Theory [AdP 47, 879 (1915)]]. Annalen der Physik (Berlin), 14(1):509–516, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590043.

Einstein:2005:ABJ

[Ein05d] A. Einstein. Antwort auf eine Bemerkung von J. Stark: ,,Über eine Anwendung des Planckschen Elementargesetzes … “ [AdP 38, 888 (1912)]. (German) [Reply to a remark by J. Stark: On an application of Planck’s elementary law … [AdP 38, 888 (1912)]]. Annalen der Physik (Berlin), 14(1):480, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/andp.200590038.

Einstein:2005:BAV

[Ein05e] A. Einstein. Bemerkung zu Abrahams vorangehender Auseinandersetzung ,,Nochmals Relativität und Gravitation“ [AdP 39, 704 (1912)]. (German) [Comment on Abraham’s previous argument Relativity and Gravity Revisited [AdP 39, 704 (1912)]]. Annalen der Physik (Berlin), 14(1):487–488, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/andp.200590040.

Einstein:2005:BGE

[Ein05f] A. Einstein. Bemerkung zu dem Gesetz von Eötvös [AdP 34, 165 (1911)]. (German) [Comment on the law of Eötvös [AdP 34, 165 (1911)]]. Annalen der Physik (Berlin), 14(1):392–397, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590027.

Einstein:2005:BAD

[Ein05g] A. Einstein. Bemerkung zu der Arbeit von D. Mirimanoff: ,,Über die Grundgleichungen … “ [AdP 28, 885 (1909)]. (German) [Comment on the work of D. Mirimanoff: On the basic equations … [AdP 28, 885 (1909)]]. Annalen der Physik (Berlin), 14(1):343–346, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590023.

Einstein:2005:BFS

[Ein05h] A. Einstein. Bemerkung zu der Franz Seletyschen Arbeit ,,Beiträge zum kosmologischen System“ [AdP 69, 436 (1922)]. (German) [Comment on the work of Franz Selety Contributions to the cosmological system [AdP 69, 436 (1922)]]. Annalen der Physik (Berlin), 14(1):582–585, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590047.

Einstein:2005:BMAb

[Ein05i] A. Einstein. Bemerkung zu meiner Arbeit: ,,Eine Beziehung zwischen dem elastischen Verhalten … “ [AdP 34, 590 (1911)]. (German) [Comment on my work: A relationship between elastic behavior … [AdP 34, 590 (1911)]]. Annalen der Physik (Berlin), 14(1):405, February 2005. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590030.

Einstein:2005:BPH

[Ein05j] A. Einstein. Bemerkungen zu den P. Hertzschen Arbeiten: ,,Über die mechanischen Grundlagen der Thermodynamik“ [AdP 34, 175 (1911)]. (German) [Remarks on P. Hertz’s works: On the Mechanical Foundations of Thermodynamics [AdP 34, 175 (1911)]]. Annalen der Physik (Berlin), 14(1):403–404, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590029.

Einstein:2005:BNH

[Ein05k] A. Einstein. Bemerkungen zu der Notiz von Hrn. Paul Ehrenfest: ,,Die Translation deformierbarer Elektronen und der Flächensatz“ [AdP 23, 206 (1907)]. (German) [Comment to the note by Mr. Paul Ehrenfest: The translation of deformable electrons and the surface set [AdP 23, 206 (1907)]]. Annalen der Physik (Berlin), 14(1):300–302, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590017.

Einstein:2005:BMAa

[Ein05l] A. Einstein. Berichtigung zu meiner Arbeit: ,,Die Plancksche Theorie der Strahlung etc.“ [AdP 22, 800 (1907)]. (German) [Correction to my work: Planck’s theory of radiation, etc. [AdP 22, 800 (1907)]]. Annalen der Physik (Berlin), 14(1):296, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/andp.200590015.

Einstein:2005:BMAc

[Ein05m] A. Einstein. Berichtigung zu meiner Arbeit: ,,Eine neue Bestimmung der Moleküldimensionen“ [AdP 34, 591 (1911)]. (German) [Correction to my work: A New Determination of Molecular Dimensions [AdP 34, 591 (1911)]]. Annalen der Physik (Berlin), 14(1):406–407, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590031.

Einstein:2005:PES

[Ein05n] A. Einstein. Das Prinzip von der Erhaltung der Schwerpunktsbewegung und die Trägheit der Energie [AdP 20, 627 (1906)]. (German) [The principle of conservation of the center of mass motion and the inertia of the energy [AdP 20, 627 (1906)]]. Annalen der Physik (Berlin), 14(1):268–274, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590011.

Einstein:2005:GAR

[Ein05o] A. Einstein. Die Grundlage der allgemeinen Relativitätstheorie [AdP 49, 769 (1916)]. (German) [The basis of general theory of relativity [AdP 49, 769 (1916)]]. Annalen der Physik (Berlin), 14(1):517–571, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590044.

Einstein:2005:PTS

[Ein05p] A. Einstein. Die Plancksche Theorie der Strahlung und die Theorie der spezifischen Wärme [AdP 22, 180 (1907)]. (German) [Planck’s theory of radiation and the theory of specific heat [AdP 22, 180 (1907)]]. Annalen der Physik (Berlin), 14(1):280– 291, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590013.

Einstein:2005:BZE

[Ein05q] A. Einstein. Eine Beziehung zwischen dem elastischen Verhalten und der spezifischen Wärme bei festen Körpern mit einatomigem Molekül [AdP 34, 170 (1911)]. (German) [A relationship between the elastic behavior and the specific heat of solid bodies with a monatomic molecule [AdP 34, 170 (1911)]]. Annalen der Physik (Berlin), 14(1):398–402, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/andp.200590028.

Einstein:2005:NBM

[Ein05r] A. Einstein. Eine neue Bestimmung der Moleküldimensionen [AdP 19, 289 (1906)]. (German) [A new determination of molecular dimensions [AdP 19, 289 (1906)]]. Annalen der Physik (Berlin), 14(1):229–247, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590008.

Einstein:2005:TGT

[Ein05s] A. Einstein. Eine Theorie der Grundlagen der Thermodynamik [AdP 11, 170 (1903)]. (German) [A theory of the foundations of thermodynamics [AdP 11, 170 (1903)]]. Annalen der Physik (Berlin), 14(1):135–153, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590002.

Einstein:2005:EBT

[Ein05t] A. Einstein. Elementare Betrachtungen über die thermische Molekularbewegung in festen Körpern [AdP 35, 679 (1911)]. (German) [Elementary considerations on thermal molecular motion in solid bodies [AdP 35, 679 (1911)]]. Annalen der Physik (Berlin), 14(1):408–424, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590032.

Einstein:2005:FCA

[Ein05u] A. Einstein. Folgerungen aus den Capillaritätserscheinungen [AdP 4, 513 (1901)]. (German) [Consequences of the capillarization phenomena [AdP 4, 513 (1901)]]. Annalen der Physik (Berlin), 14(1):87–98, February 2005. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590048.

Einstein:2005:TKS

[Ein05v] A. Einstein. Ist die Trägheit eines Körpers von seinem Energieinhalt abhängig? [AdP 18, 639 (1905)]. (German) [Is the inertia of a body dependent on its energy content? [AdP 18, 639 (1905)]]. Annalen der Physik (Berlin), 14(1):225–228, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/andp.200590007.

Einstein:2005:KTW

[Ein05w] A. Einstein. Kinetische Theorie des Wärmegleichgewichtes und des zweiten Hauptsatzes der Thermodynamik [AdP 9, 417 (1902)]. (German) [Kinetic theory of heat equilibrium and the Second Law of Thermodynamics [AdP 9, 417 (1902)]]. Annalen der Physik (Berlin), 14(1):117–134, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590001.

Einstein:2005:LSG

[Ein05x] A. Einstein. Lichtgeschwindigkeit und Statik des Gravitationsfeldes [AdP 38, 355 (1912)]. (German) [Speed of light and statics of the gravitational field [AdP 38, 355 (1912)]]. Annalen der Physik (Berlin), 14(1):444–459, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590035.

Einstein:2005:NMA

[Ein05y] A. Einstein. Nachtrag zu meiner Arbeit: ,,Thermodynamische Begründung des photochemischen Äquivalentgesetzes“ [AdP 38, 881 (1912)]. (German) [Addendum to my paper: Thermodynamic Justification of the Photochemical Equivalent Law [AdP 38, 881 (1912)]]. Annalen der Physik (Berlin), 14(1):476–479, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590037.

Einstein:2005:PPS

[Ein05z] A. Einstein. Physics, philosophy, and scientific progress. Physics Today, 58(6):46–48, June 2005. CODEN PHTOAD. ISSN 0031-9228.

Einstein:2005:PAR

[Ein05-27] A. Einstein. Prinzipielles zur allgemeinen Relativitätstheorie [AdP 55, 241 (1918)]. (German) [Principles of the General Theory of Relativity [AdP 55, 241 (1918)]]. Annalen der Physik (Berlin), 14(1):577–581, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590046.

Einstein:2005:RGE

[Ein05-28] A. Einstein. Relativität und Gravitation. Erwiderung auf eine Bemerkung von M. Abraham [AdP 38, 1059 (1912)]. (German) [Relativity and gravity. Response to a remark by M. Abraham [AdP 38, 1059 (1912)]]. Annalen der Physik (Berlin), 14(1): 481–486, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/andp.200590039.

Einstein:2005:TOH

[Ein05-29] A. Einstein. Theorie der Opaleszenz von homogenen Flüssigkeiten und Flüssigkeitsgemischen in der Nähe des kritischen Zustandes [AdP 33, 1275 (1910)]. (German) [Theory of opalescence of homogeneous liquids and liquid mixtures near the critical state [AdP 33, 1275 (1910)]]. Annalen der Physik (Berlin), 14(1): 368–391, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/andp.200590026.

Einstein:2005:TBP

[Ein05-30] A. Einstein. Thermodynamische Begründung des photochemischen Äquivalentgesetzes [AdP 37, 832 (1912)]. (German) [Thermodynamic rationale for the photochemical equivalent law [AdP 37, 832 (1912)]]. Annalen der Physik (Berlin), 14(1):436–443, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590034.

Einstein:2005:ESA

[Ein05-31] A. Einstein. Über den Einflußder Schwerkraft auf die Ausbreitung des Lichtes [AdP 35, 898 (1911)]. (German) [On the influence of gravity on the propagation of light [AdP 35, 898 (1911)]]. Annalen der Physik (Berlin), 14(1):425–435, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/andp.200590033.

Einstein:2005:GST

[Ein05-32] A. Einstein. Über die Gültigkeitsgrenze des Satzes vom thermodynamischen Gleichgewicht und über die Möglichkeit einer neuen Bestimmung der Elementarquanta [AdP 22, 569 (1907)]. (German) [On the validity limit of the thermodynamic equilibrium theorem on the possibility of a new determination of the elementary quanta [AdP 22, 569 (1907)]]. Annalen der Physik (Berlin), 14(1):292–295, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590014.

Einstein:2005:MNP

[Ein05-33] A. Einstein. Über die Möglichkeit einer neuen Prüfung des Relativitätsprinzips [AdP 23, 197 (1907)]. (German) [On the possibility of a new examination of the Principle of Relativity [AdP 23, 197 (1907)]]. Annalen der Physik (Berlin), 14(1):297–299, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590016.

Einstein:2005:RGT

[Ein05-34] A. Einstein. Über die vom Relativitätsprinzip geforderte Trägheit der Energie [AdP 23, 371 (1907)]. (German) [On the inertia of energy demanded by the Principle of Relativity [AdP 23, 371 (1907)]]. Annalen der Physik (Berlin), 14(1):303–316, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590018.

Einstein:2005:MTW

[Ein05-35] A. Einstein. Über die von der molekularkinetischen Theorie der Wärme geforderte Bewegung von in ruhenden Flüssigkeiten suspendierten Teilchen [AdP 17, 549 (1905)]. (German) [On the movement of solids suspended in dormant fluids, required by the molecular-kinetic theory of heat [AdP 17, 549 (1905)]]. Annalen der Physik (Berlin), 14(1):182–193, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590005.

Einstein:2005:MBV

[Ein05-36] A. Einstein. Über eine Methode zur Bestimmung des Verhältnisses der transversalen und longitudinalen Masse des Elektrons [AdP 21, 583 (1906)]. (German) [A method for determining the relationship of the transverse and longitudinal masses of the electron [AdP 21, 583 (1906)]]. Annalen der Physik (Berlin), 14(1):275–279, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590012.

Einstein:2005:EVL

[Ein05-37] A. Einstein. Über einen die Erzeugung und Verwandlung des Lichtes betreffenden heuristischen Gesichtspunkt [AdP 17, 132 (1905)]. (German) [On a heuristic point of view concerning the creation and transformation of light [AdP 17, 132 (1905)]]. Annalen der Physik (Berlin), 14(1):164–181, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590004.

Einstein:2005:FKA

[Ein05-38] A. Einstein. Über Friedrich Kottlers Abhandlung ,,Über Einsteins Äquivalenzhypothese und die Gravitation“ [AdP 51, 639 (1916)]. (German) [On Friedrich Kottler’s paper On Einstein’s Equivalence Hypothesis and Gravity [AdP 51, 639 (1916)]]. Annalen der Physik (Berlin), 14(1):572–576, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590045.

Einstein:2005:TTP

[Ein05-39] A. Einstein. Ueber die thermodynamische Theorie der Potentialdifferenz zwischen Metallen und vollständig dissociirten Lösungen ihrer Salze und über eine elektrische Methode zur Erforschung der Molecularkräfte [AdP 8, 798 (1902)]. (German) [On the thermodynamic theory of the potential difference between metals and fully dissociated solutions of their salts and on an electrical method for the study of molecular forces [AdP 8, 798 (1902)]]. Annalen der Physik (Berlin), 14(1):99–116, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590050.

Einstein:2005:AMT

[Ein05-40] A. Einstein. Zur allgemeinen molekularen Theorie der Wärme [AdP 14, 354 (1904)]. (German) [On the general molecular theory of heat [AdP 14, 354 (1904)]]. Annalen der Physik (Berlin), 14(1):154–163, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590003.

Einstein:2005:EBK

[Ein05-41] A. Einstein. Zur Elektrodynamik bewegter Körper [AdP 17, 891 (1905)]. (German) [Electrodynamics of moving bodies [AdP 17, 891 (1905)]]. Annalen der Physik (Berlin), 14(1):194–224, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590006.

Einstein:2005:TBB

[Ein05-42] A. Einstein. Zur Theorie der Brownschen Bewegung [AdP 19, 371 (1906)]. (German) [On the theory of Brownian motion [AdP 19, 371 (1906)]]. Annalen der Physik (Berlin), 14(1):248–258, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590009.

Einstein:2005:TLL

[Ein05-43] A. Einstein. Zur Theorie der Lichterzeugung und Lichtabsorption [AdP 20, 199 (1906)]. (German) [The theory of light generation and absorption [AdP 20, 199 (1906)]]. Annalen der Physik (Berlin), 14(1):259–267, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590010.

Einstein:2005:TSG

[Ein05-44] A. Einstein. Zur Theorie des statischen Gravitationsfeldes [AdP 38, 443 (1912)]. (German) [Theory of the static gravitational field [AdP 38, 443 (1912)]]. Annalen der Physik (Berlin), 14(1): 460–475, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/andp.200590036.

Einstein:2005:AEMb

[Ein05-45] Albert Einstein. Albert Einstein: Out of My Later Years Through His Own Word. Castle Books, Edison, NJ, USA, 2005. ISBN 0-7858-2045-0. 282 pp. LCCN QC16.E5 A3 2005.

Einstein:2005:AEA

[Ein05-46] Albert Einstein. Albert Einstein to Arnold Sommerfeld. Physics Today, 58(2):14, February 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Short letter of 5 January 1908 written in Bern, Switzerland, translated and annotated by Bertram Schwarzschild.

Einstein:2005:LMS

[Ein05-47] Albert Einstein. Lettres á Maurice Solovine: reproduites en facsimilé et traduites en français. J. Gabay, Paris, France, 2005. ISBN 2-87647-275-9 (paperback). xiii + 139 pp. LCCN ????

Einstein:2005:MRa

[Ein05-48] Albert Einstein. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Routledge, London, UK, 2005. ISBN 0-203-44953-3. 177 pp. LCCN QC173.55.

Einstein:2005:TRS

[Ein05-49] Albert Einstein. Theoretical remark on the superconductivity of metals. arXiv.org, October 27, 2005. URL http://arXiv. org/abs/physics/0510251. Translation of Albert Einstein’s article “Theoretische Bemerkungen zur Supraleitung der Metalle” [Ein22m], Gedenkboek aangeb. aan H. Kamerlingh Onnes, eaz. Leiden, E. IJdo, 1922 (page 435) Translated by Bjoern S. Schmekel with the courtesy of the Kamerlingh Onnes Laboratory, Leiden Institute of Physics, Leiden University. A typo in the citation of the original article has been corrected.

Einstein:2006:ER

[Ein06a] Albert Einstein. The Einstein reader. Citadel Press, New York, NY, USA, 2006. ISBN 0-8065-2791-9 (paperback). ix + 257 pp. LCCN QC16.E5 A25 2006.

Einstein:2006:RSGa

[Ein06b] Albert Einstein. Relativity: The Special and the General Theory. Penguin classics. Penguin, London, UK and New York, NY, USA, 2006. ISBN 0-14-303982-2, 0-486-41714-X, 1-4406-2712-6 (e-book). xxvii + 129 pp. LCCN QC173.55 .E384513 2006.

URL http://www.loc.gov/catdir/toc/fy0703/2005058696.

html. Translated by Robert W. Lawson.

Einstein:2006:WSI

[Ein06c] Albert Einstein. The world as I see it. Filiquarian Publishing, Minneapolis, MN, USA, 2006. ISBN 1-59986-824-5. 127 pp. LCCN QC71 .E5513 2006.

Einstein:2007:X

[Ein07a] Albert Einstein. E = mc^{2}. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5 .W53 2007. URL

http://store.greatbooks.org/adu-wtm.html. Foreword by

Alan Lightman.

Einstein:2007:ETR

[Ein07b] Albert Einstein. Ether and the theory of relativity. In Renn et al. [RSS^{+}07a], chapter 34, pages 613–619. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6

.G469 2007. URL https://link.springer.com/chapter/10.

1007/978-1-4020-4000-9_34. Originally published as “Äther und Relativitätstheorie at Springer, Berlin on the basis of an address delivered on 5 May 1920 at the University of Leyden.

Einstein:2007:PSP

[Ein07c] Albert Einstein. On the present state of the problem of gravitation. In Renn et al. [RSS^{+}07a], chapter 31, pages 543–566. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v.

3; QC173.6 .G469 2007. URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-9_31. Originally pub-

lished as “Zum gegenwärtigen Stande des Gravitationsproblems”, Physikalische Zeitschrift 14 1249–1262 (1913).

Einstein:2007:RP

[Ein07d] Albert Einstein. On the relativity problem. In Renn et al. [RSS^{+}07a], chapter 33, pages 605–612. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/

978-1-4020-4000-9_33. Originally published as “Zum Relativitätsproblem”, Scientia 15, 337–348 (1914).

Einstein:2007:SR

[Ein07e] Albert Einstein. The scientist’s responsibilities. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5

.W53 2007. URL http://store.greatbooks.org/adu-wtm.

html. Foreword by Alan Lightman.

Einstein:2007:STR

[Ein07f] Albert Einstein. The special theory of relativity. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN

QC7.5 .W53 2007. URL http://store.greatbooks.org/adu-

wtm.html. Foreword by Alan Lightman.

Einstein:2007:TR

[Ein07g] Albert Einstein. The theory of relativity. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5

.W53 2007. URL http://store.greatbooks.org/adu-wtm.

html. Foreword by Alan Lightman.

Einstein:2007:W

[Ein07h] Albert Einstein. The way out. In Masters and Way [MW07], page ?? ISBN 1-59558-227-4 (hardcover). LCCN UG1282.A8

O54 2007.  URL http://thenewpress.com/index.php?
option=com_title&task=view_title&metaproductid=1703;
 http://www.loc.gov/catdir/toc/ecip0718/2007020838.

htm. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Einstein:2008:NST

[Ein08a] Albert Einstein. The necessity of students & teachers as science researchers. The American Biology Teacher, 70(9):518–519, November/December 2008. ISSN 0002-7685 (print), 0002-8444 (electronic), 1938-4211, 1938-422X.

Einstein:2008:UOL

[Ein08b] Albert Einstein. Unpublished opening lecture for the course on the Theory of Relativity in Argentina, 1925 (reprinted). Science in Context, 21(3):451–459, September 2008. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Einstein:2009:CJV

[Ein09a] Albert Einstein. Comment je vois le monde. (French) [How I see the world], volume 183 of Champs. Sciences. Flammarion, Paris, France, 2009. ISBN 2-08-122904-8 (paperback). 245 pp. LCCN QC16E5 A514.2009.

Einstein:2009:EES

[Ein09b] Albert Einstein. Einstein’s essays in science. Dover, New York, NY, USA, 2009. ISBN 0-486-47011-3. xi + 114 pp. LCCN QC71 .E513 2009. URL http://store.

doverpublications.com/0486470113.html;   http://www.
loc.gov/catdir/enhancements/fy0909/2009000860-d.htm.

Einstein:2009:RET

[Ein09c] Albert Einstein. Relativity: Einstein’s theory of spacetime, time dilation, gravity, and cosmology. Red and Black Publishers, St. Petersburg, FL, USA, 2009. ISBN 1-934941-46-8. 151 pp. LCCN QC173.6 .E3613 2009.

Einstein:2009:WS

[Ein09d] Albert Einstein. Why socialism? Monthly Review — an Independent Socialist Magazine, 61(1):55–61, May 2009. ISSN 0027-0520.

Einstein:2010:CSQ

[Ein10] Albert Einstein. Comment: a skeptic of quantum theory explains his misgivings. Science News (Washington, DC), 178(11): 36, 2010. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Einstein:2011:GGF

[Ein11a] Albert Einstein. Geleitwort. (German) [foreword]. In Diels [Die11], page ?? ISBN 3-05-005392-5, 3-05-006392-0. LCCN ????

Einstein:2011:MID

[Ein11b] Albert Einstein. Maxwell’s influence on the development of the conception of physical reality. In Thomson [Tho11a], pages 66– 73. ISBN 1-107-67095-0 (paperback). LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Einstein:2013:RSG

[Ein13] Albert Einstein. Relativity: the Special and the General Theory. Routledge great minds. Routledge, London, UK, 2013. ISBN 0-415-85471-7 (paperback), 1-315-88674-X (e-book). xvii + 166 pp. LCCN QC173.55 .E384513 2013. Translation from German by Robert W. (Robert William) Lawson and Derek J. Raine.

Einstein:2015:CJV

[Ein15a] Albert Einstein. Comment je vois le monde. (French) [How I see the world]. ????, ????, 2015. ISBN ???? 147 pp. LCCN ???? French translation by Faroq Al-Hamid of Mein Weltbild.

Einstein:2015:RSG

[Ein15b] Albert Einstein. Relativity: the Special and the General theory. Princeton University Press, Princeton, NJ, USA, 100th anniversary edition, 2015. ISBN 0-691-16633-1 (hardcover). xviii + 300 pp. LCCN QC173.55 .E3843 2015. With commentaries and background material by Hanoch Gutfreund and Jürgen Renn.

Einstein:2017:RPH

[Ein17] Albert Einstein. The real problem is in the hearts of men. In Johnsen [Joh17], pages 100–106.

Einstein:2022:TDA

[Ein22] Albert Einstein. Travel Diaries of Albert Einstein: South America, 1925. Princeton University Press, Princeton, NJ, USA, 2022. ISBN 0-691-20102-1 (hardcover), 0-691-24250-X (e-book). xx + 265 pp. LCCN QC16.E5 A3 2023.

Eisenhart:1922:EES

[Eis22] Luther Pfahler Eisenhart. THE Einstein equations for the solar field from the Newtonian point of view. Science, 55(1430): 570–572, May 26, 1922. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/55/1430/570.full.pdf.

Eisenhart:1923:ES

[Eis23] L. P. Eisenhart. Einstein and Soldner. Science, 58(1512):516– 517, December 21, 1923. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/58/1512/516.2.full.pdf.

Eistein:1937:CJV

[Eis37] Albert Eistein. Comment je vois le monde. (French) [How I see the world]. Flammarion, Paris, France, 1937. vi + 254 pp. LCCN ????

Eisenstaedt:1986:RGE

[Eis86] Jean Eisenstaedt. La relativité générale à l’étiage: 1925– 1955. (French) [General Relativity at low-water level: 1925– 1955]. Archive for History of Exact Sciences, 35(2):115–185, June 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=2&spage=115.

Eisenstaedt:1993:DBB

[Eis93] Jean Eisenstaedt. Dark bodies and black holes, magic circles and Montgolfiers: Light and gravitation from Newton to Einstein. Science in Context, 6(1):83–106, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Eisner:2000:RBP

[Eis00] Werner Eisner. Rezension: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Herausgegeben von Karl von Meyenn (Band IV, Teil I: 1950–1952; Teil II: 1953–1954). Berlin/Heidelberg/New York: Springer 1996 und 1999; 968 und 1100 Seiten. (German) [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, and others. Edited by Karl of Meyenn (Volume IV, part I: 1950–1952, Part II: 1953–1954). Berlin / Heidelberg / New York: Springer 1996 and 1999, 968 and 1100 pages]. Berichte zur Wissenschaftsgeschichte, 23(1):59–62, ???? 2000. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Eisenstaedt:2002:ERG

[Eis02] Jean Eisenstaedt. Einstein et la relativité générale: les chemins de l’espace-temps. (French) [Einstein and General Relativity: the ways of space–time]. CNRS histoire des sciences. CNRS éditions, Paris, France, 2002. ISBN 2-271-05880-5. ISSN 1292-4296. 344 pp. LCCN QC173.55. EUR 29.

Eisenstaedt:2006:CHR

[Eis06] Jean Eisenstaedt. The Curious History of Relativity: How Einstein’s Theory of Gravity Was Lost and Found Again, volume 15 of CNRS editions. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11865-5. ix + 363 pp. LCCN QC173.6 EIS; 06.E06047.

Eisenstaedt:2007:NEF

[Eis07] Jean Eisenstaedt. From Newton to Einstein: A forgotten relativistic optics of moving bodies. American Journal of Physics, 75(8):741–746, August 2007. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Eisinger:2011:ER

[Eis11] Josef Eisinger. Einstein on the road. Prometheus Books, Amherst, NY, USA, 2011. ISBN 1-61614-460-2 (hardcover), 1-61614-461-0 (e-book). xxxi + 219 pp. LCCN QC16.E5 E586 2011.

Earman:1993:AGN

[EJN93] John Earman, Michel Janssen, and John D. Norton, editors. The Attraction of Gravitation: New Studies in the History of General Relativity, volume 5 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1993. ISBN 3-7643-3624-2, 0-8176-3624-2. x + 432 pp. LCCN QC173.6

.A85 1993. URL http://www.zentralblattmath.org/zmath/
en/search/?an=0846.01009.

Einstein:1944:AP

[EK44] Albert Einstein and Erich Kahler. The Arabs and Palestine. Princeton Herald, April 14, 28, 1944.

Einstein:1953:EAT

[EK53] A. Einstein and B. Kaufman. Sur l’état actuel de la théorie générale de la gravitation. (French) [On the current state of the General Theory of Gravitation]. In Louis de Broglie, physicien et penseur (French) [Louis de Broglie, physicist and thinker] [Geo53], pages 321–336. LCCN QC71 .L88.

Einstein:1954:APF

[EK54] Albert Einstein and Bruria Kaufman. Algebraic properties of the field in the relativistic theory of the asymmetric field. Annals of Mathematics (2), 59(1):230–244, January 1954. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL 2-W;   http://links.jstor.org/sici?sici=
0003-486X(195403)2:59:2<230:APOTFI>2.0.CO.

Einstein:1955:NFG

[EK55] Albert Einstein and Bruria Kaufman. A new form of the general relativistic field equations. Annals of Mathematics (2), 62(1):128–138, July 1955. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL 2-3; http:

//links.jstor.org/sici?sici=0003-486X(195507)2:62:1<
128:ANFOTG>2.0.CO.

Eisenstaedt:1992:SHG

[EK92] Jean Eisenstaedt and Anne J. Kox, editors. Studies in the history of General Relativity: based on the proceedings of the 2nd International Conference on the History of General Relativity, Luminy, France, 1988, volume 3 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1992. ISBN 0-8176-3479-7 (Boston), 3-7643-3479-7 (Berlin). LCCN QC173.6 .I572 1988.

Einstein:2003:EMS

[EK03] Albert Einstein and Martin J. Klein. Einstein’s 1912 Manuscript on the Special Theory of Relativity. George Braziller, in association with the Edmond J. Safra Philanthropic Foundation, New York, NY, USA, revised edition, 2003. ISBN 0-8076-1532-3. 174 pp. LCCN QC173.65 .E3613 2003. English and German texts appear on facing pages, with the German handwritten.

Eisenstaedt:2005:UGR

[EK05] Jean Eisenstaedt and Anne J. Kox, editors. The universe of General Relativity, volume 11 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2005. ISBN 0-8176-4380-X. xii + 500 pp. LCCN QC173.5 .U55 2005.

Ekspong:1993:P

[Eks93] Gösta Ekspong, editor. Physics 1981–1990. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-0728-X, 981-02-0729-8 (paperback). xi + 738 pp. LCCN QC6.2 .P49 1993.

Ekspong:1997:P

[Eks97] Gösta Ekspong, editor. Physics, 1991–1995. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2677-2, 981-02-2678-0 (paperback). ix + 221 pp. LCCN QC6.2 .P47 1997.

Ekspong:1999:DNL

[Eks99] Gösta Ekspong. The dual nature of light, as reflected in the Nobel Archive. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 143 (1):42–49, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181972.

Einstein:2008:CBI

[EKS^{+}08] Albert Einstein, Anne J. Kox, Tilman Sauer, Diana Kormos Buchwald, Rudy Hirschmann, Osik Moses, Benjamin Aronin, and Jennifer Stolper. Cumulative bibliography and index of citations to volumes 1–10: Collected papers of Albert Einstein. Princeton, NJ, USA, December 5, 2008. 158 pp. URL http://

press.princeton.edu/books/einstein11/c_biblio.pdf.

Einstein:1908:BAE

[EL08a] Albert Einstein and Jakob Johann Laub. Berichtigung zur Abhandlung: ,,Über die elektromagnetischen Grundgleichungen für bewegte Körper“. (German) [correction to the paper: “On the electromagnetic principal equations of moving bodies”]. Annalen der Physik (1900) (series 4), 332(11):232, 1908. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [EL08b, EL09].

Einstein:1908:EGB

[EL08b] Albert Einstein and Jakob Johann Laub. Über die elektromagnetischen Grundgleichungen für bewegte Körper. (German) [On the electromagnetic principal equations of moving bodies]. Annalen der Physik (1900) (series 4), 331(8):532–540, 1908. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See correction [EL08a] and supplement [EL09].

Einstein:1908:IEF

[EL08c] Albert Einstein and Jakob Johann Laub. Über die im elektromagnetischen Felde auf ruhende Körper ausgeübten ponderomotorischen Kräfte. (German) [On the ponderomotive forces exerted in the electromagnetic field on inertial bodies]. Annalen der Physik (1900) (series 4), 331(8):541–550, 1908. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1909:BUA

[EL09] Albert Einstein and Jakob Johann Laub. Bemerkungen zu unserer Arbeit: “Über die elektromagnetischen Grundgleichungen für bewegte Körper”. (German) [Remark on our work: “On the electromagnetic principal equations of moving bodies”]. Annalen der Physik (1900) (series 4), 333(2):445–447, 1909. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [EL08b, EL08a].

Einstein:2005:BUA

[EL05a] A. Einstein and J. Laub. Bemerkungen zu unserer Arbeit: ,,Über die elektromagnetischen Grundgleichungen für bewegte Körper“ [AdP 28, 445 (1909)]. (German) [Comments on our work: On the basic electromagnetic equations for moving bodies [AdP 28, 445 (1909)]]. Annalen der Physik (Berlin), 14(1): 339–342, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/andp.200590022.

Einstein:2005:BAE

[EL05b] A. Einstein and J. Laub. Berichtigung zur Abhandlung: ,,Über die elektromagnetischen Grundgleichungen für bewegte Körper“ [AdP 27, 232 (1908)]. (German) [Correction to the treatise: On the Electromagnetic Propositions for Moving Bodies [AdP 27, 232 (1908)]]. Annalen der Physik (Berlin), 14(1):337–338, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590021.

Einstein:2005:EGB

[EL05c] A. Einstein and J. Laub. Über die elektromagnetischen Grundgleichungen für bewegte Körper [AdP 26, 532 (1908)]. (German) [On the basic electromagnetic equations for moving bodies [AdP 26, 532 (1908)]]. Annalen der Physik (Berlin), 14(1):317–326, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590019.

Einstein:2005:IEF

[EL05d] A. Einstein and J. Laub. Über die im elektromagnetischen Felde auf ruhende Körper ausgeübten ponderomotorischen Kräfte [AdP 26, 541 (1908)]. (German) [On the ponderomotive forces exerted on resting bodies in the electromagnetic field [AdP 26, 541 (1908)]]. Annalen der Physik (Berlin), 14(1):327–336, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200590020.

Elbaz:2013:PAE

[Elb13] Claude Elbaz. Sur les programmes d’Albert Einstein et de Louis de Broglie. Une contribution. (French) [On the programs of Albert Einstein and Louis de Broglie. A contribution]. Annales de la Fondation Louis de Broglie, 38(1):195–217, 2013. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL

http://aflb.ensmp.fr/AFLB-381/aflb381m758.htm.

Einstein:2005:RSG

[ELGC05] Albert Einstein, Robert W. Lawson, Robert Geroch, and David C. Cassidy. Relativity: The Special and General Theory. Pi Press, New York, NY, USA, the masterpiece science edition, 2005. ISBN 0-13-186261-8. xxvi + 259 pp. LCCN QC173.55 .E384513 2005. URL http://catdir.loc.

gov/catdir/enhancements/fy1204/2004117390-b.html;
http://catdir.loc.gov/catdir/enhancements/fy1204/2004117390-

d.html. Translated by Robert W. Lawson. Introduction by Roger Penrose. Commentary by Robert Geroch. Historical essay by David Cassidy.

Einstein:2006:RSGb

[ELGC06] Albert Einstein, Robert W. Lawson, Robert Geroch, and David C. Cassidy. Relativity: The Special and the General Theory. Penguin classics. Plume Book, New York, NY, USA, 2006. ISBN 0-452-28784-7. xxvii + 129 pp. LCCN QC173.55

.E3613 2006. URL http://www.loc.gov/catdir/toc/fy0703/

2005058696.html. Introduction by Roger Penrose.

Eliot:1950:HB

[Eli50] George Fielding Eliot, editor. The H bomb. Didier, New York, NY, USA, 1950. 175 pp. LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Elisabeth:1992:NIS

[Eli92] Crawford Elisabeth. Nationalism and Internationalism in Science, 1880–1939. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover), 0-521-52474-1 (paperback), 0-511-59969-2 (e-book). LCCN 18 f. Four Studies of the Nobel Population.

Elkana:1982:MS

[Elk82] Yehuda Elkana. The myth of simplicity. In Holton and Elkana [HE82], pages 205–252. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Ellis:1989:EUH

[Ell89a] George Ellis. The expanding universe. A history of cosmology from 1917 to 1960. In Howard and Stachel [HS89b], pages 367–431. ISBN 0-8176-3392-8, 3-7643-3392-8. LCCN

QC173.6 .E38 1989. URL http://www.springer.com/us/
book/9780817633929.

Ellis:1989:CR

[Ell89b] George F. R. Ellis. Concepts of relativity. Physics World, 2(7): 16, July 1989. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/2/7/phwv2i7a14.pdf.

Ellis:1990:IRC

[Ell90]  George Ellis.  Innovation, resistance and change. The

transition to the expanding universe. In Bertotti et al. [BBBM90], pages 97–113. ISBN 0-521-37213-5. LCCN

QB981 .M774 1990. URL http://www.loc.gov/catdir/
description/cam024/90041803.html;   http://www.loc.
gov/catdir/toc/cam028/90041803.html.

Ellis:2005:EQU

[Ell05] John Ellis. Einstein’s quest for unification. Physics World, 18(1):56–57, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a37.
pdf.

Ellis:2010:FWA

[Ell10] Elsi Vassdal Ellis. Fall-Out: We Are All Downwinders. EVE Press, Bellingham, WA, USA, 2010. 14 pp. LCCN N7433.4.E347 F366 2010.

Elsbach:1924:KEU

[Els24] Alfred Coppel Elsbach. Kant und Einstein. Untersuchungen über das Verhältnis der modernen Erkenntnistheorie zur Relativitätstheorie. (German) [Kant and Einstein: Investigations on the relationship of the modern theory of knowledge to the Theory of Relativity]. Walter de Gruyter, Berlin, Germany, 1924. viii + 374 pp. LCCN QC6 .E6. See Einstein’s review [Ein24a].

Elsasser:1978:MPA

[Els78] Walter M. (Walter Maurice) Elsasser. Memoirs of a physicist in the atomic age. Science History Publications, New York, NY, USA, 1978. ISBN 0-88202-178-8 (Science History), 0-85274-400-5 (Adam Hilger). xii + 268 pp. LCCN QC16.E58

A35. URL http://catalog.hathitrust.org/api/volumes/
oclc/3415166.html.

Else:1995:DAT

[Els95] Jon Else. The day after Trinity: J. Robert Oppenheimer and the atomic bomb. For the record. Voyager, New York, NY, USA, 1995. ISBN 1-55940-685-2. 1 computer optical disc (89 minutes) pp. LCCN QC16.O62.

Elzinga:1995:ELN

[Elz95] Aant Elzinga. Einstein in the Land of Nobel: An episode in the interplay of science, politics, epistemology and popular culture. In Gavroglu et al. [GSW95], chapter 5, pages 73–103. ISBN 0-7923-2991-0 (set), 0-7923-2988-0. LCCN Q174 .B67 vol.

163. URL https://link.springer.com/chapter/10.1007/
978-94-017-2658-0_5.

Elzinga:2006:ENP

[Elz06] Aant Elzinga. Einstein’s Nobel Prize: a glimpse behind closed doors: the archival evidence, volume 6 of Archives of the Nobel Museum. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2006. ISBN 0-88135-283-7. xii + 228 pp. LCCN QC16.E5 E68 2006.

Einstein:1920:PRO

[EM20] Albert Einstein and H. (Hermann) Minkowski. The Principle of Relativity: original papers. The University of Calcutta, Calcutta, India, 1920. 3 + xxiii + 186 pp. LCCN QC6 .E45. Translated by Meghnad N. Saha (1893–1956) and Satyendranath N. Bose (1894–1974). with a historical introduction, by P. C. Mahalanobis.

Einstein:1923:EBK

[EM23] Albert Einstein and Hans Mühsam. Experimentelle Bestimmung der Kanalweite von Filtern. (German) [Experimental determination of the channel width of filters]. Deutsch. med. Wochenschr., 49(??):1012–1013, 1923.

Einstein:1930:ZSS

[EM30] Albert Einstein and Walter Mayer. Zwei strenge statische Lósungen der Feldgleichungen der einheitlichen Feldtheorie. (German) [Two strict static solutions of the field equations of Uniform Field Theory]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):110–120, 1930. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/GEYDB1K1.

Einstein:1931:ETG

[EM31a] Albert Einstein and Walter Mayer. Einheitliche Theorie von Gravitation und Elektrizität. (German) [Unified Theory of Gravitation and Electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):541–557, 1931. See part 2 [EM32a].

Einstein:1931:SUK

[EM31b] Albert Einstein and Walter Mayer. Systematische Untersuchung über Kompatible Feldgleichungen welche in einem Riemannschen Räume mit Fern-Parallelismus gesetzt werden können. (German) [Systematic investigation of compatible field equations which can be set in a Riemannian space with remote parallelism]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ?? (??):257–265, 1931.

Einstein:1932:ETG

[EM32a] Albert Einstein and Walter Mayer. Einheitliche Theorie von Gravitation und Elektrizität. (German) [Unified Theory of Gravitation and Electricity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):130–137, 1932. URL http:/

/einstein-annalen.mpiwg-berlin.mpg.de/related_texts/
sitzungsberichte/TSH25SM0. See part 2 [EM31a].

Einstein:1932:SVS

[EM32b] Albert Einstein and Walter Mayer. Semi-Vektoren und Spinoren. (German) [Semivectors and spinors]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, 32(??):522–550, 1932.

URL http://einstein-annalen.mpiwg-berlin.mpg.de/related_
texts/sitzungsberichte/VA2YYBC8.

Einstein:1932:PES

[EM32c] Albert Einstein and James Murphy. Prologue and epilogue: a Socratic dialogue. In Max Planck, editor, Where is Science Going?, page ?? W. W. Norton & Co., New York, NY, USA, 1932. LCCN Q175 .P57 1932. Reprinted in [Ein77a].

Einstein:1933:SNF

[EM33a] Albert Einstein and Walter Mayer. Spaltung der natürlichsten Feldgleichungen für Semi-Vektoren in Spinor-Gleichungen vom Diracschen Typus. (German) [Splitting the most natural field equations for semivectors into spinor equations of the Dirac type]. Proc. Acad. Sci. Amst., 36 (part 2)(??):615–619, 1933.

Einstein:1933:DGS

[EM33b] Albert Einstein and Walther Mayer. Dirac-Gleichung für Semi-Vektoren. (German) [Dirac equations for semivectors]. Proc. Acad. Sci. Amst., 36 (part 2)(??):497–502, 1933.

Einstein:1934:DSV

[EM34] Albert Einstein and Walter Mayer. Darstellung der Semi-Vektoren als gewöhnliche Vektoren von besonderem Differentiationscharakter. (German) [Representation of the semivectors as ordinary vectors of a special differentiation character]. Annals of Mathematics (2), 35(1):104–110, January 1934. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL 2-T;   http://links.jstor.org/sici?sici=
0003-486X(193401)2:35:1<104:DDSAGV>2.0.CO.

Einstein:1949:DAE

[EM49] Albert Einstein and Robert A. Millikan. Dr. Albert Einstein and American colleagues, 1931. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 93(7):543–545, December 30, 1949. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-
049X(19491230)93:7<543:DAEAAC>2.0.CO%3B2-U;   http:/
/www.jstor.org/stable/3143144. Theory of Relativity in

Contemporary Science. Papers Read at the Celebration of the Seventieth Birthday of Professor Albert Einstein in Princeton, March 19, 1949.

Einstein:1984:EQA

[EM84] Albert Einstein and James Clerk Maxwell. Einstein’s quotations about Maxwell. In Berger [Ber84b], pages 11–14. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Enz:1988:WPG

[EM88] Charles P. (Charles Paul) Enz and K. V. Meyenn, editors. Wolfgang Pauli: das Gewissen der Physik. (German) [Wolfgang Pauli: the conscience of physics]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1988. ISBN 3-528-08993-8. xii + 546 pp. LCCN QC16.P37 W65 1988.

Einstein:1991:EPF

[EM^{+}91] Albert Einstein, Jean-Philippe Mathieu, et al. Écrits politiques. (French [Political writings]), volume 6 of Sources du savoir; Oeuvres choisies / Albert Einstein. Éditions du Seuil, Paris, France, 1991. ISBN 2-02-010181-5, 2-222-04646-7. 289 pp. LCCN QC16.E5 A612 B34. Translation by Jacques Duvernet and others from German to French.

Eisenstein:2004:BEC

[EM04] J. P. Eisenstein and A. H. MacDonald. Bose–Einstein condensation of excitons in bilayer electron systems. Nature, 432(7018): 691–694, December 9, 2004. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v432/n7018/full/nature03081.html.

Emanuel:1985:ECN

[Ema85] Gabriel Emanuel. Einstein: Children of night: two plays. Playwrights Canada, Toronto, ON, Canada, 1985. ISBN 0-88754-409-6. 128 pp. US$5.95.

ECAS:1948:PSS

[Eme48] Emergency Committee of Atomic Scientists. A policy for survival: A statement by Emergency Committee of Atomic Scientists — April 12, 1948. Bulletin of the Atomic Scientists, 4(6): 176, 188, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1996:BSE

[EMH96] Albert Einstein, Jerry Mayer, and John P. Holms. Bitesize Einstein: Quotations on Just About Everything From the Greatest Mind of the Twentieth Century. St. Martin’s Press, New York, NY, USA, 1996. ISBN 0-312-14551-9. 92 pp. LCCN QC16.E5 A25 1996b. URL http://www.loc.gov/

catdir/bios/hol054/96018830.html;   http://www.loc.
gov/catdir/description/hol042/96018830.html.

Einstein:1979:WGF

[EMvL79] Albert Einstein, Lise Meitner, and Max von Laue. Wissenschaft unter dem Gesetz der Freiheit. (German) [Science under the law of freedom]. Physik Journal, 35(5):192–199, May 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Earman:1997:CSE

[EN97] John Earman and John D. Norton, editors. The cosmos of science: essays of exploration. Pittsburgh-Konstanz series in the philosophy and history of science. University of Pittsburgh Press, Pittsburgh, PA, USA, 1997. ISBN 0-8229-3930-4 (clothbound). xvi + 581 pp. LCCN Q175.3 .C69 1996; Q175.3 .C69 1997.

Enebakk:2012:NSP

[Ene12] Vidar Enebakk. Nobel science of peace: Norwegian neutrality, internationalism and the Nobel Peace Prize. In Lettevall et al. [LSW12], chapter 14, pages 295–315. ISBN 0-203-11679-8, 0-415-89377-1. ISSN 1404-7586. LCCN D723 .N49 2012. URL

http://libanswers.liverpool.ac.uk/faq/182315.

Einstein:1914:AEG

[ENFB14] Albert Einstein, Georg Nicolai, Wilhelm Foerster, and Otto Buek. Aufrauf an die Europäer. (German) [Manifesto to the Europeans]. This was issued in opposition to the just-started war in Europe, and was widely circulated in Germany. It was finally published as the introduction to Nicolai’s 1917 book, “The Biology of War” [Nic17, Nic18, Nic19]., October 1914.

Engler:2002:EMB

[Eng02] Gideon Engler. Einstein and the most beautiful theories in physics. International Studies in the Philosophy of Science, 16 (1):27–37, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-

9281 (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/02698590120118800.

Engler:2005:EHT

[Eng05a] Gideon Engler. Einstein, his theories, and his aesthetic considerations. International Studies in the Philosophy of Science, 19(1): 21–30, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/02698590500051068.

English:2005:EPP

[Eng05b] James F. English. The Economy of Prestige: Prizes, Awards, and the Circulation of Cultural Value. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01884-2 (paperback), 0-674-03043-5 (hardcover), 0-674-03653-0 (e-book). xii + 409 pp. LCCN AS935 .E54 2005.

Engdahl:2008:L

[Eng08] Horace Engdahl, editor. Literature, 2001–2005. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-279-436-0 (paperback), 981-279-435-2. ix + 104 pp. LCCN PN781 .L58 2008.

Englert:2014:NLB

[Eng14] François Englert. Nobel lecture: The BEH mechanism and its scalar boson. Reviews of Modern Physics, 86(3):843–??, July/ September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.843.

Engelhardt:2017:FFG

[Eng17a] W. Engelhardt. Free fall in gravitational theory, January

2017. URL https://www.researchgate.net/publication/
312118218_Free_Fall_in_Gravitational_Theory.

Engler:2017:RJC

[Eng17b] Fynn Ole Engler. Rezensionen: Jimena Canales, The Physicist & the Philosopher: Einstein, Bergson, and the Debate That Changed Our Understanding of Time, Princeton: Princeton University Press 2015. 488 S., $24,95. ISBN 978-1-4008-6577-2. Berichte zur Wissenschaftsgeschichte, 40(4):402–403, December 2017. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Enke:2018:ESE

[Enk18] Dr. Ulrike Enke. “Der erste zu sein.” [1] — Über den ersten Medizinnobelpreis für Emil von Behring im Jahr 1901. (German) [“The first to be.” [1] — on the first Nobel Prize in Medicine for Emil von Behring in 1901]. Berichte zur Wissenschaftsgeschichte, 41(1):19–46, March 2018. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
bewi.201801866.

Einstein:1960:EPa

[ENN60a] Albert Einstein, Otto Nathan, and Heinz Norden. Einstein on peace. Methuen and Co. Ltd., London, UK, 1960. xvi + 704 pp. LCCN JX 1952.E44 NAT.

Einstein:1960:EPc

[ENN60b] Albert Einstein, Otto Nathan, and Heinz Norden. Einstein on peace. Schocken Books, New York, NY, USA, 1960. xvi + 704 pp. LCCN JZ 5574 EIN. Edited by Otto Nathan and Heinz Norden. Preface by Bertrand Russell.

Einstein:1963:EP

[ENN63] Albert Einstein, Otto Nathan, and Heinz Norden. Einstein on peace. Methuen and Co. Ltd., London, UK, 1963. xvi + 704 pp. LCCN JX 1952.E44 NAT.

Einstein:1968:EP

[ENN68] Albert Einstein, Otto Nathan, and Heinz Norden. Einstein on peace. Schocken Books, New York, NY, USA, 1968. xvi + 704 pp. LCCN JZ 5574 EIN. Edited by Otto Nathan and Heinz Norden. Preface by Bertrand Russell.

Enz:2002:NTB

[Enz02] Charles P. (Charles Paul) Enz. No time to be brief: a scientific biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http://www.loc.

gov/catdir/enhancements/fy0614/2002726902-d.html;
http://www.loc.gov/catdir/enhancements/fy0614/2002726902-
t.html.

Enz:2005:PGB

[Enz05] Charles P. (Charles Paul) Enz. “Pauli hat gesagt”: eine Biografie des Nobelpreisträgers Wolfgang Pauli, 1900–1958. (German) [Pauli Said: a Biography of the Nobel Prize winner Wolfgang Pauli, 1900–1958]. Verlag Neue Zürcher Zeitung, Zürich, Switzerland, 2005. ISBN 3-03823-144-4. 167 pp. LCCN QC16.P37 E593 2005.

Elkana:1979:EEJ

[EO79] Yehuda Elkana and Adi Ophir. Einstein 1879–1979 exhibition: Jewish National and University Library, Berman Hall, Jerusalem, March, 1979. Raphael Haim Hacohen Press, Jerusalem, Israel, 1979. ISBN ???? 132 (est.) pp. LCCN ????

Eotvos:1891:AEV

[Eöt91] R. Baron Eötvös. Über die Anziehung der Erde auf verschiedene Substanzen. (German) [On the gravitation produced by the Earth on different substances]. Mathematische und Naturwissenschaftliche Berichte aus Ungarn, 8(??):65–68, ???? 1891.

URL http://www.kfki.hu/eotvos/onehund.html. Read at a

meeting of the Academy on 20 January 1889. German translation of the Hungarian original [von90]. English translation in [Eöt08].

Eotvos:1900:ESN

[Eöt00] R. Eötvös. Étude sur les surfaces de niveau et la variation de la pesanteur et du champ magnétique. (French) [Study of level surfaces and the variation of weight and magnetic field]. Rapports du congrés international de physique, pages 371–393, 1900.

URL http://www.kfki.hu/eotvos/onehund.html.

Eotvos:2008:GPE

[Eöt08] Eötvös Loránd. On the gravitation produced by the Earth on different substances. The Abraham Zelmanov Journal: The journal for General Relativity, gravitation and cosmology, 1(??): 6–9, ???? 2008. CODEN ???? ISSN 1654-9163. URL http:

//zelmanov.ptep-online.com/html/zj-2008-02.html. See

Hungarian original [von90] and concurrent German translation [Eöt91].

Einstein:1943:NER

[EP43] Albert Einstein and Wolfgang Pauli. On the non-existence of regular stationary solutions of relativistic field equations. Annals of Mathematics (2), 44(2):131–137, April 1943. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL 2-Q;   http://links.jstor.org/sici?sici=
0003-486X(194304)2:44:2<131:OTNORS>2.0.CO.

Einstein:2006:AEM

[EP06] A. Einstein and M. Planck. Antrittsrede und Erwiderung von Max Planck am Leibniztag. (German) [inaugural speech and reply from Max Planck at the Leibniz Workshop]. In Albert Einstein: Akademie-Vorträge: Sitzungsberichte der Preußischen Akademie der Wissenschaften, 1914–1932. (German) [Albert Einstein: Academy lectures: Proceedings of the Prussian Academy of Sciences, 1914–1932] [Sim06], pages 1–7. ISBN 3-527-40609-3 (hardcover), 3-527-60895-8 (e-book), 3-527-62244-6 (e-book). LCCN QC71 .E36 2006.

Einstein:1935:CQM

[EPR35] Albert Einstein, Boris Podolsky, and Nathan Rosen. Can quantum mechanical description of physical reality be considered complete? Physical Review (2), 47(10):777–780, May 15, 1935.

CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v47/i10/
p777_1.

Evans:2013:NSE

[EPW13] P. W. Evans, Huw Price, and K. B. Wharton. New slant on the EPR–Bell experiment. The British Journal for the Philosophy of Science, 64(2):297–324, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/64/2/297.full.pdf+html.

Einstein:1935:PPG

[ER35] A. Einstein and N. Rosen. The particle problem in the General Theory of Relativity. Physical Review (2), 48(1):73–77, July 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://link.aps.org/abstract/PR/v48/
p73.

Einstein:1936:TBP

[ER36] A. Einstein and N. Rosen. Two-body problem in General Relativity Theory. Physical Review, 49(5):404–405, March 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/abstract/PR/v49/p404/
s2.

Einstein:1937:GW

[ER37] A. Einstein and N. Rosen. On gravitational waves. Journal of The Franklin Institute, 223(1):43–54, January 1937. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Einstein:1955:ERM

[ER55] Albert Einstein and Bertrand Russell. The Einstein–Russell Manifesto. Issued in London, UK on July 9, 1955 (after Einstein’s death on 18 April 1955)., 1955.

Einstein:1991:PNP

[ER91] Albert Einstein and Bertrand Russell. Le pouvoir nu: propos sur la guerre et la paix, 1914–1955. (French) [Raw power: on war and peace, 1914–1955]. Collection Savoir/Cultures. Hermann, Paris, France, 1991. ISBN 2-7056-6156-4. 222 pp. LCCN ????

Engler:2013:WPM

[ER13] Fynn Ole Engler and Jürgen Renn. Wissenschaftliche Philosophie, moderne Wissenschaft und historische Epistemologie: Albert Einstein, Ludwig Fleck und Moritz Schlick im Ringen um die wissenschaftliche Rationalität. (German) [Scientific philosophy, modern science and historical epistemology: Albert Einstein, Ludwig Fleck and Moritz Schlick in the struggle for scientific rationality], volume 400 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wiss.-Geschichte, Berlin, Germany, 2013. 107 pp. LCCN ????

Erbe:2010:LED

[Erb10] Michael Erbe. Von Leibniz zu Einstein: drei Jahrhunderte Wissenschaft in Berlin. (German) [From Leibniz to Einstein: three hundred years of science in Berlin]. WJS, Wolf Jobst Siedler, Berlin, Germany, 2010. ISBN 3-937989-60-9 (paperback). 217 pp. LCCN DD866.8 .E73 2010.

Erckmann:1955:RNW

[Erc55] Rudolf Erckmann. Via regia: Nobelpreisträger auf dem Wege ins Atomzeitalter. (German) [The Regal Way: Nobel Prize winners on the road to the Atomic Age]. W. Andermann, München, West Germany, 1955. 399 pp. LCCN Q141 .E69.

Erikson:1982:PRE

[Eri82] Erik H. Erikson. Psychoanalytic reflections on Einstein’s centenary. In Holton and Elkana [HE82], pages 151–174. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Erlichson:1972:EEP

[Erl72a] H. Erlichson. Einstein and the Einstein–Podolsky–Rosen criterion of reality. American Journal of Physics, 40(2):359– 360, 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/40/2/10.1119/1.1986544.

Erlichson:1972:BEP

[Erl72b] Herman Erlichson. Bohr and the Einstein–Podolsky–Rosen paradox. American Journal of Physics, 40(4):634–636, 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/40/4/10.1119/1.1988078.

Ermenc:1989:ABS

[Erm89] Joseph J. Ermenc, editor. Atomic bomb scientists: memoirs, 1939–1945: interviews with Werner Karl Heisenberg, Paul Harteck, Lew Kowarski, Leslie R. Groves, Aristid von Grosse, C. E. Larson. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1989. ISBN 0-88736-267-2. 385 pp. LCCN QC774.A2 A86 1989. US$120.00.

Ernst:1992:LMO

[Ern92] Sabine Ernst, editor. Lise Meitner an Otto Hahn: Briefe aus den Jahren 1912 bis 1924: Edition und Kommentierung. (German) [Lise Meitner to Otto Hahn: Correspondence from the years 1912 to 1924: editing and commentary], volume 65 of Quellen und Studien zur Geschichte der Pharmazie. Wissenschaftliche Verlagsgesellschaft mbH, Stuttgart, Germany, 1992. ISBN 3-8047-1254-1. vi + 267 pp. LCCN RS61 .Q8 Bd. 65.

Einstein:1913:EAA

[ES13] Albert Einstein and Otto Stern. Einige Argumente für die Annahme einer molekular Agitation beim absoluten Nullpunkt. (German) [Some arguments for the acceptance of molecular motion at absolute zero]. Annalen der Physik (1900) (series 4), 345 (3):551–560, 1913. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Einstein:1930:R

[ES30] Albert Einstein and Leo Szilard. Refrigeration. US Patent 1,781,541., November 11, 1930. URL http://www.google.com/ patents?id=t0BRAAAAEBAJ. Application filed December 16, 1927 (serial number 240,566) and in Germany, December 16, 1926. See [Dan97a, Dan97b] for accounts of this invention.

Einstein:1945:IES

[ES45] Albert Einstein and Ernst G. Straus. The influence of the expansion of space on the gravitation fields surrounding the individual stars. Reviews of Modern Physics, 17(2–3):120–124, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/abstract/
RMP/v17/p120;   http://prola.aps.org/pdf/RMP/v17/i2-

3/p120_1. See corrections and additional remarks [ES46a].

Einstein:1946:CAR

[ES46a] A. Einstein and E. G. Straus. Corrections and additional remarks to our paper: “The influence of the expansion of space on the gravitation fields surrounding the individual stars”. Reviews of Modern Physics, 18(1):148–149, January 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/abstract/RMP/v18/p148.

See [ES45].

Einstein:1946:GRT

[ES46b] A. Einstein and E. G. Straus. A generalization of the relativistic theory of gravitation. II. Annals of Mathematics (2), 47(4):731–741, October 1946. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL 2-E; http:

//links.jstor.org/sici?sici=0003-486X(194610)2:47:4<
731:AGOTRT>2.0.CO.

Einstein:1956:LMS

[ES56] Albert Einstein and Maurice Solovine. Lettres à Maurice Solovine. Gauthier-Villars, Paris, France, 1956. xiv + 140 pp. LCCN QC 16 EIN. With an Introduction by Maurice Solovine.

Einstein:1968:BBG

[ES68] Albert Einstein and Arnold Sommerfeld. Briefwechsel. 60 Briefe aus dem goldenen Zeitalter der modernen Physik. (German) [Letter exchanges: 60 letters from the Golden Age of modern physics]. Schwabe, Stuttgart, Germany and Basel, Switzerland, 1968. 126 + 2 pp. LCCN QC16.E5 A43. With commentary by Armin Hermann.

Einstein:2005:EAA

[ES05a] A. Einstein and O. Stern. Einige Argumente für die Annahme einer molekularen Agitation beim absoluten Nullpunkt [AdP 40, 551 (1913)]. (German) [Some arguments for assuming molecular agitation at absolute zero [AdP 40, 551 (1913)]]. Annalen der Physik (Berlin), 14(1):489–499, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590041.

Einstein:2005:MW

[ES05b] Albert Einstein and Carl Seelig. Mein Weltbild. (German) [My World View], volume 36728 of Ullstein. Ullstein, Berlin, Germany, 2005. ISBN 3-548-36728-3. 231 pp. LCCN

???? URL http://www.gbv.de/dms/hebis-darmstadt/toc/
127386025.pdf.

ESA:2015:SSA

[ESA15] ESA. Satellites set for ambitious test of Einstein’s most famous theory. Scientific Computing, ??(??):??, November 11, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (elec-

tronic). URL http://www.scientificcomputing.com/news/
2015/11/satellites-set-ambitious-test-einsteins-most-
famous-theory.

Eshach:2009:NPP

[Esh09] Haim Eshach. The Nobel Prize in the physics class: Science, history, and glamour. Science & Education (Springer), 18(10): 1377–1393, October 2009. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Espinosa:1934:MES

[Esp34] Enrique Espinosa. El mundo de Einstein. (Spanish) [the world of Einstein]. La Nación (Literary Supplement) [Buenos Aires??], 3(1):??, September 16, 1934.

Espinosa:1955:MES

[Esp55] Enrique Espinosa. El mundo de Einstein, “Mein Weltbild”. (Spanish) [the world of Einstein: “My Worldview”]. Davar, 61(??):71–77, 1955.

Esterson:2006:EWD

[Est06a] Allen Esterson. Einstein’s Wife documentary: List of errors.

Web site, 2006. URL http://www.esterson.org/Einsteins_
Wife_Errors_List.htm.

Esterson:2006:EWM

[Est06b] Allen Esterson. Einstein’s Wife: Mileva Marić. Web site, 2006.

URL http://www.esterson.org/einsteinwife1.htm.

Esterson:2006:CEH

[Est06c] Allen Esterson. Critique of Evan Harris Walker’s letter in Physics Today, February 1991. Web site, 2006. URL

http://www.esterson.org/Walker_Physics_Today.htm. See

[Wal91b].

Eisen:1987:NRS

[ET87] Jonathan Eisen and Stuart Troy, editors. The Nobel reader: short fiction, poetry, and prose by Nobel laureates in literature. C. N. Potter, New York, NY, USA, 1987. ISBN 0-517-56351-7 (paperback). x + 338 pp. LCCN ???? US$12.95.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp. LCCN

QC174.12 .Q346 2007. URL http://www.loc.gov/catdir/
enhancements/fy0824/2006934045-b.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2006934045-d.html;
 http://www.loc.gov/catdir/toc/fy0708/2006934045.html.

Einstein:2020:WSS

[ET20] Albert Einstein and Edward Teller. 1950: What the scientists are saying. Bulletin of the Atomic Scientists, 76(6):350–352, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847477.

Einstein:1931:KPF

[ETP31] Albert Einstein, Richard C. Tolman, and Boris Podolsky. Knowledge of past and future in quantum mechanics. Physical Review (2), 37(6):780–781, March 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/abstract/PR/v37/p780;   http://prola.
aps.org/pdf/PR/v37/i6/p780_1.

EONR:1969:CRS

[Eur69] European Organization for Nuclear Research . Catalogue of reprints of scientific papers in the Pauli collection, volume 3 of CERN bibl. CERN, Geneva, Switzerland, 1969. v + 470 pp. LCCN Z7145 .E87.

Evans:1953:SWB

[Eva53] Medford Evans. The secret war for the A-bomb. H. Regnery Co., Chicago, IL, USA, 1953. xviii + 302 pp. LCCN HD9698.A28 E8. Introduction by James Burnham.

Evans:1995:FRR

[Eva95] M. W. Evans. The B^{(3)} field: Its role in the Rayleigh–Jeans law, Planck law, and Einstein coefficients. Foundations of Physics, 25 (2):383–389, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02055215.

Einstein:1972:JKA

[EVB^{+}72] Albert Einstein, I. N. Vesselovsky, Yu. A. Belyj, R. Plechkaitis, I. M. Rabinovich, Yu. Kh. Kopelevich, M. E. Glinka, and E. S. Shchukina. Johann Kepler (to the 400th anniversary since J. Kepler’s birthday). Istoriko-Astronomicheskie Issledovaniya, 11 (??):15–177, 1972. ISSN 0234-8632.

Eve:1939:RBL

[Eve39] A. S. (Arthur Stewart) Eve. Rutherford: being the life and letters of the Rt. Hon. Lord Rutherford, O.M. Macmillan Publishing Company, New York, NY, USA, 1939. xvi + 451 + 17 pp. LCCN QC16.R8 E9 1939a.

Eve:2013:RBL

[Eve13] A. S. (Arthur Stewart) Eve. Rutherford: being the life and letters of the Rt. Hon. Lord Rutherford, O.M. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-67881-1 (paperback). xvi + 451 pp. LCCN ????

Enz:1994:WPW

[EvM94] Charles P. (Charles Paul) Enz and K. von Meyenn. Wolfgang Pauli: Writings on physics and philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). vi + 289 pp. LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Ellis:1988:FCS

[EW88] George F. R. (George Francis Rayner) Ellis and Ruth M. Williams. Flat and curved space–times. Clarendon Press, Oxford, UK, 1988. ISBN 0-19-851169-8 (paperback), 0-19-851164-7 (hardcover). x + 351 pp. LCCN QC173.65

.E45 1988. URL http://www.loc.gov/catdir/enhancements/
fy0639/87026340-d.html;   http://www.loc.gov/catdir/
enhancements/fy0639/87026340-t.html.

Ellis:2000:FCS

[EW00] George F. R. (George Francis Rayner) Ellis and Ruth M. Williams. Flat and curved space–times. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2000. ISBN 0-19-850657-0 (hardcover), 0-19-850656-2 (paperback). x + 375 pp. LCCN QC173.65 .E45 2000. URL http://www.loc.

gov/catdir/enhancements/fy0611/2001268456-d.html;
http://www.loc.gov/catdir/enhancements/fy0611/2001268456-

t.html. Diagrams by Mauro Carfora.

Einstein:1925:RPR

[EWJ^{+}25] Albert Einstein, A. N. Whitehead, G. B. Jeffery, F. D. Murnaghan, and W. Perrett. Recent publications: Reviews: Sidelights on Relativity // The Principle of Relativity with Applications to Physical Science. American Mathematical Monthly, 32 (6):311–313, 1925. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ezawa:1979:ECSa

[Eza79a] Hiroshi Ezawa. Einstein’s contribution to statistical mechanics. In Aichelburg and Sexl [AS79], pages 69–87. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_7/.

Ezawa:1979:ECSb

[Eza79b] Hiroshi Ezawa. Einstein’s contributions to statistical mechanics, classical and quantum. Japanese Studies in the History of Science, 18(??):27–72, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Ezrahi:1982:ELR

[Ezr82] Yaron Ezrahi. Einstein and the light of reason. In Holton and Elkana [HE82], pages 253–280. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Ezzell:1987:EM

[Ezz87] Carol Ezzell. Einstein manuscript. Nature, 330(6149):597, December 23, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v330/n6149/pdf/330597c0.pdf.

Frangsmyr:1993:La

[FA93a] Tore Frängsmyr and Sture Allén, editors. Literature 1968–1980. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1174-0, 981-02-1175-9 (paperback). x + 197 pp. LCCN ????

Frangsmyr:1993:Lb

[FA93b] Tore Frängsmyr and Sture Allén, editors. Literature 1981–1990. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1176-7, 981-02-1177-5 (paperback). x + 166 pp. LCCN ????

Frangsmyr:1997:P

[FA97] Tore Frängsmyr and Irwin Abrams, editors. Peace. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-1178-3 (vol. 1, hardcover), 981-02-1179-1 (vol. 1, paperback), 981-02-1180-5 (vol. 2, hardcover), 981-02-1181-3 (vol. 2, paperback). ???? pp. LCCN JX1962.A2 P37 1997.

Fabian:1979:FRE

[Fab79] A. C. Fabian. First results from the Einstein Observatory. Nature, 279(5712):371–372, May 31, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v279/n5712/
pdf/279371a0.pdf.

College:1982:MU

[Fac82] Faculty of Lynchburg College, editor. Man and the Universe, volume 1 (series 2) of Classical selections on great issues. University Press of America, Washington, DC, USA, 1982. ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). xvi + 370 pp. LCCN Q125 .M316 1982.

Falk:2009:NIA

[Fal09] Dean Falk. New information about Albert Einstein’s brain. Frontiers in evolutionary neuroscience, 1(??):3, ???? 2009. ISSN 1663-070X.

Fant:1991:ABN

[Fan91] Kenne Fant. Alfred Bernhard Nobel. Norstedt, Stockholm, Sweden, 1991. ISBN 91-1-919072-7. 423 + 32 pp. LCCN TP268.5.N7 F36 1991.

Fant:1993:ANB

[Fan93] Kenne Fant. Alfred Nobel: a biography. Arcade, New York, NY, USA, 1993. ISBN 1-55970-222-2. viii + 342 + 16 pp. LCCN TP268.5.N7 F3613 1993. US$24.95.

Fantova:2003:IGE

[Fan03] Johanna Fantova. Introduction to “Gespräche mit Einstein”. Princeton University Library Chronicle, 65(1):51–64, Autumn

2003. ISSN 0032-8456. URL https://www.jstor.org/stable/
10.25290/prinunivlibrchro.65.1.0051.

Fang:2006:CAE

[Fan06] L. Z. Fang. China and Albert Einstein: The reception of the physicist and his theory in China, 1917–1979. China Journal, 55(??):211–212, January 2006. ISSN 1324-9347.

Farber:1953:NPW

[Far53] Eduard Farber. Nobel Prize winners in chemistry, 1901–1950, volume 31 of The Life of science library. H. Schuman, New York, NY, USA, 1953. 219 pp. LCCN QD21 .F21n 1953.

Farber:1963:NPW

[Far63] Eduard Farber. Nobel Prize winners in chemistry, 1901–1961, volume 41 of The Life of science library. Abelard-Schuman, London, UK, revised edition, 1963. vii + 341 pp. LCCN QD21 .F37 1963.

Farago:1979:BIR

[Far79] P. S. Farago. Born’s intimate reminiscences: Book review: My Life: Recollections of a Nobel Laureate, by Max Born. Nature, 282(5734):113, November 1, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v282/n5734/
pdf/282113a0.pdf.

Farmelo:2002:IMB

[Far02] Graham Farmelo, editor. It must be beautiful: great equations of modern science. Granta, London, UK, 2002. ISBN 1-86207-479-8. xvi + 283 pp. LCCN Q125 .I88 2002.

Farmelo:2003:IMB

[Far03a] Graham Farmelo, editor. It Must Be Beautiful: Great Equations of Modern Science. Granta, London, UK, 2003. ISBN 1-86207-555-7. xviii + 284 pp. LCCN Q125 .I88 2003.

Farmelo:2003:RNR

[Far03b] Graham Farmelo. A revolution with no revolutionaries: the Planck–Einstein equation for the energy of a quantum. In It Must Be Beautiful: Great Equations of Modern Science [Far03a], pages 1–27. ISBN 1-86207-555-7. LCCN Q125 .I88 2003.

Fara:2005:MTA

[Far05a] P. Fara. The maestro of time (Albert Einstein). History Today, 55(4):28–33, April 2005. ISSN 0018-2753.

Fara:2005:ME

[Far05b] Patricia Fara. Monuments to Einstein. Endeavour, 29(2):58– 59, June 2005. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Farrell:2005:DYL

[Far05c] John Farrell. The day without yesterday: Lemaître, Einstein, and the birth of modern cosmology. Thunder’s Mouth Press, New York, NY, USA, 2005. ISBN 1-56025-660-5, 1-56025-902-7 (paperback). viii + 262 pp. LCCN QB991.E94

F37 2005. URL http://www.loc.gov/catdir/enhancements/
fy0831/2006272995-d.html.

Farmelo:2009:SMH

[Far09] Graham Farmelo. The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01827-0. 539 + 8 pp. LCCN QC16.D57; QC16.D57 F37 2009. US$29.95. URL http:

//bnreview.barnesandnoble.com/t5/The-Thinking-Read/
The-Strangest-Man-The-Hidden-Life-of-Paul-Dirac-Mystic-
of-the/ba-p/1243.

Farmelo:2013:CBHa

[Far13a] Graham Farmelo. Churchill’s bomb: a hidden history of science, war and politics. Faber and Faber, London, UK, 2013. ISBN 0-571-24978-7 (hardcover). x + 554 + 12 pp. LCCN ????

Farmelo:2013:CBHb

[Far13b] Graham Farmelo. Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race. Basic Books, New York, NY, USA, 2013. ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). vi + 554 pp. LCCN UA647 .F28 2013.

Faure:1979:THB

[Fau79] J. P. Faure. Time from Henri Bergson to Albert Einstein: Relationship between Man and Cosmos. Europe — Revue Littéraire Mensuelle, 57(602–??):191–204, ???? 1979. ISSN 0014-2751.

Fayer:2010:ASH

[Fay10] Michael D. Fayer. Absolutely small: how quantum theory explains our everyday world. AMACOM, New York, NY, USA, 2010. ISBN 0-8144-1488-5. ix + 383 pp. LCCN QC174.12 .F379 2010.

Fazarinc:2015:FDB

[Faz15] Zvonko Fazarinc. Fermi–Dirac, Bose–Einstein, Maxwell– Boltzmann, and computers. Computer Applications in Engineering Education, 23(5):746–759, September 2015. CO-DEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (elec-

tronic). URL https://dblp.org/db/journals/caee/caee23.
html#Fazarinc15;   https://www.wikidata.org/entity/
Q60032924.

Freundlich:1919:FET

[FB19] Erwin Freundlich and Henry Herman Leopold Adolf Brose, editors. The foundations of Einstein’s theory of gravitation. E. P. Dutton, New York, NY, USA, 1919. xvi + 140 pp. LCCN QC6 .F73.

Freundlich:1920:FET

[FB20] Erwin Freundlich and Henry L. (Henry Leopold) Brose. The foundations of Einstein’s theory of gravitation. The University Press, Cambridge, UK, 1920. xvi + 60 + 2 pp. LCCN QC178 .F7 1920; QC178 .F7.

Freundlich:1922:FET

[FB22] Erwin Freundlich and Henry L. (Henry Leopold) Brose. The foundations of Einstein’s theory of gravitation. G. E. Stechert, New York, NY, USA, 1922. xvi + 60 pp. LCCN QC6 .F88.

Freundlich:1924:TRT

[FB24a] Erwin Freundlich and Henry Herman Leopold Adolf Brose. The Theory of Relativity: three lectures for chemists. Methuen and Co. Ltd., London, UK, 1924. xii + 98 pp. LCCN QC6 .F889; 530 F88; QC6 .F7; QC173.58 .F7; QC6 .F77. With an introduction by Viscount Haldane.

Freundlich:1924:FET

[FB24b] Erwin Freundlich and Henry L. (Henry Leopold) Brose. The foundations of Einstein’s theory of gravitation. Methuen and Co. Ltd., London, UK, second revised and enlarged edition, 1924. xvi + 140 pp. LCCN QC6 .F73 1924; QC178 .F89gE 1924a.

Feynman:2005:FTN

[FBD05] Richard P. (Richard Phillips) Feynman, Laurie M. Brown, and P. A. M. (Paul Adrien Maurice) Dirac. Feynman’s thesis: a new approach to quantum theory. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-256-366-0, 981-256-380-6 (paperback). xxii + 119 pp. LCCN QC174.12 .F48 2005.

Fawcett:1987:RB

[FC87] Carolyn R. Fawcett and Robert S. Cohen. Reason and Being, volume 17 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346. xx + 441 pp. LCCN

Q174 .B67 vol. 17. URL https://link.springer.com/book/
10.1007/978-94-009-4590-6.

Frantz:2007:MPT

[FC07] Douglas Frantz and Catherine Collins. The Man from Pakistan: The True Story of the World’s Most Dangerous Nuclear Smuggler. Grand Central, New York, NY, USA, 2007. ISBN 0-446-19958-3. xvii + 413 pp. LCCN JZ5675 .F73 2008.

Forman:2011:WCQ

[FCKT11] Paul Forman, Cathryn Carson, Alexei Kojevnikov, and Helmuth Trischler, editors. Weimar culture and quantum mechanics: selected papers by Paul Forman and contemporary perspectives on the Forman thesis. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4293-11-3. LCCN Q174.8 W45 2011. This book has its origin in the conference “The Cultural Alchemy of the Exact Sciences: Revisiting the Forman Thesis”, held at the University of British Columbia, Vancouver, on 23 to 25 March 2007.

Friedman:1985:EMM

[FD85] Alan J. Friedman and Carol C. Donley. Einstein as myth and muse. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26720-X. xiv + 224 pp. LCCN QC16.E5

F76 1985. URL http://www.loc.gov/catdir/description/
cam031/84023839.html.

Fechter:1990:AMH

[Fec90] Egidius Fechter. The Atomkeller-Museum in Haigerloch. Town of Haigerloch, Haigerloch, Germany, 1990. 20 pp.

Feis:1961:JSA

[Fei61] Herbert Feis. Japan subdued: the atomic bomb and the end of the war in the Pacific. Princeton University Press, Princeton, NJ, USA, 1961. vi + 199 pp. LCCN D767.2 .F4.

Feis:1966:ABE

[Fei66] Herbert Feis. The atomic bomb and the end of World War II. Princeton University Press, Princeton, NJ, USA, revised edition, 1966. vi + 213 pp. LCCN D767.2 .F4 1966.

Feinberg:1970:PGF

[Fei70] G. Feinberg. Particles that go faster than light. Scientific American, 222(2):69–77, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fei:2001:NPC

[Fei01] Binhai Fei. Nobel Prize: centennial 1901–2001: laureate stamps and inscriptions. Shanghai Far East Publishers, Shanghai, China, 2001. ISBN 7-80661-379-X. 777 (est.) pp. LCCN AS911.N9 F36 2001. Two volumes.

Feld:1979:EPN

[Fel79] Bernard T. Feld. Einstein and the politics of nuclear weapons. Bulletin of the Atomic Scientists, 35(3):5–16, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Feld:1982:EPN

[Fel82a] Bernard T. Feld. Einstein and the politics of nuclear weapons. In Holton and Elkana [HE82], pages 369–396. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Feld:1982:NPP

[Fel82b] Bernard T. Feld. Nobel Peace Prize: [for 1982 to Alva R. Myrdal and Alfonso Garcia Robles]. Bulletin of the Atomic Scientists, 38(10):81, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feldman:2000:NPH

[Fel00] Burton Feldman. The Nobel Prize: a history of genius, controversy, and prestige. Arcade Pubishing, New York, NY, USA, 2000. ISBN 1-55970-537-X. xi + 489 pp. LCCN AS911.N9 F38 2000.

Feldman:2011:BRB

[Fel11] Bernard J. Feldman. Book review: Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein, by Bruce J. Hunt. American Journal of Physics, 79(6):687, June 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/79/687/
1.

Fermi:1924:CSQ

[Fer24] Enrico Fermi. Considerazioni sulla quantizzazione dei sistemi che contengono degli elementi identici. (Italian) [Considerations on the quantization of systems that contain identical elements]. Il Nuovo Cimento (8), 1(1):145–152, December 1924. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://link.springer.com/article/10.1007/BF02956745;
 http://www.springerlink.com/content/97203r302q8w7unj/
;   http://www.springerlink.com/content/97203r302q8w7unj/
fulltext.pdf.

Fermi:1938:NPP

[Fer38] Enrico Fermi. The Nobel Prize in Physics 1938. Nobelprize.org,

1938. URL http://www.nobelprize.org/nobel_prizes/

physics/laureates/1938/. The Nobel Prize in Physics 1938 was awarded to Enrico Fermi “for his demonstrations of the existence of new radioactive elements produced by neutron irradiation, and for his related discovery of nuclear reactions brought about by slow neutrons”.

Fermi:1939:ARP

[Fer39] Enrico Fermi. Artificial radioactivity produced by neutron bombardment: Nobel lecture delivered at Stockholm, December 12, 1938. In Les Prix Nobel en 1938: Les Conférences Nobel, Stockholm, pages 1–8. Norstedt & Söner, Stockholm, Sweden, 1939. ISSN 0546-8175. Reprinted in [Fer70].

Fermi:1954:AFM

[Fer54] Laura Fermi. Atoms in the family: my life with Enrico Fermi. University of Chicago Press, Chicago, IL, USA and London, UK, 1954. 267 pp. LCCN QC774.F4 F4.

Fermi:1957:AW

[Fer57] Laura Fermi. Atoms for the world. University of Chicago Press, Chicago, IL, USA and London, UK, 1957. ???? pp. LCCN TK9006 .I5 1955j.

Fermi:1961:SAE

[Fer61] Laura Fermi. The story of atomic energy. Landmark books. Random House, New York, NY, USA, 1961. 184 pp. LCCN QC778 .F47.

Fermi:1970:ARP

[Fer70] Enrico Fermi. Artificial radioactivity produced by neutron bombardment: Nobel lecture delivered at Stockholm, December 12, 1938. In Enrico Fermi: physicist [Seg70], chapter A-2, pages 214–221. ISBN 0-226-74472-8. LCCN QC16.F46 S4.

Ferris:1981:AES

[Fer81] T. Ferris. Albert Einstein Special Theory of Relativity — emergence (1905) and early interpretation (1905–1911) A. I. Miller. New York Times Book Review, 86(29):11–12, ???? 1981. ISSN 0028-7806.

Fermi:1982:AFM

[Fer82a] Laura Fermi. Atoms in the family: my life with Enrico Fermi. University of New Mexico Press, Albuquerque, NM, USA, 1982. ISBN 0-8263-1060-5 (paperback). ix + 267 + 24 pp. LCCN QC774.F4 F4 1988.

Ferris:1982:SSP

[Fer82b] T. Ferris. Some strangeness in the proportion — a Centennial symposium to celebrate the achievements of Albert Einstein — H. Woolf. New York Times Book Review, 87(39):18, ???? 1982. ISSN 0028-7806.

Ferris:1982:SLS

[Fer82c] T. Ferris. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. New York Times Book Review, 87(48):9, ???? 1982. ISSN 0028-7806.

Ferris:1983:AEH

[Fer83] T. Ferris. Albert Einstein historical and cultural perspectives — G. Holton, Y. Elkana. Science Digest, 91(6):81, ???? 1983. ISSN 0036-8296.

Fermi:1987:AFM

[Fer87] Laura Fermi. Atoms in the family: my life with Enrico Fermi, volume 9 of The History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1987. ISBN 0-88318-524-5. 267 + 24 pp. LCCN QC16.F46 F47 1987. US$35.00.

Ferris:2003:CAM

[Fer03] Timothy Ferris. Coming of age in the Milky Way. Harper-Collins, New York, NY, USA and London, UK, 2003. ISBN 0-06-053595-4, 0-06-200654-1 (e-book). 510 pp. LCCN

Q125 .F425 2003.  URL http://www.loc.gov/catdir/
enhancements/fy0645/2004273573-d.html;   http://www.
loc.gov/catdir/enhancements/fy0910/2004273573-b.html.

FernandezBuey:2005:AEC

[Fer05a] Francisco Fernández Buey. Albert Einstein: ciencia y conciencia. (Spanish) [Albert Einstein: science and conscience]. Retratos del viejo topo. El Viejo Topo, Barcelona, Spain, 2005. ISBN 84-96356-21-3. 304 pp. LCCN ????

Ferreira:2005:WGW

[Fer05b] Pedro Ferreira. Why Gödel was a hero to Einstein. Physics World, 18(12):44–45, December 2005. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/12/phwv18i12a36.pdf.

Ferreira:2007:BRD

[Fer07a] Pedro G. Ferreira. Book review: David Lindley, Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for The Soul of Science. Physics Today, 60(8):57–??, August 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v60/i8/p57/s1.

Ferreira:2007:SUP

[Fer07b] Pedro G. Ferreira. The state of the universe: a primer in modern cosmology. Phoenix, London, UK, 2007. ISBN 0-7538-2256-3 (paperback). 320 pp. LCCN QB982 FER 2007; Ordered; Y; Ps; M09.G00838.

Fert:2008:NLO

[Fer08] Albert Fert. Nobel Lecture: Origin, development, and future of spintronics. Reviews of Modern Physics, 80(4):1517– 1530, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1517;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1517_1.

Ferreira:2014:PTJ

[Fer14a] Pedro G. Ferreira. Die perfekte Theorie: das Jahrhundert der Genies und der Kampf um die Relativitätstheorie. (German) [The perfect theory: the century of genius and the struggle for the Theory of Relativity]. Verlag C. H. Beck, München, Germany, 2014. ISBN 3-406-66047-9, 3-406-66138-6 (e-book). LCCN QC173.6.

Ferreira:2014:EPT

[Fer14b] Pedro G. Ferreira. Einstein’s perfect theory: General relativity rebooted. New Scientist, ??(??):??, August 4, 2014. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL

http://www.newscientist.com/article/mg22329800.500-
einsteins-perfect-theory-general-relativity-rebooted.
html.

Ferreira:2014:PTC

[Fer14c] Pedro G. Ferreira. The Perfect Theory: a Century of Geniuses and the Battle over General Relativity. Houghton Mifflin Harcourt, Boston, MA, USA, 2014. ISBN 0-547-55489-3 (hardcover), 0-544-26408-8 (paperback), 0-547-55490-7 (e-book), 1-306-36108-7 (e-book). xv + 288 pp. LCCN QC173.6 .F47 2014.

Feshbach:1956:RPM

[Fes56] Herman Feshbach. Review: Physics and Microphysics by Louis de Broglie; Martin Davidson; A. Einstein. Science, 123(3196): 550, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1750554.

Fetter:2009:RTB

[Fet09] Alexander L. Fetter. Rotating trapped Bose–Einstein condensates. Reviews of Modern Physics, 81(2):647–691, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.647;   http://rmp.aps.org/
abstract/RMP/v81/i2/p647_1.

Feuer:1971:SREa

[Feu71a] Lewis S. Feuer, B.Sc., Ph.D. The social roots of Einstein’s Theory of Relativity. Part I. Annals of Science, 27(3):277–298, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1971:SREb

[Feu71b] Lewis S. Feuer, B.Sc., Ph.D. The social roots of Einstein’s Theory of Relativity. Part II. Annals of Science, 27(4):313–344, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1974:EGS

[Feu74] Lewis Samuel Feuer. Einstein and the generations of science. Basic Books, New York, NY, USA, 1974. ISBN 0-465-01871-8. ix + 374 pp. LCCN QC173.98 .F48.

Feuer:1978:ECG

[Feu78] Lewis Samuel Feuer. Einstein et le conflit des générations. (French) [Einstein and the generational conflict]. De la science.

Éditions Complexe, Paris, France, 1978. ISBN 2-87027-023-2. 382 pp. LCCN QC173.98 F4812.

Feuer:1982:EGS

[Feu82] Lewis Samuel Feuer. Einstein and the generations of science. Transaction Books, New Brunswick, NJ, USA, second edition, 1982. ISBN 0-87855-899-3 (paperback). xlviii + 374 + 10 pp. LCCN QC16.E5 F48 1982. US$12.95.

Feynman:1964:RPF

[Fey64] Richard P. Feynman. Richard P. Feynman, The Nobel Prize in Physics 1965: Biography. In Nobel Foundation [Nob72], page ?? ISBN 0-444-40993-9. LCCN QC71.N735. URL http:/

/nobelprize.org/nobel_prizes/physics/laureates/1965/
feynman-bio.html;   http://nobelprize.org/nobel_prizes/
physics/laureates/1965/feynman-lecture.html.

Feynman:1966:DSTa

[Fey66] Richard P. Feynman. The development of the space–time view of quantum electrodynamics. In ????, editor, Les Prix Nobel 1965, pages 172–191. Imprimerie Royale P. A. Norstedt & Söner, Stockholm, Sweden, 1966. LCCN ????

URL http://www.feynmanlectures.info/other/Feynmans_
Nobel_Lecture.pdf.

Feyerabend:1974:ZE

[Fey74] P. K. Feyerabend. Zahar on Einstein. The British Journal for the Philosophy of Science, 25(1):25–28, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
25/1/25.full.pdf+html.

Feyerabend:1980:ZME

[Fey80] Paul Feyerabend. Zahar on Mach, Einstein and modern science. The British Journal for the Philosophy of Science, 31(3):273– 282, September 1980. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/31/3/273.full.pdf+html.

Feyerabend:1984:MTR

[Fey84] Paul K. Feyerabend. Mach’s theory of research and its relation to Einstein. Studies in History and Philosophy of Science Part A, 15(1):1–22, March 1984. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/003936818490027X.

Feynman:1994:CPL

[Fey94] Richard Phillips Feynman. The Character of Physical Law. Modern Library, New York, NY, USA, modern library edition, 1994. ISBN 0-679-60127-9. xx + 167 pp. LCCN QC71 .F44 1994.

Feynman:1995:SEP

[Fey95] Richard Phillips Feynman. Six easy pieces: essentials of physics, explained by its most brilliant teacher. Helix books. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40955-0, 0-201-40956-9 (set), 0-201-48308-4 (cassettes), 0-201-40825-2 (paperback). xxix + 145 pp. LCCN QC21.2 .F52 1995. US$22.00. Originally prepared for publication by Robert B. Leighton and Matthew L. (Matthew Linzee) Sands. New introduction by Paul Davies. See also [Fey97].

Feynman:1997:SEP

[Fey97] Richard Phillips Feynman. Six not-so-easy pieces: Einstein’s Relativity, symmetry, and space–time. Helix books. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-32841-0 (set), 0-201-15025-5 (hardcover), 0-201-32842-9 (paperback), 0-201-31151-8 (disc 1), 0-201-31152-6 (disc 2), 0-201-31153-4 (disc 3), 0-201-31154-2 (disc 4), 0-201-31155-0 (disc 5), 0-201-31156-9 (disc 6). xxvii + 152 pp. LCCN QC793.3.S9 F49 199. Originally prepared for publication by Robert B. Leighton and Matthew L. (Matthew Linzee) Sands. New introduction by Roger Penrose. See also [Fey95].

Feynman:1998:MIA

[Fey98] Richard Phillips Feynman. The meaning of it all: thoughts of a citizen scientist. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-7382-0166-9 (paperback). 133 pp. LCCN Q175.55 .F49 1998.

Feynman:2000:PFT

[Fey00] Richard P. Feynman. The pleasure of finding things out: the best short works of Richard P. Feynman. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0349-1 (paperback). xvi + 270 pp. LCCN Q171 .F385 1999. Edited by Jeffrey Robbins, and foreword by Freeman Dyson.

Feynman:2005:PRD

[Fey05] Michelle Feynman, editor. Perfectly Reasonable Deviations (from the Beaten Track): the Collected Letters of Richard P. Feynman. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0636-9. xxiii + 486 pp. LCCN QC16 .F49

A4 2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/toc/ecip055/

2005000049.html. Foreword by Timothy Ferris. See also interview with the editor [Ano05m].

Feynman:2006:QST

[Fey06] Richard P. (Richard Phillips) Feynman. QED: the strange theory of light and matter. Princeton science library, Alix G. Mautner memorial lectures. Princeton University Press, Princeton, NJ, USA, expanded Princeton science library edition, 2006. ISBN 0-691-12717-4, 0-691-12575-9 (paperback). xxiv + 158 pp. LCCN QC793.5.P422 F48 2006.

Feynman:2011:SEP

[Fey11] Richard P. Feynman. Six not-so-easy pieces: Einstein’s relativity, symmetry, and space–time. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02526-9 (print), 0-465-02528-5 (e-book). xxvii + 154 pp. LCCN QC793.3.S9 F49 2011.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Frangsmyr:1993:C

[FF93] Tore Frängsmyr and Sture Forsén, editors. Chemistry: 1971– 1980. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-0786-7, 981-02-0787-5 (paperback). 447 pp. LCCN ????

Faye:1994:NBC

[FF94] Jan Faye and Henry J. Folse, editors. Niels Bohr and Contemporary Philosophy, volume 153 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. xxvii + 377 pp. LCCN Q174 .B67 vol. 153 1994; QC16

1994. URL http://www.loc.gov/catdir/enhancements/
fy0823/93024825-d.html;   http://www.loc.gov/catdir/
enhancements/fy0823/93024825-t.html;   https://link.
springer.com/book/10.1007/978-94-015-8106-6.

Feynman:1995:ARP

[FF95] Richard Phillips Feynman and Michelle Feynman. The art of Richard P. Feynman: images by a curious character. GB Science Publishers SA, Basel, Switzerland, 1995. ISBN 2-88449-047-7. 173 pp. LCCN NC1075.F44 A4 1995.

Franke:1997:PLA

[FF97] Almut Franke and Fabian Franke. Paul Langevin und Albert Einstein — eine Freundschaft zwischen Relativitätstheorie und politischer Realität. (German) [Paul Langevin and Albert Einstein — a friendship between Relativity and political reality]. Berichte zur Wissenschaftsgeschichte, 20(2–3):199–215, ???? 1997. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Frize:2009:BBH

[FFF09] Monique Frize, Peter R. D. Frize, and Nadine Faulkner. The bold and the brave: a history of women in science and engineering. University of Ottawa Press, Ottawa, ON, Canada, 2009. ISBN 0-7766-0725-1. xvi + 348 pp. LCCN Q130 .F765 2009.

Fehner:2000:ONT

[FG00] Terrence R. Fehner and Francis G. (Francis George) Gosling. Origins of the Nevada Test Site. U.S. Department of Energy, Washington, DC, USA, 2000. vi + 95 pp. LCCN TK9024.N3 F44 2000.

Frey:2020:STG

[FG20] Bruno S. Frey and Anthony Gullo. Sic transit gloria mundi: What remains of famous economists after their deaths? Scientometrics, 123(1):283–298, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/s11192-020-03393-w.

Farouki:2021:EBD

[FG21] Nayla Farouki and Philippe Grangier. The Einstein–Bohr debate: Finding a common ground of understanding? Foundations of Science, 26(1):97–101, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09716-7.

Feynman:2006:FTP

[FGL^{+}06] Richard P. (Richard Phillips) Feynman, Michael A. Gottlieb, Ralph Leighton, Matthew L. (Matthew Linzee) Sands, Robert B. Leighton, and Rochus Vogt. Feynman’s tips on physics: a problem-solving supplement to the Feynman Lectures on Physics. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-8053-9063-4 (paperback). xiii + 162 pp. LCCN QC23

.F47 1989 Suppl. URL http://www.loc.gov/catdir/toc/
ecip0512/2005013077.html.

Feiveson:2014:UBF

[FGMvH14] Harold A. Feiveson, Alexander Glaser, Zia Mian, and Frank N. von Hippel. Unmaking the bomb: a fissile material approach to nuclear disarmament and nonproliferation. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02774-7 (hardcover), 0-262-31918-7 (e-book). x + 4 + 277 pp. LCCN JZ5675 .F45 2014.

Frauendiener:2006:ANA

[FGPP06] J. (Jörg) Frauendiener, D. (Domenico) Giulini, Volker Perlick, and Roger Penrose, editors. Analytical and numerical approaches to mathematical relativity, volume 692 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-31027-4, 3-540-33484-X (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). xvii + 278 pp. LCCN QC173.59.M3 A63 2006.

Feynman:1965:QMP

[FH65] Richard P. (Richard Phillips) Feynman and Albert R. Hibbs. Quantum mechanics and path integrals. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1965. xiv + 365 pp. LCCN QC174.1 .F39.

Fritzsch:2002:CSN

[FH02] Harald Fritzsch and Karin Heusch. The curvature of spacetime: Newton, Einstein, and gravitation. Columbia University Press, New York, NY, USA, 2002. ISBN 0-231-11820-1 (hardcover), 0-231-11821-X (paperback). x + 341 pp. LCCN QC173.6 .F7513

Finster:2010:QOC

[FH10] Felix Finster and Christian Hainzl. Quantum oscillations can prevent the big bang singularity in an Einstein–Dirac cosmology. Foundations of Physics, 40(1):116–124, January 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
009-9380-z.

Findlay:2011:AFD

[FH11] John M. Findlay and Bruce William Hevly. Atomic frontier days: Hanford and the American West. Emil and Kathleen Sick series in Western history and biography. Center for the Study of the Pacific Northwest in association with University of Washington Press, Seattle, WA, USA, 2011. ISBN 0-295-99097-X (paperback). xv + 368 pp. LCCN TD898.12.W2 F56 2011.

Feynman:2010:QMP

[FHS10] Richard P. (Richard Phillips) Feynman, Albert R. Hibbs, and Daniel F. Styer. Quantum mechanics and path integrals. Dover, New York, NY, USA, 2010. ISBN 0-486-47722-3. xii + 371 pp. LCCN QC174.12 .F484 2010. URL http://www.loc.gov/

catdir/enhancements/fy1006/2010004550-d.html.

Field:1988:KGC

[Fie88] Judith Veronica Field. Kepler’s geometrical cosmology. University of Chicago Press, Chicago, IL, USA and London, UK, 1988. ISBN 0-226-24823-2. xx + 243 pp. LCCN QB981

.F47 1988. URL http://www.loc.gov/catdir/enhancements/
fy0608/87010837-d.html.

Fiege:2007:ASS

[Fie07] Mark Fiege. The atomic scientists, the sense of wonder, and the bomb. Environmental History, 12(3):578–613, July 2007. ISSN 1084-5453 (print), 1930-8892 (electronic). URL http://www.

jstor.org/stable/25473133.

Fields:2015:CEC

[Fie15] Chris Fields. Close to the edge: co-authorship proximity of Nobel laureates in Physiology or Medicine, 1991–2010, to crossdisciplinary brokers. Scientometrics, 103(1):267–299, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-015-1526-5.

Fields:2016:NNT

[Fie16] Chris Fields. Nobel numbers: Time-dependent centrality measures on coauthorship graphs. Journal of the Association for Information Science and Technology, 67(9):2212–2222, September 2016. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fine:1976:YEO

[Fin76] Arthur Fine. The young Einstein and the old Einstein. In Essays in Memory of Imre Lakatos [CFW76], chapter 12, pages 145–159. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-9_12.

Fine:1981:ECQ

[Fin81] Arthur Fine. Einstein’s critique of quantum theory: The roots and significance of EPR. In Barker and Shugart [BS81], page ?? ISBN 0-87870-095-1. LCCN QC16.E5 A63. US$24.95.

Fine:1986:ER

[Fin86a] Arthur Fine. Einstein’s realism. In The shaky game: Einstein, realism, and the quantum theory [Fin86b], pages 86–111. ISBN 0-226-24946-8. LCCN QC6 .F54 1986.

Fine:1986:SGE

[Fin86b] Arthur Fine. The shaky game: Einstein, realism, and the quantum theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1986. ISBN 0-226-24946-8. xi + 186 pp. LCCN QC6 .F54 1986.

Findlay:1990:NDP

[Fin90a] Trevor Findlay. Nuclear dynamite: the peaceful nuclear explosions fiasco. Brassey’s Australia/Pergamon Press, Sydney, NSW, Australia, 1990. ISBN 0-08-034436-4. xv + 339 pp. LCCN TA748.

Fine:1990:EER

[Fin90b] A. Fine. Einstein and ensembles: Response. Foundations of Physics, 20(8):967–989, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00738375.

Fine:1993:EIQ

[Fin93] Arthur Fine. Einstein’s interpretations of the quantum theory. Science in Context, 6(1):257–273, March 1993. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fine:1997:SGE

[Fin97] Arthur Fine. The shaky game: Einstein, realism, and the quantum theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, second edition, 1997. ISBN 0-226-24948-4 (clothbound), 0-226-24949-2 (paper). xi + 220 pp. LCCN QC6 .F54 1996; QC6 .F54 1997; QC6 .F541 1996.

Finegold:2005:AEM

[Fin05] L. Finegold. Albert Einstein to Michele Besso. Physics Today, 58 (7):16, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Fine:2012:SG

[Fin12] Arthur Fine. The Shaky Game. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-92326-6, 1-280-12658-2. 233 pp. LCCN QC6 .F54 1996.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=878251;   http://www.myilibrary.com?id=353044;
 http://www.myilibrary.com?id=353044\%26ref=toc.

Fischer:1979:RJB

[Fis79] Jan Fischer. Review of Jiří Bičák, Einstein a Praha. Československý Časopis pro Fyziku, A29(??):296–297, ???? 1979.

Fisher:1985:AE

[Fis85] Phyllis Fisher. Los Alamos experience. Japan Publications, Tokyo, Japan, 1985. ISBN 0-87040-623-X. 266 pp. LCCN F804.L6 F5 1985.

Fishman:1992:FMF

[Fis92] Sylvia Barack Fishman, editor. Follow my footprints: changing images of women in American Jewish fiction. The Brandeis series in American Jewish history, culture, and life. University Press of New England [for] Brandeis University Press, Hanover, NH, USA, 1992. ISBN 0-87451-544-0, 0-87451-583-1 (paperback). xv + 506 pp. LCCN PS648.J4 F65 1992.

Fischer:2002:WHS

[Fis02] Ernst Peter Fischer. Werner Heisenberg: das selbstvergessene Genie; mit einem Nachtrag zur Taschenbuchausgabe; mit einer Tabelle. (German) [Werner Heisenberg: the forgotten genius; with a supplement to the paperback and a table], volume 3701 of Serie Piper. Piper, München, Germany, Ungekürzte Taschenbuchausgabe edition, 2002. ISBN 3-492-23701-0. 287 pp. LCCN ???? EUR 9.90.

Fischer:2010:HQE

[Fis10] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2011:EB

[Fis11] Ernst Peter Fischer. Einstein: Basics. Piper, München, Germany, 2011. ISBN 3-492-27332-7. 128 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Francis-Jones:2005:EPC

[FJTS05] Anthony Francis-Jones, John Taylor, and Francis Smith. Einstein and physics in the 21st century. Physics World, 18(3): 18, March 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/3/phwv18i3a21.pdf.

French:1968:AWN

[FK68] Warren French and Walter E. Kidd, editors. American winners of the Nobel Literary Prize. University of Oklahoma Press, Norman, OK, USA, 1968. vii + 248 pp. LCCN PS121 .F65.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Farwell:1990:EAN

[FK90] Ruth Farwell and Christopher Knee. The end of the absolute: a Nineteenth-Century contribution to General Relativity. Studies in History and Philosophy of Science Part A, 21(1):91– 121, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368190900162.

Fox:2004:EZ

[FK04] Karen C. Fox and Aries Keck. Einstein: A to Z. Wiley, New York, NY, USA, 2004. ISBN 0-471-46674-3 (paperback). x + 310 pp. LCCN QC16.E5 F68 2004. URL http://

catdir.loc.gov/catdir/bios/wiley047/2004003016.html;
 http://catdir.loc.gov/catdir/description/wiley041/
2004003016.html;   http://catdir.loc.gov/catdir/enhancements/
fy0616/2004003016-b.html;   http://catdir.loc.gov/catdir/
enhancements/fy0616/2004003016-d.html;   http://catdir.
loc.gov/catdir/toc/ecip0415/2004003016.html;   http:
//issuu.com/ammar.friendzzz/docs/a-z__a_biography_of_
albert_einstein.

Fujita:1991:BEC

[FKZ91] S. Fujita, T. Kimura, and Y. Zheng. On the Bose–Einstein condensation of free relativistic bosons with or without mass. Foundations of Physics, 21(9):1117–1130, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733389.

Franck:1949:RII

[FL49] James Franck and Robert Livingston. Remarks on intra- and inter-molecular migration of excitation energy. Reviews of Modern Physics, 21(3):505–509, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.505;
http://rmp.aps.org/abstract/RMP/v21/i3/p505_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Feynman:1988:WDY

[FL88] Richard Phillips Feynman and Ralph Leighton. What do YOU care what other people think?: Further adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 1988. ISBN 0-393-02659-0. 255 pp. LCCN QC16.F49 A3 1988. US$17.95.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: all the adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp. LCCN QC16.F49 A3 2006. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/toc/ecip0515/2005018928.html.

Flatau:2015:WAA

[Fla15] Elke Flatau. Der wissenschaftliche Autor: Aspekte seiner Typologisierung am Beispiel von Einstein, Sauerbruch, Freud und Mommsen. (German) [The scientific author: aspects of its typology using the example of Einstein, Sauerbruch, Freud and Mommsen]. Springer VS, Wiesbaden, Germany, 2015. ISBN 3-658-08140-6 (hardcover), 3-658-08141-4 (e-book). xvi + 506 pp. LCCN PN146 .F575 2015. Thesis (doctoral) — Universität, Mainz, Germany 2014.

Flexner:1933:PEI

[Fle33] Abraham Flexner. Professor Einstein and the Institute for Advanced Study. Science, 78(2019):214, September 8, 1933. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/78/2019/214.1.
full.pdf.

Fletcher:2018:ESB

[Fle18] Seth Fletcher. Einstein’s Shadow: a Black Hole, a Band of Astronomers, and the Quest to See the Unseeable. Ecco, New York, NY, 2018. ISBN 0-06-231204-9 (paperback), 0-06-231202-2 (hardcover), 0-06-231203-0 (e-book). xxvii + 255 pp. LCCN QB843.B55 F595 2018.

Fletcher:2019:RGR

[Fle19] Samuel C. Fletcher. On the reduction of general relativity to Newtonian gravitation. Studies in History and Philosophy of Modern Physics, 68(??):1–15, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301308.

Feynman:1985:SYJ

[FLH85] Richard P. Feynman, Ralph Leighton, and Edward Hutchings. “Surely You’re Joking, Mr. Feynman!”: Adventures of a Curious Character. Bantam Books, New York, NY, USA, 1985. ISBN 0-553-25649-1. xi + 322 pp. LCCN QC16.F49A37 1986.

Falk:2013:CCA

[FLN13] Dean Falk, Frederick E. Lepore, and Adrianne Noe. The cerebral cortex of Albert Einstein: a description and preliminary analysis of unpublished photographs. Brain, 136(??):1304–1327, April 2013. ISSN 0006-8950.

Flores:1998:ED

[Flo98] Francisco Flores. Einstein’s 1935 derivation of E = mc^{2}. Studies in History and Philosophy of Modern Physics, 29(2):223–243, June 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000070.

Flores:1999:ETT

[Flo99] Francisco Flores. Einstein’s theory of theories and types of theoretical explanation. International Studies in the Philosophy of Science, 13(2):123–134, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.

tandfonline.com//doi/abs/10.1080/02698599908573613.

Flores:2005:IEE

[Flo05] Francisco Flores. Interpretations of Einstein’s equation E = mc^{2}. International Studies in the Philosophy of Science, 19(3): 245–260, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-

9281 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02698590500462257. See reply [?] and response

[?].

Flores:2006:IER

[Flo06] Francisco Flores. On the interpretation of the equation E = mc^{2}: response to W. Krajewski: “On the interpretation of the equation E = mc^{2}: reply to F. Flores” [Int. Stud. Philos. Sci. 20 (2006), no. 2, 215–216; MR2253243]. International Studies in the Philosophy of Science, 20(2):217–218, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02698590600814480. See [Kra06].

Florides:2008:JLS

[Flo08] Petros S. Florides. John Lighton Synge. 23 March 1897–30 March 1995. Biographical Memoirs of Fellows of the Royal Society, 54:401–424, December 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Feynman:1963:FLP

[FLS65] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Addison-Wesley, Reading, MA, USA, 1963–1965. ???? pp. LCCN QC21.2 .F49 1963. Three volumes.

Feynman:1989:FLP

[FLS89] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-51003-0 (v. 1). ???? pp. LCCN QC21.2 .F49 1989. Three volumes.

Feynman:2006:FLP

[FLS06] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Pearson/Addison-Wesley, San Francisco, CA, USA, definitive edition, 2006. ISBN 0-8053-9046-4 (vol. 1), 0-8053-9047-2 (vol. 2), 0-8053-9049-9 (vol. 3). ???? pp. LCCN QC21.2 .F49 2006.

Flugge:1939:KEA

[Flü39] Siegfried Flügge. Kann der Energieinhalt der Atomkerne technisch nutzbar gemacht werden?. (German) [Can the energy content of atomic nuclei be made technically available?]. Die Naturwissenschaften, 27(23–24):402–410, June 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
g55p220623l6813n/.

Fluckiger:1974:AEB

[Flü74] Max Flückiger. Albert Einstein in Bern: das Ringen um ein neues Weltbild: eine dokumentarische Darstellung über den Aufstieg eines Genies. (German) [Albert Einstein in Bern: the struggle for a new world view: a documentary representation of the rise of a genius]. Verlag Paul Haupt, Bern, Switzerland, 1974. ISBN 3-258-01341-6. 219 pp. LCCN QC16.E5 F56.

Frangsmyr:1992:NLC

[FM92] Tore Frängsmyr and B. G. (Bo G.) Malmström, editors. Nobel lectures, chemistry: including presentation speeches and laureates’ biographies. –1990, volume 1981. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992. ISBN 981-02-0788-3, 981-02-0789-1 (paperback). x + 708 pp. LCCN ???? Published for the Nobel Foundation.

Freedman:2010:HC

[FM10] Wendy L. Freedman and Barry F. Madore. The Hubble constant. Annual Review of Astronomy and Astrophysics, 48:673– 710, 2010. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-

4282 (electronic). URL http://www.annualreviews.org/doi/
abs/10.1146/annurev-astro-082708-101829.

Franca:1996:MET

[FMM96] H. M. França, A. Maia, Jr., and C. P. Malta. Maxwell electromagnetic theory, Planck’s radiation law, and Bose–Einstein statistics. Foundations of Physics, 26(8):1055–1068, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.
1007/BF02061403.

Fox:1990:NLM

[FMR90] Daniel M. Fox, Marcia Meldrum, and Ira Rezak, editors. Nobel laureates in medicine or physiology: a biographical dictionary, volume 852 of Garland reference library of the humanities. Garland Publishing, New York, NY, USA, 1990. ISBN 0-8240-7892-6. xviii + 595 pp. LCCN R134 .F69 1990.

Falnes:1938:NNP

[FN38] Oscar J. Falnes and Alfred Nobel. Norway and the Nobel Peace Prize. Columbia University Press, New York, NY, USA, 1938. ISBN 020. xii + 332 pp. LCCN ????

Frandsen:2013:REC

[FN13] Tove Faber Frandsen and Jeppe Nicolaisen. The ripple effect: Citation chain reactions of a Nobel Prize. Journal of the American Society for Information Science and Technology: JASIST, 64(3):437–447, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fock:1953:AUC

[Foc53a] V. A. Fock. Against unscientific criticism of modern physical theories. Voprosy filosofii, 1(??):168–174, ???? 1953.

Fock:1953:KPS

[Foc53b] V. A. Fock. Das Kopernikanische und das Ptolemäische System im Lichte der allgemeinen Relativitätstheorie. (German) [The Copernican and the Ptolemaic system in the light of the general theory of relativity]. Sowjetwissenschaft, ??(??):805–809, ???? 1953.

Fock:1953:U

[Foc53c] V. A. Fock. [unknown]. Mysl Filosoficzna, ??(??):??, ???? 1953. Reprint of yet-to-be-identified Russian original, cited in [Inf78b, page 80].

Fock:1954:MTR

[Foc54] V. A. Fock. Die moderne Theorie von Raum und Zeit. (German) [The modern theory of space and time]. Sowjetwissenschaft, ?? (??):525–544, ???? 1954.

Fock:1955:CUC

[Foc55] V. A. Fock. The concept of uniformity, covariance and relativity in the theory of space and time. Voprosy filosofii, 4(??):131–135, ???? 1955.

Foerster:1923:RTI

[Foe23]  Ernst Foerster.  A rudder that turns itself, proof of

Einstein’s theory from the atom. Scientific American, 129(4):237, October 1923. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v129/n4/
pdf/scientificamerican1023-237.pdf.

Foer:2012:MEA

[Foe11] Joshua Foer. Moonwalking with Einstein: the art and science of remembering everything. Penguin, London, UK and New York, NY, USA, 2011. ISBN 1-59420-229-X (hardcover), 0-14-312053-0 (paperback). 307 pp. LCCN BF385 .F64 2011.

Fokker:1949:STG

[Fok49] A. D. Fokker. On the space–time geometry of a moving rigid body. Reviews of Modern Physics, 21(3):406–408, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.406;   http://rmp.aps.org/

abstract/RMP/v21/i3/p406_1. Special issue in honor of Albert Einstein on his 70th birthday.

Fokker:1953:AEI

[Fok55] A. D. Fokker. Albert Einstein, inventor of chronogeometry. Synthese, 9(1):442–445, ???? 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567431.

Fok:1959:TST

[Fok59] V. A. (Vladimir Aleksandrovich) Fok. The theory of space, time and gravitation. Pergamon Press, New York, NY, USA, 1959. 411 pp. LCCN QC6 .F573. Translated from the Russian by N. Kemmer.

Fok:1964:TSTa

[Fok64a] V. A. (Vladimir Aleksandrovich) Fok. The theory of space, time and gravitation. Macmillan Publishing Company, New York, NY, USA, second revised edition, 1964. xi + 448 pp. LCCN QC6 .F573 1964. Translated from the Russian by N. Kemmer.

Fok:1964:TSTb

[Fok64b] V. A. (Vladimir Aleksandrovich) Fok. The theory of space, time and gravitation. Pergamon, New York, NY, USA, second revised edition, 1964. xi + 448 pp. LCCN QC6 .F573 1964. Translated from the Russian by N. Kemmer.

Fokker:1965:TSW

[Fok65] A. D. (Adriaan Daniël) Fokker, editor. Time and space, weight and inertia; a chronogeometrical introduction to Einstein’s theory, volume 2 of International series of monographs in natural philosophy. Pergamon, New York, NY, USA, 1965. xv + 188 pp. LCCN QC6 .F5813 1965. Translated by D. Bijl and edited by D. Field.

Folsing:1990:KMR

[Föl90] Albrecht Fölsing. Keine “Mutter der Relativitätstheorie”. Die Zeit, 47(??):??, November 16, 1990. URL http://www.

esterson.org/Foelsing_Die_Zeit_1990.htm.

Folsing:1991:NFN

[Föl91] Ulla Fölsing. Nobel-Frauen: naturwissenschaftlerinnen im Porträt. (German) [Nobel women: natural scientists in the portrait]. Verlag C. H. Beck, München, Germany, second edition, 1991. ISBN 3-406-34018-0. 214 pp. LCCN Q141 .F55 1991.

Folsing:1993:AEB

[Föl93] Albrecht Fölsing. Albert Einstein: eine Biographie. (German) [Albert Einstein: a Biography]. Suhrkamp, Frankfurt am Main, Germany, 1993. ISBN 3-518-40489-X. 959 + 32 pp. LCCN QC16.E5 F59 1993. English translation in [Föl97].

Folse:1996:BED

[Fol96] Henry J. Folse. The Bohr–Einstein debate and the philosophers’ debate over realism versus anti-realism. In Cohen et al. [CHR96], pages 289–298. ISBN 94-015-8638-1. ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.
1007/978-94-015-8638-2_20.

Folsing:1997:AEB

[Föl97] Albrecht Fölsing. Albert Einstein: a biography. Viking, New York, NY, USA, 1997. ISBN 0-670-85545-6. xiii + 882 pp. LCCN QC16.E5 F5913 1997.

Folsing:1998:AEB

[Föl98] Albrecht Fölsing. Albert Einstein: a biography. Penguin, London, UK and New York, NY, USA, 1998. ISBN 0-670-85545-6, 0-14-023719-4 (Penguin paperback). xiii + 882 + 16 pp. LCCN QC16.E5 F5913 1997.

Folger:2015:TBH

[Fol15] Tim Folger. Theory: A brief history of time travel. Scientific American, 313(3):68–73, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-68.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-68.pdf.

Folgarait:2017:PPP

[Fol17] Leonard Folgarait. Painting 1909: Pablo Picasso, Gertrude Stein, Henri Bergson, comics, Albert Einstein, and anarchy. Yale University Press, New Haven, CT, USA, 2017. ISBN 0-300-21801-X (hardcover). xi + 226 pp. LCCN ND553.P5 F65 2017.

Foppl:2007:ARM

[Föp07] August Föppl. On absolute and relative motion. In Renn et al. [RSS^{+}07a], pages 145–152. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_15.

Forsee:1963:AET

[For63] Aylesa Forsee. Albert Einstein: theoretical physicist. Macmillan Publishing Company, New York, NY, USA, 1963. xiii + 202 pp.

LCCN QC16.E5 F6. URL http://www.loc.gov/catdir/toc/
fy0805/63014538.html.

Forman:1970:BRBe

[For70a] Paul Forman. Book review: Briefwechsel, 1916–1955, by Albert Einstein, Max Born, and Hedwig Born. Isis, 61(4):553–555, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229480;   http://www.jstor.org/
stable/pdfplus/229480.pdf.

Forman:1970:BRBc

[For70b] Paul Forman. Book review: Briefwechsel. Sechzig Briefe aus dem goldenen Zeitalter der modernen Physik by Albert Einstein; Arnold Sommerfeld; Armin Hermann. Isis, 61(2):287–288, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230001.

Forbes:1971:BRB

[For71a] Eric G. Forbes. Book review: Physics without Einstein, by Harold Aspden. Annals of Science, 27(3):302–303, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forman:1971:WCC

[For71b] Paul Forman. Weimar culture, causality, and quantum theory, 1918–1927: Adaptation by German physicists and mathematicians to a hostile intellectual environment. Historical Studies in the Physical Sciences, 3(??):1–115, ???? 1971. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757315.

Forman:1973:SIW

[For73] Paul Forman. Scientific internationalism and the Weimar physicists: The ideology and its manipulation in Germany after World War I. Isis, 64(2):150–180, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229595.

Forman:1980:BRE

[For80a] Paul Forman. Book reviews: Einstein: A Centenary Volume, edited by A. P. French; Albert Einstein’s Theory of General Relativity, edited by Gerald E. Tauber; Einstein, by Louis de Broglie and others; Einstein. Sein Leben und seine Zeit, by Philipp Frank; Documentary publications: Albert Einstein Autobiographical Notes, translated and edited by Paul Arthur Schilpp; Elie Cartan, Albert Einstein, Letters on Absolute Parallelism 1929–1932, edited by Robert Debever; Aus meinen späten Jahren, by Albert Einstein; Albert Einstein in Berlin 1918–1933, Teil I: Darstellung und Dokumente; Teil II: Spezialinventar; Images of Einstein: a Catalog, compiled by Joan N. Warnow; Albert Einstein: the Human Side. New Glimpses from his Archives, edited by Helen Dukas and Banesh Hoffman; Einstein privat. Herta W. erinnert sich an die Jahre 1927 bis 1933, by Friedrich Herneck; Centenary exhibitions: Albert Einstein, 1879–1955. A Centenary Exhibit of Manuscripts, Books, and Portraits Selected from the Humanities Research Center Collections, by Albert C. Lewis; Gedächtnisausstellung zum 100. Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner; Einstein = mc^{2}; Einstein 1879–1979. Exhibition . Bulletin of the Atomic Scientists, 36(3):49–55, March 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forman:1980:WCC

[For80b] Paul Forman. Weimar culture, causality, and quantum theory, 1918–1927: Adaptation by German physicists and mathematicians to a hostile intellectual environment. In Chant and Fauvel [CF80], page ?? ISBN 0-582-49157-6 (paperback), 0-582-49156-8. LCCN Q174.8 .D37.

Ford:1982:CAS

[For82a] Daniel F. Ford. The cult of the atom: the secret papers of the Atomic Energy Commission. Simon and Schuster, New York, NY, USA, 1982. ISBN 0-671-25301-8. 273 pp. LCCN TK9023 .F67 1982.

Ford:1982:TMI

[For82b] Daniel F. Ford. Three Mile Island: thirty minutes to meltdown. Penguin Books, Harmondsworth, Middlesex, England, 1982. ISBN 0-14-006048-0 (paperback). 271 pp. LCCN TK1345.H37 F67.

Ford:1984:CAS

[For84a] Daniel F. Ford. The cult of the atom: the secret papers of the Atomic Energy Commission. A Touchstone book. Simon and Schuster, New York, NY, USA, revised and updated edition, 1984. ISBN 0-671-25302-6 (paperback). 273 pp. LCCN TK9023 .F67 1984.

Forman:1984:SLL

[For84b] P. Forman. “Subtle Is the Lord … ”: The Science and the Life of Albert Einstein by Abraham Pais (review). Technology and Culture, 25(3):697–698, July 1984. CODEN TECUA3. ISSN 0040-165x (print), 1097-3729 (electronic).

Ford:1986:MSP

[For86] Daniel F. Ford. Meltdown: The secret papers of the Atomic Energy Commission. A Touchstone book. Simon and Schuster, New York, NY, USA, revised Touchstone edition, 1986. ISBN 0-671-63449-6 (paperback). 307 pp. LCCN TK9023 .F67 1986. US$6.95.

Ford:2001:BRJ

[For01] Kenneth W. Ford. Book review: John Canaday, The Nuclear Muse: Literature, Physics and the First Atomic Bomb. Madison: University of Wisconsin Press, 2000, xviii + 303 pages. $22.95 (paper), $60.00 (cloth). Physics in Perspective (PIP), 3 (4):492–493, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Forgan:2005:BRU

[For05] Sophie Forgan. Book review: Ulf Larsson (ed.), Cultures of Creativity: The Centennial Exhibition of the Nobel Prize. English translation by Daniel M. Olson. Nobel Museum Archives, 2. Canton, MA: Science History Publications, 2001. Pp. 228. ISBN 0-88135-288-8. $40.00 (paperback). British Journal for the History of Science, 38(2):242–243, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4028716.

Forman:2009:BRD

[For09] Paul Forman. Book review: David Lindley: Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for the Soul of Science. Isis, 100(1):180–181, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/599679.

Ford:2014:TRE

[For14] Sarah Gilbreath Ford. Of trains and relativity: Einstein and perspective in Eudora Welty’s Delta Wedding. Studies in the Novel, 46(3):354–370, Fall 2014. CODEN ???? ISSN 0039-3827 (print), 1934-1512 (electronic).

Ford:2015:BHB

[For15] Kenneth William Ford. Building the H bomb: a personal history. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4632-07-4 (hardcover), 981-4618-79-9 (paperback). xiii + 221 pp. LCCN

UG1282.A8 F67 2015. URL http://www.worldscientific.
com/worldscibooks/10.1142/9269.

Foster:2005:EHL

[Fos05] Brian Foster. Einstein and his love of music. Physics World, 18(1):34, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a28.
pdf.

Foster:2008:EU

[Fos08] Brian Foster. Einstein unmasked. Physics World, 21(9):38–39, September 2008. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/21/09/phwv21i09a34.pdf.

Fossheim:2013:SDD

[Fos13] Kristian Fossheim. Superconductivity: Discovery and Discoverers: Ten Physics Nobel Laureates Tell Their Story. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-36058-0, 3-642-36059-9 (e-book). xiii + 140 pp. LCCN ????

Fowler:2003:NGE

[Fow03] Michael Fowler. News: Galileo and Einstein: Using history to teach basic physics to nonscientists. Science & Education (Springer), 12(2):229–231, February 2003. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Fox:1965:EAE

[Fox65] J. G. Fox. Evidence against emission theories. American Journal of Physics, 33(1):1–17, January 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v33/i1/p1_s1. See [Ein09d,

Rit09b, RE09].

Fox:2015:LEO

[Fox15] Robert Fox. Lindemann and Einstein: the Oxford connexion. In Arabatzis et al. [ARS15], pages 23–31. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). LCCN Q125 .R45 2015.

URL http://www.loc.gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.loc.gov/catdir/enhancements/
fy1516/2015936523-t.html.

Fox:2018:EO

[Fox18] Robert Fox. Einstein in Oxford. Notes and Records of the Royal Society of London, 72(3):293–318, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Flerov:1940:SFU

[FP40a] Flerov and Petrjak. Spontaneous fission of uranium. Physical Review (2), 58(1):89, July 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.58.89.2.

Frisch:1940:MPR

[FP40b] Otto Robert Frisch and Rudolf Peierls. Memorandum on the properties of a radioactive “super-bomb”. Technical report, University of Birmingham, Birmingham, UK, March

1940. URL http://en.wikipedia.org/wiki/Frisch%E2%
80%93Peierls_memorandum;   http://www.atomicarchive.
com/Bios/Frisch.shtml;   http://www.atomicarchive.com/
Bios/Peierls.shtml;   http://www.atomicarchive.com/
Docs/Begin/FrischPeierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls2.shtml.

Frisch:1940:CSB

[FP40c] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium. Technical report, University of Birmingham, Birmingham, UK, March 1940. URL http://en.wikipedia.

org/wiki/Frisch%E2%80%93Peierls_memorandum;   http:
//www.atomicarchive.com/Bios/Frisch.shtml;   http:/
/www.atomicarchive.com/Bios/Peierls.shtml;   http://
www.atomicarchive.com/Docs/Begin/FrischPeierls.shtml.

Reprinted in [?].

Frisch:1940:PRS

[FP40d] Otto Robert Frisch and Rudolf Peierls. The properties of a radioactive “super-bomb”. Memorandum, University of Birmingham, Birmingham, UK, March 1940. URL

http://en.wikipedia.org/wiki/Frisch%E2%80%93Peierls_
memorandum;   http://www.atomicarchive.com/Bios/Frisch.
shtml;   http://www.atomicarchive.com/Bios/Peierls.
shtml;   http://www.atomicarchive.com/Docs/Begin/FrischPeierls.
shtml;   http://www.atomicarchive.com/Docs/Begin/FrischPeierls2.

shtml. Reprinted in [?].

Frisch:2016:RRE

[FP16] Mathias Frisch and Wolfgang Pietsch. Reassessing the Ritz–Einstein debate on the radiation asymmetry in classical electrodynamics. Studies in History and Philosophy of Modern Physics, 55(??):13–23, August 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300150.

Frisch:1959:TAP

[FPLR59] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Trends in atomic physics; essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday: Atomic physics. Interscience Publishers, New York, NY, USA, 1959. 285 pp. LCCN QC475 .B45 1959a. An identical edition is published simultaneously under the title Beiträge zur Physik und Chemie des 20. Jahrhunderts.

Fermi:1949:MRS

[FR49a] Enrico Fermi and Isidor Isaac Rabi. Minority report to the Science Advisory Committee (SAC). This report opposed the development of the hydrogen bomb on ethical grounds. It is mentioned, but not properly cited, in [Rig00, Wol09]., 1949.

Fermi:1949:ODS

[FR49b] Enrico Fermi and Isidor Isaac Rabi. An opinion on the development of the Super. This minority report to the Science Advisory Committee (SAC) opposed the development of the hydrogen bomb on ethical grounds. It is mentioned, but not properly cited, in [Rig00, Wol09]. It is reprinted in [?, pages 120–127] and [CHW91]., 1949.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896– 1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:1919:RTE

[Fra20] Philipp Frank. Zur Relativitäts-Theorie Einsteins. (German) [Towards Einstein’s relativity theory]. Lotos, 67/68:152–156, 1919/1920.

Frank:1937:WVP

[Fra37] Philipp Frank. Was versteht der Physiker unter der “Grösse” eines Körpers?: Bemerkungen zu A. Phaléns Kritik der Einsteinschen Relativitätstheorie. (German) [What does the physicist understand by the “size” of a body?: Remarks on A. Phalén’s critique of Einstein’s relativity theory]. Theoria, 3(1): 90–114, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Frank:1947:EHL

[Fra47] Philipp Frank. Einstein, his life and times. Alfred A. Knopf, New York, NY, USA, 1947. xi + 298 + xii pp. Translated by George Rosen from a German manuscript, edited and revised by Shuichi Kusaka.

Frank:1948:EHL

[Fra48] Philipp Frank. Einstein, his life and times. Jonathan Cape, London, UK, 1948. 367 pp.

Frank:1949:EML

[Fra49a] Philipp Frank. Einstein, Mach, and logical positivism. In Schilpp [Sch49b], chapter 9, pages 269–286. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Frank:1949:ESL

[Fra49b] Philipp Frank. Einstein: Sein Leben und seine Zeit. (German) [Einstein: His Life and Times]. Paul List Verlag, München, Leipzig, Freiburg i. Brg, Germany, 1949. 467 pp. LCCN ????

Frank:1949:EPS

[Fra49c] Philipp Frank. Einstein’s philosophy of science. Reviews of Modern Physics, 21(3):349–355, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.349;
http://rmp.aps.org/abstract/RMP/v21/i3/p349_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Frank:1949:ESP

[Fra49d] Philipp Frank. Einsteins Stellung zur Philosophie. (German) [Einstein’s position on philosophy]. Deutsche Beiträge, 2(??): 146–157, ???? 1949.

Frank:1950:RRT

[Fra50] Philipp Frank, editor. Relativity, a richer truth. Seeds-of-thought series. Beacon Press, Boston, MA, USA, 1950. xvi + 142 pp. LCCN QC6 F642. With a foreword by Albert Einstein.

Frank:1951:RRT

[Fra51] Philipp Frank. Relativity, a richer truth. Cape, London, UK, 1951. 124 pp. LCCN QC6 .F642 1951. With a foreword by Albert Einstein.

Frank:1952:WRA

[Fra52] Philipp Frank. Wahrheit, relativ oder absolut? (German) [Truth, relative or absolute?], volume 12 of Internationale Bibliothek für Psychologie und Soziologie. Pan-Verlag, Zürich, Switzerland, 1952. 170 pp. LCCN BD221 .F7. With a foreword by Albert Einstein. German translation by Gertrud Deuel.

Frank:1953:EHL

[Fra53] Philipp Frank. Einstein, his life and times. Alfred A. Knopf, New York, NY, USA, 1953. xxvii + 298 pp. Translated by George Rosen from a German manuscript, edited and revised by Shuichi Kusaka.

Frank:1953:E

[Fra55] Philipp G. Frank. Einstein. Synthese, 9(1):435–437, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/
10.1007/BF00567429.

Frank:1965:EHL

[Fra65] Philipp Frank. Einstein, his life and times. A Borzoi book. Alfred A. Knopf, New York, NY, USA, 1965. 298 pp. LCCN QC16 E5 F713 1965; QC 16 .E5F713 1970 PHYS; QC 16 .E5F713 1965 GERSTM. Translated by George Rosen from a German Manuscript. Edited and revised by Shuichi Kusaka. With a new introduction.

Franklin:1981:MPU

[Fra81] Allan D. Franklin. Millikan’s published and unpublished data on oil drops. Historical Studies in the Physical Sciences, 11(2): 185–201, ???? 1981. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757478.

Frank:1988:EHL

[Fra88] Philipp Frank. Einstein: his life and times. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1988. xxvii + 298 + xii + 16 pp. LCCN QC16.E5.

Fradkin:1989:FAN

[Fra89] Philip L. Fradkin. Fallout: an American nuclear tragedy. University of Arizona Press, Tucson, AZ, USA, 1989. ISBN 0-8165-1086-5. xiii + 300 pp. LCCN U264 .F72 1989.

Frank:1991:ESV

[Fra91] Philipp Frank. Einstein: sa vie et son temps. (French) [Einstein: his life and times], volume 242 of Champs. Flammarion, Paris, France, 1991. ISBN 2-08-081242-4. 473 pp. LCCN QC16 .E5F714.

Frank:1993:OEFa

[Fra93a] Sir Charles Frank, editor. Operation Epsilon: the Farm Hall transcripts. Institute of Physics Publishers, Bristol, UK, 1993. ISBN 0-7503-0274-7. ix + 313 pp. LCCN Q141 .O56 1993. US$15.00.

Frank:1993:OEFb

[Fra93b] Sir Charles Frank, OBE, FRS, editor. Operation Epsilon: the Farm Hall transcripts. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1993. ISBN 0-520-08499-3. ix + 313 pp. LCCN Q141 .O56 1993b.

Frank:1999:DEI

[Fra99] Richard B. Frank. Downfall: the end of the Imperial Japanese Empire. Random House, New York, NY, USA, 1999. ISBN 0-679-41424-X (hardcover). 484 pp. LCCN

D767.2 .F73 1999.  URL http://www.loc.gov/catdir/
bios/random059/99011838.html;   http://www.loc.gov/
catdir/description/random0411/99011838.html;   http:
//www.loc.gov/catdir/samples/random044/99011838.html.

Frayn:2000:C

[Fra00] Michael Frayn. Copenhagen. Anchor Books, New York, NY, USA, 2000. ISBN 0-385-72079-3 (paperback). 132 pp. LCCN PR6056.R3 C64 2000. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/

catdir/description/random044/00055814.html;   http:/
/www.pbs.org/hollywoodpresents/copenhagen/.

Frank:2002:EHL

[Fra02] Philipp Frank. Einstein: his life and times. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, second edition, 2002. ISBN 0-306-81109-X. xxvii + 298 + xii pp. LCCN QC16.E5.

Fradkin:2004:FAN

[Fra04] Philip L. Fradkin. Fallout: an American nuclear tragedy. Johnson Books, Boulder, CO, USA, 2004. ISBN 1-55566-331-1 (paperback). xiv + 312 pp. LCCN U264.4.N3 F73 2004. URL http:

//www.loc.gov/catdir/toc/ecip0411/2003024786.html.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2005PhP.
....7..204F;   http://www.springerlink.com/content/
1422-6960/.

Fraser:2008:CAA

[Fra08] Gordon Fraser. Cosmic anger: Abdus Salam — the first Muslim Nobel scientist. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-920846-8 (hardcover), 0-19-170895-X. LCCN QC16.S26 F73 2008.

Frankel:2011:GCI

[Fra11] Theodore Frankel. Gravitational curvature: an introduction to Einstein’s theory. Dover, New York, NY, USA, Dover edition, 2011. ISBN 0-486-48121-2 (paperback). xv + 172 pp. LCCN QC173.55 .F7 2011.

Franson:2014:EES

[Fra14] James Franson. Entanglement: Einstein’s spooky action at a distance. Philosophical Society of Washington, ??(??):??, Oc-

tober 24, 2014. URL http://www.philsoc.org/2014Fall/
2337abstract.html;   http://www.philsoc.org/2014Fall/
2337transcript.html.

Frank:2015:E

[Fra15] Philipp Frank. Einstein. Castelvecchi, Roma, Italia, 2015. ISBN 88-6944-072-9. 282 pp. LCCN ???? Translation to Italian by Fernando Rocca of [Fra02].

Freundlich:1916:GEG

[Fre16] Erwin Freundlich. Die Grundlagen der Einsteinschen Gravitationstheorie (German) [The Foundations of Einstein’s Theory of Gravitation]. Julius Springer, Berlin, Germany, 1916. 64 pp. LCCN QC178 .F67 1916.

Freundlich:1919:AES

[Fre19] Erwin Freundlich. Albert Einstein: Zum Siege seiner Relativitätstheorie. (German) [Albert Einstein: The victories of his Theory of Relativity]. Vossische Zeitung Supplement, November 30, 1919.

Freundlich:1920:GEG

[Fre20] Erwin Freundlich. Die Grundlagen der Einsteinschen Gravitationstheorie (German) [The Foundations of Einstein’s Theory of Gravitation]. J. Springer, Berlin, fourth edition, 1920. 3 + 96 pp. LCCN QC178 .F67 1920.

Freundlich:1927:GTT

[Fre27a] Erwin Freundlich. Das große Turm-Teleskop in Potsdam — Ein Besuch im Einstein Turm. (German) [The big tower telescope in Potsdam — a visit to the Einstein Tower]. Berliner Illustrierte Zeitung, 37(51):2195–2196, 1927.

Freundlich:1927:TES

[Fre27b] Erwin Freundlich. Das Turmteleskop der Einstein Stiftung. (German) [The Tower Telescope of the Einstein Foundation]. J. Springer, Berlin, Germany, 1927. ii + 43 + 1 pp. LCCN QB82.P7 F7.

Frenkel:1931:WDE

[Fre31] J. Frenkel. What does Einstein mean? Science, 74(1929):609– 618, December 18, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/74/1929/609.full.pdf.

French:1968:SR

[Fre68] A. P. (Anthony Philip) French. Special Relativity. W. W. Norton & Co., New York, NY, USA, 1968. x + 286 pp. LCCN QC6 .F68.

Frenz:1969:L

[Fre69] Horst Frenz. Literature, 1901–1967. Nobel lectures, including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1969. xxi + 640 pp. LCCN ???? Published for the Nobel Foundation.

Freedman:1979:FSP

[Fre79a] Michael I. Freedman. Frederick Soddy and the practical significance of radioactive matter. British Journal for the History of Science, 12(3):257–260, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

French:1979:ECVa

[Fre79b] A. P. (Anthony Philip) French, editor. Einstein: a centenary volume. Heinemann for the International Commission of Physics Education, London, UK, 1979. ISBN 0-435-58200-3. xx + 332 pp. LCCN QC16.E5 E37 1979b.

French:1979:ECVb

[Fre79c] A. P. (Anthony Philip) French, editor. Einstein: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1979. ISBN 0-674-24230-0. xx + 332 pp. LCCN QC16.E5 E37.

French:1979:ELD

[Fre79d] A. P. (Anthony Philip) French, editor. Einstein: le livre du centenaire. (French) [Einstein: the centenary volume]. Hier et demain, Paris, France, 1979. ISBN 2-7206-0061-X. xxiii + 333 pp. Edited by G. Delacôte and J. Souchon-Rouyer. Preface by Alfred Kastler. French translation by J.-B. Yelnick and others.

Freeman:1981:NWI

[Fre81a] Leslie J. Freeman. Nuclear witnesses: insiders speak out. Norton, New York, NY, USA, 1981. ISBN 0-393-01456-8. xxvii + 330 pp. LCCN TK9023 .N83 1981.

French:1981:TCE

[Fre81b] A. P. French. ‘and then came Einstein’. Nature, 293(5835): 766, October 29, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v293/n5835/pdf/293766a0.pdf.

French:1985:AEW

[Fre85] A. P. (Anthony Philip) French, editor. Albert Einstein: Wirkung und Nachwirkung. (German) [Albert Einstein: Effects and Aftereffects]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1985. ISBN 3-528-08497-9. xi + 392 pp.

French:1986:NBH

[Fre86] A. P. French. Niels Bohr at 100: his life and work. Physics Education, 21(4):220–226, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http:/

/stacks.iop.org/0031-9120/21/i=4/a=005.

Frenkel:1996:YIF

[Fre96] Viktor Iakovlevich Frenkel. Yakov Ilich Frenkel: his work, life, and letters. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-2741-3 (Basel), 0-8176-2741-3 (Boston). viii + 323 pp. LCCN QC16.F715 A3 1996.

Fredrickson:2006:MDA

[Fre06] D. Fredrickson. My duets with Albert Einstein. Clavier, 45(6): 6–??, July/August 2006. ISSN 0009-854X.

Freedman:2008:BRB

[Fre08] Lawrence D. Freedman. Book reviews: Hiroshima: The World’s Bomb, by Andrew J. Rotter; Atomic Tragedy: Henry L. Stimson and the Decision to Use the Bomb Against Japan, by Sean L. Malloy; The Nuclear Taboo: The United States and the Non-Use of Nuclear Weapons Since 1945, by Nina Tannenwald. Foreign Affairs, 87(6):161–162, November 2008. CO-DEN FRNAA3. ISSN 0015-7120. URL http://www.jstor.

org/stable/20699391.

Freundlich:2011:FET

[Fre11] Erwin Freundlich. The foundations of Einstein’s theory of gravitation. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-60137-1 (paperback). ???? pp. LCCN ????

Fredrikson:2013:ANO

[Fre13] Bengt Fredrikson. Alfred Nobel: den olycklige uppfinnaren. (Swedish) [Alfred Nobel: the unhappy inventor]. Fakta. LLförlaget, Stockholm, Sweden, 2013. ISBN 91-7053-440-3 (hardcover). 182 + 3 pp. LCCN TP268.5.N7 F74 2013.

Fredrikson:2015:AEM

[Fre15a] Bengt Fredrikson. Albert Einstein — Det modiga geniet. (Swedish) [Albert Einstein — The brave genius]. LL-förlaget, ????, 2015. ISBN 91-7053-525-6 (hardcover). ???? pp. LCCN ????

FreireJunior:2015:QDR

[Fre15b] Olival Freire Junior. The Quantum Dissidents: Rebuilding the Foundations of Quantum Mechanics (1950–1990). Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-662-44661-8, 3-662-44662-6 (e-book). xvi + 356 pp. LCCN QC173.98 .F74 2015.

Frenkel:2018:AEP

[Fre18] Edward Frenkel. AMS Einstein public lecture: imagination and knowledge. Notices of the American Mathematical Society, 65 (4):410–412, 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Frisch:1939:RSP

[Fri39a] Otto R. Frisch. Radioactivity and subatomic phenomena. Annual Reports on the Progress of Chemistry, 36:7–24, 1939. CO-DEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Frisch:1939:PED

[Fri39b] Otto Robert Frisch. Physical evidence for the division of heavy nuclei under neutron bombardment. Nature, 143(3616):276, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.chemteam.info/
Chem-History/Frisch-Fission-1939.html;   http://www.
nature.com/nature/journal/v143/n3616/pdf/143276a0.pdf.

Friendly:1950:QD

[Fri50]  Fred Friendly.  The quick and the dead.  Four-part

NBC radio miniseries on the creation of the atomic bomb, narrated by American entertainer Bob Hope, with Helen Hayes playing Lise Meitner and with Enrico Fermi and other leading scientists from the Manhattan Project receiving some mention., 1950. URL https://www.discogs.

com/National-Broadcasting-Company-The-The-Quick-And-
The-Dead-Volume-1-The-Atom-Bomb-/release/5350974;
https://www.questia.com/library/journal/1P3-1718351551/
living-with-the-bomb-fred-friendly-s-the-quick-and.

Frick:1964:PME

[Fri64] Martin Frick. On a possibility of measuring the Einsteinshift photoelectrically. Nature, 204(4963):1076, December 12, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v204/n4963/pdf/2041076a0.pdf.

Frisch:1968:LMD

[Fri68] Otto R. Frisch. Lise Meitner dies; nuclear-physics pioneer. Physics Today, 21(12):101, December 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

scitation.aip.org/content/aip/magazine/physicstoday/
article/21/12/10.1063/1.3034630.

Frisch:1973:BRA

[Fri73] O. R. Frisch. Book review: Albert Einstein: Creator and Rebel, by B. Hoffman. London: Hart-Davis, MacGibbon, 1973. pp xv + 272. Physics Bulletin, 24(11):679, November 1973. CODEN PHSBB4. ISSN 0031-9112.

Frisch:1979:GRH

[Fri79a] Otto Robert Frisch. Great relativist and humanitarian — life and times of Albert Einstein 1879–1955. New Scientist, 81 (1145):752–755, ???? 1979. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Frisch:1979:WLR

[Fri79b] Otto Robert Frisch. What Little I Remember. Cambridge University Press, Cambridge, UK, 1979. ISBN 0-521-22297-4. xi + 227 pp. LCCN QC16.F75 A38. URL http://www.loc.

gov/catdir/description/cam022/78018096.html;   http:
//www.loc.gov/catdir/toc/cam026/78018096.html.

Friedman:1981:NPP

[Fri81] Robert Marc Friedman. Nobel Physics Prize in perspective. Nature, 292(5826):793–798, August 27, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Friedman:1983:FST

[Fri83] Michael Friedman. Foundations of space–time theories: relativistic physics and philosophy of science. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-07239-6. xvi + 385 pp. LCCN QC173.55 .F74 1983. US$35.00.

Friedman:1989:TCQ

[Fri89] Robert Marc Friedman. Text, context, and quicksand: Method and understanding in studying the Nobel Science Prizes. Historical Studies in the Physical and Biological Sciences, 20(1): 63–77, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print),

1533-8355 (electronic). URL http://www.jstor.org/stable/
27757635.

Friedman:2001:PEB

[Fri01] Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. Times Books, New York, NY, USA, 2001. ISBN 0-7167-3103-7. xv + 379 pp. LCCN QC49 .F75 2001.

Friedlaender:2007:ARM

[Fri07] Benedict Immanuel Friedlaender. Absolute or relative motion? In Renn et al. [RSS^{+}07a], pages 127–144. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/
978-1-4020-4000-9_14.

Friedrich:2009:BEC

[Fri09] Bretislav Friedrich. Bose–Einstein condensates: BECs from the fridge. Nature Physics, 5(10):712–714, October 2009. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic).

URL http://www.nature.com/nphys/journal/v5/n10/full/
nphys1417.html.

Fritzsch:2011:YWM

[Fri11] Harald Fritzsch. You are wrong, Mr. Einstein!: Newton, Einstein, Heisenberg, and Feynman discussing quantum mechanics.

World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4324-99-X. xxi + 178 pp. LCCN QC174.12 .F755 2011.

Friedman:2012:SAS

[Fri12] Robert Marc Friedman. “Has the Swedish Academy of Sciences seen nothing, heard nothing, and understood nothing?”: the First World War, biased neutrality, and the Nobel Prizes in science. In Lettevall et al. [LSW12], chapter 5, pages 90–114. ISBN 0-203-11679-8, 0-415-89377-1. ISSN 1404-7586. LCCN

D723 .N49 2012. URL http://libanswers.liverpool.ac.uk/
faq/182315.

Friedman:2014:STG

[Fri14] Michael Friedman. Space, time, and geometry. In Janssen and Lehner [JL14b], chapter 13, pages 398–420. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Friege:2017:MME

[Fri17] Johanna Friege. Meitner’s Matilda effect. Varsity, ??(??):??, Oc-

tober 12, 2017. URL https://www.varsity.co.uk/science/
13682.

Fruton:1982:ICB

[Fru82] Joseph S. Fruton. The interplay of chemistry and biology at the turn of the century. In Bernhard et al. [BCS82], pages 74–96. ISBN 0-08-027939-2. LCCN Q124.6 .N63

1981. US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500104.

Feld:1972:CWL

[FS87] Bernard T. Feld and Gertrud Weiss Szilard, editors. The collected works of Leo Szilard: Scientific papers, volume 1. MIT Press, Cambridge, MA, USA, 1972–1987. ISBN 0-262-06039-6 (vol. 1: 1972), 0-262-19168-7 (vol. 2: 1978), 0-262-19260-8 (vol. 3: 1987). xxii + 737 pp. LCCN QC3 .S97. With Kathleen R. Winsor. Foreword by Jacques Monod. Introductory essays by Carl Eckart, Bernard T. Feld, Maurice Goldhaber, Aaron Novick, and Julius Tabin.

Fermi:1995:PBP

[FS95] Rachel Fermi and Esther Samra. Picturing the Bomb: Photographs from the Secret World of the Manhattan Project. H.

N. Abrams, New York, NY, USA, 1995. ISBN 0-8109-3735-2. 232 pp. LCCN QC773.A1 F47 1995. Introduction by Richard Rhodes.

Feoli:1998:BMW

[FS98] A. Feoli and G. Scarpetta. De Broglie matter waves from the linearized Einstein field equations. Foundations of Physics Letters, 11(4):395–403, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Folman:2001:BEC

[FS01] Ron Folman and Jörg Schmiedmayer. Bose–Einstein condensates: Mastering the language of atoms. Nature, 413(6855): 466–467, October 4, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v413/n6855/full/413466a0.html.

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-1. LCCN ???? URL http://deposit.

ddb.de/cgi-bin/dokserv?id=2758637\%26prov=M\%26dok\_
var=1\%26dok\_ext=htm. 4 CDs.

Ferreira:2009:ETG

[FS09] Pedro G. Ferreira and Glenn D. Starkman. Einstein’s theory of gravity and the problem of missing mass. Science, 326(5954): 812–815, November 6, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/326/5954/812.abstract.

Freundlich:1931:NPF

[FSE31] E. E. Freundlich, C. E. St. John, and A. Einstein. New proofs found of an Einstein idea: He is told findings show rays of star light bend in passing near sun. New York Times, page 3 (column 3), February 21, 1931. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Fischbach:1986:ERE

[FSS^{+}86a] Ephraim Fischbach, Daniel Sudarsky, Aaron Szafer, Carrick Talmadge, and S. H. Aronson. Errata: Reanalysis of the Eötvös Experiment. Physical Review Letters, 56(13):1427, March 1986.

CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.56.1427. See [FSS^{+}86b].

Fischbach:1986:REE

[FSS^{+}86b] Ephraim Fischbach, Daniel Sudarsky, Aaron Szafer, Carrick Talmadge, and S. H. Aronson. Reanalysis of the Eötvös experiment. Physical Review Letters, 56:3–6, January 1986. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/

PhysRevLett.56.3. See minor erratum [FSS^{+}86a].

Friedlaender:2008:KGE

[FTG08] Salomo Friedlaender, Detlef Thiel, and Hartmut Geerken. Kant gegen Einstein: Fragelehrbuch (nach Immanuel Kant und Ernst Marcus) zum Unterricht in den vernunftwissenschaftlichen Vorbedingungen der Naturwissenschaft. (German) [Kant to Einstein: Question textbook (by Immanuel Kant and Ernst Marcus) for instruction in the preconditions of rational scientific natural sciences]. Waitawhile: Books on Demand, Herrsching and Norderstedt, Germany, second edition, 2008. ISBN 3-8370-0052-4. 197 pp. LCCN QC6 .F79 2008.

Fuwa:2015:EPN

[FTZ^{+}15] Maria Fuwa, Shuntaro Takeda, Marcin Zwierz, Howard M. Wiseman, and Akira Furusawa. Experimental proof of nonlocal wavefunction collapse for a single particle using homodyne measurements. Nature Communications, 6(??):6665–??, March 24, 2015. CODEN NCAOBW. ISSN 2041-1723 (elec-

tronic). URL http://www.nature.com/ncomms/2015/150324/
ncomms7665/full/ncomms7665.html.

Fuller:1984:DWBb

[Ful84a] John G. Fuller. The day we bombed Utah: America’s most lethal secret. Signet Books, New York, NY, USA, 1984. ISBN 0-451-13482-6. 268 pp. LCCN CPB Box no. 1342 vol. 7.

Fuller:1984:DWBa

[Ful84b] John G. (John Grant) Fuller. The day we bombed Utah: America’s most lethal secret. New American Library, New York, NY, USA, 1984. ISBN 0-453-00457-1. 268 pp. LCCN U264 .F84

Furley:1987:GC

[Fur87] David J. Furley. The Greek cosmologists. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33328-8 (vol. 1). ???? pp. LCCN BD495 .F87 1987. URL http://www.loc.

gov/catdir/description/cam023/86026384.html;   http:
//www.loc.gov/catdir/enhancements/fy0632/86026384-t.
html.

Fussell:1988:TGA

[Fus88] Paul Fussell. Thank God for the Atom Bomb, and Other Essays. Summit Books, New York, NY, USA, 1988. ISBN 0-671-63866-1. 298 pp. LCCN AC8 .F94 1988.

Fussell:1990:TGA

[Fus90] Paul Fussell. Thank God for the atom bomb, and other essays. Ballantine Books, New York, NY, USA, 1990. ISBN 0-345-36135-0 (paperback). 257 pp. LCCN AC8 .F94 1990. US$4.95.

Freedman:1985:HDS

[Fv85] D. Z. Freedman and P. van Nieuwenhuizen. The hidden dimensions of spacetime. Scientific American, 252(3):62–69, March 1985. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fetter-Vorm:2013:TGH

[FV13] Jonathan Fetter-Vorm. Trinity: a graphic history of the first atomic bomb. Novel graphics. Hill and Wang: a division of Farrar, Straus and Giroux, New York, NY, USA, 2013. ISBN 0-8090-9355-3. 154 pp. LCCN QC773.3.U5 F47 2013.

Freedman:1978:SUL

[FvN78] Daniel Z. Freedman and Peter van Nieuwenhuizen. Supergravity and the unification of the laws of Physics. Scientific American, 238(2):126–143, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v238/n2/
pdf/scientificamerican0278-126.pdf.

Fierz:1960:TPT

[FW60] Markus Fierz and Victor Frederick Weisskopf. Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli. Interscience Publishers, New York, NY, USA, 1960. x + 328 pp. LCCN QC3 .F52. See [?] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Frisch:1967:DFH

[FW67] Otto Robert Frisch and John A. Wheeler. The discovery of fission: How it all began and mechanism of fission. Physics Today, 20(11):43–52, November 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v20/i11/p43_s1. Reprinted in [WP85, pages 272–281].

Frisch:1985:DF

[FW85] Otto R. Frisch and John A. Wheeler. The discovery of fission. In Weart and Phillips [WP85], pages 272–281. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [FW67].

Feldman:2007:MSE

[FW07] Burton Feldman and Katherine Williams. 112 Mercer Street: Einstein, Russell, Gödel, Pauli, and the end of innocence in science. Arcade Publishing, New York, NY, USA, 2007. ISBN 1-55970-704-6. xx + 243 pp. LCCN Q141 .F345 2007. URL http:

//www.loc.gov/catdir/toc/ecip078/2007001194.html.

Farys:2017:MCG

[FW17] Rudolf Farys and Tobias Wolbring. Matched control groups for modeling events in citation data: an illustration of Nobel Prize effects in citation networks. Journal of the Association for Information Science and Technology, 68(9):2201–2210, September 2017. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

G:1920:ENE

[G.20] W. G. Euclid, Newton, and Einstein. Nature, 104(2624):627– 630, February 12, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v104/n2624/pdf/104627a0.pdf.

G:1933:CAE

[G.33] T. G. The case against Einstein. Nature, 131(3304):260, February 25, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v131/n3304/pdf/131260a0.pdf.

Gumbrecht:2011:GMW

[G^{+}11] Hans Ulrich Gumbrecht et al., editors. Geist und Materie. What is life?: the intellectual pertinence of Erwin Schrödinger. Stanford University Press, Stanford, CA, USA, 2011. ISBN 0-8047-6915-X (cloth), 0-8047-6916-8 (paperback). ???? pp. LCCN QC16.S265 W4313 2011.

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gowing:1979:AB

[GA79] Margaret Gowing and Lorna Arnold. The atomic bomb. Science in a social context. Butterworths, London, UK; Boston, MA, USA, 1979. ISBN 0-408-71311-9. 56 pp. LCCN QC773 .G66.

Graetzer:1981:DNF

[GA81] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission. The Development of science. Arno Press, New York, NY, USA, 1981. ISBN 0-405-13846-6. viii + 120 + 2 pp. LCCN QC790 .G68 1981.

Gembillo:2002:WHS

[GA02] Giuseppe Gembillo and Costanza Altavilla, editors. Werner Heisenberg: scienzato e filosofo. (Italian) [Werner Heisenberg: Scientist and Philosopher], volume 14 of Interazioni. A. Siciliano, Messina, Italy, 2002. 291 pp. LCCN QC16.H35 W473 2002.

Gabor:1995:EWW

[Gab95] Andrea Gabor. Einstein’s wife: work and marriage in the lives of five great twentieth-century women. Viking, New York, NY, USA, 1995. ISBN 0-670-84210-9 (hardcover), 0-14-015993-2 (paperback). xxiv + 341 pp. LCCN HQ759 .G3 1995.

Gundlach:1996:NTM

[GAHS96] J. H. Gundlach, E. G. Adelberger, B. R. Heckel, and H. E. Swanson. New technique for measuring Newton’s constant G. Physical Review D (Particles and Fields), 54(2):R1256–R1259, ???? 1996. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500. URL http://lheawww.gsfc.

nasa.gov/users/merk/G/Big_G.html;   http://link.aps.
org/doi/10.1103/PhysRevD.54.R1256;   http://prd.aps.
org/abstract/PRD/v54/i2/pR1256_1.

Gallant:1979:SGA

[Gal79] R. A. Gallant. Spinoza, God, and Albert Einstein. Omni, 1(11): 136, ???? 1979. ISSN 0149-8711.

Galison:1982:TPE

[Gal82] Peter Galison. Theoretical predispositions in experimental physics: Einstein and the gyromagnetic experiments, 1915– 1925. Historical Studies in the Physical Sciences, 12(2):285–323, ???? 1982. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757498.

Gallagher:1993:AGZ

[Gal93] Carole Gallagher. American ground zero: the secret nuclear war. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-07146-0. xxxiii + 427 pp. LCCN U264.4.N3 G35 1993.

Galbraith:1994:KES

[Gal94] James K. Galbraith. Keynes, Einstein and scientific revolution. The American Prospect, 5(16):62–67, Winter 1994. CODEN APROEY. ISSN 1049-7285. URL http://virtualschool.

edu/mon/Economics/GalbraithKeynesAndEinstein.html.

Galaburda:1999:AEB

[Gal99] A. M. Galaburda. Albert Einstein’s brain. The Lancet (London, England), 354(9192):1821, November 20, 1999. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547X (electronic).

Galison:2000:ECP

[Gal00] Peter Galison. Einstein’s clocks: The place of time. Critical Inquiry, 26(2):355–389, Winter 2000. ISSN 0093-1896 (print),

1539-7858 (electronic). URL http://www.jstor.org/stable/
1344127.

Galison:2003:ECP

[Gal03] Peter Galison. Einstein’s Clocks and Poincaré’s Maps: Empires of Time. W. W. Norton & Co., New York, NY, USA, 2003. ISBN 0-393-02001-0 (hardcover), 0-393-32604-7 (paperback). 389 pp. LCCN QB209 .G35 2003. US$23.95.

Galison:2004:ECP

[Gal04a] Peter Galison. Einstein’s clocks, Poincaré’s maps: empires of time. W. W. Norton & Co., New York, NY, USA, 2004. ISBN 0-393-32604-7. 389 pp. LCCN QB209 .G35 2003.

Galison:2004:EC

[Gal04b] Peter Galison. Einstein’s compass. Scientific American, 291(3):66–69, September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v291/n3/
pdf/scientificamerican0904-66.pdf.

GalindoTixaire:2007:CEA

[Gal07] Alberto Galindo Tixaire. Cosmology and Einstein (Albert Einstein). Arbor — Ciencia Pensamiento y Cultura, 183(728):869– 876, November/December 2007. ISSN 0210-1963.

Gamow:1942:LHT

[Gam42a] G. Gamow. A layman’s history of the theory of Relativity. (books on science for laymen: From Copernicus to Einstein). The Scientific Monthly, 55:569–??, December 1942. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://adsabs.harvard.edu/abs/1942SciMo..55..569R.

Gamow:1942:RLH

[Gam42b] George Gamow. Review: a layman’s history of the theory of relativity: From Copernicus to Einstein by Hans Reichenbach. The Scientific Monthly, 55(6):569–??, December 1942. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/pdfplus/17860.pdf.

Gamow:1949:RC

[Gam49] G. Gamow. On relativistic cosmogony. Reviews of Modern Physics, 21(3):367–373, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.367;
http://rmp.aps.org/abstract/RMP/v21/i3/p367_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Gamow:1961:BPa

[Gam61a] George Gamow. Biography of Physics, volume TB567 of Harper torchbooks. The Cloister library. Harper, New York, NY, USA, 1961. 338 pp. LCCN QC7 .G263 1964.

Gamow:1961:BP

[Gam61b] George Gamow. Biography of physics, volume TB567 of Harper torchbooks. The Cloister library. Harper, New York, NY, USA, 1961. 338 pp. LCCN QC7 .G263 1964.

Gamow:1966:TYSa

[Gam66a] George Gamow. Thirty years that shook physics; the story of quantum theory. Doubleday, Garden City, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3 1966b.

Gamow:1966:TYSb

[Gam66b] George Gamow. Thirty years that shook physics; the story of quantum theory, volume S45 of Science study series. Anchor Books, New York, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3.

Gamow:1972:TYS

[Gam72] George Gamow. Thirty years that shook physics; the story of quantum theory, volume 38 of Science study series. Heinemann Educational, London, UK, 1972. ISBN 0-435-55071-3. viii + 224 + 12 pp. LCCN QC173.98 .G35 1972.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Gamow:1988:GPG

[Gam88] George Gamow. The great physicists from Galileo to Einstein. Dover, New York, NY, USA, 1988. ISBN 0-486-25767-3. vi + 338 pp. LCCN QC7 .G27 1988; QC7 .G14b 1988.

Ganley:1997:NIC

[Gan97] W. T. Ganley. A note on the intellectual connection between Albert Einstein and Thorstein Veblen. Journal of Economic Issues, 31(1):245–251, March 1997. ISSN 0021-3624.

Garbedian:1939:AEM

[Gar39] H. Gordon (Haig Gordon) Garbedian. Albert Einstein, maker of universes. Funk and Wagnalls Company, New York, NY, USA, 1939. x + 1 + 328 pp. LCCN QC16.E5 G3.

Garbedian:1940:EHU

[Gar40] Haig Gordon Garbedian. Einstein: hacedor de universos. (Spanish) [Albert Einstein, maker of universes]. Editorial Losada, S.A., Buenos Aires, Argentina, 1940. 2 + 7–292 pp. LCCN QC16.E5 G36. Spanish translation by Felipe Jiménez de Asúa of [Gar39].

Garbedian:1941:AES

[Gar41] Haig Gordon Garbedian. Albert Einstein: skaparen av en ny världsbild. (Swedish) [Albert Einstein: Maker of a new world view]. Natur och kultur, Stockholm, Sweden, 1941. ???? pp. LCCN QC16.E5 G37. Swedish translation by Oskar Klein of [Gar39].

Garbedian:1942:ECU

[Gar42] Haig Gordon Garbedian. Einstein, o criador de universos. (Portuguese) [Albert Einstein, maker of universes]. J. Olympio, Rio de Janeiro, Brazil, 1942. 2 + 3–367 + 1 pp. LCCN QC16.E5 G35. Portuguese translation by Julio Cesar de Mello e Souza of [Gar39].

Gardner:1962:RMC

[Gar62] Martin Gardner. Relativity for the million. Macmillan Publishing Company, New York, NY, USA, 1962. 182 pp. LCCN QC6 .G28. Illustrated by Anthony Ravielli.

Garber:1970:CMK

[Gar70] Elizabeth Wolfe Garber. Clausius and Maxwell’s Kinetic Theory of Gases. Historical Studies in the Physical Sciences, 2(??):299– 319, ???? 1970. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757309.

Gardner:1972:QTR

[Gar72] Michael R. Gardner. Quantum-theoretical realism: Popper and Einstein v. Kochen and Specker. The British Journal for the Philosophy of Science, 23(1):13–23, February 1972.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
23/1/13.full.pdf+html.

Gardner:1976:RE

[Gar76] Martin Gardner. The Relativity explosion. Vintage Books, New York, NY, USA, 1976. ISBN 0-394-72104-7. xi + 198 pp. LCCN QC173.57 .G37 1976.

GarciaCamarero:1977:AE

[Gar77] Ernesto García Camarero, editor. Albert Einstein. Caminos abiertos. Editorial Hernando, Madrid, Spain, 1977. ISBN 84-7155-245-0. 160 pp. LCCN QC16.E5 A64. 300ptas.

Gardner:1993:CMA

[Gar93] Howard Gardner. Creating minds: an anatomy of creativity seen through the lives of Freud, Einstein, Picasso, Stravinsky, Eliot, Graham, and Gandhi. Basic Books, New York, NY, USA, 1993. ISBN 0-465-01455-0, 0-465-01454-2. xvi + 464 pp. LCCN BF408 .G33 1993. URL http://www.loc.gov/

catdir/enhancements/fy0830/92056172-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0830/92056172-d.
html.

Garwin:2008:LNW

[Gar08] Richard L. Garwin. Living with nuclear weapons: Sixty years and counting. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 152(1):69– 82, March 2008. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://www.jstor.org/stable/
25478469.

Gaspar:1954:AHF

[Gás54] R. Gáspár. Über eine Approximation des Hartree–Fockschen Potentials Durch eine Universelle Potentialfunktion. (German) [On an approximation of the Hartree–Fock potential by a universal potential function]. Acta physica Academiae Scientiarum Hungaricae, 3(3–4):263–286, April 1954. CODEN APA-

HAQ. ISSN 0001-6705. URL http://www.springerlink.
com/content/w24337x7040846x5/. In this paper, the Hartree–

Fock–Slater X-α equations [Sla51, Sla72] are rederived by an independent method, with the prediction of a value α = 2/3, instead of Slater’s original α = 1. The same factor was later rediscovered in independent work [KS65].

Gasco:2005:PMP

[Gas05] E. Gasco. Il Principio di Mach: le prime considerazioni di Einstein (1907–12). (Italian) [Mach’s Principle: Einstein’s main consideration (1907–12)]. Quaderni di Storia della fisica, 13: 75–92, 2005. CODEN ???? ISSN 1594-9974 (print), 1827-

6164 (electronic). URL http://en.sif.it/journals/qsf/
econtents.

Gates:2009:ETH

[Gat09] Evalyn Gates. Einstein’s telescope: the hunt for dark matter and dark energy in the universe. W. W. Norton & Co., New York, NY, USA, 2009. ISBN 0-393-06238-4. xiv + 305 pp. LCCN QB791.3 .G38 2009.

Gavroglou:1995:FLS

[Gav95] Kostas Gavroğlou. Fritz London: a scientific biography. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-43273-1. xxiv + 299 pp. LCCN QC16.L645 G38

1995.  URL http://www.loc.gov/catdir/description/
cam026/94000691.html;   http://www.loc.gov/catdir/toc/
cam021/94000691.html.

Gottfried:1988:CSN

[GB88] Kurt Gottfried and Bruce G. Blair, editors. Crisis, Stability and Nuclear War. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-505146-7, 0-19-505147-5 (paperback). xi + 354 pp. LCCN U263

.C75 1988. URL http://www.loc.gov/catdir/enhancements/
fy0636/88012552-d.html;   http://www.loc.gov/catdir/
enhancements/fy0636/88012552-t.html;   http://www.loc.
gov/catdir/enhancements/fy0724/88012552-b.html.

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2):267– 347, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print),

1533-8355 (electronic). URL http://www.jstor.org/stable/
27757627.

Galison:2003:EPM

[GB03] Peter L. Galison and D. Graham Burnett. Einstein, Poincaré & modernity: A conversation. Dædalus, 132(2):41–55, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.org/stable/20027839.

Goldsmith:2006:EHL

[GB06] Donald Goldsmith and Marcia Bartusiak, editors. E = Einstein: his life, his thought and his influence on our culture. Sterling Pub. Co., New York, NY, USA, 2006. ISBN 1-4027-3787-4. 352 pp. LCCN QC16.E5 E18 2006. URL http://www.loc.

gov/catdir/enhancements/fy0709/2007271614-d.html;
http://www.loc.gov/catdir/toc/fy0706/2007271614.html.

Goldsmith:2008:EHL

[GB08] Donald Goldsmith and Marcia Bartusiak, editors. E = Einstein: his life, his thought and his influence on our culture. Sterling Pub. Co., New York, NY, USA, 2008. ISBN 1-4027-6319-0. 352 pp. LCCN QC16.E5 E18 2008. URL http://catdir.loc.

gov/catdir/enhancements/fy0709/2007271614-d.html;
http://catdir.loc.gov/catdir/toc/fy0706/2007271614.
html.

Geoffrey-Brumfiel:2011:ESY

[GB11] Geoffrey-Brumfiel. Einstein starts your car. Nature News Blog, January 12, 2011. CODEN ???? ISSN ???? URL http:/

/blogs.nature.com/news/2011/01/einstein_starts_your_
car_1.html.

Ghatak:2015:HDQ

[GB15] Kamakhya Prasad Ghatak and Sitangshu Bhattacharya. Heavily-doped 2D-quantized structures and the Einstein relation, volume 260 of Springer tracts in modern physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-08379-1, 3-319-08380-5 (e-book). xl + 347 pp. LCCN QC176-176.9. URL http://link.springer.

com/openurl?genre=book\%26isbn=978-3-319-08379-7.

Garber:1995:MHS

[GBE95] Elizabeth Garber, Stephen G. Brush, and C. W. F. (C. W. Francis) Everitt. Maxwell on heat and statistical mechanics: on “avoiding all personal enquiries” of molecules. Lehigh University Press, Bethlehem, PA, USA, 1995. ISBN 0-934223-34-3. 550 pp. LCCN QC310.2 .M39 1995.

Goenner:1996:AEP

[GC96a] H. Goenner and G. Castagnetti. Albert Einstein as pacifist and Democrat during World War I. Science in Context, 9(4):325– 386, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goenner:1996:EPD

[GC96b] Hubert Goenner and Giuseppe Castagnetti. Einstein as pacifist and democrat during World War I. Science in Context, 9(4): 325–386, December 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Garwin:2002:MMF

[GC02] Richard L. Garwin and Georges Charpak. Megawatts and megatons: the future of nuclear power and nuclear weapons. University of Chicago Press, Chicago, IL, USA and London, UK, 2002. ISBN 0-226-28427-1 (paperback). xvii + 412 pp. LCCN QC792 .C4713 2002. URL http://catdir.

loc.gov/catdir/bios/uchi051/2002027143.html;   http:
//catdir.loc.gov/catdir/description/uchi051/2002027143.
html;   http://catdir.loc.gov/catdir/toc/uchi051/2002027143.
html.

Goenner:2020:EQP

[GC20] Hubert Goenner and Giuseppe Castagnetti. Establishing quantum physics in Berlin: Einstein and the Kaiser Wilhelm Institute for Physics, 1917–1922. SpringerBriefs in history of science and technology. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-63121-4, 3-030-63122-2 (e-book), 3-030-63123-0. ISSN 2211-4564. vi + 127 pp. LCCN QC16.E5 G64 2020.

Goldhaber:2006:VEM

[GCO^{+}06] Alfred Scharff Goldhaber, Tom Cornsweet, Hans C. Ohanian, Ravi Gomatam, Ron Larson, Brian C. Hall, Roger G. Newton, Bob Eisenberg, Robert E. Becker, and Steven Weinberg. The value of Einstein’s mistakes. Physics Today, 59(4):10–16, April 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://scitation.aip.org/journals/
doc/PHTOAD-ft/vol_59/iss_4/10_1.shtml. See [Wei05a].

Glasstone:1977:ENW

[GD77] Samuel Glasstone and Philip J. Dolan, editors. The effects of nuclear weapons. U.S. Dept. of Defense, Washington, DC, USA, third edition, 1977. 653 pp. LCCN UF767 .E33 1977.

Guy:1990:FEE

[GD90] Reed Guy and Robert Deltete. Fine, Einstein, and ensembles. Foundations of Physics, 20(8):943–965, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00738374.

Gehlen:2015:EFC

[GdS15] Juliana Marques Tejkowski Gehlen and Antônio V. dos Santos. Einstein e Freud: contribuiçäoes para o ensino da Física Moderna: Uma nova mecânica de aula. (Portuguese) [Einstein and Freud: contributions to the teaching of modern physics: A new mechanics class]. Novas Ediçäoes Acadêmicas, Saarbrücken, Germany, 2015. ISBN 3-639-75770-X. ???? pp. LCCN ????

Goldman:1997:EGA

[GE97] Robert N. Goldman and Albert Einstein. Einstein’s God: Albert Einstein’s quest as a scientist and as a Jew to replace a forsaken God. Jason Aronson, Northvale, NJ, USA, 1997. ISBN 1-56821-983-0. xviii + 166 pp. LCCN QC16.E5 G67 1997.

Gearhart:1990:EBE

[Gea90] Clayton A. Gearhart. Einstein before 1905: The early papers on statistical mechanics. American Journal of Physics, 58(5): 468–480, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/58/5/10.1119/1.16478.

Gearhart:1994:BRB

[Gea94] Clayton A. Gearhart. Book review: The Private Lives of Albert Einstein, by Roger Highfield and Paul Carter. American Journal of Physics, 62(11):1054–1055, November 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/62/1054/1.

Gearhart:2002:PQH

[Gea02] Clayton A. Gearhart. Planck, the quantum, and the historians. Physics in Perspective (PIP), 4(2):170–215, May 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-

002-8363-7. See also [Bad09].

Gefter:2002:TEL

[Gef02] Amanda Gefter. Throwing Einstein for a loop. Scientific American, 287(6):40–41, December 2002. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v287/n6/
pdf/scientificamerican1202-40.pdf.

Gefter:2014:TEL

[Gef14] Amanda Gefter. Trespassing on Einstein’s lawn: a father, a daughter, the meaning of nothing, and the beginning of everything. Bantam Books, New York, NY, USA, 2014. ISBN 0-345-53143-4 (hardcover), 0-345-53963-X (e-book). 418 pp. LCCN QC6 .G34 2014.

Geim:2011:NLR

[Gei11] Andre K. Geim. Nobel Lecture: Random walk to graphene. Reviews of Modern Physics, 83(3):851–862, July 2011. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.83.851;   http://rmp.aps.org/abstract/RMP/
v83/i3/p851_1.

Geisler:2015:WLG

[Gei15] Harald Geisler. Write like a genius: New font released on centennial of Einstein’s general theory of relativity. Scientific Computing, ??(??):??, May 7, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.scientificcomputing.com/news/2015/05/
write-genius-new-font-released-centennial-einsteins-
general-theory-relativity.

Genese:1923:EPb

[Gen23a] R. W. Genese. An Einstein paradox. Nature, 111(2796): 742, June 2, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v111/n2796/pdf/111742a0.pdf.

Genese:1923:EPc

[Gen23b] R. W. Genese. An Einstein paradox. Nature, 111(2800): 880, June 30, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v111/n2800/pdf/111880a0.pdf.

Genese:1923:EPA

[Gen23c] R. W. Genese. An Einstein paradox: an apology. Nature, 112 (2808):283, August 25, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v112/n2808/pdf/112283a0.pdf.

Genovesi:1998:LHB

[Gen98] A. Genovesi. Letters from Henri Bergson to Albert Einstein — introduction, transcription, Italian translation and notes. Filosofia, 49(1):3–41, January/April 1998. ISSN 0015-1823.

Genzel:2005:GMS

[Gen05] Reinhard Genzel. Galaxien und massive Schwarze Löcher. (German) [Galaxies and massive black holes]. In Steiner [Ste05a], pages 123–132. ISBN 3-540-21060-1. LCCN QC16.E5 A659

2005. URL http://link.springer.com/chapter/10.1007/3-
540-30595-5_5/.

Genzel:2022:NLF

[Gen22] Reinhard Genzel. Nobel Lecture: A forty-year journey. Reviews of Modern Physics, 94(2):020501–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.020501.

George:1953:LBP

[Geo53] André George. Louis de Broglie, physicien et penseur (French) [Louis de Broglie, physicist and thinker]. Editions Albin Michel, Paris, France, 1953. xi + 497 pp. LCCN QC71 .L88.

Gerlach:1969:OHF

[Ger69] Walther Gerlach. Otto Hahn, ein Forscherleben unserer Zeit. (German) [Otto Hahn: a researcher’s life in our time]. Deutsches Museum. Abhandlungen und Berichte, 37. Jahrg., 1969, Heft 3. R. Oldenbourg, München, West Germany, 1969. 84 pp. LCCN AM101 .M9743.

Geroch:1978:GRB

[Ger78] Robert Geroch. General relativity from A to B. Phoenix book. University of Chicago Press, Chicago, IL, USA and London, UK, 1978. ISBN 0-226-28864-1 (paperback), 0-226-28863-3. xi + 225 pp. LCCN ????

Gerlach:1979:RAE

[Ger79] Walther Gerlach. Reminiscences of Albert Einstein from 1908 to 1930. In Aichelburg and Sexl [AS79], pages 189–200. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_14/.

Gerber:1992:HFC

[Ger92] Michele Stenehjem Gerber. On the home front: the Cold War legacy of the Hanford Nuclear Site. University of Nebraska Press, Lincoln, NE, USA, 1992. ISBN 0-8032-2145-2. 312 pp. LCCN TD898.12.W2 G47 1992. URL http://www.loc.gov/

catdir/enhancements/fy0728/92008746-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0728/92008746-d.
html.

Gerber:1997:HFC

[Ger97] Michele Stenehjem Gerber. On the home front: the Cold War legacy of the Hanford Nuclear Site. University of Nebraska Press, Lincoln, NE, USA, 1997. ISBN 0-8032-7068-2 (paperback), 0-585-25712-4 (e-book). 334 + 12 pp. LCCN TD898.12.W2 G47 1992.

Gerber:2002:HFC

[Ger02]  Michele Stenehjem Gerber.  On the home front: the

Cold War legacy of the Hanford Nuclear Site. University of Nebraska Press, Lincoln, NE, USA, second edition, 2002. ISBN 0-8032-7101-8 (paperback). 363 pp. LCCN

TD898.12.W2 G47 2002. URL http://www.loc.gov/catdir/
enhancements/fy0727/2001027956-b.html;   http://www.
loc.gov/catdir/enhancements/fy0727/2001027956-d.html.

Gerber:2007:HFC

[Ger07] Michele Stenehjem Gerber. On the Home Front: the Cold War Legacy of the Hanford Nuclear Site. University of Nebraska Press, Lincoln, NE, USA, third edition, 2007. ISBN 0-8032-5995-6 (paperback). x + 391 pp. LCCN

TD898.12.W2 G47 2007. URL http://www.loc.gov/catdir/
enhancements/fy0708/2007004538-b.html;   http://www.
loc.gov/catdir/enhancements/fy0708/2007004538-d.html;

 http://www.loc.gov/catdir/toc/ecip0710/2007004538.
html.

Germann:2015:BRE

[Ger15] C. B. Germann. Book review: Einstein and the Quantum: The Quest of the Valiant Swabian by by A. Douglas Stone. Leonardo (Oxford, England), 28(2):208–209, ???? 2015. CO-DEN LEONDP. ISSN 0024-094X (print), 1530-9282 (elec-

tronic). URL https://muse.jhu.edu/article/577891.

Getler:2006:EWR

[Get06] Michael Getler. Einstein’s wife: The relative motion of “facts”. PBS Ombudsman column, December 15, 2006. URL

http://www.pbs.org/ombudsman/2006/12/einsteins_wife_
the_relative_motion_of_facts.html.

Giovannitti:1965:DDB

[GF65] Len Giovannitti and Fred Freed. The decision to drop the bomb. Coward-McCann, New York, NY, USA, 1965. 348 pp. LCCN UA23 .G62.

Giovannitti:1967:DDB

[GF67] Len Giovannitti and Fred Freed. The decision to drop the bomb. Methuen and Co. Ltd., London, UK, 1967. 384 pp. LCCN UA23 .G62 1967.

Galvagno:2005:PPC

[GG05a] Mariano Galvagno and Gastón Giribet. The particle problem in classical gravity: a historical note on 1941. European Journal of Physics, 26(6):S97, 2005. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.iop.org/ 0143-0807/26/i=6/a=S03. Comments on [Ein41a, EP43].

Gribbin:2005:AMA

[GG05b] John R. Gribbin and Mary Gribbin. Annus mirabilis: 1905, Albert Einstein, and the Theory of Relativity. Chamberlain Brothers, New York, NY, USA, 2005. ISBN 1-59609-144-4. viii + 310 pp. LCCN QC16.E5 G73 2005.

Gribbin:2015:EMG

[GG15] John Gribbin and Mary Gribbin. Einstein’s masterwork: 1915 and the general theory of relativity. Icon, London, UK, 2015. ISBN 1-84831-852-9 (hardcover), 1-84831-854-5 (e-book). xiii + 226 pp. LCCN QC173.6 .G74 2015.

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge, London, UK, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Galison:2013:RSR

[GGK13] Peter Galison, Michael D. Gordin, and David Kaiser. The Roots of Special Relativity: Science and Society. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2013. ISBN 1-136-

70916-9. 418 pp. URL http://books.google.com/books?id=
ocnbAAAAQBAJ.

Gnaiger:1994:WCL

[GGW94] E. (Erich) Gnaiger, Frank N. Gellerich, and Markus Wyss, editors. What is controlling life?: 50 years after Erwin Schrödinger’s “What is Life?”, volume 205 of Modern trends in biothermokinetics (Innsbruck, Austria) (vol. 3). Innsbruck University Press, Innsbruck, Austria, 1994. ISBN 3-901249-17-6. 335 pp. LCCN QP517.B54 W48 1994.

Gerlach:1984:OHF

[GH84] Walther Gerlach and Dietrich Hahn. Otto Hahn, 1879–1968: ein Forscherleben unserer Zeit. (German) [Otto Hahn, 1879–1968: a researcher’s life in our time], volume 45 of Grosse Naturforscher. Wissenschaftliche Verlagsgesellschaft, Stuttgart, West Germany, 1984. ISBN 3-8047-0757-2. ISSN 0072-7741. 267 pp. LCCN QD22.H2 G47 1984.

Gibbons:1993:EQG

[GH93] G. W. Gibbons and S. W. (Stephen W.) Hawking, editors. Euclidean quantum gravity. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-0515-5, 981-02-0516-3 (paperback). xiii + 586 pp.

LCCN QC178 .E93 1993. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0874.53056.

Glick:2001:SPR

[GH01] Thomas F. Glick and Mark G. Henderson. The scientific and popular receptions of Darwin, Freud, and Einstein. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 14, pages 229–238. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-010-0602-6_14.

Gron:2007:EGT

[GH07] Øyvind Grøn and Sigbjørn Hervik. Einstein’s General theory of Relativity: with modern applications in cosmology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-69199-5, 0-387-69200-2 (e-book). xx + 538 pp. LCCN QC173.55 .O85 2007.

Ghani:1982:ASN

[Gha82] Abdul Ghani. Abdus Salam: a Nobel laureate from a Muslim country: a biographical sketch. Maaref [??] Printers, Karachi, Pakistan, 1982. ISBN ???? xv + 234 + 8 pp. LCCN QC16.S26 G45 1982.

Ghatak:2014:EP

[Gha14] Kamakhya Prasad Ghatak. Einstein’s photoemission. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-11187-6, 3-319-11188-4 (e-book). xxxviii + 495? pp. LCCN ???? URL http://link.springer.

com/openurl?genre=book\%26isbn=978-3-319-11187-2;
http://www.loc.gov/catdir/enhancements/fy1501/2014953237-
d.html;   http://www.loc.gov/catdir/enhancements/fy1501/
2014953237-t.html.

Ghassib:2017:RME

[Gha17] Hisham Ghassib. Reason and method in Einstein’s relativity. International Studies in the Philosophy of Science, 31(4):331– 342, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565204.

Ghirardi:2005:SLG

[Ghi05] G. C. Ghirardi. Sneaking a Look at God’s Cards: Unraveling the Mysteries of Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, revised edition, 2005. ISBN 0-691-04934-3, 0-691-12139-7, 0-691-13037-X. xix + 488 pp. LCCN Q173 .G4813 2004.

Goldberg:1996:EKJ

[GHK^{+}96] Stanley Goldberg, Dieter Hoffmann, Alexei B. Kojewnikow, Fritz Krafft, and Helmut Rechenberg. Die erste Kernwaffendetonation am 16. Juli 1945 in Alamogordo, New Mexico — Vorgeschichte, Ereignis, Wirkungen, Öffentliche Podiumsdiskussion. (German) [The first nuclear weapon detonation on 16 July 1945 in Alamogordo, New Mexico — history, event, effects: public panel discussion]. Berichte zur Wissenschaftsgeschichte, 19(2–3):157–182, ???? 1996. CO-DEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/bewi.

19960190214/abstract. See [Gol96] for Part II.

Gabrielse:2006:NDF

[GHK^{+}06] G. Gabrielse, D. Hanneke, T. Kinoshita, M. Nio, and B. Odom. New determination of the fine structure constant from the electron g value and QED. Physical Review Letters, 97(3):030802:1– 030802:4, July 2006. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.97.030802. See erratum

[GHK^{+}07] that revises the value of the fine-structure constant.

Gabrielse:2007:END

[GHK^{+}07] G. Gabrielse, D. Hanneke, T. Kinoshita, M. Nio, and B. Odom. Erratum: New Determination of the Fine Structure Constant from the Electron g Value and QED [Phys. Rev. Lett. 97, 030802 (2006)]. Physical Review Letters, 99(3):039902:1– 039902:2, July 2007. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:/

/link.aps.org/doi/10.1103/PhysRevLett.99.039902. See

[GHK^{+}06].

Ghose:2017:ETN

[Gho17] Partha Ghose, editor. Einstein, Tagore, and the nature of reality, volume 2 of Routledge studies in the philosophy of mathematics and physics. Routledge, Taylor and Francis Group, London, UK, 2017. ISBN 1-138-68524-0 (hardcover), 1-134-85934-1 (e-book). xviii + 237 pp. LCCN BD331 .E425 2017.

Galison:2008:ECH

[GHS08] Peter Galison, Gerald James Holton, and S. S. (Silvan S.) Schweber, editors. Einstein for the 21st Century: His Legacy in Science, Art, and Modern Culture. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13520-7 (hardcover). xviii + 363 + 16 pp. LCCN QC16.E5 E446 2008.

Galison:2018:ECH

[GHS18] Peter Galison, Gerald James Holton, and S. S. (Silvan S.) Schweber, editors. Einstein for the 21st Century: His Legacy in Science, Art, and Modern Culture. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17790-2 (paperback). xviii + 363 + 16 pp. LCCN QC16.E5 E446 2008.

Gomori:2021:MER

[GHS21] Márton Gömöri and Gábor Hofer-Szabó. On the meaning of EPR’s reality criterion. Synthese, 199(5–6):13441–13469, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/
article/10.1007/s11229-021-03382-3.

Greenberger:1990:BTI

[GHSZ90] Daniel M. Greenberger, Michael A. Horne, Abner Shimony, and Anton Zeilinger. Bell’s theorem without inequalities. American Journal of Physics, 58(12):1131–143, December 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/ajpias/v58/i12/p1131/s1.

Greenberger:2009:CQP

[GHW09] Daniel M. Greenberger, Klaus Hentschel, and Friedel Weinert, editors. Compendium of quantum physics: concepts, experiments, history and philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-70622-4, 3-540-70626-7 (e-book). xvi + 901 pp. LCCN C174.12 .C66 2009.

Giannoni:1970:ELP

[Gia70] Carlo Giannoni. Einstein and the Lorentz–Poincaré Theory of Relativity. PSA: Proceedings of the Biennial Meeting of the Philosophy of Science Association, 1970(??):575–589, 1970. ISSN 0270-8647 (print), 2327-9486 (electronic). URL http://www.

jstor.org/stable/495789.

Giannoni:1971:ELP

[Gia71] Carlo Giannoni. Einstein and the Lorentz–Poincaré theory of relativity. In Buck and Cohen [BC71], pages 575–589. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-010-3142-4_39.

Giacconi:1980:EXR

[Gia80] Riccardo Giacconi. The Einstein X-Ray Observatory. Scientific American, 242(2):80–102, February 1980. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v242/n2/
pdf/scientificamerican0280-80.pdf.

Giacconi:2003:NLD

[Gia03] Riccardo Giacconi. Nobel Lecture: The dawn of X-ray astronomy. Reviews of Modern Physics, 75(3):995–1010, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.995;   http://rmp.aps.org/
abstract/RMP/v75/i3/p995_1.

Giangreco:2009:HPO

[Gia09] D. M. Giangreco. Hell to pay: Operation Downfall and the invasion of Japan, 1945–47. Naval Institute Press, Annapolis, MD, USA, 2009. ISBN 1-59114-316-0. xxiii + 362 pp.

LCCN D767.2. URL http://www.gbv.de/dms/bowker/toc/
9781591143161.pdf.

Gibbons:1991:PET

[Gib91a] Ann Gibbons. Putting Einstein to the test — in space: Gravity Probe B — an audacious experiment conceived 30 years ago to test General Relativity — nears the launching pad. Science, 254 (5034):939–941, November 15, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/254/5034/939.full.pdf.

Gibilisco:1991:UET

[Gib91b] Stan Gibilisco. Understanding Einstein’s theories of relativity: man’s new perspective on the cosmos. Dover, New York, NY, USA, 1991. ISBN 0-486-26659-1 (paperback). viii + 200 pp. LCCN QC173.57 .G5 1991. US$6.95. URL http://www.loc.

gov/catdir/description/dover033/90020570.html.

Gibson:1996:NWU

[Gib96] James N. Gibson. Nuclear weapons of the United States: an illustrated history. Schiffer military history. Schiffer Publishing, Atglen, PA, USA, 1996. ISBN 0-7643-0063-6. 236 pp. LCCN U264.3 .G53 1996.

Gibbs:1997:NE

[Gib97] W. Wayt Gibbs. Newton 1, Einstein 0. Scientific American, 277(6):40–42, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v277/n6/
pdf/scientificamerican1297-40.pdf.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Gieser:2005:IKD

[Gie05] Suzanne Gieser. The innermost kernel: depth psychology and quantum physics: Wolfgang Pauli’s dialogue with C. G. Jung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-20856-9 (hardcover), 3-540-26986-X. xii + 378 pp. LCCN QC15 .G52 2005. URL

http://www.loc.gov/catdir/toc/fy0614/2004113652.html.

Gilpin:1962:ASN

[Gil62] Robert Gilpin. American scientists and nuclear weapons policy. Princeton University Press, Princeton, NJ, USA, 1962. viii + 352 pp. LCCN UA23 .G58.

Gilbert:1980:EE

[Gil80] Felix Gilbert. Einstein’s Europe. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Gilbert:1984:AEH

[Gil84a] F. Gilbert. Albert Einstein, historical and cultural perspectives — the Centennial Symposium in Jerusalem — G. Holton, Y. Elkana. The Journal of Modern History, 56(1):129–133, ???? 1984. ISSN 0022-2801 (print), 1537-5358 (electronic).

Gilbert:1984:SLS

[Gil84b] F. Gilbert. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. The Journal of Modern History, 56(1):129–133, ???? 1984. ISSN 0022-2801 (print), 1537-5358 (electronic).

Gilman:2005:EVJ

[Gil05] S. L. Gilman. Einstein’s violin: Jews and the performance of identity. Modern Judaism, 25(3):219–236, October 2005. CO-DEN ???? ISSN 0276-1114 (print), 1086-3273 (electronic). URL

http://muse.jhu.edu/article/187438.

Gilder:2008:AEW

[Gil08] Louisa Gilder. The Age of Entanglement: When Quantum Physics Was Reborn. Alfred A. Knopf, New York, NY, USA, 2008. ISBN 1-4000-9526-3, 1-4000-4417-0. xvi + 443 pp. LCCN QC174.12 .G528 2008.

Gilder:2009:AEW

[Gil09] Louisa Gilder. The Age of Entanglement: When Quantum Physics Was Reborn. Vintage Books, New York, NY, USA, 2009. ISBN 1-4000-9526-3. xviii + 443 pp. LCCN QC174.12 .G528 2009.

Gillies:2012:BRE

[Gil12] Donald Gillies. Book review: Erling Norrby, Nobel Prizes and Life Sciences. London and Singapore: World Scientific Press, 2010. Pp. xvi + 317. ISBN 978-981-4299-37-4. £25.00 (paperback). British Journal for the History of Science, 45(1):142–143, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Nor10].

Gilmore:2019:BRF

[Gil19] Gerard Gilmore. Book review: The formative years of relativity: the history and meaning of Einstein’s Princeton lectures. Contemporary Physics, 60(2):208–209, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gimbel:2012:EJS

[Gim12] Steven Gimbel. Einstein’s Jewish science: physics at the intersection of politics and religion. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0554-7 (hardcover), 1-4214-0575-X (e-book). viii + 245 pp. LCCN QC173.55 .G55 2012.

Gimbel:2015:EHS

[Gim15] Steven Gimbel. Einstein: his space and times. Jewish lives. Yale University Press, New Haven, CT, USA, 2015. ISBN 0-300-19671-7 (hardcover), 0-300-21361-1 (e-book). ix + 191 pp. LCCN QC16.E5 G53 2015.

Ginzburg:1959:AST

[Gin59] V. L. Ginzburg. Artificial satellites and the Theory of Relativity. Scientific American, 200(5):149–160, May 1959. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

https://www.math.utah.edu/pub/tex/bib/einstein.bib;
 http://www.nature.com/scientificamerican/journal/v200/
n5/pdf/scientificamerican0559-149.pdf.

Ginzburg:1962:EVG

[Gin62] V. L. Ginzburg. Experimental verifications of the general theory of relativity. In Anonymous [Ano62], pages 57–71. LCCN QC 6 .R295. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Gingerich:1982:BRE

[Gin82] Owen Gingerich. Book review: Einstein’s special theory analysed: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911). Journal for the History of Astronomy, 13(2):135, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ginzburg:2001:PLR

[Gin01] V. L. (Vitaliĭ Lazarevich) Ginzburg. The Physics of a Lifetime: Reflections on the Problems and Personalities of 20th Century Physics. Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67534-5, 3-642-08699-3. xiii + 513

pp. LCCN QC7 .G59 2001. URL http://catdir.loc.gov/
catdir/enhancements/fy0812/00066112-d.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy0812/00066112-
t.html.

Ginzburg:2004:NLS

[Gin04] Vitaly L. Ginzburg. Nobel Lecture: On superconductivity and superfluidity (what I have and have not managed to do) as well as on the “physical minimum” at the beginning of the XXI century. Reviews of Modern Physics, 76(3):981–998, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.981;   http://rmp.aps.org/
abstract/RMP/v76/i3/p981_1.

Gingras:2008:CCS

[Gin08] Yves Gingras. The collective construction of scientific memory: The Einstein–Poincaré connection and its discontents, 1905– 2005. History of Science (UK), 46(1):75–114, March 2008. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/46/1/75.full.pdf+
html.

Gingras:2015:CPF

[Gin15] Yves Gingras. The creative power of formal analogies in physics: The case of Albert Einstein. Science & Education (Springer), 24(5–6):529–541, July 2015. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Ginoux:2022:AED

[Gin22] Jean-Marc Ginoux. Albert Einstein and the doubling of the deflection of light. Foundations of Science, 27(3):829–850, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.
1007/s10699-021-09783-4.

Giovanelli:2013:LAV

[Gio13a] Marco Giovanelli. Leibniz-Äquivalenz vs. Einstein-Äquivalenz. Was man von der Logisch-Empiristischen(Fehl-)Interpretation des Punkt-Koinzidenz-Arguments lernen kann. (German) [Leibniz equivalence vs. Einstein equivalence. What can be learned from the logical-empirical (mis) interpretation of the pointcoincidence argument]. Philosophia Naturalis, 50(1):115–164, ???? 2013. CODEN ???? ISSN 0031-8027.

Giovanelli:2013:TCP

[Gio13b] Marco Giovanelli. Talking at cross-purposes: how Einstein and the logical empiricists never agreed on what they were disagreeing about. Synthese, 190(17):3819–3863, November 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/s11229-
012-0229-1;   http://link.springer.com/content/pdf/
10.1007/s11229-012-0229-1.pdf.

Giovanelli:2014:OML

[Gio14] Marco Giovanelli. ‘But one must not legalize the mentioned sin’: Phenomenological vs. dynamical treatments of rods and clocks in Einstein’s thought. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):20–44, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000963.

Giovanelli:2016:SCG

[Gio16] Marco Giovanelli. ‘… But I still can’t get rid of a sense of artificiality’: the Reichenbach–Einstein debate on the geometrization of the electromagnetic field. Studies in History and Philosophy of Modern Physics, 54(??):35–51, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300642.

Giovanelli:2020:LTB

[Gio20] Marco Giovanelli. ‘Like thermodynamics before Boltzmann.’ on the emergence of Einstein’s distinction between constructive and principle theories. Studies in History and Philosophy of Modern Physics, 71(??):118–157, August 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219820300174.

Giovanelli:2022:GVU

[Gio22] Marco Giovanelli. Geometrization vs. unification: the Reichenbach–Einstein quarrel about the Fernparallelismus field theory. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03531-2.

Giroud:1986:MCL

[Gir86] Françoise Giroud. Marie Curie, a life. Holmes and Meier, New York, NY, USA, 1986. ISBN 0-8419-0977-6, 0-8419-0978-4 (paperback). 291 + 16 pp. LCCN QD22.C8 G4813 1986.

Gisin:2012:LHN

[Gis12]  N. (Nicolas) Gisin.  L’impensable hasard: non-localité,

téléportation et autres merveilles quantiques. Sciences. Odile Jacob, Paris, France, 2012. ISBN 2-7381-2831-9. 170 pp. LCCN QC174.17.E58 G57 2012.

Gisin:2014:QCH

[Gis14] Nicolas Gisin. Quantum Chance: Nonlocality, Teleportation and Other Quantum Marvels. Springer International Publishing, Cham, Switzerland, 2014. ISBN 3-319-05472-4, 3-319-05473-2 (e-book). xxiii + 109 pp. LCCN QC173.96-174.52. Foreword by Alain Aspect. Translation from French by Stephen Lyle.

Gittus:1963:U

[Git63] J. H. Gittus. Uranium. Butterworths, London, UK, 1963. xiii + 623 pp. LCCN TN799.U7 G5.

Giulini:2005:SRF

[Giu05] D. (Domenico) Giulini. Special Relativity: a First Encounter, 100 Years Since Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856746-4 (hardcover), 0-19-856747-2. viii + 168 pp. LCCN QC173.65 .G5813

2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/toc/ecip052/
2004024330.html.

Guth:2005:ICE

[GK05] Alan H. Guth and David I. Kaiser. Inflationary cosmology: Exploring the universe from the smallest to the largest scales. Science, 307(5711):884–890, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Glassner:2012:RAE

[GKP12a] Edwin Glassner and Heidi König-Porstner. [Rezension von:] Albert Einstein, Geometrie und Erfahrung. (German) [[Review of:] Albert Einstein, geometry and experience]. In Stadler and Wendel [SW12], pages 459–467. ISBN 3-211-32769-X (print), 3-211-69443-9 (electronic). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-211-69443-5_20.

Glassner:2012:REI

[GKP12b] Edwin Glassner and Heidi König-Porstner. Rostocker Ehrendoktoren. III. Albert Einstein. (German) [Rostock honorary doctors. III. Albert Einstein]. In Stadler and Wendel [SW12], pages 207–215. ISBN 3-211-32769-X (print), 3-211-69443-9

(electronic). LCCN ???? URL http://link.springer.com/
chapter/10.1007/978-3-211-69443-5_12.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Glasstone:1957:ENW

[Gla57] Samuel Glasstone, editor. The effects of nuclear weapons. U.S. Atomic Energy Commission, Washington, DC, USA, 1957. xii + 579 pp. LCCN UF767 .U522 1957.

Glasstone:1962:ENW

[Gla62] Samuel Glasstone, editor. The effects of nuclear weapons. U.S. Atomic Energy Commission, Washington, DC, USA, second edition, 1962. xvi + 730 pp. LCCN UF767 .U5216 1962.

Glanz:2000:DAP

[Gla00] James Glanz. Dr. Abraham Pais, 82, physicist and science historian, dies: the writer of one of the most acclaimed of all biographies of Einstein. New York Times, page B7, July 31, 2000. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://search.proquest.com/docview/91585630?.

Glauber:2006:NLO

[Gla06] Roy J. Glauber. Nobel Lecture: One hundred years of light quanta. Reviews of Modern Physics, 78(4):1267– 1278, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1267;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1267_1.

Glanz:2007:MML

[Gla07] Christian Glanz. Making music, loving and keeping one’s mouth shut! Albert Einstein’s relations with music. Österreichische Musikzeitschrift, 62(6):81, ???? 2007. ISSN 0029-9316.

Gleick:1992:GLS

[Gle92] James Gleick. Genius: The Life and Science of Richard Feynman. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-40836-3. x + 531 pp. LCCN QC16.F49G54 1992. US$27.50.

Gleiser:2010:TEC

[Gle10] Marcelo Gleiser. A tear at the edge of creation: a radical new vision for life in an imperfect universe. Free Press, New York, NY, USA, 2010. ISBN 1-4391-0832-3. xvii + 285 pp. LCCN QB981

.G575 2010. US$25.00. URL http://www.loc.gov/catdir/
enhancements/fy1009/2009046247-s.html;   http://www.
loc.gov/catdir/enhancements/fy1011/2009046247-b.html;
 http://www.loc.gov/catdir/enhancements/fy1011/2009046247-
d.html.

Gopalakrishnan:2009:ECF

[GLG09] Sarang Gopalakrishnan, Benjamin L. Lev, and Paul M. Goldbart. Emergent crystallinity and frustration with Bose–Einstein condensates in multimode cavities. Nature Physics, 5(11):845– 850, October 4, 2009. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.nature.com/

nphys/journal/v5/n11/full/nphys1403.html.

Glick:1984:BRS

[Gli84] Thomas F. Glick. Book review: Societat Catalana de Ciències Físiques, Químiques i Matemátiques: Centenari de la naixença d’Albert Einstein (Catalan) [Centenary of the Birth of Albert Einstein]. Isis, 75(2):428–429, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231892.

Glick:1987:CRR

[Gli87] Thomas F. Glick, editor. The Comparative Reception of Relativity, volume 103 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2498-9. vii + 412 pp. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US).

Glick:1988:ESR

[Gli88] Thomas F. Glick. Einstein in Spain: Relativity and the recovery of science. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-05507-6. xiii + 391 pp. LCCN Q175.52.S7

G48 1988. US$49.95. URL http://www.loc.gov/catdir/
description/prin031/87018946.html.

Gover:2015:PEH

[GLW15] A. Rod Gover, Emanuele Latini, and Andrew (Andrew Kenneth) Waldron. Poincaré–Einstein holography for forms via conformal geometry in the bulk, volume 235(1106) of Memoirs of the American Mathematical Society. American Mathematical Society, Providence, RI, USA, May 2015. ISBN 1-4704-1092-3, 1-4704-2224-7 (e-book). ISSN 0065-9266 (print), 1947-6221 (electronic). v + 95 pp. LCCN QA641 .G68 2015.

Glymour:1990:BRB

[Gly90] Clark Glymour. Book review: Einstein and the History of General Relativity. Don Howard and John Stachel, Eds. Birkhäuser Boston, Cambridge, MA, 1989. xii, 445 pp. $69. Einstein Studies, vol. 1. From a conference, North Andover, MA, May 1986. Science, 248(4957):879, May 18, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/248/4957/879.full.pdf.

Gruber:1980:SS

[GM80] Bruno Gruber and Richard S. Millman, editors. Symmetries in science. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

Gundlach:2000:MNC

[GM00] Jens H. Gundlach and Stephen M. Merkowitz. Measurement of Newton’s constant using a torsion balance with angular acceleration feedback. Physical Review Letters, 85(14):2869–2872, October 2000. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/doi/

10.1103/PhysRevLett.85.2869.

Galluzzi:2004:BMN

[GM04] Paolo Galluzzi and Laura Manetti. Beautiful minds: i Nobel italiani. Giunti, Firenze, Italia, 2004. ISBN 88-09-03814-2. 127 pp. LCCN AS911.N9 B43 2004.

Goodman:2006:TDU

[GM06] Barak Goodman and John Maggio. Ten days that unexpectedly changed America: Einstein’s letter. A&E Home Video, P.O. Box 2284, South Burlington, VT 05407, 2006. URL http://

www.store.aetv.com/. 60-minute video.

Gompert:1977:NWW

[GMG^{+}77] David C. Gompert, Michael Mandelbaum, Richard L. Garwin, John H. Barton, and Sidney D. Drell. Nuclear weapons and world politics: alternatives for the future. 1980s project/Council on Foreign Relations. McGraw-Hill, New York, NY, USA, 1977. ISBN 0-07-023713-1, 0-07-023714-X (paperback). xii + 370 pp. LCCN JX1974.7 .N85.

Greiner:2002:CRM

[GMHB02] Markus Greiner, Olaf Mandel, Theodor W. Hänsch, and Immanuel Bloch. Collapse and revival of the matter wave field of a Bose–Einstein condensate. Nature, 419(6902):51–54, September 5, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v419/n6902/full/nature00968.html.

Grandin:2012:WWO

[GMOP12] Karl Grandin, Piero Mazzinghi, Nils Olander, and Giuseppe Pelosi, editors. A wireless world: one hundred years since the Nobel Prize to Guglielmo Marconi, volume 42 of Bidrag till Kunglige Vetenskapsakademiens historia. Center for History of Science, The Royal Swedish Academy of Sciences, Stockholm, Sweden, 2012. ISBN 91-7190-178-7 (paperback). ISSN 0081-9956. xx + 370 pp. LCCN Q64 .S78 no.42.

Goldsmith:1980:EFH

[GMW80] Maurice Goldsmith, Alan L. (Alan Lindsay) Mackay, and James Woudhuysen, editors. Einstein, the first hundred years. Pergamon, New York, NY, USA, 1980. ISBN 0-08-025019-X. xiii + 200 pp. LCCN QC16.E5 E52 1980.

Grene:1986:SS

[GN86] Marjorie Grene and Debra Nails. Spinoza and the Sciences, volume 91 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8511-0, 94-009-4514-0 (ebook). ISSN 0068-0346. xix + 339 pp. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-009-4514-2.

Gron:2011:ETR

[GN11] Øyvind Grøn and Arne Næss. Einstein’s Theory: a Rigorous Introduction for the Mathematically Untrained. Springer Science+Business Media, LLC, New York, NY, USA, 2011. ISBN 1-4614-0705-2, 1-4614-0706-0. xvii + 341 pp. LCCN QC173.55 .G76 2011.

Gibbs:1999:NPS

[GNS99] W. Wayt Gibbs, Sasha Nemecek, and Gary Stix. The 1998 Nobel Prizes in Science. Scientific American, 280(1):16–19, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Goodman:1981:SBD

[GO81] David Goodman and Robert Olby. Science and belief: from Darwin to Einstein. Block 5: Unit 10–11, The mystery of life. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11004-5. 53 pp. LCCN QH341 .G66 1981.

Gangui:2008:EUO

[GO08] Alejandro Gangui and Eduardo L. Ortiz. Einstein’s unpublished opening lecture for his course on Relativity theory in Argentina, 1925. Science in Context, 21(3):435–450, August 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Godel:1949:ENT

[Göd49a] Kurt Gödel. An example of a new type of cosmological solutions of Einstein’s field equations of gravitation. Reviews of Modern Physics, 21(3):447–450, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.447;
http://rmp.aps.org/abstract/RMP/v21/i3/p447_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Godel:1949:RAR

[Göd49b] Kurt Gödel. A remark about the relationship between Relativity Theory and idealistic philosophy. In Schilpp [Sch49c], chapter 21, pages 555–562. CODEN 21. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

GoeppertMayer:1949:CSN

[Goe49] Maria Goeppert Mayer. On closed shells in nuclei. II. Physical Review (2), 75(12):1969–1970, June 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.75.1969;   http:
//prola.aps.org/abstract/PR/v75/i12/p1969_1;   http:
//www.nobelprize.org/nobel_prizes/physics/laureates/

Goenner:1970:MPE

[Goe70] Hubert Goenner. Mach’s Principle and Einstein’s theory of gravitation. In Ernst Mach: Physicist and Philosopher [CS70], chapter 10, pages 200–215. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-017-1462-4_10.

Goenner:1993:RRT

[Goe93] Hubert Goenner. The reaction to Relativity theory I: The anti-Einstein campaign in Germany in 1920. Science in Context, 6 (1):107–133, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goenner:2003:AEF

[Goe03] Hubert Goenner. Albert Einstein and Friedrich Dessauer: Political views and political practice. Physics in Perspective (PIP), 5(1):21–66, March 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/jkg195fabv2pjdcq/fulltext.
pdf.

Goenner:2004:HUF

[Goe04] Hubert F. M. Goenner. On the history of unified field theories. Living Reviews in Relativity, 7(1):2:1–2:153, 2004. ISSN 1433-8351.

Goenner:2005:EB

[Goe05] Hubert Goenner. Einstein in Berlin 1914–1933. Beck, München, Germany, 2005. ISBN 3-406-52731-0 (clothbound). 367 pp. LCCN QC16.E5 G59 2005.

Goenner:2015:AE

[Goe15] Hubert Goenner. Albert Einstein. Verlag C. H. Beck oHG, München, Germany, 2015. ISBN 3-406-67592-1, 3-406-67593-X (e-book). 128 pp. LCCN QC16.E5. URL http:/

/alltitles.ebrary.com/Doc?id=11031926;   http://lib.
myilibrary.com?id=750950;   http://public.eblib.com/
choice/PublicFullRecord.aspx?p=1988059;   http://swb.
eblib.com/patron/FullRecord.aspx?p=1988059.

Goenner:2017:GRG

[Goe17] Hubert Goenner. General relativity and the growth of a subdiscipline “gravitation” in Germany. European Physical Journal H, 42(3):395–430, August 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

Goin:1991:NL

[Goi91] Peter Goin. Nuclear landscapes. Creating the North American landscape. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4077-5, 0-8018-4078-3 (paperback). xxii + 151 pp. LCCN U264.3 .G65 1991.

Goldberg:1967:HPE

[Gol67] Stanley Goldberg. Henri Poincaré and Einstein’s Theory of Relativity. American Journal of Physics, 35(10):934, October 1967. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/35/10/10.1119/1.1973643.

Goldberg:1968:ERE

[Gol68] Stanley Goldberg. The Early response to Einstein’s Special Theory of Relativity, 1905–1911: a Case Study in National Differences. Ed.D. thesis, Graduate School of Education, Harvard University, Cambridge, MA, USA, 1968. 424 pp. URL http:/

/books.google.com/books?id=NDc9AQAAIAAJ.

Goldberg:1969:LTE

[Gol69] Stanley Goldberg. The Lorentz theory of electrons and Einstein’s theory of relativity. American Journal of Physics, 37 (10):982–994, October 1969. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Goldberg:1970:DEB

[Gol70a] Stanley Goldberg. In defense of ether: The British response to Einstein’s Special Theory of Relativity, 1905–1911. Historical Studies in the Physical Sciences, 2(??):89–125, ???? 1970. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.

org/stable/27757305.

Goldberg:1970:PSE

[Gol70b] Stanley Goldberg. Poincaré’s silence and Einstein’s Relativity: The role of theory and experiment in Poincaré’s physics. British Journal for the History of Science, 5(1):73–84, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goldhaber:1972:RPD

[Gol72] Maurice Goldhaber. Remarks on the prehistory of the discovery of slow neutrons. Proceedings - Royal Society of Edinburgh, Section A: Mathematical and Physical Sciences, 70A(??):191– 195, January 1972. CODEN PREAAY. ISSN 0080-4541 (print), 2053-5902 (electronic).

Goldberg:1976:MPP

[Gol76] Stanley Goldberg. Max Planck’s philosophy of nature and his elaboration of the Special Theory of Relativity. Historical Studies in the Physical Sciences, 7(??):125–160, ???? 1976. CODEN HSPSAS. ISBN 0-691-08169-7. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757355.

Goldschmidt:1982:ACW

[Gol82] Bertrand Goldschmidt. The Atomic Complex: a Worldwide Political History of Nuclear Energy. American Nuclear Society, La Grange Park, IL, USA, 1982. ISBN 0-89448-550-4 (hardcover), 0-89448-551-2 (paperback). xiv + 479 pp. LCCN QC773 .G6313 1982.

Goldberg:1983:AEC

[Gol83a] Stanley Goldberg. Albert Einstein and the creative act: the case of Special Relativity. In Aris et al. [ADS83], pages 232– 253. ISBN 0-8166-1087-8. LCCN QC7 .S77 1983.

Goldman:1983:DAS

[Gol83b] Martin Goldman. The demon in the aether: the story of James Clerk Maxwell. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-026-5. 224 + 12 pp. LCCN QC16.M4 G65 1983. US$15.00.

Goldberg:1984:BRB

[Gol84a] Stanley Goldberg. Book review: “Subtle is the Lord… ”: The Science and the Life of Albert Einstein, by Abraham Pais. American Journal of Physics, 52(10):951–953, October 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://link.aip.org/link/?AJP/52/951/1.

Goldberg:1984:BRH

[Gol84b] Stanley Goldberg. Book reviews: Harry Woolf: Some Strangeness in the Proportion: A Centennial Symposium to Celebrate the Achievements of Albert Einstein; Gerald Holton and Yehuda Elkana: Albert Einstein: Historical and Cultural Perspectives. Isis, 75(2):429–430, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231893.

Goldberg:1984:URO

[Gol84c] Stanley Goldberg. Understanding relativity: origin and impact of a scientific revolution. Birkhäuser Boston Inc., Cambridge, MA, USA, 1984. ISBN 0-8176-3150-X. xvii + 494 pp. LCCN QC173.65 .G65 1984.

Goldberg:1987:UAO

[Gol87] S. Goldberg. As usual, anything but ordinary: The Collected Papers of Albert Einstein, Vol 1, the Early Years, 1879– 1902 — J. Stachel. The Scientist (Philadelphia, PA), 1(21):21, September 21, 1987. ISSN 0890-3670 (print), 1945-5127 (elec-

tronic). URL http://www.the-scientist.com/?articles.
view/articleNo/8957/.

Goldberg:1995:GSC

[Gol95a] Stanley Goldberg. Groves and the scientists: Compartmentalization and the building of the bomb. Physics Today, 48(8):38– 43, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v48/i8/p38_s1.

Goldsmith:1995:EGB

[Gol95b] Donald Goldsmith. Einstein’s greatest blunder?: the cosmological constant and other fudge factors in the physics of the Universe. Harvard University Press, Cambridge, MA, USA, 1995. ISBN 0-674-24241-6. vi + 216 + 28 pp. LCCN QB981 .G594 1995.

Goldberg:1996:GGB

[Gol96] Stanley Goldberg. General Groves and the bombing of Hiroshima and Nagasaki. (Part II). Berichte zur Wissenschaftsgeschichte, 19(4):207–217, ???? 1996. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/bewi.

19960190402/abstract. See [GHK^{+}96] for Part I.

Goldsmith:1997:EBY

[Gol97] Donald Goldsmith. Einstein”n büyük yanılgısı. (Turkish) [Einstein’s greatest blunder]. Harvard University Press, Cambridge, MA, USA, 1997. ISBN 0-674-24241-6. 176 + 12 pp. LCCN ???? Translation of [?] to Turkish by Fatma Esin.

Goldsmith:1998:NPE

[Gol98a] Donald Goldsmith. Najwiȩksza pomyłka Einsteina?: stała kosmologiczna i inne niewiadome w fizyce Wszechsświata. (Polish) [Einstein’s greatest blunder?: The cosmological constant and other uncertainties in the physics of the Universe]. Prośzyński i S-ka, Warszawa, Poland, 1998. ISBN 83-7180-069-X. 205 + 24 pp. LCCN ???? Translation of [Gol95b] to Polish by Bogumił Bieniok and Ewa L. Łokas.

Goldsmith:1998:UNS

[Gol98b] Donald Goldsmith. Uchū-no-shōtai Ainshutain-no-ōinaru misu? (Japanese) [Einstein’s greatest blunder?]. Tōkyō Aotsuchisha, Tokyo, Japan, 1998. ISBN 4-7917-5548-0. 295 + 5 + 24 pp. LCCN ????

Goldberg:1999:EGA

[Gol99] Stanley Goldberg. The Enola Gay affair: What evidence counts when we commemorate historical events? Osiris (Series 2), 14 (??):176–186, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301967.

Goldberg:2004:AEE

[Gol04] S. Goldberg. Albert Einstein, Esq. Georgetown Law Journal, 93(1):319–333, November 2004. ISSN 0016-8092.

Goldsmith:2005:OGI

[Gol05] Barbara Goldsmith. Obsessive genius: the inner world of Marie Curie. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2005. ISBN 0-393-05137-4 (hardcover), 0-393-32748-5 (paperback). 256 pp. LCCN QD22.C8 G56 2005. URL http:

//catdir.loc.gov/catdir/toc/ecip0419/2004015027.html.

Goldenstern:2015:AEG

[Gol15] Joyce Goldenstern. Albert Einstein: genius of the theory of relativity. Genius scientists and their genius ideas. Enslow Publishers, Inc., Berkeley Heights, NJ, USA, 2015. ISBN 0-7660-6515-4 (hardcover), 0-7660-6516-2 (paperback), 0-7660-6517-0 (e-book). 96 pp. LCCN QC16.E5 G65 2015.

Goncharov:1996:ASH

[Gon96a] German A. Goncharov. American and Soviet H-bomb development programmes: historical background. Physics-Uspekhi, 39 (10):1033–1044, October 1996. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL https://nuke.fas.

org/guide/russia/nuke/goncharov-h-bomb.pdf.

Goncharov:1996:TM

[Gon96b] German A. Goncharov. Thermonuclear milestones. Physics Today, 49(11):44, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMA

[Gon96c] German A. Goncharov. Thermonuclear milestones: (1) the American effort. Physics Today, 49(11):45–48, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMB

[Gon96d] German A. Goncharov. Thermonuclear milestones: (2) beginnings of the Soviet H-bomb program. Physics Today, 49(11): 50–54, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMR

[Gon96e] German A. Goncharov. Thermonuclear milestones: (3) the race accelerates. Physics Today, 49(11):56–61, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Gonzales:2000:MPA

[Gon00] Doreen Gonzales. The Manhattan Project and the atomic bomb in American history. In American history. Enslow Publishers, Berkeley Heights, NJ, USA, 2000. ISBN 0-89490-879-0. 128 pp. LCCN QC773.3.U5 G65 2000.

Gonzalez:2005:BRS

[Gon05] Cristina Gonzalez. Book review: Shakespeare, Einstein, and the Bottom Line: The Marketing of Higher Education. The Review of Higher Education, 28(3):433–434, 2005. CODEN ???? ISSN 0162-5748 (print), 1090-7009 (electronic). URL https://muse.

jhu.edu/article/179947.

Gonzales:2012:SMP

[Gon12] Doreen Gonzales. The secret of the Manhattan Project. Stories in American history. Enslow Publishers, Berkeley Heights, NJ, USA, 2012. ISBN 0-7660-3954-4 (hardcover), 1-4644-0024-5 (paperback). 128 pp. LCCN QC773.3.U5 G65 2012.

Goodstein:1975:LCA

[Goo75] J. R. Goodstein. Levi-Civita, Albert Einstein, and relativity in Italy. Atti dei Convegni Lincei, 8(??):43–51, ???? 1975.

Gooding:1985:BRC

[Goo85] David Gooding. Book review: Carl Gustaf Bernhard, Elisabeth Crawford and Per Sörbom (eds), Science, Technology and Society in the Time of Alfred Nobel. Oxford and New York: Pergamon Press (for the Nobel Foundation), 1982. Pp. xvii + 426. £29.75, $59.50. British Journal for the History of Science, 18(2):239–240, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4026338.

Goodman:1990:MMW

[Goo90a] Ellen Goodman. The myth of Mileva: When it came to marriage, Einstein was no genius. Chicago Tribune, ??(??):??, March 18, 1990. ISSN 1085-6706 (print), 2165-171X (electronic).

Goodman:1990:SGM

[Goo90b] Ellen Goodman. Out of the shadows of ‘great’ men. Boston Globe, ??(??):15, March 15, 1990. ISSN 0743-1791.

Goodheart:2000:DMA

[Goo00] A. Goodheart. Driving Mr. Albert — a trip across America with Einstein’s brain. New York Times Book Review, ??(??):6, July 30, 2000. ISSN 0028-7806.

Goodstein:2002:BRB

[Goo02] David Goodstein. Book review: Einstein, Picasso. Space, time, and the beauty that causes havoc. Physics in Perspective (PIP), 4(2):247–248, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodman:2003:GHB

[Goo03a] Michael S. Goodman. The grandfather of the hydrogen bomb?: Anglo–American intelligence and Klaus Fuchs. Historical Studies in the Physical and Biological Sciences, 34(1):1–22, September 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Goodstein:2003:RSN

[Goo03b] D. Goodstein. The road to Stockholm: Nobel prizes, science and scientists. Physics in Perspective (PIP), 5(4):473, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodchild:2004:ETRa

[Goo04a] Peter Goodchild. Edward Teller: the real Dr Strangelove. Weidenfeld and Nicolson, London, UK, 2004. ISBN 0-297-60734-0. 467 pp. LCCN QC16.T37 G66 2004b.

Goodchild:2004:ETRb

[Goo04b] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004.

URL http://alsos.wlu.edu/information.aspx?id=2237.

Gordin:2002:BRP

[Gor02] Michael D. Gordin. Book review: Playing dice with Einstein: M. Jammer, Einstein and Religion: Physics and Theology. (268 pp.) Princeton University Press, Princeton, 1999, hardcover, US $26.95, UK £18.95, ISBN 0-691-00699-7. Studies in History and Philosophy of Modern Physics, 33(1):95–100, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000399. Translation and revision of

Einstein und die Religion published by Universitätsverlag Konstanz. See [JAvWE95a, Jam99, Jam02].

Gordin:2007:BRB

[Gor07a] Michael D. Gordin. Book review: Einstein: His Life and Universe. By Walter Isaacson. (New York: Simon & Schuster, 2007. xxii, 675 pp. $32.00, ISBN 978-0-7432-6473-0). The Journal of American History, 94(3):964, 2007. ISSN 0021-8723 (print),

1945-2314 (electronic). URL http://jah.oxfordjournals.
org/content/94/3/964.1.short.

Gordin:2007:FDA

[Gor07b] Michael D. Gordin. Five days in August: how World War II became a nuclear war. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12818-9. xv + 209 pp. LCCN

D767.25.H6 G67 2007. URL http://www.loc.gov/catdir/
enhancements/fy0668/2006049337-d.html;   http://www.
loc.gov/catdir/enhancements/fy0708/2006049337-t.html;
 http://www.loc.gov/catdir/enhancements/fy0734/2006049337-
b.html.

Gordin:2009:RCD

[Gor09] Michael D. Gordin. Red cloud at dawn: Truman, Stalin, and the end of the atomic monopoly. Farrar, Strauss, and Giroux, New York, NY, USA, 2009. ISBN 0-374-25682-9. xii + 402 pp. LCCN U264 .G67 2009.

Gordon:2012:PDE

[Gor12] Jeremy Gordon. Playing dice with Einstein: Reflections on God and the nature of suffering. Conservative Judaism, 64(1):14– 21, 2012. CODEN ???? ISSN 0010-6542 (print), 1947-4717

(electronic). URL https://muse.jhu.edu/article/494236.

Gorroochurn:2017:GDP

[Gor17] Prakash Gorroochurn. God does not play dice: Revisiting Einstein’s rejection of probability in quantum mechanics. The Mathematical Scientist, 42(2):61–73, August 2017. ISSN 0312-3685 (print), 1475-6080 (electronic). URL http:

//www.appliedprobability.org/content.aspx?Group=tms&
Page=TMS422.

Gordin:2020:EB

[Gor20] Michael D. Gordin. Einstein in Bohemia. Princeton University Press, Princeton, NJ, USA, 2020. ISBN 0-691-17737-6 (hardcover), 0-691-19984-1 (e-book). xi + 343 pp. LCCN QC16.E5 G6794 2020.

Gordin:2022:ELM

[Gor22] Michael D. Gordin. Einsteinian language: Max Talmey, Benjamin Lee Whorf and linguistic relativity. British Journal for the History of Science, 55(2):145–165, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL https://www.cambridge.org/core/journals/
british-journal-for-the-history-of-science/article/
einsteinian-language-max-talmey-benjamin-lee-whorf-
and-linguistic-relativity/ECE744AA2067BA1EC96296C2E89DAC18.

Gosling:1994:MPH

[Gos94] Francis G. (Francis George) Gosling. The Manhattan Project: making the atomic bomb. Energy history series. History Division, Executive Secretariat, Human Resources and Administration, Department of Energy, Washington, DC, USA, September 1994. vii + 66 pp.

Gosling:1999:MPM

[Gos99] Francis G. (Francis George) Gosling. The Manhattan Project: making the atomic bomb. US Dept. of Energy, Washington, DC, USA, January 1999. vii + 66 pp.

Gosling:2001:MPM

[Gos01] Francis G. (Francis George) Gosling. The Manhattan Project: making the atomic bomb. US Department of Energy, Washington, DC, USA, 2001 edition, October 2001. vi + 66 pp.

Gottlieb:1980:AEP

[Got80] I. Gottlieb. Albert Einstein — a pathfinder in physics. Revue Roumaine de Physique, 25(3):215–219, ???? 1980. ISSN 0035-4090.

Gotschl:1992:ESW

[Göt92] Johann Götschl, editor. Erwin Schrödinger’s world view: the dynamics of knowledge and reality, volume 16 of Theory and decision library. Series A, Philosophy and methodology of the social sciences. Kluwer Academic, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1694-0. viii + 237 pp. LCCN QC16.S265 E78 1992.

Goto:2000:EQC

[Got00] Kunio Goto. Einstein’s quantum conditions in early quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):148– 155, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Gott:2001:TTE

[Got01] J. Richard Gott. Time Travel in Einstein’s Universe: the Physical Possibilities of Travel Through Time. Houghton-Mifflin, Boston, MA, USA, 2001. ISBN 0-395-95563-7. xii + 291 pp. LCCN QC173.59.S65 G67 2001. URL http:

//catalog.hathitrust.org/api/volumes/oclc/45603941.
html;   http://catdir.loc.gov/catdir/description/hm021/
00054243.html;   http://catdir.loc.gov/catdir/toc/hm021/
00054243.html.

Gott:2002:TTE

[Got02] J. Richard Gott. Time Travel in Einstein’s Universe: The Physical Possibilities of Travel Through Time. Mariner Books, Boston, MA, USA, 2002. ISBN 0-618-25735-7. xii + 291 pp. LCCN ????

Goudsmit:1947:HGN

[Gou47a] Samuel A. Goudsmit. Heisenberg on the German Nuclear Power Project. Bulletin of the Atomic Scientists, 3(11):343, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Hei47b].

Goudsmit:1947:A

[Gou47b] Samuel Abraham Goudsmit. Alsos. H. Schuman, New York, NY, USA, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947.

Goudsmit:1947:AFG

[Gou47c] Samuel Abraham Goudsmit. Alsos: the failure in German science. Sigma Books, London, UK, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947a.

Goudsmit:1948:LSA

[Gou48] Samuel Abraham Goudsmit. L’Allemagne et le secret atomique (la mission Alsos). (French) [Germany and the atomic secret (Alsos mission)]. A. Fayard, Paris, France, 1948. 250 + 1 pp. LCCN D810.S2 G615.

Goudsmit:1983:A

[Gou83] Samuel Abraham Goudsmit. Alsos. History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-09-9. xxx + 259 pp. LCCN D810.S2 G6 1983.

Goudsmit:1996:A

[Gou96] Samuel Abraham Goudsmit. Alsos. History of modern physics and astronomy. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-415-5. xxxviii + 259 pp. LCCN D810.S2 G6 1996.

Gourgoulhon:2010:RRH

[Gou10] Éric Gourgoulhon. Relativité restreinte: des particules à l’astrophysique. (French) [Constrained Relativity: from particles to astrophysics]. Savoirs actuels. Physique. EDP sciences, Les Ulis, France, 2010. ISBN 2-7598-0067-9 (EDP sciences), 2-271-07018-X (CNRS éd). xxvi + 776 pp. LCCN QC173.65 G68 2010.

URL http://catalogue.bnf.fr/ark:/12148/cb414117131.

Gourgoulhon:2013:SRG

[Gou13] Éric Gourgoulhon. Special relativity in general frames: from particles to astrophysics. Graduate texts in physics, 1868-4513. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-37275-9 , 3-642-37276-7 (e-book). xxx + 784 pp. LCCN QC173.65 .G6813 2013.

Gowing:1964:BAEa

[Gow64a] Margaret Gowing. Britain and atomic energy, 1939–1945. Macmillan, London, UK, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964a.

Gowing:1964:BAEb

[Gow64b] Margaret Gowing. Britain and Atomic Energy, 1939–1945. St. Martin’s Press, New York, NY, USA, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Gowing:1988:BB

[Gow88] Margaret Gowing. Britain and the bomb: The origin of Britain’s determination to be a nuclear power. Contemporary Record: the journal of the Institute of Contemporary British History, 2(2): 36–40, June 1988. ISSN 0950-9224.

Gowing:1993:JCA

[Gow93] Margaret Gowing, F.R.S., F.B.A. James Chadwick and the atomic bomb. Notes and Records of the Royal Society of London, 47(1):79–92, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531395.

Gurgel:2011:EDS

[GP11] Iva Gurgel and Mauricio Pietrocola. An epistemological discussion of scientific imagination: The construction of knowledge across Albert Einstein’s views. Revista Brasileira de Ensino de Fisica, 33(1):??, January/March 2011. ISSN 1806-1117.

Gates:2019:PER

[GP19] S. James Gates, Jr. and Cathie Pelletier. Proving Einstein Right: the Daring Expeditions That Changed How We Look at the Universe. PublicAffairs, New York, NY, USA, 2019. ISBN 1-5417-6223-1 (e-book), 1-5417-6225-8 (hardcover). viii + 356 + 16 pp. LCCN QB544.19 .G38 2019. URL https://

historynewsnetwork.org/article/173408.

Glick:2001:RDI

[GPSR01] Thomas F. Glick, Miguel Angel Puig-Samper, and Rosaura Ruiz. The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil, volume 221 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346.

282 (est.) pp. LCCN ???? URL https://link.springer.com/
book/10.1007/978-94-010-0602-6.

Gogate:1944:SFL

[GR44] D. V. Gogate and R. N. Rai. Surface flow of liquid helium II and Bose–Einstein degeneracy. Nature, 153(3881):342, March 18, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v153/n3881/pdf/153342a0.pdf.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Grodzins:1965:AAS

[GR65] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Simon and Schuster, New York, NY, USA, 1965. xvii + 616 pp. LCCN D842 .B78.

Giere:1996:OLE

[GR96] Ronald N. Giere and Alan W. Richardson, editors. Origins of logical empiricism, volume 16 of Minnesota Studies in the Philosophy of Science. University of Minnesota Press, Minneapolis, MN, USA and London, UK, 1996. ISBN 0-8166-2834-3 (hardcover). vii + 392 pp. LCCN Q1 .M665; B824.6 .O75 1996; Q175.A1 M64; Q175 .M48; Q175 .M64.

Glick:2006:SFE

[GR06] Thomas F. Glick and José M. Sánchez Ron. Science frustrated: The ‘Einstein Institute’ in Madrid. Minerva [Washington, DC], 44(4):355–378, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-9013-z.

Gavroglu:2007:PHS

[GR07] Kostas Gavroglu and Jürgen Renn. Positioning the History of Science, volume 248 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5195-6, 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346. vii + 188 pp. LCCN Q125 .P77 2007. URL https://link.

springer.com/book/10.1007/1-4020-5420-3.

Gutfreund:2015:RRH

[GR15] Hanoch Gutfreund and Jürgen Renn. The road to relativity: the history and meaning of Einstein’s “The foundation of general relativity” featuring the original manuscript of Einstein’s masterpiece. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16253-0 (hardcover). xviii + 237 pp. LCCN QC173.6 .G88 2015.

Gutfreund:2017:FYR

[GR17] Hanoch Gutfreund and Jürgen Renn. The formative years of relativity: the history and meaning of Einstein’s Princeton lectures: featuring Einstein’s classic text The meaning of relativity in its historical context. Princeton University Press, Princeton, NJ, USA, 2017. ISBN 0-691-17463-6 (hardcover), 1-4008-8868-9 (e-book). xiv + 415 pp. LCCN QC173.6 .G878 2017. URL

http://www.jstor.org/stable/j.ctt1vxm7ts.

Gutfreund:2020:EEb

[GR20a] Hanoch Gutfreund and Jürgen Renn. Einstein on Einstein.

Web site, May 18, 2020. URL https://press.princeton.edu/
ideas/hanoch-gutfreund-and-jurgen-renn-on-einstein-
on-einstein.

Gutfreund:2020:EEA

[GR20b] Hanoch Gutfreund and Jürgen Renn. Einstein on Einstein: autobiographical and scientific reflections. Princeton University Press, Princeton, NJ, USA, 2020. ISBN 0-691-18360-0 (hardcover), 0-691-20011-4 (e-book). xi + 197 pp. LCCN QC16.E5 G88 2020.

Gutfreund:2023:ERH

[GR23] Hanoch Gutfreund and Jürgen Renn. The Einsteinian Revolution: the Historical Roots of His Breakthroughs. Princeton University Press, Princeton, NJ, USA, 2023. ISBN 0-691-16876-8 (hardcover), 0-691-25649-7 (e-book). xviii + 249 pp. LCCN QC16.E5 G88 2023; QC16.E5 G885 2023.

Gray:1949:NP

[Gra49] George W. Gray. The Nobel prizes. Scientific American, 181(6):11–17, December 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v181/n6/
pdf/scientificamerican1249-11.pdf.

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39c, Bey49].

Graves:1971:CFC

[Gra71] John Cowperthwaite Graves. The Conceptual Foundations of Contemporary Relativity Theory. MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-07040-5. ix + 361 pp. LCCN QC6 .G687. Foreword by John Archibald Wheeler.

Graham:1982:REI

[Gra82] Loren R. Graham. The reception of Einstein’s ideas: two examples from contrasting political cultures. In Holton and Elkana [HE82], pages 107–138. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Grandy:1994:LSS

[Gra94a] David Grandy. Leo Szilard: Science as a mode of being. Ph.D. thesis, Indiana University, Bloomington, IN, USA, 1994. 312

pp. URL http://search.proquest.com/docview/304128624.

Grant:1994:PSO

[Gra94b] Edward Grant. Planets, stars, and orbs: the medieval cosmos, 1200–1687. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43344-4 (hardcover). xxiii + 816 pp. LCCN QB981 .G664 1994. URL http://

www.loc.gov/catdir/description/cam025/93025899.html;
 http://www.loc.gov/catdir/toc/cam025/93025899.html;
 http://www.zentralblattmath.org/zmath/en/search/?an=
0862.01006.

Grandy:1996:LSS

[Gra96] David Grandy. Leo Szilard: science as a mode of being. University Press of America, Lanham, MD, USA, 1996. ISBN 0-7618-0308-4 (hardcover). xiii + 189 pp. LCCN QC16.S95 G73 1996. Based on the author’s doctoral dissertation [Gra94a].

Graefrath:1999:EPE

[Gra99] R. Graefrath. The Einsteinturm in Potsdam (Erich Mendelssohn, renovation, Germany, Albert Einstein, scientific architecture). Denkmalpflege, 57(2):116–125, ???? 1999. ISSN 0947-031X.

Grandy:2001:OLE

[Gra01] David Grandy. The otherness of light: Einstein and Levinas. Postmodern Culture, 12(1):??, September 2001. CODEN ???? ISSN 1053-1920.

Grant:2005:H

[Gra05] R. G. Grant. Hiroshima. How did it happen? Lucent Books, Farmington Hills, MI, USA, 2005. ISBN 1-59018-602-8. 48 pp. LCCN D767.25.H6 G728 2005. URL http://www.loc.

gov/catdir/toc/ecip052/2004024458.html.

Grassmann:2007:RNE

[Gra07] Hermann Grassmann. On the relation of non–Euclidean geometry to extension theory. In Renn et al. [RSS^{+}07b], pages 1079–1080. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (ebook). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL https://link.springer.com/chapter/10.
1007/978-1-4020-4000-9_47. Originally published as Ap-

pendix 1 (1877) to “A New Branch of Mathematics: The ‘Ausdehnungslehre’ of 1844 and other works” (Chicago: Open Court, 1995), 279–280.

Gray:2018:SGG

[Gra18] Norman Gray. A Student’s Guide to General Relativity. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-107-18346-4 (hardcover), 1-316-63479-5 (paperback). xii + 151 pp. LCCN QC173.6 .G732 2018.

Greca:2018:BRE

[GRC18] Ileana M. Greca and Luis Rodríguez-Cano. Book review: An Einstein encyclopaedia: Alice Calaprice, Daniel Kennefick and Robert Schulmann (2015). An Einstein Encyclopaedia. Princeton University Press, Princeton. ISBN: 978-0-691-14174-9, 347 pages, price: $24.95 (paperback). Science & Education (Springer), 27(9–10):1021–1023, December 2018. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Greenberger:1996:EMB

[Gre96] Daniel M. Greenberger. Einstein: the man behind the myth. Physics World, 9(9):59–60, September 1996. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
9/9/phwv9i9a26.pdf.

Greene:1999:EUSa

[Gre99a] B. (Brian) Greene. The Elegant Universe: Superstrings, Hidden Dimensions, and the Quest for the Ultimate Theory. Jonathan Cape, London, UK, 1999. ISBN 0-224-05299-3. xiii + 448 pp. LCCN QC794.6.S85 G75 1999b.

Greene:1999:EUSb

[Gre99b] B. (Brian) Greene. The Elegant Universe: Superstrings, Hidden Dimensions, and the Quest for the Ultimate Theory. W. W. Norton & Co., New York, NY, USA, 1999. ISBN 0-393-04688-5. xiii + 448 pp. LCCN QC794.6.S85 G75 1999.

Greene:2000:EUS

[Gre00] B. (Brian) Greene. The Elegant Universe: Superstrings, Hidden Dimensions, and the Quest for the Ultimate Theory. Vintage Books, New York, NY, USA, 2000. ISBN 0-375-70811-1. xvi + 447 pp. LCCN QC794.6.S85 G75 2000.

Gornitz:2001:BSM

[GRE^{+}01] Thomas Görnitz, Helmut Rechenberg, Berthold-Georg Englert, Thomas Walcher, Katja Bammel, Thomas W. Beneke, Wolfgang W. Schwippert, and Filip Floecel. Bücher und Software: Meyenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg U. A./Kraus: Von der Uranspaltung zur Göttinger Erklärung. Otto Hahn, Werner Heisenberg und Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers/Dürr: Bohmsche Mechanik als Grundlage der Quantenmechanik/Thomas/Weise: The Structure of the Nucleon/Sube: Langenscheidts Fachwörterbuch Physik. Deutsch-Englisch/ConceptDraw 1.6 Zeichnen mit intelligenten Objekten /Rath: Quantenphysik für Windows Version 1.5. Physikalische Blätter, 57(11):82–86, November 2001. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
20010571123/abstract.

Green:2003:AE

[Gre03a] Jim Green, editor. Albert Einstein. Rebel lives. Ocean Press, Melbourne, Victoria, Australia, 2003. ISBN 1-876175-63-X (paperback). 88 + 1 pp. LCCN QC16.E5 A645 2003. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/cons041/2003105753.
html.

Grenthe:2003:C

[Gre03b] Ingmar Grenthe, editor. Chemistry, 1996–2000. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-02-4958-6, 981-02-4959-4 (paperback). x + 462 pp. LCCN QD39 .C4862 2003.

Green:2004:AES

[Gre04a] Jim Green. Albert Einstein: selected writings. LeftWord Books, New Delhi, India, 2004. ISBN 81-87496-43-6 (paperback). 88 + 1 pp. LCCN ????

Greene:2004:FCS

[Gre04b] Brian Greene. The Fabric of the Cosmos: Space, Time, and the Texture of Reality. Vintage Books, New York, NY, USA, 2004. ISBN 0-375-41288-3, 0-375-72720-5 (paperback). xii + 569 pp. LCCN QB982 .G74 2004.

Greiner:2004:RPM

[Gre04c] Walter Greiner. Relativity Principle and Michelson–Morley Experiment, chapter 29, pages 362–369. Classical Theoretical Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-95586-0 (print), 0-387-21851-3 (e-book). LCCN QC125.2 .G7413 2004eb. URL

http://www.myilibrary.com?id=18940.

Greene:2005:FCS

[Gre05a] Brian Greene. The Fabric of the Cosmos: Space, Time, and the Texture of Reality. Vintage Books, New York, NY, USA, 2005. ISBN 0-375-72720-5. xii + 569 pp. LCCN QB982 .G74 2004.

Greenspan:2005:ECW

[Gre05b] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Gregory:2005:FHU

[Gre05c] Jane Gregory. Fred Hoyle’s universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-850791-7 (hardcover), 0-19-151358-X (e-book). x + 406 + 16 pp. LCCN

QB36.H75 G74 2005. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005279804-d.html;   http://www.
loc.gov/catdir/enhancements/fy0636/2005279804-t.html;

 http://www.loc.gov/catdir/enhancements/fy0724/2005279804-
b.html.

Greene:2011:HRP

[Gre11] B. (Brian) Greene. The Hidden Reality: Parallel Universes and the Deep Laws of the Cosmos. Alfred A. Knopf, New York, NY, USA, 2011. ISBN 0-307-26563-3. xi + 370 pp. LCCN QC6 .G6885 2011.

Greene:2012:MM

[Gre12] Brian Greene. The mystery of the multiverse. Newsweek, pages 20–25, May 28, 2012. ISSN 0028-9604.

Greene:2015:JER

[Gre15a] Brian Greene. 100 Jahre Einsteins Raumzeit: Der Glanz des Genies. (German) [100 years of Einstein’s space–time: The splendor of genius]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, October 2015. CODEN SPEKDI. ISSN 0170-2971. URL http://www.spektrum.de/

magazin/einsteins-glanz-und-wirkung/1362266.

Greene:2015:EWM

[Gre15b] Brian Greene. Essay: Why He matters. Scientific American, 313(3):34–37, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v313/n3/full/
scientificamerican0915-34.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
34.pdf.

Greenberg:2019:BRPa

[Gre19a] Arthur Greenberg. Book review: The Posthumous Nobel Prize in Chemistry. Volume 1. Bulletin for the History of Chemistry, 44(1):62–68, ???? 2019. CODEN BHCHET. ISSN 1053-

4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v44-1/v44-1%20p62-68.pdf.

Greenberg:2019:BRPb

[Gre19b] Arthur Greenberg. Book review: The Posthumous Nobel Prize in Chemistry. Volume 2. Bulletin for the History of Chemistry, 44(2):139–147, ???? 2019. CODEN BHCHET. ISSN 1053-

4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v44-2/v44-2%20p139-147.pdf.

Gregis:2019:AAM

[Gré19c] Fabien Grégis. Assessing accuracy in measurement: the dilemma of safety versus precision in the adjustment of the fundamental physical constants. Studies in History and Philosophy of Science Part A, 74(??):42–55, April 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0039368117302972.

Gribanov:1979:PVAa

[Gri79a] D. P. Gribanov. Philosophical views of Albert Einstein. Voprosy Filosofii, ??(2):15–27, ???? 1979. ISSN 0042-8744.

Gribanov:1979:PVAb

[Gri79b] D. P. Gribanov. Philosophical views of Albert Einstein. Soviet Studies in Philosophy, 18(2):72–94, ???? 1979. ISSN 0038-5883.

Gribanov:1981:AEA

[Gri81] D. P. Gribanov. Albert Einstein attitude toward idealistic philosophy. Voprosy Filosofii, ??(7):64–75, ???? 1981. ISSN 0042-8744.

Gribbin:1984:SSCb

[Gri84a] John R. Gribbin. In search of Schrödinger’s cat: quantum physics and reality. Bantam Books, New York, NY, USA, 1984. ISBN 0-553-34103-0 (paperback). xvi + 302 pp. LCCN QC173.98 .G75 1984.

Gribbin:1984:SSCa

[Gri84b] John R. Gribbin. In search of Schrödinger’s cat: the startling world of quantum physics explained. Wildwood House, London, UK, 1984. ISBN 0-7045-3071-6. xvi + 302 pp. LCCN QC173.98 .G75x 1984b. US$12.50.

Gribbin:1993:PAA

[Gri93] J. Gribbin. Pay attention, Albert Einstein. New Scientist, 137 (1854):28–31, January 2, 1993. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Gribbin:1995:SKSa

[Gri95a] John R. Gribbin. Schrödinger’s kittens and the search for reality: solving the quantum mysteries. Little, Brown and Co., Boston, MA, USA, 1995. ISBN 0-316-32838-3. ix + 261 pp. LCCN QC174.12 .G747 1995.

Gribbin:1995:SKS

[Gri95b] John R. Gribbin. Schrödinger’s kittens and the search for reality: solving the quantum mysteries. Little, Brown and Co., Boston, MA, USA, 1995. ISBN 0-316-32838-3. ix + 261 pp. LCCN QC174.12 .G747 1995. US$23.95.

Griffiths:1995:IQM

[Gri95c] David J. (David Jeffery) Griffiths. Introduction to quantum mechanics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-124405-1. ix + 394 pp. LCCN QC174.12 .G75

1995. US$40.00. URL http://www.zentralblattmath.org/
zmath/en/search/?an=0818.00001.

Griffiths:2005:IQM

[Gri05] David J. (David Jeffery) Griffiths. Introduction to quantum mechanics. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2005. ISBN 0-13-111892-7. ix + 468 pp. LCCN QC174.12 .G75 2005.

Gribbin:2010:SMP

[Gri10] John R. Gribbin. In search of the multiverse: parallel worlds, hidden dimensions, and the ultimate quest for the frontiers of reality. Wiley, New York, NY, USA, 2010. ISBN 0-470-61352-1 (hardcover), 0-470-92656-2 (e-book). xii + 227 pp. LCCN

QB981 .G7566 2010. URL http://catalogimages.wiley.
com/images/db/jimages/9780470613528.jpg.

Gribbin:2012:ESQ

[Gri12] John R. Gribbin. Erwin Schrödinger and the quantum revolution. Bantam Press, London, UK, 2012. ISBN 0-593-06776-2 (hardcover), 0-593-06865-3 (paperback), 1-4464-6571-3 (ebook). xii + 322 + 16 pp. LCCN QC16.S265 G75 2012.

Grimberg:2015:ERG

[Gri15] Markus Grimberg. Einstein — Ein Roman. (German) [Einstein — a novel]. neobooks Self-Publishing, München, Germany, 2015. ISBN 3-7380-3017-4. ???? pp. LCCN ????

Greiner:2003:EMB

[GRJ03] Markus Greiner, Cindy A. Regal, and Deborah S. Jin. Emergence of a molecular Bose–Einstein condensate from a Fermi gas. Nature, 426(6966):537–540, November 26, 2003. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v426/n6966/
full/nature02199.html.

Grodzins:1949:ABP

[Gro49] Morton Grodzins. Americans betrayed: politics and the Japanese evacuation. University of Chicago Press, Chicago, IL, USA and London, UK, 1949. xvii + 444 pp. LCCN D769.8.A6 G7.

Grodzins:1956:LDS

[Gro56] Morton Grodzins. The loyal and the disloyal: social boundaries of patriotism and treason. University of Chicago Press, Chicago, IL, USA and London, UK, 1956. x + 319 pp. LCCN JC328 .G7.

Groves:1962:NIC

[Gro62] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Harper & Row, New York, NY, USA, 1962. xiv + 464 pp. LCCN QC773.A1 G7.

Groueff:1967:MPUa

[Gro67a] Stéphane Groueff. Manhattan Project: the untold story of the making of the atomic bomb. Collins, London, UK, 1967. 416 pp. LCCN QC773.3.U5 G7 1967b.

Groueff:1967:MPUb

[Gro67b] Stéphane Groueff. Manhattan Project; the untold story of the making of the atomic bomb. Little, Brown and Co., Boston, MA, USA, 1967. xii + 372 pp. LCCN QC773.3.U5 G7.

Groueff:1968:MPU

[Gro68] Stéphane Groueff. Manhattan Project: the untold story of the making of the atomic bomb, volume Q3643. Bantam Books, New York, NY, USA, June 1968. xii + 429 pp. LCCN QC773.3.U5 G7 1968.

Gron:1982:MSC

[Grø82] Ø. Grøn. The metric in a static cylindrical elastic medium and in an empty rotating frame as solutions of Einstein’s field equations. Foundations of Physics, 12(5):509–520, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00729998.

Groves:1983:NIC

[Gro83] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1983. ISBN 0-306-80189-2, 0-7867-4822-2 (e-book). xiv + 464 + 8 pp. LCCN QC773.A1 G7.

Groueff:2000:MPU

[Gro00] Stéphane Groueff. Manhattan Project: the untold story of the making of the atomic bomb. Authors Guild backinprint.com edition. iUniverse.com, Inc., Lincoln, NE, USA, 2000. ISBN 0-595-09238-1 (paperback). xii + 372 pp. LCCN QC773.3.U6 G76 2000.

Gross:2005:NLD

[Gro05] David J. Gross. Nobel Lecture: The discovery of asymptotic freedom and the emergence of QCD. Reviews of Modern Physics, 77(3):837–849, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.837;   http://rmp.
aps.org/abstract/RMP/v77/i3/p837_1.

Grondin:2012:WBB

[Gro12] David Grondin, editor. War beyond the battlefield. Routledge, London, UK, 2012. ISBN 0-415-52368-0. v + 220 pp. LCCN U21.5 .W37 2012.

Goenner:1999:EWG

[GRRS99a] Hubert Goenner, Jürgen Renn, Jim Ritter, and Tilman Sauer, editors. The Expanding Worlds of General Relativity, volume 7 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4060-6 (hardcover), 3-7643-4060-6. xv + 512 pp. LCCN QC173.6 .E97 1999.

Goenner:1999:HGR

[GRRS99b] Hubert Goenner, Jürgen Renn, Jim Ritter, and Tilman Sauer, editors. History of General Relativity IV: The Expanding Worlds of General Relativity. Based upon the Fourth International Conference, Berlin, Germany, 31 July–3 August 1995, volume 7 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4060-6 (hardcover), 3-7643-4060-6. LCCN QC173.6 .E97 1999.

Gok:2016:IRF

[GRS16] Abdullah Gök, John Rigby, and Philip Shapira. The impact of research funding on scientific outputs: Evidence from six smaller European countries. Journal of the Association for Information Science and Technology, 67(3):715–730, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Giamarchi:2008:BEC

[GRT08] Thierry Giamarchi, Christian Rüegg, and Oleg Tchernyshyov. Bose–Einstein condensation in magnetic insulators. Nature Physics, 4(3):198–204, March 3, 2008. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.

nature.com/nphys/journal/v4/n3/full/nphys893.html.

Grunbaum:1957:PRA

[Gru57] Adolf Grunbaum. The philosophical retention of absolute space in Einstein’s General Theory of Relativity. The Philosophical Review, 66(4):525–534, October 1957. URL http://www.

jstor.org/stable/2182748.

Grunbaum:1962:RPH

[Grü62] Adolf Grünbaum. The relevance of philosophy to the history of the Special Theory of Relativity. The Journal of Philosophy, 59(21):561–574, October 1962. ISSN 0022-362X (print),

1939-8549 (electronic). URL http://www.jstor.org/stable/

2023278. American Philosophical Association Eastern Division: Symposium Papers to be presented at the Fifty-ninth Annual Meeting, New York City, December 27–29, 1962.

Grunbaum:1973:CEP

[Grü73a] Adolf Grünbaum. Critique of Einstein’s philosophy of geometry. In Philosophical Problems of Space and Time [Grü73c], chapter 4, pages 106–151. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346. LCCN ???? URL https://link.

springer.com/chapter/10.1007/978-94-010-2622-2_4.

Grunbaum:1973:PAM

[Grü73b] Adolf Grünbaum. Philosophical appraisal of E. A. Milne’s alternative to Einstein’s STR. In Philosophical Problems of Space and Time [Grü73c], chapter 13, pages 410–417. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-010-2622-2_13.

Grunbaum:1973:PPS

[Grü73c] Adolf Grünbaum. Philosophical Problems of Space and Time, volume 12 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, second, enlarged edition, 1973. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346. xxiv

+ 884 pp. LCCN ???? URL https://link.springer.com/
book/10.1007/978-94-010-2622-2.

Gruning:1990:HAE

[Grü90] Michael Grüning. Ein Haus für Albert Einstein: Erinnerungen, Briefe, Dokumente. (German) [A house for Albert Einstein: Recollections, Letters, Documents]. Verlag der Nation, Berlin, Germany, 1990. ISBN 3-373-00324-5. 583 pp. LCCN QC16.E5 H38 1990.

Grundmann:1997:EAE

[Gru97] Siegfried Grundmann. Einsteins Akte: Einsteins Jahre in Deutschland aus der Sicht der deutschen Politik. (German) [Einstein’s Acts: Einstein’s Years in Germany from the Perspective of German Politics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63197-6. xvi + 535 pp. LCCN QC16.E5.

Grujic:2000:ED

[Gru00] Petar Grujic. Einstein 1, Dirac 1. Physics World, 13(5):19– 20, May 2000. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/13/5/phwv13i5a19.pdf.

Grujic:2005:EG

[Gru05a] Petar Grujic. Einstein and God. Physics World, 18(4):21, April 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/18/4/phwv18i4a30.pdf.

Grunden:2005:SWW

[Gru05b] Walter E. Grunden. Secret weapons and World War II: Japan in the shadow of big science. Modern war studies. University Press of Kansas, Lawrence, KS, USA, 2005. ISBN 0-7006-1383-8 (hardcover). xi + 335 pp. LCCN UF505.J3

G78 2005. URL http://www.loc.gov/catdir/toc/ecip053/
2004026485.html.

Grundmann:2005:EDS

[Gru05c] Siegfried Grundmann. The Einstein dossiers: science and politics — Einstein’s Berlin period with an appendix on Einstein’s FBI file. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-25661-X (hardcover), 3-540-31104-1. xix + 459 pp. LCCN

QC16.E5 G7513 2005. URL http://www.loc.gov/catdir/
enhancements/fy0663/2005930169-d.html;   http://www.
loc.gov/catdir/toc/fy0611/2005930169.html.

Grundmann:2005:IFP

[Gru05d] Siegfried Grundmann. Im Fadenkreuz von politischer Polizei und Geheimdiensten: Albert Einstein. (German) [In the crosshairs of political police and intelligence agencies: Albert Einstein]. In Steiner [Ste05a], pages 151–169. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http://link.springer.

com/chapter/10.1007/3-540-30595-5_7/.

Gruschka:2007:TSB

[Gru07] Roland Gruschka. Tuvia Schalit’s Di spetsyele relativitets-teorye of 1927 and other introductions to the Theory of Relativity in Yiddish. Science in Context, 20(2):317–339, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Grunberg:2008:NLS

[Grü08] Peter A. Grünberg. Nobel Lecture: From spin waves to giant magnetoresistance and beyond. Reviews of Modern Physics, 80 (4):1531–1540, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1531;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1531_1.

Greenberger:1999:EEP

[GRZ99] Daniel M. Greenberger, Wolfgang L. Reiter, and Anton Zeilinger, editors. Epistemological and experimental perspectives on quantum physics, volume 7 of Vienna Circle Institute yearbook. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6063-X. ISSN 0929-6328. x + 377 pp. LCCN QC174.13 .E65 1999.

Gerstell:1950:HSA

[GS50] Richard Gerstell and Leo Szilard, former owner. How to survive an atomic bomb, volume 845 of A Bantam book. Bantam Books, New York, NY, USA, 1950. 149 pp. LCCN UF767 .G39.

Gordon:1959:ASR

[GS59] Isabel S. Gordon and Sophie Sorkin, editors. The armchair science reader. Simon and Schuster, New York, NY, USA, 1959. 832 pp. LCCN PN6110.S37 G6.

Garuccio:1980:SDA

[GS80] A. Garuccio and F. Selleri. Systematic derivation of all the inequalities of Einstein locality. Foundations of Physics, 10(3– 4):209–216, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715068.

Giulini:2005:EIP

[GS05] D. Giulini and N. Straumann. Einstein’s impact on the physics of the twentieth century. ArXiv Physics e-prints, July

2005. URL http://adsabs.harvard.edu/abs/2005physics..
.7107G;   http://arxiv.org/abs/physics/0507107.

Giulini:2006:EIP

[GS06] Domenico Giulini and Norbert Straumann. Einstein’s impact on the physics of the twentieth century. Studies in History and Philosophy of Modern Physics, 37(1):115–173, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000924.

Ghojogh:2022:GDR

[GS22] Benyamin Ghojogh and Smriti Sharma. Gravitational dimensionality reduction using Newtonian gravity and Einstein’s general relativity. CoRR, ??(??):??, ???? 2022. CODEN ???? ISSN

???? URL https://dblp.org/db/journals/corr/corr2211.
html#abs-2211-01369.

Gerton:2000:DOG

[GSPH00] Jordan M. Gerton, Dmitry Strekalov, Ionut Prodan, and Randall G. Hulet. Direct observation of growth and collapse of a Bose–Einstein condensate with attractive interactions. Nature, 408(6813):692–695, December 7, 2000. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v408/n6813/
full/408692a0.html.

Gavroglu:1995:PPS

[GSW95] Kōstas Gavroglu, John J. Stachel, and Marx W. Wartofsky, editors. Physics, Philosophy, and the Scientific Community: Essays in the Philosophy and History of the Natural Sciences and Mathematics in Honor of Robert S. Cohen, volume 163 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2991-0 (set), 0-7923-2988-0. xxvii + 383 pp. LCCN Q174 .B67 vol. 163. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0823/94022250-
d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0823/94022250-t.html.

Gamow:1938:LER

[GT38] George Gamow and Edward Teller. Letter to editor: The rate of selective thermonuclear reactions. Physical Review, 53(7): 608–609, April 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v53/i7/p608_1.

Giacconi:1980:EON

[GT80] Riccardo Giacconi and Harvey Tananbaum. The Einstein Observatory: New perspectives in astronomy. Science, 209(4459): 865–876, August 22, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/209/4459/865.full.pdf.

Gilmore:2022:ERV

[GTP22] Gerard Gilmore and Gudrun Tausch-Pebody. The 1919 eclipse results that verified general relativity and their later detractors: a story re-told. Notes and Records of the Royal Society of London, 76(1):155–180, March 20, 2022. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0040.

Gubskij:1992:LNP

[GUB92] E. F. (Evgeniĭ Fedorovich) Gubskij, O. Uitmen, and K. G. Bernhard. Laureaty Nobelevskoj premii: Jenciklopediĭ. Kniga 1, A–L. Progress, Moskva, 1992. ISBN 5-01-002539-6. xxxiv + 740 + 1 pp. LCCN ????

Guillen:1995:FEC

[Gui95] Michael Guillen. Five equations that changed the world: the power and poetry of mathematics. Hyperion, New York, NY, USA, 1995. ISBN 0-7868-6103-7. viii + 277 pp. LCCN QC24.5 .G85 1995.

Guichelaar:2009:SAV

[Gui09] Jan Guichelaar. De Sitter: een alternatief voor Einsteins heelalmodel. (Dutch) [De Sitter: an alternative to Einstein’s model of the universe], volume 31 of Wetenschappelijke biografie. Natuurwetenschap and Techniek, Amsterdam, The Netherlands, 2009. ISBN 90-8571-181-9. 159 pp. LCCN 124.6-127.2.

Guichelaar:2018:WSE

[Gui18] Jan Guichelaar. Willem de Sitter: Einstein’s Friend and Opponent. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-98336-9 (hardcover), 3-319-98337-7 (e-book), 3-319-98338-5. ISSN 2365-0613 (print), 2365-0621 (electronic). xxii + 278 + 71 pp. LCCN QB36.S6

G85 2018. URL http://link.springer.com/10.1007/978-3-
319-98337-0.

Guntercomp:1969:BEP

[Gun69] P. A. Y. (Pete Addison Y.) Gunter, editor. Bergson and the evolution of physics. University of Tennessee Press, Knoxville, TN, USA, 1969. ISBN 0-87049-092-3. xi + 348 pp. LCCN QC16.B45 G8; B2430.B43 F69.

Gunther:2011:GDS

[Gun11] Ralph Gunther. A guided dream: selected data on the Nobel Prize, volume 166 of Scripta Humanistica. Scripta Humanistica, Potomac, MD, USA, 2011. ISBN 1-882528-57-3. 229 pp. LCCN AS911.N9 G85 2011.

Gunderman:2015:WSG

[Gun15] Richard Gunderman. When science gets ugly: The story of Philipp Lenard and Albert Einstein. Scientific Computing, ??(??):??, August 24, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic). URL http:

//www.scientificcomputing.com/articles/2015/08/when-
science-gets-ugly-story-philipp-lenard-and-albert-einstein.

Gupta:1957:EOT

[Gup57] Suraj N. Gupta. Einstein’s and other theories of gravitation. Reviews of Modern Physics, 29(3):334–336, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.29.334;   http://rmp.aps.org/
abstract/RMP/v29/i3/p334_1.

Guth:1975:ESC

[Gut75] Zdeněk Guth. Einsteinovi a Smíchov. (Czech) [Einstein in Smichov]. Neděle s Lidová demokracie, 31(15):9–10, April 12, 1975.

Guterl:1994:AEK

[Gut94] F. Guterl. Albert Einstein — keyhole view of a genius. Scientific American, 270(1):26, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Guth:1997:IUQ

[Gut97] Alan H. Guth. The inflationary universe: the quest for a new theory of cosmic origins. Perseus Publishers, Cambridge, MA, USA, 1997. ISBN 0-201-32840-2 (paperback), 0-201-14942-7 (hardcover). xv + 358 pp.

Gutzwiller:1998:MES

[Gut98] Martin C. Gutzwiller. Moon–Earth–Sun: The oldest threebody problem. Reviews of Modern Physics, 70(2):589–639, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.70.589;   http://rmp.aps.org/
abstract/RMP/v70/i2/p589_1.

Gutfreund:2015:ZGG

[Gut15] Hanoch Gutfreund. Zwei der Glänzendsten Gestirne: Max Planck und Albert Einstein. (German) [Two of the most shining stars: Max Planck and Albert Einstein]. In Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy] [LMOS15], pages 310–343. ISBN 3-11-037598-2 (hardcover). LCCN AS182.B38 B477 2015.

Greenberg:1983:AHT

[GW83] Martin Harry Greenberg and Charles G. Waugh, editors. The Arbor House treasury of Nobel Prize winners. Arbor House, New York, NY, USA, 1983. ISBN 0-87795-613-8 (paperback), 0-87795-511-5 (hard). 319 pp. LCCN PS648.S5 A7 1983. US$7.95.

Glass:1993:EB

[GW93] Philip Glass and Robert Wilson. Einstein on the beach. Elektra Nonesuch (New York) sound recording (201 minutes)., 1993.

Goldstine:1996:SWJ

[GW96] H. H. Goldstine and E. P. Wigner. The scientific work of John von Neumann. In Wigner [Wig96m], pages 123–126. ISBN 3-540-56972-3. Reprinted from [?].

Gimbel:2006:DEH

[GW06] Steven Gimbel and Anke Walz, editors. Defending Einstein: Hans Reichenbach’s writings on space, time, and motion. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85958-1 (hardcover). vi + 216 pp. LCCN QC173.55 .R439 2006.

Gingras:2010:WIB

[GW10] Yves Gingras and Matthew L. Wallace. Why it has become more difficult to predict Nobel Prize winners: a bibliometric analysis of nominees and winners of the chemistry and physics prizes (1901–2007). Scientometrics, 82(2):401–412, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-009-0035-9.

Gwynne:2002:FVA

[Gwy02] Peter Gwynne. The FBI vs Albert Einstein. Physics World, 15 (6):13, June 2002. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/15/6/phwv15i6a22.pdf.

Gwynne:2010:GMS

[Gwy10] Peter Gwynne. General Motors sued for ‘denigrating’ Einstein’s image. Physics World, 23(7):7, July 2010. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
23/07/phwv23i07a12.pdf.

Gyllensten:1978:NPL

[Gyl78] Lars Gyllensten. The Nobel Prize in literature. Swedish Academy, Stockholm, Sweden, 1978. 15 + 1 pp. LCCN ????

Greenstein:1997:QCM

[GZ97] George Greenstein and Arthur Zajonc. The quantum challenge: modern research on the foundations of quantum mechanics. The challenge series: The Jones and Bartlett series in physics and astronomy. Jones and Bartlett, Boston, MA, USA, 1997. ISBN 0-7637-0467-9 (hardcover), 0-7637-0216-1 (paperback), 0-585-25206-8 (e-book). xvi + 224 pp. LCCN QC174.12 .G73 1997.

Greenstein:2006:QCM

[GZ06] George Greenstein and Arthur Zajonc. The quantum challenge: modern research on the foundations of quantum mechanics. Jones and Bartlett, Boston, MA, USA, second edition, 2006. ISBN 0-7637-2470-X. xviii + 300 pp. LCCN QC174.12 .G73 2006.

Heisenberg:1961:MPb

[H^{+}61] Werner Heisenberg et al., editors. On modern physics. C. N. Potter, New York, NY, USA, 1961. 108 pp. LCCN QC71 .O5. English translation by M. Goodman and J. W. Binns.

H:1989:NP

[H.89] J. H. Nobel Prizes. Scientific American, 261(6):34–35, December 1989. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v261/n6/pdf/scientificamerican1289-
34b.pdf.

Howard:2017:GE

[H^{+}17] Ron Howard et al. Genius: Einstein. 10 episodes of television series, 2017. Developed for TV by Noah Pink and Ken Biller. National Geographic Channel.

Hewlett:1962:HUS

[HA62] Richard G. Hewlett and Oscar E. Anderson. A history of the United States Atomic Energy Commission. Volume 1, The New World, 1939–1946. Pennsylvania State University Press, University Park, PA, USA, 1962. ???? pp. LCCN HD9698.U58.

Haberman:1999:P

[HA99] Frederick W. Haberman and Irwin Abrams, editors. Peace. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3414-7 (1901–1925), 981-02-3415-5 (1926–1950), 981-02-3416-3 (1951–1970). ???? pp. LCCN JX1937 .P38

1999. URL http://nobelprize.org/nobel_prizes/peace/
video_lectures.html.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Haber:1982:CIP

[Hab82] L. F. Haber. Chemical innovation in peace and in war. In Bernhard et al. [BCS82], pages 271–282. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500220.

Habashi:2002:AAB

[Hab02] Fathi Habashi. From alchemy to atomic bombs: history of chemistry, metallurgy, and civilization. Métallurgie Extractive Québec, Sainte Foy, QC, Canada, 2002. ISBN 2-922686-00-0s. x + 357 pp. LCCN QD11 .H23 2002.

Habashi:2009:INM

[Hab09] Fathi Habashi. Ida Noddack and the missing elements. Education in Chemistry, 46(2):48–51, March 2009. CO-DEN EDCHAU. ISSN 0013-1350 (print), 1749-5326 (elec-

tronic). URL http://www.rsc.org/education/eic/issues/
2009March/ida-noddack-rhenium-nuclear-fission.asp.

Hacker:1987:DTR

[Hac87] Barton C. Hacker. The dragon’s tail: radiation safety in the Manhattan Project, 1942–1946. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1987. ISBN 0-520-05852-6. x + 258 pp. LCCN U264 .H33 1987.

Hacker:1994:ECA

[Hac94] Barton C. Hacker. Elements of controversy: the Atomic Energy Commission and radiation safety in nuclear weapons testing, 1947–1974. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1994. ISBN 0-520-08323-7. xxiv + 614 pp. LCCN U264.3 .H333 1994. URL http://

hdl.handle.net/2027/heb.00533;   http://www.loc.gov/
catdir/bios/ucal051/93041611.html;   http://www.loc.
gov/catdir/description/ucal041/93041611.html.

Hacker:1998:NPW

[Hac98] Carlotta Hacker. Nobel Prize winners. Women in profile. Crabtree Pub., New York, NY, USA, 1998. ISBN 0-7787-0007-0, 0-7787-0029-1 (paperback). 48 pp. LCCN AS911.N9 H29 1998.

Hadlington:1987:ES

[Had87] Simon Hadlington. Einstein for sale. Nature, 330(6144):97, November 18, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v330/n6144/pdf/330097c0.pdf.

Hewlett:1990:AS

[HAD90] Richard G. Hewlett, Oscar E. Anderson, and Francis Duncan. Atomic shield, 1947–1952, volume 2 of A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1990. ISBN 0-520-07187-5. xviii + 718 pp. LCCN HD9698.U52 H55 1989 vol. 2; QC792.7.

Hagerstrom:1946:EVS

[Häg46] Axel Hägerström. Erkenntnistheoretische Vorausseteungen der speziellen Relativitätstheorie Einsteins. (German) [Theoretical predictions of Einstein’s theory of special relativity]. Theoria, 12(1–2):1–68, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hagner:2005:EBP

[Hag05] Michael Hagner. Einstein on the Beach: der Physiker als Phänomen. (German) [Einstein on the Beach: the physicist as a phenomenon]. Fischer Taschenbuch Verlag, Frankfurt am Main, Germany, 2005. ISBN 3-596-16515-6 (paperback). 326 pp. LCCN QC16.E5 E516 2005.

Hagar:2008:LME

[Hag08] Amit Hagar. Length matters: The Einstein–Swann correspondence and the constructive approach to the Special Theory of Relativity. Studies in History and Philosophy of Modern Physics, 39(3):532–556, September 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000178.

Hagar:2014:DCQ

[Hag14] Amit Hagar. Discrete or continuous? The quest for fundamental length in modern physics. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06280-2, 1-139-95816-X (e-book), 1-139-96027-X (e-book), 1-107-47734-4 (e-book), 1-139-95921-2 (e-book). xi + 267 pp. LCCN QC173.59.S65 H34 2014.

Hahn:1942:EEE

[Hah42] Otto Hahn. Einiges über die experimentelle Entwirrung der bei der Spaltung des Urans auftretenden Elemente und Atomarten. (German) [Something on the experimental disentanglement of the fission of uranium occurring in the elements and types of atoms], volume 3 of Abhandlungen der Preussischen Akademie der Wissenschaften, Jahrg. 1942. Mathematischnaturwissenschaftliche Klasse. Akademie der Wissenschaften, in Kommission bei W. de Gruyter, Berlin, Germany, 1942. 30 pp. LCCN AS182 .B335 1942, Nr. 3.

Hahn:1948:KUI

[Hah48] Otto Hahn. Die Kettenreaktion des Urans und ihre Bedeutung. (German) [The chain reaction of uranium and its significance]. Deutscher Ingenieur-Verlag, Düsselfdorf, West Germany, 1948. 46 pp. LCCN QC780 .H3.

Hahn:1962:RUW

[Hah62] Otto Hahn. Vom Radiothor zur Uranspaltung: Eine wissenschaftliche Selbstbiographie. (German) [From radiothorium to uranium fission: a scientific autobiography]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1962. 156 + 47 pp. LCCN QD22.H2 A3.

Hahn:1966:OHS

[Hah66] Otto Hahn. Otto Hahn: a scientific autobiography. Charles Scribner’s Sons, New York, NY, USA, 1966. xxiv + 296 + 16 pp. LCCN QD22.H2 1966. Translated and edited by Willy Ley. With an introduction by Glenn T. Seaborg.

Hahn:1967:OHS

[Hah67] Otto Hahn. Otto Hahn: a scientific autobiography. MacGibbon and Kee, London, UK, 1967. xxiv + 296 + 16 pp. LCCN QD22.H2 1967. Translated and edited by Willy Ley. With an introduction by Glenn T. Seaborg.

Hahn:1968:MLG

[Hah68] Otto Hahn. Mein Leben. (German) [My Life]. Bruckmann, München, West Germany, 1968. 271 + 8 pp. LCCN QD22.H2 A28.

Hahn:1970:MLA

[Hah70] Otto Hahn. My life: the autobiography of a scientist. Herder and Herder, New York, NY, USA, 1970. 240 pp. LCCN QD22.H2 A2813.

Hahn:1973:BRBb

[Hah73] Roger Hahn. Book review: Les Prix Nobel en 1970 by The Nobel Foundation. Isis, 64(3):436–437, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229768.

Hahn:1975:EEG

[Hah75] Otto Hahn. Erlebnisse und Erkenntnisse. (German) [Experiences and insights]. Econ-Verlag, Düsseldorf, West Germany, 1975. ISBN 3-430-13732-2. 320 pp. LCCN QD22.H2 A26. DM36.00. With an introduction by Karl-Erik Zimen. Edited by Dietrich Hahn.

Hahn:1979:OHB

[Hah79] Dietrich Hahn. Otto Hahn: Begründer des Atomzeitalters: eine Biographie in Bildern und Dokumenten. (German) [Otto Hahn: Founder of the Atomic Age: a biography in pictures and documents]. List, München, West Germany, 1979. ISBN 3-471-77841-1. 357 pp. LCCN QD22.H2 O87. DM98.00. With a foreword by Reimar Lüst, a foreword by Paul Matussek, and an introduction by Walther Gerlach.

Hahn:1981:OHK

[Hah81] Dietrich Hahn. Otto Hahn in der Kritik: eine Auswahl deutschsprachiger Rezensionen und persönlicher Stellungnahmen zu den Büchern von und über Otto Hahn seit 1948. (German) [Otto Hahn in criticism: a selection of German reviews and personal opinions in the books by and about Otto Hahn since 1948]. H. Moos, München, West Germany, 1981. ISBN 3-7879-0198-1 (paperback). 116 pp. LCCN QD22.H2 O88 1981.

Hahn:1988:OHL

[Hah88] Dietrich Hahn. Otto Hahn: Leben und Werk in Texten und Bildern. (German) [Otto Hahn: life and work in texts and pictures], volume 1089 of Insel Taschenbuch. Insel, Frankfurt am Main, West Germany, 1988. ISBN 3-458-32789-4 (paperback). 361 pp. LCCN QD22.H2 O88 1988.

Haisch:2010:PGU

[Hai10] Bernard Haisch. The purpose-guided universe: believing in Einstein, Darwin, and God. New Page Books, Franklin Lakes, NJ, USA, 2010. ISBN 1-60163-122-7, 1-60163-733-0 (e-book). 222 pp. LCCN BL240.3 .H355 2010.

Hakim:1999:IRG

[Hak99]  Rémi Hakim.  An Introduction to Relativistic Gravita-

tion. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-45312-7, 0-521-45930-3 (paperback). ix + 271 pp. LCCN QC173.6 .H35 1999. URL http://www.loc.

gov/catdir/description/cam021/97017392.html;   http:
//www.loc.gov/catdir/toc/cam021/97017392.html.

Hakim:2007:EAN

[Hak07] Joy Hakim. Einstein adds a new dimension, volume 3 of The Story of science. Smithsonian Books, Washington, DC, 2007. ISBN 1-58834-162-3. xi + 468 pp. LCCN Q125 .H263 2007.

Haldane:1921:RR

[Hal21] Viscount R. B. (Richard Burdon) Haldane. The Reign of Relativity. John Murray, London, UK, 1921. xix + 434 pp. LCCN BD221 .H3 1921b.

Halla:1955:ATA

[Hal55] Franz Halla. Anläßlich des Todes von Albert Einstein. (German) [On the occasion of the death of Albert Einstein]. Mitteilungen aus der anthroposophischen Arbeit in Deutschland, 32(??):74– 75, June 1955.

Halasz:1959:NBA

[Hal59] Nicholas Halasz. Nobel: a biography of Alfred Nobel. Orion Press, New York, NY, USA, 1959. 281 pp. LCCN TP268.5.N7 H3.

Halperin:1965:CB

[Hal65] Morton H. Halperin. China and the bomb. Pall Mall Press, London, UK, 1965. LCCN ????

Halevi:1980:EM

[Hal80] P. Halevi. Einstein and Maxwell. Physics Today, 33(2):15, February 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Halliwell:1991:QCC

[Hal91] Jonathan J. Halliwell. Quantum cosmology and the creation of the universe. Scientific American, 265(6):76–?? (Intl. ed. 28–??), December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hales:1997:ASL

[Hal97] Peter B. (Peter Bacon) Hales. Atomic spaces: living on the Manhattan Project. University of Illinois Press, Urbana, IL, USA, 1997. ISBN 0-252-02296-3 (hardcover). 447 pp. LCCN QC773.3.U5 H35 1997.

Hallengren:2004:NLS

[Hal04a] Anders Hallengren, editor. Nobel laureates in search of identity and integrity: voices of different cultures. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-038-6, 981-256-074-2 (paperback). ???? pp. LCCN ????

Hallyn:2004:SRS

[Hal04b] Fernand Hallyn. Les structures rhétoriques de la science: de Kepler à Maxwell. (French). [The rhetorical structures of science: from Kepler to Maxwell]. Des travaux. Seuil, Paris, France, 2004. ISBN 2-02-063249-7. 321 pp. LCCN Q174.8 .H35 2004.

Hall:2006:NLD

[Hal06] John L. Hall. Nobel Lecture: Defining and measuring optical frequencies. Reviews of Modern Physics, 78(4):1279– 1295, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1279;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1279_1.

Halpern:2007:KEF

[Hal07] Paul Halpern. Klein, Einstein, and five-dimensional unification. Physics in Perspective (PIP), 9(4):390–405, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://www.springerlink.com/content/
j97p35877401w357/fulltext.pdf.

Halpern:2009:CSW

[Hal09] Paul Halpern. Collider: the search for the world’s smallest particles. John Wiley and Sons, Hoboken, NJ, USA, 2009. ISBN 0-470-28620-2 (cloth). xii + 260 pp. LCCN QC787.P73 H35 2009. US$27.95.

Hall:2012:DMF

[Hal12] Alice Hall. Disability and modern fiction: Faulkner, Morrison, Coetzee and the Nobel Prize for Literature. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2012. ISBN 0-230-29209-7 (hardcover). ix + 220 pp. LCCN PN56.5.H35 H35

2012. URL http://www.netread.com/jcusers2/bk1388/093/
9780230292093/image/lgcover.9780230292093.jpg.

Halpern:2015:EDS

[Hal15] Paul Halpern. Einstein’s dice and Schrödinger’s cat: how two great minds battled quantum randomness to create a unified theory of physics. Basic Books, New York, NY, USA, 2015. ISBN 0-465-07571-1 (hardcover). x + 271 pp. LCCN QC174.17.C45 H35 2015.

Halpern:2017:QLH

[Hal17] Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Basic Books, New York, NY, USA, 2017. ISBN 0-465-09758-8 (hardcover), 0-465-09759-6 (e-book). ix + 311 pp. LCCN QC174.12 .H347 2017.

Halpern:2018:CPH

[Hal18] Paul Halpern. Celebrity physicist: How the press sensationalized Einstein’s search for a unified field theory. Physics in Perspective (PIP), 20(3):254–271, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hamilton:1973:AE

[Ham73] Peter N. (Peter Napier) Hamilton. Albert Einstein. Makers of modern thought. Judson Press, Valley Forge, PA, USA, 1973. ISBN 0-8170-0595-1. 59 pp. LCCN QC6 .H275.

Hamburg:1985:SAE

[Ham85] D. A. Hamburg. The spirit of Albert Einstein toward free inquiry and judicious use of technology. Einstein Quarterly Journal of Biology and Medicine, 3(4):160–163, ???? 1985. ISSN 0724-6706.

Hammontree:1986:AEY

[Ham86] Marie Hammontree. Albert Einstein, young thinker. The Childhood of famous Americans series. Aladdin Books, New York, NY, USA, 1986. ISBN 0-02-041860-4 (paperback). 192 pp. LCCN QC16.E5 H33 1986. US$3.95. Illustrated by Robert Doremus.

Hamby:1997:BRB

[Ham97] Alonzo L. Hamby. Book reviews: The Decision to Use the Atomic Bomb and the Architecture of an American Myth, by Gar Alperovitz; The Decision to Drop the Atomic Bomb, by Dennis D. Wainstock; The Last Great Victory: The End of World War II, July/August 1945, by Stanley Weintraub; Harry S. Truman and the Bomb: A Documentary History, by Robert H. Ferrell; History Wars: The Enola Gay and Other Battles for the American Past, by Edward T. Linenthal and Tom Engelhardt. The Journal of American History, 84(2):609–614, September 1997. CODEN ???? ISSN 0021-8723 (print), 1945-2314 (elec-

tronic). URL http://jah.oxfordjournals.org/content/
84/2/609.full.pdf+html;   http://jah.oxfordjournals.
org/content/84/2/609.short;   http://www.jstor.org/
stable/2952578.

Hamblin:2008:PWR

[Ham08] Jacob Darwin Hamblin. Poison in the well: radioactive waste in the oceans at the dawn of the nuclear age. Rutgers University Press, New Brunswick, NJ, USA, 2008. ISBN 0-8135-4220-0 (hardcover). x + 311 pp. LCCN TD898 .H35 2008. URL http:

//www.loc.gov/catdir/toc/ecip0715/2007015498.html.

Hamblin:2011:BRB

[Ham11a] Jacob Darwin Hamblin. Book review: Bruce J. Hunt, Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Baltimore: Johns Hopkins University Press, 2010, 182 pages. $20.00 (paper). Physics in Perspective (PIP), 13(1):117–118, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hamblin:2011:PPL

[Ham11b] Jacob Darwin Hamblin. Pursuing power and light: Technology and physics from James Watt to Albert Einstein. Physics in Perspective (PIP), 13(1):117–118, March 2011. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hanle:1979:IBH

[Han79a] Paul A. Hanle. Indeterminacy before Heisenberg: The case of Franz Exner and Erwin Schrödinger. Historical Studies in the Physical Sciences, 10(??):225–269, ???? 1979. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757391.

Hanle:1979:SCS

[Han79b] Paul A. Hanle. The Schrödinger–Einstein correspondence and the sources of wave mechanics. American Journal of Physics, 47(7):644–648, 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.aip.

org/content/aapt/journal/ajp/47/7/10.1119/1.11950.

Hanrion:1979:PAE

[Han79c] R. Hanrion. Poor Albert Einstein. Histoire, ??(11):77–78, ????

1979. ISSN 0182-2411.

Hanle:1980:BRB

[Han80a] Paul A. Hanle. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. Isis, 71(2):354–355, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230243;
http://www.jstor.org/stable/pdfplus/230243.pdf.

Hansen:1980:ECE

[Han80b] James R. Hansen. Einstein: A Centenary Exhibition ed. by the National Museum of History and Technology (review). Technology and Culture, 21(4):693–695, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890783/pdf.

Hansen:1988:UNW

[Han88] Chuck Hansen. US nuclear weapons: the secret history. Orion Books, New York, NY, USA, 1988. ISBN 0-517-56740-7. 232 pp. LCCN QC773.3.U.

Hansen:1992:LUS

[Han92] Chuck Hansen. Letter: Ulam’s shock waves. Bulletin of the Atomic Scientists, 48(1):45, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BHM91].

Hansen:1997:LMF

[Han97] Chuck Hansen. Letter: More fallout on fallout. Bulletin of the Atomic Scientists, 53(5):3, 58, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM97].

Hansch:2006:NLP

[Hän06] Theodor W. Hänsch. Nobel Lecture: Passion for precision. Reviews of Modern Physics, 78(4):1297–1309, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.78.1297;   http://rmp.aps.org/abstract/RMP/
v78/i4/p1297_1.

Hansson:2018:AWN

[Han18] Dr. Nils Hansson. Anmerkungen zur wissenschaftshistorischen Nobelpreisforschung. (German) [Notes on science–historical Nobel Prize research]. Berichte zur Wissenschaftsgeschichte, 41(1): 7–18, March 2018. CODEN BEWID8. ISSN 0170-6233 (print),

1522-2365 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/bewi.201801869.

Hanson:2022:CDE

[Han22] Jason Hanson. Coupling the Dirac and Einstein equations through geometry. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.
1007/s10701-021-00519-9.

Harre:1969:STSa

[Har69] Rom Harré, editor. Scientific Thought 1900–1960: a Selective Survey. Clarendon Press, Oxford, UK, 1969. ISBN 0-19-858125-4, 0-19-858126-4. viii + 277 pp. LCCN Q175 .H327.

Harder:1972:CCE

[Har72] Allen Harder, M.A., Ph.D. The Copernican character of Einstein’s cosmology. Annals of Science, 29(4):339–347, 1972. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hartline:1979:EEH

[Har79a] Beverly Karplus Hartline. Einstein explores high energy astrophysics. Science, 204(4400):1399–1400, June 29, 1979. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/204/4400/1399.
full.pdf.

Hartline:1979:EPX

[Har79b] Beverly Karplus Hartline. Einstein pictures the X-ray sky. Science, 205(4401):31–33, July 6, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/205/4401/31.full.pdf.

Hartman:1979:CRE

[Har79c] Hyman Hartman. Commentary: Reflections on Einstein’s 100th birthday. Bulletin of the Atomic Scientists, 35(10):9, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harwood:1979:BRN

[Har79d] Jonathan Harwood. Book review: Nineteenth and Twentieth Centuries: Scientific Elite: Nobel Laureates in the United States. By Harriet Zuckerman. New York: The Free Press, 1977. Pp. xv + 335. $14.95. British Journal for the History of Science, 12(2):233–234, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4025767.

Harrison:1981:CSU

[Har81] Edward Robert Harrison. Cosmology, the science of the universe. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-22981-2. xi + 430 pp. LCCN QB981 .H32.

Harman:1985:WPS

[Har85] Peter Michael Harman, editor. Wranglers and physicists: studies on Cambridge physics in the nineteenth century, Studies on Cambridge mathematical physics in the nineteenth century. Manchester University Press, Manchester, UK, 1985. ISBN 0-7190-1756-4. LCCN QC9.G7 W73 1985. US$27.00 (US).

Harris:1989:ESC

[Har89] Sidney Harris. Einstein simplified: cartoons on science. Rutgers University Press, New Brunswick, NJ, USA, 1989. ISBN 0-8135-1410-X. 162 pp. LCCN NC1429.H33315 A4 1989.

US$9.95. URL http://www.loc.gov/catdir/enhancements/
fy1206/88031284-d.html.

Harman:1991:BRA

[Har91] P. M. Harman. Book reviews: Albert Einstein and John Stachel: The Collected Papers of Albert Einstein. Volume II: The Swiss Years: Writings, 1900–1909. Albert Einstein and Anna Beck: The Collected Papers of Albert Einstein. Volume II: The Swiss Years: Writings, 1990–1909. English Translation. Isis, 82(4): 768–769, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233384.

Harwit:1995:BHT

[Har95] Martin Harwit. Black Holes and Time Warps: Einstein’s Outrageous Legacy by Kip S. Thorne (review). Technology and Culture, 36(1):208–209, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888031/pdf.

Harman:1996:BRA

[Har96a] P. M. Harman. Book review: Albert Einstein, Martin J. Klein, A. J. Kox, Jurgen Renn, and Robert Schulmann: The Collected Papers of Albert Einstein. Volume 4: The Swiss Years: Writings, 1912–1914. Isis, 87(3):568–569, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236042.

Harman:1996:BRC

[Har96b] P. M. Harman. Book review: The collected papers of Albert Einstein, vol 5, The Swiss years: Correspondence, 1902–1914 — M. J. Klein, A. J. Kox, R. Schulmann. Annals of Science, 53 (1):93–94, January 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Harwit:1996:EDL

[Har96c] Martin Harwit. An exhibit denied: lobbying the history of Enola Gay. Copernicus, New York, NY, USA, 1996. ISBN 0-387-94797-3 (hardcover). xxv + 477 pp. LCCN D767.25.H6 H348 1996.

Hardy:1997:EPR

[Har97] Lucien Hardy. Einstein–Podolsky–Rosen reasoning in nonlocality theorems. In Cohen et al. [CHS97], pages 105–111. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346.

LCCN QC173.96-174.52. URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-7_8.

Harman:1998:NPJ

[Har98] P. M. (Peter Michael) Harman. The natural philosophy of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56102-7, 0-521-00585-X (paperback). xiv + 232 pp. LCCN QC16.M4 H37 1998. URL http://

www.loc.gov/catdir/description/cam028/97035703.html;
 http://www.loc.gov/catdir/toc/cam021/97035703.html;
 http://www.zentralblattmath.org/zmath/en/search/?an=
1005.01006.

Harrison:2000:CSU

[Har00] Edward Robert Harrison. Cosmology: the science of the universe. Cambridge University Press, Cambridge, UK, second edition, 2000. ISBN 0-521-66148-X. x + 567 pp. LCCN QB981 .H32 2000. URL http://www.loc.

gov/catdir/description/cam029/99010172.html;   http:
//www.loc.gov/catdir/samples/cam034/99010172.html;
http://www.loc.gov/catdir/toc/cam021/99010172.html.

Hargittai:2002:RSN

[Har02a] István Hargittai. The road to Stockholm: Nobel Prizes, science, and scientists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850912-X. xvii + 342 + 24 pp.

LCCN Q141 .H267 2002. URL http://www.loc.gov/catdir/
enhancements/fy0614/2002283888-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2002283888-b.html;
 http://www.loc.gov/catdir/toc/fy033/2002283888.html.

Harman:2002:BRB

[Har02b] P. M. Harman. Book review: Electrodynamics from Ampère to Einstein: Olivier Darrigol, Oxford University Press, Oxford and New York, 2000, xix + 532 pp., UK £75, ISBN 0-19-850594-9. Studies in History and Philosophy of Modern Physics, 33(2): 371–373, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219802000138. See [Dar00].

Harman:1990:SLP

[Har02c] P. M. (Peter Michael) Harman, editor. The scientific letters and papers of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1990–2002. ISBN 0-521-25625-9 (vol. 1), 0-521-25626-7 (vol. 2), 0-521-25627-5 (vol. 3). xxviii + 748 (vol. 1), xxx + 999 + 15 (vol. 2), xxvii + 932 pp. LCCN QC670 .M385 1990. URL http://www.loc.

gov/catdir/description/cam023/89000452.html;   http:
//www.loc.gov/catdir/enhancements/fy0732/89000452-b.
html;   http://www.loc.gov/catdir/toc/cam021/89000452.
html.

Hartcup:2003:ESS

[Har03a] Guy Hartcup. The effect of science on the Second World War. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2003. ISBN 1-4039-0643-2 (paperback), 0-333-67061-2 (hardcover). xv + 214 + 8 pp. LCCN D810.S2 H37 2003. URL http://www.

loc.gov/catdir/bios/hol052/2002042819.html;   http:
//www.loc.gov/catdir/description/hol032/2002042819.
html;   http://www.loc.gov/catdir/toc/hol031/2002042819.
html.

Hartle:2003:GIE

[Har03b] J. B. (James B.) Hartle. Gravity: an introduction to Einstein’s General Relativity. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-8053-8662-9. xxii + 582 pp. LCCN QC173.6

.H38 2003. URL http://wps.aw.com/aw_hartle_gravity_1/
0,6533,512494-,00.html.

Hargittai:2006:OVM

[Har06a] István Hargittai. Az öt világformáló marslakó. Vince, Budapest, Hungary, 2006. ISBN 963-9552-77-1. 397 + 32 pp. LCCN QC15 .H27155 2006.

Hargittai:2006:MSF

[Har06b] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp.

LCCN QC15 .H27 2006. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005029427-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2005029427-b.html;
 http://www.loc.gov/catdir/toc/ecip061/2005029427.html.

Harle:2007:BRS

[Har07] R. Harle. Book review: Sex, Drugs, Einstein, & Elves: Sushi, Psychedelics, Parallel Universes, and the Quest for Transcendence. Leonardo (Oxford, England), 40(1):93–94, ???? 2007. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (elec-

tronic). URL https://muse.jhu.edu/article/209702.

Haroche:2013:NLC

[Har13a] Serge Haroche. Nobel Lecture: Controlling photons in a box and exploring the quantum to classical boundary. Reviews of Modern Physics, 85(3):1083–1102, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.85.1083;
 http://rmp.aps.org/abstract/RMP/v85/i3/p1083_1.

Harwit:2013:STU

[Har13b] Martin Harwit. In search of the true universe: the tools, shaping, and cost of cosmological thought. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-04406-5 (hardcover), 1-107-62025-2 (paperback). xvii + 393 pp. LCCN QB32 .H27 2013.

Harzing:2013:PTG

[Har13c] Anne-Wil Harzing. A preliminary test of Google Scholar as a source for citation data: a longitudinal study of Nobel Prize winners. Scientometrics, 94(3):1057–1075, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/s11192-
012-0777-7.

Hasday:2004:AEG

[Has04] Judy L. Hasday. Albert Einstein: the Giant of 20th Century Science. Nobel Prize-winning scientists. Enslow Publishers, Berkeley Heights, NJ, USA, 2004. ISBN 0-7660-2185-8. 128 pp. LCCN

QC16.E5 H37 2004. URL http://www.loc.gov/catdir/toc/
ecip046/2003015335.html.

Hasegawa:2005:RES

[Has05] Tsuyoshi Hasegawa. Racing the enemy: Stalin, Truman, and the surrender of Japan. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2005. ISBN 0-674-01693-9 (hardcover), 0-674-03840-1 (e-book). ix + 382 pp. LCCN D813.J3 H37 2005.

Hasegawa:2007:EPW

[Has07a] Tsuyoshi Hasegawa, editor. The end of the Pacific war: reappraisals. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 2007. ISBN 0-8047-5427-6. xvi + 331 pp.

LCCN D813.J3 E64 2007. URL http://www.loc.gov/catdir/
toc/ecip0620/2006028407.html.

Hashi:2007:SET

[Has07b] Hisaki. Hashi. The significance of Einstein’s theory of relativity in Nishida’s “Logic of Field”. Philosophy East and West, 57(4):457–481, 2007. CODEN ???? ISSN 0031-8221 (print), 1529-1898 (electronic). URL https://muse.jhu.edu/

article/222654.

Hasler:2009:EPR

[Has09] Ludwig Hasler. End with planning rage. thanks to Albert Einstein. Du — die Zeitschrift der Kultur, ??(801):118, November 2009. ISSN 0012-6837.

Haukelid:1947:DDN

[Hau47] Knut Haukelid. Det demrer en dag. (Norwegian) [It dawns one day]. Nasjonalforlaget, Oslo, Norway, 1947. 189 pp. LCCN D802.N7 H36 1947.

Haukelid:1948:CKH

[Hau48] Knut Haukelid. Capitaine Knut Haukelid. L’Epopée de l’eau lourde. (French) [Captain Knut Haukelid: the epic of heavy water]. Éditions de l’Elan, Paris, France, 1948. 206 pp. LCCN ???? French translation by Georges Charbonnier of [Hau47].

Haukelid:1949:KOT

[Hau49] Knut Haukelid. Kampen om det tunge vand. (Danish) [Fight about heavy water]. Aschehoug Dansk Forlag, København, Danmark, 1949. ???? pp. LCCN ???? Danish translation by Grete Juel Jørgensen of [Hau47].

Haukelid:1950:DDN

[Hau50] Knut Haukelid. Det demrer en dag. (Norwegian) [It dawns one day]. Ponni-bok. Nasjonalforlaget, Oslo, Norway, 1950. 247 pp. LCCN D802.N7 H36 1947.

Haukelid:1953:KOT

[Hau53] Knut Haukelid. Kampen om tungtvannet. (Norwegian) [Fight about heavy water]. Essforlagene, Oslo, Norway, 1953. 205 + 16 pp. LCCN D802.N7 H36x.

Haukelid:1954:SAA

[Hau54] Knut Haukelid. Skis against the Atom: An account of the sabotage of German heavy water plant and supplies in Norway, 1942–44. William Kimber, London, UK, 1954. 201 pp. LCCN ???? Translation by Francis Hamilton Lyon of the original Norwegian Kampen om tungtvannet [Fight about heavy water].

Haukelid:1963:KOT

[Hau63] Knut Haukelid. Kampen om det tunge vand. (Danish) [Fight about heavy water]. Hirschsprung, København, Danmark, second edition, 1963. 159 pp. LCCN ???? Danish translation by Grete Juel Jørgensen of [Hau47].

Haukelid:1964:KOT

[Hau64] Knut Haukelid. Kampen om det tunge vand. (Danish) [Fight about heavy water]. Aschehoug Dansk Forlag, København, Danmark, second edition, 1964. 159 pp. LCCN ???? Danish translation by Grete Juel Jørgensen of [Hau47].

Haukelid:1965:KOT

[Hau65] Knut Haukelid. Kampen om tungtvannet. (Norwegian) [Fight about heavy water]. Krigsbokserien. Cappelens Forlag, Oslo, Norway, 1965. 192 pp. LCCN D802.N7 H36.

Haukelid:1974:ATS

[Hau74] Knut Haukelid. Attack on Telemark: Skis against the atom. Ballantine Books, New York, NY, USA, 1974. ISBN 0-345-23879-6 (hardcover). 160 pp. LCCN ????

Haukelid:1983:KOT

[Hau83] Knut Haukelid. Kampen om tungtvannet. (Norwegian) [Fight about heavy water]. Nordmenn i krig. Mortensen, Oslo, Norway, 1983. ISBN 82-527-0715-7. 205 + 12 pp. LCCN ????

Hau:2008:OIP

[Hau08] Lene Vestergaard Hau. Optical information processing in Bose– Einstein condensates. Nature Photonics, 2(8):451–453, August 2008. CODEN NPAHBY. ISSN 1749-4885 (print), 1749-

4893 (electronic). URL http://www.nature.com/nphoton/
journal/v2/n8/full/nphoton.2008.140.html.

Havranek:1977:JAE

[Hav77] Jan Havránek. Ke jmenováni Alberta Einsteina profesorem v Praze. (Czech) [On the appointment of Albert Einstein as a professor in Prague]. Acta Universitatis Carolinae — Historia Universitatis Carolinae Pragensis, 17(2):105–130, ???? 1977.

Havranek:1980:MKE

[Hav80] Jan Havránek. Materiály k Einsteinovu Pražskěmu půisobení z Archivu Univerzity Karlovy. (Czech) [Einstein’s materials: Prague work from the Archives of Charles University]. Acta Universitatis Carolinae — Historia Universitatis Carolinae Pragensis, 22(1):109–134, ???? 1980.

Havas:1988:TEP

[Hav88] Peter Havas. Translation of Einstein papers. Science, 240 (4860):1716, June 24, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/240/4860/1716.2.full.pdf.

Hawking:1988:BHT

[Haw88] S. W. (Stephen W.) Hawking. A brief history of time: from the Big Bang to black holes. Bantam Books, New York, NY, USA, 1988. ISBN 0-553-05243-8. x + 198 pp. LCCN QB981 .H377

1988. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
random0410/87033333.html.

Hawking:1992:SHB

[Haw92] S. W. (Stephen W.) Hawking, editor. Stephen Hawking’s A brief history of time: a reader’s companion. Bantam Books, New York, NY, USA, 1992. ISBN 0-553-07772-4. ix + 194 pp. LCCN QB981.H377 S74 1992. US$25.00.

Hawking:1998:BHT

[Haw98] S. W. (Stephen W.) Hawking. A brief history of time. Bantam Books, New York, NY, USA, updated and expanded tenth anniversary edition, 1998. ISBN 0-553-10953-7, 0-553-38016-8. ix + 212 pp. LCCN QB981 .H377 1998. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/description/random043/98021874.html.

Hawking:2001:UN

[Haw01] S. W. (Stephen W.) Hawking. The universe in a nutshell. Bantam Books, New York, NY, USA, 2001. ISBN 0-553-80202-X. viii + 216 pp. LCCN QC174.12 .H39 2001. URL http://

www.loc.gov/catdir/bios/random053/2001035757.html;
http://www.loc.gov/catdir/description/random046/2001035757.
html;   http://www.loc.gov/catdir/samples/random042/
2001035757.html.

Haw:2002:CSB

[Haw02a] M. D. Haw. Colloidal suspensions, Brownian motion, molecular reality: a short history. Journal of Physics: Condensed Matter, 14(33):7769–7779, August 26, 2002. CODEN JCOMEL. ISSN 0953-8984 (print), 1361-648X (electronic). URL http:

//iopscience.iop.org/0953-8984/14/33/315;   http://
www.physik.uni-augsburg.de/theo1/hanggi/Colloidal.pdf.

Hawking:2002:FS

[Haw02b] S. W. (Stephen W.) Hawking, editor. The future of spacetime. W. W. Norton & Co., New York, NY, USA, 2002. ISBN 0-393-02022-3. 220 pp. LCCN QC173.59.S65 F87 2002.

Hawking:2002:SGG

[Haw02c] S. W. (Stephen W.) Hawking, editor. On the shoulders of giants: the great works of physics and astronomy. Running Press, Philadelphia, PA, USA, 2002. ISBN 0-7624-1348-4. xiii + 1264 pp. LCCN QC6.2 .O5 2002. URL http://www.loc.gov/

catdir/toc/fy036/2002100441.html.

Hawking:2002:SPR

[Haw02d] S. W. (Stephen W.) Hawking, editor. Selections from the Principle of Relativity. On the shoulders of giants. Running Press, Philadelphia, PA, USA, 2002. ISBN 0-7624-2023-5. xvi + 100 pp. LCCN ????

Haw:2005:ERW

[Haw05] Mark Haw. Einstein’s random walk. Physics World, 18(1):19– 22, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/1/phwv18i1a25.pdf;   http:/
/users-phys.au.dk/fogedby/statphysII/notes/Einstein'
s-random-walk.pdf.

Hawking:2007:SPI

[Haw07] Stephen Hawking, editor. A stubbornly persistent illusion: the essential scientific works of Albert Einstein. Running, Philadelphia, PA, USA, 2007. ISBN 0-7624-3003-6 (hardcover). xi + 468 pp. LCCN QC16.E5 A2 2007. URL http://www.loc.

gov/catdir/enhancements/fy0832/2007935658-b.html;
http://www.loc.gov/catdir/enhancements/fy0832/2007935658-
d.html;   http://www.loc.gov/catdir/toc/fy0805/2007935658.
html.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Hayano:1991:AEA

[Hay91] T. Hayano. Albert Einstein achievement life and psychopathology time series analysis applied in the number of scientific reports. Journal of the Wakayama Medical Society, 42(1):79–92, ???? 1991. ISSN 0043-0013.

Hayden:2002:IE

[Hay02] T. Hayden. The inner Einstein. US News and World Report, 133(22):60–66, 68–69, December 9, 2002. CODEN XNWRAV. ISSN 0041-5537.

Heisenberg:1963:KDQ

[HB63] Werner Heisenberg and Niels Bohr, editors. Die Kopenhagener Deutung der Quantentheorie. (German) [The Copenhagen Interpretation of Quantum Theory]. E. Battenberg, Stuttgart, Germany, 1963. 67 pp. LCCN QC174.1 .K65.

Holton:2001:PHA

[HB01] Gerald Holton and Stephen G. Brush. Physics, the human adventure: from Copernicus to Einstein and beyond. Rutgers University Press, New Brunswick, NJ, USA, third edition, 2001. ISBN 0-8135-2907-7, 0-8135-2908-5. xv + 582 pp. LCCN Q160 .H654 2001.

Hull:2005:RPH

[HB05] McAllister H. Hull and Amy Bianco. Rider of the pale horse: a memoir of Los Alamos and beyond. University of New Mexico Press, Albuquerque, NM, USA, 2005. ISBN 0-8263-3553-5. 158 pp. LCCN QC16.H85 A2 2005. URL http://www.loc.gov/

catdir/toc/ecip0511/2005011952.html.

Heisenberg:1971:QFP

[HBD71] Werner Heisenberg, Erich Bagge, and H. Peter Dürr, editors. Quanten und Felder; physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg. (German) [Quanta and fields: physical and philosophical views. On the 70th Birthday of Werner Heisenberg]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1971. 366 pp. LCCN QC174.45 .Q35.

Halling:2018:LNK

[HBFH18] Thorsten Halling, Ragnar Björk, Heiner Fangerau, and Nils Hansson. Leopoldina: Ein Netzwerk für künftige Nobelpreisträger für Physiologie oder Medizin? / Leopoldina: a network for future Nobel Laureates in Physiology or Medicine? Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 102 (2):211–233, 2018. CODEN SUARAH. ISSN 0039-4564. URL

https://www.jstor.org/stable/44988538.

Hoyle:2000:DAC

[HBN00] Sir Fred Hoyle, Geoffrey R. Burbidge, and Jayant Vishnu Narlikar. A Different Approach to Cosmology: From a Static Universe Through the Big Bang Towards Reality. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66223-0 (hardcover), 0-521-01926-5 (paperback). xi + 357 pp. LCCN QB981 .H754 2000. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/description/cam029/99015821.html;   http:
//www.loc.gov/catdir/samples/cam032/99015821.html;
http://www.loc.gov/catdir/toc/cam021/99015821.html.

Hilgartner:1982:NNL

[HBO82] Stephen Hilgartner, Richard C. Bell, and Rory O’Connor. Nukespeak: nuclear language, visions, and mindset. Sierra Club Books, San Francisco, CA, USA, 1982. ISBN 0-87156-307-X. xiv + 282 pp. LCCN QC773.3.U5 H54 1982.

Heylighen:1999:ECV

[HBR99] Francis Heylighen, Johan Bollen, and Alexander Riegler, editors. The evolution of complexity: the violet book of “Einstein meets Magritte”, volume 8 of Einstein meets Magritte. VUB University Press, Brussels, Belgium, 1999. ISBN 0-7923-5764-7 (hardcover). LCCN Q172.5.C45 E945 1999. See other volumes in this series [Aer99a, AMM99, Aer99b, ABW99, AVvdV99, BAM99, CSvB99].

Hoddeson:1997:RSM

[HBRD97] Lillian Hoddeson, Laurie Brown, Michael Riordan, and Max Dresden, editors. The rise of the Standard Model: particle physics in the 1960s and 1970s. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-57082-4 (hardcover), 0-521-57816-7 (paperback). xxx + 714 pp. LCCN QC794.6.S75 R57 1997.

Highfield:1993:PLA

[HC93] J. R. L. (John Roger Loxdale) Highfield and Paul Carter. The private lives of Albert Einstein. Faber, London, UK, 1993. ISBN 0-571-16744-6, 0-571-17170-2s. xii + 355 + 8 pp. LCCN ????

Highfield:1994:GLA

[HC94a] Roger Highfield and Paul Carter. Die geheimen Leben des Albert Einstein: eine Biographie. (German) [The private lives of Albert Einstein]. Byblos-Verlag, Berlin, Germany, 1994. ISBN 3-929029-33-2. 409 pp. LCCN ???? URL http://d-nb.info/ 940315955/0. German translation by Anita Ehlers of [HC93].

Highfield:1994:VLV

[HC94b] Roger Highfield and Paul Carter. Het verborgen leven van Albert Einstein. (Dutch) [The private lives of Albert Einstein]. Aramith, Bloemendaal, The Netherlands, 1994. ISBN 90-6834-148-0. 368 pp. LCCN ???? Dutch translation of [HC93] by Jan van de Craats.

Highfield:1994:VSA

[HC94c] Roger Highfield and Paul Carter. Le vite segrete di Albert Einstein. (Italian) [The private lives of Albert Einstein]. Muzzio biblioteca. F. Muzzio, Padova, Italy, 1994. ISBN 88-7021-712-4s. xiii + 331 pp. LCCN ???? Italian translation of [HC93] by Giovanna Mannino.

Highfield:1994:PLA

[HC94d] Roger Highfield and Paul Carter. The private lives of Albert Einstein. St. Martin’s Press, New York, NY, USA, 1994. ISBN 0-312-11047-2. xii + 353 + 8 pp. LCCN QC16.E5 H54 1994. US$23.95; CAN$33.50.

Highfield:1995:PCA

[HC95] Roger Highfield and Paul Carter. Prywatne çzycie Alberta Einsteina. (Polish) [The private lives of Albert Einstein]. Na ścieçzkach nauki. Prószyński, Warszawa, Poland, 1995. ISBN 83-86669-55-1. 416 pp. LCCN ???? Polish translation of [HC93] by Marek Krośniak.

Highfield:1996:GLA

[HC96a] Roger Highfield and Paul Carter. Die geheimen Leben des Albert Einstein: eine Biographie. (German) [The private lives of Albert Einstein: a biography], volume 30561 of Dtv. Dt. Taschenbuch-Verlag, München, Germany, 1996. ISBN 3-423-30561-4. 409 pp.

LCCN ???? URL http://d-nb.info/948704667/0. German

translation of [HC93].

Highfield:1996:VLV

[HC96b] Roger Highfield and Paul Carter. Het verborgen leven van Albert Einstein. (Dutch) [The private lives of Albert Einstein], volume 296 of Rainbow pocketboeken. Muntinga, Amsterdam, The Netherlands, 1996. ISBN 90-417-0044-7. 520 pp. LCCN ???? Dutch translation of [HC93] by Jan van de Craats.

Highfield:1996:VPE

[HC96c] Roger Highfield and Paul Carter. Las vidas privadas de Einstein. (Spanish) [The private lives of Albert Einstein]. Biografías Espasa. Espasa Calpe, Madrid, Spain, 1996. ISBN 84-239-2282-0. 356 pp. LCCN ???? Spanish translation of [HC93] by Javier Alfaya.

Highfield:1998:ECZ

[HC98] Roger Highfield and Paul Carter. Einshtein: chastnaìa zhizn’. (Russian) [The private lives of Albert Einstein]. Zakharov, Moskva, Russia, 1998. ISBN 5-8159-0009-5 (hardcover). 367 + 16 pp. LCCN QC16.E5 H5417 1998x. US$12.50.

Highfield:2004:GLA

[HC04] Roger Highfield and Paul Carter. Die geheimen Leben des Albert Einstein: eine Biographie. (German) [The private lives of Albert Einstein: a biography]. Marixverlag, Wiesbaden, Germany, 2004. ISBN 3-937715-43-6. 409 pp. LCCN ???? German translation of [HC93].

Hebel:1978:RAP

[HCD^{+}78] L. Charles Hebel, Eldon L. Christensen, Fred A. Donath, Warren E. Falconer, Leon J. Lidofsky, Ernest J. Moniz, Thomas H. Moss, Robert L. Pigford, Thomas H. Pigford, Gene I. Rochlin, Robert H. Silsbee, Mcdonald E. Wrenn, Hans Frauenfelder, Theodore L. Cairns, W. K. H. Panofsky, and M. Gene Simmons (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society by the study group on nuclear fuel cycles and waste management. Reviews of Modern Physics, 50(1):S1–S176, January 1978. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
50.S1;   http://rmp.aps.org/abstract/RMP/v50/i1/pS1_
1.

Hewlett:1969:HUS

[HD69] Richard G. Hewlett and Francis Duncan. A history of the United States Atomic Energy Commission. Volume 2. Atomic Shield, 1947–1952. Pennsylvania State University Press, University Park, PA, USA, 1969. 718 pp. LCCN QC791.948.

Hoffmann:1972:AEC

[HD72] Banesh Hoffmann and Helen Dukas. Albert Einstein, creator and rebel. Viking, New York, NY, USA, 1972. ISBN 0-670-11181-3. xv + 272 pp. LCCN QC16.E5 H64 1972.

Hoffmann:1973:AEC

[HD73] Banesh Hoffmann and Helen Dukas. Albert Einstein, creator and rebel. Hart-Davis, MacGibbon, London, UK, 1973. ISBN 0-246-10606-9. xv + 272 pp. LCCN QC16.E5 H64 1972.

Hoffman:1975:AEC

[HD75] Banesh Hoffman and Helen Dukas. Albert Einstein, créateur et rebelle. (French) [Albert Einstein, creator and rebel]. Éditions du Seuil, Paris, France, 1975. 298 pp. LCCN QC16.E5 H6312.

Hoddeson:2002:TGL

[HD02] Lillian Hoddeson and Vicki Daitch. True genius: the life and science of John Bardeen: the only winner of two Nobel Prizes in physics. Joseph Henry Press, Washington, DC, USA, 2002. ISBN 0-309-08408-3, 0-309-09511-5 (paperback). xi + 467 + 8 pp. LCCN QC16.B27 H63 2002. URL http://www.loc.gov/

catdir/toc/fy037/2002007967.html.

Hammontree:2014:AE

[HD14] Marie Hammontree and Robert Doremus. Albert Einstein. History’s all-stars. Aladdin, New York, NY, USA, Aladdin edition, 2014. ISBN 1-4814-1497-6 (hardcover), 1-4814-1496-8 (paperback), 1-4814-1498-4 (e-book). 195 pp. LCCN PZ7.H186 Ak 2014.

Hewlett:1989:HUS

[HDHA89] Richard G. Hewlett, Francis Duncan, Jack M. Holl, and Oscar E. Anderson, editors. A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1989. ISBN 0-520-07187-5 (vol. 1), 0-520-07187-5 (vol. 2), 0-520-06018-0 (vol. 3). ???? pp. LCCN HD9698.U52 H55 1989.

Holton:1982:AEH

[HE82] Gerald James Holton and Yehuda Elkana, editors. Albert Einstein, historical and cultural perspectives: the centennial symposium in Jerusalem. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Hawking:2008:LSS

[HE08] S. W. (Stephen W.) Hawking and George Francis Rayner Ellis. The large scale structure of space–time. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-09906-4 (paperback). xi + 391 pp. LCCN QC173.59.S65 H38 2008.

Heathcote:1964:BRN

[Hea64] N. H. De V. Heathcote. Book review: Nobel: The Man and his Prizes. Edited by the Nobel Foundation. Pp. xii + 690. Frontispiece. Amsterdam, London and New York: Elsevier Publishing Company, 1962. 60s. net. British Journal for the History of Science, 2(1):79–80, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4025091.

Healey:2011:REB

[Hea11] Richard Healey. Reduction and emergence in Bose–Einstein condensates. Foundations of Physics, 41(6):1007–1030, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-010-9481-8.

Heafner:2013:ETR

[Hea13] Joe Heafner. Book review: The Einstein Theory of Relativity: A Trip to the Fourth Dimension, by Lillian R. Lieber, with illustrations by Hugh Gray Lieber, edited and with a new Forward by David Derbes and Robert Jantzen, and published by Paul Dry Books, Philadelphia, pp. xviii + 350 (2008) $10.00 paperback, ISBN 978-1-58988-044-3. The Physics Teacher, 51(3): 191, March 2013. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Heacox:2015:EUP

[Hea15] William D. Heacox. The expanding universe: a primer on relativistic cosmology. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-11752-6 (hardcover). xx + 272 pp. LCCN QB991.E94 H43 2015. URL http:

//www.cambridge.org/us/academic/subjects/astronomy/
cosmology-and-relativity/expanding-universe-primer-
relativistic-cosmology.

Haensel:1980:BFK

[HEB^{+}80] R. Haensel, Karl Ehrlich, K. Bethge, H. Soffel, W. von Witsch, and H. Rechenberg. Bücher Filme: Kunz: Synchrotron Radiation Techniques and Applications/Lehmann: Interaction of Radiation with Solids and Elementary Defect Production/ Olson u. Schumacher: Collective Ion Acceleration/Brosche u Sündermann: Tidal Friction and the Earth’s Rotation/ Haas u. Koch: Physik-Lehrbuch für Pharmazeuten und Mediziner/Hermann, Meyenn u. Weisskopf: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Bd. I: 1919–1929. Physikalische Blätter, 36(10):318–320, October 1980. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19800361013/abstract.

Hecht:2011:HEC

[Hec11] Eugene Hecht. How Einstein confirmed E_{0} = mc^{2}. American Journal of Physics, 79(6):591–600, June 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v79/i6/p591_s1.

Hecht:2012:BNA

[Hec12] Gabrielle Hecht. Being Nuclear: Africans and the Global Uranium Trade. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01726-1 (hardcover). xx + 451 pp. LCCN HD9539.U72

A43 2012. URL https://en.wikipedia.org/wiki/Uranium_
in_Africa.

Hecker:2016:DCH

[Hec16] Siegfried S. Hecker. Doomed to Cooperate: How American and Russian Scientists Joined Forces to Avert Some of the Greatest Post-Cold War Dangers. Bathtub Row Press, Los Alamos, NM, USA, 2016. ISBN 0-941232-44-1 (hardcover). xxviii + 534 + xii (vol. 1), 436 (vol. 2) pp. LCCN JX1428.R8 D66 2016.

Heeger:2001:NLS

[Hee01] Alan J. Heeger. Nobel Lecture: Semiconducting and metallic polymers: The fourth generation of polymeric materials. Reviews of Modern Physics, 73(3):681–700, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.681;   http://rmp.aps.org/
abstract/RMP/v73/i3/p681_1.

Heffermehl:2010:NPP

[Hef10] Fredrik S. Heffermehl. The Nobel Peace Prize: what Nobel really wanted. Praeger, Santa Barbara, CA, USA, 2010. ISBN 0-313-38744-3 (hardcover), 0-313-38745-1 (e-book). xii + 239 pp. LCCN JZ5537 .H44 2010.

Hehr:1998:NER

[Heh98] G. S. Hehr. Nobel episode revisited. Nyoom-Aneesh, Toronto, ON, Canada, 1998. ISBN 0-9683556-0-9. xi + 99 pp. LCCN ????

Heisenberg:1932:NPP

[Hei32] Werner Heisenberg. The Nobel Prize in Physics 1932.

Nobelprize.org, 1932. URL http://www.nobelprize.org/
nobel_prizes/physics/laureates/1932/. The Nobel Prize

in Physics 1932 was awarded to Werner Heisenberg “for the creation of quantum mechanics, the application of which has, inter alia, led to the discovery of the allotropic forms of hydrogen”.

Heisenberg:1934:EQG

[Hei34] Werner Heisenberg. Die Entwicklung der Quantenmechanik. (German) [The development of quantum mechanics]. In Die moderne Atomtheorie: Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Lectures on modern atomic theory: the acceptance of the Nobel Prize in 1933 in Stockholm], pages 1–18. S. Hirzel, Leipzig, 1934. LCCN ????

Heisenberg:1935:WHA

[Hei35] Werner Heisenberg. Werner Heisenberg [autobiographical note]. In ????, editor, Les Prix Nobel en 1933. (French) [The Nobel Prizes in 1933], page 85. P. A. Norstedt & Söner, Stockholm, Sweden, 1935. LCCN ????

Heisenberg:1946:ATA

[Hei46] Werner Heisenberg. Über die Arbeiten zur Technischen Ausnutzung der Atomkernenergie in Deutschland. (German) [On the work toward the technical use of atomic energy in Germany]. Die Naturwissenschaften, 33(11):325–329, 1946. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1947:RGTa

[Hei47a] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Nature, 160(4059):211–215, August 16, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v160/n4059/pdf/160211a0.pdf. Abridged translation of Ger-

man original [Hei46]. See also [Hei53b, Hei53c], and response [Gou47a].

Heisenberg:1947:RGT

[Hei47b] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Nature, 160(4059):211–215, August 16, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v160/n4059/pdf/160211a0.pdf. Abridged translation of Ger-

man original [Hei46]. See also [Hei53b, Hei53c], and response [Gou47a].

Heisenberg:1949:AEJ

[Hei49a] Werner Heisenberg. Albert Einstein 70 Jahre alt. (German) [Albert Einstein 70 years old]. Welt am Sonntag, 11(??):5, March 13, 1949.

Heitler:1949:DCT

[Hei49b] Walter Heitler. The departure from classical thought in modern physics. In Schilpp [Sch49b], chapter 6, pages 179–198. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Heisenberg:1950:PPQ

[Hei50] Werner Heisenberg. The physical principles of the quantum theory. Dover, New York, NY, USA, 1950. ISBN 0-486-60113-7. 183 pp. LCCN QC174.1 .H4 1950. Translated into English by Carl Eckart and Frank C. Hoyt.

Heisenberg:1953:MSJ

[Hei53a] Werner Heisenberg. Aus meinen späten Jahren, von Albert Einstein. (Übersetzung von Dr. Hildegard Blomeyer, nach der amerikanischen Ausgabe “Out of my later years”, Philosophical Library, New York 1950). Deutsche Verlagsanstalt, Stuttgart. 1952. 280 S., Leinen DM 13.80. (German) [Out of my later years, by Albert Einstein. (Translation by Dr. Hildegard Blomeyer, according to the American edition “Out of my later years”, Philosophical Library, New York, 1950). German Publishing house, Stuttgart. 1952, 280 pp, Linen DM 13.80]. Angewandte Chemie, 65(1):43, 1953. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1953:NPa

[Hei53b] Werner Heisenberg. Nuclear Physics. Methuen and Co. Ltd., London, UK, 1953. ???? pp. LCCN ???? English translation of [HW49].

Heisenberg:1953:NPb

[Hei53c] Werner Heisenberg. Nuclear Physics. Philosophical Library, New York, NY, USA, 1953. ix + 225 pp. LCCN QC173 .H3854.

URL http://alsos.wlu.edu/information.aspx?id=95. En-

glish translation of [HW49]. See also [Hei47b].

Heisenberg:1955:AEW

[Hei55] Werner Heisenberg. Albert Einsteins wissenschaftliches Werk. (German) [Albert Einstein’s scientific work]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 10(??):897–902, ???? 1955. CODEN UNIVA8. ISSN 0041-9079.

Heisenberg:1961:MPa

[Hei61a] Werner Heisenberg, editor. On modern physics. Orion Press, London, UK, 1961. 108 pp. LCCN QC71 .O5 1961.

Heitler:1961:ES

[Hei61b] Walter Heitler. Erwin Schrödinger. 1887–1961. Biographical Memoirs of Fellows of the Royal Society, 7(??):221–228, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769408.

Heisenberg:1965:EQG

[Hei65a] Werner Heisenberg. Die Entwicklung der Quantenmechanik. (German) [The development of quantum mechanics]. In ????, editor, Nobel Lectures Including Presentation Speeches and Laureates: Biographies: Physics 1922–1941, pages 290–301. Elsevier, Amsterdam, The Netherlands, 1965. LCCN ???? Reprint of [Hei34].

Heisenberg:1965:GNP

[Hei65b] Werner Heisenberg. Gespräch mit Nobelpreisträger Professor Dr. Werner Heisenberg Physik. (German) [Interview with Nobel Laureate Professor Dr. Werner Heisenberg, physics]. In ????, editor, Wohin führt die Wissenschaft? Neun Gespräche mit deutschen Gelehrten. Aus einer Sendereihe von Radio Bremen. Bremer Beitrage. (German) [Where is the science? Nine interviews with German scholars. From a series of Radio Bremen. Bremen contributions], pages 39–54. Verlag B. C. Heye, Bremen, West Germany, 1965.

Heisenberg:1969:TGG

[Hei69a] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. R. Piper, München, Germany, 1969. ISBN 3-492-01791-6. 333 pp. LCCN ????

Heisenberg:1969:NP

[Hei69b] Werner Heisenberg. Nuclear Physics. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1969. ISBN 0-8371-2089-6. viii + 224 pp. LCCN QC173 .H3854. Reprint of [Hei53b]. See also [Hei47b].

Heisenberg:1971:PBEb

[Hei71a] Werner Heisenberg. Physics and beyond: encounters and conversations, volume 42 of World perspectives. Harper & Row, New York, NY, USA, 1971. xviii + 247 pp. LCCN QC173 .H38613 1971. Translated from the German by Arnold J. Pomerans.

Heisenberg:1971:PBE

[Hei71b] Werner Heisenberg. Physics and beyond; encounters and conversations, volume 42 of World perspectives. Harper & Row, New York, NY, USA, 1971. xviii + 247 pp. LCCN QC173 .H38613 1971. Translated from the German by Arnold J. Pomerans.

Heisenberg:1972:TGG

[Hei72] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. R. Piper, München, Germany, fourth edition, 1972. ISBN 3-492-01791-6. 333 pp. LCCN ????

Heilbron:1974:BRBa

[Hei74a] J. L. Heilbron. Book review: Albert Einstein Creator and Rebel by Banesh Hoffmann; Helen Dukas. Isis, 65(1):128–129, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228917.

Heisenberg:1974:AF

[Hei74b] Werner Heisenberg. Across the Frontiers. Harper & Row, New York, NY, USA, 1974. ISBN 0-06-011823-7. ???? pp. LCCN

???? URL http://alsos.wlu.edu/information.aspx?id=
941.

Heisenberg:1974:SWA

[Hei74c] Werner Heisenberg. The scientific work of Albert Einstein. In Across the Frontiers [Hei74b], pages 1–7. ISBN 0-06-011823-

7. LCCN ???? URL http://alsos.wlu.edu/information.
aspx?id=941.

Heisenberg:1975:BGA

[Hei75a] Werner Heisenberg. Begegnungen und Gesprache mit Albert Einstein. (German) [Encounters and conversations with Albert Einstein]. In Tradition in der Wissenschaft. (German) [Tradition in science] [Hei77], pages 111–125. ISBN 3-492-00454-7. LCCN ???? English translation in [Hei83b].

Heisenberg:1975:BZG

[Hei75b] Werner Heisenberg. Die Beziehungen zwischen Gesellschaft und Wissenschaft im Spiegel der Lindauer Tagungen. (German) [The relationship between society and science in the Lindau Meetings]. In A. Dées de Steno, editor, Nobel führte sie zusammen. Begegnungen in Lindau. (German) [Nobel brought them together. Meetings at Lindau], pages 84–99. Belser Presse, Stuttgart, West Germany, 1975. LCCN ????

Heisenberg:1977:TWG

[Hei77] Werner Heisenberg. Tradition in der Wissenschaft. (German) [Tradition in science]. Piper, Munich, West Germany, 1977. ISBN 3-492-00454-7. 145 pp. LCCN ????

Heilbron:1979:BRB

[Hei79a] J. L. Heilbron. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. Science (New Series), 204(4394):740– 741, May 18, 1979. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1748293.pdf.

Heilbron:1979:MBB

[Hei79b] J. L. Heilbron. Max Born: Book review: My Life: Recollections of a Nobel Laureate. Science, 204(4394):740–741, May 18, 1979. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/204/4394/740.full.pdf.

Heilbron:1982:FSP

[Hei82] J. L. Heilbron. Fin-de-siècle physics. In Bernhard et al. [BCS82], pages 51–73. ISBN 0-08-027939-2. LCCN Q124.6 .N63

1981. US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500098.

Heims:1983:SLS

[Hei83a] S. J. Heims. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. Technology Review (M.I.T.), 86(5): 67–68, ???? 1983. CODEN TEREAU. ISSN 0040-1692.

Heisenberg:1983:ECA

[Hei83b] Werner Heisenberg. Encounters and conversations with Albert Einstein. In Tradition in Science [Hei83c], pages 107–122. ISBN 0-8164-2488-8. LCCN Q175 .H3932 1981. English translation of [Hei75a].

Heisenberg:1983:TS

[Hei83c] Werner Heisenberg. Tradition in Science. Seabury Press, New York, NY, USA, 1983. ISBN 0-8164-2488-8. 141 pp. LCCN Q175 .H3932 1981. All German chapters translated by Peter Heath.

Heilbron:1986:DUM

[Hei86] J. L. Heilbron. The dilemmas of an upright man: Max Planck as spokesman for German science. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1986. ISBN 0-520-05710-4. xiii + 238 + 16 pp. LCCN QC16.P6 H45 1986.

Heisenberg:1989:EEO

[Hei89] Werner Heisenberg. Encounters with Einstein: And other essays on people, places, and particles. Princeton Science Library. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-02433-2. vi + 141 pp. Reprint of the 1983 original, With an epilogue by Hans-Peter Dürr.

Heisenberg:1993:FSQ

[Hei93] Werner Heisenberg. Fortschritte und Schwierigkeiten in der Quantentheorie der Elementarteilchen (Lindau 1953). (German) [Advances and difficulties in the quantum theory of elementary particles (Lindau 1953)]. In Blum et al. [BDR93], pages 309– 319. ISBN 3-540-13848-X. LCCN ???? Lecture presented 2 July 1953 at the Nobel Prize Winners’ Meeting in Lindau, West Germany.

Heisenberg:1995:LBP

[Hei95] Werner Heisenberg. A lecture on bomb physics: February 1942. Physics Today, 48(6):27–30, June 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Translated by William Sweet, with an introduction by David Cassidy.

Heilbron:2000:DUM

[Hei00] J. L. Heilbron. The dilemmas of an upright man: Max Planck and the fortunes of German science. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00439-6. xiii + 254 pp. LCCN QC16.P6 H45 2000.

Heilbron:2005:OGH

[Hei05] J. L. Heilbron, editor. The Oxford guide to the history of physics and astronomy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517198-5. xxi + 358 pp.

LCCN QC7 .O94 2005. URL http://www.loc.gov/catdir/
enhancements/fy0639/2004020388-d.html;   http://www.
loc.gov/catdir/enhancements/fy0725/2004020388-b.html.

Heighway:2010:EAV

[Hei10] Jack Heighway. Einstein, the Aether and Variable Rest Mass. Lulu.com, ????, 2010. ISBN 1-61658-620-6. 174 pp. LCCN ????

Held:1980:GRG

[Hel80] Alan Held, editor. General Relativity and gravitation: one hundred years after the birth of Albert Einstein. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40265-3 (vol. 1), 0-306-40266-1 (vol. 2). xix + 598 (vol. 1), vii + 540 (vol. 2) pp. LCCN QC173.6 G464.

Hellman:1982:EBS

[Hel82a] Geoffrey Hellman. Einstein and Bell: Strengthening the case for microphysical randomness. Synthese, 53(3):445–460, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/article/10.
1007/BF00486161.

Hellman:1982:SEL

[Hel82b] Geoffrey Hellman. Stochastic Einstein-locality and the Bell theorems. Synthese, 53(3):461–503, December 1982. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486162.

Hellberg:1983:AN

[Hel83] Thomas Hellberg. Alfred Nobel. Alno, Stockholm, Sweden, 1983. ISBN 91-7260-977-X. 141 pp. LCCN ????

Heller:1984:AES

[Hel84] M. Heller. Albert Einstein Special Relativity, emergence (1905) and early interpretation (1905–1911) — A. I. Miller. Review of Metaphysics, 37(3):642–643, ???? 1984. ISSN 0034-6632.

Held:1998:BED

[Hel98] Carsten Held. Die Bohr–Einstein-Debatte: Quantenmechanik und physikalische Wirklichkeit. (German) [The Bohr–Einstein Debate: Quantum Mechanics and Physical Reality]. Ferdinand Schöningh, Paderborn, Germany, 1998. ISBN 3-506-73823-2 (paperback). 292 pp. LCCN 0.2hel a0160 a0165 a0170 a0365.

Heller:2009:CCV

[Hel09a] Michael Heller. Continuous creation versus a beginning. In Ultimate Explanations of the Universe [Hel09c], chapter 5, pages 43–55. ISBN 3-642-02102-6 (hardcover), 3-642-02103-4 (e-book). LCCN QB981 .H4513 2009.

Heller:2009:PEU

[Hel09b] Michael Heller. Problems with the eternity of the universe. In Ultimate Explanations of the Universe [Hel09c], chapter 2, pages 15–22. ISBN 3-642-02102-6 (hardcover), 3-642-02103-4 (e-book). LCCN QB981 .H4513 2009.

Heller:2009:UEU

[Hel09c] Michael Heller. Ultimate Explanations of the Universe. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-02102-6 (hardcover), 3-642-02103-4 (e-book). xii + 216 pp. LCCN QB981 .H4513 2009.

Held:2011:SGS

[Hel11] E. B. Held. A spy’s guide to Santa Fe and Albuquerque. University of New Mexico Press, Albuquerque, NM, USA, 2011. ISBN 0-8263-4935-8 (paper). xxvii + 95 pp. LCCN E743.5 .H415 2011.

Helsvig:2014:APM

[Hel14] Kim G. Helsvig. The Abel prize: The missing Nobel in mathematics? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):1–30, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Held:2015:EBI

[Hel15] Carsten Held. Einstein’s boxes: Incompleteness of quantum mechanics without a separation principle. Foundations of Physics, 45(9):1002–1018, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-014-9845-6.

Henry:1920:CE

[Hen20] M. A. Henry. Checking up Einstein. Scientific American, 122(24):646–647, June 12, 1920. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v122/n24/
pdf/scientificamerican06121920-646.pdf.

Hendry:1981:PP

[Hen81] John Hendry. Pauli as philosopher. The British Journal for the Philosophy of Science, 32(3):277–282, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/32/3/277.
full.pdf+html.

Hendry:1982:AEF

[Hen82a] J. Hendry. Albert Einstein — four commemorative lectures — L. S. Swenson, C. P. Snow, H. Stein, and I. Prigogine. British Journal for the History of Science, 15(50):200–201, ???? 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Hendry:1982:BRNa

[Hen82b] John Hendry. Book review: The New Physics Loyd S. Swenson Jr, C. P. Snow, Howard Stein and Ilya Prigogine, Albert Einstein: four commemorative lectures. Austin: The Humanities Research Center, The University of Texas at Austin, 1979. Pp. 64. $3.50. A. P. French (ed.), Einstein. A centenary volume. London: Heinemann Educational Books, 1979. Pp. xx + 332. £10.50. Colette M. Kinnon with A. N. Kholodinin and J. G. Richardson, The impact of modern scientific ideas on society: in commemoration of Einstein. Dordrecht, Boston & London: D. Reidel, 1981. Pp. xiv + 203. Df150.00/$26.50. British Journal for the History of Science, 15(2):200–201, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4025973.

Hendry:1984:CQM

[Hen84] John Hendry. The creation of quantum mechanics and the Bohr– Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Hentschel:1985:FVM

[Hen85] Klaus Hentschel. On Feyerabend’s version of ‘Mach’s theory of research and its relation to Einstein’. Studies in History and Philosophy of Science Part A, 16(4):387–394, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900196.

Hendry:1986:BRJ

[Hen86a] John Hendry. Book reviews: Jagdish Mehra & Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 1, The Quantum Theory of Planck, Einstein, Bohr and Sommerfeld: Its Foundations and the Rise of its Difficulties 1900– 1925. Volume 2, The Discovery of the Quantum Mechanics 1925. Volume 3, The Formulation of Matrix Mechanics and its Modifications 1925–1926. Volume 4, The Fundamental Equations of Quantum Mechanics 1925–1926 and The Reception of the New Quantum Mechanics 1925–1926. New York, Heidelberg and Berlin: Springer-Verlag, 1982. Volume 1 in two parts, pp. xlvii + 372, vi + 506. ISBN 3-540-90642-8, 3-540-90667-3. DM 75, DM 85. Volume 2, pp. vii + 355. ISBN 3-540-90674-6. DM 65. Volume 3, pp. vii + 334. ISBN 3-540-90675-4. DM 75. Volume 4, pp. viii + 322. ISBN 3-540-90680-0. DM 75. Andrew Pickering. Constructing Quarks. A Sociological History of Particle Physics. Edinburgh: Edinburgh University Press, 1984. Pp. xi + 468. ISBN 0-85224-458-4 £20. British Journal for the History of Science, 19(2):206–208, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026596.

Hendry:1986:JCM

[Hen86b] John Hendry. James Clerk Maxwell and the theory of the electromagnetic field. Adam Hilger Ltd., Bristol, UK, 1986. ISBN 0-85274-563-X. xix + 305 pp. LCCN QC665.E4

H46 1986. URL http://www.loc.gov/catdir/enhancements/
fy0745/86196851-d.html.

Hentschel:1986:KES

[Hen86c] Klaus Hentschel. Die Korrespondenz Einstein–Schlick: Zum Verhältnis der Physik zur Philosophie. (German) [The Einstein– Schlick correspondence: On the relationship of physics to philosophy]. Annals of Science, 43(5):475–488, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:1987:ENT

[Hen87] Klaus Hentschel. Einstein, Neokantianismus und Theorienholismus. (German) [Einstein, neokantianism, and holistic theory]. Kant-Studien: Philosophische Zeitschrift der Kant-Gesellschaft, 78(1–4):459–470, January 1987. ISSN 0022-8877 (print), 1613-

1134 (electronic). URL http://elib.uni-stuttgart.de/
opus/volltexte/2010/5111.

Hendry:1988:BRC

[Hen88a] J. Hendry. Book review: The Collected Papers of Albert Einstein, Vol 1 — the Early Years, 1879–1902 — J. Stachel. British Journal for the History of Science, 21(69):262, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Hendry:1988:BRJa

[Hen88b] John Hendry. Book review: John Stachel (ed.). The Collected Papers of Albert Einstein. Volume 1. The Early Years, 1879– 1902. Princeton: Princeton University Press, 1987. Pp. lxvi + 433. ISBN 0-691-08407-6. £35. English translation by Anna Beck, £14.10 (paperback), £6.25 (microfiche), available only to purchasers of main volume. British Journal for the History of Science, 21(2):262, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4026992.

Hentschel:1990:IFS

[Hen90] Klaus Hentschel. Interpretationen und Fehlinterpretationen der speziellen und der allgemeinen Relativitätstheorie durch Zeitgenossen Albert Einsteins. (German) [Interpretations and misinterpretations of the Special and the General Theory of Relativity by Albert Einstein’s contemporaries], volume 6 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1990. ISBN 3-7643-2438-4. xvii + 574 + cxi pp. LCCN QC173.52 .H46 1990. URL

http://www.gbv.de/dms/ilmenau/toc/025365711.PDF.

Henderson:1991:DBS

[Hen91] Janice Adrienne Henderson. On the distances between sun, moon, and Earth according to Ptolemy, Copernicus, and Reinhold, volume 1(30) of Studia Copernicana; Brill’s series. Brill, Leiden, The Netherlands, 1991. ISBN 90-04-09378-8. ISSN 0925-6806. xiv + 220 pp. LCCN QB991.C66 H46 1991.

Hentschel:1992:ETE

[Hen92a] Klaus Hentschel. Der Einstein-Turm: Erwin F. Freundlich und die Relativitätstheorie; Ansätze zu einer “dichten Beschreibung” von institutionellen, biographischen und theoriengeschichtlichen Aspekten. (German) [The Einstein Tower: Erwin F. friendly and the Theory of Relativity; Dense approaches to a description of institutional, biographical and theoretical-historical aspects]. Spektrum Akademischer Verlag, Heidelberg, Germany, 1992. ISBN 3-86025-025-6. 192 pp. LCCN QB36.F73 H46 1992. See also English translation [Hen97].

Hentschel:1992:EAT

[Hen92b] Klaus Hentschel. Einstein’s attitude towards experiments: Testing Relativity Theory 1907–1927. Studies in History and Philosophy of Science Part A, 23(4):593–624, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290014W.

Hentschel:1994:EFF

[Hen94] Klaus Hentschel. Erwin Finlay Freundlich and testing Einstein’s Theory of Relativity. Archive for History of Exact Sciences, 47(2):143–201, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=2&spage=143.

Hentschel:1996:PNS

[Hen96] Klaus Hentschel, editor. Physics and National Socialism: an anthology of primary sources, volume 18 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-5312-0 (Basel), 0-8176-5312-0 (Boston). ci + 406 + civ pp. LCCN Q127.G3 P48 1996.

Hentschel:1997:ETI

[Hen97] Klaus Hentschel. The Einstein Tower: an intertexture of dynamic construction, Relativity Theory, and astronomy. Writing science. Stanford University Press, Stanford, CA, USA, 1997. ISBN 0-8047-2824-0. xiv + 226 pp. LCCN QB462.65. URL

http://www.gbv.de/dms/goettingen/220404755.pdf. En-

glish translation by Ann M. Hentschel of the German original [Hen92a].

Henshall:1998:VHN

[Hen98] Philip Henshall. Vengeance: Hitler’s nuclear weapon: fact or fiction? Sutton Publishing Ltd., Stroud, Gloucestershire, UK, 1998. ISBN 0-7509-2051-3 (paperback). xii + 180 pp. LCCN UA710 .H46 1998.

Henshall:2000:NAG

[Hen00a] Philip Henshall. The Nuclear Axis: Germany, Japan and the atom bomb race, 1939–1945. Sutton Publishing Ltd., Stroud, Gloucestershire, UK, 2000. ISBN 0-7509-2293-1. viii + 230 pp. LCCN UG1282.A8 H45 2000.

Hentschel:2000:BRF

[Hen00b] Klaus Hentschel. Book review: Fritz Stern, Einstein’s German World, Princeton, Princeton University Press, 1999, ix + 335 pp. (ISBN 0-691-05939-X). Nuncius, 15(1):449–452, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00641.

Henriksen:2003:SHC

[Hen03] Jens H. Henriksen. Starling, his contemporaries and the Nobel Prize: one hundred years with hormones. Scandinavian journal of clinical and laboratory investigation. Supplement, 238:iii + 64, 2003.

Hentschel:2005:PHB

[Hen05] Ann M. Hentschel. Peripatetic highlights in Bern. Physics in Perspective (PIP), 7(1):107–129, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
004-0232-0.

Hentschel:2007:MAM

[Hen07] Klaus Hentschel. The mental aftermath: the mentality of German physicists 1945–1949. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920566-3 (hardcover). 205 pp. LCCN QC9.G3 H46 2007. URL http://www.

loc.gov/catdir/toc/fy0805/2007280005.html.

Hentschel:2009:RBG

[Hen09] Klaus Hentschel. Rezension: Einstein’s Generation: The Origins of the Relativity Revolution von Richard Staley. Berichte zur Wissenschaftsgeschichte, 32(3):304, September 2009. CO-DEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Henchman:2011:BRT

[Hen11a] A. Henchman. Book reviews: ThermoPoetics: Energy in Victorian Literature and Science, and: Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Victorian Studies, 54(1):142–145, Autumn 2011. CODEN ???? ISSN 0042-5222 (print), 1527-2052 (electronic).

Henchman:2011:PPL

[Hen11b] Anna Henchman. Pursuing power and light: Technology and physics from James Watt to Albert Einstein. Victorian Studies, 54(1):142–145, Fall 2011. ISSN 0042-5222 (print), 1527-2052 (electronic).

Hentschel:2011:BRH

[Hen11c] Klaus Hentschel. Book review: Helmut Rechenberg, Werner Heisenberg — Die Sprache der Atome. Leben und Wirken — Eine wissenschaftliche Biographie. Die ‘Fröhliche Wissenschaft’ (Jugend bis Nobelpreis). 2 vols. Heidelberg: Springer, 2010. Pp. xxi + 603. ISBN 978-3-540-69221-8. £168.00 (hardback). British Journal for the History of Science, 44(4):612–614, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heneberg:2013:SUA

[Hen13] Petr Heneberg. Supposedly uncited articles of Nobel laureates and Fields medalists can be prevalently attributed to the errors of omission and commission. Journal of the American Society for Information Science and Technology: JASIST, 64(3):448– 454, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hentschel:2016:BRM

[Hen16] Klaus Hentschel. Book review: Milena Wazeck; Geoffrey S. Koby. Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s. Isis, 107(1):193–195, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hentschel:2021:BRJ

[Hen21] Klaus Hentschel. Book review: Jost Lemmerich. Max von Laue-Furchtlos und treu: Eine Biographie des Nobelpreisträgers für Physik. Isis, 112(3):622–623, ???? 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Herneck:1933:AEL

[Her33] Friedrich Herneck. Albert Einstein. Ein Leben für Wahrheit, Menschlichkeit und Frieden. (German) [Albert Einstein: a life for truth, humanity, and peace]. Buchverlag Der Morgen, Berlin, Germany, 1933. ???? pp. LCCN QC16.E5 H4.

Hersey:1946:H

[Her46a] John Hersey. Hiroshima. Penguin, London, UK and New York, NY, USA, 1946. ix + 119 pp. LCCN D767.25.H6 H4 1946.

Hersey:1946:RLH

[Her46b] John Hersey. A reporter at large: Hiroshima. The New Yorker, ??(??):15–26, 28, 30–33, 36–43, 46–52, 54–58, 61–68, August 31,

1946. ISSN 0028-792X. URL http://archives.newyorker.
com/?i=1946-08-31#folio=015.

Herzfeld:1949:NSM

[Her49] Karl F. Herzfeld. Nodal surfaces in molecular wave functions. Reviews of Modern Physics, 21(3):527–530, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.527;   http://rmp.aps.org/

abstract/RMP/v21/i3/p527_1. Special issue in honor of Albert Einstein on his 70th birthday.

Herneck:1963:AEL

[Her63a] Friedrich Herneck. Albert Einstein. Ein Leben für Wahrheit, Menschlichkeit und Frieden. (German) [Albert Einstein: a life for truth, humanity, and peace]. Buchverlag Der Morgen, Berlin, Germany, 1963. ???? pp. LCCN QC16.E5 H4.

Herneck:1963:BAE

[Her63b] Friedrich Herneck. Zum Briefwechsel Albert Einsteins mit Ernst Mach. (German) [On Albert Einstein’s correspondence with Ernst Mach]. Forschungen und Fortschritte, 37(8):239–243, August 1963. CODEN FOFOAI. ISSN 0367-2794.

Hermann:1966:AEJ

[Her66a] Armin Hermann. Albert Einstein und Johannes Stark: Briefwechsel und Verhältnis der beiden Nobelpreisträger. (German) [Albert Einstein and Johannes Stark: Correspondence and relation of the two Nobel Prize Laureates]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 50(3):267–285, ???? 1966. CODEN SUARAH. ISSN 0039-4564. URL http:/

/www.jstor.org/stable/20775517.

Herneck:1966:BZE

[Her66b] Friedrich Herneck. Die Beziehungen zwischen Einstein und Mach, dokumentarisch dargestellt. (German) [The relationships between Einstein and Mach, depicted in a documentary]. Wissenschaftliche Zeitschrift der Friedrich-Schiller-Universität Jena, 15(??):1–14, ???? 1966.

Herneck:1967:AEL

[Her67] Friedrich Herneck. Albert Einstein. Ein Leben für Wahrheit, Menschlichkeit und Frieden. (German) [Albert Einstein: a life for truth, humanity, and peace]. Buchverlag Der Morgan, Berlin, Germany, third edition, 1967. 269 pp. LCCN QC16.E5 H4 1967.

Hermann:1971:GQT

[Her71] Armin Hermann. The Genesis of Quantum Theory (1899–1913). MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-08047-8. ix + 165 pp. LCCN QC174 .H4613 1971.

Herneck:1974:AE

[Her74] Friedrich Herneck. Albert Einstein. G. Teubner, Leipzig, Germany, 1974. ???? pp. LCCN QC16.E5 H397.

Herneck:1975:AEL

[Her75] Friedrich Herneck. Albert Einstein: ein Leben für Wahrheit, Menschlichkeit und Frieden. (German) [Albert Einstein: a life for truth, humanity, and peace], volume 14 of Biographien hervorragender Naturwissenschaftler, Techniker und Mediziner. BSB Teubner, Leipzig, Germany, second edition, 1975. 124 pp. LCCN QC16.E5 H397 1975. 5.00M.

Hermann:1976:WH

[Her76a] Armin Hermann. Werner Heisenberg, 1901–1976. Inter Nationes, Bonn-Bad Godesberg, West Germany, 1976. 146 pp. LCCN QC16.H35 H4813.

Hermann:1976:WHS

[Her76b] Armin Hermann. Werner Heisenberg in Selbstzeugnissen und Bilddokumenten. (German) [Werner Heisenberg: in self certifications and pictures], volume 240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, 1976. 152 pp. LCCN QC16.H35 .H47; QC16.H35 H47.

Herneck:1976:ESW

[Her76c] Friedrich Herneck. Einstein und sein Weltbild: Aufsätze und Vorträge. (German) [Einstein and his World View: Essays and Lectures]. Buchverlag Der Morgen, Berlin, Germany, 1976. 365 pp. LCCN QC16.E5 H42. 10.80M.

Hermann:1977:SNA

[Her77] Armin Hermann. Die Suche nach dem Absoluten: Max Planck und die Platonische Philosophie. (German) [The search for the absolute: Max Planck and Platonic philosophy]. ????, ????, 1977. ???? pp. LCCN ????

Hermann:1978:AEU

[Her78] Armin Hermann. Albert Einstein: Umsturz im Weltbild der Physik. (German) [Albert Einstein: upheaval in the world of physics]. ????, ????, 1978. 14–33 pp. LCCN ????

Hermann:1979:NPR

[Her79a] Armin Hermann. The new physics: the route into the atomic age: in memory of Albert Einstein, Max von Laue, Otto Hahn, Lise Meitner. Inter Nationes, Bonn-Bad Godesberg, West Germany, 1979. 175 pp. LCCN QC773 .H4813.

Herneck:1979:AE

[Her79b] Friedrich Herneck. Albert Einstein. Biographien hervorragender Naturwissenschaftler, Techniker und Mediziner; 14. Farrar, Strauss, and Giroux, New York, NY, USA, fourth edition, 1979. 116 pp. LCCN QC16.E5 H397 1979. 5.00M.

Herken:1980:WWA

[Her80] Gregg Herken. The winning weapon: the atomic bomb in the Cold War, 1945–1950. Knopf, New York, NY, USA, 1980. ISBN 0-394-50394-5. x + 425 pp. LCCN D843 .H438 1980.

Herken:1982:WWA

[Her82a] Gregg Herken. The winning weapon: the atomic bomb in the Cold War, 1945–1950. Vintage Books, New York, NY, USA, 1982. ISBN 0-394-75160-4. x + 425 pp. LCCN D843 .H438 1982.

Hershberg:1982:EVM

[Her82b] James Hershberg. Ends versus means: James B. Conant and American atomic policy, 1939–47. Thesis (a.b., honors), Harvard University, Cambridge, MA, USA, 1982.

Hermann:1984:WHS

[Her84] Armin Hermann. Werner Heisenberg mit Selbstzeugnissen und Bilddokumenten. (German) [Werner Heisenberg: in self certifications and pictures], volume 240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, 1984. ISBN 3-499-50240-2. 152 pp. LCCN QC16.H37 H47 1984.

Hersey:1985:H

[Her85a] John Hersey. Hiroshima. Alfred A. Knopf, New York, NY, USA, 1985. ISBN 0-394-54844-2. 196 pp. LCCN D767.25.H6

H4 1985. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
random045/85040346.html.

Hersey:1985:RLH

[Her85b] John Hersey. A reporter at large: Hiroshima: The aftermath. The New Yorker, ??(??):37–63, July 15, 1985. ISSN 0028-

792X. URL http://www.newyorker.com/magazine/1985/07/
15/hiroshima-the-aftermath.

Herken:1988:WWA

[Her88] Gregg Herken. The winning weapon: the atomic bomb in the Cold War, 1945–1950: with a new preface. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-02286-0 (paperback). xvi + 425 pp. LCCN D843 .H438 1988. US$12.50.

URL http://www.jstor.org/stable/j.ctt7ztvzm.

Hersey:1989:H

[Her89a] John Hersey. Hiroshima. Vintage Books, New York, NY, USA, 1989. ISBN 0-679-72103-7. 152 pp. LCCN D767.25.H6

H4 1989. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
random045/85040346.html.

Hershberg:1989:JBC

[Her89b] James Gordon Hershberg. James B. Conant, nuclear weapons, and the Cold War, 1945–1950. Thesis (Ph.D.), Department of Philosophy, Tufts University, Medford, MA, USA, 1989. Three volumes.

Hershberg:1993:JBC

[Her93] James G. (James Gordon) Hershberg. James B. Conant: Harvard to Hiroshima and the Making of the Nuclear Age. Alfred A. Knopf, New York, NY, USA, 1993. ISBN 0-394-57966-6. ix + 948 pp. LCCN CT275.C757 H46 1993. US$35.00 (US$46.00 Can.).

Hermann:1994:EWS

[Her94a] Armin Hermann. Einstein der Weltweise und sein Jahrhundert: eine Biographie. (German) [Einstein the worldly wise man and his century: a biography], volume 2303 of Piper. Piper, München, Germany, 1994. ISBN 3-492-22303-6. 635 pp. LCCN ????

Hersey:1994:H

[Her94b] John Hersey. Hiroshima. G. K. Hall, Thorndike, Me., 1994. ISBN 0-8161-5978-5. 204 pp. LCCN D767.25.H6

H4 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
random045/85040346.html.

Herken:2000:CCP

[Her00] Gregg Herken. Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, revised and expanded edition, 2000. ISBN 0-8047-3966-8, 0-8047-3770-3 (paperback). xv + 358 pp. LCCN Q127.U6 H394

2000. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
cam021/00026546.html;   http://www.loc.gov/catdir/toc/
cam027/00026546.html.

Herbert:2001:VRR

[Her01a] Christopher Herbert. Victorian relativity: radical thought and scientific discovery. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-32732-9, 0-226-32733-7 (paperback). xv + 302 pp. LCCN

BD221 .H47 2001. URL http://www.loc.gov/catdir/bios/
uchi051/00012177.html;   http://www.loc.gov/catdir/
description/uchi051/00012177.html;   http://www.loc.
gov/catdir/toc/uchi051/00012177.html.

Hermann:2001:WHS

[Her01b] Armin Hermann. Werner Heisenberg: mit Selbstzeugnissen und Bilddokumenten dargestellt. (German) [Werner Heisenberg: in self certifications and pictures], volume 50240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, seventh edition, 2001. ISBN 3-499-50240-2. 151 pp. LCCN QC16.H35 .H47.

Herken:2002:BBT

[Her02] Gregg Herken. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Henry Holt and Company, Incorporated, New York, NY, USA, 2002. ISBN 0-8050-6588-1 (hardcover). xiv + 448 pp. LCCN QC16.O62 H47 2002. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:

//alsos.wlu.edu/information.aspx?id=1264;   http://
www.loc.gov/catdir/bios/hol051/2002017219.html;   http:
//www.loc.gov/catdir/description/hol021/2002017219.
html.

Herken:2003:CBR

[Her03] Gregg Herken. Comment on book review of Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken [Am. J. Phys. 71(4), 411–415 (2003)]. American Journal of Physics, 71(7):647–648, July 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v71/i7/p647_s1;   http://link.aip.

org/link/?AJP/71/647/1. See [Ber03]. Includes a response by reviewer Jeremy Bernstein.

Hermann:2004:EWS

[Her04] Armin Hermann. Einstein der Weltweise und sein Jahrhundert: eine Biographie. (German) [Einstein the worldly wise man and his century: a biography], volume 4036 of Piper. Piper, München, Germany, 2004. ISBN 3-492-24036-4. 635 pp. LCCN ????

Herschbach:2006:ES

[Her06] Dudley Herschbach. Einstein as a student. Report, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, Massachusetts, USA, June 27, 2006. URL http://

www.chem.harvard.edu/herschbach/Einstein_Student.pdf.

Herweck:2011:AEV

[Her11] Don Herweck. Albert Einstein: een vriendelijk genie. (Dutch) [Albert Einstein: a gentle genius]. Schoolsupport, Noordhorn, The Netherlands, 2011. ISBN 90-8664-284-5. 31 pp. LCCN ???? Translation to Dutch by Roel Gille of English original.

Herring:2017:BRP

[Her17] Emily Herring. Book review: The Physicist and the Philosopher: Einstein, Bergson and the Debate that Changed our Understanding of Time. Annals of Science, 74(1):86–87, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hersch:2021:BRR

[Her21] Matthew Hersch. Book review: Robert Serber; introduction by Richard Rhodes. The Los Alamos Primer: The First Lectures on How to Build an Atomic Bomb. Isis, 112(1):209–210, ???? 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hessenbruch:2004:BRL

[Hes04] Arne Hessenbruch. Book review: Lillian Hoddeson and Vicki Daitch, True Genius: The Life and Science of John Bardeen, the Only Winner of Two Nobel Prizes in Physics. Washington, DC: Joseph Henry Press, 2002. Pp. xi + 467. ISBN 0-309-08408-3. £20.95, $27.95 (hardback). British Journal for the History of Science, 37(2):230–231, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028358.

Handchen:2012:MSN

[HES^{+}12] Vitus Händchen, Tobias Eberle, Sebastian Steinlechner, Aiko Samblowski, Torsten Franz, Reinhard F. Werner, Roman Schnabel, et al. Materials science and nanotechnology — news & features: Bose, Einstein and chips. Nature Photonics, 6(9):598– 601, August 19, 2012. CODEN NPAHBY. ISSN 1749-4885 (print), 1749-4893 (electronic). URL http://www.nature.com/

materials/news/news/011004/portal/6855-1.html. Origi-

nal published 4 October 2001.

Hess:2015:EWR

[Hes15a] Karl Hess. Einstein was right! Pan Stanford Publishing, Singapore, 2015. ISBN 981-4463-69-8 (hardcover), 981-4463-70-1 (e-book). xii + 205 pp. LCCN QC173.6 .H47 2015.

Hesse:2015:WES

[Hes15b] Christian Hesse. Was Einstein seinem Papagei erzählte Die besten Witze aus der Wissenschaft. (German) [What Einstein told his parrot: The best jokes from science], volume 6084 of C. H. Beck Paperback. C. H. Beck, München, Germany, third edition, 2015. ISBN 3-406-67917-X. 233 pp. LCCN ????

Hillman:2015:MWS

[HEWW15] Bruce J. Hillman, Birgit Ertl-Wagner, and Bernd C. Wagner. The man who stalked Einstein: how Nazi scientist Philipp Lenard changed the course of history. Lyons Press, Guilford, CT, USA, 2015. ISBN 1-4930-1001-8 (hardcover). ???? pp. LCCN QC16.L4 H55 2015.

Hey:1999:FCE

[Hey99] Anthony J. G. Hey. Feynman and computation: exploring the limits of computers. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0057-3. xxiii + 438 pp. LCCN QA76 .F46 1999, QC52 .F49 199. US$50.00.

Hevly:1998:AW

[HF98] Bruce William Hevly and John M. Findlay, editors. The Atomic West, volume 7 of The Emil and Kathleen Sick lecture-book series in Western history and biography. University of Washington Press, Seattle, WA, USA, 1998. ISBN 0-295-97749-3 (hardcover), 0-295-97716-7 (paperback). x + 286 pp. LCCN QC773.3.U5 A85 1998.

Hsu:2005:YGA

[HF05] J. P. (Jong-Ping) Hsu and Dana Fine, editors. 100 Years of Gravity and Accelerated Frames: The Deepest Insights of Einstein and Yang–Mills, volume 9 of Advanced series on theoretical physical science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-256-335-0 (hardcover), 981-270-340-3 (e-book). xxxvii + 623 pp. LCCN QC178 .A15 2005.

Heinze:2022:NOP

[HF22] Thomas Heinze and Joel Emanuel Fuchs. National and organizational patterns of Nobel laureate careers in physiology/ medicine, physics, and chemistry. Scientometrics, 127(12):7273– 7288, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04250-0.

Henley:2006:IAP

[HFN06] Ernest M. Henley, E. Norval Fortson, and Warren G. Nagourney, editors. An isolated atomic particle at rest in free space: a tribute to Hans Dehmelt, Nobel Laureate. Alpha Science International, Oxford, UK, 2006. ISBN 1-84265-161-7. v + 151 pp. LCCN QC793.28 .I86 2006.

Heinemann-Gruder:1992:SAG

[HG92] Andreas Heinemann-Grüder. Die sowjetische Atombombe. (German) [The Soviet Atomic Bomb]. Verlag Westfälisches Dampfboot, Münster, Germany, 1992. ISBN 3-924550-65-4. 168 pp. LCCN 0.2hei a0165 a0175 a2000 b7900.

Hentschel:2005:AEH

[HG05a] Ann Hentschel and Gerd Grasshoff. Albert Einstein: “Those Happy Bernese Years”. Stämpfli, Bern, Switzerland, 2005. ISBN 3-7272-1177-6. 180 pp. LCCN QC16.E5 H38413 2005.

Hon:2005:HEM

[HG05b] Giora Hon and Bernard R. Goldstein. How Einstein made asymmetry disappear: Symmetry and Relativity in 1905. Archive for History of Exact Sciences, 59(5):437–544, July 2005. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=59&issue=5&spage=437.

Hon:2006:AVE

[HG06a] Giora Hon and Bernard R. Goldstein. Adding velocities without exceeding the velocity of light: Wilhelm Wien’s algorithm (1904) and Albert Einstein’s light postulate (1905). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(2):89–113, April 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2006:SAE

[HG06b] Giora Hon and Bernard R. Goldstein. Symmetry and asymmetry in electrodynamics from Rowland to Einstein. Studies in History and Philosophy of Modern Physics, 37(4):635–660, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000487.

Harmon:2007:SLG

[HG07] Joseph E. Harmon and Alan G. Gross, editors. The scientific literature: a guided tour. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-31655-6 (hardcover), 0-226-31656-4 (paperback). xxiv + 327 pp. LCCN Q225.5 .S35 2007. URL http://catdir.loc.

gov/catdir/enhancements/fy0701/2006016547-t.html;
http://catdir.loc.gov/catdir/enhancements/fy0707/2006016547-
b.html;   http://catdir.loc.gov/catdir/enhancements/
fy0707/2006016547-d.html.

Hawkins:1987:TLW

[HGS87] Helen S. Hawkins, G. Allen Greb, and Gertrud Weiss Szilard. Toward a livable world: Leo Szilard and the crusade for nuclear arms control, volume 3 of Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19260-8, 0-262-08162-8. lxxiv + 499 pp. LCCN QC3 .S97 vol. 3.

Habicht:1910:EPN

[HH10] Conrad Habicht and Paul Habicht. Elektrostatischer Potentialmultiplikator nach A. Einstein. (German) [Electrostatic potential multiplier according to A. Einstein]. Physikalische Zeitschrift, 11(??):532–535, ???? 1910. CODEN PHZTAO. ISSN 0369-982X. This article describes joint work between the Habicht brothers and Albert Einstein on the design (by Einstein; see [Ein08a]) and construction (by the Habichts) of a commercial device for accurate measurements of small voltages and currents. See [Maa07] for the history and impact of that effort. Only three such original devices survive in museums, but a modern reconstruction and evaluation has been carried out [SU06, SWU09].

Hewlett:1989:APW

[HH89] Richard G. Hewlett and Jack M. Holl. Atoms for peace and war, 1953–1961: Eisenhower and the Atomic Energy Commission, volume 3 of California studies in the history of science; A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1989. ISBN 0-520-06018-0. xxix + 696 pp. LCCN HD9698.U52 H55 1989 vol. 3; QC792.7.

Hetherington:1996:HCG

[HH96] Norriss S. Hetherington and Edwin Powell Hubble. Hubble’s cosmology: a guided study of selected texts, volume 11 of Pachart history of astronomy series. Pachart Publishing House, Tucson, AZ, USA, 1996. ISBN 0-88126-287-0. xx + 218 pp. LCCN QB855 .H47 1996.

Hargittai:2000:CSC

[HH00] István Hargittai and Magdolna Hargittai, editors. Candid science: conversations with famous chemists. Imperial College Press, London, UK, 2000. ISBN 1-86094-151-6, 1-86094-228-8 (paperback). xii + 516 pp. LCCN QD21 .H294 2000.

Hargittai:2002:CSI

[HH02] István Hargittai and Magdolna Hargittai, editors. Candid science II: conversations with famous biomedical scientists. Imperial College Press, London, UK, 2002. ISBN 1-86094-280-6, 1-86094-288-1 (paperback). x + 604 pp. LCCN QP511.7 .H37 2002.

Hargittai:2003:CSI

[HH03] István Hargittai and Magdolna Hargittai, editors. Candid science III: more conversations with famous chemists. Imperial College Press, London, UK, 2003. ISBN 1-86094-337-3 (paperback), 1-86094-336-5 (hardcover). x + 507 pp. LCCN QD21 .H295 2003.

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711 pp. LCCN QC15 .H295 2004. URL

http://www.worldscibooks.com/physics/p304.html.

Hargittai:2004:EPW

[HH04b] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1–19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). LCCN QC15 .H295 2004. URL http://www.

worldscibooks.com/physics/p304.html.

Hargittai:2005:CSV

[HH05] Balazs Hargittai and István Hargittai. Candid science V: conversations with famous scientists. Imperial College Press, London, UK, 2005. ISBN 1-86094-505-8 , 1-86094-506-6 (paperback). xvi + 695 pp. LCCN Q141 .H264 2005. URL http://

www.loc.gov/catdir/toc/fy0602/2004062538.html.

Hargittai:2006:CSV

[HH06] István Hargittai and Magdolna Hargittai. Candid science VI: more conversations with famous scientists. Imperial College Press, London, UK, 2006. ISBN 1-86094-693-3, 1-86094-694-1 (paperback). x + 885 pp. LCCN Q141 .H264 2006.

Hentschel:2014:CFW

[HH14] Klaus Hentschel and Dieter Hoffmann, editors. Carl Friedrich von Weizsäcker: Physik — Philosophie — Friedensforschung: Leopoldina-Symposium, vom 20. bis 22. Juni 2012 in Halle (Saale): mit 110 Abbildungen und 1 Tabelle, volume 63 of Acta Historica Leopoldina. Deutsche Akademie der Naturforscher Leopoldina — Nationale Akademie der Wissenschafte, Halle (Saale), Germany, 2014. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Hanson-Harding:2016:AEI

[HH16] Alexandra Hanson-Harding. Albert Einstein: ingenious physicist and father of relativity. Britannica beginner bios. Britannica Educational Publishing in association with Rosen Educational Services, New York, NY, USA, 2016. ISBN 1-68048-255-6 (hardcover), 1-5081-0060-8 (paperback). ???? pp. LCCN QC16.E5 H36 2016.

Hansel:2001:BEC

[HHHR01] W. Hänsel, P. Hommelhoff, T. W. Hänsch, and J. Reichel. Bose–Einstein condensation on a microelectronic chip. Nature, 413(6855):498–501, October 4, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v413/n6855/
full/413498a0.html;   http://www.nature.com/physics/
highlights/6855-1.html.

Hendricks:2006:IMP

[HHJ^{+}06] Vincent F. Hendricks, David Jalal Hyder, Klaus Frovin Jørgensen, Jesper Lützen, and Stig Andur Pedersen, editors. Interactions: mathematics, physics and philosophy, 1860–1930, volume 251 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346. xiv + 326 pp. LCCN Q175 .B69 v.251 2006;

Q175 .I51377 2006. URL http://www.loc.gov/catdir/
enhancements/fy0825/2007425756-d.html;   http://www.
loc.gov/catdir/toc/fy0713/2007425756.html.

Hoddeson:1993:CAT

[HHMW93] Lillian Hoddeson, Paul W. Henriksen, Roger A. Meade, and Catherine L. Westfall, editors. Critical assembly: a technical history of Los Alamos during the Oppenheimer years, 1943–1945. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44132-3 (hardcover). xv + 509 pp. LCCN QC773.3.U5 C75 1993. URL http://www.loc.

gov/catdir/description/cam025/92036611.html;   http:
//www.loc.gov/catdir/toc/cam029/92036611.html. With

contributions by Gordon Baym, Richard Hewlett, Alison Kerr, Robert Penneman, Leslie Redman, and Robert Seidel.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hawking:1979:GRE

[HI79] S. W. (Stephen W.) Hawking and W. Israel, editors. General Relativity: an Einstein centenary survey. Cambridge University Press, Cambridge, UK, 1979. ISBN 0-521-22285-0. xviii + 919 pp. LCCN QC173.6 .G46.

Hawking:1987:THY

[HI87] S. W. (Stephen W.) Hawking and W. (Werner) Israel, editors. Three hundred years of gravitation. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34312-7, 0-521-37976-8 (paperback). xiii + 690 pp. LCCN QC178 .T47 1987.

Hiebert:1968:BRB

[Hie68] Erwin N. Hiebert. Book review: Letters on Wave Mechanics: Schrödinger, Planck, Einstein, Lorentz. K. Przibram, Ed. Translated from the German with an introduction by Martin J. Klein. Philosophical Library, New York, 1967. xx + 75 pp., illus. $6. Science, 160(3828):664–666, May 10, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/160/
3828/664.full.pdf.

Hiebert:1982:DPC

[Hie82] Erwin N. Hiebert. Developments in physical chemistry at the turn of the century. In Bernhard et al. [BCS82], pages 97–115. ISBN 0-08-027939-2. LCCN Q124.6 .N63

1981. US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500116.

Hiedemann:1984:BRE

[Hie84] Robert E. Hiedemann. Book review: Einstein and Modern French Drama: An Anthology by Kenneth S. White. Modern Drama, 27(3), Fall 1984. CODEN ???? ISSN 0026-7694 (print), 1712-5286 (electronic). URL https://muse.jhu.edu/

article/498615/summary.

Higgs:2014:NLE

[Hig14] Peter W. Higgs. Nobel lecture: Evading the Goldstone theorem. Reviews of Modern Physics, 86(3):851–??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.851.

Hildebrand:1939:KSV

[Hil39] Bengt Hildebrand. Kungl. Svenska Vetenskapsakademien: Förhistoria, Grundläggning och Första Organisation. (Swedish) [Royal Swedish Academy of Science: Prehistory, Foundation, and First Organization]. Kungel. Vetenskapsakademien, Stockholm, Sweden, 1939. xv + 885 pp. LCCN ????

Hilton:2003:EWL

[Hil03] Geraldine Hilton. Einstein’s wife: The life of Mileva Marić Einstein. Film produced by Melsa Films in conjunction with Oregon Public Broadcasting and Australian Broadcasting Corporation., 2003.

Hilbert:2007:FPFa

[Hil07a] David Hilbert. The foundations of physics (first communication). In Renn et al. [RSS^{+}07b], pages 989–1001. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_43. Originally published 20 November 2015.

Hilbert:2007:FPFb

[Hil07b] David Hilbert. The foundations of physics (first communication). In Renn et al. [RSS^{+}07b], pages 1003–1015. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_44. Originally published as “Die Grundlagen der Physik (Erste Mitteilung)”, Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Math.-phys. Klasse, (8) 395–407 (1916).

Hilbert:2007:FPS

[Hil07c] David Hilbert. The foundations of physics (second communication). In Renn et al. [RSS^{+}07b], pages 1017–1038. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_45. Originally published as “Die Grundlagen der Physik (Zweite Mitteilung)”, Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Math.-phys. Klasse, 53–76 (1917).

Hinshaw:1949:ESP

[Hin49] Virgil G. Hinshaw, Jr. Einstein’s social philosophy. In Schilpp [Sch49c], chapter 25, pages 647–661. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Hines:1998:FEB

[Hin98] Terence Hines. Further on Einstein’s brain. Experimental Neurology, 150(2):343–344, April 1998. CODEN EXNEAC. ISSN 0014-4886 (print), 1090-2430 (electronic).

Hirosige:1969:OLT

[Hir69] Tetu Hirosige. Origins of Lorentz’ Theory of Electrons and the concept of the electromagnetic field. Historical Studies in the Physical Sciences, 1(??):151–209, ???? 1969. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757298.

Hirondel:1974:SRG

[Hir74] D. Hirondel. Special Relativity in General Relativity? Journal of Mathematical Physics, 15(9):1471–1473, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hirosige:1976:EPM

[Hir76] Tetu Hirosige. The ether problem, the mechanistic worldview, and the origins of the Theory of Relativity. Historical Studies in the Physical Sciences, 7(??):3–82, ???? 1976. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757354.

Hirsch:1997:GOS

[Hir97] Henry R. Hirsch. Gerontological observations supporting Einstein and Mao. Perspectives in Biology and Medicine, 40(4): 562–563, 1997. CODEN PBMEA8. ISSN 0031-5982 (print), 1529-8795 (electronic).

Hiroiwa:1998:HWP

[Hir98] Chikahiro Hiroiwa. Hirochima witness for peace: testimony of A-bomb survivor Suzuko Numata. Soeisha/books sanseido, ????, Japan, 1998. ISBN 4-88142-208-1. 285 pp. LCCN ????

Harrison:2016:NTF

[HJ16] Adam P. Harrison and Dileepan Joseph. Numeric tensor framework: Exploiting and extending Einstein notation. Journal of Computational Science, 16:128–139, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1877750316300813.

Heilbron:1969:GBA

[HK69] John L. Heilbron and Thomas S. Kuhn. The genesis of the Bohr atom. Historical Studies in the Physical Sciences, 1(??): vi, 211–290, ???? 1969. CODEN HSPSAS. ISSN 0073-2672.

URL http://www.jstor.org/stable/27757291.

Hawking:1994:ETE

[HK94] Stephen William Hawking and Hainer Kober. Einsteins Traum: Expeditionen an die Grenzen der Raumzeit. Rowohlt Verlag GmbH, Reinbek bei Hamburg, Germany, 1994. ISBN 3-498-02919-3. 190 pp. LCCN ????

Hagen:2010:ADE

[HKS10] Margareth Hagen, Randi Koppen, and Margery Vibe Skagen, editors. The art of discovery: encounters in literature and science. Aarhus University Press, Århus, Denmark, 2010. ISBN 87-7934-501-8. 275 pp. LCCN PN55 .A78 2010.

Hoddeson:2008:FPF

[HKW08] Lillian Hoddeson, Adrienne W. Kolb, and Catherine Westfall. Fermilab: physics, the frontier, and megascience. University of Chicago Press, Chicago, IL, USA and London, UK, 2008. ISBN 0-226-34623-4 (cloth). xi + 497 pp. LCCN

QC789.2.U62 F474 2008. URL http://www.loc.gov/catdir/
toc/ecip0810/2008006254.html.

Huebener:2008:FD

[HL08a] R. P. (Rudolf Peter) Huebener and H. Lübbig. A focus of discoveries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-279-034-9. x + 185 pp. LCCN QC51.G62 P495 2008.

Huebener:2008:IIA

[HL08b] R. P. (Rudolf Peter) Huebener and H. Lübbig. The Imperial Institute and Albert Einstein. In A focus of discoveries [HL08a], chapter 9, pages 105–126. ISBN 981-279-034-9. LCCN QC51.G62 P495 2008.

Hong:2010:LXM

[HL10] Bin Hong and Xiaobo Liu. Liu Xiaobo mian mian guan: 2010 nian Nuobei’er he ping jiang de zhu. (Mandarin) [Dr. Liu Xiaobo: visions and perspectives as the winner of 2010 Nobel Peace Prize]. Po si chu ban gong si, Xianggang, China, 2010. ISBN 1-921815-05-1. 347 pp. LCCN PL2879.X53 Z75 2010.

Huebener:2012:FD

[HL12a] R. P. (Rudolf Peter) Huebener and H. Lübbig. A focus of discoveries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2012. ISBN 981-4390-49-6. x + 185 pp. LCCN ????

Huebener:2012:IIA

[HL12b] R. P. (Rudolf Peter) Huebener and H. Lübbig. The Imperial Institute and Albert Einstein. In A focus of discoveries [HL12a], chapter 9, pages 105–126. ISBN 981-4390-49-6. LCCN ????

Hansen:1991:LCT

[HLZJ91] Jean-Pierre Hansen, D. Levesque, and Jean Zinn-Justin, editors. Liquides, cristallisation et transition vitreuse = Liquids, freezing and glass transition: Les Houches, session LI, 3–28 juillet 1989. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-88927-2 (part I), 0-444-88928-0 (part II). LCCN QC138 .E36 1989.

Hirsch:1990:HBW

[HM90] Daniel Hirsch and William G. Mathews. The H-bomb: Who really gave away the secret? Bulletin of the Atomic Scientists, 46(1):22–30, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kno90, MHM90, BHM91, HM97].

Hirsch:1997:LHB

[HM97] Daniel Hirsch and William G. Mathews. Letter: H-bomb secrets. Bulletin of the Atomic Scientists, 53(4):3, 58, July/ August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90, KAS96] and comment [Han97].

Hawking:2005:BHT

[HM05a] S. W. (Stephen W.) Hawking and Leonard Mlodinow. A briefer history of time. Bantam Books, New York, NY, USA, 2005. ISBN 0-553-80436-7. xi + 162 pp. LCCN

QB981 .H3773 2005. URL http://www.loc.gov/catdir/
enhancements/fy0624/2005042949-b.htm;   http://www.
loc.gov/catdir/enhancements/fy0624/2005042949-d.htm;
 http://www.loc.gov/catdir/enhancements/fy0624/2005042949-
s.htm.

Hora:2005:ETL

[HM05b] Heinrich Hora and George H. (George Hunter) Miley, editors. Edward Teller lectures: lasers and inertial fusion energy. Imperial College Press, London, UK, 2005. ISBN 1-86094-468-X, 1-86094-727-1. xi + 365 pp. LCCN QC791.775

.L37 E39 2005. URL http://ebooks.worldscinet.com/ISBN/
9781860947278/9781860947278.shtml.

Hawking:2010:GD

[HM10] Stephen Hawking and Leonard Mlodinow. The Grand Design. Bantam Books, New York, NY, USA, 2010. ISBN 0-553-80537-1 (hardcover), 0-593-05829-1 (hardcover), 0-593-05830-5 (paperback), 0-553-84088-6 (paperback), 0-553-90707-7 (e-book). 198 pp. LCCN QC794.6.G7 H385 2010.

Hamsun:2011:NNK

[HM11] Marie Hamsun and Elmer T. Magnuson. The Nobel novelist Knut Hamsun during the Nazi occupation of Norway: the final chapter that was omitted from Marie Hamsun’s autobiography of their life together. Edwin Mellen Press, Lewiston, NY, USA, 2011. ISBN 0-7734-3944-7 (hardcover). v + 145 pp. LCCN PT8950 .H3 Z642513 2011; PT8950.H3 Z642513 2011. An authorized translation from the Norwegian and with an introduction.

Haunschild:2024:HCR

[HMW24] Robin Haunschild, Werner Marx, and Jürgen Weis. How can revivals of scientific publications be explained using bibliometric methods? A case study discovering booster papers for the 1985 Physics Nobel Prize paper. Scientometrics, 129(2):1079– 1095, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04906-z.

Hnizdo:1998:CME

[Hni98] V. Hnizdo. Common misrepresentation of the Einstein– Podolsky–Rosen argument. Foundations of Physics Letters, 11 (4):359–369, ???? 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hoag:2003:EPG

[Hoa03] Hannah Hoag. Einstein’s papers go online. Nature News, May 20, 2003. CODEN ???? ISSN ???? URL http://www.nature.

com/news/2003/030520/full/news030519-1.html.

Hoch:1983:RCE

[Hoc83] Paul K. Hoch. The reception of central European refugee physicists of the 1930s: U.S.S.R., U.K., U.S.A. Annals of Science, 40(3):217–246, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoch:1993:BRE

[Hoc93] Paul K. Hoch. Book review: Elisabeth Crawford, Nationalism and Internationalism in Science, 1880–1939: Four Studies of the Nobel Population. Cambridge: Cambridge University Press, 1992. Pp. xii + 157. ISBN 0-521-40386-3. £27.95, $44.95. British Journal for the History of Science, 26(1):121–122, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.jstor.org/stable/4027383.

Hoefer:1994:ESM

[Hoe94] Carl Hoefer. Einstein’s struggle for a Machian gravitation theory. Studies in History and Philosophy of Science Part A, 25(3): 287–335, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368194900566.

Hoeneveld:2018:BRM

[Hoe18] Friso Hoeneveld. Book review: Martijn van Calmthout. Sam Goudsmit: Zijn jacht op de atoombom van Hitler. Isis, 109(1): 207–208, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hoffmann:1949:IAE

[Hof49a] Banesh Hoffmann. The influence of Albert Einstein. Scientific American, 180(3):52–55, March 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v180/n3/
pdf/scientificamerican0349-52.pdf.

Hoffmann:1949:KNR

[Hof49b] Banesh Hoffmann. Kron’s non-Riemannian electrodynamics. Reviews of Modern Physics, 21(3):535–540, July 1949. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.21.535;   http://rmp.aps.org/abstract/RMP/

v21/i3/p535_1. Special issue in honor of Albert Einstein on his 70th birthday.

Hoffmann:1955:HWM

[Hof55] Heinrich Hoffmann. Hitler was my friend. Burke, London, UK, 1955. 256 pp. LCCN DD247.H5 H635 1955. Translated by R. H. Stevens.

Hoffmann:1979:EZ

[Hof79a] Banesh Hoffmann. Einstein and Zionism. In Aichelburg and Sexl [AS79], pages 169–175. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5

A66. DM48.00. URL http://link.springer.com/chapter/
10.1007/978-3-322-91080-6_12/.

Hoffmann:1979:EC

[Hof79b] Banesh Hoffmann. Einstein the catalyst. Physics Today, 32(3): 36–40, March 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995457.

Hoffmann:1982:RAE

[Hof82a] Banesh Hoffmann. Reminiscences [of Albert Einstein]. In Holton and Elkana [HE82], pages 401–404. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Hoffmann:1982:SEA

[Hof82b] Banesh Hoffmann. Some Einstein anomalies. In Holton and Elkana [HE82], pages 91–106. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Hoffman:1997:AEC

[Hof97a] A. Hoffman. Albert Einstein at Caltech. California History, 76 (4):109–10+, Winter 1997. ISSN 0162-2897.

Hoffmann:1997:EIR

[Hof97b] Banesh Hoffmann. Einsteins Ideen: das Relativitätsprinzip und seine historischen Wurzeln. (German) [Einstein’s Ideas: the Principle of Relativity and its historical roots]. Spektrum Akademischer Verlag, Heidelberg, Germany, 1997. ISBN 3-8274-0252-2. 198 pp. LCCN ???? Translation by Hajo Suhr from the American edition.

Hoffmann:1999:RR

[Hof99] Banesh Hoffmann. Relativity and its roots. Dover, New York, NY, USA, 1999. ISBN 0-486-40676-8 (paperback). 161 pp. LCCN QC173.52 .H63 1999. URL http://www.loc.gov/

catdir/description/dover032/98048651.html.

Hoffmann:2003:EK

[Hof03] Dieter Hoffmann. Einstein und die Kunst. (German) [Einstein and art]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 11(3):196, September 2003. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL

http://link.springer.com/article/10.1007/BF03004118.

Hoffmann:2006:EBS

[Hof06] Dieter Hoffmann. Einsteins Berlin: Auf den Spuren eines Genies. (German) [Einstein’s Berlin: in the footsteps of a genius]. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40596-8. viii + 224 pp. LCCN QC16.E5 H644 2006.

Hoffman:2008:QIE

[Hof08a] Jascha Hoffman. Q & A: Insight into Einstein. Nature, 453 (7198):987, June 18, 2008. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v453/n7198/full/453987a.html.

Hoffmann:2008:MPK

[Hof08b] Dieter Hoffmann. Max Planck: Kaum eine Sitzung die ich versäumt habe. Max Planck und die Physikalische Gesellschaft. (German) [Max Planck: Hardly a meeting that I had neglected. Max Planck and the Physical Society]. Physik Journal, 7(3):27–33, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.

pro-physik.de/details/articlePdf/1104315/issue.html;
 http://www.pro-physik.de/details/news/1121445/Kaum_
eine_Sitzung_die_ich_versaeumt_habe.html.

Hoffmann:2008:ZWP

[Hof08c] Dieter Hoffmann. Zwischen Wissenschaft und Politik: Max Planck und Albert Einstein:. Kollegen im Widerstreit. (German) [Between science and politics: Max Planck and Albert Einstein: Colleagues in the conflict]. Spektrum der Wissenschaft (German translation of Scientific American), 5 (??):32–39, May 25, 2008. CODEN SPEKDI. ISSN 0170-

2971. URL http://www.spektrum.de/magazin/max-planck-
und-albert-einstein-kollegen-im-widerstreit/947196.

Hoffman:2010:DHU

[Hof10] David E. (David Emanuel) Hoffman. The dead hand: the untold story of the Cold War arms race and its dangerous legacy. Anchor Books, New York, NY, USA, 2010. ISBN 0-307-38784-4 (paperback). viii + 577 pp. LCCN U264 .H645 2010.

Hoffmann:2013:EBF

[Hof13] Dieter Hoffmann. Einstein’s Berlin: in the footsteps of a genius. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1040-0 (paperback). xiii + 175 pp. LCCN QC16.E5 H64413 2013.

Hofer:2018:NLB

[Hof18] Prof. Dr. Hans-Georg Hofer. In Netzwerken von Laureaten. Beobachtungen und Überlegungen zur historischen Nobelpreisforschung. (German) [In networks of laureates. Observations and reflections on historical Nobel Prize research]. Berichte zur Wissenschaftsgeschichte, 41(1):98–103, March 2018. CO-DEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
bewi.201801870.

Hogan:1996:HHM

[Hog96] Michael J. Hogan, editor. Hiroshima in history and memory. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56206-6 (hardcover), 0-521-56682-7 (paperback). xiv + 238 pp. LCCN D767.25.H6 H6456 1996. URL http://catdir.loc.

gov/catdir/description/cam027/95045850.html;   http:
//catdir.loc.gov/catdir/toc/cam025/95045850.html.

Hohenberg:2013:BRA

[Hoh13] Pierre Hohenberg. Book review: From the Atomic Bomb to the Landau Institute: Autobiography. Top Non-Secret, Isaak M. Khalatnikov, (translated from Russian by Sergei Esipov and Yurii Morozov) Springer, 2012. $69.95 (214 pp.). ISBN 978-3-642-27560-9. Physics Today, 66(11):52, November 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hoiberg:2001:YNL

[Hoi01] Dale Hoiberg, editor. 100 years with Nobel laureates. Encyclopaedia Britannica (India) and I.K. International, New Delhi, India, 2001. ISBN 81-88237-00-0. xiii + 1098 pp. LCCN AS911.N9 A15 2001.

Holmes:1930:EEJ

[Hol30] John Haynes Holmes. Einstein, Eddington and Jeans: Is science vindicating religion?, volume 7 of Community pulpit. Ser. 1930– 31. The Community Church, New York, NY, USA, 1930. 22 pp. LCCN BL241 .H6.

Holton:1967:IEE

[Hol68a] Gerald Holton. Influences on Einstein’s early work in Relativity Theory. American Scholar, 37(1):59–79, Winter 1967–1968. ISSN 0003-0937 (print), 2162-2892 (electronic). URL http://

www.jstor.org/stable/41210232.

Holton:1968:MES

[Hol68b] Gerald Holton. Mach, Einstein, and the search for reality. Dædalus, 97(2):636–673, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www. jstor.org/stable/20023833. Reprinted in [Hol73, pp. 219– 259].

Holton:1969:ECE

[Hol69a] Gerald Holton. Einstein and the “crucial” experiment. American Journal of Physics, 37(10):968–982, October 1969. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Holton:1969:EMC

[Hol69b] Gerald Holton. Einstein, Michelson, and the “crucial” experiment. Isis, 60(2):133–197, Summer 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229907.

Holton:1970:MES

[Hol70] Gerald J. Holton. Mach, Einstein, and the search for reality. In Ernst Mach: Physicist and Philosopher [CS70], chapter 9, pages 165–199. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (ebook). ISSN 0068-0346. LCCN ???? URL https://link.

springer.com/chapter/10.1007/978-94-017-1462-4_9.

Holton:1973:TOS

[Hol73] Gerald James Holton. Thematic origins of scientific thought: Kepler to Einstein. Harvard University Press, Cambridge, MA, USA, 1973. ISBN 0-674-87745-4. 495 pp. LCCN Q125 .H722.

Holton:1978:SPM

[Hol78a] Gerald Holton. Subelectrons, presuppositions, and the Millikan– Ehrenhaft dispute. Historical Studies in the Physical Sciences, 9(??):161–224, ???? 1978. CODEN HSPSAS. ISSN 0073-2672.

URL http://www.jstor.org/stable/27757378.

Holton:1978:SIC

[Hol78b] Gerald James Holton. The scientific imagination: case studies. Cambridge University Press, Cambridge, UK, 1978. ISBN 0-521-21700-8, 0-521-29237-9 (paperback). xvi + 382 pp. LCCN Q175 .H775.

Holton:1979:EHP

[Hol79a] Gerald Holton. Einstein and his perception of order in the universe, volume 1980 of Gerhard Herzberg lecture. Carleton University, Ottawa, ON, Canada, 1979. 20 pp.

Holton:1979:EMC

[Hol79b] Gerald Holton. Einstein’s model for constructing a scientific theory. In Aichelburg and Sexl [AS79], pages 109–136. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_9/.

Holton:1980:ESP

[Hol80] Gerald Holton. Einstein’s scientific program: The formative years. In Woolf [Woo80b], pages 49–?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Holloway:1981:ENA

[Hol81a] David Holloway. Entering the nuclear arms race: The Soviet decision to build the atomic bomb, 1939–45. Social Studies of Science, 11(2):159–197, May 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.

jstor.org/stable/284865;   https://journals.sagepub.
com/doi/pdf/10.1177/030631278101100201.

Holton:1981:FAP

[Hol81b] G. Holton. The formation of the American physics community in the 1920s and the coming of Albert Einstein. Minerva [Washington, DC], 19(4):569–581, ???? 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Holton:1982:IES

[Hol82] Gerald Holton. Introduction: Einstein and the shaping of our imagination. In Holton and Elkana [HE82], pages vii–?? ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Holton:1984:MPU

[Hol84] Gerald Holton. The migration of physicists to the United States. Bulletin of the Atomic Scientists, 40(4):18–24, April 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holton:1986:ASB

[Hol86] Gerald James Holton. The advancement of science, and its burdens: the Jefferson Lecture and other essays. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25244-X, 0-521-27243-2 (paperback). xi + 351 pp. LCCN Q173

.H734 1986. URL http://www.loc.gov/catdir/description/
cam032/86009722.html.

Holton:1988:EMC

[Hol88a] Gerald James Holton. Einstein, Michelson and the “crucial” experiment. In Thematic origins of scientific thought: Kepler to Einstein [Hol88d], page ?? ISBN 0-674-87747-0, 0-674-87748-9 (paperback). LCCN Q125 .H722 1988.

Holton:1988:MES

[Hol88b] Gerald James Holton. Mach, Einstein, and the search for reality. In Thematic origins of scientific thought: Kepler to Einstein [Hol88d], page ?? ISBN 0-674-87747-0, 0-674-87748-9 (paperback). LCCN Q125 .H722 1988.

Holton:1988:OST

[Hol88c] Gerald James Holton. On the origins of the Special Theory of Relativity. In Thematic origins of scientific thought: Kepler to Einstein [Hol88d], page ?? ISBN 0-674-87747-0, 0-674-87748-9 (paperback). LCCN Q125 .H722 1988.

Holton:1988:TOS

[Hol88d] Gerald James Holton. Thematic origins of scientific thought: Kepler to Einstein. Harvard University Press, Cambridge, MA, USA, revised edition, 1988. ISBN 0-674-87747-0, 0-674-87748-9 (paperback). 499 pp. LCCN Q125 .H722 1988.

Holtzman:1988:NSC

[Hol88e] Jack M. Holtzman. A note on Schrödinger’s Cat and the unexpected hanging paradox. The British Journal for the Philosophy of Science, 39(3):397–401, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/39/3/397.
full.pdf+html.

Hollingdale:1989:MM

[Hol89a] Stuart H. Hollingdale. Makers of Mathematics. Penguin, London, UK and New York, NY, USA, 1989. ISBN 0-14-022732-6. xiii + 433 pp. LCCN QA21 .H74 1989.

Holton:1989:MME

[Hol89b] Gerald Holton. More on Mach and Einstein: Investigation of an important phase in the rise of logical positivism/empiricism. Methodology and Science, 22(2):67–81, ???? 1989.

Holton:1991:ESR

[Hol91a] Gerald Holton. Einstein in Spain: Relativity and the Recovery of Science by Thomas F. Glick (review). Technology and Culture, 32(1):179–182, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888824/pdf.

Holton:1991:SEC

[Hol91b] Gerald Holton. Spengler, Einstein and the controversy over the end of science. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 28(2):543–556, 1991. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Holton:1992:MME

[Hol92] Gerald Holton. More on Mach and Einstein. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92], chapter 14, pages 263–276. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN

C5.53; QC6.9. URL https://link.springer.com/chapter/
10.1007/978-94-011-2771-4_14.

Holloway:1994:SBS

[Hol94] David Holloway. Stalin and the bomb: the Soviet Union and atomic energy, 1939–1956. Yale University Press, New Haven, CT, USA, 1994. ISBN 0-300-06056-4. xvi + 464 pp. LCCN UA770 .H632 1994.

Holton:1995:EB

[Hol95a] Gerald Holton. Einstein and books. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 11, pages 273–279. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. LCCN ???? URL https://link.springer.

com/chapter/10.1007/978-94-011-0217-9_11.

Holton:1995:EHO

[Hol95b] Gerald James Holton. Einstein, history, and other passions, volume 16 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-333-7, 1-56396-417-1. ISSN 1057-4964. xiii + 312 pp. LCCN Q173 .H7342 1995.

Holdren:1996:PBP

[Hol96a] John P. Holdren. Peace-building in the post-Cold War world. Bulletin of the Atomic Scientists, 52(2):29–32, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holton:1996:EHO

[Hol96b] Gerald James Holton. Einstein, history, and other passions: the rebellion against science at the end of the twentieth century. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-40716-7. xii + 240 + 8 pp. LCCN Q173 .H7342 1996.

Holton:1998:ECR

[Hol98a] Gerald Holton. Einstein and the cultural roots of modern science. Dædalus, 127(1):1–44, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027475.

Holton:1998:ASB

[Hol98b] Gerald James Holton. The advancement of science, and its burdens: with a new introduction. Harvard University Press, Cambridge, MA, USA, 1998. ISBN 0-674-00530-9 (paperback). xlix + 352 pp. LCCN QK596.C8 Y46.

Holton:1998:SIN

[Hol98c] Gerald James Holton. The scientific imagination: with a new introduction. Harvard University Press, Cambridge, MA, USA, 1998. ISBN 0-674-79488-5. xlii + 382 pp. LCCN Q175 .H775 1998.

Holton:1999:RMS

[Hol99] Gerald Holton. R. A. Millikan’s struggle with the meaning of Planck’s constant. Physics in Perspective (PIP), 1(3):231–237, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160050020.

Holton:2000:EHO

[Hol00a] Gerald Holton. Einstein, history, and other passions: the rebellion against science at the end of the Twentieth Century.

Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00433-7. xii + 240 + 4 pp. LCCN ????

Holton:2000:WHA

[Hol00b] Gerald Holton. Werner Heisenberg and Albert Einstein. Physics Today, 53(7):38–42, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v53/i7/p38/s1.

Holbrow:2003:ACC

[Hol03a] Charles H. Holbrow. In appreciation: Charles C. Lauritsen: a reasonable man in an unreasonable world. Physics in Perspective (PIP), 5(4):419–472, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0178-7.

Holton:2003:ETP

[Hol03b] Gerald Holton. Einstein’s third paradise. Dædalus, 132(4):26– 34, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print),

1548-6192 (electronic). URL http://www.aip.org/history/
einstein/essay-Einsteins-Third-Paradise.pdf;   http:
//www.jstor.org/stable/20027878.

Holland:2005:WWE

[Hol05a] Peter Holland. What’s wrong with Einstein’s 1927 hiddenvariable interpretation of quantum mechanics? Foundations of Physics, 35(2):177–196, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/s10701-004-1940-7.

Holton:2005:BRC

[Hol05b] G. Holton. Book review: The collected papers of Albert Einstein, vol 9. The Berlin years: Correspondence, January 1919–April 1920. Nature, 433(7023):195–196, January 20, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Holton:2005:VVS

[Hol05c] Gerald James Holton. Victory and vexation in science: Einstein, Bohr, Heisenberg, and others. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01519-3. xi + 229 pp. LCCN Q180.A3 H65 2005.

Holton:2007:WES

[Hol07] Gerald Holton. The woman in Einstein’s shadow. In Renn and Gavroglou [RG07], chapter 17, pages 95–98. ISBN 1-4020-5420-3, 1-4020-5419-X. ISSN 0068-0346. LCCN Q125

.P77 2007eb. URL https://link.springer.com/chapter/10.
1007/1-4020-5420-3_17.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (ebook). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Homolova:2004:EOJ

[Hom04] Marie Homolová. Einstein očima Johanny z Čech. (Czech) [Einstein through the eyes of Johanna from the Czech Republic]. Lidové noviny, ??(??):20, May 15, 2004. See note about errors in this article [Gor20, note 6, page 330].

Home:2005:WSS

[Hom05] Roderic W. Home. William Sutherland and the ‘Sutherland– Einstein’ diffusion relation: theoretical physics in a colonial setting. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):125–138, November 2005. CODEN HISCDU. ISSN 0285-4821.

Home:2006:RAE

[Hom06] R. W. Home. The rush to accelerate: Early stages of nuclear physics research in Australia. Historical Studies in the Physical and Biological Sciences, 36(2):213–241, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hon:1989:FHV

[Hon89] Giora Hon. Franck and Hertz versus Townsend: a study of two types of experimental error. Historical Studies in the Physical and Biological Sciences, 20(1):79–106, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL

http://www.jstor.org/stable/27757636.

Hon:1996:DSD

[Hon96] Giora Hon. Disturbing, but not surprising: Did Gödel surprise Einstein with a rotating universe and time travel? Foundations of Physics, 26(4):501–521, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02071218. See comments

[You98].

Hon:2004:GEM

[Hon04] Giora Hon. Gödel, Einstein, Mach: Casting constraints on allembracing concepts. Foundations of Science, 9(1):25–64, March 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.
1023/B:FODA.0000014880.52117.f0.

Hooker:1970:CEP

[Hoo70] C. A. Hooker. Concerning Einstein’s, Podolsky’s, and Rosen’s objection to quantum theory. American Journal of Physics, 38 (7):851–857, 1970. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://scitation.aip.org/
content/aapt/journal/ajp/38/7/10.1119/1.1976483. See

comments [Sch71].

Hoover:1977:LEB

[Hoo77] P. Hoover. Letter to Einstein beginning Dear Albert’. Chicago Review, 28(3):156–157, ???? 1977. ISSN 0009-3696 (print), 2327-5804 (electronic).

Hook:1988:PHS

[HOO88] Sidney Hook, William L. O’Neill, and Roger O’Toole. Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer, volume 107 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 94-010-7793-2, 94-009-2873-4 (ebook). ISSN 0068-0346. 494 (est.) pp. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-009-2873-2.

Hooker:1991:PPI

[Hoo91] C. A. Hooker. Projection, physical intelligibility, objectivity and completeness: the divergent ideals of Bohr and Einstein. The British Journal for the Philosophy of Science, 42(4):491–511, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/42/4/491.full.pdf+html.

Hoppe:2007:RBN

[Hop07] Brigitte Hoppe. Rezension: Bittere Nobelpreise von Alfred Neubauer. Berichte zur Wissenschaftsgeschichte, 30(2):171–172, June 2007. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Horz:1979:AEP

[Hor79a] H. Horz. Albert Einstein and philosophy. Deutsche Zeitschrift für Philosophie, 27(2):149–161, ???? 1979. ISSN 0012-1045.

Horz:1979:AEB

[Hor79b] H. Horz. Albert Einstein in Berlin, 1913–1933. 1. description and documents. 2 Einstein bibliography of works available in GDR — German — C. Kirsten and H. Treder. Deutsche Zeitschrift für Philosophie, 27(2):241–244, ???? 1979. ISSN 0012-1045.

Horvitz:2000:QSW

[Hor00] Leslie Alan Horvitz. The quotable scientist: words of wisdom from Charles Darwin, Albert Einstein, Richard Feynman, Galileo, Marie Curie, and more. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-136063-8. x + 169 pp. LCCN Q173 .H739 2000. URL http://www.loc.

gov/catdir/description/mh021/00709492.html;   http:
//www.loc.gov/catdir/enhancements/fy0707/00709492-b.
html;   http://www.loc.gov/catdir/toc/mh021/00709492.
html.

Horn:2009:RWC

[Hor09] Karen Ilse Horn. Roads to wisdom, conversations with ten Nobel laureates in economics. Edward Elgar, Cheltenham, UK, 2009. ISBN 1-84844-670-5 (hardcover), 1-84844-921-6 (e-book). vii + 369 pp. LCCN HB76 .H67 2009.

Hossenfelder:2015:TEH

[Hos15] Sabine Hossenfelder. Thought experiments: Head trip. Scientific American, 313(3):46–49, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-46.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-46.pdf.

Hotakainen:2015:ETK

[Hot15] Markus Hotakainen. Einstein on täysi kaheli: [review of] Täydellinen teoria — sata vuotta neroutta ja taistelu yleisestä suhteellisuusteoriasta (A complete theory — one hundred years of genius and general relativity theory). (Finnish) [Einstein is a full fruitcake]. Tähdet ja avaruus [Stars and space], 45(1):52–??, 2015. ISSN 0355-9467.

Houston:1965:BRA

[Hou65] William V. Houston. Book review: Albert Einstein and the cosmic world order, by Cornelius Lanczos, Interscience Publishers, vi + 139 pp. The Physics Teacher, 3(5):231, ???? 1965. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Houston:2016:BWS

[Hou16] Louis M. Houston. A de Broglie wave solution to the Maxwell– Schrödinger–Einstein equations. Theor. Math. Appl., 6(3):1–13, 2016. ISSN 1792-9687 (print), 1792-9709 (electronic).

Howard:1921:PTY

[How21] Ralph Howard. Prices — today, yesterday and before the War, the Einstein Award, and more. Scientific American, 124(5):81–83, January 29, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v124/n5/
pdf/scientificamerican01291921-81.pdf.

Howey:1950:FGS

[How50] Walter Howey, editor. The faith of great scientists: a collection of “My faith” articles from the American weekly. Hearst Publishing Company, New York, NY, USA, 1950. 63 pp. LCCN BL240 .A643.

Howson:1976:MAP

[How76] Colin Howson, editor. Method and appraisal in the physical sciences: the critical background to modern science, 1800–1905. Cambridge University Press, Cambridge, UK, 1976. ISBN 0-521-21110-7. vii + 344 pp. LCCN Q175 .M5417. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0906/75044580-
d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0906/75044580-t.html;   http://www.gbv.de/dms/bowker/
toc/9780521211109.pdf.

Howard:1984:RCE

[How84] Don Howard. Realism and conventionalism in Einstein’s philosophy of science: The Einstein–Schlick correspondence. Philosophia Naturalis, 21(??):618–629, 1984.

Howard:1985:ELS

[How85] Don Howard. Einstein on locality and separability. Studies in History and Philosophy of Science Part A, 16(3):171–201, September 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368185900019.

Howard:1989:HSM

[How89] Don Howard. Holism, separability, and the metaphysical implications of the Bell experiments. In Cushing and Mc-Mullin [CM89], pages 224–253. ISBN 0-268-01578-3, 0-268-01579-1 (paperback). LCCN QC174.12 .P43 1989. US$39.95,

US$19.95. URL http://en.wikipedia.org/wiki/Bell%27s_

theorem. See [Bel64].

Howard:1990:ED

[How90a] Don Howard. Einstein and Duhem. Synthese, 83(??):363–384, 1990. ISSN 0039-7857(print) 1573-0964 (electronic).

Howard:1990:SKW

[How90b] Don Howard. ’Nicht sein kann was nicht sein darf,” or the prehistory of EPR, 1909–1935: Einstein’s early worries about the quantum mechanics of composite systems. In Miller [Mil90], pages 61–111. ISBN 0-306-43608-6. LCCN QC173.96 .N385 1989.

Howard:1992:EEC

[How92] Don Howard. Einstein and Eindeutigkeit [clarity]: a neglected theme in the philosophical background to general relativity. In Eisenstaedt and Kox [EK92], pages 154–243. ISBN 0-8176-3479-7 (Boston), 3-7643-3479-7 (Berlin). LCCN QC173.6 .I572 1988.

Howard:1993:LEH

[How93a] Don Howard. Out of the labyrinth? Einstein, Hertz, and the Göttingen answer to the hole argument. In Earman et al. [EJN93], pages 30–62. ISBN 3-7643-3624-2, 0-8176-3624-2. LCCN QC173.6 .A85 1993. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0846.01009.

Howard:1993:WER

[How93b] Don Howard. Was Einstein really a realist? Perspectives on Science, 1(2):204–251, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Howard:1994:KVW

[How94a] Don Howard. ‘A kind of vessel in which the struggle for eternal truth is played out’ — Albert Einstein and the role of personality in science. In Langdon and McGann [LM94], pages 111–138. LCCN BD225 .N38 1993. URL https:

//www3.nd.edu/~dhoward1/Albert%2520Einstein%2520and%
2520the%2520Role%2520of%2520Personality%2520in%2520Science.
pdf.

Howard:1994:EKO

[How94b] Don Howard. Einstein, Kant, and the origins of logical empiricism. In Salmon and Wolters [SW94], pages 45–105. ISBN 3-87940-477-1, 0-8229-3740-9 (USA). LCCN Q174 L64 1994; Gd.CAR.B5; M95.E09243.

Howard:1996:REM

[How96a] Don Howard. Relativity, Eindeutigkeit, and monomorphism: Rudolf Carnap and the development of the categoricity concept in formal semantics. In Giere and Richardson [GR96], pages 115–164. ISBN 0-8166-2834-3 (hardcover). LCCN Q1 .M665; B824.6 .O75 1996; Q175.A1 M64; Q175 .M48; Q175 .M64.

Howarth:1996:SHT

[How96b] Richard J. Howarth. Sources for a history of the ternary diagram. British Journal for the History of Science, 29(3):337–356, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4027737.

Howard:1997:AEB

[How97a] D. Howard. Albert Einstein. a biography — A. Fölsing. Science, 278(5341):1241–1242, November 14, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Howard:1997:PBV

[How97b] Don Howard. A peek behind the Veil of Maya: Einstein, Schopenhauer, and the historical background of the conception of space as a ground for the individuation of physical systems. In Earman and Norton [EN97], pages 87–150. ISBN 0-8229-3930-4 (clothbound). LCCN Q175.3 .C69 1996; Q175.3 .C69 1997.

Howard:1998:ADL

[How98] Don Howard. Astride the divided line: Platonism, empiricism, and Einstein’s epistemological opportunism. In Shanks [Sha98], pages 143–163. ISBN 90-420-0642-0. ISSN 0303-8157. LCCN QC6 .I33 1998.

Howard:1999:PCP

[How99a] Don Howard. Point coincidences and pointer coincidences: Einstein on invariant structure in spacetime theories. In Goenner et al. [GRRS99b], pages 463–500. ISBN 0-8176-4060-6 (hardcover), 3-7643-4060-6. LCCN QC173.6 .E97 1999.

Howe:1999:GE

[How99b] Michael J. A. Howe. Genius explained. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-64018-0 (hardcover), 0-521-64968-4 (paperback). ix + 221 pp. LCCN BF416.A1 H68 1999.

Howard:2004:EPS

[How04] Don A. Howard. Einstein’s philosophy of science. In Edward N. Zalta, editor, The Stanford Encyclopedia of Philosophy (Spring 2004 Edition). Stanford University Press, Stanford, CA, USA,

2004. URL http://plato.stanford.edu/archives/spr2004/
entries/einstein-philscience/.

Howard:2005:AEP

[How05a] D. A. Howard. Albert Einstein as a philosopher of science. Physics Today, 58(12):34–40, December 2005. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Howard:2005:EPS

[How05b] Don Howard. Einstein as a philosopher of science. World-Wide Web lecture slides., February 20, 2005. URL http://www.

nd.edu/~dhoward1/Einstein2.pdf. Presented at the meeting

of the American Association for the Advancement of Science, Washington, DC, February 20, 2005.

Howard:2014:EDT

[How14] Don Howard. Einstein and the development of twentiethcentury philosophy of science. In Janssen and Lehner [JL14b], chapter 11, pages 354–376. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014. URL

http://www.nd.edu/~dhoward1/Einstein's%20Philosophy%
20of%20Science-Cambridge%20Companion-Final%20V.pdf.

Hoyle:1946:SEH

[Hoy46] Fred Hoyle. The synthesis of the elements from hydrogen. Monthly Notices of the Royal Astronomical Society, 106(5):343– 383, ???? 1946. CODEN MNRAA4. ISSN 0035-8711 (print),

1365-2966 (electronic). URL http://adsabs.harvard.edu/
abs/1946MNRAS.106..343H.

Hoyle:1954:NRO

[Hoy54] Fred Hoyle. On nuclear reactions occurring in very hot stars. I. The synthesis of elements from carbon to nickel. Astrophysical Journal. Supplement Series, 1(??):121–146, September 1954. CODEN APJSA2. ISSN 0067-0049 (print), 1538-

4365 (electronic). URL http://labs.adsabs.harvard.edu/
ui/abs/1954ApJS....1..121H. Reprinted in the centennial

edition Astrophysical Journal 525, 571 (1999) with a modern commentary by David Arnett.

Hoyle:1977:EEC

[Hoy77] Fred Hoyle. Energy or extinction?: The case for nuclear energy. Heinemann, London, UK, 1977. ISBN 0-435-54430-6. vii + 81 pp. LCCN TK9153 .H69.

Hoyle:1979:EEC

[Hoy79] Fred Hoyle. Energy or extinction?: the case for nuclear energy. Open University set book. Heinemann Educational, London, UK, second edition, 1979. ISBN 0-435-54431-4 (paperback). xx + 80 pp. LCCN TK9153 .H69 1979.

Hoyle:1994:HWW

[Hoy94] Fred Hoyle. Home is where the wind blows: chapters from a cosmologist’s life. University Science Books, Mill Valley, CA, USA, 1994. ISBN 0-935702-27-X (hardcover). xi + 443 + 24 pp. LCCN QB460.72.H69 A3 1994.

Heisenberg:2002:WHG

[HR02] Werner Heisenberg and Helmut Rechenberg, editors. Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). (German) [Werner Heisenberg: Appraisal and test certificates for graduations and Habilitations (1929–1942)], volume 29 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS-Verl., Berlin, Germany, second edition, 2002. ISBN 3-928577-43-3. 266 pp. LCCN ???? EUR 27.50.

He:2013:GME

[HR13] Q. Y. He and M. D. Reid. Genuine multipartite Einstein– Podolsky–Rosen steering. Physical Review Letters, 111(25): 250403, December 2013. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.111.250403;   http:/
/www.scientificcomputing.com/news/2014/03/einsteins-
entanglement-produces-quantum-encryption;   http://
www.swinburne.edu.au/engineering/caous/news_and_events/
mulitipartite%20EPR%20steering%20paper.htm.

Hu:2017:NPW

[HR17] Xiaojun Hu and Ronald Rousseau. Nobel Prize winners 2016: Igniting or sparking foundational publications? Scientometrics, 110(2):1053–1063, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11192-
016-2205-x.

Hentschel:1990:KPR

[HRP90] Klaus Hentschel, Hans Reichenbach, and Joseph Petzoldt. Die Korrespondenz Petzoldt–Reichenbach: zur Entwicklung der “wissenschaftlichen Philosophie” in Berlin. (German) [The Petzoldt–Reichenbach correspondence: development “of scientific philosophy” in Berlin], volume 12 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. SIGMA, Berlin, Germany, 1990. ISBN 3-928068-03-2. 91 pp. LCCN Q175.3 .H46 1990.

Holmes:2003:RBR

[HRR03] Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors. Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1039-7 (hardcover), 0-306-48152-9 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xv + 325 pp. LCCN

Q225.5 .R475 2003. URL http://www.loc.gov/catdir/
enhancements/fy0814/2002041330-d.html;   http://www.
loc.gov/catdir/enhancements/fy0814/2002041330-t.html.

Hahn:1939:NEA

[HS39a] Otto Hahn and Fritz Strassmann. Nachweis der Entstehung aktiver Bariumisotope aus Uran und Thorium durch Neutronenbestrahlung; Nachweis weiterer aktiver Bruchstücke bei der Uranspaltung. (German) [Evidence of formation of active barium isotopes of uranium and thorium by neutron irradiation: further evidence of active fragments from the fission of uranium]. Die Naturwissenschaften, 27(6):89–95, February 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
w406757166152l83/.

Hahn:1939:ZUD

[HS39b] Otto Hahn and Fritz Strassmann. Über das Zerplatzen des Urankernes durch langsame Neutronen. (German) [On the bursting of the uranium nucleus by slow neutrons], volume 12 of Abhandlungen der Preussischen Akademie der Wissenschaften. Jahrg. 1939. Mathematisch-naturwissenschaftliche Klasse. Akademie der Wissenschaften, in Kommission bei W. de Gruyter u. Co., Berlin, Germany, 1939. 20 pp. LCCN AS182 .B335 1939, nr. 12.

Hahn:1939:NVB

[HS39c] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Die Naturwissenschaften, 27(1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87–91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hahn:1944:CAS

[HS44] Otto Hahn and Fritz Strassmann. Die chemische Abscheidung der bei der Spaltung des Urans entstehenden Elemente und Atomarten (Allgemeiner Teil). (German) [The deposition of the chemical in the fission of uranium: resulting elements and atom types (general part)], volume 12 of Abhandlungen der Preussischen Akademie der Wissenschaften. Jahrg. 1944. Mathematisch-naturwissenschaftliche Klasse. Verlag der Akademie der Wissenschaften, in Kommission bei W. de Gruyter, Berlin, Germany, 1944. 14 pp. LCCN AS182 .B335 1944, Nr. 12.

Hittmair:1951:SGT

[HS51] O. Hittmair and Erwin Schrödinger. Studies in the generalized theory of gravitation II: the velocity of light, volume A8 of Communications of the Dublin Institute for Advanced Studies. Institiúid Árd-Léinn Bhaile Átha Cliath (Dublin Institute for Advanced Studies), Dublin, Ireland, 1951. 15 pp. LCCN QC407 .H56.

Heilbron:1989:LHL

[HS89a] John L. Heilbron and Robert W. Seidel. Lawrence and his laboratory: a history of the Lawrence Berkeley Laboratory. California studies in the history of science. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1989. ISBN 0-520-06426-7 (vol. 1). xxvii + 586 pp. LCCN QC789.2.U62 L384 1989. URL

http://ark.cdlib.org/ark:/13030/ft5s200764;   http:/
/publishing.cdlib.org/ucpressebooks/view?docId=ft5s200764;
 http://www.loc.gov/catdir/bios/ucal051/89004820.html;
 http://www.loc.gov/catdir/description/ucal041/89004820.
html.

Howard:1989:EHG

[HS89b] Don Howard and John J. Stachel, editors. Einstein and the History of General Relativity: Based on the Proceedings of the 1986 Osgood Hill Conference, North Andover, Massachusetts, 8–11 May 1986, volume 1 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1989. ISBN 0-8176-3392-8, 3-7643-3392-8. LCCN QC173.6 .E38 1989.

URL http://www.springer.com/us/book/9780817633929.

Howard:2000:EFY

[HS00] Don Howard and John J. Stachel, editors. Einstein: the formative years, 1879–1909, volume 8 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 0-8176-4030-4, 3-7643-4030-4. viii + 258 pp. LCCN QC16.E5 E524 2000. URL http://www.loc.gov/

catdir/enhancements/fy0812/98006996-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0812/98006996-t.
html.

Harvey:2005:SP

[HS05] Alex Harvey and Engelbert Schucking. A small puzzle from 1905. Physics Today, 58(3):34–36, March 2005. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). In his first Relativity paper, Einstein made one erroneous prediction. Although it should have been withdrawn when he generalized the theory to include gravity, the original error has received surprisingly little attention.

Harrison:2006:EBM

[HS06] Suzanne S. Harrison and Patrick H. Sullivan. Einstein in the boardroom: moving beyond intellectual capital to I-stuff. Wiley, New York, NY, USA, 2006. ISBN 0-471-70332-X (hardcover). xii + 226 pp. LCCN HD53 .H376 2006. URL http://www.loc.

gov/catdir/enhancements/fy0623/2005031932-d.html;
http://www.loc.gov/catdir/enhancements/fy0740/2005031932-
b.html;   http://www.loc.gov/catdir/toc/ecip063/2005031932.
html.

Harrigan:2010:EIE

[HS10] Nicholas Harrigan and Robert W. Spekkens. Einstein, incompleteness, and the epistemic view of quantum states. Foundations of Physics, 40(2):125–157, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
009-9347-0.

Hansson:2014:GSM

[HS14] Nils Hansson and Udo Schagen. ,,In Stockholm hatte man offenbar irgendwelche Gegenbewegung” — Ferdinand Sauerbruch (1875–1951) und der Nobelpreis. (German) [“In Stockholm there was apparently any backlash” — Ferdinand Sauerbruch (1875– 1951) and the Nobel Prize]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 22(3):133–161, September 2014. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-

9144 (electronic). URL http://link.springer.com/article/
10.1007/s00048-014-0114-8.

Heisenberg:1934:MAE

[HSD34a] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Die moderne Atomtheorie: die bei der entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Modern Atomic Theory: the 1933 Nobel Prize in Physics lecture in Stockholm]. Verlag von S. Hirzel, Leipzig, Germany, 1934. 2 + 45 + 1 pp. LCCN QC173 .H38.

Heisenberg:1934:MQM

[HSD34b] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Modern Quantum Mechanics: Three Nobel Reports. Technico-Theoretical State Press, Leningrad and Moscow, USSR, 1934. ???? pp.

Hsu:2000:ERB

[Hsu00] J. P. (Jong-Ping) Hsu. Einstein’s Relativity and beyond: new symmetry approaches, volume 7 of Advanced series on theoretical physical science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-3888-6. xxi + 418 pp. LCCN QC173.65 .H78 2000.

Heilbron:1981:LHL

[HSW81] J. L. Heilbron, Robert W. Seidel, and Bruce R. Wheaton. Lawrence and his laboratory: A historian’s view of the Lawrence years. Lawrence Berkeley Laboratory Research Review Magazine, 6(3):??, Fall 1981. ISBN 0-918102-09-X (pa-

perback). URL http://www2.lbl.gov/Science-Articles/
Research-Review/Magazine/1981/.

Heim:2009:KWS

[HSW09] Susanne Heim, Carola Sachse, and Mark Walker, editors. The Kaiser Wilhelm Society under National Socialism. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-87906-X (hardcover). xxiv + 477 pp. LCCN

Q49.K14174 H45 2009. URL http://www.loc.gov/catdir/
enhancements/fy0914/2008054455-b.htm;   http://www.
loc.gov/catdir/enhancements/fy0914/2008054455-d.htm;

 http://www.loc.gov/catdir/enhancements/fy0914/2008054455-
t.htm.

Hazen:2010:GIS

[HT10] Robert M. Hazen and James Trefil, editors. Great ideas of science: a reader in the classic literature of science. Cognella, San Diego, CA, USA, 2010. 256 pp. LCCN Q161.2 .G74 2010.

Hawkins:1983:MDH

[HTS83] David Hawkins, Edith C. Truslow, and Ralph Carlisle Smith, editors. Manhattan District history, Project Y, the Los Alamos story, volume 2 of History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-08-0. xxvi + 506 + 9 pp. LCCN QC773.3.U5 M25 1983.

Hewitt:1988:URS

[HTS^{+}88] J. N. Hewitt, E. L. Turner, D. P. Schneider, B. F. Burke, G. I. Langston, C. R. Lawrence, et al. Unusual radio source MG1131+0456: a possible Einstein ring. Nature, 333(6173): 537–540, June 9, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v333/n6173/pdf/333537a0.pdf.

Hu:2004:OCE

[Hu04] Danian Hu. Organized criticism of Einstein and Relativity in China, 1949–1989. Historical Studies in the Physical and Biological Sciences, 34(2):311–338, March 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hu:2005:CAE

[Hu05] Danian Hu. China and Albert Einstein: the reception of the physicist and his theory in China 1917–1979. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01538-X. xiv + 257 pp. LCCN QC16.E5 H79 2005. URL http://www.

loc.gov/catdir/toc/fy053/2004059690.html.

Hubble:1929:RBD

[Hub29] Edwin Hubble. A relation between distance and radial velocity among extra-galactic nebulae. Proceedings of the National Academy of Sciences of the United States of America, 15(3): 168–173, March 1929. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://hubblesite.org/

;   http://www.jstor.org/stable/85225.

Hubisz:2013:MBR

[Hub13] John L. Hubisz. MicroReviews by the Book Review Editor: Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics: Fulvio Melia. The Physics Teacher, 51(5):319, May 2013. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Hughes:1954:ERB

[Hug54] A. J. Hughes. The Einstein relation between relative viscosity and volume concentration of suspensions of spheres. Nature, 173 (4414):1089–1090, June 5, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v173/n4414/pdf/1731089a0.pdf.

Hughes:1989:SIQ

[Hug89] R. I. G. Hughes. The Structure and Interpretation of Quantum Mechanics. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-84391-6 (hardcover), 0-674-84392-4 (paperback). xiii + 369 pp. LCCN QC174.12 .H82 1989.

Hughes:1993:EII

[Hug93] Thomas P. Hughes. Einstein, inventors, and invention. Science in Context, 6(1):25–42, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Huggett:1999:SZE

[Hug99] Nick Huggett. Space from Zeno to Einstein: classic readings with a contemporary commentary. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-08271-3 (hardcover), 0-262-58169-8 (paperback). xi + 274 pp. LCCN QC6 .S6625 1997.

Hughes:2004:DBR

[Hug04] J. Hughes. Deconstructing the bomb: recent perspectives on nuclear history. British Journal for the History of Science, 37(4):455–464, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/4028151;   http://www.jstor.
org/stable/4028643.

Hulet:2001:BEC

[Hul01] Randall G. Hulet. Bose–Einstein condensation: Getting excited about helium. Nature, 411(6836):425–428, May 24, 2001. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v411/n6836/
full/411425a0.html.

Humphreys:1912:WPR

[Hum12] W. J. Humphreys. What is the Principle of Relativity? Scientific American, 106(23):525–526, June 8, 1912. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v106/n23/pdf/scientificamerican06081912-525.
pdf.

Humason:1936:ARV

[Hum36] M. L. (Milton L.) Humason. The apparent radial velocities of 100 extra-galactic nebulae. Astrophysical Journal, 83(10):10– 23, January 1936. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Humbert:1951:AON

[Hum51] Pierre Humbert. Analyses d’ouvrages: Nobel, Savants et découvertes, (Collection ^{≪}Les Savants et le Monde ^{≫}) par Louis de Broglie; André George. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):369–370, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905033.

Humiston:2013:D

[Hum13] Terisa Humiston. Downwinders. Master of Fine Arts, Pacific University, Forest Grove, OR 97116, USA, 2013. 104 pp.

Hunt:1991:M

[Hun91] Bruce J. Hunt. The Maxwellians. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2641-3. x + 266 pp. LCCN QC670 .H84 1991.

Hunziker:2001:AEM

[Hun01] Herbert Hunziker. Albert Einstein: Maturitätsprüfung in Mathematik 1896. (German) [Albert Einstein: matriculation in mathematics 1896]. Elemente der Mathematik, 56(2):45– 54, ???? 2001. ISSN 0013-6018 (print), 1420-8962 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000170050087.

Hunner:2007:IAG

[Hun07] Jon Hunner. Inventing Los Alamos: The Growth of an Atomic Community. University of Oklahoma Press, Norman, OK, USA, 2007. ISBN 0-8061-3891-2. LCCN ????

Hunt:2010:PPL

[Hun10] Bruce J. Hunt. Pursuing power and light: technology and physics from James Watt to Albert Einstein. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9358-5 (hardcover), 0-8018-9359-3 (paperback). 182 pp. LCCN T173.8 .H92 2010.

Hunziker:2015:PTA

[Hun15] Herbert Hunziker. The physical tourist: Albert Einstein’s magic mountain: An Aarau education. Physics in Perspective (PIP), 17(1):55–69, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0153-5.

Huntelmann:2018:PEN

[Hün18] Dr. Axel Cäsar Hüntelmann. Paul Ehrlich und der Nobelpreis. Die Konstruktion wissenschaftlicher Exzellenz. (German) [Paul Ehrlich and the Nobel Prize. The construction of scientific excellence]. Berichte zur Wissenschaftsgeschichte, 41(1):47–72, March 2018. CODEN BEWID8. ISSN 0170-6233 (print), 1522-

2365 (electronic). URL https://onlinelibrary.wiley.com/
doi/abs/10.1002/bewi.201801867.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hussain:2015:BRM

[Hus15] Rawaa Mahmoud Hussain. Book reviews: Michel Janssen and Christoph Lehner (eds.), The Cambridge Companion to Einstein. New York: Cambridge University Press, 2014. Pp. xvi + 562. ISBN 978-0-521-82834-5. £65.00 (hardback). British Journal for the History of Science, 48(2):374–375, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Holland:1988:QPS

[HV88] P. R. Holland and J. P. Vigier. The quantum potential and signalling in the Einstein–Podolsky–Rosen experiment. Foundations of Physics, 18(7):741–750, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00734154.

Hermann:1979:WPW

[HvMW79] Armin Hermann, Karl von Meyenn, and Victor F. (Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band I: 1919–1929. (German) [Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume I: 1919–1929], volume 2 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-08962-4, 3-540-08962-4. xlvii + 577 pp. LCCN QC16.P37 A34; QC16.P37

W64. URL http://www.springer.com/physics/book/978-3-

540-08962-9. Edited by Armin Hermann, K. von Meyenn and Victor F. Weisskopf, With a foreword by Victor F. Weisskopf, With an introduction by Armin Hermann.

Heisenberg:1949:PAG

[HW49] Werner Heisenberg and Wilhelm Westphal. Die Physik der Atomkerne. (German) [Physics of the atomic nucleus]. Friedrich Vieweg und Sohn, Braunschweig, Germany, third extended edition, 1949. viii + 192 pp. LCCN ???? English translation in [Hei53b, Hei53c].

Hachiya:1955:HDJ

[HW55] Michihiko Hachiya and Warner Wells, editors. Hiroshima diary; the journal of a Japanese physician, August 6–September 30, 1945. University of North Carolina Press, Chapel Hill, NC, USA, 1955. 238 pp. LCCN D767.25.H6 H3.

Herneck:1978:EPH

[HW78] Friedrich Herneck and Herta Waldow. Einstein privat: Herta W. erinnert sich an die Jahre 1927 bis 1933. (German) [Einstein privately: Herta W. remembers the years 1927 to 1933]. Der Morgen, Berlin, Germany, 1978. 169 + 8 pp. LCCN QC16.E5 H412. 7.50M.

Hey:1996:EM

[HW96] Anthony J. G. Hey and Patrick Walters. Einstein’s mirror. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43504-8 (hardcover), 0-521-43532-3 (paperback). xii + 291 pp. LCCN QC173.55 .H49 1996. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0968.83002.

Hoffmann:2004:GPS

[HW04] Dieter Hoffmann and Mark Walker. The German Physical Society under National Socialism. Physics Today, 57(12):52– 58, December 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/57/12/
10.1063/1.1878335.

Home:2007:ESQ

[HW07a] Dipankar Home and Andrew Whitaker. Einstein’s Struggles with Quantum Theory: a Reappraisal. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71520-7. 300 pp. LCCN QC16.E5 2007.

Home:2007:PBE

[HW07b] Dipankar Home and Andrew Whitaker. The philosophical background: Einstein and Mach. In Einstein’s Struggles with Quantum Theory: a Reappraisal [HW07a], page 300. ISBN 0-387-71520-7. LCCN QC16.E5 2007.

Hodge:2008:NFV

[HW08] Nathan Hodge and Sharon Weinberger. A nuclear family vacation: travels in the world of atomic weaponry. Bloomsbury USA, New York, NY, USA, 2008. ISBN 1-59691-378-9. viii + 324 pp.

LCCN U264 .H635 2008. URL http://www.loc.gov/catdir/
enhancements/fy0809/2008002013-b.html;   http://www.
loc.gov/catdir/enhancements/fy0809/2008002013-d.html.

Hoffmann:2012:GPS

[HW12] Dieter Hoffmann and Mark Walker, editors. The German Physical Society in the Third Reich: physicists between autonomy and accommodation. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-00684-8 (hardcover). xxiii + 458 pp. LCCN Q49 .G225 2012.

Haubold:1986:JIT

[HY86] Hans Joachim Haubold and Eiichi Yasui. Jun Ishiwaras Text über Albert Einsteins Gastvortrag an der Universität zu Kyoto am 14. Dezember 1922. (German) [Jun Ishiwara’s text of Albert Einstein’s guest lecture at Kyoto University on 14 December 1922]. Archive for History of Exact Sciences, 36(3):271–279, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=3&spage=271. Translated from

the Japanese by Hans Joachim Haubold and Eiichi Yasui.

Higuchi:2015:ITO

[HY15] Toshihiro Higuchi and Masakatsu Yamazaki. Introduction: Transnational origins of the Russell–Einstein Manifesto and the radiological dimensions of the nuclear arms race. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):1–7, August 2015. CODEN HISCDU. ISSN 0285-4821.

Hymans:2012:ANA

[Hym12] Jacques E. C. Hymans. Achieving nuclear ambitions: scientists, politicians and proliferation. Cambridge University Press, Cambridge, UK, 2012. ISBN 0-521-76700-8 (hardcover), 0-521-13225-8 (paperback). xii + 315 pp. LCCN JZ5665 .H95

2012. URL http://assets.cambridge.org/97805217/67002/
cover/9780521767002.jpg.

Hsu:2001:LPI

[HZ01] J. P. (Jong-Ping) Hsu and Yuanzhong Zhang, editors. Lorentz and Poincaré invariance: 100 years of Relativity, volume 8 of Advanced series on theoretical physical science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2001. ISBN 981-02-4721-4. xxxi + 583 pp. LCCN

QC174.17.S9 H78 2001. URL http://www.loc.gov/catdir/
toc/fy031/2002280428.html.

Ianniello:1993:ENB

[Ian93] Maria Grazia Ianniello. Elastic Nachwirkung, Brownian motion and the tide against determinism: 1835–1920. Historical Studies in the Physical and Biological Sciences, 24(1):41–100, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Inouye:1998:OFR

[IAS^{+}98] S. Inouye, M. R. Andrews, J. Stenger, H.-J. Miesner, D. M. Stamper-Kurn, W. Ketterle, et al. Observation of Feshbach resonances in a Bose–Einstein condensate. Nature, 392(6672): 151–154, March 12, 1998. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v392/n6672/full/392151a0.html.

Ibuse:1979:BR

[Ibu79] Masuji Ibuse. Black rain. Kodansha International, Tokyo, Japan, 1979. ISBN 0-87011-364-X, 4-7700-0695-0 (Japan). 300 pp. LCCN PL830.B8 K813 1979.

Iglewicz:2007:EFP

[Igl07] Boris Iglewicz. Einstein’s first published paper. The American Statistician, 61(4):339–342, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ihde:1974:BRBb

[Ihd74] Aaron J. Ihde. Book reviews: Les Prix Nobel en 1971 by Institutions Nobel; Nobel Lectures, Chemistry, 1963–1970 by Nobel Foundation; Nobel Lectures, Physics, 1963–1970 by Nobel Foundation. Isis, 65(4):544, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229366.

Ilic:2003:HHM

[IK03] Mirjana Ilić and Andreas Kleinert. “Herzallerliebstes helenchen”. Mileva Einsteins briefe an Helene Savić. (German) [“Dearest little Helen”. Mileva Einstein’s letters to Helene Savić]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 11(1):29–33, March 2003. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL

http://link.springer.com/article/10.1007/BF02908584.

Illy:1979:AEP

[Ill79] József Illy. Albert Einstein in Prague. Isis, 70(1):76–84, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230879.

Illy:1989:EEE

[Ill89a] J. Illy. Einstein and the Eötvös experiment — a letter from Albert Einstein to Willy Wien. Annals of Science, 46(4):417– 422, July 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Illy:1989:ETL

[Ill89b] József Illy. Einstein teaches Lorentz, Lorentz teaches Einstein: their collaboration in General Relativity, 1913–1920. Archive for History of Exact Sciences, 39(3):247–289, September 1989. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=39&issue=3&spage=247.

Illy:1989:EEV

[Ill89c] József Illy. Einstein und der Eötvös-Versuch: Ein Brief Albert Einsteins an Willy Wien. (German) [Einstein and the Eötvös experiment: a letter between Albert Einstein and Willy Wien]. Annals of Science, 46(4):417–422, ???? 1989. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00033798900200311.

Illy:2006:AMA

[Ill06] József Illy. Albert meets America: how journalists treated genius during Einstein’s 1921 travel. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8457-8. xvii + 345 pp. LCCN QC16.E5 A68 2006. URL http://catdir.loc.

gov/catdir/enhancements/fy0664/2006005266-b.html;
http://catdir.loc.gov/catdir/enhancements/fy0664/2006005266-
d.html;   http://catdir.loc.gov/catdir/toc/ecip069/
2006005266.html.

Illy:2012:PEE

[Ill12] József Illy. The practical Einstein: experiments, patents, inventions. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0457-5 (hardcover), 1-4214-0533-4 (ebook). xi + 202 pp. LCCN QC16.E5 I45 2012.

Illy:2019:EG

[Ill19] József Illy. Einstein’s gyros. Physics in Perspective (PIP), 21(4): 274–295, December 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-019-00248-7.pdf. See

correction [Ill20].

Illy:2020:CEG

[Ill20] József Illy. Correction to: Einstein’s gyros. Physics in Perspective (PIP), 22(1):52, March 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-020-00250-4.

pdf. See [Ill19].

Impey:2019:EML

[Imp19] Chris Impey. Einstein’s Monsters: the Life and Times of Black Holes. W. W. Norton & Co., New York, NY, USA, 2019. ISBN 1-324-00093-7 (hardcover), 1-324-00094-5 (e-book). xix + 295 pp. LCCN QB843.B55 I47 2019.

Infeld:1931:INE

[Inf31] Leopold Infeld. Über eine Interpretation der neuen Einsteinschen Weltgeometrie auf dem Boden der klassischen Mechanik. (German) [On an interpretation of the new Einstein’s world geometry on the ground of classical mechanics]. Physikalische Zeitschrift, 32:110–112, 1931. CODEN PHZTAO. ISSN 0369-982X.

Infeld:1933:NDN

[Inf33] Leopold Infeld. Nowe drogi nauki: kwanty i materja (Polish) [New pathways of science: quanta and matter], volume 2 of Z Dziedziny Nauki i Techniki. nakł. Mathesis Polskiej, Warszawa, Poland, 1933. x + 284 pp. LCCN ????

Infeld:1934:WMS

[Inf34] Leopold Infeld. The world of modern science: matter and quanta. G. P. Putnam’s Sons, New York, NY, USA, 1934. 287 pp. LCCN ???? Translation by Louis Infield of [Inf33]. Introduction by Albert Einstein.

Infeld:1941:QESa

[Inf41a] Leopold Infeld. Quest: the evolution of a scientist. Doubleday, Doran, and Co., New York, NY, USA, 1941. iv + 342 pp. LCCN QC16.I6 A3 1941.

Infeld:1941:QESb

[Inf41b] Leopold Infeld. Quest: the evolution of a scientist. Scientific Book Club, London, UK, 1941. 288 pp. LCCN ????

Infeld:1942:QA

[Inf42] Leopold Infeld. Quest: an autobiography. Readers Union, London, UK, 1942. viii + 272 pp. LCCN QC16.I6 A3 1942.

Infeld:1947:PE

[Inf47] Leopold Infeld. Portrait: Einstein. American Scholar, 16(3): 337–341, 1947. ISSN 0003-0937 (print), 2162-2892 (electronic).

URL http://www.jstor.org/stable/41204928.

Infeld:1949:GRS

[Inf49] Leopold Infeld. On the structure of our universe. In Schilpp [Sch49c], chapter 18, pages 477–499. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Infeld:1950:AEH

[Inf50a] Leopold Infeld. Albert Einstein: his work and its influence on our world. Scribner, New York, NY, USA, revised edition, 1950. v + 132 pp. LCCN ????

Infeld:1950:AES

[Inf50b] Leopold Infeld. Albert Einstein speaks his mind: Out of my Later Years, by Albert Einstein. New York Times, ??(??):146, May 28, 1950. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/111436294/.

Infeld:1950:LNE

[Inf50c] Leopold Infeld. Letter: The new Einstein theory and the equations of motion. Nature, 166(4234):1075, December 23, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1950:NET

[Inf50d] Leopold Infeld. The new Einstein theory and the equations of motion. Nature, 166(4234):1075, December 23, 1950. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v166/n4234/
pdf/1661075a0.pdf.

Infeld:1950:ENT

[Inf50e] Leopold Infeld. On Einstein’s new theory. American Scholar, 19 (4):423–433, Autumn 1950. ISSN 0003-0937 (print), 2162-2892

(electronic). URL http://www.jstor.org/stable/41206693.

Infeld:1951:NET

[Inf51a] Leopold Infeld. The new Einstein theory and the equations of motion. Acta Physica Polonica, 10:284–293, 1951. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1951:ENT

[Inf51b] Leopold Infeld. On Einstein’s new theory. In Annual report of the Board of Regents of the Smithsonian Institution, pages 189–197. Smithsonian Institution, Washington, DC, USA, 1951.

LCCN QC 6 .I54 1952. URL https://babel.hathitrust.org/
cgi/pt?id=uc1.31210022604829. Reprint of [Inf50e].

Infeld:1953:AES

[Inf53a] Leopold Infeld. Albert Einstein: sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein: his work and his influence on our world]. Schönbrunn-Verlag, Wien, Austria, 1953. 173 pp. LCCN ????

Infeld:1953:RVF

[Inf53b] Leopold Infeld. [rebuttal to V. A. Fock’s view of Albert Einstein]. Mysl Filosoficzna, ??(??):??, ???? 1953. Cited in [Inf78b, page 80].

Infeld:1954:DAE

[Inf54] Leopold Infeld. [In defense of Albert Einstein]. Mysl Filosoficzna, ??(??):??, ???? 1954. Cited in [Inf78b, page 85].

Infeld:1955:AE

[Inf55a] Leopold Infeld. [Albert Einstein]. Twórczość, ??(??):??, ???? 1955.

Infeld:1955:DAE

[Inf55b] Leopold Infeld. [Death of Albert Einstein]. L’Humanité, ?? (??):??, ???? 1955. Cited in [Inf78b, page 85], and reprinted in several Polish newspapers. Later enlarged and published in [Inf55a].

Infeld:1955:GRG

[Inf55c] Leopold Infeld. Die Geschichte der Relativitätstheorie. (German) [The history of the Theory of Relativity]. Die Naturwissenschaften, 42(15):431–436, August 1, 1955. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Infeld:1955:OKD

[Inf55d] Leopold Infeld. Od Kopernika do Einsteina. Polish [From Copernicus to Einstein]. Kosmos, Series B, 1(??):209–226, 1955.

Infeld:1956:AEJ

[Inf56a] Leopold Infeld. Albert Einstein, jego dzieło i rola w nauce. (Polish) [Albert Einstein his work and its influence on our world]. Państwowe wydawnictwo naukowe, Warszawa, Poland, 1956. 192 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1956:AES

[Inf56b] Leopold Infeld. Albert Einstein. Sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein. His work and his influence on our world]. Akademie-Verlag, Berlin, West Germany, 1956. 174 pp. LCCN ????

Infeld:1956:MVO

[Inf56c] Leopold Infeld. Moi Vospominaniya ob Einshteine. (Russian) [My memories of Einstein]. Uspekhi Fizicheskikh Nauk, 59(1): 135–184, 1956. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL http://ufn.ru/ru/articles/
1956/5/i/.

Infeld:1956:MWE

[Inf56d] Leopold Infeld. Moje wspomnienia o Einsteinie. (Polish) [My Memories of Einstein]. Iskry, Warszawa, Poland, 1956. 148 + 3 pp. LCCN ????

Infeld:1956:EMG

[Inf56e] Leopold Infeld. On equations of motion in general relativity theory. Helv. Phys. Acta Supp., 4(??):206–209, ???? 1956.

Infeld:1957:AESb

[Inf57a] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, 1957. 181 + 11 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1957:AESa

[Inf57b] Leopold Infeld. Albert Einstein: sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein: his work and his influence on our world]. Akademie-Verlag, Berlin, West Germany, 1957. 173 pp. LCCN ????

Infeld:1958:AEJ

[Inf58] Leopold Infeld. Albert Einstein : jeho dílo a vliv na ná sv t. (Czech) [Albert Einstein: His work and influence on our world], volume 74 of Universita vojáka. Naše vojsko, Praha, Czechoslovakia, 1958. 117 + 4 pp. LCCN ???? Translation to Czech by Lenka Teigová-Stachová of the Polish original.

Infeld:1959:MVN

[Inf59] Leopold Infeld. Mé vzpomínky na Einstein. (Czech) [My Memories of Einstein], volume 171 of Knihovna eskoslovenské Spolenosti pro íení politických a vdeckých znalostí. Orbis, Praha, Czechoslovakia, 1959. 79 pp. LCCN ???? Translation to Czech by Bojana Fiedlerová of the Polish original.

Infeld:1960:AES

[Inf60] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, third edition, 1960. 187 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1961:AES

[Inf61a] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, third edition, 1961. 181 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1961:AEH

[Inf61b] Leopold Infeld. Albert Einstein: his work and its influence on our world. Scribner, New York, NY, USA, revised edition, 1961. v + 134 pp. LCCN ????

Infeld:1961:ESO

[Inf61c] Leopold Infeld. Einstein, su obra y su influencia en nuestro mundo. (Spanish) [Einstein, his work and his influence in our world]. Lautaro, Buenos Aires, Argentina, 1961. 174 pp. LCCN ????

Infeld:1963:AEU

[Inf63] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world], volume 23 of Piccola biblioteca Einaudi. Giulio Einaud, Torino, Italia, fourth edition, 1963. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1968:AEU

[Inf68] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world]. Giulio Einaud, Torino, Italia, sixth edition, 1968. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1969:LEK

[Inf69] Leopold Infeld. Leben mit Einstein : Kontur einer Erinnerung. (German) [Living with Einstein: Contours of a memory]. Europa-Verlag, Wien, Austria, 1969. 223 pp. LCCN ???? Translation to German of ???? by Walter Hacker.

Infeld:1973:CE

[Inf73] Leopold Infeld. From Copernicus to Einstein. In Barbara Bieńkowska, editor, The Scientific World of Copernicus on the occasion of the 500th anniversary of his birth 1473–1973, pages 66–83. D. Reidel Publishing Company, Dordrecht, The Netherlands / Boston, MA, USA, 1973. ISBN 94-010-2618-1. LCCN

QB36.C8. URL http://adsabs.harvard.edu/abs/1973swc..

conf...66I. Foreword by Zdeněk Kopal.

Infeld:1974:AEU

[Inf74] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world]. Giulio Einaud, Torino, Italia, 1974. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1978:O

[Inf78a] Leopold Infeld. Oppenheimer. In Why I left Canada: reflections on science and politics [Inf78b], pages 160–180. ISBN 0-7735-0272-6. LCCN QC16.I6 A3213. Infeld makes a strong rebuttal in support of Albert Einstein against a speech made by J. Robert Oppenheimer on 14 December 1965. He also records harsh criticism by Philip Morrison of Oppenheimer.

Infeld:1978:WLC

[Inf78b] Leopold Infeld. Why I left Canada: reflections on science and politics. McGill-Queen’s University Press, Montréal, Québec, Canada, 1978. ISBN 0-7735-0272-6. xii + 212 pp. LCCN QC16.I6 A3213. Translations from Polish to English by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild (1921–1977).

Infeld:1979:AAN

[Inf79a] Leopold Infeld. Albert Ajnštajn: njegova dela i njihov uticaj na naš svet. (Serbian) [Albert Einstein: his work and its impact on our world], volume 7 of Biblioteka Zanimljiva nauk. Nolit, Beograd, Yugoslavia, 1979. 131 pp. LCCN ???? Translation to Serbian by Branislav Lalović of [Inf50a].

Infeld:1979:AE

[Inf79b] Leopold Infeld. Albert Einstein. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, second edition, 1979. ISBN 83-01-00079-1. 188 + 4 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1980:QA

[Inf80] Léopold Infeld. Quest: an autobiography. Chelsea Pub. Co., New York, NY, USA, second edition, 1980. ISBN 0-8284-0309-0 (hardcover), 0-8218-4073-8. 361 pp. LCCN QC16.I6 A3 1980.

Infeld:1983:ESO

[Inf83] Leopold Infeld. Einstein, su obra y su influencia en nuestro mundo. (Spanish) [Einstein, his work and his influence in our world]. Editorial Leviatán, Buenos Aires, Argentina, 1983. 197 pp. LCCN ????

Infeld:1984:AE

[Inf84] Leopold Infeld. Albert Einstein. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, third edition, 1984. ISBN 83-01-05210-4. 188 + 4 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1989:QES

[Inf89] Leopold Infeld. Quest: the evolution of a scientist. Transaction Publishers, New Brunswick, NJ, USA, 1989. ISBN 0-88738-796-9. ???? pp. LCCN QC16.I6 A318 1989. With a new introduction by Peter Bergmann.

Infeld:1992:AEU

[Inf92] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world], volume 23 of Piccola biblioteca Einaudi. Giulio Einaudi, Torino, Italia, thirteenth edition, 1992. ISBN 88-06-04416-8. 145 pp. LCCN ????

Infeld:1999:MNE

[Inf99] Leopold Infeld. Der Mann neben Einstein : ein Leben zwischen Raum und Zeit. (German) [The man next to Einstein: a life between space and time]. Weymann-Bauer-Verlag, Rostock, Germany, 1999. ISBN 3-929395-42-8. 234 pp. LCCN ???? Translation to German by Kurt Kelm of original Polish edition [?].

Infeld:2006:QA

[Inf06] Leopold Infeld. Quest: an autobiography. AMS Chelsea Publishing, Providence, RI, USA, second edition, 2006. ISBN 0-8218-4073-8, 0-8284-0309-0. 361 pp. LCCN QC16.I6 A3 1980.

Infeld:2015:MGE

[Inf15] Eva Infeld. My grandfather and Einstein. Web site, January

6, 2015. URL https://www.fiatphysica.com/blog/people/
einstein-theory-of-relativity-evolution-of-physics.

Ingber:2005:ACL

[Ing05] Marc Ingber, editor. Alles über Champignons; Ein Lehrpfad durch die Erdgeschichte; Mit Albert Einstein ins Universum; In Klees Gesicht schauen. (German) [All about mushrooms; A trail through the geological history; With Albert Einstein into the universe; looking at Klees’ face], volume Heft 8/2005 of Die neue Schulpraxis. Zollikofer, St. Gallen, Switzerland, 2005. 63 pp. LCCN ????

Inhaber:1976:QRN

[IP76] H. Inhaber and K. Przednowek. Quality of research and the Nobel Prizes. Social Studies of Science, 6(1):33–50, February 1, 1976. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600102.

Israelit:1985:EDP

[IR85] Mark Israelit and Nathan Rosen. Einstein: Distant parallelism and electromagnetism. Foundations of Physics, 15(3):365–377, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00737324.

Ireland:1989:AE

[Ire89] Karin Ireland. Albert Einstein. Pioneers in change. Silver Burdett Press, Englewood Cliffs, NJ, USA, 1989. ISBN 0-382-09523-5 (lib. bdg.). 108 pp. LCCN QC16.E5 I74 1989.

Irons:2004:REA

[Iro04] F. E. Irons. Reappraising Einstein’s 1905 application of thermodynamics and statistics to radiation. European Journal of Physics, 25(??):269–277, 2004. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://ej.

iop.org/links/rsy0lFzw0/nj_s3lPp2xG8KOK4av5vpA/ejp4_
2_013.pdf;   http://stacks.iop.org/EJP/25/269.

Irving:1967:GAB

[Irv67a] David John Cawdell Irving. The German atomic bomb; the history of nuclear research in Nazi Germany. Simon and Schuster, New York, NY, USA, 1967. 329 pp. LCCN QC773.A1 I69.

Irving:1967:VHG

[Irv67b] David John Cawdell Irving. The virus house: [Germany’s atomic research and Allied counter-measures]. William Kimber, London, UK, 1967. 288 + 12 pp. LCCN QC773.A1 I7.

Irving:1983:GAB

[Irv83] David John Cawdell Irving. The German atomic bomb: the history of nuclear research in Nazi Germany. A Da Capo paperback. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1983. ISBN 0-306-80198-1 (paperback). 329 + 16 pp. LCCN QC773.3.G3 I78 1983. US$9.95.

Israel:1931:HAG

[IRW31] Hans Israël, Erich Ruckhaber, and Rudolf Weinmann, editors. Hundert Autoren gegen Einstein. (German) [One hundred authors against Einstein]. R. Voigtländer Verlag, Leipzig, Germany, 1931. 104 pp. LCCN ???? URL https:/

/www.google.com/books/edition/Hundert_Autoren_gegen_
Einstein/kmi4AAAAIAAJ.

Infeld:1949:MTP

[IS49] L. Infeld and A. Schild. On the motion of test particles in General Relativity. Reviews of Modern Physics, 21(3):408– 413, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.408;   http://rmp.aps.
org/abstract/RMP/v21/i3/p408_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Isaksson:1985:FPG

[Isa85] Eva Isaksson. Der finnische Physiker Gunnar Nordström und sein Beitrag zur Entstehung der allgemeinen Relativitätstheorie A. Einsteins. (German) [The Finnish physicist Gunnar Nordström and his contribution to the development of A. Einstein’s General Theory of Relativity]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 22(??):29–52, ???? 1985. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic).

Isaacson:2007:EHL

[Isa07] Walter Isaacson. Einstein: his life and universe. Simon and Schuster, New York, NY, USA, 2007. ISBN 0-7432-6473-8. xxii + 675 + 16 pp. LCCN QC16.E5 I76 2007. URL http://www.

loc.gov/catdir/enhancements/fy0708/2006051264-d.html.

Isaacson:2009:HED

[Isa09] Walter Isaacson. How Einstein divided America’s Jews. Atlantic Monthly, ??(??):??, December 2009. ISSN 1072-7825 (print),

2151-9463 (electronic). URL http://www.theatlantic.com/
magazine/archive/2009/12/how-einstein-divided-americas-
jews/307763/.

Isaacson:2015:HHE

[Isa15] Walter Isaacson. History: How Einstein reinvented reality. Scientific American, 313(3):38–45, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-38.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-38.pdf.

Ishiwara:1923:REL

[Ish23] Jun Ishiwara. The Record of Einstein’s Lectures. Kaizosha, Tokyo, Japan, 1923. ???? pp. LCCN ???? Reprinted in 1971 by Tosho in Tokyo.

Ishiwara:1977:EKR

[Ish77] Jun Ishiwara. Einstein Kōen-Roku. Tokyo-Toshu, Tokyo, Japan, 1977. Japanese translation of German lecture that Einstein gave in Kyoto in December 1922. See also English translation [Oga79], and its note on Einstein’s knowledge in 1905 of the Michelson–Morley experiment of 1887.

Israel:1973:RAC

[Isr73] W. Israel, editor. Relativity, astrophysics and cosmology. Proceedings of the summer school held, 14–26 August, 1972 at the Banff Centre, Banff Alberta, volume 38 of Astrophysics and space science library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0369-8. LCCN QB460 .R44.

Israel:1998:EPR

[Isr98] Werner Israel. 1905: Einstein’s paper revolution. Physics World, 11(10):47–48, October 1998. CODEN PHWOEW.

ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/11/10/phwv11i10a33.
pdf.

Itagaki:1982:WDM

[Ita82] Ryoichi Itagaki. Why did Mach reject Einstein’s theory of relativity? Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):81–95, 1982. CO-DEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.

jp/naid/110009837277.

Itagaki:1999:EKL

[Ita99] Ryoichi Itagaki. Einstein’s ‘Kyoto Lecture’: The Michelson– Morley experiment. Science, 283(5407):1455, March 5, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/283/

5407/1455.4.full. The author disputes earlier translations from German to Japanese to English [Oga79, Ein82b], and concludes that Einstein did not know of the Michelson–Morley experiment.

Ivry:1999:BRE

[Ivr99] Sara Ivry. Book review: Einstein’s Daughter: The Search for Lieserl. By Michele Zackheim. Riverhead, $26.95. New York Times, ??(??):BR21, November 28, 1999. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
110111525/.

Jabr:2011:BNC

[Jab11] Ferris Jabr. Biology: A Nobel celebration. Scientific American, 304(6):54–63, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v304/n6/full/
scientificamerican0611-54.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
54.pdf.

Jabr:2013:CNG

[Jab13] Ferris Jabr. Chemistry: a Nobel gathering. Scientific American, 309(1):68–79, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v309/n1/full/

scientificamerican0713-68.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
68.pdf.

Jackson:1987:ISR

[Jac87] J. David Jackson. The impact of Special Relativity on theoretical physics. Physics Today, 40(5):34–42, May 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://scitation.aip.org/content/aip/magazine/
physicstoday/article/40/5/10.1063/1.881108.

Jacobi:2008:SNA

[Jac08] Manfred Jacobi. Die Suche nach dem Absoluten. Zum 150. Geburtstag von Max Planck. (German) [The search for the absolute. On the 150th birthday of Max Planck]. Physik in unserer Zeit, 39(2):64–70, 2008. CODEN PHUZAH. ISSN 1521-3943.

Jacobs:2010:EG

[Jac10] Howy Jacobs. Einstein’s grandchildren. EMBO reports, 11(4): 239, April 2010. CODEN ???? ISSN ???? URL http://www.

nature.com/embor/journal/v11/n4/full/embor201038.html.

Jacobsen:2011:AUH

[Jac11] Annie Jacobsen. Area 51: an uncensored history of America’s top secret military base. Little, Brown and Co., Boston, MA, USA, 2011. ISBN 0-316-13294-2. xv + 523 + 32 pp. LCCN UG634.5.A74 J33 2011. US$27.99.

Jacob:2012:HDP

[Jac12] Antoine Jacob. Histoire du prix Nobel. Bourin, Paris, France, 2012. ISBN 2-84941-338-0 (paperback). 256 pp. LCCN AS911.N9 J33 2012. 22,00 EUR.

Jaeger:2006:EKR

[Jae06] Lydia Jaeger. Einstein und die kosmische Religion. (German) [Einstein and the cosmic religion]. Philosophia Naturalis, 43(2): 313–327, ???? 2006. CODEN ???? ISSN 0031-8027.

Jagerman:2001:MRR

[Jag01] Louis S. Jagerman. The mathematics of relativity for the rest of us: all you wanted to know about relativity but thought you couldn’t possibly understand. Trafford, Victoria, BC, Canada, 2001. ISBN 1-55212-567-X. ii + 447 pp. LCCN QC173.57.

Jaki:1978:JGS

[Jak78] Stanley L. Jaki. Johann Georg von Soldner and the gravitational bending of light, with an English translation of his essay on it published in 1801. Foundations of Physics, 8(11–12): 927–950, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/BF00715064.

Jakobson:1982:ESL

[Jak82] Roman Jakobson. Einstein and the science of language. In Holton and Elkana [HE82], pages 139–150. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Jaksch:2008:BEC

[Jak08] Dieter Jaksch. Bose–Einstein condensates: a peek and a poke. Nature Physics, 4(12):906–908, December 2008. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic).

URL http://www.nature.com/nphys/journal/v4/n12/full/
nphys1143.html.

Jakobsen:2013:DBB

[Jak13] Knud Jakobsen. Danskeren bag bomben: Om tømrerlærlingen fra Holstebro, der hjalp atombomben til verden og bestemte, hvor den skulle kastes. (Danish) [The Dane behind the bomb: about the carpenter’s apprentice from Holstebro who helped develop the atomic bomb and decided where it should be dropped]. Berlingske Media Forlag, København, Danmark, 2013. ISBN 87-7108-983-7. 279 pp. LCCN C774.L39 .J35 2013.

James:1930:ENG

[Jam30] Edwin L. James. Einstein near goal of a unified cosmos: Scientists now keenly await the equations to unite all physical phenomena. text of English lecture. Physicist traces evolution of views of the universe — sees solution of space next step. New York Times, ??(??):??, June 8, 1930. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/98927700/.

Jammer:1966:CDQ

[Jam66] Max Jammer. The conceptual development of quantum mechanics. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1966. xii + 399 pp. LCCN QC174.1 .J26.

Jammer:1970:BRB

[Jam70] Max Jammer. Book review: Space through the Ages. The Evolution of Geometrical Ideas from Pythagoras to Hilbert and Einstein. Cornelius Lanczos. Academic Press, New York, 1970. x, 322 pp., illus. $11.50. Science, 170(3963):1183, December 11, 1970. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/170/3963/1183.1.full.pdf.

Jammer:1974:BED

[Jam74a] Max Jammer. The Bohr–Einstein debate. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 5, pages 108–158. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:EVC

[Jam74b] Max Jammer. Early versions of the complementarity interpretation. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 4, pages 85–107. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:IOL

[Jam74c] Max Jammer. The incompleteness objection and later versions of the complementarity interpretation. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 6, pages 159–252. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74d] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. Wiley, New York, NY, USA, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

Jammer:1982:EQP

[Jam82] Max Jammer. Einstein and quantum physics. In Holton and Elkana [HE82], pages 59–78. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Jammer:1989:CDQ

[Jam89] Max Jammer. The Conceptual Development of Quantum Mechanics, volume 12 of The History of modern physics, 1800-1950. Tomash Publishers, Los Angeles, CA, USA, second edition, 1989. ISBN 0-88318-617-9. xvii + 436 pp. LCCN QC174.12 .J36 1989.

James:1993:NLC

[Jam93] Laylin K. James, editor. Nobel laureates in chemistry, 1901– 1992. History of modern chemical sciences. American Chemical Society, Washington, DC, USA, 1993. ISBN 0-8412-2459-5 (hardcover), 0-8412-2690-3 (paperback). xvii + 798 pp. LCCN QD21 .N63 1993.

Jammer:1999:ERP

[Jam99] Max Jammer. Einstein and religion: physics and theology. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00699-7 (harcover), 0-691-10297-X (paperback). 279 pp. LCCN QC16.E5 J36 1999. URL http://site.ebrary.

com/lib/academiccompletetitles/home.action;   http:/
/site.ebrary.com/lib/alltitles/docDetail.action?docID=
10496617;   http://www.gbv.de/dms/bowker/toc/9780691006994.
pdf;   http://www.loc.gov/catdir/description/prin021/
99024124.html;   http://www.loc.gov/catdir/samples/
prin031/99024124.html;   http://www.loc.gov/catdir/
toc/prin032/99024124.html.

Jammer:2002:ERP

[Jam02] Max Jammer. Einstein and religion: physics and theology. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-10297-X (paperback), 1-4008-4087-2 (e-book). viii + 279 pp. LCCN QC16.E5 J36 2002. UK£11.95. URL http://www.

gbv.de/dms/bowker/toc/9780691102979.pdf.

James:2006:ASH

[Jam06] I. M. (Ioan Mackenzie) James. Asperger’s syndrome and high achievement: some very remarkable people. Jessica Kingsley, London, UK, 2006. ISBN 1-84310-388-5 (paperback). 224 pp. LCCN RC553.A88 J35 2006.

Janis:1973:BRB

[Jan73] Allen I. Janis. Book review: The Conceptual Foundations of Contemporary Relativity Theory: John Cowperthwaite Graves, Cambridge, Mass.: The MIT Press, 1971. xi + 361 pp. $15.00. Studies in History and Philosophy of Science Part A, 4(3):300– 306, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0039368173900137. See [Gra71].

Janis:1977:FRB

[Jan77] Allen I. Janis. Film review: Doctor Einstein Before Lunch, produced by NBC. American Journal of Physics, 45(1):116, January 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/45/1/10.1119/1.10897.

Janssen:1995:CBL

[Jan95] Michael Heinrich Paul Janssen. A Comparison Between Lorentz’s Ether Theory and Special Relativity in the Light of the Experiments of Trouton and Noble. Ph.D. dissertation in history and philosophy of science, University of Pittsburgh, Pittsburgh, PA, USA, November 1995. xv + 312 pp. URL http://

pittcat.pitt.edu/cgi-bin/Pwebrecon.cgi?BBID=1570240;
 http://search.proquest.com/docview/304219554.

Janssen:2002:RSR

[Jan02] Michel Janssen. Reconsidering a scientific revolution: The case of Einstein versus Lorentz. Physics in Perspective (PIP), 4(4):421–446, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/53vxdwrfnm99h8l3/fulltext.
pdf.

Janssen:2003:TES

[Jan03] Michel Janssen. The Trouton experiment, E = mc^{2}, and a slice of Minkowski space–time. In Ashtekar et al. [ACH^{+}03], pages 27–54. ISBN 1-4020-1284-5. LCCN

Q174 .B67 vol. 234. URL http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-d.html;   http://www.
loc.gov/catdir/enhancements/fy0813/2003046757-t.html.

Janssen:2005:PHE

[Jan05] Michel Janssen. Of pots and holes: Einstein’s bumpy road to general relativity. Annalen der Physik (Berlin), 14(1):58–85, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/andp.200410130.

Janssen:2006:CEA

[Jan06] Michel Janssen. 2005: The centenary of Einstein’s annus mirabilis. Studies in History and Philosophy of Modern Physics, 37(1):1–4, March 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S135521980500095X.

Janssen:2007:WDE

[Jan07] Michel Janssen. What did Einstein know and when did he know it? A Besso memo dated August 1913. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 785–837. ISBN 1-4020-3942-5. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 2; QC173.6

.G469 2007. URL https://link.springer.com/chapter/10.
1007/978-1-4020-4000-9_8.

Janis:2008:SRC

[Jan08] Allen I. Janis. Simultaneity, Relativity and conventionality. Studies in History and Philosophy of Modern Physics, 39(1): 217–224, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219807000986.

Janssen:2012:TBH

[Jan12] Michel Janssen. The twins and the bucket: How Einstein made gravity rather than motion relative in general relativity. Studies in History and Philosophy of Modern Physics, 43(3):159–175, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000056.

Janssen:2014:ASR

[Jan14a] Michel Janssen. Appendix: Special relativity. In Janssen and Lehner [JL14b], pages 455–506. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Janssen:2014:NSL

[Jan14b] Michel Janssen. “No success like failure … ”: Einstein’s quest for General Relativity, 1907–1920. In Janssen and Lehner [JL14b], chapter 6, pages 167–227. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Jarlskog:2008:LRN

[Jar08a] Cecilia Jarlskog. Lord Rutherford of Nelson, his 1908 Nobel Prize in Chemistry, and why he didn’t get a second prize. Journal of Physics: Conference Series, 136(1):012001:1–012001:16, 2008. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596

(electronic). URL http://stacks.iop.org/1742-6596/136/

i=1/a=012001. Presented at the XXIII Conference on Neutrino Physics and Astrophysics.

Jarlskog:2008:RNP

[Jar08b] Cecilia Jarlskog. Rutherford’s Nobel Prize and the one he didn’t get. CERN Courier: International Journal of High-Energy Physics, 59(??):??, ???? 2008. CODEN CE-COA2. ISSN 0304-288X (print), 2077-9550 (electronic).

URL https://cerncourier.com/rutherfords-nobel-prize-
and-the-one-he-didnt-get/.

Jammer:1995:ERG

[JAvWE95a] Max Jammer, Jürgen Audretsch, Carl Friedrich von Weizsäcker, and Albert Einstein. Einstein und die Religion. (German) [Einstein and Religion]. Universitätsverlag Konstanz, Konstanz, Germany, 1995. ISBN 3-87940-484-4. 125 pp. LCCN BL240.2. DM 24.80; SFR 24.80; S 194.00. URL http://www.gbv.de/

dms/goettingen/148710794.pdf.

Jammer:1995:ER

[JAvWE95b] Max Jammer, Jürgen Audretsch, Carl Friedrich von Weizsäcker, and Albert Einstein. Einstein und die Religion. (German) [Einstein and Religion]. Universitätsverlag Konstanz, Konstanz, Germany, 1995. ISBN 3-87940-484-4. 125 pp. LCCN BL240.2. DM 24.80; SFR 24.80; S 194.00. URL http://www.gbv.de/

dms/goettingen/148710794.pdf.

Jayaprakash:1990:MHN

[Jay90] Nallukunnel Damodaran Jayaprakash. The meaning of Hiroshima Nagasaki: the decision to use atomic bombs on Japan and its implications for humankind. Delhi Science Forum, New Delhi, India, 1990. 150 pp. LCCN D767.25.H6.

Jayaraman:2005:AER

[Jay05] T. Jayaraman. Albert Einstein: Radical pacifist and democrat. Current Science, 89(12):2141–2145, December 25, 2005. CO-DEN CUSCAM. ISSN 0011-3891.

Jorda:1997:PNE

[JDB^{+}97] S. Jorda, E. Dreisigacker, Gunnar Berg, R. Scharf, I. Peschel, R. Bock, I. Hofmann, J. Redfearn, S. Jorda, and R. Scharf. Physik-Nobelpreis für die Entwicklung von Laserkühlung und Atomfallen/kurzgefaßt/Max Planck zum Gedenken/WE-Heraeus-Programm Weiterbildung in den Grundlagen unternehmerischen Denkens und Handelns jetzt auch für Fachhochschulen/Einstein im ADLON/kurzgefaßt/Auf Lebenszeit Studierender? Oder: Beitragsvermeidung durch Stillhalten/Frühjahrstagungen: Internet-Adressen/Blick in die USA/ Großbritannien: Diamond und JET/JET setzt Weltrekordmarken/Trägheitsfusion mit schweren Ionen im Aufwind/ Trägheitsfusion/Erstes Bose–Einstein-Kondensat in Europa/ Quasiteilchen mit gebrochenzahliger Ladung beobachtet/ Laserlicht verwandelt sich in Materie. (German) [Nobel Prize in Physics for the development of laser cooling and atom traps/in brief/Max Planck commemoration/we-heraeus-program training in the basics of entrepreneurship now for polytechnics/ Einstein in ADLON/in brief/student’s life? or: Post prevention by keeping quiet/Spring Meetings: Internet addresses/looking at the United States/UK: Diamond and JET/JET sets world record marks/inertial confinement fusion with heavy ions on the rise/inertial confinement fusion/first Bose–Einstein condensate in Europe/quasiparticles with the fractional charge observed/ laser light turns into matter]. Physikalische Blätter, 53(11): 1078–1088, 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Jeans:1930:MUa

[Jea30a] Sir James Hopwood Jeans. The Mysterious Universe. Cambridge University Press, Cambridge, UK, 1930. viii + 154 pp. LCCN Q171 .J37.

Jeans:1930:MUb

[Jea30b] Sir James Hopwood Jeans. The Mysterious Universe. The Macmillan Company, New York, NY, USA, 1930. viii + 2 + 163 pp. LCCN Q171 .J37.

Jeans:1932:MU

[Jea32] Sir James Hopwood Jeans. The Mysterious Universe. Cambridge University Press, Cambridge, UK, second edition, 1932. viii + 142 pp. LCCN Q171 .J37.

Jeans:1938:MU

[Jea38] Sir James Hopwood Jeans. The Mysterious Universe. Penguin, London, UK and New York, NY, USA, second edition, 1938. ix + 192 pp. LCCN Q171 .J37.

Jeans:1943:MU

[Jea43] Sir James Hopwood Jeans. The Mysterious Universe. Cambridge University Press, Cambridge, UK, second edition, 1943. viii + 142 pp. LCCN Q171 .J37 1948.

Jeans:1948:MU

[Jea48] Sir James Hopwood Jeans. The Mysterious Universe. Cambridge University Press, Cambridge, UK, 1948. 142 pp. LCCN Q171 .J37 1948.

Jeans:1976:MU

[Jea76] Sir James Hopwood Jeans. The Mysterious Universe. AMS Press, New York, NY, USA, 1976. ISBN 0-404-14742-9. xi + 192 + 1 pp. LCCN Q171 .J37 1976.

Jeandron:2008:RBE

[Jea08] Michelle Jeandron. Record bid for Einstein letter. Physics World, 21(6):13, June 2008. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/21/06/phwv21i06a22.
pdf.

Jehle:1951:BRB

[Jeh51] Herbert Jehle. Book review: Albert Einstein—Philosopher-Scientist, by Paul Arthur Schilpp. American Journal of Physics, 19(4):252–253, April 1951. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/?AJP/19/252/2.

Jenkin:2001:UPW

[Jen01] John Jenkin. A unique partnership: William and Lawrence Bragg and the 1915 Nobel Prize in Physics. Minerva [Washington, DC], 39(4):373–392, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012783802528.

Jennings:2002:HWB

[Jen02] Peter Jennings. Hiroshima: Why the bomb was dropped. MPI Home Video, Orland Park, IL, USA, 2002. ISBN 0-7886-0444-9. LCCN D767.25 .H6 H57 2002.

Jerome:2002:EFJ

[Jer02] Fred Jerome. The Einstein file: J. Edgar Hoover’s secret war against the world’s most famous scientist. St. Martin’s Press, New York, NY, USA, 2002. ISBN 0-312-28856-5. xxii + 358 pp. LCCN QC16.E5 J46 2002. URL http://www.loc.gov/

catdir/toc/fy022/2001058850.html.

Jette:1977:IB

[Jet77] Eleanor Jette. Inside Box 1663. Los Alamos Historical Society, Los Alamos, NM, USA, 1977. 132 pp. LCCN QC773.3.U5 J47 1977.

Jette:2007:IB

[Jet07] Eleanor Jette. Inside Box 1663. Los Alamos Historical Society, Los Alamos, NM, USA, second edition, 2007. ISBN 0-941232-02-6 (paperback). 253 pp. LCCN QC773.3.U5

J47 2007. URL http://www.loc.gov/catdir/toc/ecip0719/
2007022927.html.

Jin:1996:EK

[Jin96] Doo Jung Jin. Einstein for kg? Nature, 383(6602):662, October 24, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v383/n6602/pdf/383662c0.pdf.

Johnson:1981:CBF

[JJ81] Charles W. Johnson and Charles O. Jackson. City behind a fence: Oak Ridge, Tennessee, 1942–1946. University of Tennessee Press, Knoxville, TN, USA, 1981. ISBN 0-87049-303-5, 0-87049-309-4 (paperback). xxiii + 248 pp. LCCN F444.O3 J63. US$14.50.

Janssen:2014:ICC

[JL14a] Michel Janssen and Christoph Lehner. Introduction: [the Cambridge companion to Einstein]. In Janssen and Lehner [JL14b], pages 1–37. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Janssen:2014:CCE

[JL14b] Michel Janssen and Christopher Lehner, editors. The Cambridge Companion to Einstein. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). xvi + 562 pp. LCCN QC16.E5 C36 2014.

Jungnickel:1986:IMN

[JM86] Christa Jungnickel and Russell McCormmach. Intellectual mastery of nature: theoretical physics from Ohm to Einstein. University of Chicago Press, Chicago, IL, USA and London, UK, 1986. ISBN 0-226-41581-3 (vol. 1), 0-226-41584-8 (vol. 2). xxiii + 350 pp. LCCN QC7.J86 1986. US$55.00.

Jungnickel:2017:SPH

[JM17] Christa Jungnickel and Russell McCormmach. The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-49564-X (hardcover), 3-319-49565-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxi + 460 pp. LCCN QC9.G3 J86 2017.

Johnstone:2005:EBI

[JMH^{+}05] Gary Johnstone, Aidan McArdle, Shirley Henderson, John Lithgow, and David Bodanis. Einstein’s big idea. WGBH Boston Video, Boston, MA, USA, letterbox edition, 2005. ISBN 1-59375-317-9. LCCN QC173.6 .N68 2005. 1 videodisc (ca. 112 min.).

Jentschura:2014:ADF

[JN14] U. D. Jentschura and I. Nándori. Attempts at a determination of the fine-structure constant from first principles: a brief historical overview. European Physical Journal H, 39(5):591–613, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2014-50044-7.

Janssen:2007:IVZ

[JNR^{+}07a] Michel Janssen, John Norton, Jürgen Renn, Sauer Tilmann, and John Stachel. Introduction to volumes 1 and 2: The Zurich Notebook and the genesis of General Relativity. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 7–20. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_1.

Janssen:2007:GGRa

[JNR^{+}07b] Michel Janssen, John Norton, Jürgen Renn, Sauer Tilmann, John Stachel, and Lindy Divarci. The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source, volume 250(1) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. 487 pp. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

Joel:1919:UAE

[Joe19] Kurt Joel. Unterredung mit Albert Einstein. (German) [Interview with Albert Einstein]. Neues Wiener Journal, ??(??):??, December 25, 1919.

Joffe:1955:PAE

[Jof55] Abraham F. Joffe. Pamyati Alberta Eynshtina (Russian) [Memories of Albert Einstein]. Uspekhi Fizicheskikh Nauk, 57(2):188– 192, ???? 1955. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL https://ufn.ru/ru/articles/
1955/10/b/.

Joffe:1962:VFM

[Jof62] Abraham Joffe. Vstrechi s fizikami, moi vospominaniia o zarubezhnykh fizikah. (Russian) [Meetings with Physicists, My Reminiscences of Physics Abroad]. Gusudarstvennoye Izdatelstvo Fiziko-Matematitsheskoi Literatury, Moscow, USSR, 1962. 141 pp.

Joffe:1967:BPG

[Jof67] Abraham Joffe. Begegnungen mit Physikern. (German) [Meetings with physicists]. Teubner Verlag, Leipzig, East Germany, second edition, 1967. 141 pp. German translation by Konrad Werner of Russian original [Jof62].

Johnsen:1946:AB

[Joh46] Julia E. Johnsen. The Atomic Bomb. H. W. Wilson Company, New York, NY, USA, 1946. 335 pp. LCCN HD9698.A2 J6.

Johnson:1995:BRBb

[Joh95] Jeffrey A. Johnson. Book review: Fritz Haber: Chemiker, Nobelpreistrager, Deutscher, Jude by Dietrich Stoltzenberg. Isis, 86 (3):518–519, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235088.

Johansen:2004:PNP

[Joh04] T. K. Johansen. Plato’s natural philosophy: a study of the Timaeus–Critias. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-79067-0. vi + 218 pp.

LCCN B387 .J64 2004. URL http://www.loc.gov/catdir/
enhancements/fy0668/2003070331-d.html;   http://www.
loc.gov/catdir/enhancements/fy0668/2003070331-t.html;
 http://www.loc.gov/catdir/enhancements/fy0732/2003070331-
b.html.

Johnson:2005:CE

[Joh05a] M. Alex Johnson. The culture of Einstein. NBC News,

April 18, 2005. URL http://www.nbcnews.com/id/7406337/#
.UpygZOK7SOt.

Johnston:2005:WEN

[Joh05b] Sean F. Johnston. From white elephant to Nobel Prize: Dennis Gabor’s wavefront reconstruction. Historical Studies in the Physical and Biological Sciences, 36(1):35–70, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Johnson:2006:MMR

[Joh06] Jeffrey Allan Johnson. Master Mind: The Rise and Fall of Fritz Haber, the Nobel Laureate Who Launched the Age of Chemical Warfare (review). Technology and Culture, 47(4):835–837, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204984.

Johnston:2007:HLH

[Joh07] Barbara Rose Johnston, editor. Half-Lives and Half-Truths: Confronting the Radioactive Legacies of the Cold War. Resident scholar series. School for Advanced Research Press, Santa Fe, NM, USA, 2007. ISBN 1-930618-82-4 (paperback). x + 326 pp. LCCN RA393 .H33 2007. URL http://www.loc.gov/

catdir/toc/ecip077/2006102477.html.

Johnson:2011:ASS

[Joh11a] Amy (Poet) Johnson. Aung San Suu Kyi: a Nobel Peace Prize winner and a female Burmese President-Elect under seige: a treatise, sonnets for a modern trailblazer. AuthorHouse, Central Milton Keynes, UK, 2011. ISBN 1-4520-7876-9 (paperback). x + 65 pp. LCCN ????

Johnston:2011:PPL

[Joh11b] Sean F. Johnston. Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein, by Bruce J. Hunt. Technology and Culture, 52(2):403–404, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

http://muse.jhu.edu/article/431377.

Johnson:2013:BEC

[Joh13] Dexter Johnson. Bose–Einstein condensate made at room temperature for first time. IEEE Spectrum, ??(??):??, December 10, 2013. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340

(electronic). URL http://spectrum.ieee.org/nanoclast/
semiconductors/materials/bose-einstein-condensate-made-
at-room-temperature-for-first-time. See [PSM^{+}13] for the

original research.

Johnston:2014:WTP

[Joh14] Hamish Johnston. What type of physics should you do if you want to bag a Nobel prize? Physics World, 27(10):??, October 2, 2014. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058

(electronic). URL https://physicsworld.com/a/what-type-
of-physics-should-you-do-if-you-want-to-bag-a-nobel-
prize/.

Johnsen:2017:AB

[Joh17] Julia E. Johnsen, editor. The atomic bomb. Deutsche Nationalbibliothek, Leipzig and Frankfurt am Main, Germany, 2017. 335 pp.

Johns:2019:VEH

[Joh19] Oliver Davis Johns. Validity of the Einstein hole argument. Studies in History and Philosophy of Modern Physics, 68(??): 62–70, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S135521981830025X.

Jones:1944:CE

[Jon44] H. Spencer Jones. From Copernicus to Einstein. Nature, 153 (3872):67–68, January 15, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v153/n3872/pdf/153067a0.pdf.

Jones:1985:MAA

[Jon85] Vincent C. Jones. Manhattan, the Army and the atomic bomb. United States Army in World War II. Special studies. Center of Military History, U.S. Army, Washington, DC, USA, 1985. ISBN 99941-62-44-6. xx + 660 + 4 pp. LCCN QC773.3.U5 J65 1985.

Jones:1989:WMS

[Jon89] D. E. H. Jones. The world of mathematics — a small library of the literature of mathematics from Ah-Mose the Scribe to Albert Einstein — J. R. Newman. Nature, 337(6206):420, February 2, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Jones:2001:TSB

[Jon01] Peter Jones. The true story of Britain’s H-bomb tests: [book review:] Britain and the H-Bomb, Lorna Arnold, 2001 Palgrave 288pp £15.99pb/$69/95hb. Physics World, 14(9):54–55, September 2001. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://stacks.iop.org/2058-
7058/14/i=9/a=34.

Jones:2006:TTE

[Jon06] Nicola Jones. Testing times for Einstein’s theory. In The Field, February 21, 2006. URL http://blogs.nature.

com/inthefield/2006/02/testing_times_for_einsteins_
th.html.

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Jones:2017:PCF

[Jon17] B. J. T. (Bernard Jean Trefor) Jones. Precision cosmology: the first half million years. Cambridge University Press, Cambridge, UK, 2017. ISBN 0-521-55433-0. xiv + 761 pp. LCCN QB981 .J664 2017.

Jordan:1969:AES

[Jor69] Pascual Jordan. Albert Einstein. Sein Lebenswerk und die Zukunft der Physik. (German) [Albert Einstein. His life’s work and the future of physics], volume 6 of Wirkung und Gestalt. Huber, Frauenfeld, Switzerland and Stuttgart, Germany, 1969. 302 pp. LCCN QC16.E5.

Jost:1979:EZZ

[Jos79] Res Jost. Einstein und Zürich, Zürich und Einstein. Vierteljahrsschrift der Naturforschenden Gesellschaft in Zürich, 124 (??):7–23, ???? 1979. CODEN VNGZAL. ISSN 0042-5672.

Jost:1980:CEP

[Jos80] Res Jost. Comment on “Einstein on particles, fields, and the quantum theory”. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Josephson:1995:BRB

[Jos95a] Paul R. Josephson. Book review: Alexander A. Friedmann: The Man Who Made the Universe Expand, by Eduard A. Tropp, Artur D. Chernin, Alexander Dron, Michael Burov, and Victor Ya. Frenkel. Journal of Interdisciplinary History, 26(2):318– 320, ???? 1995. ISSN 0022-1953 (print), 1530-9169 (electronic).

URL http://www.jstor.org/stable/206645.

Jost:1995:MET

[Jos95b] Res Jost. Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers], volume 34 of Lecture notes in physics. New series M, monographs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. 286 pp. LCCN

QC16.J67 J67 1995. URL http://www.springerlink.com/
openurl.asp?genre=issue\%26issn=0940-7677\%26volume=

34. Edited posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Joseph:2009:HPM

[Jos09] Timothy W. Joseph, editor. Historic photos of the Manhattan Project. Turner Publishing Company, Nashville, TN, USA, 2009. ISBN 1-59652-521-5. x + 205 pp. LCCN QC773.A1

J67 2009. URL http://www.loc.gov/catdir/enhancements/
fy1204/2008910973-d.html.

Jung:1952:NPS

[JP52] C. G. (Carl Gustav) Jung and Wolfgang Pauli, editors. Naturerklärung und Psyche Synchronizität als ein Prinzip akausaler Zusammenhänge [von C. G. Jung]. Der Einfluss archetypischer Vorsstellungen auf die Bildung naturwissenschaftlicher Theorien bei Kepler [von W. Pauli]: Synchronizität als ein Prinzip akausaler Zusammenhänge, volume 4 of Studien aus dem C. G. Jung-Institut, Zürich. Rascher, Zürich, Switzerland, 1952. 194 pp. LCCN BF1033 .N38.

Jung:1955:INP

[JP55] C. G. (Carl Gustav) Jung and Wolfgang Pauli, editors. The Interpretation of nature and the psyche. Synchronicity: an acausal connecting principle [by C. G. Jung]. The influence of archetypal ideas on the scientific theories of Kepler [by W. Pauli], volume 51 of Bollingen series. Pantheon Books, New York, NY, USA, 1955. vii + 247 pp. LCCN BF1033 .N385.

Janta:1989:TPP

[JP89] J. Janta and J. Pantoflíček, editors. Trends in Physics, 1984: Proceedings of the 6th General Conference of the European Physical Society, 27–31 August 1984, Prague, Czechoslovakia. Polygrafia [for the Union of Czechoslovak Mathematicians and Physicists], Prague, Czechoslovakia, 1989. LCCN QC1 .E857 1984.

Janke:2008:PIN

[JP08] W. (Wolfhard) Janke and Axel Pelster, editors. Path integrals: new trends and perspectives: proceedings of the 9th International Conference: Dresden, Germany, September 23–28, 2007. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-283-726-4, 981-283-727-2 (ebook). LCCN QC174.17.P27 I57 2007eb.

Janssen:2007:UKH

[JR07] Michel Janssen and Jürgen Renn. Untying the knot: How Einstein found his way back to field equations discarded in the Zurich Notebook. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 839–925. ISBN 1-4020-3942-5. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_9.

Janssen:2015:ASH

[JR15] Michel Janssen and Jürgen Renn. Arch and scaffold: How Einstein found his field equations. Physics Today, 68(11):30– 36, November 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/full/10.1063/PT.3.2979.

Jantzen:2002:NMG

[JRG02] Robert T. Jantzen, Remo Ruffini, and V. G. Gurzadyan, editors. The Ninth Marcel Grossmann Meeting: on recent developments in theoretical and experimental general relativity, gravitation, and relativistic field theories: proceedings of the MGIX MM meeting held at the University of Rome “La Sapienza”, 2-8 July 2000. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2002. ISBN 981-238-010-8 (set), 981-238-995-4 (pt. A), 981-238-994-6 (pt. B), 981-238-993-8 (pt. C). LCCN QC173.6 M37 2000.

Janssen:2007:CNG

[JRN^{+}07a] Michel Janssen, Jürgen Renn, John Norton, Sauer Tilmann, and John Stachel. A commentary on the notes on gravity in the Zurich Notebook. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 489–714. ISBN 1-4020-3942-5. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_6.

Janssen:2007:GGRb

[JRN^{+}07b] Michel Janssen, Jürgen Renn, John Norton, Sauer Tilmann, and John Stachel. The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays, volume 250(2) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3942-5. ISSN 0068-0346. 489–938 pp. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

Jauncey:1991:USE

[JRT^{+}91] D. L. Jauncey, J. E. Reynolds, A. K. Tzioumis, T. W. B. Muxlow, R. A. Perley, D. W. Murphy, R. A. Preston, E. A. King, A. R. Patnaik, D. L. Jones, D. L. Meier, D. J. Bird, D. G. Blair, J. D. Bunton, R. W. Clay, M. E. Costa, R. A. Duncan, R. H. Ferris, R. G. Gough, P. A. Hamilton, D. W. Hoard, A. Kemball, M. J. Kesteven, E. T. Lobdell, A. N. Lurten, P. M. McCulloch, J. D. Murray, G. D. Nicolson, A. P. Rao, A. Savage, M. W. Sinclair, L. Skjerve, L. Taaffe, R. M. Wark, G. L. White, et al. An unusually strong Einstein ring in the radio source PKS1830-211. Nature, 352(6331):132–134, July 11, 1991. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v352/n6331/
pdf/352132a0.pdf.

Janossy:1987:MLJ

[JS87] Lajos Jánossy and Erwin Schrödinger. In memoriam: Lajos Jánossy 75, Erwin Schrödinger 100. MTA Központi Fizikai Kutató Intézete, Budapest, Hungary, 1987. ISBN 963-372-404-X. xx + 148 pp. LCCN QC16.J36 I5 1987.

Janssen:2002:BYWa

[JSI^{+}02a] Michel Janssen, Robert Schulmann, József Illy, Christoph Lehner, and Diana Kormos Buchwald. The Berlin years: writings, 1918–1921, volume 7 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05717-6. xlviii + 689 pp. URL http://

einsteinpapers.press.princeton.edu/vol7-doc;   http:
//einsteinpapers.press.princeton.edu/vol7-trans;   http:
//press.princeton.edu/titles/7184.html;   http://press.
princeton.edu/titles/7185.html.

Janssen:2002:BYWb

[JSI^{+}02b] Michel Janssen, Robert Schulmann, József Illy, Christoph Lehner, and Diana Kormos Buchwald, editors. The Berlin Years: Writings, 1918–1921. English Translation Supplement, volume 7 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05718-4. xvi + 383 pp. LCCN QC16.E5 A2 1987. URL http://press.

princeton.edu/titles/7185.html.

Jeffries:1987:GWO

[JSSZ87] Andrew D. Jeffries, Peter R. Saulson, Robert E. Spero, and Michael E. Zucker. Gravitational wave observatories. Scientific American, 256(6):50–57, 60, June 1987. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jerome:2005:ERR

[JT05] Fred Jerome and Rodger Taylor. Einstein on race and racism. Rutgers University Press, New Brunswick, NJ, USA, 2005. ISBN 0-8135-3617-0 (hardcover). 206 pp. LCCN QC16.E5

J466 2005. URL http://www.loc.gov/catdir/toc/ecip053/
2004023485.html.

Jucovy:2004:BRP

[Juc04] Peter M. Jucovy. Book review: Possessing Genius: The Bizarre Odyssey of Einstein’s Brain. Perspectives in Biology and Medicine, 47(1):148–150, 2004. CODEN PBMEA8. ISSN 0031-5982 (print), 1529-8795 (electronic).

Jungk:1958:BTT

[Jun58] Robert Jungk. Brighter than a thousand suns: a personal history of the atomic scientists. Harcourt Brace, New York, NY, USA, 1958. 369 pp. LCCN QC773 .J813. Translated by James Cleugh.

Jungk:1982:BTT

[Jun82] Robert Jungk. Brighter than a thousand suns: a personal history of the atomic scientists. Penguin, London, UK and New York, NY, USA, 1982. ISBN 0-14-020667-1 (paperback). 329 pp.

Jung:2005:RBZ

[Jun05a] Tobias Jung. Rezension: “Zwei wirkliche Kerle”. Neues zur Entdeckung der Gravitationsgleichungen der Allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert von Daniela Wuensch. (German) [Review: “Two real guys.” News of the discovery of the gravitational equations of General Relativity by Albert Einstein and David Hilbert by Daniela Wuensch]. Berichte zur Wissenschaftsgeschichte, 28(4):356–357, December 2005. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Jung:2005:RBE

[Jun05b] Tobias Jung. Rezensionen: Gödel, Einstein und die Folgen. Vermächtnis einer ungewöhnlichen Freundschaft von Palle Yourgrau. (German) [Reviews: Gödel, Einstein, and the consequences. Legacy of an unusual friendship by Palle Yourgrau]. Berichte zur Wissenschaftsgeschichte, 28(2):187, June 2005. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Jung:2008:AER

[Jun08] Tobias Jung. Albert Einstein: Revolutionär oder “Bewahrer des Alten”?. (German) [Albert Einstein: Revolutionary or “Preserver of the Old?”]. Berichte zur Wissenschaftsgeschichte, 31 (3):264–281, September 2008. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Jundt:2014:GRW

[Jun14] Thomas Jundt. Greening the Red, White, and Blue: the Bomb, Big Business, and Consumer Resistance in Postwar America. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-979120-1 (hardcover). xii + 306 pp. LCCN GE197 .J86 2014.

Juttner:1921:RBE

[Jüt21] F. Jüttner. [Rezensionen:] Die Entstehung und bisherige Entwicklung der Quantentheorie. Nobel-Vortrag, gehalten vor der Königlich Schwedischen Akademie der Wissenschaften zu Stockholm am 2. Juni 1920 von Max Planck. 32 Seiten. Verlag von Joh. Ambr. Barth, Leipzig 1920. Preis geh. 4 Mk. Zeitschrift für Elektrochemie und angewandte physikalische Chemie, 27(23– 24):583, 1921. ISSN 0005-9021.

Johnson:2006:SMA

[JW06] G. W. Johnson and Mark E. Walker. Sir Michael Atiyah’s Einstein Lecture: “The Nature of Space”. Notices of the American Mathematical Society, 53(6):674–678, June/July 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.org/notices/200606/comm-
walker.pdf.

Jones:2011:ADS

[JW11] Benjamin F. Jones and Bruce A. Weinberg. Age dynamics in scientific creativity. Proceedings of the National Academy of Sciences of the United States of America, 108(47):18910–18914, September 2011. CODEN PNASA6. ISSN 0027-8424 (print),

1091-6490 (electronic). URL http://www.pnas.org/content/
early/2011/11/03/1102895108.abstract;   http://www.
pnas.org/lookup/suppl/doi:10.1073/pnas.1102895108/-/
DCSupplemental.

K:1912:EP

[K.12] R. K. Einstein in Prag. Prager Tagblatt, 37(144):2, May 26, 1912.

Kacser:1996:LMM

[Kac96] Claude Kacser. Lise Meitner, Manne Siegbahn, and Adolf Hitler. American Journal of Physics, 63(2):106–107, February 1996. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/63/2/10.1119/1.17992. See [Sim94] and

rebuttal [Tru96].

Kaempffert:1931:HEU

[Kae31] Waldemar Kaempffert. How to explain the universe? Science in a quandary: One after another the theories put out by the scientists have been exploded and now science in its uncertainty has been forced to become idealistic and to drop the idea of a mechanical universe. New York Times, ???(??):120, January 11, 1931. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/99177884/13F390B7A1F783DA798.

Kaempffert:1938:BRE

[Kae38] Waldemar Kaempffert. Book review: Einstein Himself Explains: With Remarkable Clarity and Simplicity He and His Collaborator Trace the Ideas Behind Relativity. The Evolution of Physics. The Growth of Ideas Front Early Concepts to Relativity and Quanta. By Albert Einstein and Leopold Infeld. x + 319 pp.

New York: Simon & Schuster. $2.50. New York Times, ??(??): 86, April 10, 1938. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/102650329/.

Kaempffert:1939:WSW

[Kae39] Waldemar Kaempffert. This week in science: When uranium splits: Doubtful source of power; cancer and X-rays; neutron possibilities. New York Times, ??(??):D9, March 5, 1939. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/102937542.

Kaempffert:1945:SSB

[Kae45] Waldemar Kaempffert. Story of scientists’ ‘battle’ for atom bomb secret revealed in Smyth Report. New York Times, ??(??): 8, August 16, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107074385/.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kaempffert:2013:DAE

[Kae13a] Waldemar Kaempffert. Discovery of the antiproton ends a long search, confirms Einstein’s equation. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Kaempffert:2013:EEH

[Kae13b] Waldemar Kaempffert. Einstein expounds his new theory — lights all askew in the heavens — science seeks secret of life in star rays. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Kahn:1960:TW

[Kah60] Herman Kahn. On thermonuclear war. Princeton University Press, Princeton, NJ, USA, 1960. 651 pp. LCCN UF767 .K25 1960.

Kahn:1969:TWT

[Kah69] Herman Kahn. On thermonuclear war; [three lectures and several suggestions]. Free Press paperback. Free Press, New York, NY, USA, second edition, 1969. xxxvi + 668 pp. LCCN UF767 .K25 1969.

Kaiser:1981:RAM

[Kai81] Walter Kaiser. Rezensionen: Arthur I. Miller: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911). Reading, Massachusetts: Addison-Wesley 1981. XXVIII, 466 SS., Geb. $39,50; Kart. $27,50. Berichte zur Wissenschaftsgeschichte, 4(3–4):283–284, ???? 1981. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kaiser:1994:BHA

[Kai94] David Kaiser. Bringing the human actors back on stage: the personal context of the Einstein–Bohr debate. British Journal for the History of Science, 27(2):129–152, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4027432.

Kaiser:2002:EOF

[Kai02] David Kaiser. Einstein, Oppenheimer, Feynman: Physics in the 20th Century. Technical Report MIT OpenCourseWare STS.042J / 8.225J, MIT, Cambridge, MA, USA, 2002. Free, independent study course that explores the changing roles of physics and physicists during the 20th Century.

Kaiser:2005:BRB

[Kai05a] David Kaiser. Book review: Masters of Theory: Cambridge and the Rise of Mathematical Physics by Andrew Warwick. Journal of Interdisciplinary History, 35(4):644–645, ???? 2005. ISSN 0022-1953 (print), 1530-9169 (electronic). URL http://www.

jstor.org/stable/3656384.

Kaiser:2005:DTA

[Kai05b] David Kaiser. Drawing theories apart: the dispersion of Feynman diagrams in postwar physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4 K35 2005. URL http://www.

loc.gov/catdir/bios/uchi051/2004023335.html;   http:
//www.loc.gov/catdir/enhancements/fy0617/2004023335-
d.html;   http://www.loc.gov/catdir/toc/ecip051/2004023335.
html.

Kaku:2004:ECH

[Kak04] Michio Kaku. Einstein’s cosmos: how Albert Einstein’s vision transformed our understanding of space and time. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2004. ISBN 0-393-05165-X (hardcover). 251 pp. LCCN QC173.59.S65 K356 2004. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip0411/2003025580.html.

Kaku:2005:ECH

[Kak05] Michio Kaku. Einstein’s cosmos: how Albert Einstein’s vision transformed our understanding of space and time. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2005. ISBN 0-393-32700-0 (paperback). 268 pp. LCCN ???? URL http://

www.loc.gov/catdir/toc/ecip0411/2003025580.html.

Kalla:1935:AER

[Kal35] Elno Kalla. Albert Einsteins religion. Theoria, 1(1–2):58–67, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaloyerou:1992:WZD

[Kal92] P. N. Kaloyerou. On the Wootters–Zurek development of Einstein’s two-slit experiment. Foundations of Physics, 22(11): 1345–1377, November 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883665.

Kalai:2013:FHP

[Kal13] Ehud Kalai. Foreword: The high priest of game theory. American Mathematical Monthly, 120(5):384–385, May 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.05.384.pdf.

Kallstrand:2022:SNC

[Käl22] Gustav Källstrand. Science by Nobel committee: decision making and norms of scientific practice in the early physics and chemistry prizes. British Journal for the History of Science, 55(2):187–205, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.org/core/journals/british-
journal-for-the-history-of-science/article/science-
by-nobel-committee-decision-making-and-norms-of-scientific-
practice-in-the-early-physics-and-chemistry-prizes/
48B80DE55848512740AE7E05A5C77717.

Kanitscheider:1979:ETT

[Kan79] Bernulf Kanitscheider. Einstein’s treatment of theoretical concepts. In Aichelburg and Sexl [AS79], pages 137–158. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_10/.

Kaneko:1984:EVJ

[Kan84] Tsutomu Kaneko. Einstein’s view of Japan’s culture. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):51–76, September 1984. CODEN HIS-

CDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837320.

Kaneko:1987:EIJ

[Kan87] Tsutomu Kaneko. Einstein’s impact on Japanese intellectuals. In Glick [Gli87], chapter 10, pages 351–379. ISBN 90-277-2498-9. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.com/chapter/10.1007/
978-94-009-3875-5_10.

Kantha:1989:EZ

[Kan89] Sachi Sri Kantha. Einstein in Zurich. Nature, 337(6203): 110, January 12, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v337/n6203/pdf/337110b0.pdf.

Kanigel:1992:AEM

[Kan92a] R. Kanigel. Albert Einstein, Mileva Marić — the Love letters — J. Renn, R. Schulmann, Editors. New York Times Book Review, ??(??):15, July 19, 1992. ISSN 0028-7806.

Kantha:1992:AED

[Kan92b] S. S. Kantha. Albert Einstein dyslexia and the significance of Brodmann Area 39 of his left cerebral cortex. Medical Hypotheses, 37(2):119–122, February 1992. CODEN MEHYDY. ISSN 0306-9877 (print), 1532-2777 (electronic).

Kanigel:1994:PLA

[Kan94a] R. Kanigel. The private lives of Albert Einstein — R. Highfield, P. Carter. New York Times Book Review, ??(??):11–12, September 18, 1994. ISSN 0028-7806.

Kantha:1994:AAE

[Kan94b] S. S. Kantha. An appraisal of Albert Einstein chronic illness. Medical Hypotheses, 42(5):340–346, May 1994. CODEN MEHYDY. ISSN 0306-9877 (print), 1532-2777 (electronic).

Kane:1995:PGO

[Kan95a] G. L. Kane. The particle garden: our universe as understood by particle physicists. Helix books. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40780-9 (hardcover), 0-201-40826-0 (paperback). xvi + 224 pp. LCCN QC793.2 K365 1995; 539.72; 97.E04260; 528.

Kantha:1995:EL

[Kan95b] Sachi Sri Kantha. Einstein and Lorentz. Nature, 376(6536): 111, July 13, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v376/n6536/pdf/376111b0.pdf.

Kant:1996:AEM

[Kan96a] Horst Kant. Albert Einstein, Max von Laue, Peter Debye und das Kaiser-Wilhelm-Institut für Physik in Dahlem (1917–1939). (German) [Albert Einstein, Max von Laue, Peter Debye and Kaiser Wilhelm Institute of Physics in Dahlem (1917–1939)]. In vom Brocke and Laitko [vBL96a], pages 227–244. ISBN 3-11-015483-8 (vol. 1). LCCN

Q180.G42 K35 1996. URL http://www.degruyter.com/
viewbooktoc/product/146831;   http://www.reference-
global.com/doi/book/10.1515/9783110802443.

Kantha:1996:EMF

[Kan96b] S. Sri Kantha. Einstein’s medical friends and their influence on his life. Medical Hypotheses, 45(3):257–260, March 1996. CODEN MEHYDY. ISSN 0306-9877 (print), 1532-2777 (electronic).

Kantha:1996:ED

[Kan96c] Sachi Sri Kantha. An Einstein dictionary. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1996. ISBN 0-313-28350-8. 298 pp. LCCN QC16.E5 K36 1996.

Kanon:1997:AN

[Kan97] Joseph Kanon. Los Alamos: a novel. Broadway Books, New York, NY, USA, 1997. ISBN 0-553-06224-7. 403 pp. LCCN PS3561.A476 L6 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/bios/random057/96044055.html;   http://www.loc.
gov/catdir/description/random041/96044055.html;   http:
//www.loc.gov/catdir/samples/random041/96044055.html.

Kant:2002:WHG

[Kan02] Horst Kant. Werner Heisenberg and the German Uranium Project, volume 203 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wissenschaftsgeschichte, Berlin, Germany, 2002. 40 pp. LCCN ????

Kao:2010:NLS

[Kao10] Charles K. Kao. Nobel Lecture: Sand from centuries past: Send future voices fast. Reviews of Modern Physics, 82(3): 2299–2303, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2299;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2299_1.

Kapur:1940:DMO

[Kap40] P. L. Kapur. Does the mesotron obey Bose–Einstein or Fermi– Dirac statistics? Nature, 145(3663):69, January 13, 1940. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v145/n3663/
pdf/145069a0.pdf.

Kaplan:1961:BBP

[Kap61] Martin M. Kaplan. Books: Einstein on Peace, by Otto Nathan and Heinz Norden, with a preface by Bertrand Russell. Bulletin of the Atomic Scientists, 17(7):291–292, September 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitsa:1980:AE

[Kap80] P. L. Kapitsa. Albert Einstein. Voprosy Filosofii, ??(6):29–31, ???? 1980. ISSN 0042-8744.

Kapitsa:1981:AE

[Kap81] P. L. Kapitsa. Albert Einstein. Soviet Studies in Philosophy, 19 (3):3–10, ???? 1981. ISSN 0038-5883.

Kaplan:1991:WAn

[Kap91a] Fred M. Kaplan. The Wizards of Armageddon. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1991. ISBN 0-8047-1884-9 (paperback), 0-8047-9617-3 (e-book). 452 pp. LCCN U263 .K36 1991. URL http://catdir.loc.gov/

catdir/enhancements/fy1413/90071236-d.html.

Kaplan:1991:AEN

[Kap91b] L. Kaplan. Albert Einstein, a necrological approach. Centennial Review, 35(3):591–606, Fall 1991. ISSN 0162-0177.

Kaplan:2020:BPG

[Kap20] Fred M. Kaplan. The Bomb: Presidents, Generals, and the Secret History of Nuclear War. Simon and Schuster Paperbacks, New York, NY, USA, 2020. ISBN 1-982107-29-4, 1-982107-30-8 (paperback), 1-982107-31-6 (e-book). 372 + 8 pp. LCCN U264.3 .K37 2021.

Karl:2005:BRE

[Kar05] Gabriel Karl. Book reviews: Einstein Defiant: Genius Versus Genius in the Quantum Revolution, Edmund Blair Bolles, Joseph Henry Press, Washington, DC, 2004. $27.95 (348 pp.). ISBN 0-309-08998-0. Einstein’s Cosmos: How Albert Einstein’s Vision Transformed Our Understanding of Space and Time, Michio Kaku, Atlas Books / W. W. Norton, New York, 2004. $22.95 (251 pp.). ISBN 0-393-05165-X. Physics Today, 58(4):

59–60, April 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.1955481.

Karier:2010:ICF

[Kar10] Thomas Mark Karier. Intellectual capital: forty years of the Nobel Prize in economics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76326-6 (hardcover). xiii + 351 pp. LCCN HD53 .K37 2010.

Kasner:1920:NET

[Kas20] Edward Kasner. Note on Einstein’s theory of gravitation and light. Science, 52(1348):413–414, October 29, 1920. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/52/
1348/413.full.pdf.

Kasner:1921:ESF

[Kas21a] Edward Kasner. The Einstein solar field and space of six dimensions. Science, 53(1367):238–239, March 11, 1921. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/53/1367/238.
full.pdf.

Kasner:1921:ECE

[Kas21b] Edward Kasner. Einstein’s cosmological equations. Science, 54 (1396):304–305, September 30, 1921. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/54/1396/304.full.pdf.

Khariton:1996:WIW

[KAS96] Yuli Khariton, Viktor Adamskii, and Yuri Smirnov. The way it was. Bulletin of the Atomic Scientists, 52(9):54–59, November/ December 1996. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/

1G1-18855733.html. See comment [HM97].

KaSartaj:2015:SEB

[KaS15] Cartoon KaSartaj. The secrets of Einstein’s brain. Web 7minute video, August 18, 2015. URL https://www.youtube.

com/watch?v=wJOMtIL--Kw.

Katzin:1952:BRBb

[Kat52] Leonard I. Katzin. Book review: Nobel — The Man and His Prizes, edited by the Nobel Foundation, and written by H. Schück, R. Sohlman, A. Osterling, G. Liljestrand, A. Westgren, M. Siegbahn, A. Schou, and N. K. Stähl . Bulletin of the Atomic Scientists, 8(5):164, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1989:NLE

[Kat89] Bernard S. Katz, editor. Nobel laureates in economic sciences: a biographical dictionary, volume 850 of Garland reference library of the humanities. Garland, New York, NY, USA, 1989. ISBN 0-8240-5742-2. xvii + 339 pp. LCCN HB76.N63 1989.

Katsumori:1992:TRE

[Kat92] Makoto Katsumori. The theories of Relativity and Einstein’s philosophical turn. Studies in History and Philosophy of Science Part A, 23(4):557–592, December 1992. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290013V.

Katz:2004:BRZ

[Kat04] Shaul Katz. Berlin roots — Zionist incarnation: The ethos of pure mathematics and the beginnings of the Einstein Institute of Mathematics at the Hebrew University of Jerusalem. Science in Context, 17(1–2):199–234, May 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Katzir:2005:PRP

[Kat05] Shaul Katzir. Poincaré’s relativistic physics: Its origins and nature. Physics in Perspective (PIP), 7(3):268–292, Septem-

ber 2005. URL http://www.springerlink.com/content/
w40237nj121084w3/.

Kaufherr:1985:NAE

[Kau85] Tirzah Kaufherr. A new approach to the Einstein–Podolsky– Rosen paradox. Foundations of Physics, 15(10):1043–1051, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00732846.

Kauffman:2014:BRU

[Kau14] George B. Kauffman. Book review: Ulf Lagerkvist: Erling Norrby (ed.): The periodic table and a missed Nobel Prize. Foundations of Chemistry, 16(3):249–251, October 2014. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

URL http://link.springer.com/article/10.1007/s10698-
014-9205-9.

Kawaki:1934:BE

[Kaw34] Ayao Kawaki. A Biography of Einstein. Kaizosha, Tokyo, Japan, 1934. ???? pp. LCCN ????

Kayser:1946:SPS

[Kay46] Rudolf Kayser. Spinoza, portrait of a spiritual hero. Philosophical Library, New York, NY, USA, 1946. xix + 11 + 326 pp. LCCN B3997 .K33. With an introduction by Albert Einstein. Translated by Amy Allen and Maxim Newmark.

Kayser:1968:SPS

[Kay68] Rudolf Kayser. Spinoza, portrait of a spiritual hero. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1968. xix + 326 pp. LCCN B3997 .K33 1968. With an introduction by Albert Einstein. Translated by Amy Allen and Maxim Newmark. Reprint of 1946 edition.

Knebel:1960:NHG

[KB60] Fletcher Knebel and Charles W. (Charles Waldo) Bailey III. No high ground. Harper, New York, 1960. 272 pp. LCCN D767.25.H6 K55.

Knebel:1983:NHG

[KB83] Fletcher Knebel and Charles W. (Charles Waldo) Bailey III. No high ground. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1983. ISBN 0-313-24221-6. ix + 272 pp. LCCN D767.25.H6 K55 1983.

Kerber:1987:ESD

[KDK87] Gabriele Kerber, Auguste Dick, and Wolfgang Kerber. Erwin Schrödinger 1887–1961: Dokumente, Materialien und Bilder. Österreichischen Zentralbibliothek für Physik, Vienna, Austria, 1987. ISBN 3-900538-09-3. 158 pp. LCCN QC16.S265 D65 1987. An exhibition of the Österreichischen Zentralbibliothek für Physik. With contributions by Nándor Balázs and others.

Kox:2005:UGR

[KE05] Anne J. Kox and Jean Eisenstaedt, editors. The universe of General Relativity, volume 11 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2005. ISBN 0-8176-4380-X, 0-8176-4454-7 (e-book). x + 383 pp. LCCN QC173.5 .U55 2005. URL http://www.loc.

gov/catdir/enhancements/fy0663/2005047817-d.html;
http://www.loc.gov/catdir/toc/fy0602/2005047817.html.

Kean:2012:VTO

[Kea12] Sam Kean. The Violinist’s Thumb: and Other Lost Tales of Love, War, and Genius, as Written by Our Genetic Code. Little, Brown and Co., Boston, MA, USA, 2012. ISBN 0-316-18231-1 (hardcover). ix + 403 pp. LCCN QH431 .K24 2012.

Keating:2018:LNP

[Kea18] Brian (Brian Gregory) Keating. Losing the Nobel Prize: a story of cosmology, ambition, and the perils of science’s highest honor. W. W. Norton & Co., New York, NY, USA, 2018. ISBN 1-324-00091-0 (hardcover). xxi + 326 pp. LCCN QB991.B54 K43 2018.

Keeney:2004:ANL

[Kee04] John Keeney, editor. Australia’s Nobel Laureates: adventures in innovation. ETN communications for ABIE, Roseville, NSW, Australia, 2004. ISBN 0-9580207-2-8. 310 pp. LCCN AS911.N9 A89 2004.

Keel:2005:SEF

[Kee05] W. Keel. The sky at Einstein’s feet. Springer Praxis books in astrophysics and astronomy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-26130-3. xiii + 246 pp. LCCN QC173.57 .K44 2006.

Keeney:2011:MGC

[Kee11] L. Douglas Keeney. 15 minutes: General Curtis LeMay and the countdown to nuclear annihilation. St. Martin’s Press, New York, NY, USA, 2011. ISBN 0-312-61156-0. x + 372 + 16 pp. LCCN U264.3 .K43 2011.

Keller:1958:CBS

[Kel58] Joseph B. Keller. Corrected Bohr–Sommerfeld quantum conditions for nonseparable systems. Annals of Physics, 4:180– 188, 1958. CODEN APNYA6. ISSN 0003-4916 (print),

1096-035X (electronic). URL http://www.maths.ed.ac.uk/
~aar/papers/keller1.pdf;   http://www.sciencedirect.
com/science/article/pii/0003491658900320.

Keller:1985:SM

[Kel85] Joseph B. Keller. Semiclassical mechanics. SIAM Review, 27(4): 485–504, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:2004:RMP

[Kel04] Cynthia C. Kelly, editor. Remembering the Manhattan Project: Perspectives on the Making of the Atomic Bomb and Its Legacy. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Kelly:2006:OMP

[Kel06a] Cynthia C. Kelly, editor. Oppenheimer and the Manhattan Project: insights into J. Robert Oppenheimer, “Father of the atomic bomb”. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-418-7 (hardcover), 981-256-599-X (paperback). vii + 173 pp. LCCN QC16.O62 O66 2006. URL http://www.loc.

gov/catdir/toc/fy0702/2006283849.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5897.

Kelly:2006:TMS

[Kel06b] Kate Kelly. That’s not in my science book: a compilation of little-known facts. Taylor Trade Pub., Lanham, UK, 2006. ISBN 1-58979-290-4 (paperback). xi + 226 pp. LCCN Q162 .K44 2006.

Kelly:2007:MPB

[Kel07] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb by its creators, eyewitnesses, and historians. Black Dog and Leventhal Publishers, New York, NY, USA, 2007. ISBN 1-57912-747-9, 1-57912-808-4 (paperback), 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U5

M27 2007. URL http://www.loc.gov/catdir/toc/ecip0720/
2007023984.html.

Kelly:2009:MPB

[Kel09] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb in the words of its creators, eyewitnesses, and historians. Black Dog and Leventhal, New York, NY, USA, 2009. ISBN 1-57912-808-4 (paperback), 1-57912-747-9, 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U6 M27 2009. US$10.95.

Kelly:2012:GMP

[Kel12] Cynthia C. Kelly. A guide to the Manhattan Project sites in Manhattan. Atomic Heritage Foundation, Washington, DC, USA, 2012. ISBN 0-9859448-0-3. 64 pp. LCCN ????

Keller:2014:LPP

[Kel14] Ole Keller. Light: the physics of the photon. Series in optics and optoelectronics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4398-4043-1 (hardcover). xix + 462 pp. LCCN QC793.5.P42 .K45

2014. URL http://www.loc.gov/catdir/enhancements/
fy1413/2013362573-b.html.

Kemp:2012:EMH

[Kem12] Scott Kemp. The end of Manhattan: How the gas centrifuge changed the quest for nuclear weapons. Technology and Culture, 53(2):272–305, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL http://muse.jhu.edu/

journals/technology_and_culture/v053/53.2.kemp.pdf.

Kennard:1925:RSR

[Ken25] E. H. Kennard. Review: Sidelights on Relativity, by A. Einstein, G. B. Jeffery, and W. Perrett. The Philosophical Review, 34 (2):204–205, March 1925. URL 2-M; http://links.jstor.

org/sici?sici=0031-8108(192503)34:2<204:SOR>2.0.CO.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of Hbomb at Pacific Proving Ground. New York Times, ??(??): E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/113104238/.

Kennedy:2003:STE

[Ken03] J. B. (John Bernard) Kennedy. Space, time and Einstein: an introduction. Acumen, Chesham, UK, 2003. ISBN 1-902683-66-8 (hardcover), 1-902683-67-6 (paperback). x + 244 pp.

Kennedy:2005:EEP

[Ken05a] W. L. Kennedy. On Einstein’s 1905 electrodynamics paper. Studies in History and Philosophy of Modern Physics, 36(1): 61–65, March 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219804000784.

Kennefick:2005:EVP

[Ken05b] Daniel Kennefick. Einstein versus the Physical Review. Physics Today, 58(9):43–48, September 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/physicstoday.org/resource/1/phtoad/v58/i9/p43_s1.

Kent:2005:EOC

[Ken05c] Paul W. Kent. Einstein in Oxford: celebrating the centenary of the 1905 publications. Report, Department of Physics, University of Oxford, Oxford, UK, 2005. 19 pp.

Kennefick:2007:TST

[Ken07] Daniel Kennefick. Traveling at the speed of thought: Einstein and the quest for gravitational waves. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-11727-6 (hardcover). xii + 319 pp. LCCN QC179 .K46 2007. URL http://

www.loc.gov/catdir/enhancements/fy0834/2006938366-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2006938366-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2006938366-t.html.

Kennefick:2011:RBU

[Ken11] Daniel Kennefick. Review of Einstein’s Unification by Jeroen van Dongen. Foundations of Physics, 41(2):278–280, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9532-1.

Kennedy:2012:SGE

[Ken12] Robert E. Kennedy. A student’s guide to Einstein’s major papers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-969403-6 (hardcover). xxi + 303 pp. LCCN QC16.E5 K466 2012.

Kent:2013:MQI

[Ken13] Adrian Kent. Might quantum-induced deviations from the Einstein equations detectably affect gravitational wave propagation? Foundations of Physics, 43(6):707–718, June 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
013-9716-6.

Kennefick:2014:EGW

[Ken14] Daniel J. Kennefick. Einstein, gravitational waves, and the theoretician’s regress. In Janssen and Lehner [JL14b], chapter 8, pages 270–280. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Kennefick:2019:NSD

[Ken19] Daniel Kennefick. No Shadow of a Doubt: the 1919 Eclipse That Confirmed Einstein’s Theory of Relativity. Princeton University Press, Princeton, NJ, USA, 2019. ISBN 0-691-18386-4. viii + 403 pp. LCCN ????

Kent:2022:EEE

[Ken22] Deborah Kent. Einstein, Eddington, e o/and the eclipse: Impressões de viagem/travel impressions. British Journal for the History of Mathematics, 37(2):162–163, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/26375451.2022.
2056968.

Kerszberg:1987:RRE

[Ker87] Pierre Kerszberg. The relativity of rotation in the early foundations of General Relativity. Studies in History and Philosophy of Science Part A, 18(1):53–79, March 1987. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900112.

Kerker:1989:CCC

[Ker89a] M. Kerker. Classics and classicists of colloid and interface science. 8 Albert Einstein. Journal of Colloid and Interface Science, 129(1):291–295, April 1989. CODEN JCISA5. ISSN 0021-9797 (print), 1095-7103 (electronic).

Kerszberg:1989:BRC

[Ker89b] P. Kerszberg. Book review: The Collected Papers of Albert Einstein, Vol 1, the Early Years, 1879–1902 J. Stachel. Annals of Science, 46(1):98–99, January 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Kerszberg:1989:IUE

[Ker89c] Pierre Kerszberg. The invented universe: the Einstein–de Sitter controversy (1916–17) and the rise of relativistic cosmology. Oxford science publications. Clarendon Press, Oxford, UK, 1989. ISBN 0-19-851876-5. ix + 403 pp. LCCN QB981 .K47 1989.

US$37.50. URL http://www.loc.gov/catdir/enhancements/
fy0638/88038563-d.html;   http://www.loc.gov/catdir/
enhancements/fy0638/88038563-t.html.

Keswani:1965:OCRa

[Kes65a] G. H. Keswani. Origin and concept of Relativity (I). The British Journal for the Philosophy of Science, XV (60):286–306, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/content/XV/60/286.full.pdf+
html;   http://links.jstor.org/sici?sici=0007-0882%
28196502%2915%3A60%3C286%3AOACOR%28%3E2.0.CO%3B2-F;
 http://www.jstor.org/stable/686536. See note [Din65].

Keswani:1965:OCRb

[Kes65b] G. H. Keswani. Origin and concept of Relativity (II). The British Journal for the Philosophy of Science, XVI(61):19–32, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/XVI/61/19.full.pdf+html;   http://www.jstor.
org/stable/686136.

Ketterle:1999:ESB

[Ket99] Wolfgang Ketterle. Experimental studies of Bose–Einstein condensation. Physics Today, 52(12):30–35, December 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.882898.

Ketterle:2002:NLW

[Ket02] Wolfgang Ketterle. Nobel Lecture: When atoms behave as waves: Bose–Einstein condensation and the atom laser. Reviews of Modern Physics, 74(4):1131–1151, October 2002. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.74.1131;   http://rmp.aps.org/abstract/RMP/
v74/i4/p1131_1.

Kevles:2005:BRC

[Kev05] D. J. Kevles. Book review: The collected papers of Albert Einstein, vol 9, The Berlin years correspondence, January 1919– April 1920. TLS — The Times Literary Supplement, ??(5328): 3–4, May 13, 2005. ISSN 0307-661X.

Keynes:1933:EB

[Key33] John Maynard Keynes. Essays in Biography. Macmillan Publishing Company, New York, NY, USA, 1933. x + 318 pp. LCCN DA574.A1 K4.

Keynes:1989:CWJ

[Key89a] John Maynard Keynes. The Collected Writings of John Maynard Keynes: Volume X: Essays in Biography. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-333-10721-7. xix + 460 pp. LCCN HB171 .K44 v.10. With a new introduction by Donald Winch.

Keynes:1989:E

[Key89b] John Maynard Keynes. Einstein. In The Collected Writings of John Maynard Keynes: Volume X: Essays in Biography [Key89a], chapter 37, pages 382–388. ISBN 0-333-10721-7. LCCN HB171 .K44 v.10. With a new introduction by Donald Winch.

Keynes:2013:EB

[Key13] John Maynard Keynes. Essays in biography, volume 10 of The collected writings of John Maynard Keynes. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-65643-5 (paperback). xlv + 460 pp. LCCN HB171. With a new introduction by Donald Winch.

Kouznetsov:1974:CEB

[KF74] B. Kouznetsov and A. Frenk. La correspondance Einstein– Besso. Revue d’Histoire des Sciences, 27(1):77–82, January 1974. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632081.

Ketcham:2003:VBE

[KF03] Peter M. Ketcham and David L. Feder. Visualizing Bose– Einstein condensates. Computing in Science and Engineering, 5(1):86–89, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://csdl.

computer.org/dl/mags/cs/2003/01/c1086.htm;   http://
csdl.computer.org/dl/mags/cs/2003/01/c1086.pdf.

Kirby:2005:AEM

[KH05] K. Kirby and F. Houle. Albert Einstein to Michele Besso — reply. Physics Today, 58(7):16, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Khalatnikov:2012:ABL

[Kha12] Isaak M. Khalatnikov. From the atomic bomb to the Landau Institute: autobiography. Top non-secret. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-27560-5 (hard cover). xii + 214 pp. LCCN QC16.K43 K43 2012.

Khan:2020:TMN

[Kha20] Sieeka Khan. Top 10 most notable science anniversaries in 2020. Web site., January 8, 2020. URL https://www.sciencetimes.

com/articles/24639/20200107/science.htm.

Khrennikov:2008:RNL

[Khr08] Andrei Khrennikov. The role of von Neumann and Lüders postulates in the Einstein, Podolsky, and Rosen considerations: Comparing measurements with degenerate and nondegenerate spectra. Journal of Mathematical Physics, 49(5):052102, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v49/i5/p052102_s1.

Kiang:1946:DLG

[Kia46] Su-Ching Kiang. Deflexion of light in the gravitational field without using Einstein geometry. Nature, 157(3999):842, June 22, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v157/n3999/pdf/157842c0.pdf.

Kiefer:1982:EEW

[Kie82] Klaus H. Kiefer. Einstein & Einstein: Wechselseitige Erhellung der Künste und Wissenschaften um 1915. (German) [Albert] Einstein and [Carl] Einstein, reciprocal influences between arts and sciences in 1915. Komparatistische Hefte, 5–6:181–194, 1982. ISSN 0172-9209. See later skepticism about this article [Kie94] [from remark in [van12, page 176, reference 130]].

Kiefer:1994:DIW

[Kie94] Klaus H. Kiefer. Diskurswandel im Werk Carl Einsteins: ein Beitrag zur Theorie und Geschichte der europäischen Avantgarde. (German) [Discourse of change in the work of Carl Einstein: a contribution to the theory and history of European avant-garde], volume 7 of Communicatio. Niemeyer, Tübingen, Germany, 1994. ISBN 3-484-63007-8. ISSN 0941-1704. vii + 631 pp. LCCN PT2609.I6 Z74 1994.

Kiernan:2013:GAC

[Kie13] Denise Kiernan. The girls of Atomic City: the untold story of the women who helped win World War II. Simon and Schuster, New York, NY, USA, 2013. ISBN 1-4516-1752-6 (hardcover), 1-4516-1754-2 (e-book). xvii + 373 + 16 pp. LCCN F444.O3 K54 2013.

Kiefer:2015:AEB

[Kie15] Claus Kiefer. Albert Einstein, Boris Podolsky, Nathan Rosen: Kann die quantenmechanische Beschreibung der physikalischen Realität als vollständig betrachtet werden?. (German) [Albert Einstein, Boris Podolsky, Nathan Rosen: can the quantum mechanical description of physical reality be considered complete?]. Klassische Texte der Wissenschaft. Springer Spektrum, Berlin, Germany, 2015. ISBN 3-642-41998-4, 3-642-41999-2. 125 pp. LCCN ????

Kifner:1995:HCR

[Kif95] John Kifner. Hiroshima: The controversy that refuses to die. New York Times, ??(??):A16, January 31 (or 13??), 1995. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Kilmister:1965:RBW

[Kil65] C. W. Kilmister. Review: Briefe zur Wellenmechanik, by von Schrödinger, Planck, Einstein, and Lorentz, edited by K. Przibram. Mathematical Gazette, 49(369):346–347, October 1965.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL http://www.jstor.org/stable/3612915.

Kiley:1969:EAR

[Kil70] John F. Kiley. Einstein and Aquinas: a rapprochement. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1969–1970. xli + 124 pp. LCCN QC6 .K437.

Kilmister:1980:ECA

[Kil80] C. W. Kilmister. Elie Cartan and Albert Einstein — letters on absolute parallelism 1929–1932 — R. Debever. Annals of Science, 37(2):247, ???? 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Kilmister:1984:AEH

[Kil84a] C. W. Kilmister. Albert Einstein — historical and cultural perspective — G. Holton, Y. Elkana. British Journal for the History of Science, 17(56):227, ???? 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Kilmister:1984:BRPb

[Kil84b] C. W. Kilmister. Book review: Physics and its concepts: Gerald Holton and Yehuda Elkana (ed), Albert Einstein: historical and cultural perspective. Princeton: University Press, 1982. Pp. xxxii + 439. ISBN 0-691-08299-5. £24.70. British Journal for the History of Science, 17(2):227, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026561.

Kilmister:1987:SCC

[Kil87] C. W. (Clive William) Kilmister, editor. Schrödinger: centenary celebration of a polymath. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://www.loc.gov/

catdir/enhancements/fy0901/86028338-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0901/86028338-t.

html. Papers presented at a conference, March 31–April 3, 1987, at Imperial College, London.

Kilmister:1989:SCC

[Kil89] C. W. (Clive William) Kilmister, editor. Schrödinger, centenary celebration of a polymath. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37929-6 (paperback), 0-521-34017-9. LCCN QC16.S265 S36 1989.

Kilmister:1994:ESF

[Kil94] C. W. (Clive William) Kilmister. Eddington’s search for a fundamental theory: a key to the universe. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-37165-1 (hardcover). xii + 256 pp. LCCN QC6 .K439 1994. URL http://www.loc.

gov/catdir/description/cam026/94009035.html;   http:
//www.loc.gov/catdir/toc/cam021/94009035.html.

Kilmister:1995:BRC

[Kil95] C. W. Kilmister. Book review: The Collected Papers of Albert Einstein, Vol 3, the Swiss Years — Writings, 1909–11 — M. J. Klein, Anne J. Kox, J. Renn, R. Schulmann. Annals of Science, 52(2):198–199, March 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Kilmister:1998:BRC

[Kil98] C. W. Kilmister. Book review: The collected papers of Albert Einstein, vol 6, The Berlin years: 1914–17. Annals of Science, 55(3):320–321, July 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

King:1995:EAL

[Kin95] B. Tom King. An Einstein addition law for nonparallel boosts using the geometric algebra of space–time. Foundations of Physics, 25(12):1741–1755, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF02057886.

King:2011:ET

[Kin11] Ritchie S. King. The Einstein telescope. IEEE Spectrum, 48 (7):14, July 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kirkpatrick:1980:CPE

[Kir80] Paul Kirkpatrick. Confirming the Planck–Einstein equation hν = (1/2)mv^{2}. American Journal of Physics, 48(10):803– 806, 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/48/10/10.1119/1.12228.

Kirshner:2002:EUE

[Kir02] Robert P. Kirshner. The Extravagant Universe: Exploding Stars, Dark Energy, and the Accelerating Cosmos. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05862-

8. xii + 282 pp. LCCN QB843.S95 K57 2002. URL http://www.

loc.gov/catdir/description/prin022/2002029268.html;
 http://www.loc.gov/catdir/samples/prin031/2002029268.
html;   http://www.loc.gov/catdir/toc/prin031/2002029268.
html.

Kirsch:2005:PGP

[Kir05] Scott Kirsch. Proving grounds: Project Plowshare and the unrealized dream of nuclear earthmoving. Rutgers University Press, New Brunswick, NJ, USA, 2005. ISBN 0-8135-3666-9 (hardcover). xi + 257 pp. LCCN TA748 .K57 2005. URL http://

www.loc.gov/catdir/toc/ecip057/2005002579.html.

Kirshner:2013:AUN

[Kir13] Robert P. Kirshner. The accelerating universe: A Nobel surprise. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 157(4):438– 456, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640226.

Kistiakowsky:1980:TR

[Kis80] George B. Kistiakowsky. Trinity — a reminiscence. Bulletin of the Atomic Scientists, 36(6):19–22, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1982:FA

[Kis82] George B. Kistiakowsky. The four anniversaries. Bulletin of the Atomic Scientists, 38(10):2–3, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kitfield:2000:DNS

[Kit00] James A. Kitfield. The decline of the nuclear stockpile. Air Force magazine, 82(2):54–57, February 2000. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782

(electronic).  URL http://www.airforce-magazine.com/
MagazineArchive/Pages/2000/February%202000/0200nuclear.
aspx.

Kahn:2007:TW

[KJ07] Herman Kahn and Evan Jones. On thermonuclear war. Transaction Publishers, New Brunswick, NJ, USA, transaction edition, 2007. ISBN 1-4128-0664-X. xxi + 668 pp. LCCN U263

.K324 2007. URL http://www.loc.gov/catdir/toc/fy0802/
2006050043.html.

Ko:2024:UPR

[KJY24] Byoung-Kwon Ko, Yeongkyun Jang, and Jae-Suk Yang. Universalism and particularism in the recommendations of the Nobel Prize for science. Scientometrics, 129(2):847–868, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-023-04921-0.

Kahn:1975:LES

[KK75] Carla Kahn and Franz Kahn. Letters from Einstein to de Sitter on the nature of the Universe. Nature, 257(5526):451– 454, October 9, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v257/n5526/pdf/257451a0.pdf.

Keswani:1983:IRR

[KK83] G. H. Keswani and C. W. Kilmister. Intimations of Relativity: Relativity before Einstein. The British Journal for the Philosophy of Science, 34(4):343–354, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/34/4/343.
full.pdf+html.

Kaplan:1989:PAE

[KK89] Rita J. Kaplan and Stanley H. Kaplan. Problems at Albert Einstein. Nature, 339(6219):10, May 4, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v339/n6219/
pdf/339010c0.pdf.

Krauss:2005:RHC

[KK05] Robert Krauss and Amelia Krauss, editors. The 509th remembered: a history of the 509th Composite Group as told by the veterans themselves, 509th anniversary reunion, Wichita, Kansas October 7–10, 2004. 509th Press, Buchanan, MI, USA, second edition, 2005. ISBN 0-923568-66-2. 346 pp. LCCN D790.253 509th .A16 2005. URL http:/

/www.enolagay509th.com;   http://www.loc.gov/catdir/
toc/ecip0510/2005010355.html.

Krull:2009:AE

[KK09] Kathleen Krull and Boris Kulikov. Albert Einstein. Giants of science. Viking, New York, NY, USA, 2009. ISBN 0-670-06332-0 (hardcover). 141 pp. LCCN QC16.E5 K75 2009.

Kar:2006:SRC

[KKC06] Guruprasad Kar, Samir Kunkri, and Sujit K. Choudhary. Special Relativity, causality and quantum mechanics — 1. Resonance, 11(8):41–52, November 2006. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.

springerlink.com/content/07m702671008720l/.

Kademani:1999:SPN

[KKJ99] B. S. Kademani, V. L. Kalyane, and Suresh Jange. Scientometric portrait of Nobel laureate Dorothy Crowfoot Hodgkin. Scientometrics, 45(2):233–250, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02458435.

Kademani:2005:NLT

[KKKM05] B. S. Kademani, V. L. Kalyane, Vijai Kumar, and Lalit Mohan. Nobel laureates: Their publication productivity, collaboration and authorship status. Scientometrics, 62(2):261–268, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-005-0019-3.

Klein:1993:SYW

[KKRS93] Martin J. Klein, Anne J. Kox, Jürgen Renn, and Robert Schulmann, editors. The Swiss Years: Writings, 1909–1911, volume 3 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-08772-5. xxxvi + 644 pp.

Klein:1995:SYW

[KKRS95] Martin J. Klein, Anne J. Kox, Jürgen Renn, and Robert Schulmann. The Swiss Years: Writings, 1912–1914, volume 4 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-02610-

6. xii + 314 pp. LCCN QC16.E5A2 1995. URL http://

einsteinpapers.press.princeton.edu/vol4-doc;   http:
//einsteinpapers.press.princeton.edu/vol4-trans;   http:
//press.princeton.edu/titles/5603.html;   http://press.
princeton.edu/titles/5837.html;   http://press.princeton.
edu/TOCs/c5603.html;   http://press.princeton.edu/TOCs/

c5837.html. The Swiss years: writings, 1912–1914, English translation by Anna Beck in consultation with Don Howard, with a preface by Beck and Howard.

Klein:1993:SYC

[KKS93] Martin J. Klein, Anne J. Kox, and Robert Schulmann, editors. The Swiss years: correspondence, 1902–1914, volume 5 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-03322-6. xlxx + 724 pp. LCCN QC16.E5 A2 1987. URL http://

einsteinpapers.press.princeton.edu/vol5-doc;   http:
//press.princeton.edu/titles/5340.html.

Kox:1996:BYW

[KKS96] Anne J. Kox, Martin J. Klein, and Robert Schulmann. The Berlin years: writings, 1914–1917, volume 6 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01086-

2. xxv + 626 pp. LCCN QC 7 084. URL http://

einsteinpapers.press.princeton.edu/vol6-doc;   http:
//einsteinpapers.press.princeton.edu/vol6-trans;   http:
//press.princeton.edu/titles/5836.html;   http://press.
princeton.edu/titles/6161.html;   http://press.princeton.
edu/TOCs/c5836.html.

Klein:1949:TEF

[Kle49] O. Klein. On the thermodynamical equilibrium of fluids in gravitational fields. Reviews of Modern Physics, 21(3):531– 533, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.531;   http://rmp.aps.
org/abstract/RMP/v21/i3/p531_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Klein:1959:PEC

[Kle59] Martin J. Klein, editor. Paul Ehrenfest: Collected scientific papers. North-Holland, Amsterdam, The Netherlands, 1959. xii + 632 pp. LCCN QC3 .E4. With an introduction by H. B. G. Casimir.

Klein:1963:EFP

[Kle63] Martin J. Klein. Einstein’s first paper on quanta. The Natural Philosopher, 2:59–86, 1963. CODEN NPHLAC. ISSN 0547-9592.

Klein:1964:EWP

[Kle64] Martin J. Klein. Einstein and the wave–particle duality. The Natural Philosopher, 3:27–??, 1964. CODEN NPHLAC. ISSN 0547-9592.

Klein:1965:ESC

[Kle65a] Martin J. Klein. Einstein and some civilized discontents. Physics Today, 18(1):38–44, January 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Klein:1965:ESH

[Kle65b] Martin J. Klein. Einstein, specific heats, and the early quantum theory. Science, 148(3667):173–180, April 9, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/148/
3667/173.extract.

Klein:1967:TET

[Kle67] Martin J. Klein. Thermodynamics in Einstein’s thought. Science, 157(3788):509–516, August 4, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/157/3788/509.full.pdf.

Klein:1970:BRBb

[Kle70a] Martin J. Klein. Book review: Briefwechsel, 1916–1955. Albert Einstein and Hedwig and Max Born. Commentary by Max Born. Nymphenburger, Munich, 1969. 332 pp. DM 24.80. Science, 169(3943):360–361, July 24, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/169/3943/360.3.full.pdf.

Klein:1970:BRBa

[Kle70b] Martin J. Klein. Book review: Briefwechsel, 1916-1955. Albert Einstein and Hedwig and Max Born. Commentary by Max Born. Nymphenburger, Munich, 1969. 332 pp. DM 24.80. Science, 169(3943):360–361, July 24, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/169/3943/360.3.full.pdf.

Klein:1970:FPB

[Kle70c] Martin J. Klein. The first phase of the Bohr–Einstein dialogue. Historical Studies in the Physical Sciences, 2(??):iv + 1–39, ???? 1970. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757302.

Klein:1970:PE

[Kle70d] Martin J. Klein. Paul Ehrenfest. North-Holland, Amsterdam, The Netherlands, 1970. ISBN 0-7204-0163-1 (v. 1). xvi + 330 + 5 pp. LCCN QC16.E45 K58.

Klein:1973:BRB

[Kle73] Martin J. Klein. Book review: Correspondance 1903–1955 by Albert Einstein; Michele Besso; Pierre Speziali. Isis, 64(3):435– 436, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
229767.

Klein:1974:PEE

[Kle74] Oskar Klein. Le principe d’équivalence d’Einstein utilisé pour une alternative de la cosmologie relativiste en regardant le système des galaxies comme limité et non comme l’univers. (French) [Einstein’s Principle of Equivalence used for an alternative of relativistic cosmology by looking at the system of galaxies as limited and not as the universe]. Kongelige Danske Videnskabernes Selskab: Matematisk-fysiske meddelelser, 39(2): 18, 1974.

Klein:1975:BAE

[Kle75a] Martin J. Klein. “Albert Einstein: The Education of a Genius” produced, written, and directed by Harold Mantell. American Journal of Physics, 43(5):473, May 1975. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/43/473/1.

Kleinert:1975:ALA

[Kle75b] Andreas Kleinert. Anton Lampa und Albert Einstein. Die Neubesetzung der physikalischen Lehrstuhle an der deutschen Universität Prag 1909 und 1910. (German) [Anton Lampa and Albert Einstein. The call to the Chair in Physics at the German University at Prague 1909 and 1910]. Gesnerus: Swiss Journal of the history of medicine and sciences, 32(3–4):284–292, ????

1975. ISSN 0016-9161. URL https://www.e-periodica.ch/
digbib/view?pid=ges-001:1975:32::314.

Klein:1980:NFF

[Kle80] Martin J. Klein. No firm foundation: Einstein and the early quantum theory. In Woolf [Woo80b], pages 161–185. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Klein:1981:DAC

[Kle81] Martin J. Klein. Not by discoveries alone: The centennial of Paul Ehrenfest. Physica A, 106(1–2):3–14, 1981. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378437181902016.

Klein:1982:AEF

[Kle82a] M. J. Klein. Albert Einstein — four commemorative lectures — Loyd S. Swenson, Jr., C. P. Snow, Howard Stein, and Ilya Prigogine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):344–345, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klein:1982:FSP

[Kle82b] Martin J. Klein. Fluctuations and statistical physics in Einstein’s early work. In Holton and Elkana [HE82], pages 39–58. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Klein:1984:BRA

[Kle84] Martin J. Klein. Book review: Abraham Pais: “Subtle Is the Lord … ”: The Science and the Life of Albert Einstein. Isis, 75(2):377–379, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231838.

Kleinert:1985:NES

[Kle85] Andreas Kleinert. Noch einmal:Sommerfeld und Einstein. (German) [Once again: Sommerfeld and Einstein]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 69(1):96–97, ???? 1985. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.

org/stable/20776958.

Kleinert:1993:PWB

[Kle93] Andreas Kleinert. Paul Weyland, der Berliner Einstein-Töter. (German) [Paul Weyland, the Berlin Einstein-killer]. In Albrecht Helmuth, editor, Naturwissenschaft und Technik in der Geschichte. 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaften und Technik am Historischen Institut der Universität Stuttgart, pages 198–232. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. LCCN ????

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres … (French) [The seven-times revolution: Albert Einstein and the others …]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Klein:2010:PEN

[Kle10] Martin J. Klein. Paul Ehrenfest, Niels Bohr, and Albert Einstein: Colleagues and friends. Physics in Perspective (PIP), 12(3):307–337, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0025-6;
 http://www.springerlink.com/content/g08710644434v387/
fulltext.pdf.

Kleppner:2014:BRE

[Kle14a] Daniel Kleppner. Book review: Einstein and the Quantum: The Quest of the Valiant Swabian, A. Douglas Stone, Princeton U. Press, 2013. $29.95 (332 pp.). ISBN 978-0-691-13968-5. Physics Today, 67(4):48, April 1, 2014. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/PT.3.2349.

Klevgard:2014:EFE

[Kle14b] Paul A. Klevgard. Einstein and the formal equivalence of mass and energy. arxiv.org, December 1, 2014. URL http://arxiv.

org/pdf/1412.2060.pdf.

Klein:2016:PQA

[Kle16] Étienne Klein. Le pays qu’habitait Albert Einstein. (French) [The country that Albert Einstein lived in]. Actes Sud, Arles, France, 2016. ISBN 2-330-06662-7. 243 pp. LCCN ????

Klein:2017:BEI

[Kle17a] Étienne Klein. La bicicletta di Einstein. (Italian) [Einstein’s bicycle], volume 55. Ponte alle Grazie, Milano, Italy, 2017. ISBN 88-6833-693-6. 225 pp.

Kleinknecht:2017:EHB

[Kle17b] Konrad Kleinknecht. Einstein und Heisenberg: Begründer der modernen Physik. (German) [Einstein and Heisenberg: Founders of Modern Physics]. Verlag W. Kohlhammer, Stuttgart, Germany, 2017. ISBN 3-17-032385-7 (paperback), 3-17-032388-1 (e-book). 221 pp. LCCN QC15 .K54 2017.

Klickstein:1962:AEA

[Kli62a] H. S. Klickstein. Albert Einstein and the Albert medical center. The Journal of the Albert Einstein Medical Center, Philadelphia, 10(??):150–154, July 1962. ISSN 0002-4708.

Klickstein:1962:CRB

[Kli62b] H. S. Klickstein. A cumulative review of bibliographies of the published writings by Albert Einstein. The Journal of the Albert Einstein Medical Center, Philadelphia, 10(??):141–149, July 1962. ISSN 0002-4708.

Klose:1963:QBR

[Klo63] W. Klose. Quantentheorie: [book review:] E. Schrödinger, M. Planck, A. Einstein, and H. A. Lorentz, Briefe zur Wellenmechanik, Herausgegeben von K. Przibram. Wien: Springer-Verlag 1963. 67 S. ????, ??(??):454, ???? 1963.

Klotz:1985:EV

[Klo85] Irving M. Klotz. Einstein and the Vatican. Physics Today, 38 (7):13–15, July 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2814628.

Klotz:1997:CTF

[Klo97] Irving M. Klotz. Captives of their fantasies: The German atomic bomb scientists. Journal of Chemical Education, 74(2):204–209, February 1997. CODEN JCEDA8. ISSN

0021-9584. URL http://jchemed.chem.wisc.edu/Journal/
issues/1997/feb/abs204.html;   http://jchemed.chem.
wisc.edu/Journal/issues/1997/feb/PlusSub/V74N02/p204.
pdf.

Karazija:2004:NPP

[KM04] Romualdas Karazija and Alina Momkauskaite. The Nobel Prize in Physics — regularities and tendencies. Scientometrics, 61(2): 191–205, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1023/B%3ASCIE.0000041648.87075.de.

Kuran:1997:TBA

[KMN^{+}97] Peter Kuran, Alan Munro, Scott Narrie, Don Pugsley, William Shatner, Edward Teller, Frank H. Shelton, Barbu Marian, William T. Stromberg, Jacqueline Zietlow, and William S. Conner. Trinity and beyond: the atomic bomb movie. Goldhil Video, Thousand Oaks, CA, USA, 1997. ISBN ???? 1 videocassette (95 minutes) pp. LCCN ????

Khavinson:2008:FTA

[KN08] Dmitry Khavinson and Genevra Neumann. From the fundamental theorem of algebra to astrophysics: A “harmonious” path. Notices of the American Mathematical Society, 55(6):666–675, June/July 2008. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/journals/
notices/200806/noti-jun08-cov1.jpg. See follow-on article

[Pet10].

Kroker:2015:ET

[KN15] Stefanie Kroker and Ronny Nawrodt. The Einstein telescope. IEEE Instrumentation & Measurement Magazine, 18(3):4–8, June 2015. CODEN ???? ISSN 1094-6969 (print), 1941-

0123 (electronic). URL https://dblp.org/db/journals/imm/
imm18.html#KrokerN15.

Kim:2019:NPE

[KN19] Y. S. Kim and Marilyn E. Noz. New perspectives on Einstein’s E = mc^{2}. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2019. ISBN 981-323-770-8 (hardcover), 981-323-771-6 (e-book). xi + 192 pp. LCCN QC173.59.S65 K56 2018.

Knight:2012:BRB

[Kni12a] David Knight. Book review: Science Secrets: The Truth about Darwin’s Finches, Einstein’s Wife, and Other Myths. Isis, 103 (2):387, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/667476.

Knight:2012:BRS

[Kni12b] David Knight. Book review: Science Secrets: The Truth about Darwin’s Finches, Einstein’s Wife, and Other Myths. Isis, 103 (2):387, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/667476.

Knispel:2013:EDP

[Kni13] B. Knispel. Einstein@Home discovery of 24 pulsars in the Parkes multi-beam pulsar survey. Astrophysical Journal, 774(2):93–??, ???? 2013. CODEN ASJOAB. ISSN 0004-637X (print), 1538-

4357 (electronic). URL http://stacks.iop.org/0004-637X/
774/i=2/a=93.

Knippel:2016:EIC

[Kni16] Jean-Michel Knippel. Éditorial : Information, communication, James Clerk Maxwell et Albert Einstein. (French). Bull. d’Informatique Approfondie et Appl., 103(??):9–10, ???? 2016.

CODEN ???? ISSN ???? URL http://biaa.eu/-upload/
editono103.pdf;   https://dblp.org/db/journals/biaa/
biaa103.html#Knippel16.

Knoll:1990:LHB

[Kno90] Erwin Knoll. Letter: H-bomb secret never was. Bulletin of the Atomic Scientists, 46(3):53–54, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Knox:1997:BYW

[Kno97] A. J. Knox, editor. The Berlin Years: Writings, 1914–1917. English Translation Supplement, volume 6 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-01734-4 (paperback). xii + 449 pp. LCCN QC16.E5 A2 1987. URL http://press.princeton. edu/titles/6161.html. English translation of selected texts by Alfred Engel in consultation with Engelbert Schucking, with a preface by Engel and Schucking.

Kobayashi:2009:NLC

[Kob09] Makoto Kobayashi. Nobel Lecture: CP violation and flavor mixing. Reviews of Modern Physics, 81(3):1019– 1025, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1019;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1019_1.

Koch:2001:NLB

[Koc01] Richard Koch. The natural laws of business: applying the theories of Darwin, Einstein, and Newton to achieve business success. Currency/Doubleday, New York, NY, USA, 2001. ISBN 0-385-50159-5. 275 pp. LCCN HF5386 .K763 2001eb. URL

http://site.ebrary.com/lib/byuprovo/Doc?id=10035368.

Koenig:1995:KAS

[Koe95] Robert Koenig. Keeping alive the spirit of Einstein. Science, 267 (5204):1598, March 17, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/267/5204/1598.full.pdf.

Kohn:1995:HCW

[Koh95] Richard H. Kohn. History and the culture wars: The case of the Smithsonian Institution’s Enola Gay exhibition. The Journal of American History, 82(3):1036–1063, ???? 1995. ISSN 0021-8723 (print), 1945-2314 (electronic).

Kohn:1999:NLE

[Koh99] W. Kohn. Nobel Lecture: Electronic structure of matter-wave functions and density functionals. Reviews of Modern Physics, 71(5):1253–1266, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1253;   http:
//rmp.aps.org/abstract/RMP/v71/i5/p1253_1.

Kojevnikov:2008:BRD

[Koj08] Alexei Kojevnikov. Book review: David E. Rowe, Robert Schulmann (eds.), Einstein on Politics. His Private Thoughts and Public Stands on Nationalism, Zionism, War, Peace, and the Bomb. Princeton: Princeton University Press, 2007. xxix + 523 pp., ISBN 13: 978-0-691-12094-2. Nuncius, 23(2):441–442, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01093.

Kolman:1940:KVE

[Kol40] E. Kol’man. K vystupleniiu Einshteina po voprosu o sovremennoi fizike (Russian) [On Einstein’s lecture on modern physics]. Pod znamenem marksizma, ??(12):100–105, ???? 1940.

Kollros:1956:AES

[Kol56] Louis Kollros. Albert Einstein en Suisse : souvenirs. (French) [Albert Einstein in Switzerland: memories]. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1956. 281 (est.) pp.

Kollros:1958:AES

[Kol58] Louis Kollros. Alberto Einstein in Svizzera. (Italian) [Albert Einstein in Switzerland], volume 9 of Filosofia della scienza. Ed. di “Filosofia”, Torino, Italia, 1958. 12 pp. Translation of [Kol56] to Italian by Cordelia Guzzo.

Kolossvary:1959:EB

[Kol59] Bela G. Kolossvary. Eötvös balance. American Journal of Physics, 27(5):336–343, May 1959. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Kolomy:1972:AEJ

[Kol72] Rudolf Kolomý. Albert Einstein a jeho vztah v Praze. (Czech) [Albert Einstein and his relationship in Prague]. Pokroky matematiky, fysiky a astronomie, 17(??):265–272, ???? 1972.

Kolomy:1974:AEV

[Kol74] Rudolf Kolomý. Albert Einstein v Praze. (Czech) [Albert Einstein in Prague]. Vesmí, 53(4):112–115, ???? 1974.

Kolata:1977:ESE

[Kol77] Gina Bari Kolata. Einstein skeptical of ESP after all. Science, 197(4311):1349, September 30, 1977. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/197/4311/1349.full.pdf.

Kolb:2007:UTE

[Kol07] Rocky Kolb. A unified theory of Einstein. Bulletin of the Atomic Scientists, 63(6):58–59, November 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//bos.sagepub.com/content/63/6/58.abstract. Review of

[Isa07].

Konitzer:2015:BSE

[Kon15] Franziska Konitzer. Bohr schlägt Einstein im Doppelspalt: mit Sauerstoff-Molekülen haben Forscher ein Gedankenexperiment von Albert Einstein realisiert und die spukhaften Ergebnisse der Quantenwelt erneut bestätigt. (German) [Bohr proposes Einstein in the double slit: with oxygen molecules, researchers have realized a thought experiment by Albert Einstein and reconfirmed the spooky results of the quantum world]. Bild der Wissenschaft, 52:40–41, 2015. ISSN 0006-2375.

Kopff:1921:GER

[Kop21] August Kopff. Grundzüge der Einsteinschen Relativitätstheorie. (German) [The basis of Einstein’s theory of relativity]. Hirzel, Leipzig, Germany, 1921. viii + 198 pp. LCCN ????

Kopff:1923:MTR

[Kop23a] August Kopff. The mathematical theory of relativity. Methuen and Co. Ltd., London, UK, 1923. viii + 14 pp. LCCN ????

Kopff:1923:FDR

[Kop23b] Augusto Kopff. I fondamenti della relatività Einsteiniana: valore e interpretazione della teoria. (Italian) [The foundations of Einsteinian relativity: value and interpretation of the theory]. Ulrico Hoepli, Milano, Italy, 1923. xxx + 455 pp. LCCN ???? Italian translation of German original Grundzüge der Einsteinschen Relativitätstheorie (1921 and 1923). Edited by R. Contu and T. Bembo.

Kopal:1964:OPF

[Kop64] Zdenĕk Kopal. Obituary: Prof. E. F. Freundlich [1885– 1964]. Nature, 204(4960):727, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v204/n4960/
pdf/204727a0.pdf.

Korch:1979:AEA

[Kor79] H. Korch. Albert Einstein about philosophic problems of natural scientific cognition. Deutsche Zeitschrift für Philosophie, 27(2): 167–179, ???? 1979. ISSN 0012-1045.

KormosBarkan:1994:SMC

[Kor94] Diana Kormos Barkan. Simply a matter of chemistry? The Nobel Prize for 1920. Perspectives on Science, 2(4):357–395, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Korner:1996:PC

[Kör96] T. W. (Thomas William) Körner. The Pleasures of Counting. Cambridge University Press, Cambridge, UK, 1996. ISBN 1-107-05056-1 (e-book), 0-521-56823-4, 0-521-56087-X, 1-316-02367-2. x + 534 pp. LCCN QA93 .K65 1996. URL http:

//catdir.loc.gov/catdir/description/cam029/97108334.
html;   http://catdir.loc.gov/catdir/toc/cam027/97108334.
html.

Kostro:1998:AEH

[Kos98] Ludwik Kostro. Albert Einstein’s hypothetism. Science & Education (Springer), 7(3):317–322, May 1998. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). URL http://

link.springer.com/article/10.1023/A:1008661622142.

Kostro:2000:EE

[Kos00] Ludwik Kostro. Einstein and the ether. Apeiron, Montréal, Québec, Canada, 2000. ISBN 0-9683689-4-8 (paperback). iv + 242 pp.

Koshiba:2003:NLB

[Kos03] Masatoshi Koshiba. Nobel Lecture: Birth of neutrino astrophysics. Reviews of Modern Physics, 75(3):1011–1020, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.75.1011;   http://rmp.aps.org/abstract/
RMP/v75/i3/p1011_1.

Kosmulski:2020:NLH

[Kos20] Marek Kosmulski. Nobel laureates are not hot. Scientometrics, 123(1):487–495, April 2020. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/s11192-
020-03378-9;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03378-9.pdf.

Kotecky:1993:KEM

[Kot93] Roman Kotecký. Korespondence Einsteina s Masarykem: Domněle ztracený dopis a pozapomenutá historie jedné intervence. (Czech). [Einstein’s correspondence with Masaryk: Supposedly lost letter and forgotten history of one intervention]. Vesmír, 72(??):566–568, ???? 1993.

Kovacs:2001:GBN

[Kov01] László Kovács. Georg von Békésy, Nobel Laureate in physiology, experimental physicist and art collector was born 100 years ago. Science & Education (Springer), 10(1–2):149–152, January 2001. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Kovacevic:2020:SRO

[Kov20] Mladen Kovacevic. Signaling to relativistic observers: an Einstein–Shannon–Riemann encounter. Problems of Information Transmission, 56(4):303–308, ???? 2020. CODEN PRITA9. ISSN 0032-9460 (print), 1608-3253 (electronic). URL https://

dblp.org/db/journals/poit/poit56.html#Kovacevic20.

Kox:1988:HAL

[Kox88] Anne J. Kox. Hendrik Antoon Lorentz, the ether, and the General Theory of Relativity. Archive for History of Exact Sciences, 38(1):67–78, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=1&spage=67.

Kox:1993:ELM

[Kox93] Anne J. Kox. Einstein and Lorentz: More than just good colleagues. Science in Context, 6(1):43–56, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kox:1995:ESH

[Kox95] A. J. Kox. Einstein, specific heats, and residual rays: The history of a retracted paper. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 9, pages 245–257. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-9_9.

Kox:2006:CCA

[Kox06] Anne J. Kox. Confusion and clarification: Albert Einstein and Walther Nernst’s Heat Theorem, 1911–1916. Studies in History and Philosophy of Modern Physics, 37(1):101–114, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000687.

Kox:2014:ESP

[Kox14] A. J. Kox. Einstein on statistical physics: fluctuations and atomism. In Janssen and Lehner [JL14b], chapter 3, pages 103– 116. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Kallen:1955:MSD

[KP55] Gunnar Källén and Wolfgang Pauli. On the mathematical structure of T. D. Lee’s model of a renormalizable field theory, volume 30(7) of Det Kongelige Danske videnskabernes selskab. Matematisk-fysiske meddelelser. I kommission hos Munksgaard, København, Danmark, 1955. 23 pp. LCCN AS281 .D215 bd. 30, no. 7.

Kraus:1919:DRT

[Kra20a] Oskar Kraus. Ueber die Deutung der Relativitäts-Theorie Einsteins. (German) [On the interpretation of Einstein’s relativity theory]. Lotos, 67/68:146–152, 1919/1920.

Kraus:1920:EGE

[Kra20b] Oskar Kraus. Einwendungen gegen Einstein: Philosophische Betrachtungen gegen die Relativitätstheorie. (German) [Objections to Einstein: Philosophical considerations against the theory of relativity]. Neue Freie Presse, (20130):2–3, September 11, 1920.

Kraus:1920:KGE

[Kra20c] Oskar Kraus. Zum Kampfe gegen Einstein und die Relativitätstheorie. Deutsche Zeitung Bohemia, 93(208):5, September 3, 1920.

Kraus:1921:VBB

[Kra21] Oskar Kraus. Die Verwechselungen von ‘Beschreibungsmittel’ und ‘Beschreibungsobjekt’ in der Einsteinschen speziellen und allgemeinen Relativitätstheorie. (German) [The confusion of ‘means of description’ and object of description’ in Einstein’s special and general theory of relativity]. Kant-Studien, 26(??): 454–486, 1921.

Kraus:1923:GSRa

[Kra23a] Oskar Kraus. Der gegenwärtige Stand der Relativitätstheorie, Erste Teil. (German) [The current state of relativity theory, part one]. Neue Freie Presse, (21238):16–17, October 25, 1923.

Kraus:1923:GSRb

[Kra23b] Oskar Kraus. Der gegenwärtige Stand der Relativitätstheorie, Zweite Teil. (German) [The current state of relativity theory, part two]. Neue Freie Presse, (21240):16, October 27, 1923.

Kraus:1925:OBA

[Kra25] Oskar Kraus. Offene Briefe an Albert Einstein und Max von Laue über die gedanklichen Grundlagen der speziellen und allgemeinen Relativitätstheorie. (German) [Open letters to Albert Einstein and Max von Laue on the conceptual foundations of the special and general theory of relativity]. Wilhelm Braumüller, Wien, Austria, 1925. xvi + 104 pp.

Kragh:1979:NBS

[Kra79] Helge Kragh. Niels Bohr’s second atomic theory. Historical Studies in the Physical Sciences, 10(??):123–186, ???? 1979. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.

org/stable/27757389.

Krafft:1981:ISS

[Kra81] Fritz Krafft. Im Schatten der Sensation: Leben und Wirken von Fritz Strassmann. (German) [In the shadow of sensation: the life and work of Fritz Strassmann]. Verlag Chemie, Weinheim, West Germany, 1981. ISBN 3-527-25818-3. xvii + 541 pp. LCCN QD22.S77 K7.

Kragh:1982:CPT

[Kra82a] Helge Kragh. Cosmo-physics in the thirties: Towards a history of Dirac cosmology. Historical Studies in the Physical Sciences, 13(1):69–108, ???? 1982. CODEN HSPSAS. ISSN 0073-2672.

URL http://www.jstor.org/stable/27757506.

Kranzberg:1982:IWS

[Kra82b] Melvin Kranzberg. The industrialization of Western society, 1860–1914. In Bernhard et al. [BCS82], pages 209– 230. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500189.

Krafft:1988:WBI

[Kra88] Prof.Dr. Fritz Krafft. Wissenschaft und Bildung im Deutschland Des 19. Jahrhunderts: An der Schwelle zum Atomzeitalter. Die Vorgeschichte der Entdeckung der Kernspaltung im Dezember 1938. (German) [Science and education in Germany in the 19th Century: On the threshold of the atomic age. The history of the discovery of nuclear fission, December 1938]. Berichte zur Wissenschaftsgeschichte, 11(4):227–251, ???? 1988. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kragh:1990:DSB

[Kra90] Helge Kragh. Dirac: a scientific biography. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-01756-4 (paperback), 0-521-38089-8 (hardcover). x + 389 pp. LCCN QC16.D57 K73 1990. URL http://www.loc.

gov/catdir/description/cam023/89017257.html;   http:
//www.loc.gov/catdir/toc/cam028/89017257.html.

Krauss:1992:UNE

[Kra92] Lawrence M. Krauss. Universe from nothing: Einstein, the Belgian priest and the puzzle of the Big Bang. Scientific American, ??(??):??, February 10, 1992. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

https://www.scientificamerican.com/article/universe-
from-nothing/.

Kragh:1996:CCH

[Kra96a] Helge Kragh. Cosmology and Controversy: the Historical Development of Two Theories of the Universe. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02623-8 (hardcover). xiii + 500 pp. LCCN QB981

.K73 1996. URL http://www.loc.gov/catdir/description/
prin031/96005612.html;   http://www.loc.gov/catdir/
toc/prin031/96005612.html.

Kragh:1996:GGR

[Kra96b] Helge Kragh. Gamow’s game: The road to the Hot Big Bang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):335–361, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1996Cent.
..38..335K;   http://onlinelibrary.wiley.com/doi/10.
1111/j.1600-0498.1996.tb00020.x/abstract.

Kragh:1999:QGH

[Kra99] Helge Kragh. Quantum generations: a history of physics in the twentieth century. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-01206-7 (hardcover), 0-691-09552-3 (paperback). xiv + 494 pp. LCCN QC7 .K7 2002. URL

http://press.princeton.edu/titles/6683.html.

Kragh:2000:MPR

[Kra00] Helge Kragh. Max Planck: the reluctant revolutionary. Physics World, 13(12):31–36, December 2000. CODEN PHWOEW.

ISSN 0953-8585. URL http://physicsworldarchive.iop.
org/full/pwa-pdf/13/12/phwv13i12a34.pdf. This paper de-

scribes Max Planck’s steps in 1900 to the notion of quantization, which Albert Einstein exploited so well in 1905.

Kragh:2002:BRP

[Kra02] Helge Kragh. Book review: The Politics of Excellence: Behind the Nobel Prize in Science, Robert Marc Friedman. W. H. Freeman, New York, 2001. $30.00 (379 pp.) ISBN 0-7167-3103-7. Physics Today, 55(3):63, March 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Kragh:2005:GGF

[Kra05] Helge Kragh. George Gamow and the ‘factual approach’ to relativistic cosmology. In Kox and Eisenstaedt [KE05], pages 175–188. ISBN 0-8176-4380-X, 0-8176-4454-7 (e-book). LCCN

QC173.5 .U55 2005. URL http://adsabs.harvard.edu/abs/
2005ugr..book..175K.

Krajewski:2006:IER

[Kra06] Władysław Krajewski. On the interpretation of the equation E = mc^{2}: reply to F. Flores: “Interpretations of Einstein’s equation E = mc^{2}” [Int. Stud. Philos. Sci. 19 (2005), no. 3, 245–260; MR2201925]. International Studies in the Philosophy of Science, 20(2):215–216, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02698590600814449. See [Flo05, Flo06].

Kragh:2007:CCM

[Kra07a] Helge Kragh. Conceptions of cosmos: from myths to the accelerating universe: a history of cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920916-2. 276 pp. LCCN QB981 .K729 2007.

URL http://www.loc.gov/catdir/enhancements/fy0723/
2006027692-b.html;   http://www.loc.gov/catdir/toc/
ecip0619/2006027692.html.

Kragh:2007:CEC

[Kra07b] Helge S. Kragh. Cosmology and the entropic creation argument. Historical Studies in the Physical and Biological Sciences, 37 (2):369–382, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kragh:2008:ECR

[Kra08] Helge Kragh. Entropic creation: religious contexts of thermodynamics and cosmology. Science, technology and culture, 1700– 1945. Ashgate, Aldershot, Hampshire, England, 2008. ISBN 0-7546-6414-7. 272 pp. LCCN QB981 .K732 2008. URL http:

//www.loc.gov/catdir/toc/ecip0727/2007038440.html.

Krauss:2011:QMR

[Kra11] Lawrence Maxwell Krauss. Quantum Man: Richard Feynman’s Life in Science. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-06471-9. xvii + 350 pp. LCCN QC16.F49 K73 2011. URL http:

//www.scientificcomputing.com/news-DS-Illuminating-
the-Life-and-Legacy-of-Richard-Feynman-032211.aspx.

Kragh:2013:SAC

[Kra13] Helge Kragh. Svante Arrhenius, cosmical physicist and auroral theorist. History of Geo- and Space Sciences, 4(2):61–69, ???? 2013. CODEN ???? ISSN 2190-5010 (print), 2190-5029

(electronic). URL https://www.hist-geo-space-sci.net/4/
61/2013/.

Krause:2014:AE

[Kra14a] Michael Krause. Albert Einstein (1879–1955). In CERN: how we found the Higgs boson [Kra14b], chapter 11, page ?? ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). LCCN QC793 .K73 2014.

Krause:2014:CHW

[Kra14b] Michael Krause. CERN: how we found the Higgs boson. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). xiii + 243 pp. LCCN QC793 .K73 2014.

Krause:2014:EQ

[Kra14c] Michael Krause. Einstein quotes. In CERN: how we found the Higgs boson [Kra14b], chapter 11, page ?? ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (ebook). LCCN QC793 .K73 2014.

Krause:2014:ECC

[Kra14d] Michael Krause. Einstein’s cosmological constant. In CERN: how we found the Higgs boson [Kra14b], chapter 12, page ?? ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). LCCN QC793 .K73 2014.

Kramer:2015:NPW

[Kra15a] David Kramer. New park will honor US atomic heritage. Physics Today, 68(3):24–25, March 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Krauss:2015:CWE

[Kra15b] Lawrence M. Krauss. Cosmology: What Einstein got wrong. Scientific American, 313(3):50–55, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-50.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-50.pdf.

Kragh:2017:NPS

[Kra17] Helge Kragh. The Nobel Prize system and the astronomical sciences. Journal for the History of Astronomy, 48(3):257–280, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-

8556 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0021828617721574.

Kragh:2018:GLP

[Kra18] Helge Kragh. Georges Lemaître, pioneer of modern theoretical cosmology. Foundations of Physics, 48(10):1333–1348, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kreider:1958:IEC

[Kre58] Marlin Books Kreider. Identifying Einstein’s “creative force”. In The evidence of God in an expanding universe: forty American scientists declare their affirmative views on religion [Mon58], page ?? LCCN BL240 .M715.

Krefetz:1972:CDE

[Kre72] Elliott Krefetz. Comment on: On the Derivation of Einstein’s Field Equations. American Journal of Physics, 40(9): 1347, 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/40/9/10.1119/1.1986837. See [Cha72].

Krim:1999:BRA

[Kri99] Jacqueline Krim. Book review: Against all odds: The triumphs and defeats of Lise Meitner: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. The Physics Teacher, 37(4):202, April 1999. CODEN PHTEAH. ISSN 0031-921X (print), 1943-

4928 (electronic). URL http://scitation.aip.org/content/
aapt/journal/tpt/37/4/10.1119/1.880229.

Krige:2001:NPP

[Kri01] John Krige. The 1984 Nobel Physics Prize for heterogeneous engineering. Minerva [Washington, DC], 39(4):425–443, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012735919366.

Krige:2008:BRK

[Kri08] John Krige. Book review: Klaus Hentschel, The Mental Aftermath. The Mentality of German Physicists 1945–1949. Oxford: Oxford University Press, 2007. 205 pp., ISBN 978-0-19-920566-0. Nuncius, 23(1):176–177, ???? 2008. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00445.

Krivit:2019:NFR

[Kri19a] Steven B. Krivit. The Nobel Foundation’s retraction of the Rutherford transmutation claim. Web site., May 19,

2019. URL http://news.newenergytimes.net/2019/05/19/
the-nobel-foundations-retraction-of-the-rutherford-
transmutation-claim/.

Krivit:2019:RRR

[Kri19b] Steven B. Krivit. Rutherford’s reluctant role in nuclear transmutation. Web site., May 18, 2019. URL http://news.

newenergytimes.net/2019/05/18/rutherfords-reluctant-
role-in-nuclear-transmutation/.

Krivit:2019:WFS

[Kri19c] Steven B. Krivit. The world’s first successful alchemist (it wasn’t Rutherford). Web site., May 14, 2019. URL http://

news.newenergytimes.net/2019/05/14/the-worlds-first-
successful-alchemist-it-wasnt-rutherford/.

Kasprzak:2006:BEC

[KRK^{+}06] J. Kasprzak, M. Richard, S. Kundermann, A. Baas, P. Jeambrun, J. M. J. Keeling, F. M. Marchetti, M. H. Szymaśka, R. André, J. L. Staehli, V. Savona, P. B. Littlewood, B. Deveaud, Le Si Dang, et al. Bose–Einstein condensation of exciton polaritons. Nature, 443(7110):409–414, September 28, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v443/n7110/full/nature05131.html.

Kroner:1953:BAE

[Krö53] F. Kröner. Buchbesprechung: Albert Einstein, philosopherscientist, Herausgegeben von P. A. Schilpp. The Library of Living Philosophers, Inc., Evanston, Illinois, 1949. Dialectica: International Review of Philosophy of Knowledge, 7(1):61–69, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Krotkov:1980:NP

[Kro80] Yuri Krotkov. The Nobel Prize. Simon and Schuster, New York, NY, USA, 1980. ISBN 0-671-24255-5. 348 pp. LCCN PZ4.K9377 PG3482.8.R66 No.

Kromphardt:1981:UE

[Kro81] Karl Kromphardt. Unsöld on Einstein. Nature, 291(5814): 374, June 4, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v291/n5814/pdf/291374d0.pdf.

Kroemer:2001:NLQ

[Kro01] Herbert Kroemer. Nobel Lecture: Quasielectric fields and band offsets: teaching electrons new tricks. Reviews of Modern Physics, 73(3):783–793, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.73.783;
http://rmp.aps.org/abstract/RMP/v73/i3/p783_1.

Krstic:1991:AME

[Krs91] Dord Krstić. Appendix A: Mileva Einstein-Marić. In Hans Albert Einstein: Reminiscences of His Life and Our Life Together [Ein91g], pages 85–99. ISBN 0-87414-083-8. LCCN TC140.E38 E38 1991. With a foreword by Robert W. Clark.

Krstic:2004:MAE

[Krs04] Dord Krstić. Mileva & Albert Einstein: Their Love and Scientific Collaboration. Didakta, Radovljica, Slovenia, 2004. ISBN 961-6530-08-9. 247 pp. LCCN ????

Krull:1994:AES

[Kru94] Frank Krull. Albert Einstein in seinen erkenntnistheoretischen Äußerungen. (German) [Albert Einstein in his epistemic theory pronouncements]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 78(2):154–170, ???? 1994. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.org/

stable/20777461.

Kruger:2000:TDU

[Krü00] Thomas Krüger. Towards a deeper understanding of the Einstein–Podolsky–Rosen problem. Foundations of Physics, 30 (11):1869–1890, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A:1003758305043.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN QC16.H35 W474 2005.

URL http://www.loc.gov/catdir/toc/fy0613/2006402966.
html.

Kraus:2005:WES

[KRZ^{+}05] Ute Kraus, Hanns Ruder, Corvin Zahn, Marc Borchers, and Daniel Weiskopf. Was Einstein sicher auch gern gesehen hätte? Visualisierung relativistischer Effekte. (German) [What Einstein would certainly like to see — visualization of relativistic effects]. In Steiner [Ste05a], pages 133–150. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http://link.springer.

com/chapter/10.1007/3-540-30595-5_6/.

Kohn:1965:SCE

[KS65] W. Kohn and L. J. Sham. Self-consistent equations including exchange and correlation effects. Physical Review (2), 140(4A):A1133–A1138, November 1965. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.
140.A1133;   http://www.nobelprize.org/nobel_prizes/
chemistry/laureates/1998/. The authors of this paper re-

discover an earlier result [Gás54], and obtain the same value α = 2/3, compared to Slater’s original α = 1 [Sla51, Sla72].

Kramer:1980:ESE

[KS80] D. (Dietrich) Kramer and Ernst Schmutzer, editors. Exact solutions of Einstein’s field equations. Cambridge monographs on mathematical physics. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-23041-1. 425 pp. LCCN QC173.6 .E96.

Kurylo:1981:FBL

[KS81] Friedrich Kurylo and Charles Susskind. Ferdinand Braun, a life of the Nobel Prize winner and inventor of the cathode-ray oscilloscope. MIT Press, Cambridge, MA, USA, 1981. ISBN 0-262-11077-6. xiv + 289 pp. LCCN QC16.B68 K813.

Khariton:1993:KV

[KS93] Yuli Khariton and Yuri Smirnov. The Khariton version. Bulletin of the Atomic Scientists, 49(4):20–31, May 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.thebulletin.org/article.php?
art_ofn=may93khariton.

Kox:1995:NTE

[KS95] A. J. Kox and Daniel M. Siegel. No Truth Except in the Details: Essays in Honor of Martin J. Klein, volume 167 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. xxv + 382 + 13 + 1 pp. LCCN ???? URL https://

link.springer.com/book/10.1007/978-94-011-0217-9.

Knapp:2012:KHB

[KS12] Jane F. Knapp and Robert D. Schremmer. Kansas horse & buggy doctor receives letter from Albert Einstein. Missouri medicine, 109(1):46, January/February 2012. ISSN 0026-6620.

Kox:2009:CIB

[KSB^{+}09] Anne J. Kox, Tilman Sauer, Diana Kormos Buchwald, Rudy Hirschmann, Osik Moses, Benjamin Aronin, and Jennifer Stolper. Cumulative Index, Bibliography, List of Correspondence, Chronology, and Errata to Volumes 1–10, volume 11 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14187-8. xv + 619 pp. LCCN QC16.E5 A2 2009.

URL http://einsteinpapers.press.princeton.edu/vol11;
 http://press.princeton.edu/titles/8953.html.

Kubrick:1963:DSH

[KSG63] Stanley Kubrick, Terry Southern, and Peter George. Dr. Strangelove, or, how I learned to stop worrying and love the bomb. Motion picture., 1963.

Kincheloe:1999:SGE

[KST99] Joe L. Kincheloe, Shirley R. Steinberg, and Deborah J. Tippins. The stigma of genius: Einstein, consciousness, and education, volume 111 of Counterpoints. P. Lang, New York, NY, USA, 1999. ISBN 0-8204-4431-6 (paperback). xxvi + 218 pp. LCCN LB875.E562 K56 1999.

Klaers:2010:BEC

[KSVW10] Jan Klaers, Julian Schmitt, Frank Vewinger, and Martin Weitz. Bose–Einstein condensation of photons in an optical microcavity. Nature, 468(7323):545–548, November 24, 2010. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v468/n7323/
full/nature09567.html.

Kirsten:1979:AEB

[KT79] Christa Kirsten and Hans-Jürgen Treder. Albert Einstein in Berlin: 1913–1933, volume 6–7 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie-Verlag, Berlin, Germany, 1979. ISBN ???? 287 + 8 (vol. 1), 295 (vol.

2) pp. LCCN QC16.E5 A67. 42.00M (vol. 1).

Kaku:1995:BEC

[KT95] Michio Kaku and Jennifer Trainer Thompson. Beyond Einstein: the cosmic quest for the theory of the universe. Anchor Books, New York, NY, USA, 1995. ISBN 0-385-47781-3 (paperback). xi + 224 pp. LCCN QC173.7 .K35 1995. URL http://www.

loc.gov/catdir/description/random042/95001815.html.

Kaku:2015:TDT

[KT15] Michio Kaku and Jennifer Trainer Thompson. La teoria del tutto : Einstein e le nuove vie della fisica. (Italian) [The theory of everything: Einstein and the new ways of physics]. Castelvecchi, Roma, Italia, 2015. ISBN 88-6944-118-0. 221 pp. LCCN ????

Kahn:1938:TC

[KU38] B. Kahn and G. E. Uhlenbeck. On the theory of condensation. Physica, 5(5):399–416, May 1938. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Kubo:1986:BMN

[Kub86] Ryogo Kubo. Brownian motion and nonequilibrium statistical mechanics. Science, 233(4761):330–334, July 18, 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/233/
4761/330.abstract.

Kuchment:2011:WNT

[Kuc11] Anna Kuchment. For whom the Nobel tolls. Scientific American, 304(1):27, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v304/n1/full/
scientificamerican0111-27.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
27.pdf.

Kuhn:1954:BRB

[Kuh54] Thomas S. Kuhn. Book review: Dialogue Concerning the Two Chief World Systems — Ptolemaic & Copernican. Galileo Galilei. Translated by Stillman Drake, foreword by Albert Einstein. Univ. California Press, Berkeley, 1953. 496 pp. Illus. $10. Science, 119(3095):546–547, April 23, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/119/3095/546.3.full.pdf.

Kuhn:1980:ECP

[Kuh80] Thomas S. Kuhn. Einstein’s critique of Planck. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Kuhn:1984:RP

[Kuh84] Thomas S. Kuhn. Revisiting Planck. Historical Studies in the Physical Sciences, 14(2):231–252, ???? 1984. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757534.

Kumar:2010:QEB

[Kum10] Manjit Kumar. Quantum: Einstein, Bohr and the great debate about the nature of reality. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-07829-9 (hardcover). xvi + 448 + 16 pp. LCCN QC173.98 .K86 2010; QC173.98.K86.

Kumar:2011:GRP

[Kum11] Manjit Kumar. Le grand roman de la physique quantique: Einstein, Bohr et le dêbat sur la nature de la réalité. (French) [Quantum: Einstein, Bohr and the great debate about the nature of reality]. J. C. Lattés, Paris, France, 2011. ISBN 2-7096-2465-6. 524 + 16 pp. LCCN ????

Kuna:1976:KSC

[Kun76] Franz Kuna, editor. On Kafka: Semi-centenary perspective. Elek, London, UK, 1976. ISBN 0-236-40050-9. xi + 195 pp. Proceedings of the Kafka Symposium, University of East Anglia, Norwich, 7–10 July 1974.

Kursunoglu:1979:NTH

[Kur79] Behram Kursunoglu. A non-technical history of the generalized theory of gravitation dedicated to the Albert Einstein centennial. In Perlmutter and Scott [PS79], pages 15–37. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178 .O72 1979. URL http://link.

springer.com/chapter/10.1007/978-1-4684-3596-2_2/.

Kurzman:1986:DBC

[Kur86] Dan Kurzman. Day of the bomb: countdown to Hiroshima. Mc-Graw-Hill, New York, NY, USA, 1986. ISBN 0-07-035683-1. xiv + 546 + 24 pp. LCCN D767.25.H6 K865 1986.

Kurian:2002:NSB

[Kur02] George Thomas Kurian. The Nobel scientists: a biographical encyclopedia. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-927-1. 420 pp. LCCN Q141 .K78 2002.

Kuran:2006:HPA

[Kur06] Peter Kuran. How to photograph an atomic bomb. VCE, Santa Clarita, CA, USA, 2006. ISBN 1-889054-19-4 (paperback), 1-889054-11-9. 141 pp. LCCN UG476 .K87 2006. US$24.95.

Kurzweil:2012:HCM

[Kur12] Ray Kurzweil. How to Create a Mind: the Secret of Human Thought Revealed. Penguin, London, UK and New York, NY, USA, 2012. ISBN 0-14-312404-8. xi + 336 pp. LCCN QP385 .K87 2013.

Kutcher:2009:CMC

[Kut09] Gerald Kutcher. Contested medicine: cancer research and the military. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-46531-4 (hardcover). x + 247 pp. LCCN RC267 .K88 2009.

Kuppers:1982:ANP

[KUW82] Günter Küppers, Norbert Ulitzka, and Peter Weingart. The awarding of the Nobel Prize: Decisions about significance in science. In Bernhard et al. [BCS82], pages 332–351. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B978008027939850027X.

Kuzneetisov:1972:ED

[Kuz72] B. G. (Boris Grigorevich) Kuzneetisov. Einstein and Dostoyevsky. Hutchinson, London, UK, 1972. ISBN 0-09-106660-3. 112 pp. LCCN QC6 .K78413.

Kuznetsov:1979:CPS

[Kuz79] B. Kuznetsov. Coincidental parallels — some more about Albert Einstein and Fedor Mikhailovich Dostoevski. Novyi Mir, ??(3): 224–235, ???? 1979. ISSN 0130-7673.

Kuznetsov:1987:EE

[Kuz87a] B. G. Kuznetsov. Einstein and Epicurus. In Reason and Being [FC87], chapter 4, pages 67–108. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346. LCCN Q174

.B67 vol. 17. URL https://link.springer.com/chapter/10.
1007/978-94-009-4590-6_4.

Kuznetsov:1987:SE

[Kuz87b] B. G. Kuznetsov. Spinoza and Einstein. In Reason and Being [FC87], chapter 7, pages 178–210. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346. LCCN Q174

.B67 vol. 17. URL https://link.springer.com/chapter/10.
1007/978-94-009-4590-6_7.

Krause:2011:BSP

[KV11] Décio Krause and Antonio Videira. Brazilian Studies in Philosophy and History of Science: An account of recent works, volume 290 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346. xiii + 352 pp. LCCN Q175.3 .B73

2011. URL https://link.springer.com/book/10.1007/978-
90-481-9422-3.

Kvasnica:1979:NRW

[Kva79] J. Kvasnica. Non-relativistic work of Albert Einstein. Ceskoslovensky Časopis pro Fyziku Sekce A, 29(3):212–221, ???? 1979. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN

QC16.D57 R46 1987. URL http://www.loc.gov/catdir/
description/cam023/86033409.html;   http://www.loc.
gov/catdir/toc/cam028/86033409.html.

Kleint:1993:WHL

[KW93] Christian Kleint and Gerald Wiemers, editors. Werner Heisenberg in Leipzig, 1927–1942, volume 58(2) of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse, 0365-6470. Akademie-Verlag, Berlin, Germany, 1993. ISBN 3-05-501585-1. 263 pp. LCCN QC16.H35 W48 1933. EUR$607 (!).

Krahl:2002:ANS

[KWB^{+}02] Hilde Krahl, Mathias Wieman, Dieter Borsche, Werner Hinz, Dorothea Wieck, and Harald Braun. The Alfred Nobel story. Wellspring, New York, NY, USA, 2002. ISBN 0-7942-0140-7. LCCN PN1995.9.P42 A43 2002. 1 videodisc (108 min.).

Kigar:1997:ECN

[KWGH97] D. L. Kigar, S. F. Witelson, I. I. Glezer, and T. Harvey. Estimates of cell number in temporal neocortex in the brain of Albert Einstein. Society for Neuroscience Abstracts, 23(1–2): 213, ???? 1997. ISSN 0190-5295.

Kremens:1957:REC

[KWM57] V. Kremens, A. Woldow, and R. S. Monheit. Recent experience with congenital heart disease at the Albert Einstein medical center. The Journal of the Albert Einstein Medical Center, Philadelphia, 5(2):141–152, March 1957. ISSN 0002-4708.

Kikot:2005:AS

[KWMS05] Thom Kikot, Leyna Juliet Weber, Jean Meltzer, and Megan Selheim. The actinide series. Periodic table of the elements. Films for the Humanities and Sciences, Hamilton, NJ, USA, 2005. LCCN QD172.A3 A28 2005. 56-minute video film, originally produced as part of the television program Assignment Discovery.

Kuran:1999:AJW

[KZS99] Peter Kuran, Jacqueline Zietlow, and William Shatner, editors. Atomic journeys: welcome to ground zero. Goldhil Video, Thousand Oaks, CA, USA, 1999. ISBN 1-58565-141-9. 1 videocassette (53 min.) pp. LCCN ????

L:1958:RLB

[L.58] A.-L. L. Review: Louis de Broglie, physicien et penseur by Albert Einstein, Louis de Broglie, and 44 others. Revue Philosophique de la France et de l’Étranger, 148:102, 1958. ISSN 0035-3833 (print), 2104-385X (electronic). URL http://www.

jstor.org/stable/41089668.

L:1989:SOB

[L.89] K. L. Some other books of interest: Book review: Einstein Simplified: Cartoons in Science. Sidney Harris. Rutgers University Press, New Brunswick, NJ, 1989. Science, 244(4900):90– 91, April 7, 1989. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/244/4900/90.3.full.pdf.

Laporte:1988:CRL

[LA88] Paul M. Laporte and Rudolf Arnheim. Cubism and Relativity with a letter of Albert Einstein. Leonardo (Oxford, England), 21(3):313–315, July 1988. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL https://muse.jhu.edu/

article/600680/summary.

Lacki:2008:BRB

[Lac08] Jan Lacki. Book review: The Collected Papers of Albert Einstein, by Albert Einstein; Diana Kormos Buchwald; Tilman Sauer; Ze’ev Rosenkranz; József Illy; Virginia Iris Holmes. Volume 10: The Berlin Years: Correspondence, May-December 1920, and Supplementary Correspondence, 1909–1920. Isis, 99 (4):850–851, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597714.

Lacayo:2011:EEL

[Lac11] Richard Lacayo, editor. Einstein: the enduring legacy of a modern genius. Time Books, New York, NY, USA, 2011. ISBN 1-60320-173-4. 96 pp. LCCN QC16.E5

L23 2011. URL http://www.loc.gov/catdir/enhancements/
fy1502/2010941157-d.html.

Lagerkvist:2003:PMN

[Lag03] Ulf Lagerkvist. Pioneers of Microbiology and the Nobel Prize. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 1-281-92824-0, 981-238-233-X (hardcover), 981-238-234-8 (paperback), 981-277-556-0 (e-book). ix + 178 pp. LCCN QR21 .L34 2003. URL https:/

/www.worldscientific.com/worldscibooks/10.1142/5161.

Lagerkvist:2012:PTM

[Lag12] Ulf Lagerkvist, editor. The periodic table and a missed Nobel Prize. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2012. ISBN 981-4295-95-7 (paperback), 981-4295-96-5 (e-book). xii + 122 pp. LCCN AS911.N9 .L34 2012; QD11 .L34 2012; QD467 .L34 2012.

URL http://www.worldscientific.com/worldscibooks/10.

1142/7658. Edited by Erling Norrby.

Lahav:2020:DEI

[Lah20] Ofer Lahav. Dark energy: is it ‘just’ Einstein’s cosmological constant Λ? Contemporary Physics, 61(2):132–145, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Laird:1920:ETS

[Lai20] Elizabeth R. Laird. Einstein’s theory and shift of spectrum lines. Science, 52(1331):14–15, July 2, 1920. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/52/1331/14.full.pdf.

Lake:1985:SLS

[Lak85] K. W. Lake. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. Queens Quarterly, 92(4):829–830, Winter 1985. ISSN 0033-6041.

Lamson:1930:RPR

[Lam30] K. W. Lamson. Recent publications: Reviews: A Simplified Presentation of Einstein’s Unified Field Equations, by Tullio Levi-Civita and John Dougall. American Mathematical Monthly, 37 (3):148, March 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lammert:1979:WZE

[Lam79] Eberhard Lammert. Der Wissenschaftler zwischen Erkenntnis und Verantwortung: Zur Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner und Max von Laue am 1. März 1979 in Berlin. (German) [The scientist between knowledge and responsibility: Celebrating the 100th birthdays of Albert Einstein, Otto Hahn, Lise Meitner and Max von Laue on March 1, 1979 in Berlin]. Physik Journal, 35(11):526–530, November 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Lamy:2007:AOB

[Lam07] Jérôme Lamy. Analyses d’ouvrages: L’Empire du temps: Les horloges d’Einstein et les cartes de Poincaré par Peter Galison. Revue d’Histoire des Sciences, 60(1):258–260, January 2007. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23634263.

Lamphere:2021:CNC

[Lam21] Robert L. Lamphere. 105.15 Could Newton have calculated the deflection angle of starlight? Mathematical Gazette, 105(562): 154–158, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10515-

could-newton-have-calculated-the-deflection-angle-of-
starlight/32D1E2D1183594AD676C3B47403C93EB.

Langevin:1908:TDM

[Lan08] Paul Langevin. Sur la théorie du mouvement brownien. (French) [On the theory of Brownian motion]. Comptes rendus de l’Académie des sciences, Paris, 146(??):530–533, ???? 1908. English translation, with historical remarks, in [LG97].

Lanczos:1921:FBM

[Lan21] Cornelius Lanczos. Die funktionentheoretischen Beziehungen der Maxwellsche Aethergleichungen — Ein Beitrag zur Relativitäts und Elektronentheorie. (German) [The functiontheoretical relationships of the Maxwellian aether-equations — a contribution to the Theory of Relativity and electrons]. Ph.D. thesis, University of Szeged, Szeged, Hungary, 1921. Dedicated, with permission, to Albert Einstein.

Lanczos:1937:PAR

[Lan37] Cornelius Lanczos. The philosophical aspects of Relativity. Two lectures delivered by invitation on March 2 and 4, 1937 in the course Life Views of Great Men of Science, Philosophy 143b at

Indiana University, 1937. URL http://catalog.hathitrust.
org/Record/000483261.

Lang:1948:ETA

[Lan48] Daniel Lang. Early Tales of the Atomic Age. Doubleday, Garden City, NY, USA, 1948. 223 pp. LCCN QC173 .L35.

Lanczos:1949:LMR

[Lan49] Cornelius Lanczos. Lagrangian multiplier and Riemannian spaces. Reviews of Modern Physics, 21(3):497–502, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.497;   http://rmp.aps.org/

abstract/RMP/v21/i3/p497_1. Special issue in honor of Albert Einstein on his 70th birthday.

Lanczos:1955:AET

[Lan55] C. Lanczos. Albert Einstein and the Theory of Relativity. Il Nuovo Cimento (10), 2(Supplement 5):1193–1220, 1955. CO-DEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Lanczos:1959:AER

[Lan59] Cornelius Lanczos. Albert Einstein and the role of theory in contemporary physics. American Scientist, 47(??):41–59, March 1959. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Lanczos:1965:AEC

[Lan65] Cornelius Lanczos. Albert Einstein and the cosmic world order: six lectures delivered at the University of Michigan in the Spring of 1962. Interscience Publishers, New York, NY, USA, 1965. vi + 139 pp. LCCN QC 16 .E35 L252 1965. URL http://

catalog.hathitrust.org/Record/001477802.

Lanczos:1967:AES

[Lan67a] Cornelius Lanczos. Albert Einshtein i stroenie kosmosa: Shest lektsii, Prochitannykh vesnoi 1962 G. v Michiganskom universitete. (Russian) [Albert Einstein and the cosmic world order: six lectures delivered at the University of Michigan in the Spring of 1962]. Nauka, Moscow, USSR, 1967. 156 pp. LCCN ???? Translated from English to Russian by V. A. Ugarova.

Lanczos:1967:AEP

[Lan67b] Cornelius Lanczos. Albert Einstein i porza̧dek wszechświata, volume 92 of Omega: współczesna biblioteka naukowa. Państwowe Wydawnictwo Naukowe, Warszawa, Poland, 1967. 146 + 4 pp. LCCN ???? Translated from English to Polish by Barbara Stanosz.

Lanczos:1967:CCV

[Lan67c] Cornelius Lanczos. Che cosa ha veramente detto Einstein?. (Italian) [What did Einstein really say?]. Ubaldini, Rome, Italy, 1967. 154 pp. LCCN ????

Lanczos:1967:EEE

[Lan67d] Cornelius Lanczos. Einstein equations and electromagnetism. Journal of Mathematical Physics, 8(4):829–836, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lanczos:1970:STA

[Lan70] Cornelius Lanczos. Space through the ages: the evolution of geometrical ideas from Pythagoras to Hilbert and Einstein. Academic Press, New York, NY, USA, 1970. x + 320 pp. URL

http://catalog.hathitrust.org/Record/000164850.

Lanczos:1972:EPS

[Lan72] C. Lanczos. Einstein’s path from Special to General Relativity. In Synge and O’Raifeartaigh, ed [SO72], pages 5–19. ISBN 0-19-851126-4. LCCN QC6 .G358. URL http://adsabs.harvard.

edu/abs/1972grec.conf....5L. Edited for the Royal Irish

Academy by L. O’Raifeartaigh.

Lanczos:1974:EDa

[Lan74a] Cornelius Lanczos. The Einstein decade (1905–1915). Histories of science series. Elek Science, London, UK, 1974. ISBN 0-236-17632-3. xiv + 230 pp. LCCN QC16.E5 L33

1974b. URL http://catalog.hathitrust.org/api/volumes/
oclc/1113008.html;   http://catalog.hathitrust.org/
Record/000033210.

Lanczos:1974:EDb

[Lan74b] Cornelius Lanczos. The Einstein decade, 1905–1915. Academic Press, New York, NY, USA, 1974. ISBN 0-12-435840-3. xiv + 230 pp. LCCN QC16.E5 L33 1974.

URL http://catalog.hathitrust.org/api/volumes/oclc/
1113008.html;   http://catalog.hathitrust.org/Record/
000033210.

Lande:1974:AEQ

[Lan74c] Alfred Landé. Albert Einstein and the quantum riddle. American Journal of Physics, 42(6):459–464, June 1974. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/42/459/1.

Lanouette:1994:AS

[Lan94] William Lanouette. Atomic spies. Debate with the authors of Special Tasks by Soviet spymaster Pavel Sudoplatov [SSSS95], which falsely alleged that Szilard, Niels Bohr, J. Robert Oppenheimer and Enrico Fermi were Soviet agents within the Manhattan Project. MacNeil/Lehrer NewsHour, April 26, 1994.

Lancaster:2000:TDS

[Lan00] Kurt Lancaster. Theatrical deconstructionists: The social “gests” of Peter Sellars’s Ajax and Robert Wilson’s Einstein on the Beach. Modern Drama, 43(3):461–468, Fall 2000. CODEN ???? ISSN 0026-7694 (print), 1712-5286 (electronic).

Langridge:2002:EOM

[Lan02] Robert Langridge. Einstein and the orbit of Mercury. Science, 297(5580):336, July 19, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/297/5580/336.2.full.pdf.

Landsman:2006:WCM

[Lan06] N. P. Landsman. When champions meet: Rethinking the Bohr–Einstein debate. Studies in History and Philosophy of Modern Physics, 37(1):212–242, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000869.

Landsman:2007:BCQ

[Lan07a] N. P. Landsman. Between classical and quantum. In Butterfield and Earman [BE07], pages 417–553 (part A). ISBN 0-444-51560-7 (set), 0-444-53001-0 (part A), 0-444-53002-9 (part B).

LCCN QC6 .P44 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006050854-d.html.

Lanouette:2007:EE

[Lan07b] William Lanouette. Enlisting Einstein. In Kelly [Kel07], pages 38–41. ISBN 1-57912-747-9, 1-57912-808-4 (paperback), 1-60376-206-X (e-book). LCCN QC773.3.U5 M27

2007. URL http://www.loc.gov/catdir/toc/ecip0720/
2007023984.html.

Lanouette:2008:BRP

[Lan08] William Lanouette. Book review: The political Einstein: Einstein on Politics: His Private Thoughts and Public Stands on Nationalism, Zionism, War, Peace, and the Bomb eds. David E. Rowe and Robert Schulmann. Princeton and Oxford: Princeton University Press, 2007, 523 pp. Issues in Science and Technology, 24(2):??, Winter 2008. ISSN 0748-5492 (print), 1938-1557

(electronic). URL http://issues.org/24-2/br_lanouette-
2/.

Lange:2014:DER

[Lan14] M. Lange. Did Einstein really believe that principle theories are explanatorily powerless? Perspectives on Science, 22(4):449– 463, ???? 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/article/
561191.

Lanouette:2015:BRB

[Lan15] William Lanouette. Book review: Bomb control: Thermonuclear Monarchy: Choosing Between Democracy and Doom, by Elaine Scarry. New York and London: W. W. Norton & Company, 2014, 582 pp. Issues in Science and Technology, 31(4):??, Summer 2015. ISSN 0748-5492 (print), 1938-1557 (electronic).

URL http://issues.org/31-4/book-review-bomb-control/

.

Largeault:1984:AEH

[Lar84a] J. Largeault. Albert Einstein historical and cultural perspectives, the Centennial Symposium in Jerusalem (1979) — G. Holton, Y. Elkana. Archives de Philosophie, 47(1):139–141, ???? 1984. ISSN 0003-9632.

Largeault:1984:SLS

[Lar84b] J. Largeault. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. Archives de Philosophie, 47(1):141– 144, ???? 1984. ISSN 0003-9632.

Larsson:2001:MMO

[Lar01] Ulf Larsson, editor. Människor, miljöer och kreativitet. Cultures of creativity: the centennial exhibition of the Nobel prize, volume 2 of Nobel Museum Archives. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2001. ISBN 0-88135-288-8. 228 pp. LCCN AS911.N9 M34 2001.

Larsson:2008:ANN

[Lar08] Ulf Larsson. Alfred Nobel: networks of innovation, volume 10 of Archives of the Nobel Museum. Nobel Museum, Stockholm, Sweden, 2008. ISBN 0-88135-399-X (paperback). 216 pp. LCCN TP268.5.N7 L37 2008.

Laroche:2012:PNS

[Lar12] Josepha Laroche. Les prix Nobel: sociologie d’une élite transnationale. (French) [The Nobel Prizes: sociology of a transnational elite]. Liber, Montréal, QC, Canada, 2012. ISBN 2-89578-378-0 (paperback). 184 + 1 pp. LCCN AS911.N9 L37 2012.

Larkoski:2019:EPP

[Lar19] Andrew J. Larkoski. Elementary Particle Physics: an Intuitive Introduction. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-49698-9 (hardcover). 540 (est.) pp. LCCN QC793.2 .L37 2019.

Latour:1988:RAE

[Lat88] Bruno Latour. A relativistic account of Einstein’s relativity. Social Studies of Science, 18(1):3–44, February 1, 1988. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL http://www.jstor.org/stable/285375;   https://
journals.sagepub.com/doi/pdf/10.1177/030631288018001001.

Laurence:1945:UAB

[Lau45] William Leonard Laurence. U.S. atom bomb site belies Tokyo tales: Tests on New Mexico range confirm that blast, and not radiation, took toll. New York Times, ??(??):1, 4, September 12, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/107140479/.

Laurence:1946:DZS

[Lau46] William Leonard Laurence. Dawn over zero; the story of the atomic bomb. Alfred A. Knopf, New York, NY, USA, 1946. xii + 274 + viii pp. LCCN UF767 .L3.

Laurence:1950:EPH

[Lau50] William L. Laurence. Einstein publishes his ‘master theory’: Long-awaited new chapter to relativity volume is product of 30 years of labor. Revised at last minute. Text and formulae presented first in typewritten form condensed and altered. New York Times, ??(??):29, February 15, 1950. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
111520708/.

Laurence:1951:DHR

[Lau51] William L. Laurence. Day of Hiroshima recalls atom race: Two leaders in atomic work at Columbia University. New York Times, ??(??):3, August 6, 1951. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112188459/fulltextPDF.

Laurence:1955:FRO

[Lau55a] William L. Laurence. Father of relativity and outstanding pacifist inspired U.S. to make atomic bomb: Warning stirred president to act; Einstein Relativity theory, one of history’s boldest, proved by nuclear fission; he sought cosmic law; devoted his life to develop all-embracing equations of forces in the Universe. New York Times, ??(??):24, April 19, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

search.proquest.com/docview/113433792.

Laurence:1955:KCS

[Lau55b] William L. Laurence. Key clue sought in Einstein brain; son asked study of Einstein brain; class prize note from Dr. Einstein thanking it for 76th birthday gift; Einstein letter hailed trend on atom power. New York Times, ??(??):24, April 20, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Laurence:1959:MAD

[Lau59] William Leonard Laurence. Men and atoms; the discovery, the uses, and the future of atomic energy. Simon and Schuster, New York, NY, USA, 1959. 302 pp. LCCN QC773 .L3.

Laub:1962:AEA

[Lau62] Jakob Johann Laub. Albert Einstein und Albert Gockel. (German) [Albert Einstein and Albert Gockel]. Academia Friburgensis, 20(??):30–33, ???? 1962.

Laurence:1972:DZS

[Lau72] William Leonard Laurence. Dawn over zero; the story of the atomic bomb. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, second edition, 1972. ISBN 0-8371-6064-2. xv + 289 + viii pp. LCCN UF767 .L3 1972.

Laurikainen:1988:BAP

[Lau88] Kalervo Vihtori Laurikainen. Beyond the atom: the philosophical thought of Wolfgang Pauli. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-19456-8 (US). xix + 234 pp. LCCN QC16.P37 L3813

Laudisa:1995:EBN

[Lau95] Federico Laudisa. Einstein, Bell, and nonseparable realism. The British Journal for the Philosophy of Science, 46(3):309–329, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/46/3/309.full.pdf+html.

Laurikainen:1997:MAE

[Lau97] Kalervo Vihtori Laurikainen. The message of the atoms: essays on Wolfgang Pauli and the unspeakable. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61754-X, 3-642-64457-0, 3-642-60560-5 (e-book). ix + 203 pp. LCCN QC174.12 .L39 1997.

Laudisa:1998:GOM

[Lau98] Federico Laudisa. The great old man’s secrets: Einstein and the foundations of quantum mechanics. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 35(1):125–145, 1998. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Laughlin:1999:NLF

[Lau99] R. B. Laughlin. Nobel Lecture: Fractional quantization. Reviews of Modern Physics, 71(4):863–874, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.71.863;   http://rmp.aps.org/
abstract/RMP/v71/i4/p863_1.

Laughlin:2005:DUR

[Lau05] Robert B. Laughlin. A different universe: reinventing physics from the bottom down. Basic Books, New York, NY, USA, 2005. ISBN 0-465-03828-X (hardcover). xviii + 254 pp. LCCN QC24.5

.L38 2005. URL http://www.loc.gov/catdir/toc/ecip054/
2004028059.html.

Laucht:2012:EGK

[Lau12] Christoph Laucht. Elemental Germans: Klaus Fuchs, Rudolf Peierls, and the making of British nuclear culture 1939–59. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2012. ISBN 0-230-35487-4 (hardcover). xiv + 274 pp. LCCN QC773.3.G7 L38 2012.

Laurence:2013:FMS

[Lau13] William L. Laurence. Fermi measures speed of neutron / Bohr and Einstein at odds. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Laudisa:2015:AEI

[Lau15] Federico Laudisa. Albert Einstein e l’immagine scientifica del mondo. (Italian) [Albert Einstein and the scientific picture of the world], volume 455 of Quality paperbacks. Carocci editore, Roma, Italy, 2015. ISBN 88-430-7721-X. 131 pp. LCCN QC16.E5 L2584 2015. 13.00 EUR. URL http://www.loc.gov/

catdir/toc/casalini14/3027275.pdf.

Lavenda:1985:BM

[Lav85] Bernard H. Lavenda. Brownian motion. Scientific American, 252(2):70–?? (Intl. ed. 56–67), February 1985. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Law:1921:ECI

[Law21] Frank E. Law. From the Einstein Contest — III. Scientific American, 124(10):186–187, March 5, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v124/n10/
pdf/scientificamerican03051921-186.pdf.

Lawler:2000:EPR

[Law00] Andrew Lawler. Einstein probe remains earthbound. Science, 287(5459):1726, March 10, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/287/5459/1726.1.full.

Lawler:2007:BES

[Law07] Andrew Lawler. Beyond Einstein should start with dark energy probe, says panel. Science, 317(5844):1480–1481, September 14, 2007. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/317/5844/1480.full.pdf.

Lazzarotto:2015:EGR

[Laz15] Quentin Lazzarotto, Director. Einstein and General Relativity, a singular story. 52-minute documentary film., November

23, 2015. URL http://ihp.fr/en/brief/documentaryIHP/
Einstein.

London:1939:TOM

[LB39] Fritz London and Edmond Bauer. La théorie de l’observation en mécanique quantique. (French) [The theory of observation in quantum mechanics], volume 775 of Exposés de physique générale, pub. sous la direction de Paul Langevin … III; Actualités scientifiques et industrielles. Hermann and Cie, Paris, France, 1939. 51 + 1 pp. LCCN Q111 .A3 no. 775.

Ladenburg:1949:LTB

[LB49] R. Ladenburg and D. Bershader. On laminar and turbulent boundary layer in supersonic flow. Reviews of Modern Physics, 21(3):510–515, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.510;   http://rmp.
aps.org/abstract/RMP/v21/i3/p510_1. Special issue in honor

of Albert Einstein on his 70th birthday.

Lightman:1990:OLW

[LB90] Alan Paige Lightman and Roberta Brawer. Origins: the lives and worlds of modern cosmologists. Harvard University Press, Cambridge, MA, USA, 1990. ISBN 0-674-64470-0, 0-674-64471-9. xi + 563 pp. LCCN QB981 .L54 1990.

Ljunggren:2002:NPM

[LB02] Bengt Ljunggren and G. W. Bruyn. The Nobel Prize in Medicine and the Karolinska Institute: the story of Axel Key and Alfred Nobel. Karger, Basel, Switzerland, 2002. ISBN 3-8055-7297-2. xi + 232 pp. LCCN R484 .L586 2002.

Levi-Civita:1929:PME

[LC29] T. Levi-Civita. A proposed modification of Einstein’s fieldtheory. Nature, 123(3105):678–679, May 4, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v123/n3105/
pdf/123678a0.pdf.

Levi-Civita:2007:NPG

[LC07] Tullio Levi-Civita. Notion of parallelism on a general manifold and consequent geometrical specification of the Riemannian curvature (excerpts). In Renn et al. [RSS^{+}07b], pages 1081–1088.

ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_48. Originally published as “Nozione di parallelismo in una varietà qualunque e conseguente specificazione geometrica della curvatura riemanniana”, Circolo Mathematico di Palermo, Rendiconti 42, 173–204 (1916).

Langston:1990:GMD

[LCL^{+}90] G. I. Langston, S. R. Conner, J. Lehar, B. F. Burke, and K. W. Weiler. Galaxy mass deduced from the structure of Einstein ring MG1654+1346. Nature, 344(6261):43–45, March 1, 1990. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v344/n6261/
pdf/344043a0.pdf.

Larson:2008:CRF

[LD08] Drake Larson and Nora DeCaprio. The cults of relativity: finding Einstein, Twain and a universe beyond E = mc^{2}. Hellgate Press, Ashland, OR, USA, 2008. ISBN 1-55571-651-2. vii + 193 pp. LCCN QC173.57 .L37 2008. URL http://catdir.loc.

gov/catdir/toc/ecip0727/2007039603.html.

Langevin:1912:GWH

[LdB12a] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911 / sous les auspices de m. E. Solvay. Gauthier-Villars,

Paris, France, 1912. URL https://archive.org/details/
lathoriedurayo00inst.

Langevin:1912:TDR

[LdB12b] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911 / sous les auspices de m. E. Solvay. (French) [Radiation theory and quanta: reports and discussions of the meeting held in Brussels from 30 October to 3 November 1911 / under the auspices of Mr. E. Solvay]. Gauthier-Villars, Paris, France, 1912. URL https:

//archive.org/details/lathoriedurayo00inst.

LeLionnais:1952:DE

[Le 52] F. Le Lionnais. Descartes et Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 5(2):139–154, avril–juin 1952. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23905404.

LePoidevin:2003:TFD

[Le 03] Robin Le Poidevin. Travels in four dimensions: the enigmas of space and time. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-875254-7 (hardcover), 0-19-875255-5 (paperback). xviii + 275 pp. LCCN BD632 .L46 2003; M03.F01152.

Lyon:1984:AFF

[LE84] Fern Lyon and Jacob Evans, editors. Los Alamos, the first forty years. Los Alamos Historical Society, Los Alamos, NM, USA, 1984. ISBN 0-941232-06-9 (paperback). 176 pp. LCCN F804.L6 L67 1984. US$12.95.

Linenthal:1996:HWE

[LE96] Edward Tabor Linenthal and Tom Engelhardt, editors. History wars: the Enola Gay and other battles for the American past. Metropolitan Books, New York, NY, USA, 1996. ISBN 0-8050-4387-X (paperback), 0-8050-4386-1 (hardcover). vi + 295 pp. LCCN E840.4 .H57 1996. URL http://www.loc.gov/

catdir/bios/hol056/96000886.html;   http://www.loc.
gov/catdir/description/hol055/96000886.html.

Leadstone:2007:RA

[Lea07] Stuart Leadstone. Relativity in action. Physics World, 20(8): 19, August 2007. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/20/8/phwv20i8a31.pdf.

Leclercq:1986:NA

[Lec86] Jacques Leclercq, editor. The nuclear age. Le Chêne, Paris, France, 1986. ISBN 2-85108-439-9. 417 pp. LCCN TK1078 .L43 1986.

Ledger:2014:WCM

[Led14] James Ledger. When Chaplin met Einstein: for mixed chamber ensemble. Australian Music Centre, Grosvenor Place, NSW, Australia, 2014. 21-page musical score.

Lee:1993:SGV

[Lee93] You Qin Lee. Science gets the vote, but who was Einstein? Nature, 365(6443):200, September 16, 1993. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v365/n6443/
pdf/365200c0.pdf.

Lee:2006:NSV

[Lee06] Sabine Lee. ‘In no sense vital and actually not even important’? Reality and perception of Britain’s contribution to the development of nuclear weapons. Contemporary British History, 20(2):159–185, June 2006. ISSN 1361-9462 (print), 1743-7997

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/13619460600600680.

Leffler:1995:TDD

[Lef95] Melvyn P. Leffler. Truman’s decision to drop the atomic bomb. IHJ Bulletin: A Quarterly Publication of the International House of Japan, 15(3):1–7, Summer 1995. ISSN 0285-2608.

Leggett:2001:BEC

[Leg01] Anthony J. Leggett. Bose–Einstein condensation in the alkali gases: Some fundamental concepts. Reviews of Modern Physics, 73(2):307–356, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.307;   http://rmp.
aps.org/abstract/RMP/v73/i2/p307_1. See erratum [Leg03].

Leggett:2003:EBE

[Leg03] Anthony J. Leggett. Erratum: Bose–Einstein condensation in the alkali gases: Some fundamental concepts [Rev. Mod. Phys. 73, 307 (2001)]. Reviews of Modern Physics, 75 (3):1083, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1083;   http:
//rmp.aps.org/abstract/RMP/v75/i3/p1083_1. See [Leg01].

Leggett:2004:NLS

[Leg04] Anthony J. Leggett. Nobel Lecture: Superfluid ^{3}He: the early days as seen by a theorist. Reviews of Modern Physics, 76 (3):999–1011, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.76.999;   http://rmp.
aps.org/abstract/RMP/v76/i3/p999_1.

Leggett:2005:QMP

[Leg05] A. J. Leggett. The quantum measurement problem. Science, 307(5711):871–872, February 11, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Leggett:2015:BRB

[Leg15] Don Leggett. Book reviews: A Vision of Modern Science: John Tyndall and the Role of the Scientist in Victorian Culture; ThermoPoetics: Energy in Victorian Literature and Science; Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein; Communicating Physics: The Production, Circulation and Appropriation of Ganot’s Textbooks in France and England, 1851–1887. Historical Studies in the Natural Sciences, 45(3):501–510, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.jstor.org/

stable/10.1525/hsns.2015.45.3.501.

Lehmkuhl:2014:WED

[Leh14a] Dennis Lehmkuhl. Why Einstein did not believe that general relativity geometrizes gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):316–326, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000695.

Lehner:2014:ERH

[Leh14b] Christoph Lehner. Einstein’s realism and his critique of quantum mechanics. In Janssen and Lehner [JL14b], chapter 10, pages 306–353. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Lehmkuhl:2019:GRH

[Leh19] Dennis Lehmkuhl. General relativity as a hybrid theory: the genesis of Einstein’s work on the problem of motion. Studies in History and Philosophy of Modern Physics, 67(??):176–190, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301314.

Leighton:2000:TBR

[Lei00] Ralph Leighton. Tuva Or Bust!: Richard Feynman’s Last Journey. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-32069-3 (paperback), 0-393-02953-0. 260 pp. LCCN QC16.F49 L45 2000.

Leibowitz:2008:HHC

[Lei08] J. R. (Jack R.) Leibowitz. Hidden harmony: the connected worlds of physics and art. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8866-2 (hardcover). xi + 148 pp. LCCN N70 .L454 2008. URL http://www.loc.

gov/catdir/enhancements/fy0805/2007043958-b.html;
http://www.loc.gov/catdir/enhancements/fy0805/2007043958-
d.html;   http://www.loc.gov/catdir/toc/ecip083/2007043958.
html.

Lorentz:1913:RSA

[LEM13] Hendrik Antoon Lorentz, Albert Einstein, and Hermann Minkowski. Das Relativitätsprinzip. Eine Sammlung von Abhandlungen. (German) [The Relativity Principle. A collection of essays], volume 2 of Fortschritte der mathematischen Wissenschaften in Monographien. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1913. 89 pp. Mit Anmerkungen von Arnold Sommerfeld und Vorwort von Otto Blumenthal. [With comments from Arnold Sommerfeld and preface by Otto Blumenthal.].

Lorentz:1923:PRC

[LEM^{+}23] H. A. (Hendrik Antoon) Lorentz, Albert Einstein, H. (Hermann) Minkowski, Hermann Weyl, and Arnold Sommerfeld. The Principle of Relativity: a collection of original memoirs on the Special and General Theory of Relativity. Methuen and Co. Ltd., London, UK, 1923. viii + 216 pp. LCCN QC6 .L63. Translated by W. Perrett and G. B. (George Barker) Jeffery.

Lemaitre:1949:CAR

[Lem49a] Georges Lemaître. Cosmological application of Relativity. Reviews of Modern Physics, 21(3):357–366, July 1949. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.21.357;   http://rmp.aps.org/abstract/RMP/

v21/i3/p357_1. Special issue in honor of Albert Einstein on his 70th birthday.

Lemaitre:1949:CC

[Lem49b] Georges Edward Lemaître. The cosmological constant. In Schilpp [Sch49c], chapter 16, pages 437–456. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Lorentz:1990:RG

[LEM90] H. A. Lorentz, A. Einstein, and H. Minkowski. Das Relativitätsprinzip. (German) [The Principle of Relativity]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, ninth edition, 1990. ISBN 3-519-07306-4. vi + 159 pp. Eine Sammlung von Abhandlungen. [A collection of essays]. With a contribution by H. Weyl, annotations by A. Sommerfeld and forewords by O. Blumenthal.

Lemonick:2009:GSH

[Lem09] Michael D. Lemonick. The Georgian star: how William and Caroline Herschel revolutionized our understanding of the cosmos. Great discoveries. Atlas and Company, New York, NY, USA, 2009. ISBN 0-393-06574-X (hardcover). 199 + 4 pp. LCCN QB35 .L36 2009.

Lorentz:1952:PRC

[LEMW52] H. A. (Hendrik Antoon) Lorentz, A. (Albert) Einstein, H. (Hermann) Minkowski, and H. (Hermann) Weyl. The Principle of Relativity: a Collection of Original Memoirs on the Special and General Theory of Relativity. Dover, New York, NY, USA, 1952. ISBN 0-486-60081-5, 0-486-31840-0 (e-book). viii + 216 pp. LCCN QC 173.55 L67 1952. With notes by A. Sommerfeld. Translated By W. Perrett and G. B. Jeffery.

Lenzen:1939:BRE

[Len39] V. F. Lenzen. Book review: The Evolution of Physics by Albert Einstein; Leopold Infeld. Isis, 30(1):124–125, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225593.

Lenzen:1949:ETK

[Len49] Victor F. Lenzen. Einstein’s theory of knowledge. In Schilpp [Sch49c], chapter 13, pages 355–384. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Lenz:2017:EUE

[Len17] Siegfried Lenz. Einstein überquert die Elbe bei Hamburg: Erzählungen. (German) [Einstein crosses the Elbe near Hamburg: stories]. Atlantik Verlag, Hamburg, Germany, 2017. ISBN 3-455-00049-5. 208 pp. LCCN ????

Leopold:2009:URC

[Leo09] Ellen Leopold. Under the radar: cancer and the Cold War. Critical issues in health and medicine. Rutgers University Press, New Brunswick, NJ, USA, 2009. ISBN 0-8135-4404-1. xi + 284 pp. LCCN RC276 .L46 2009. URL http://www.loc.gov/

catdir/toc/ecip0811/2008007751.html.

Lepore:2018:FEB

[Lep18] Frederick E. Lepore. Finding Einstein’s Brain. Rutgers University Press, New Brunswick, NJ, USA, 2018. ISBN 0-8135-8039-0 (hardcover), 0-8135-8041-2 (e-book). ???? pp. LCCN QC16.E5 L378 2018.

Lerner:1973:ENC

[Ler73] Aaron Bunsen Lerner. Einstein and Newton: a comparison of the two greatest scientists. Lerner Publications Co., Minneapolis, MN, USA, 1973. ISBN 0-8225-0752-8. 234 + 24 pp. LCCN QC16.E5 L39 1973.

Lerner:1991:BBN

[Ler91] Eric J. Lerner. The Big Bang never happened: a startling refutation of the dominant theory of the origin of the universe. Times Books/Random House, New York, NY, USA, 1991. ISBN 0-8129-1853-3. xiii + 466 pp. LCCN QB991.B54 L47 1991.

Leroy:2003:CNP

[Ler03] Francis Leroy, editor. A century of Nobel Prizes recipients: chemistry, physics, and medicine. Marcel Dekker, Inc., New York, NY, USA, 2003. ISBN 0-8247-0876-8. 380 pp. LCCN

Q141 .C252 2003. URL http://www.crcnetbase.com/isbn/
9780824708764.

Lessen:1974:DEP

[Les74] Martin Lessen. Derivation of Einstein’s postulate regarding the energy of light quanta from a more elementary postulate and special relativity. Foundations of Physics, 4(3):417– 419, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708547.

Levy:1980:E

[Lév80] François Lévy. Einstein. Biographies travelling. Duculot, Paris-Gembloux, France, 1980. ISBN 2-8011-0273-3. 143 pp. LCCN QC/16/E5/L48/1980 SS; QC16.E35 L668 1980.

Levi:1998:LLB

[Lev98] Barbara Goss Levi. At long last, a Bose–Einstein condensate is formed in hydrogen. Physics Today, 51(10):17–19, October 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.882393.

Levi:2001:CKW

[Lev01a] Barbara G. Levi. Cornell, Ketterle, and Wieman share Nobel Prize for Bose–Einstein condensates. Physics Today, 54(12):14– 16, December 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.1445529.

Levi:2001:MFN

[Lev01b]  Barbara Goss Levi.  Magnetic forces need not apply:

Bose–Einstein condensates can be made in an optical trap. Physics Today, 54(7):20–22, July 1, 2001. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://physicstoday.scitation.org/doi/full/10.
1063/1.1397387.

Levenson:2003:EB

[Lev03] Thomas Levenson. Einstein in Berlin. Bantam Books, New York, NY, USA, 2003. ISBN 0-553-10344-X. viii + 486 pp. LCCN QC16.E5 L395 2003.

Levenson:2004:EB

[Lev04a] Thomas Levenson. Einstein in Berlin. Bantam Books, New York, NY, USA, 2004. ISBN 0-553-37844-9 (paperback). viii

+ 486 pp. LCCN ???? URL http://www.randomhouse.com/
bantamdell/catalog/display.pperl?isbn=9780553378443.

Leveugle:2004:RPE

[Lev04b] Jules Leveugle. La relativité, Poincaré et Einstein, Planck, Hilbert: Histoire véridique de la théorie de la relativité. (French) [Relativity, Poincaré and Einstein, Planck, Hilbert: the true history of the Theory of Relativity]. Harmattan, Paris, France, 2004. ISBN 2-7475-6862-8. xi + 330 pp. LCCN QC173.52 .L48 2004.

Levin:2007:CCH

[Lev07a] Frank S. Levin, editor. Calibrating the Cosmos: How Cosmology Explains Our Big Bang Universe. Astronomers and #x2019; Universe Series, 1614-659X. Springer Science+Business Media, LLC, New York, NY, USA, 2007. ISBN 0-387-49768-4, 0-387-30778-8. vii + 301 + 4 pp. LCCN QB982 .L48 2007.

Levin:2007:HIU

[Lev07b] Frank S. Levin. Homogeneous, isotropic universes. In Calibrating the Cosmos: How Cosmology Explains Our Big Bang Universe [Lev07a], chapter 6, pages 121–139. ISBN 0-387-49768-4, 0-387-30778-8. LCCN QB982 .L48 2007.

Levenson:2015:HVL

[Lev15] Thomas Levenson. The hunt for Vulcan: … and how Albert Einstein destroyed a planet, discovered relativity, and deciphered the universe. Random House, New York, NY, USA, 2015. ISBN 0-8129-9898-7 (hardcover),, 0-8129-8829-9 (e-book). xv + 229 pp. LCCN BF1724.2.V84 T46 2015.

Levin:2016:BHB

[Lev16] Janna Levin. Black hole blues: and other songs from outer space. Alfred A. Knopf, New York, NY, USA, 2016. ISBN 0-307-95819-1 (hardcover), 0-307-94848-X (paperback). 241 pp.

LCCN QC179 .L48 2016. URL http://www.loc.gov/catdir/
enhancements/fy1605/2015046692-b.html;   http://www.
loc.gov/catdir/enhancements/fy1605/2015046692-d.html.

Lewis:1926:CP

[Lew26] Gilbert N. Lewis. The conservation of photons. Nature, 118(2981):874–875, December 18, 1926. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v118/n2981/pdf/
118874a0.pdf;   http://www.nobeliefs.com/photon.htm.

This paper introduces the name ‘photon’ for Einstein’s ‘light particle’.

Lewis:1979:AEC

[Lew79] Albert C. Lewis. Albert Einstein, 1879–1955: a centenary exhibit of manuscripts, books, and portraits selected from the Humanities Research Center collections, 5 February–31 March 1979, Humanities Research Center, the University of Texas at Austin. Humanities Research Center, the University of Texas at Austin, Austin, TX, USA, 1979. ISBN ???? 38 pp. LCCN QC16.E5 U55 1979. Preface by John A. Wheeler.

Lopez-Fernandez:2006:AAE

[LFKW06] Xose M. Lopez-Fernandez, Andrzej Krawczyk, and Slawomir Wiak. 100th anniversary of Albert Einstein’s papers with regard to field theory. Compel — the International Journal for Computation and Mathematics in Electrical and Electronic Engineering, 25(3):546–550, ???? 2006. ISSN 0332-1649.

Lemons:1997:PLP

[LG97] Don S. Lemons and Anthony Gythiel. Paul Langevin’s 1908 paper “On the Theory of Brownian Motion’ [“Sur la théorie du mouvement brownien,” C. R. Acad. Sci. (Paris) 146, 530 533 (1908)]. American Journal of Physics, 65 (11):1079–1081, November 1997. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

adsabs.harvard.edu/abs/1997AmJPh..65.1079L;   http:/
/ajp.aapt.org/resource/1/ajpias/v65/i11/p1079_s1. See

[Lan08].

Liang:2021:KRB

[LHDH21] Guoqiang Liang, Haiyan Hou, Ying Ding, and Zhigang Hu. Knowledge recency to the birth of Nobel Prizewinning articles: Gender, career stage, and country. Journal of Informetrics, 14(3):Article 101053, ???? 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719301804.

Luo:1998:DNG

[LHF^{+}98] Jun Luo, Zhong-Kun Hu, Xiang-Hui Fu, Shu-Hua Fan, and Meng-Xi Tang. Determination of the Newtonian gravitational constant G with a nonlinear fitting method. Physical Review D (Particles and Fields), 59(4):042001:1–042001:6, December 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918

(electronic), 1538-4500. URL http://link.aps.org/doi/10.
1103/PhysRevD.59.042001.

Lou:2023:SBS

[LHZ^{+}23] Wen Lou, Jiangen He, Lingxin Zhang, Zhijie Zhu, and Yongjun Zhu. Support behind the scenes: the relationship between acknowledgement, coauthor, and citation in Nobel articles. Scientometrics, 128(10):5767–5790, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-023-04803-5.

Luo:2020:DDS

[LHZA20] Yufeng Luo, Roland Haas, Qian Zhang, and Gabrielle Allen. DataVault: a data storage infrastructure for the Einstein Toolkit. CoRR, ??(??):??, ???? 2020. CODEN ????

ISSN ???? URL https://arxiv.org/abs/2012.06635;
https://dblp.org/db/journals/corr/corr2012.html#abs-
2012-06635.

Li:2007:CAE

[Li07] Shang-Jen Li. China and Albert Einstein: The reception of the physicist and his theory in China, 1917–1979. Annals of Science, 64(1):115–118, January 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Libby:1979:UP

[Lib79] Leona Marshall Libby. The Uranium People. Crane Russak, New York, NY, USA, 1979. ISBN 0-8448-1300-1 (Crane Russak), 0-684-16242-3 (Scribners). x + 341 + 16 pp. LCCN QC773.3.U5 L52.

Lide:1967:CSE

[Lid67] David R. Lide. Correction of some erroneous calculations of the Einstein A coefficient for the 18 cm transition of OH. Nature, 213(5077):694–695, February 18, 1967. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v213/n5077/
pdf/213694a0.pdf.

Lieber:1945:ETR

[Lie45] Lillian R. Lieber. The Einstein Theory of Relativity. Farrar & Rinehart, Inc., Toronto, ON, Canada, 1945. x + 324 pp. LCCN ????

Liebscher:1978:DEV

[Lie78] Dierck-Ekkehard Liebscher. Derivation of Einstein’s velocity addition theorem through use of the invariant double ratio. Foundations of Physics, 8(1–2):131–135, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708493.

Lieber:2008:ETR

[Lie08] Lillian R. (Lillian Rosanoff) Lieber, editor. The Einstein theory of relativity: a trip to the fourth dimension. Paul Dry Books, Philadelphia, PA, USA, 2008. ISBN 1-58988-044-7 (paperback). xvii + 350 pp. LCCN QC173.55 .L517 2008. Illustrations by Hugh Gray Lieber. Edited and with a new foreword by David Derbes and Robert Jantzen.

Lightman:1994:ED

[Lig94] Alan Lightman. Einstein’s Dreams. Warner Books, New York, NY, USA, 1994. ISBN 0-446-67011-1. 179 pp. LCCN PS3562.I45397 E38 1994.

Lightman:1997:CGB

[Lig97] Alan Lightman. The contradictory genius: Albert Einstein: A Biography by Albrecht Fölsing, translated by Ewald Osers. New York Review of Books, 44(6):14, April 10, 1997. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/1997/apr/10/the-contradictory-
genius/.

Lightman:2004:ENG

[Lig04] Alan Lightman. Einstein and Newton: Genius compared. Scientific American, 291(3):108–109, September 2004. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v291/n3/pdf/scientificamerican0904-108.pdf.

Lightman:2005:DGBb

[Lig05] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55.D57 .L53. URL http://www.loc.

gov/catdir/enhancements/fy0623/2005040854-b.html;
http://www.loc.gov/catdir/enhancements/fy0623/2005040854-
d.html;   http://www.loc.gov/catdir/enhancements/fy0623/
2005040854-t.html.

Lightman:2018:SSI

[Lig18] Alan P. Lightman. Searching for Stars on an Island in Maine. Pantheon Books, New York, NY, USA, 2018. ISBN 1-101-87186-5 (hardcover), 1-101-87187-3 (e-book). viii + 226 pp. LCCN QB981 .L545 2017.

Lilienthal:1963:CHB

[Lil63] David Eli Lilienthal. Change, hope and the bomb. Princeton University Press, Princeton, NJ, USA, 1963. viii + 168 pp. LCCN HD9698.U52 L5.

Lilburn:1986:AEB

[Lil86] T. Lilburn. ‘Albert Einstein, Berne Patent Office, 1905’. Poetry Australia, ??(105):43, ???? 1986. ISSN 0032-2059.

Lindsay:1956:RBM

[Lin56] R. B. Lindsay. Review: Biographical Memoirs of Fellows of the Royal Society: Vol. 1. Physics Today, 9(10):36, October 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lindsay:1960:BRT

[Lin60] R. Bruce Lindsay. Book review: Trends in Atomic Physics: essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday. Edited by O. R. Frisch, F. A. Paneth, F. Laves, P. Rosbaud. 285 pp. (Vieweg & Sohn, Germany) Interscience Publishers, Inc., New York, 1959. $7.50. Physics Today, 13(9):45, September 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

scitation.aip.org/content/aip/magazine/physicstoday/
article/13/9/10.1063/1.3057115.

Lindbeck:1999:PES

[Lin99] Assar Lindbeck. The prize in economic sciences in memory of Alfred Nobel — 1969–1998, volume 668 of Seminar paper / Institute for International Economic Studies, University of Stockholm. The Institute for International Economic Studies, University of Stockholm, Stockholm, Swedish, 1999. 20 + vi pp. LCCN ????

Lindqvist:2001:NEC

[Lin01] Svante Lindqvist. The Nobel exhibition cultures of creativity: the centennial exhibition of the Nobel Prize, 1901–2001. Minerva [Washington, DC], 39(4):461–465, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012733001619.

Lindorff:2004:PJM

[Lin04a] David P. Lindorff. Pauli and Jung: the meeting of two great minds. Quest Books, Wheaton, IL, USA, 2004. ISBN 0-8356-0837-9. xiv + 299 pp. LCCN BF109.J8 L55 2004.

Linton:2004:EEH

[Lin04b] C. M. (Christopher M.) Linton. From Eudoxus to Einstein: a history of mathematical astronomy. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-82750-7 (hardcover), 0-511-21109-0, 0-511-21286-0, 0-511-20751-4 (e-book), 0-511-21467-7 (e-book). xii + 516 pp. LCCN QB47 .L56 2004; QB47 LIN. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://catdir.loc.
gov/catdir/enhancements/fy0732/2003067569-b.html;
http://www.loc.gov/catdir/description/cam041/2003067569.
html;   http://www.loc.gov/catdir/toc/cam041/2003067569.
html.

Lindley:2006:UU

[Lin06] D. V. (Dennis Victor) Lindley. Understanding uncertainty. Wiley, New York, NY, USA, 2006. ISBN 0-470-04383-0. xv + 250 pp. LCCN QA273 .L534 2006. URL http://www.loc.

gov/catdir/enhancements/fy0741/2006046183-b.html;
http://www.loc.gov/catdir/enhancements/fy0741/2006046183-
d.html;   http://www.loc.gov/catdir/toc/fy0706/2006046183.
html.

Lindley:2007:UEH

[Lin07] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Doubleday, Garden City, NY, USA, 2007. ISBN 0-385-51506-5. vii + 257 pp. LCCN

QC174.17.H4 L56 2007. URL http://www.loc.gov/catdir/
toc/ecip0614/2006017029.html.

Lindley:2008:UEH

[Lin08] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Anchor Books, New York, NY, USA, 2008. ISBN 1-4000-7996-9. viii + 257 pp. LCCN

QC174.17.H4 L56 2007. URL http://www.loc.gov/catdir/
toc/ecip0614/2006017029.html.

Lincoln:2009:QFL

[Lin09] Don Lincoln. The Quantum Frontier: the Large Hadron Collider. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9144-2 (hardcover). xiv + 172 pp. LCCN QC793.5.B62 L56 2009. URL http://www.loc.gov/

catdir/toc/ecip0819/2008022647.html.

Lindley:2013:UU

[Lin13] D. V. (Dennis Victor) Lindley. Understanding uncertainty. Wiley, New York, NY, USA, revised edition, 2013. ISBN 1-118-65012-3 (hardcover), 1-118-65011-5 (e-book). xvi + 393 pp. LCCN QA273 .L534 2013.

Lincoln:2015:YAE

[Lin15] Don Lincoln. 100 years after Einstein’s breakthrough, tensions remain with quantum gravity. Discovered in November 1915, Albert Einstein’s general theory of relativity can predict with astonishing precision the movements of stars and planets. But it remains unable to describe behavior in the subatomic realm. The Christian Science Monitor, ??(??):??, November 29, 2015. ISSN 0882-7729 (print), 1540-4617 (electronic).

URL https://www.csmonitor.com/Science/2015/1129/100-
years-after-Einstein-s-breakthrough-tensions-remain-
with-quantum-gravity.

Lingenhohl:2017:EQD

[Lin17] Daniel Trösch Lingenhöhl. Einstein für Quanten-Dilettanten — Kalender 2018: Ein vergnüglicher Crashkurs in Sachen Naturwissenschaften. (German) [Einstein for quantum dilettantes — calendar 2018: A pleasant crash course in matters of the natural sciences]. KV&H Verlag, Unterhaching, Germany, 2017. ISBN 3-8400-1713-0. 648 pp. LCCN ????

Lipkin:2000:ECR

[Lip00] Harry Lipkin. Einstein could read. Physics World, 13(2):18–19, February 2000. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/13/2/phwv13i2a21.pdf.

Lisle:2016:ES

[Lis16] John Lisle. Einstein up in smoke. Physics in Perspective (PIP), 17(4):354–360, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-015-0171-y.

Litten:1991:ENA

[Lit91] Frederick S. Litten. Einstein and the Noulens Affair. British Journal for the History of Science, 24(4):465–467, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027131.

Little:2009:BRB

[Lit09] Monroe H. Little. Book review: Atomic Tragedy: Henry L. Stimson and the Decision to Use the Bomb against Japan, by Sean L. Malloy. The Journal of American History, 96(1):279–280, June 2009. CODEN ???? ISSN 0021-8723 (print), 1945-2314 (electronic). URL http://jah.

oxfordjournals.org/content/96/1/279.full.pdf+html;
http://jah.oxfordjournals.org/content/96/1/279.short.

Liu:1992:ERT

[Liu92] Chuang Liu. Einstein and relativistic thermodynamics in 1952: a historical and critical study of a strange episode in the history of modern physics. British Journal for the History of Science, 25(2):185–206, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4027298.

Livio:2013:BBD

[Liv13] Mario Livio. Brilliant blunders: from Darwin to Einstein — colossal mistakes by great scientists that changed our understanding of life and the universe. Simon and Schuster, New York, NY, USA, 2013. ISBN 1-4391-9236-7 (hardcover), 1-4391-9238-3 (e-book). 341 pp. LCCN Q172.5.E77 L58 2013.

Lin:2011:SOC

[LJGS11] Y.-J. Lin, K. Jiménez-García, and I. B. Spielman. Spin– orbit-coupled Bose–Einstein condensates. Nature, 471(7336): 83–86, March 2, 2011. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v471/n7336/full/nature09887.html.

Levy-Leboyer:1982:CFS

[LL82] Maurice Levy-Leboyer. The contribution of French scientists and engineers to the development of modern managerial structures in the early part of the twentieth century. In Bernhard et al. [BCS82], pages 283–297. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500232.

Leon:1980:TRR

[LLH^{+}80] X. Leon, P. Langevin, J. Hadamard, A. Einstein, E. Cartan, P. Painleve, P. Levy, J. Perrin, J. Becquerel, L. Brunschvieg, E. Leroy, H. Bergson, E. Meyerson, and H. Pieron. Theory of Relativity + re-publication of debate with Albert Einstein — meeting of April 6, 1922 from Bulletin de la Société Française de Philosophie. Pensée, ??(210):12–29, ???? 1980. ISSN 0031-4773.

Liang:2022:RDI

[LLH22] Guoqiang Liang, Ying Lou, and Haiyan Hou. Revisiting the disruptive index: evidence from the Nobel Prizewinning articles. Scientometrics, 127(10):5721–5730, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-022-04499-z.

Liang:2023:MDM

[LLSH23] Guoqiang Liang, Yaqin Li, Lurui Song, and Chaoguang Huo. Magnitude decrease of the Matthew effect in citations: a study based on Nobel Prize articles. Scientometrics, 128(12):6357– 6371, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04874-4.

Levy:2007:CDC

[LLSS07] S. Levy, E. Lahoud, I. Shomroni, and J. Steinhauer. The a.c. and d.c. Josephson effects in a Bose–Einstein condensate. Nature, 449(7162):579–583, October 4, 2007. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v449/n7162/
full/nature06186.html.

Lifton:1990:GMN

[LM90] Robert Jay Lifton and Eric Markusen. The genocidal mentality: Nazi Holocaust and nuclear threat. Basic Books, New York, NY, USA, 1990. ISBN 0-465-02662-1. xiii + 346 pp. LCCN U263 .L53 1990.

Langdon:1994:NHP

[LM94] John H. (John Howard) Langdon and Mary E. McGann, editors. The natural history of paradigms: science and the process of intellectual evolution. University of Indianapolis Press, Indianapolis, IN, USA, 1994. LCCN BD225 .N38 1993.

Lifton:1995:HAF

[LM95] Robert Jay Lifton and Greg Mitchell. Hiroshima in America: fifty years of denial. Putnam’s Sons, New York, NY, USA, 1995. ISBN 0-399-14072-7. xviii + 425 pp. LCCN D767.25.H6 L42 1995.

Lorentz:1923:RSA

[LME23] H. A. (Hendrik Antoon) Lorentz, H. (Hermann) Minkowski, and Albert Einstein. Das Relativitätsprinzip: eine Sammlung von Abhandlungen. (German) [The Principle of Relativity: a collection of works]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fifth edition, 1923. 89 pp. LCCN ????

Lerager:1988:SCP

[LML88] Jim Lerager, K. Z. (Karl Ziegler) Morgan, and Susan D. Lambert. In the shadow of the cloud: photographs and histories of America’s atomic veterans. Fulcrum, Golden, CO, USA, 1988. ISBN 1-55591-030-0. 116 pp. LCCN UB369 .L47 1988.

Leibfried:2015:BWE

[LMOS15] Stephan Leibfried, Christoph Markschies, Ernst Osterkamp, and Günter Stock. Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy]. De Gruyter Akademie Forschung, Berlin, Germany, 2015. ISBN 3-11-037598-2 (hardcover). 517 pp. LCCN AS182.B38 B477 2015.

Loach:1998:CCU

[Loa98] Judi Loach. Le Corbusier and the creative use of mathematics. British Journal for the History of Science, 31(2):185–215, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027762.

Lodge:1920:ESS

[Lod20a] Oliver Lodge. Einstein’s shift of spectral lines. Nature, 106 (2664):373, November 18, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v106/n2664/pdf/106373a0.pdf.

Lodge:1920:TES

[Lod20b] Oliver Lodge. Testing Einstein’s shift of spectral lines. Nature, 106(2661):280, October 28, 1920. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v106/n2661/pdf/
106280a0.pdf.

Lodge:1925:ESD

[Lod25] Oliver Lodge. Einstein shift and Doppler shift. Nature, 116 (2930):938, December 26, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v116/n2930/pdf/116938d0.pdf.

Loewenberg:1988:BRC

[Loe88a] P. Loewenberg. Book review: The Collected Papers of Albert Einstein, Vol 1, the Early Years, 1879–1902 J. Stachel. Science, 239(4839):510–512, January 29, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Loewenberg:1988:EHY

[Loe88b] Peter Loewenberg. Einstein in his youth: Book review: The Collected Papers of Albert Einstein. Vol. 1, The Early Years 1879–1902. Science, 239(4839):510–512, January 29, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/239/
4839/510.full.pdf.

Logan:1996:CM

[Log96] Jonothan Logan. The critical mass: As Allied scientists established the feasibility of an atomic bomb, Germany’s leading theorist came to a mistaken conclusion. American Scientist, 84(3):

263–277, May/June 1996. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/29775672.

Lombardo:2005:NSS

[Lom05] D. J. Cirilo Lombardo. New spherically symmetric monopole and regular solutions in Einstein–Born–Infeld theories. Journal of Mathematical Physics, 46(4):042501, April 2005. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p042501_s1.

London:1937:CNS

[Lon37] Fritz London. Une conception nouvelle de la supraconductibilité. (French) [A new concept of superconductivity]. Hermann and cie, Paris, France, 1937. 2 + 7–80 + 1 pp. LCCN QC611 .L6 1937. Translated by Edmond Bauer and J. Winter.

London:1938:PLH

[Lon38a] F. London. The λ-phenomenon of liquid helium and the Bose–Einstein degeneracy. Nature, 141(3571):643–644, April 9, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v141/n3571/pdf/141643a0.pdf.

London:1938:BEC

[Lon38b] F. London. On the Bose–Einstein condensation. Physical Review, 54(11):947–954, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

London:1950:S

[Lon54] Fritz London. Superfluids. Structure of matter series. Wiley, New York, NY, USA, 1950–1954. ???? pp. LCCN QC611 .L58.

London:1961:S

[Lon61] Fritz London. Superfluids. Dover, New York, NY, USA, second edition, 1961. ???? pp. LCCN QC611 .L582.

Longair:1984:TCP

[Lon84] Malcolm S. Longair. Theoretical concepts in physics: an alternative view of theoretical reasoning in physics for final-year undergraduates. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25550-3, 0-521-27553-9 (paperback). xiii + 366 pp. LCCN QC20 .L66 1984. URL http://www.loc.

gov/catdir/description/cam022/83018928.html;   http:
//www.loc.gov/catdir/toc/cam021/83018928.html.

Longair:2003:TCP

[Lon03] Malcolm S. Longair. Theoretical concepts in physics: an alternative view of theoretical reasoning in physics. Cambridge University Press, Cambridge, UK, second edition, 2003. ISBN 0-521-52878-X (paperback), 0-521-82126-6. xvii + 569 pp. LCCN QC20 .L64 2003. URL http://www.loc.gov/

catdir/description/cam031/2002073612.html;   http://
www.loc.gov/catdir/samples/cam041/2002073612.html;
http://www.loc.gov/catdir/toc/cam031/2002073612.html.

Longair:2013:CCH

[Lon13] M. S. Longair. The cosmic century: a history of astrophysics and cosmology. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-66936-7. xvi + 545 pp. LCCN QB461 .L66 2013.

Lorentz:1914:RDV

[Lor14] H. A. Lorentz. Das relativitätsprinzip: drei Vorlesungen gehalten in Teylers Stiftung zu Haarlem. (German) [The Principle of Relativity: three lectures held in Teylers Stiftung in Haarlem]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1914. 52 pp. LCCN ???? Edited by W. H. Keesom.

Lorentz:1917:ETG

[Lor17] H. A. Lorentz. On Einstein’s Theory of Gravitation. Proc. Amsterdams Acad., XIX:1341–??, 1917.

Lorentz:1920:ETR

[Lor20] H. A. (Hendrik Antoon) Lorentz. The Einstein theory of relativity: a concise statement. Brentano’s, New York, NY, USA, 1920. 64 pp. LCCN QC6 .L5.

Lorentz:1921:MME

[Lor21] Hendrik A. Lorentz. The Michelson–Morley experiment and the dimensions of moving bodies. Nature, 106(2677):793–795, February 17, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). Translated by Robert W. Lawson from the German manuscript. Special issue on Relativity. This is one of only three papers by foreign authors in this issue; the others are [Ein21b, Wey21].

Lorentz:1927:LTP

[Lor31] H. A. (Hendrik Antoon) Lorentz. Lectures on theoretical physics, delivered at the University of Leiden. Macmillan Publishing Company, New York, NY, USA, 1927–1931. ???? pp. LCCN QC3 .L78. Translated by Silberstein, Ludwik, b 1872 and Trivelli, A. P. H. (Adrian Peter Herman), b 1879, Bremekamp, Hendrik, b 1880, Bruins, Eva Dina, b 1885, Johanna Reudler, Tettje Klazina (Jolles) Clay, Claude August Crommelin, A. D. (Adriaan Daniël) Fokker and Geertruida Luberta de Haas-Lorentz.

Lorentz:2007:CG

[Lor07a] Hendrik A. Lorentz. Considerations on gravitation. In Renn et al. [RSS^{+}07a], pages 113–126. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_13.

Lorentz:2007:ONQ

[Lor07b] Hendrik A. Lorentz. Old and new questions in physics (excerpt). In Renn et al. [RSS^{+}07a], pages 287–301. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/

978-1-4020-4000-9_21. Originally published as “Alte unde neue Fragen der Physik”, Physikalische Zeitschrift 11, 1234– 1257 (1910).

Lorge:2008:AMR

[Lor08] Peter Allan Lorge. The Asian military revolution: from gunpowder to the bomb. New approaches to Asian history. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-84682-X (hardcover), 0-521-60954-2 (paperback). xi + 188 pp. LCCN M08.E09475; DS33.7 .L66 2008. URL http://www.loc.

gov/catdir/enhancements/fy0806/2007051673-b.html;
http://www.loc.gov/catdir/enhancements/fy0806/2007051673-
d.html;   http://www.loc.gov/catdir/enhancements/fy0806/
2007051673-t.html.

Lorentz:2009:ETR

[Lor09] H. A. (Hendrik Antoon) Lorentz. The Einstein theory of relativity. WLC, ????, 2009. ISBN 1-55742-712-7 (paperback). 28 pp. LCCN QC6 .L5 1920.

Lorentz:2010:ETRb

[Lor10a] H. A. Lorentz. The Einstein Theory of Relativity: a Concise Statement. Mobi Classics. MobileReference.com, Boston, MA, USA, 2010. ISBN 1-60778-566-8. 64 pp. LCCN QC173.55.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=543089.

Lorentz:2010:ETRa

[Lor10b] H. A. (Hendrik Antoon) Lorentz. The Einstein theory of relativity: a concise statement. Kessinger Pub., LaVergne, TN, USA, 2010. ISBN 1-104-38753-0. 64 pp. LCCN ????

Lorentz:2010:ETRc

[Lor10c] Hendrik Antoon Lorentz. Einstein theory of relativity: a concise statement. General Books, ????, 2010. ISBN 1-154-44781-2. ???? pp. LCCN ????

Lorenzen:1976:RER

[Lor77] P. Lorenzen. Eine Revision der Einsteinschen Revision. (German) [A revision of the Einsteinian revision]. Philosophia Naturalis, 16(4):383–391, 1976/77. CODEN ???? ISSN 0031-8027.

Loudon:1983:QTL

[Lou83] Rodney Loudon. The quantum theory of light. Oxford science publications. Clarendon Press, Oxford, UK, second edition, 1983. ISBN 0-19-851155-8 (paperback), 0-19-851152-3 (hardcover). xiv + 393 pp. LCCN QC446.2 .L68 1983.

Lovejoy:1986:SLS

[Lov86] D. Lovejoy. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. Science & Society, 50(1):117–121, Spring 1986. ISSN 0036-8237 (print), 1943-2801 (electronic).

Annette:2012:BWM

[LOvT12] Annette Lykknes, Donald Opitz, and Brigitte van Tiggelen, editors. For better or for worse? Collaborative couples in the sciences, volume 44 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 3-0348-0285-4, 3-0348-0286-2 (e-book). LCCN ????

Lowdin:1978:LTP

[Löw78] Per-Olov Löwdin. “Low Temperature Physics and Cosmic Microwave Background Radiation”: the 1978 Nobel prizes in physics. In Enciclopedia della Scienca e della Technica (EST) Yearbook, page ?? Ed. Arnoldo Mondadow, Milano, Italy, 1978. ISBN ???? LCCN ????

Lowdin:1998:SCF

[Löw98] Per-Olov Löwdin. Some comments on the foundations of physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-2913-5. xvi + 121 pp. LCCN QC173.58 .L69 1998.

Low:2001:ESN

[Low01] Morris Low. From Einstein to Shirakawa: the Nobel Prize in Japan. Minerva [Washington, DC], 39(4):445–460, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012740020275.

Leffler:2005:OCW

[LP05] Melvyn P. Leffler and David S. Painter, editors. Origins of the Cold War: an international history. Rewriting histories. Routledge, London, UK, second edition, 2005. ISBN 0-415-34109-4 (hardcover), 0-415-34110-8 (paperback). xv + 352 pp.

LCCN D842 .O86 2005. URL http://www.loc.gov/catdir/
enhancements/fy0651/2004051306-d.html.

Labbe:2011:E

[LP11] Brigitte Labbé and Michel Puech. Einstein. Milan jeunesse, Toulouse, France, 2011. ISBN 2-7459-4888-1. ???? pp. LCCN ????

Lightman:1975:PBR

[LPPT75] Alan P. Lightman, William H. Press, Richard H. Price, and Saul A. Teukolsky, editors. Problem book in Relativity and Gravitation. Princeton University Press, Princeton, NJ, USA, 1975. ISBN 0-691-08160-3, 0-691-08162-X (paperback). xiv + 599 pp. LCCN QC173.55 .P76.

Levinovitz:2001:NPF

[LR01] Agneta Wallin Levinovitz and Nils Ringertz, editors. The Nobel Prize: the first 100 years. Imperial College Press, London, UK, 2001. ISBN 981-02-4854-7 (e-book), 981-02-4664-1, 981-02-4665-X (paperback). viii + 235 pp. LCCN AS911.N9 N586 2001.

Levinovitz:2002:BNN

[LR02] Agneta Wallin Levinovitz and Nils Ringertz. Bai nian Nuobei’er jiang. Shi jie ke ji chu ban gong si, Singapore, 2002. ISBN 981-02-4757-5. viii + 227 pp. LCCN AS911.N9 N586125 2002.

Liu:2014:CAD

[LR14] Yuxian Liu and Ronald Rousseau. Citation analysis and the development of science: a case study using articles by some Nobel prize winners. Journal of the Association for Information Science and Technology, 65(2):281–289, February 2014. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Logan:1991:LEH

[LRD^{+}91] Jonothan L. Logan, Helmut Rechenberg, Max Dresden, A. Van Der Ziel, and Mark Walker. Letters to the Editor: Heisenberg, Goudsmit and the German “A-bomb”. Physics Today, 44(5): 13, 15, 90–92, 94–96, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v44/i5/p13_s1. Let-

ters in response to [?], including Walker’s rebuttal.

Li:2022:LIR

[LRH22] Xian Li, Ronald Rousseau, and Xiaojun Hu. Is low interdisciplinarity of references an unexpected characteristic of Nobel Prize winning research? Scientometrics, 127(4):2105– 2122, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04290-0.

Lachieze-Rey:2015:EPR

[LRL15] Marc Lachièze-Rey and Ludovic Ligot. Einstein à la plage: la relativité dans un transat. (French) [Einstein on the beach: Relativity in a deckchair]. Dunod, Paris, France, 2015. ISBN 2-10-072223-9 (paperback). 159 pp. LCCN ????

Lachieze-Rey:2017:EPR

[LRL17] Marc Lachièze-Rey and Ludovic Ligot. Einstein à la plage: la relativité dans un transat. (French) [Einstein at the beach: relativity in a deckchair]. ????, ????, nouvelle edition, 2017. ISBN 2-10-076339-3 (paperback). 174 pp. LCCN ????

Lehner:2012:ECW

[LRS12] Christoph Lehner, Jürgen Renn, and Matthias Schemmel, editors. Einstein and the changing worldviews of physics, volume 12 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 0-8176-4939-5, 0-8176-4940-9 (e-book). xii + 363 pp. LCCN QC173.6 .E46 2012.

Lanouette:1992:GSB

[LS92] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. C. Scribner’s Sons, New York, NY, USA, 1992. ISBN 0-684-19011-7. xix + 587 + 16 pp. LCCN QC16.S95 L36 1992. US$35.00 (US$44.50 Can.).

URL http://adsabs.harvard.edu/abs/1994gsbl.book.....
L.

Logan:1993:SCH

[LS93] Jonothan L. Logan and Robert Serber. Scientific correspondence: Heisenberg and the bomb. Nature, 362(6416):117, March 11, 1993. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v362/n6416/full/362117a0.html.

Lanouette:1994:GSB

[LS94] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-46888-7 (paperback). xix + 587 pp. LCCN QC16.S95

L36 1994. URL http://www.loc.gov/catdir/description/
uchi053/94012738.html.

Lohmeier:2005:EAK

[LS05] Dieter Lohmeier and Bernhardt Schell, editors. Einstein, Anschütz und der Kieler Kreiselkompass der Briefwechsel zwischen Albert Einstein und Hermann Anschütz-Kaempfe und andere Dokumente. (German) [Einstein, Anschütz and the Kiel gyro compass]. Raytheon Marine, Kiel, Germany, second edition, 2005. ISBN 3-00-016598-3. 259 pp. LCCN QC16.E5 A4 2005b. With a contribution by Jobst Broelmann.

Linguerri:2008:EPI

[LS08] Sandra Linguerri and Raffaella Simili. Einstein parla italiano: itinerari e polemiche. (Italian) [Einstein speaks Italian: itineraries and controversy]. Pendragon, Bologna, Italia, 2008.

ISBN 88-8342-530-8. x + 359 pp. LCCN ???? 20.00 EUR. URL

http://www.gbv.de/dms/bsz/toc/bsz283148705inh.pdf.

Lindley:1994:AUT

[LSF94] D. V. (Dennis Victor) Lindley, Adrian F. M. Smith, and P. R. (Peter R.) Freeman, editors. Aspects of uncertainty: a tribute to D. V. Lindley. Wiley series in probability and mathematical statistics. Wiley, New York, NY, USA, 1994. ISBN 0-471-94347-9. xviii + 392 pp. LCCN QA279.4 .A87 1994.

US$63.95. URL http://www.zentralblatt-math.org/zmath/
en/search/?an=0827.00022.

Lehmkuhl:2016:TTS

[LSS16] Dennis Lehmkuhl, Gregor Schiemann, and Erhard Scholz, editors. Towards a theory of spacetime theories. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 1-4939-3209-8 (hardcover), 1-4939-3210-1 (e-book). LCCN QC173.59.S65 T69 2017.

Levine:2009:TNM

[LSW09] Frances Levine, Louise Stiver, and Marta Weigle, editors. Telling New Mexico: a new history. Museum of New Mexico Press, Santa Fe, NM, 2009. ISBN 0-89013-552-5 (hardcover), 0-89013-556-8 (paperback). 483 pp. LCCN F796.

Lettevall:2012:NTC

[LSW12] Rebecka Lettevall, Geert Somsen, and Sven Widmalm, editors. Neutrality in twentieth-century Europe: intersections of science, culture, and politics after the First World War, volume 18 of Routledge studies in cultural history. Routledge, London, UK, 2012. ISBN 0-203-11679-8, 0-415-89377-1. ISSN 1404-7586. ix + 351 pp. LCCN D723 .N49 2012. URL http://libanswers.

liverpool.ac.uk/faq/182315.

Lingenhohl:2015:EFQ

[LT15] Daniel Lingenhöhl and Thomas Trösch. Einstein für Quanten-Dilettanten 2016: Ein vergnüglicher Crashkurs in Sachen Naturwissenschaften. (German) [Einstein for quantum dilettantes 2016: A more fun crash course in natural sciences]. KV&H Verlag, ????, 2015. ISBN 3-8400-1139-6. 648 pp. LCCN ????

Lecuyer:2013:LMI

[LU13] Christophe Lécuyer and Takahiro Ueyama. The logics of materials innovation: The case of gallium nitride and blue light emitting diodes. Historical Studies in the Natural Sciences, 43 (3):243–280, June 2013. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/
pdfplus/10.1525/hsns.2013.43.3.243.pdf.

Lucas:2011:BS

[Luc11] Amand Lucas. The Bomb and the Swastika. CreateSpace Independent Publishing, ????, 2011. 68 pp. URL http://

www.amazon.com/The-Bomb-Swastikahistorys-scientists/

dp/1466426675. A play in four acts.

Lucas:20xx:LEA

[Lucxx] Amand A. Lucas. La Lettre d’Einstein au Président Roosevelt. ????, ????, 20xx. ???? pp. Scénario d’une pièce de théâtre.

Ludovici:1957:NPW

[Lud57] Laurence James Ludovici. Nobel Prize winners. Arco Publishers, London, UK, 1957. 226 pp. LCCN ????

Ludyk:2013:EMF

[Lud13] Günter Ludyk. Einstein in matrix form: exact derivation of the theory of Special and General Relativity without tensors. Graduate texts in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-35797-0, 3-642-35798-9 (e-book). LCCN QC173.6 .L83 2013.

Luiggi:2011:QAG

[Lui11] Cristina Luiggi. Q&A: Aging geniuses: A new study shows that over the past century, the age at which scientists produce their most valuable work is increasing. The Scientist (Philadelphia, PA), ??(??):??, November 8, 2011. ISSN 0890-3670 (print), 1945-5127 (electronic). URL http://www.the-

scientist.com/?articles.view/articleNo/31369/. News

report on [JW11].

Lukoschik:2002:RII

[Luk02] Bernd Lukoschik. Realismus und Instrumentalismus im Weltbild des frühen Einstein. (German) [Realism and instrumentalism in the world view of the early Einstein]. Philosophia Naturalis, 39(1):111–140, 2002. CODEN ???? ISSN 0031-8027.

Luminet:2011:GOE

[Lum11] Jean-Pierre Luminet. Golden oldie editorial: Editorial note to: Georges Lemaître, the beginning of the world from the point of view of quantum theory. General Relativity and Gravitation, 43(10):2911–2928, ???? 2011. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL http://www.

springerlink.com/content/l596v5992421238r/.

Lundstrom:1974:ANS

[Lun74] Ragnhild Lundström. Alfred Nobel som internationell företagare. Den nobelska sprängämnesindustrin 1864–1886. (Swedish) [Alfred Nobel’s international business. The Nobel explosives industry 1864–1886], volume 10 of Uppsala studies in economic history. Acta Universitatis Upsaliensis. University of Upsala, Uppsala, Sweden, 1974. ISBN 91-554-0197-X. 272 pp. LCCN HD9663.S82 L85.

Lundqvist:1992:NLP

[Lun92] Stig Lundqvist, editor. Nobel lectures. Physics, 1971–1980: including presentation speeches and laureates’ biographies. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992. ISBN 981-02-0726-3. xi + 603 pp. LCCN QC6.2 .P46 1992.

Lutz:2005:AES

[Lut05] Barbara Anna Lutz. Albert Einstein’s summer house by Konrad Wachsmann in Caputh. Architectura — Zeitschrift für Geschichte der Baukunst, 35(2):178–198, ???? 2005. ISSN 0044-863X.

Ladenburg:1934:ANP

[LW34] Rudolf Ladenburg and Eugene Wigner. Award of the Nobel Prizes in Physics to Professors Heisenberg, Schrödinger and Dirac. The Scientific Monthly, 38(1):86–91, January 1934. CO-DEN SCMOAA. ISSN 0096-3771. URL http://www.jstor.

org/stable/15534.

Lewis:1971:APT

[LW71] Richard S. Lewis and Jane Wilson, editors. Alamogordo plus twenty-five years; the impact of atomic energy on science, technology, and world politics. Viking, New York, NY, USA, 1971.

ISBN 0-670-11151-1. vi + 281 pp. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

Lewis:1988:CBB

[LX88] John Wilson Lewis and Litai Xue. China Builds the Bomb, volume 280 of ISIS Studies in International Security and Arms Control. Stanford University Press, Stanford, CA, USA, 1988. ISBN 0-8047-1452-5, 0-8047-1841-5 (paperback). xviii + 329 + 16 pp. LCCN U263.L49 1.

Liu:1995:CNY

[LY95] C. S. (Chao Shiuan) Liu and Shing-Tung Yau, editors. Chen Ning Yang: a great physicist of the twentieth century. International Press, Cambridge, MA, USA, 1995. ISBN 1-57146-001-2. LCCN QC16.Y364 A3 1995.

Lynch:1933:CAE

[Lyn33] Arthur Lynch. The case against Einstein. Nature, 131(3313): 622, April 29, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v131/n3313/pdf/131622b0.pdf.

M:1925:TTE

[M.25] D. B. M. The tyranny of time: Einstein or Bergson? Nature, 116(2907):91–92, July 18, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v116/n2907/pdf/116091a0.pdf.

MendezCarpio:2008:AHN

[M08] Jorge Méndez Carpio. Azcona Hoyo, “Nobel de la paz y la democracia”. J. A. Méndez Carpio, Tegucigalpa, Honduras, 2008. ISBN 99926-45-24-5, 99926-11-13-8. 234 pp. LCCN F1508.33.A97 M36 2008.

Millikan:1931:RRD

[MA31] Robert A. Millikan and Walter S. Adams. The reason and the results of Dr. Einstein’s visit to the California Institute of Technology. Science, 73(1893):380–381, April 10, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/73/
1893/380.full.pdf.

Maas:2007:EEC

[Maa07] Ad Maas. Einstein as engineer: The case of the little machine. Physics in Perspective (PIP), 9(3):305–328, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0308-0;   http://www.
springerlink.com/content/3x7341xrt45047gx/fulltext.

pdf. See also [SU06, SWU09].

Maas:2013:BRB

[Maa13a] Ad Maas. Book review: József Illy, The Practical Einstein: Experiments, Patents, Inventions. Isis, 104(3):627–628, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674492.pdf.

Maas:2013:BRJ

[Maa13b] Ad Maas. Book review: József Illy, The Practical Einstein: Experiments, Patents, Inventions. Isis, 104(3):627–628, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674492.pdf.

MacKinnon:1977:HMR

[Mac77] Edward MacKinnon. Heisenberg, models, and the rise of matrix mechanics. Historical Studies in the Physical Sciences, 8(??): 137–188, ???? 1977. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757370.

MacPherson:1985:STF

[Mac85] Malcolm MacPherson. A small town with a footnote to history: Hidden below, the Nazi A-bomb lab Haigerloch. New York Times, ??(??):10, 47, November 3, 1985. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/111210585.

Macrossan:1986:NRB

[Mac86] M. N. Macrossan. A note on Relativity before Einstein. The British Journal for the Philosophy of Science, 37(2):232–234, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

MacKenzie:1993:IAH

[Mac93] Donald A. MacKenzie. Inventing accuracy: a historical sociology of nuclear missile guidance. Inside technology. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-63147-4 (paperback), 0-262-13258-3. xiii + 464 pp. LCCN UG1312.B34 M33 1993.

MacDiarmid:2001:NLS

[Mac01] Alan G. MacDiarmid. Nobel Lecture: “synthetic metals”: a novel role for organic polymers. Reviews of Modern Physics, 73 (3):701–712, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.701;   http://rmp.
aps.org/abstract/RMP/v73/i3/p701_1.

Macalister:2008:EGW

[Mac08] Todd Macalister. Einstein’s god: a way of being spiritual without the supernatural. Apocryphile, Berkeley, CA, USA, 2008. ISBN 1-933993-65-0. iv + 76 pp. LCCN ????

MacCallum:2013:ESE

[Mac13] Malcolm MacCallum. Exact solutions of Einstein’s equations. Scholarpedia, 8(12):8584, ???? 2013. CODEN ???? ISSN 1941-

6016. URL https://dblp.org/db/journals/scholarpedia/
scholarpedia8.html#MacCallum13.

Maddox:1983:WWH

[Mad83] John Maddox. Where was Hoyle? Nature, 305(5837):750, October 27, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/nature/
journal/v305/n5937/pdf/305750b0.pdf.

Maddox:1989:EMD

[Mad89] John Maddox. Einstein’s much delayed little joke. Nature, 337 (6209):685, February 23, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v337/n6209/pdf/337685a0.pdf.

Maddox:1995:BET

[Mad95a] John Maddox. Beyond Einstein’s theory of gravitation? Nature, 374(6525):759, April 27, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v374/n6525/pdf/374759a0.pdf.

Maddox:1995:WVH

[Mad95b] Robert James Maddox. Weapons for victory: the Hiroshima decision fifty years later. University of Missouri Press, Columbia, MO, USA, 1995. ISBN 0-8262-1037-6. 215 pp. LCCN

D767.25.H6 M23 1995. URL http://www.h-net.org/review/
hrev-a0a0m2-aa.

Maddox:2004:WVH

[Mad04] Robert James Maddox. Weapons for victory: the Hiroshima decision. University of Missouri Press, Columbia, MO, USA, 2004. ISBN 0-8262-1562-9 (paperback). xvii + 215 pp. LCCN D767.25.H6 M23 2004.

Maddox:2007:HHM

[Mad07] Robert James Maddox, editor. Hiroshima in history: the myths of revisionism. University of Missouri Press, Columbia, MO, USA, 2007. ISBN 0-8262-1732-X (hardcover). vi + 213 pp. LCCN D767.25.H6 H6335 2007. URL http://www.loc.gov/

catdir/toc/ecip076/2006100945.html.

Magill:1989:NPW

[Mag89] Frank Northen Magill. The Nobel prize winners. Physics. Salem Press, Pasadena, CA, USA, 1989. ISBN 0-89356-557-1 (vol. 1), 0-89356-559-8 (vol. 2), 0-89356-560-1 (vol. 3). il + 1364 +

xix pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/
12148/cb374181682. Three volumes.

Maglic:1995:ZSS

[Mag95] Rastko Maglić, editor. Zbornik sa savetovanja doprinos Mileve Ajnštajn-Marić nauci: 13.-14. maj 1994. Matica srpska. (Serbian) [Proceedings from the conference on Mileva Einstein-Marić’s science: 13.-14. May 1994. Matica srpska]. Univerzitet u Novom Sadu, Novi Sad, Serbia, 1995. ISBN ???? 182 pp. LCCN ????

Magueijo:2003:FTS

[Mag03] João Magueijo. Faster than the speed of light: the story of a scientific speculation. Perseus Publishers, Cambridge, MA, USA, 2003. ISBN 0-7382-0525-7 (Perseus), 0-14-200361-1 (Penguin). vi + 279 pp. LCCN QC407 .M34 2003. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/toc/fy044/2002112394.html.

Magueijo:2009:BDE

[Mag09] João Magueijo. A brilliant darkness: the extraordinary life and disappearance of Ettore Majorana, the troubled genius of the nuclear age. Basic Books, New York, NY, USA, 2009. ISBN 0-465-00903-4. xxi + 280 pp. LCCN QC774.M34 M35 2009.

Mahner:2001:SRS

[Mah01] Martin Mahner, editor. Scientific realism: selected essays of Mario Bunge. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-892-5. 438 pp. LCCN Q175.32 .R42 B87 2001.

Mahon:2003:MWC

[Mah03] Basil Mahon. The man who changed everything: the life of James Clerk Maxwell. Wiley, New York, NY, USA, 2003. ISBN 0-470-86088-X, 0-470-86171-1 (paperback). xx + 226 + 16 pp. LCCN QC16.M4 M34 2003. URL http://www.loc.gov/

catdir/bios/wiley047/2004426225.html;   http://www.
loc.gov/catdir/description/wiley042/2004426225.html;
 http://www.loc.gov/catdir/toc/wiley041/2004426225.
html.

Mahaffey:2009:AAN

[Mah09] James A. Mahaffey. Atomic Awakening: a New Look at the History and Future of Nuclear Power. Pegasus Books, New York, NY, USA, 2009. ISBN 1-60598-040-4. xxiii + 344 pp. LCCN TK9145 .M34 2009; TK9145.M34.

Mahaffey:2014:AAH

[Mah14] James Mahaffey. Atomic Accidents: a History of Nuclear Meltdowns and Disasters: from the Ozark Mountains to Fukushima. Pegasus, New York, NY, USA, 2014. ISBN 1-60598-680-1. xxi + 442 + 16 pp. LCCN ????

Mahaffey:2015:AAH

[Mah15] James A. Mahaffey. Atomic accidents: a history of nuclear meltdowns and disasters: from the Ozark Mountains to Fukushima. Pegasus Books, New York, NY, USA, 2015. ISBN 1-60598-680-1. xxi + 442 + 16 pp. LCCN TK9152 .M34 2015; TK9152 .M284 2015.

Mahaffey:2017:AAS

[Mah17] James A. Mahaffey. Atomic Adventures: Secret Islands, Forgotten N-rays, and Isotopic Murder — a Journey Into the Wild World of Nuclear Science. Pegasus Books Ltd., New York, NY, 2017. ISBN 1-68177-421-6 (hardcover). xxxiii + 363 + 24 pp. LCCN TK9145 .M315 2017.

Maiocchi:1990:CBM

[Mai90] Roberto Maiocchi. The case of Brownian motion. British Journal for the History of Science, 23(3):257–283, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.jstor.org/stable/4026755.

MaierRainwater:1999:SEH

[Mai99] Terese Sian Maier Rainwater. 60 shakes: an educational history of the making of the atomic bomb, Los Alamos, New Mexico 1943–1946. Ph.D. dissertation, University of Kansas, 1999. vi + 187 pp.

Malamud:1965:BRA

[Mal65] Herbert Malamud. Book review: Albert Einstein and the cosmic world order, by Cornelius Lanczos, Interscience Publishers, vi + 139 pp. Physics Today, 18(6):54, June 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Mallett:1979:HMI

[Mal79] Ronald L. Mallett. Higgs mechanism and the inverse Einstein– Infeld–Hoffman problem. Journal of Mathematical Physics, 20 (5):921–922, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i5/p921_s1.

Malik:1985:YHN

[Mal85] John Malik. The yields of the Hiroshima and Nagasaki nuclear explosion. Technical report LA-8819, Los Alamos National Laboratory, Los Alamos, NM, USA, 1985.

Maltese:1991:RRT

[Mal91] Giulio Maltese. The rejection of the Ricci tensor in Einstein’s first tensorial Theory of Gravitation. Archive for History of Exact Sciences, 41(4):363–381, December 1991. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=41&issue=4&spage=363.

Malmstrom:1997:C

[Mal97] B. G. (Bo G.) Malmström, editor. Chemistry, 1991–1995. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2679-9, 981-02-2680-2 (paperback). ix + 296 pp. LCCN QD39 .C4862 1997.

Maltese:2000:LER

[Mal00] Guilio Maltese. The late entrance of Relativity into Italian scientific community (1906–1930). Historical Studies in the Physical and Biological Sciences, 31(1):125–173, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Malloy:2008:ATH

[Mal08] Sean L. (Sean Langdon) Malloy. Atomic tragedy: Henry L. Stimson and the decision to use the bomb against Japan. Cornell University Press, Ithaca, NY, USA, 2008. ISBN 0-8014-4654-6 (hardcover), 0-8014-7629-1 (paperback). xi + 233 pp. LCCN D767.25.H6 M26 2008.

Malloy:2009:CSN

[Mal09] Sean L. Mallo. Contemporary scholarship and new light on the A-bomb decision. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2): 91–104, December 2009. CODEN HISCDU. ISSN 0285-4821.

Malloy:2012:VPW

[Mal12] Sean L. Malloy. ‘A very pleasant way to die’: Radiation effects and the decision to use the atomic bomb against Japan. Diplomatic History, 33(3):515–545, April 2012. ISSN 0145-2096 (print), 1467-7709 (electronic). URL http://dh.

oxfordjournals.org/content/36/3/515.abstract.

Mannoury:1953:CPA

[Man55] G. Mannoury. The cultural phenomenon Albert Einstein. Synthese, 9(1):438–441, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00567430.

Manchester:1964:AK

[Man64] William Raymond Manchester. The arms of Krupp, 1587–1968. Michael Joseph, London, UK, 1964. ISBN 0-7181-0553-2. 1053 pp. LCCN HD9523.9.

Manchester:1968:AKC

[Man68a] William Raymond Manchester. The arms of Krupp, 1587–1968. Little, Brown and Co., Boston, MA, USA, 1968. xvi + 976 pp. LCCN HD9523.9.K7 M35.

Manchester:1968:KCF

[Man68b] William Raymond Manchester. Krupp: Chronik einer Familie. (German) [Krupp: The Chronicles of a Family]. Heyne, München, Germany, 1968. ISBN 3-453-55045-5. 907 + 8 pp. LCCN HD9523.9.K7 M35.

Manchester:1970:AK

[Man70] William Raymond Manchester. The arms of Krupp, 1587–1968. Bantam Books, New York, NY, USA, 1970. ISBN 0-553-13149-4, 0-553-25992-X (paperback). 1068 (est.) pp. LCCN HD9523.9.

Mann:1980:PSR

[Man80] Robert Mann. On playing with scientists: remarks at the Einstein centennial celebration concert by the Juilliard quartet. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Manger:1986:SIT

[Man86] J. Manger. Study in indifference — on the trans-Atlantic contrasts in the reception accorded to a biography of Albert Einstein. Nature, 324(6093):169, November 13, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Manchester:2003:AK

[Man03] William Raymond Manchester. The arms of Krupp, 1587–1968. Back Bay Books, Boston, US, USA, 2003. ISBN 0-316-54490-6 (hardcover); 0-316-52940-0 (paperback). xvi + 976 pp. LCCN HD9523.9.K7 M35 2003.

Mansfield:2013:LRS

[Man13] P. Mansfield. The long road to Stockholm: the story of magnetic resonance imaging (MRI): an autobiography. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966454-4. x + 241 pp. LCCN QC16.M346 A3 2013;

Q143.M2265 A3 2013. URL http://www.loc.gov/catdir/
enhancements/fy1403/2012946411-b.html;   http://www.
loc.gov/catdir/enhancements/fy1403/2012946411-d.html;
 http://www.loc.gov/catdir/enhancements/fy1403/2012946411-
t.html.

MED:1945:ABH

[Man5 a] Manhattan Engineer District. The Atomic Bombings of Hiroshima and Nagasaki. ????, ????, 1945 (??). ii + 45 pp. LCCN ????

MED:1945:PAB

[Man5 b] Manhattan Engineer District. Photographs of the Atomic Bombings of Hiroshima and Nagasaki. ????, ????, 1945 (??). 102 pp. LCCN ????

Maor:2007:PTY

[Mao07] Eli Maor. The Pythagorean Theorem: a 4,000-year History. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12526-0. xvi + 259 pp. LCCN

QA460.P8 M36 2007. URL http://www.loc.gov/catdir/
enhancements/fy0704/2006050969-d.html;   http://www.
loc.gov/catdir/enhancements/fy0726/2006050969-t.html;
 http://www.loc.gov/catdir/enhancements/fy0734/2006050969-
b.html.

Mardles:1940:VSE

[Mar40] E. W. J. Mardles. Viscosity of suspensions and the Einstein equation. Nature, 145(3686):970, June 22, 1940. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v145/n3686/
pdf/145970a0.pdf.

Margenau:1949:ECR

[Mar49] Henry Margenau. Einstein’s conception of reality. In Schilpp [Sch49b], chapter 8, pages 243–268. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Marianoff:1951:EI

[Mar51] Dimitri Marianoff. Einstein dans l’intimité. (French) [Einstein: an intimate study of a great man]. Jeheber, Genève, Switzerland, 1951. 304 pp. LCCN QC16.E5 M314.

Margolin:1973:BRB

[Mar73] Jean-Claude Margolin. Book review: Correspondance 1916-1955 by Albert Einstein, Max Born, Bertrand Russell, Werner Heisenberg, and Pierre Leccia. Les Études philosophiques, 1 (??):75, January/March 1973. CODEN ???? ISSN 0014-2166 (print), 2101-0056 (electronic). URL http://www.jstor.

org/stable/20846392;   http://www.jstor.org/stable/
pdfplus/20846392.pdf.

Markov:1979:RSS

[Mar79] M. A. Markov. Reflections of a Soviet scientist on Einstein. Bulletin of the Atomic Scientists, 35(3):27–33, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Marshall:1984:EW

[Mar84] Trevor W. Marshall. From Einstein to where? Nature, 308 (5960):669, April 12, 1984. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v308/n5960/pdf/308669a0.pdf.

Marshak:1988:PHB

[Mar88] Robert E. Marshak. The pragmatic humanism of Bohr, Einstein and Sakharov. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 132(3): 268–275, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/3143854.

Maroist:1991:NCP

[Mar91] Richard Maroist. Nobel, la colombe de la paix mondiale et la petite Jessica. (French) [Nobel, the dove of world peace and little Jessica]. Roman jeunesse. Éditions R. Marois, ????, QC, Canada, 1991. ISBN 2-9802518-0-1. 86 pp. LCCN ????

Margaritondo:1995:RCE

[Mar95a] Giorgio Margaritondo. Röntgen, Curie and Einstein: the beginning of trouble for physics? Physics World, 8(9):15–16, September 1995. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/8/9/phwv8i9a11.pdf.

Markovic:1995:SSN

[Mar95b] Živko Marković. Sećanje starih novosadana na ajnštajnove (Serbian) [Recollections of old Novi Sad residents about the Einsteins]. In Rastko Maglić, editor, Dorpinos Mileve Ajinštajn-Marić nauci [Contributions of Mileva Einstein-Marić to Science], page ???? MST Gajic, Belgrade, Serbia, 1995. ISBN ????

Mariscotti:1996:SAH

[Mar96a] Mario Mariscotti. El Secreto Atómico de Huemul. (Spanish) [The Atomic Secret of Huemul]. Estudio Sigma, Buenos Aires, Argentina, third edition, 1996. ISBN 950-9446-24-6. 286 pp. LCCN QC773.3.A7 M37 1996.

Martin:1996:PLA

[Mar96b] Steve Martin. Picasso at the Lapin Agile and other plays. Grove Press, New York, NY, USA, 1996. ISBN 0-8021-1595-0. 150 pp. LCCN PS3563.A7293 P53 1996.

Marmet:1997:ETR

[Mar97] Paul Marmet. Einstein’s Theory of Relativity versus classical mechanics. Newton Physics Books, Gloucester, ON, Canada, 1997. ISBN 0-921272-18-9. 200 (est.) pp. LCCN QC6 .M356 1997.

Martin:2000:QSS

[Mar00] Craig Martin. Quads, shoeboxes, and sunken living rooms: a history of Los Alamos housing, volume Monograph 4 of The Los Alamos story. Los Alamos Historical Society, Los Alamos, NM, USA, 2000. ISBN 0-941232-24-7 (paperback). 149 pp. LCCN NA7235.N62 L676 2000.

Margaritondo:2001:TTA

[Mar01] Giorgio Margaritondo. Telling the truth about Einstein. Physics World, 14(4):17–18, April 2001. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
14/4/phwv14i4a19.pdf.

Martinez:2003:RRR

[Mar03] Alberto A. Martínez. Railways and the roots of relativity. Physics World, 16(11):51, November 2003. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/

16/11/phwv16i11a40.pdf. Review of Einstein’s Clocks, Poincaré’s Maps: Empires of Time, by Peter Galison, 2003 Sceptre / Norton, 384pp, £16.99 / $23.95hb.

Martinez:2004:AAE

[Mar04a] Alberto A. Martínez. Arguing about Einstein’s wife. Physics World, 17(4):14, April 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/17/4/phwv17i4a22.
pdf.

Martinez:2004:EW

[Mar04b] Alberto A. Martínez. Einstein’s wife. Physics World, 17(9):20, September 2004. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/17/9/phwv17i9a30.pdf.

Martinez:2004:MHI

[Mar04c]  Alberto A. Martínez.  Material history and imaginary

clocks: Poincaré, Einstein, and Galison on simultaneity. Physics in Perspective (PIP), 6(2):224–240, June 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic).  URL http://www.springerlink.com/content/
qd3k029w8d2lfp1m/fulltext.pdf.

Martinez:2004:REE

[Mar04d] Alberto A. Martínez. Ritz, Einstein, and the Emission Hypothesis. Physics in Perspective (PIP), 6(1):4–28, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://www.springerlink.com/content/
jq5dekc9hntg2qte/.

Martinez:2005:HEH

[Mar05] Alberto A. Martínez. Handling evidence in history: The case of Einstein’s wife. School Science Review, 86(316):49–56, March 2005. ISSN 0036-6811. URL https://www.ase.org.uk/

system/files/journal-issue/.

Marris:2006:TTE

[Mar06] Emma Marris. Testing times for Einstein’s theory. Nature News, February 21, 2006. CODEN ???? ISSN ???? URL http://www.

nature.com/news/2006/060221/full/news060220-7.html.

Martinez:2007:MNW

[Mar07] Alberto A. Martínez. Martínez writings: Einstein’s wife. Web

site, 2007. URL http://martinezwritings.com/m/Maric.
html..

Marsh:2008:CGE

[Mar08] Gerald E. Marsh. Charge, geometry, and effective mass in the Kerr–Newman solution to the Einstein field equations. Foundations of Physics, 38(10):959–968, October 2008. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
008-9245-x.

Martinez:2009:KLO

[Mar09] Alberto A. Martínez. Kinematics: the Lost Origins of Einstein’s Relativity. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9135-3 (hardcover). xix + 464 pp. LCCN QA841 .M37 2009.

Martinez:2010:BRB

[Mar10] Alberto A. Martínez. Book review: Einstein’s Generation: The Origins of the Relativity Revolution, Richard Staley. University of Chicago Press, Chicago (2009), pp. x + 494. US $38.00 PB, ISBN-13 978-0-226-77057-4. Studies in History and Philosophy of Modern Physics, 41(4):366–367, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000365. See [Sta08a].

Marriott:2011:MMN

[Mar11a] Barbara Marriott. Myths and mysteries of New Mexico: true stories of the unsolved and unexplained. Myths and mysteries series. Globe Pequot Press, Guilford, CT, USA, 2011. ISBN 0-7627-5873-2, 0-7627-6745-6 (e-book). x + 188 pp. LCCN F796.6 .M37 2011.

Martinez:2011:BRC

[Mar11b] Alberto A. Martinez. Book review: The Collected Papers of Albert Einstein. Annals of Science, 68(2):267–280, ???? 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Martinez:2011:ERM

[Mar11c] Alberto A. Martínez. Essay review: The myriad pieces of Einstein’s remains: John Stachel, Robert Schulmann, Diana Kormos Buchwald, et al, editors, The Collected Papers of Albert Einstein. Annals of Science, 68(2):267–280, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinez:2011:SST

[Mar11d] Alberto Antonio Martínez. Science secrets: the truth about Darwin’s finches, Einstein’s wife, and other myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2011. ISBN 0-8229-4407-3 (hardcover). xviii + 324 pp. LCCN Q173 .M316 2011.

Marcus:2015:SSE

[Mar15] Solomon Marcus. Starting from the scenario Euclid–Bolyai– Einstein. Synthese, 192(7):2139–2149, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/s11229-
013-0346-5;   http://link.springer.com/content/pdf/
10.1007/s11229-013-0346-5.pdf.

Marquardt:2017:EWG

[Mar17] Siegfried Marquardt. Einstein war der größte Plagiator in der modernen Wissenschaftsgeschichte: ein Wissenschaftskrimi ohne Leichen. (German) [Einstein was the greatest plagiarizer in the modern history of science: a science fiction without corpses]. Wissenschaften, Königs Wusterhausen, Germany, 2017. 68 pp.

Marsh:2022:PFN

[Mar22] Allison Marsh. Past forward: The Nobelist’s first laser. IEEE Spectrum, 59(3):68, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mashhoon:1974:CET

[Mas74] Bahram Mashhoon. Can Einstein’s theory of gravitation be tested beyond the geometrical optics limit? Nature, 250(5464): 316–317, July 26, 1974. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v250/n5464/pdf/250316a0.pdf.

Mason:1995:CAO

[Mas95] Katrina R. Mason. Children of Los Alamos: An Oral History of the Town Where the Atomic Age Began, volume 19 of Twayne’s oral history series. Twayne Publishers, New York, NY, USA, 1995. ISBN 0-8057-9138-8 (hardcover), 0-8057-9139-6 (paperback). xiv + 204 + 8 pp. LCCN QC773.3.U5 M29 1995.

Mastrobisi:1999:ANC

[Mas99] Giorgio Jules Mastrobisi. Aporie di una nuova concezione dell’etere nel manoscritto einsteiniano del 1920. (Italian) [History of a new concept in ether. The 1920 Einsteinian manuscript]. Nuncius, 14(2):629–649, ???? 1999. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00094.

Mastrobisi:2001:BRM

[Mas01] Giorgio Jules Mastrobisi. Book review: Max Jammer, Einstein and Religion: Physics and Theology, Princeton, Princeton University Press, 1999, pp. 279 (ISBN 0-691-00699-7). Nuncius, 16 (2):850–854, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00866.

Mastrobisi:2002:SAE

[Mas02] Giorgio Jules Mastrobisi. Il ≪manoscritto di Singapore≫ (1923) di Albert Einstein. Per una teoria del ≪campo unificato≫ tra possibilità fisica e necessità matematica (Italian) [The ≪Singapore manuscript≫ (1923) of Albert Einstein. On a theory of the ≪unified field≫ between physical possibility and mathematical need]. Nuncius, 17(1):269–305, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00694.

Massey:2005:AEA

[Mas05a] Robert U. Massey. Albert Einstein’s Annus Mirabilis. Connecticut Medicine, 69(5):307–308, May 2005. CODEN CNMEAH. ISSN 0010-6178.

Massimi:2005:PEP

[Mas05b] Michela Massimi. Pauli’s exclusion principle: the origin and validation of a scientific principle. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83911-4 (hardcover). xiv + 211 pp. LCCN QC174.17.P3 M37 2005. URL http://www.loc.

gov/catdir/enhancements/fy0632/2005296620-d.html;
http://www.loc.gov/catdir/enhancements/fy0632/2005296620-
t.html;   http://www.loc.gov/catdir/enhancements/fy0733/
2005296620-b.html.

Maskawa:2009:NLW

[Mas09] Toshihide Maskawa. Nobel Lecture: What does CP violation tell us? Reviews of Modern Physics, 81(3):1027– 1030, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1027;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1027_1.

Matukuma:1940:DEE

[Mat40] T. Matukuma. Derivation of Einstein effect from eclipse observations of 1936. Nature, 146(3695):264–265, August 24, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v146/n3695/pdf/146264b0.pdf.

Mathieu:1986:AOBd

[Mat86a] J.-P. Mathieu. Analyses d’ouvrages: Galilée, Newton lus par Einstein. Espace et relativité, (Coll. ^{≪}Philosophies ^{≫}) par Françoise Balibar. Revue d’Histoire des Sciences, 39(2):183, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23632723.

Mathieu:1986:AOBc

[Mat86b] J.-P. Mathieu. Analyses d’ouvrages: The beginnings of the Nobel institution. The Science Prizes, 1901–1915 par Elisabeth Crawford. Revue d’Histoire des Sciences, 39(1):88, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632585.

Matzner:1991:NSR

[Mat91] Richard Matzner. Naked singularities raise questions over general relativity. Physics World, 4(6):25–26, June 1991. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
4/6/phwv4i6a22.pdf.

Mather:2007:NLB

[Mat07] John C. Mather. Nobel Lecture: From the Big Bang to the Nobel Prize and beyond. Reviews of Modern Physics, 79(4): 1331–1348, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1331;   http:
//rmp.aps.org/abstract/RMP/v79/i4/p1331_1.

Mattsson:2010:DEM

[Mat10] Teppo Mattsson. Dark energy as a mirage. General Relativity and Gravitation, 42(3):567–599, ???? 2010. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL http://

www.springerlink.com/content/8287361563158411/.

Maudlin:1990:SST

[Mau90] Tim Maudlin. Substances and space–time: What Aristotle would have said to Einstein. Studies in History and Philosophy of Science Part A, 21(4):531–561, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900324.

Maurer:1996:WSK

[Mau96] Margarete Maurer. “Weil nicht sein kann was nicht sein darf … ‘Die Eltern’ oder ‘der Vater’ der Relativitätstheorie?” German [Because one can not be what one can not be … the ‘parents’ or ‘the father’ of the theory of relativity?]. PC News (Vienna) 48, 11(3):20–27, June 1996.

Mauskopf:2009:BRU

[Mau09] Seymour Mauskopf. Book review: Ulf Larsson, Alfred Nobel. Networks of Innovation. Sagamore Beach: Science History Publications, 2008. 216 pp., ISBN 978-0-88135-399-0. Nuncius, 24 (1):242–243, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00471.

Maxwell:1993:ISRa

[Max93a] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. part one: How to solve the problem of induction. The British Journal for the Philosophy of Science, 44(1):61–79, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL

http://bjps.oxfordjournals.org/content/44/1/61.full.
pdf+html;   http://www.jstor.org/stable/687850.

Maxwell:1993:ISRc

[Max93b] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. part three: Einstein, aim-oriented empiricism and the discovery of Special and General Relativity. The British Journal for the Philosophy of Science, 44(2):275–305, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/44/2/275.full.pdf+html;
http://www.jstor.org/stable/687649.

Maxwell:1993:ISRb

[Max93c] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. Part two: Aim-oriented empiricism and scientific essentialism. The British Journal for the Philosophy of Science, 44(1):81–101, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL

http://bjps.oxfordjournals.org/content/44/1/81.full.
pdf+html;   http://www.jstor.org/stable/687851.

May:1995:FSR

[May95] Michael M. May. Fearsome security: The role of nuclear weapons. Brookings Review, 13(3):24–27, ???? 1995.

Moszkowski:1921:ESH

[MB21] Alexander Moszkowski and Henry L. (Henry Leopold) Brose. Einstein, the searcher: his work explained from dialogues with Einstein. Methuen and Co. Ltd., London, UK, 1921. xi + 246 pp. LCCN QC6 .M85eE; QC .M6; QC6 .M916e 1921.

Marshak:1966:PMP

[MB66] R. E. Marshak and J. Warren Blaker, editors. Perspectives in modern physics. Essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July, 1966. Interscience Publishers, New York, NY, USA, 1966. xii + 673 pp. LCCN QC774.B4 M3.

McGuire:1992:NPW

[MB^{+}92] Paula McGuire, Gert H. Brieger, et al., editors. Nobel Prize winners. Supplement, 1987–1991: an H. W. Wilson biographical dictionary. H. W. Wilson Co., New York, NY, USA, 1992. ISBN 0-8242-0834-X. 143 pp. LCCN AS911.N9 N59 1987 Suppl.

McGrayne:1994:NPW

[MB94] Sharon Bertsch McGrayne and Nina Byers. Nobel Prize women in science. Physics Today, 47(7):63, July 1994. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Montwill:2011:QAD

[MB11] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (ebook). ix + 248 pp. LCCN QC174.12 .M66 2012.

Michels:1939:EEE

[MBD39] A. Michels, A. Bijl, and J. De Boer. Effect of an excitation energy on the specific heat of liquid helium II and its relation to the exchange effect in a non-ideal Bose–Einstein gas. Nature, 144(3648):594–595, September 30, 1939. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v144/n3648/
pdf/144594b0.pdf.

Mehlberg:1980:LAC

[MC80] Henry Mehlberg and Robert S. Cohen. Lechalas’ adaptation of the causal theory of time to the laws of pre-Einsteinian physics. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 4, pages 70–90. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-009-8935-1_4.

McClintock:1948:NPT

[McC48] Marshall McClintock. The Nobel Prize treasury. Doubleday, Garden City, NY, USA, 1948. xvi + 612 pp. LCCN ????

Mccrea:1955:PAE

[Mcc55] W. H. Mccrea. Prof. Albert Einstein, For.Mem.R.S. Nature, 175 (4465):925–926, May 28, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v175/n4465/pdf/175925a0.pdf.

McCrea:1957:OEW

[McC57] W. H. McCrea. On the objective of Einstein’s work. The British Journal for the Philosophy of Science, VIII(29):18–29, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/VIII/29/18.full.pdf+html.

McCrea:1968:CAH

[McC68] W. H. McCrea. Cosmology after half a century: Fifty years after Einstein’s paper of 1917 cosmology is in a supremely interesting state. Science, 160(3834):1295–1299, June 21, 1968. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/160/3834/1295.
full.pdf.

McCormmach:1970:ELE

[McC70] Russell McCormmach. Einstein, Lorentz, and the electron theory. Historical Studies in the Physical Sciences, 2(??):41–87, ???? 1970. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757304.

McCausland:1977:DAU

[McC77] Ian McCausland. The Dingle Affair: An unresolved scientific controversy. Web document, June 8, 1977. URL http:/

/www.gsjournal.net/Science-Journals/Essays/View/2567.

Refutation of the anti-relativity views of [Din72].

Mccrea:1978:AEP

[Mcc78] W. H. Mccrea. Albert Einstein physicist (1879–1955). Interdisciplinary Science Reviews, 3(4):275–293, ???? 1978. ISSN 0308-0188.

McCausland:1984:ESR

[McC84] Ian McCausland. Einstein and Special Relativity: Who wrote the added footnotes? The British Journal for the Philosophy of Science, 35(1):60–61, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/35/1/60.full.pdf+html.

Mccormmach:1987:BRC

[Mcc87] R. Mccormmach. Book review: The Collected Papers of Albert Einstein, Vol 1, the Early Years, 1879–1902, A. Beck, Translator. New York Times Book Review, ??(??):33, September 27, 1987. ISSN 0028-7806.

McCrea:1991:ASE

[McC91] Sir William McCrea. Arthur Stanley Eddington. Scientific American, 264(6):92–?? (Intl. ed. 66–??), June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McCarty:2001:NLH

[McC01a] Marilu Hurt McCarty. The Nobel laureates: how the world’s greatest economic minds shaped modern thought. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-135614-2. xvi + 397 pp. LCCN HB87 .M337 2001.

McCook:2001:NSN

[McC01b] Alison McCook. News scan: The Nobel Prizes for 2001. Scientific American, 285(6):29, December 2001. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/2001/1201issue/1201inbrief.html.

McCook:2001:NP

[McC01c] Alison McCook. The Nobel Prizes for 2001. Scientific American, 285(6):29, December 2001. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v285/n6/
pdf/scientificamerican1201-29a.pdf.

McCall:2011:CMN

[McC11] Martin W. McCall. Classical mechanics: from Newton to Einstein: a modern introduction. Wiley, New York, NY, USA, second edition, 2011. ISBN 0-470-71574-X (hardcover), 0-470-71572-3 (paperback). xiv + 235 pp. LCCN QC125.2. UK£100.00; UK£37.50.

McCormick:2015:AE

[McC15] Lisa Wade McCormick. Albert Einstein. Great science writers. Rosen Publishing, New York, NY, USA, 2015. ISBN 1-4777-7687-7 (hardcover). 112 pp. LCCN QC16.E5 M3514 2015.

McGrayne:1993:NPW

[McG93] Sharon Bertsch McGrayne. Nobel Prize women in science: their lives, struggles, and momentous discoveries. Carol Publishing Group, Secaucus, NJ, USA, 1993. ISBN 1-55972-146-4. xi + 419 pp. LCCN Q141 .B42 1993.

McGrayne:1998:NPW

[McG98] Sharon Bertsch McGrayne. Nobel Prize women in science: their lives, struggles, and momentous discoveries. Joseph Henry Press, Washington, DC, USA, second edition, 1998. ISBN 0-309-07270-0 (paperback), 0-8065-2025-6 (paperback), 0-9702256-0-1. xi + 451 pp. LCCN Q141 .M358 1998. URL http://www.

nap.edu/catalog/10016.html.

McGrath:2019:TEM

[McG19] Alister McGrath. Theory of Everything (That Matters): a Brief Guide to Einstein, Relativity, and His Surprising Thoughts on God. Tyndale House Publishers, Carol Stream, IL, USA, 2019. ISBN 1-4964-3807-8. xiii + 217 pp. LCCN QC16.E5 M3625 2019.

McHarris:2007:CMQ

[McH07] Wm. C. McHarris. Chaos meets quantum mechanics: Possible nonlinear vindication of Einstein’s arguments: Paradoxes of the Copenhagen interpretation: Nonlinear parallels. Complexity, 12 (4):12–18, March /April 2007. CODEN COMPFS. ISSN 1076-2787 (print), 1099-0526 (electronic). URL https://dblp.org/

db/journals/complexity/complexity12.html#McHarris07.

McKay:1984:MAA

[McK84] H. A. C. (Herbert Alwyn Cochrane) McKay. The making of the atomic age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-289174-X (paperback), 0-19-219193-4 (hardcover). xii + 153 + 8 pp. LCCN QC773 .M35 1984.

McKain:2003:MUA

[McK03a] Mark McKain, editor. Making and using the atomic bomb. History firsthand. Greenhaven Press / Thomson, San Diego, CA, USA, 2003. ISBN 0-7377-1413-1 (paperback), 0-7377-1412-3. 240 pp. LCCN QC773.3.U5 M23 2003.

McKeown:2003:IFU

[McK03b] William (William Thomas) McKeown. Idaho Falls: the untold story of America’s first nuclear accident. ECW Press, Toronto, ON, Canada, 2003. ISBN 1-55022-562-6, 1-55490-562-1 (e-book), 1-55490-543-5 (e-book). 269 pp.

LCCN TK1345.I2 M38 2003. URL http://catdir.loc.gov/

catdir/bios/ipg051/2003467097.html;   http://catdir.
loc.gov/catdir/description/ipg051/2003467097.html;
http://catdir.loc.gov/catdir/enhancements/fy1204/2003467097-
s.html.

McLay:1999:LME

[McL99] David B. McLay. Lise Meitner and Erwin Schrödinger: Biographies of two Austrian physicists of Nobel stature. Minerva [Washington, DC], 37(1):75–94, ???? 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004536223625.

McMillan:1951:TEE

[McM51] Edwin M. McMillan. The transuranium elements: Early history. Report UCRL-1619, Radiation Laboratory, University of California, Berkeley, Berkeley, CA, USA, Decem-

ber 1951. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0361.pdf. Nobel Lecture given at

Stockholm on December 12, 1951.

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

McManus:2014:EBP

[McM14] Joanna McManus. Einstein in Britain: a portrait. Notes and Records of the Royal Society of London, 68(3):311–315, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
43287734.

McPhee:1974:CBE

[McP74] John A. McPhee. The curve of binding energy. Farrar, Strauss, and Giroux, New York, NY, USA, 1974. ISBN 0-374-13373-5. 232 pp. LCCN UF767 .M215 1974. URL http://www.loc.

gov/catdir/bios/hol059/74001226.html;   http://www.
loc.gov/catdir/description/hol054/74001226.html.

McPherson:1995:OGS

[McP95] Stephanie Sammartino McPherson. Ordinary genius: the story of Albert Einstein. Carolrhoda Books, Minneapolis, MN, USA, 1995. ISBN 0-87614-788-0. 95 pp. LCCN QC16.E5 M38 1995.

McPherson:1997:OGS

[McP97] Stephanie Sammartino McPherson. Ordinary Genius: The Story of Albert Einstein. Twenty-First Century Books, Minneapolis, 1997. ISBN 1-57505-067-6. 95 (est.) pp. LCCN ????

Mota:2009:EPE

[MCS09] Elsa Mota, Paulo Crawford, and Ana Simões. Einstein in Portugal: Eddington’s expedition to Principe and the reactions of Portuguese astronomers (1917–1925). British Journal for the History of Science, 42(2):245–273, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mcvittie:1948:E

[Mcv48] G. C. Mcvittie. Einstein. Nature, 162(4120):591, October 16, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v162/n4120/pdf/162591a0.pdf.

Miller:1947:BEB

[MCW47] Leslie Miller, Robert Considine, and Frank Wead. The Beginning or the End: Book of the Film. ????, ????, 1947. LCCN ????

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http:/

/library.ucsd.edu/dc/object/bb0103915g. This is a reason-

ably accurate 83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Mermin:2015:ZER

[MD15] David N. Mermin and Matthias Delbrück. Es ist an der Zeit Einsteins Relativitätstheorie verstehen. (German) [It is time to understand Einstein’s relativity theory]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-662-47151-5. 348 pp. LCCN ????

Mackay:2003:NEL

[MDT^{+}03] Michael E. Mackay, Tien T. Dao, Anish Tuteja, Derek L. Ho, Brooke Van Horn, Ho-Cheol Kim, and Craig J. Hawker. Nanoscale effects leading to non-Einstein-like decrease in viscosity. Nature Materials, 2(11):762–766, October 19, 2003. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (elec-

tronic). URL http://www.nature.com/nmat/journal/v2/
n11/full/nmat999.html.

Mamedov:2014:PNE

[ME14a] B. A. Mamedov and M. Y. Esmer. On the philosophical nature of Einstein’s mass–energy equivalence formula E = mc^{2}. Foundations of Science, 19(4):319–329, November 2014. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic).

URL http://link.springer.com/article/10.1007/s10699-
013-9339-6.

Meltzer:2014:AEd

[ME14b] Brad Meltzer and Chris Eliopoulos. I am Albert Einstein. Ordinary people change the world. Dial Books for Young Readers, an imprint of Penguin Group (USA) LLC, New York, NY, USA, 2014. ISBN 0-8037-4084-0 (hardcover). 39 pp. LCCN QC16.E5 M448 2014.

Meadows:1973:E

[Mea73] A. J. Meadows. Einstein. Nature, 246(5433):431–432, December 14, 1973. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v246/n5433/pdf/246431a0.pdf.

Mealli:2021:ACQ

[Mea21] Fabrizia Mealli. Answering causal questions: Angrist, Imbens and the Nobel prize. Significance (Oxford, England), 18(6):4–5, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Medicus:1976:BRB

[Med76] Heinrich A. Medicus. Book review: Albert Einstein in Bern by Max Flückiger. American Journal of Physics, 44(9):897, September 1976. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?
AJP/44/897/1.

Medicus:1994:FAT

[Med94] Heinrich A. Medicus. The friendship among three singular men: Einstein and his Swiss friends Besso and Zangger. Isis, 85(3): 456–478, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235463.

Meeropol:2003:EFO

[Mee03] Robert Meeropol. An execution in the family: one son’s journey. St. Martin’s Press, New York, NY, USA, 2003. ISBN 0-312-30636-9, 0-312-30637-7 (paperback). xii + 273 pp. LCCN HV28.M365 A3 2003. URL http://catdir.loc.gov/

catdir/bios/hol052/2003043117.html;   http://catdir.
loc.gov/catdir/description/hol032/2003043117.html;
http://catdir.loc.gov/catdir/samples/hol051/2003043117.
html.

Mehra:1973:EHT

[Meh73a] Jagdish Mehra. Einstein, Hilbert, and the theory of gravitation. In The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972 [Meh73b], chapter 7, pages 92–178. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Mehra:1973:PCN

[Meh73b] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Mehra:1975:EFS

[Meh75a] Jagdish Mehra. Einstein and the foundation of statistical mechanics. Physica A, 79(5):447–477, March 1975. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Mehra:1975:SB

[Meh75b] Jagdish Mehra. Satyendra Bose. Biographical Memoirs of Fellows of the Royal Society, 7(??):117–138, 1975. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mehra:1975:SCP

[Meh75c] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mehra:1987:NBD

[Meh87] Jagdish Mehra. Niels Bohr discussions with Albert Einstein, Werner Heisenberg, and Erwin Schrödinger — the origins of the Principles of Uncertainty and Complementarity. Foundations of Physics, 17(5):461–506, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01559698.

Mehra:1998:OMH

[Meh98] Jagdish Mehra. One month in the history of the discovery of general relativity theory. Foundations of Physics Letters, 11 (1):41–60, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Meitner:1939:NPF

[Mei39] Lise Meitner. New products of the fission of the thorium nucleus. Nature, 143(3624):637, April 15, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v143/n3624/
pdf/143637a0.pdf.

Meier:1954:ANN

[Mei54] Ernst Meier. Alfred Nobel; Nobelstiftung; Nobelpreise. ????, Berlin, 1954. 300 pp. LCCN ????

Meitner:1962:RWR

[Mei62] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. International Atomic Energy Agency Bulletin, 4(0):4–6, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004790608su.pdf.

Meitner:1963:RWR

[Mei63a] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. Advancement of Science, 19(??):363–365, ???? 1963. CODEN ADSCAH. ISSN 0001-866X. English translation of [Mei63b].

Meitner:1963:WIK

[Mei63b] Lise Meitner. Wege und Irrwege zur Kernenergie. (German) [right and wrong roads to the discovery of nuclear energy]. Naturwissenschaftliche Rundschau, 16(??):167–169, ???? 1963. CODEN NARSAC. ISSN 0028-1050. English translation in [Mei63a].

Meitner:1964:LMLa

[Mei64a] Lise Meitner. Lise Meitner looks back. Advancement of Science, 21(??):39–46, ???? 1964. CODEN ADSCAH. ISSN 0001-866X.

Meitner:1964:LMLb

[Mei64b] Lise Meitner. Lise Meitner looks back. International Atomic Energy Agency Bulletin, 6(1):4–12, ???? 1964. CODEN IAE-BAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL

http://www.iaea.org/Publications/Magazines/Bulletin/
Bull061/06101400412.pdf.

Meitner:2005:LME

[Mei05] Lise Meitner. Lise Meitner: Erinnerungen an Otto Hahn; mit Beiträgen von Mitarbeitern und Weggefährten. (German) [Lise Meitner: Memories of Otto Hahn; with contributions from employees and companions]. S. Hirzel Verlag, Stuttgart, Germany, 2005. ISBN 3-7776-1380-0 (paperback). 167 pp. LCCN

???? URL http://catalog.hathitrust.org/api/volumes/

oclc/71048128.html. Edited by Dietrich Hahn.

Meilinger:2012:EAA

[Mei12] Phillip S. Meilinger. Early atomic air. Air Force magazine, 95 (3):74–77, March 2012. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.airforce-

magazine.com/MagazineArchive/Pages/2012/March%202012/
0312air.aspx.

Melehy:1963:FEE

[Mel63] M. A. Melehy. A fundamental extension of Einstein’s diffusion — mobility relationship. Nature, 198(4884):980–981, June 8, 1963. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v198/n4884/pdf/198980b0.pdf.

Melcher:1978:AEP

[Mel78] Horst Melcher. Albert Einsteins Patente. (German) [Albert Einstein’s patents]. Spektrum, 9(??):23 26, ???? 1978.

Melcher:1979:AEE

[Mel79a] Horst Melcher. Albert Einstein und die experimentelle Physik. (German) [Albert Einstein and experimental physics]. Physik in der Schule, 17(??):1–20, ???? 1979.

Melcher:1979:AEW

[Mel79b] Horst Melcher. Albert Einstein wider Vorurteile und Denkgewohnheiten. (German) [Albert Einstein against prejudices and habits of thought]. Reihe Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1979. ISBN 3-528-06849-3. 107 pp. LCCN QC16.E5 M44. DM14.80.

Melcher:1982:BRA

[Mel82] Horst Melcher. Book review: Arthur I. Miller: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–11). Isis, 73(3):483–484, September 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231513.

Melzer:2000:BHS

[Mel00] Richard Melzer. Breakdown: how the secret of the atomic bomb was stolen during World War II. Sunstone Press, Santa Fe, NM, 2000. ISBN 0-86534-304-7. 160 pp. LCCN QC789.2.U62 M45 2000.

Melia:2009:CEC

[Mel09] Fulvio Melia. Cracking the Einstein code: Relativity and the birth of black hole physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-51951-1. xi + 137 pp. LCCN QC173.6 .M434 2009. With an afterword by Roy Kerr.

Menges:1926:ETR

[Men26] Charles L. R. E. Menges. On Einstein’s Theory of Relativity. Science, 63(1634):427–428, April 23, 1926. CODEN SCIEAS.

ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/63/1634/427.full.pdf.

Menger:1949:TRG

[Men49] Karl Menger. The theory of relativity and geometry. In Schilpp [Sch49c], chapter 17, pages 457–474. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Merton:1938:SSO

[Mer38] Robert K. Merton. Science and the social order. Philosophy of Science, 5(3):321–337, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/184838.

Mercier:1955:FYT

[Mer55] André Mercier. Fifty years of the theory of Relativity. Nature, 175(4465):919–921, May 28, 1955. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v175/n4465/pdf/
175919a0.pdf.

Mercier:1979:BRG

[Mer79] André Mercier. Birth and rôle of the GRG-organization and the cultivation of international relations among scientists in the field. In Aichelburg and Sexl [AS79], pages 177–188. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_13/.

Mermin:1985:BBP

[Mer85a] N. David Mermin. A bolt from the blue: the E-P-R paradox. In French and Kennedy [FK85], pages 141–147. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Mermin:1985:MTW

[Mer85b] N. David Mermin. Is the Moon there when nobody looks? Reality and the quantum theory. Physics Today, 38(4):38–47, April 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Merlan:1997:TE

[Mer97] Thomas Merlan. The Trinity experiments, volume 9701 of Human systems research report. Human Systems Research, Tularosa, NM, USA, 1997. ISBN 1-887523-15-4 (paperback). ???? pp. LCCN 9709 BOOK NOT YET IN LC.

Merlan:2001:LTB

[Mer01] Thomas Merlan. Life at Trinity base camp. Report HSR 9831, Human Systems Research, Tularosa, NM, USA, 2001. Also issued as White Sands Missile Range Archaeological Research Report No. 01-07.

Mermin:2005:ATU

[Mer05] N. David Mermin. It’s about time: understanding Einstein’s Relativity. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12201-6 (clothbound). xv + 192 pp. LCCN QC173.65 .M47 2005. URL http://people.ccmr.

cornell.edu/~mermin/about-time/errata.html;   http:/
/www.pupress.princeton.edu/titles/8112.html.

Merali:2010:EPC

[Mer10] Zeeya Merali. Einstein passes cosmic test. Nature News, March 10, 2010. CODEN ???? ISSN ???? URL http://www.nature.

com/news/2010/100310/full/news.2010.113.html.

Mermin:2011:UED

[Mer11] N. David Mermin. Understanding Einstein’s 1905 derivation of E = Mc^{2}. Studies in History and Philosophy of Modern Physics, 42(1):1–2, February 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000754. See comment [Oha12b] and

reply [Mer12].

Mermin:2012:ROC

[Mer12] N. David Mermin. Reply to Ohanian’s comment. Studies in History and Philosophy of Modern Physics, 43(3):218–219, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000044. See [Mer11, Oha12b].

Mermin:2016:WQR

[Mer16] N. David Mermin. Why quark rhymes with pork, and other scientific diversions. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-02430-7 (hardcover), 1-139-16257-8 (ebook), 1-316-47617-0 (e-book). ???? pp. LCCN QC71 .M373 2016.

Message:1983:MBE

[Mes83] P. J. Message. Mostly before Einstein: explanations for the anomalous orbit. Nature, 301(5899):445, February 3, 1983. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v301/n5899/
pdf/301445a0.pdf.

Metzler:1995:AEB

[Met95] G. Metzler. Albert Einstein — a biography — German — A. Fölsing. Historische Zeitschrift, 260(2):492–494, April 1995. ISSN 0018-2613.

Meyerson:1925:DR

[Mey25] Émile Meyerson. La déeduction relativiste. Payot, Paris, France, 1925. xvi + 396 pp. LCCN ????

Meyerson:1985:RDE

[Mey85] Émile Meyerson. The relativistic deduction: epistemological implications of the Theory of Relativity, volume 83 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1699-4. liii + 268 pp. LCCN QC173.585; Q174 .B67 vol. 83. With a review by Albert Einstein.

Meyer:1992:ME

[Mey92a] Stefan Meyer. Mach and Einstein. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92], chapter 7, pages 151–164. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN C5.53; QC6.9. URL https://link.

springer.com/chapter/10.1007/978-94-011-2771-4_7.

Meyerson:1992:DR

[Mey92b] Émile Meyerson. La déduction relativiste. J. Gabay, Sceaux, France, 1992. ISBN 2-87647-088-8. xvi + 396 pp. LCCN ????

Meyers:2007:HAS

[Mey07] Morton A. Meyers. Happy accidents: serendipity in modern medical breakthroughs. Arcade Pub., New York, NY, USA, 2007. ISBN 1-55970-819-0, 1-55970-845-X (paperback). xiv + 390 pp.

LCCN R149 .M49 2007. URL http://www.loc.gov/catdir/
toc/ecip076/2006100551.html.

Meitner:1939:DUN

[MF39a] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction. Nature, 143(3615): 239–240, February 11, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v143/n3615/pdf/143239a0.pdf. This

paper, and [MF39c], both submitted 16 January 1939 (see [Mei62]), provided the first published explanation of nuclear disintegration, called ‘nuclear fission’ by Frisch, that was first observed experimentally by Hahn and Strassmann in December 1939 [HS39c]. Frisch’s paper [Fri39b] describes the first experimental confirmation. It was these results that Niels Bohr intended to hold confidential until their journal publication during his January 1939 trip to the USA, but his traveling companion Léon Rosenfeld [Ros72] wasn’t informed of that intent, and the news escaped and spread quickly.

Meitner:1939:PFUa

[MF39b] Lise Meitner and Otto Robert Frisch. On the products of the fission of uranium and thorium under neutron bombardment. Math.-fys. Meddr, 17(5):1–13, 1939.

Meitner:1939:PFUb

[MF39c] Lise Meitner and Otto Robert Frisch. Products of the fission of the uranium nucleus. Nature, 143(3620):471–472, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v143/n3620/pdf/143471a0.pdf. See note in [MF39a].

Martinez-Frias:2006:RCA

[MFHR06] J. Martinez-Frias, D. Hochberg, and F. Rull. A review of the contributions of Albert Einstein to earth sciences — in commemoration of the World Year of Physics. Die Naturwissenschaften, 93(2):66–71, February 2006. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Mendelsson:2014:AEA

[MFO14] Dalia Mendelsson, Edith Falk, and Amalya L. Oliver. The Albert Einstein archives digitization project: opening hidden treasures. Library Hi Tech, 32(2):318–335, ???? 2014. CODEN ???? ISSN 0737-8831 (print), 2054-166X (electronic). URL https:/

/dblp.org/db/journals/lht/lht32.html#MendelssonFO14.

Marconi:196x:VPN

[MFS^{+}6x] Guglielmo Marconi, Enrico Fermi, Emilio Segrè, Daniele Bovet, Giulio Natta, Grazia Deledda, Luigi Pirandello, Salvatore Quadimodo, and Angelo Monteverdi, editors. Le Voci dei Premi Nobel italiani. (Italian). [The voices of Italian Nobel Prize winners]. Discoteca di Stato DSM 251, ????, 196x. Sound recording (33 1/3 rpm).

Murrell:2002:CEF

[MG02] John N. Murrell and Nicole Grobert. The centenary of Einstein’s first scientific paper. Notes and Records of the Royal Society of London, 56(1):89–94, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://

www.jstor.org/stable/532124.

Martienssen:1995:BAF

[MGG^{+}95] W. Martienssen, S. Großmann, P. Grassberger, D. Meissner, W. Sandhas, H. Nicolai, D. Hoffmann, H. Kant, P. Richter, H. Atmanspacher, M. Eckert, H. Rechenberg, H. Maier-Leibnitz, and D. Haarer. Buchbesprechungen: Argyris/Faust: Die Erforschung des Chaos/Plaschko/Brod: Nichtlineare Dynamik, Bifurkation und Chaotische Systeme/Bunde/Havlin: Fractals in Science/Goetzberger/Voß/Knobloch: Sonnenenergie: Photovoltaik/Lindner: Grundkurs Theoretische Physik/Vilenkin/Shellard: Cosmic Strings and other Topological Defects/Weiss: Großforschung in Deutschland Geschichte des Hahn-Meitner-Instituts/Fölsing: Wilhelm Conrad Röntgen Aufbruch ins Innere der Materie/Brandmüller: Galilei und die Kirche Ein Fall und seine Lösung/Meÿenn: Wolfgang Pauli Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. Band 3: 1940 49/Atmanspacher/Primas: Der Pauli-Jung-Dialog und seine Bedeutung für die moderne Wissenschaft/Cassidy: Werner Heisenberg Leben und Werk/Weinberg: The First Nuclear Era The Life and Times of a Technological Fixer/Wenske: Wörterbuch Chemie/Dictionary of Chemistry. Physikalische Blätter, 51(11):1101–1106, November 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19950511117/abstract.

Menke:1971:WWT

[MH71] H. Menke and G. Herrmann. Was waren die ‘Transurane’ der dreißiger Jahre in Wirklichkeit?. (German) [What were the ‘transuranics’ in the Thirties in reality?]. Radiochimica Acta, 16 (??):119–123, ???? 1971. CODEN RAACAP. ISSN 0033-8230.

Maas:2009:SRW

[MH09] Ad Maas and Hans Hooijmaijers. Scientific research in World War II: what scientists did in the war. Routledge studies in modern history. Routledge/Taylor and Francis, New York, NY, 2009. ISBN 0-203-88318-7 (e-book), 0-7103-1340-3 (hardcover). xii + 240 pp. LCCN Q141 .H195 2009. URL http://www.loc.

gov/catdir/toc/ecip0824/2008033118.html.

Margo:2012:AGA

[MH12] Curtis E. Margo and Lynn E. Harman. Allvar Gullstrand, Albert Einstein, and a Nobel dilemma revisited. The Pharos of Alpha Omega Alpha—Honor Medical Society, 75(2):14–19, ???? 2012. ISSN 0031-7179.

Moss:1990:LWF

[MHM90] Norman Moss, Daniel Hirsch, and William G. Mathews. Letters: What Fuchs didn’t know, and when he didn’t know it. Bulletin of the Atomic Scientists, 46(4):51–52, May 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Makhijani:1995:NWG

[MHY95] Arjun Makhijani, Howard Hu, and Katherine Yih, editors. Nuclear wastelands: a global guide to nuclear weapons production and its health and environmental effects. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-13307-5. xxiv + 666 pp. LCCN TD195.N85 N83 1995.

Michelmore:1962:EPM

[Mic62] Peter Michelmore. Einstein: Profile of the Man. Dodd, Mead & Company, New York, NY, USA, 1962. xii + 269 pp. LCCN QC16.E5 M623.

Michelmore:1963:EPM

[Mic63] Peter Michelmore. Einstein: Profile of the Man. Frederick Muller, London, UK, 1963. xi + 241 pp.

Michels:1968:BRB

[Mic68] Walter C. Michels. Book review: Letters on Wave Mechanics; Albert Einstein, Erwin Schroedinger, Max Planck, H. A. Lorentz, translated by Martin J. Klein from a collection edited by K. Przibram. American Journal of Physics, 36(8):766–767, August 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/?AJP/36/
766/2.

Michaelis:1978:AES

[Mic78] A. R. Michaelis. Albert Einstein as seen by artists. Interdisciplinary Science Reviews, 3(4):263–264, ???? 1978. ISSN 0308-0188.

Mie:2007:FTM

[Mie07a] Gustav Mie. Foundations of a theory of matter (excerpts). In Renn et al. [RSS^{+}07b], pages 633–697. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_36. Originally published as “Grundlagen einer Theorie der Materie”, Annalen der Physik, 37, 511–534 (1912), 40, 1–65 (1913).

Mie:2007:PRG

[Mie07b] Gustav Mie. The principle of the relativity of the gravitational potential. In Renn et al. [RSS^{+}07b], pages 729–743. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_38. Originally published as “Das Prinzip von der Relativität der Gravitationspotentials” in Arbeiten aus den Gebieten der Physik, Mathematik, Chemie, Festschrift Julius Elster und Hans Geitel zum sechzigsten Geburtstag, Braunschweig, Friedrich Vieweg und Sohn, 251–268 (1915).

Mie:2007:RCE

[Mie07c] Gustav Mie. Remarks concerning Einstein’s theory of gravitation. In Renn et al. [RSS^{+}07b], pages 699–728. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/

978-1-4020-4000-9_37. Originally published as “Bemerkungen zu der Einsteinschen Gravitationstheorie”, Physikalische Zeitschrift 15(3), 115–122 (1914) and 15(3), 169–176 (1914).

Mikulak:1965:RTS

[Mik65] Maxim William Mikulak. Relativity theory and Soviet communist philosophy, 1922–1960. ????, ????, 1965. 367 pp. LCCN HX541 .M538 1965a.

Millikan:1913:EEC

[Mil13] Robert Andrews Millikan. On the elementary electric charge and the Avogadro constant. Physical Review, 2(2): 109–143, August 1913. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.2.109;   http://physics.
nist.gov/cgi-bin/cuu/Value?e;   http://prola.aps.org/

abstract/PR/v2/i2/p109_1. In this paper, Millikan reported his famous oil-drop experiments of 1909 that led to a measurement of the unit of charge as e = 4.774 ± 0.009 × 10^{−10} esu, which corresponds to e = 1.592 ± 0.003 × 10^{−19} Coulomb. The best modern value (2011 CODATA Value: elementary charge) is e = 1.602 176 565(35) × 10^{−19} Coulomb. Millikan also reports Avogadro’s number as N = 6.062 ± 0.012 × 10^{23}. The 2010 CODATA value is N = 6.022 141 29(27) × 10^{23}.

Millikan:1916:EPE

[Mil16] Robert A. Millikan. Einstein’s photoelectric equation and contact electromotive force. Physical Review, 7(1):18–32, January 1916. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.7.18;   http://prola.aps.org/abstract/PR/v7/
i1/p18_1.

Millikan:1924:NPP

[Mil24] Robert Andrews Millikan. The Nobel Prize in Physics 1923 Lecture: May 23, 1924: The electron and the light-quant from the experimental point of view, 1924. URL http:

//en.wikipedia.org/wiki/Planck%27s_constant;   http:
//www.nobelprize.org/nobel_prizes/physics/laureates/

1923/. Millikan received the 1923 Nobel Prize in Physics “for his work on the elementary charge of electricity and on the photoelectric effect”. The latter work confirmed Einstein’s predictions of 1905 [Ein05d] about the quantum nature of that effect.

Millikan:1949:AEH

[Mil49a] Robert A. Millikan. Albert Einstein on his seventieth birthday. Reviews of Modern Physics, 21(3):343–345, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.343;   http://rmp.aps.org/

abstract/RMP/v21/i3/p343_1. Special issue in honor of Albert Einstein on his 70th birthday.

Milne:1949:GGR

[Mil49b] E. A. Milne. Gravitation without general relativity. In Schilpp [Sch49c], chapter 15, pages 409–435. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Miller:1975:AEM

[Mil75] Arthur I. Miller. Albert Einstein and Max Wertheimer: a gestalt psychologist’s view of the genesis of Special Relativity theory. History of Science (UK), 13(2):75–103, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/13/2/75.full.pdf+html.

Miller:1979:HSR

[Mil79] Arthur I. Miller. “On the History of the Special Relativity Theory”. In Aichelburg and Sexl [AS79], pages 89–108. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66. DM48.00. URL http://link.springer.

com/chapter/10.1007/978-3-322-91080-6_8/.

Miller:1981:AES

[Mil81] Arthur I. Miller. Albert Einstein’s Special Theory of Relativity: emergence (1905) and early interpretation, 1905–1911. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-04680-6, 0-201-04679-2 (paperback). xxviii + 466 pp. LCCN QC173.52 .M54.

Miller:1982:SRT

[Mil82] Arthur I. Miller. The Special Relativity Theory: Einstein’s response to the physics of 1905. In Holton and Elkana [HE82], pages 3–26. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Miller:1990:STY

[Mil90] Arthur I. Miller, editor. Sixty-two years of uncertainty: historical, philosophical, and physical inquiries into the foundations of quantum mechanics, volume 226 of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1990. ISBN 0-306-43608-6. LCCN QC173.96 .N385 1989.

Miller:1991:UCD

[Mil91] Richard L. (Richard Lee) Miller. Under the cloud: the decades of nuclear testing. Two-Sixty Press, The Woodlands, TX, USA, 1991. ISBN 1-881043-05-3. xii + 547 pp. LCCN U264 .M55 1991.

Miller:1992:AEJ

[Mil92] Arthur I. Miller. Albert Einstein’s 1907 Jahrbuch paper: The first step from SRT to GRT. In Eisenstaedt and Kox [EK92], pages 319–335. ISBN 0-8176-3479-7 (Boston), 3-7643-3479-7 (Berlin). LCCN QC173.6 .I572 1988.

Miller:1993:PLA

[Mil93] A. I. Miller. The private lives of Albert Einstein — R. Highfield, P. Carter. Nature, 366(6453):371–372, November 25, 1993. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Mills:1994:STQ

[Mil94] Robert Mills. Space, time, and quanta: an introduction to contemporary physics. W. H. Freeman and Company, New York, NY, USA, 1994. ISBN 0-7167-2436-7. xiv + 482 pp. LCCN QC21.2 .M57 1994.

Miller:1998:AES

[Mil98] Arthur I. Miller. Albert Einstein’s Special Theory of Relativity: emergence (1905) and early interpretation (1905–1911). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94870-8 (softcover). xix + 446 pp. LCCN QC173.52 .M54 1998.

Miller:1999:EFS

[Mil99] A. I. Miller. Einstein’s first steps toward General Relativity: Gedanken experiments and axiomatics. Physics in Perspective (PIP), 1(1):85–104, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/qx2xj7m80nx0ph79/fulltext.
pdf.

Miller:2001:EPS

[Mil01] Arthur I. Miller. Einstein, Picasso: space, time, and the beauty that causes havoc. Basic Books, New York, NY, USA, 2001. ISBN 0-465-01859-9 (paperback). x + 357 pp. LCCN N72.S3

M55 2001. URL http://www.loc.gov/catdir/enhancements/
fy0831/00065130-d.html.

Miller:2005:ESF

[Mil05a] Arthur I. Miller. Empire of the stars: friendship, obsession and betrayal in the quest for black holes. Little, Brown, London, UK, 2005. ISBN 0-316-72555-2 (hardcover). 400 pp. LCCN QB35 .M552 2005.

Miller:2005:ESO

[Mil05b] Arthur I. Miller. Empire of the stars: obsession, friendship, and betrayal in the quest for black holes. Houghton Mifflin, Boston, MA, USA, 2005. ISBN 0-618-34151-X. xx + 364 pp.

LCCN QB35 .M55 2005. URL http://www.loc.gov/catdir/
enhancements/fy0618/2004060909-d.html;   http://www.
loc.gov/catdir/enhancements/fy0618/2004060909-s.html;
 http://www.loc.gov/catdir/enhancements/fy0618/2004060909-
t.html.

Millet:2005:OPR

[Mil05c] Lydia Millet. Oh pure and radiant heart. Soft Skull Press, Brooklyn, NY, USA, 2005. ISBN 1-932360-85-9. 489 pp. LCCN

PS3563.I42175 O37 2005. URL http://www.loc.gov/catdir/
toc/ecip056/2005001028.html.

Millet:2006:OPR

[Mil06] Lydia Millet. Oh pure and radiant heart. Harcourt, Orlando, FL, USA, 2006. ISBN 0-15-603103-5 (paperback). 532 pp. LCCN

PS3563.I42175 O37 2006. URL http://www.loc.gov/catdir/
enhancements/fy0623/2005023808-b.html;   http://www.
loc.gov/catdir/enhancements/fy0623/2005023808-d.html;
 http://www.loc.gov/catdir/enhancements/fy0662/2005023808-
s.html;   http://www.loc.gov/catdir/enhancements/fy0662/
2005023808-t.html.

Miles:2008:PFI

[Mil08] Daniel Miles. The Phantom Fallout-Induced Cancer Epidemic in Southwestern Utah: Downwinders Deluded and Waiting to Die. BookSurge, Charleston, SC, USA, 2008. ISBN 1-4392-0647-3. xii + 130 pp. LCCN RA569 .M54 2008.

Milentijevic:2012:MMAa

[Mil10a] Radmila Milentijević. Mileva Marić Ajnštajn: Život sa Albertom Ajnštajnom. (Serbian) [Mileva Marić Einstein: life with Albert Einstein]. Matica Srpska, Novi Sad, Serbia, 2010. ISBN 86-7946-066-4. ???? pp. LCCN ????

Miller:2010:JPP

[Mil10b] Arthur I. Miller. 137: Jung, Pauli, and the Pursuit of a Scientific Obsession. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33864-9. xxiii + 336 pp. LCCN QC16.P37 M55 2010.

Miller:2011:BRB

[Mil11a] David Philip Miller. Book review: Bruce J. Hunt, Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Baltimore: Johns Hopkins University Press, 2010. Pp. x + 182. ISBN 978-0-8018-9359-9. £10.50 (paperback). British Journal for the History of Science, 44(4):609–610, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Miller:2011:PPL

[Mil11b] David Philip Miller. Pursuing power and light: Technology and physics from James Watt to Albert Einstein. British Journal for the History of Science, 44(163):609–610, December 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Milentijevic:2012:MMAb

[Mil12a] Radmila Milentijević. Mileva Marić Ajnštajn: Život sa Albertom Ajnštajnom. (Serbian) [Mileva Marić Einstein: life with Albert Einstein]. Prosveta, Belgrade, Serbia, 2012. ???? pp.

Miller:2012:YWM

[Mil12b] D. J. Miller. You are Wrong Mr Einstein! Newton, Einstein, Heisenberg and Feynman Discussing Quantum Mechanics, by Harald Fritzsch, Scope: general interest. Level: undergraduates and/or general readers. Contemporary Physics, 53(5):437–439, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milentijevic:2013:MME

[Mil13a] Radmila Milentijević. Mileva Marić Einstein: Vivre avec Albert Einstein. (French) [Mileva Marić Einstein: Life with Albert Einstein]. Ed. Âge d’Homme, Lausanne, Switzerland, 2013. ISBN 2-8251-4232-8. 491 pp. LCCN QC16.E52 M5414 2013.

Miles:2013:RCD

[Mil13b] Daniel W. Miles. Radioactive Clouds of Death over Utah: Downwinders’ Fallout Cancer Epidemic Updated. Trafford Publishing, ????, 2013. ISBN 1-4907-1097-3. 151 pp. LCCN RA569 .M55 2013.

Milentijevic:2015:MME

[Mil15] Radmila Milentijević. Mileva Marić Einstein: Life with Albert Einstein. United World Press, New York, NY, USA, 2015. ???? pp.

Minor:1984:AEW

[Min84] D. E. Minor. Albert Einstein on writing. Journal of Technical Writing and Communication, 14(1):13–18, ???? 1984. ISSN 0047-2816.

Minsky:1989:PAE

[Min89] Leonard Minsky. Problems at Albert Einstein. Nature, 339 (6219):10, May 4, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v339/n6219/pdf/339010b0.pdf.

Minkowski:2007:MRP

[Min07] Hermann Minkowski. Mechanics and the relativity postulate. In Renn et al. [RSS^{+}07a], pages 275–285. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/

978-1-4020-4000-9_20. Originally published as “Mechanik und Relativitätspostulat” appended to “Die Grundgleichungen für elektromagnetischen Vorgänge in bewegten Körpern” in Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Mathematisch-Physikalische Klasse, 53–111 (1908).

Mirsky:1994:AIN

[Mir94] Steve Mirsky. The annual Ig Nobel Prizes. Scientific American, 271(6):22–26, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v271/n6/
pdf/scientificamerican1294-22.pdf.

Mirsky:2002:EHT

[Mir02] Steve Mirsky. Einstein’s hot time. Scientific American, 287(3):102, September 2002. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v287/n3/
pdf/scientificamerican0902-102.pdf.

Mirsky:2004:EP

[Mir04] Steve Mirsky. Einstein’s parrot. Scientific American, 291 (1):118, July 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v291/n1/pdf/scientificamerican0704-
118.pdf.

Mirsky:2011:AGN

[Mir11] Steve Mirsky. Anti gravity: Noble Nobel faces. Scientific American, 305(3):94, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v305/n3/full/
scientificamerican0911-94.html;   http://www.nature.
com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
94.pdf.

Missner:1985:WEB

[Mis85] Marshall Missner. Why Einstein became famous in America. Social Studies of Science, 15(2):267–291, May 1, 1985. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic).

URL http://www.jstor.org/stable/285389;   https://
journals.sagepub.com/doi/pdf/10.1177/030631285015002003.

Mitchell:1987:ECN

[Mit87] Ralph G. Mitchell. Einstein and Christ: a new approach to the defence of the Christian religion, volume 5 of Theology and science at the frontiers of knowledge. Scottish Academic Press, Edinburgh, Scotland, 1987. ISBN 0-7073-0453-9. 231 pp. LCCN BL240.2.

Mitchell:2002:BBB

[Mit02] William C. (William Carl) Mitchell. Bye Bye Big Bang: Hello Reality. Cosmic Sense Books, Carson City, NV, USA, 2002. ISBN 0-9643188-1-4. xii + 444 pp. LCCN QB991.B54 M57 2002.

Mitton:2005:Fhl

[Mit05] Simon Mitton. Fred Hoyle: a life in science. Aurum, London, UK, 2005. ISBN 1-85410-961-8 (hardcover). xi + 369 + 8 pp. LCCN QB36.H75 M584 2005.

Mars-Jones:1990:VE

[MJ90] Adam Mars-Jones. Venus Envy, volume 14 of Chatto counterblasts. Chatto and Windus, London, UK, 1990. ISBN 0-7011-3585-9. 39 pp. LCCN PR6051.M5 E3636 1990.

Matson:2012:PNP

[MJ12] John Matson and Ferris Jabr. Physics: Nobel pursuits. Scientific American, 307(1):62–73, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v307/n1/
full/scientificamerican0712-62.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
62.pdf.

Mercier:1956:FJR

[MK56] Andre Mercier and Michel Kervaire, editors. Fünfzig Jahre Relativitätstheorie = Cinquantenaire de la théorie de la relativité = Jubilee of relativity theory: Bern, 11.-16. Juli 1955: Verhandlungen — actes — proceedings. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1956. LCCN ????

McHugh:1930:BADe

[MKBF30a] F. D. McHugh, Alexander Klemin, A. E. Buchanan, Jr., and Morris Fishbein. Scientific American digest: Huge lodestone specimen, Einstein, living immortal, and more. Scientific American, 143(6):466–484, December 1930. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v143/n6/pdf/scientificamerican1230-466.pdf.

McHugh:1930:SAD

[MKBF30b] F. D. McHugh, Alexander Klemin, A. E. Buchanan, Jr., and Morris Fishbein. Scientific American digest: Huge lodestone specimen, Einstein, living immortal, and more. Scientific American, 143(6):466–484, December 1930. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v143/n6/pdf/scientificamerican1230-466.pdf.

Mlodinow:2001:EWS

[Mlo01] Leonard Mlodinow. Euclid’s Window: the Story of Geometry from Parallel Lines to Hyperspace. Free Press, New York, NY, USA, 2001. ISBN 0-684-86523-8, 0-684-86524-6 (paperback), 0-7139-9634-X. xii + 306 pp. LCCN QA443.5.M56 2001.

Moore:1958:RSE

[MM58] Joan W. Moore and Burton M. Moore. The role of the scientific elite in the decision to use the atomic bomb. Social Problems, 6(1):78–85, ???? 1958. CODEN SOPRAG. ISSN 0037-7791 (print), 1533-8533 (electronic). URL http://www.jstor.org/

stable/798998.

Muller-Markus:1960:ESK

[MM66] Siegfried Müller-Markus. Einstein und die Sowjetphilosophie; Krisis einer Lehre. (German) [Einstein and the Soviet philosophy; a crisis in teaching]. Sovietica: Abhandlungen des Osteuropa-Instituts, Universität Freiburg/Schweiz [1]. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1960–1966. ???? pp. LCCN QC6 .M83.

Mullen:1983:UWC

[MM83] Peter Mullen and David Martin, editors. Unholy warfare: the church and the bomb. Blackwell, Oxford, UK, 1983. ISBN 0-631-13453-0, 0-631-13454-9. xvii + 247 pp. LCCN BR115.A85 U54 1983.

Mehra:2000:CMS

[MM00] Jagdish Mehra and Kimball A. Milton. Climbing the mountain: the scientific biography of Julian Schwinger. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850658-9 (hardcover), 0-19-852745-4 (paperback). xii + 677 pp. LCCN QC16.S29 M45

2000. URL http://www.loc.gov/catdir/enhancements/
fy0606/99087323-d.html;   http://www.loc.gov/catdir/
enhancements/fy0606/99087323-t.html.

Millar:1996:CDS

[MMMM96] David Millar, Ian Millar, John Millar, and Margaret Millar. The Cambridge Dictionary of Scientists. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56185-X (hardcover), 0-521-56718-1 (paperback). xii + 387 pp. LCCN Q141 .C128 1996. URL http://www.loc.

gov/catdir/description/cam027/95038471.html;   http:
//www.loc.gov/catdir/toc/cam021/95038471.html.

Mouton:2000:PTI

[MNOW00] W. G. Mouton, M. Naef, K. T. Otten, and H. E. Wagner. The progress of the treatment of intrarenal aortic aneurysms throughout the last century with the example of the death of Albert Einstein. International Journal of Artificial Organs, 23 (8):588, August 2000. ISSN 0391-3988.

MacAndrew:2000:LPT

[MNT^{+}00] Tim MacAndrew, Robert H. Norman, Jeff Templon, Kevin W. Wall, Shari Lawrence Pfleeger, Joseph C. Sligo, Christopher Jack, and Terry Ritter. Letters: Probability theory and software engineering; food for thought; uncovering erroneous assumptions; Einstein’s Nobel Prize; small-project process improvement; the truth about cryptography. Computer, 33(2): 4, 6–8, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.computer.

org/co/books/co2000/pdf/r2004.pdf.

Murphy:1995:WLN

[MO95] Michael P. (Michael Patrick) Murphy and Luke A. J. O’Neill, editors. What is life?: the next fifty years: speculations on the future of biology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-45509-X (hardcover). LCCN

QH331 .W465 1995. URL http://www.loc.gov/catdir/
description/cam026/94049438.html;   http://www.loc.
gov/catdir/toc/cam026/94049438.html. Contains most of

the contributions presented at a conference held at Trinity College, Dublin, from September 20–22, 1993.

Morsch:2006:DBE

[MO06] Oliver Morsch and Markus Oberthaler. Dynamics of Bose– Einstein condensates in optical lattices. Reviews of Modern Physics, 78(1):179–215, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.78.179;
http://rmp.aps.org/abstract/RMP/v78/i1/p179_1.

Martin:2013:CVA

[MO13] Walt Martin and Magda Ott, editors. The cosmic view of Albert Einstein. Sterling, New York, NY, USA, 2013. ISBN 1-4549-0776-2 (hardcover). xxiii + 151 pp. LCCN QC16.E5 A25

2013. URL http://catdir.loc.gov/catdir/enhancements/
fy1412/2014395622-b.html;   http://catdir.loc.gov/catdir/
enhancements/fy1412/2014395622-d.html.

Moffat:2008:RGP

[Mof08] John W. Moffat. Reinventing gravity: a physicist goes beyond Einstein. Smithsonian Books/Collins, New York, NY, USA, 2008. ISBN 0-06-117088-7. xiv + 272 pp. LCCN QC178 .M64 2008.

Moffat:2010:EWB

[Mof10] John W. Moffat. Einstein wrote back: my life in physics. Thomas Allen Publishers, Toronto, ON, Canada, 2010. ISBN 0-88762-615-7 (hardcover). 244 pp. LCCN QC16.M64 A3 2010.

Mogren:2002:WSU

[Mog02] Eric W. (Eric William) Mogren. Warm Sands: Uranium Mill Tailings Policy in the Atomic West. University of New Mexico Press, Albuquerque, NM, USA, 2002. ISBN 0-8263-2280-8 (hardcover). x + 241 pp. LCCN TD899.U73 M64 2002. URL

http://www.h-net.org/review/hrev-a0d4j6-aa.

Molina:1963:EES

[Mol63] Antonio Ma. Molina. Einstein’s epistemology of the scientific method. The Thomist: A Speculative Quarterly Review, 26(1): 100–110, January 1963. CODEN ???? ISSN 2473-3725.

Moldauer:1974:RAE

[Mol74] P. A. Moldauer. Reexamination of the arguments of Einstein, Podolsky, and Rosen. Foundations of Physics, 4(2):195–205, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00712686.

Moller:1984:WEQ

[Möl84] Horst Möller. Wissenschaft in der Emigration — Quantitative und geographische Aspekte. (German) [Science in exile — quantitative and geographical aspects]. Berichte zur Wissenschaftsgeschichte, 7(1):1–9, ???? 1984. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/bewi.
19840070102/abstract;   http://onlinelibrary.wiley.
com/journal/10.1002/%28ISSN%291522-2365/issues.

Monsma:1958:EGE

[Mon58] John Clover Monsma. The evidence of God in an expanding universe: forty American scientists declare their affirmative views on religion. Putnam, New York, NY, USA, 1958. 250 pp. LCCN BL240 .M715.

Monod:2004:EHK

[Mon04] Emmanuel Monod. Einstein, Heisenberg, Kant: methodological distinction and conditions of possibilities. Information and Organization, 14(2):105–121, April 2004. CODEN ???? ISSN 1471-7727 (print), 1873-7919 (electronic). URL https://dblp.

org/db/journals/iando/iando14.html#Monod04.

Monneret:2012:NVA

[Mon12] Claude Monneret. Nobel, vous avez dit Nobel. (French) [Nobel, you said Nobel]. Éditions Bénévent, Nice, France, 2012. ISBN 2-7563-2388-8 (paperback). 135 pp. LCCN ????

Monaldi:2019:SSR

[Mon19] Daniela Monaldi. The statistical style of reasoning and the invention of Bose–Einstein statistics. Berichte zur Wissenschaftsgeschichte, 42(4):307–337, December 2019. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Moore:1981:SBDa

[Moo81a] James R. Moore. Science and belief: from Darwin to Einstein. Block 1 (Unit 1), Beliefs in science: an introduction. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11000-2. 44 pp. LCCN Q175 .M7844 1981.

Moore:1981:SBDc

[Moo81b] James R. Moore. Science and belief: from Darwin to Einstein. Block 2: Unit 2–3, Science and metaphysics in Victorian Britain. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11001-0. 73 pp. LCCN BD111 .S34 1981.

Moore:1981:SBDb

[Moo81c] James R. Moore. Science and belief: from Darwin to Einstein. Block 7 (Unit 15), The future of science and belief: theological views in the twentieth century. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11006-1. 58 pp. LCCN BL240.2 .M63 1981.

Moore:1989:SLT

[Moo89] Walter John Moore. Schrödinger: life and thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35434-X. xi + 513 pp. LCCN QC16.S265 M66 1989. URL http://

www.loc.gov/catdir/description/cam023/88025807.html;
 http://www.loc.gov/catdir/toc/cam021/88025807.html.

Moore:1994:LES

[Moo94] Walter John Moore. A life of Erwin Schrödinger. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-46934-1 (paperback). xiii + 349 pp. LCCN QC16.S265

M65 1994. URL http://www.loc.gov/catdir/description/
cam026/93050147.html;   http://www.loc.gov/catdir/toc/
cam028/93050147.html.

Moore:1996:ENSb

[Moo96a] Mike Moore. Editor’s note: Shackling the genie. Bulletin of the Atomic Scientists, 52(2):2, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:ESV

[Moo96b] Walter John Moore. Erwin Schrödinger: una vida. (Spanish) [Erwin Schrödinger: a Life]. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55593-0 (paperback). ix + 444 pp. LCCN QC16.S265 M6518 1996. URL http://www.loc.

gov/catdir/description/cam027/95051783.html;   http:
//www.loc.gov/catdir/toc/cam023/95051783.html. Spanish

translation of [Moo94].

Moore:2002:EGI

[Moo02] Pete Moore. E = mc^{2}: the great ideas that shaped our world. Friedman/Fairfax Publishers, New York, NY, USA, 2002. ISBN 1-58663-699-5, 1-84222-554-5 (paperback). 192 pp. LCCN Q162

.M82 2002. URL http://www.loc.gov/catdir/description/
ste031/2003430735.html.

Moore:2010:NIN

[Moo10] Richard Moore. Nuclear illusion, nuclear reality: Britain, the United States and nuclear weapons, 1958–64. Nuclear weapons and international security since 1945. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23067-9 (hardcover). xv + 332 + 8 pp. LCCN UA647 .M66 2010.

Moore:2016:RGT

[Moo16] Kate Moore. The Radium Girls: They Paid with Their Lives, Their Final Fight Was for Justice. Simon and Schuster, New York, NY, USA, 2016. ISBN 1-4711-4757-6 (paperback), 1-4711-5387-8 (hardcover), 1-4711-5389-4 (e-book), 1-4926-4937-6 (ebook). xiii + 465 pp. LCCN HD6067.2.U6 M66 2016.

Moore:2017:RGD

[Moo17] Kate Moore. The Radium Girls: the Dark Story of America’s Shining Women. Sourcebooks, Inc., Naperville, IL, USA, 2017. ISBN 1-4926-4935-X (hardcover), 1-4926-5095-1 (paperback), 1-4926-4936-8 (e-book), 1-4926-4937-6 (e-book). xvi + 479 + 8 pp. LCCN HD6067.2.U6 M66 2017.

Morton:1957:DUA

[Mor57] Louis Morton. The decision to use the atomic bomb. Foreign Affairs, 35(2):334–353, January 1957. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20031230.

Morland:1979:HBS

[Mor79a] Howard Morland. The H-bomb secret: To know how is to ask why. The Progressive, 43(11):3–12, November 1979. CODEN

???? ISSN 0033-0736. URL http://progressive.org/?q=
node/2252.

Morrison:1979:AEM

[Mor79b] R. D. Morrison. Albert Einstein — the methodological unity underlying science and religion. Zygon, 14(3):255–266, ???? 1979. ISSN 0591-2385 (print), 1467-9744 (electronic).

Morrill:1995:NPM

[Mor95a] John E. Morrill. A Nobel Prize in mathematics. American Mathematical Monthly, 102(10):888–891, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morrison:1995:NTW

[Mor95b] Philip Morrison. Nothing is too wonderful to be true, volume 11 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-363-9 (hardcover). xi + 446 pp. LCCN Q173.

Morrison:1995:RNW

[Mor95c] Philip Morrison. Recollections of a nuclear war. Scientific American, 273(2):42–46, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v273/n2/
pdf/scientificamerican0895-42.pdf.

Morgan:1996:REN

[Mor96] P. Morgan. Reason embodied in nature — some notes towards the ultimate reality and meaning of Albert Einstein. Ultimate Reality and Meaning, 19(1):16–21, March 1996. ISSN 0709-549X.

Moring:2000:CIG

[Mor00] Gary Moring. The complete idiot’s guide to understanding Einstein. Alpha Books, Indianapolis, IN, USA, 2000. ISBN 0-585-22698-9 (e-book), 0-02-863180-3 (paperback). xix + 376 pp.

LCCN QC16.E5 M82 2000. URL http://site.ebrary.com/
id/10048483;   http://www.netLibrary.com/urlapi.asp?
action=summary\%26v=1\%26bookid=25368.

Moring:2004:CIG

[Mor04] Gary Moring. The complete idiot’s guide to understanding Einstein. Alpha Books, Indianapolis, IN, USA, second edition, 2004. ISBN 1-59257-185-9. xxii + 404 pp. LCCN QC16.E5 M67 2004.

Morita:2016:ECQ

[Mor16] Kunihisa Morita. Einstein’s criticism of quantum mechanics and Humean philosophy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):65–74, March 2016. CODEN HISCDU. ISSN 0285-4821.

Moszkowski:1921:EES

[Mos21] Alexander Moszkowski. Einstein: Einblicke in seine Gedankenwelt. Gemeinverstandliche Betrachtungen über die Relativitätstheorie und ein neues Weltsystem (German) [Einstein: Insights into his Thought World. Common Considerations on the Theory of Relativity and a New World System]. Hoffmann und Campe, Hamburg and Berlin, Germany, 1921. 240 (est.) pp.

Moszkowski:1922:EES

[Mos22] Alexander Moszkowski. Einstein, einblicke in seine gedankenwelt. (German) [Einstein, Views of his thought world]. F. Fontane and Co., Berlin, Germany, 1922. 240 pp. LCCN QC6 .M59; Microfilm 21570 QC. English translation in [Mos72].

Moss:1968:MWP

[Mos68] Norman Moss. Men who play God: the story of the H-bomb and how the world came to live with it. Harper & Row, New York, NY, USA, 1968. 352 pp. LCCN UF767 .M68 1968b.

Moszkowski:1970:CE

[Mos70] Alexander Moszkowski. Conversations with Einstein. Horizon Press, New York, NY, USA, 1970. ISBN 0-8180-0215-8. xxiii + 246 pp. LCCN QC6 .M5913. Translation by Henry L. Brose of [Mos22]. Introduction by Henry LeRoy Finch.

Moszkowski:1972:CE

[Mos72] Alexander Moszkowski. Conversations with Einstein. Sidgwick and Jackson, London, UK, 1972. ISBN 0-283-97924-0. xxiii + 246 pp. LCCN QC6 .M5913 1972. Translation by Henry L. Brose of [Mos22]. Introduction by Henry LeRoy Finch. Foreword by C. P. Snow.

Moss:1987:KFMa

[Mos87a] Norman Moss. Klaus Fuchs: the man who stole the atom bomb. Grafton, London, UK, 1987. ISBN 0-246-13158-6, 0-586-20299-4. 216 + 8 pp. LCCN UB271.R92F8.

URL http://catalog.hathitrust.org/api/volumes/oclc/
15195592.html;   https://www.math.utah.edu/pub/tex/
bib/einstein.bib.

Moss:1987:KFMb

[Mos87b] Norman Moss. Klaus Fuchs: the man who stole the atom bomb. St. Martin’s Press, New York, NY, USA, 1987. ISBN 0-312-01349-3. 216 + 8 pp. LCCN UB271.R92 F836 1987.

Moskowitz:2013:PMB

[Mos13a] Clara Moskowitz. Puzzling measurement of “Big G” gravitational constant ignites debate. Scientific American, ?? (??):??, September 18, 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.scientificamerican.com/article.cfm?id=puzzling-
measurement-of-big-g-gravitational-constant-ignites-
debate-slide-show.

Moskowitz:2013:SA

[Mos13b] Clara Moskowitz. Puzzling measurement of “Big G” gravitational constant ignites debate. Scientific American, ?? (??):??, September 18, 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.scientificamerican.com/article.cfm?id=puzzling-
measurement-of-big-g-gravitational-constant-ignites-
debate-slide-show.

Moses:2016:BDS

[Mos16a] George Moses. BBC documentary 2016: Secrets inside Einstein’s brain. Web video (98 minutes), June 12, 2016. URL

https://www.youtube.com/watch?v=Gi8NZpiC1M8.

Moszkowski:2016:EES

[Mos16b] Alexander Moszkowski. Einstein: Einblicke in seine Gedankenwelt. BookRix, München, Germany, 2016. ISBN 3-7368-3290-7 (e-book). ???? pp. LCCN ???? URL http://d-nb.info/

1088613195/34;   http://nbn-resolving.de/urn:nbn:de:
101:1-2016030922553.

Moskvitch:2018:TTA

[Mos18] Katia Moskvitch. Troubled times for alternatives to Einstein’s theory of gravity. Quanta Newsletter Web story.,

April 30, 2018. URL https://www.quantamagazine.org/
troubled-times-for-alternatives-to-einsteins-theory-
of-gravity-20180430.

Moskowitz:2021:CC

[Mos21] Clara Moskowitz. Cosmic conundrum. Scientific American, 324(2):24–29, February 2021. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://

www.scientificamerican.com/article/the-cosmological-
constant-is-physics-most-embarrassing-problem/.

Mott:1979:NPS

[Mot79] Sir Nevill Mott. Nobel prizes in science. Contemporary Physics, 20(2):227–228, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moulton:1920:ETS

[Mou20] H. Fletcher Moulton. The Einstein theory and spectral displacement. Nature, 104(2621):532, January 22, 1920. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v104/n2621/
pdf/104532c0.pdf.

Mourou:2019:NLE

[Mou19] Gerard Mourou. Nobel Lecture: Extreme light physics and application. Reviews of Modern Physics, 91(3):030501–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.030501.

Moyer:1979:RSE

[Moy79] Donald Franklin Moyer. Revolution in science: The 1919 eclipse test of general relativity. In Perlmutter and Scott [PS79], pages 55–101. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178 .O72

1979. URL http://link.springer.com/chapter/10.1007/
978-1-4684-3596-2_4/.

Moyer:1981:BRB

[Moy81] Albert E. Moyer. Book review: Albert Einstein, the Human Side: New Glimpses from His Archives by Helen Dukas; Banesh Hoffmann. Isis, 72(1):147–148, March 1981. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231238.

Merleau-Ponty:1974:LGT

[MP74] Jacques Merleau-Ponty. Leçons sur la genèse des théories physiques: Galilée, Ampère, Einstein. (French) [Lectures on the foundations of physical theories: Galileo, Ampère, Einstein]. Sciences de l’homme. J. Vrin, Paris, France, 1974. ISBN ???? 172 pp. LCCN ????

Medawar:1999:HGS

[MP99a] Jean S. Medawar and David Pyke. Hitler’s Gift: Scientists Who Fled Nazi Germany. Richard Cohen Books and the European Jewish Publication Society, London, UK, 1999. ISBN 1-86066-172-6. xx + 268 pp. LCCN Q141 .M385 1999.

Morgan:1999:AGO

[MP99b] K. Z. (Karl Ziegler) Morgan and Ken M. Peterson. The angry genie: one man’s walk through the nuclear age. University of Oklahoma Press, Norman, OK, USA, 1999. ISBN 0-8061-3122-5. xvii + 218 pp. LCCN R895.6.U6 M67 1999.

Medawar:2001:HGT

[MP01] Jean S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Medawar:2012:HGT

[MP12] Jean S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Books, New York, NY, USA, 2012. ISBN 1-61145-709-2. xx + 268 pp. LCCN Q141 .M42 2012. Foreword by Max Perutz.

Marino:2009:TDE

[MPBL09] A. M. Marino, R. C. Pooser, V. Boyer, and P. D. Lett. Tunable delay of Einstein–Podolsky–Rosen entanglement. Nature, 457(7231):859–862, February 12, 2009. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v457/n7231/
full/nature07751.html.

Martin:2008:EE

[MPM^{+}08] Philip Martin, Mark Pybus, Peter Moffat, David Tennant, and Andy Serkis. Einstein and Eddington. BBC Video film., 2008. A look at the evolution of Albert Einstein’s theory of relativity, and Einstein’s relationship with British scientist Sir Arthur Eddington, the first physicist to understand his ideas.

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82d, MR82e, MR87b, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The formulation of matrix mechanics and its modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44

vol. 3. URL http://www.loc.gov/catdir/enhancements/
fy1005/82003253-d.html;   http://www.loc.gov/catdir/
enhancements/fy1005/82003253-t.html. See other volumes

in this series [MR82a, MR82c, MR82d, MR82e, MR87b, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4. See other volumes in this series [MR82a, MR82b, MR82d, MR82e, MR87b, MR00].

Mehra:1982:QTPa

[MR82d] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR87b, MR00].

Mehra:1982:QTPb

[MR82e] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR87b, MR00].

Mehra:1987:ESRc

[MR87a] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol.

5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR00].

Mehra:1987:ESR

[MR87b] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol.

5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR00].

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The completion of quantum mechanics, 1926–1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp. LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR87b].

Morlot:2017:IAE

[MRS17] Frédéric Morlot, Anne-Margot Ramstein, and Anna Street. Les illuminations d’Albert Einstein. (French) [Albert Einstein’s bright ideas]. Plato and Co. Diaphanes, Zurich, Switzerland, 2017. ISBN 3-03734-935-2, 3-03734-936-0. 63 pp. LCCN ????

Mermin:1982:JDL

[MS82] N. D. Mermin and Gina M. Schwarz. Joint distributions and local realism in the higher-spin Einstein–Podolsky–Rosen experiment. Foundations of Physics, 12(2):101–135, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00736844.

Moore:2009:NPD

[MS09] Samuel K. Moore and Neil Savage. The Nobel Prize and its discontents. IEEE Spectrum, 46(12):11–12, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morton-Smith:2015:O

[MS15] Tom Morton-Smith. Oppenheimer. Oberon Books Ltd., London, UK, 2015. ISBN 1-78319-198-8. ???? pp. LCCN ????

URL http://stratford-upon-avon-theatre.blogspot.com/
2015/01/rsc-oppenheimer-review-stratford-upon.html.

Maier:2017:MFE

[MS17] Corinne Maier and Anne Simon. Marx, Freud, and Einstein: heroes of the mind. Nobrow, London, UK, 2017. ISBN 1-910620-31-9. 189 pp. LCCN HX39.5 .M253313 2017.

Maier:2015:EGN

[MSK15] Corinne Maier, Anne Simon, and Anja Kootz. Einstein: Ein Graphic Novel. (German) [Einstein: a graphic novel]. Knesebeck, München, Germany, 2015. ISBN 3-86873-809-6. 64 pp. LCCN ????

Madarasz:2017:TDF

[MSS17] Judit X. Madarász, Gergely Székely, and Mike Stannett. Three different formalisations of Einstein’s relativity principle. The Review of Symbolic Logic, 10(3):530–548, ???? 2017. CODEN ???? ISSN 1755-0203 (print), 1755-0211 (elec-

tronic). URL https://dblp.org/db/journals/rsl/rsl10.
html#MadaraszSS17.

Madarasz:2022:GWT

[MSS22] Judit X. Madarász, Mike Stannett, and Gergely Székely. Groups of WorldView transformations implied by Einstein’s special principle of relativity over arbitrary ordered fields. The Review of Symbolic Logic, 15(2):334–361, ???? 2022. CODEN

???? ISSN ???? URL https://dblp.org/db/journals/rsl/
rsl15.html#MadaraszSS22.

Muller:2007:RTC

[MST^{+}07a] Holger Müller, Paul Louis Stanwix, Michael Edmund Tobar, Eugene Ivanov, Peter Wolf, Sven Herrmann, Alexander Senger, Evgeny Kovalchuk, and Achim Peters. Relativity tests by complementary rotating Michelson–Morley experiments. arXiv.org,

2007. URL http://arxiv.org/abs/0706.2031v1. Published

in [MST^{+}07b].

Muller:2007:TRC

[MST^{+}07b] Holger Müller, Paul Louis Stanwix, Michael Edmund Tobar, Eugene Ivanov, Peter Wolf, Sven Herrmann, Alexander Senger, Evgeny Kovalchuk, and Achim Peters. Tests of Relativity by complementary rotating Michelson–Morley experiments. Physical Review Letters, 99(5):050401:1–050401:4, August 3, 2007. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://arxiv.org/abs/

0706.2031v1;   http://prl.aps.org/abstract/PRL/v99/
i5/e050401.

Mendelsohn:1988:STM

[MSW88] Everett Mendelsohn, Merritt Roe Smith, and Peter Weingart, editors. Science, technology, and the military, volume 12/1, 12/2 of Sociology of the sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2780-5 (vol. 1), 90-277-2783-X (vol. 2). xxx + 288 (vol. 1), vii + 292–562 (vol. 2) pp. LCCN U42 .S35

1988. URL http://link.springer.com/book/10.1007/978-
94-017-2958-1.

Ma:2012:PWN

[MSYW12] Caifeng Ma, Cheng Su, Junpeng Yuan, and Yishan Wu. Papers written by Nobel Prize winners in physics before they won the prize: an analysis of their language and journal of publication. Scientometrics, 93(3):1151–1163, December 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1007/s11192-
012-0748-z.

Morrison:1998:REH

[MT98] Philip Morrison and Kosta Tsipis. Reason enough to hope: America and the world of the twenty-first century. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-13344-X (hardcover). xv + 210 pp. LCCN JZ5675 .M67 1998.

Meyer-Ter-Vehn:1988:ESP

[MTVB88] J. Meyer-Ter-Vehn and Th. Bührke. EIR und SIN zum Paul-Scherrer-Institut zusammengelegt/Stuttgarter Max-Planck-Institute erweitert/Geheimer Fortschritt bei der Kernfusion/ Nobelpreis: Preisgeld erhöht/Océ-van der Grinten-Preis 1988/ Grünes Licht für den Bau eines Infrarot-Observatoriums/Neu. (German) [EIR and SIN merged to form the Paul Scherrer Institute/Stuttgart Max Planck Institute expanded/Secret advances in nuclear fusion/Nobel Prize: prize money increased/Océ van der Grinten Prize 1988/Green light for the construction of an infrared observatory/news]. Physik Journal, 44(5):128–154, May 1988. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Misner:1973:GCC

[MTW73] Charles W. Misner, Kip S. Thorne, and John Archibald Wheeler. Gravitation. W. H. Freeman and Company, New York, NY, USA, 1973. ISBN 0-7167-0334-3, 0-7167-0344-0 (paperback). xxvi + 1279 pp. LCCN QC178 .M57.

Mercier:1979:GRA

[MTY79] André Mercier, Hans Jürgen Treder, and Wolfgang Yourgrau. On General Relativity: an analysis of the fundamentals of the Theory of General Relativity and Gravitation. Akademie-Verlag, Berlin, Germany, 1979. 154 pp. LCCN QC173.6 .M45.

Muhling:2011:ERW

[Müh11] Markus Mühling. Einstein und die Religion: das Wechselverhältnis zwischen religiös-weltanschaulichen Gehalten und naturwissenschaftlicher Theoriebildung Albert Einsteins in seiner Entwicklung. (German) [Einstein and Religion: the changing relationship between religious-ideological contents and Albert Einstein’s scientific theory in its development], volume 23 of Religion, Theologie und Naturwissenschaft. Vandenhoeck and Ruprecht, Göttingen, Germany, 2011. ISBN 3-525-56989-0. 392 pp. LCCN QC16.E5 M84 2011. EUR 49.95.

Muirden:1994:HDW

[Mui94] James Muirden. How do we know Einstein was right? Simon and Schuster Young Books, Hemel Hempstead, UK, 1994. ISBN 0-7500-1521-7. 45 pp.

Mukunda:2005:CAE

[Muk05] N. Mukunda. A celebration of Albert Einstein. Current Science, 89(12):1973–1974, December 25, 2005. CODEN CUS-CAM. ISSN 0011-3891.

Mullner:1999:DGR

[Mul99] Ross M. Mullner. Deadly Glow: the Radium Dial Worker Tragedy. American Public Health Association, Washington, DC, 1999. ISBN 0-87553-245-4. xii + 175 pp. LCCN RC965.R25 M85 1999.

Muldoon:2002:DPK

[Mul02] Ciara Muldoon. Did Picasso know about Einstein? Physics World, 15(11):30–31, November 2002. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http:/

/physicsworld.com/cws/article/print/2002/nov/01/did-
picasso-know-about-einstein;   http://physicsworldarchive.
iop.org/full/pwa-pdf/15/11/phwv15i11a33.pdf.

Muldoon:2005:EFN

[Mul05] Ciara Muldoon. Einstein and Freud: novel thoughts. Physics World, 18(6):42–43, June 2005. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/6/phwv18i6a40.pdf.

Muller:2010:PTF

[Mul10] R. (Richard) Muller. Physics and technology for future presidents: an introduction to the essential physics every world leader needs to know. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13504-5 (hardcover). xi + 517 pp. LCCN QC24.5 .M85 2010.

Mundlak:1936:CPR

[Mun36] Max Mundlak. The consequences of philosophy: a reply to Planck and Einstein. J. Bale, Sons and Danielsson, Ltd., London, UK, 1936. 86 pp. LCCN Q175 .M98.

Murnaghan:1925:RPR

[Mur25] F. D. Murnaghan. Recent publications: Reviews: Sidelights on Relativity, by Albert Einstein, G. B. Jeffery, and W. Perrett; The Principle of Relativity with Applications to Physical Science, by A. N. Whitehead. American Mathematical Monthly, 32(6):311–313, June 1925. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL 2-T; http:/

/links.jstor.org/sici?sici=0002-9890(192506/07)32:6<
311:SOR>2.0.CO.

Murdoch:1990:NBP

[Mur90] Dugald Murdoch. Niels Bohr’s Philosophy of Physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-33320-2 (hardcover), 0-521-37927-X (paperback). x + 294 pp. LCCN QC16.B6 M87 1987. US$16.95.

Murdoch:1994:BED

[Mur94] Dugald Murdoch. The Bohr–Einstein dispute. In Faye and Folse [FF94], chapter 14, pages 303–324. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://link.

springer.com/chapter/10.1007/978-94-015-8106-6_14.

Murray:2011:BEE

[Mur11] Lydia Murray. Between Elmo and Einstein. Nature, 477(7362): 1, August 31, 2011. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

naturejobs/2011/110901/full/nj7362-1.html.

Musser:2004:WER

[Mus04]  George Musser. Was Einstein right?  Scientific Amer-

ican, 291(3):88–91, September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v291/n3/
pdf/scientificamerican0904-88.pdf.

Musser:2015:QPC

[Mus15a] George Musser. Quantum physics: Is the Cosmos random? Scientific American, 313(3):88–93, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-88.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-88.pdf.

Musser:2015:SAD

[Mus15b] George Musser. Spooky action at a distance: the phenomenon that reimagines space and time — and what it means for black holes, the big bang, and theories of everything. Scientific American/Farrar, Straus and Giroux, New York, NY, USA, 2015. ISBN 0-374-29851-3 (hardcover), 0-374-71355-3 (e-book). ???? pp. LCCN QC173.59.S65 M88 2015.

Marianoff:1944:EIS

[MW44] Dimitri Marianoff and Palma Wayne. Einstein: an intimate study of a great man. Doubleday, Doran, Garden City, NY, USA, 1944. x + 211 pp. LCCN QC16.E5 M3.

Masters:1946:OWN

[MW46a] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Masters:1946:VEI

[MW46b] Dexter Masters and Katherine Way, editors. Een Verden eller ingen. (Danish) [One world or none]. ????, København, Danmark, 1946. 220 pp. LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Marage:1999:SCB

[MW99] Pierre Marage and Grégoire Wallenborn, editors. The Solvay councils and the birth of modern physics, volume 22 of Science networks, historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-5705-3 (Basel), 0-8176-5705-3 (Boston). xiii + 224 pp. LCCN QC7.

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8

O54 2007.  URL http://thenewpress.com/index.php?
option=com_title&task=view_title&metaproductid=1703;
 http://www.loc.gov/catdir/toc/ecip0718/2007020838.

htm. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Muller:2011:SRV

[MW11] Thomas Muller and Daniel Weiskopf. Special-relativistic visualization. Computing in Science and Engineering, 13(4):85–93, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Myridis:2021:BRL

[Myr21a] N. E. Myridis. Book review: The large-scale structure of the Universe, by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton University Press, 2020, 448 pp., £50.00 (paperback), ISBN: 978-0-691-20983-8. Contemporary Physics, 62 (1):66–67, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2021:BRC

[Myr21b] Nikolaos E. Myridis. Book review: Cosmology’s century: an inside history of our modern understanding of the universe Physics & Astronomy, by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton, NJ, Princeton University Press, 2020, 440 pp., $28.00 (hardback), ISBN (e-book) 978-0-691-20166-5. Scope: monograph. Level: non-specialists, undergraduate, advanced undergraduate, postgraduate, researcher, teacher, specialist, scientist. Contemporary Physics, 62(4):239–240, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2021:BRP

[Myr21c] Nikolaos E. Myridis. Book review: Principles of physical Cosmology by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton, NJ, Princeton University Press, 2020, 774 pp., £62.00 (e-book), ISBN 978-0-691-20672-1. Princeton series in physics & astronomy. Scope: reference. Level: nonspecialists, advanced undergraduate, postgraduate, researcher, teacher, specialist, scientist. Contemporary Physics, 62(4):243– 244, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

N:1938:RPA

[N.38] E. N. Book review: Physik als Abenteuer der Erkenntnis by Albert Einstein, Leopold Infeld. The Journal of Philosophy, 35(18):500–501, September 1, 1938. ISSN 0022-362X (print), 1939-8549 (electronic). URL http://links.jstor.org/

sici?sici=0022-362X(19380901)35:18<500:PAADE>2.0.CO%
3B2-O;   http://www.jstor.org/stable/2017661.

NunesdosSantos:1992:EEV

[NA92] A. M. Nunes dos Santos and Christopher Auretta. Eddington e Einstein: verificão experimental da teoria da relatividade generalizada na Ilha do Príncipe, volume 17 of Panfletos Gradiva. Gradiva, Lisboa, Portugal, 1992. ISBN 972-662-252-2s. 144 pp. LCCN QC173.57 .E33 1992.

Nachtikal:1924:EPS

[Nac24] František Nachtikal. Einsteinův posuv spektrálních car slunečních. (Czech) [Einstein’s shift of solar spectral lines]. Časopis pro pěstování matematiky a fysiky, 53(4):414–416, ???? 1924.

Nackoney:1973:GIB

[Nac73] R. W. Nackoney. The gravitational influence of a beam of light. I. Journal of Mathematical Physics, 14(9):1239–1247, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nackoney:1977:GIB

[Nac77] R. W. Nackoney. The gravitational influence of a beam of light of variable flux. Journal of Mathematical Physics, 18(11):2146– 2152, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nachmansohn:1979:GJP

[Nac79] David Nachmansohn. German–Jewish pioneers in science, 1900–1933: highlights in atomic physics, chemistry, and biochemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-90402-6, 3-540-90402-6. xx + 388 pp. LCCN QD21 .N33. URL http://

catalog.hathitrust.org/api/volumes/oclc/4804322.html.

Nachtigall:19xx:IPM

[Nacxx] D. K. Nachtigall, editor. Internalizing physics: making physics part of one’s life: eleven essays of Nobel Laureates, volume 48 of Science and technology education document series. Education Sector, UNESCO, Paris, France, 19xx. iv + 114 pp. LCCN ????

Nader:1979:BRB

[Nad79] Laura Nader. Book reviews: Scientific Elite: Nobel Laureates in the United States, by Harriet Zuckerman. Bulletin of the Atomic Scientists, 35(1):40–41, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadis:2002:NSJ

[Nad02] Steve Nadis. News scan: Joke hunter of science: Funnyman Marc Abrahams tackles an improbable role and an Ig Nobel cause. Scientific American, 286(4):28, April 2002. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/2002/0402issue/0402scicit6.html.

Nagel:1950:SRC

[Nag50] Ernest Nagel. Species and relativity: Charles Darwin. The Naturalist as a Cultural Force, by Paul B. Sears, and Albert Einstein. His Work and Its Influence on our World, by Leopold Infeld. New York Times, ??(??):BR4, February 26, 1950. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/111510278.

Nagasaka:1971:EPR

[Nag71] Gen-Ichiro Nagasaka. The Einstein–Podolsky–Rosen Paradox reexamined. In Buck and Cohen [BC71], pages 437–445. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-010-3142-4_27.

Nagel:1982:DCN

[Nag82] Bengt Nagel. The discussion concerning the Nobel Prize for Max Planck. In Bernhard et al. [BCS82], pages 352– 376. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500281.

Nahin:2001:TTE

[Nah01] P. J. Nahin. Time travel in Einstein’s universe, J. Richart Gott, Houghton Mifflin, New York, 2001. ISBN: 0-395-95563-7, 291 pp., $25 [books]. IEEE Spectrum, 38(7):79–80, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nambu:2009:NLS

[Nam09] Yoichiro Nambu. Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization. Reviews of Modern Physics, 81(3):1015–1018, July 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1015;   http://rmp.aps.org/abstract/RMP/v81/i3/p1015_

1. See note [Nam10].

Nambu:2010:PNN

[Nam10] Yoichiro Nambu. Publisher’s note: Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization [Rev. Mod. Phys. 81, 1015 (2009)]. Reviews of Modern Physics, 82(4):3199, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.3199;
 http://rmp.aps.org/abstract/RMP/v82/i4/p3199_1. See

[Nam09].

Nansen:1978:P

[Nan78] Fridtjof Nansen. Perspective: 1922. Bulletin of the Atomic Scientists, 34(7):1, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the author’s Nobel Prize Lecture, 29 December 1922, Oslo, Norway.

Narlikar:2010:IR

[Nar10] Jayant Vishnu Narlikar. An Introduction to Relativity. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51497-5 (hardback), 0-521-73561-0 (paperback). viii + 363 pp. LCCN QC173.6 .N369 2010; for. URL http://assets.

cambridge.org/97805215/14972/cover/9780521514972.jpg.

NMHT:1979:ECEc

[Nat79] National Museum of History and Technology, editor. Einstein, a centenary exhibition [March 1979–March 1980]. Smithsonian Institution Press, Washington, DC, USA, 1979. ISBN 87-88104-62-1. 48 pp. LCCN QC16.E5 N37 1979.

Nathan:1981:EPb

[Nat81] Otto Nathan. The Einstein papers. Science, 213(4514):1313– 1314, September 18, 1981. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/213/4514/1313.2.full.pdf.

Natarajan:2005:EAD

[Nat05] Vasant Natarajan. Einstein as armchair detective: The case of stimulated radiation. Resonance, 10(12):148–162, December 2005. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X

(electronic). URL http://www.springerlink.com/content/
n3877601p514w672/.

Nauenberg:2016:EEP

[Nau16] Michael Nauenberg. Einstein’s equivalence principle in quantum mechanics revisited. American Journal of Physics, 84(11):879– 882, November 2016. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Nauenberg:2017:JSB

[Nau17] Michael Nauenberg. John Stewart Bell and Twentieth-Century Physics: Vision and Integrity. Whitaker, Andrew. 476 pp. Oxford U.P., New York, 2016. Price: $44.95 (hardcover). ISBN 978-0-19-874299-9. American Journal of Physics, 85(11):880, November 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Navon:1982:EPI

[Nav82] Yitzhak Navon. On Einstein and the Presidency of Israel. In Holton and Elkana [HE82], pages 293–296. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Navarro:1991:ESM

[Nav91] Luis Navarro. On Einstein’s statistical-mechanical approach to the early quantum theory (1904–1916). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):39–58, June 1991. CODEN HISCDU. ISSN 0285-4821.

Navarroveguillas:1994:AEL

[Nav94] L. Navarroveguillas. Albert Einstein — life, work and philosophy — Spanish — J. Merleauponty, A. L. Sanzsaenz, translator. Arbor — Ciencia Pensamiento y Cultura, 149(588):112– 116, December 1994. ISSN 0210-1963.

Navarro:1998:GEF

[Nav98] Luis Navarro. Gibbs, Einstein and the foundations of statistical mechanics. Archive for History of Exact Sciences, 53(2):147– 180, June 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=2&spage=147.

Navarro:2015:BRM

[Nav15] Jaume Navarro. Book reviews: Milena Wazeck, Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s. Translated by Geoffrey S. Koby. Cambridge: Cambridge University Press, 2014. ISBN 978-1-107-01744-3. £65.00/$99.00 (hardback). British Journal for the History of Science, 48(2):375–377, 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Navarro:2021:WEH

[Nav21] Jaume Navarro. Whittaker, Einstein, and the history of the aether: Alternative interpretation, blunder, or bigotry? History of Science (UK), 59(3):287–314, September 1, 2021. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275320968408.

Noel-Baker:1979:PNP

[NB79] Philip Noel-Baker. Perspective: Nobel Peace Laureates: 1959. Bulletin of the Atomic Scientists, 35(5):1, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 10 December 1959, Oslo, Norway.

Nussbaumer:2009:DEU

[NB09] Harry Nussbaumer and Lydia Bieri. Discovering the expanding universe. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51484-3 (hardcover). xvii + 226 pp. LCCN QB981 .N865 2009.

Nowik-Boltyk:2012:SNU

[NBDD^{+}12] P. Nowik-Boltyk, O. Dzyapko, V. E. Demidov, N. G. Berloff, and S. O. Demokritov. Spatially non-uniform ground state and quantized vortices in a two-component Bose–Einstein condensate of magnons open. Scientific Reports, 2, June 29, 2012. CODEN SRCEC3. ISSN ???? URL http://www.nature.com/

srep/2012/120629/srep00482/full/srep00482.html.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57(6):546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.57.546. This paper reports experi-

ments that confirm the Bohr–Wheeler prediction [BW39b] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Ndiaye:2007:NBD

[Ndi07] Pap Ndiaye. Nylon and bombs: DuPont and the march of modern America. Studies in industry and society. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8444-6 (hardcover). 289 + 8 pp. LCCN

HD9651.9.D8 N3513 2007. URL http://www.loc.gov/catdir/
enhancements/fy0703/2006004143-b.html;   http://www.
loc.gov/catdir/enhancements/fy0703/2006004143-d.html;
 http://www.loc.gov/catdir/toc/ecip068/2006004143.html.

Neemann:2006:OET

[Ne’06] Yuval Ne’emann. One Einstein, two scientific revolutions, three routes to unification, four very dark clouds. AIP Conference Proceedings, 861(1):245–246, 2006. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

URL http://scitation.aip.org/content/aip/proceeding/
aipcp/10.1063/1.2399580.

Neeman:1979:CAP

[Nee79] Y. Neeman. Coherence, abstraction, and personal involvement — Albert Einstein, physicist and humanist. Impact of Science on Society, 29(1):17–25, ???? 1979. ISSN 0019-2872.

Neffe:2005:EBG

[Nef05] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography]. Rowohlt, Reinbek bei Hamburg, Germany, second edition, 2005. ISBN 3-498-04685-3. 490 + 14 pp. LCCN QC16.E5 N44 2005.

Neffe:2006:EBG

[Nef06] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography], volume 61937 of rororo. Rowohlt Taschenbuch Verlag, Reinbek bei Hamburg, Germany, 2006. ISBN 3-499-61937-7 (paperback). 490 + 8 pp. LCCN ???? EUR 9.90, SFR 17.90.

Neffe:2007:EB

[Nef07] Jürgen Neffe. Einstein: a biography. Farrar, Strauss, and Giroux, New York, NY, USA, 2007. ISBN 0-374-14664-0 (hardcover). x + 461 + 16 pp. LCCN QC16.E5 N4413

2007. URL http://www.loc.gov/catdir/enhancements/

fy0705/2006026136-b.. Translated by Shelley Frisch from the original German edition Einstein: eine Biographie (ISBN 3-498-04685-3, 3-499-61937-7) [Nef05, Nef06].

Neffe:2011:AEA

[Nef11] Jürgen Neffe. Albert Einstein and the atomic bomb. History Reader, ??(??):??, August 6, 2011. URL http://www.

thehistoryreader.com/modern-history/albert-einstein-
atomic-bomb/.

Nelson:1977:CPN

[Nel77] Bryce Nelson. Corn patch Nobel laureate. Bulletin of the Atomic Scientists, 33(8):48–50, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:2014:ARE

[Nel14] Craig Nelson. The age of radiance: the epic rise and dramatic fall of the atomic era. Scribner, New York, NY, USA, 2014. ISBN 1-4516-6043-X (hardcover), 1-4516-6044-8 (paperback), 1-4516-6045-6 (e-book). 4a37 pp. LCCN QC773 .N45 2014.

Nersessian:1984:FEC

[Ner84] Nancy J. Nersessian. Faraday to Einstein: constructing meaning in scientific theories. Science and philosophy. M. Nijhoff, Dordrecht, The Netherlands, 1984. ISBN 90-247-2997-1, 0-7923-0950-2 (paperback). xiv + 196 pp. LCCN Q175 .N3885 1984.

Nersessian:1986:WWL

[Ner86] Nancy J. Nersessian. “Why wasn’t Lorentz Einstein?” An examination of the scientific method of H. A. Lorentz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3):205–242, September 1986. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.
1111/j.1600-0498.1986.tb00857.x/abstract.

Nesvadba:1960:EUM

[Nes60] Josef Nesvadba. Einsteintův mozek. (Czech) [Einstein’s brain]. Mladá Fronta, Prague, Czechoslovakia, 1960. 162 pp.

Nesbit:2014:WANc

[Nes14] Tarashea Nesbit. The wives of Los Alamos: a novel. Bloomsbury, New York, NY, USA, 2014. ISBN 1-62040-503-2 (hardcover). 233 pp. LCCN PS3614.E467 W58 2014.

Neuenschwande:2006:TEE

[Neu06] Dwight E. Neuenschwande. Taking Einstein’s ethics into the 21st century: “remember your humanity”. The SPS Observer,

23(2):??, Summer 2006. URL http://www.spsobserver.org/
2006/observer_einstein.pdf.

Neve:1992:AEM

[Nev92] M. Neve. Albert Einstein, Mileva Marić — the Love letters — J. Renn, R. Schulmann, Editors. TLS — The Times Literary Supplement, ??(4659):11, July 17, 1992. ISSN 0307-661X.

Newman:1956:WMS

[New56] James R. Newman. The world of mathematics: a small library of the literature of mathematics from A’h-mose the Scribe to Albert Einstein. G. Allen, London, UK, 1956. xviii + 2535 pp. LCCN ???? Four volumes.

Newman:1960:WMS

[New60] James Roy Newman, editor. The World of Mathematics: a small library of the literature of mathematics from A’h-mosé the scribe to Albert Einstein: presented with commentaries and notes. Allen and Unwin, London, UK, 1960. xviii + 2535 pp. LCCN ???? Four volumes.

Times:1965:HP

[New65] New York Times. Hiroshima Plus 20. Delacorte Press, New York, NY, USA, 1965. vi + 211 pp. LCCN UF767 N448. Introduction by John W. Finney. Special quotes selected and compiled by Ruth Block.

Newhouse:1988:WPN

[New88] John Newhouse. War and peace in the nuclear age. Vintage Books, New York, NY, USA, 1988. ISBN 0-394-72645-4. xii + 486 pp. LCCN U263 .N48 1990.

Newhouse:1989:NAH

[New89] John Newhouse. The nuclear age: from Hiroshima to Star Wars. Michael Joseph, London, UK, 1989. ISBN 0-7181-3263-7. xii + 486 + 48 pp. LCCN D842 .N38x 1989.

Newton:2009:HPC

[New09] Roger G. Newton. How Physics Confronts Reality: Einstein Was Correct, But Bohr Won the Game. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2009. ISBN 981-4277-02-9 (hardcover), 981-4277-03-7 (paperback). ix + 147 pp. LCCN QC173.98 .N49 2009.

Noah-Gray:2007:WKE

[NG07]  Noah-Gray. What kept Einstein smart?  Action Po-

tential, August 22, 2007. CODEN ???? ISSN ????

URL http://blogs.nature.com/actionpotential/2007/08/
what_kept_einstein_smart_1.html.

Ng:2011:CME

[Ng11] Yew-Kwang Ng. Common mistakes in economics by the public, students, economists, and Nobel Laureates. Economic issues, problems and perspectives. Nova Science Publishers, New York, NY, USA, 2011. ISBN 1-61761-606-0. xv + 171 pp. LCCN HB171 .N56 2011.

Nieto:1989:AED

[NHG89] Michael Martin Nieto, Richard J. Hughes, and T. Goldman. Actually, eötvös did publish his results in 1910, it s just that no one knows about it …. American Journal of Physics, 57(5): 397–404, May 1989. CODEN AJPIAS. ISSN 1943-2909.

Nelkowski:1979:ESB

[NHP^{+}79] H. Nelkowski, A. Hermann, H. Poser, R. Schrader, and R. Seiler, editors. Einstein Symposium Berlin, aus Anlaß der 100. Wiederkehr seines Geburtstages, 25. bis 30. März 1979. (German) [Einstein Symposium Berlin, on the occasion of the hundredth anniversary of his birthday, 25–30 March, 1979], volume 100 of Lecture Notes in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. viiii + 550 pp. DM49.50, US$27.80.

Nicolai:1917:BKB

[Nic17] Georg Friedrich Nicolai. Die Biologie des Krieges: Betrachtungen eines deutschen Naturforschers. (German) [The Biology of War: Reflections of a German Scientist]. Orell Füssli, Zürich, Switzerland, 1917. 463 pp. LCCN U21 .N4 1917. The introduction contains the Einstein manifesto [ENFB14].

Nicolai:1918:BW

[Nic18] Georg Friedrich Nicolai. The Biology of War. The Century Company, New York, NY, USA, 1918. xxx + 553 pp. LCCN U21 .N5 FT MEADE. Translated by Constance A. Grande and Julian Grande from the 1917 Swiss German edition [Nic17, Nic19]. The introduction contains the Einstein manifesto [ENFB14].

Nicolai:1919:BKB

[Nic19] Georg Friedrich Nicolai. Die Biologie des Krieges: Betrachtungen eines deutschen Naturforschers. (German) [The Biology of War: Reflections of a German Scientist]. Orell Füssli, Zürich, Switzerland, second edition, 1919. 552 pp. LCCN U21 .N4 1919. The introduction contains the Einstein manifesto [ENFB14].

Nichols:1934:YOC

[Nic34] Joseph B. Nichols. You have one chance in a hundred to understand Einstein. Scientific American, 150 (2):72–73, February 1934. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v150/n2/
pdf/scientificamerican0234-72.pdf.

Nichols:1987:RTP

[Nic87] Major General Kenneth D. (Kenneth David) Nichols. The road to Trinity: a personal account of how America’s nuclear policies were made. Morrow, New York, NY, USA, 1987. ISBN 0-688-06910-X. 401 pp. LCCN QC774.N45 A3 1987.

Nichols:2020:BRD

[Nic20] Tiffany Nichols. Book review: Daniel Kennefick. No Shadow of a Doubt: The 1919 Eclipse That Confirmed Einstein’s Theory of Relativity. Isis, 111(2):417–418, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nier:1989:SRM

[Nie89] Alfred O. Nier. Some reminiscences of mass spectrometry and the Manhattan Project. Journal of Chemical Education, 66(5): 385–388, 1989. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Nikbin:2006:NEM

[Nik06] Darius Nikbin. Newton, Einstein and the monstrous calf. Physics World, 19(10):8, October 2006. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
19/10/phwv19i10a16.pdf.

Nikolic:2012:EBE

[Nik12] Hrvoje Nikolić. EPR before EPR: a 1930 Einstein–Bohr thought experiment revisited. European Journal of Physics, 33(5):1089, 2012. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404

(electronic). URL http://stacks.iop.org/0143-0807/33/i=
5/a=1089.

Nirenstein:1925:EBA

[Nir25] Mauricio Nirenstein. Einstein en Buenos Aires. (Spanish) [Einstein in Buenos Aires]. Verbvm: Revista del Centro de Estudiantes de Filosofía y Letras de la Universidad de Buenos Aires, 18(??):167–178, ???? 1925.

Nisio:1979:TEW

[Nis79] Sigeko Nisio. The transmission of Einstein’s work to Japan. Japanese Studies in the History of Science, 18(??):1–8, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Newton-John:2005:AEM

[NJ05] I. B. Newton-John. Albert Einstein to Max Born. Physics Today, 58(6):16, June 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Newman:1992:RAG

[NJP92] Ted Newman, Allen I. (Allen Ira) Janis, and John R. (John Robert) Porter, editors. Recent advances in General Relativity: essays in honor of Ted Newman, volume 4 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1992. ISBN 0-8176-3541-6, 3-7643-3541-6. xi + 266 pp. LCCN QC173.6 .R43 1991. URL http://www.

zentralblatt-math.org/zmath/en/search/?an=0840.53076.

Papers from the Discussion Conference on Recent Advances in General Relativity, held at the University of Pittsburgh, May 3–5, 1990.

Nightingale:1984:NEF

[NK84] J. D. Nightingale and R. E. Kelly. Note on Einstein’s first paper. American Journal of Physics, 52(6):560–561, 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/52/6/10.1119/1.13621.

Nakao:2015:YNR

[NKY15] Maika Nakao, Takeshi Kurihara, and Masakatsu Yamazaki. Yasushi Nishiwaki, radiation biophysics, and peril and hope in the nuclear age. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):8–35, August 2015. CODEN HISCDU. ISSN 0285-4821.

Niaz:2012:RWP

[NM12] Mansoor Niaz and Cecilia Marcano. Reconstruction of Waveparticle Duality and Its Implications for General Chemistry Textbooks. Springer briefs in education. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4395-5 (paperback). viii + 46 pp. LCCN

QC476.W38 N53 2012. URL http://www.springerlink.com/
content/978-94-007-4395-3.

Neumann:1981:UE

[NMA^{+}81] Holger Neumann, Olaf Melsheimer, Wolfgang Adamczak, Reinhard Eckhorn, Jürgen Altmann, Wolfgang Bayer, Ralf Beckmann, Norbert Stelte, Reinhard Brandt, Ludwig Schweitzer, Gustav Sauer, Peter Thomas, et al. Unsöld on Einstein. Nature, 293(5832):422, October 8, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v293/n5832/pdf/293422a0.pdf.

Nathan:1960:EPa

[NN60a] Otto Nathan and Heinz Norden, editors. Einstein on peace. Methuen and Co. Ltd., London, UK, 1960. 704 pp. LCCN JX1952 .E44. Preface by Bertrand Russell.

Nathan:1960:EPb

[NN60b] Otto Nathan and Heinz Norden, editors. Einstein on peace. Simon and Schuster, New York, NY, USA, 1960. xvi + 704 pp. LCCN JX1952 .E44. Preface by Bertrand Russell.

Nathan:1975:EP

[NN75a] Otto Nathan and Heinz Norden. Einstein on Peace. Schocken Books, New York, NY, USA, 1975. ISBN 0-8052-0191-2. 704 pp.

Nathan:1975:FWW

[NN75b] Otto Nathan and Heinz Norden. Über den Frieden: Weltordnung oder Weltuntergang? (German) [On Freedom: World Order or World Collapse?]. Land und Cie AG, Bern, Switzerland, 1975. ISBN 3-261-01384-2. 675 pp. LCCN JX1952 .E44 1975. Forward by Bertrand Russell. German translation of [NN60a].

Nathan:1981:EPa

[NN81] Otto Nathan and Heinz Norden, editors. Einstein on peace. Avenel Books, New York, NY, USA, 1981 edition, 1981. ISBN 0-517-34580-3. xiv + 704 pp. LCCN JX1952 .E44 1981.

Nielsen:2001:NTN

[NN01] Keld Nielsen and Henry Nielsen, editors. Nabo til Nobel: historien om tretten danske Nobelpriser. (Danish) [Neighbor of Nobel: the story of thirteen Danish Nobel Prize winners]. Aarhus Universitetsforlag, Århus, Danmark, 2001. ISBN 87-7288-900-4. 570 pp. LCCN AS911.N9 N33 2001. DKR 398,00. URL

http://da.unipress.dk/udgivelser/n/nabo-til-nobel/.

Nobelstiftelsen:1964:CNL

[Nob64a] Nobelstiftelsen. Chemistry: Nobel lectures in chemistry. Nobel lectures, including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1964. ????

pp. LCCN QD39 .N735. URL http://nobelprize.org/nobel_
prizes/chemistry/video_lectures.html. Published for the

Nobel Foundation.

Nobelstiftelsen:1964:NLC

[Nob64b] Nobelstiftelsen. Nobel lectures, chemistry 1942–1962: including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1964. 710 pp. LCCN ????

Nobelstiftelsen:1966:NLCa

[Nob66a] Nobelstiftelsen. Nobel lectures, chemistry 1901–1921: including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1966. 409 pp. LCCN ????

Nobelstiftelsen:1966:NLCb

[Nob66b] Nobelstiftelsen. Nobel lectures, chemistry 1922–1941: including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1966. 506 pp. LCCN ????

Nobelstiftelsen:1967:P

[Nob67] Nobelstiftelsen. Physics. Nobel lectures, including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1967. ???? pp. LCCN QC71 .N64. Published for the Nobel Foundation.

Nobel:1972:NLI

[Nob72] Nobel Foundation, editor. Nobel lectures, including presentation speeches and Laureates’ biographies, 1963–1970. Elsevier, Amsterdam, The Netherlands, 1972. ISBN 0-444-40993-9. xi + 349 pp. LCCN QC71.N735.

Nobile:1995:JS

[Nob95] Philip Nobile, editor. Judgment at the Smithsonian. Marlowe and Co., New York, NY, USA, 1995. ISBN 1-56924-841-9. xcvii + 270 pp. LCCN D767.25.H6 J83 1995. Smithsonian script by the curators at the National Air and Space Museum. Afterword by Barton J. Bernstein.

Nobelstiftelsen:1964:PM

[Nob03] Nobelstiftelsen. Physiology or medicine. Nobel lectures, including presentation speeches and laureates’ biographies. Elsevier, Amsterdam, The Netherlands, 1964–2003. ISBN 981-238-005-1 (vol. 8). ???? pp. LCCN QH311 .N74p 1964-. Published for the Nobel Foundation.

Noddack:1934:EGE

[Nod34] Ida Noddack. Über das Element 93. (German) [On element 93 [neptunium]]. Zeitschrift für Angewandte Chemie, 47(37):653–655, September 15, 1934. ISSN 0932-2132.

URL http://en.wikipedia.org/wiki/Neptunium;   http:/
/onlinelibrary.wiley.com/doi/10.1002/ange.19340473707/
abstract;   http://www.chemteam.info/Chem-History/Noddack-

1934.html. According to Frisch [FW67, page 48], “Ida Noddack, a German chemist, quite rightly pointed out that they might be lighter elements [after bombardment of uranium by neutrons]; but her comments (published in a journal not much read by chemists and hardly at all by physicists) were regarded as mere pedantry. She did not indicate how such light elements could be formed; her paper had probably no effect whatever on later work.”. English translation in [GA71, pages 16–20].

Noer:1993:BRB

[Noe93] Richard J. Noer. Book review: Albert Einstein and Mileva Marić: The Love Letters, by Jürgen Renn and Robert Schulmann. American Journal of Physics, 61(5):479, May 1993. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/61/479/1.

Noer:2011:BRJ

[Noe11] Richard Noer. Book review: Jennifer Coopersmith, Energy, the Subtle Concept: The Discovery of Feynman’s Blocks from Leibniz to Einstein. Oxford: Oxford University Press, 2010, xiv + 400 pages. $55.00 (cloth). Physics in Perspective (PIP), 13(3): 379–380, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Nola:1988:RRS

[Nol88] Robert Nola, editor. Relativism and realism in science, volume 6 of Australasian studies in history and philosophy of science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2647-7. ix + 299 pp. LCCN Q175 .R385 1988.

Noll:2005:EIM

[Nol05] Roger Noll. Einstein’s interoffice memo? Science, 309(5740): 1490–1491, September 2, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/309/5740/1490.full.pdf.

Nordmann:1921:EUL

[Nor21] Charles Nordmann. Einstein et l’Univers: une lueur dans le mystére des choses. (French) [Einstein and the Universe: a glow in the mystery of things]. Hachette, Paris, France, 1921. 221 pp.

Nordmann:1922:EWG

[Nor22a] Charles Nordmann. Einstein und das Weltall. (German) [Einstein and the Universe]. Hoffmann, Stuttgart, Germany, 1922. 199 pp.

Nordmann:1922:EBF

[Nor22b] Charles Nordmann. With Einstein on the battle fields. Literary Digest, 73(??):586–592, June 3, 1922.

Northrop:1949:ECS

[Nor49] Filmer S. C. Northrop. Einstein’s conception of science. In Schilpp [Sch49c], chapter 14, pages 385–408. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

North:1965:MUH

[Nor65] John David North. The measure of the universe: a history of modern cosmology. Clarendon Press, Oxford, UK, 1965. xxviii + 436 pp. LCCN QB981 .N77.

Nordtvedt:1972:GTE

[Nor72] Kenneth L. Nordtvedt, Jr. Gravitation theory: Empirical status from solar system experiments: All observations to date are consistent with Einstein’s General Relativity theory of gravity. Science, 178(4066):1157–1164, December 15, 1972. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/178/4066/1157.
full.pdf.

Norton:1984:HEF

[Nor84] John Norton. How Einstein found his field equations: 1912– 1915. Historical Studies in the Physical Sciences, 14(2):253–316, ???? 1984. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757535. Reprinted in [HS89b, 101–

159].

Norton:1985:WWE

[Nor85] John Norton. What was Einstein’s Principle of Equivalence? Studies in History and Philosophy of Science Part A, 16(3):203–246, September 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900020.

Norton:1989:CCE

[Nor89] John Norton. Coordinates and covariance: Einstein’s view of space–time and the modern view. Foundations of Physics, 19(10):1215–1263, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731880.

North:1990:MUH

[Nor90] John David North. The measure of the universe: a history of modern cosmology. Dover, New York, NY, USA, 1990. ISBN 0-486-66517-8 (paperback). xxviii + 436 pp. LCCN QB981 .N77 1990. US$12.95.

Norton:1992:ENE

[Nor92] John D. Norton. Einstein, Nordström and the early demise of scalar, Lorentz-covariant theories of gravitation. Archive for History of Exact Sciences, 45(1):17–94, March 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=45&issue=1&spage=17.

Norton:1993:GCF

[Nor93] John D. Norton. General covariance and the foundations of general relativity: eight decades of dispute. Reports on Progress in Physics, 56(7):791–858, July 1993. CO-DEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (elec-

tronic). URL http://iopscience.iop.org/0034-4885/56/7/
001;   http://stacks.iop.org/0034-4885/56/i=7/a=001.

North:1994:FHA

[Nor94] John David North. The Fontana history of astronomy and cosmology. Fontana history of science. Fontana Press, London, UK, 1994. ISBN 0-00-686177-6 (paperback). xxvii + 697 + 8 pp. LCCN ????

Norton:2000:NRS

[Nor00] John D. Norton. ‘Nature is the Realisation of the Simplest Conceivable Mathematical Ideas’: Einstein and the canon of mathematical simplicity. Studies in History and Philosophy of Modern Physics, 31(2):135–170, June 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000350.

North:2001:EM

[Nor01] Joseph Roy D. North. Einstein’s motivation. Science, 291(5505): 828, February 2, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/291/5505/828.2.full.

Norrby:2002:CNP

[Nor02a] Erling Norrby. A century of Nobel Prizes. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 146(4):323–336, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558309.

Norris:2002:RBG

[Nor02b] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Steerforth Press, South Royalton, VT, USA, 2002. ISBN 1-58642-039-9. xxi + 722 pp. LCCN UG128.G76 N67 2002. URL

http://www.loc.gov/catdir/toc/fy036/2001057629.html.

Norris:2003:RBG

[Nor03a] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Turnaround, London, UK, 2003. ISBN 1-58642-067-4. 228 pp. LCCN UG128.G76 N67 2002. URL http://www.loc.

gov/catdir/toc/fy036/2001057629.html.

Norton:2003:SF

[Nor03b] John D. Norton. The N-Stein family. In Ashtekar et al. [ACH^{+}03], pages 55–68. ISBN 1-4020-1284-5. LCCN

Q174 .B67 vol. 234. URL http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-d.html;   http://www.
loc.gov/catdir/enhancements/fy0813/2003046757-t.html.

Norton:2004:EIG

[Nor04] John D. Norton. Einstein’s investigations of Galilean covariant electrodynamics prior to 1905. Archive for History of Exact Sciences, 59(1):45–105, November 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=1&spage=45.

Norton:2006:AEQ

[Nor06] John D. Norton. Atoms, entropy, quanta: Einstein’s miraculous argument of 1905. Studies in History and Philosophy of Modern Physics, 37(1):71–100, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000699.

Nordstrom:2007:IGM

[Nor07a] Gunnar Nordström. Inertial and gravitational mass in relativistic mechanics. In Renn et al. [RSS^{+}07a], pages 499–521. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v.

3; QC173.6 .G469 2007. URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-9_29.

Nordstrom:2007:TGS

[Nor07b] Gunnar Nordström. On the theory of gravitation from the standpoint of the Principle of Relativity. In Renn et al. [RSS^{+}07a], pages 523–542. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_30.

Nordstrom:2007:PRG

[Nor07c] Gunnar Nordström. The Principle of Relativity and gravitation. In Renn et al. [RSS^{+}07a], pages 489–497. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/
978-1-4020-4000-9_28.

Norton:2007:ENE

[Nor07d] John D. Norton. Einstein, Nordström, and the early demise of scalar, Lorentz covariant theories of gravitation. In Renn et al. [RSS^{+}07a], pages 413–487. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_27.

Norton:2007:WWE

[Nor07e] John D. Norton. What was Einstein’s “fateful prejudice”? In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 715–783. ISBN 1-4020-3942-5. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007. URL https://link.springer.

com/chapter/10.1007/978-1-4020-4000-9_7.

Norrby:2010:NPL

[Nor10] Erling Norrby. Nobel Prizes and life sciences. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010. ISBN 981-4299-36-7 (hardcover), 981-4299-37-5 (paperback). xv + 317 pp. LCCN QH315 .N667 2010.

Nordlinger:2012:PTS

[Nor12] Jay Nordlinger. Peace, they say: a history of the Nobel Peace Prize, the most famous and controversial prize in the world. Encounter Books, New York, NY, USA, 2012. ISBN 1-59403-598-9 (hardcover). viii + 459 pp. LCCN JZ5537 .N67 2012.

Norrby:2013:NPN

[Nor13] Erling Norrby. Nobel Prizes and nature’s surprises. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4520-98-5 (hardcover), 981-4520-99-3 (paperback). 400 (est.) pp. LCCN QH315. URL http://

www.worldscientific.com/worldscibooks/10.1142/8881.

Norton:2014:EST

[Nor14] John D. Norton. Einstein’s Special Theory of Relativity and the problems in the electrodynamics of moving bodies that led him to it. In Janssen and Lehner [JL14b], chapter 2, pages 72–102. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Norden:2015:NLC

[Nor15] Bengt Norden. Nobel Lectures in Chemistry 2006–2010. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4630-16-0 (hardcover), 981-4630-17-9 (paperback). x + 433 pp. LCCN QD39 .N67 2014.

Norton:2016:HED

[Nor16] John D. Norton. How Einstein did not discover. Physics in Perspective (PIP), 18(3):249–282, September 2016. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
016-0186-z.

Novak:1992:AEH

[Nov92] L. Novak. Albert Einstein halala. (Hungarian) [The death of Albert Einstein]. Orvosi hetilap, 133(52):3318–3319, December 27, 1992. ISSN 0030-6002 (print), 1788-6120 (electronic).

Novikov:1998:RT

[Nov98] I. D. (Igor Dmitrievich) Novikov. The River of Time. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-46177-4, 0-521-46737-3 (paperback). xxii + 275 pp.

LCCN QB209 .N68 1998. URL http://www.loc.gov/catdir/
description/cam028/97043010.html;   http://www.loc.
gov/catdir/toc/cam021/97043010.html;   http://www.zentralblattmath.
org/zmath/en/search/?an=0997.83500.

Novoselov:2011:NLG

[Nov11] K. S. Novoselov. Nobel Lecture: Graphene: Materials in the Flatland. Reviews of Modern Physics, 83(3):837–849, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.837;   http://rmp.aps.org/
abstract/RMP/v83/i3/p837_1.

Novosyadlyj:2018:C

[Nov18] Bohdan Novosyadlyj. Century of Λ. European Physical Journal H, 43(3):267–280, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2018-90007-y.

Nielsen:1988:NEM

[NP88] Birger Nielsen and Hendrik Pedersen. A note on Einstein metrics — a simple application of a symbolic algebra system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):7–11, January 1988. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nachmansohn:1988:GAW

[NS88] David Nachmansohn and Roswitha Schmid. Die grosse Aera der Wissenschaft in Deutschland 1900 bis 1930: jüdische und nichtjüdische Pioniere in der Atomphysik, Chemie und Biochemie. (German) [The great era of science in Germany 1900 to 1930: Jewish and non-Jewish pioneers in atomic physics, chemistry and biochemistry]. Wissenschaftliche Verlagsgesellschaft, Stuttgart, Germany, 1988. ISBN 3-8047-0662-2. 398 pp. LCCN ????

Naumann:1996:EAC

[NS96] Robert Naumann and Henry Stroke. Einstein and the atomic clock. Physics World, 9(4):76, April 1996. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
9/4/phwv9i4a34.pdf.

Natarajan:2005:STR

[NS05] Vasant Natarajan and Diptiman Sen. The Special Theory of Relativity. Resonance, 10(4):32–50, April 2005. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X

(electronic). URL http://www.springerlink.com/content/
r352522639232765/.

Nugayev:1987:GSM

[Nug88] R. M. Nugayev. The genesis and structure of models in the modern theory of gravity. International Studies in the Philosophy of Science, 2(1):84–104, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.

tandfonline.com//doi/abs/10.1080/02698598708573304.

Nugayev:1999:ERC

[Nug99] Rinat M. Nugayev. Einstein’s revolution: a case study in communicative rationality. Foundations of Science, 4(2):155–204, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL http://link.springer.com/article/
10.1023/A:1009619128932.

Nugayev:2000:ERR

[Nug00] Rinat M. Nugayev. Einstein’s revolution: reconciliation of mechanics, electrodynamics and thermodynamics. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 37(1):181– 207, 2000. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Numbers:2009:GGJ

[Num09] Ronald L. Numbers, editor. Galileo goes to jail: and other myths about science and religion. Harvard University Press, Cambridge, MA, USA, 2009. ISBN 0-674-03327-2, 0-674-05439-3 (e-book). x + 302 pp. LCCN Q126.8 .G35 2009.

Nunes:2022:SSS

[Nun22] Rafael Nunes. Search for sub-solar mass binaries with Einstein Telescope and Cosmic Explorer. Entropy, 24(2):262, ???? 2022. CODEN ENTRFG. ISSN 1099-4300. URL https://dblp.org/

db/journals/entropy/entropy24.html#Nunes22.

Nussbaumer:2014:ECH

[Nus14] Harry Nussbaumer. Einstein’s conversion from his static to an expanding universe. European Physical Journal H, 39(1): 37–62, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2013-40037-6.

Nuvolari:2011:RBP

[Nuv11] Alessandro Nuvolari. Reviews: Pursuing Power and Light. Technology and Physics from James Watt and Albert Einstein — by Bruce J. Hunt. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):251–252, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Newton:1949:LSE

[NW49] T. D. Newton and E. P. Wigner. Localized states for elementary systems. Reviews of Modern Physics, 21(3):400–406, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.400;   http://rmp.aps.org/

abstract/RMP/v21/i3/p400_1. Special issue in honor of Albert Einstein on his 70th birthday.

Nye:1972:MRP

[Nye72] Mary Jo Nye. Molecular reality: a perspective on the scientific work of Jean Perrin. History of science library. Macdonald, London, 1972. ISBN 0-444-19596-3 (American Elsevier). xi + 201 pp. LCCN QC16.P4 N9.

Nye:1988:BRBb

[Nye88] Mary Jo Nye. Book review: The Nobel Population, 1901–1937: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry by Elisabeth Crawford; J. L. Heilbron; Rebecca Ullrich. Isis, 79(4):684, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234760.

Nye:1999:PCP

[Nye99] M. J. Nye. A physicist in the corridors of power: P. M. S. Blackett’s opposition to atomic weapons following the War. Physics in Perspective (PIP), 1(2):136–156, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050013.

Nye:2004:BPW

[Nye04] Mary Jo Nye. Blackett: physics, war, and politics in the Twentieth Century. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01548-7 (hardcover). x + 255 pp. LCCN QC16.B59 N94 2004.

Nye:2019:STM

[Nye19] Mary Jo Nye. Shifting trends in modern physics, Nobel recognition, and the histories that we write. Physics in Perspective (PIP), 21(1):3–22, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Oppenheimer:19xx:JRO

[OBxx] J. Robert Oppenheimer and Niels Bohr. J. Robert Oppenheimer papers (1799–1980) (bulk 1947–1967) and Niels Henrik David Bohr papers. ????, ????, 19xx. ???? pp. LCCN 0456G; Vault 0202A; Microfilm 16,646-1P. URL http://hdl.loc.gov/loc.

mss/eadmss.ms998007.

Oliveira:2005:HDD

[OB05] João Gama Oliveira and Albert-László Barabási. Human dynamics: Darwin and Einstein correspondence patterns. Nature, 437(7063):1251, October 26, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v437/n7063/
full/4371251a.html.

Oberheim:2016:REL

[Obe16] Eric Oberheim. Rediscovering Einstein’s legacy: How Einstein anticipates Kuhn and Feyerabend on the nature of science. Studies in History and Philosophy of Science Part A, 57(??):17– 26, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S003936811500165X.

Nobelstiftelsen:1972:NMH

[Ode72a] Wilhelm Odelberg, editor. Nobel, the man and his prizes. American Elsevier Pub. Co., New York, third edition, 1972. ISBN 0-444-00117-4. x + 659 pp. LCCN AS911 .N7553 1972.

Odelberg:1972:NMH

[Ode72b] Wilhelm Odelberg, editor. Nobel, the man and his prizes. American Elsevier Pub. Co., New York, NY, USA, third edition, 1972. ISBN 0-444-00117-4. x + 659 pp. LCCN AS911 .N7553 1972. Individual sections written by H. Schück and others.

Oe:1981:HN

[Ōe81] Kenzaburō Ōe. Hiroshima notes. YMCA Press, Tokyo, Japan, 1981. 181 pp. LCCN D767.25.H6 O3513 1982. Y1200.

Oe:1995:DES

[Ōe95a] Kenzaburō Ōe. The day the Emperor spoke in a human voice. The New York Times Magazine, ??(??):103–105, May 7, 1995. ISSN 0362-1308.

Oe:1995:HN

[Ōe95b] Kenzaburō Ōe. Hiroshima notes. Marion Boyars, New York, NY, USA, 1995. ISBN 0-7145-3007-7 (hardcover). 192 pp. LCCN D767.25.H6 O3513 1995.

Oe:1996:HN

[Ōe96] Kenzaburō Ōe. Hiroshima notes. Grove Press, New York, NY, USA, 1996. ISBN 0-8021-3464-5. 192 pp. LCCN

D767.25.H6 O3513 1996. URL http://www.loc.gov/catdir/
enhancements/fy1311/96005812-d.html.

Ogawa:1979:JEE

[Oga79] Tsuyoshi Ogawa. Japanese evidence for Einstein’s knowledge of the Michelson–Morley experiment. Japanese Studies in the History of Science, 18(??):73–81, ???? 1979. CODEN JSHIAE. ISSN 0090-0176. This article contains a partial translation of Einstein’s lecture How I created the theory of relativity at Kyoto University on 14 December 1922. A full translation is available in [Ein82b].

Ohanian:1976:GS

[Oha76] Hans C. Ohanian. Gravitation and spacetime. W. W. Norton & Co., New York, NY, USA, 1976. ISBN 0-393-09198-8. xiv + 461 pp. LCCN QC178 .O35.

Ohanian:1977:WPE

[Oha77] Hans C. Ohanian. What is the principle of equivalence? American Journal of Physics, 45(10):903–909, October 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Ohanian:1979:BRB

[Oha79] Hans Ohanian. Book review: Albert Einstein’s Theory of General Relativity: 60 Years of its Influence on Man and the Universe, edited by Gerald Tauber. American Journal of Physics, 47(12):1111, December 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/47/1111/1.

Ohanian:2008:EMH

[Oha08] Hans C. Ohanian. Einstein’s Mistakes: the Human Failings of Genius. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-06293-7. xix + 394 pp. LCCN QC7 .O33 2008.

Ohanian:2009:DEP

[Oha09] Hans C. Ohanian. Did Einstein prove E = mc^{2}? Studies in History and Philosophy of Modern Physics, 40(2):167–173, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000112.

Ohanian:2012:CMS

[Oha12a] Hans C. Ohanian. A comment on Mermin’s “Understanding Einstein’s 1905 derivation of E = mc^{2}”. Studies in History and Philosophy of Modern Physics, 43(3):215–217, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000202. See [Mer11] and reply

[Mer12].

Ohanian:2012:CMU

[Oha12b] Hans C. Ohanian. A comment on Mermin’s “Understanding Einstein’s 1905 derivation of E = mc^{2}”. Studies in History and Philosophy of Modern Physics, 43(3):215–217, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000202. See [Mer11] and reply

[Mer12].

Ohanian:2019:CEB

[Oha19] Hans C. Ohanian. Comedy of errors boosted 1920s Einstein mania. Physics Today, 72(11):10–11, November 2019. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/
full/10.1063/PT.3.4332.

Oliphant:1947:SSA

[OJCP^{+}47] M. L. Oliphant, F. Joliot-Curie, R. E. Peierls, Albert Einstein, Federation of American Scientists, and General Leslie Groves. Statements on the second anniversary of Hiroshima. Bulletin of the Atomic Scientists, 3(9):235–236, 252, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Okamoto:1981:AEJ

[OK81] Ippei Okamoto and Kenkichiro Koizumi. Albert Einstein in Japan — 1922. American Journal of Physics, 49(10):930–940, October 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/?AJP/49/
930/1.

Ozeri:2005:CBB

[OKSD05] R. Ozeri, N. Katz, J. Steinhauer, and N. Davidson. Colloquium: Bulk Bogoliubov excitations in a Bose–Einstein condensate. Reviews of Modern Physics, 77(1):187–205, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.187;   http://rmp.aps.org/
abstract/RMP/v77/i1/p187_1.

Okun:2005:MEM

[Oku05] Lev Okun. Mass, energy and the meaning of relativity. Physics World, 18(10):20, October 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/10/phwv18i10a29.pdf.

OLeary:1964:RDP

[O’L64] Austin J. O’Leary. Redshift and deflection of photons by gravitation: a comparison of relativistic and Newtonian treatments. American Journal of Physics, 32(1):52–55, January 1964. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Oliphant:1972:RRC

[Oli72] Sir Mark Oliphant. Rutherford: recollections of the Cambridge days. Elsevier, Amsterdam, The Netherlands, 1972. ISBN 0-444-40968-8. xii + 162 pp. LCCN QC16.R8 O54.

Olivi:1999:DNM

[Oli99] Fred J. Olivi. Decision at Nagasaki: the mission that almost failed. ????, ????, 1999. ISBN ???? ???? pp. LCCN ????

Olwell:1997:AWA

[Olw97] Russell Brian Olwell. Atomic workers, atomic city: labor and community in Oak Ridge, Tennessee, 1942–1950. Thesis (Ph.D.), Massachusetts Institute of Technology, Program in Science, Technology, and Society, Cambridge, MA, USA, 1997. 83 pp.

Olwell:2004:WAC

[Olw04] Russell B. Olwell. At work in the atomic city: a labor and social history of Oak Ridge, Tennessee. University of Tennessee Press, Knoxville, TN, USA, 2004. ISBN 1-57233-324-3 (hardcover). ix + 165 pp. LCCN HD8039.A62 U434 2004. URL http://www.

loc.gov/catdir/toc/ecip0416/2004005732.html.

ORaifeartaigh:2014:ECM

[OM14] C. O’Raifeartaigh and B. McCann. Einstein’s cosmic model of 1931 revisited: an analysis and translation of a forgotten model of the universe. European Physical Journal H, 39(1): 63–85, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2013-40038-x. Contains En-

glish translation on pages 82–85 of [Ein31o]. See correction [OM22].

ORaifeartaigh:2018:ILE

[OM18] Cormac O’Raifeartaigh and Simon Mitton. Interrogating the legend of Einstein’s ”biggest blunder”. Physics in Perspective (PIP), 20(4):318–341, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

ORaifeartaigh:2022:CEC

[OM22] C. O’Raifeartaigh and B. McCann. Correction to: Einstein’s cosmic model of 1931 revisited: an analysis and translation of a forgotten model of the universe. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-022-00040-1.

See [OM14].

ORaifeartaigh:2014:ESS

[OMNM14] Cormac O’Raifeartaigh, Brendan McCann, Werner Nahm, and Simon Mitton. Einstein’s steady-state theory: an abandoned model of the cosmos. European Physical Journal H, 39(3):353– 367, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2014-50011-x.

Oberhansly:2001:DAT

[OO01] Curtis Oberhansly and Dianne Nelson Oberhansly. Downwinders: an Atomic Tale. Black Ledge Press, Salt Lake City, UT, USA, 2001. ISBN 0-9707965-9-5. 425 pp. LCCN PS3615.B47 D69 2001.

ORaifeartaigh:2021:HPR

[OOM21] Cormac O’Raifeartaigh, Michael O’Keeffe, and Simon Mitton. Historical and philosophical reflections on the Einstein–de Sitter model. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL https://link.springer.com/article/10.
1140/epjh/s13129-021-00007-8.

ORaifeartaigh:2015:ECR

[OONM15] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. Einstein’s cosmology review of 1933: a new perspective on the Einstein–de Sitter model of the cosmos. European Physical Journal H, 40(3):301–335, September 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-50061-y.

ORaifeartaigh:2017:ESM

[OONM17] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. Einstein’s 1917 static model of the universe: a centennial review. European Physical Journal H, 42(3):431– 474, August 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

ORaifeartaigh:2018:OHY

[OONM18] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. One hundred years of the cosmological constant: from ‘superfluous stunt” to dark energy. European Physical Journal H, 43(1):73–117, April 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2017-80061-7.

OHara:1987:HMS

[OP87] J. G. (James G.) O’Hara and W. (Willibald) Pricha. Hertz and the Maxwellians: a study and documentation of the discovery of electromagnetic wave radiation, 1873–1894, volume 8 of IEE History of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1987. ISBN 0-86341-101-0. xiv + 154 pp. LCCN QC670 .O47 1987.

Opfell:1978:LLW

[Opf78] Olga S. Opfell. The Lady Laureates: Women Who Have Won the Nobel Prize. Scarecrow Press, Metuchen, NJ, USA, 1978. ISBN 0-8108-1161-8. xvii + 267 pp. LCCN AS911.N9 O63.

Oppenheimer:1939:CSB

[Opp39] J. R. Oppenheimer. Celebration of the sixtieth birthday of Albert Einstein. Science, 89(2311):335–336, April 14, 1939. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/89/2311/335.2.
extract.

Oppenheimer:1955:OM

[Opp55] J. Robert Oppenheimer. The open mind. Simon and Schuster, New York, NY, USA, 1955. 146 pp. LCCN QC780.

Oppenheimer:1956:E

[Opp56] J. Robert Oppenheimer. Einstein. Reviews of Modern Physics, 28(1):1–2, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.28.1;   http://
rmp.aps.org/abstract/RMP/v28/i1/p1_1. In memoriam trib-

ute.

Oppenheimer:1963:OM

[Opp63] J. Robert Oppenheimer. The open mind. Simon and Schuster, New York, NY, USA, 1963. viii + 152 pp. LCCN QC780.

Oppenheimer:1965:AE

[Opp65] J. R. Oppenheimer. On Albert Einstein. UNESCO, Paris. Published in [Opp66]., 1965.

Oppenheimer:1966:AE

[Opp66] J. Robert Oppenheimer. On Albert Einstein. New York Review of Books, ??(??):4–5, March 17, 1966. CODEN ???? ISSN 0028-7504 (print), 1944-7744 (electronic).

Oppenheim:1979:CNP

[Opp79a] Charles Oppenheim. Could the 1978 Nobel Prizewinner in Chemistry have been predicted? Social Studies of Science, 9 (4):507–508, November 1, 1979. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631277900900407.

Oppenheimer:1979:AE

[Opp79b] J. R. Oppenheimer. On Albert Einstein. In French [Fre79c], page ?? ISBN 0-674-24230-0. LCCN QC16.E5 E37.

Oppenheimer:1979:CSB

[Opp79c] J. Robert Oppenheimer. Celebration of the sixtieth birthday of Albert Einstein. Bulletin of the Atomic Scientists, 35(3):39, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Oppenheimer:1979:OE

[Opp79d] J. Robert Oppenheimer. Oppenheimer on Einstein. Bulletin of the Atomic Scientists, 35(3):36–38, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Ohanian:1994:GS

[OR94] Hans C. Ohanian and Remo Ruffini. Gravitation and spacetime. W. W. Norton & Co., New York, NY, USA, second edition, 1994. ISBN 0-393-96501-5. xv + 679 pp. LCCN QC178

.O35 1994. URL http://www.zentralblattmath.org/zmath/
en/search/?an=0846.53052.

ORaifeartaigh:1997:DGT

[O’R97] L. (Lochlainn) O’Raifeartaigh. The dawning of gauge theory. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-02978-4, 0-691-02977-6 (paperback). ix + 249 pp. LCCN QC174.45

.O7 1997. URL http://www.loc.gov/catdir/description/
prin021/96043337.html;   http://www.loc.gov/catdir/
toc/prin031/96043337.html.

Orrman-Rossiter:2021:PID

[OR21] Kevin Orrman-Rossiter. Places of ‘invention and discovery’ and the Nobel Prize in Physics. Notes and Records of the Royal Society of London, 75(3):439–460, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0014.

Orr:1978:PNP

[Orr78] John Boyd Orr. Perspective: Nobel Peace Laureates: 1949. Bulletin of the Atomic Scientists, 34(9):1, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 12 December 1949, Oslo, Norway.

Ortiz:1995:CIE

[Ort95] Eduardo L. Ortiz. A convergence of interests: Einstein’s visit to Argentina in 1925. Ibero-Americanisches Archiv, 20(??):67–126, ???? 1995.

Orzel:2012:HTR

[Orz12] Chad Orzel. How to teach relativity to your dog. Basic Books, New York, NY, USA, 2012. ISBN 0-465-02331-2 (paperback), 0-465-02937-X (e-book). vi + 327 pp. LCCN QC173.57 .O79 2012.

Oppenheimer:1939:CGC

[OS39] J. R. Oppenheimer and H. Snyder. On continued gravitational contraction. Physical Review (2), 56(5):455–459, September 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p455_1.

OSullivan:2001:HDN

[O’S01] Abigail O’Sullivan. Henry Dale’s Nobel Prize winning ‘discovery’. Minerva [Washington, DC], 39(4):409–424, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012787903437.

Oshikawa:2006:BEC

[Osh06] Masaki Oshikawa. Bose–Einstein condensates: Fluctuating fringes. Nature Physics, 2(10):663–664, October 2006. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic).

URL http://www.nature.com/nphys/journal/v2/n10/full/
nphys430.html.

Ostermann:2015:UEE

[Ost15] Peter Ostermann. Unterwegs mit Einstein und dem Esel Zur Entfesselung der Urknall-Kosmologie. (German) [Traveling with Einstein and the donkey: liberation from the shackles of bigbang cosmology]. digIT Verlag GmbH, Bruttig-Fankel, Germany, 2015. ISBN 3-941550-26-8. 632 pp. LCCN ????

Oechsner:1986:BKP

[OSWR86] H. Oechsner, A. Schlachetzki, H. Walther, and H. Rechenberg. Buchbesprechungen: Kirschner: Polarized Electrons at Surfaces/Paul, Teubner: Optoelektronische Halbleiterbauelemente/Stitch, Bass: Laser Handbook, Vol. 4/Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.: Teil II: 1930–1939. Physikalische Blätter, 42(7):252–253, July 1986. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1002/phbl.19860420726/abstract.

Obenhous:1986:EBC

[OV86] Mark Obenhous and Chrisann Verges. Einstein on the beach — the changing image of opera. Obenhous Films, Inc. (New York), January 31, 1986.

Overholt:1987:OVS

[Ove87] James Overholt, editor. These are our voices: the story of Oak Ridge, 1942–1970. Children’s Museum of Oak Ridge, Oak Ridge, TN, USA, 1987. ISBN 0-9606832-4-0. xxvi + 550 pp. LCCN F444.O3 T48 1987. US$19.95.

Ovendale:1996:BRB

[Ove96] Ritchie Ovendale. Book review: The Decision to Use the Atomic Bomb and the Architecture of an American Myth, by Gar Alperovitz [and others]. International Affairs (Royal Institute of International Affairs 1944–), 72(1):187, January 1996. CODEN ???? ISSN 0020-5850 (print), 1468-2346 (electronic). URL

http://www.jstor.org/stable/2624796.

Overbye:2000:ELS

[Ove00] Dennis Overbye. Einstein in love: a scientific romance. Viking, New York, NY, USA, 2000. ISBN 0-670-89430-3. xv + 416 pp. LCCN QC16.E5 O9 2000.

Overbye:2001:ELS

[Ove01] Dennis Overbye. Einstein in love: a scientific romance. Penguin, London, UK and New York, NY, USA, 2001. ISBN 0-670-89430-3 (hardcover), 0-14-100221-2 (paperback). xv + 416 pp. LCCN QC16.E5 O9 2001. US$15.00.

Overbye:2005:QTT

[Ove05] Dennis Overbye. Quantum trickery: Testing Einstein’s strangest theory. New York Times, ??(??):F1, F4, December 27, 2005. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/92999357/.

Overbye:2013:TFT

[Ove13] Dennis Overbye. A trip forward in time. your travel agent: Einstein. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Oliver:1954:ERB

[OW54] Donald R. Oliver and Stacey G. Ward. The Einstein relation between relative viscosity and volume concentration of suspensions of spheres. Nature, 173(4414):1090, June 5, 1954. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v173/n4414/
pdf/1731090a0.pdf.

Owens:2012:TRE

[Owe12] Trevor Owens. Tripadvisor rates Einstein: using the social web to unpack the public meanings of a cultural heritage site. International Journal of Web Based Communities, 8(1):40–56, ???? 2012. CODEN ???? ISSN 1477-8394 (print), 1741-8216 (electronic). URL https://dblp.

org/db/journals/ijwbc/ijwbc8.html#Owens12;   https:/
/www.wikidata.org/entity/Q56028324.

Oxlade:2003:AE

[Oxl03] Chris Oxlade. Albert Einstein. Twentieth-century history makers. Raintree Steck-Vaughn, Austin, TX, USA, 2003. ISBN 0-7398-5259-0. 112 pp. LCCN QC16.E5 O94 2003.

Oyibo:2004:GUT

[Oyi04] G. A. Oyibo. Grand unified theorem: discovery of the theory of everything and the fundamental building block of quantum theory. Nova Science Publishers, New York, NY, USA, 2004. ISBN 1-59033-835-9 (hardcover), 1-62808-163-5 (e-book). xiii + 126 pp. LCCN QC794.6.G7 O947 2004. URL http://catdir.

loc.gov/catdir/toc/ecip047/2003017395.html.

P:1930:SPE

[P.30] H. T. H. P. A simplified presentation of Einstein’s unified field equations. Nature, 125(3161):813, May 31, 1930. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v125/n3161/
pdf/125813b0.pdf.

P:1950:AOBa

[P.50] R. B. P. Analyses d’ouvrages: Les Prix Nobel en 1946. Revue d’Histoire des Sciences et de Leurs Applications, 3(2):185–186, avril–juin 1950. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23904653.

P:1990:TE

[P.90] R. P. Tilting at Einstein. Science, 250(4985):1208, November 30, 1990. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/250/4985/1208.full.pdf.

Primack:2006:VCU

[PA06] J. R. (Joel R.) Primack and Nancy Ellen Abrams. The view from the center of the universe: discovering our extraordinary place in the cosmos. Riverhead Books, New York, NY, USA, 2006. ISBN 1-59448-914-9. 386 pp. LCCN QB981 .P85 2006.

Panoutsopoulos:2021:CBA

[PA21] Grigoris Panoutsopoulos and Theodore Arabatzis. CERN’s balancing act between unity and disunity: The ”Sister Experiments” UA1 and UA2 and CERN’s first Nobel Prize. Physics in Perspective (PIP), 23(4):181–201, December 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

PWRS:1972:DML

[Pac72] Pacific War Research Society. The Day Man Lost: Hiroshima, 6 August 1945. Kodansha International, Tokyo, Japan, 1972. ISBN 0-87011-174-4. 312 + 32 pp. LCCN D767.25.H6 B85.

PWRS:1981:DML

[Pac81] Pacific War Research Society. The Day Man Lost: Hiroshima, 6 August 1945. Kodansha International, Tokyo, Japan, 1981. ISBN 0-87011-471-9. 312 + 32 pp. LCCN D767.25.H6 D39

Page:1921:RRD

[Pag21] Leigh Page. Relativity and a rotating disk. Scientific American, 124(13):247, March 26, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v124/n13/
pdf/scientificamerican03261921-247.pdf.

Pais:1979:EQT

[Pai79] A. Pais. Einstein and the quantum theory. Reviews of Modern Physics, 51(4):863–914, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/abstract/RMP/v51/p863;   http:/
/link.aps.org/doi/10.1103/RevModPhys.51.863;   http:
//prola.aps.org/abstract/RMP/v51/i4/p863_1;   http:
//prola.aps.org/pdf/RMP/v51/i4/p863_1;   http://rmp.
aps.org/abstract/RMP/v51/i4/p863_1.

Pais:1980:EPF

[Pai80] Abraham Pais. Einstein on particles, fields, and the quantum theory. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Pais:1982:SLS

[Pai82] Abraham Pais. “Subtle is the Lord”: the science and the life of Albert Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1982. ISBN 0-19-853907-X. xvi + 552 + 9 pp. LCCN QC16.E5 P26 1982.

Pais:1986:IBM

[Pai86] Abraham Pais. Inward bound: of matter and forces in the physical world. Clarendon Press, Oxford, UK, 1986. ISBN 0-19-851971-0 , 0-19-851997-4 (paperback). xiv + 666 pp.

LCCN QC7 .P27 1986. URL http://books.google.com/
books?id=jtsPAQAAMAAJ;   http://catalog.hathitrust.
org/api/volumes/oclc/12581763.html.

Pais:1988:KB

[Pai88] Abraham Pais. Knowledge and belief: The impact of Einstein’s Relativity Theory. American Scientist, 76(2):154–158, March/April 1988. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27855069.

Pais:1991:NBT

[Pai91] Abraham Pais. Niels Bohr’s times: in physics, philosophy, and polity. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773

.P35 1991. US$35.00. URL http://www.loc.gov/catdir/
enhancements/fy0635/90027248-d.html.

Pais:1994:ELH

[Pai94] Abraham Pais. Einstein lived here: essays for the layman. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853994-0. xvi + 282 pp. LCCN QC16.E5 P25 1994. US$22.42.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp.

LCCN Q141 .P29 2000.  URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/fy02/99046603.html.

Palter:1983:AES

[Pal83a] R. M. Palter. Albert Einstein Special Theory of Relativity — emergence (1905) and early interpretation (1905–11) — A. Miller. Annals of Science, 40(6):657–662, ???? 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Palter:1983:ERH

[Pal83b] Robert M. Palter. Essay review: History, philosophy, and physics in Einstein’s Special Relativity paper: Arthur Miller, Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911). Annals of Science, 40(6): 657–662, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Mil81].

PalevskyGranados:1995:TQW

[Pal95] Mary Palevsky Granados. The tough question will always remain: Did we have to use the bomb? Los Angeles Times Magazine, ??(??):10–11, 28–30, June 25, 1995.

PalevskyGranados:1998:AFC

[Pal98] Mary Palevsky Granados. Atomic fragments: conversations with seven Manhattan Project veterans fifty years after the making of the bomb. Thesis (Ph.D.), Fielding Institute, Fielding Graduate University, Santa Barbara, CA, USA, 1998. vii + 356 pp.

PalladinoBosia:1999:GAE

[Pal99] M. (Maritza) Palladino Bosia. Da Galileo ad Einstein: corso di fisica per i licei scientifici. (Italian) [From Galileo to Einstein: physics course for scientific high schools]. Petrini, Torino, Italia, 1999. ISBN 88-494-0551-0. 300 pp. LCCN ????

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp. LCCN

QC773.3.U5 P35 2000. URL http://catdir.loc.gov/catdir/
bios/ucal052/99087422.html;   http://catdir.loc.gov/
catdir/description/ucal042/99087422.html.

Panofsky:1994:PP

[Pan94] Wolfgang K. H. (Wolfgang Kurt Hermann) Panofsky. Particles and policy. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1994. ISBN 1-56396-247-0. xii + 232 pp. LCCN QC793.28 .P35 1994.

Pang:2002:ESV

[Pan02] Alex Soojung-Kim Pang. Empire and the Sun: Victorian solar eclipse expeditions. Writing science. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-3925-0, 0-8047-3926-9 (paperback). x + 196 + 12 pp.

LCCN QB541 .P36 2002. URL http://www.loc.gov/catdir/
enhancements/fy0710/2001057661-b.html;   http://www.
loc.gov/catdir/enhancements/fy0710/2001057661-d.html;
 http://www.loc.gov/catdir/toc/fy032/2001057661.html.

Panek:2004:ICE

[Pan04] Richard Panek. The invisible century: Einstein, Freud, and the search for hidden universes. Viking, New York, NY, USA, 2004. ISBN 0-670-03074-0. 258 pp. LCCN Q175 .P3325 2004.

Panek:2005:YAE

[Pan05] R. Panek. The year of Albert Einstein. Smithsonian, 36(3): 108–??, June 2005. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Paolucci:1984:HCM

[Pao84] Henry Paolucci. Hegel and the celestial mechanics of Newton and Einstein. In Hegel and the Sciences [CW84], chapter 5, pages 55–85. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-009-6233-0_5.

Papantonopoulos:2015:MET

[Pap15] E. (Eleftherios) Papantonopoulos, editor. Modifications of Einstein’s theory of gravity at large distances: review talks given in the Seventh Aegean Summer School on Beyond Einstein’s Theory of Gravity, held in Parikia on Paros Island, Greece, from 23 to 28 September 2013, volume 892 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-10069-6, 3-319-10070-X (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). LCCN QC178 .M63 2015.

Parker:1979:AQG

[Par79] Leonard Parker. Applied quantum gravity: Applications of the semiclassical theory. In Perlmutter and Scott [PS79], pages 145–166. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178 .O72

1979. URL http://link.springer.com/chapter/10.1007/
978-1-4684-3596-2_7/.

Parker:1986:EDS

[Par86] Barry R. Parker. Einstein’s dream: the search for a unified theory of the universe. Plenum Press, New York, NY, USA; London, UK, 1986. ISBN 0-306-42343-X. ix + 287 pp. LCCN QC173.7 .P36 1986. URL http://www.loc.gov/

catdir/enhancements/fy0830/86015139-d.html.

Parker:1991:ER

[Par91] Bècky Parker. Down-to-earth relativity. Physics World, 4(9):83, September 1991. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/4/9/phwv4i9a37.pdf.  Review of

Black Holes and Uncle Albert, Russell Stannard, 1991 Faber and Faber 145pp £8.99hb.

Park:1999:CTP

[Par99] Buhm Soon Park. Chemical translators: Pauling, Wheland and their strategies for teaching the theory of resonance. British Journal for the History of Science, 32(1):21–46, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.jstor.org/stable/4027968.

Parker:2000:EBR

[Par00] Barry R. Parker. Einstein’s brainchild: Relativity made relatively easy! Prometheus Books, Amherst, NY, USA, 2000. ISBN 1-57392-857-7. 280 pp. LCCN QC16.E5 P37 2000.

Parker:2003:EPS

[Par03] Barry R. Parker. Einstein: the passions of a scientist. Prometheus Books, Amherst, NY, USA, 2003. ISBN 1-59102-063-8. 297 pp. LCCN QC16.E5 P367 2003. URL http://www.

loc.gov/catdir/toc/fy037/2002036942.html.

Parker:2004:AEV

[Par04] Barry Parker. Albert Einstein’s vision: remarkable discoveries that shaped modern science. Prometheus Books, Amherst, NY, USA, 2004. ISBN 1-59102-186-3 (hardcover). 286 pp. LCCN QC16.E5 P365 2004. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip0415/2004003990.html.

Paret:2005:EFP

[Par05] Peter Paret. Einstein and Freud’s pamphlet Why War? Historically Speaking, 6(6):14–19, 2005. CODEN ???? ISSN 1941-4188 (print), 1944-6438 (electronic).

Park:2006:BRG

[Par06] David Park. Book review: Gerald Holton, Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Cambridge, Mass. and London: Harvard University Press, 2005. xiv + 229 pages. $35.00 (cloth). Physics in Perspective (PIP), 8(4): 481–483, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Parkinson:2009:SRC

[Par09] Frank Parkinson. Science and religion at the crossroads. Imprint Academic, Exeter, UK, 2009. ISBN 1-84540-151-4 (paperback). vii + 162 pp. LCCN BL240.3 .P375 2009.

Parsons:2011:MED

[Par11] Paul Parsons. 3-Minute Einstein: Digesting his life, theories, and influence in 3-minute morsels. Ivy Press, Lewes, East Sussex BN7 2NS, UK, 2011. ISBN 1-4351-3002-2. 160 pp. LCCN ????

Pash:1969:AM

[Pas69] Colonel Boris T. Pash. The Alsos Mission. Award House, New York, NY, USA, 1969. 256 pp. LCCN D810.S2 P28.

Pash:1970:AM

[Pas70] Colonel Boris T. Pash. The Alsos Mission. Universal Publishing and Distributing Corporation, New York, NY, USA, 1970. 256 pp.

Pasternak:2010:YDA

[Pas10] Judy Pasternak. Yellow dirt: an American story of a poisoned land and a people betrayed. Free Press, New York, NY, USA, 2010. ISBN 1-4165-9482-5, 1-4391-0046-2 (e-book). xiii + 317 + 16 pp. LCCN E99.N3 P378X 2010 (LC); E99.N3

P378 2010. US$26.00. URL http://www.loc.gov/catdir/
enhancements/fy1006/2010005546-b.html;   http://www.
loc.gov/catdir/enhancements/fy1006/2010005546-d.html.

Paty:1979:PDA

[Pát79a] Libor Pátý. Pamětni deska Albertu Einsteinovi v Prazi v Lesnické ulici (Czech) [Memorial plaque to Albert Einstein in Prague in Lesnické street]. Československý časopis pro fyziku, A29(??):311, ???? 1979.

Paty:1979:RAE

[Pat79b] M. Paty. Realism of Albert Einstein. Pensée, ??(204):18–37, ???? 1979. ISSN 0031-4773.

Paty:1985:ECA

[Pat85] Michel Paty. Einstein et la complémentarité au sens de Bohr: du retrait dans le tumulte aux arguments d’incomplétude. (French) [Einstein and complementarity according to Bohr: withdrawal in the tumult of the incompleteness arguments]. Revue d’Histoire des Sciences, 38(3–4):325–351, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:/

/www.jstor.org/stable/23632502;   http://www.persee.
fr/web/revues/home/prescript/article/rhs_0151-4105_
1985_num_38_3_4010.

Paty:1986:ES

[Pat86] Michel Paty. Einstein and Spinoza. In Spinoza and the Sciences [GN86], chapter 12, pages 267–302. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-009-4514-2_12.

Paty:1993:HVQ

[Pat93] M. Paty. On the hidden-variables of quantum-mechanics — Albert Einstein, David Bohm and Louis Debroglie. Pensée, ?? (292):93–116, March/April 1993. ISSN 0031-4773.

Paty:1995:NEO

[Pat95] Michel Paty. The nature of Einstein’s objections to the Copenhagen interpretation of quantum mechanics. Foundations of Physics, 25(1):183–204, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02054665.

Paty:1999:RTR

[Pat99] Michel Paty. La réception de la théorie de la relativité au Brésil et l’influence des traditions scientifiques européennes. (French) [The reception of the Theory of Relativity in Brazil and the influence of European scientific traditions]. Archives Internationales d’Histoire des Sciences, 49(143):331–368, 1999.

Paterniti:2000:DMA

[Pat00] Michael Paterniti. Driving Mr. Albert: a trip across America with Einstein’s brain. Dial Press, New York, NY, USA, 2000. ISBN 0-385-33300-5. xi + 211 pp. LCCN RB17.H365 P38 2000.

Paterniti:2001:CAE

[Pat01a] Michael Paterniti. Chauffør for Albert Einstein: gennem USA med geniets hjerne i en plasticboks. (Danish) [Driver for Albert Einstein: through the USA with the genius’ brain in a plastic box]. Høst, København, Denmark, 2001. ISBN 87-14-29721-3. 224 pp.

Paterniti:2001:DMA

[Pat01b] Michael Paterniti. Driving Mr. Albert: a trip across America with Einstein’s brain. Delta, New York, NY, USA, 2001. ISBN 0-385-33303-X (paperback). xi + 211 pp. LCCN RB17.H365 P38 2000c. US$10.95.

Pathria:2003:TR

[Pat03] R. K. Pathria. The Theory of Relativity. Dover, New York, NY, USA, 2003. ISBN 0-486-42819-2 (paperback). x + 314 pp. LCCN QC173.55 .P37 2003. URL http://www.loc.gov/

catdir/description/dover031/2003043452.html;   http:
//www.loc.gov/catdir/toc/dover031/2003043452.html.

Pauli:1921:RGT

[Pau21] Wolfgang Pauli. Relativitätstheorie. (German) [Theory of Relativity]. In Arnold Sommerfeld, editor, Encyklopädie der mathematischen Wissenschaften, pages iv + 540–775. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1921. LCCN ????

Pauli:1925:ZAE

[Pau25] Wolfgang Pauli. Über den Zusammenhang des Abschlusses der Elektronengruppen im Atom mit der Komplexstruktur der Spektren. (German) [On the relation of the completion of groups of electrons in the atom with the complex structure of spectra]. Zeitschrift für Physik, 31(1):765–783, February 1925. CODEN ZEPYAA. ISSN 0044-3328. URL

http://link.springer.com/article/10.1007/BF02980631;
 http://www.springerlink.com/content/20w1m0vr050j033r/

fulltext.pdf. This is the paper in which Pauli introduced the famous Exclusion Principle, for which he received the Nobel Prize in Physics 1945 [Pau45].

Pauli:1942:AND

[Pau42] Hertha Ernestine Pauli. Alfred Nobel, dynamite king, architect of peace. L. B. Fischer, New York, NY, USA, 1942. 5 + 3–325 pp. LCCN ????

Pauli:1945:NPP

[Pau45] Wolfgang Pauli. The Nobel Prize in Physics 1945. Nobel-

prize.org, 1945. URL http://www-groups.dcs.st-and.ac.
uk/~history/Extras/Pauli_Exclusion_Principle.html;

http://www.nobelprize.org/nobel_prizes/physics/laureates/

1945/pauli-bio.html. The Nobel Prize in Physics 1945 was awarded to Wolfgang Pauli “for the discovery of the Exclusion Principle, also called the Pauli Principle”.

Pauli:1947:AND

[Pau47a] Hertha Ernestine Pauli. Alfred Nobel, dynamite king — architect of peace. Nicholson and Watson, London, UK, 1947. 323 pp. LCCN TP268.5.N7 P3 1947.

Pauli:1947:EPQ

[Pau47b] Wolfgang Pauli. Exclusion Principle and quantum mechanics: Discours prononcé à la réception du prix Nobel de physique 1945. (French) [Lecture given on receiving the 1945 Nobel Prize in Physics]. Dialectica: International Review of Philosophy of Knowledge, 1(2):204, 1947. ISSN 1746-8361.

Pauli:1949:ECQ

[Pau49] Wolfgang Pauli. Einstein’s contributions to quantum theory. In Schilpp [Sch49b], chapter 4, pages 147–160. ISBN 0-87548-286-

4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Pauli:1955:EBQ

[Pau55] Wolfgang Pauli. Einsteins Beitrag zur Quantentheorie. (German) [Einstein’s contribution to quantum theory]. In ????, editor, Albert Einstein als Philosoph und Naturforscher. (German) [Albert Einstein as philosopher and scientist], pages 74–83. ????, Stuttgart, West Germany, 1955. LCCN ????

Pauli:1958:AEE

[Pau58] Wolfgang Pauli. Albert Einstein in der Entwicklung der Physik. (German) [Albert Einstein in the development of physics]. Neue Zürcher Zeitung, 89(??):??, January 12, 1958.

Pauli:1961:AVP

[Pau61a] Wolfgang Pauli. Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology], volume 115 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961. 183 pp. LCCN ????

Pauli:1961:IAE

[Pau61b] Wolfgang Pauli. Impressionen über Albert Einstein. (German) [Impressions of Albert Einstein]. In Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology] [Pau61a], pages 81–?? LCCN ????

Pauling:1979:PNP

[Pau79] Linus Pauling. Perspective: Nobel Peace Laureates: 1959. Bulletin of the Atomic Scientists, 35(6):1, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 10 December 1963, Oslo, Norway.

Pauli:1981:TR

[Pau81] Wolfgang Pauli. Theory of Relativity. Dover, New York, NY, USA, 1981. ISBN 0-486-64152-X. xiv + 241 pp. LCCN QC173.55 .P3813 1981. Translated from the German [Pau21] by G. Field.

Paul:1993:MWG

[Pau93] Erich Robert Paul. The Milky Way galaxy and statistical cosmology, 1890–1924. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35363-7 (hardcover). xiv + 262 pp. LCCN QB857.7 .P38 1993. URL http://www.loc.

gov/catdir/description/cam025/92036530.html;   http:
//www.loc.gov/catdir/toc/cam029/92036530.html.

Pauli:1994:AED

[Pau94a] Wolfgang Pauli. Albert Einstein and the development of physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 13, pages 117–124. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:ECQ

[Pau94b] Wolfgang Pauli. Einstein’s contribution to quantum theory. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 9, pages 85–94. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:IAE

[Pau94c] Wolfgang Pauli. Impressions of Albert Einstein. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 12, pages 113–116. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:STC

[Pau94d] Wolfgang Pauli. Space, time and causality in modern physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 10, pages 95–106. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0812/94015098-d.
htm.

Pauli:1994:TRS

[Pau94e] Wolfgang Pauli. The theory of Relativity and science. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 11, pages 107–112. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0812/94015098-d.
htm.

Paz:2010:PPC

[Paz10] Octavio Paz. Pasado y presente en claro: 20 años del Premio Nobel. (Spanish) [Past and clear present: 20 years of the Nobel Prize]. Dirección de Publicaciones del Consejo Nacional para la Cultura y las Artes, México, DF, México, 2010. ISBN 607-455-512-5. 45 pp. LCCN PQ7297.P285 A6 2010.

Porzio:1972:PCM

[PB72] Domenico Porzio and Heinrich Böll. La provocazione: [da Cristo a Mao, da Socrate a Marcuse, da Galileo a Einstein]. (Italian) [The provocation: [from Christ to Mao, from Socrates to Marcuse, from Galileo to Einstein]]. Ferro, Milano, Italia, 1972. 733 pp. LCCN ????

Povh:1979:ESP

[PB79] I. L. Povh and A. D. Barinberg. From the Einstein–Szilard patent to modern magnetohydrodynamics. Impact of Science on Society, 29(1):49–60, ???? 1979. ISSN 0019-2872.

Provost:2006:RPT

[PB06] Jean-Pierre Provost and Christian Bracco. La Relativité de Poincaré de 1905 et les transformations actives. (French) [Poincaré’s Relativity of 1905 and active transformations]. Archive for History of Exact Sciences, 60(3):337–351, May 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0003-9519&volume=60&issue=3&
spage=337.

Perrin:2012:HBT

[PBM^{+}12] A. Perrin, R. Bücker, S. Manz, T. Betz, C. Koller, T. Plisson, T. Schumm, J. Schmiedmayer, et al. Hanbury Brown and Twiss correlations across the Bose–Einstein condensation threshold. Nature Physics, 8(3):195–198, January 29, 2012. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (elec-

tronic). URL http://www.nature.com/nphys/journal/v8/
n3/full/nphys2212.html.

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN

QC16.O62 P35 2006. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005002173-d.html;   http://www.
loc.gov/catdir/enhancements/fy0723/2005002173-b.html;
 http://www.loc.gov/catdir/toc/ecip056/2005002173.html.

Price:1977:UPF

[PCH77] William Charles Price, Seymour S. Chissick, and Werner Heisenberg, editors. The Uncertainty Principle and foundations of quantum mechanics: a fifty years’ survey. Wiley, New York, NY, USA, 1977. ISBN 0-471-99414-6. xvii + 572 pp. LCCN QC174.125 .U5.

Psaltis:2015:ABH

[PD15] Dimitrios Psaltis and Sheperd S. Doeleman. Astronomy: The black hole test. Scientific American, 313(3):74–79, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
74.html;   http://www.nature.com/scientificamerican/
journal/v313/n3/pdf/scientificamerican0915-74.pdf.

Pais:1949:NRP

[PE49] A. Pais and S. T. Epstein. Note on relativistic properties of self-energies. Reviews of Modern Physics, 21(3):445–446, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.445;   http://rmp.aps.org/

abstract/RMP/v21/i3/p445_1. Special issue in honor of Albert Einstein on his 70th birthday.

Peare:1949:AEB

[Pea49] Catherine Owens Peare. Albert Einstein: a biography for young people. Henry Holt and Company, Incorporated, New York, NY, USA, 1949. 152 pp. LCCN QC16.E5 P4.

Pearson:1958:PP

[Pea58] Lester Pearson. Peace and prosperity. Bulletin of the Atomic Scientists, 14(6):214, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peacock:1990:ERU

[Pea90] John Peacock. Einstein rings unveil the invisible. Physics World, 3(5):20–21, May 1990. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/3/5/phwv3i5a18.
pdf.

Pearl:2006:DMA

[Pea06] Nancy Pearl. Driving Mr. Albert: A trip across America with Einstein’s brain. Library Journal, 131(20):183, December 2006. CODEN LIBJA7. ISSN 0363-0277.

Peacock:2008:QRH

[Pea08] Kent A. Peacock. The Quantum Revolution: a Historical Perspective. Greenwood guides to great ideas in science. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2008. ISBN 0-313-33448-X. ISSN 1559-5374. xviii + 220 pp. LCCN

QC173.98 .P43 2008. URL http://www.loc.gov/catdir/toc/
ecip081/2007039786.html.

Pechenkin:2004:BRB

[Pec04] A. A. Pechenkin. Book review: Einstein studies in Russia: Yuri Balashov and Vladimir Vizgin (Eds.); Birkhäuser, Basel, 2002, 315pp, US $59.95, ISBN 0-8176-4263-3. Studies in History and Philosophy of Modern Physics, 35(1):120–124, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000844. See [BV02].

Peebles:1993:PPC

[Pee93] P. J. E. (Phillip James Edwin) Peebles. Principles of Physical Cosmology. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-07428-3, 0-691-01933-9 (paperback). xviii + 718 pp. LCCN QB981 .P424

1993. US$79.50US$29.95. URL http://www.loc.gov/catdir/
description/prin021/92033370.html.

Peebles:2014:DHB

[Pee14] Phillip James Edwin Peebles. Discovery of the hot Big Bang: What happened in 1948. European Physical Journal H, 39 (2):205–223, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2014-50002-y.

Pegram:1939:EFN

[Peg39] George B. Pegram. Enrico Fermi — Nobel Prize man in physics for 1938. The Scientific Monthly, 49(2):182–184, August 1939. CODEN SCMOAA. ISSN 0096-3771. URL http://www.jstor.

org/stable/17044.

Peierls:1939:CCN

[Pei39] Rudolf Peierls. Critical conditions in neutron multiplication. Mathematical proceedings of the Cambridge Philosophical Society, 35(4):610–615, November 1939. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic). This is the first paper with an estimate of the critical mass of uranium needed to sustain a chain reaction. It contains a note added in proof to the fission model of Bohr and Wheeler [BW39b] that appeared after this paper was submitted.

Peierls:1955:AE

[Pei55] Rudolf Peierls. Albert Einstein. In Physical Society Year Book 1955, pages 69–71. ????, ????, 1955.

Peierls:1971:SIB

[Pei71] Rudolf Peierls. Strong interactions [Born–Einstein correspondence]. New Scientist, ??(??):339, May 6, 1971. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Peierls:1983:WED

[Pei83] Rudolf Peierls. What Einstein did: Subtle Is The Lord… the Science and the Life of Albert Einstein by Abraham Pais. New York Review of Books, 30(7):21–23, April 28, 1983. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:/

/www.nybooks.com/articles/archives/1983/apr/28/what-
einstein-did/.

Peierls:1985:BPR

[Pei85] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). xii + 350 + 12 pp. LCCN QC16.P375

A32 1985. URL http://site.ebrary.com/id/10897402;
https://www.jstor.org/stable/j.ctt7ztn4b.

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Penrose:1979:BH

[Pen79a] Roger Penrose. Black holes. In Aichelburg and Sexl [AS79], pages 33–50. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66.

DM48.00. URL http://link.springer.com/chapter/10.
1007/978-3-322-91080-6_4/.

Penzias:1979:OE

[Pen79b] Arno A. Penzias. The origin of the elements. Reviews of Modern Physics, 51(3):425–431, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.51.425;
http://rmp.aps.org/abstract/RMP/v51/i3/p425_1.

Penrose:2003:RGE

[Pen03] Roger Penrose. The rediscovery of gravity: the Einstein equation of general relativity. In Farmelo [Far03a], pages 180–212. ISBN 1-86207-555-7. LCCN Q125 .I88 2003.

Penrose:2005:RRC

[Pen05] Roger Penrose. The road to reality: a complete guide to the laws of the universe. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-679-45443-8. xxviii + 1099 pp. LCCN QC20 .P366 2005. URL http://www.loc.

gov/catdir/bios/random051/2004061543.html;   http://
www.loc.gov/catdir/description/random051/2004061543.
html;   http://www.loc.gov/catdir/enhancements/fy0619/
2004061543-s.html;   http://www.loc.gov/catdir/enhancements/
fy0619/2004061543-t.html.

Peplow:2004:GPT

[Pep04a] Mark Peplow. Gravity probe to test Einstein. Nature News, April 19, 2004. CODEN ???? ISSN ???? URL http://www.

nature.com/news/2004/040419/full/news040412-11.html;
 http://www.nature.com/physics/highlights/6985-2.html.

Pepperell:2004:BRC

[Pep04b] Robert Pepperell. Book review: The Curvature of Spacetime: Newton, Einstein, and Gravitation. Leonardo (Oxford, England), 37(3):255–256, June 2004. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL http://muse.

jhu.edu/article/169882.

Pepperell:2004:BRE

[Pep04c] Robert Pepperell. Book reviews: Einstein, Picasso: Space, Time, and the Beauty that Causes Havoc, and: Inner Vision: An Exploration of Art and the Brain. Leonardo (Oxford, England), 37(1):73–75, February 2004. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL http://muse.

jhu.edu/article/52535.

Pepin:2005:ELR

[Pep05] Jerry Pepin. Einstein’s light relief? Physics World, 18(12):20, December 2005. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/18/12/phwv18i12a28.pdf.

Perrin:1910:BMM

[Per10] Jean Perrin. Brownian movement and molecular reality. Taylor and Francis, London, UK and Boca Raton, FL, USA, 1910. 93 pp. LCCN ???? Reprinted in [Per05b].

Perrin:1913:AFA

[Per13] Jean Baptiste Perrin. Les atomes. (French) [Atoms]. Félix Alcan, Paris, France, 1913. xvi + 299 pp. LCCN ???? See also English translations [Per23a, Per90].

Perrin:1923:A

[Per23a] Jean Baptiste Perrin. Atoms. Constable, London, UK, second edition, 1923. xiv + 231 pp. LCCN QC173 P45. Authorised translation by D. Ll. Hammick. See original edition [Per13]. Reprinted in [Per90].

Perrine:1923:CHA

[Per23b] C. D. Perrine. Contribution to the history of attempts to test the Theory of Relativity by means of astronomical observations. Astronomische Nachrichten, 219(17–18):281–284, ???? 1923. CO-DEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/asna.
19232191706.

Perrin:1939:CRC

[Per39] Francis Perrin. Calcul relatif aux conditions éventuelles de transmutation en chaîne de l’uranium. (French) [Calculation of possible conditions on the uranium chain transmutation]. Comptes rendus de l’Académie des sciences, Paris, 208(18):1394–1396, May 1, 1939. URL http://

gallica.bnf.fr/ark:/12148/bpt6k6238836c/f26.image.r=
perrin%20calcul%20relative%20aux%20conditions#.

Peres:1985:EGB

[Per85] Asher Peres. Einstein, Gödel, Bohr. Foundations of Physics, 15 (2):201–205, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00735292.

Perrin:1990:A

[Per90] Jean Perrin. Atoms. Ox Bow Press, Woodbridge, CT, USA, 1990. ISBN 0-918024-78-1, 0-918024-79-X (paperback). xiv + 231 pp. LCCN QC173 .P4213 1990.

Pereboom:1991:MEP

[Per91] Derk Pereboom. Mathematical expressibility, perceptual relativity, and secondary qualities. Studies in History and Philosophy of Science Part A, 22(1):63–88, March 1991. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190015K.

Persson:1997:NLI

[Per97] Torsten Persson, editor. Nobel lectures including presentation speeches and laureates’ biographies: 1991–1995. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-3059-1, 981-02-3060-5 (paperback), 981-02-2677-2 (hardcover), 981-02-2678-0 (paperback). 220 pp. LCCN ????

Perlick:2000:ROF

[Per00] Volker Perlick. Ray Optics, Fermat’s Principle, and Applications to General Relativity, volume 61 of Lecture notes in physics: New series m, Monographs; Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66898-5. x + 220 pp. LCCN QC389 .P37 2000.

Peres:2005:EPR

[Per05a] A. Peres. Einstein, Podolsky, Rosen, and Shannon. Foundations of Physics, 35(3):511–514, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/s10701-004-1986-6.

Perrin:2005:BMM

[Per05b] Jean Perrin. Brownian movement and molecular reality. Dover phoenix editions. Dover, New York, NY, USA, 2005. ISBN 0-486-44257-8. 93 pp. LCCN QC184 .P47 2005. URL http://

www.loc.gov/catdir/enhancements/fy0619/2004061764-d.

html. Reprint of [Per10].

Perers:2009:EAK

[Per09] Karin Perers, editor. Erik Axel Karlfeldt: the road to the Nobel Prize, volume 41 of Karlfeldtsamfundet’s publications. Karlfeldtsamfundet, Stockholm, Sweden, 2009. ISBN 91-978554-0-5. 96 pp. LCCN PT9875.K2 Z65 2009; PT9875 .K2 Z65 2009.

Peruzzi:2010:BRS

[Per10] Giulio Peruzzi. Book review: Sandra Linguerri e Raffaella Simili (a cura di), Einstein parla italiano. Itinerari e polemiche. Bologna: Edizioni Pendragon, 2008. x + 359 pp., ISBN 978-88-8342-530-1. Nuncius, 25(1):159–160, ???? 2010. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00271.

Perlmutter:2012:NLM

[Per12] Saul Perlmutter. Nobel Lecture: Measuring the acceleration of the cosmic expansion using supernovae. Reviews of Modern Physics, 84(3):1127–1149, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.1127;
 http://rmp.aps.org/abstract/RMP/v84/i3/p1127_1;   http:
//www.nobelprize.org/nobel_prizes/physics/laureates/
2011/.

Perovic:2021:DQN

[Per21] Slobodan Perović. From Data to Quanta : Niels Bohr’s Vision of Physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2021. ISBN 0-226-79847-X. 251 pp. LCCN QC16.

Pesic:2013:ERH

[Pes13] Peter Pesic. Essay review: Hermann Weyl’s neighborhood: Umgebungen: Symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus, by Norman Sieroka; Chronos Verlag, Zurich, 2010, pp. 411, Price EUR 43,00, hardcover, ISBN 978-3-0340-1006-1. Studies in History and Philosophy of Science Part A, 44(1):150–153, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000751. See [Sie10].

Petzoldt:1921:SRG

[Pet21] Joseph Petzoldt. Die Stellung der Relativitätstheorie in der geistigen Entwicklung der Menschheit. (German) [The position of the theory of relativity in the spiritual development of mankind]. Sibyllen-Verlag, Dresden, Germany, 1921. 131 pp.

Petzoldt:1923:SRG

[Pet23] Joseph Petzoldt. Die Stellung der Relativitätstheorie in der geistigen Entwicklung der Menschheit. (German) [The position of the theory of relativity in the spiritual development of mankind]. Johann Ambrosius Barth, Leipzig, Germany, second edition, 1923. 131 pp.

Pettitt:1972:ABD

[Pet72] Roland A. Pettitt. Los Alamos before the dawn. Pajarito Publications, Los Alamos, NM, USA, 1972. 64 pp. LCCN F804.L6 P47.

Peter:2008:NLE

[Pet08] Englund Peter. Nobel lectures in economic sciences 2001–2005. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-279-438-7, 981-279-439-5. 525 pp. LCCN HB171 .E2443 2008. pbk.

Petters:2010:GAL

[Pet10] Arlie O. Petters. Gravity’s action on light. Notices of the American Mathematical Society, 57(11):1392–1409, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/journals/
notices/201011/201011-full-issue.pdf. See preceding arti-

cle [KN08], and book-length treatment [Per00].

Pfalz:2012:BVR

[Pfa12] Maike Pfalz. Einsteins Versprechen, von À. Rovira, F. Miralles. Physik Journal, ??(??):??, January 16, 2012. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL

http://www.pro-physik.de/details/rezension/1442143/
Einsteins_Versprechen.html. See [RMHD11].

Pfeiffer:1977:DUW

[Pfe77] Heinrich Pfeiffer, editor. Denken und Umdenken: zu Werk und Wirkung von Werner Heisenberg. (German) [Thinking and rethinking: the work and influence of Werner Heisenberg]. Piper, München, West Germany, 1977. ISBN 3-492-02275-8. 279 + 10 pp. LCCN QC16.H35 D46. DM28.00.

Pfleeger:1999:AEE

[Pfl99] Shari Lawrence Pfleeger. Albert Einstein and empirical software engineering. Computer, 32(10):32–37, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/co/books/co1999/pdf/rx032.
pdf;   http://www.computer.org/computer/co1999/rx032abs.
htm.

Podosenov:2017:DIE

[PFM17] Stanislav Podosenov, Jaykov Foukzon, and Elena Men’kova. Difficulties in the Interpretation of the Einstein’s Relativity Theory: Basics, Concepts, Methods. LAP Lambert Academic Publishing, Saarbrücken, Germany, 2017. ISBN 3-330-06799-3. 104 pp. LCCN ????

Payne-Gaposchkin:1984:AOR

[PGH84] Cecilia Helena Payne-Gaposchkin and Katherine Haramundanis. An autobiography and other recollections. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25752-2. vi + 269 pp. LCCN ????

Payne-Gaposchkin:1996:CPG

[PGH96] Cecilia Payne-Gaposchkin and Katherine Haramundanis, editors. Cecilia Payne-Gaposchkin: an autobiography and other recollections. Cambridge University Press, Cambridge, UK, second edition, 1996. ISBN 0-521-48251-8 (hardcover), 0-521-48390-5 (paperback). xxii + 277 pp. LCCN QB36.G37 A33 1996. URL http://catdir.loc.

gov/catdir/description/cam027/95044001.html;   http:
//catdir.loc.gov/catdir/toc/cam021/95044001.html.

Phelps:1921:SCT

[Phe21] Charles C. Phelps. Stopping the chimney thief, how the Einstein Contest worked, and more. Scientific American, 124(6):101–103, February 5, 1921. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v124/n6/
pdf/scientificamerican02051921-101.pdf.

Phillips:1922:NET

[Phi22] H. B. Phillips. Note on Einstein’s theory of gravitation. Journal of Mathematics and Physics (MIT), 1(3): 177–190, April 1922. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/
10.1002/sapm192213177.

Philips:1974:MAE

[Phi74] J. Philips. El mundo de Albert Einstein. (Spanish) [The world of Albert Einstein]. Revista del Colegio Nacional de Enfermeras, 21(1):24–26, July/September 1974. ISSN 0045-7329.

Phipps:1997:NVE

[Phi97] Thomas E. Phipps, Jr. Newton versus Einstein: How matter interacts with matter. Foundations of Physics, 27(10):1457–1460, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF02551523.

Phillips:1998:NLL

[Phi98] William D. Phillips. Nobel Lecture: Laser cooling and trapping of neutral atoms. Reviews of Modern Physics, 70(3):721– 741, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.70.721;   http://rmp.aps.
org/abstract/RMP/v70/i3/p721_1.

Pauli:1979:WPW

[PHvMW93] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor F. (Victor Frederick) Weisskopf. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–1993. ISBN 0-387-08962-4, 3-540-08962-4, 0-387-13609-6, 3-540-13609-6, 0-387-54911-0, 3-540-54911-0. ???? pp. LCCN QC16.P37 W64.

Pauli:1979:WBB

[PHvMW01a] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor Frederick Weisskopf, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, ua. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11, 15, 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–2001. ISBN 0-387-08962-4 (vol. 1). ???? pp. LCCN QC16.P37 A34.

Pauli:2001:WBBa

[PHvMW01b] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor Frederick Weisskopf, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, ua. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11, 15, 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–2001. ISBN 0-387-08962-4 (vol. 1). ???? pp. LCCN QC16.P37 A34.

Piaggio:1929:EOUa

[Pia29a] H. T. H. Piaggio. Einstein’s and other unitary field theories: An explanation for the general reader. Nature, 123(3109): 839–841, June 1, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v123/n3109/pdf/123839a0.pdf.

Piaggio:1929:EOUb

[Pia29b] H. T. H. Piaggio. Einstein’s and other unitary field theories: An explanation for the general reader. Nature, 123(3110): 877–879, June 8, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v123/n3110/pdf/123877a0.pdf.

Piaggio:1930:CS

[Pia30] H. T. H. Piaggio. Concept of space. Nature, 125(3163):897–898, June 14, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Piaggio:1931:CMM

[Pia31] H. T. H. Piaggio. (1) Cours de mécanique: mécanique des solides indéformables, mécanique des milieux continus déformables, théorie sommaire des machines et de l’aviation, les mécaniques de Newton et d’Einstein; (2) Cours de géométrie;

(3) Cours d’analyse professé à l’École polytechnique; (4) Cours de mathématiques générales (analyse et géométrie); (5) Erreurs et moindres carrés; (6) Principes géométriques d’analyse;

(7) Leçons sur les ensembles analytiques et leurs applications. Nature, 127(3215):883–884, June 13, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v127/n3215/
full/127883a0.html.

Pickering:1921:ET

[Pic21] William H. Pickering. The Einstein theories. Scientific American, 3:292–295, April 1, 1921. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v3/
n4msupp/pdf/scientificamerican04011921-292supp.pdf.

Pickett:1993:HE

[Pic93] Maria-Elena Pickett. Haunted by Einstein. Physics World, 6(1): 76, January 1993. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/6/1/phwv6i1a42.pdf.

Pickover:2005:SDE

[Pic05] Clifford A. Pickover. Sex, Drugs, Einstein and Elves: Sushi, Psychedelics, Parallel Universes, and the Quest for Transcendence. Smart Publications, Petaluma, CA, USA, 2005. ISBN 1-890572-17-9. xxxii + 318 pp. LCCN BF408 .P534 2005.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN Q175.32.R45 P53 2008. URL

http://catdir.loc.gov/catdir/toc/ecip087/2007051167.
html;   http://www.gbv.de/dms/goettingen/555633241.
pdf;   http://www.oup.com/us/catalog/general/subject/
Physics/?view=usa&ci=9780195336115.

Pierpont:1923:GRE

[Pie23] James Pierpont. The geometry of Riemann and Einstein. American Mathematical Monthly, 30(8):425–438, December 1923. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pierpont:1924:GRE

[Pie24a] James Pierpont. The geometry of Riemann and Einstein. American Mathematical Monthly, 31(1):26–39, January 1924. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pierpont:1924:RPR

[Pie24b] James Pierpont. Recent publications: Reviews: Relativity, a Systematic Treatment of Einstein’s Theory, by J. Rice. American Mathematical Monthly, 31(8):395–398, October 1924. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pike:1963:FEE

[Pik63] E. R. Pike. A fundamental extension of Einstein’s diffusion — mobility relationship. Nature, 199(4897):994, September 7, 1963. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v199/n4897/pdf/199994b0.pdf.

Piner:2000:CSCj

[Pin00a] Mary-Louise G. Piner. Computer Society connection: Second annual design competition seeks student participants; Society presents new award for undergraduate teachers: [Joseph Zachary recognized for innovative educational materials; Timothy Long and Ohio State colleague Bruce Weide honored]; IEEE Fellow Jack Kilby receives Nobel Prize in Physics. Computer, 33 (11):70–73, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.computer.

org/co/books/co2000/pdf/ry070.pdf.

Pinsent:2000:GLT

[Pin00b] Andrew Pinsent. God’s library through Einstein’s eyes. Physics World, 13(6):55, June 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/13/6/phwv13i6a31.
pdf.

Pines:2015:WWD

[Pin15] David Pines. What we don’t understand, we explain to each other. The Physics Teacher, 53(9):526–531, December 2015. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Pippard:1983:SLS

[Pip83] B. Pippard. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. TLS — The Times Literary Supplement, ??(4174):315, ???? 1983. ISSN 0307-661X.

Pippard:1997:AEB

[Pip97] B. Pippard. Albert Einstein, a biography — A. Fölsing. TLS — The Times Literary Supplement, ??(4928):5, September 12, 1997. ISSN 0307-661X.

Piron:2005:NEG

[Pir05] C. Piron. New Einstein gravitation*. Foundations of Physics, 35(9):1643–1647, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-005-6486-9. See

clarification [Pir07].

Piron:2007:FCN

[Pir07] Constantin Piron. Further clarification of “New Einstein Gravitation”. Foundations of Physics, 37(4–5):761–762, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-007-9128-6. See [Pir05].

Pistone:2015:EP

[Pis15] Massimo Pistone. Einstein & Parmenide. Armando, Roma, Italia, 2015. ISBN 88-6677-866-4. 79 pp. LCCN ????

Pitts:2016:EPS

[Pit16] J. Brian Pitts. Einstein’s physical strategy, energy conservation, symmetries, and stability: “But Grossmann & I believed that the conservation laws were not satisfied”. Studies in History and Philosophy of Modern Physics, 54(??):52–72, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300733.

Pitici:2017:BWM

[Pit17] Mircea Pitici, editor. The Best Writing on Mathematics, volume 2016. Princeton University Press, Princeton, NJ, USA, 2017. ISBN 0-691-17529-2 (paperback). xxii + 377 pp. LCCN QA1 .B337; QA93 .B476 2016.

Pyenson:2011:BSR

[PJMS11] Lewis Pyenson, Sean F. Johnston, Alberto A. Martínez, and Richard Staley. Book symposium: Revisiting the history of Relativity: Richard Staley: Einstein’s Generation: The Origins of the Relativity Revolution. Chicago: University of Chicago Press, 2008, x + 494 pp, $38 PB, $98 HB. Metascience, 20(1):53–73,

March 2011. URL http://www.springerlink.com/content/
w08j885701u325t3/.

Plebanski:2006:IGR

[PK06] Jerzy Plebański and Andrzej Krasiński. An Introduction to General Relativity and Cosmology. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85623-X (hardcover), 1-139-45840-X. xix + 534 pp. LCCN

QC173.6 .P64 2006. URL http://www.loc.gov/catdir/
enhancements/fy0618/2006299685-d.html;   http://www.
loc.gov/catdir/enhancements/fy0618/2006299685-t.html.

Petrov:2017:MTG

[PKLT17] Alexander N. Petrov, Sergei Kopeikin, Robert R. Lompay, and Bayram Tekin. Metric theories of gravity: perturbations and conservation laws, volume 38 of De Gruyter studies in mathematical physics. Walter de Gruyter GmbH, Berlin, Germany, 2017. ISBN 3-11-035173-0, 3-11-035178-1 (e-book). xxiii + 595 pp. LCCN QB337 .P48 2017.

Potters:2017:SST

[PL17] Jan Potters and Bert Leuridan. Studying scientific thought experiments in their context: Albert Einstein and electromagnetic induction. Studies in History and Philosophy of Modern Physics, 58(??):1–11, May 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219816301915.

Planck:1900:VWS

[Pla00a] Max Planck. Über eine Verbesserung der Wienschen Spektralgleichung. (German) [On an improvement of Wien’s equation for the spectrum]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(13):202–204, October 19, 1900. CO-DEN VDPEAZ. ISSN 0372-5448. URL http://web.

ihep.su/dbserv/compas/src/planck00/eng.pdf;   http:
//web.ihep.su/owa/dbserv/hw.part2?s_c=PLANCK+1900;
https://hdl.handle.net/2027/coo.31924056107224?urlappend=

%3Bseq=520. English translation in [?].

Planck:1900:TGE

[Pla00b] Max Planck. Zur Theorie das Gesetzes der Energieverteilung im Normalspektrum. (German) [On the law of distribution of energy in the normal spectrum]. Verhandl. Dtsch. phys. Ges., 2(??):237–??, ???? 1900. URL http://web.

ihep.su/dbserv/compas/src/planck00b/eng.pdf;   http:
//web.ihep.su/owa/dbserv/hw.part2?s_c=PLANCK+1900B.

Planck:1901:GEI

[Pla01] Max Planck. Über das Gesetz der Energieverteilung im Normalspektrum. (German) [On the law of distribution of energy in the normal spectrum]. Annalen der Physik (1900) (series 4), 309(3):553–563, ???? 1901. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/andp.
19013090310/abstract;   http://web.ihep.su/dbserv/compas/
src/planck01/eng.pdf;   http://web.ihep.su/owa/dbserv/
hw.part2?s_c=PLANCK+1901.

Planck:1920:EBEa

[Pla20a] Max Planck. Die Entstehung und bisherige Entwicklung der Quantentheorie. Nobel-Vortrag, (mit Lebenslauf). Gehalten vor der königl. Schwed. Akad. d. Wiss. zu Stockholm am 2.6.1920. (German) [The emergence and current development of quantum theory. Nobel Lecture, (with curriculum vitae). Held before the Royal Swedish Academy of Sciences in Stockholm on 2.6.1920]. J. A. Barth, Leipzig, Germany, 1920. 32 pp. LCCN ????

Planck:1920:EBEb

[Pla20b] Max Planck. Entstehung und bisherige Entwicklung der Quantentheorie: Nobelvortrag, gehalten vor der Königlich Schwedischen Akademie der Wissenschaften zu Stockholm am 2. Juni 1920. (German) [Emergence and current development of Quantum Theory: Nobel Lecture, held in front of the Royal Swedish Academy of Sciences in Stockholm on 2 June 1920]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1920. 14 pp. LCCN ????

Planck:1920:GPS

[Pla20c] Max Planck. The genesis and present state of development of the quantum theory. Nobel Lecture, June 2, 1920, June

2, 1920. URL http://www.nobelprize.org/nobel_prizes/
physics/laureates/1918/planck-lecture.html. The Nobel

Prize in Physics for 1918 was award to Max Karl Ernst Ludwig Planck “in recognition of the services he rendered to the advancement of Physics by his discovery of energy quanta”.

Planck:1922:EBE

[Pla22a] Max Planck. Die Entstehung und bisherige Entwicklung der Quantentheorie. (German) [The emergence and current development of quantum theory]. Nörstedt, Stockholm, Sweden, 1922. 1–14 pp. Nobel Prize in Physics lecture, 2 June 1920.

Planck:1922:ODQ

[Pla22b] Max Planck. The Origin and Development of the Quantum Theory. Clarendon Press, Oxford, UK, 1922. 23 + 1 pp. LCCN QC174 .P8. English translation by Hans Thacher Clarke and Ludwik Silberstein of the Nobel Prize address delivered before the Royal Swedish Academy of Sciences at Stockholm, 2 June, 1920.

Planck:1932:WSG

[Pla32] Max Planck. Where is Science Going? W. W. Norton & Co., New York, NY, USA, 1932. 221 pp. LCCN Q175 .P57. Prologue by Albert Einstein. Translation and biographical note by James Murphy.

Planck:1933:WSG

[Pla33] Max Planck. Where is Science Going? George Allen and Unwin, London, UK, 1933. 224 pp. LCCN Q175 .P57. Prologue by Albert Einstein. Translation and biographical note by James Murphy.

Planck:1959:NSC

[Pla59] Max Planck. The new science. 3 complete works: Where is science going? The universe in the light of modern physics; The philosophy of physics. Greenwich editions. Meridian Books, New York, NY, USA, 1959. 328 pp. LCCN QC6 .P623. Translated from the German by James Murphy and W. H. Johnston.

Planck:1977:WSG

[Pla77] Max Planck. Where is Science Going? American Mathematical Society, Providence, RI, USA, 1977. ISBN 0-404-14696-1. 221 pp. LCCN Q175 .P57 1977. Prologue by Albert Einstein. English translation and biographical note by James Murphy.

Planck:1981:WSG

[Pla81] Max Planck, editor. Where is science going? Ox Bow Press, Woodbridge, CT, USA, 1981. ISBN 0-918024-21-8, 0-918024-22-6 (paperback). 224 pp. LCCN Q175 .P436x 1981. Edited by James Murphy, with a preface by Albert Einstein.

Planck:2010:ODQ

[Pla10] Max Planck. The Origin and Development of the Quantum Theory. Project Gutenberg, ????, 2010. 23 pp. URL https://www.

gutenberg.org/ebooks/33663. English translation by Hans

Thacher Clarke and Ludwik Silberstein of the Nobel Prize address delivered before the Royal Swedish Academy of Sciences at Stockholm, 2 June, 1920.

Piscopo:2000:NTD

[PLM^{+}0 ] Filippo. Piscopo, Lorena. Luciano, Giovanni. Maggi, Dario. Fo, and Franca Rame, editors. A Nobel for two: a documentary on Dario Fo and Franca Rame. Distributed by Contemporary Arts Media, South Fremantle, WA, Australia, 2000 (??). LCCN PQ4866.O2 Z462 2000 DVD. 1 DVD (55 min.).

Plotnitsky:2006:BRB

[Plo06] Arkady Plotnitsky. Book review: It’s About Time: Understanding Einstein’s Relativity. By N. David Mermin, Princeton University Press, Princeton, NJ, USA, 2005, xv, 192 pp. (Hardcover). Foundations of Physics, 36(8):1286–1290, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-006-9063-y.

Papenfuss:2002:YWH

[PLS02] Dietrich Papenfuss, D. (Dieter) Lüst, and Wolfgang Schleich, editors. 100 Years Werner Heisenberg: works and impact. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2002. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). ix + 299 pp. LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.loc.

gov/catdir/bios/wiley046/2003268171.html;   http://
www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Pendergrass:1987:MCD

[PNT87] Gurli Pendergrass, Lorelle Nelson, and A. Costandina Titus. The Mushroom Cloud and the Downwinders. Forlaget Futurum, Nykøbing Mors, Denmark, 1987. ISBN 87-88497-13-5. 72 pp. LCCN ????

Pockley:1998:GRA

[Poc98] Peter Pockley. General Relativity: Australia plans gravitational telescope. Physics World, 11(9):9, September 1998. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
11/9/phwv11i9a10.pdf.

Poincare:1902:SHF

[Poi02] Henri Poincaré. La Science et l’Hypothèse. (French) [Science and Hypothesis]. Flammarion, Paris, France, 1902. ???? pp.

Poincare:2007:DEE

[Poi07] Henri Poincaré. On the dynamics of the electron (excerpts). In Renn et al. [RSS^{+}07a], pages 253–271. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6

.G469 2007. URL https://link.springer.com/chapter/10.

1007/978-1-4020-4000-9_19. Originally published as “Sur la dynamique de l’électron” in Rendiconti del Circolo Matematico di Palermo 21, 129–175 (1906), dated Paris July 1905.

Polkinghorne:1983:ERS

[Pol83] Revd. Dr. J. C. Polkinghorne, F.R.S. Essay review: The science and the life of Albert Einstein: Subtle Is The Lord: the Science and the Life of Albert Einstein, by A. Pais. Contemporary Physics, 24(3):302–305, ???? 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Polikarov:1989:CEP

[Pol89] Asaria Polikarov. Über den Charakter von Einsteins philosophischem Realismus. (German) [On the character of Einstein’s philosophical realism]. Philosophia Naturalis, 26(1):135–158, 1989. CODEN ???? ISSN 0031-8027.

Polverini:1991:AEF

[Pol91] L. Polverini. Albert Einstein and the Fascist Oath of 1931. Rivista Storica Italiana, 103(1):268–280, April 1991. ISSN 0035-

Polenberg:2002:MJR

[Pol02] Richard Polenberg, editor. In the matter of J. Robert Oppenheimer: the security clearance hearing. Cornell paperbacks. Cornell University Press, Ithaca, NY, USA, 2002. ISBN 0-8014-8661-0 (paperback), 0-8014-3783-0 (cloth). xxxii + 409 pp. LCCN QC16.O62 I5 2002. URL http://www.h-net.org/

review/hrev-a0c5p6-aa.

Politzer:2005:NLD

[Pol05] David Politzer. Nobel Lecture: The dilemma of attribution. Reviews of Modern Physics, 77(3):851–856, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.851;   http://rmp.aps.org/
abstract/RMP/v77/i3/p851_1.

Poor:1921:PME

[Poo21] Charles Lane Poor. Planetary motions and the Einstein theories. Scientific American, 3:484–486, June 1, 1921. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v3/
n6msupp/pdf/scientificamerican06011921-484supp.pdf.

Pool:1990:CES

[Poo90] Robert Pool. Closing in on Einstein’s Special Relativity Theory: With exquisitely precise measurements not possible even a few years ago, physicists are testing the theory explicitly. Science, 250(4985):1207–1208, November 30, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/250/4985/1207.full.pdf.

Popovici:1957:AE

[Pop57] Andrei Popovici. Albert Einstein. Gaz. Mat. Fiz. Ser. A, 9(62): 207–213, 1957.

Popper:1966:NDB

[Pop66] Karl R. Popper. A note on the difference between the Lorentz– Fitzgerald contraction and the Einstein contraction. The British Journal for the Philosophy of Science, XVI(64):332–333, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/XVI/64/332.full.pdf+html.

Popovic:1998:JPP

[Pop98] Milan Popović, editor. Jedno prijateljstvo: pisma Mileve i Alberta Ajnštajna Heleni Savić. (Serbian) [A Friendship: Letters from Mileva and Albert Einstein to Helene Savić]. CID, Podgorića, Montenegro, 1998. ISBN 86-495-0056-0. 324 (est.) pp. LCCN ???? Marić’s original German letters with parallel Serbian translation.

Pople:1999:NLQ

[Pop99] John A. Pople. Nobel Lecture: Quantum chemical models. Reviews of Modern Physics, 71(5):1267–1274, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1267;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1267_1.

Popovic:2003:ASL

[Pop03] Milan Popović, editor. In Albert’s shadow: the life and letters of Mileva Marić, Einstein’s first wife. The Johns Hopkins University Press, Baltimore, MD, USA, 2003. ISBN 0-8018-7856-X (hardcover). xiii + 182 pp. LCCN QC16.E5213

A4 2003. URL http://www.loc.gov/catdir/bios/jhu051/
2003010638.html;   http://www.loc.gov/catdir/description/
jhu051/2003010638.html;   http://www.loc.gov/catdir/
toc/ecip043/2003010638.html.

Popovic:2005:JPP

[Pop05] Milan Popović. Jedno prijateljstvo: pisma Mileve i Alberta Ajnštajna Heleni Savić. (Serbian) [A Friendship: Letters from Mileva and Albert Einstein to Helene Savić]. Plato, Beograd, Serbian, 2005. ISBN 86-447-0280-7. 324 pp. LCCN ???? Marić’s original German letters with parallel Serbian translation.

Popper:2007:LDS

[Pop07] Karl Raimund Popper. La logique de la découverte scientifique. (French) [The logic of scientific discovery]. Bibliothèque scientifique Payot. Payot, Paris, France, 2007. ISBN 2-228-90201-2 (paperback). 480 pp. LCCN Q175. Translation to French by Nicole Thyssen-Rutten Philippe Devaux. Preface by Jacques Monod. In an appendix, an unpublished letter from Albert Einstein to the author, 11 September 1935.

Porter:1981:SSG

[Por81] Theodore M. Porter. A statistical survey of gases: Maxwell’s social physics. Historical Studies in the Physical Sciences, 12(1): 77–116, ???? 1981. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757490.

Posey:2011:WAS

[Pos11] Carl Posey. Wendover’s atomic secret: How B-29 crews trained to drop the bomb. Air & Space Smithsonian, ?? (??):??, March 2011. ISSN 0886-2257 (print), 1942-6488

(electronic). URL http://www.airspacemag.com/military-
aviation/wendovers-atomic-secret-78620832/.

Potenza:2011:DNG

[Pot11] Marco Potenza. Dio non gioca a dadi: l’eredità di Einstein; spazio, tempo e materia. (Italian) [God does not play dice: the legacy of Einstein, space, time and matter]. U. Hoepli, Milano, Italy, 2011. ISBN 88-203-4702-4. x + 205 pp. LCCN ????

Poundstone:1993:PDJ

[Pou93] William Poundstone. Prisoner’s dilemma: John von Neumann, game theory, and the puzzle of the bomb. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-286162-X (paperback), 0-385-41580-X. xi + 294 + 8 pp. LCCN QA29.V66 P68 1993. US$12.95 (US$15.95 Can.).

Powers:1993:HWS

[Pow93] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Knopf, New York, NY, USA, 1993. ISBN 0-394-51411-4. xi + 607 pp. LCCN QC16.H35 P69 1993. US$27.50, CAN$34.50.

Powell:1994:SCN

[Pow94a] Corey S. Powell. Science and the citizen: Nobel notes. Scientific American, 270(2):16–20, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v270/n2/
pdf/scientificamerican0294-16.pdf.

Powers:1994:HWS

[Pow94b] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Back Bay Books, Boston, MA, USA, 1994.

ISBN 0-316-71623-5. xi + 607 pp. LCCN QC16.H35 P69 1994. US$16.95.

Powers:2000:HWS

[Pow00] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 2000. ISBN 0-306-81011-5 (paperback). xi + 607 pp. LCCN QC16.H35 P69 2000.

Powell:2002:GEH

[Pow02] Corey S. Powell. God in the equation: how Einstein became the prophet of the new religious era. Free Press, New York, NY, USA, 2002. ISBN 0-684-86348-0 (hardcover), 0-684-86349-9. 277 (est.) pp. LCCN BL240.3 .P69 2002.

Powell:2015:FPC

[Pow15] Corey S. Powell. Fundamental physics: Cleaning up after Einstein. Scientific American, 313(3):60–67, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-60.html;
 http://www.nature.com/scientificamerican/journal/v313/
n3/pdf/scientificamerican0915-60.pdf.

Plesch:1995:SRA

[PP95] Janos Plesch and Peter H. Plesch. Some reminiscences of Albert Einstein. Notes and Records of the Royal Society of London, 49 (2):303–328, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532017.

Priwer:2003:EEB

[PP03] Shana Priwer and Cynthia Phillips. The everything Einstein book: from matter and energy to space and time, all you need to understand the man and his theories. The everything series. Adams Media Corporation, Avon, MA, USA, 2003. ISBN 1-58062-961-X. xii + 289 pp. LCCN QC16.E5

P75 2003. URL http://www.loc.gov/catdir/toc/ecip042/
2003009085.html.

Phillips:2005:TYD

[PP05] Cynthia Phillips and Shana Priwer. 101 Things You Didn’t Know about Einstein: Sex, Science, and the Secrets of the Universe. F+W Publications Company, Avon, MA, USA, 2005.

ISBN 1-59337-388-0 (paperback), 0-7607-9272-0 (hardcover). x + 244 pp. LCCN QC16.E5P45 2005.

Priwer:2006:EEE

[PP06] Shana Priwer and Cynthia Phillips. Essential Einstein: everything you need to know about the world’s most acclaimed genius. Adams Media, Avon, MA, USA, 2006. ISBN 1-59869-071-X (paperback). xii + 289 pp. LCCN QC16.E5 P75 2006.

Phillips:2007:EEE

[PP07] Cynthia Phillips and Shana Priwer. Essential Einstein: everything you need to know about the world’s most acclaimed genius. David and Charles, Newton Abbot, UK, 2007. ISBN 0-7153-2736-4 (paperback). xii + 289 pp. LCCN ???? Reprint of [PP06].

Priwer:2005:TSE

[PPC05] Shana Priwer, Cynthia Phillips, and Pedro Crespo. Todo sobre Einstein. (Spanish) [Everything about Einstein]. Ediciones Robinbook, Barcelona, España, 2005. ISBN 84-96222-35-7. 316 pp. LCCN ???? Spanish translation by Pedro Crespo of [PP03].

Penrose:1984:SST

[PR86] Roger Penrose and Wolfgang Rindler. Spinors and space– time. Cambridge monographs on mathematical physics. Cambridge University Press, Cambridge, UK, 1984/1986. ISBN 0-521-24527-3 (vol. 1.), 0-521-25267-9 (vol. 2). x + 458 pp. LCCN QC20.7.S65 P46 1984. URL http://www.loc.

gov/catdir/description/cam022/82019861.html;   http:
//www.loc.gov/catdir/samples/cam031/82019861.html;
http://www.loc.gov/catdir/toc/cam022/82019861.html.

Peruzzi:2011:QSU

[PR11] Giulio Peruzzi and Matteo Realdi. The quest for the size of the universe in early relativistic cosmology (1917–1930). Archive for History of Exact Sciences, 65(6):659–689, November 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=65&issue=6&spage=659.

Pratt:2006:ELP

[Pra06] Simon M. Pratt. Einstein’s landmark paper — 100 years on: 100 years after publication, why is Einstein’s “The Electrodynamic moving body” still so influential? Report, Thomson Scientific, New York, NY, USA, November 2006. 8

pp. URL http://ip-science.thomsonreuters.com/m/pdfs/
klnl/2006-11/einstein.pdf.

Pratt:2012:WTE

[Pra12] David Pratt. 1,000 wisest things ever said: wisdom of the Nobel Prize winners. Biteback Publishing [Robson Press], London, UK, 2012. ISBN 1-84954-398-4. xv + 270 pp. LCCN AS911.N9 P738 2012.

Preston:2005:BFM

[Pre05] Diana Preston. Before the fallout: from Marie Curie to Hiroshima. Berkley Books, New York, NY, USA, 2005. ISBN 0-425-20789-7. xiv + 400 pp. LCCN ????

Price:1982:GRP

[Pri82] Richard H. Price. General relativity primer. American Journal of Physics, 50(4):300–329, 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/scitation.aip.org/content/aapt/journal/ajp/50/4/10.
1119/1.12853.

Pribic:1990:NLL

[Pri90] Rado Pribić, editor. Nobel laureates in literature: a biographical dictionary, volume 849 of Garland reference library of the humanities. Garland Publishing, New York, NY, USA, 1990. ISBN 0-8240-5741-4. xxiv + 473 pp. LCCN ????

Price:1999:BRN

[Pri99] Richard H. Price. Book review: Nick Huggett, Space from Zeno to Einstein. Cambridge, Mass.: MIT Press, 1999. xi + 274 pages. $45.00 (cloth), $22.50 (paper). Physics in Perspective (PIP), 1(4):457–458, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pritscher:2010:EZL

[Pri10] Conrad P. Pritscher. Einstein and Zen: learning to learn, volume 384 of Counterpoints: studies in the postmodern theory of education, 1058-1634. Peter Lang, New York, NY, USA, 2010. ISBN 1-4331-0871-2 (hardcover), 1-4331-0870-4 (paperback). xiv + 242 pp. LCCN LB875.E562 P74 2010.

Price:2012:LFT

[Pri12a] Katy Price. Loving faster than light: romance and readers in Einstein’s universe. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-68073-8 (hardcover), 0-226-68075-4 (e-book). xi + 261 pp. LCCN PR478.S26 P75 2012.

Priestley:2012:MRN

[Pri12b] Rebecca Priestley. Mad on radium: New Zealand in the Atomic Age. Auckland University Press, Auckland, NZ, 2012. ISBN 1-86940-727-X (paperback), 1-77558-586-7 (e-book). xii + 284 pp. LCCN TK9122.5 .P75 2012.

Prokhovnik:1974:DEP

[Pro74] S. J. Prokhovnik. Did Einstein’s programme supersede Lorentz’s? The British Journal for the Philosophy of Science, 25(4):336–340, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/25/4/336.full.pdf+html;
http://links.jstor.org/sici?sici=0007-0882%28197412%
2925%3A4%3C336%3ADEPSL%3E2.0.CO%3B2-U.

Przibram:1963:SPE

[Prz63] K. Przibram, editor. Schrödinger–Planck–Einstein–Lorentz: Briefe zur Wellenmechanik. (German) [Schrödinger–Planck– Einstein–Lorentz: Letters on Wave Mechanics]. Herausgegeben im Auftrage der österreichischen Akademie der Wissenschaften von K. Przibram. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1963. vi + 68 pp.

Przibram:1967:LWM

[Prz67] Karl Przibram. Letters on Wave Mechanics: Schrödinger, Planck, Einstein, Lorentz. Philosophical Library, New York, NY, USA, 1967. ISBN 0-85478-271-0. xv + 75 pp. LCCN QC174.2 .P73 1968. Translation by Martin J. Klein of the German original [Prz63].

Przibram:1986:LWM

[Prz86] Karl Przibram, editor. Letters on wave mechanics: Albert Einstein, Erwin Schrödinger, Max Planck and Hendrik A. Lorentz. Philosophical Library, New York, NY, USA, 1986. ISBN ???? ???? pp. LCCN ????

Przibram:2011:AEL

[Prz11] Karl Przibram, editor. Albert Einstein: Letters on wave mechanics: correspondence with H. A. Lorentz, Max Planck, and Erwin Schrödinger. Open Road Integrated Media, New York, NY, USA, 2011. ISBN 1-4532-0468-7. xv + 101 pp. LCCN QC174.2.

Perlmutter:1979:PAE

[PS79] Arnold Perlmutter and Linda F. Scott, editors. On the Path of Albert Einstein, volume 15 of Studies in the Natural Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178

.O72 1979. URL http://www.springerlink.com/content/
978-1-4684-3596-2.

Peskin:1995:IQF

[PS95] Michael Edward Peskin and Daniel V. Schroeder. An introduction to quantum field theory. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-50397-2, 0-201-50934-2 (paperback). xxii + 842 pp. LCCN QC174.45 .P465 1995b; QC174.45 .P465 1995; QC174.45 .P48 1995.

Pitts:2003:NFR

[PS03] J. Brian Pitts and W. C. Schieve. Nonsingularity of flat Robertson–Walker models in the special relativistic approach to Einstein’s equations. Foundations of Physics, 33(9):1315– 1321, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A:1025641327666.

Pitts:2004:NCE

[PS04] J. Brian Pitts and W. C. Schieve. Null cones and Einstein’s equations in Minkowski spacetime. Foundations of Physics, 34 (2):211–238, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B:FOOP.0000019582.44548.6a.

Perez:2010:EQT

[PS10] Enric Pérez and Tilman Sauer. Einstein’s quantum theory of the monatomic ideal gas: non-statistical arguments for a new statistics. Archive for History of Exact Sciences, 64(5):561–612, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Parikh:2016:GEE

[PS16] Maulik Parikh and Sudipta Sarkar. Generalized Einstein’s equations from Wald entropy. Entropy, 18(4):119, ???? 2016. CO-DEN ENTRFG. ISSN 1099-4300. URL https://dblp.org/

db/journals/entropy/entropy18.html#ParikhS16.

Polemis:2022:WSD

[PS22] Michael L. Polemis and Thanasis Stengos. What shapes the delay in the Nobel Prize discoveries? a research note. Scientometrics, 127(2):803–811, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://

link.springer.com/article/10.1007/s11192-021-04241-1.

Pycior:1996:CCS

[PSAA96] Helena M. (Helena Mary) Pycior, Nancy G. Slack, and Pnina G. Abir-Am, editors. Creative couples in the sciences. Lives of women in science. Rutgers University Press, New Brunswick, NJ, USA, 1996. ISBN 0-8135-2187-4, 0-8135-2188-2 (paperback). xi + 369 + 16 pp. LCCN Q141 .C68 1996.

Penrose:1997:LSH

[PSC^{+}97] Roger Penrose, Abner Shimony, Nancy Cartwright, M. S. Longair, and S. W. (Stephen W.) Hawking. The large, the small, and the human mind. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56330-5 (hardcover). xviii + 185 pp. LCCN Q335 .L36 1997.

Plumhof:2013:LRT

[PSM^{+}13] Johannes D. Plumhof, Thilo Stöferle, Lijian Mai, Ullrich Scherf, and Rainer F. Mahrt. Letter: Room-temperature Bose–Einstein condensation of cavity exciton polaritons in a polymer. Nature Materials, ??(??):??, 2013. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (electronic). URL http://www.

nature.com/nmat/journal/vaop/ncurrent/full/nmat3825.

html. See [Joh13] for the news story.

Pulte:2006:SBH

[Pul06] Helmut Pulte. The space between Helmholtz and Einstein: Moritz Schlick on spatial intuition and the foundations of geometry. In Hendricks et al. [HHJ^{+}06], pages 185–206. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (ebook). ISSN 0068-0346. LCCN Q175 .B69 v.251 2006;

Q175 .I51377 2006. URL http://www.loc.gov/catdir/
enhancements/fy0825/2007425756-d.html;   http://www.
loc.gov/catdir/toc/fy0713/2007425756.html.

Pullin:2011:BRB

[Pul11] Jorge Pullin. Book review: Numerical Relativity: Solving Einstein’s Equations on the Computer, by Thomas W. Baumgarte and Stuart L. Shapiro. Physics Today, 64(2):49–50, February 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.org/resource/
1/phtoad/v64/i2/p49_s2. See [BS10].

Purcell:1963:BKSd

[Pur63] John Francis Purcell. The best-kept secret; the story of the atomic bomb. Vanguard Press, New York, 1963. 188 pp. LCCN QC773.A1 P8.

Pusey:1991:CS

[Pus91] P. M. Pusey. Colloidal suspensions. In Hansen et al. [HLZJ91], pages 767–942. ISBN 0-444-88927-2 (part I), 0-444-88928-0 (part II). LCCN QC138 .E36 1989.

Putnam:1925:VET

[Put25] William Lowell Putnam. Velocities in the Einstein theory. American Mathematical Monthly, 32(9):450–455, November 1925. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Puttaswamaiah:2009:FMP

[Put09a] K. Puttaswamaiah, editor. Franco Modigliani: peerless Twentieth-Century macroeconomist: Nobel laureate in economics. Isle Pub., Enfield, NH, USA, 2009. ISBN 0-9823895-3-1. vi + 116 pp. LCCN HB109.M63 F73 2009.

Puttaswamaiah:2009:MFN

[Put09b] K. Puttaswamaiah, editor. Milton Friedman: Nobel monetary economist: a review of his theories and policies. Isle Pub. Co., Enfield, NH, 2009. ISBN 0-9823895-0-7. xiv + 285 pp. LCCN HB119.F84 M557 2009.

Pauli:1949:IRR

[PV49] W. Pauli and F. Villars. On the invariant regularization in relativistic quantum theory. Reviews of Modern Physics, 21 (3):434–444, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.434;   http://rmp.
aps.org/abstract/RMP/v21/i3/p434_1. Special issue in honor

of Albert Einstein on his 70th birthday.

Pauli:1993:WBB

[PvM93] Wolfgang Pauli and Karl von Meyenn, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band III: 1940–1949. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others. Volume 3: 1940–1949]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-78802-6. ???? pp. LCCN ????

Poisson:2014:GNP

[PW14] Eric Poisson and Clifford M. Will. Gravity: Newtonian, post-Newtonian, relativistic. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03286-5 (hardcover). xiv + 780 pp. LCCN QC173.59.M3 P65 2014.

URL http://www.cambridge.org/us/academic/subjects/
physics/cosmology-relativity-and-gravitation/gravity-
newtonian-post-newtonian-relativistic.

Price:2017:DQS

[PW17] Huw Price and Ken Wharton. Dispelling the quantum spooks: A clue that Einstein missed? In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17], chapter 7, pages 123–137. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346. LCCN

QC173.59.S65 T56 2017. URL https://link.springer.com/
chapter/10.1007/978-3-319-53725-2_7.

Pyenson:1973:GRE

[Pye73] Lewis Robert Pyenson. The Goettingen reception of Einstein’s General Theory of Relativity. Ph.D. thesis, Department of Physics, The Johns Hopkins University, Baltimore, MD, USA, November 1973. xxii + 413 pp. URL http://search.

proquest.com/docview/288263009.

Pyenson:1975:RRG

[Pye75] Lewis Pyenson. La réception de la relativité généralisée : disciplinarité et institutionalisation en physique. (French) [The reception of General Relativity: discipline and institutionalization in physics]. Revue d’Histoire des Sciences, 28(1):61–73, January 1975. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23631814.

Pyenson:1976:EES

[Pye76] Lewis Pyenson. Einstein’s early scientific collaboration. Historical Studies in the Physical Sciences, 7(??):iv + 83–123, ???? 1976. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757352.

Pyenson:1977:HME

[Pye77] Lewis Pyenson. Hermann Minkowski and Einstein’s Special Theory of Relativity. Archive for History of Exact Sciences, 17(1):71–95, May 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=1&spage=71. With an appendix

of Minkowski’s “Funktiontheorie” manuscript.

Pyenson:1980:BRBb

[Pye80] Lewis Pyenson. Book reviews: Einstein. A Centenary Exhibition by Paul Forman; Paul Hanle; Albert Einstein in Berlin 1913–1933 by Christa Kirsten; Hans-Jurgen Treder; Images of Einstein: A Catalogue by Joan Warnow; Einstein 1897–1979. Exhibition by Yehuda Elkana; Adi Ophir; Einstein. A Centenary Volume by A. P. French; Albert Einstein. His Influence on Physics, Philosophy and Politics by P. C. Aichelburg; R. U. Sexl. Isis, 71(2):356–359, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230245.

Pyenson:1982:AEE

[Pye82a] Lewis Pyenson. Audacious enterprise: The Einsteins and electrotechnology in late Nineteenth-Century Munich. Historical Studies in the Physical Sciences, 12(2):373–392, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757500.

Pyenson:1982:RLW

[Pye82b] Lewis Pyenson. Relativity in late Wilhelmian Germany: The appeal to a preestablished harmony between mathematics and physics. Archive for History of Exact Sciences, 27(2):137– 155, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=2&spage=137.

Pyenson:1985:YEA

[Pye85] Lewis Pyenson. The young Einstein: the advent of relativity. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-779-9. xiv + 255 + 9 pp. LCCN QC16.E5 P93 1985. URL http://www.

loc.gov/catdir/enhancements/fy0745/85120280-d.html.

Pyenson:1986:BRBa

[Pye86] Lewis Pyenson. Book review: El secreto atomico de Huemul: Cronica del origen de la energia atomica en la Argentina by Mario Mariscotti. Isis, 77(2):389–391, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232723.

Pyenson:1989:BRJ

[Pye89] Lewis Pyenson. Book reviews: Just the facts: The Collected Papers of Albert Einstein, by Albert Einstein; John Stachel; David C. Cassidy; Robert Schulmann; Jürgen Renn; Olga Griminger; Gary Smith; Robert Summerfield; The Collected Papers of Albert Einstein. Volume I: The Early Years, 1879–1902. English Translation by Albert Einstein; Anna Beck; Peter Havas. Volume I: The Early Years, 1879–1902. Isis, 80(1):129–135, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234378.

Pyenson:1990:END

[Pye90] Lewis Pyenson. Einstein’s natural daughter. History of Science (UK), 28(4):365–379, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/28/4/365.full.pdf+html.

Pyenson:1993:BRK

[Pye93] Lewis Pyenson. Book review: Klaus Hentschel: Interpretationen und Fehlinterpretationen der speziellen und der allgemeinen Relativitätstheorie durch Zeitgenossen Albert Einsteins. Isis, 84(2):404–405, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236293.

Pyenson:1995:BRA

[Pye95] Lewis Pyenson. Book reviews: Albert Einstein, Martin J. Klein, A. J. Kox, Jurgen Renn, Jed Buchwald, Jean Eisenstaedt, Don Howard, John Norton, Tilman Sauer: The Collected Papers of Albert Einstein. Volume 3: The Swiss Years: Writings, 1909– 1911. Albert Einstein and Anna Beck: The Collected Papers of Albert Einstein. Volume 3: The Swiss Years: Writings, 1909– 1911. Albert Einstein, Martin J. Klein, A. J. Kox, Robert Schulmann, Paolo Brenni, Klaus Hentschel, Jurgen Renn, and Laura Ruetsche:The Collected Papers of Albert Einstein. Volume 5: The Swiss Years: Correspondence, 1902–1914. Isis, 86(3):514– 515, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
235083.

Pyenson:1997:BRA

[Pye97] Lewis Pyenson. Book review: Albert Einstein, A. J. Kox, Martin J. Klein, Robert Schulmann, József Illy, Jean Eisenstaedt, Rita Fountain, and Annette Pringle: The Collected Papers of Albert Einstein. Volume 6: The Berlin Years: Writings, 1914–1917. Isis, 88(3):562–564, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236206.

Pyenson:2006:BRA

[Pye06] Lewis Pyenson. Book reviews: Albert Einstein: The Collected Papers of Albert Einstein, edited by Michel Janssen, Robert Schulmann, József Illy, Christoph Lehner, Diana Kormos Buchwald, Daniel Kennefick, A. J. Kox, David Rowe, R. Hirschmann, O. Moses, A. Mynttinen, A. Pringle, and R. Fountain. Albert Einstein. The Collected Papers of Albert Einstein: Volume 7: The Berlin Years: Writings, 1918–1921. Isis, 97(4):766–767, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
pdfplus/10.1086/512897.pdf.

Pyenson:2010:BRC

[Pye10] Lewis Pyenson. Book review: Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics by Fulvio Melia. Journal of Interdisciplinary History, 41(2):274–276, Autumn 2010. CODEN ???? ISSN 0022-1953 (print), 1530-9169 (electronic).

URL http://muse.jhu.edu/article/391013.

Pyenson:2013:EPQ

[Pye13] Lewis Pyenson. The Einstein–Picasso question: Neo-idealist abstraction in the decorative arts and manufactures. Historical Studies in the Natural Sciences, 43(3):281–333, June 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X

(electronic). URL http://www.jstor.org/stable/pdfplus/
10.1525/hsns.2013.43.3.281.pdf.

Quinn:1996:MC

[QM96] Susan Quinn and Laurent Muhleisen. Marie Curie. O. Jacob, Paris, France, 1996. ISBN 2-7381-0388-X. 483 pp. LCCN

???? URL http://opac.nebis.ch/cgi-bin/showAbstract.
pl?u20=273810388X.

Quinn:2008:MMA

[QN08] Helen R. Quinn and Yossi Nir. The Mystery of the Missing Antimatter. Science essentials. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13309-3. xii + 278 pp. LCCN

QC173.3 .Q856 2008. URL http://www.loc.gov/catdir/
enhancements/fy0835/2007934402-b.html;   http://www.
loc.gov/catdir/enhancements/fy0835/2007934402-d.html;
 http://www.loc.gov/catdir/enhancements/fy0835/2007934402-
t.html.

Quinn:2001:MCE

[QP01] Susan Quinn and Kimmo Pietiläinen. Marie Curie: elämä. Terra cognita, Helsinki, Finland, 2001. ISBN 952-5202-52-6. 534 + 16 pp. LCCN ????

Quinn:2013:IDU

[QPSD13] Terry Quinn, Harold Parks, Clive Speake, and Richard Davis. Improved determination of G using two methods. Physical Review Letters, 111:101102, September 5, 2013. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.111.101102.

Quill:1938:TE

[Qui38] Laurence L. Quill. The transuranium elements. Chemical Reviews, 23(1):87–155, August 1938. CODEN CHREAY. ISSN 0009-2665 (print), 1520-6890 (electronic). URL http://pubs.

acs.org/doi/abs/10.1021/cr60074a007.

Quinn:1940:BRE

[Qui40] R. B. Quinn. Book review: The Evolution of Physics, by Albert Einstein and Leopold Infeld. Popular Astronomy, 48:

286, December 1940. URL http://adsabs.harvard.edu/abs/
1940PA.....48..286E;   http://adsabs.harvard.edu/abs/
1940PA.....48..286Q.

Quinn:1995:MCL

[Qui95] Susan Quinn. Marie Curie: a life. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-671-67542-7 (hardcover). 509 pp. LCCN QD22.C8 Q56 1995.

Quinn:1996:MCL

[Qui96] Susan Quinn. Marie Curie: a life. Radcliffe biography series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-88794-0 (paperback). 509 pp. LCCN QD22.C8 Q56

1996. URL http://www.loc.gov/catdir/enhancements/
fy0831/96000167-b.html;   http://www.loc.gov/catdir/
enhancements/fy0831/96000167-d.html.

Quinn:1998:MCV

[Qui98] Susan Quinn. Marie Curie: una vita. Le Vite / [Bollati Boringhieri]. Bollati Boringhieri, Torino, Italy, 1998. ISBN 88-339-1116-0. 547 pp. LCCN ????

Quinn:2005:NCE

[Qui05] Terry Quinn, FRS. Newton’s calendar, Einstein and 340 years of Philosophical Transactions. Notes and Records of the Royal Society of London, 59(3):219–222, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041500.

Rhodes:2000:JPB

[RA00] Richard Rhodes and Piotr Tø. Amsterdamski. Jak powstała bomba atomowa. (Polish) [The Making of the Atomic Bomb]. Prószyński i S-ka, Warszawa, Poland, 2000. ISBN 83-7255-131-6. 781 + 42 pp.

Ruiz:2001:DHC

[RA01] Rosaura Ruiz and Francisco J. Ayala. Darwinism: Its hard core. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 15, pages 239–261. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-010-0602-6_15.

Rauscher:2011:OMP

[RA11] Elizabeth A. Rauscher and Richard L. Amoroso. Orbiting the moons of Pluto: complex solutions to the Einstein, Maxwell, Schrödinger, and Dirac equations, volume 45 of Series on knots and everything. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4324-24-8 (hardcover), 981-4324-25-6 (e-book). xvii + 391 pp. LCCN QC173.55 .R34 2011.

Reif-Acherman:2013:HKO

[RA13] Simón Reif-Acherman. Heike Kamerlingh Onnes and the Nobel Prize in Physics for 1913: The highest honor for the lowest temperatures. Physics in Perspective (PIP), 15(4):415–450, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-013-0118-0.

Rabinowitch:1963:DNA

[Rab63] Eugene Rabinowitch. The dawn of a new age: reflections on science and human affairs. University of Chicago Press, Chicago, IL, USA and London, UK, 1963. viii + 332 pp. LCCN CB151 .R25.

Rabi:1970:SCC

[Rab70] Isidor I. Rabi. Science: the center of culture. World Publishing Company, New York, NY, USA, 1970. xx + 155 pp. LCCN Q175.

Rabi:19xx:RPB

[Rabxx] I. I. (Isidor Isaac) Rabi. I. I. Rabi papers, 1899–1989 (bulk 1945–1968). ????, ????, 19xx. ???? pp. LCCN 0639F; LCA (Unprocessed ac. 22,752); Vault 0202A; Oversize 6:10. URL

http://hdl.loc.gov/loc.mss/eadmss.ms998009.

Racah:1949:DTC

[Rac49] Giulio Racah. On the decomposition of tensors by contraction. Reviews of Modern Physics, 21(3):494–496, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.494;   http://rmp.aps.org/

abstract/RMP/v21/i3/p494_1. Special issue in honor of Albert Einstein on his 70th birthday.

Radcliff:2008:AMP

[Rad08] Matthew Radcliff. Absent-minded professor or romantic artist? The depiction of creativity in documentary biographies of Albert Einstein. Journal of Popular Film and Television, 36(2):62–71, Summer 2008. ISSN 0195-6051.

Rafelski:2017:RMEb

[Raf17] Johann Rafelski. Relativity Matters: From Einstein’s E = mc^{2} to Laser Particle Acceleration and Quark–Gluon Plasma. SpringerLink: Bücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-51231-5. xxv + 468 + 85 pp. LCCN QC178 QC173.5-173.65.

Rainich:1946:RPR

[Rai46] G. Y. Rainich. Recent publications: Reviews: The Meaning of Relativity, by Albert Einstein. American Mathematical Monthly, 53(2):93–94, 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL 2-Y; http://links.jstor.

org/sici?sici=0002-9890(194602)53:2<93:TMOR>2.0.CO.

Raine:1975:AER

[Rai75] Derek J. Raine. Albert Einstein and relativity. Pioneers of science and discovery. Priory Press, London, UK, 1975. ISBN 0-85078-209-0. 96 pp. LCCN QC16.E5 R34.

Rainwater:2014:MEW

[Rai14] C. Rainwater. “Maybe Einstein was part Yaqui: Deposing thought in works by Endrezze and Silko. Studies in American Indian Literatures, 26(1):1–28, ???? 2014. CODEN ???? ISSN 0730-3238 (print), 1548-9590 (electronic). URL http://

muse.jhu.edu/article/539872..

Raman:1922:EAE

[Ram22] C. V. Raman. Einstein’s aberration experiment. Nature, 109 (2737):477–478, April 15, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v109/n2737/pdf/109477b0.pdf.

Raman:1979:BRBc

[Ram79a] V. V. Raman. Book review: Albert Einstein, 1879–1955: A centenary exhibit of manuscripts, books, and portraits, selected from the Humanities Research Center Collections, by Albert C. Lewis. American Journal of Physics, 47(12):1109, December 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/?AJP/47/
1109/1.

Raman:1979:BRBa

[Ram79b] V. V. Raman. Book review: Albert Einstein Autobigraphical Notes, A Centennial Edition, edited by Paul Arthur Schilpp. American Journal of Physics, 47(12):1107, December 1979. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/47/1107/1.

Raman:1979:BRBb

[Ram79c] V. V. Raman. Book review: Albert Einstein: The Human Side, edited by Helen Dukas and Banesh Hoffmann. American Journal of Physics, 47(12):1107, December 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/47/1107/2.

Raman:1981:BRB

[Ram81] V. V. Raman. Book review: Relativity, Quanta, and Cosmology in the Development of the Scientific Thought of Albert Einstein, edited by Francesco de Finis. American Journal of Physics, 49 (11):1082–1084, November 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/49/1082/1.

Ramunni:1982:AOB

[Ram82] Girolamo Ramunni. Analyses d’ouvrages: Pioneers of Science, Nobel Prize Winners in Physics par Robert L. Weber; J. M. A. Lenihan. Revue d’Histoire des Sciences, 35(4):361, October 1982. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23632191.

Ramunni:1983:AOBa

[Ram83] Girolamo Ramunni. Analyses d’ouvrages: Albert Einstein’s Special Theory of Relativity. Emergence (1905) and Early Interpretation (1905–1911) par Arthur I. Miller. Revue d’Histoire des Sciences, 36(2):191–192, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632381.

Raman:1990:BRB

[Ram90] Varadaraja V. Raman. Book review: The Collected Papers of Albert Einstein. Vol. 2. The Swiss Years: Writings, 1900– 1909, edited by John Stachel. American Journal of Physics, 58(7):700–701, July 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/?AJP/58/700/1.

Randall:2005:WPB

[Ran05] Lisa Randall. Warped passages: brane-worlds, particles, strings, and the Universe’s hidden dimensions. Ecco, New York, NY, USA, 2005. ISBN 0-06-053108-8. xii + 499 pp. LCCN QC6 .R26 2005.

Rasetti:1968:EFF

[Ras68] Franco Rasetti, editor. Enrico Fermi e la fisica italiana. Accademia nazionale dei Lincei. Celebrazioni lincee. Accademia nazionale dei Lincei, Roma, Italia, 1968. 18 pp. LCCN QC16.F4 E5.

Ravin:1997:AEH

[Rav97] James G. Ravin. Albert Einstein and his mentor Max Talmey — the Seventh Charles B Snyder Lecture. Documenta Ophthalmologica, 94(1–2):1–17, ???? 1997. ISSN 0012-4486 (print), 1573-

2622 (electronic). URL http://link.springer.com/article/

10.1007/BF02629677. Presented in part at the tenth annual meeting of the Cogan Ophthalmic History Society, Philadelphia College of Physicians, March 8 and 9, 1997.

Rayski:1971:LCF

[Ray71] Jerzy Rayski. Limitations of the concept of free fields in Einstein’s theory of gravitation. Foundations of Physics, 1(3): 203–209, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708607.

Ray:1987:ER

[Ray87] Christopher Ray. The evolution of Relativity. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-423-4. x + 211 pp. LCCN

QC173.52 .R39 1987. URL http://www.loc.gov/catdir/
enhancements/fy0745/87021137-d.html.

Ray:1990:CCE

[Ray90] Christopher Ray. The cosmological constant: Einstein’s greatest mistake? Studies in History and Philosophy of Science Part A, 21(4):589–604, December 1990. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900346.

Ray:1992:FLA

[Ray92] Christopher Ray. Fundamental laws and ad hoc decisions: a reply to Curry. Studies in History and Philosophy of Science Part A, 23(4):661–664, December 1992. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290017Z. See [Ray90, Cur92].

Ray:1995:TEC

[Ray95] Sitansu Ray. The Tagore–Einstein conversations: Reality and the human world, causality and chance. In Tymieniecka [Tym95], pages xvi + 340. ISBN 0-7923-3373-X. LCCN B3279.H94 A129 vol. 47 B829.5.

Razavy:2011:HQM

[Raz11] Mohsen Razavy. Heisenberg’s quantum mechanics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4304-11-5 (paperback), 981-4304-10-7. xix + 657 pp. LCCN QC174.12 .R39 2011.

Raz:2016:GTN

[Räz16] Tim Räz. Gone till November: A disagreement in Einstein scholarship. In Sauer and Scholl [SS16], chapter 9, pages 179–200. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346.

LCCN Q126.8 .P45 2016. URL https://link.springer.com/
chapter/10.1007/978-3-319-30229-4_9.

Roberts:1981:GLE

[RB81] David H. Roberts and Bernard F. Burke. Gravitational lenses: From Einstein to the double quasar. Technology Review (M.I.T.), 83(2):68–??, February/March 1981. CO-DEN TEREAU. ISSN 0040-1692. URL http://www.

technologyreview.com/magazine/1981/02/pdf/.

Reed:2004:AIH

[RB04] Thomas C. Reed and George Bush. At the abyss: an insider’s history of the Cold War. Presidio Press/Ballantine Books, New York, NY, USA, 2004. ISBN 0-89141-821-0. ix + 368 pp. LCCN D843 .R37 2004. URL http://www.loc.

gov/catdir/bios/random056/2004300901.html;   http://
www.loc.gov/catdir/description/random052/2004300901.
html;   http://www.loc.gov/catdir/enhancements/fy0620/
2004300901-s.html.

Rogers:2005:SVW

[RB05] Everett M. Rogers and Nancy R. Bartlit. Silent voices of World War II: when sons of the Land of Enchantment met sons of the Land of the Rising Sun. Sunstone Press, Santa Fe, NM, USA, 2005. ISBN 0-86534-423-X (hardcover). 348 pp. LCCN

D769.85.N33 R64 2005. URL http://www.loc.gov/catdir/
toc/ecip052/2004025033.html.

Raychaudhuri:2003:GRA

[RBB03] Amal K. Raychaudhuri, Sriranjan Banerji, and Asit Banerjee. General Relativity, Astrophysics, and Cosmology. Astronomy and astrophysics library. Springer-Verlag, Springer-Verlag, 2003. ISBN 0-387-40628-X. xi + 296 pp. LCCN QB461 .R28 2003.

Richter:2008:MST

[RBS^{+}08] Jörg Richter, Tycho Brake, Antje Samiralow, Werner Kiefer, Christian Sichau, Christian Hainzinger, and Lorenz Kloska. Milestones in Science and Technology. Planet Wissen. AKSA, Aarau, Switzerland, 2008. LCCN ????

Reader:2013:WYN

[RC13] Joseph Reader and Charles W. Clark. 1932, a watershed year in nuclear physics. Physics Today, 66(3):44–49, March 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See also [CR12], and corrections and comments [AG13].

Renn:1998:AEA

[RCD98] Jürgen Renn, Giuseppe Castagnetti, and Peter Damerow. Albert Einstein: alte und neue Kontexte in Berlin. (German) [Albert Einstein: old and new contexts in Berlin]. Report, Max-Planck-Institut für Wissenschaftsgeschichte, Berlin, Germany, 1998. 30 pp.

Rayner-Canham:1997:DTS

[RCRC97] Marelene F. Rayner-Canham and Geoffrey Rayner-Canham, editors. A devotion to their science: pioneer women of radioactivity. McGill-Queen’s University Press and Chemical Heritage Foundation, Montréal, Québec, Canada and Philadelphia, PA, USA, 1997. ISBN 0-7735-1608-5 (McGill hardcover), 0-7735-1642-5 (McGill paper), 0-941901-16-5 (Chemical Heritage hardcover), 0-941901-15-7 (Chemical Heritage paper). xiv + 307 pp. LCCN QC15 .R39 1997.

Renn:2005:DLA

[RD05] Jürgen Renn and Peter Damerow, editors. Dokumente eines Lebensweges: zur Ausstellung Albert Einstein. (German) [Documents of a life’s pathway: for the Albert Einstein exhibition]. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2005. ISBN 3-527-40571-2. 560 pp. LCCN ????

Reid:2009:CEP

[RDB^{+}09] M. D. Reid, P. D. Drummond, W. P. Bowen, E. G. Cavalcanti, P. K. Lam, H. A. Bachor, U. L. Andersen, and G. Leuchs. Colloquium: The Einstein–Podolsky–Rosen paradox: From concepts to applications. Reviews of Modern Physics, 81(4): 1727–1751, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1727;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1727_1.

Roati:2008:ALN

[RDF^{+}08] Giacomo Roati, Chiara D’Errico, Leonardo Fallani, Marco Fattori, Chiara Fort, Matteo Zaccanti, Giovanni Modugno, Michele Modugno, Massimo Inguscio, et al. Anderson localization of a non-interacting Bose–Einstein condensate. Nature, 453(7197): 895–898, June 12, 2008. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v453/n7197/full/nature07071.html.

Rodrigues:2007:MFM

[RdO07] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 722 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71292-5 (hardcover), 3-540-71293-3 (e-book). ISSN 0075-8450. xiv + 445 pp. LCCN

QC173.59.S65 R63 2007. URL http://www.loc.gov/catdir/
enhancements/fy0826/2007923174-d.html;   http://www.
loc.gov/catdir/enhancements/fy0826/2007923174-t.html.

Rodrigues:2016:MFM

[RdO16] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 922 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2016. ISBN 3-319-27636-0 (print), 3-319-27637-9 (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). xvi + 587 pp. LCCN QC173.59.S65. URL http://

link.springer.com/10.1007/978-3-319-27637-3.

Renn:2003:RFR

[RDS03] Jürgen Renn, Lindy Divarci, and Petra Schröter, editors. Revisiting the Foundations of Relativistic Physics: Festschrift in Honor of John Stachel, volume 234 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346. xxii + 649 pp. LCCN

Q174 .B67 vol. 234. URL http://www.gbv.de/dms/bowker/
toc/9781402012846.pdf;   http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-d.html;   http://www.
loc.gov/catdir/enhancements/fy0813/2003046757-t.html;
 https://link.springer.com/book/10.1007/978-94-010-
0111-3.

Ritz:1909:GSSb

[RE09] W. Ritz and A. Einstein. Zum gegenwärtigen Stande des Strahlungsproblems. (German) [On the current state of the radiation problem]. Physikalische Zeitschrift, 10(??):323–324, ???? 1909. CODEN PHZTAO. ISSN 0369-982X. See [Ein09d, Rit09b, Fox65].

Russell:1963:AAW

[RE63] Bertrand Russell and Albert Einstein. An appeal for the abolition of war. In Grodzins and Rabinowitch [GR63], pages 539– 541. LCCN D842 .B78. Reprint of [BEI^{+}55b].

Rhodes:1990:SVH

[RE90] Richard Rhodes and Lennart Edberg. Det sista vapnet: hur atombomben kom till. (Swedish) [The Last Weapon: How the Atomic Bomb was Created]. Hammarström and Åberg, Johanneshov, Sweden, 1990. ISBN 91-7638-080-7 (paperback). 714 + 42 pp.

Rechenberg:1986:BRE

[Rec86] Helmut Rechenberg. Book review: Elizabeth Crawford. The Beginning of the Nobel Institution. The Science Prizes, 1901– 1915. Cambridge: Cambridge University Press, 1985. Pp. ix + 281. ISBN 0-521-26584-3. £22.50. British Journal for the History of Science, 19(2):229, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4026612.

Rechenberg:2002:WHS

[Rec02] Helmut Rechenberg, editor. Werner Heisenberg (1901–1976): Schritte in die neue Physik; [Begleitbuch zur Ausstellung zum 100. Geburtstag von Werner Heisenberg]. (German) [Werner Heisenberg (1901–1976): Steps into new physics [Introductory book for the exhibition for the 100th birthday of Werner Heisenberg]], volume 2 of Veröffentlichungen des Universitätsarchivs Leipzig. Sax-Verl., Beucha, Germany, second edition, 2002. ISBN 3-934544-25-8c. 152 pp. LCCN ???? EUR 10.00,

SFR 19.00. URL http://www.gbv.de/du/services/agi/
DADD826AB2E19CF5C1256D2E004F95CC/420000071853.

Recami:2007:ERD

[Rec07] Erasmo Recami. Einstein e il rinnovamento delle scienze. (Italian) [Einstein and the renewal of science]. arxiv.org, ??(??):??,

September 18, 2007. URL http://arxiv.org/abs/0709.2758.

Rechenberg:2009:WHS

[Rec09] Helmut Rechenberg, editor. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis).

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-69221-5 (print), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35 2009.

URL http://ebooks.ciando.com/book/index.cfm/bok\_id/
46829;   http://www.ciando.com/img/books/width167/3540692223\
_k.jpg;   http://www.ciando.com/pictures/bib/3540692223bib\
_t\_1.jpg.

Rechenberg:2010:WHS

[Rec10] Helmut Rechenberg. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die ,,fröhliche Wissenschaft“ (Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)]. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35

R434 2010. URL http://ebooks.ciando.com/book/index.
cfm/bok_id/46829;   http://www.ciando.com/img/books/
width167/3540692223_k.jpg;   http://www.ciando.com/
pictures/bib/3540692223bib_t_1.jpg.

Redman:1926:EDO

[Red26] Leander A. Redman. The Einstein delusion and other essays. A. M. Robertson, San Francisco, CA, USA, 1926. 3 pp. LCCN Q171 .R33.

Redhead:1984:SLS

[Red84a] M. L. G. Redhead. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. British Journal for the History of Science, 17(56):226–227, ???? 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Redhead:1984:BRP

[Red84b] Michael L. G. Redhead. Book review: Physics and its concepts: Abraham Pais, ‘Subtle is the Lord’: the science and the life of Albert Einstein. Oxford: Clarendon Press, 1982. Pp. xvi + 552. ISBN 0-19-853907-X. £15. British Journal for the History of Science, 17(2):226–227, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4026560.

Redniss:2010:RMP

[Red10] Lauren Redniss. Radioactive: Marie and Pierre Curie, a tale of love and fallout. !t Books/Harper Collins, New York, NY, USA, 2010. ISBN 0-06-135132-6. 205 pp. LCCN

QD22.C8 R395 2010. URL http://www.loc.gov/catdir/
enhancements/fy1111/2011281920-b.html;   http://www.
loc.gov/catdir/enhancements/fy1111/2011281920-d.html.

Reeves:1987:EPE

[Ree87] Barbara J. Reeves. Einstein politicized: The early reception of relativity in Italy. In Glick [Gli87], chapter 6, pages 189– 229. ISBN 90-277-2498-9. LCCN Q174 .B67 vol. 103; QC173.5.

US$79.00 (US). URL https://link.springer.com/chapter/
10.1007/978-94-009-3875-5_6.

Rees:2000:JSN

[Ree00] Martin J. Rees. Just six numbers: the deep forces that shape the universe. Basic Books, New York, NY, USA, 2000. ISBN 0-465-03673-2. xi + 195 pp. LCCN QB 981 R434 2000.

Reed:2005:RLM

[Ree05a] B. Cameron Reed. Resource letter MP-1: The Manhattan Project and related nuclear research. American Journal of Physics, 73(9):805–811, September 2005. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/scitation.aip.org/content/aapt/journal/ajp/73/9/10.
1119/1.1949629.

Reeder:2005:SPN

[Ree05b] Carolyn Reeder. The secret project notebook. Los Alamos Historical Society, Los Alamos, NM, USA, 2005. ISBN 0-941232-33-6. 247 pp. LCCN PZ7.R235237 Sec 2005.

Reed:2006:SLV

[Ree06] B. C. Reed. Seeing the light: Visibility of the July ’45 Trinity atomic bomb test from the inner solar system. The Physics Teacher, 44:604–606, December 2006. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL http:

//adsabs.harvard.edu/abs/2006PhTea..44..604R.

Reed:2009:PMP

[Ree09] Bruce Cameron Reed. Physics of the Manhattan Project. Trafford Publishing, Victoria, BC, Canada, 2009. ISBN 1-4269-0079-1, 1-4269-0081-3 (e-book). 182 pp. LCCN QC790 .R44 2009.

Reed:2011:RLM

[Ree11a] B. Cameron Reed. Resource letter MP-2: The Manhattan Project and related nuclear research. American Journal of Physics, 79(2):151–163, February 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://aapt.

scitation.org/doi/full/10.1119/1.3533209.

Reed:2011:PMP

[Ree11b] Bruce Cameron Reed. The physics of the Manhattan Project. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2011. ISBN 3-642-14708-9 (hardcover), 3-642-14709-7 (e-book). xiii + 170 pp. LCCN QC790 .R44 2011. URL http://d-

nb.info/100407090X/04;   http://www.loc.gov/catdir/
enhancements/fy1113/2010937572-d.html;   http://www.
loc.gov/catdir/enhancements/fy1113/2010937572-t.html.

Reed:2011:TVM

[Ree11c] Cameron Reed. From Treasury vault to the Manhattan Project. American Scientist, 99(1):40–47, January/February 2011. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/25766759.

Reed:2014:FMP

[Ree14a] B. Cameron Reed. The feed materials program of the Manhattan Project: A foundational component of the nuclear weapons complex. Physics in Perspective (PIP), 16(4):461–479, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-014-0146-4.

Reed:2014:MP

[Ree14b] B. Cameron Reed. The Manhattan Project. Physica Scripta, 89 (10):108003:1–108003:26, 2014. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). URL http://stacks.iop.

org/1402-4896/89/i=10/a=108003.

Reed:2014:HSM

[Ree14c] Bruce Cameron Reed. The history and science of the Manhattan Project. Undergraduate Lecture Notes in Physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-40296-8 (hardcover), 3-642-40297-6 (ebook). ISSN 2192-4791. xvi + 451 pp. LCCN QC702.7.H42; QC770-798; QC773.A1; QC793.5.H32-793.5.H329.

Reed:2014:PMP

[Ree14d] Bruce Cameron Reed. The physics of the Manhattan Project. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2014. ISBN 3-662-43532-2 (hardcover), 3-662-43533-0 (e-book). xvii + 222 pp. LCCN QC790 .R44 2015.

Reed:2015:NMC

[Ree15a] B. Cameron Reed. Note on the minimum critical mass for a tamped fission bomb core. American Journal of Physics, 83 (11):969–971, November 2015. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Reed:2015:NWR

[Ree15b] B. Cameron Reed. Nuclear weapons at 70: reflections on the context and legacy of the Manhattan Project. Physica Scripta, 90(8):088001:1–088001:20, 2015. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). URL http://

stacks.iop.org/1402-4896/90/i=8/a=088001.

Reed:2015:ABS

[Ree15c] Bruce Cameron Reed. The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). 239 (est.) pp. LCCN QC773.3.U5 R443 2015eb. URL

http://iopscience.iop.org/book/978-1-6270-5991-6.

Reed:2015:BRB

[Ree15d] Cameron Reed. Book review: Building the H Bomb: A Personal History. Kenneth W. Ford. 234 pp. World Scientific, Singapore, 2015. Price $24 (paper). ISBN 978-981-4618-79-3. American Journal of Physics, 83(10):902–903, October 2015. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Reed:2016:EPG

[Ree16a] B. Cameron Reed. Erratum: A physicist’s guide to The Los Alamos Primer (2016 Physica Scripta 91 [https://doi.org/10.1088/0031-8949/91/11/113002] 113002). Physica Scripta, 91(12):129601:1, 2016. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896

(electronic). URL http://stacks.iop.org/1402-4896/91/i=

12/a=129601. See [Ree16b].

Reed:2016:PGA

[Ree16b] B. Cameron Reed. A physicist’s guide to The Los Alamos Primer. Physica Scripta, 91(11):113002:1–113002:30, 2016. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (elec-

tronic). URL http://stacks.iop.org/1402-4896/91/i=11/

a=113002. See erratum [Ree16a].

Reed:2016:RLM

[Ree16c] B. Cameron Reed. Resource letter MP-3: The Manhattan Project and related nuclear research. American Journal of Physics, 84(10):734–745, October 2016. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://aapt.

scitation.org/doi/full/10.1119/1.4961499.

Reed:2017:EPF

[Ree17a] B. Cameron Reed. An examination of the potential fission-bomb weaponizability of nuclides other than ^{235}U and ^{239}Pu. American Journal of Physics, 85(1):38–44, January 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Reed:2017:RAP

[Ree17b] B. Cameron Reed. Revisiting The Los Alamos Primer. Physics Today, 70(9):42–49, September 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Reed:2017:DCH

[Ree17c] Cameron Reed. Book review: Doomed to Cooperate: How American and Russian Scientists Joined Forces to Avert Some of the Greatest Post-Cold War Dangers. Hecker, Siegfried S., Ed. Vol. I. 568 pp., Vol. II. 447 pp. Bathtub Row Press, Los Alamos, New Mexico, 2016. Price: $80 (hardcover). ISBN 978-0-941232-44-9. American Journal of Physics, 85(7):558–559, July 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Reed:2017:BRT

[Ree17d] Cameron Reed. Book review: True Genius. The Life and Work of Richard Garwin, The Most Influential Scientist You’ve Never Heard Of. Shurkin, Joel N. 308 pp. Prometheus Books, Amherst, New York, 2017. Price: $25 (hardcover). ISBN 978-1-63388-223-2. American Journal of Physics, 85(10):802–803, October 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Reed:2020:MPS

[Ree20] Bruce Cameron Reed. Manhattan Project: The Story of the Century. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-45733-84b (hardcover), 3-030-45734-6 (e-book). xiv + 553 pp. LCCN QC773.3.U5 R4436

2020. URL http://www.springerlink.com/content/978-3-
030-45734-1.

Regis:1987:WGE

[Reg87] Edward Regis. Who got Einstein’s office?: eccentricity and genius at the Institute for Advanced Study. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-12065-8. xv + 316 pp. LCCN Q180.U5 R35 1987. US$17.95.

Reichenbach:1920:REA

[Rei20] Hans Reichenbach. Relativitätstheorie und Erkenntnis Apriori. (German) [The Theory of Relativity and Realization Apriori]. Julius Springer, Berlin, Germany, 1920. ???? pp. LCCN ???? See English translation in [RR65].

Reichenbach:1921:EBG

[Rei21] Hans Reichenbach. Die Einsteinsche Bewegungslehre. (German) [Einstein’s theory of motion]. Die Umschau, 25(35):501–505, August 27, 1921.

Reichenbach:1924:ARR

[Rei24] Hans Reichenbach. Axiomatik der relativistischen Raum–Zeit-Lehre. (German) [Axiomatization of the theory of relativistic space–time], volume 72 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1924. ???? pp. English translation in [RR69].

Reichenbach:1927:PRZ

[Rei27] Hans Reichenbach. Die Philosophie der Raum–Zeit-Lehre. (German) [The Philosophy of Space–Time Theory]. Walter de Gruyter, Berlin, Germany, 1927. vi + 380 pp. LCCN ????

Reichenbach:1928:PRZa

[Rei28a] Hans Reichenbach. Die Philosophie der Raum–Zeit-Lehre. (German) [The Philosophy of Space–Time Theory]. Julius Springer, Berlin, Germany, 1928. ???? pp. LCCN ???? English translation in [RF57].

Reichenbach:1928:PRZb

[Rei28b] Hans Reichenbach. Die Philosophie der Raum–Zeit-Lehre. (German) [The Philosophy of Space–Time Theory]. Walter de Gruyter, Berlin, Germany, 1928. vi + 380 pp. LCCN BD632 .R4.

Reiser:1930:AEB

[Rei30] Anton Reiser. Albert Einstein: a Biographical Portrait. Albert and Charles Boni, New York, NY, USA, 1930. xiv + 225 + 8 pp. LCCN ????

Reiser:1931:AEB

[Rei31] Anton Reiser. Albert Einstein: a Biographical Portrait. Thornton Butterworth, London, UK, 1931. 223 pp. LCCN ????

Reichinstein:1934:AEP

[Rei34] David Reichinstein. Albert Einstein: a picture of his life and his conception of the world. Stella Publishing House, Prague, Czechoslovakia, 1934. 255 pp. Translation by M. Juers and D. Sigmund from the German original.

Reichinstein:1941:RGR

[Rei41] D. Reichinstein. Die Religion des Gebildeten (Religion von Giordano Bruno, Spinoza, Kant, Goethe, Tolstoi, Einstein, Gandhi, Vivekananda etc.). Allgemeinverständlich: mit einem Anhang für Fachleute der Naturwissenschaft: Gott, das Geheimnisvolle und die Wissenschaft. (German) [The religion of the educated (religion of Giordano Bruno, Spinoza, Kant, Goethe, Tolstoy, Einstein, Gandhi, Vivekananda, etc.). General course with an appendix for professionals in the science, God, the mystery and the science]. Aristoteles Verlag, Zürich, Switzerland, second edition, 1941. 142 pp. LCCN ????

Reichenbach:1942:CE

[Rei42] Hans Reichenbach. From Copernicus to Einstein. Philosophical library, Alliance Book Corporation, New York, NY, USA, 1942. 123 pp. LCCN QC6 .R3853. Translation by Ralph Bubrich Winn of German original Von Kopernikus bis Einstein.

Reichenbach:1949:PST

[Rei49] Hans Reichenbach. The philosophical significance of the Theory of Relativity. In Schilpp [Sch49b], chapter 10, pages 287–311. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Reither:1950:RVL

[Rei50] William H. Reither. Review: Voices of Liberalism: 2, by Albert Einstein and George D. Stoddard. Philosophy and Phenomenological Research, 11(1):120–122, September 1950. ISSN 0031-

8205. URL 2-C;   http://links.jstor.org/sici?sici=
0031-8205(195009)11:1<120:VOL2>2.0.CO.

Reither:1951:RWS

[Rei51] William H. Reither. Review: The World as I See it, by Albert Einstein. Philosophy and Phenomenological Research, 11(3): 447–448, March 1951. ISSN 0031-8205. URL 2-G; http://

links.jstor.org/sici?sici=0031-8205(195103)11:3<447:
TWAISI>2.0.CO.

Reichenbach:1957:CE

[Rei57] Hans Reichenbach. From Copernicus to Einstein. Wisdom library. Philosophical Library, New York, NY, USA, 1957. 93 pp. LCCN QC173.55 .R4413 1957. Translation by Ralph Bubrich Winn of German original Von Kopernikus bis Einstein.

Reichenbach:1958:PST

[Rei58] Hans Reichenbach. The philosophy of space and time. Dover, New York, NY, USA, 1958. ISBN 0-486-60443-8. xvi + 295 pp. LCCN BD632 .R4; BD632 .R413; BD632 .R27pE; BD632 .R45 1958; BD632 .R413 1958; QA9 .R27pE. Translated by Maria Reichenbach and John Freund. With introductory remarks by Rudolf Carnap.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reichenbach:1980:CE

[Rei80] Hans Reichenbach. From Copernicus to Einstein. Dover, New York, NY, USA, 1980. ISBN 0-486-23940-3 (paperback). 123 pp. LCCN QC173.55 .R4413 1970. URL http://www.loc.gov/

catdir/description/dover032/79055911.html. Translation

by Ralph Bubrich Winn of German original Von Kopernikus bis Einstein.

Reiser:1982:ESM

[Rei82] Stanley Joel Reiser. The emergence of scientific medicine: a view from the bedside. In Bernhard et al. [BCS82], pages 121–134. ISBN 0-08-027939-2. LCCN Q124.6 .N63

1981. US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B978008027939850013X.

Reingold:1986:BRB

[Rei86] Nathan Reingold. Book reviews: The Beginnings of the Nobel Institution: The Science Prizes, 1901–1915 by Elisabeth Crawford; The Nobel Prize by Peter Wilhelm. Isis, 77(1):145– 146, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
232536.

Reichhardt:1996:EPP

[Rei96a] Tony Reichhardt. Einstein paper on the path to riches. Nature, 379(6568):755, February 29, 1996. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v379/n6568/pdf/
379755b0.pdf.

Reines:1996:NPP

[Rei96b] F. Reines. The neutrino: from poltergeist to particle. Reviews of Modern Physics, 68(2):317–327, April 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
68.317;   http://rmp.aps.org/abstract/RMP/v68/i2/p317_
1;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/1995/.

Reid:2001:AE

[Rei01] Struan Reid. Albert Einstein. Groundbreakers. Heinemann Library, Oxford, UK, 2001. ISBN 0-431-10456-5 (paperback), 0-431-10444-1 (casebound). 48 pp.

Reimer:2005:SPR

[Rei05] Albert Reimer, editor. Spacetime physics research trends, volume 248 of Horizons in world physics. Nova Science Publishers, New York, NY, USA, 2005. ISBN 1-59454-322-4. xi + 169 pp. LCCN QC1 .A4114 vol. 248; QC173.59.S65. URL http://www.

loc.gov/catdir/toc/ecip058/2005005502.html.

Reichenbach:2006:DEH

[Rei06] Hans Reichenbach. Defending Einstein: Hans Reichenbach’s Writings on Space, Time and Motion. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85958-1. vi + 216 pp. LCCN QC173.55 .R439 2006. UK£40.00.

URL http://www.cambridge.org/uk/catalogue/catalogue.

asp?isbn=0521859581. Edited and translated by Steven Gimbel.

Reisman:2007:JRN

[Rei07] Arnold Reisman. Jewish refugees from Nazism, Albert Einstein, and the modernization of higher education in Turkey (1933– 1945). Aleph: Historical Studies in Science and Judaism, 7(??): 253–281, ???? 2007. CODEN ???? ISSN 1553-3956 (print),

1565-1525 (electronic). URL https://muse.jhu.edu/pub/3/
article/221073.

Reich:2011:TPU

[Rei11a] Eugenie Samuel Reich. Troubled probe upholds Einstein. Nature, 473(7346):131–132, May 10, 2011. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/news/2011/110510/full/473131a.html.

Reichenbach:2011:DEH

[Rei11b] Hans Reichenbach. Defending Einstein: Hans Reichenbach’s writing on space, time and motion. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-37116-3. ???? pp. LCCN ????

Rennie:1991:INP

[Ren91] John Rennie. The Ig Nobel Prizes. Scientific American, 265(6):26, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v265/n6/
pdf/scientificamerican1291-26.pdf.

Renn:1993:EDG

[Ren93] Jürgen Renn. Einstein as a disciple of Galileo: A comparative study of concept development in physics. Science in Context, 6 (1):311–341, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:1997:ECD

[Ren97a] Jürgen Renn. Einstein’s controversy with Drude and the origin of statistical mechanics: a new glimpse from the “Love Letters”. Archive for History of Exact Sciences, 51(4):315–354, December 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=4&spage=315.

Renn:1997:KTD

[Ren97b] Jürgen Renn. Von der klassischen Trägheit zur dynamischen Raumzeit: Albert Einstein und Ernst Mach. (German) [From classic inertia to dynamic space–time: Albert Einstein and Ernst Mach]. Berichte zur Wissenschaftsgeschichte, 20(2–3): 189–198, ???? 1997. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Renn:2005:EIB

[Ren05a] J. Renn. Einstein’s invention of Brownian motion. Annalen der Physik (Berlin), 14(1):23–37, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200410131.

Renn:2005:AEIb

[Ren05b] Jürgen Renn. Albert Einstein — Ingenieur des Universums. Hundert Autoren für Einstein. Albert Einstein — chief engineer of the universe. One hundred authors for Einstein. History of knowledge. Wiley-VCH, Weinheim, Germany, 2005. ISBN 3-527-40574-7, 3-527-40579-8. 472 pp. LCCN

QC16.E5 A67713 2005. URL http://www.loc.gov/catdir/
enhancements/fy0645/2006530450-b.html;   http://www.
loc.gov/catdir/enhancements/fy0645/2006530450-d.html;
 http://www.loc.gov/catdir/enhancements/fy0645/2006530450-
t.html.

Renn:2005:AEIa

[Ren05c] Jürgen Renn, editor. Albert Einstein: Ingenieur des Universums. Hundert Autoren für Einstein. (German) [Albert Einstein: Chief Engineer of the Universe. One hundred authors for Einstein]. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2005. ISBN 3-527-40569-0, 3-527-40571-2. 254 pp. LCCN QC16.E5 A6759 2005.

Renn:2005:EAP

[Ren05d] Jürgen Renn. Einstein’s Annalen papers: the complete collection 1901–1922. Wiley-VCH, Weinheim, Germany, 2005. ISBN 3-527-40564-X. 585 pp. LCCN QC174.12 .E334 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0626/2005277124-
b.html;   http://www.loc.gov/catdir/enhancements/fy0626/
2005277124-d.html;   http://www.loc.gov/catdir/enhancements/

fy0626/2005277124-t.html. Text in German, with introductory essays in English.

Renn:2005:ELW

[Ren05e] Jürgen Renn, editor. Einsteins Leben und Werk im Kontext zur Ausstellung Albert Einstein — Ingenieur des Universums, die vom Max-Planck-Institut für Wissenschaften anlässlich des Einsteinjahres 2005 entwickelt wurde; [Ausstellung im Kronprinzenpalais, Berlin, vom 16. Mai bis 30. September 2005]. (German) [Einstein’s Life and Work in Context …]. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2005. ISBN 3-527-40573-9. 254 pp. LCCN ????

Renn:2005:IEA

[Ren05f] Jürgen Renn. Introduction [to Einstein’s Annalen papers]. Annalen der Physik (Berlin), 14(1):9–14, February 2005. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
andp.200590049.

Renn:2005:WER

[Ren05g] Jürgen Renn. Wie Einstein die Relativitätstheorie fand. (German) [How Einstein found the Theory of Relativity]. In Steiner [Ste05a], pages 41–78. ISBN 3-540-21060-1. LCCN QC16.E5

A659 2005. URL http://link.springer.com/chapter/10.
1007/3-540-30595-5_2/.

Renn:2006:SRZ

[Ren06] Jürgen Renn. Auf den Schultern von Riesen und Zwergen: Einsteins unvollendete Revolution. (German) [On the Shoulders of Giants and Dwarfs: Einstein’s unfinished revolution], volume 1 of Abenteuer Wissensgeschichte. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40595-X (hardcover). 364 pp.

LCCN ???? URL http://catalogue.bnf.fr/ark:/12148/
cb39976846k.

Renn:2007:CPD

[Ren07a] Jürgen Renn. Classical physics in disarray: The emergence of the riddle of gravitation. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 21–80. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https://link.springer.

com/chapter/10.1007/978-1-4020-4000-9_2.

Renn:2007:P

[Ren07b] Jürgen Renn. Preface. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 1–6. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

Renn:2007:SAS

[Ren07c] Jürgen Renn. The summit almost scaled: Max Abraham as a pioneer of a relativistic theory of gravitation. In Renn et al. [RSS^{+}07a], pages 305–330. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_22.

Renn:2007:TWG

[Ren07d] Jürgen Renn. The third way to General Relativity: Einstein and Mach in context. In Renn et al. [RSS^{+}07a], pages 21–75.

ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https://link.springer.

com/chapter/10.1007/978-1-4020-4000-9_11.

Rentetzi:2008:TMG

[Ren08] Maria Rentetzi. Trafficking materials and gendered experimental practices: radium research in early 20th Century Vienna. Columbia University Press, New York, NY, USA, 2008. ISBN 0-231-13558-0 (hardcover), 0-231-50959-6 (e-book). xxiii + 279 pp. LCCN QC795.34 .R46 2008. URL http://www.loc.gov/

catdir/toc/ecip0828/2008040581.html.

Renn:2013:EMS

[Ren13] Jürgen Renn. Einstein as a missionary of science. Science & Education (Springer), 22(10):2569–2591, October 2013. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Renn:2021:BRS

[Ren21] Jürgen Renn. Book review: Simões, Ana, Sousa, Ana Matilde. The global adventure of science: Einstein, Eddington and the eclipse. Lisbon, Portugal: Chili com Carne, 2019, 245 pp. ISBN: 978-989-8363-41-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):616–617, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reston:1950:UHB

[Res50] James Reston. U.S. hydrogen bomb delay urged pending bid to Soviet: Officials, said to include Lilienthal, appeal to President to seek agreement and thus avert making weapon of terrible power. New York Times, ??(??):1, 11, January 17, 1950. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/

docview/111352790. According to [BO76, page 227], this report, which appears on the top front page, is the first printed public story in the USA about the hydrogen bomb. It followed a radio broadcast by Drew Pearson on 15-Jan-1950 on that subject. However, it was not the first in this newspaper: there were at about 20 earlier stories from 8-Aug-1945 to 10-Jan-1950 that contain “hydrogen” and “bomb”. The story names President Harry S Truman, Atomic Energy Commission (AEC) Chairman David Lilienthal, Major-General Leslie R. Groves, AEC Commissioner Lewis L. Strauss, but names no scientists whatever.

Reynolds:1925:NGA

[Rey25] C. N. Reynolds. Note on the geometric aspects of Einstein’s theory. American Mathematical Monthly, 32(2):74–75, February 1925. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rabinowitch:1955:AE

[RF55] Eugene Rabinowitch and James Franck. Albert Einstein. Bulletin of the Atomic Scientists, 11(6):202–203, June 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reichenbach:1957:PST

[RF57] Maria Reichenbach and John Freund, editors. The Philosophy of Space and Time. Dover, New York, NY, USA, 1957. ???? pp. LCCN ???? English translation of [Rei28a, Rei28b].

Raman:1969:WWI

[RF69] V. V. Raman and Paul Forman. Why was it Schrödinger who developed de Broglie’s ideas? Historical Studies in the Physical Sciences, 1(??):291–314, ???? 1969. CODEN HSPSAS. ISSN

0073-2672. URL http://www.jstor.org/stable/27757299.

Rabinowitch:1979:MAE

[RF79] Eugene Rabinowitch and James Franck. In memoriam [Albert Einstein]. Bulletin of the Atomic Scientists, 35(3):40, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Rifkind:2005:NPW

[RF05] David Rifkind and Geraldine L. Freeman. The Nobel Prize winning discoveries in infectious diseases. Elsevier/Academic Press, Amsterdam, The Netherlands, 2005. ISBN 0-12-369353-5 (paperback), 0-08-045957-9. LCCN RC111 .R54 2005. URL http:

//site.ebrary.com/lib/royallibrary/Doc?id=10138142.

Robert:1945:LTP

[RFP^{+}45] Owen J. Robert, J. W. Fulbright, Claude Pepper, Elbert D. Thomas, the Right Reverent Henry St. George Tucker, Edward A. Conway, S.J., Louis Finkelstein, Mortimer J. Adler, Charles G. Bolte, Gardner Cowles, Jr., Albert Einstein, Dorothy Canfield Fisher, Albert D. Lasker, Thomas Mann, Cord Meyer, Christopher Morley, Carl Van Doren, Mark Van Doren, Walter F. Wanger, and Robert J. Watt. Letters to The Times: Peace by law our one hope: San Francisco Charter must be made to work to avoid another war. New York Times, ??(??):20, October 10, 1945. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/107222982/5497E3A766954244PQ/1.

Renn:2007:PHS

[RG07] Jürgen Renn and Kåostas Gavroglou, editors. Positioning the history of science, volume 248 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5420-3, 1-4020-5419-X. ISSN 0068-0346. vii + 188 pp. LCCN Q125 .P77 2007eb.

Reiprich:2002:WHG

[RH02] Kurt Reiprich and Werner Heisenberg, editors. Werner Heisenberg: zum 100. Geburtstag. (German) [Werner Heisenberg: 100th Birthday]. Sonderheft // Rohrbacher Kreis. Rosa-Luxemburg-Stiftung Sachsen, Leipzig, Germany, 2002. ISBN 3-89819-116-8. 64 pp. LCCN ????

Rhodes:1986:MAB

[Rho86] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 1986. ISBN 0-671-44133-7 (paperback), 0-671-65719-4 (hardcover). 886 + 42 pp. LCCN QC773 .R46 1986.

Rhodes:1988:AGS

[Rho88a] Richard Rhodes. Die Atombombe oder Die Geschichte des 8. Schöpfungstages. (German) [The Atomic Bomb, or The History of the Eighth Day of Creation]. Greno, Nördlingen, Germany, 1988. ISBN 3-89190-522-X. 915 pp.

Rhodes:1988:MABa

[Rho88b] Richard Rhodes. The Making of the Atomic Bomb. Touchstone, New York, NY, USA, 1988. ISBN 0-684-81378-5 (paperback), 0-671-44133-7 (cased). 886 + 42 pp. LCCN QC773 .R46 1988.

Rhodes:1988:MABb

[Rho88c] Richard Rhodes. The making of the atomic bomb. A Penguin book History, archaeology. Penguin, London, UK and New York, NY, USA, 1988. ISBN 0-14-011667-2 (paperback), 0-14-014997-X (hardcover). 886 pp. LCCN QC773 .R46 1986.

Rhodes:1993:TAF

[Rho93] Richard Rhodes. Trademark Los Alamos: the first Soviet bomb. Video recording at Director’s Colloquium, 10 June 1993, Los Alamos National Laboratory., June 10, 1993. Video recording that details how the first Soviet bomb was a copy of the Los Alamos Trinity device based upon Klaus Fuch’s spying intelligence.

Rhodes:1995:DSM

[Rho95] Richard Rhodes. Dark Sun: The Making of the Hydrogen Bomb. Sloan technology series. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-684-80400-X, 0-684-82414-0 (paperback). 731 + 32 pp. LCCN UG1282.A8 R46 1995.

Rhodes:2007:AFM

[Rho07] Richard Rhodes. Arsenals of folly: the making of the nuclear arms race. Alfred A. Knopf, New York, NY, USA, 2007. ISBN 0-375-41413-4. 386 pp. LCCN U264 .R48 2007. URL http://

www.loc.gov/catdir/toc/ecip0716/2007017613.html.

Rhodes:2012:MABa

[Rho12] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 25th anniversary edition, 2012. ISBN 1-4516-7761-8 (paperback), 1-4391-2622-4 (e-book). 838 + 42 pp. LCCN QC773 .R46 2012.

Rich:1983:EMV

[Ric83] Vera Rich. Einstein manuscripts on view. Nature, 301(5897): 190, January 20, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v301/n5897/pdf/301190b0.pdf.

Richards:1993:TAM

[Ric93] Hugh T. (Hugh Taylor) Richards. Through Los Alamos, 1945: memoirs of a nuclear physicist. Arlington Press, Madison, WI, USA, 1993. ISBN 0-9637521-1-1. 95 pp. LCCN QC774.R52 A3

Rich:2001:FC

[Ric01] James Rich. Fundamentals of cosmology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41350-2. xi + 302 pp. LCCN

QB981 .R486 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/2001020334-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/2001020334-t.html.

Richelson:2006:SBA

[Ric06a] Jeffrey Richelson. Spying on the bomb: American nuclear intelligence from Nazi Germany to Iran and North Korea. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-05383-0 (hardcover). 702 pp. LCCN UB271.U5 R53 2006.

Richter:2006:TPP

[Ric06b] Burton Richter. Theory in particle physics: Theological speculation versus practical knowledge. Physics Today, 59(10):8–??, October 11, 2006. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
vol-59/iss-10/p8.html.

Ridenour:1950:HB

[Rid50] Louis N. Ridenour. The hydrogen bomb. Scientific American, 182(3):11–15, March 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v182/n3/
pdf/scientificamerican0350-11.pdf. See also Parts II

[Bet50a] and III [Bac50].

Riess:2012:NLM

[Rie12] Adam G. Riess. Nobel Lecture: My path to the accelerating Universe. Reviews of Modern Physics, 84(3):1165– 1175, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1165;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1165_1;   http:/
/www.nobelprize.org/nobel_prizes/physics/laureates/
2011/.

Rigden:2000:IIR

[Rig00] John S. Rigden. Isidor Isaac Rabi: 29 July 1898–11 January 1988. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 144(1):113–118, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp. LCCN QD181.H1 R54 2002. URL http://catdir.loc.

gov/catdir/toc/fy022/2001051708.html.

Rigden:2005:ESG

[Rig05a] John S. Rigden. Einstein 1905: the standard of greatness. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-02104-5 (paperback), 0-674-04275-1 (e-book). ix + 173 pp. LCCN QC16.E5 R54 2005.

Rigden:2005:ERP

[Rig05b] John S. Rigden. Einstein’s revolutionary paper. Physics World, 18(4):18–19, April 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/4/phwv18i4a25.
pdf.

Rindler:1960:SR

[Rin60] Wolfgang Rindler. Special relativity. University mathematical texts. Oliver and Boyd, Edinburgh, UK; London, UK, 1960. 186 pp. LCCN QC6 .R48.

Rindler:1966:SRC

[Rin66] Wolfgang Rindler. Special relativity, volume 23 of University mathematical texts. Oliver and Boyd, Edinburgh, UK; London, UK, second edition, 1966. xii + 196 pp. LCCN QC6 .R48 1966.

Rindler:1969:ERS

[Rin69] Wolfgang Rindler. Essential relativity; special, general, and cosmological. Van Nostrand Reinhold, New York, NY, USA, 1969. xi + 319 pp. LCCN QC6 .R477.

Rindler:1977:ERS

[Rin77] Wolfgang Rindler. Essential relativity: special, general, and cosmological. Texts and monographs in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1977. ISBN 0-387-07970-X. xiv + 284 pp. LCCN QC173.55 .R56 1977.

Rindler:1982:ISR

[Rin82] Wolfgang Rindler. Introduction to special relativity. Clarendon Press, Oxford, UK, 1982. ISBN 0-19-853182-6 (paperback), 0-19-853181-8 (paperback). x + 185 pp. LCCN QC173.65 .R56 1982. US$12.95; US$29.50.

Rindler:1991:ISR

[Rin91] Wolfgang Rindler. Introduction to special relativity. Clarendon Press, Oxford, UK, second edition, 1991. ISBN 0-19-853953-3, 0-19-853952-5 (paperback). x + 169 pp. LCCN QC173.65 .R56 1991. US$37.50US$18.25. URL http://www.loc.gov/

catdir/enhancements/fy0603/90048748-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0603/90048748-t.
html.

Rindler:2001:RSG

[Rin01] Wolfgang Rindler. Relativity: special, general, and cosmological. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850836-0 (paperback), 0-19-850835-2 (hardcover). xiii + 428 pp. LCCN QC173.55 .R563

2001. URL http://www.loc.gov/catdir/enhancements/
fy0610/00067605-d.html;   http://www.loc.gov/catdir/
enhancements/fy0610/00067605-t.html.

Ringholz:2002:UFS

[Rin02] Raye Carleson Ringholz. Uranium Frenzy: Saga of the Nuclear West. Utah State University Press, Logan, UT, USA, revised and expanded edition, 2002. ISBN 0-87421-432-7 (paperback). xiii + 344 pp. LCCN HD9539.U72 U5366 2002.

URL http://www.loc.gov/catdir/enhancements/fy0809/
2002009421-d.html;   http://www.loc.gov/catdir/toc/
fy033/2002009421.html.

Ring:2003:TST

[Rin03] R. Ring. Toy story — toward a general theory of perpetual motion (a reader describes his personal encounter with Albert Einstein in 1935). American Heritage, 54(4):70, September 2003. ISSN 0002-8738.

Rindler:2006:RSG

[Rin06] Wolfgang Rindler. Relativity: special, general, and cosmological. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2006. ISBN 0-19-856732-4 (paperback), 0-19-856731-6 (hardcover). xv + 430 pp. LCCN

QC173.55 .R563 2006. URL http://www.loc.gov/catdir/
enhancements/fy0725/2006277893-b.html;   http://www.
loc.gov/catdir/enhancements/fy0725/2006277893-d.html;
 http://www.loc.gov/catdir/toc/fy0611/2006277893.html.

Rindler:2009:GEMa

[Rin09] Wolfgang Rindler. Gödel, Einstein, Mach, Gamow, and Lanczos: Gödel’s remarkable excursion into cosmology. American Journal of Physics, 77(6):498–510, June 2009. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://adsabs.harvard.edu/abs/2009AmJPh..77..498R;
http://ajp.aapt.org/resource/1/ajpias/v77/i6/p498_s1.

Reprinted in [BPP^{+}11, Chapter 9].

Rindler:2011:GEMb

[Rin11a] Wolfgang Rindler. Gödel, Einstein, Mach, Gamow, and Lanczos: Gödel’s remarkable excursion into cosmology. In Baaz et al. [BPP^{+}11], pages 185–212. ISBN 0-521-76144-1 (print), 0-511-97423-X (electronic). LCCN QA9.65 .K87 2011.

Rindler:2011:GEM

[Rin11b] Wolfgang Rindler. Gödel, Einstein, Mach, Gamow, and Lanczos: Gödel’s remarkable excursion into cosmology. In Baaz et al. [BPP^{+}11], pages 185–212. ISBN 0-521-76144-1 (print), 0-511-97423-X (electronic). LCCN QA9.65 .K87 2011.

Riordan:1987:HQT

[Rio87] Michael Riordan. The hunting of the quark: a true story of modern physics. A Touchstone book. Simon and Schuster, New York, NY, USA, 1987. ISBN 0-671-50466-5, 0-671-64884-5 (paperback). 399 pp. LCCN QC793.5.Q252 R56 1987.

Rispens:2006:ENI

[Ris06] Sybe Izaak Rispens. Einstein in Nederland: een intellectuele biografie. (Dutch) [Einstein in The Netherlands: an intellectual biography]. Ambo, Amsterdam, The Netherlands, 2006. ISBN 90-263-1903-7. 242 pp. LCCN ????

Ritz:1909:GGG

[Rit09a] Walter Ritz. Die Gravitation. (German) [Gravitation]. Rivista di Scienza “Scientia”, 5(??):241–255, ???? 1909.

Ritz:1909:GSSa

[Rit09b] Walter Ritz. Zum gegenwärtigen Stand des Strahlungsproblems; Erwiderung auf der Aufsatz des Herrn. A Einstein. (German) [On the current state of the radiation problem: response to the essay of Mr. A. Einstein]. Physikalische Zeitschrift, 10(??):224– 225, ???? 1909. CODEN PHZTAO. ISSN 0369-982X.

Renn:2007:GGR

[RJS07] Jürgen Renn, M. (Michael) Janssen, and Matthias Schemmel, editors. The Genesis of General Relativity, volume 250 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. ???? pp. LCCN

QC173.6 .G469 2007. URL http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html.

Roberts:1939:UAP

[RK39] R. B. Roberts and J. B. H. Kuper. Uranium and atomic power. Journal of Applied Physics, 10(9):612–614, September 1, 1939. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://scitation.aip.

org/content/aip/journal/jap/10/9/10.1063/1.1707351.

Reichenbach:1977:PRZ

[RK77] Hans Reichenbach and Andreas Kamlah, editors. Philosophie der Raum–Zeit-Lehre. (German) [The Philosophy of Space– Time Theory], volume 2 of Gesammelte Werke in 9 Bänden / Hans Reichenbach. Hrsg. von Andreas Kamlah …. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1977. ISBN 3-528-08362-X. 442 pp. LCCN ????

Rubel-Kolb:1986:EEE

[RK86] Eduard Rübel-Kolb. Eduard Einstein: Erinnerungen ehemaliger Klassenkameraden am Zürcher Gymnasium. (German) [Eduard Einstein: Memoirs of former classmates at the Zurich School]. P. Haupt, Bern, Switzerland, 1986. ISBN 3-258-03555-5. 119 pp. LCCN CT1398.E39 E38 1986.

Rose-Kobe:1994:REC

[RK94] Almut Rose-Kobe. Rezension: Elisabeth Crawford: Nationalism and internationalism in science, 1880–1939 — Four studies of the Nobel population. Cambridge, Cambridge University Press 1992. 157 Seiten. Berichte zur Wissenschaftsgeschichte, 17(3):216–218, ???? 1994. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Rosenblum:2011:QEP

[RK11] Bruce Rosenblum and Fred Kuttner. Quantum Enigma: Physics Encounters Consciousness. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2011. ISBN 0-19-975381-4 (paperback). x + 287 pp. LCCN QC174.13 .R67 2011.

Roll:1964:EIP

[RKD64] P. G. Roll, R. Krotkov, and R. H. Dicke. The equivalence of inertial and passive gravitational mass. Annals of Physics, 26 (3):442–517, February 20, 1964. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

Richter:2008:MNT

[RKS08] Jörg Richter, Werner Kiefer, and Antje Samiralow. Meilensteine der Naturwissenschaft und Technik. (German) [Milestones of Science and Technology]. Planet Wissen. AKSA, Aarau, Switzerland, 2008. LCCN ????

Rechenberg:2005:WHI

[RKW05] Helmut Rechenberg, Christian Kleint, and Gerald Wiemers, editors. Werner Heisenberg im Spiegel seiner Leipziger Schüler und Kollegen. (German) [Werner Heisenberg as seen by his Leipzig pupils and colleagues]. Leipziger Uni-Vlg, Leipzig, Germany, 2005. ISBN 3-86583-079-X (paperback). 235 pp. LCCN ???? EUR 24.00.

Raner:1992:EB

[RL92] Guy H. Raner and Lawrence S. Lerner. Einstein’s beliefs. Nature, 358(6382):102, July 9, 1992. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v358/n6382/pdf/
358102e0.pdf.

Riess:2016:CPD

[RL16] Adam G. Riess and Mario Livio. Cosmology: The puzzle of dark energy. Scientific American, 314(3):38–43, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-38.html;
 http://www.nature.com/scientificamerican/journal/v314/
n3/pdf/scientificamerican0316-38.pdf.

Radosh:1997:RF

[RM97] Ronald Radosh and Joyce Milton. The Rosenberg file. Yale University Press, New Haven, CT, USA, second edition, 1997. ISBN 0-300-07205-8 (paperback). xxx + 616 pp. LCCN

KF224.R6 R32 1997. URL http://www.h-net.org/review/

hrev-a0a7x6-aa. With a new introduction containing revelations from National Security Agency and Soviet sources.

Rovira:2011:EVR

[RMHD11] Àlex Rovira, Francesc Miralles Contijoch, and Maria Hoffmann-Dartevelle. Einsteins Versprechen: Roman. (German) [Einstein’s Promise: Novel]. List, Berlin, Germany, 2011. ISBN 3-471-35051-9. 382 pp. LCCN ???? EUR 18.00 (DE); EUR 18.50 (AT); SFR 29.90.

RMITU:2021:PUL

[RMI21a] RMIT University. Previously unknown letter reveals Einstein’s thinking on bees, birds and physics: A recently discovered letter written by Albert Einstein discusses a link between physics and biology, seven decades before evidence emerges. Press release, May 13, 2021. URL https://www.rmit.edu.

au/news/media-releases-and-expert-comments/2021/may/
einstein-letter. See [DGG^{+}21].

RMITU:2021:RDL

[RMI21b] RMIT University. Recently discovered letter written by Albert Einstein discusses link between physics and biology seven decades before evidence emerges, June 26, 2021. URL https://

scitechdaily.com/recently-discovered-letter-written-
by-albert-einstein-discusses-link-between-physics-and-
biology-seven-decades-before-evidence-emerges/. See

[DGG^{+}21].

Rodgers:1950:TMR

[RMR^{+}50] Douglas Rodgers, Roger Muir, Eleanor Roosevelt, Albert Einstein, David Eli Lilienthal, J. Robert Oppenheimer, Brien McMahon, Hans A. (Hans Albrecht) Bethe, Harry Winne, Allan Kline, and Detlev W. (Detlev Wulf) Bronk. Today with Mrs. Roosevelt. NBC Television Network, United States, February 12, 1950. LCCN FSA 1920 (arch neg); FSA 1922 (mag trk); FCB 4264 (ref print); FTA 3029 (pos trk); FTA 3028 (masterpos); FCB 4367 (ref print, copy 2); FSA 1921 (neg trk). Seven tape reels.

Rosen:1996:DEP

[RMR96] Nathan Rosen, A. (Ady) Mann, and M. Revzen, editors. The dilemma of Einstein, Podolsky and Rosen, 60 years later: an international symposium in honour of Nathan Rosen — Haifa, March 1995, volume 12 of Annals of the Israel Physical Society, 0309-8710. Institute of Physics Publishers, Bristol, UK, 1996. ISBN 0-7503-0394-8. LCCN QC174.12 .D55 1996.

Robertson:1968:RC

[RN68] H. P. Robertson and Thomas W. Noonan. Relativity and Cosmology. W. B. Saunders Company, Philadelphia, PA, USA, 1968. xxxiii + 456 pp. LCCN QC6 .R635.

Rodriguez-Navarro:2016:RAB

[RN16] Alonso Rodríguez-Navarro. Research assessment based on infrequent achievements: a comparison of the United States and Europe in terms of highly cited papers and Nobel Prizes. Journal of the Association for Information Science and Technology, 67(3):731–740, March 2016. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Robertson:1938:NPP

[Rob38] H. P. Robertson. Note on the preceding paper (“The gravitational equations and the problem of motion” [MR1503389] by A. Einstein, L. Infeld and B. Hoffmann): the two body problem in General Relativity. Annals of Mathematics (2), 39(1): 101–104, 1938. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Robertson:1939:RYAa

[Rob39a] H. P. Robertson. Relativity — 20 years after. Scientific American, 160(6):358–359, June 1939. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v160/n6/
pdf/scientificamerican0639-358.pdf.

Robertson:1939:RYAb

[Rob39b] H. P. Robertson. Relativity — 20 years after. Scientific American, 161(1):22–24, July 1939. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v161/n1/
pdf/scientificamerican0739-22.pdf.

Robson:1948:RSP

[Rob48] J. W. Robson. Review: Spinoza: Portrait of a Spiritual Hero, by Rudolf Kayser, Albert Einstein, Amy Allen, and Maxim Newmark. Philosophy and Phenomenological Research, 8(4): 725–726, June 1948. ISSN 0031-8205. URL 2-L; http:/

/links.jstor.org/sici?sici=0031-8205(194806)8:4<725:
SPOASH>2.0.CO.

Robertson:1949:GBP

[Rob49a] H. P. Robertson. Geometry as a branch of physics. In Schilpp [Sch49b], chapter 11, pages 313–332. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Robertson:1949:PVO

[Rob49b] H. P. Robertson. Postulate versus observation in the Special Theory of Relativity. Reviews of Modern Physics, 21(3):378– 382, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.21.378;   http://rmp.aps.
org/abstract/RMP/v21/i3/p378_1. Special issue in honor of

Albert Einstein on his 70th birthday.

Robinson:1950:ORS

[Rob50] George O. Robinson. The Oak Ridge story; the saga of a people who share in history. Southern Publishers, Kingsport, TN, USA, 1950. 181 pp. LCCN F444.O3 R6.

Robinson:1958:WTH

[Rob58] G. O. (George Oscar) Robinson. And what of tomorrow: the human drama in the atomic revolution and the promise of a golden age. Comet Press Books, New York, NY, USA, second edition, 1958. 178 pp. LCCN QC773.3.U5 R6 1958.

Robinson:1975:REI

[Rob75] Arthur L. Robinson. Relativity: Experiments increase confidence in Einstein. Science, 188(4193):1099–1101, June 13, 1975. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/188/
4193/1099.full.pdf.

Roberts:2001:BUS

[Rob01] Sam Roberts. The brother: the untold story of atomic spy David Greenglass and how he sent his sister, Ethel Rosenberg, to the electric chair. Random House, New York, NY, USA, 2001. ISBN 0-375-50013-8. x + 543 + 16 pp. LCCN

KF224.R6 R63 2001. URL http://www.loc.gov/catdir/
bios/random052/00054781.html;   http://www.loc.gov/
catdir/description/random0415/00054781.html;   http:
//www.loc.gov/catdir/samples/random045/00054781.html.

Robinson:2005:EHY

[Rob05] Andrew Robinson. Einstein: a hundred years of Relativity. Harry N. Abrams, New York, NY, USA, 2005. ISBN 0-8109-5923-2 (hardcover). 256 pp. LCCN QC16.E5 R63

2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/toc/ecip059/
2005006593.html.

Robinson:2010:SGG

[Rob10] Andrew Robinson. Sudden genius?: the gradual path to creative breakthroughs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-956995-9 (hardcover). xxxv + 371 + 8 pp. LCCN BF408 .R628 2010.

Robinson:2012:SED

[Rob12] Andrew Robinson, editor. The scientists: an epic of discovery. Thames and Hudson, New York, NY, USA, 2012. ISBN 0-500-25191-6. 304 pp. LCCN Q141 .S3712 2012.

Roberts:2014:BUS

[Rob14] Sam Roberts. The brother: the untold story of the Rosenberg case: with a new epilogue. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4767-4738-5 (paperback), 1-4767-4739-3 (ebook). 574 + 16 pp. LCCN KF224.R6 R63 2014.

Robinson:2015:EHG

[Rob15a] Andrew Robinson. Einstein : un homme, un génie. (French) [Einstein: a man, a genius]. Flammarion, Paris, France, 2015. ISBN 2-08-135443-8. ???? pp. LCCN ????

Robinson:2015:EHY

[Rob15b] Andrew Robinson. Einstein: a Hundred Years of Relativity. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-8109-5923-2 (hardcover), 0-691-16989-6 (paperback). 256 pp. LCCN QC16.E5 R63 2015.

Rockwell:2003:CNW

[Roc03] Theodore Rockwell. Creating the new world: stories and images from the dawn of the Atomic Age. 1st Books Library, Bloomington, IN, USA, 2003. ISBN 1-4033-9087-8 (paperback), 1-4033-9086-X (e-book), 1-4107-0333-9 (dust jacket). x + 373 pp. LCCN TK9023 .R63 2003.

Rodgers:1992:MRE

[Rod92] Peter Rodgers. Music: Roll over Einstein …. Physics World, 5(10):8, October 1992. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/5/10/phwv5i10a8.
pdf.

Rodgers:2004:RGL

[Rod04] Peter Rodgers. Relativity gets lunar test. Physics World, 17 (6):9, June 2004. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/17/6/phwv17i6a16.pdf.

Rodgers:2005:ECG

[Rod05] Peter Rodgers. Einstein celebrations get off to a flyer. Physics World, 18(2):5, February 2005. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/2/phwv18i2a5.pdf.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (ebook). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rodriguez:2022:MMW

[Rod22] Juan Gabriel Rodríguez. Making the most of world talent for science? The Nobel Prize and Fields Medal experience. Scientometrics, 127(2):813–847, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://

link.springer.com/article/10.1007/s11192-021-04236-y.

Roensch:1993:LWL

[Roe93] Eleanor (Jerry) Stone Roensch. Life within limits: glimpses of everyday life at Los Alamos, New Mexico, seen through the experiences of a young female soldier while on military service there, May 1944 to April 1946. Los Alamos Historical Society, Los Alamos, NM, USA, 1993. ISBN 1-941232-14-X. 66 pp. LCCN D811 .R597 1993.

Rogers:2006:NLC

[Rog06] David Rogers. Nobel laureate contributions to 20th Century chemistry. Royal Society of Chemistry, Cambridge, UK, 2006. ISBN 0-85404-356-X (hardcover). xii + 656 pp. LCCN

QD21 .R73 2006. URL http://www.rsc.org/shop/books/
2006/9780854043569.asp.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rohrlich:1960:SES

[Roh60] F. Rohrlich. Self-energy and stability of the classical electron. American Journal of Physics, 28(7):639–643, October 1960. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/

v28/i7/p639_s1. See also [TN03, But68, Jan95, Teu96, Jan03].

Rohrlich:1987:PRO

[Roh87] Fritz Rohrlich. From Paradox to Reality: Our Basic Concepts of the Physical World. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-37605-X (paperback) , 0-521-30749-X (hardcover). x + 227 pp. LCCN QC173.55 .R64 1987. URL

http://www.cambridge.org/9780521376051.

Rolland:1915:ADM

[Rol15] Romain Rolland. Au-dessus de la Mêlée. (French) [Above the fray]. P. Ollendorff, Paris, France, 1915. 2 + 163 + 1 pp. LCCN D523 .R748 1915v.

Rolland:1919:VSO

[Rol19] Romain Rolland. V storonye ot skhvatki = Au-dessus de la mêlée. (Russian/French) [Above the fray]. Izdatelstvo Petrogradskovo soveieta rabochikh i krasniye deputatov [Petrograd publisher of the Soviet of Workers and Red Deputies], Petrograd, USSR, 1919. 160 pp. LCCN Cyr4 D 210.

Rollins:1947:HPE

[Rol47] Eleanor Rollins. The Hanford Project. An evaluative analysis of how administration, especially public welfare met the unusual situation created in the community by the making of the atomic bomb in Hanford, Washington, during World War II. Masters essay, Social Work, Columbia University, New York, NY, USA, 1947. ???? pp.

Rolland:1971:LLA

[Rol71] Romain Rolland. L’esprit libre: Au-dessus de la mêlée [et] Les précurseurs. (French) [The free spirit: Above the fray and its precursors]. Les Chefs d’Œuvres de Romain Rolland. Cercle du bibliophile, Évreux, France, 1971. xxii + 341 pp. LCCN D523 .R748 1971. Preface by Jean Albertini. Original illustrations by Pietro Sarto. Reprint of [Rol15].

Rolbein:1989:NCR

[Rol89] Seth Rolbein. Nobel Costa Rica: a timely report on our peaceful pro-Yankee, Central American neighbor. St. Martin’s Press, New York, NY, USA, 1989. ISBN 0-312-02262-X. 253 pp. LCCN F1548.2.R6 1988.

Rolnik:2010:TIP

[Rol10] Eran J. Rolnik. Therapy and ideology: Psychoanalysis and its vicissitudes in pre-state Israel (including some hitherto unpublished letters by Sigmund Freud and Albert Einstein). Science in Context, 23(4):473–506, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Romer:1991:EMP

[Rom91] Robert H. Romer. Editorial: Memorable papers from the American Journal of Physics, 1933–1990. American Journal of Physics, 59(3):201–207, March 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Ron:2007:EXC

[Ron07] Jose Manuel Sanchez Ron. Einstein and XX century philosophy (Albert Einstein). Arbor — Ciencia Pensamiento y Cultura, 183(728):833–853, November/December 2007. ISSN 0210-1963.

Ronald:2010:FF

[Ron10] Ann Ronald. Friendly fallout 1953. University of Nevada Press, Reno, NV, USA, 2010. ISBN 0-87417-825-8. x + 231 pp. LCCN PS3568.O56583 F75 2010.

Roosevelt:1939:LPA

[Roo39] Franklin D. Roosevelt. Letter to Professor Albert Einstein. Web document, October 10, 1939. URL http:

//www.pugetsound.edu/files/resources/7579_Roosevelt-
einstein-letter.png.

Rosen:1937:PPW

[Ros37] Nathan Rosen. Plane polarized waves in the General Theory of Relativity. Physikalische Zeitschrift der Sowjetunion, 12(??):366–372, ???? 1937. CODEN PHZSAL. ISSN 0369-

9811. URL http://relativity.livingreviews.org/refdb/
record/7613.

Rosen:1949:PRS

[Ros49] Nathan Rosen. A particle at rest in a static gravitational field. Reviews of Modern Physics, 21(3):503–505, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.503;   http://rmp.aps.org/

abstract/RMP/v21/i3/p503_1. Special issue in honor of Albert Einstein on his 70th birthday.

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72], pages 289–299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http:

//adsabs.harvard.edu/abs/1972cht..conf..289R.

Rosen:1979:CQM

[Ros79a] Nathan Rosen. Can quantum-mechanical description of physical reality be considered complete? In Aichelburg and Sexl [AS79], pages 57–67. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5

A66. DM48.00. URL http://link.springer.com/chapter/
10.1007/978-3-322-91080-6_6/.

Rosenberg:1979:AAS

[Ros79b] David Alan Rosenberg. American atomic strategy and the hydrogen bomb decision. The Journal of American History, 66 (1):62–87, June 1979. ISSN 0021-8723 (print), 1945-2314 (elec-

tronic). URL http://jah.oxfordjournals.org/content/66/
1/62.full.pdf+html.

Rosenfeld:1979:ECB

[Ros79c] Léon Rosenfeld. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday). In Cohen and Stachel [CS79a], chapter II.10, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7.

URL http://link.springer.com/chapter/10.1007/978-94-
009-9349-5_35.

Rossi:1980:MEE

[Ros80] Arcangelo Rossi. Mach ed Einstein: Influenza della Meccanica di Mach sul pensiero di Einstein. (Italian) [Mach and Einstein: Influence of the Mach’s Mechanics on Einstein’s thought]. Physis: Rivista Internazionale di Storia della Scienza, 22(2):279–292, 1980. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Rosen:1982:RAE

[Ros82a] Nathan Rosen. Reminiscences [of Albert Einstein]. In Holton and Elkana [HE82], pages 405–408. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Rosenberg:1982:GRS

[Ros82b] Nathan Rosenberg. The growing role of science in the innovation process. In Bernhard et al. [BCS82], pages 231–246. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500190.

Roseveare:1982:MPV

[Ros82c] N. T. Roseveare. Mercury’s perihelion, from Le Verrier to Einstein. Oxford science publications. Clarendon Press, Oxford, UK, 1982. ISBN 0-19-858174-2. viii + 208 pp. LCCN QB611 .R67 1982. US$20.00.

Rose:1986:RFK

[Ros86] Paul Lawrence Rose. Rezensionen: Fritz Krafft: Im Schatten der Sensation. Leben und Wirken von Fritz Strassmann. Weinheim/Deerfield Beach, Florida/Basel: Verlag Chemie 1981. XVII und 541 Seiten, DM 150. Berichte zur Wissenschaftsgeschichte, 9(2):131, ???? 1986. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Roseberg:1995:DTJ

[Rös95] Ulrich Röseberg. Did they just misunderstood each other? Logical empiricists and Bohr’s complementarity argument. In Gavroglu et al. [GSW95], pages 105–124. ISBN 0-7923-2991-0 (set), 0-7923-2988-0. LCCN Q174 .B67 vol. 163. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0823/94022250-
d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0823/94022250-t.html.

Rosenkranz:1998:ATL

[Ros98] Ze’ev Rosenkranz. Albert through the looking-glass: the personal papers of Albert Einstein, volume 59 of JNUL publications. Albert Einstein Archives, Jewish National and University Library, Hebrew University of Jerusalem, Jerusalem, 1998. ISBN 965-90170-0-6. xv + 156 pp. LCCN QC16.E5 R67 1998. URL http:

//www.loc.gov/catdir/toc/fy046/98196114.html. See also

revised and enlarged edition [RW07].

Rosen:2002:ACF

[Ros02a] Terry L. Rosen. The Atomic City: a firsthand account by a son of Los Alamos. Sunbelt Eakin, Austin, TX, USA, 2002. ISBN 1-57168-752-1, 1-57168-722-6 (paperback). xiii + 193 pp. LCCN F804.L6 R67 2002.

Rosenkranz:2002:ES

[Ros02b] Ze’ev Rosenkranz. The Einstein scrapbook. Albert Einstein Archives, Jewish National and University Library, Hebrew University of Jerusalem, Jerusalem, Israel, 2002. ISBN 0-8018-7203-0 (hardcover). xviii + 199 pp. LCCN QC16.E5

R674 2002. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.loc.gov/catdir/bios/
jhu052/2002005379.html;   http://www.loc.gov/catdir/
description/jhu051/2002005379.html.

Rosenkranz:2005:AEP

[Ros05] Ze’ev Rosenkranz. Albert Einstein: privat und ganz persönlich. (German) [Albert Einstein: private and rather personable]. Albert-Einstein-Archiv, Jüdische National- und Universitätsbibliothek, Hebräische Universität Jerusalem, Jerusalem, Israel, second edition, 2005. ISBN 3-03823-185-1. 235 pp. LCCN QC16.E5 R6715 2005.

Rosenkranz:2011:EBI

[Ros11] Ze’ev Rosenkranz. Einstein before Israel: Zionist icon or iconoclast? Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14412-5 (hardcover). xviii + 344 pp. LCCN QC16.E5 R673 2011.

Rosen:2012:HYI

[Ros12] Rebecca J. Rosen. And here you have it, ladies and gentlemen, E = mc^{2} and other Einstein Archive treasures: A newly unveiled digitization project is a feast for Einstein fans. Atlantic Monthly, ??(??):??, March 19, 2012. ISSN 1072-7825 (print), 2151-9463 (electronic). URL http://www.

theatlantic.com/technology/archive/2012/03/and-here-
you-have-it-ladies-and-gentlemen-e-mc-sup-2-sup-and-
other-einstein-archive-treasures/254754/.

Rosenkranz:2018:TDA

[Ros18] Ze’ev Rosenkranz, editor. The travel diaries of Albert Einstein. The Far East, Palestine, and Spain, 1922–1923. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17441-5. 256 pp. LCCN Q124.6-127.2.

Rotblat:1979:EPW

[Rot79] Joseph Rotblat. Einstein the pacifist warrior. Bulletin of the Atomic Scientists, 35(3):21–26, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Rotenstreich:1982:RR

[Rot82] Nathan Rotenstreich. Relativity and relativism. In Holton and Elkana [HE82], pages 175–204. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Rothenberg:1987:EBC

[Rot87] Albert Rothenberg. Einstein, Bohr, and creative thinking in science. History of Science (UK), 25(2):147–166, June 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/25/2/147.
full.pdf+html.

Rothman:1992:RRA

[Rot92] Hal Rothman. On rims and ridges: the Los Alamos area since 1880. University of Nebraska Press, Lincoln, NE, USA, 1992. ISBN 0-8032-3901-7. xiii + 376 + 13 pp. LCCN F804.L6

R68 1992. URL http://www.loc.gov/catdir/enhancements/
fy0727/91024418-b.html;   http://www.loc.gov/catdir/
enhancements/fy0727/91024418-d.html.

Rotblat:1995:RFA

[Rot95] Joseph Rotblat. Reminiscences on the fortieth anniversary of the Russell–Einstein manifesto. Presidential address at the 45th Pugwash Conference on Science and World Affairs, Hiroshima, Japan, July 1995., 1995.

Rotblat:1996:RYH

[Rot96] Joseph Rotblat. Remember your humanity. Bulletin of the Atomic Scientists, 52(2):26–28, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothman:1997:RRA

[Rot97] Hal Rothman. On rims and ridges: the Los Alamos area since 1880. University of Nebraska Press, Lincoln, NE, USA, 1997. ISBN 0-8032-8966-9 (paperback). xiii + 384 + 13 pp.

LCCN F804.L6 R68 1997. URL http://www.loc.gov/catdir/
enhancements/fy0728/97211154-d.html.

Rotter:2008:WB

[Rot08] Andrew Jon Rotter. The world’s bomb. Making of the modern world. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-280437-5 (hardcover), 0-19-956976-2 (paperback). viii + 4 + 371 + 16 pp. LCCN QC773 .R67 2008.

Rouze:1964:ROM

[Rou64] Michel Rouzé. Robert Oppenheimer, the man and his theories. A Profile in science. Souvenir Press, London, UK, 1964. 189 pp. LCCN QC16.O62 R613 1964.

Rouze:1965:ROM

[Rou65] Michel Rouzé. Robert Oppenheimer, the man and his theories. A Profile in science. P. S. Eriksson, New York, 1965. 192 pp. LCCN QC16.O62 R613 1965.

Rowe:1985:LB

[Row85] Dorothy Rowe. Living with the bomb. Routledge and Keagan Paul, London, UK, 1985. ISBN 0-7102-0477-9. xi + 243 pp. LCCN U264 .R68 1985.

Rowe:2000:BRA

[Row00] David E. Rowe. Book review: Albert Einstein, Robert Schulmann, A. J. Kox, Michel Janssen, József Illy, Ann M. Hentschel, and Klaus Hentschel: The Collected Papers of Albert Einstein: The Berlin Years: Correspondence, 1914–1918. Isis, 91(1):183– 185, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
237620.

Rowe:2001:EMH

[Row01] David E. Rowe. Einstein meets Hilbert: At the crossroads of physics and mathematics. Physics in Perspective (PIP), 3(4):379–424, November 2001. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/PL00000538;
 http://www.springerlink.com/content/5byc4pv8q4gq5qy1/
fulltext.pdf.

Rowlinson:2005:ECP

[Row05] J. S. Rowlinson. Einstein: The classical physicist. Notes and Records of the Royal Society of London, 59(3):255–271, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
30041502.

Rowe:2006:BRB

[Row06a] David E. Rowe. Book review: “Zwei wirkliche Kerle”: Neues zur Entdeckung der Gravitationsgleichungen der Allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert, Wuensch Daniela. Termessos, Göttingen (2005), ISBN 3-938016-04-3. Historia Mathematica, 33(4):500–508, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000509.

Rowe:2006:BRZ

[Row06b] David E. Rowe. Book review: “Zwei wirkliche Kerle”: Neues zur Entdeckung der Gravitationsgleichungen der Allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert, Wuensch Daniela. Termessos, Göttingen (2005), ISBN 3-938016-04-3. Historia Mathematica, 33(4):500–508, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000509.

Rowe:2006:EAE

[Row06c] David E. Rowe. Einstein’s allies and enemies: debating Relativity in Germany, 1916–1920. In Hendricks et al. [HHJ^{+}06], pages 231–280. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346. LCCN Q175 .B69 v.251

2006; Q175 .I51377 2006. URL http://www.loc.gov/catdir/
enhancements/fy0825/2007425756-d.html;   http://www.
loc.gov/catdir/toc/fy0713/2007425756.html.

Rowe:2006:TGG

[Row06d] David E. Rowe. “Two genuine guys”: New information on the discovery of the gravitation equations of the General Relativity Theory through Albert Einstein and David Hilbert. Historia Mathematica, 33(4):500–508, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Rowe:2008:BRE

[Row08] David E. Rowe. Book review: Einstein studies, volume 11: a retrospective review [The universe of General Relativity]. Studies in History and Philosophy of Modern Physics, 39(3):667– 686, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S135521980800018X. See [EK05].

Rowe:2012:ERW

[Row12] David E. Rowe. Einstein and Relativity: What price fame? Science in Context, 25(2):197–246, May 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rowe:2017:BRS

[Row17] David E. Rowe. Book review: Steven Gimbel. Einstein: His Space and Times. Isis, 108(1):207–208, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roy:1919:DTR

[Roy19] M. N. Roy. Discussion on the theory of relativity. Astron. Soc., LXXX(2):96–??, December 1919.

Realdi:2009:ESB

[RP09] Matteo Realdi and Giulio Peruzzi. Einstein, de Sitter and the beginning of relativistic cosmology in 1917. General Relativity and Gravitation, 41(2):225–247, ???? 2009. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL http://

www.springerlink.com/content/h8jlm02x0p03n212/.

Reichenbach:1965:TRP

[RR65] Hans Reichenbach and Maria Reichenbach, editors. The Theory of Relativity and A Priori Knowledge. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1965. xliv + 116 pp. LCCN QC 6 .R384.

Reichenbach:1969:ATR

[RR69] Hans Reichenbach and Maria Reichenbach. Axiomatization of the Theory of Relativity. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1969. xix + 208 pp. LCCN QC6 .R3813. English translation of [Rei24].

Rassias:1982:SSP

[RR82] Themistocles M. Rassias and George M. Rassias, editors. Selected studies, physics-astrophysics, mathematics, history of science: a volume dedicated to the memory of Albert Einstein. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86161-0. xii + 392 pp. LCCN QB461 .S374 1982.

Rodrigues:1989:MTT

[RR89] Waldyr A. Rodrigues, Jr. and Marcio A. F. Rosa. The meaning of time in the Theory of Relativity and “Einstein’s later view of the Twin Paradox”. Foundations of Physics, 19(6):705–724, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00731906. See [Sac85].

Rynasiewicz:2006:TPE

[RR06] Robert Rynasiewicz and Jürgen Renn. The turning point for Einstein’s Annus mirabilis. Studies in History and Philosophy of Modern Physics, 37(1):5–35, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000961.

Renn:2014:ECR

[RR14] Jürgen Renn and Robert Rynasiewicz. Einstein’s Copernican revolution. In Janssen and Lehner [JL14b], chapter 1, pages 38–71. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Ruter:2017:EFK

[RR17] Martina Rüter and Simon Rüter. Einstein für Kids — Kalender 2018: Forschen, staunen und entdecken. (German) [Einstein for Kids Calendar 2018: Research, astonish, and discover]. KV&H Verlag, Unterhaching, Germany, 2017. ISBN 3-8400-1729-7. 640 pp. LCCN ????

Rosenfeld:1964:NBH

[RRK^{+}64] Léon Rosenfeld, Erik Rüdinger, Oskar Klein, Werner Heisenberg Hendrik G. B. Casimir, Otto Robert Frisch, Stefan Rozental, Aage Bohr, Abraham Pais, Jørgen Kalckar, Christian Møller, Mogens Pihl, Viktor F. Weisskopf, Johannes Pedersen, Viggo Kampmann, Richard Courant, Paul A. M. Dirac, Hans Henrik Koch, William Scharff, Mogens Andersen, Hans Bohr, and Niels Bohr. Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers]. J. H. Schultz Forlag, Copenhagen, DK, 1964. 341 pp. LCCN ????

Recami:2007:ESM

[RRR07] Erasmo Umberto M. Recami, Umberto Victor G. Recami, and Erasmo Recami. Einstein, Sciascia, Majorana, Amaldi, e il rapporto tra intellettuali e potere. (Italian) [Einstein, Sciascia, Majorana, Amaldi, and the relationship of intellectuals with the power]. arxiv.org, ??(??):??, September 28, 2007. URL http:/

/arxiv.org/abs/0709.4672.

Rosenthal-Schneider:1949:PAE

[RS49] Ilse Rosenthal-Schneider. Presuppositions and anticipations in Einstein’s physics. In Schilpp [Sch49b], chapter 3, pages 129– 146. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Rosenthal-Schneider:1980:RE

[RS80a] Else Rosenthal-Schneider. Reminiscenses of Einstein. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Rosenthal-Schneider:1980:RST

[RS80b] Ilse Rosenthal-Schneider. Reality and scientific truth: discussions with Einstein, von Laue, and Planck. Wayne State University Press, Detroit, MI, USA, 1980. ISBN 0-8143-1650-6. 148 pp. LCCN QC6 .R5696. Edited by Thomas Braun. Foreword by Arthur I. Miller.

Renn:1992:AEM

[RS92] Jürgen Renn and Robert J. Schulmann, editors. Albert Einstein/Mileva Marić: the love letters. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-08886-1 (paperback). xxxi + 107 pp. LCCN 03.E01958. URL http:

//press.princeton.edu/titles/4988.html. Translated by

Shawn Smith.

Renn:1996:EZN

[RS96] Jürgen Renn and Tilman Sauer. Einsteins Züricher Notizbuch. Die Entdeckung der Feldgleichungen der Gravitation im Jahre 1912. (German) [Einstein’s Zurich notebook. The discovery of the gravitational field equations in 1912]. Physikalische Blätter, 52(9):865–872, September 1996. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19960520907/abstract.

Renn:1999:HMR

[RS99] Jürgen Renn and Tilman Sauer. Heuristics and mathematical representation in Einstein’s search for a gravitational field equation. In Goenner et al. [GRRS99b], page ?? ISBN 0-8176-4060-6 (hardcover), 3-7643-4060-6. LCCN QC173.6 .E97 1999.

Renn:2003:ESM

[RS03a] Jürgen Renn and Tilman Sauer. Eclipses of the stars: Mandl, Einstein, and the early history of gravitational lensing. In Ashtekar et al. [ACH^{+}03], pages 69–92. ISBN 1-4020-1284-5.

LCCN Q174 .B67 vol. 234. URL http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-d.html;   http://www.
loc.gov/catdir/enhancements/fy0813/2003046757-t.html.

Renn:2003:EIR

[RS03b] Jürgen Renn and Tilman Sauer. Errors and insights: Reconstructing the genesis of general relativity from Einstein’s Zürich Notebook. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 253–268. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.com/chapter/10.1007/0-306-
48152-9_12.

Raussen:2006:IPD

[RS06] Martin Raussen and Christian Skau. Interview with Peter D. Lax. Notices of the American Mathematical Society, 53(2):223– 229, February 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.org/

notices/200602/comm-lax.pdf. Peter D. Lax is the recipient

of the 2005 Abel Prize of the Norwegian Academy of Science and Letters. On May 24, 2005, prior to the Abel Prize celebrations in Oslo, Lax was interviewed by Martin Raussen of Aalborg University and Christian Skau of the Norwegian University of Science and Technology. This interview originally appeared in the European Mathematical Society Newsletter, September 2005, pages 24–31.

Renn:2007:PCP

[RS07a] Jürgen Renn and Tilman Sauer. Pathways out of classical physics: Einstein’s double strategy in his search for the gravitational field equation. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 113–312. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https://link.springer.

com/chapter/10.1007/978-1-4020-4000-9_4.

Renn:2007:GTC

[RS07b] Jürgen Renn and Matthias Schemmel. Gravitation in the twilight of classical physics: An introduction. In Renn et al. [RSS^{+}07a], pages 1–18. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_10.

Renn:2007:HFP

[RS07c] Jürgen Renn and John Stachel. Hilbert’s foundation of physics: From a theory of everything to a constituent of General Relativity. In Renn et al. [RSS^{+}07b], pages 857–973. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_41.

Rigden:2007:EF

[RS07d] John S. Rigden and Roger H. Stuewer. Editorial: FAPP. Physics in Perspective (PIP), 9(1):1–3, January 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0334-6.

Rojo:2007:AEP

[RS07e] Jose Antonio Rojo and Leandro Sequeiros. Albert Einstein (1879–1955): Physics, philosophy and mystic. Pensamiento, 63(238):659–692, ???? 2007. ISSN 0031-4749.

Rowe:2007:EPH

[RS07f] David E. Rowe and Robert J. Schulmann, editors. Einstein on politics: his private thoughts and public stands on nationalism, Zionism, war, peace, and the bomb. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12094-3. xxxiv + 523 pp. LCCN QC16.E5 E5157 2007. URL http://www.loc.

gov/catdir/enhancements/fy0704/2006100303-b.html;
http://www.loc.gov/catdir/enhancements/fy0704/2006100303-
d.html;   http://www.loc.gov/catdir/toc/ecip076/2006100303.
html.

Reed:2009:NEP

[RS09] Thomas C. Reed and Danny B. Stillman. The nuclear express: a political history of the bomb and its proliferation. Zenith Press, Minneapolis, MN, USA, 2009. ISBN 0-7603-3502-8. viii + 392 pp. LCCN U264 .R44 2009.

Richie:2011:HSR

[RS11a] Jackson Richie and Jean Sherrod. The house by the side of the road. University of Alabama Press, Tuscaloosa, AL, USA, 2011. ISBN 0-8173-1694-9 (hardcover), 0-8173-8326-3

(e-book). LCCN ???? URL http://site.ebrary.com/lib/
royallibrary/Doc?id=10527717.

Rigden:2011:BNG

[RS11b] John S. Rigden and Roger H. Stuewer. Book notes [Galileo’s Two Chief World Systems and Two New Sciences; Harald Fritzsch, You Are Wrong, Mr Einstein!]. Physics in Perspective (PIP), 13(3):373–375, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0067-4.

Rosenthal-Schneider:2013:BEL

[RS13] Ilse Rosenthal-Schneider. Begegnungen mit Einstein, von Laue und Planck Realität und Wissenschaftliche Wahrheit. (German) [Discussions with Einstein, von Laue und Planck: Reality and Scientific Truth]. Vieweg + Teubner Verlag, 2013. ISBN 3-663-01885-7. ???? pp. LCCN ????

Rukop:1949:KMA

[RSPS49] H. Rukop, Arnold Sommerfeld, K. Philipp, and H. Schimank. Karl Mey/Albert Einstein 70 Jahre/Otto Hahn 70 Jahre/Georg Christoph Lichtenberg. (German) [Karl Mey slash Albert Einstein 70 years slash Otto Hahn 70 years slash Georg Christoph Lichtenberg]. Physik Journal, 5(3):124–133, March 1949. CO-DEN PJHOB2. ISSN 1521-3722.

Renn:1997:OGL

[RSS97] Jürgen Renn, Tilman Sauer, and John Stachel. The origin of gravitational lensing: a postscript to Einstein’s 1936 Science paper. Science, 275(5297):184–186, January 10, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/275/
5297/184.full.pdf. See [Ein36e].

Renn:2007:GGRa

[RSS^{+}07a] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 3. Gravitation in the Twilight of Classical Physics: Between Mechanics, Field Theory, and Astronomy, volume 250(3) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9. ISSN 0068-0346. 619 pp. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Renn:2007:GGRb

[RSS^{+}07b] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 4. Gravitation in the Twilight of Classical Physics: the Promise of Mathematics, volume 250(4) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. 621–1152 pp. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL http://

d-nb.info/985765933/34;   http://www.loc.gov/catdir/
enhancements/fy0824/2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/fy0824/2007440325-t.html;
 https://link.springer.com/book/10.1007/978-1-4020-
4000-9.

Rotblat:1993:NWF

[RSU93] Joseph Rotblat, J. Steinberger, and B. M. Udgaonkar, editors. A nuclear-weapon-free world: desirable?, feasible? A Pugwash monograph. Westview Press, Boulder, CO, USA, 1993. ISBN 0-8133-8718-3. xiv + 228 pp. LCCN JX1974.7 .N847

1993. URL http://www.loc.gov/catdir/enhancements/
fy0831/92021168-b.html.

Rhodes:1990:AGS

[RT90] Richard Rhodes and Peter Torberg. Die Atombombe oder Die Geschichte des 8. Schöpfungstages. (German) [The Atomic Bomb, or The History of the Eighth Day of Creation]. Verlag Volk und Welt, Berlin, Germany, 1990. ISBN 3-353-00717-2 (clothbound). 915 pp.

Rindler:1987:GGV

[RTR87] Wolfgang Rindler, A. (Andrzej) Trautman, and Ivor Robinson, editors. Gravitation and geometry: a volume in honour of Ivor Robinson, volume 4 of Monographs and textbooks in physical science. Bibliopolis, Napoli, Italy, 1987. ISBN 88-7088-142-3. 505 pp. LCCN QC178 .G62 1987.

Reid:1965:ESM

[RTS^{+}65] R. W. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. 58-minute black-and-white film from BBC Television, London, UK., 1965. URL http://explore. bfi.org.uk/4ce2b6d1c3ab0. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Reid:1969:ESM

[RTS^{+}69] R. W. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. 42-minute black-and-white motion picture from Time-Life Films, Paramus, NJ, USA., 1969. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Reid:20xx:ESM

[RTS^{+}xx] R. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. Motion picture, 20xx. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Rudolph:2017:QQ

[Rud17] Terry Rudolph. Q is for quantum. Amazon Fulfillment, Wroclaw, Poland, 2017. ISBN 0-9990635-0-2. x + 153 pp. LCCN ????

Ruger:1988:ACE

[Rüg88] Alexander Rüger. Atomism from cosmology: Erwin Schrödinger’s work on wave mechanics and space–time structure. Historical Studies in the Physical and Biological Sciences, 18(2):377– 401, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print),

1533-8355 (electronic). URL http://www.jstor.org/stable/
27757607.

Rupke:2009:ELT

[Rup09] Nicolaas A. Rupke, editor. Eminent Lives in Twentieth-Century Science and Religion. Lang, Frankfurt am Main, Germany, second edition, 2009. ISBN 3-631-58120-3. 371 pp. LCCN BL240.3 .E48 2009. EUR 49.00. With chapters on Rachel Carson, Charles A. Coulson, Theodosius Dobzhansky, Arthur S. Eddington, Albert Einstein, Ronald A. Fisher, Julian Huxley, Pascual Jordan, Robert A. Millikan, Ivan P. Pavlov, Michael I. Pupin, Abdus Salam, and Edward O. Wilson.

Russell:1925:RNT

[Rus25] Henry Norris Russell. Remarkable new tests favor the Einstein theory. Scientific American, 133(2):88, August 1925. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v133/n2/pdf/scientificamerican0825-88.pdf.

Russell:1931:AR

[Rus31] Bertrand Russell. The ABC of Relativity. K. Paul, Trench, Trubner, London, UK, 1931. 231 pp. LCCN QC6 .R8 1931.

Russell:1937:RE

[Rus37] Henry Norris Russell. A Relativistic eclipse. Scientific American, 156(2):76–77, February 1937. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v156/n2/
pdf/scientificamerican0237-76.pdf.

Russell:1943:BCR

[Rus43] Bertrand Russell. El A. B. C. de le relatividad. (Spanish) [The A. B. C. of Relativity]. Ediciones Imán, Buenos Aires, Argentina, 1943. 194 pp. LCCN QC6 .R817. Spanish translation by Ernesto R. Sábato.

Russell:1958:AR

[Rus58] Bertrand Russell. The ABC of Relativity. G. Allen and Unwin, London, UK, second edition, 1958. 139 pp. LCCN QC6 .R8 1958. Revision edited by Felix Pirani.

Russell:1969:ARC

[Rus69] Bertrand Russell. The ABC of Relativity. Allen and Unwin, London, UK, third edition, 1969. ISBN 0-04-521001-2. 139 pp. LCCN QC6 .R8 1969. Revised and edited by F. A. E. (Felix Arnold Edward) Pirani.

Russell:1981:SBD

[Rus81] Colin A. Russell. Science and belief: from Darwin to Einstein. Block 3: Unit 4–5, Time, chance and thermodynamics. The Open University Press, Milton Keynes, UK, 1981. ISBN 0-335-11002-9. 74 pp. LCCN QC311 .R784 1981.

Russell:1985:PLP

[Rus85a] B. R. Russell. Preparation-limited predictions in Einstein– Podolsky–Rosen experiments. Foundations of Physics, 15(8): 861–869, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738319.

Russell:1985:AR

[Rus85b] Bertrand Russell. The ABC of Relativity. New American Library, New York, NY, USA, fourth edition, 1985. ISBN 0-451-62381-9 (paperback). 220 pp. LCCN QC173.55 .R85 1985. US$4.50. Edited by F. A. E. (Felix Arnold Edward) Pirani.

Russ:1990:PAP

[Rus90] Harlow W. (Harlow Wilson) Russ. Project Alberta: the preparation of atomic bombs for use in World War II. Exceptional Books, Los Alamos, NM, USA, 1990. ISBN 0-944482-01-5. (various) pp. LCCN QC773.3.M36 R87 1990.

Russell:1991:ELG

[Rus91] Bertrand Russell. Einstein’s law of gravitation. In Ferris and Fadiman [FF91], pages 194–202. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Russell:1997:AR

[Rus97] Bertrand Russell. ABC of Relativity. Routledge, London, UK, fifth edition, 1997. ISBN 0-415-15429-4. xvi + 9–155 pp. LCCN QC173.55 .R85 1999. URL http://www.loc.gov/

catdir/enhancements/fy0908/2001278032-d.htm. With an

introduction by Peter Clark.

Russo:1998:MME

[Rus98] Frank P. Russo. The Michelson–Morley experiment: the final solution? Speculations in Science and Technology, 21(2):73–78,

August 1998. URL http://www.springerlink.com/content/
h17j267710137137/.

Russell:2009:AR

[Rus09] Bertrand Russell. ABC of Relativity. Routledge classics. Routledge, London, UK, 2009. ISBN 0-415-47382-9. xvii + 150 pp. LCCN QC6 .R8 2009.

Ruskin:2017:HVH

[Rus17] Steve Ruskin. Book review: The Hunt for Vulcan: … and How Albert Einstein Destroyed a Planet, Discovered Relativity, and Deciphered the Universe, Thomas Levenson, 244 pp. Random House, New York, 2015. Price $26 (hardcover). ISBN 978-0-8129-9898-6. American Journal of Physics, 85(2):159– 160, February 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3 S36 1994. URL http://www.loc.

gov/catdir/description/cam025/92041633.html;   http:
//www.loc.gov/catdir/samples/cam031/92041633.html;
http://www.loc.gov/catdir/toc/cam028/92041633.html.

Rechenberg:2002:WHG

[RW02] Helmut Rechenberg and Gerald Wiemers, editors. Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). (German) [Werner Heisenberg: Appraisal and test certificates for graduations and habilitations], volume 29 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS Verlag, Berlin, Germany, second edition, 2002. ISBN 3-928577-43-3. 266 pp. LCCN QC16.H35 W475 2002.

Rosenkranz:2007:AEP

[RW07] Ze’ev Rosenkranz and Barbara Wolff. Albert Einstein: the persistent illusion of transience. Albert Einstein Archives, Jerusalem, Israel, 2007. ISBN 965-493-325-X (hardcover). 263 pp. LCCN QC16.E5 A69 2007.

Ren:2023:DCI

[RWL23] Jingjing Ren, Fang Wang, and Minglu Li. Dynamics and characteristics of interdisciplinary research in scientific breakthroughs: case studies of Nobel-winning research in the past 120 years. Scientometrics, 128(8):4383–4419, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-023-04762-x.

Rennie:1997:SBN

[RWY97] John Rennie, Paul Wallich, and Philip Yam. Special briefing: The 1996 Nobel Prizes in science. Scientific American, 276(1):14–18, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v276/n1/
pdf/scientificamerican0197-14.pdf.

Ryan:2006:RLT

[Rya06] Eugene E. Ryan. Reflections on light and time in the philosophy of Franciscus Patricius and in the 1905 paper of Albert Einstein “The electrodynamics of moving bodies”. Synthesis Philosophica, 21(2):195–208, ???? 2006. ISSN 0352-7875.

Ryckman:1992:POC

[Ryc92] T. A. Ryckman. “P(oint)-c(oincidence) thinking”: The ironical attachment of logical empiricism to General Relativity (and some lingering consequences). Studies in History and Philosophy of Science Part A, 23(3):471–497, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290005Q.

Ryckman:1999:ECG

[Ryc99] T. A. Ryckman. Einstein, Cassirer, and general covariance — then and now. Science in Context, 12(4):585–619, December 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ryckman:2005:RRP

[Ryc05] Thomas Ryckman. The Reign of Relativity: Philosophy in Physics, 1915–1925. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517717-7. ix + 317 pp. LCCN

QC173.52 .R93 2005. URL http://www.loc.gov/catdir/
enhancements/fy0617/2004041576-d.html;   http://www.
loc.gov/catdir/enhancements/fy0725/2004041576-b.html.

Ryckman:2014:BRE

[Ryc14] Thomas Ryckman. A believing rationalist”: Einstein and “the truly valuable” in Kant. In Janssen and Lehner [JL14b], chapter 12, pages 377–397. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Ryden:2017:IC

[Ryd17] Barbara Sue Ryden. Introduction to cosmology. Cambridge University Press, Cambridge, UK, second edition, 2017. ISBN 1-107-15483-9. xii + 264 pp. LCCN QB981 .R93 2017.

Ryff:2014:NRD

[Ryf14] Luiz Carlos Ryff. Null-result detection and Einstein–Podolsky– Rosen correlations. Foundations of Physics, 44(1):58–70, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-013-9762-0.

Rynasiewicz:2005:OEE

[Ryn05] R. Rynasiewicz. The optics and electrodynamics of On the Electrodynamics of Moving Bodies. Annalen der Physik (Berlin), 14 (1):38–57, February 2005. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/andp.200410135.

Slater:1956:ELG

[S^{+}57] John C. (John Clarke) Slater et al., editors. Elektrische Leitungsphänomene. (German) [Electrical conductivity], volume 19–20 of Handbuch der Physik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1956/1957. ???? pp. LCCN QC21 .H327 Bd. 19-20.

Schuck:1962:NMH

[S^{+}62] H. Schück et al. Nobel, the man and his prizes. Elsevier, Amsterdam, The Netherlands, second edition, 1962. x + 690 pp. LCCN ????

Schellnhuber:2010:GSN

[S^{+}10] Hans-Joachim Schellnhuber et al., editors. Global sustainability: a Nobel cause. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76934-5. LCCN GE140 .G59 2010.

URL http://www.loc.gov/catdir/enhancements/fy1010/
2010292978-d.html;   http://www.loc.gov/catdir/enhancements/
fy1010/2010292978-t.html.

Spielberg:1985:SIS

[SA85] Nathan Spielberg and Bryon D. Anderson. Seven ideas that shook the universe. Wiley, New York, NY, USA, 1985. ISBN 0-471-81477-6. viii + 291 pp. LCCN QC21.2 .S65

1985. URL http://www.loc.gov/catdir/enhancements/
fy0607/85006442-d.html;   http://www.loc.gov/catdir/
enhancements/fy0607/85006442-t.html.

Spielberg:1987:SIS

[SA87] Nathan Spielberg and Bryon D. Anderson. Seven ideas that shook the universe. Wiley, New York, NY, USA, 1987. ISBN 0-471-85974-5, 0-471-84816-6 (paperback). viii + 263 pp. LCCN QC21.2 .S65 1987.

Spielberg:1995:SIS

[SA95] Nathan Spielberg and Bryon D. Anderson. Seven ideas that shook the universe. Wiley, New York, NY, USA, second edition, 1995. ISBN 0-471-30606-1. xi + 355 pp. LCCN QC21.2

.S65 1995. URL http://www.loc.gov/catdir/description/
wiley032/94032986.html;   http://www.loc.gov/catdir/
toc/onix02/94032986.html.

Sachs:1981:IMA

[Sac81] Mendel Sachs. Ideas of matter: from ancient times to Bohr and Einstein. University Press of America, Washington, DC, USA, 1981. ISBN 0-8191-1615-7, 0-8191-1616-5 (paperback). xi + 322 pp. LCCN QC171.2 .S23 1981.

Sachs:1985:ELV

[Sac85] Mendel Sachs. On Einstein’s later view of the Twin Paradox. Foundations of Physics, 15(9):977–980, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00739038. See comments [RR89].

Sachs:1988:EVB

[Sac88] Mendel Sachs. Einstein versus Bohr: the continuing controversies in physics. Open Court, La Salle, IL, USA, 1988. ISBN 0-8126-9064-8, 0-8126-9065-6 (paperback). xxiii + 296 pp. LCCN QC7 .S2 1988.

Sachs:2010:PU

[Sac10] Mendel Sachs. Physics of the universe. Imperial College Press, London, UK, 2010. ISBN 1-84816-532-3 (hardcover), 1-84816-604-4 (paperback). xvii + 128 pp. LCCN QB981 .S253 2010.

Shapiro:1971:FTG

[SAI^{+}71] Irwin I. Shapiro, Michael E. Ash, Richard P. Ingalls, William B. Smith, Donald B. Campbell, Rolf B. Dyce, Raymond F. Jurgens, and Gordon H. Pettengill. Fourth test of General Relativity: New radar result. Physical Review Letters, 26:1132–1135, May 3, 1971. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.26.1132.

Salzer:1974:TLE

[Sal74] Herbert E. Salzer. Two letters from Einstein concerning his distant parallelism field theory. Archive for History of Exact Sciences, 12(1):89–96, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=1&spage=89.

Salewski:1995:ZBG

[Sal95] Michael Salewski. Das Zeitalter der Bombe: die Geschichte der atomaren Bedrohung von Hiroshima bis heute. (German) [The age of the bomb: the story of the atomic threat of Hiroshima until today], volume 1103 of Beck’sche Reihe C. H. Beck Wissen. Beck, München, Germany, 1995. ISBN 3-406-39203-2. 334 pp. LCCN ????

Salvatori:1999:AEB

[Sal99] Roberto Salvatori. Albert Einstein’s brain. The Lancet (London, England), 354(9192):1821–1822, November 20, 1999. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547X (electronic).

Salam:2005:ELD

[Sal05] Abdus Salam. Einstein’s last dream: The space–time unification of fundamental forces. Resonance, 10(12):246–253, December 2005. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X

(electronic). URL http://www.springerlink.com/content/
a3075114r14mq262/.

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947–1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sample:2010:MMP

[Sam10] Ian Sample. Massive: the missing particle that sparked the greatest hunt in science. Basic Books, New York, NY, USA, 2010. ISBN 0-465-01947-1. xi + 260 pp. LCCN QC793.5.B62 S26

Sanesi:1977:TLA

[San77] Elena Sanesi. Three letters by Albert Einstein and some information on Einstein’s stay at Pavia. Physis: Rivista Internazionale di Storia della Scienza, 18 (1976)(??):174–178 (1977), 1977. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Sanger:1981:BF

[San81a] Penny Sanger. Blind Faith. McGraw-Hill Ryerson, Toronto, ON, Canada, 1981. ISBN 0-07-092423-6 (paperback). viii + 182 pp. LCCN TD899.U73 S26 1981.

Santilli:1981:ILE

[San81b] Ruggero Maria Santilli. An intriguing legacy of Einstein, Fermi, Jordan, and others: The possible invalidation of quark conjectures. Foundations of Physics, 11(5–6):383–472, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00727075.

Sanger:1995:WBO

[San95] Stephen L. Sanger. Working on the Bomb: an Oral History of WWII Hanford. Portland State University, Continuing Education Press, Portland, OR, USA, 1995. ISBN 0-87678-115-6 (paperback). xiii + 264 pp. LCCN QC773.3.U5 S265 1995. Edited by Craig Wollner.

Sanders:2007:BRE

[San07] Barry C. Sanders. Book review: Einstein: a miraculous year one century later. Science and Public Policy, 34(6): 449–450, July 2007. ISSN 0302-3427 (print), 1471-5430 (elec-

tronic). URL http://spp.oxfordjournals.org/content/34/
6/449.full.pdf+html.

Santos:2011:DMC

[San11] Gildo Magalhães Santos. A debate on magnetic current: the troubled Einstein–Ehrenhaft correspondence. British Journal for the History of Science, 44(3):371–400, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sangwal:2015:GDC

[San15] Keshra Sangwal. On the growth dynamics of citations of articles by some Nobel Prize winners. Journal of Informetrics, 9 (3):466–476, July 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715000449.

Sarton:1935:BRBe

[Sar35] George Sarton. Book review: Comment je vois le monde by Albert Einstein; Colonel Cros. Isis, 23(1):278–280, June 1935. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225238.

Sartori:1996:URS

[Sar96] Leo Sartori. Understanding relativity: a simplified approach to Einstein’s theories. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1996. ISBN 0-520-07986-8 (cloth), 0-520-20029-2 (paper). xiii + 367 pp. LCCN QC173.55 .S367 1996. URL http://www.loc.gov/

catdir/bios/ucal051/94049358.html;   http://www.loc.
gov/catdir/description/ucal041/94049358.html.

Sarkar:2008:EDE

[Sar08] Subir Sarkar. Is the evidence for dark energy secure? General Relativity and Gravitation, 40(2–3):269–284, ???? 2008. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532

(electronic). URL http://www.springerlink.com/content/
c712v10751434048/.

Sauer:2004:CEEa

[Sau04a] Tilman Sauer. The challenge of editing Einstein’s scientific manuscripts. arXiv.org, December 22, 2004. URL http://

arxiv.org/abs/physics/0412142.

Sauer:2004:CEEb

[Sau04b] Tilman Sauer. The challenge of editing Einstein’s scientific manuscripts. Documentary editing, 26(??):145–165, ???? 2004. ISSN 0196-7134 (print), 2167-1451 (electronic).

Sauer:2005:EEH

[Sau05] Tilman Sauer. Einstein equations and Hilbert action: What is missing on page 8 of the proofs for Hilbert’s First Communication on the Foundations of Physics? Archive for History of Exact Sciences, 59(6):577–590, October 2005. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=6&spage=577.

Sauer:2006:FET

[Sau06] Tilman Sauer. Field equations in teleparallel space-time: Einstein’s Fernparallelismus approach toward unified field theory. Historia Mathematica, 33(4):399–439, November 2006. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005001060.

Sauer:2007:EET

[Sau07a] Tilman Sauer. Einstein and the early Theory of Superconductivity, 1919–1922. Archive for History of Exact Sciences, 61(2):159–211, March 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=2&spage=159.

Sauer:2007:EEH

[Sau07b] Tilman Sauer. Einstein equations and Hilbert action: What is missing on page 8 of the proofs for Hilbert’s first communication on the foundations of physics? In Renn et al. [RSS^{+}07b], pages 975–988. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (ebook). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL https://link.springer.com/chapter/10.
1007/978-1-4020-4000-9_42.

Sauer:2007:EME

[Sau07c] Tilman Sauer. An Einstein manuscript on the EPR paradox for spin observables. Studies in History and Philosophy of Modern Physics, 38(4):879–887, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000330.

Sauer:2008:ROP

[Sau08] Tilman Sauer. Remarks on the origin of path integration: Einstein and Feynman. In Janke and Pelster [JP08], pages 3–13. ISBN 981-283-726-4, 981-283-727-2 (e-book). LCCN QC174.17.P27 I57 2007eb.

Sauer:2014:EUF

[Sau14] Tilman Sauer. Einstein’s unified field theory program. In Janssen and Lehner [JL14b], chapter 9, pages 281–305. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN

QC16.E5 C36 2014. URL http://philsci-archive.pitt.
edu/3293/1/uft.pdf.

Sauer:2016:BRP

[Sau16] Tilman Sauer. Book review: Paul Halpern. Einstein’s Dice and Schrödinger’s Cat: How Two Great Minds Battled Quantum Randomness to Create a Unified Theory of Physics. Isis, 107(2): 427–428, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sauer:2019:EWS

[Sau19] Tilman Sauer. Einstein’s working sheets and his search for a unified field theory. European Physical Journal H, 44(4–5): 371–379, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2019-100019-6.

Sauer:2021:SEG

[Sau21] Tilman Sauer. Soldner, Einstein, gravitational light deflection and factors of two. Annalen der Physik (1900), 533(8):2100203, August 2021. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Singer:1982:PTN

[SAY^{+}82] I. B. Singer, K. J. Arrow, R. S. Yalow, A. Penzias, R. Feynman, and D. Baltimore. Psychology tomorrow — the Nobel view. Psychology Today, 16(12):21–??, ???? 1982. ISSN 0033-3107.

Sayen:1985:EAS

[Say85] Jamie Sayen. Einstein in America: the scientist’s conscience in the age of Hitler and Hiroshima. Crown, New York, NY, USA, 1985. ISBN 0-517-55604-9. ix + 340 pp. LCCN QC16.E5 S285 1985.

Schlick:1920:STC

[SB20] Moritz Schlick and Henry L. (Henry Leopold) Brose. Space and time in contemporary physics, an introduction to the Theory of Relativity and Gravitation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1920. x + 89 pp. LCCN QC6 .S3 1920a.

Salomon-Bayet:1982:BNP

[SB82] Claire Salomon-Bayet. Bacteriology and Nobel Prize selections, 1901–1920. In Bernhard et al. [BCS82], pages 377–400. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500293.

Schlick:2006:STC

[SB06a] Moritz Schlick and Henry L. (Henry Leopold) Brose. Space and time in contemporary physics: an introduction to the Theory of Relativity and Gravitation. Prometheus Books, Amherst, NY, USA, 2006. ISBN 1-59102-417-X (paperback). 98 pp. LCCN

QC173.59.S65 S35 2006. URL http://www.loc.gov/catdir/
toc/ecip0610/2006008266.html.

Schlick:2006:RZG

[SB06b] Moritz Schlick and Henry Leopold Brose. Raum und Zeit in der gegenwärtigen Physik. (German) [Space and time in contemporary physics: an introduction to the Theory of Relativity and Gravitation]. Prometheus Books, Amherst, NY, USA, 2006. ISBN 1-59102-417-X (paperback). 98 pp.

LCCN QC173.59.S65. URL http://www.loc.gov/catdir/
toc/ecip0610/2006008266.html.

Schlagberger:2016:WID

[SBB16] Elisabeth Maria Schlagberger, Lutz Bornmann, and Johann Bauer. At what institutions did Nobel laureates do their prizewinning work? An analysis of biographical information on Nobel laureates from 1994 to 2014. Scientometrics, 109(2):723– 767, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11192-016-2059-2.pdf.

Stephenson:2000:UUI

[SBF00] Bruce Stephenson, Marvin Bolt, and Anna Felicity Friedman. The universe unveiled: instruments and images through history. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-79143-X (hardcover). 152 pp.

LCCN QB86 .S74 2000. URL http://www.loc.gov/catdir/
description/cam021/00059883.html;   http://www.loc.

gov/catdir/enhancements/fy0731/00059883-b.html;   http:
//www.loc.gov/catdir/toc/cam021/00059883.html.

Szilard:1960:SWSb

[SBSW60] Leo Szilard, Hans A. (Hans Albrecht) Bethe, Lewis L. Strauss, and Jerome B. (Jerome Bert) Wiesner, editors. Small world: Szilard, Bethe, Strauss, Wiesner [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? 28 minute 16mm film.

Scarry:2014:TMC

[Sca14] Elaine Scarry. Thermonuclear monarchy: choosing between democracy and doom. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-08008-0 (hardcover). 582 pp. LCCN JX1974.7 .S2177 2014.

Schlick:1915:PBR

[Sch15] Moritz Schlick. Die philosophische Bedeutung des Relativitätsprinzips. (German) [The philosophical meaning of the Principle of Relativity]. Zeitschrift für Philosophie und philosophische Kritik, 159(??):129–175, 1915. English translation in [SMvdVS79].

Schlick:1917:RZG

[Sch17] Moritz Schlick. Raum und Zeit in den gegenwärtigen Physik. Zur Einführung in das Verständnis der allgemeinen Relativitätstheorie. (German) [Space and time in contemporary physics. On the introduction to the understanding of the Theory of General Relativity]. Julius Springer, Berlin, Germany, 1917. ???? pp. LCCN ???? English translation of the third edition in [Bro20a].

Schlick:1919:RZG

[Sch19] Moritz Schlick. Raum und Zeit in der gegenwärtigen Physik: zur Einführung in das Verständnis der Relativitäts- und Gravitationstheorie. (German) [Space and time in contemporary physics. On the introduction to the understanding of the Theory of Relativity and Gravitation]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1919. iv + 85 pp. LCCN QC173.59.S65.

Schlick:1920:RZG

[Sch20] Moritz Schlick. Raum und Zeit in der gegenwärtigen Physik: zur Einführun in das Verständnis der Relativitäts- und Gravitationstheorie. (German) [Space and time in contemporary physics. On the introduction to the understanding of the Theory of Relativity and Gravitation]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1920. v + 89 pp. LCCN ????

Schlick:1921:KED

[Sch21] Moritz Schlick. Kritizistische oder empiristische Deutung der neuen Physik (German) [Critical or empirical interpretation of the new physics]. Kant-Studien, 26(??):96–111, 1921. English translation in [SMvdVS79, vol. 1, pp. 322–334].

Schlick:1922:RZG

[Sch22a] Moritz Schlick. Raum und Zeit in der gegenwärtigen Physik: zur Einführun in das Verständnis der Relativitäts- und Gravitationstheorie. (German) [Space and time in contemporary physics. On the introduction to the understanding of the Theory of Relativity and Gravitation]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1922. iv + 106 pp. LCCN ????

Schott:1922:STE

[Sch22b] G. A. Schott. Some terrestrial experiments on gravitation and Einstein’s theory. Nature, 109(2726):106, January 26, 1922. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v109/n2726/
pdf/109106b0.pdf.

Schalit:1927:SRT

[Sch27] Tuvia Schalit. Di spetsyele relativitets-teorye. Aynshtayns shite un Minkovskis ‘velt’. (Yiddish) [The Special Theory of Relativity: Einstein’s Theory and Minkowski’s ‘World’]. Eygener farlag, Berlin, Germany, 1927. ?? pp. LCCN ???? Self published. Preface by Albert Einstein.

Schrodinger:1928:FLW

[Sch28] Erwin Schrödinger. Four lectures on wave mechanics: delivered at the Royal Institution, London, on 5th, 7th, 12th, and 14th March, 1928. Blackie, Glasgow, Scotland, 1928. viii + 53 pp. LCCN QA927 .S33 1928a.

Schrodinger:1934:GWG

[Sch34] Erwin Schrödinger. Der Grundgedanke der Wellenmechanik. (German) [Der Grundgedanke der Wellenmechanik], pages 19– 36. Verlag von S. Hirzel, Leipzig, Germany, 1934.

Schrodinger:1935:GWG

[Sch35a] Erwin Schrödinger. Der Grundgedanke der Wellenmechanik. (German) [The basic idea of wave mechanics]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 1–13. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1935:DPR

[Sch35b] Erwin Schrödinger. Discussion of probability relations between separated systems. Mathematical proceedings of the Cambridge Philosophical Society, 31(4):555–563, October 1, 1935. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064

(electronic). URL http://www.jiscjournalarchives.ac.uk/
openurl.html?ref=cup/PSP31_04/S0305004100013554a.pdf.

Schrodinger:1935:ESA

[Sch35c] Erwin Schrödinger. Erwin Schrödinger [autobiographical sketch]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 86–88. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1935:TGT

[Sch35d] Erwin Schrödinger. Trinkspruch. (German) [Toast]. In Les Prix Nobel en 1933. (French) [The 1933 Nobel Prizes], pages 79–81. Norstedt & Söner, Stockholm, Sweden, 1935.

Schrodinger:1936:PRB

[Sch36] Erwin Schrödinger. Probability relations between separated systems. Mathematical proceedings of the Cambridge Philosophical Society, 32(3):446–452, October 1, 1936. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Schrodinger:1944:WLP

[Sch44] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1944. viii + 91 pp. LCCN QH331 .S355 1945. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schrodinger:1945:WLP

[Sch45] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1945. viii + 91 pp. LCCN QH331 .S355 1945. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schilpp:1949:AEPd

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein als Philosoph und Naturforscher: eine Auswahl. (German) [Albert Einstein as philosopher and scientist: a selection]. Kohlhammer, Stuttgart, West Germany, 1949. xv + 539 pp. LCCN ????

Schilpp:1949:AEPa

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49c] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schilpp:1949:AEPc

[Sch49d] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 3. Library of Living Philosophers, Evanston, IL, USA, 1949. xvi + 781 pp. LCCN QC16.E5 S3.

Schilpp:1949:CLP

[Sch49e] Paul Arthur Schilpp. Chronological list of principal works. In Albert Einstein: Philosopher-Scientist [Sch49c], pages 759–760. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schouten:1949:MFC

[Sch49f] J. A. Schouten. On meson fields and conformal transformations. Reviews of Modern Physics, 21(3):421–424, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.421;   http://rmp.aps.org/

abstract/RMP/v21/i3/p421_1. Special issue in honor of Albert Einstein on his 70th birthday.

Schlegel:1950:BRL

[Sch50a] Richard Schlegel. Book review: Leopold Infeld, Albert Einstein: His Work and Its Influence On Our World. American Journal of Physics, 18(5):327, May 1950. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/18/327/1.

Schuck:1950:NAF

[Sch50b]  Henrik Schück.  Nobelprisen 50 år: forskare, diktare,

fredskämpar. Nobelstiftelsen, Stockholm, Sweden, 1950. 512 + 11 pp. LCCN ????

Schilpp:1951:AED

[Sch51a] Paul Arthur Schilpp, editor. Albert Einstein als Philosoph und Naturforscher: eine Auswahl. (German) [Albert Einstein as philosopher and scientist: a selection]. Philosophen des 20. Jahrhunderts. Kohlhammer, Stuttgart, Germany, 1951. xv + 539 pp. LCCN ????

Schilpp:1951:AEP

[Sch51b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist. Tudor Publishing Co., New York, NY, USA, second edition, 1951. xvi + 781 pp. LCCN QC16.E5 S3 1951.

Schrodinger:1952:TQJb

[Sch52a] E. Schrödinger. Are there quantum jumps ?: Part II. The British Journal for the Philosophy of Science, III(11):233–242, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/III/11/233.full.pdf+html.

Schrodinger:1952:TQJa

[Sch52b] E. Schrödinger. Are there quantum jumps?: Part I. The British Journal for the Philosophy of Science, III(10):109–123, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/III/10/109.full.pdf+html.

Schrodinger:1953:RQT

[Sch53] E. Schrödinger. Relativistic quantum theory. The British Journal for the Philosophy of Science, IV(16):328–329, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/IV/16/328.full.pdf+html.

Schwartz:1954:RBB

[Sch54] Harry Schwartz. Red bloc building own atomic ‘pool’: Soviet has a virtual monopoly and benefits most — China may have reactor plant. New York Times, ??(??):3, September 7, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112937767.

Schilpp:1955:AEP

[Sch55] Paul Arthur Schilpp, editor. Albert Einstein als Philosoph und Naturforscher: eine Auswahl. (German) [Albert Einstein as philosopher and scientist: a selection], volume 1 of Philosophen des 20. Jahrhunderts. Kohlhammer, Stuttgart, West Germany, 1955. xv + 539 pp. LCCN ???? URL http://www.gbv.de/

dms/faz-rez/560630_FAZ_0037_5_0001.pdf.

Schrodinger:1956:EU

[Sch56a] Erwin Schrödinger. Expanding universes. Cambridge University Press, Cambridge, UK, 1956. viii + 93 pp. LCCN QB500 .S36.

Schrodinger:1956:WLOa

[Sch56b] Erwin Schrödinger. What is life? and other scientific essays. Cambridge University Press, Cambridge, UK, 1956. viii + 263 pp. LCCN QH331 .S355 1956.

Schrodinger:1956:WLOb

[Sch56c] Erwin Schrödinger. What is life? and other scientific essays. Doubleday, Garden City, NY, 1956. ???? pp. LCCN QH331 .S355 1956.

Schrodinger:1956:WLO

[Sch56d] Erwin Schrödinger. What is life? and other scientific essays. Cambridge University Press, Cambridge, UK, 1956. viii + 263 pp. LCCN QH331 .S355 1956.

Schrodinger:1957:STMa

[Sch57] Erwin Schrödinger. Science theory and man. George Allen and Unwin ltd., London, UK, 1957. xxiv + 2 + 27–223 pp.

LCCN Q171 .S3 1957a. URL http://www.zbp.univie.ac.at/
schrodinger/ebibliographie/publications.htm. Formerly

published under the title Science and the human temperament. Translated from German by James Vincent Murphy W. H. Jonston. Foreword by Ernest Rutherford.

Schilling:1961:HBD

[Sch61] Warner Schilling. The H-bomb decision: How to decide without actually choosing. Political science quarterly, 76(??):24–46, March 1961.

Schrodinger:1962:WLP

[Sch62a] Erwin Schrödinger. What is life? The physical aspect of the living cell. Cambridge University Press, Cambridge, UK, 1962. viii + 91 pp. LCCN QH331 .S355 1962. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Nobelstiftelsen:1962:NMH

[Sch62b] Henrik Schück. Nobel, the man and his prizes. Elsevier, Amsterdam, The Netherlands, second edition, 1962. x + 690 pp. LCCN AS911 .N7553 1962.

Schlick:1963:STC

[Sch63] Moritz Schlick. Space and time in contemporary physics; an introduction to the Theory of Relativity and Gravitation. Dover, New York, NY, USA, 1963. 89 pp. LCCN QC6 .S3 1963. Rendered into English by Henry L. Brose, with an introd. by F. A. Lindemann.

Scheinman:1965:AEP

[Sch65] Lawrence Scheinman. Atomic energy policy in France under the Fourth Republic. Princeton University Press, Princeton, NJ, USA, 1965. xxiv + 259 pp. LCCN HD9698.F72 S3.

Schrodinger:1967:WLP

[Sch67] Erwin Schrödinger. What is life? The physical aspect of the living cell & Mind and matter. Cambridge University Press, Cambridge, UK, 1967. 178 pp. LCCN QH331 .S355 1967. Based on lectures delivered under the auspices of the Institute at Trinity College, Dublin, in February 1943.

Schilpp:1970:AEP

[Sch70] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher– Scientist. The Library of Living Philosophers. Open Court Publishing Co., La Salle, IL, USA, third edition, 1970. xviii + 781 pp. LCCN ????

Schlegel:1971:EPR

[Sch71] Richard Schlegel. The Einstein–Podolsky–Rosen paradox. American Journal of Physics, 39(4):458, 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://scitation.aip.org/content/aapt/journal/
ajp/39/4/10.1119/1.1986185. See [Hoo70].

Schaffner:1974:EVL

[Sch74] Kenneth F. Schaffner. Einstein versus Lorentz: Research programmes and the logic of comparative theory evaluation. The British Journal for the Philosophy of Science, 25(1):45–78, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/25/1/45.full.pdf+html;
http://links.jstor.org/sici?sici=0007-0882%28197403%
2925%3A1%3C45%3AEVLRPA%3E2.0.CO%3B2-B.

Schwartz:1975:LFS

[Sch75] H. M. Schwartz. On the logical foundations of special relativity. American Journal of Physics, 43(4):362–364, 1975. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/43/4/10.1119/1.10083.

Schwartz:1977:ECEa

[Sch77a] H. M. Schwartz. Einstein’s comprehensive 1907 essay on relativity, Part I. American Journal of Physics, 45(6):512– 517, 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/45/6/10.1119/1.10949.

Schwartz:1977:ECEb

[Sch77b] H. M. Schwartz. Einstein’s comprehensive 1907 essay on relativity, Part II. American Journal of Physics, 45(9):811– 817, 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/45/9/10.1119/1.11053.

Schwartz:1977:ECEc

[Sch77c] H. M. Schwartz. Einstein’s comprehensive 1907 essay on relativity, Part III. American Journal of Physics, 45(10):899– 902, 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/45/10/10.1119/1.10743.

Schwartz:1977:EFP

[Sch77d] H. M. Schwartz. Einstein’s first paper on relativity. American Journal of Physics, 45(1):18–25, 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/scitation.aip.org/content/aapt/journal/ajp/45/1/10.
1119/1.10927.

Schwartz:1978:FRB

[Sch78] John Schwartz. Film review: A Portrait of Albert Einstein, produced by Francis S. Lestingi. American Journal of Physics, 46 (7):780, July 1978. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?
AJP/46/780/1.

Schilpp:1979:AEP

[Sch79a] Paul Arthur Schilpp, editor. Albert Einstein als Philosoph und Naturforscher: eine Auswahl. (German) [Albert Einstein as philosopher and scientist: a selection]. Philosophen des 20. Jahrhunderts. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1979. ISBN 3-528-08427-8. xv + 539 pp. LCCN ????

Schilpp:1979:ANA

[Sch79b] Paul Arthur Schilpp, editor. Autobiographical notes: [Albert Einstein]. Open Court Publishing Company, La Salle, IL, USA, 1979. ISBN 0-87548-352-6, 0-8126-9179-2 (paperback). 95 pp. LCCN QC16.E5 A3 1991.

Schlegel:1979:BRB

[Sch79c] Richard Schlegel. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. American Scientist, 67(5):623, September 1979. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
pdfplus/27849521.pdf.

Schwartz:1979:EIP

[Sch79d] Joe Schwartz. Einstein’s influence on politics and philosophy. Nature, 281(5729):324–325, September 27, 1979. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v281/n5729/
pdf/281324a0.pdf.

Schlicker:1981:AEP

[Sch81a] Wolfgang Schlicker. Albert Einstein: Physiker und Humanist, volume 26 of Illustrierte historische Hefte. Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1981. ISBN ???? 43 pp. LCCN ????

Schwartz:1981:EPD

[Sch81b] Joe Schwartz. Einstein pour débutants. (French) [Einstein for beginners]. Boréal express, Montréal, Québec, Canada, 1981. ISBN 2-89052-027-7. 169 pp. US$9.95.

Schrodinger:1982:CPW

[Sch82a] Erwin Schrödinger. Collected papers on wave mechanics: together with his four lectures on wave mechanics. Chelsea, New York, NY, USA, third edition, 1982. ISBN 0-8284-1302-9. xiii + 207 pp. LCCN QC174.2.

Schwarz:1982:MPR

[Sch82b] Boris Schwarz. Musical and personal reminiscences of Albert Einstein. In Holton and Elkana [HE82], pages 409–416. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Schilpp:1983:AEP

[Sch83a] Paul Arthur Schilpp, editor. Albert Einstein als Philosoph und Naturforscher: eine Auswahl. (German) [Albert Einstein as philosopher and scientist: a selection], volume 8 of Facetten der Physik. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1983. ISBN 3-528-08538-X. 249 pp. LCCN ???? URL http:/

/www.gbv.de/dms/bs/toc/025261541.pdf.

Schmutzer:1983:PNI

[Sch83b] Ernst Schmutzer, editor. Proceedings of the Ninth International Conference on General Relativity and Gravitation, Jena, 14–16 July 1980. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24669-5. LCCN QC173.6 .I57 1980.

Schilpp:1984:NAA

[Sch84a] Paul Arthur Schilpp, editor. Notas autobiográficas: [Albert Einstein], volume 1005 of El libro de Bolsillo. Humanidades. Alianza, Madrid, Spain, 1984. ISBN 84-206-0005-9. 87 pp. LCCN QC16.E5 S3418 1995. Translation of [Sch79a] to Spanish by Miguel Paredes.

Schroder:1984:AEH

[Sch84b] W. Schroder. Albert Einstein — historical and cultural perspectives — G. Holton, Y. Elkana. Historische Zeitschrift, 238(1): 108–109, ???? 1984. ISSN 0018-2613.

Schutz:1985:FCG

[Sch85] Bernard F. Schutz. A First Course in General Relativity. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-25770-0, 0-521-27703-5 (paperback). xiv + 376 pp. LCCN QC173.6 .S38 1985. URL http://www.loc.

gov/catdir/description/cam022/83023205.html;   http:
//www.loc.gov/catdir/enhancements/fy0729/83023205-b.
html;   http://www.loc.gov/catdir/toc/cam028/83023205.
html.

Schwinger:1986:ELU

[Sch86] Julian Seymour Schwinger. Einstein’s legacy: the unity of space and time, volume 16 of Scientific American Library series. Scientific American Library, New York, NY, USA, 1986. ISBN 0-7167-5011-2. xii + 250 pp. LCCN QC173.59.S65 S39 1986.

Schabas:1987:BRB

[Sch87a] Margaret Schabas. Book review: Lives of the Laureates: Seven Nobel Economists by William Breit; Roger W. Spencer. Isis, 78 (3):464–465, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232040.

Schoenberg:1987:ASA

[Sch87b] E. R. Schoenberg. Arnold Schoenberg and Albert Einstein, their relationship and views on Zionism. Journal of the Arnold Schoenberg Institute, 10(2):134–187, November 1987. ISSN 0146-5856.

Schrodinger:1987:ENG

[Sch87c] Annemarie Schrödinger. Erinnerungen an den Nobelpreis. (German) [Memories of the Nobel Prize]. In Erwin Schrödinger 1887–1961: Dokumente, Materialien und Bilder [KDK87], pages 84–?? ISBN 3-900538-09-3. LCCN QC16.S265 D65 1987. An exhibition of the Österreichischen Zentralbibliothek für Physik. With contributions by Nándor Balázs and others.

Schroder:1988:AER

[Sch88a] W. Schroder. Albert Einstein relationship with members of the Göttingen Gesellschaft der Wissenschaften. Archive for History of Exact Sciences, 39(2):157–171, ???? 1988. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Schroder:1988:AES

[Sch88b] Wilfried Schröder. Albert Einstein in seinen Beziehungen zu Mitgliedern der Gesellschaft der Wissenschaften in Göttingen. (German) [Albert Einstein and his relations with members of the Göttingen Scientific Society]. Archive for History of Exact Sciences, 39(2):157–171, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=157.

Schutz:1988:SWR

[Sch88c] Bernard Schutz. Strong and weak relativity. Physics World, 1(12):38, December 1988. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/1/12/phwv1i12a27.

pdf. Review of Flat and Curved Space–times, by G. F. R. Ellis and R. M. Williams, 1988, Oxford University Press, 368pp, £40.00 hb £17.50 pb. See rebuttal [Ell89b].

Schrodinger:1989:ST

[Sch89] Erwin Schrödinger. Statistical thermodynamics. Dover, New York, NY, USA, 1989. ISBN 0-486-66101-6. 95 pp. LCCN QC311.5 .S36 1989. US$3.95. URL http://www.loc.gov/

catdir/description/dover032/89016929.html.

Schweber:1990:YJC

[Sch90] S. S. Schweber. The young John Clarke Slater and the development of quantum chemistry. Historical Studies in the Physical and Biological Sciences, 20(2):339–406, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL

http://www.jstor.org/stable/27757647.

Schutz:1991:VSG

[Sch91] Bernard Schutz. Violin strings and general relativity. Physics World, 4(8):24, August 1991. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/4/8/phwv4i8a24.
pdf.

Schrodinger:1992:WLP

[Sch92a] Erwin Schrödinger. What is life?: the physical aspect of the living cell; with, Mind and matter; and Autobiographical sketches. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-42708-8 (paperback). viii + 184 pp. LCCN QH331 .S357 1992. URL http://www.loc.

gov/catdir/description/cam024/91028415.htm;   http:/
/www.loc.gov/catdir/toc/cam028/91028415.htm.

Schwartz:1992:AEM

[Sch92b] J. Schwartz. Albert Einstein and Mileva Marić — the Love letters — J. Renn, R. Schulmann. Nature, 360(6402):377–378, November 26, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Schwartz:1992:ELB

[Sch92c] Joseph Schwartz. Einstein as lover: Book review: Albert Einstein and Mileva Marić: The Love Letters. Nature, 360(6402): 377–378, November 26, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v360/n6402/pdf/360377a0.pdf.

Schmid:1993:AEL

[Sch93a] B. Schmid. Albert Einstein letters to Thomas Mann. Musik in Bayern, ??(46):5–16, ???? 1993. ISSN 0937-583X.

Schulmann:1993:EPO

[Sch93b] Robert Schulmann. Einstein at the Patent Office: Exile, salvation, or tactical retreat? Science in Context, 6(1):17–24, March 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schutz:1993:CIS

[Sch93c] Bernard Schutz. Choice introduction to special relativity. Physics World, 6(6):57–58, June 1993. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/

6/6/phwv6i6a26.pdf. Review of Spacetime Physics, by E. F. Taylor and J. A. Wheeler, (2nd edn), 1992, W. H. Freeman, 312pp, £14.95pb.

Schweber:1994:QMW

[Sch94] S. S. (Silvan S.) Schweber. QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). xxviii + 732 pp. LCCN QC680 .S34 1994. US$72.50. URL http://www.

loc.gov/catdir/description/prin021/93033550.html;
http://www.loc.gov/catdir/toc/prin031/93033550.html.

Schon:1995:PLA

[Sch95a] I. Schon. The pursuer of light — Albert Einstein — B. Deswaan. Journal of Adolescent & Adult Literacy, 39(3):261, November 1995. ISSN 1081-3004.

Schulmann:1995:PCE

[Sch95b] Robert Schulmann. From periphery to center: Einstein’s path from Bern to Berlin (1902–1914). In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 10, pages 259–271. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. LCCN ???? URL https://link.springer.

com/chapter/10.1007/978-94-011-0217-9_10.

Schirmacher:1996:GES

[Sch96a] Wolfgang Schirmacher, editor. German essays on science in the 20th century, volume 82 of German library. Continuum, New York, NY, USA, 1996. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). xviii + 314 pp. LCCN QC9.G3 G47 1996.

Schon:1996:PLA

[Sch96b] I. Schon. The pursuer of light: Albert Einstein — Spanish — B. deSwaan. Journal of Adolescent & Adult Literacy, 40(1):77, September 1996. ISSN 1081-3004.

Schramm:1996:BBO

[Sch96c] David N. Schramm. The Big Bang and other explosions in nuclear and particle astrophysics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2024-3, 981-02-2025-1 (paperback). xix + 710 pp. LCCN QB464 .S37 1995.

Schrodinger:1999:FIW

[Sch99a] Erwin Schrödinger. The fundamental idea of wave mechanics. Resonance, 4(2):92–103, ???? 1999. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.

springerlink.com/content/x25558467pq78238/fulltext.
pdf.

Schulman:1999:BHT

[Sch99b] Eric Schulman. A briefer history of time: from the Big Bang to the Big Mac. W. H. Freeman and Company, New York, NY, USA, 1999. ISBN 0-7167-3389-7. xiv + 171 pp. LCCN QB982 .S38 1999.

Schapiro:2000:UPA

[Sch00a] Meyer Schapiro. The unity of Picasso’s art. George Braziller, New York, NY, USA, 2000. ISBN 0-8076-1479-3. 199 pp. LCCN

N6853.P5 S33 2000. URL http://www.loc.gov/catdir/toc/
fy034/00034318.html.

Scheibe:2000:OSR

[Sch00b] Erhard Scheibe. The origin of scientific realism: Boltzmann, Planck, Einstein. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00], chapter 2, pages 31–44. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-015-9391-5_2.

Schmidt:2000:EFE

[Sch00c] Bernd G. Schmidt. Einstein’s field equations and their physical implications: selected essays in honour of Jürgen Ehlers, volume 540 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67073-4, 3-540-46580-4 (e-book). ISSN 0075-8450. xiii + 433 pp. LCCN QC173.6 .E39 2000.

URL http://rave.ohiolink.edu/ebooks/ebc/3540465804;
 http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-67073-5;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=1616-6361&volume=540.

Schweber:2000:SBB

[Sch00d] S. S. (Silvan S.) Schweber. In the shadow of the bomb: Bethe, Oppenheimer, and the moral responsibility of the scientist. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-04989-0 (hardcover). xviii + 260 + 8 pp. LCCN QC774.O56 S32 2000.

Scheideler:2002:SMA

[Sch02a] Britta Scheideler. The scientist as moral authority: Albert Einstein between elitism and democracy, 1914–1933. Historical Studies in the Physical and Biological Sciences, 32(2):319– 346, March 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Scheubel:2002:AER

[Sch02b] R. S. Scheubel. Albert Einstein and Riemann’s ‘Musiklexicon’, 1938–1952. Musikforschung, 55(4):386–394, October/December 2002. ISSN 0027-4801.

Schwinger:2002:ELU

[Sch02c] Julian Seymour Schwinger. Einstein’s legacy: the unity of space and time. Dover, New York, NY, USA, 2002. ISBN 0-486-41974-6 (paperback). xii + 250 pp. LCCN QC173.59.S65

S38 2002. URL http://www.loc.gov/catdir/description/
dover031/2002073813.html.

Schirrmacher:2003:ETP

[Sch03a] Arne Schirrmacher. Experimenting theory: The proofs of Kirchhoff’s radiation law before and after Planck. Historical Studies in the Physical and Biological Sciences, 33(2):299–335, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Schrodinger:2003:CPW

[Sch03b] Erwin Schrödinger. Collected papers on wave mechanics: together with his four lectures on wave mechanics. AMS Chelsea Publishing, Providence, RI, USA, third edition, 2003. ISBN 0-8218-3524-6. xiii + 207 pp. LCCN QC174.2 .S3213 2003.

Schweber:2003:BRB

[Sch03c] Silvan Schweber. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. Physics Today, 56(5):59–60, May 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

org/resource/1/phtoad/v56/i5/p59_s1.

Schweber:2003:AEF

[Sch03d] Silvan S. Schweber. Albert Einstein and the founding of Brandeis University. In Ashtekar et al. [ACH^{+}03], pages 615–?? ISBN 1-4020-1284-5. LCCN Q174 .B67 vol. 234. URL http://

www.loc.gov/catdir/enhancements/fy0813/2003046757-d.
html;   http://www.loc.gov/catdir/enhancements/fy0813/
2003046757-t.html.

Schlenoff:2004:CE

[Sch04a] Daniel C. Schlenoff. A century of Einstein. Scientific American, 291(3):102–105, September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v291/n3/
pdf/scientificamerican0904-102.pdf.

Schultz:2004:NDD

[Sch04b] Steven Schultz. Newly discovered diary chronicles Einstein’s last years. Princeton Weekly Bulletin, 93(25):??, April 26, 2004.

URL https://pr.princeton.edu/pwb/04/0426/.

Schumm:2004:DTB

[Sch04c] Bruce A. Schumm. Deep down things: the breathtaking beauty of particle physics. The Johns Hopkins University Press, Baltimore, MD, USA, 2004. ISBN 0-8018-7971-X (hardcover). x + 378 pp. LCCN QC793.2 .S35 2004. URL http://www.

loc.gov/catdir/bios/jhu052/2004050710.html;   http:
//www.loc.gov/catdir/description/jhu051/2004050710.
html.

Scheideler:2005:DET

[Sch05a] B. Scheideler. Democratic and elitist thinking in Albert Einstein. On the political biography of a moralist in the Weimar Republic. Vierteljahrshefte für Zeitgeschichte, 53(3):381–38+, July 2005. ISSN 0042-5702.

Schemmel:2005:ARG

[Sch05b] Matthias Schemmel. An astronomical road to General Relativity: The continuity between classical and relativistic cosmology in the work of Karl Schwarzschild. Science in Context, 18(3): 451–478, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schirrmacher:2005:DMA

[Sch05c] Arne Schirrmacher, editor. Dreier Männer Arbeit in der frühen Bundesrepublik: Max Born, Werner Heisenberg und Pascual Jordan als politische Grenzgänger. (German) [Three men’s work in the early Federal Republic: Max Born, Werner Heisenberg and Pascual Jordan as political border crossers], volume 296 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wissenschaftsgeschichte, Berlin, Germany, 2005. 50 + 6 pp. LCCN ????

Schlick:2005:STC

[Sch05d] Moritz Schlick. Space and time in contemporary physics: an introduction to the Theory of Relativity and Gravitation. Dover, New York, NY, USA, 2005. ISBN 0-486-44283-7 (paperback). x + 87 pp. LCCN QC173.59.S65 S3513 2005. URL http://www.

loc.gov/catdir/enhancements/fy0618/2004059121-d.html.

Schwarzenbach:2005:VGA

[Sch05e] Alexis Schwarzenbach. Das verschmähte Genie: Albert Einstein und die Schweiz / Alexis Schwarzenbach. (German) [The spurned genius: Albert Einstein and Switzerland]. Deutsche Verlags-Anstalt, Munich, Germany, 2005. 215 (est.) pp. LCCN ????

Schwarzenbach:2005:GDA

[Sch05f] Alexis Schwarzenbach. Le génie dédaigné: Albert Einstein et la Suisse. (French) [The despised genius: Albert Einstein and Switzerland]. Metropolis, Genève, Switzerland, 2005. ISBN 2-88340-158-6. 257 + 8 pp. LCCN QC16.E5 S32814 2005.

Schweber:2006:EOI

[Sch06] Silvan S. Schweber. Einstein and Oppenheimer: Interactions and intersections. Science in Context, 19(4):513–559, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schell:2007:SDN

[Sch07a] Jonathan Schell. The seventh decade: the new shape of nuclear danger. The American empire project. Metropolitan Books/Henry Holt, New York, NY, USA, 2007. ISBN 0-8050-8129-1. 251 pp. LCCN U264.3 .S43 2007. US$24.00. URL http:

//www.loc.gov/catdir/enhancements/fy0711/2007014238-
b;   http://www.loc.gov/catdir/enhancements/fy0711/
2007014238-b.html;   http://www.loc.gov/catdir/enhancements/
fy0711/2007014238-d;   http://www.loc.gov/catdir/enhancements/
fy0711/2007014238-d.html;   http://www.loc.gov/catdir/
enhancements/fy0809/2007014238-s;   http://www.loc.
gov/catdir/toc/ecip0715/2007014238;   http://www.loc.
gov/catdir/toc/ecip0715/2007014238.html.

Schemmel:2007:CBC

[Sch07b] Matthias Schemmel. The Continuity Between Classical and Relativistic Cosmology in the Work of Karl Schwarzschild, pages 1081–1108. Volume 250(3) of Renn et al. [RJS07], 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN QC173.6 .G469 2007. URL http://www.

springerlink.com/content/v64669k31r250854/.

Schucking:2007:BRE

[Sch07c] Engelbert L. Schucking. Book review: Einstein: His Life and Universe, Walter Isaacson, Simon & Schuster, New York, 2007. $32.00 (675 pp.). ISBN 978-0-7432-6473-0. Physics Today, 60(11):59–61, November 1, 2007. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://physicstoday.scitation.org/doi/full/10.
1063/1.2812127.

Schwarzschild:2007:TRU

[Sch07d] Karl Schwarzschild. Things at rest in the universe. In Renn et al. [RSS^{+}07a], pages 183–190. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_17. Originally published as “Was in der Welt ruht” in Die Zeit, Vienna 11, No. 142, 181–183, 19 June 1897.

Schweber:2008:EOM

[Sch08] S. S. (Silvan S.) Schweber. Einstein and Oppenheimer: the meaning of genius. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02828-7. xiv + 412 pp. LCCN

QC16.E5 S3285 2008. URL http://www.loc.gov/catdir/toc/
ecip082/2007043108.html.

Schroder:2009:BRE

[Sch09a] Wilfried Schröder. Book review: Einsteins Kosmos. Untersuchungen zur Geschichte der Kosmologie, Relativitätstheorie und zu Einsteins Wirken und Nachwirken. (= Acta His. Astron. Vol. 27) by Hilmar W. Duerbeck, Wolfgang R. Dick. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 93(2):241–242, ???? 2009. CODEN SUARAH. ISSN 0039-4564. URL http:/

/www.jstor.org/stable/20778418.

Schutz:2009:FCG

[Sch09b] Bernard F. Schutz. A First Course in General Relativity. Cambridge University Press, Cambridge, UK, second edition, 2009. ISBN 0-521-88705-4. xv + 393 pp. LCCN QC173.6

.S38 1985. URL http://www.cambridge.org/catalogue/
catalogue.asp?isbn=9780521887052.

Scheubel:2010:CSS

[Sch10a] Robert Schmitt Scheubel, editor. “Chronique scandaleuse”: Schönberg, dieser Einstein der Musik: Kritiken zu Arnold Schönbergs Werken. (German) [Scandalous chronicle: Schoenberg, this Einstein of music: reviews of Arnold Schoenberg’s works]. Consassis.de, Berlin, Germany, 2010. ISBN 3-937416-23-4 (paperback). 241 pp. LCCN ML410.S283 C48 2010.

Schneir:2010:FVW

[Sch10b] Walter Schneir. Final verdict: what really happened in the Rosenberg case. Melville House, Brooklyn, NY, USA, 2010. ISBN 1-935554-16-6. 203 pp. LCCN HX84.R6 S27 2010. Preface and afterword by Miriam Schneir.

Schrodinger:2010:EU

[Sch10c] Erwin Schrödinger. Expanding universes. Cambridge University Press, Cambridge, UK, 2010. ???? pp. LCCN QB500 .S36.

Scharf:2012:APN

[Sch12a] Rainer Scharf. Ausgezeichnete Physik: der Nobelpreis und die Geschichte einer Wissenschaft. (German) [Excellent physics: the Nobel Prize and the History of Science]. Nobel und die Folgen. Bückle and Böhm, Regensburg, Germany, 2012. ISBN 3-941530-09-7 (hardcover), 3-941530-10-0. 303 pp. LCCN QC7 .S28 2012.

Anonymous:2012:APN

[Sch12b] Rainer Scharf, editor. Ausgezeichnete Physik: der Nobelpreis und die Geschichte einer Wissenschaft. (German) [Excellent physics: the Nobel Prize and the History of Science]. Bückle and Böhm, Regensburg, Germany, 2012. ISBN 3-941530-09-7 (hardcover). 303 pp. LCCN QC7 .S28 2012.

Schils:2012:AE

[Sch12c] René Schils. Albert Einstein. In How James Watt invented the copier: forgotten inventions of our great scientists [Sch12d], pages 123–129. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book). LCCN T15 .S35513 2012. URL http://link.

springer.com/chapter/10.1007/978-1-4614-0860-4_20.

Schils:2012:HJW

[Sch12d] René Schils. How James Watt invented the copier: forgotten inventions of our great scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book). vii + 170 pp. LCCN T15 .S35513 2012.

Schmidt:2012:NLA

[Sch12e] Brian P. Schmidt. Nobel Lecture: Accelerating expansion of the universe through observations of distant supernovae. Reviews of Modern Physics, 84(3):1151–1163, July 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.1151;   http://rmp.aps.org/abstract/RMP/v84/i3/p1151_
1;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/2011/.

Schwarzenbach:2012:VGA

[Sch12f] Alexis Schwarzenbach. Das verschmähte Genie: Albert Einstein und die Schweiz / Alexis Schwarzenbach. (German) [The spurned genius: Albert Einstein and Switzerland]. Heyne, Rolf, Collection, Munich, Germany, 2012. ISBN 3-89910-531-1. 220 (est.) pp. LCCN ????

Schewe:2013:MGP

[Sch13a] Phillip F. Schewe. Maverick Genius: the Pioneering Odyssey of Freeman Dyson. Thomas Dunn Books, St. Martin’s Press, New York, NY, USA, 2013. ISBN 0-312-64235-0 (hardcover), 1-250-02101-4 (e-book). x + 339 pp. LCCN QC16.D95 S34 2013.

Schroer:2013:EJC

[Sch13b] Bert Schroer. The Einstein–Jordan conundrum and its relation to ongoing foundational research in local quantum physics. European Physical Journal H, 38(2):137–173, March 2013. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30059-x.

Schaaf:2014:WBN

[Sch14a] Michael Schaaf. Weizsäcker, Bethe und der Nobelpreis. (German) [Weizsäcker, Bethe and the Nobel Prize]. In Hentschel and Hoffmann [HH14], pages 145–158. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Schulmann:2014:EP

[Sch14b] Robert Schulmann. Einstein’s politics. In Janssen and Lehner [JL14b], chapter 14, pages 421–454. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Schwarcz:2015:RCH

[Sch15a] Joe Schwarcz. The right chemistry: How the Manhattan Project and ‘the bomb’ came to be. The Montreal Gazette, ??(??):??, May 10, 2015. ISSN 0384-1294. URL http:

//montrealgazette.com/technology/science/the-right-
chemistry-how-the-manhattan-project-and-the-bomb-came-
to-be.

Schwichtenberg:2015:PS

[Sch15b] Jakob Schwichtenberg. Physics from Symmetry. Undergraduate lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-19200-0, 3-319-19201-9 (e-book). xix + 279 pp. LCCN QC174.17.S9 S4 2015.

Schick:2017:HNP

[Sch17a] Michael Schick. Historical note on 2016 Physics Nobel. Physics Today, 70(8):13, August 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schiesser:2017:SCM

[Sch17b] William E. Schiesser. Spline Collocation Methods for Partial Differential Equations: with Applications in R. Wiley, New York, NY, USA, 2017. ISBN 1-119-30103-3 (hardcover), 1-119-30105-X (PDF), 1-119-30104-1 (ePub), 1-119-30106-8 (online). xv + 549 pp. LCCN QA377 .S355 2017.

Scholten:2017:EAA

[Sch17c] Hans Scholten. Die Entfesselung der Abendländer Von Albertus Magnus bis Albert Einstein. (German) [The Unleashing of the Westerners From Albertus Magnus to Albert Einstein]. neobooks, München, Germany, 2017. ISBN 3-7380-9984-0. 182 pp. LCCN ????

Scholz:2019:CAB

[Sch19] Erhard Scholz. E. Cartan’s attempt at bridge-building between Einstein and the Cosserats — or how translational curvature became to be known as torsion. European Physical Journal H, 44 (1):47–75, February 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2018-90059-x.

Schneider:2020:WPC

[Sch20] Mike D. Schneider. What’s the problem with the cosmological constant? Philosophy of Science, 87(1):1–20, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schultz:2021:AYC

[Sch21] Isaac Schultz. After 69 years, chemists finally get a good look at einsteinium. Web site., February 3, 2021. URL https:

//gizmodo.com/after-69-years-chemists-finally-get-a-
good-look-at-ein-1846188962.

Sciama:1953:OI

[Sci53] D. W. Sciama. On the origin of inertia. Monthly Notices of the Royal Astronomical Society, 113(1):34–42, ???? 1953. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic).

Sciama:1961:IES

[Sci61] D. W. Sciama. On the interpretation of the Einstein– Schrödinger unified field theory. Journal of Mathematical Physics, 2(4):472–477, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i4/p472_s1.

Sciama:1979:C

[Sci79] Dennis W. Sciama. Cosmology. In Aichelburg and Sexl [AS79], pages 17–23. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66.

DM48.00. URL http://link.springer.com/chapter/10.
1007/978-3-322-91080-6_2/.

Sciama:1980:GRS

[Sci80] D. W. Sciama. General Relativity since Einstein: Book review: General Relativity:: An Einstein Centenary Survey, S. W. Hawking and W. Israel, eds. Cambridge University Press, New York, 1979. xviii, 920 pp., illus. $74.50. Science, 207(4431): 631–632, February 8, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/207/4431/631.full.pdf.

Sclove:1989:AAW

[Scl89] Richard E. Sclove. From alchemy to atomic war: Frederick Soddy’s “technology assessment” of atomic energy, 1900–1915. Science, Technology and Human Values, 14(2):163–194, Spring 1989. ISSN 0162-2439 (print),1552-8251 (electronic). URL

http://www.jstor.org/stable/690079.

Scotti:2015:EE

[Sco15] Camillo Maurizio Scotti. Einstein Evolution. Casa Editrice Kimerik, Patti, Italia, 2015. ISBN 88-6884-512-1. 172 pp. LCCN ????

Stachel:1989:SYW

[SCRS89] John Stachel, David C. Cassidy, Jürgen Renn, and Robert Schulmann, editors. The Swiss Years: Writings, 1900–1909, volume 2 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08526-9. xxxvi + 656 pp. LCCN QC16.E5 A2 1987. Translations from the German by Anna Beck.

Stachel:1987:EY

[SCS87] John J. Stachel, David C. Cassidy, and Robert Schulmann, editors. The Early Years, 1879–1902, volume 1 of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08407-6. lxvi + 433 pp. LCCN QC16.E5 A2 1987. US$52.50. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://einsteinpapers.press.princeton.edu/vol1-doc;
 http://press.princeton.edu/titles/2515.html;   http:
//www.loc.gov/catdir/description/prin021/86043132.html;
 http://www.loc.gov/catdir/toc/prin032/86043132.html.

Published with a translation into English by Anna Beck, ISBN 0-691-08475-0 [?], which is not available separately. Consultant for the translation: Peter Havas.

Smoot:1993:WT

[SD93] George Smoot and Keay Davidson. Wrinkles in time. W. Morrow, New York, NY, USA, 1993. ISBN 0-688-12330-9. vi + 331 pp. LCCN QB981 .S695 1993. US$25.00.

Sorensen:2001:MPE

[SDCZ01] A. Sørensen, L.-M. Duan, J. I. Cirac, and P. Zoller. Manyparticle entanglement with Bose–Einstein condensates. Nature, 409(6816):63–66, January 4, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v409/n6816/full/409063a0.
html.

Szilard:1928:EMF

[SE28] Leo Szilard and Albert Einstein. Electrodynamic movement of fluid metals particularly for refrigerating machines. British Patent No. 303,065 (filed December 24, 1928, issued May 26, 1930), December 24, 1928. Reprinted in [FS87, pages 540–542].

Szilard:1926:CBL

[SE45] Leo Szilard and Albert Einstein. [correspondence between Leo Szilard and Albert Einstein]. Personal and scientific correspondence between the authors., 1926–1945. URL http://library.

ucsd.edu/dc/object/bb24247539.

Samuel:1951:EP

[SE51] Viscount Herbert Louis Samuel and Albert Einstein. Essay in physics. Blackwell, Oxford, UK, 1951. 154 pp. LCCN QC 6 SAM.

Sass:1979:EWK

[SE79] Hans-Martin Sass and A. Einstein. Einstein über “wahre Kultur” und die Stellung der Geometrie im Wissenschaftssystem:

Ein Brief Albert Einsteins an Hans Vaihinger vom Jahr 1919. (German) [Einstein on “true culture” and the position of geometry in the science system: A letter from Albert Einstein to Hans Vaihinger of 1919]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):316– 319, ???? 1979. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/
25170513.

Stavis:1990:CBB

[SE90] Barrie Stavis and Albert Einstein. Correspondence between Barrie Stavis and Albert Einstein. Cardozo Studies in Law and Literature, 2(2):??, Autumn–Winter 1990. ISSN 1043-1500 (print), 2324-6510 (electronic). URL http://www.jstor.org/

stable/743328.

Seaborg:1958:TE

[Sea58] Glenn Theodore Seaborg. The transuranium elements. Mrs. Hepsa Ely Silliman memorial lectures, Yale University. Yale University Press, New Haven, CT, USA, 1958. xx + 328 pp. LCCN QD172.T7 S38.

Seaborg:1972:NMC

[Sea72] Glenn Theodore Seaborg. Nuclear milestones: a collection of speeches. W. H. Freeman and Company, New York, NY, USA, 1972. ISBN 0-7167-0342-4. 390 pp. LCCN QC773.3.U5 S42.

Seaborg:1979:AER

[Sea79] Glenn T. Seaborg. Albert Einstein – a reflection. Bulletin of the Atomic Scientists, 35(3):20, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

See:1925:RNE

[See25] T. J. J. (Thomas Jefferson Jackson) See. Researches in non-Euclidian geometry and the theory of relativity: a systematic study of twenty fallacies in the geometry of Riemann, including the so-called curvature of space and radius of World curvature, and of eighty errors in the physical theories of Einstein and Eddington, showing the complete collapse of the theory of relativity. Naval Observatory, Mare Island, CA, USA, 1925. 237 pp. LCCN QA685 .S44 1925a.

Seelig:1952:AES

[See52] Carl Seelig. Albert Einstein und die Schweiz. (German) [Albert Einstein and Switzerland]. Europa-Verlag, Zurich, Switzerland and Vienna, Austria, 1952. 253 pp.

Seelig:1953:AEM

[See53] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View], volume 65 of Ullstein Buch. Europa Verlag, München, Germany and Zürich, Switzerland, 1953. 275 pp. LCCN ???? First published in 1934 in Amsterdam.

Seelig:1954:AED

[See54] Carl Seelig. Albert Einstein: eine dokumentarische Biographie. (German) [Albert Einstein: a documentary biography]. Europa Verlag, München, Germany and Zürich, Switzerland, 1954. 304 pp. LCCN ????

Seelig:1955:AEM

[See55] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View], volume 65 of Ullstein Buch. Verlag “Das goldene Vlies”, Zürich, Switzerland, 1955. 201 pp. LCCN ???? First published in 1934 in Amsterdam.

Seelig:1956:AED

[See56a] Carl Seelig. Albert Einstein: a documentary biography. Staples Press, London, UK, 1956. 240 pp. LCCN QC16.E5 S333 1956.

Seelig:1956:HZD

[See56b] Carl Seelig, editor. Helle Zeit — dunkle Zeit: in memoriam Albert Einstein. (German) [Bright Times — Dark Times: in memory of Albert Einstein]. Europa Verlag, München, Germany and Zürich, Switzerland, 1956. ISBN 3-322-84225-8. 171 pp. LCCN ???? With contributions and letters from Hedwig Born.

Seelig:1960:AEL

[See60] Carl Seelig. Albert Einstein: Leben und Werk: eines Genies unserer Zeit. (German) [Albert Einstein: Life and Work: a Genius of Our Time]. Europa Verlag, München, Germany and Zürich, Switzerland, 1960. 436 pp. LCCN ????

Seelig:1961:CVM

[See61] Carl Seelig. Como vejo o mundo [Albert Einstein]. (Portuguese) [As I See the World]. Emp. Nac. de Publicidade, Lisboa, Portugal, 1961. 328 pp. LCCN ???? Portuguese translation by Ruth San Payo Araújo of [See55].

Seelig:1964:AEM

[See64] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View], volume 65 of Ullstein Buch. Ullstein, Frankfurt a. M., Germany, 1964. ???? pp. LCCN ???? First published in 1934 in Amsterdam.

Seeger:1965:BRB

[See65a] Raymond J. Seeger. Book review: Albert Einstein and the Cosmic World Order, by Cornelius Lanczos. American Journal of Physics, 33(11):972, November 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/?AJP/33/972/1.

Seeger:1965:BRA

[See65b] Raymond J. Seeger. Book review: Albert Einstein and the cosmic world order, by Cornelius Lanczos, Interscience Publishers, vi + 139 pp. American Journal of Physics, 33(11):972, November 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Seelig:1980:AEM

[See80] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View], volume 65 of Ullstein Materialien. Ullstein, Frankfurt a. M., Germany, 1980. ISBN 3-548-35024-0. 200 pp. LCCN QC71 .E55 1980. First published in 1934 in Amsterdam.

Seelig:1986:AEM

[See86] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View]. Ullstein Materialien, Frankfurt a. M., Germany, 1986. ISBN 3-548-35024-0. 200 pp. LCCN ???? First published in 1934 in Amsterdam.

Seelig:1991:AEM

[See91] Carl Seelig, editor. Albert Einstein: Mein Weltbild. (German) [Albert Einstein: My World View]. Ullstein Materialien, Frankfurt a. M., Germany, 1991. ISBN 3-548-40138-4. 330 pp. LCCN ???? First published in 1934 in Amsterdam.

Segre:1970:EFPb

[Seg70] Emilio Segrè. Enrico Fermi: physicist. University of Chicago Press, Chicago, IL, USA and London, UK, 1970. ISBN 0-226-74472-8. x + 276 pp. LCCN QC16.F46 S4.

Segre:1976:PSN

[Seg76] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea. (Italian) [Personalities and discoveries in contemporary physics]. Biblioteca della EST. Edizioni scientifiche e tecniche Mondadori, Milano, Italia, 1976. 297 pp. LCCN QC7 .S44. Ciclo di lezioni tenute dal nov. 1972 fino al marzo 1973.

Segre:1980:XRQ

[Seg80] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. W. H. Freeman and Company, New York, NY, USA, 1980. ISBN 0-7167-1146-X, 0-7167-1147-8 (paperback). ix + 337 pp. LCCN QC7 .S4413. English translation of [Seg76].

Segre:1983:PSN

[Seg83] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea: dai raggi X ai quark. Biblioteca della EST, 0303-2752. Edizioni scientifiche e tecniche Mondadori (IS), Milano, Italia, second edition, 1983. 297 pp. LCCN ????

Segre:1984:PML

[Seg84] Emilio Segrè. Les physiciens modernes et leurs découvertes: des rayons X aux quarks. (French) [Modern physicists and their discoveries: from X-rays to quarks]. Temps des sciences. Fayard, Paris, France, 1984. 456 pp. LCCN QC7 .S4414 1984. French translation of [Seg76].

Segre:1986:MKR

[Seg86a] Emilio Segre. Mi-karne Ranotgen ove-ad kvarkim: fisikaim modernim ove-tagliyotehe. Keter, Yerushalayim, Israel, 1986. 331 pp. LCCN ???? Hebrew translation of [Seg76].

Segre:1986:WLM

[Seg86b] Emilio Segrè. Wu li ming ren he wu li fa xian. Zhi shi chu ban she, Shanghai, People’s Republic of China, 1986. iii + 366 pp.

LCCN QC7 .S4412. Mandarin Chinese translation by Zuwei Liu of [Seg76].

Segre:1987:RXA

[Seg87] Emilio Segre. Dos raios X aos quarks: fisicos modernos e suas descobertas. (Portuguese) [From X-rays to quarks: modern physicists and their discoveries], volume 24 of Pensamento cientifico. Ed. UnB, Brasilia, Brazil, 1987. ISBN 85-230-0078-X. 345 pp. LCCN ????

Segre:1989:DNF

[Seg89] Emilio G. Segrè. The discovery of nuclear fission. Physics Today, 42(7):43–48, July 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v42/i7/p38_s1.

Segala:1991:BRL

[Seg91] Marco Segala. Book review: Lewis Samuel Feuer, Einstein e la sua generazione. Nascita e sviluppo di teorie scientifiche, Bologna, Il Mulino, 1990 (≪Le occasioni≫ 35) ISBN 88-15-02809-9 (tr. it. di Gianfranco Ceccarelli di Einstein and the Generations of Science, New Brunswick, Transaction, Inc., 1982). Nuncius, 6(2):449–450, ???? 1991. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01541.

Segre:1995:ABE

[Seg95] Claudio G. Segrè. Atoms, bombs, and Eskimo kisses: a memoir of father and son. Viking, New York, NY, USA, 1995. ISBN 0-670-86307-6. xi + 287 pp. LCCN DG465.7.S43 A3 1995.

Segal:2003:MUN

[Seg03] Sanford L. Segal. Mathematicians under the Nazis. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-00451-X (hardcover). xxii + 530 pp. LCCN QA28

.S44 2003. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
prin031/2002070399.html;   http://www.pupress.princeton.
edu/titles/7558.html.

Segre:2007:XRQ

[Seg07a] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. Dover classics of science and mathematics. Dover, New York, NY, USA, 2007. ISBN 0-486-45783-4. ix + 339 pp. LCCN QC7 .S4413 2007. URL http://www.loc.gov/

catdir/enhancements/fy0702/2006102450-d.html.

Segre:2007:FCS

[Seg07b] Gino Segrè. Faust in Copenhagen: a struggle for the soul of physics. Viking, New York, NY, USA, 2007. ISBN 0-670-03858-X. x + 310 pp. LCCN QC15 .S427 2007. URL http://www.loc.

gov/catdir/enhancements/fy0743/2006052807-b.html;
http://www.loc.gov/catdir/enhancements/fy0743/2006052807-
d.html.

Segre:2011:OGM

[Seg11] Gino Segrè. Ordinary geniuses: Max Delbrück, George Gamow, and the origins of genomics and Big Bang cosmology. Viking, New York, NY, USA, 2011. ISBN 0-670-02276-4. xxi + 330 pp. LCCN QH31.D434 S44 2011.

Seidel:1979:BRB

[Sei79] Robert W. Seidel. Book review: Scientific Elite. Nobel Laureates in the United States by Harriet Zuckerman. Isis, 70(2):283–284, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230804.

Seidel:1984:BRB

[Sei84] Robert W. Seidel. Book review: Science, Technology, and Society in the Time of Alfred Nobel by Carl Gustaf Bernhard; Elisabeth Crawford; Per Sorbom. Isis, 75(4):727–728, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232424.

Seidel:1994:SEE

[Sei94] Edward Seidel. Solving Einstein’s equations on supercomputers. In Brown et al. [BCEP94], pages 529–531. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Seidel:1995:ADA

[Sei95] Robert W. Seidel. Los Alamos and the development of the atomic bomb. Otowi Crossing Press, Los Alamos, NM, USA, 1995. ISBN 0-9645703-0-0 (paperback). iv + 109 + 12 pp. LCCN TN22.5.L6 S45 1995.

Seidel:1999:GJL

[Sei99a] Robert W. Seidel. The Golden Jubilees of Lawrence Berkeley and Los Alamos National Laboratories. Osiris (Series 2), 14 (??):187–202, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301968.

Seitz:1999:AEB

[Sei99b] Jay A. Seitz. Albert Einstein’s brain. The Lancet (London, England), 354(9192):1822–1823, November 20, 1999. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547X (electronic).

Seife:2003:AOS

[Sei03a] Charles Seife. Alpha and Omega: the search for the beginning and end of the universe. Viking, New York, NY, USA, 2003. ISBN 0-670-03179-8. viii + 294 + 2 pp.

LCCN QB981 .S446 2003. URL http://www.loc.gov/catdir/
enhancements/fy0719/2002044853-b.html;   http://www.
loc.gov/catdir/enhancements/fy0719/2002044853-d.html;
 http://www.loc.gov/catdir/enhancements/fy0719/2002044853-
s.html;   http://www.loc.gov/catdir/enhancements/fy0719/
2002044853-t.html.

Seife:2003:RGW

[Sei03b] Charles Seife. Relativity goes where Einstein sneered to tread. Science, 299(5604):185, January 10, 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/299/5604/185.1.full.pdf.

Seife:2004:GPG

[Sei04] Charles Seife. Gravity probe to give Einstein a pricey highprecision test. Science, 304(5669):385, April 16, 2004. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/304/
5669/385.full.pdf.

Selleri:1978:CEL

[Sel78] F. Selleri. On the consequences of Einstein locality. Foundations of Physics, 8(1–2):103–116, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00708489.

Selleri:1988:QMV

[Sel88] Franco Selleri, editor. Quantum mechanics versus local realism: the Einstein–Podolsky–Rosen paradox. Physics of atoms and molecules. Plenum Press, New York, NY, USA; London, UK, 1988. ISBN 0-306-42739-7. xviii + 461 pp. LCCN QC174.12 .Q36 1988.

Schroeder:2003:EDP

[SEM03] Hans G. Schroeder, Albert Einstein, and Hans Mushham. Experimental determination of the pore size of filters. 1923. PDA journal of pharmaceutical science and technology, 57(5):331– 332, September/October 2003. ISSN 1079-7440.

Semendeferi:2008:LNC

[Sem08] Ioanna Semendeferi. Legitimating a nuclear critic: John Gofman, radiation safety, and cancer risks. Historical Studies in the Natural Sciences, 38(2):259–301, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Sen:2021:EFH

[Sen21] Paul Sen. Einstein’s fridge: how the difference between hot and cold explains the universe. Scribner, New York, NY, USA, 2021. ISBN 1-5011-8130-0 (hardcover), 1-5011-8132-7 (e-book). xii + 305 pp. LCCN QC311 .S5118 2021.

Serber:1943:AP

[Ser43] Robert Serber. The Los Alamos primer. Report LA-1, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April

1943. 24 pp. URL http://www.fas.org/sgp/othergov/doe/

lanl/docs1/00349710.pdf. Unclassified and publicly released on 19 December 1963.

Serwer:1977:UZP

[Ser77] Daniel Serwer. Unmechanischer Zwang: Pauli, Heisenberg, and the rejection of the mechanical atom, 1923–1925. Historical Studies in the Physical Sciences, 8(??):189–256, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757371.

Servos:1988:BRB

[Ser88] John W. Servos. Book review: Intellectual Mastery of Nature: Theoretical Physics from Ohm to Einstein. I. The Torch of Mathematics, 1800–1870. II. The Now Mighty Theoretical Physics, 1870–1925, by Christa Jungnickel and Russell McCormmach. Journal of Interdisciplinary History, 19(1):106–108, ???? 1988. ISSN 0022-1953 (print), 1530-9169 (electronic). URL

http://www.jstor.org/stable/204229.

Seth:2003:PPC

[Set03] Suman Seth. Principles and problems: Constructions of theoretical physics in Germany, 1890–1918. Ph.D. dissertation, Department of History, Princeton University, Princeton, NJ, USA, November 2003. xiii + 315 pp. URL https://search.

proquest.com/pqdtglobal/docview/288241277/.

Seth:2010:BRE

[Set10a] S. Seth. Book review: Einstein’s Generation: The Origins of the Relativity Revolution. Journal of Interdisciplinary History, 41(2):273–274, ???? 2010. CODEN ???? ISSN 0022-1953 (print), 1530-9169 (electronic). URL https://muse.jhu.edu/

article/391012.

Seth:2010:BRB

[Set10b] Suman Seth. Book review: Einstein’s Generation: The Origins of the Relativity Revolution by Richard Staley. Journal of Interdisciplinary History, 41(2):273–274, ???? 2010. ISSN 0022-1953 (print), 1530-9169 (electronic).

Seth:2011:BRB

[Set11] Suman Seth. Book review: Bruce J. Hunt: Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Isis, 102(4):782–783, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664874.

Severance:1999:EVS

[Sev99] John B. Severance. Einstein: visionary scientist. Clarion Books, New York, NY, USA, 1999. ISBN 0-618-15317-9 (electronic), 0-395-93100-2. 144 pp. LCCN QC16.E5 S394 1999. URL http:/

/www.loc.gov/catdir/description/hm022/98051396.html;
 http://www.netLibrary.com/urlapi.asp?action=summary&
v=1&bookid=56240.

Slater:1933:ITP

[SF33] John C. (John Clarke) Slater and Nathaniel Herman Frank. Introduction to theoretical physics. International series in physics.

McGraw-Hill, New York, NY, USA, 1933. xx + 576 pp. LCCN QC21 .S595 1933c.

Slater:1947:E

[SF47a] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1947. xiii + 240 pp. LCCN QC760 .S55.

Slater:1947:ECJ

[SF47b] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1947. xiii + 240 pp. LCCN QC760 .S55.

Slater:1947:MCJ

[SF47c] John C. (John Clarke) Slater and Nathaniel Herman Frank. Mechanics. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1947. xiii + 297 pp. LCCN QA805 .S632.

Slater:1969:E

[SF69] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. Dover, New York, NY, USA, 1969. ISBN 0-486-62263-0. xiii + 240 pp. LCCN QC760 .S55 1969.

URL http://www.loc.gov/catdir/description/dover032/
69017476.html.

Slater:1983:M

[SF83] John C. (John Clarke) Slater and Nathaniel Herman Frank. Mechanics. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1983. ISBN 0-313-24064-7 (hardcover). xiii + 297 pp. LCCN QA805 .S632 1983.

Stoff:1991:MPD

[SFW91] Michael B. Stoff, Jonathan F. Fanton, and R. Hal (Richard Hal) Williams, editors. The Manhattan Project: a documentary introduction to the Atomic Age. Temple University Press, Philadelphia, PA, USA, 1991. ISBN 0-87722-787-X. xxi + 290 pp. LCCN QC773.A1 M36 1991.

Schroeder-Gudehus:1982:DLC

[SG82] Brigitte Schroeder-Gudehus. Division of labour and the common good: The international association of academies, 1899–1914.

In Bernhard et al. [BCS82], pages 3–20. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500062.

Snyder:2011:APP

[SG11] Sharon Snyder and Toni Michnovicz Gibson. Los Alamos and the Pajarito Plateau. Images of America. Arcadia Publishing, Charleston, SC, USA, 2011. ISBN 0-7385-8483-5 (paperback). 127 pp. LCCN F804.L6 S67 2011.

Schulmann:1995:EW

[SH95] Robert Schulmann and Gerald Holton. Einstein’s wife. New York Times, ??(??):??, October 8, 1995. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Schucking:1999:EPE

[SH99] E. L. (Engelbert L.) Schucking and Alex Harvey, editors. On Einstein’s path: essays in honor of Engelbert Schucking. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98564-6 (hardcover). LCCN QC173.5 .O6 1999. URL http://www.loc.gov/

catdir/enhancements/fy0816/98020292-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0816/98020292-t.
html.

Stober:2001:CSW

[SH01] Dan Stober and Ian Hoffman. A convenient spy: Wen Ho Lee and the politics of nuclear espionage. Simon and Schuster, New York, NY, USA, 2001. ISBN 0-7432-2378-0. 384 + 8 pp. LCCN

UB271.C62 L47 2001. URL http://www.loc.gov/catdir/
bios/simon052/2001054945.html;   http://www.loc.gov/
catdir/description/simon032/2001054945.html;   http:
//www.loc.gov/catdir/toc/fy02/2001054945.html.

Shapley:1954:BBG

[Sha54] Harlow Shapley. Books: Galileo Galilei: Dialogue Concerning the Two Chief World Systems — Ptolemaic and Copernican, translated by Stillman Drake with a foreword by Albert Einstein; Galileo Galilei: Dialogue on the Great World Systems, revised by Giorgio de Santillana. Bulletin of the Atomic Scientists, 10(5):162, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shankland:1963:CAE

[Sha63] R. S. Shankland. Conversations with Albert Einstein. American Journal of Physics, 31(1):47–57, January 1963. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/

v31/i1/p47_s1. Transcriptions of five conversations between 4 February 1950 and 11 December 1954. See also [Hol69b, Sha73a, Sha75].

Shankland:1964:MMEa

[Sha64a] R. S. Shankland. Michelson–Morley Experiment. American Journal of Physics, 32(1):16–35, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Shankland:1964:MMEb

[Sha64b] R. S. Shankland. The Michelson–Morley Experiment. Scientific American, 211(5):107–114, November 1964. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v211/n5/pdf/scientificamerican1164-107.pdf.

Shapiro:1964:FTG

[Sha64c] Irwin I. Shapiro. Fourth test of general relativity. Physical Review Letters, 13(26):789–791, December 28, 1964. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/

PhysRevLett.13.789. See experimental results in [SPA^{+}68b, SPA^{+}68a, SAI^{+}71].

Shadowitz:1968:SR

[Sha68] Albert Shadowitz. Special Relativity. Saunders, Philadelphia, PA, USA, 1968. xiii + 203 pp. LCCN ???? Reprinted as [Sha88].

Shankland:1973:CAE

[Sha73a] R. S. Shankland. Conversations with Albert Einstein. II. American Journal of Physics, 41(7):895–901, July 1973. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/41/895/1. Further clarifi-

cation of the transcriptions of five conversations between 4 February 1950 and 11 December 1954. See also [Sha63, Hol69b, Sha75].

Shapley:1973:IAS

[Sha73b] Deborah Shapley. Institute for Advanced Study: Einstein is a hard act to follow. Science, 179(4079):1209–1211, March 23, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/
179/4079/1209.full.pdf.

Shankland:1975:CCA

[Sha75] R. S. Shankland. Comment on “Conversations with Albert Einstein. II”. American Journal of Physics, 43(5):464, May 1975. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/43/464/

1. See also earlier reports [Sha63, Hol69b, Sha73a].

Shankland:1979:AER

[Sha79a] R. S. Shankland. Albert Einstein — in remembrance. Biography (Honolulu), 2(3):190–200, Summer 1979. CODEN BGPYE2. ISSN 0162-4962 (print), 1529-1456 (electronic). URL https://

muse.jhu.edu/article/371229/summary.

Shapley:1979:BYE

[Sha79b] Deborah Shapley. It’s a big year for Einstein memorabilia. Science, 203(4378):342–344, January 26, 1979. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1747761;
http://www.sciencemag.org/content/203/4378/342.full.
pdf.

Shamos:1987:GEP

[Sha87] Morris H. (Morris Herbert) Shamos, editor. Great experiments in physics: firsthand accounts from Galileo to Einstein. Dover, New York, NY, USA, 1987. ISBN 0-486-25346-5 (paperback). viii + 370 pp. LCCN QC7 .G74 1987. URL http://store.

doverpublications.com/0486253465.html.

Shadowitz:1988:SR

[Sha88] Albert Shadowitz. Special relativity. Dover, New York, NY, USA, 1988. ISBN 0-486-65743-4. xiii + 203 pp. LCCN QC173.65 .S53 1988. US$5.95. URL http://www.loc.gov/

catdir/description/dover031/88011125.html.

Shadowitz:1992:ITS

[Sha92] S. Shadowitz. I’m told that I sat on Einstein’s knee when I was 5 years old. Albert Shadowitz espionage and Communism accusations. Queens Quarterly, 99(4):824–836, Winter 1992. ISSN 0033-6041.

Shanks:1998:III

[Sha98] Niall Shanks, editor. Idealization IX: idealization in contemporary physics, volume 63 of Poznań studies in the philosophy of the sciences and the humanities. Rodopi, Amsterdam, The Netherlands and Atlanta, GA, USA, 1998. ISBN 90-420-0642-0. ISSN 0303-8157. 238 pp. LCCN QC6 .I33 1998.

Shachtman:1999:AZC

[Sha99] Tom Shachtman. Absolute Zero and the Conquest of Cold. Houghton-Mifflin, Boston, MA, USA, 1999. ISBN 0-395-93888-0, 0-547-52595-8, 0-618-08239-5 (paperback), 1-299-88289-7. 261 pp. LCCN QC278 .S48 1999. URL http://www.loc.

gov/catdir/description/hm022/99033305.html;   http:
//www.loc.gov/catdir/enhancements/fy0735/99033305-b.
html;   http://www.loc.gov/catdir/samples/hm051/99033305.
html.

Shah:2014:EAT

[Sha14] Huma Shah. The emotions of Alan Turing: The boy who explained Einstein’s theory of relativity aged 15\frac{1}{2} for his mother. International Journal of Synthetic Emotions, 5(1):23–30, ???? 2014. CODEN ???? ISSN 1947-9093 (print), 1947-9107 (elec-

tronic). URL https://dblp.org/db/journals/ijse/ijse5.
html#Shah14.

Sherwin:1975:WDA

[She75] Martin J. Sherwin. A world destroyed: the atomic bomb and the Grand Alliance. Alfred A. Knopf, New York, NY, USA, 1975. ISBN 0-394-49794-5. xvi + 315 + xi pp. LCCN D753 .S48 1975. US$10.00.

Shea:1983:OHRa

[She83a] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario Series in Philosophy of Science, A Series of Books in Philosophy of Science, Methodology, Epistemology, Logic, History of Science, and Related Fields. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 94-009-7133-8 (e-book), 94-009-7135-4. ISSN 1566-659X. 268 pp. LCCN D1-DX301.

Shea:1983:OHRb

[She83b] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1584-X. 252 pp. LCCN QC773 .O87 1983. URL http:

//link.springer.com/book/10.1007/978-94-009-7133-2.

Sherwin:1987:WDH

[She87] Martin J. Sherwin. A world destroyed: Hiroshima and the origins of the arms race. Vintage Books, New York, NY, USA, 1987. ISBN 0-394-75204-X (paperback). xxx + 375 pp. LCCN

D842 .S49 1987. US$9.95. URL http://www.loc.gov/catdir/
description/random044/86040523.html.

Shelton:1988:ROS

[She88] Jim Shelton. The role of observation and simplicity in Einstein’s epistemology. Studies in History and Philosophy of Science Part A, 19(1):103–118, March 1988. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900210.

Sherwin:2003:WDH

[She03] Martin J. Sherwin. A world destroyed: Hiroshima and its legacies. Stanford University Press, Stanford, CA, USA, third edition, 2003. ISBN 0-8047-3957-9. xl + 375 pp.

LCCN D842 .S49 2003. URL http://www.loc.gov/catdir/
enhancements/fy0710/2003014107-b.html;   http://www.
loc.gov/catdir/enhancements/fy0710/2003014107-d.html.

Sheldon:2015:ERN

[She15] Eric Sheldon. Essay review: Nobel lectures: Physics (2006– 2010) [review of Nobel Lectures: Physics (2006–2010), edited by Lars Brink, Singapore, NJ, and London, World Scientific 2014, xii + 371 pp. ISBN 978-981-4612-67-8 (hardback), 978-981-4612-68-5 (paperback), 978-9-81-461270-08 (e-book)]. Contemporary Physics, 56(2):225–232, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2016:AER

[She16a] Eric Sheldon. Albert Einstein’s relativistic road ‘less traveled’: A review The road to relativity: the history and meaning of Einstein’s ‘The Foundation of General Relativity’, by Albert Einstein (Facsimile German Manuscript) and foreword by John Stachel, with annotations and commentary by Hanoch Gutfreund and Jürgen Renn. Contemporary Physics, 57(2):250–254, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2016:RSG

[She16b] Eric Sheldon. Relativity: The Special and the General Theory, 100th Anniversary Edition, by Albert Einstein, Scope: general interest, edited book. Level: general readership, non-specialists, pre-university, undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 57(1):147–149, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shields:1949:BWA

[Shi49] Margaret C. Shields. Bibliography of the writings of Albert Einstein to October 1949. In Schilpp [Sch49c], pages 689–758. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Shipman:1975:IHT

[Shi75] H. L. Shipman. The implausible history of triple star models for Cygnus X-1: Evidence for a black hole. Astrophysical Letters, 16 (1):9–12, January 1, 1975. CODEN ASTLAI. ISSN 0004-6388.

Shirakawa:2001:NLD

[Shi01] Hideki Shirakawa. Nobel Lecture: The discovery of polyacetylene film — the dawning of an era of conducting polymers. Reviews of Modern Physics, 73(3):713–718, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.713;   http://rmp.aps.org/
abstract/RMP/v73/i3/p713_1.

Shifman:2017:STT

[Shi17] Misha Shifman. Standing Together In Troubled Times. World Scientific Publishing Company, Singapore, 2017. ISBN 981-320-102-9, 981-320-102-9 (e-book). xxii + 337 pp. LCCN

QC16.P37 A4 201. URL http://public.eblib.com/choice/
PublicFullRecord.aspx?p=4800883.

Sadler:2006:SSB

[SHL^{+}06] L. E. Sadler, J. M. Higbie, S. R. Leslie, M. Vengalattore, and D. M. Stamper-Kurn. Spontaneous symmetry breaking in a quenched ferromagnetic spinor Bose–Einstein condensate. Nature, 443(7109):312–315, September 21, 2006. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v443/n7109/
full/nature05094.html.

Shorter:2021:BRP

[Sho21] R. S. Shorter. Book review: The physics of the Manhattan Project by Bruce Cameron Reed, New York, Springer, Nature Switzerland AG, 2021, 256 pp., £43.99 (e-book), ISBN: 978-3-030-61372-3. Scope: general interest, textbook. Level: general readership, undergraduate, teacher. Contemporary Physics, 62(2):118, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shurkin:2017:TGL

[Shu17] Joel N. Shurkin. True genius: the life and work of Richard Garwin, the most influential scientist you’ve never heard of. Prometheus Books, Amherst, NY, USA, 2017. ISBN 1-63388-223-3 (hardcover), 1-63388-224-1 (e-book). LCCN QC16.G38.

URL http://lib.myilibrary.com?id=994754.

Spitzer:1949:PM

[SHW^{+}49] Lyman Spitzer, Jr., William A. Higinbotham, Arthur S. Wightman, Donald R. Hamilton, David Bohm, Roy Britten, Robert R. Bush, Elmer G. Butler, Albert Einstein, Luther P. Eisenhart, Samuel A. Goudsmit, M. Stanley Livingston, Stuart Mudd, David Pines, Oswald Veblen, and Irving Wolff. To prevent misunderstanding. Physics Today, 2(8):3–??, August 1, 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.3066618.

Sichau:2009:EIA

[Sic09] Christian Sichau. Einstein, interaktiv und zum Anfassen. (German) [Einstein, interactive and hands-on]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 17 (1):85–92, February 2009. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00048-008-0327-9.pdf.

Sida:1976:GRE

[Sid76] D. W. Sida. A geometrical relationship for the Einstein and Ricci tensors. Foundations of Physics, 6(4):477–483, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.
1007/BF00715036.

Siegel:1991:IME

[Sie91] Daniel M. Siegel. Innovation in Maxwell’s electromagnetic theory: molecular vortices, displacement current, and light. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-35365-3. x + 225 pp. LCCN QC670 .S48 1991. URL http://

www.loc.gov/catdir/description/cam024/90042511.html;
 http://www.loc.gov/catdir/toc/cam029/90042511.html.

Siegel:1997:BRC

[Sie97] D. M. Siegel. Book review: The collected papers of Albert Einstein, vol 4, The Swiss years: Writings, 1912–1914 — M. J. Klein, Anne J. Kox, J. Renn, R. Schulmann, S. Bergia, J. Illy, M. Janssen, J. D. Norton, T. Sauer. Annals of Science, 54 (2):207–208, March 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Siegel:2007:BRC

[Sie07] Daniel Siegel. Book review: The collected papers of Albert Einstein. Isis, 98(4):800, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sieroka:2010:USK

[Sie10] Norman Sieroka. Umgebungen: symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus. (German) [Environments: symbolic constructivism following Hermann Weyl and Fritz Medicus], volume 8 of Legierungen. Chronos Verlag, Zürich, Switzerland, 2010. ISBN 3-0340-1006-0. 411 pp. LCCN ???? EUR 43.00; SFR 58.00. URL http:/

/www.chronos-verlag.ch/php/book_latest-new.php?book=
978-3-0340-1006-1.

Sigal:1978:BPT

[Sig78] Leon V. Sigal. Bureaucratic politics & tactical use of committees: The Interim Committee & the decision to drop the atomic bomb. Polity, 10(3):326–364, Spring 1978. ISSN 0032-3497 (print), 1744-1684 (electronic). URL http://www.jstor.org/

stable/3234412.

Sigurdsson:1992:IMS

[Sig92] S. Sigurdsson. Interpretation and misinterpretation of the Special and the General Theory of Relativity by Albert Einstein contemporaries — German — K. Hentschel. Annals of Science, 49(6):577–583, November 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Silberstein:1922:TGR

[Sil22] Ludwik Silberstein. The theory of General Relativity and gravitation. D. Van Nostrand Company, New York, NY, USA, second edition, 1922. 3 + 141 pp. LCCN QC6 .S55 1924. Based on a course of lectures delivered at the Conference on Recent Advances in Physics held at the University of Toronto, in January, 1921.

Silberstein:1923:TRE

[Sil23] Ludwik Silberstein. The true relation of Einstein’s to Newton’s equations of motion. Nature, 112(2822):788–789, December 1, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v112/n2822/pdf/112788b0.pdf.

Silberstein:1924:TR

[Sil24] Ludwik Silberstein. The Theory of Relativity. Macmillan Publishing Company, New York, NY, USA, second edition, 1924. x + 563 pp. LCCN QC6 .S55 1924.

Silberstein:1930:SUA

[Sil30] Ludwik Silberstein. The size of the universe: attempts at a determination of the curvature radius of spacetime. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1930. viii + 215 + 1 pp. LCCN QC6 .S48.

Silberstein:1936:MLG

[Sil36] Ludwik Silberstein. Minimal lines and geodesics within matter: a fundamental difficulty of Einstein’s theory. Nature, 138(3502):

1012, December 12, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v138/n3502/pdf/1381012a0.pdf.

Simha:1980:AER

[Sim80] R. Simha. Albert Einstein and rheology. Journal of Rheology, 24(6):935, ???? 1980. ISSN 0148-6055.

Sime:1990:LME

[Sim90] Ruth Lewin Sime. Lise Meitner’s escape from Germany. American Journal of Physics, 58(3):262–267, March 1990. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/58/3/10.1119/1.16196. See also [Sim94].

Sime:1994:LMS

[Sim94] Ruth Lewin Sime. Lise Meitner in Sweden 1938–1960: Exile from physics. American Journal of Physics, 62(8):695– 701, August 1994. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://scitation.aip.org/
content/aapt/journal/ajp/62/8/10.1119/1.17498.  See

[Sim90, Kac96, Tru96].

Sime:1996:LML

[Sim96] Ruth Lewin Sime. Lise Meitner: a life in physics, volume 13 of California studies in the history of science. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1996. ISBN 0-520-08906-5, 0-520-20860-9 (paperback), 0-585-05524-6 (e-book). xiii + 526 pp. LCCN QC774.M4 S56 1996.

Simon:2005:AER

[Sim05] J. J. Simon. Albert Einstein, radical — a political profile. Monthly Review — an Independent Socialist Magazine, 57(1): 1–12, May 2005. ISSN 0027-0520.

Simon:2006:AEA

[Sim06] Dieter Simon. Albert Einstein: Akademie-Vorträge: Sitzungsberichte der Preußischen Akademie der Wissenschaften, 1914– 1932. (German) [Albert Einstein: Academy lectures: Proceedings of the Prussian Academy of Sciences, 1914–1932]. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40609-3 (hardcover), 3-527-60895-8 (e-book), 3-527-62244-6 (e-book). xiv + 431 pp. LCCN QC71 .E36 2006.

Simoes:2022:STS

[Sim22] Ana Simões. In the shadow of the 1919 total solar eclipse: The two British expeditions and the politics of invisibility. Berichte zur Wissenschaftsgeschichte, 45(4):581–601, December 2022. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Singh:1961:GIT

[Sin61] Jagjit Singh. Great ideas and theories of modern cosmology. Dover, New York, NY, USA, 1961. 276 pp. LCCN BD511 .S5 1961.

Singh:1970:GIT

[Sin70a] Jagjit Singh. Great ideas and theories of modern cosmology. Dover, New York, NY, USA, revised and enlarged edition, 1970. ISBN 0-486-20925-3. 416 pp. LCCN BD511 .S5 1970b.

Singh:1970:MC

[Sin70b] Jagjit Singh. Modern cosmology. Pelican books. Penguin, London, UK and New York, NY, USA, revised edition, 1970. ISBN 0-14-021060-1. 416 + 12 pp. LCCN BD511 .S5 1970.

Sinha:1975:NLL

[Sin75] Shri Murari Sinha. Nobel laureates of literature, 1901–1973 (with a note on the laureates of 1974). S. Chand, New Delhi, India, 1975. xlviii + 397 + 1 pp. LCCN ????

Sinha:2000:EST

[Sin00a] Supurna Sinha. Einstein and the Special Theory of Relativity. Resonance, 5(3):6–15, March 2000. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.

springerlink.com/content/a72403552x885872/.

Sinha:2000:PST

[Sin00b] Supurna Sinha. Poincaré and the Special Theory of Relativity. Resonance, 5(2):12–15, February 2000. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://

www.springerlink.com/content/h0331238v8010741/.

Singer:2001:EDI

[Sin01a] W. Singer. ‘Endless dawns of imagination’ (Albert Einstein and Rabindranath Tagore, a portfolio). Kenyon Review, 23(2):7–33, Spring 2001. ISSN 0163-075X.

Singer:2001:PET

[Sin01b] Wendy Singer. Portfolio: Einstein and Tagore. The Kenyon Review, 23(2):7–33, Spring 2001. CODEN ???? ISSN 0163-075X.

Singh:2002:CVR

[Sin02] Rajinder Singh. C. V. Raman and the discovery of the Raman Effect. Physics in Perspective (PIP), 4(4):399–420, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160200002;   http://www.nobelprize.org/
nobel_prizes/physics/laureates/1930/.

Singh:2004:NLC

[Sin04a] Rajinder Singh. Nobel laureate C. V. Raman’s work on light scattering: historical contributions to a scientific biography. Logos, Berlin, Germany, 2004. ISBN 3-8325-0567-9 (paperback). xii + 282 pp. LCCN ????

Singh:2004:BBO

[Sin04b] Simon Singh. Big Bang: the origin of the universe. Fourth Estate, New York, NY, USA, 2004. ISBN 0-00-716220-0, 0-00-715251-5 (hardcover). 532 pp. LCCN

QB991.B54 S56 2004. URL http://www.simonsingh.com/
;   http://www.simonsingh.com/Big_Bang.html;   http:
//www.simonsingh.com/Big_Bang_Reviews.html.

Sinha:2005:EST

[Sin05] Supurna Sinha. Einstein and the Special Theory of Relativity. Resonance, 10(12):96–105, December 2005. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Singh:2007:IPC

[Sin07] Rajinder Singh. India’s Physics and Chemistry Nobel Prize nominators and nominees in colonial and international context. Notes and Records of the Royal Society of London, 61(3):333– 345, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462636.

Singh:2010:OSA

[Sin10a] Rajinder Singh. The other side of Albert Einstein. Current Science, 99(7):863, October 10, 2010. CODEN CUSCAM. ISSN 0011-3891.

Sington:2010:EG

[Sin10b] Philip Sington. The Einstein girl. Vintage Books, New York, NY, USA, 2010. ISBN 0-09-953579-3 (paperback), 0-8466-5290-0. 391 pp. LCCN ????

Sington:2012:EMG

[Sin12] Philip Sington. Das Einstein-Mädchen. (German) [The Einstein girl]. dtv, Munich, Germany, 2012. ISBN 3-423-21399-X. 464 pp. LCCN ???? Translation of [Sin10b] to German by Sophie Zeitz.

Stenger:1998:SDG

[SISK^{+}98] J. Stenger, S. Inouye, D. M. Stamper-Kurn, H.-J. Miesner, A. P. Chikkatur, W. Ketterle, et al. Spin domains in groundstate Bose–Einstein condensates. Nature, 396(6709):345–348, November 26, 1998. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v396/n6709/full/396345a0.html.

Saffer:1982:CZGa

[SK82a] Thomas H. Saffer and Orville E. Kelly. Countdown zero: GI Victims of U.S. Atomic Testing. Putnam, New York, NY, USA, 1982. ISBN 0-399-12685-6. 351 pp. LCCN UF767 .S218 1982. US$15.95. Introduction by Former Secretary of the Interior, Stewart L. Udall.

Saffer:1982:CZGb

[SK82b] Thomas H. Saffer and Orville E. Kelly. Countdown zero: GI Victims of U.S. Atomic Testing. Penguin, London, UK and New York, NY, USA, 1982. ISBN 0-14-006724-8 (paperback). 350 pp. LCCN U264 .S23 1983. Introduction by Former Secretary of the Interior, Stewart L. Udall.

Seaborg:1994:PSJ

[SK94] Glenn Theodore Seaborg and Ronald L. Kathren. The plutonium story: the journals of Professor Glenn T. Seaborg 1939– 1946. Battelle Press, Columbus, OH, USA, 1994. ISBN 0-935470-75-1. ix + 920 pp. LCCN QC773.3.U5 S424 1994.

Seeger:2009:SMN

[SK09] Petra. Seeger and Eric R. Kandel. In search of memory: the neuroscientist Eric Kandel, winner of the Nobel Prize. Icarus Films., Brooklyn, NY, 2009. LCCN RC339.52.K362 A3 2008. 1 videodisc (95 min.).

Skates:1994:IJA

[Ska94] John Ray Skates. The invasion of Japan: alternative to the bomb. University of South Carolina Press, Columbia, SC, USA, 1994. ISBN 0-87249-972-3. xii + 276 pp. LCCN D767.2 .S56 1994.

Schrodinger:1987:DMB

[SKD^{+}87] Erwin Schrödinger, Gabriele Kerber, Auguste Dick, Wolfgang Kerber, et al., editors. Dokumente, Materialien und Bilder zur 100. Wiederkehr des Geburtstages von Erwin Schrödinger. (German) [Documents, materials and pictures on the 100th return of the birthday of Erwin Schrödinger]. Fassbaender, Wien, Austria, 1987. ISBN 3-900538-09-3. 158 pp. LCCN QC16.S265 D65 1987.

Skidmore:2004:AWH

[Ski04] Max J. Skidmore. After the White House: former presidents as private citizens. Palgrave Macmillan, New York, NY, USA, 2004. ISBN 0-312-29559-6. ix + pp. LCCN E176.1

.S613 2004. URL http://www.loc.gov/catdir/description/
hol042/2003068939.html;   http://www.loc.gov/catdir/
toc/hol041/2003068939.html.

Schulmann:1998:BYCa

[SKJI98a] Robert Schulmann, Anne J. Kox, Michel Janssen, and József Illy, editors. The Berlin Years: Correspondence, 1914–1917, volume 8A of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-04849-5 (hardcover). lxx + 590 + 8 pp. LCCN QC16.E5 A2 1987.

URL http://einsteinpapers.press.princeton.edu/vol8-
trans;   http://einsteinpapers.press.princeton.edu/
vol8a-doc;   http://press.princeton.edu/titles/6174.
html;   http://press.princeton.edu/titles/6206.html.

Schulmann:1998:BYCc

[SKJI98b] Robert Schulmann, Anne J. Kox, Michel Janssen, and József Illy, editors. The Berlin Years: Correspondence, 1914–1918.

English Translation Supplement, volume 8. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-04841-X. xxii

+ 714 pp. URL http://einsteinpapers.press.princeton.
edu/vol8-trans;   http://press.princeton.edu/titles/

6174.html. English translation by Ann M. Hentschel in consultation with Klaus Hentschel.

Schulmann:1998:BYCb

[SKJI98c] Robert Schulmann, Anne J. Kox, Michel Janssen, and József Illy, editors. The Berlin Years: Correspondence, 1918, volume 8B of The collected papers of Albert Einstein. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-04849-5 (hardcover). lxx + 590 + 8 pp. LCCN QC16.E5 A2 1987.

URL http://einsteinpapers.press.princeton.edu/vol8-
trans;   http://einsteinpapers.press.princeton.edu/
vol8b-doc;   http://press.princeton.edu/titles/6174.
html;   http://press.princeton.edu/titles/6206.html.

Sklar:1985:PSP

[Skl85] Lawrence Sklar. Philosophy and spacetime physics. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1985. ISBN 0-520-05374-5. x + 335 pp. LCCN QC173.59.S65 S54 1985.

Stephani:2003:ESE

[SKM^{+}03] Hans Stephani, Dietrich Kramer, Malcolm MacCallum, Cornelius Hoenselaers, and Eduard Herlt, editors. Exact solutions of Einstein’s field equations. Cambridge monographs on mathematical physics. Cambridge University Press, Cambridge, UK, second edition, 2003. ISBN 0-521-46136-7. xxix + 701 pp. LCCN QC173.6 .E96 2003. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/description/cam031/2002071495.html;   http://
www.loc.gov/catdir/toc/cam031/2002071495.html.

Seaborg:1990:EBU

[SL90] Glenn Theodore Seaborg and Walter D. Loveland. The Elements Beyond Uranium. Wiley, New York, NY, USA, 1990. ISBN 0-471-89062-6. xiii + 359 pp. LCCN QD172.T7

S35 1990. URL http://www.loc.gov/catdir/description/
wiley031/90012643.html;   http://www.loc.gov/catdir/
toc/onix02/90012643.html.

Stannard:1991:BHU

[SL91] Russell Stannard and John Levers. Black holes and Uncle Albert. Faber, London, UK, 1991. ISBN 0-571-14452-7, 0-571-14453-5 (paperback). 145 pp. LCCN ????

Stephan:1993:ANP

[SL93] Paula E. Stephan and Sharon G. Levin. Age and the Nobel Prize revisited. Scientometrics, 28(3):387–399, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026517.

Seife:2005:WSU

[SL05a] Charles Seife and Andrew Lawler. We’re so sorry, Uncle Albert. Science, 307(5711):869–870, February 11, 2005. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Special Issue: Einstein’s Legacy.

Stannard:2005:BHU

[SL05b] Russell Stannard and John Levers. Black holes and Uncle Albert. Faber, London, UK, 2005. ISBN 0-571-22614-0 (paperback). 145 pp. LCCN ????

Slater:1939:ICP

[Sla39] John C. (John Clarke) Slater. Introduction to chemical physics. International series in physics; F. K. Richtmeyer, consulting editor. McGraw-Hill, New York, NY, USA, 1939. xiv + 521 pp. LCCN QC171 .S55.

Slater:1942:MT

[Sla42] John C. (John Clarke) Slater. Microwave Transmission. International series in physics. McGraw-Hill, New York, NY, USA, 1942. x + 309 pp. LCCN QC661 .S635.

Slater:1947:PUR

[Sla47] John C. (John Clarke) Slater. Peredacha ulstrakorotkikh radiovola. ????, ????, 1947. 344 pp. LCCN QC661 .S639 1947.

Slater:1950:ME

[Sla50] John C. (John Clarke) Slater. Microwave electronics. The Bell Telephone Laboratories series. Van Nostrand, New York, NY, USA, 1950. xiv + 406 pp. LCCN QC721 .S6.

Slater:1951:SHF

[Sla51] J. C. Slater. A simplification of the Hartree–Fock method. Physical Review (2), 81(3):385–390, February 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v81/i3/p385_1.

Slater:1953:ESA

[Sla53] John C. (John Clarke) Slater. Electronic structure of atoms and molecules, volume 3 of Massachusetts Institute of Technology. Solid-State and Molecular Theory Group. Technical report. Solid-State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, 1953. iv + 212 pp. LCCN QC173 .M4 no. 3.

Slater:1953:ESS

[Sla54] John C. (John Clarke) Slater. Electronic structure of solids. Solid-State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, 1953/1954. ???? pp. LCCN QC173 .M4 no. 4-6.

Slater:1955:MP

[Sla55] John C. (John Clarke) Slater. Modern physics. McGraw-Hill, New York, NY, USA, 1955. 322 pp. LCCN QC7 .S57.

Slater:1956:ESS

[Sla57] John C. (John Clarke) Slater. The electronic structure of solids. In Siegfried Flügge, editor, Handbuch der Physik: Elektrische Leitungsphänomene. (German) [Electrical conductivity phenomena], volume 19–20, pages 1–136. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1956/1957. ISBN 0-387-02040-3, 3-540-02040-3. LCCN QC21 .H327.

Slater:1960:QTA

[Sla60] John C. (John Clarke) Slater. Quantum theory of atomic structure. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1960. ???? pp. LCCN QC174.1 .S53.

Slater:1963:QTM

[Sla63] John C. Slater. Quantum Theory of Molecules and Solids: Electronic-Structure of Molecules, volume 1. McGraw-Hill, New York, NY, USA, 1963. xv + 485 pp. LCCN QC174.1 .S551.

Slater:1969:CDQ

[Sla69a] John C. (John Clarke) Slater. Concepts and development of quantum physics. Dover, New York, NY, USA, 1969. ISBN 0-486-62265-7. xi + 322 pp. LCCN QC174.1 .S528 1969.

Slater:1969:ME

[Sla69b] John C. (John Clarke) Slater. Microwave electronics. Dover, New York, NY, USA, 1969. ISBN 0-486-62264-9. xiv + 406 pp. LCCN TK7876 .S58 1969.

Slater:1970:ICP

[Sla70] John C. (John Clarke) Slater. Introduction to chemical physics. Dover, New York, NY, USA, 1970. ISBN 0-486-62562-1. xiv + 521 pp. LCCN QC171 .S55 1970.

Slater:1972:SEC

[Sla72] John C. Slater. Statistical exchange-correlation in the selfconsistent field. Advances in Quantum Chemistry, 6:1–92, 1972. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327608605419.

Slater:1979:CMO

[Sla79] John C. (John Clarke) Slater. The calculation of molecular orbitals. Wiley, New York, NY, USA, 1979. ISBN 0-471-03181-X. vi + 104 pp. LCCN QD461 .S49 1979.

Slayton:2013:ACP

[Sla13] Rebecca Slayton. Arguments that count: physics, computing, and missile defense, 1949–2012. Inside technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01944-2 (hardcover). xi + 325 pp. LCCN UA22 .S57 2013.

Slavov:2022:KIE

[Sla22] Matias Slavov. Kaila’s interpretation of Einstein–Minkowski invariance theory. Studies in History and Philosophy of Science Part A, 93(??):57–65, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368122000474.

Shomroni:2009:EOS

[SLLS09] I. Shomroni, E. Lahoud, S. Levy, and J. Steinhauer. Evidence for an oscillating soliton/vortex ring by density engineering of a Bose–Einstein condensate. Nature Physics, 5(3):193–197, January 18, 2009. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v5/n3/full/nphys1177.html.

Slosson:1922:TE

[Slo22] Edwin E. Slosson. On translating Einstein. Science, 56(1461): 752–754, December 29, 1922. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/56/1461/752.full.pdf.

Slowik:2011:MAE

[Slo11] Edward S. Slowik. Mechanics from Aristotle to Einstein. Annals of Science, 68(1):142–144, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stark:1941:JDP

[SM41] Johannes Stark and Wilhelm Müller, editors. Jüdische und deutsche Physik: Vorträge zur Eröffnung des Kollegiums für theoretische Physik an der Universität München. (German) [Jewish and German Physics: Lectures at the opening of the College of Theoretical Physics at the University of Munich]. Heling, Leipzig, Germany, 1941. 55 pp. LCCN ????

Schwartz:1979:EB

[SM79] Joseph Schwartz and Michael McGuinness. Einstein for beginners. Pantheon Books, New York, NY, USA, 1979. ISBN 0-394-50588-3, 0-394-73801-2 (paperback). 173 pp. LCCN QC16.E5 S32 1979. US$8.95, US$2.95.

Schwartz:1992:EB

[SM92] Joseph Schwartz and Michael McGuinness. Einstein for beginners. Icon Books, Cambridge, UK, 1992. ISBN 1-874166-02-1. 173 pp.

Schwartz:1999:IE

[SM99] Joseph Schwartz and Michael McGuinness. Introducing Einstein. Icon, Trumpington, UK, 1999. ISBN 1-84046-060-1, 1-874166-02-1. 169 pp. LCCN QC16.E5 S39 1999.

Smeenk:2007:MTM

[SM07] Christopher Smeenk and Christopher Martin. Mie’s theories of matter and gravitation. In Renn et al. [RSS^{+}07b], pages 623– 632. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/
978-1-4020-4000-9_35.

Spangenburg:2008:NBA

[SM08] Ray Spangenburg and Diane Kit Moser, editors. Niels Bohr: Atomic Theorist. Makers of modern science. Chelsea House, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6178-5. xvi + 141 pp. LCCN QC16.B63 S63 2008.

Smeenk:2014:ERC

[Sme14] Christopher Smeenk. Einstein’s role in the creation of Relativistic Cosmology. In Janssen and Lehner [JL14b], chapter 7, pages 228–269. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Spiechowicz:2023:DCB

[SMHL23] Jakub Spiechowicz, Ivan G. Marchenko, Peter Hänggi, and Jerzy Luczka. Diffusion coefficient of a Brownian particle in equilibrium and nonequilibrium: Einstein model and beyond. Entropy, 25(1):42, January 2023. CODEN ENTRFG. ISSN

1099-4300. URL https://dblp.org/db/journals/entropy/
entropy25.html#SpiechowiczMHL23.

Smith:1951:BRA

[Smi51] Vincent Edward Smith. Book review: Albert Einstein: Philosopher–Scientist ed. by Paul Arthur Schilpp. The Thomist: A Speculative Quarterly Review, 14(3):404–407, 1951. CODEN

???? ISSN 2473-3725. URL https://muse.jhu.edu/article/
644417/summary.

Smith:1958:BDU

[Smi58] Alice Kimball Smith. Behind the decision to use the atomic bomb: Chicago 1944–45. Bulletin of the Atomic Scientists, 14 (8):288–312, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1965:PHS

[Smi65] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. University of Chicago Press, Chicago, IL, USA and London, UK, 1965. xiv + 591 pp. LCCN Q127.U6 S6.

Smith:1970:PHS

[Smi70] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-69026-8. xvi + 398 pp. LCCN Q127.U6 S6 1971.

Smith:1982:EUA

[Smi82] Robert W. (Robert William) Smith. The expanding universe: astronomy’s “Great Debate,” 1900–1931. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23212-0. xiv + 220 pp. LCCN QB981 .S69 1982.

Smith:1986:SG

[Smi86] Martin Cruz Smith. Stallion Gate. Random House, New York, NY, USA, 1986. ISBN 0-394-53006-3. 321 pp. LCCN PS3569.M5377 S7 1986. US$17.95. URL http://www.loc.gov/

catdir/description/random0412/85024444.html.

Smith:1989:BRC

[Smi89a] C. Smith. Book review: The Collected Papers of Albert Einstein, Part 1, the Early Years, 1879–1902 J. Stachel. Journal of Interdisciplinary History, 19(3):488–490, Winter 1989. ISSN 0022-1953 (print), 1530-9169 (electronic).

Smith:1989:BRB

[Smi89b] Crosbie Smith. Book review: The Collected Papers of Albert Einstein. I. The Early Years, 1879–1902, by John Stachel. Journal of Interdisciplinary History, 19(3):488–490, ???? 1989. ISSN 0022-1953 (print), 1530-9169 (electronic). URL http://www.

jstor.org/stable/204367.

Smith:1992:NC

[Smi92] Marie Smith. Nobel crimes. Carroll and Graf Pub., New York, NY, USA, 1992. ISBN 0-88184-914-6 (paperback). viii + 263 pp. LCCN PN6071.D45 N63 1992.

Smith:1996:DSE

[Smi96] Dinitia Smith. Dark side of Einstein emerges in his letters. New York Times, ??(??):C11, November 6, 1996. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
109673807/.

Smith:1997:BET

[Smi97] Nicholas J. J. Smith. Bananas enough for time travel? The British Journal for the Philosophy of Science, 48(3):363–389, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/48/3/363.full.pdf+html.

Smith:2002:BRB

[Smi02] Robert W. Smith. Book review: The Einstein tower: an intertexture of dynamic construction, Relativity Theory, and astronomy: Klaus Hentschel and Ann M. Hentschel (Trans.); Stanford University Press, 270 pp., US $51, ISBN 0-8047-2824-0. Studies in History and Philosophy of Modern Physics, 33(3):591– 599, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219802000278. See [Hen97].

Smith:2003:E

[Smi03] Peter D. (Peter Daniel) Smith. Einstein. Haus, London, UK, 2003. ISBN 1-904341-14-4, 1-904341-15-2 (paperback). 170 pp.

LCCN QC16.E5 S65 2003. URL http://catalog.hathitrust.
org/api/volumes/oclc/51964351.html.

Smith:2004:ASL

[Smi04] P. D. Smith. In Albert’s shadow — the life and letters of Mileva Marić, Einstein’s first wife. TLS — The Times Literary Supplement, ??(5259):28, January 16, 2004. ISSN 0307-661X.

Smil:2005:BRD

[Smi05a] Vaclav Smil. Book review: Dietrich Stoltzenberg: Fritz Haber: Chemist, Nobel Laureate, German, Jew. Isis, 96(2):310–312, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/
491536.

Smith:2005:AEM

[Smi05b] P. D. Smith. Albert Einstein — man of the century. TLS — The Times Literary Supplement, ??(5354):17, November 11, 2005. ISSN 0307-661X.

Smith:2007:AMA

[Smi07a] P. D. Smith. Albert meets America — how journalists treated genius Einstein’s 1921 travels. TLS — The Times Literary Supplement, ??(5425):33, March 23, 2007. ISSN 0307-661X.

Smith:2007:DMR

[Smi07b] Peter D. (Peter Daniel) Smith. Doomsday men: the real Dr. Strangelove and the dream of the superweapon. St. Martin’s Press, New York, NY, USA, 2007. ISBN 0-312-37397-X. xxii + 552 pp. LCCN QC773.A1 S65 2007. URL http://www.loc.

gov/catdir/enhancements/fy0801/2007032786-b.html;
http://www.loc.gov/catdir/enhancements/fy0801/2007032786-
d.html.

Smith:2010:NLI

[Smi10] George E. Smith. Nobel Lecture: The invention and early history of the CCD. Reviews of Modern Physics, 82(3): 2307–2312, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2307;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2307_1.

Smith:2015:CPC

[Smi15a] Daniel Smith. Comment penser comme Einstein. (French) [How to think like Einstein]. Marabout, Paris, France, 2015. ISBN 2-501-09968-0 (hardcover). xxii + 213 pp. LCCN ???? URL

http://catalogue.bnf.fr/ark:/12148/cb44265922m.

Smith:2015:DWE

[Smi15b] Daniel Smith. Denken wie Einstein. (German) [Thinking like Einstein]. Anaconda Verlag, Köln, Germany, 2015. ISBN 3-7306-0269-1. 256 pp. LCCN ????

Smith:2021:LEM

[Smi21]  Meg Weston Smith.  Away from the limelight: Ein-

stein and Milne. Notes and Records of the Royal Society of London, 75(1):159–165, March ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2019.0047.

Seaborg:1946:RED

[SMKW46] G. T. Seaborg, E. M. Mcmillan, J. W. Kennedy, and A. C. Wahl. Radioactive element 94 from deuterons on uranium. Physical Review, 69(7–8):366–367, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.69.366.2.

Szilard:1960:SWSa

[SMM^{+}60] Leo Szilard, Thomas Edward Murray, John B. (John Bruce) Medaris, Edward Teller, and Jerome B. (Jerome Bert) Wiesner, editors. Small world: Szilard, Murray, Medaris, Teller, Wiesner [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? 28 minute 16mm film.

Smolin:2006:TPR

[Smo06] Lee Smolin. The trouble with physics: the rise of string theory, the fall of a science, and what comes next. Houghton-Mifflin, Boston, MA, USA, 2006. ISBN 0-618-55105-0. xxiii + 392 pp.

LCCN QC6 .S6535 2006. URL http://www.loc.gov/catdir/
enhancements/fy0634/2006007235-d.html;   http://www.
loc.gov/catdir/enhancements/fy0666/2006007235-s.html;
 http://www.loc.gov/catdir/toc/ecip069/2006007235.html.

Smolin:2007:OEB

[Smo07a] Lee Smolin. The other Einstein: Einstein: His Life and Universe by Walter Isaacson. Einstein: A Biography by Jürgen Neffe, translated from the German by Shelley Frisch. Subtle Is the Lord: The Science and the Life of Albert Einstein by Abraham Pais. The Private Lives of Albert Einstein by Roger Highfield and Paul Carter. Einstein in Love: A Scientific Romance by Dennis Overbye. Einstein’s Clocks, Poincaré’s Maps: Empires of Time by Peter Galison. Einstein on Politics edited by David Rowe and Robert Schulmann. Einstein on Race and Racism by Fred Jerome and Rodger Taylor. The Collected Papers of Albert Einstein by Albert Einstein. New York Review of Books, 54(10):76–??, June 14, 2007. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2007/jun/14/the-other-einstein/.

Smolin:2007:TPR

[Smo07b]  Lee Smolin.  The Trouble with Physics: the Rise of

String Theory, the Fall of a Science, and What Comes Next. Mariner Books, Boston, MA, USA, 2007. ISBN

0-618-91868-X (paperback).  xxiii + 392 pp.  LCCN

QC6 .S6535 2007.  URL http://www.loc.gov/catdir/
enhancements/fy0634/2006007235-d.html;   http://www.
loc.gov/catdir/enhancements/fy0737/2006007235-b.html;
 http://www.loc.gov/catdir/toc/ecip069/2006007235.html.

Smoot:2007:NLC

[Smo07c] George F. Smoot. Nobel Lecture: Cosmic microwave background radiation anisotropies: Their discovery and utilization. Reviews of Modern Physics, 79(4):1349–1379, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.79.1349;   http://rmp.aps.org/abstract/RMP/
v79/i4/p1349_1.

Smolin:2013:TRC

[Smo14] Lee Smolin. Time Reborn: From the Crisis in Physics to the Future of the Universe. Mariner Books, Boston, MA, USA, 2014. ISBN 0-544-24559-8. xxxi + 319 pp. LCCN QB991.Q36 S66 2014.

Smolin:2019:CRB

[Smo19a] Lee Smolin. The challenge of realism: de Broglie and Einstein. In Einstein’s Unfinished Revolution: the Search for What Lies Beyond the Quantum [Smo19b], chapter 7, pages 97–106. ISBN 1-59420-619-8 (hardcover), 0-698-16913-1 (e-book). LCCN QC174.13 .S6545 2019. URL https:

//www.penguinrandomhouse.com/books/316818/einsteins-
unfinished-revolution-by-lee-smolin/9781594206191. Il-

lustrations by Kaca Bradonjic.

Smolin:2019:EUR

[Smo19b] Lee Smolin. Einstein’s Unfinished Revolution: the Search for What Lies Beyond the Quantum. Penguin Press, New York, NY, USA, 2019. ISBN 1-59420-619-8 (hardcover), 0-698-16913-1 (e-book). xxix + 322 pp. LCCN QC174.13

.S6545 2019. URL https://www.penguinrandomhouse.com/
books/316818/einsteins-unfinished-revolution-by-lee-

smolin/9781594206191. Illustrations by Kaca Bradonjic.

Schlick:1979:PP

[SMvdVS79] Moritz Schlick, Henk L. Mulder, and Barbara F. B. van de Velde-Schlick. Philosophical papers, volume 11 of Vienna circle collection. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0314-0 (vol. 1). various pp. LCCN B3329.S4872 E5 1979. English translation by Peter Heath of [Sch15].

Shikibu:2009:L

[SMW^{+}09] Murasaki Shikibu, Jack McLaughlin, Richard White, Marge Piercy, Jacob Riis, John Berryman, Ronald Blythe, Susan B. Anthony, Henry Mayhew, Albert Einstein, and Henry David Thoreau. Labor. American Scholar, 78(4):118–119, Fall 2009. ISSN 0003-0937.

Smyth:1945:AEMb

[Smy45a] H. D. Smyth. Atomic energy for military purposes. Reviews of Modern Physics, 17(4):351–471, October 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.351;   http://rmp.aps.org/abstract/RMP/v17/i4/p351_
1.

Smyth:1945:AEMa

[Smy45b] Henry DeWolf Smyth. Atomic energy for military purposes; the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. Princeton University Press, Princeton, NJ, USA, 1945. ix + 1 + 264 pp. LCCN QC173 .S4735 1945a. URL http:/

/www.archive.org/details/atomicenergyform00smytrich;
 http://www.atomicarchive.com/Docs/SmythReport/index.
shtml;   https://en.wikipedia.org/wiki/Smyth_Report.

Smyth:1945:GADa

[Smy45c] Henry DeWolf Smyth. A general account of the development of methods of using atomic energy for military purposes under the auspices of the United States Government, 1940–1945. ????, ????, 1945. (various) pp. LCCN QC173 .S474 1945a.

Smyth:1945:GADb

[Smy45d] Henry DeWolf Smyth. A general account of the development of methods of using atomic energy for military purposes under the auspices of the United States Government, 1940–1945. H. M. Stationery Office, London, UK, 1945. iv + 143 + 1 pp. LCCN QC173 .S474 1945b.

Smyth:1945:GADc

[Smy45e] Henry DeWolf Smyth. A general account of the development of methods of using atomic energy for military purposes under the auspices of the United States Government, 1940–1945. United States Government Printing Office, Washington, DC, USA, 1945. vii + 182 pp. LCCN QC173 .S474 1945.

Smyth:1946:ACD

[Smy46] Henry DeWolf Smyth. Atomnaëiìa çenergiëiìa dlëiìa voennykh ëtìseleĭ; ofiëtìsialsnyi otchet o razrabotke atomnoi bomby pod nablëiìudeniem pravitelstva SShA. (Russian) [Atomic energy for military purposes]. Gos. transp. zhel-dor. izd-vo, Moskva, USSR, 1946. 276 pp. LCCN QC173 .S47477. Translated from English by G. N. Ivanova.

Smyth:1948:AEM

[Smy48] Henry DeWolf Smyth. Atomic Energy for Military Purposes. Princeton University Press, Princeton, NJ, USA, 1948. vii + 308 pp. LCCN QC173 .S4735 1948.

Smyth:1976:AEM

[Smy76a] Henry DeWolf Smyth. Atomic energy for military purposes: the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. The Politics and strategy of World War II. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1976. ISBN 0-306-70767-5. ix + 264 + 4 pp. LCCN QC773.A1 S69 1976.

Smyth:1976:SR

[Smy76b] Henry DeWolf Smyth. The “Smyth Report”. The Princeton University Library Chronicle, 37(3):173–190, Spring 1976. ISSN

0032-8456. URL http://libweb5.princeton.edu/visual_
materials/pulc/pulc_v_37_n_3.pdf. This is a historical ret-

rospective on the author’s famous report [Smy45b, Smy45a] that described the Manhattan Project to the public. The periodical issue in which this article appears (the whole issue is in the PDF file) also contains other articles about the production and dissemination of the report. It is believed to have been translated into at least 40 languages.

Smyth:1978:AEM

[Smy78] Henry DeWolf Smyth. Atomic energy for military purposes: the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. AMS Press, New York, NY, USA, 1978. ISBN 0-404-14703-8. vii + 308 + 4 pp. LCCN QC773.3.U5 S67 1978.

Smyth:1989:AEM

[Smy89] Henry DeWolf Smyth. Atomic energy for military purposes: the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1721-4 (clothbound), 0-8047-1722-2 (paperback). xv + 324 pp. LCCN QC773.A1 S69 1989.

Sharov:1993:EHD

[SN93] A. S. (Aleksandr Sergeevich) Sharov and I. D. (Igor Dmitrievich) Novikov. Edwin Hubble, the discoverer of the big bang universe. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41617-5. xv + 187 pp. LCCN

QB36.H83 S5313 1993. URL http://www.loc.gov/catdir/
description/cam025/92024683.html;   http://www.loc.
gov/catdir/toc/cam028/92024683.html.

Schreiber:2006:GSE

[SNA^{+}06] Liev Schreiber, Liam Neeson, F. Murray Abraham, Peter Thomas, John Lithgow, Simon Callow, Scott Handy, Aidan McArdle, David R. Axelrod, Peter Jones, David Espar, Susan Kopman Lewis, Alastair Reid, Allan Cubitt, Christopher Oxley, Christopher Sykes, Gary Johnstone, Dava. Sobel, and David Bodanis. Genius: the science of Einstein, Feynman, Newton, Darwin, and Galileo. WGBH Boston Video, S. Burlington, VT, USA, widescreen and standard formats. edition, 2006. ISBN 1-59375-575-9, 1-59375-573-2 (Galileo’s battle for the heavens), 1-57807-858-X (Darwin’s dangerous idea), 1-59375-323-3 (Newton’s dark secrets), 1-59375-576-7 (The best mind since Einstein), 1-59375-317-9 (Einstein’s big idea). LCCN Q158.5 .G46 2006. 5 videodiscs (ca. 480 min.).

Snell:1982:GSH

[Sne82] Arthur H. Snell. Graveyard shift, Hanford, 28 September 1944 — Henry W. Newson. American Journal of Physics, 50 (4):343–348, April 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.aip.

org/content/aapt/journal/ajp/50/4/10.1119/1.13039.

Snow:1961:SG

[Sno61] C. P. (Charles Percy) Snow. Science and government, volume 1960 of The Godkin lectures at Harvard University. Harvard University Press, Cambridge, MA, USA, 1961. 88 pp. LCCN Q127.G4 S62 1961.

Snow:1981:P

[Sno81] C. P. (Charles Percy) Snow. The physicists. Macmillan Publishing Company, New York, NY, USA, 1981. ISBN 0-333-32228-2. 192 pp. LCCN ???? Introduction by William Cooper.

Synge:1972:GRP

[SO72] J. L. (John Lighton) Synge and L. (Lochlainn) O’Raifeartaigh, ed, editors. General relativity; papers in honour of J. L. Synge. Clarendon Press, Oxford, UK, 1972. ISBN 0-19-851126-4. ix + 277 pp. LCCN QC6 .G358. Edited for the Royal Irish Academy by L. O’Raifeartaigh.

Sherby:2002:WWN

[SO02] Louise S. Sherby and Wilhelm Odelberg, editors. The Who’s Who of Nobel Prize winners, 1901–2000. Oryx Press, Westport, CT, USA, fourth edition, 2002. ISBN 1-57356-414-1. xxiii + 277 pp. LCCN AS911.N9 W53 2002. URL http://www.e-streams.

com/es0805/es0805_2063.htm.

Sodei:1995:HNH

[Sod95] Rinjiro Sodei. Hiroshima/Nagasaki as history and politics. The Journal of American History, 82(3):1119–1123, ???? 1995.

Soderlind:2010:NBC

[Söd10] Ulrica Söderlind. The Nobel banquets: a century of culinary history (1901–2001). World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010. ISBN 981-4313-11-4 (hardcover), 981-4317-97-7 (paperback). xviii + 319 pp. LCCN AS911.N9 S5813 2010.

Solovine:1956:AEL

[Sol56] Maurice Solovine, editor. Albert Einstein: Lettres à Maurice Solovine: 1906–1966. Gauthier-Villars, Paris, France, 1956. xiii + 139 pp. LCCN ????

Solovine:1959:FAE

[Sol59] Maurice Solovine. Freundschaft mit Albert Einstein. (German) [Friendship with Albert Einstein]. Physikalische Blätter, 15(3): 97–103, ???? 1959. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Solonche:1987:PCC

[Sol87] J. R. Solonche. ‘On a photograph of Charlie Chaplin and Albert Einstein’. American Scholar, 56(1):70, Winter 1987. ISSN 0003-0937 (print), 2162-2892 (electronic).

Solomon:2016:CBE

[Sol16] Adam Ross Solomon. Cosmology beyond Einstein. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-46620-8 (hardcover), 3-319-46621-6 (ebook). 239 pp. LCCN ????

Sommerfeld:1949:AEJ

[Som49a] Arnold Sommerfeld. Albert Einstein 70 Jahre. (German) [Albert Einstein 70 years]. Physikalische Blätter, 5(??):127–??, ???? 1949. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Sommerfeld:1949:AEG

[Som49b] Arnold Sommerfeld. Albert Einstein zum 70. Geburtstag. (German) [Albert Einstein’s 70th birthday]. Dtsch. Beiträge Jahrg., 3(??):3–??, ???? 1949.

Sommerfeld:1949:AES

[Som49c] Arnold Sommerfeld. To Albert Einstein’s seventieth birthday. In Schilpp [Sch49b], chapter 1, pages 97–105. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Sommerfeld:1955:AE

[Som55a] Arnold Sommerfeld. Albert Einstein. In Schilpp [Sch55], pages

37–42. LCCN ???? URL http://www.gbv.de/dms/faz-rez/
560630_FAZ_0037_5_0001.pdf.

Sommerfeld:1955:AEP

[Som55b] Arnold Sommerfeld. Albert Einstein als Philosoph und Naturforscher. (German) [Albert Einstein as philosopher and natural science researcher]. In ????, pages 37–42. Kohlhammer, Stuttgart, West Germany, 1955.

Sommer:2005:DBH

[Som05a] Klaus P. Sommer. In das Deutschland ,,von Hilbert und Einstein”. Briefe von Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl und Althoff an David Hilbert, gefunden auf einem Göttinger Dachboden. (German) [In the Germany of “Hilbert and Einstein”. Letters of Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl and Althoff to David Hilbert, found in a Goettingen loft]. Berichte zur Wissenschaftsgeschichte, 28(4):283–303, December 2005. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Sommer:2005:DHE

[Som05b] Klaus P. Sommer. In das Deutschland ,,von Hilbert und Einstein”. Briefe von Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl und Althoff an David Hilbert, gefunden auf einem Göttinger Dachboden. (German) [In the Germany “of Hilbert and Einstein.” Letters of Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl and Althoff to David Hilbert, found in a loft in Göttingen]. Berichte zur Wissenschaftsgeschichte, 28(4), December 2005. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Sopka:1988:QPA

[Sop88] Katherine Russell Sopka. Quantum physics in America: the years through 1935, volume 10 of The history of modern physics, 1800-1950. Tomash, Los Angeles, CA, USA, 1988. ISBN 0-88318-553-9. xxvi + 376 pp. LCCN QC173.98 .S67 1988. Foreword by Abraham Pais. Introduction by Gerald Holton.

Sourkes:1966:NPW

[Sou66] Theodore L. (Theodore Lionel) Sourkes. Nobel Prize winners in medicine and physiology, 1901–1965, volume 45 of The Life of science library. Abelard-Schuman, London, UK, 1966. ix + 464 pp. LCCN ????

Sousa:2019:EEE

[Sou19] Ana Matilde Sousa. Einstein Eddington e o eclipse: impressäoes de viagem = Einstein, Eddington and the eclipse: travel impressions. Associaçäao Chili com Carne, Lisboa, Portugal, 2019. ISBN 989-8363-41-X (paperback). 245 pp. LCCN QB544.19 .S56 2019. Translation to English by Ana Simäoes.

Sowell:2001:ESB

[Sow01] Thomas Sowell. The Einstein syndrome: bright children who talk late. Basic Books, New York, NY, USA, 2001. ISBN 0-465-08141-X. xiv + 217 pp.

Singh:1959:BES

[SP59] A. D. Singh and R. K. Pathria. Bose–Einstein statistics and helium films. Nature, 183(4662):668, March 7, 1959. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v183/n4662/
pdf/183668a0.pdf.

Schilling:2012:BMI

[SP12] René L. Schilling and Lothar Partzsch. Brownian Motion: An Introduction to Stochastic Processes. Walter de Gruyter, Berlin, Germany, 2012. ISBN 3-11-027889-8 (paperback), 3-11-027898-7 (e-book). xiv + 380 pp. LCCN QA274.75 .S35 2012. URL http:

//www.degruyter.com/doi/book/10.1515/9783110278989;
 http://www.degruyter.com/viewbooktoc/product/181486.

Sparberg:1958:AEB

[Spa58] Esther B. Sparberg. The atomic energy breakthrough. Thesis Ph.D., Science Manpower Project, Teachers College, Columbia University, New York, NY, USA, 1958. 230–444 pp.

Sparberg:1964:SDF

[Spa64] Esther B. Sparberg. A study of the discovery of fission. American Journal of Physics, 32(1):2–8, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Shapiro:1968:EFT

[SPA^{+}68a] Irwin I. Shapiro, Gordon H. Pettengill, Michael E. Ash, Melvin L. Stone, William B. Smith, Richard P. Ingalls, and Richard A. Brockelman. Erratum: Fourth test of General Relativity: Preliminary results. Physical Review Letters, 21(4):266, July 22, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/doi/

10.1103/PhysRevLett.21.266.3.

Shapiro:1968:FTG

[SPA^{+}68b] Irwin I. Shapiro, Gordon H. Pettengill, Michael E. Ash, Melvin L. Stone, William B. Smith, Richard P. Ingalls, and Richard A. Brockelman. Fourth test of general relativity: Preliminary results. Physical Review Letters, 20(22):1265–1269, May 3, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/doi/

10.1103/PhysRevLett.20.1265. See erratum [SPA^{+}68a].

Spagnou:2017:MDT

[Spa17] Pierre Spagnou. Les mystères du temps: de Galilée à Einstein. (French) [The mysteries of time: from Galileo to Einstein]. Le Banquet scientifique. CNRS Éditions, Paris, France, 2017. ISBN 2-271-08911-5 (paperback). ISSN 2109-8638. 277 pp. LCCN ????

Speziali:1962:AOB

[Spe62] Pierre Speziali. Analyses d’ouvrages: Lettres à Maurice Solovine, reproduites en facsimilé et traduites en français avec une introduction et trois photographies par Albert Einstein. Revue d’Histoire des Sciences et de Leurs Applications, 15(1):84– 85, January 1962. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23904814.

Speziali:1967:AOBc

[Spe67] Pierre Speziali. Analyses d’ouvrages: La théorie de la relativité restreinte, exposé élémentaire par Albert Einstein; Henri Montias. Revue d’Histoire des Sciences et de Leurs Applications, 20 (4):393–394, October 1967. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23904717.

Speziali:1968:AOBc

[Spe68] Pierre Speziali. Analyses d’ouvrages: Relativitätstheorie und Weltanschauung. Zur philosophischen und wissenschaftlichen Wirkung Albert Einsteins (Von einem Autorenkollektiv). Revue d’Histoire des Sciences et de Leurs Applications, 21(3):268–271, juillet–septembre 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23909921.

Speziali:1972:AEM

[Spe72] Pierre Speziali, editor. Albert Einstein et Michele Besso: correspondance 1903–1955, volume 17 of Centre de recherches historiques. Hermann, Paris, France, 1972. lxiii + 559 pp. LCCN ????

Specker:1979:EUF

[Spe79] Hans Eugen Specker. Einstein und Ulm: Festakt, Schülerwettbewerb und Ausstellung zum 100. Geburtstag von Albert Einstein. (German) [Einstein and Ulm: ceremony, student competition and exhibition for the 100th birthday of Albert Einstein], volume 1 of Forschungen zur Geschichte der Stadt Ulm: Reihe Dokumentation. Kohlhammer, Stuttgart, Germany, 1979. ISBN 3-17-005522-4. 135 pp. LCCN QC16.E5 E53.

Specter:1995:ESQ

[Spe95] Michael Specter. Einstein’s son? It’s a question of relativity. New York Times, ??(??):1, July 22, 1995. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

https://search.proquest.com/hnpnewyorktimes/docview/
109446895/.

Speake:2005:NCT

[Spe05] Clive C. Speake. Newton’s constant and the twenty-first century laboratory. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 363(1834): 2265–2287, September 15, 2005. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Spitz:1996:CVI

[Spi96] H. H. Spitz. Commentary on Vandervert’s “From idiots savants to Albert Einstein”. New Ideas in Psychology, 14(1):93– 96, March 1996. ISSN 0732-118X. See [Van96].

Spielberg:2005:SIS

[Spi05] Nathan Spielberg. Seven ideas that shook the universe. MJF Books/Fine Communications, New York, NY, 2005. ISBN 1-56731-707-3. ???? pp. LCCN ????

Sponsel:2002:CRS

[Spo02] Alistair Sponsel. Constructing a “revolution in science”: the campaign to promote a favourable reception for the 1919 solar eclipse experiments. British Journal for the History of Science, 35(4):439–467, December 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4028276.

Spotts:2016:EVM

[Spo16] Pete Spotts. Einstein verified: ‘monumental’ gravitational waves discovery was 100 years in making. An international team of researchers announced Thursday that they became the first group to detect a cosmic phenomenon that Albert Einstein first predicted 100 years ago. The Christian Science Monitor, ??(??):??, February 11, 2016. ISSN 0882-7729 (print), 1540-4617 (electronic). URL https://www.csmonitor.

com/Science/Spacebound/2016/0211/Einstein-verified-
Monumental-gravitational-waves-discovery-was-100-years-
in-making.

Serber:1992:APF

[SR92] R. (Robert) Serber and Richard Rhodes. The Los Alamos primer: the first lectures on how to build an atomic bomb. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1992. ISBN 0-520-07576-5. xxxiii + 98 + 8 pp. LCCN QC773.A1 S47 1992. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/bios/ucal051/91014068.html;   http://www.loc.
gov/catdir/description/ucal041/91014068.html.

Singh:1999:CVR

[SR99] Rajinder Singh and Falk Riess. C. V. Raman, M. N. Saha and the Nobel Prize for the year 1930. Indian Journal of the History of Science, 34(??):61–75, ???? 1999. CODEN IJHSA4.

ISSN 0019-5235. URL http://insa.nic.in/writereaddata/
UpLoadedFiles/IJHS/Vol34_1_5_RSingh.pdf.

Singh:2001:NPP

[SR01] R. Singh and F. Riess. The 1930 Nobel Prize for Physics: A close decision? Notes and Records of the Royal Society of London, 55 (2):267–283, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532100.

Singh:2004:NLS

[SR04] Rajinder Singh and Falk Riess. The Nobel laureate Sir Chandrasekhara Venkata Raman FRS and his contacts with the British scientific community in a social and political context. Notes and Records of the Royal Society of London, 58(1):47–64, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
4142032.

Simkin:2011:REA

[SR11] Mikhail Simkin and Vwani Roychowdhury. Von Richthofen, Einstein and the AGA: Estimating achievement from fame. Significance (Oxford, England), 8(1):22–26, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sretenovic:1929:LRS

[Sre29] Misa Sretenović. Ljubavni roman slavnog naučnika Alberta Ajnstajna i Srpkinje, Mileve Marić. (Serbian) [Romance novel by famous scientist Albert Einstein and Serb woman, Mileva Marić]. Politika (Belgrade), ??(??):??, May 23, 1929.

SriKantha:1994:AAE

[Sri94] S. Sri Kantha. An appraisal of Albert Einstein’s chronic illness. Medical Hypotheses, 42(5):340–346, ???? 1994. CODEN MEHYDY. ISSN 0306-9877 (print), 1532-2777 (electronic).

Srinivasan:1999:WDB

[Sri99] Ganesan Srinivasan, editor. From white dwarfs to black holes: the legacy of S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-76996-8. 240 pp. LCCN QB51.

Sohlman:1929:NDP

[SS29] Ragnar Sohlman and Henrik Schück. Nobel, dynamite and peace. Cosmopolitan Book Corporation, New York, NY, USA, 1929. 353 pp. LCCN TP268.5.N7 N73. Translation by Brian Lunn and Beatrix Lunn of the 1926 Swedish original Alfred Nobel och hans släkt [Alfred Nobel and his family].

Savage:1965:RUT

[SS65] John Savage and Barbara Storms. Reach to the unknown: The Trinity Story, July 16, 1945. Atom, 2(8):??, ???? 1965.

Schiavulli:1979:FCE

[SS79] L. Schiavulli and F. Selleri. Further consequences of Einstein locality. Foundations of Physics, 9(5–6):339–352, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00708526.

Schlessinger:1986:WWN

[SS86] Bernard S. Schlessinger and June H. Schlessinger, editors. The Who’s Who of Nobel Prize winners. Oryx Press, Phoenix, AZ, USA, 1986. ISBN 0-89774-136-6 , 0-89774-193-5. xii + 212 pp. LCCN AS911.N9 W53 1986.

Schlessinger:1991:WWN

[SS91] Bernard S. Schlessinger and June H. Schlessinger, editors. The Who’s Who of Nobel Prize winners, 1901–1990. Oryx Press, Phoenix, AZ, USA, second edition, 1991. ISBN 0-89774-599-X. xxii + 234 pp. LCCN AS911.N9 W53 1991.

Schlessinger:1996:WWN

[SS96] Bernard S. Schlessinger and June H. Schlessinger, editors. The Who’s Who of Nobel Prize winners, 1901–1995. Oryx Press, Phoenix, AZ, USA, third edition, 1996. ISBN 0-89774-899-9. xx + 251 pp. LCCN AS911.N9 W53 1996.

Schlogl:1997:WTP

[SS97] Reinhard Schlögl and Erwin Schrödinger. Wellen und Teilchen — Pionierleistungen der modernen Physik: Portrait des Nobelpreisträgers Erwin Schrödinger zum 110. Geburtstag. (German) [Waves and particles — pioneering achievements of modern physics: Portrait of the Nobel laureate Erwin Schrödinger on his 110th birthday]. Radiosendung 19.8.1997 [Radio broadcast 19.8.1997]., August 19, 1997.

Samuelson:1998:NLI

[SS98a] Bengt Samuelson and Michael Sohlman. Nobel lectures: including presentation speeches and laureates’ biographies, physics: 1942–1962. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3403-1. ISSN 0169-006X. xi + 619 pp. LCCN ????

Samuelsson:1998:NLP

[SS98b] Bengt Samuelsson and Michael Sohlman, editors. Nobel Lectures in Physics (1963–1970), volume 4 of Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3404-X. ix + 349 pp. LCCN QC71 .P455 1998.

Samuelsson:1998:NLI

[SS98c] Bengt Samuelsson and Michael Sohlman, editors. Nobel Lectures: Including Presentation Speeches and Laureates’ Biographies, Physics: 1922–1941. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3402-3. ISSN 0169-006X. x + 456 pp. LCCN ????

Seaborg:2001:AAA

[SS01] Glenn Theodore Seaborg and Eric Seaborg. Adventures in the atomic age: from Watts to Washington. Farrar, Straus and Giroux, New York, NY, USA, 2001. ISBN 0-374-29991-9 (hardcover). viii + 312 + 24 pp. LCCN QD22.S436 A3 2001. URL

http://www.loc.gov/catdir/bios/hol054/00049522.html;
 http://www.loc.gov/catdir/description/hol042/00049522.
html;   http://www.loc.gov/catdir/toc/fy022/00049522.
html.

Schwarz:2004:SRE

[SS04] Patricia M. (Patricia Margaret) Schwarz and John H. Schwarz. Special Relativity: from Einstein to strings. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81260-7. xii + 376 pp. LCCN QC173.65 SCH.

Schucking:2015:EAH

[SS15] E. L. (Engelbert L.) Schucking and Eugene J. Surowitz. Einstein’s apple: homogeneous Einstein fields. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4630-07-1 (hardcover), 981-4630-08-X (ebook). xii + 301 pp. LCCN QC178 .S35 2015.

Sauer:2016:PHC

[SS16] Tilman Sauer and Raphael Scholl, editors. The Philosophy of Historical Case Studies, volume 319 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346. viii + 296 pp. LCCN Q126.8 .P45 2016. URL https://link.

springer.com/book/10.1007/978-3-319-30229-4.

Sauer:2021:EIP

[SS21a] Tilman Sauer and Tobias Schütz. Einstein on involutions in projective geometry. Archive for History of Exact Sciences, 75(5): 523–555, September 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-020-00270-z.

Sauer:2021:EWM

[SS21b] Tilman Sauer and Tobias Schütz. Einstein’s Washington manuscript on unified field theory. Berichte zur Wissenschaftsgeschichte, 44(1):94–105, March 2021. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Sanchez:2013:NMH

[SSG13] Joseph P. Sánchez, Robert L. S. Spude, and Art Gómez. New Mexico: a history. University of Oklahoma Press, Norman, OK, USA, 2013. ISBN 0-8061-4256-1, 1-4619-4445-7 (e-book). LCCN F796 .S26 2013.

Steger:2013:SSP

[SSH13] Volker Steger, Adam Smith, and Tim Hunt. Sketches of science: photo sessions with Nobel laureates, volume 18 of Archives of the Nobel Museum, 1404-7586. Nobelmuseum, Stockholm, Sweden, 2013. ISBN 3-939201-04-9 (hardcover). 232 pp. LCCN ???? Issued on the occasion of the Traveling Exhibition “Sketches of Science” by the Nobel Museum: Stockholm, Nobel Museum 8 June–30 September 2012: Frankfurt, Airport Terminal A, 5 December 2012–9 January 2013: Berlin, Landesvertretung Baden-Württemberg, 18 January–7 February 2013: Heidelberg, Carl-Bosch-Museum, 22 February–2 June 2013: Mainau Island, Mainau Castel, 24 June–31 August 2013.

Silberstein:2021:BCE

[SSM21] Michael Silberstein, William Stuckey, and Timothy McDevitt. Beyond causal explanation: Einstein’s principle not Reichenbach’s. Entropy, 23(1):114, ???? 2021. CODEN ENTRFG. ISSN 1099-4300. URL https://dblp.org/db/

journals/entropy/entropy23.html#SilbersteinSM21.

Soles:1998:ELP

[SSS^{+}98] Paul Soles, LaTaye Studwood, Raymond Storey, David Devine, and Richard Mozer. Einstein: light to the power of 2, 1998. ISBN 1-894449-05-3. 1 videocassette (55:05 min.).

Sudoplatov:1995:STM

[SSSS95] Pavel Sudoplatov, Anatolii Pavlovich Sudoplatov, Jerrold L. Schecter, and Leona Schecter. Special tasks: the memoirs of an unwanted witness, a Soviet spymaster. Little, Brown and Co., Boston, MA, USA, updated edition, 1995. ISBN 0-316-77352-2 (hardcover), 0-316-82115-2 (paperback). xxxi + 527 pp. LCCN JN6529.I6 S83 1995.

See:1923:SFE

[ST23] T. J. J. See and Robert Trumpler. Soldner, Foucault and Einstein. Science, 58(1506):372, November 9, 1923. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/58/1506/372.1.
full.pdf.

Schroder:1994:ELD

[ST94a] W. Schröder and H.-J. Treder. The ‘Einstein–Laue’ discussion. British Journal for the History of Science, 27(1):113–114, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027585.

Schroder:1994:ELV

[ST94b] Wilfried Schröder and Hans-Jürgen Treder. Zu Einsteins letzter Vorlesung — Beobachtbarkeit, Realität und Vollständigkeit in Quanten- und Relativitätstheorie. (German) [On Einstein’s last lecture — observability, reality, and completeness in quantum and Relativity theory]. Archive for History of Exact Sciences, 48(2):149–154, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=48&issue=2&spage=149.

Schroder:1998:ECP

[ST98] Wilfried Schröder and Hans-Jürgen Treder. Einstein and cosmic physics. Foundations of Physics, 28(6):1013–1020, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1018829616282.

Schroder:2006:AME

[ST06] W. Schröder and H.-J. Treder. Aspects of the Mach–Einstein doctrine and geophysical application (a historical review). Foundations of Physics, 36(6):883–901, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
006-9050-3.

Stark:1936:PLD

[Sta36] Johannes Stark. Philipp Lenard als deutscher Naturforscher. (German) [Philipp Lenard as German natural scientist]. Nationalsozialistische Monatshefte, 71(??):106–111, January 1936.

Stang:1979:PNP

[Sta79] Fredrik Stang. Perspective: Nobel Peace Laureates: Carl von Ossietzky, 1935. Bulletin of the Atomic Scientists, 35(2):1, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Von Ossietzky was awarded the Prize, but the Nazi government refused to release him from a concentration camp, and decreed that no German in the future could accept any Nobel Prize. Von Ossietzky died in hospital under Nazi surveillance in May 1938.

Stachel:1980:ERR

[Sta80a] John Stachel. Einstein and the rigidly rotating disk. In Held [Hel80], pages 1–15. ISBN 0-306-40265-3 (vol. 1), 0-306-40266-1 (vol. 2). LCCN QC173.6 G464. Reprinted in [Sta89a].

Stachel:1980:ESG

[Sta80b] John Stachel. Einstein’s search for general covariance, 1912– 1915. In Schmutzer [Sch83b], page 312. ISBN 0-521-24669-5. LCCN QC173.6 .I57 1980. Reprinted in [HS89b].

Stachel:1982:E

[Sta82a] John Stachel. Einstein. Science, 218(4576):989–990, December 3, 1982. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/218/4576/989.full.pdf.

Stachel:1982:EMC

[Sta82b] John Stachel. Einstein and Michelson: The context of discovery and the context of justification. Astronomische Nachrichten, 303 (1):47–53, 1982. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Stachel:1986:EE

[Sta86] John Stachel. Eddington and Einstein. In Ullmann-Margalit [UM86], pages 225–250. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. LCCN

Q174 .B67 vol. 95 Q175. URL http://link.springer.com/
chapter/10.1007/978-94-009-4566-1_18/.

Stachel:1987:MMT

[Sta87] John Stachel. “A man of my type”—editing the Einstein papers. British Journal for the History of Science, 20(1):57–66, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). Conference on documentary editing (Tarrytown, NY, 1985).

Stachel:1989:ERR

[Sta89a] John Stachel. Einstein and the rigidly rotating disk. In Howard and Stachel [HS89b], pages 48–62. ISBN 0-8176-3392-8, 3-7643-3392-8. LCCN QC173.6 .E38 1989. URL http://www.

springer.com/us/book/9780817633929. Reprint of [Sta80a].

Stahle:1989:ANN

[Sta89b] Nils K. Stahle. Alfred Nobel and the Nobel Prizes. Swedish Institute, Stockholm, Sweden, 1989. ISBN 91-520-0241-1. 31 pp. LCCN MLCS 92/11407 (A).

Stannard:1989:ICR

[Sta89c] Russell Stannard. Introducing children to relativity. Physics World, 2(5):23–24, May 1989. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/2/5/phwv2i5a22.
pdf.

Staguhn:1992:GLM

[Sta92a] Gerhard Staguhn. God’s laughter: man and his cosmos. Harper-Collins, New York, NY, USA and London, UK, 1992. ISBN 0-06-019004-3. viii + 255 pp. LCCN BD513 .S7913 1992. US$23.00.

Stannard:1992:TNO

[Sta92b] Russell Stannard. Les trous noirs et l’oncle Albert. École des loisirs, Paris, France, 1992. ISBN 2-211-02756-3. 252 pp. LCCN ???? French translation by Raphaëlle Desplechin and Fabrice Desplechin of [SL91].

Stachel:1993:NPE

[Sta93a] John Stachel. Nobel Prize: Einstein was right. Physics World, 14(6):7, June 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/6/11/phwv6i11a5.
pdf.

Stachel:1993:WER

[Sta93b] John Stachel. Why Einstein reinvented the ether. Physics World, 14(6):55–56, June 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/14/6/phwv14i6a33.
pdf.

Stawstrom:1993:EWR

[Sta93c] Carl-Olov Stawström. Einstein at work, rest and play. Physics World, 6(12):53–54, December 1993. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
6/12/phwv6i12a35.pdf.

Stachel:1994:LEC

[Sta94] John Stachel. Lanczos’s early contributions to Relativity and his relationship with Einstein. In Brown et al. [BCEP94], pages 201–221. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Stachel:1995:HR

[Sta95a] John Stachel. History of relativity. In Brown et al. [BPP95], pages 249–256. ISBN 1-56396-314-0 (US: set), 1-56396-047-8 (US: vol. 1), 1-56396-048-6 (US: vol. 2), 1-56396-049-4 (US: vol. 3), 0-7503-0310-7 (UK: set), 0-7503-0353-0 (UK: vol. 1), 0-7503-0354-9 (UK: vol. 2), 0-7503-0355-7 (UK: vol. 3). LCCN QC7

.T84 1995. URL http://www.loc.gov/catdir/enhancements/
fy0668/95041186-d.html.

Staley:1995:IMS

[Sta95b] R. Staley. Interpretations and misinterpretations of special and general relativity theory by contemporaries of Albert Einstein

— German — K. Hentschel. British Journal for the History of Science, 28(99):482–483, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Stannard:1995:OAU

[Sta95c] Russell Stannard. Onkel Albert und der Urknall: eine neue Geschichte um Einstein und seine Theorie. (German) [Uncle Albert and the Big Bang: a new story about Einstein and his theory], volume 80055: Fischer Schatzinsel of Fischer. Fischer-Taschenbuch-Verl., Frankfurt am Main, Germany, 1995. ISBN 3-596-80055-2. 164 pp. LCCN ????

Stachel:1998:EMY

[Sta98a] John Stachel, editor. Einstein’s miraculous year: five papers that changed the face of physics. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-05938-1. xv + 198 pp. LCCN QC7 .E52 1998. US$19.95. With the assistance of Trevor Lipscombe, Alice Calaprice, and Sam Elworthy, and with a foreword by Roger Penrose.

Stachel:1998:OEE

[Sta98b] John Stachel. The other Einstein: Einstein contra field theory. Science in Context, 6(1):275–290, March 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Staley:1998:HRP

[Sta98c] Richard Staley. On the histories of Relativity: The propagation and elaboration of Relativity Theory in participant histories in Germany, 1905–1911. Isis, 89(2):263–299, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237756.

Stange:1998:KAR

[Sta98d] Thomas Stange. Die kernphysikalischen Ambitionen des Reichspostministers Ohnesorge. (German) [The nuclear physics ambitions of Empire Postmaster Ohnesorge]. Berichte zur Wissenschaftsgeschichte, 21(2–3):159–174, ???? 1998. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Stachel:1999:EIS

[Sta99a] John Stachel. Einstein and Infeld, seen through their correspondence. Acta Physica Polonica B, 30(10):2879–2908, October 1999. CODEN APOBBB. ISSN 0587-4254 (print),

1509-5770 (electronic). URL http://www.actaphys.uj.edu.
pl/fulltext?series=Reg&vol=30&page=2879. The appendix

Scientific Writings of Leopold Infeld is an shortened version of the annotated bibliography in [?].

Staley:1999:BRD

[Sta99b] Richard Staley. Book review: Denis Brian, Einstein: A Life, New York, John Wiley, 1996, 509 pp. (ISBN 0-471-11459-6) $30.00 USA, $36.95 Can, £18.99 UK. Nuncius, 14(2):713–715, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00265.

Stannard:1999:NWM

[Sta99c] Russell Stannard. The New World of Mr. Tompkins: George Gamow’s Classic Mr. Tompkins in Paperback. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-63009-6 (hardcover), 0-521-63992-1 (paperback), 0-511-00800-7 (e-book), 0-511-58169-6 (e-book). xii + 258 pp. LCCN QC71.S775 1999. US$24.99. URL http://www.loc.

gov/catdir/description/cam029/98050379.html;   http:
//www.loc.gov/catdir/enhancements/fy0732/98050379-b.
html;   http://www.loc.gov/catdir/toc/cam021/98050379.
html.

Staley:2000:BRC

[Sta00] R. Staley. Book review: The collected papers of Albert Einstein, vol 6, the Berlin years: writings, 1914–1917. Annals of Science, 57(3):313–315, July 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505x (electronic).

Stachel:2002:EBZ

[Sta02] John J. Stachel. Einstein from “B” to “Z”, volume 9 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002. ISBN 0-8176-4143-2, 3-7643-4143-2 (Basel). xi + 556 pp. LCCN QC16.E5 S69 2002.

Stairs:2003:TGR

[Sta03a] Ingrid H. Stairs. Testing General Relativity with pulsar timing. Living Reviews in Relativity, 6:5–49, 2003. CODEN ????

ISSN 1433-8351. URL http://relativity.livingreviews.
org/Articles/lrr-2003-5/.

Stanley:2003:EHW

[Sta03b] Matthew Stanley. “An expedition to heal the wounds of war”: The 1919 eclipse and Eddington as Quaker adventurer. Isis, 94(1):57–89, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376099.

Stachel:2005:BRB

[Sta05a] John Stachel. Book review: Einstein’s clocks, Poincaré’s maps; Empires of time. Studies in History and Philosophy of Modern Physics, 36(1):202–210, March 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000796. See [Gal04a].

Stachel:2005:EMY

[Sta05b] John Stachel, editor. Einstein’s miraculous year: five papers that changed the face of physics. Princeton University Press, Princeton, NJ, USA, centenary edition, 2005. ISBN 0-691-12228-8 (paperback). lxxiv + 198 pp. LCCN QC7 .E52x 2005; QC7

.E52 2005. US$16.95. URL http://press.princeton.edu/
titles/6272.html.

Stadler:2005:MOM

[Sta05c] Ulrich Stadler. Moving objects, moved observers: On the treatment of the problem of Relativity in poetic texts and scientific prose. Science in Context, 18(4):607–627, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stanley:2005:ECB

[Sta05d] Matthew Stanley. The 1919 eclipse: a celebrity is born. Physics World, 18(1):25–26, January 2005. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/1/phwv18i1a26.pdf.

Stapp:2006:QID

[Sta06] Henry P. Stapp. Quantum interactive dualism, II: The Libet and Einstein–Podolsky–Rosen causal anomalies (Benjamin Libet, Albert Einstein, Boris Podolsky, Nathan Rosen). Erkenntnis, 65(1):117–142, ???? 2006. ISSN 0165-0106.

Stachel:2007:FTA

[Sta07a] John Stachel. The first two acts. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 81–111. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https://link.

springer.com/chapter/10.1007/978-1-4020-4000-9_3.

Stachel:2007:SNG

[Sta07b] John Stachel. The story of Newstein or: Is gravity just another pretty force? In Renn et al. [RSS^{+}07b], pages 1041–1078. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-
4020-4000-9_46.

Stanley:2007:PMR

[Sta07c] Matthew Stanley. Practical mystic: religion, science, and A. S. Eddington. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-77097-4 (cloth). x + 313 pp. LCCN BL240.3 .S725 2007. URL http://www.loc.

gov/catdir/enhancements/fy0809/2007005482-d.html;
http://www.loc.gov/catdir/enhancements/fy0809/2007005482-
t.html;   http://www.loc.gov/catdir/enhancements/fy0827/
2007005482-b.html.

Stanley:2007:STS

[Sta07d] Matthew Stanley. So simple a thing as a star: the Eddington– Jeans debate over astrophysical phenomenology. British Journal for the History of Science, 40(1):53–82, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4500682.

Staley:2008:EGO

[Sta08a] Richard Staley. Einstein’s generation: the origins of the Relativity revolution. University of Chicago Press, Chicago, IL, USA and London, UK, 2008. ISBN 0-226-77056-7 (cloth), 0-226-77057-5 (paperback). x + 494 pp. LCCN QC7 .S777 2008.

URL http://www.loc.gov/catdir/enhancements/fy0829/
2008014155-d.html;   http://www.loc.gov/catdir/toc/
ecip0815/2008014155.html;   http://www.press.uchicago.
edu/presssite/metadata.epl?mode=synopsis&bookkey=306867.

Stannard:2008:RVS

[Sta08b] Russell Stannard. Relativity: a very short introduction, volume 190 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-923622-4 (paperback). 114 pp. LCCN QC173.55 .S73 2008.

Stachel:2009:GC

[Sta09a] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Stapp:2009:EL

[Sta09b] Henry Stapp. Einstein locality. In Greenberger et al. [GHW09], pages 182–188. ISBN 3-540-70622-4, 3-540-70626-7 (e-book). LCCN C174.12 .C66 2009.

Stachel:2015:EH

[Sta15] John Stachel. Einstein and Hilbert. In Arabatzis et al. [ARS15], pages 33–39. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). LCCN Q125 .R45 2015. URL http://www.loc.

gov/catdir/enhancements/fy1516/2015936523-d.html;
http://www.loc.gov/catdir/enhancements/fy1516/2015936523-
t.html.

Stamenkovic:2016:NUE

[Sta16] Philippe Stamenkovic. On the “negative utility” of Ernst Cassirer’s philosophy of physics: an application to the EPR argument. Studies in History and Philosophy of Modern Physics, 55 (??):34–42, August 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219816300296.

Stanley:2019:EWB

[Sta19] Matthew Stanley. Einstein’s War: the Birth of Relativity Amid the Vicious Nationalism of World War I. Dutton, New York, NY, USA, 2019. ISBN 1-5247-4541-3 (hardcover), 1-5247-4543-X (e-book). vi + 390 pp. LCCN QC173.52 .S73 2019.

Stanley:2020:BRJ

[Sta20] Matthew Stanley. Book review: Jan Guichelaar. Willem de Sitter: Einstein’s Friend and Opponent. Isis, 111(1):200–201, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stanley:2021:BRE

[Sta21] Matthew Stanley. Book review: Einstein in Bohemia, by M. Gordin, Princeton, NJ, Princeton University Press, 2020, xi + 343 pp., $29.95. Hardback. ISBN 978-0-691-17737-3. Annals of Science, 78(1):131–132, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stern:1949:TE

[Ste49] Otto Stern. On the term k ln n! in the entropy. Reviews of Modern Physics, 21(3):534–535, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.534;
http://rmp.aps.org/abstract/RMP/v21/i3/p534_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Stevenson:1953:NPW

[Ste53] Lloyd G. Stevenson. Nobel Prize winners in medicine and physiology, 1901–1950, volume 29 of Life of science library. H. Schuman, New York, NY, USA, 1953. ix + 291 pp. LCCN ????

Steahle:1960:ANN

[Stê60] Nils K. Stêahle. Alfred Nobel and the Nobel Prizes. Nobel Foundation and Swedish Institute, Stockholm, Sweden, 1960. 86 pp. LCCN ????

Stern:1982:EG

[Ste82] Fritz Stern. Einstein’s Germany. In Holton and Elkana [HE82], pages 319–346. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Stern:1986:EG

[Ste86] Fritz Stern. Einstein and Germany. Physics Today, 39(2):40– 49, February 1986. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v39/i2/p40_s1.

Stephan:1989:SBF

[Ste89] Inge Stephan. Das Schicksal der begabten Frau: im Schatten berühmter Männer. (German) [The Fate of the Gifted Woman:

In the Shadows of Famous Men]. Kreuz Verlag, Stuttgart, Germany, third edition, 1989. ISBN 3-7831-0987-6. 217 pp. LCCN CT3203 .S740 1989; HQ759 .S688 1989.

Stewart:1990:BRC

[Ste90] I. Stewart. Book review: The Collected Papers of Albert Einstein, Paperback Supplement to Volume 2 — A. Einstein, A. Beck, Translator. TLS — The Times Literary Supplement, ?? (4549):615, June 8, 1990. ISSN 0307-661X.

Steinsmith:1995:EL

[Ste95a] William Steinsmith. Einstein and Lorentz. Nature, 376(6536): 111, July 13, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v376/n6536/pdf/376111a0.pdf.

Steneck:1995:BRB

[Ste95b] Nicholas Steneck. Book review: Science on Trial: The Whistle-Blower, the Accused, and the Nobel Laureate by Judy Sarasohn. Isis, 86(1):143–144, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/236479.

Sterrett:1998:SLL

[Ste98] Susan G. Sterrett. Sounds like light: Einstein’s Special Theory of Relativity and Mach’s work in acoustics and aerodynamics. Studies in History and Philosophy of Modern Physics, 29(1):1– 35, March 11, 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219897000270.

Stern:1999:EGW

[Ste99] Fritz Richard Stern. Einstein’s German world. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-05939-x. ix + 335 pp. LCCN DD239 .S74 1999. URL http://www.gbv.de/

dms/bowker/toc/9780691059396.pdf;   http://www.loc.
gov/catdir/description/prin021/99020128.htm;   http:
//www.loc.gov/catdir/description/prin021/99020128.html;
 http://www.loc.gov/catdir/toc/prin032/99020128.htm;
 http://www.loc.gov/catdir/toc/prin032/99020128.html.

Stern:2001:EGW

[Ste01] Fritz Richard Stern. Einstein’s German world. Princeton University Press, Princeton, NJ, USA, 2001. ISBN 0-691-07458-5 (paperback). ix + 335 pp. LCCN DD239 .S74 2001.

URL http://www.loc.gov/catdir/description/prin021/
99020128.htm;   http://www.loc.gov/catdir/toc/prin032/
99020128.htm.

Stephani:2004:RIS

[Ste04] Hans Stephani. Relativity: an introduction to Special and General Relativity. Cambridge University Press, Cambridge, UK, third edition, 2004. ISBN 0-521-81185-6 (hardcover), 0-521-01069-1 (paperback). xx + 396 pp. LCCN QC17 .S7413

2004.  URL http://www.loc.gov/catdir/description/
cam032/2003053217.html;   http://www.loc.gov/catdir/
toc/cam032/2003053217.html.

Steiner:2005:AEG

[Ste05a] F. (Frank) Steiner, editor. Albert Einstein: Genie, Visionär und Legende. (German) [Albert Einstein: Genius, Visionary, and Legend]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-21060-1. xii + 222 pp. LCCN QC16.E5 A659 2005.

Steiner:2005:AEU

[Ste05b] Frank Steiner. Albert Einstein: Von Ulm nach Princeton. (German) [Albert Einstein: From Ulm to Princeton]. In Steiner [Ste05a], pages 1–40. ISBN 3-540-21060-1. LCCN QC16.E5

A659 2005. URL http://link.springer.com/chapter/10.
1007/3-540-30595-5_1/.

Steiner:2005:EKR

[Ste05c] Frank Steiner. Einsteins kosmische Religiosität. (German) [Einstein’s cosmic religiosity]. In Steiner [Ste05a], pages 191–217. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http:/

/link.springer.com/chapter/10.1007/3-540-30595-5_9/.

Stephens:2005:BRE

[Ste05d] Carlene E. Stephens. Book review: Einstein’s Clocks, Poincaré’s Maps: Empires of Time. Technology and Culture, 46(1):241– 243, 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

Stewart:2007:CDM

[Ste07] Philip J. Stewart. A century on from Dmitrii Mendeleev: tables and spirals, noble gases and Nobel prizes. Foundations of Chemistry, 9(3):235–245, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-007-9038-x.

Stewart:2016:CCH

[Ste16] Ian Stewart. Calculating the Cosmos: How Mathematics Unveils the Universe. Basic Books, New York, NY, USA, 2016. ISBN 1-5416-1725-8 (paperback). 346 + 8 pp. LCCN QB981 .S82 2018.

Stenhouse:2022:BRA

[Ste22] Brigitte Stenhouse. Book reviews: Allan Esterson and David C. Cassidy, Einstein’s Wife: The Real Story of Mileva Einstein-Marić Cambridge, MA: MIT Press, 2019. Pp. 336. ISBN: 978-0-2625-3897-8. $19.95 (paperback). ISBN: 978-0-2620-3961-1. $29.95 (hardback). British Journal for the History of Science, 55:531–533, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/british-journal-for-the-
history-of-science/article/allan-esterson-and-david-
c-cassidy-einsteins-wife-the-real-story-of-mileva-einsteinmaric-
cambridge-ma-mit-press-2019-pp-336-isbn-9780262538978-
1995-paperback-isbn-9780262039611-2995-hardback/EC290FBBFB3E58DC60E32E6AE4B6CD49.

Stimson:1947:DUB

[Sti47] H. L. Stimson. The decision to use the bomb. Harper’s Magazine, 194(2):97–107, February 1947. CODEN HAMAA3.

ISSN 1045-7143. URL http://www.harpers.org/archive/

1947/02/0032863. Reprinted in [Sti76]. See also [Lef95] for a contrary view based on a half-century of historical scholarship, and access to previously-secret US documents.

Stimson:1976:DUB

[Sti76] H. L. Stimson. The decision to use the bomb. In Baker [Bak76b], pages 14–28. ISBN 0-03-089873-0. LCCN ???? Reprinted from [Sti47]. See also [Lef95] for a contrary view based on a halfcentury of historical scholarship, and access to previously-secret US documents.

Stix:1998:SBN

[Sti98] Gary Stix. Special briefing: The 1997 Nobel Prizes in science. Scientific American, 278(1):14–18, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-14.pdf.

Stiefelhagen:2005:TAE

[Sti05] Peter Stiefelhagen. Zum 50. Todestag von Albert Einstein (1879–1955). “Der Tod ist eine alte Schuld”. (German) [The 50th death anniversary of Albert Einstein (1879–1955). “Death is an old debt”]. MMW Fortschritte der Medizin, 147(10):56, March 10, 2005. ISSN 1438-3276.

Stiehm:2013:CPW

[Sti13a] Judith Hicks Stiehm. Champions for Peace: Women Winners of the Nobel Peace Prize. Rowman and Littlefield, Lanham, MD, USA, second edition, 2013. ISBN 1-4422-2151-8 (hardcover), 1-4422-2152-6 (e-book). LCCN JZ5540 .S74 2014.

Stix:2013:NRG

[Sti13b] Gary Stix. Neuroscience: Roots of genius. Scientific American, 308(2):23, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:2013:RGN

[Sti13c] Gary Stix. Roots of genius: A new image collection of Albert Einstein’s brain may provide insight into the physicist’s profound ability to visualize space and time. Scientific American, 308(2):23, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stoffel:1996:MGL

[Sto96] Jean-François Stoffel, editor. Mgr. Georges Lemaître, savant et croyant: actes du colloque commémoratif du centième anniversaire de sa naissance (Louvain-la-Neuve, le 4 novembre 1994): suivi de La physique d’Einstein: texte inédit. (French) [Monsignor Georges Lemaître, scholar and believer: proceedings of the commemorative conference on the hundredth anniversary of his birth (Louvain-la-Neuve, 4 November 1994), followed by the physics of Einstein: unpublished text]. Centre Interfacultaire d’Étude en Histoire des Sciences, Louvain-la-Neuve, Belgium, 1996. ISBN 2-930175-02-8 (paperback). LCCN ????

Stoltzenberg:1998:FHC

[Sto98] Dietrich Stoltzenberg. Fritz Haber: Chemiker, Nobelpreisträger, Deutscher, Jude: eine Biographie. (German) [Fritz Haber: Chemist, Nobel Laureate, German, Jew: a biography]. VCH, Weinheim, Germany, second edition, 1998. ISBN 3-527-29573-9 (paperback). xiv + 669 pp. LCCN ????

Stormer:1999:NLF

[Sto99] Horst L. Stormer. Nobel Lecture: The fractional quantum Hall effect. Reviews of Modern Physics, 71(4):875–889, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.71.875;   http://rmp.aps.org/
abstract/RMP/v71/i4/p875_1.

Stoof:2002:BEC

[Sto02] Henk T. C. Stoof. Bose–Einstein condensation: Breaking up a superfluid. Nature, 415(6867):25–26, January 3, 2002. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v415/n6867/
full/415025a.html.

Stone:2005:EUI

[Sto05] A. Douglas Stone. Einstein’s unknown insight and the problem of quantizing chaos. Physics Today, 58(8):37–43, August 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/
1/phtoad/v58/i8/p37_s1. See [Ein17h, Kel58].

Stoddart:2012:LEF

[Sto12] Kristan Stoddart. Losing an empire and finding a role: Britain, the USA, NATO and nuclear weapons, 1964–70. Nuclear weapons and international security since 1945. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2012. ISBN 0-230-30088-X (hardcover). xv + 327 pp. LCCN U264.5.G7 S76 2012.

Stone:2013:EQQ

[Sto13] A. Douglas Stone. Einstein and the quantum: the quest of the valiant Swabian. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-13968-7 (hardcover), 1-4008-4834-2 (e-book), 1-299-87423-1 (e-book). x + 332 pp. LCCN

QC16.E5 S76 2013. URL http://www.npr.org/2013/11/01/
242356997/einsteins-real-breakthrough-quantum-theory.

Stone:2017:EGM

[Sto17] A. Douglas Stone. Book review: Einstein’s Greatest Mistake: A Biography, David Bodanis, Houghton Mifflin Harcourt, 2016, 304 p, $27.00, ISBN 978-0-544-80856-0. Physics Today, 70(5): 58–59, May 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Straus:1949:SRE

[Str49] E. G. Straus. Some results in Einstein’s Unified Field Theory. Reviews of Modern Physics, 21(3):414–420, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.414;   http://rmp.aps.org/

abstract/RMP/v21/i3/p414_1. Special issue in honor of Albert Einstein on his 70th birthday.

Strauss:1968:ETC

[Str68a] M. Strauss. Einstein’s theories and the critics of Newton. Synthese, 18(2–3):251–284, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413778.

Strickland:1968:SPA

[Str68b] Donald A. Strickland. Scientists in politics; the Atomic Scientists Movement, 1945–46,: Atomic Scientists Movement, 1945– 46. Purdue University Studies, Lafayette, IN, USA, 1968. xi + 149 pp. LCCN QC773 .S8.

Straus:1982:RAE

[Str82] Ernst G. Straus. Reminiscences [of Albert Einstein]. In Holton and Elkana [HE82], pages 417–425. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Strandh:1983:ANM

[Str83] Sigvard Strandh. Alfred Nobel: mannen, verket, samtiden. (Swedish) [Alfred Nobel: the man, his work, and his times]. Natur och kultur, Stockholm, Sweden, 1983. ISBN 91-27-01283-2. 339 + 8 pp. LCCN TP268.5.N7 S77 1983.

Straumann:1990:BRC

[Str90] N. Straumann. Book review: The Collected Papers of Albert Einstein, Vol 2, the Swiss Years — Writings, 1900–1909 — J. Stachel. Nature, 344(6269):899, April 26, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Straumann:1994:AEW

[Str94] Norbert Straumann. Albert Einstein: On the way to the theory of gravitation. Vierteljahrsschrift der Naturforschenden Gesellschaft in Zuerich, 139(3):103–112, ???? 1994. ISSN 0042-5672.

Strathern:1997:EOR

[Str97a] Paul Strathern. Einstein og relativiteten. (Danish) [Einstein and Relativity]. Polyteknisk, Lyngby, Denmark, 1997. ISBN 87-502-0798-9. 82 pp.

Strobach:1997:EZL

[Str97b] Niko Strobach. Einsteins Zug und logische Gesetze. (German) [Einstein’s train and logical laws]. Philosophia Naturalis, 34(1): 21–31, 1997. CODEN ???? ISSN 0031-8027.

Strathern:1999:ERB

[Str99] Paul Strathern. Einstein and Relativity: The Big Idea. The big idea. Anchor Books, New York, NY, USA, 1999. ISBN 0-385-49244-8. 102 pp. LCCN QC16.E5 S77 1999.

Stranges:2004:BRA

[Str04] Anthony N. Stranges. Book review: Ahmed Zewail: Voyage through Time: Walks of Life to the Nobel Prize. Isis, 95(1): 164–165, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/423589.

Stranges:2005:FHC

[Str05a] Anthony N. Stranges. Fritz Haber: Chemist, Nobel Laureate, German, Jew (review). Technology and Culture, 46(2):448– 449, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184345.

Straumann:2005:WZN

[Str05b] Norbert Straumann. Weiße Zwerge, Neutronensterne und Schwarze Löcher. (German) [White dwarfs, neutron stars, and black holes]. In Steiner [Ste05a], pages 93–122. ISBN 3-540-21060-1. LCCN QC16.E5 A659 2005. URL http://link.

springer.com/chapter/10.1007/3-540-30595-5_4/.

Straumann:2011:FSC

[Str11a] Norbert Straumann. On the first Solvay Congress in 1911. European Physical Journal H, 36(3):379–399, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20043-9.

Strickland:2011:WSC

[Str11b] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Straumann:2013:GR

[Str13a] Norbert Straumann. General Relativity. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 94-007-5409-4 (hardcover), 94-007-5410-8 (e-book). xix + 735 pp. LCCN QC173.6 .S912 2013.

Stryker:2013:MPM

[Str13b] James Stryker. Manhattan Project: The Making of the Bomb. CreateSpace Independent Publishing Platform, ????, 2013. ISBN 1-4943-6101-9. 270 (est.) pp. LCCN ????

Strogatz:2017:EFP

[Str17] Steven Strogatz. Einstein’s first proof. In Pitici [Pit17], page ?? ISBN 0-691-17529-2 (paperback). LCCN QA1 .B337; QA93 .B476 2016.

Strickland:2019:NLG

[Str19] Donna Strickland. Nobel Lecture: Generating high-intensity ultrashort optical pulses. Reviews of Modern Physics, 91(3): 030502–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
030502.

Study:1914:RWL

[Stu14] Eduard Study. Die realistische Weltansicht und die Lehre vom Räume. Geometrie, Anschauung und Erfahrung. (German) [The realistic world view and the science of space. Geometry, opinion and experience], volume 54 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1914. ix + 145 pp. LCCN ????

Study:1923:RWL

[Stu23] Eduard Study. Die realistische Weltansicht und die Lehre vom Räume. Geometrie, Anschauung und Erfahrung. (German) [The realistic world view and the science of space. Geometry, opinion and experience]. Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1923. LCCN ????

Stuewer:1975:CET

[Stu75] Roger H. Stuewer. The Compton effect: turning point in physics. Science History Publications, New York, NY, USA, 1975. ISBN 0-88202-012-9. xii + 367 pp. LCCN QC794.6.S3 S88 1975.

Stuewer:1986:RSM

[Stu86] Roger H. Stuewer. Rutherford’s satellite model of the nucleus. Historical Studies in the Physical and Biological Sciences, 16 (2):321–352, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.jstor.org/

stable/27757568.

Stuhlhofer:1987:AEN

[Stu87] Dr. Franz Stuhlhofer. Anregungen und Empfelhlungen: Nobelpreise und Nationalismus. Berichte zur Wissenschaftsgeschichte, 10(2):114–116, ???? 1987. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Stuewer:1993:MEN

[Stu93] Roger H. Stuewer. Mass-energy and the neutron in the Early Thirties. Science in Context, 6(1):195–238, Spring 1993. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2(1):76–129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://alsos.wlu.edu/

information.aspx?id=1736. See errata [Ano94b].

Stubhaug:2010:GML

[Stu10] Arild Stubhaug. Gösta Mittag-Leffler: a man of conviction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11671-X (hardcover), 3-642-11672-8 (e-book). x + 733 pp. LCCN QA29.M5233 S8913 2010. Translated by Tiina Nunnally.

Stuewer:2014:ECL

[Stu14] Roger H. Stuewer. The experimental challenge of light quanta. In Janssen and Lehner [JL14b], chapter 5, pages 143–166. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Stykes:1994:OIN

[Sty94] Mervin Stykes. And the other 1994 Ig Nobel Prize winners are. Scientific American, 271(6):26, December 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-26a.pdf.

Styer:2000:SWQ

[Sty00] Daniel F. Styer. The strange world of quantum mechanics. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66104-8 (hardcover), 0-521-66780-1 (paperback). xiv + 154 pp. LCCN QC174.12 .S879 2000. URL http://www.loc.

gov/catdir/description/cam029/99013559.html;   http:
//www.loc.gov/catdir/samples/cam034/99013559.html;
http://www.loc.gov/catdir/toc/cam021/99013559.html.

Segers:2006:ELM

[SU06] Danny Segers and Jos Uyttenhove. Einstein’s ‘little machine’ as an example of charging by induction. American Journal of Physics, 74(8):670–676, August 2006. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v74/i8/p670_s1. See also

[Maa07, SWU09].

Subramanian:2003:GOV

[Sub03] V. K. Subramanian. The great ones, volume 1. Abhinav Publications, New Delhi, India, 2003. ISBN 81-7017-421-X. 233 pp. LCCN CT105 .S83 2003.

Sudbery:2011:QEB

[Sud11] Tony Sudbery. Quantum: Einstein, Bohr and the Great Debate About the Nature of Reality, by Manjit Kumar, Scope: general interest. Level: general readership. Contemporary Physics, 52 (3):251–254, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sugimoto:1989:AEP

[Sug89] Kenji Sugimoto. Albert Einstein: a photographic biography. Schocken Books, New York, NY, USA, 1989. ISBN 0-8052-4047-0. 202 pp. LCCN QC16.E5 S84131 1989; QC16.E5 S8413 1989.

Sullivan:1964:SRE

[Sul64] Walter Sullivan. Scientist revises Einstein’s theory: Hoyle’s proposal realigns mass–gravity relation. New York Times, ??(??): 84, June 21, 1964. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/115858143/.

Sullivan:2013:ERK

[Sul13] Walter Sullivan. Einstein: Relativity in the kitchen. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Susskind:2006:CLS

[Sus06] Leonard Susskind. The Cosmic Landscape: String Theory and the Illusion of Intelligent Design. Little, Brown and Co., Boston, MA, USA, 2006. ISBN 0-316-15579-9 (hardcover). xii + 403 pp. LCCN QB981 .S886 2005; QB981 .S886 2006. URL http:

//www.loc.gov/catdir/toc/ecip0515/2005018796.html.

Susskind:2008:BHW

[Sus08] Leonard Susskind. The Black Hole War: My Battle with Stephen Hawking to Make the World Safe for Quantum Mechanics. Little, Brown and Co., Boston, MA, USA, 2008. ISBN 0-316-01640-3. viii + 470 pp. LCCN

QC174.12 .S896 2008. URL http://www.loc.gov/catdir/
enhancements/fy0804/2007048355-b.html;   http://www.
loc.gov/catdir/enhancements/fy0804/2007048355-d.html.

Suter:1997:NHE

[Sut97] Keith Suter. Nobel hears East Timor. Bulletin of the Atomic Scientists, 53(1):47–48, January/February 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1987:RFE

[SW87] Gerrit J. Smith and Robert Weingard. A relativistic formulation of the Einstein–Podolsky–Rosen paradox. Foundations of Physics, 17(2):149–171, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00733205.

Salmon:1994:LLS

[SW94] Wesley C. Salmon and Gereon Wolters, editors. Logic, language, and the structure of scientific theories: proceedings of the Carnap–Reichenbach centennial, University of Konstanz, 21–24 May 1991, volume 2 of Pittsburgh-Konstanz series in the philosophy and history of science. University of Pittsburgh Press, Pittsburgh, PA, USA, 1994. ISBN 3-87940-477-1, 0-8229-3740-9 (USA). LCCN Q174 L64 1994; Gd.CAR.B5; M95.E09243.

Smith:1995:ROL

[SW95] Alice Kimball Smith and Charles Weiner. Robert Oppenheimer: Letters and Recollections. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1995. ISBN 0-8047-2620-5. xxii + 376 pp. LCCN QC16.O62 A4 1995. URL http://

www.loc.gov/catdir/description/cam027/95067508.html.

Stadler:2012:MSR

[SW12] Friedrich Stadler and Hans Jürgen Wendel, editors. Moritz Schlick Rostock, Kiel, Wien: Aufsätze, Beiträge, Rezensionen 1919-1925. (German) [Moritz Schlick: Rostock, Kiel, Vienna: essays, contributions, reviews, 1919–1925], volume 5. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-211-32769-X (print), 3-211-69443-9 (electronic). xii + 839 pp. LCCN ???? URL http://link.

springer.com/bookseries/7287.

Swann:1930:SRP

[Swa30] W. F. G. Swann. Statement in regard to Professor Einstein’s publications. Science, 71(1841):390–391, April 11, 1930. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/71/
1841/390.full.pdf.

Swartley:1992:NMA

[Swa92] Ron Swartley. New Mexico’s atomic tour: a guided trip through the birth and flowering of the Atomic Age in New Mexico. Frontier Image Press, Albuquerque, NM, USA, 1992. ISBN 0-9634309-0-4. 52 pp. LCCN QC773.3.U5 S9 1992.

Swartley:1995:NMA

[Swa95] Ron Swartley. New Mexico’s atomic tour: a guided trip through the birth and flowering of the Atomic Age in New Mexico. Frontier Image Press, Las Cruces, NM, USA, second edition, 1995. ISBN 0-9634309-5-5. 63 pp. LCCN QC773.3.U5 S9 1995.

Swanson:2002:BRR

[Swa02] Eleanor Swanson. Book reviews: Radium Girls, and: The Laboratory at Night, and: Marie Curie and Albert Einstein Hike in Engadine, and: Dr. Tobé Attends Mme. Curie on Her Deathbed. The Missouri Review, 25(1):27–33, 2002. CODEN ???? ISSN 0191-1961 (print), 1548-9930 (electronic). URL https://muse.

jhu.edu/article/409684/summary.

Swenson:1979:GRE

[Swe79] Loyd S. Swenson, Jr. Genesis of relativity: Einstein in context, volume 5 of Studies in the history of science. B. Franklin, New York, NY, USA, 1979. ISBN 0-89102-101-9. xv + 266 + 4 pp. LCCN QC16.E5 S95 1979. US$18.95.

Swenson:2013:EAH

[Swe13] Loyd S. Swenson, Jr. The Ethereal Aether: A History of the Michelson–Morley–Miller Aether-drift Experiments, 1880–1930.

University of Texas Press, Austin, TX, USA, 2013. ISBN 0-292-75836-7. ???? pp. LCCN ????

Swing:1945:RSB

[Swi45] Raymond Swing. Raymond Swing’s broadcast[: History of atomic bomb]. Script for radio broadcast on station WMAL on ABC network., September 7, 1945. URL http://library.

ucsd.edu/dc/object/bb0718253r.

Swinbanks:1991:EBS

[Swi91] David Swinbanks. Einstein brought down to size. Nature, 349 (6310):556, February 14, 1991. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v349/n6310/pdf/349556c0.pdf.

Segers:2009:CTR

[SWU09] Danny Segers, Kristel Wautier, and Jos Uyttenhove. Construction and testing of a replica of Einstein’s “little machine”. American Journal of Physics, 77(1):10–12, January 2009. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/

ajpias/v77/i1/p10_s1. See also [SU06, Maa07].

Symborska:1998:P

[Sym98] Wisława Symborska. PI. In Poems, New and Collected, 1957–1997, pages 174–175. Harcourt Brace, New York, NY, USA, 1998. ISBN 0-15-100353-X. LCCN PG7178.Z9 A222

1998. URL http://www.nobelprize.org/nobel_prizes/
literature/laureates/1996/;   http://www.nobelprize.
org/nobel_prizes/literature/laureates/1996/szymborska.

html. Translated from the Polish by Stanisław Barańczak and Clare Cavanagh.

Synge:1965:ES

[Syn65] J. L. Synge. Einstein spaces. Nature, 207(4998):673, August 14, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v207/n4998/pdf/207673a0.pdf.

Selleri:1988:LDD

[SZ88] F. Selleri and A. Zeilinger. Local deterministic description of Einstein–Podolsky–Rosen experiments. Foundations of Physics, 18(12):1141–1158, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889429.

Szasz:1984:DSR

[Sza84] Ferenc Morton Szasz. The day the sun rose twice: the story of the Trinity Site nuclear explosion, July 16, 1945. University of New Mexico Press, Albuquerque, NM, USA, 1984. ISBN 0-8263-0768-X (paperback), 0-8263-0767-1. xi + 233 pp. LCCN QC773.A1 S93 1995; QC773.A1 S93 1984.

Szasz:1992:BSMa

[Sza92a] Ferenc Morton Szasz. British scientists and the Manhattan Project: the Los Alamos years. Macmillan, Basingstoke, UK, 1992. ISBN 1-349-12731-0, 0-333-56597-5. xx + 167 pp. LCCN

???? URL http://link.springer.com/10.1007/978-1-349-
12731-3.

Szasz:1992:BSMb

[Sza92b] Ferenc Morton Szasz. British scientists and the Manhattan Project: the Los Alamos years. St. Martin’s Press, New York, NY, USA, 1992. ISBN 0-312-06167-6. xx + 167 pp. LCCN

QC773.3.U5 S97 1991. URL http://www.loc.gov/catdir/
enhancements/fy0809/91019904-d.html.

Szilard:1945:ABP

[Szi45] Leo Szilard. Atomic bombs and the postwar position of the United States. Unpublished memo to US President Franklin D. Roosevelt, who died before the memo could be delivered., March 1945. Reproduced in [WS78, pages 196–204], and with an introduction to the President by Albert Einstein reproduced in [LS92, pages 261–261].

Szilard:1971:CLC

[Szi71] Gertrud Weiss Szilard. Communications: Lunar craters. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1992:VDO

[Szi92] Leo Szilard. The voice of the dolphins and other stories. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, expanded edition, 1992. ISBN 0-8047-1753-2, 0-8047-1754-0. vi + 182 pp. LCCN PS3569.Z5 V65 1992. US$20.00, US$8.95.

Szilard:2005:DHH

[Szi05] Szilárd Leó. A delfinek hangja. (Hungarian) [The voice of dolphins]. Kairosz, Budapest, Hungary, 2005. ISBN 963-9568-95-3. 124 pp. LCCN PS3569.Z5 V6515 2005.

tHooft:1997:SUB

[’t 97] Gerard ’t Hooft. In search of the ultimate building blocks. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-55083-1 (hardcover), 0-521-57883-3 (paperback). xiii + 191 pp. LCCN QC794.6.S75 H66 1997. URL http://www.loc.

gov/catdir/description/cam027/96031468.html;   http:
//www.loc.gov/catdir/toc/cam021/96031468.html.

tHooft:2000:NLC

[’t 00] Gerard ’t Hooft. Nobel Lecture: a confrontation with infinity. Reviews of Modern Physics, 72(2):333–339, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.333;   http://rmp.aps.org/
abstract/RMP/v72/i2/p333_1. See erratum [’t 02].

tHooft:2002:ENL

[’t 02] Gerard ’t Hooft. Erratum: Nobel Lecture: A confrontation with infinity [Rev. Mod. Phys. 72, 333 (2000)]. Reviews of Modern Physics, 74(4):1343, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.74.1343;
 http://rmp.aps.org/abstract/RMP/v74/i4/p1343_1. See

[’t 00].

Trbuhovic-Duric:1969:USA

[T– D69] Desanka Trbuhović-– Durić. U senci Alberta Ajnštajna. (Serbian) [In the shadow of Albert Einstein]. Bagdala Press, Kruševac, Serbia, 1969. 271 pp. LCCN ????

Trbuhovic-Duric:1982:ISA

[T– D82a] Desanka Trbuhović-– Durić. Im Schatten Albert Einsteins: das tragische Leben der Mileva Einstein-Marić. (German) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Verlag Paul Haupt, Bern, Switzerland, 1982. ISBN ???? ???? pp. LCCN QC16.E5 D815 1982.

Trbuhovic-Duric:1982:MME

[T– D82b] Desanka Trbuhović-– Durić. Mileva Marić-Einstein 1875–1948. Verlag Paul Haupt, Bern, Switzerland, 1982. ISBN ???? ???? pp. LCCN ???? German translation of [T– D69]. With an epilogue (Nachwort) by Werner G. Zimmermann.

Trbuhovic-Duric:1983:ISA

[T– D83] Desanka Trbuhović-– Durić. Im Schatten Albert Einsteins: das tragische Leben der Mileva Einstein-Marić. (German) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Verlag Paul Haupt, Bern, Switzerland, second edition, 1983. ISBN 3-258-03288-2. 189 pp. LCCN QC16.E5 D815 1983.

Trbuhovic-Duric:1985:ISA

[T– D85] Desanka Trbuhović-– Durić. Im Schatten Albert Einsteins: das tragische Leben der Mileva Einstein-Marić. (German) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Verlag Paul Haupt, Bern, Switzerland, third edition, 1985. ISBN 3-258-03499-0. 192 pp. LCCN QC16.E5 D815 1988.

Trbuhovic-Duric:1988:ISA

[T– D88] Desanka Trbuhović-– Durić. Im Schatten Albert Einsteins: das tragische Leben der Mileva Einstein-Marić. (German) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Verlag Paul Haupt, Bern, Switzerland, fourth edition, 1988. ISBN 3-258-03973-9. 213 pp. LCCN QC16.E5 D815 1988.

Trbuhovic-Duric:1992:SAE

[T– D92] Desanka Trbuhović-– Durić. A la sombra de Albert Einstein: la trágica vida de Mileva Einstein Marić. (Spanish) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. La Tempestad, Barcelona, Spain, 1992. ISBN 84-7948-014-9. 198 + 16 pp. LCCN ???? Spanish translation by José Antonio Alemany.

Trbuhovic-Duric:1993:ISA

[T– D93] Desanka Trbuhović-– Durić. Im Schatten Albert Einsteins: das tragische Leben der Mileva Einstein-Marić. (German) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Verlag Paul Haupt, Bern, Switzerland, fifth edition, 1993. ISBN 3-258-04700-6. 213 pp. LCCN QC16.E5 D815

1992. URL http://www.zentralblatt-math.org/zmath/en/
search/?an=0771.01022.

Trbuhovic-Duric:1995:USA

[T– D95] Desanka Trbuhović-– Durić. U senci Alberta Ajnštajna. (Serbian) [In the shadow of Albert Einstein]. Klub NT, Belgrade, Serbia, 1995. ISBN 86-82167-32-8. 181 pp. LCCN QC16.E5 D8 1995.

Trbuhovic-Duric:2007:VSA

[T– D07] Desanka Trbuhović-– Durić. Ve stínu Alberta Einsteina tragický život Milevy Einsteinové-Marićové. (Czech) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Edice Žena a věda. Academia, Prague, Czechoslovakia, 2007. ISBN 80-200-1464-0. 173 pp. LCCN ????

Trbuhovic-Duric:2012:OAE

[T– D12] Desanka Trbuhović-– Durić. En la ombro de Alberto Ejnštejno. (Esperanto) [In the shadow of Albert Einstein: the tragic life of Mileva Einstein-Marić]. Interpress, Beograd, Serbia, 2012. ISBN 86-7561-128-5. 188 pp. LCCN ????

Tanner:1997:PLN

[TA97] P. G. Tanner and T. E. Allibone, F.Eng., F.R.S. The patent literature of Nobel laureate Dennis Gabor (1900–1979). Notes and Records of the Royal Society of London, 51(1):105–120, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
532039.

Tong:2017:ETN

[TA17] Sichao Tong and Per Ahlgren. Evolution of three Nobel Prize themes and a Nobel snub theme in chemistry: a bibliometric study with focus on international collaboration. Scientometrics, 112(1):75–90, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11192-017-2377-z.pdf.

Takaki:1995:HWA

[Tak95] Ronald T. Takaki. Hiroshima: why America dropped the atomic bomb. Little, Brown and Co., Boston, MA, USA, 1995. ISBN 0-316-83122-0. 193 + 8 pp. LCCN D769.2 T35 1995; D769.2 .T35 1995.

Talmey:1932:PRE

[Tal32a] Max Talmey. Personal recollections of Einstein’s boyhood and youth. Scripta Mathematica, 1(??):68–71, September 1932.

ISSN 0036-9713. URL https://en.wikipedia.org/wiki/
Max_Talmey;   https://en.wikipedia.org/wiki/Scripta_
Mathematica.

Talmey:1932:RTS

[Tal32b] Max Talmey. The Relativity Theory simplified and the formative period of its inventor. Falcon Press, New York, NY, USA, 1932. xi + 186 pp. With an introduction by George B. Pegram.

Tal:1982:JUS

[Tal82] Uriel Tal. Jewish and universal social ethics in the life and thought of Albert Einstein. In Holton and Elkana [HE82], pages 297–318. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Tamborrino:2013:AFF

[Tam14] E. Tamborrino. Aleksander Friedmann e la ^{≪}rivoluzione copernicana^{≫}in cosmologia: l’impatto sulla cosmologia relativistica einsteiniana. (Italian) [Aleksander Friedmann and the “Copernican revolution” in cosmology: the impact on the Einsteinian relativistic cosmology]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 49(1–2):??, 2013–2014. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Tannebaum:1966:AE

[Tan66] S. Tannebaum. Albert Einstein. 2. Journal of the American Optometric Association, 37(10):971, October 1966. ISSN 0003-0244.

Tanaka:1992:MEK

[Tan92] Setsuko Tanaka. Mach, Einstein, and Kuwaki. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92], chapter 16, pages 297–332. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN

C5.53; QC6.9. URL https://link.springer.com/chapter/
10.1007/978-94-011-2771-4_16.

Tansey:2011:EEN

[Tan11] E. M. Tansey. The early education of a Nobel laureate: Henry Dale’s schooldays. Notes and Records of the Royal Society of London, 65(4):379–391, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://

www.jstor.org/stable/23056305.

Taton:1964:AOBm

[Tat64] René Taton. Analyses d’ouvrages: Les Prix Nobel de physique et de chimie par Pierre Auger. Revue d’Histoire des Sciences et de Leurs Applications, 17(2):180–181, avril–juin 1964. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23904688.

Taub:1949:SMS

[Tau49] A. H. Taub. A special method for solving the Dirac equations. Reviews of Modern Physics, 21(3):388–392, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.388;   http://rmp.aps.org/

abstract/RMP/v21/i3/p388_1. Special issue in honor of Albert Einstein on his 70th birthday.

Taub:1963:JNCb

[Tau63] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon, New York, NY, USA, 1963. x + 538 pp. LCCN ????

Tauber:1979:AET

[Tau79] Gerald E. Tauber, editor. Albert Einstein’s Theory of General Relativity: 60 Years of its Influence on Man and the Universe. Crown Publishers, New York, NY, USA, 1979. ISBN 0-517-53661-7. 352 pp. LCCN QC173.6 .A4 1979.

Taubes:1986:NDP

[Tau86] Gary Taubes. Nobel dreams: power, deceit, and the ultimate experiment. Random House, New York, NY, USA, 1986. ISBN 0-394-54503-6. xxiv + 261 pp. LCCN QC793.4 .T38 1986. US$19.95.

Taubes:1988:NDP

[Tau88] Gary Taubes. Nobel dreams: power, deceit, and the ultimate experiment. Tempus Books of Microsoft Press, Redmond, WA, USA, 1988. ISBN 1-55615-112-8 (paperback). xxiv + 261 pp. LCCN QC793.4 .T38 1988. US$8.95.

Taubman:2012:PFC

[Tau12] Philip Taubman. The partnership: five cold warriors and their quest to ban the bomb. Harper, New York, NY, USA, 2012. ISBN 0-06-174400-X (hardcover), 0-06-174407-7 (paperback). xviii + 478 pp. LCCN U264.3 .T38 2012.

Taylor:1921:RRS

[Tay21] A. E. Taylor. Reviews: Relativity, the Special and the General Theory: A Popular Exposition by Albert Einstein and Robert W. Lawson, Space, Time, and Gravitation: An Outline of the General Theory of Relativity by A. S. Eddington, and The Concept of Nature by A. N. Whitehead. Mind (New Series), 30(117):

76–83, January 1921. URL 2-Z;   http://links.jstor.org/
sici?sici=0026-4423(192101)2:30:117<76:RTSATG>2.0.CO.

Taylor:1949:SHS

[Tay49] F. Sherwood (Frank Sherwood) Taylor. A short history of science and scientific thought, with readings from the great scientists from the Babylonians to Einstein. W. W. Norton & Co., New York, NY, USA, 1949. 368 pp. LCCN Q125.T28 1949.

Taylor:1995:B

[Tay95] Theodore Taylor. The bomb. Harcourt Brace and Co., San Diego, CA, USA, 1995. ISBN 0-15-200867-5. 197 pp. LCCN PZ7.T2186 Bo 1995. URL http://www.loc.

gov/catdir/description/har041/95010683.html;   http:
//www.loc.gov/catdir/enhancements/fy0733/95010683-b.
html.

Taylor:2007:B

[Tay07] Theodore Taylor. The bomb. Harcourt, Orlando, FL, USA, 2007. ISBN 0-15-206165-7s. 200 pp. LCCN PZ7.T2186 Bo 1995.

Teller:2002:HSU

[TB02] Teller Ede and Bekk Mária. Huszadik századi utazás: Tudományban és politikában. (Hungarian) [Twentieth Century Travel: Science and politics]. Huszadik Század Intézet, Budapest, Hungary, 2002. ISBN 963-9406-55-4. 595 pp. LCCN QC16.T37 M5516 2001.

Toro:1994:EA

[TC94] Tibor Toró and Paul L. Csonka. Einstein anticipated. Physics World, 7(12):17, December 1994. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
7/12/phwv7i12a17.pdf.

Tagore:1931:NR

[TE31a] Rabindranath Tagore and Albert Einstein. The nature of reality. Modern Review (Calcutta), 49(??):42–43, 1931.

Tagore:1931:TTE

[TE31b] Rabindranath Tagore and Albert Einstein. Tagore talks with Einstein. Asia, 31(??):138–142, 1931.

Teevan:2002:AEB

[Tee02] D. Teevan. Albert Einstein and Bernard Lonergan on empirical method. Zygon, 37(4):873–890, December 2002. ISSN 0591-2385 (print), 1467-9744 (electronic).

Teller:1976:OWH

[Tel76] Edward Teller. Obituary: Werner Heisenberg. Nature, 260 (5552):657–658, April 15, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v260/n5552/pdf/260657a0.pdf.

Teller:1978:AES

[Tel78] E. Teller. Albert Einstein — 3 successes, 3 failures — unconventional thoughts about an unconventional man. Interdisciplinary Science Reviews, 3(4):265–266, ???? 1978. ISSN 0308-0188.

Teller:1995:IIQ

[Tel95] Paul Teller. An interpretive introduction to quantum field theory. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-07408-9. x + 176 pp. LCCN QC174.45

.T45 1995. US$35.00. URL http://www.loc.gov/catdir/
description/prin031/94013846.html;   http://www.loc.
gov/catdir/toc/prin031/94013846.html.

Tonomura:1989:DSE

[TEM^{+}89] A. Tonomura, J. Endo, T. Matsuda, T. Kawasaki, and H. Ezawa. Demonstration of single electron buildup of an interference pattern. American Journal of Physics, 57(2):117–120, February 1989. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v57/i2/p117_s1.

Teske:1969:ESG

[Tes69] Armin Teske. Einstein und Smoluchowski: Zur Geschichte der Brownschen Bewegung und der Opaleszenz. (German) [Einstein and Smoluchowski: On the history of Brownian movement and opalescence]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 53(3):292–305, ???? 1969. CODEN SUARAH. ISSN

0039-4564. URL http://www.jstor.org/stable/20775771.

Tessman:1999:ED

[Tes99] Irwin Tessman. Einstein’s diploma. Science, 285(5426):337, July 16, 1999. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/285/5426/337.6.full.

Tesinska:2012:PTF

[Těš12] Emilie Těšinská. Profilováni teoretické fyziky na pražské univerzitě a vazby s pražským působením A. Einsteina před 100 lety. (Czech). Pokroky Matematiky, Fyziky a Astronomie, 57 (2):146–168, ???? 2012. CODEN PMFAA4. ISSN 0032-2423.

Teukolsky:1996:ETN

[Teu96] Saul A. Teukolsky. The explanation of the Trouton–Noble experiment revisited. American Journal of Physics, 64(9):1104– 1109, September 1996. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v64/i9/p1104_s1.

Teweles:1919:E

[Tew19] Heinrich Teweles. Einstein. Prager Tagblatt, 44(??):??, December 18, 1919.

Teichmann:2002:PPM

[TEW02] Jürgen Teichmann, Michael Eckert, and Stefan Wolff. Physicists and physics in Munich. Physics in Perspective (PIP), 4 (3):333–359, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8372-6.

Tromel:1999:LBL

[TF99] Martin Trömel and Rudolf Fleischmann. Leserbriefe zu: ,,Lise Meitner und der Nobelpreis”. (German) [Readers’ letters on: “Lise Meitner and the Nobel Prize”]. Physik in unserer Zeit, 30 (2):81, 1999. CODEN PHUZAH. ISSN 0031-9252.

Tropp:1993:AFM

[TFC93] E. A. (Eduard Abramovich) Tropp, Viktor Iakovlevich Frenkel, and A. D. (Artur Davidovich) Chernin. Alexander A. Friedmann: the man who made the universe expand. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38470-2. x + 267 pp. LCCN QC16.F73 T7613 1993. URL http://

www.loc.gov/catdir/description/cam025/92028315.html;
 http://www.loc.gov/catdir/toc/cam028/92028315.html.

Trbuhovic-Gjuric:1991:MEV

[TGC91] Desanka Trbuhović-Gjurić and Nicole Casanova. Mileva Einstein: une vie: [biographie]. Biographie, Biographie. Des femmes, Paris, France, 1991. ISBN 2-7210-0407-7. ISSN 0297-3128. 244 pp. LCCN ????

Turner:1979:EMI

[TH79] Dean Turner and Richard Hazelett, editors. The Einstein myth and the Ives papers: a counter-revolution in physics: with excerpts from Ives’ correspondence, The Einstein myth by Dean Turner, a condensation of Euclid or Einstein by J. J. (Jeremiah Joseph) Callahan, and papers and comments by others. Devin-Adair Co., Old Greenwich, CT, USA, 1979. ISBN 0-8159-5823-4. vii + 313 pp. LCCN QC173.59.S65 I93 1979.

tenHagen:2020:LVG

[tH20] Sjang L. ten Hagen. The Local versus the Global in the history of relativity: The case of Belgium. Science in Context, 33(3): 227–250, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/local-
versus-the-global-in-the-history-of-relativity-the-
case-of-belgium/FE3225A6006BBBADECDAE7519CB451C7.

Thackray:1977:BRB

[Tha77] Arnold Thackray. Book review: Method and Appraisal in the Physical Sciences: The Critical Background to Modern Science, 1800-1905 by Colin Howson. Journal of Interdisciplinary History, 8(2):353–354, ???? 1977. ISSN 0022-1953 (print),

1530-9169 (electronic). URL http://www.jstor.org/stable/
202795.

PP:1946:NFF

[The46] The Plutonium Project. Nuclei formed in fission: Decay characteristics, fission yields, and chain relationships. Journal of the American Chemical Society, 68(11):2411–2442, 1946. CO-DEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Thee:1995:PNP

[The95] Marek Thee, editor. Peace!: by the Nobel Peace Prize laureates: an anthology. Cultures of peace. UNESCO Publishing, Paris, France, 1995. ISBN 92-3-103193-7 (paperback). 570 pp. LCCN JX1937 .P42 1995.

Theaker:2012:BRE

[The12] Martin Theaker. Book review: Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939–59, Christoph Laucht, Basingstoke, Palgrave Macmillan, 2012, xiv + 274 pp., ISBN 978-0-230-35487-6 (hbk) (£55.00). Contemporary British History, 26(4):553–555, December 2012. ISSN 1361-9462 (print), 1743-7997 (electronic).

Thirring:1921:IET

[Thi21] Hans Thirring. The Ideas of Einstein’s Theory: the Theory of Relativity in Simple Language. Methuen and Co. Ltd., London, UK, 1921. xiv + 167 pp. LCCN QC6.T5. URL

http://hdl.handle.net/2027/mdp.39015004968445. Trans-

lation from 1921 German edition to English by Rhoda A. B. Russell.

Thirring:1946:GAE

[Thi46] Hans Thirring. Die Geschichte der Atombombe; mit einer elementaren Einführung in die Atomphysik auf Grund der Originalliteratur gemeinverständlich dargestellt. (German) [The history of the atomic bomb, with an elementary introduction to atomic physics according to the original literature, understandably presented]. Wissenschaft für jedermann; Phönix-Bücherel. “Neues Österreich” Zeitungs- und Verlagsgesellschaft, Wien, Austria, 1946. 149 pp. LCCN QC778 .T47.

Thiessen:2003:EG

[Thi03] Vern Thiessen. Einstein’s gift. Playwrights Canada Press, Toronto, ON, Canada, 2003. ISBN 0-88754-678-1. vii + 108 pp. LCCN PR9199.3.T4486 E38 2003. URL http://www.loc.

gov/catdir/enhancements/fy0621/2004401001-d.html.

Thomson:1931:JCMb

[Tho31] Sir Joseph John Thomson, editor. James Clerk Maxwell; a commemoration volume, 1831–1931. Cambridge University Press, Cambridge, UK, 1931. 3 + 146 pp. LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Thomsen:1983:MDG

[Tho83] D. E. Thomsen. Many dimensions in gravity theory — quantizing General Relativity means rolling dice on Albert Einstein field. Science News (Washington, DC), 124(4):60–61, ???? 1983. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Thomas:1985:DSR

[Tho85] Donald Serrell Thomas. The day the sun rose twice. Macmillan Publishing Company, New York, NY, USA, 1985. ISBN 0-333-38604-3. 285 pp. LCCN PR6070.H6 D3 1985.

Thompson:1988:ANL

[Tho88a] Larry Thompson. America’s Nobel laureates in medicine, physiology, and chemistry: a tribute to America’s living Nobel Prize winners. Friends of the National Library of Medicine, Washington, DC, USA, second edition, 1988. 62 pp. LCCN ???? URL

http://resource.nlm.nih.gov/8909073.

Thouart:1988:FPF

[Tho88b] D. Thouart. Film portrayals of famous scientists: Galileo, Louis Pasteur, Rene Laennec, Jean Henri Fabre, Thomas Edison, Albert Einstein, Marie Curie, and others. Historia, ??(504):96–99, December 1988. ISSN 0018-2281.

Thorne:1993:BHT

[Tho93] Kip S. Thorne. Black holes and time warps: Einstein’s outrageous legacy. Commonwealth Fund Book Program. W. W. Norton & Co., New York, NY, USA, 1993. ISBN 0-393-03505-0. 619 pp. LCCN QC6.T526 1993; QC6 .T565 1995; QC6 .T526 1994. Foreword by Stephen Hawking.

Thorne:1994:BHT

[Tho94] Kip S. Thorne. Black holes and time warps: Einstein’s outrageous legacy. The Commonwealth Fund Book Program. W. W. Norton & Co., New York, NY, USA, 1994. ISBN 0-393-03505-0, 0-393-31276-3. 619 pp. LCCN QC6 .T526 1994. URL http://

www.gbv.de/dms/bowker/toc/9780393035056.pdf.

Thompson:1997:NPW

[Tho97] Clifford Thompson, editor. Nobel Prize winners. 1992–1996, Supplement: an H. W. Wilson biographical dictionary. H. W. Wilson Co., New York, NY, USA, 1997. ISBN 0-8242-0906-0. 165 pp. LCCN AS911.N9 N59 1992 Suppl.

Thomas:2000:AEL

[Tho00] M. Thomas. Albert Einstein and LD: An evaluation of the evidence. Journal of Learning Disabilities, 33(2):149–157, March/ April 2000. ISSN 0022-2194.

Thomsen:2005:AEM

[Tho05] M. Thomsen. Albert Einstein to Michele Besso. Physics Today, 58(7):14–15, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Thorpe:2006:OTI

[Tho06] Charles Thorpe. Oppenheimer: the tragic intellect. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-79845-3 (hardcover), 0-226-79846-1 (paperback). xx + 413 + 12 pp. LCCN QC16.O62 T56 2006. URL http://www.loc.

gov/catdir/enhancements/fy0666/2006015223-b.html;
http://www.loc.gov/catdir/enhancements/fy0666/2006015223-
d.html;   http://www.loc.gov/catdir/enhancements/fy0666/
2006015223-t.html.

Thomson:2011:JCM

[Tho11a] Joseph John Thomson, editor. James Clerk Maxwell: a commemoration volume 1831–1931. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-67095-0 (paperback). 6 + 146 pp. LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Thorpe:2011:BRS

[Tho11b] Charles Thorpe. Book review: Silvan S. Schweber, Einstein and Oppenheimer, ISBN-13 978-0-674-02828-9. Annals of Science, 68(4):558–561, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thompson:2012:NPC

[Tho12] Gilbert Thompson, editor. Nobel Prizes that changed medicine. Imperial College Press, London, UK, 2012. ISBN 1-84816-825-X (hardcover). 329 pp. LCCN ????

Thomas:2013:AFE

[Tho13] Kelly Devine Thomas. The advent and fallout of EPR. The Institute Letter, Institute for Advanced Study, ??(??):??, Fall

2013. URL http://www.ias.edu/ideas/2013/epr-fallout.

Thorpe:2015:HTL

[Tho15] Scott Thorpe. How to think like Einstein: simple ways to break the rules and discover your hidden genius. Sourcebooks, Inc., Naperville, Illinois, second edition, 2015. ISBN 1-4926-2627-9. ???? pp. LCCN BF449 .T48 2015.

Thomas:2017:PPM

[Tho17] Linda Carrick Thomas. Polonium in the Playhouse: the Manhattan Project’s secret chemistry work in Dayton, Ohio. Trillium, an imprint of The Ohio State University Press, Columbus, OH, USA, 2017. ISBN 0-8142-1338-3 (hardcover). xxi + 247 pp. LCCN QC773.3.U5 T46 2017.

Thorne:2018:NLL

[Tho18] Kip S. Thorne. Nobel lecture: LIGO and gravitational waves III. Reviews of Modern Physics, 90(4):040503–??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL https://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.040503.

Throesch:2017:BEF

[Thr17] Elizabeth Throesch. Before Einstein: the fourth dimension in fin-de-siècle in literature and culture. Anthem nineteenthcentury series. Anthem Press, New York, NY, USA, 2017. ISBN 1-78308-623-8, 1-78308-625-4 (e-book), 1-78308-624-6 (ebook). LCCN PR149.S75. URL http://www.jstor.org/

stable/10.2307/j.ctt1jktqh1;   http://www.oapen.org/
search?identifier=626390.

Tibbets:2005:REG

[Tib05] Paul W. (Paul Warfield) Tibbets. Return of the Enola Gay. Enola Gay Remembered, Inc., New Hope, PA, 60th anniversary restoration edition, 2005. ISBN 0-9703666-0-4. 339 pp. LCCN D767.25.H6 T533 2004.

Tibbetts:2013:HGS

[Tib13] Gary G. Tibbetts. How the Great Scientists Reasoned: the Scientific Method in Action. Elsevier insights. Elsevier, Amsterdam, The Netherlands, 2013. ISBN 0-12-398498-X (hardcover). viii + 148 pp. LCCN Q175 .T547 2013.

Tiozzo:2009:LIP

[Tio09] Enrico Tiozzo. La letteratura italiana e il premio Nobel: storia critica e documenti. (Italian) [Italian literature and the Nobel Prize: a critical history and documents], volume 351 of Biblioteca dell’“Archivum Romanicum.” Serie I, Storia, letteratura, paleografia. Leo S. Olschki, Firenze, Italia, 2009. ISBN 88-222-5817-7. vii + 355 pp. LCCN PQ4087 .T675 2009; PQ4087 .T56 2009.

Tippett:2010:EGC

[Tip10] Krista Tippett. Einstein’s God: conversations about science and the human spirit. Penguin, London, UK and New York, NY, USA, 2010. ISBN 0-14-311677-0 (paperback), 1-4487-2130-X, 1-101-19541-X (e-book), 1-101-19583-5 (e-book). 286 pp. LCCN BL240.3 .T57 2010.

Tippett:2012:ETB

[Tip12] Krista Tippett. Einstein’ın Tanrısı: bilim ve insan ruhu üzerine sohbetler, volume 7 of h20 kitap; Bilim ve Din dizisi; 1. h2o Kitap, Istanbul, Turkey, 2012. ISBN 605-62473-6-8. 271 pp. LCCN BL240.3 .T5720 2013. Turkish translation by Gizem Aldoğan of [Tip10].

Tiselius:1952:PSN

[Tis52] Arne Tiselius. Protest from Swedish Nobel Laureate. Bulletin of the Atomic Scientists, 8(7):256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Tu:2010:NDG

[TLW^{+}10] Liang-Cheng Tu, Qing Li, Qing-Lan Wang, Cheng-Gang Shao, Shan-Qing Yang, Lin-Xia Liu, Qi Liu, and Jun Luo. New determination of the gravitational constant G with time-of-swing method. Physical Review D (Particles and Fields), 82 (2):022001, July 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500. URL http://link.

aps.org/doi/10.1103/PhysRevD.82.022001.

Thomas:1977:RAA

[TM77] Gordon Thomas and Max Morgan Witts. Ruin from the air: the atomic mission to Hiroshima. Hamilton, London, UK, 1977. ISBN 0-241-89726-2. xvii + 386 + 4 pp. LCCN D767.25.H6 T53 1977.

Thomas:1990:RAE

[TM90] Gordon Thomas and Max Morgan Witts. Ruin from the air: the Enola Gay’s atomic mission to Hiroshima. Scarborough House, Chelsea, MI, 1990. ISBN 0-8128-8509-0. xvii + 386 + 16 pp. LCCN D767.25.H6 T53 1990.

Trouton:1903:MFA

[TN03] Frederick Thomas Trouton and H. R. Noble. The mechanical forces acting on a charged electric condenser moving through space. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 202(??):165–181, ???? 1903. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://en.wikipedia.

org/wiki/Trouton_Noble_experiment.

Toader:2023:ECC

[Toa23] Iulian D. Toader. Einstein completeness as categoricity. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.1007/s10701-023-00675-0.

Tobies:1994:AEF

[Tob94] Renate Tobies. Albert Einstein und Felix Klein. Naturwissenschaftliche Rundschau, 47(??):345–354, ???? 1994. CODEN NARSAC. ISSN 0028-1050.

Todd:2014:BHL

[Tod14] Neil Todd. A brief history of Lord Rutherford’s radium. Notes and Records of the Royal Society of London, 68(3): 279–300, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://rsnr.

royalsocietypublishing.org/content/68/3/279;   http:
//www.jstor.org/stable/43287732.

Tolman:1934:RTC

[Tol34] Richard Chace Tolman. Relativity, thermodynamics and cosmology. The International series of monographs on physics. Clarendon Press, Oxford, UK, 1934. xv + 501 + 1 pp. LCCN QC6 .T57.

Tolf:1976:RRS

[Tol76] Robert W. Tolf. The Russian Rockefellers: the saga of the Nobel family and the Russian oil industry, volume 158 of Hoover Institution publication. Hoover Institution Press, Stanford University, Stanford, CA, USA, 1976. ISBN 0-8179-6581-5. xv + 269 + 12 pp. LCCN HD9575.R82 T64.

Tolman:1987:RTC

[Tol87] Richard Chace Tolman. Relativity, thermodynamics, and cosmology. Dover, New York, NY, USA, 1987. ISBN 0-486-65383-8 (paperback). xv + 501 pp. LCCN QC173.65

.T65 1987. URL http://www.loc.gov/catdir/description/
dover031/87006728.html.

Tol:2023:NBN

[Tol23] Richard S. J. Tol. Nobel begets Nobel in economics. Journal of Informetrics, 17(4):??, November 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000822.

Tol:2024:NF

[Tol24] Richard S. J. Tol. The Nobel family. Scientometrics, 129(3): 1329–1346, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04936-1.

Tomlin:1999:BEC

[Tom99] Sarah Tomlin. Bose–Einstein condensates: Visions of vortices. Nature, 397(6717):301, January 28, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v397/n6717/
full/397301b0.html.

Tom:2010:ICM

[Tom10] Karier Tom. Intellectual Capital, Forty years of the Nobel Prize in Economics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76326-6, 0-511-77897-X (e-book). xiii + 351 pp. LCCN 18 f.

Tomczyk:2022:DEP

[Tom22] Hannah Tomczyk. Did Einstein predict Bose–Einstein condensation? Studies in History and Philosophy of Science Part A, 93(??):30–38, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368122000450.

Tonietti:1997:AEA

[Ton97] Tito M. Tonietti. Albert Einstein and Arnold Schönberg correspondence. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 5(1):1–22, December 1997. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL

http://link.springer.com/article/10.1007/BF02913641.

Topper:1989:BRT

[Top89] David Topper. Book review: Thematic Origins of Scientific Thought: Kepler to Einstein, by Gerald Holton. Leonardo (Oxford, England), 22(2):267, April 1989. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL https://muse.

jhu.edu/article/602266/summary.

Topper:2007:QSS

[Top07a] David Topper. Quirky sides of scientists: true tales of ingenuity and error from physics and astronomy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71019-1. xiii + 210 pp. LCCN Q172.5.E77 T67

Topper:2007:WND

[Top07b] David Topper. A well-nigh discovery: Einstein and the expanding universe. In Quirky sides of scientists: true tales of ingenuity and error from physics and astronomy [Top07a], chapter 12, pages 201–202. ISBN 0-387-71019-1. LCCN Q172.5.E77 T67 2007.

Tornqvist:1981:SEP

[Tör81] Nils A. Törnqvist. Suggestion for Einstein–Podolsky–Rosen experiments using reactions like e^{+}e^{−} → ΛΛ̄ → π^{−}pπ^{+}p̄. Foundations of Physics, 11(1–2):171–177, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715204.

Torstendahl:1982:EIP

[Tor82] Rolf Torstendahl. Engineers in industry, 1850–1910: Professional men and new bureaucrats. A comparative approach. In Bernhard et al. [BCS82], pages 253–270. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500219.

Torsten:2003:NLE

[Tor03] Persson Torsten. Nobel lectures in economic sciences 1996–2000. Nobel lectures including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-02-4960-8, 981-02-4961-6. xii + 365 pp. LCCN 9.5 nob. pbk. A collection of the Nobel lectures delivered by the prizewinners in the economic sciences, together with their biographies, portraits and the presentation speeches at the award ceremonies for the period 1996-2000. Each lecture is based on the work for which the laureate was awarded the prize.

Tousignant:2005:BRM

[Tou05] N. Tousignant. Book review: Milan Popovic (editor), In Albert’s Shadow: The Life and letters of Mileva Marić, Einstein’s First Wife. European Review of History = Revue européenne d’histoire, 12(3):522–523, ???? 2005. ISSN 1350-7486 (print),

1469-8293 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/13507480500492298.

Tourville:2014:AFW

[Tou14] Jacqueline Tourville. Albie’s first word: a tale inspired by Albert Einstein’s childhood. Schwartz and Wade Books, New York, NY, USA, 2014. ISBN 1-322-22990-2, 0-385-39265-6, 0-307-97893-1, 0-307-97894-X, 0-307-97895-8 (e-book). 40 (est.) pp. LCCN PZ7.T6477.

Townes:1965:NLP

[Tow65] Charles H. Townes. 1964 Nobel lecture: Production of coherent radiation by atoms and molecules. IEEE Spectrum, 2(8):30– 43, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tozser:2015:ETO

[Tőz15] Árpád Tőzsér. Einstein a teremtést olvassa: naplók (2005–2007) naplója. (Hungarian) [The creation of Einstein: diary (2005– 2007)]. Kalligram, Pozsony, Slovakia [??], 2015. ISBN 80-8101-828-X. 280 + 2 pp. LCCN ????

Troemel-Ploetz:1990:MEM

[TP90] Senta Troemel-Ploetz. Mileva Einstein-Marić: The woman who did Einstein’s mathematics. Women’s Studies International Forum, 13(5):415–432, ???? 1990. CODEN WSINDA. ISSN 0277-5395 (print), 1879-243X (electronic).

URL https://www.sciencedirect.com/science/article/
abs/pii/027753959090094E.

Thorne:1986:BHM

[TPM86] Kip S. Thorne, Richard H. Price, and Douglas A. (Douglas Alan) Macdonald, editors. Black holes: the membrane paradigm. Yale University Press, New Haven, CT, USA, 1986. ISBN 0-300-03769-4, 0-300-03770-8 (paperback). xii + 367 pp. LCCN QB843.B55 B59 1986.

Travis:2000:BRB

[Tra00] Anthony S. Travis. Book review: Nobel Laureates in Chemistry, 1901–1992 by Laylin K. James. Isis, 91(3):641–642, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237994.

Tracy:2016:HIM

[Tra16] Suzanne Tracy. Haunting interactive map shows every nuclear detonation since 1945. Scientific Computing, ??(??):??, March 7, 2016. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156

(electronic). URL http://www.scientificcomputing.com/
news/2016/03/haunting-interactive-map-shows-every-nuclear-
detonation-1945.

Trevor:1921:RRS

[Tre21] J. E. Trevor. Review: Relativity. The Special and General Theory, by Albert Einstein; Robert W. Lawson. The Philosophical Review, 30(2):213–214, March 1921. URL 2-O; http://

links.jstor.org/sici?sici=0031-8108(192103)30:2<213:
RTSAGT>2.0.CO.

Treder:1966:ESE

[Tre66] Hans-Jürgen Treder, editor. Einstein-Symposium “Entstehung, Entwicklung und Perspektiven der Einsteinschen Gravitationstheorie”: vom 2.–5. November 1965 in Berlin aus Anlaß des

50. Jahrestages der Entdeckung der allgemeinen relativistischen Gravitationsgleichungen und ihrer Vorlage in der Gesamtsitzung der Berliner Akademie am 4. November 1915 durch Albert Einstein: Vorträge und Diskussionen. Akademie-Verlag, Berlin, Germany, 1966. LCCN ????

Treder:1971:GAL

[Tre71] Hans-Jürgen Treder, editor. Gravitationstheorie und Äquivalenzprinzip: Lorentz-Gruppe, Einstein-Gruppe und Raumstruktur. (German) [Gravitational theory and the Principle of Equivalence: Lorentz Group, Einstein Group, and the structure of space], volume 26 of Mathematische Lehrbücher und Monographien Abt. 2, Mathematische Monographien. Akademie-Verlag, Berlin, Germany, 1971. viii + 122 pp. LCCN ????

Treder:1975:ESE

[Tre75a] H.-J. Treder. The Einstein shift in Einstein’s box experiment. Foundations of Physics, 5(1):135–142, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01100322.

Trenn:1975:RRA

[Tre75b] Thaddeus J. Trenn. Rutherford and recoil atoms: The metamorphosis and success of a once stillborn theory. Historical Studies in the Physical Sciences, 6(??):513–547, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757348.

Treder:1979:AEB

[Tre79a] Hans Jürgen Treder, editor. Albert Einstein: Beiträge zum 100. Geburtstag A. Einsteins. (German) [Albert Einstein: Contributions for Einstein’s 100th Birthday], volume 14 of Potsdamer Forschungen. Reihe B. Pädagogische Hochschule “Karl Liebknecht,” Wissenschaftlich-Technisches Zentrum, Potsdam, Germany, 1979. 100 pp. LCCN QC16.E5 A658 1979.

Trenn:1979:CRE

[Tre79b] Thaddeus J. Trenn. The central role of energy in Soddy’s holistic and critical approach to nuclear science, economics, and social responsibility. British Journal for the History of Science, 12(3): 261–276, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Trenn:1988:BRE

[Tre88] Thaddeus J. Trenn. Book review: Elisabeth Crawford, J. L. Heilbron, Rebecca Ullrich. The Nobel Population (1901–1937): A census of the Nominators and Nominees for the Prizes in Physics and Chemistry. Berkeley: Office for History of Science and Technology, 1987. Pp. vii + 337. ISBN 0-918102-15-4. $20.00. British Journal for the History of Science, 21(4):497, December 1988. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4026969.

Trulock:1996:MHH

[Tru96] David W. Trulock. Meitner, Hahn, Hitler, and Siegbahn. American Journal of Physics, 64(5):523–524, May 1996. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/64/5/10.1119/1.18271. See [Sim94, Kac96].

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001.

URL http://www.loc.gov/catdir/enhancements/fy0833/
2001097880-d.html;   http://www.loc.gov/catdir/toc/
fy02/2001097880.html.

Tsipis:2008:DRR

[Tsi08] Kosta Tsipis. David Rowe and Robert Schulmann, Einstein on Politics. Journal of Cold War Studies, 10(3):140–143, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/241504;
 http://muse.jhu.edu/article/241504/pdf;   https://
direct.mit.edu/jcws/article-pdf/10/3/140/697109/jcws.
2008.10.3.140.pdf.

Tsui:1999:NLI

[Tsu99] Daniel C. Tsui. Nobel Lecture: Interplay of disorder and interaction in two-dimensional electron gas in intense magnetic fields. Reviews of Modern Physics, 71(4):891–895, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.71.891;   http://rmp.aps.org/
abstract/RMP/v71/i4/p891_1.

Taylor:1970:UFN

[TT70] Raymond W. Taylor and Samuel Woolley Taylor. Uranium fever; or, No talk under $1 million. Macmillan Publishing Company, New York, NY, USA, 1970. xi + 400 pp. LCCN F830 .T3.

Turki:2020:FCW

[TTB20] Houcemeddine Turki, Mohamed Ali Hadj Taieb, and Mohamed Ben Aouicha. Facts to consider when analyzing the references of Nobel Prize scientific background. Scientometrics, 124(1):787– 790, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03456-y. See [Bjø20b] and re-

sponse [Bjø20c, Bjø20a].

Teller:1968:CUN

[TTHJ68] Edward Teller, Wilson K. Talley, Gary H. Higgins, and Gerald W. Johnson, editors. The constructive uses of nuclear explosives. McGraw-Hill, New York, NY, USA, 1968. xiv + 320 pp. LCCN TK9153 .C6.

Tugnoli:1998:BRH

[Tug98] Claudio Tugnoli. Book review: Henri Bergson, Durata e simultaneità (a proposito della teoria di Einstein), a cura di Paolo Taroni, Pitagora editrice, Collana di Storia delle idee, sezione Fonti e Documenti, Bologna, 1997, 209 pp. Nuncius, 13(1):338–341, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00293.

Tulkes:1979:AE

[Tul79] Samuel G. Tulkes. Albert Einstein. Molecular and Cellular Biochemistry, 27(1):67, ???? 1979. CODEN MCBIB8. ISSN 0300-8177 (print), 1573-4919 (electronic). URL http://link.

springer.com/article/10.1007/BF00217794.

Turner:1940:MHN

[Tur40a] Louis A. Turner. The missing heavy nuclei. Physical Review, 57 (11):950–957, June 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.57.950.

Turner:1940:NTE

[Tur40b] Louis A. Turner. The nonexistence of transuranic elements. Physical Review, 57(2):157, January 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.57.157.

Turner:1940:NF

[Tur40c] Louis A. Turner. Nuclear fission. Reviews of Modern Physics, 12(1):1–29, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.12.1;
 http://rmp.aps.org/abstract/RMP/v12/i1/p1_1. This pa-

per surveys the early work from 1934 to 1939 on nuclear fission, including more than 140 from 1939 alone following the news of the Hahn and Strassmann work that Niels Bohr brought to the USA in January 1939.

Turner:1946:AE

[Tur46] Louis A. Turner. Atomic energy from U^{238}. Physical Review, 69(7–8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.69.366.

Turner:1966:ECD

[Tur66] Barry E. Turner. Einstein A coefficient for the Λ doublet transitions of the ground state of OH. Nature, 212(5058):184– 185, October 8, 1966. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v212/n5058/pdf/212184b0.pdf.

Turner:1967:CEC

[Tur67] Barry E. Turner. Calculations of the Einstein A coefficient for the 18 cm transition of OH. Nature, 214(5086):379, April 22, 1967. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v214/n5086/pdf/214379a0.pdf.

Turner:1988:GL

[Tur88] Edwin L. Turner. Gravitational lenses. Scientific American, 259(1):26–32, July 1988. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Turchetti:2003:ASG

[Tur03] Simone Turchetti. Atomic secrets and governmental lies: nuclear science, politics and security in the Pontecorvo case. British Journal for the History of Science, 36(4):389–415, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4028213.

Turchetti:2005:BRR

[Tur05] Simone Turchetti. Book review: Robert Marc Friedman, The Politics of Excellence. Behind the Nobel Prize in Science. New York: Times Books, 2001. xv + 379 pp., ISBN 0-7167-3103-7. Nuncius, 20(1):272–273, ???? 2005. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01096.

Turchetti:2006:BRG

[Tur06a] Simone Turchetti. Book review: Gerald Holton, Victory and Vexation in Science. Einstein, Bohr, Heisenberg and Others. Cambridge, Mass.: Harvard University Press, 2005. xiv + 230 pp., ill., ISBN 0-674-01519-3. Nuncius, 21(1):191–192, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00401.

Turchetti:2006:IBN

[Tur06b] Simone Turchetti. The invisible businessman: Nuclear physics, patenting practices, and trading activities in the 1930s. Historical Studies in the Physical and Biological Sciences, 37(1): 153–172, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Topper:2000:PEQ

[TV00] David Topper and Dwight E. Vincent. Posing Einstein’s question: Questioning Einstein’s pose. The Physics Teacher, 38 (5):278–??, May 2000. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL http://tpt.aapt.org/

resource/1/phteah/v38/i5/p278_s1.

Topper:2007:ETB

[TV07] David Topper and Dwight Vincent. Einstein’s 1934 twoblackboard derivation of energy-mass equivalence. American Journal of Physics, 75(11):978–983, November 2007. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://m.ajp.aapt.org/resource/1/ajpias/v75/i11/
p978_s1.

Teerikorpi:2009:EUO

[TVL^{+}09a] Pekka Teerikorpi, Mauri Valtonen, Kirsi Lehto, Harry Lehto, Gene Byrd, and Arthur Chernin. The Evolving Universe and the Origin of Life: The Search for Our Cosmic Roots. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-09534-9 (print), 0-387-09534-9 (e-book). xvii + 508 + 12 pp. LCCN QB981 .E845 2009.

Teerikorpi:2009:FIU

[TVL^{+}09b] Pekka Teerikorpi, Mauri Valtonen, Kirsi Lehto, Harry Lehto, Gene Byrd, and Arthur Chernin. Finite or infinite universe: Cosmological models. In The Evolving Universe and the Origin of Life: The Search for Our Cosmic Roots [TVL^{+}09a], chapter 23, pages 291–307. ISBN 0-387-09534-9 (print), 0-387-09534-9 (e-book). LCCN QB981 .E845 2009.

Taylor:1966:SP

[TW66] Edwin F. Taylor and John Archibald Wheeler. Spacetime physics. A Series of books in physics. W. H. Freeman and Company, New York, NY, USA, 1966. 208 pp. LCCN QC6 .T35.

Thomas:1977:EG

[TW77] Gordon Thomas and Max Morgan Witts. Enola Gay. Stein and Day, New York, NY, USA, 1977. ISBN 0-8128-2150-1. 327 + 11 pp. LCCN D767.25.H6 T5.

Taylor:1992:SPI

[TW92] Edwin F. Taylor and John Archibald Wheeler. Spacetime physics: introduction to special relativity. W. H. Freeman and Company, New York, NY, USA, second edition, 1992. ISBN 0-7167-2327-1. vii + 312 pp. LCCN QC173.65 T37

1991. URL http://www.loc.gov/catdir/enhancements/
fy1011/92000722-d.html;   http://www.loc.gov/catdir/
enhancements/fy1011/92000722-t.html.

Taylor:2000:EBH

[TW00] Edwin F. Taylor and John Archibald Wheeler. Exploring black holes: introduction to general relativity. Addison-Wesley Longman, Reading, MA, USA, 2000. ISBN 0-201-38423-X. 330 (est.) pp. LCCN QC173.6 .T39 2000.

Terhal:2003:QEM

[TWD03] Barbara M. Terhal, Michael M. Wolf, and Andrew C. Doherty. Quantum entanglement: a modern perspective. Physics Today, 56(4):46–52, April 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Tymieniecka:1995:HEB

[Tym95] Anna-Teresa Tymieniecka, editor. Heaven, earth, and the inbetween in the harmony of life, or, Phenomenology in the continuing oriental/occidental dialogue: Phenomenology in the continuing oriental, volume 47 of Analecta Husserliana. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3373-X. LCCN B3279.H94 A129 vol. 47 B829.5.

Udall:1994:MAP

[Uda94] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Pantheon Books, New York, NY, USA, 1994. ISBN 0-679-43364-3. xii + 399 pp. LCCN E840 .U33 1994.

Udall:1998:MAP

[Uda98] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Rutgers University Press, New Brunswick, NJ, USA, 1998. ISBN 0-8135-2546-2. xii + 399 pp. LCCN E840 .U33 1998. URL http://www.loc.gov/

catdir/enhancements/fy1502/97053053-d.html.

Uffink:2006:IDE

[Uff06] Jos Uffink. Insuperable difficulties: Einstein’s statistical road to molecular physics. Studies in History and Philosophy of Modern Physics, 37(1):36–70, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000948.

Uhlenbeck:1980:SRA

[Uhl80] George E. Uhlenbeck. Some reminiscenses about Einstein’s visits to Leiden. In Woolf [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Renn:2015:EWA

[uJR15] Michel Janssen und Jürgen Renn. Einsteins Weg zur allgemeinen Relativitätstheorie. (German) [Einstein’s road to the General Theory of Relativity]. Spektrum der Wissenschaft (German translation of Scientific American), ?? (??):??, October 2015. CODEN SPEKDI. ISSN 0170-

2971. URL http://www.spektrum.de/magazin/einsteins-
weg-zur-allgemeinen-relativitaetstheorie/1362267.

Upadhye:2004:SAS

[UKKP04] Rekha P. Upadhye, V. L. Kalyane, Vijai Kumar, and E. R. Prakasan. Scientometric analysis of synchronous references in the Physics Nobel lectures, 1981–1985: A pilot study. Scientometrics, 61(1):55–68, September 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3ASCIE.0000037362.11986.42.

Ulam:1976:AM

[Ula76] Stanisław M. Ulam. Adventures of a mathematician. Charles Scribner’s Sons, New York, NY, USA, 1976. ISBN 0-684-14391-7. xi + 317 + 25 pp. LCCN QA29.U4 A33; QA29.U4 A33 1976; QA29.U4 A331. Paperback editions published in 1977 and 1983. Translated into Japanese (1979).

Ullmann-Margalit:1986:PSI

[UM86] Edna Ullmann-Margalit, editor. The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2, volume 95(2) of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. ix + 250 pp. LCCN Q174 .B67

vol. 95 Q175. URL http://www.springerlink.com/content/
978-94-009-4566-1;   https://link.springer.com/book/
10.1007/978-94-009-4566-1.

Ullmann-Margalit:1992:SEB

[UM92] Edna Ullmann-Margalit. The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4, volume 146 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. ix + 302 pp. LCCN Q174 .S354 1992. URL https://link.

springer.com/book/10.1007/978-94-011-2688-5.

Ungar:1998:PEH

[Ung98] Abraham A. Ungar. From Pythagoras to Einstein: The hyperbolic Pythagorean theorem. Foundations of Physics, 28(8): 1283–1321, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A:1018874826277.

Ungar:2007:EVA

[Ung07] Abraham Albert Ungar. Einstein’s velocity addition law and its hyperbolic geometry. Computers and Mathematics with Applications, 53(8):1228–1250, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL https://

dblp.org/db/journals/cma/cma53.html#Ungar07.

USACHRE:1996:FRA

[Uni96] United States.Advisory Committee on Human Radiation Experiments. Final report of the Advisory Committee on Human Radiation Experiments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-510792-6. xxxii + 620 pp. LCCN R853.H8 U53 1996. URL http://www.loc.gov/

catdir/enhancements/fy0604/96190400-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0604/96190400-t.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
96190400-b.html.

USFBI:2000:AE

[Uni00] United States Federal Bureau of Investigation. Albert Einstein. Federal Bureau of Investigation, Washington, DC, USA, 2000. ISBN ???? 1427 pp. LCCN ???? URL http://foia.fbi.gov/

foiaindex/einstein.htm.

Unruh:1999:LTG

[Unr99] Bill Unruh. Lives, times and general relativity. Physics World, 12(3):42–43, March 1999. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/12/3/phwv12i3a30.

pdf. Review of The River of Time, 1998, Cambridge University Press, 275pp, £30.00 / $49.95hb, £9.95 / $15.95pb.

Unsold:1964:PKE

[Uns64] A. Unsöld. Ptolemäus — Kopernikus — Einstein. Physikalische Blätter, 20(5):204–211, May 1964. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/phbl.19640200502/

.

Unsold:1981:UE

[Uns81] Albrecht Unsöld. Unsöld on Einstein. Nature, 291(5814): 374, June 4, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v291/n5814/pdf/291374c0.pdf.

Unser:2000:WPM

[Uns00] Margit Unser, editor. Wolfgang Pauli and modern physics = Wolfgang Pauli und die moderne Physik. ETH-Bibliothek, Zürich, Switzerland, 2000. ISBN 3-9521386-2-2. 124 pp. LCCN QC16.P37 W635 2000.

Urbach:1922:KBP

[Urb22] Benno Urbach. Kritische Bemerkungen zur philosophischen Bekampfung der Einsteinschen Relativitätstheorie durch Prof.

Dr. O. Kraus. (German) [Critical remarks on the philosophical fight against Einstein’s theory of relativity by Prof. Dr. O. Kraus]. Lotos, 70(??):309–332, 1922.

Ushenko:1949:EIC

[Ush49] Andrew Paul Ushenko. Einstein’s influence on contemporary philosophy. In Schilpp [Sch49c], chapter 24, pages 607–645. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Uvnas:1982:RPD

[Uvn82] Börje Uvnäs. The rise of physiology during the nineteenth century. In Bernhard et al. [BCS82], pages 135– 145. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981.

US$49.50 (est.). URL http://www.sciencedirect.com/
science/article/pii/B9780080279398500141.

Vaas:2017:EEG

[Vaa17] Rüdiger Vaas. Einfach Einstein! Geniale Gedanken schwerelos verständlich. (German) [Simply Einstein! Ingenious thoughts, weightless, understandable]. Kosmos, Stuttgart, Germany, 2017. ISBN 3-440-15836-5. 300 pp. LCCN ????

Vail:1989:AEI

[Vai89] A. Vail. Albert Einstein introduction to Diels’ translation of Lucretius. Classical World, 82(6):435–436, July/August 1989. ISSN 0009-8418.

Vallarta:1949:GRE

[Val49] M. S. Vallarta. Galactic rotation effect and the origin of cosmic radiation. Reviews of Modern Physics, 21(3):356, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.356;   http://rmp.aps.org/

abstract/RMP/v21/i3/p356_1. Special issue in honor of Albert Einstein on his 70th birthday.

Vallentin:1954:DAE

[Val54a] Antonina Vallentin. The drama of Albert Einstein. Doubleday, Garden City, NY, USA, 1954. 312 pp. LCCN WZ E35V 1954; QC16.E5 V314 1954a; QC16.E5 V314 1954; QC16.E5 V1b 1954; QC16.E5 V1833 1954. Translated by Moura Budberg.

Vallentin:1954:EBC

[Val54b] Antonina Vallentin. Einstein, a biography. Weidenfeld and Nicolson, London, UK, 1954. 219 pp. LCCN QC16.E3 V2; QC16.E5 V314 1954; QC16.E5 V35; QC16.E5 V314; QC16.E5 V3 1954.

Vallentin:1955:AEG

[Val55] Antonina Vallentin. Albert Einstein: geniet och människan. (Swedish) [Albert Einstein: genius and human]. Natur och kultur, Stockholm, Sweden, 1955. 233 pp.

Vallentin:1964:E

[Val64] Antonina Vallentin. Einstein. In Rapport and Wright [RW64], pages 275–297.

Valdevit:2006:OFS

[Val06] Giampaolo Valdevit. Oppenheimer fra scienza e potere. Una storia americana. (Italian) [Oppenheimer between science and power. An American story]. Studi Storici, 47(1):115–142, January/March 2006. ISSN 0039-3037 (print), 2036-458X (elec-

tronic). URL http://www.jstor.org/stable/20567340.

VanderPost:1970:NNM

[Van70] Laurens Van der Post. The night of the new moon. Hogarth, London, UK, 1970. ISBN 0-7012-0340-4. 157 + 6 pp. LCCN D805.J4 V34.

VanderPost:1971:PB

[Van71] Laurens Van der Post. The prisoner and the bomb. Morrow, New York, NY, USA, 1971. 152 pp. LCCN D805.J4 V34 1971.

VanFraassen:1974:EPR

[Van74] Bas C. Van Fraassen. The Einstein–Podolsky–Rosen paradox. Synthese, 29(1–4):291–309, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484962.

Vanzandt:1979:AET

[Van79] J. D. Vanzandt. Albert Einstein theory of general-relativity — G. Tauber. Auslegung — a Journal of Philosophy, 7(1):88–96, ???? 1979. ISSN 0733-4311.

VanHelden:1985:MUC

[Van85] Albert Van Helden. Measuring the universe: cosmic dimensions from Aristarchus to Halley. University of Chicago Press, Chicago, IL, USA and London, UK, 1985. ISBN 0-226-84881-7, 0-226-84882-5 (paperback). viii + 203 pp. LCCN QB15

.V33 1985. URL http://www.loc.gov/catdir/description/
uchi052/84016397.html;   http://www.loc.gov/catdir/
enhancements/fy0608/84016397-t.html;   http://www.loc.
gov/catdir/enhancements/fy0609/84016397-b.html.

VanAssche:1988:IPF

[Van88] Pieter Van Assche. Ignored priorities: First fission fragment (1925) and first mention of fission (1934). Nuclear Europe, 6–7 (??):24–25, 1988. CODEN NUEUDS. ISSN 0254-8348.

Vandervert:1996:ISA

[Van96] L. R. Vandervert. From idiots savants to Albert Einstein: A brain-algorithmic explanation of savant and everyday performance. New Ideas in Psychology, 14(1):81–92, March 1996. ISSN 0732-118X. See commentary [Spi96].

VanFlandern:1998:DMM

[Van98] Tom Van Flandern. Dark matter, missing planets, and new comets: paradoxes resolved, origins illuminated. North Atlantic Books, Berkeley, CA, USA, revised edition, 1998. ISBN 1-55643-268-2. xxxvii + 514 pp. LCCN QB43.2.

VanRiper:2002:SPC

[Van02] A. Bowdoin Van Riper. Science in popular culture: a reference guide. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2002. ISBN 0-313-31822-0. xvii + 314 pp. LCCN Q172.5.P65 V36 2002.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2

V36 2003. URL http://www.loc.gov/catdir/toc/fy045/
2002043646.html.

VanderKloot:2004:AFF

[Van04] W. Van der Kloot. April 1915: Five future Nobel Prize-winners inaugurate weapons of mass destruction and the academic– industrial–military complex. Notes and Records of the Royal Society of London, 58(2):149–160, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142047.

vanDelft:2006:AEL

[van06] D. van Delft. Albert Einstein in Leiden. Physics Today, 59(4): 57–62, April 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

VanDongen:2007:ERA

[Van07a] Jeroen Van Dongen. Emil Rupp, Albert Einstein, and the canal ray experiments on wave-particle duality: Scientific fraud and theoretical bias. Historical Studies in the Physical and Biological Sciences, 37(S1):73–120, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://

www.jstor.org/stable/10.1525.

VanDongen:2007:IER

[Van07b] Jeroen Van Dongen. The interpretation of the Einstein– Rupp experiments and their influence on the history of quantum mechanics. Historical Studies in the Physical and Biological Sciences, 37(S1):121–131, September 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/10.
1525/hsps.2007.37.S.121.pdf.

VanDongen:2007:REF

[Van07c] Jeroen Van Dongen. Reactionaries and Einstein’s fame: “German scientists for the preservation of pure science,” Relativity, and the Bad Nauheim Meeting. Physics in Perspective (PIP), 9(2):212–230, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0318-y;   http://www.
springerlink.com/content/l6518246k7367437/fulltext.
pdf.

Vankov:2010:GRP

[Van10] Anatoli Andrei Vankov. General Relativity problem of Mercury’s perihelion advance revisited. arxiv.org, pages 1–46, Au-

gust 10, 2010. URL http://arxiv.org/abs/1008.1811.

Vankov:2011:EPE

[Van11] Anatoli Andrei Vankov. Einstein’s paper: “Explanation of the Perihelion Motion of Mercury from General Relativity The-

ory”, January 7, 2011. URL http://www.gsjournal.net/old/
eeuro/vankov.pdf.

vanDongen:2012:MIM

[van12] Jeroen van Dongen. Mistaken identity and mirror images: Albert and Carl Einstein, Leiden and Berlin, Relativity and revolution. Physics in Perspective (PIP), 14(2):126–177, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00016-012-0084-y.pdf;   http://
www.springerlink.com/content/w55417550px35577/fulltext.
pdf.

VanTiggelen:2022:BRN

[Van22] Brigitte Van Tiggelen. Book review: A Nobel Affair: The Correspondence Between Alfred Nobel and Sofie Hess. Edited and translated by Erika Rummel. Pp. 304, illus., index. University of Toronto Press: Toronto. 2017. £62.00. ISBN: 978-1-48-750177-8. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):336–337, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00026980.2022.2066253.

vonBrill:1912:RET

[vB12] Alexander von Brill. Das Relativitätsprinzip: eine Einführung in die Theorie. (German) [The Principle of Relativity: an Introduction to the Theory]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1912. 28 pp. LCCN ????

vonBrill:1914:RET

[vB14] Alexander von Brill. Das Relativitätsprinzip: eine Einführung in die Theorie. (German) [The Principle of Relativity: an Introduction to the Theory]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, second edition, 1914. 33 pp. LCCN ????

vonBrill:1918:RET

[vB18] Alexander von Brill. Das Relativitätsprinzip: eine Einführung in die Theorie. (German) [The Principle of Relativity: an Introduction to the Theory], volume 3 of Abhandlungen und Vorträge aus dem Gebiete der Mathematik, Naturwissenschaft und Technik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, third edition, 1918. 48 pp. LCCN ????

vonBrunn:1919:HEB

[vB19] A. von Brunn. Zu Hrn. Einsteins Bemerkung über die unregelmäßigen Schwankungen der Mondlänge von der genährten Periode des Umlaufs der Mondknoten. (German) [On Mr. Einstein’s remark about the irregular fluctuations of the length of the lunar period of the orbit of the Moon nodes]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, 38(??):710–711, July 24, 1919. CODEN SPWPAI. ISSN 0371-2435. See remark [Ein19c].

vonBrill:1920:RET

[vB20] Alexander von Brill. Das Relativitätsprinzip: eine Einführung in die Theorie. (German) [The Principle of Relativity: an Introduction to the Theory]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fourth edition, 1920. 43 pp. LCCN ????

vonBrunn:1973:AAE

[vB73] A. von Brunn. 100 authors against Einstein. In Weber [Web73], page ?? ISBN 0-85498-027-X. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Villa:1996:UIG

[VB96] Brian L. Villa and John Bonnett. Understanding indignation: Gar Alperovitz, Robert Maddox, and the decision to use the atomic bomb: Book reviews: The Decision to Use the Atomic Bomb, and the Architecture of an American Myth, by Gar Alperovitz [and others], and Weapons for Victory: The Hiroshima Decision Fifty Years Later, by Robert James Maddox. Reviews in American History, 24(3):529–536, September 1996. CODEN ???? ISSN 0048-7511 (print), 1080-6628 (electronic). URL

http://www.jstor.org/stable/30030699.

vonBaeyer:2001:BRD

[vB01] Hans Christian von Baeyer. Book review: Dennis Overbye, Einstein in Love: A Scientific Romance. New York: Viking, 2000, xv + 416 pages. $27.95 (cloth). Physics in Perspective (PIP), 3(3):375–376, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2002:BRG

[vB02] Hans Christian von Baeyer. Book review: Gregg Herken, Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI, revised and expanded edition. Stanford: Stanford University Press, 2000, xv + 358 pages. $22.95 (paper). Physics in Perspective (PIP), 4(1):119–120, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBoehm:2005:EWW

[vB05] Gero von Boehm. E = mc^{2} wer war Albert Einstein? (German) [E = mc^{2}: who was Albert Einstein?]. Collection Rolf Heyne, München, Germany, 2005. ISBN 3-89910-251-7. 175 pp. LCCN QC16.E5 B64 2005.

vomBrocke:1996:KWM

[vBL96a] Bernhard vom Brocke and Hubert Laitko, editors. Die Kaiser-Wilhelm-/Max-Planck-Gesellschaft und ihre Institute: Studien zu ihrer Geschichte. Walter de Gruyter, Berlin, Germany, 1996. ISBN 3-11-015483-8 (vol. 1). xiii + 672 pp. LCCN

Q180.G42 K35 1996. URL http://www.degruyter.com/
viewbooktoc/product/146831;   http://www.reference-
global.com/doi/book/10.1515/9783110802443.

vomBrocke:1996:KWS

[vBL96b] Bernhard vom Brocke and Hubert Laitko, editors. Die Kaiser-Wilhelm-/Max-Planck-Gesellschaft und ihre Institute: Studien zu ihrer Geschichte. (German) [The Kaiser Wilhelm/Max Planck Society and their institutions: studies on their history]. Walter de Gruyter, Berlin, Germany, 1996. ISBN 3-11-015483-8 (vol. 1). xiii + 672 pp. LCCN

Q180.G42 K35 1996. URL http://www.degruyter.com/
viewbooktoc/product/146831;   http://www.reference-
global.com/doi/book/10.1515/9783110802443.

Rucker:1977:GRF

[vBRvBR77] Rudy v. B. (Rudy von Bitter) Rucker. Geometry, Relativity, and the Fourth Dimension. Dover, New York, NY, USA, 1977. ISBN 0-486-23400-2. 133 pp. LCCN QA699

.R8 1977. URL http://www.loc.gov/catdir/description/
dover031/76022240.html;   http://www.loc.gov/catdir/
enhancements/fy1318/76022240-t.html.

Borzeszkowski:1982:QTE

[vBT82] H.-H. v. Borzeszkowski and H.-J. Treder. Quantum theory and Einstein’s General Relativity. Foundations of Physics, 12(11): 1113–1129, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01300549.

Borzeszkowski:1994:EEF

[vBT94] H.-H. v. Borzeszkowski and H.-J. Treder. Einstein equations and Fierz–Pauli equations with self-interaction in quantum gravity. Foundations of Physics, 24(6):949–962, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02067656.

vonBorzeszkowski:1996:MED

[vBT96] H.-H. von Borzeszkowski and H.-J. Treder. Mach–Einstein doctrine and general relativity. Foundations of Physics, 26 (7):929–942, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02148835.

vanCalmthout:2005:ELLd

[vC05a] Martijn van Calmthout. Einsteins licht: een leven met relativiteit. (Dutch) [Einstein’s light: a life with relativity]. Contact, Amsterdam, The Netherlands, 2005. ISBN 90-254-2841-X. 204 pp. LCCN QC16.E5 C35 2005.

vanCalmthout:2005:MRO

[vC05b] Martijn van Calmthout. Moet die rooie onruststoker Einstein hier prof worden? (Dutch) [should the redheaded troublemaker Einstein be a professor here?]. De Volkskrant, ??(??):35, May 14, 2005.

vanCalmthout:2016:SGJ

[vC16] Martijn van Calmthout. Sam Goudsmit: de jacht op de atoombom van Hitler. Meulenhoff, Amsterdam, The Netherlands, 2016. ISBN 90-290-8958-X (paperback), 94-023-0744-3 (e-book). 284 + 16 pp. LCCN DS135.N6 G656 2016.

vanCalmthout:2018:SGH

[vCH18] Martijn van Calmthout and Michiel Horn, editors. Sam Goudsmit and the hunt for Hitler’s atom bomb. Prometheus Books, Amherst, NY, USA, 2018. ISBN 1-63388-450-3 (hardcover), 1-63388-451-1 (e-book). 243 + 14 pp. LCCN DS135.N6 G65613 2018.

vanCalmthout:2017:RQSc

[vCK17] Martijn van Calmthout and Leo P. Kouwenhoven. Real Quanta: Simplifying Quantum Physics for Einstein and Bohr. Dundurn Press Limited, Toronto, ON, Canada, 2017. ISBN 1-4597-4049-1 (paperback), 1-4597-4050-5 (PDF), 1-4597-4051-3 (epub). 192 pp. LCCN QC174.123 .C3513 2017.

vanDongen:2002:EKK

[vD02a] Jeroen van Dongen. Einstein and the Kaluza–Klein particle. Studies in History and Philosophy of Modern Physics, 33(2): 185–210, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219802000175.

vanDongen:2002:EUG

[vD02b] Jeroen van Dongen. Einstein’s Unification: General Relativity and the Quest for Mathematical Naturalness. Ph.D. dissertation, University of Amsterdam, Amsterdam, The Netherlands, 2002.

vanDongen:2004:EMS

[vD04] Jeroen van Dongen. Einstein’s methodology, semivectors and the unification of electrons and protons. Archive for History of Exact Sciences, 58(3):219–254, March 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=3&spage=219.

vanDongen:2009:RMM

[vD09] Jeroen van Dongen. On the role of the Michelson–Morley Experiment: Einstein in Chicago. Archive for History of Exact Sciences, 63(6):655–663, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/content/86387g62p7jukj80/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=6&spage=655.

vanDongen:2010:EU

[vD10a] Jeroen van Dongen. Einstein’s Unification. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88346-6. x + 213 pp. LCCN QC173.6 .D66 2010. URL http://assets.

cambridge.org/97805218/83467/cover/9780521883467.jpg;
 http://www.loc.gov/catdir/enhancements/fy1009/2010017820-
b.html;   http://www.loc.gov/catdir/enhancements/fy1009/
2010017820-d.html;   http://www.loc.gov/catdir/enhancements/
fy1009/2010017820-t.html.

vanDongen:2010:EOO

[vD10b] Jeroen van Dongen. On Einstein’s opponents, and other crackpots. Studies in History and Philosophy of Modern Physics, 41 (1):78–80, January 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219809000598.

vanDelft:2014:PEF

[vD14] Dirk van Delft. Paul Ehrenfest’s final years. Physics Today, 67 (1):41–47, January 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

vanDongen:2015:ERA

[vD15] Jeroen van Dongen. Essay review: Albert Einstein and the History and Philosophy of Science, The Cambridge Companion to Einstein, edited by Michel Janssen and Christoph Lehner. Isis, 106(3):684–689, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/683458.

vanDongen:2017:EVV

[vD17] Jeroen van Dongen. The epistemic virtues of the virtuous theorist: On Albert Einstein and his autobiography. In van Dongen and Paul [vDP17], chapter 5, pages 63–77. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https://link.

springer.com/chapter/10.1007/978-3-319-48893-6_5.

vanDongen:2021:STE

[vD21] Jeroen van Dongen. String theory, Einstein, and the identity of physics: Theory assessment in absence of the empirical. Studies in History and Philosophy of Science Part A, 89(??):164– 176, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0039368121001023.

vandenBrink:1995:ED

[vdB95] Michaël W. J. van den Brink. Einstein’s debt. Nature, 375 (6529):272, May 25, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v375/n6529/pdf/375272c0.pdf.

vanDongen:2017:EVS

[vDP17] Jeroen van Dongen and Herman Paul, editors. Epistemic Virtues in the Sciences and the Humanities, volume 321 of Boston Studies in the Philosophy and History of Science. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346. vi + 198 pp. LCCN Q174 .B67 v. 321; Q175.32.V57

E65 2017. URL https://link.springer.com/book/10.1007/
978-3-319-48893-6.

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland, Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:1968:SQM

[vdW68] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics, volume 5 of Classics of science. Dover, New York, NY, USA, 1968. ISBN 0-486-61881-1. xi + 430 pp. LCCN QC174.1 .W3 1968.

vanderWaerden:2007:SQM

[vdW07] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics. Dover, New York, NY, USA, 2007. ISBN 0-486-45892-X (paperback). xi + 430 pp. LCCN

QC174.12 .W34 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006050791-d.html.

vonEotvos:1896:UGE

[vE96] R. von Eötvös. Untersuchungen über Gravitation und Erdmagnetismus. (German) [Investigations about gravity and geomagnetism]. Wiedemann Ann., 59:354–400, 1896. URL http://

www.kfki.hu/eotvos/onehund.html.

vonEotvos:1910:BBD

[vE10] R. von Eötvös. Bericht über Beobachtungen mit der Drehwaage. (German) [Report about observations made with the torsion balance]. In ????, editor, Verhandlungen der vom 21. Bis 29. September 1909 in London und Cambridge abgehaltenen sechzenten allgemeinen Konferenz der Internationalen Erdmessung. (German) [Proceedings of the 16th General Conference of International Geodesy, London–Cambridge, 21–29 September 1909], volume 16 (??), pages 319–350. G. Reimer, Berlin,

Germany, 1910. URL http://www.kfki.hu/eotvos/onehund.

html. See [AHH05] for discussion of recent results with the torsion balance.

vanErkelens:1995:SVW

[vE95] Herbert van Erkelens. Het spel van de wijsheid: Pauli, Jung en de menswording van God. (Dutch) [The game of wisdom: Pauli, Jung and the gestation (??) of God]. Kok Agora, Kampen, The Netherlands, 1995. ISBN 90-391-0668-1. 182 pp. LCCN BL245 .E75 1995.

Veale:2001:DMA

[Vea01] S. Veale. Driving Mr. Albert: A trip across America with Einstein’s brain. New York Times Book Review, ??(??):28, June 17, 2001. ISSN 0028-7806.

vanErkelens:2002:WPG

[vEA02] Herbert van Erkelens and Thomas Arzt. Wolfgang Pauli und der Geist der Materie. (German) [Wolfgang Pauli and the spirit of matter], volume 7 of Studien aus der existential-psychologischen Bildungs- und Begegnungsstätte Todtmoos-Rütte. Königshausen and Neumann, Würzburg, Germany, 2002. ISBN 3-8260-2222-X. 278 pp. LCCN QC16.P37 E75 2002.

Vecchiato:2017:VAG

[Vec17] Alberto Vecchiato. Variational Approach to Gravity Field Theories: from Newton to Einstein and beyond. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-51209-9, 3-319-51211-0 (e-book). 300 pp. LCCN ????

Veltman:2000:NLW

[Vel00] Martinus J. G. Veltman. Nobel Lecture: From weak interactions to gravitation. Reviews of Modern Physics, 72(2):341– 349, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.341;   http://rmp.aps.
org/abstract/RMP/v72/i2/p341_1.

Veltman:2003:FME

[Vel03] Martinus Veltman. Facts and mysteries in elementary particle physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-148-1, 981-238-149-X (paperback). viii + 340 pp. LCCN QC793.2 .V45 2003.

Venkataraman:1988:JLL

[Ven88] G. (Ganesan) Venkataraman. Journey into light: life and science of C. V. Raman. Published by Indian Academy of Sciences in co-operation with Indian National Science Academy, Bangalore, India, 1988. ISBN 81-85324-00-X. xviii + 570 + 24 pp. LCCN QC16.R27 V46 1988. URL http://www.loc.gov/

catdir/enhancements/fy0639/89903531-d.html.

Venkataraman:1994:QRB

[Ven94a] G. Venkataraman. Quantum revolution I: The breakthrough. Vignettes in physics. Sangam Books, London, UK, 1994. ISBN 0-86311-452-0. ix + 189 pp. LCCN ????

Venkataraman:1994:QRI

[Ven94b] G. Venkataraman. Quantum revolution. II. QED: the jewel of physics. Sangam, London, UK, 1994. ISBN 0-86311-453-9 (paperback). viii + 136 pp. LCCN QC680 .V46 1994.

Venkataraman:1994:QRP

[Ven94c] Ganesan Venkataraman. Quantum revolution. III. What is reality? Vignettes in physics. Sangam Books, London, UK, 1994. ISBN 0-86311-454-7 (paperback). ix + 128 pp. LCCN ????

Veneziano:2012:MBT

[Ven12] Gabriele Veneziano. The myth of the beginning of time. Scientific American, 306(2):78–89, February 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v21/n1s/full/scientificamericantime0212-78.
html.

vonEotvos:1909:BGP

[vEPF09] R. von Eötvös, D. Pekár, and E. Fekete. Beiträge zum Gesetz der Proportionalität von Trägheit and Gravität. (German) [Contributions to the law of proportionality of inertia and gravity]. With motto “Ars longa, vita brevis”. Submitted to the Beneke Foundation in Göttingen, but the original text has been

lost., 1909. URL http://www.kfki.hu/eotvos/onehund.html.

vonEotvos:1922:BGP

[vEPF22] R. von Eötvös, D. Pekár, and E. Fekete. Beiträge zur Gesetze der Proportionalität von Trägheit und Gravität. (German) [Contributions to the law of proportionality of inertia and gravity]. Annalen der Physik (Leipzig), 68(??):11–16, ????

1922. URL http://www.kfki.hu/eotvos/onehund.html. En-

glish translation for the U. S. Department of Energy by J. Achzenter, M. Bickeböller, K. Bräuer, P. Buck, E. Fischbach, G. Lubeck, C. Talmadge, University of Washington preprint 40048-13-N6. More complete English text reprinted earlier in Annales Universitatis Scientiarium Budapestiensis de Rolando Eötvös Nominate, Sectio Geologica, 7, 111, 1963. See also [NHG89].

VergaraCaffarelli:2001:EFI

[Ver01] Roberto Vergara Caffarelli, editor. Enrico Fermi: immagini e documenti. (Italian) [Enrico Fermi: images and documents]. Limonaia, Pisa, Italia, 2001. ISBN 88-8492-092-2. 111 pp. LCCN QC16.F46 E56 2001.

Vesaas:1982:BAH

[Ves82] Halldis Moren Vesaas, editor. Barn av Hiroshima. Samlaget, Oslo, Norway, second edition, 1982. ISBN 82-521-2079-2. 112 pp. LCCN ????

vonFoerster:1979:ELP

[vF79] Thomas von Foerster. Einstein’s life in pictures. Physics Today, 32(3):26–33, March 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Vavilov:1948:OLD

[VFIS48] Sergei Vavilov, A. N. Frumkin, A. F. Ioffe, and N. N. Semyonov. Open letter to Dr. Einstein — from four Soviet scientists. Bulletin of the Atomic Scientists, 4(2):34, 37–38, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Ein48m]. Reprinted in [VFIS63].

Vavilov:1963:AES

[VFIS63] Sergei Vavilov, A. N. Frumkin, A. F. Ioffe, and N. N. Semyonov. On Albert Einstein’s support of world government. In Grodzins and Rabinowitch [GR63], pages 125–129. LCCN D842 .B78. See reply [Ein63b]. Reprint of [VFIS48].

vonGleich:1930:ERP

[vG30] Gerald von Gleich. Einsteins Relativitätstheorie und physikalische Wirklichkeit. (German) [Einstein’s Theory of Relativity and physical reality]. Johann Ambrosius Barth, Leipzig, Germany, 1930. viii + 142 pp. LCCN ????

vonHevesy:1962:ARR

[vH62] Georg von Hevesy. Adventures in radioisotope research: the collected papers of George Hevesy in two volumes. Pergamon, New York, NY, USA, 1962. vii + 1–515 (vol. 1), vii + 517–1047 (vol. 2) pp. LCCN QC 796.I7; QH324 .H43.

vonHumboldt:1997:CSP

[vH97] Alexander von Humboldt. Cosmos: a sketch of a physical description of the universe. Foundations of natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1997. ISBN 0-8018-5502-0 (vol. 1: paperback), 0-8018-5503-9 (vol. 2: paperback). 375 (vol. 1), 367 (vol. 2) pp. LCCN QB981 .H8613 1997. URL http://www.loc.gov/

catdir/bios/jhu051/96036421.html;   http://www.loc.
gov/catdir/description/jhu052/96036421.html.

vonHalban:1939:ENL

[vHJK39a] Hans von Halban, Frédéric Joliot, and Lew Kowarski. Energy of neutrons liberated in the nuclear fission of uranium induced by thermal neutrons. Nature, 143(3631):939, June 3, 1939. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v143/n3631/
pdf/143939a0.pdf.

vonHalban:1939:LNN

[vHJK39b] Hans von Halban, Frédéric Joliot, and Lew Kowarski. Liberation of neutrons in the nuclear explosion of uranium. Nature, 143 (3620):470–471, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v143/n3620/pdf/143470a0.pdf.

vonHalban:1939:NNL

[vHJK39c] Hans von Halban, Frédéric Joliot, and Lew Kowarski. Number of neutrons liberated in the nuclear fission of uranium. Nature, 143 (3625):680, April 22, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v143/n3625/pdf/143680a0.pdf. According

to [BHT86, page 215], this paper ‘galvanized the German [nuclear] scientists into action’, because it demonstrated that neutrons were produced in sufficient numbers in uranium fission to make a chain reaction possible. See also its two companions [vHJK39b, vHJK39a].

Ignatowsky:1910:RGP

[vI10a] W. v. Ignatowsky. Das Relativitätsprinzip. (German) [The Principle of Relativity]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1910. 40 pp. LCCN ????

Ignatowsky:1910:SKR

[vI10b] W. v. Ignatowsky. Der starre Körper und das Relativitätsprinzip. (German) [The rigid body and the Principle of Relativity]. Annalen der Physik (1900) (series 4), 33(??):607– 630, ???? 1910. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Ignatowsky:1911:BAS

[vI11a] W. v. Ignatowsky. Bemerkung zu der Arbeit: “Der starre Körper und das Relativitätsprinzip”. (German) [Remark on my work: The rigid body and the Principle of Relativity]. Annalen der Physik (1900) (series 4), 34(??):373–375, ???? 1911. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Ignatowsky:1911:BMA

[vI11b] W. v. Ignatowsky. Eine Bemerkung zu meiner Arbeit: Einige allgemeine Bemerkungen zum Relativitätsprinzip. (German) [A note on my work: Some general comments on the Principle of Relativity]. Physikalische Zeitschrift, 12(18):779, September 15, 1911. CODEN PHZTAO. ISSN 0369-982X. URL http://hdl.

handle.net/2027/mdp.39015011417709?urlappend=%3Bseq=
837.

Videira:2017:NEN

[Vid17] Antonio A. P. Videira. Neither Einstein nor Bergson, but both! Science & Education (Springer), 26(3–4):425–428, May 2017.

CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Viereck:1929:IE

[Vie29] George Sylvester Viereck. Interview with A. Einstein. Saturday Evening Post, ??(??):??, October 26, 1929.

Vigeveno:1983:BESa

[Vig83a] Guido Vigeveno. The bomb and European security. Hurst, London, UK, 1983. ISBN 0-905838-89-0. xii + 131 pp. LCCN UA646 .V5 1983b. With a foreword by Eugene V. Rostow.

Vigeveno:1983:BESb

[Vig83b] Guido Vigeveno. The bomb and European security. Indiana University Press, Bloomington, IN, USA, 1983. ISBN 0-253-31208-6, 0-253-21220-0. xii + 131 pp. LCCN ???? With a foreword by Eugene V. Rostow.

Vigier:1993:DNE

[Vig93] J. P. Vigier. From Descartes and Newton to Einstein and de Broglie. Foundations of Physics, 23(1):1–4, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01883986.

Vigier:1994:PTR

[Vig94] J. P. Vigier. Possible test of the reality of superluminal phase waves and particle phase space motions in the Einstein–de Broglie–Bohm causal stochastic interpretation of quantum mechanics. Foundations of Physics, 24(1):61–83, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02053908.

Vigier:1995:DIF

[Vig95] Jean-Pierre Vigier. Derivation of inertial forces from the Einstein–de Broglie–Bohm (E.d.B.B.) causal stochastic interpretation of quantum mechanics. Foundations of Physics, 25(10):1461–1494, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02057462.

Vilenkin:2006:MWO

[Vil06] A. (Alexander) Vilenkin. Many worlds in one: the search for other universes. Hill and Wang (a division of Farrar, Straus and Giroux), New York, NY, USA, 2006. ISBN 0-8090-9523-8 (hardcover). viii + 235 pp. LCCN

QB981 .V526 2006. URL http://www.loc.gov/catdir/
enhancements/fy0623/2005027057-b.html;   http://www.
loc.gov/catdir/enhancements/fy0623/2005027057-d.html;
 http://www.loc.gov/catdir/enhancements/fy0701/2005027057-
s.html;   http://www.loc.gov/catdir/toc/ecip0519/2005027057.
html.

Virilio:1992:IME

[Vir92] Paul Virilio. “Das irreale Monument”: [der Einstein-Turm]. (German) [“The Unreal Monument”: [The Einstein Tower]], volume 165 of Internationaler Merve-Diskurs. Merve, Berlin, Germany, 1992. ISBN 3-88396-091-8. 49 pp. LCCN ????

Vista:1994:BRL

[Vis94] Michele Vista. Book review: Laylin K. James (ed.), Nobel Laureates in Chemistry, 1901–1992. Washington, American Chemical Society and the Chemical Heritage Foundation, 1993, xviii + 798 pp. (History of Modern Chemical Sciences). Nuncius, 9(2):940–942, ???? 1994. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01747.

Vista:1995:BRD

[Vis95] Michele Vista. Book review: Dietrich Stoltzenberg, Fritz Haber: Chemiker, Nobelpreisträger, Deutscher Jude, Weinheim, VCH, 1994, xiv + 670 pp., ill. Nuncius, 10(1):410–413, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00440.

Vitouch:2015:EZZ

[Vit15] Anatol Vitouch. Einstein in Zürich: Zwölf kurze Geschichten. (German) [Einstein in Zürich: Twelve short stories]. Labor edition, Wien, Austria, 2015. ISBN 3-902800-14-3. 176 pp.

LCCN ???? URL http://d-nb.info/1064106161/04.

Vizgin:1979:AEP

[Viz79] V. P. Vizgin. Albert Einstein and the problem of constructing a scientific theory. Voprosy Filosofii, ??(10):56–64, ???? 1979. ISSN 0042-8744.

vonKluber:1960:DEL

[vK60] H. von Klüber. The determination of Einstein’s light-deflection in the gravitational field of the sun. Vistas in Astronomy, 3 (1):47–77, January 1960. CODEN VASTA6. ISSN 0083-6656 (print), 1872-9207 (electronic).

vonKluber:1965:EFF

[vK65a] Harald von Klüber. Erwin Finlay-Freundlich [1885–1964]. Astronomische Nachrichten, 288(??):281–286, December 1965. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1965AN....
288..281V.

vonKluber:1965:ONE

[vK65b] Harald von Klüber. Obituary notices: Erwin Finlay-Freundlich [1885–1964]. Quarterly Journal of the Royal Astronomical Society, 6(??):82–84, March 1965. CODEN QJRAAK. ISSN 0035-

8738. URL http://adsabs.harvard.edu/abs/1965QJRAS...
6...82..

vonKarman:1949:CST

[vKL49] Theodore von Kármán and C. C. Lin. On the concept of similiarity in the theory of isotropic turbulence. Reviews of Modern Physics, 21(3):516–519, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.516;
http://rmp.aps.org/abstract/RMP/v21/i3/p516_1. Special

issue in honor of Albert Einstein on his 70th birthday.

vonLaue:1909:RGP

[vL09] Max Theodor Felix von Laue. Das Relativitätsprinzip. (German) [The Principle of Relativity]. Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1909. 208 pp. LCCN ???? According to [Del11a, 265], this is believed to be the first textbook on Special Relativity.

vonLaue:1911:RGP

[vL11] Max Theodor Felix von Laue. Das Relativitätsprinzip. (German) [The Principle of Relativity], volume 38 of Die Wissenschaft: Sammlung naturwissenschaftlicher und mathematischer Monographien. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1911. x + 208 pp. LCCN ????

vonLaue:1913:RGP

[vL13] Max Theodor Felix von Laue. Das Relativitätsprinzip. (German) [The Principle of Relativity], volume 38 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1913. xii + 272 pp. LCCN ????

vonLaue:1919:RBR

[vL19] Max Theodor Felix von Laue. Das Relativitätstheorie. Band 1. Das Relativitätsprinzip der Lorentztransformation. (German) [The Theory of Relativity. Volume 1. The Principle of Relativity of the Lorentz Transformation]. Friedrich Vieweg und Sohn, Braunschweig, Germany, third edition, 1919. xii + 272 pp. LCCN ????

vonLaue:1921:RBR

[vL21] Max von Laue. Die Relativitätstheorie. Band 1. Das Relativitätsprinzip der Lorentztransformation. (German) [The Theory of Relativity. Volume 1. The Principle of Relativity of the Lorentz Transformation]. Friedrich Vieweg und Sohn, Braunschweig, Germany, fourth edition, 1921. ???? pp. LCCN ????

vonLaue:1923:RBA

[vL23] Max von Laue. Die Relativitätstheorie. Band 2. Die Allgemeine Relativitätstheorie und Einsteins Lehre von der Schwerkraft. (German) [The theory of relativity. Volume 2. The General Theory of Relativity and Einstein’s Theory of Gravity]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1923. ???? pp. LCCN ????

vonLaue:1949:AET

[vL49a] M. von Laue. Zu Albert Einsteins 70-tem Geburtstag. Reviews of Modern Physics, 21(3):348–349, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.348;   http://rmp.aps.org/abstract/RMP/v21/i3/p348_

1. Special issue in honor of Albert Einstein on his 70th birthday.

vonLaue:1949:IE

[vL49b] Max von Laue. Inertia and energy. In Schilpp [Sch49c], chapter 19, pages 501–533. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

vonLaue:1952:RBR

[vL52] Max von Laue. Die Relativitätstheorie. Band 1. Das Relativitätsprinzip der Lorentztransformation. (German) [The Theory of Relativity. Volume 1. The Principle of Relativity of the Lorentz Transformation]. Friedrich Vieweg und Sohn, Braunschweig, Germany, fifth edition, 1952. ???? pp. LCCN ????

vonLaue:1955:AE

[vL55a] Max von Laue. Albert Einstein. Frankfurter Allgemeine Zeitung, ??(??):??, April 23, 1955.

vonLaue:1955:AEJ

[vL55b] Max von Laue. Albert Einstein: 50 Jahre Relativitätstheorie und Lichtquanten (Eröffnungsansprache). (German) [Albert Einstein: 50 years of relativity and light quanta (opening address)]. Physikalische Blätter, 11(??):228–230, 1955. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

vonLaue:1955:AES

[vL55c] Max von Laue. Albert Einstein, der Schöpfer der Relativitätstheorie und Entdecker der Lichtquanten. (German) [Albert Einstein, the creator of the theory of relativity and discoverer of light quanta]. In Forscher und Wissenschaftler im heutigen Europa (und Erde). (German) [Researchers and Scientists in Europe (and Earth) Today], pages 47–55. Stalling Verlag, Oldenburg, West Germany, 1955.

vonLaue:1955:RBR

[vL56a] Max von Laue. Die Relativitätstheorie. Band 1. Das Relativitätsprinzip der Lorentztransformation. (German) [The Theory of Relativity. Volume 1. The Principle of Relativity of the Lorentz Transformation], volume 68 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, fourth edition, 1955–1956. ???? pp. LCCN ????

vonLaue:1956:ER

[vL56b] Max von Laue. Einstein und die Relativitätstheorie. Naturwissenschaften, 43(??):1–8, ???? 1956. Vortrag gehalten am 23 September 1955 auf dem Deutschen Physikertag in Wiesbaden.

vonLaue:1956:KBEa

[vL56c] Max von Laue. Von Kopernikus bis Einstein. Jahrbuch der Max Planck Gesellschaft, ??(??):150–172, 1956.

vonLaue:1956:KBEb

[vL56d] Max von Laue. Von Kopernikus bis Einstein. Naturwissenschaftliche Rundschau, 3(??):83–89, 1956. CODEN NARSAC. ISSN 0028-1050.

vonLaue:1957:AE

[vL57] Max von Laue. Albert Einstein. In Die grossen Deutschen, volume IV, pages 386–387. Propyläenverlag, ????, 1957.

vonLaue:1959:AE

[vL59] Max von Laue. Albert Einstein. Neue Deutsche Biographie, 4 (??):404–408, ???? 1959.

vonLaue:1965:RBR

[vL65] Max von Laue. Die Relativitätstheorie. Band 1. Das Relativitätsprinzip der Lorentztransformation. (German) [The Theory of Relativity. Volume 1. The Principle of Relativity of the Lorentz Transformation], volume 68 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, fourth edition, 1965. ???? pp. LCCN ????

Vlachy:1979:NP

[Vla79] J. Vlachý. Nobel Prizes. Scientometrics, 1(3):295–301, March 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016314.

Einstein:1990:AER

[vM90] Karl von Meyenn. Albert Einsteins Relativitätstheorie: die grundlegenden Arbeiten. (German) [Albert Einstein’s Theory of Relativity: the basic work]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1990. ISBN 3-528-06336-X, 3-322-83770-X. vi + 331 pp. LCCN QC173.55 .E36 1990. URL http://link.

springer.com/book/10.1007/978-3-322-83770-7.

vonMeyenn:1993:WPW

[vM93] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band III: 1940–1949. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume III: 1940–1949, volume 11 of Sources in the history of mathematics and physical sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54911-0. ISSN 0172-6315. lxiv + 1070

pp. LCCN QC16.P37 A34. URL http://www.springer.com/
physics/book/978-3-540-54911-6.

vonMeyenn:1996:WPW

[vM96] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil I: 1950–1952. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part I: 1950–1952, volume 14 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-59442-6. ISSN 0172-6315. xxxvii + 968 pp. LCCN QC16.P37 A34. URL http://

www.springerlink.com/content/978-3-540-78803-4.

vonMettenheim:1998:PVE

[vM98] Christoph von Mettenheim. Popper versus Einstein: on the philosophical foundations of physics. Mohr Siebeck, Tübingen, Germany, 1998. ISBN 3-16-146910-0. 238 pp. LCCN MLCM

99/02554 (Q). URL http://www.zentralblattmath.org/
zmath/en/search/?an=0949.83005.

Vargish:1999:IMR

[VM99a] Thomas Vargish and Delo E. Mook. Inside modernism: Relativity theory, cubism, narrative. Yale University Press, New Haven, CT, USA, 1999. ISBN 0-300-07613-4 (hardcover). xii + 185 + 16 pp. LCCN BH301.M54 V37 1999. See [van12] for a discussion of Einstein’s influence on modern art.

vonMeyenn:1999:WPW

[vM99b] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil II: 1953–1954. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part II: 1953–1954, volume 15 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64312-5. ISSN 0172-6315. xxxv + 1100 pp. URL http://www.springer.com/

physics/book/978-3-540-64312-8.

vanMeijgaard:2000:WPC

[vM00] Harry van Meijgaard. Wolfgang Pauli centennial 1900–2000: an extension of the dissertation 1998. Emmy Memorial Center, Oldenzaal, The Netherlands, 2000. ISBN 90-805493-1-2. 184 pp. LCCN QC16.P37 M45 2000.

vonMeyenn:2001:WPW

[vM01] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil III: 1955–1956 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part III: 1955–1956, volume 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. lxv + 994 pp. LCCN QC16.P37 A34.

Vane:2005:NML

[VM05a] Howard R. Vane and Chris Mulhearn. The Nobel Memorial laureates in economics: an introduction to their careers and main published works. Edward Elgar, Cheltenham, UK, 2005. ISBN 1-84376-600-0. xii + 362 pp. LCCN HB76 .V36 2005.

vonMettenheim:2005:AEI

[vM05b] Christoph von Mettenheim. Albert Einstein oder Der Irrtum eines Jahrhunderts. (German) [Albert Einstein or the mistake of a century]. ????, ????, 2005. ISBN ???? 217 pp.

LCCN ???? URL http://www.christoph.mettenheim.de/
eap-o5web2.pdf.

vonMeyenn:2005:WPW

[vM05c] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil IV: 1957–1958 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part IV: A:1957, B:1958, volume 18 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-40296-9. ISSN 0172-6315. xl + 1585 pp. LCCN QC16.P37 A34. URL http:/

/www.springerlink.com/content/978-3-540-26832-1.

vonMettenheim:2009:AEI

[vM09] Christoph von Mettenheim. Albert Einstein oder Der Irrtum eines Jahrhunderts. (German) [Albert Einstein or the mistake of a century]. Books on Demand, Norderstedt, Germany, 2009. ISBN 3-8370-3361-9. 379 pp. LCCN ???? URL http://www.

christoph.mettenheim.de/eap-o5web2.pdf.

vonMettenheim:2012:AEI

[vM12] Christoph von Mettenheim. Albert Einstein oder Der Irrtum eines Jahrhunderts. (German) [Albert Einstein or the mistake of a century]. Books on Demand, Norderstedt, Germany, second edition, 2012. ISBN 3-8448-8976-0, 3-8448-2367-0. 365 pp. LCCN ????

vonMeyenn:1985:WPW

[vMHW85] Karl von Meyenn, Armin Hermann, and Victor F. (Victor Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band II, 1930– 1939. (German) [Wolfgang Pauli: scientific correspondence with Bohr, Einstein, Heisenberg and others. Volume II, 1930–1939], volume 6 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

vonNeumann:1937:OGV

[vN37] John von Neumann. Über ein ökonomisches Gleichungssystem und eine Verallgemeinerung des Brouwerschen Fixpunktsatzes. (German) [On an economics equation system and a generalization of the Brouwer Fixed-Point Theorem]. Erg. eines Math. Kolloqu., 8(??):73–83, ???? 1937. Reprinted in [Tau63, Paper 3]. English translation in [vN45].

vonNeumann:1945:MGE

[vN45] John von Neumann. A model of general economic equilibrium. Review of Economic Studies, 13(1):1–9, ???? 1945. CODEN ???? ISSN 0034-6527 (print), 1467-937X (electronic). English translation of German original [vN37].

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974–1975. Gift, Nicholas A. Vonneuman, 1993.

Vogel:2020:BRN

[Vog20] Manuel Vogel. Book review: New perspectives on Einstein’s E = mc^{2}, World scientific, by Y. S. Kim and M. E. Noz, Singapore, World Scientific, 2018, 200 pp., $78.00 (hardback), ISBN 978-981-3237-70-4. Scope: monograph. Level: researcher, specialist, scientist. Contemporary Physics, 61(1):62–63, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2021:BRR

[Vog21] Dr. Manuel Vogel. Book review: Relativity and cosmology: volume 5 of modern classical physics by Kip S. Thorne and Roger D. Blandford, Princeton, Princeton University Press, 2021, 416 pp., £48 (softcover), ISBN: 978-0-691-20739-1. Scope: textbook. Level: advanced undergraduates. Contemporary Physics, 62(2): 122, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Voigt:1986:ERA

[Voi86] A. Voigt. Entwurf einer nicht-relativistischen Alternative zur Einsteinschen speziellen und allgemeinen Relativitätstheorie. (German) [Draft of a non-relativistic alternative to Einstein’s special and general relativity theory]. Philosophia Naturalis, 23 (4):438–462, 1986. CODEN ???? ISSN 0031-8027.

Volkman:2002:SGW

[Vol02] Ernest Volkman. Science goes to war: the search for the ultimate weapon, from Greek fire to Star Wars. Wiley, New York, NY, USA, 2002. ISBN 0-471-41007-1. ix + 278 pp.

LCCN U27 .V65 2002. URL http://www.loc.gov/catdir/
bios/wiley044/2002284341.html;   http://www.loc.gov/
catdir/description/wiley037/2002284341.html;   http:
//www.loc.gov/catdir/toc/wiley023/2002284341.html.

vonEotvos:1890:FVK

[von90] von Eötvös Loránd. A föld vonzása különböző anyagokra. (Hungarian) [The Earth’s pull of different materials]. Akadém iai Értesitő, 2(??):108–110, February 1890. Read at a meeting of the Academy on 20 January 1889. See concurrent German translation [Eöt91] and later English translation [Eöt08].

Vitale:2006:GAE

[VR06] Antonio Vitale and Carlo Rubbia, editors. Da Galileo ad Einstein. (German) [From Galileo to Einstein]. Progetto Leonardo; Collana “Marche scienza”. Esculapio, Bologna, Italia, 2006. ISBN 88-7488-127-4. xvi + 224 pp. LCCN ????

Vrana:1957:AE

[Vrá57] Jiří Vrána. Albert Einstein. Pokroky matematiky, fysiky a astronomie, 2(??):320–333, ???? 1957.

vonSmoluchowski:1906:KTB

[vS06] Marian von Smoluchowski. Zur kinetischen Theorie der Brownschen Molekularbewegung und der Suspensionen. (German) [On the kinetic theory of Brownian motion and suspensions]. Annalen der Physik (1900) (series 4), 21(14): 756–780, March 15, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

en.wikipedia.org/wiki/Marian_Smoluchowski;   http://
www.physik.uni-augsburg.de/annalen/history/historic-
papers/1906_326_756-780.pdf.

Soldner:1921:ALS

[vS21a] J. (Johann Georg von) Soldner. Über die Ablenkung eines Lichtstrahls von seiner geradlinigen Bewegung durch die Attraktion eines Weltkörpers, an welchem er nähe vorbeigeht; von J. Soldner, 1801. (German) [On the deviation of a light ray from its straight-line motion due to the attraction of a heavenly body that it passes by closely]. Annalen der Physik (1900) (series 4), 65(15):593–604, 1921. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL http://en.wikipedia.org/wiki/Johann_Georg_von_
Soldner;   http://en.wikisource.org/?curid=755966;
http://onlinelibrary.wiley.com/doi/10.1002/andp.19213701503.

With a foreword, and considerable editing and abridging, by Philipp Lenard. An English translation of the complete original 1801 article is available in Wikisource.

Soldner:1921:BAL

[vS21b] J. (Johann Georg von) Soldner. Über die Ablenkung eines Lichtstrahls von seiner geradlinigen Bewegung durch die Attraktion eines Weltkörpers, an welchem er nähe vorbeigeht; von J. Soldner, 1801. (German) [On the deviation of a light ray from its straight-line motion due to the attraction of a heavenly body that it passes by closely]. Annalen der Physik (1900) (series 4), 65(15):593–604, 1921. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

URL http://en.wikipedia.org/wiki/Johann_Georg_von_
Soldner;   http://en.wikisource.org/?curid=755966;
http://onlinelibrary.wiley.com/doi/10.1002/andp.19213701503.

With a foreword, and considerable editing and abridging, by Philipp Lenard. An English translation of the complete original 1801 article is available in Wikisource.

vonSuttner:1978:PNP

[vS78] Bertha von Suttner. Perspective: Nobel Peace Laureates: 1905. Bulletin of the Atomic Scientists, 34(8):1, October 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 18 April 1906, Oslo, Norway.

vonSchirach:2014:NPH

[vS14] Richard von Schirach. Die Nacht der Physiker: Heisenberg, Hahn, Weizsäcker und die deutsche Bombe. (German) [The night of the physicists: Heisenberg, Hahn, Weizsäcker and the German bomb], volume 61642 of rororo sachbuch. Rowohlt Taschenbuch, Reinbek, Germany, 2014. ISBN 3-499-61642-4.

256 pp. LCCN ???? URL http://d-nb.info/1044314834/04.

Vargas:1999:CEU

[VT99] Jose G. Vargas and Douglas G. Torr. The Cartan–Einstein unification with teleparallelism and the discrepant measurements of Newton’s constant G. Foundations of Physics, 29(2): 145–200, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A:1018840720961.

Vincent:2022:EOC

[VT22] Dwight E. Vincent and David R. Topper. Einstein’s Oxford cosmology blackboards: open portals to 1931. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-022-00046-9.

Vucinich:2001:ESI

[Vuc01] Alexander Vucinich. Einstein and Soviet ideology. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 2001. ISBN 0-8047-4209-X. viii + 291 pp. LCCN HX514

.V83 2001. URL http://www.loc.gov/catdir/description/
cam021/2001032266.html;   http://www.loc.gov/catdir/
enhancements/fy0710/2001032266-b.html.

Weizsacker:1979:EIP

[vW79] Carl Friedrich v. Weizsäcker. Einstein’s importance to physics, philosophy, and politics. In Aichelburg and Sexl [AS79], pages 159–168. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66.

DM48.00. URL http://link.springer.com/chapter/10.
1007/978-3-322-91080-6_11/.

Vonborzeszkowski:1986:PMR

[VW86] H. H. Vonborzeszkowski and R. Wahsner. Physical means of recognition and physical reality — a discussion between Niels Bohr and Albert Einstein on the status of quantum mechanics. Deutsche Zeitschrift für Philosophie, 34(12):1098–1106, ???? 1986. ISSN 0012-1045.

vonWeizsacker:2002:GPA

[vW02] Carl Friedrich von Weizsäcker. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg], volume 33078 of dtv. Deutscher Taschenbuch-Verlag, München, Germany, unabridged edition, 2002. ISBN 3-423-33078-3. 376 pp. LCCN QC6. EUR-D 12.50.

Vance-Watkins:1995:WFB

[VWA95] Lequita Vance-Watkins and Mariko Aratani, editors. White flash, black rain: women of Japan relive the bomb. Milkweed Editions, Minneapolis, MN, USA, 1995. ISBN 1-57131-402-4. x + 104 pp. LCCN D767.25.H6 W45 1995.

vonWeizsacker:2004:GPA

[vWR04] Carl Friedrich von Weizsäcker and Helmut Rechenberg. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg]. Marix, Wiesbaden, Germany, 2004. ISBN 3-937715-46-0. 375 pp.

LCCN ????  URL http://www.zentralblattmath.org/
zmath/en/search/?an=1104.83008.

W:1925:RVE

[W.25] D. M. W. (1) Relativity: a very Elementary Exposition; (2) Relativity, Meaning, and Motion; (3) The Common Sense of the Theory of Relativity; (4) La relativité dégagée d’hypothèses métaphysiques: exposé des théories d’Einstein, discussion de ces théories, essai d’une théorie nouvelle construite dans l’espace et le temps classiques. Nature, 116(2929):895, December 19, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v116/n2929/pdf/116895a0.pdf.

Wilber:1987:CCE

[W^{+}87] Ken Wilber et al., editors. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, 1987. ISBN 84-7245-172-0. 298 pp. LCCN ????

Wirth:2003:ARS

[WA03] John D. Wirth and Linda K. (Linda Kathleen) Aldrich, editors. Los Alamos: the Ranch School years, 1917–1943. University of New Mexico Press, Albuquerque, NM, USA, 2003. ISBN 0-8263-2883-0. xiii + 306 + 32 pp. LCCN LD7501.L7298

W57 2003. URL http://www.loc.gov/catdir/toc/ecip042/
2003007243.html.

Wachhorst:2006:MRB

[Wac06] Wyn Wachhorst. Movie review: Ten Days That Unexpectedly Changed America: Einstein’s Letter, Directed and produced by Barak Goodman and John Maggio. History Channel, 2006. 60 mins. (A&E Home Video, P.O. Box 2284, South Burlington, VT 05407; 888-423-1212). The Journal of American History, 93(3):983–984, 2006. ISSN 0021-8723 (print), 1945-2314 (elec-

tronic). URL http://jah.oxfordjournals.org/content/93/
3/983.short;   http://www.store.aetv.com.

Wade:1981:BEC

[Wad81a] Nicholas Wade. Brain of Einstein continues peregrinations. Science, 213(4507):521, July 31, 1981. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/213/4507/521.full.pdf.

Wade:1981:NDT

[Wad81b] Nicholas Wade. The Nobel duel: two scientists’ 21-year race to win the world’s most coveted research prize. Anchor Press/Doubleday, Garden City, NY, USA, 1981. ISBN 0-385-14981-6. xi + 321 + 4 pp. LCCN QP572.H9 W32 1981.

Wadia:2005:LAE

[Wad05] S. R. Wadia. The legacy of Albert Einstein (1879–1955). Current Science, 89(12):1971–1972, December 25, 2005. CODEN CUSCAM. ISSN 0011-3891.

Wadia:2007:LAE

[Wad07] S. R. (Spenta R.) Wadia, editor. The legacy of Albert Einstein: a collection of essays in celebration of the Year of Physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-277-271-5 (e-book), 981-270-049-8. xviii + 260 pp. LCCN QC16.E5 L36 2007eb.

Walker:1921:DEG

[Wal21] George W. Walker. A difficulty in Einstein’s gravitational theory. Nature, 107(2684):169–170, April 7, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v107/n2684/
pdf/107169c0.pdf.

Wallace:1961:STA

[Wal61] William A. Wallace. St. Thomas Aquinas, Galileo, and Einstein. The Thomist: A Speculative Quarterly Review, 24(1): 1–22, January 1961. CODEN ???? ISSN 2473-3725.

Wallace:1962:P

[Wal62] Irving Wallace. The Prize. Simon and Schuster, New York, NY, USA, 1962. 768 pp. LCCN PZ4.W1875 1962.

Wallace:1972:EAR

[Wal72] William A. Wallace. Einstein and Aquinas: A rapprochement by John F. Kiley. The Thomist: A Speculative Quarterly Review, 36(3):527–528, July 1972. CODEN ???? ISSN 2473-3725.

Wald:1977:STG

[Wal77a] Robert M. Wald. Space, time, and gravity: the theory of the big bang and black holes. University of Chicago Press, Chicago, IL, USA and London, UK, 1977. ISBN 0-226-87030-8. viii + 131 pp. LCCN QB981 .W24. US$10.95.

Walzer:1977:JUW

[Wal77b] Michael Walzer. Just and unjust wars: a moral argument with historical illustrations. Basic Books, New York, NY, USA, 1977. ISBN 0-465-03704-6. xx + 361 pp. LCCN U21.2

.W345. URL http://www.loc.gov/catdir/enhancements/
fy0830/77075252-d.html.

Wald:1979:GRA

[Wal79] Robert Wald. General Relativity after Einstein. Nature, 280(5724):703–704, August 23, 1979. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v280/n5724/pdf/
280703a0.pdf.

Walsh:1981:WEP

[Wal81] John Walsh. Waiting for the Einstein papers: Scholars now have a duplicate archive to turn to at Princeton, but conflict between estate and press delays publication of papers. Science, 213(4505):309–311, July 17, 1981. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/213/4505/309.full.pdf. See the

project editor’s account [Sta87] for a review, and the opposing view of the literary executor of the Einstein estate [Nat81]. See also [Sau04a, Sau04b] on the challenges of editing Einstein’s works.

Walsh:1982:EPC

[Wal82] John Walsh. Einstein papers coming on. Science, 218(4573): 664, November 12, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/218/4573/664.full.pdf.

Walsh:1983:EPP

[Wal83] John Walsh. Einstein Papers Project gets NSF interim grant. Science, 219(4580):39, January 7, 1983. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/219/4580/39.1.full.pdf.

Wald:1984:GR

[Wal84] Robert M. Wald. General Relativity. University of Chicago Press, Chicago, IL, USA and London, UK, 1984. ISBN 0-226-87032-4, 0-226-87033-2 (paperback). xiii + 491 pp. LCCN

QC173.6 .W35 1984. URL http://www.loc.gov/catdir/
description/uchi051/83017969.html;   http://www.loc.
gov/catdir/enhancements/fy0609/83017969-b.html;   http:
//www.loc.gov/catdir/toc/uchi051/83017969.html.

Waldrop:1988:EIR

[Wal88a] M. Mitchell Waldrop. Einstein’s impossible ring: Found. Science, 240(4860):1733, June 24, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/240/4860/1733.full.pdf.

Walsh:1988:ECE

[Wal88b] John Walsh. Editorial changes for Einstein papers. Science, 240 (4850):278, April 15, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/240/4850/278.full.pdf.

Walker:1989:GNS

[Wal89] Mark Walker. German National Socialism and the quest for nuclear power, 1939–1949. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36413-2. x + 290 pp. LCCN TK1078 .W35 1989. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/description/cam032/88036458.html;   http:
//www.loc.gov/catdir/toc/cam031/88036458.html.

Walker:1990:ME

[Wal90a] Evan Harris Walker. Ms. Einstein. Unpublished paper delivered to the AAAS session on “The Young Einstein”, New Orleans., February 18, 1990.

Walker:1990:DUB

[Wal90b] J. Samuel Walker. The decision to use the bomb: A historiographical update. Diplomatic History, 14(1):97–114, January 1990. CODEN ???? ISSN 0145-2096 (print), 1467-7709 (electronic).

Wali:1991:CBC

[Wal91a] K. C. (Kameshwar C.) Wali. Chandra: a biography of S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA and London, UK, 1991. ISBN 0-226-87054-5, 0-226-87055-3 (paperback). x + 341 + 32 pp. LCCN

QB36.C46 W35 1991. URL http://catdir.loc.gov/catdir/
description/uchi052/90010845.html;   http://catdir.
loc.gov/catdir/enhancements/fy0608/90010845-t.html;
 http://catdir.loc.gov/catdir/enhancements/fy0609/90010845-
b.html;   http://www.gbv.de/dms/bowker/toc/9780226870540.
pdf;   http://www.loc.gov/catdir/description/uchi052/
90010845.html;   http://www.loc.gov/catdir/enhancements/
fy0608/90010845-t.html;   http://www.loc.gov/catdir/
enhancements/fy0609/90010845-b.html;   http://www.nobelprize.

org/nobel_prizes/physics/laureates/1983/chandrasekhar-
or.html;   http://www.zentralblatt-math.org/zmath/en/
search/?an=0755.01045.

Walker:1991:LEM

[Wal91b] E. H. Walker. Letter to the Editor: Mileva Marić’s relativistic role. Physics Today, 44(2):122, February 1991. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See comments [Est06c].

Wald:1992:STG

[Wal92a] Robert M. Wald. Space, time, and gravity: the theory of the big bang and black holes. University of Chicago Press, Chicago, IL, USA and London, UK, second edition, 1992. ISBN 0-226-87028-6 (hardcover), 0-226-87029-4 (paper). x + 153 pp. LCCN QB981

.W24 1992. URL http://www.loc.gov/catdir/description/
uchi051/91028034.html;   http://www.loc.gov/catdir/
enhancements/fy0608/91028034-t.html;   http://www.loc.
gov/catdir/enhancements/fy0609/91028034-b.html.

Walker:1992:MGA

[Wal92b] Mark Walker. Myths of the German atom bomb. Nature, 359 (6395):473–474, October 8, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v359/n6395/full/359473a0.
html.

Walker:1992:PPW

[Wal92c] Mark Walker. Physics and propaganda: Werner Heisenberg’s Foreign Lectures under National Socialism. Historical Studies in the Physical and Biological Sciences, 22(2):339–389, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/27757685.

Walzer:1992:JUW

[Wal92d] Michael Walzer. Just and unjust wars: a moral argument with historical illustrations. Basic Books, New York, NY, USA, second edition, 1992. ISBN 0-465-03701-1 (paperback). xxxiv + 361 pp. LCCN U21.2 .W345 1992. URL http://www.loc.gov/

catdir/enhancements/fy0832/92245978-d.html.

Wald:1994:QFT

[Wal94a] Robert M. Wald. Quantum field theory in curved spacetime and black hole thermodynamics. Chicago lectures in physics. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-87025-1 (hardcover), 0-226-87027-8 (paper). xiii + 205 pp. LCCN QC174.45

.W35 1994. URL http://www.loc.gov/catdir/description/
uchi051/94011065.html;   http://www.loc.gov/catdir/
enhancements/fy0609/94011065-b.html;   http://www.loc.
gov/catdir/toc/uchi051/94011065.html;   http://www.
zentralblattmath.org/zmath/en/search/?an=0842.53052.

Walker:1994:BRBb

[Wal94b] Mark Walker. Book review: Nationalism and Internationalism in Science, 1880–1939: Four Studies of the Nobel Population by Elisabeth Crawford. Isis, 85(2):355–356, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236554.

Walker:1995:NSM

[Wal95] Mark Walker. Nazi science: myth, truth, and the German atomic bomb. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-44941-2. viii + 325 pp. LCCN Q127.G3 W35 1995.

Walker:1997:PUD

[Wal97] J. Samuel Walker. Prompt and utter destruction: Truman and the use of atomic bombs against Japan. University of North Carolina Press, Chapel Hill, NC, USA, 1997. ISBN 0-8078-2361-9 (hardcover), 0-8078-4662-7 (paperback), 0-8078-6618-0 (e-book). xiii + 142 pp. LCCN D767.25.H6 W355 1997. URL

http://www.h-net.org/review/hrev-a0b3w6-aa;   http:/
/www.jstor.org/stable/1408325; http://www.jstor.org/
stable/25054041;   http://www.jstor.org/stable/27551824;
 http://www.jstor.org/stable/3642279.

Wald:1998:BHR

[Wal98] Robert M. Wald, editor. Black holes and relativistic stars. University of Chicago Press, Chicago, IL, USA and London, UK, 1998. ISBN 0-226-87034-0 (hardcover). xii + 278 pp. LCCN

QB843.B55 B585 1998. URL http://www.loc.gov/catdir/
description/uchi051/97043102.html;   http://www.loc.
gov/catdir/toc/uchi052/97043102.html.

Walter:1999:AOB

[Wal99] Scott Walter. Analyses d’ouvrages: The Collected Papers, vol. 6 : The Berlin years : writings, 1914–1917 par Albert Einstein; A. J. Kox; Martin J. Klein; Robert Schulman. Revue d’Histoire des Sciences, 52(1):163–164, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://

www.jstor.org/stable/23633577.

Waller:2002:ELT

[Wal02a] John Waller. Einstein’s luck: the truth behind some of the greatest scientific discoveries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-860719-9. xi + 308 pp. LCCN Q125 .W266 2002. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/toc/fy034/2002033326.html.

Walter:2002:BRB

[Wal02b] Scott Walter. Book review: Beyond Einstein’s Velocity Addition Law. By Abraham A. Ungar. Fundamental Theories of Physics

117. Kluwer Academic, Dordrecht, The Netherlands, 2001, xlii + 413 pp., $138.00 (hardcover). Foundations of Physics, 32(2): 327–330, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A:1014469428383.

Walker:2003:OHV

[Wal03] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm ,,Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/
KWG/Ergebnisse/Ergebnisse10.pdf.

Walker:2004:PUD

[Wal04a] J. Samuel Walker. Prompt and utter destruction: Truman and the use of atomic bombs against Japan. University of North Carolina Press, Chapel Hill, NC, USA, 2004. ISBN 0-8078-5607-X (paperback). xv + 142 pp. LCCN D767.25.H6.

Walker:2004:TMI

[Wal04b] J. Samuel Walker. Three Mile Island: Nuclear Crisis in Historical Perspective. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 2004. ISBN 0-520-24683-7 (paperback). xi + 303 pp. LCCN TK1345.H37 W35 2005.

Walker:2005:HCK

[Wal05a] Stephen Walker. Hiroshima: Countdown der Katastrophe. (German) [Hiroshima: Countdown of the Catastrophe]. Bertelsmann, München, Germany, 2005. ISBN 3-570-00844-4. 399 pp. LCCN ???? URL http://swbplus.

bsz-bw.de/bsz118746170inh.htm;   http://swbplus.bsz-
bw.de/bsz118746170vlg.htm.

Walker:2005:SCHa

[Wal05b] Stephen Walker. Shockwave: countdown to Hiroshima. Harper-Collins, New York, NY, USA and London, UK, 2005. ISBN 0-06-074284-4 (hardcover), 0-7195-6625-8. 352 + 16 pp. LCCN D767.25.H6 W35 2005; D767.25.H6 W36 2005; D767.25.H6 W357 2005.

Walker:2005:SCHb

[Wal05c] Stephen Walker. Shockwave: the countdown to Hiroshima. John Murray, London, UK, 2005. ISBN 0-7195-6625-8s (hardcover), 0-7195-6773-4 (paperback). xiv + 352 pp. LCCN D767.25.H6 W36 2005.

Wali:2006:MBB

[Wal06a] Kameshwar Wali. The man behind Bose statistics. Physics Today, 59(10):46–52, October 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Walker:2006:OHR

[Wal06b] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0277-3.

Walker:2006:SCHa

[Wal06c] Stephen Walker. Shockwave: countdown to Hiroshima. Harper Perennial, New York, NY, 2006. ISBN 0-06-074285-2 s(paperback). xiv + 352 + 16 + 16 pp. LCCN D767.25.H6 W36 2006. US$14.95.

Walker:2006:SCHb

[Wal06d] Stephen Walker. Shockwave: the countdown to Hiroshima. John Murray, London, UK, 2006. ISBN 0-7195-6626-6s (paperback). xiv + 352 + 16 pp. LCCN D767.25.H6 W36 2006.

Walzer:2006:JUW

[Wal06e] Michael Walzer. Just and unjust wars: a moral argument with historical illustrations. Basic Books, New York, NY, USA, fourth edition, 2006. ISBN 0-465-03707-0 (paperback). xxviii + 361 pp. LCCN U21.2 .W345 2006.

Walter:2007:BVF

[Wal07] Scott Walter. Breaking in the 4-vectors: The four-dimensional movement in gravitation, 1905–1910. In Renn et al. [RSS^{+}07a], pages 193–252. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https://

link.springer.com/chapter/10.1007/978-1-4020-4000-9_
18.

Walzer:2015:JUW

[Wal15] Michael Walzer. Just and unjust wars: a moral argument with historical illustrations. Basic Books, New York, NY, USA, fifth edition, 2015. ISBN 0-465-05271-1 (hardcover), 0-465-05270-3 (e-book). xxxii + 381 pp. LCCN U21.2 .W345 2015.

Waldinger:2016:BBS

[Wal16] Fabian Waldinger. Bombs, brains, and science: The role of human and physical capital for the creation of scientific knowledge. Review of Economics and Statistics, 98(5):811–831, December 2016. CODEN RECSA9. ISSN 0034-6535 (print), 1530-9142

(electronic). URL https://direct.mit.edu/rest/article/
98/5/811/58619/Bombs-Brains-and-Science-The-Role-of-
Human-and.

Walker:2017:PHG

[Wal17] Prof. Dr. Mark Walker. Physics, history, and the German atomic Bomb. Berichte zur Wissenschaftsgeschichte, 40(3):271–288, September 2017. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Wang:1995:TPP

[Wan95] Hao Wang. Time in philosophy and in physics: From Kant and Einstein to Gödel. Synthese, 102(2):215–234, February 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01089801.

Wang:2000:SEG

[Wan00] Steve C. Wang. In search of Einstein’s genius. Science, 289 (5484):1477, September 1, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/289/5484/1477.full.

Warnow:1979:IEC

[War79] Joan Nelson Warnow, editor. Images of Einstein: a catalog. American Institute of Physics, Woodbury, NY, USA, revised edition, 1979. ISBN 0-521-41923-9, 0-88318-248-3. 77 pp. LCCN QC16.E5 I46 1979.

Ward:1984:SLS

[War84] C. Ward. Subtle is the lord… the science and the life of Albert Einstein — A. Pais. Southern Humanities Review, 18(2):175– 178, ???? 1984. ISSN 0038-4186.

Warwick:1989:IRE

[War89] Andrew Warwick. International relativity: The establishment of a theoretical discipline. Studies in History and Philosophy of Science Part A, 20(1):139–149, March 1989. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900381.

Warwick:1991:RFL

[War91] Andrew Warwick. On the role of the FitzGerald–Lorentz contraction hypothesis in the development of Joseph Larmor’s electronic theory of matter. Archive for History of Exact Sciences, 43(1):29–91, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=43&issue=1&spage=29.

Warwick:1992:CMC

[War92] Andrew Warwick. Cambridge mathematics and Cavendish physics: Cunningham, Campbell and Einstein’s Relativity 1905–1911: Part I: The uses of theory. Studies in History and Philosophy of Science Part A, 23(4):625–656, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290015X. See also Part II [War93].

Warwick:1993:CMC

[War93] Andrew Warwick. Cambridge mathematics and Cavendish physics: Cunningham, Campbell and Einstein’s Relativity 1905–1911: Part II: Comparing traditions in Cambridge physics. Studies in History and Philosophy of Science Part A, 24(1):1– 25, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.
com/science/article/pii/003936819390022C. See also Part

I [War92].

Warwick:1995:BRB

[War95] Andrew Warwick. Book review: The collected papers of Albert Einstein, vol 3, The Swiss years: Writings, 1909–1911 — M. J. Klein, A. J. Kox, J. Renn, R. Schulmann. British Journal for the History of Science, 28(4):480–482, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4027563.

Warwick:1996:BRC

[War96a] A. Warwick. Book review: The collected papers of Albert Einstein, vol 5, The Swiss years: Correspondence, 1902–1914 — M. J. Klein, A. J. Kox, J. Renn, R. Schulmann. British Journal for the History of Science, 29(100):109–110, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001x (electronic).

Warwick:1996:BRM

[War96b] Andrew Warwick. Book review: Martin J. Klein, Anne J. Kox, Juergen Renn and Robert Schulmann (eds.), The Collected Papers of Albert Einstein. Volume 5: The Swiss Years: Correspondence, 1902–1914. English translation by Anna Beck and Don Howard. Princeton: Princeton University Press, 1995. Pp. xvii + 384. ISBN 0-691-00099-9. £29.95, $19.50. British Journal for the History of Science, 29(1):109–110, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4027530.

Ward:1999:CRL

[War99] Chip Ward. Canaries on the Rim: Living Downwind in the West. The haymarket series. Verso, London, UK, 1999. ISBN 1-85984-750-1 (hardcover). 238 pp. LCCN GE155.G74 W37

Warwick:2003:BRR

[War03a] Andrew Warwick. Book review: Righting the history of electrodynamics: Olivier Darrigol, Electrodynamics from Ampère to Einstein. Oxford University Press, 2000. Pp. xi + 532, £79.50 (hardback). ISBN 0-19-850594-9. Notes and Records of the Royal Society of London, 57(2):253–254, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557707.

Warwick:2003:MTC

[War03b] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2003. ISBN 0-226-87374-9 (cloth), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN

QC19.6 .W37 2003. URL http://www.loc.gov/catdir/
bios/uchi051/2002153732.html;   http://www.loc.gov/
catdir/description/uchi051/2002153732.html;   http:/
/www.loc.gov/catdir/toc/fy037/2002153732.html.

Wasson:1992:LNP

[Was92] Tyler Wasson, editor. Laureaty Nobelevskoj premii: enciklopedia. (Russian) [Nobel Prize laureates: Encyclopaedia]. Progress, Moscow, Russia, 1992. ISBN 5-01-002539-6. LCCN ???? Two volumes.

Waters:1966:WOC

[Wat66] Frank Waters. The woman at Otowi Crossing: a novel. Swallow Press, Denver, CO, USA, 1966. 300 pp. LCCN ????

Waters:1970:WOC

[Wat70] Frank Waters. The woman at Otowi Crossing: a novel. Swallow Press, Chicago, IL, USA, 1970. 300 pp. LCCN

????  URL http://www.loc.gov/catdir/enhancements/
fy0666/86023820-b.html;   http://www.loc.gov/catdir/
enhancements/fy0666/86023820-d.html.

Waters:1988:WOC

[Wat88] Frank Waters. The woman at Otowi Crossing: a novel. Sage Books, Athens, OH, USA, revised edition, 1988. ISBN 0-8040-0893-0 (paperback). 314 pp. LCCN PS3545.A82

W6 1988. URL http://www.loc.gov/catdir/enhancements/
fy0666/86023820-b.html;   http://www.loc.gov/catdir/
enhancements/fy0666/86023820-d.html.

Watkins:1997:LMF

[Wat97a] Sallie A. Watkins. Lise Meitner: the foiled Nobelist. In Rayner-Canham and Rayner-Canham [RCRC97], chapter 16, pages 163–191. ISBN 0-7735-1608-5 (McGill hardcover), 0-7735-1642-5 (McGill paper), 0-941901-16-5 (Chemical Heritage hardcover), 0-941901-15-7 (Chemical Heritage paper). LCCN QC15 .R39

1997. URL http://www.jstor.org/stable/j.ctt809w2.

Watson:1997:QSW

[Wat97b] Andrew Watson. Quantum spookiness wins, Einstein loses in photon test. Science, 277(5325):481, July 25, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.sciencemag.org/content/277/5325/481.full.

Watson:1998:ETR

[Wat98] Andrew Watson. Einstein’s theory rings true. Science, 280 (5361):205, April 10, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/280/5361/205.full.

Watson:1999:GGV

[Wat99] Andrew Watson. Gravity’s gravity vindicates Einstein. Science, 286(5442):1065–1067, November 5, 1999. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.sciencemag.org/content/286/5442/1065.full.

Watkins:2000:BRB

[Wat00] Sallie A. Watkins. Book review: Nobel Prize Women in Science: Their Lives, Struggles, and Momentous Discoveries by Sharon Bertsch McGrayne. Isis, 91(2):390–391, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237002.

Wazeck:2009:EGO

[Waz09] Milena Wazeck. Einsteins Gegner: die öffentliche Kontroverse um die Relativitätstheorie in den 1920er Jahren. (German) [Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s]. Campus, Frankfurt am Main, Germany, 2009. ISBN 3-593-38914-2. 429 pp. LCCN QC173.585 .W39 2009. See also English translation in [Waz14].

Wazeck:2010:EFC

[Waz10] Milena Wazeck. The 1922 Einstein Film: Cinematic innovation and public controversy. Physics in Perspective (PIP), 12(2): 163–179, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://www.springerlink.com/
content/f625752811824q46/fulltext.pdf.

Wazeck:2014:EOP

[Waz14] Milena Wazeck. Einstein’s opponents: the public controversy about the Theory of Relativity in the 1920s. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-01744-0 (hardcover). xxi + 355 pp. LCCN QC173.585 .W3913

2014. URL http://assets.cambridge.org/97811070/17443/

cover/9781107017443.jpg. English translation by Geoffrey S. Koby of [Waz09].

Wasson:1987:NPW

[WB^{+}87] Tyler Wasson, Gert H. Brieger, et al., editors. Nobel Prize winners: an H. W. Wilson biographical dictionary. H. W. Wilson, New York, NY, USA, 1987. ISBN 0-8242-0756-4. xxxiv + 1165 pp. LCCN AS911.N9 N59 1987.

Wolfson:2000:ERQ

[WB00] Richard Wolfson and Andreas Burgstaller. Einstein’s relativity and the quantum revolution: modern physics for non-scientists. Great courses, Science and mathematics. Teaching Computers, Chantilly, VA, USA, second edition, 2000. ISBN 1-56585-565-5. LCCN QC173.58 .W65 2000d. US$254.95. 4 videodiscs (720 min.).

Warner:2001:SAS

[WB01] Edith Warner and Patrick Burns, editors. In the shadow of Los Alamos: selected writings of Edith Warner. University of New Mexico Press, Albuquerque, NM, USA, 2001. ISBN 0-8263-1974-2. xiv + 226 pp. LCCN

F804.L6 W37 2001. URL http://www.loc.gov/catdir/
enhancements/fy1115/2001002383-b.html;   http://www.
loc.gov/catdir/enhancements/fy1115/2001002383-d.html.

Weeks:2000:TDD

[WCL^{+}00] Eric R. Weeks, J. C. Crocker, Andrew C. Levitt, Andrew Schofield, and D. A. Weitz. Three-dimensional direct imaging of structural relaxation near the colloidal glass transition. Science, 287(5454):626–631, January 28, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/287/5453/627.abstract.

Wainer:1996:VRN

[WD96] Howard Wainer and John W. Durso. Visual revelations: A Nobel graph. Chance, 9(1):12–16, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wilczek:2006:FRM

[WD06] Frank Wilczek and Betsy Devine. Fantastic realities: 49 mind journeys and a trip to Stockholm. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-649-X (hardcover), 981-256-655-4 (paperback), 981-277-430-0 (e-book). ix + 522 pp. LCCN QC75 .W55 2006.

Wilber:2002:CCE

[WdC02] Ken Wilber and Pedro de Casso. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, seventh edition, 2002. ISBN 84-7245-172-0. 296 pp. LCCN ????

Winteler-Einstein:1924:AEB

[WE24] Maja Winteler-Einstein. Albert Einstein —- Beitrag für sein Lebensbild. Draft for a biography of the author’s brother., February 15, 1924. Reprinted in [SCS87].

Winteler-Einstein:1987:AEB

[WE87] Maya Winteler-Einstein. Albert Einstein—Beitrag für sein Lebensbild. (German) [Albert Einstein: Contribution to his life view]. In Stachel et al. [SCS87], page ?? ISBN 0-691-08407-6. LCCN QC16.E5 A2 1987. US$52.50. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://einsteinpapers.press.princeton.edu/vol1-doc;
 http://press.princeton.edu/titles/2515.html;   http:
//www.loc.gov/catdir/description/prin021/86043132.html;
 http://www.loc.gov/catdir/toc/prin032/86043132.html.

Based on a 1924 manuscript.

Weart:1976:SS

[Wea76] Spencer R. Weart. Scientists with a secret. Physics Today, 29(2):23–30, February 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v29/i2/p23_s1.

Weart:1988:NFH

[Wea88] Spencer R. Weart. Nuclear fear: a history of images. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-62835-7. 535 pp. LCCN QC773 .W43 1988.

Weart:2012:RNF

[Wea12] Spencer R. Weart. The rise of nuclear fear. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-05233-1. ix + 367 pp. LCCN QC773 .W44 2012.

Weber:1961:GRG

[Web61] J. (Joseph) Weber. General Relativity and Gravitational Waves, volume 10 of Interscience tracts on physics and astronomy. Interscience Publishers, New York, NY, USA, 1961. 200 pp. LCCN QC6 .W4.

Weber:1973:RWS

[Web73] Robert L. Weber, editor. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weber:1979:GR

[Web79] Joseph Weber. Gravitational radiation. In Aichelburg and Sexl [AS79], pages 25–31. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5

A66. DM48.00. URL http://link.springer.com/chapter/
10.1007/978-3-322-91080-6_3/.

Weber:1993:INE

[Web93] Tullio Weber. Indeterminism, nonseparability and the Einstein– Podolsky–Rosen Paradox. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 10, pages 199–209. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-011-2496-6_10.

Webb:1999:MUC

[Web99] Stephen Webb. Measuring the universe: the cosmological distance ladder. Springer-Praxis series in astronomy and astrophysics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-85233-106-2 (paperback). xvi + 342 pp. LCCN QB991.C66 W43

1999. URL http://www.loc.gov/catdir/enhancements/
fy0816/98043764-d.html;   http://www.loc.gov/catdir/
enhancements/fy0816/98043764-t.html.

Weber:2004:GRG

[Web04] J. (Joseph) Weber. General Relativity and Gravitational Waves. Dover, New York, NY, USA, 2004. ISBN 0-486-43887-2 (paperback). viii + 200 pp. LCCN QC173.6 .W43 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0618/2004056006-
d.html.

Weeramantry:1987:NWS

[Wee87] C. G. Weeramantry. Nuclear weapons and scientific responsibility. Longwood Academic, Wolfeboro, NH, USA, 1987. ISBN 0-89341-542-1. ???? pp. LCCN JX5133.A7 W44 1987. US$25.00.

Wegener:2008:BRB

[Weg08] Daan Wegener. Book review: Einstein: His Life and Universe, Walter Isaacson. Simon, Schuster (2007). 704 pp., $32, ISBN-13 978-0-7432-6473-0. Studies in History and Philosophy of Modern Physics, 39(4):914–915, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000609. See [Isa07].

Weinberg:1972:GCP

[Wei72] Steven Weinberg. Gravitation and cosmology: principles and applications of the General Theory of Relativity. Wiley, New York, NY, USA, 1972. ISBN 0-471-92567-5. xxviii + 657 pp.

LCCN QC6 .W47. URL http://www.loc.gov/catdir/toc/
onix04/78037175.html.

Weiner:1988:OHS

[Wei88] Charles Weiner. Oral history of science: A mushrooming cloud? The Journal of American History, 75(2):548–559, September 1988. ISSN 0021-8723 (print), 1945-2314 (electronic). URL

http://www.jstor.org/stable/1887871.

Weinberg:1989:CCP

[Wei89] Steven Weinberg. The cosmological constant problem. Reviews of Modern Physics, 61(1):1–23, January 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
61.1.

Weindling:1994:BRE

[Wei94] Paul Weindling. Book review: Elisabeth Crawford, Nationalism and internationalism in science, 1880–1939. Four Studies of the Nobel Population, Cambridge, Cambridge University Press, 1992, xii + 157 p. Nuncius, 9(2):880–881, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01297.

Weintraub:1995:LGV

[Wei95] Stanley Weintraub. The last great victory: the end of World War II, July/August 1945. Truman Talley Books/Dutton, New York, NY, USA, 1995. ISBN 0-525-93687-4. xii + 730 + 16 pp. LCCN D755.7 .W44 1995.

Weinberg:1995:QTF

[Wei00] Steven Weinberg. The quantum theory of fields. Cambridge University Press, Cambridge, UK, 1995–2000. ISBN 0-521-58555-4 (set: hardcover), 0-521-67056-X (set: paperback), 0-521-55001-7 (vol. 1: hardcover), 0-521-67053-5 (vol. 1: paperback), 0-521-55002-5 (vol. 2: hardcover), 0-521-67054-3 (vol. 2: paperback), 0-521-66000-9 (vol. 3: hardcover), 0-521-67055-1 (vol. 3: paperback). xxvi + 609 pp. LCCN 27; BFJ; QC174.45 .W45 1995. URL http://www.loc.

gov/catdir/description/cam026/95002782.html;   http:
//www.loc.gov/catdir/toc/cam021/95002782.html.

Weiner:2003:BRA

[Wei03] Douglas R. Weiner. Book review: Alexander Vucinich, Einstein and Soviet Ideology. Common Knowledge, 9(2):352, Spring 2003.

URL http://muse.jhu.edu/article/41461.

Weigel:2004:ECH

[Wei04] J. W. Weigel. Einstein’s cosmos: How Albert Einstein’s vision transformed our understanding of space and time. Library Journal, 129(7):119, April 15, 2004. CODEN LIBJA7. ISSN 0363-0277.

Weinberg:2005:EM

[Wei05a] Steven Weinberg. Einstein’s mistakes. Physics Today, 58(11): 31–35, November 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See letters [GCO^{+}06].

Weinberg:2005:QTF

[Wei05b] Steven Weinberg. The quantum theory of fields. 1. Foundations, 2. Modern applications, 3. Supersymmetry. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-67053-5 (paperback). xxvi + 609 pp. LCCN ???? URL http://

www.gbv.de/dms/ohb-opac/481933859.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=1069.81500.

Weinfurter:2005:PE

[Wei05c] Harald Weinfurter. The power of entanglement. Physics World, 18(1):47–51, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a34.
pdf.

Weinberg:2006:VEM

[Wei06a] Steven Weinberg. The value of Einstein’s mistakes. Physics Today, 59(4):15–16, April 1, 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.4797363.

Weinberger:2006:IWJ

[Wei06b] Sharon Weinberger. Imaginary Weapons: a Journey Through the Pentagon’s Scientific Underworld. Nation Books, New York, NY, USA, 2006. ISBN 1-56025-849-7. xxviii + 276 pp. LCCN U393 .W45 2006. US$26.00. URL http://www.npr.org/

templates/story/story.php?storyId=5478373.

Weinberg:2008:C

[Wei08] Steven Weinberg. Cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-852682-2. xvii + 593 pp. LCCN QB981 .W475 2008. URL http://www.loc.

gov/catdir/enhancements/fy0907/2007049371-b.html;
http://www.loc.gov/catdir/enhancements/fy0907/2007049371-
d.html;   http://www.loc.gov/catdir/enhancements/fy0907/
2007049371-t.html.

Weiche:2012:IMI

[Wei12a] Iris Weiche. “Imagination is more important than knowledge, for knowledge is limited” (Albert Einstein). Phlebologie, 41(6): 285, December 2012. ISSN 0939-978X.

Weinstein:2012:AEF

[Wei12b] Galina Weinstein. Albert Einstein and the Fizeau 1851 water tube experiment. arxiv.org, ??(??):??, April 16, 2012. URL

http://arxiv.org/abs/1204.3390.

Weinstein:2012:AEZ

[Wei12c] Galina Weinstein. Albert Einstein at the Zürich Polytechnic: a rare mastery of Maxwell’s electromagnetic theory. arxiv.org,

??(??):??, May 19, 2012. URL http://arxiv.org/abs/1205.
4335.

Weinstein:2012:AER

[Wei12d] Galina Weinstein. Albert Einstein: Rebellious wunderkind. arxiv.org, ??(??):??, May 21, 2012. URL https://arxiv.org/

abs/1205.4509.

Weinstein:2012:AEC

[Wei12e] Galina Weinstein. Albert Einstein’s close friends and colleagues from the Patent Office. arxiv.org, ??(??):??, May 17, 2012. URL

http://arxiv.org/abs/1205.3904.

Weinstein:2012:AEM

[Wei12f] Galina Weinstein. Albert Einstein’s methodology. arxiv.org, ??(??):??, September 24, 2012. URL http://arxiv.org/abs/

1209.5181.

Weinstein:2012:BAE

[Wei12g] Galina Weinstein. Biographies of Albert Einstein — mastermind of theoretical physics. arxiv.org, ??(??):??, May 24, 2012. URL

http://arxiv.org/abs/1205.5539.

Weinstein:2012:BHP

[Wei12h] Galina Weinstein. A biography of Henri Poincaré — 2012 centenary of the death of Poincaré. arxiv.org, ??(??):??, July 3, 2012.

URL http://arxiv.org/abs/1207.0759.

Weinstein:2012:DMM

[Wei12i] Galina Weinstein. Did Mileva Marić assist Einstein in writing his 1905 path breaking papers? arxiv.org, ??(??):??, April 16,

2012. URL http://arxiv.org/abs/1204.3551.

Weinstein:2012:DPE

[Wei12j] Galina Weinstein. Did Poincaré explore the inertial mass–energy equivalence? arxiv.org, ??(??):??, April 30, 2012. URL http:/

/arxiv.org/abs/1204.6575.

Weinstein:2012:DSR

[Wei12k] Galina Weinstein. A discussion of special relativity. arxiv.org,

??(??):??, May 1, 2012. URL http://arxiv.org/abs/1205.
0221.

Weinstein:2012:ENT

[Wei12l] Galina Weinstein. The Einstein–Nordström theory. arxiv.org,

??(??):??, May 27, 2012. URL http://arxiv.org/abs/1205.
5966.

Weinstein:2012:ECLa

[Wei12m] Galina Weinstein. Einstein chases a light beam. arxiv.org, ??

(??):1–53, April 9, 2012. URL http://arxiv.org/abs/1204.
1833.

Weinstein:2012:EIS

[Wei12n] Galina Weinstein. Einstein on the impossibility of superluminal velocities. arxiv.org, ??(??):??, March 22, 2012. URL http://

arxiv.org/abs/1203.4954.

Weinstein:2012:ESC

[Wei12o] Galina Weinstein. Einstein the stubborn: Correspondence between Einstein and Levi-Civita. arxiv.org, ??(??):??, February

20, 2012. URL http://arxiv.org/abs/1202.4305.

Weinstein:2012:ESG

[Wei12p] Galina Weinstein. Einstein’s 1912–1913 struggles with gravitation theory: Importance of static gravitational fields theory. arxiv.org, ??(??):??, February 13, 2012. URL http://arxiv.

org/abs/1202.2791.

Weinstein:2012:ECLb

[Wei12q] Galina Weinstein. Einstein’s clocks and Langevin’s twins. arxiv.org, ??(??):??, May 4, 2012. URL http://arxiv.org/

abs/1205.0922.

Weinstein:2012:EPE

[Wei12r] Galina Weinstein. Einstein’s pathway to the equivalence principle 1905–1907. arxiv.org, ??(??):??, August 25, 2012. URL

http://arxiv.org/abs/1208.5137.

Weinstein:2012:BEFa

[Wei12s] Galina Weinstein. From the Berlin “Entwurf” field equations to the Einstein tensor I: October 1914 until beginning of November 1915. arxiv.org, ??(??):??, January 25, 2012. URL http://

arxiv.org/abs/1201.5352.

Weinstein:2012:BEFb

[Wei12t] Galina Weinstein. From the Berlin “Entwurf” field equations to the Einstein tensor II: November 1915 until March 1916. arxiv.org, ??(??):??, January 25, 2012. URL http://arxiv.

org/abs/1201.5353.

Weinstein:2012:BEFc

[Wei12u] Galina Weinstein. From the Berlin “Entwurf” field equations to the Einstein tensor III: March 1916. arxiv.org, ??(??):??,

January 25, 2012. URL http://arxiv.org/abs/1201.5358.

Weinstein:2012:GGR

[Wei12v] Galina Weinstein. Genesis of general relativity — discovery of general relativity. arxiv.org, ??(??):??, April 16, 2012. URL

http://arxiv.org/abs/1204.3386.

Weinstein:2012:MBA

[Wei12w] Galina Weinstein. Max Born, Albert Einstein and Hermann Minkowski’s space–time formalism of special relativity. arxiv.org, ??(??):??, October 25, 2012. URL http://arxiv.

org/abs/1210.6929.

Weinstein:2012:PDE

[Wei12x] Galina Weinstein. Poincaré’s dynamics of the electron — a theory of relativity? arxiv.org, ??(??):??, April 30, 2012. URL

http://arxiv.org/abs/1204.6576.

Weinstein:2012:VMV

[Wei12y] Galina Weinstein. Variation of mass with velocity: “Kugeltheorie” or “Relativtheorie”. arxiv.org, ??(??):??, May 27, 2012.

URL http://arxiv.org/abs/1205.5951.

Weinert:2013:EPT

[Wei13a] Friedel Weinert. EPR and the ‘passage’ of time. Philosophia Naturalis, 50(2):173–199, ???? 2013. CODEN ???? ISSN 0031-8027.

Weinstein:2013:RPL

[Wei13b] Galina Weinstein. The 1905 relativity paper and the “light quantum”. arxiv.org, ??(??):??, July 8, 2013. URL http://

arxiv.org/abs/1307.2132.

Weinstein:2013:ECE

[Wei13c] Galina Weinstein. Einstein and the conservation of energymomentum in general relativity. arxiv.org, ??(??):??, October

10, 2013. URL http://arxiv.org/abs/1310.2890.

Weinstein:2013:EDF

[Wei13d] Galina Weinstein. Einstein’s 1916 derivation of the field equations. arxiv.org, ??(??):??, October 24, 2013. URL http://

arxiv.org/abs/1310.6541.

Weinstein:2013:GGA

[Wei13e] Galina Weinstein. George Gamow and Albert Einstein: Did Einstein say the cosmological constant was the “biggest blunder” he ever made in his life? arxiv.org, ??(??):??, October 3,

2013. URL http://arxiv.org/abs/1310.1033.

Weinstein:2013:MSW

[Wei13f] Galina Weinstein. The mythical snake which swallows its tail: Einstein’s matter world. arxiv.org, ??(??):??, September 25,

2013. URL http://arxiv.org/abs/1309.6590.

Weinstein:2013:PTE

[Wei13g] Galina Weinstein. PBR theorem and Einstein’s quantum hole argument. arxiv.org, ??(??):??, January 31, 2013. URL http:/

/arxiv.org/abs/1301.7526.

Weinstein:2014:DEN

[Wei14a] Galina Weinstein. Did Einstein “nostrify” Hilbert’s final form of the field equations for general relativity? arxiv.org, ??(??):??,

December 24, 2014. URL http://arxiv.org/abs/1412.1816.

Weinstein:2014:ESP

[Wei14b] Galina Weinstein. Einstein, Schwarzschild, the perihelion motion of Mercury and the rotating disk story. arxiv.org, ??(??):??,

November 26, 2014. URL http://arxiv.org/abs/1411.7370.

Weinstein:2015:EPS

[Wei15a] Galina Weinstein. Einstein’s pathway to the Special Theory of Relativity. Cambridge Scholars Publishing, Cambridge, UK, 2015. ISBN 1-4438-7434-5 (print), 1-4438-7889-8. x + 384 pp. LCCN QC173.65 .W45 2015.

Weinstein:2015:EUR

[Wei15b] Galina Weinstein. Einstein’s uniformly rotating disk and the hole argument. arxiv.org, ??(??):??, April 15, 2015. URL http:

//arxiv.org/abs/1504.03989.

Weiss:2015:TEDb

[Wei15c] Eugène-Henri Weiss. La théorie d’Einstein dévoilée aux profanes: adaptation de l’ouvrage de Rudolf Lämmel Wege zur Relativitäts Théorie. (French) [The theory of Einstein unveiled to the profane: adaptation of the book by Rudolf Lämmel The Roads to Relativity Theory]. Éditions des Traboules, Brignais, France, 2015. ISBN 2-35916-075-3, 2-35916-032-X. 144 pp.

LCCN ???? URL http://catalogue.bnf.fr/ark:/12148/
cb44275019f.

Weinstein:2018:WDE

[Wei18a] Galina Weinstein. Why did Einstein reject the November tensor in 1912–1913, only to come back to it in November 1915? Studies in History and Philosophy of Modern Physics, 62(??):98–122, May 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300266.

Weiss:2018:NLL

[Wei18b] Rainer Weiss. Nobel lecture: LIGO and the discovery of gravitational waves I. Reviews of Modern Physics, 90(4):040501– ??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL https://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.040501.

Wells:1914:WSF

[Wel14] H. G. (Herbert George) Wells. The world set free. A story of mankind. Macmillan, London, UK, 1914. viii +

286 pp. URL http://www.nobelprize.org/nobel_prizes/

chemistry/laureates/1921/. This [then science-fiction] book was substantially influenced by the writings of Frederick Soddy (Nobel Prize in Chemistry 1921 [awarded in fall 1922] “for his contributions to our knowledge of the chemistry of radioactive substances, and his investigations into the origin and nature of isotopes”) on nuclear transmutations and the possibility of extracting enormous amounts of energy from them; see [Scl89].

Wells:1976:WSF

[Wel76] H. G. (Herbert George) Wells. The world set free. Corgi science fiction. Corgi, London, UK, 1976. ISBN 0-552-10258-X (paperback). 192 pp. LCCN PR5774 .W65 1976.

Wells:1988:WSF

[Wel88] H. G. (Herbert George) Wells. The world set free. Hogarth, London, UK, 1988. ISBN 0-7012-0576-8. 191 pp. LCCN ???? With a new introduction by Brian Wilson Aldiss.

Welsh:1993:NSS

[Wel93] Ian Welsh. The NIMBY syndrome: its significance in the history of the nuclear debate in Britain. British Journal for the History of Science, 26(1):15–32, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027338.

Welsome:1999:PFA

[Wel99] Eileen Welsome. The Plutonium Files: America’s secret medical experiments in the Cold War. Dial Press, New York, NY, USA, 1999. ISBN 0-385-31402-7. ix + 580 pp. LCCN RA1231.R2 W45 1999. URL http:/

/catdir.loc.gov/catdir/bios/random059/99010991.html;

 http://catdir.loc.gov/catdir/description/random047/
99010991.html;   http://catdir.loc.gov/catdir/samples/
random043/99010991.html;   http://lcweb.loc.gov/catdir/
toc/99010991.html.

Welsome:2000:PFA

[Wel00] Eileen Welsome. The Plutonium Files: America’s Secret Medical Experiments in the Cold War. Delta, New York, NY, USA, 2000. ISBN 0-385-31954-1, 1-299-00836-4 (e-book). ix + 580 + 16 pp. LCCN RA1231.R2 W45 2000.

Weller:2006:FNC

[Wel06] George Weller. First into Nagasaki: the censored eyewitness dispatches on post-atomic Japan and its prisoners of war. Crown Publishers, New York, NY, USA, 2006. ISBN 0-307-34201-8. x + 320 + 8 pp. LCCN

D767.25.N3 W45 2006. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006011345-b.html;   http://www.
loc.gov/catdir/enhancements/fy0702/2006011345-d.html;
 http://www.loc.gov/catdir/enhancements/fy0809/2006011345-

s.html. Edited with an essay by Anthony Weller, and a foreword by Walter Cronkite.

Wellerstein:2008:PBN

[Wel08] Alex Wellerstein. Patenting the bomb: Nuclear weapons, intellectual property, and technological control. Isis, 99(1):57– 87, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/587556.

Wenzl:1949:ETR

[Wen49] Aloys Wenzl. Einstein’s theory of relativity, viewed from the standpoint of critical realism, and its significance for philosophy. In Schilpp [Sch49c], chapter 23, pages 581–606. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Wesley:1980:EDS

[Wes80] J. P. Wesley. Einstein dynamics without special-relativistic kinematics. Foundations of Physics, 10(5–6):503–511, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00708747.

West:2001:VTN

[Wes01] Thomas G. West. Visual thinkers and Nobel Prizes. Computer Graphics, 35(1):14–15, February 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2004:MCG

[Wes04] Nigel West. Mortal crimes: the greatest theft in history: Soviet penetration of the Manhattan Project. Enigma Books, New York, NY, USA, 2004. ISBN 1-929631-29-4 (hardcover), 1-936274-82-5 (e-book). xxii + 279 + 16 pp. LCCN UB271.R9 W47 2004.

West:2005:AEM

[Wes05] J. O. West. Albert Einstein to Michele Besso. Physics Today, 58(7):15–16, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Westfall:2014:GIG

[Wes14] Catherine Westfall. Gaining inspiration from Galileo, Einstein and Oppenheimer. History of Physics Newsletter, 12(5):5, 7, 9, Fall 2014. CODEN ???? ISSN ???? URL http://www.aps.

org/units/fhp/newsletters/fall2014/index.cfm.

West:2015:EC

[Wes15] Tom West. The Einstein code. Pan Books, London, UK, 2015. ISBN 1-4472-1034-4 (paperback), 1-4472-4660-8 (ePub e-book). 400 pp. LCCN ????

Wetzels:1980:TRG

[Wet80] W. Wetzels. The Theory of Relativity — generally comprehensible — Albert Einstein techniques of popular scientific didactics. Lili — Zeitschrift für Literaturwissenschaft und Linguistik, 10 (40):14–24, ???? 1980. ISSN 0049-8653.

Weyl:1918:RZM

[Wey18] Hermann Weyl. Raum, Zeit, Materie: Vorlesungen über allgemeine Relativitätstheorie. (German) [Space, Time, Matter: Lectures on General Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1918. viii + 234 pp. LCCN ????

Weyland:1920:ERW

[Wey20] Paul Weyland. Einsteins Relativitätstheorie: Eine wissenschaftliche Massensuggestion. (German) [Einstein’s Theory of Relativity: a scientific mass suggestion]. Tägliche Rundschau, ??(??):??, August 6, 1920.

Weyl:1921:EG

[Wey21] Hermann Weyl. Electricity and gravitation. Nature, 106(2677): 800–802, February 17, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). Translated by Robert W. Lawson from the German manuscript. Special issue on Relativity. This is one of only three papers by foreign authors in this issue; the others are [Ein21b, Lor21].

Weyl:1922:STMa

[Wey22a] Hermann Weyl. Space–time–matter. Methuen and Co. Ltd., London, UK, 1922. xi + 330 pp. LCCN QC6 .W54rE; QC6 .W55; QC6 .W5. English translation by Henry Leopold Brose of [Wey18].

Weyl:1922:STMb

[Wey22b] Hermann Weyl. Space–time–matter. Dover, New York, NY, USA, fourth edition, 1922. xvi + 330 pp. LCCN QC6 .W5413 1922; QC6 .W4 1920; QC6 .W4 1922. English translation by Henry Leopold Brose of [Wey18].

Weyl:1923:ARG

[Wey23] Hermann Weyl. Zur allgemeinen Relativitätstheorie. (German) [On the General Relativity Theory]. Physikalische Zeitschrift, 24 (??):230–232, ???? 1923. CODEN PHZTAO. ISSN 0369-982X.

Weyl:1950:STM

[Wey50] Hermann Weyl. Space, time, matter. Dover, New York, NY, USA, 1950. xvi + 330 pp. LCCN QC6 .W55 1952; QC6 .W421 1950; QC6 .W54s 1950; QC6 .W54rE 1952; QC6 .W5 1950; QC6 .W4. Translated from the German by Henry L. Brose.

Weyl:2007:PIG

[Wey07] Hermann Weyl. Purely infinitesimal geometry (excerpt). In Renn et al. [RSS^{+}07b], pages 1089–1105. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/chapter/10.1007/978-1-

4020-4000-9_49. Originally published as “Reine Infinitesimalgeometrie”, Mathematisch Zeitschrift 2, 384–411 (1918).

Weyl:2009:GOR

[Wey09] H. Weyl. Golden oldie: Republication of: On the General Relativity Theory. General Relativity and Gravitation, 41(7):1661– 1666, ???? 2009. CODEN GRGVA8. ISSN 0001-7701 (print),

1572-9532 (electronic). URL http://www.springerlink.com/

content/107735q12g501356/. English translation of [Wey23]. See also editorial note [Ehl09].

Wheeler:1949:CET

[WF49] John Archibald Wheeler and Richard Phillips Feynman. Classical electrodynamics in terms of direct interparticle action. Reviews of Modern Physics, 21(3):425–433, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.425;   http://rmp.aps.org/

abstract/RMP/v21/i3/p425_1. Special issue in honor of Albert Einstein on his 70th birthday.

Wheeler:1998:GBH

[WF98] John Archibald Wheeler and Kenneth William Ford. Geons, black holes, and quantum foam: a life in physics. W. W. Norton & Co., New York, NY, USA, 1998. ISBN 0-393-04642-7. 380 pp. LCCN QC16.W48 A3 1998.

White:1994:ELSb

[WG94] Michael White and John Gribbin. Einstein: a life in science. Dutton, New York, NY, USA, 1994. ISBN 0-525-93750-1. viii + 279 pp. LCCN 0.2whi a0160.

White:1995:ELS

[WG95] Michael White and John R. Gribbin. Einstein: a life in science. Plume, New York, NY, USA, 1995. ISBN 0-452-27146-0 (paperback). viii + 279 pp. LCCN QC16.E5 W47 1995.

Wigner:1996:SWJ

[WG96] E. P. Wigner and H. H. Goldstine. The scientific work of John von Neumann. In Wigner [Wig96m], pages 123–126. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wellerstein:2017:SSH

[WG17] Alex Wellerstein and Edward Geist. The secret of the Soviet hydrogen bomb. Physics Today, 70(4):40–47, April 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Williams:1999:PTS

[WH99] J. E. Williams and M. J. Holland. Preparing topological states of a Bose–Einstein condensate. Nature, 401(6753):568– 572, October 7, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v401/n6753/full/401568a0.html.

Whitfield:2007:WMR

[WH07] Donald Whitfield and James L. Hicks, editors. What’s the matter?: readings in physics. Great Books Foundation, Chicago, IL, USA, 2007. ISBN 1-880323-91-5. xx + 540 pp. LCCN

QC7.5 .W53 2007. URL http://store.greatbooks.org/adu-

wtm.html. Foreword by Alan Lightman.

Wheeler:1968:EV

[Whe68a] John A. Wheeler. Einstein’s vision. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. vii + 108 pp.

Wheeler:1968:EVW

[Whe68b] John A. Wheeler. Einsteins vision: Wie steht es heute mit Einsteins Vision, alles als Geometrie aufzufassen?. (German) [Einstein’s vision: What about Einstein’s vision today of understanding everything as geometry?]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. vii + 108 pp.

Wheeler:1970:PER

[Whe70] John A. Wheeler. Predvidenie Ejnštejna. (Russian) [Einstein’s vision]. Mir, Moscow, USSR, 1970. 112 pp.

Wheeler:1979:BHI

[Whe79a] John A. Wheeler. The black hole: an imaginary conversation with Albert Einstein. In Aichelburg and Sexl [AS79], pages 51–56. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5 A66.

DM48.00. URL http://link.springer.com/chapter/10.
1007/978-3-322-91080-6_5/.

Wheeler:1979:MSO

[Whe79b] John A. Wheeler. Mercer Street and other memories. In Aichelburg and Sexl [AS79], pages 201–211. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN QC16.E5

A66. DM48.00. URL http://link.springer.com/chapter/
10.1007/978-3-322-91080-6_15/.

Wheeler:1979:EOS

[Whe79c] John Archibald Wheeler. Einstein and other seekers of the larger view. Science and Public Policy, 6(6):396–404, December 1979. ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

spp.oxfordjournals.org/content/6/6/396.full.pdf+html.

Wheeler:1980:AEM

[Whe80] John Archibald Wheeler. Albert Einstein: March 14, 1879– April 18, 1955. Biographical memoirs — National Academy of Sciences of the United States of America, 51:96–117, 1980. CODEN BMNSAC. ISBN 0-309-59901-6. ISSN 0077-

2933. URL http://www.nap.edu/catalog.php?record_id=
574;   http://www.nap.edu/catalog/574.html.

Wheeler:1990:JGS

[Whe90] John Archibald Wheeler. A journey into gravity and spacetime, volume 31 of Scientific American Library series. Scientific American Library, New York, NY, USA, 1990. ISBN 0-7167-5016-3. xii + 257 pp. LCCN QB334 .W49 1990. US$10.95.

Wheeler:1991:AE

[Whe91] John Archibald Wheeler. Albert Einstein. In Ferris and Fadiman [FF91], pages 563–576. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Wheeler:1992:HU

[Whe92] John Archibald Wheeler. At Home in the Universe, volume 9 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1992. ISBN 0-88318-862-7. ix + 371 pp. LCCN Q158.5 .W44 1992.

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11– 83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitehead:1922:PRA

[Whi22] Alfred North Whitehead. The Principle of Relativity with applications to physical science. Cambridge University Press, Cambridge, UK, 1922. xii + 190 pp. LCCN QC173.55 .W55 1922.

Whittaker:1927:OPR

[Whi27] E. T. Whittaker. The outstanding problems of Relativity. Science, 66(1706):223–229, September 9, 1927. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/66/1706/223.

Whittaker:1943:ANEa

[Whi43a] E. T. Whittaker. Aristotle, Newton, Einstein. Science, 98 (2542):249–254, September 17, 1943. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/98/2542/249.full.pdf.

Whittaker:1943:ANEb

[Whi43b] E. T. Whittaker. Aristotle, Newton, Einstein. II. Science, 98 (2543):267–270, September 24, 1943. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/98/2543/267.full.pdf.

Whitney:1955:NNS

[Whi55a] Peter D. Whitney. Nine noted scientists urge war ban: Scientists ask abolition of war; Einstein among signers of plea. Warning on nuclear peril was signed by Einstein. New York Times, ?? (??):1, July 1955. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/113191041.

Whittaker:1955:AE

[Whi55b] Edmund Whittaker. Albert Einstein. 1879–1955. Biographical Memoirs of Fellows of the Royal Society, 1(??):37–67, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://links.jstor.org/sici?
sici=0080-4606(195511)1<37:AE1>2.0.CO%3B2-2;   http:
//www.jstor.org/stable/769242.

Whitrow:1967:EMH

[Whi67] G. J. Whitrow. Einstein, the man and his achievement: a series of broadcast talks under the general editorship of G. J. Whitrow. British Broadcasting Corporation, London, UK, 1967. xv + 94 pp.

White:1983:EMF

[Whi83a] Kenneth Steele White. Einstein and modern French drama: an analogy. University Press of America, Washington, DC, USA, 1983. ISBN 0-8191-2942-9, 0-8191-2943-7 (paperback). xvi + 116 pp. LCCN PQ558 .W48 1983.

Whitrow:1983:AES

[Whi83b] G. J. Whitrow. Albert Einstein Special Theory of Relativity — emergence (1905) and early interpretation (1905–1911) A. I. Miller. The British Journal for the Philosophy of Science, 34(1): 78–84, ???? 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

White:1993:ELS

[Whi93] Michael White. Einstein: a life in science. Simon and Schuster, New York, NY, USA, 1993. ISBN 0-671-71170-9. 279 pp. LCCN QC16.E5 W47 1993; QC 16 .E5W47 1993.

White:1994:ELSa

[Whi94] Michael White. Einstein: a life in science. Pocket Books, New York, NY, USA, revised edition, 1994. ISBN 0-671-01044-1 (paperback). viii + 279 pp. UK£96.99.

Whitaker:1996:EBQ

[Whi96] Andrew Whitaker. Einstein, Bohr, and the quantum dilemma. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48220-8 (hardcover), 0-521-48428-6 (paperback). xvii + 349 pp. LCCN QC174.12 .W48 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:
//www.cambridge.org/uk/catalogue/catalogue.asp?isbn=
0521484286;   http://www.loc.gov/catdir/description/
cam027/95018270.html;   http://www.loc.gov/catdir/toc/
cam021/95018270.html.

Whitworth:2001:EWR

[Whi01] Michael H. Whitworth. Einstein’s wake: relativity, metaphor, and modernist literature. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-818640-1. ix + 254 pp. LCCN PR478.S26 W47 2002. URL http://www.loc.

gov/catdir/enhancements/fy0613/2002277601-d.html;
http://www.loc.gov/catdir/toc/fy033/2002277601.html.

Whitaker:2004:EPB

[Whi04a] M. A. B. Whitaker. The EPR paper and Bohr’s response: a reassessment. Foundations of Physics, 34(9):1305–1340, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Whitehead:2004:PRA

[Whi04b] Alfred North Whitehead. The Principle of Relativity with applications to physical science. Dover phoenix editions. Dover, New York, NY, USA, 2004. ISBN 0-486-43888-0. xii + 190 pp. LCCN QC173.55 .W55 2004. URL http://www.loc.gov/

catdir/enhancements/fy0618/2004050197-d.html.

Whitbeck:2005:AEM

[Whi05] C. Whitbeck. Albert Einstein to Michele Besso — reply. Physics Today, 58(7):16–17, July 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitaker:2006:EBQ

[Whi06] Andrew Whitaker. Einstein, Bohr and the Quantum Dilemma: From Quantum Theory to Quantum Information. Cambridge University Press, Cambridge, UK, second edition, 2006. ISBN 0-521-67102-7. xviii + 461 pp. LCCN QC174.12 .W48 2006.

UK£27.99. URL http://www.cambridge.org/uk/catalogue/
catalogue.asp?isbn=0521671027.

Whitaker:2016:JSB

[Whi16] Andrew Whitaker. John Stewart Bell and twentieth century physics: vision and integrity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-874299-1. xvi + 460 pp. LCCN QC16.B442 W45 2016.

Wildermuth:2005:BEC

[WHL^{+}05] Stephan Wildermuth, Sebastian Hofferberth, Igor Lesanovsky, Elmar Haller, L. Mauritz Andersson, Sönke Groth, Israel Bar-Joseph, Peter Krüger, Jörg Schmiedmayer, et al. Bose–Einstein condensates: Microscopic magnetic-field imaging. Nature, 435 (7041):440, May 25, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v435/n7041/full/435440a.html.

Wick:1995:IBS

[Wic95] David Wick. The infamous boundary: seven decades of controversy in quantum physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3785-0, 3-7643-3785-0. xii + 244 pp. LCCN QC173.98 .W53 1995.

Widmalm:1995:SNN

[Wid95] Sven Widmalm. Science and neutrality: The Nobel prizes of 1919 and scientific internationalism in Sweden. Minerva [Washington, DC], 33(4):339–360, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096517.

Widmalm:2007:BRA

[Wid07] Sven Widmalm. Book review: Aant Elzinga: Einstein’s Nobel Prize: A Glimpse behind Closed Doors: The Archival Evidence. Isis, 98(3):644–645, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/524259.

Wien:1921:RSP

[Wie21] Wilhelm Wien. Die Relativitätstheorie vom Standpunkt der Physik und Erkenntnislehre. (German) [The Theory of Relativity from the standpoint of physics and epistemology]. Johann Ambrosius Barth, Berlin, Germany, 1921. 36 pp. LCCN ????

Wiener:1929:DEE

[Wie29] Norbert Wiener. Dirac equations and Einstein theory. Nature, 123(3112):944–945, June 22, 1929. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v123/n3112/pdf/
123944c0.pdf.

Wiechert:1992:MRV

[Wie92] Emil Wiechert. Mach’s relativity vs. Einstein’s relativity. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92], chapter 8, pages 165–182. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN

C5.53; QC6.9. URL https://link.springer.com/chapter/
10.1007/978-94-011-2771-4_8.

Wiederkehr:2006:PEL

[Wie06] Karl Heinrich Wiederkehr. Photoeffekte, Einsteins Lichtquanten und die Geschichte ihrer Akzeptanz: Zum 125 jährigen Geburtstag Albert Einsteins. (German) [Photoelectric effect, Einstein’s light quanta, and the history of their acceptance: On the 125th birthday of Albert Einstein]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 90(2):132–142, ???? 2006. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.org/

stable/20778025.

Wigner:1949:AES

[Wig49] E. P. Wigner. Albert Einstein seventy. Physics Today, 2 (7):24–??, July 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v2/i7/p24/s1.

Wigner:1964:ENI

[Wig64a] Eugene P. Wigner. Ereignisse, Naturgesetze und Invarianzprinzipien: Nobel-Vortrag am 12. Dezember 1963. (German) [Events, laws of nature, and invariance principles: Nobel Lecture, 12 December 1963]. Angewandte Chemie, 76(17), September 7, 1964. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Wigner:1964:ELNa

[Wig64b] Eugene P. Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. In Anonymous, editor, Les Prix Nobel en 1963, pages 120–132. Nobel Foundation, Stockholm, Sweden, 1964. LCCN ????

Wigner:1964:ELNb

[Wig64c] Eugene P. Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. Science, 145(3636): 995–999, September 4, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1714068;   http://www.sciencemag.org/content/
145/3636/995.full.pdf.

Wigner:1965:ETI

[Wig65] Wigner Jenő. Események, természettörvények és invarianciaelvek: 1963. évi nobel-előadás. (Hungarian) [Events, laws of nature and invariance: 1963 annual Nobel Lecture]. Fizikai Szemle (Budapest), 15(1):1–6, ???? 1965. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1967:ELN

[Wig67a] Eugene Wigner. Events, laws of nature, and invariance principles: Nobel lecture, December 12, 1963. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67b], pages 38–50. LCCN Q171 .W65.

Wigner:1967:SRS

[Wig67b] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Wigner:1979:EIR

[Wig79] Eugene P. Wigner. Einstein’s ideals: Review of Albert Einstein, The human side: New glimpses from his archives, Selected and edited by Helen Dukas and Banesh Hoffman. Pp. 168. Princeton University, 1979. Nature, 282(5735):179–180, November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v282/n5735/pdf/282179a0.pdf.

Wigner:1980:TYK

[Wig80] Eugene P. Wigner. Thirty years of knowing Einstein. In Woolf [Woo80b], pages 461–468. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Wigner:1982:FSN

[Wig82] Eugene P. Wigner. Family, science, Nobel physicists gift to world. Press Republican, 3(??):??, May 15, 1982. CODEN ???? ISSN ????

Wiggan:1986:OFR

[Wig86] Richard Wiggan. Operation Freshman: the Rjukan heavy water raid, 1942. W. Kimber, London, UK, 1986. ISBN 0-7183-0571-X. 176 + 16 pp. LCCN D794.5 .W53 1986.

Wigner:1996:ADM

[Wig96a] E. P. Wigner. Address delivered to the Memorial Meeting [for Paul Dirac]. In Wigner [Wig96m], pages 214–215. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:ABE

[Wig96b] E. P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Wigner [Wig96m], pages 133–138. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:BJN

[Wig96c] E. P. Wigner. Biography of John von Neumann. In Wigner [Wig96m], pages 209–210. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CRA

[Wig96d] E. P. Wigner. Concluding remarks (address at the Dirac Symposium). In Wigner [Wig96m], pages 195–196. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EM

[Wig96e] E. P. Wigner. Einstein — a memoir. In Wigner [Wig96m], page 197. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EF

[Wig96f] E. P. Wigner. Enrico Fermi (1901–1954). In Wigner [Wig96m], pages 115–119. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EAE

[Wig96g] E. P. Wigner. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. In Wigner [Wig96m], pages 198– 200. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:JN

[Wig96h] E. P. Wigner. John von Neumann (1903–1957). In Wigner [Wig96m], pages 127–130. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:LS

[Wig96i] E. P. Wigner. Leo Szilard (1898–1964). In Wigner [Wig96m], pages 139–149. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:NLE

[Wig96j] E. P. Wigner. New light on Einstein letter — an interview with E. P. Wigner and A. M. Weinberg by D. Sundberg. In Wigner [Wig96m], pages 216–218. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:RPD

[Wig96k] E. P. Wigner. Remembering Paul Dirac. In Wigner [Wig96m], pages 219–230. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:TYK

[Wig96l] E. P. Wigner. Thirty years of knowing Albert Einstein. In Wigner [Wig96m], pages 201–208. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96m] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Willers:1936:BWHb

[Wil36] R. A. Willers. Buchbesprechung: W. Heisenberg, E. Schrödinger, P. A. M. Dirac, Die moderne Atomtheorie. Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. 45 S. u. 6 Fig. Leipzig 1934, Verlag von S. Hirzel. Preis kart. 2,50 M. Zeitschrift für Angewandte Mathematik und Mechanik, 16(2):126, 1936. CODEN ZAMMAX. ISSN 1521-4001.

Wilson:1946:RUF

[Wil46] Robert R. Wilson. Radiological use of fast protons. Radiology, 47(5):487–491, November 1946. CODEN RADLAX. ISSN 0033-8419 (print), 1527-1315 (electronic). URL http://pubs.rsna.

org/doi/10.1148/47.5.487.

Wilson:1969:HNP

[Wil69] Mitchell Wilson. How Nobel prizewinners get that way. Atlantic Monthly, ??(??):69–74, December 1969. ISSN 1072-7825 (print),

2151-9463 (electronic). URL http://www.theatlantic.com/
author/mitchell-wilson/.

Will:1972:EFL

[Wil72] Clifford M. Will. Einstein on the firing line. Physics Today, 25 (10):23–29, October 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Williams:1974:ANP

[Wil74] Trevor Illtyd Williams. Alfred Nobel pioneer of high explosives. Pioneers of science and discovery. Priory Press, London, UK, 1974. ISBN 0-85078-128-0. 96 pp. LCCN TP268.5.N7.

Wilson:1975:AOT

[Wil75] Jane Wilson, editor. All in our time: the reminiscences of twelve nuclear pioneers. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1975. 236 pp. LCCN QC773.A1 A44. Reprinted by the Educational Foundations for Nuclear Science, Chicago, IL, USA (1975).

Williamson:1977:LEE

[Wil77] Robert B. Williamson. Logical economy in Einstein’s “On the Electrodynamics of Moving Bodies”. Studies in History and Philosophy of Science Part A, 8(1):49–60, March 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900188.

Wilson:1979:EQM

[Wil79] E. Bright Wilson. Einstein and quantum mechanics. International Journal of Quantum Chemistry, 16(S13):1–4, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Will:1981:TEG

[Wil81] Clifford M. Will. Theory and experiment in gravitational physics. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-23237-6 (hardcover), 0-521-31710-X (paperback). x + 342 pp. LCCN QC178 .W47. URL http://www.loc.

gov/catdir/description/cam031/80039642.html;   http:
//www.loc.gov/catdir/enhancements/fy1106/80039642-t.
html.

Wilson:1983:RSG

[Wil83] David Wilson. Rutherford, simple genius. Hodder and Stoughton, London, UK, 1983. ISBN 0-340-23805-4. 638 + 8 pp. LCCN QC16.R8 W54 1983. US$14.95. URL http://catalog.

hathitrust.org/api/volumes/oclc/10560773.html.

Williams:1984:AEH

[Wil84] L. P. Williams. Albert Einstein, historical and cultural perspectives — the Centennial Symposium in Jerusalem — G. Holton, Y. Elkana. Technology and Culture, 25(3):699–700, ???? 1984. CODEN TECUA3. ISSN 0040-165x (print), 1097-3729 (electronic).

Will:1986:WER

[Wil86] Clifford M. Will. Was Einstein Right?: Putting General Relativity to the Test. Basic Books, New York, NY, USA, 1986. ISBN 0-465-09088-5. xii + 274 + 8 pp. LCCN QC173.6 .W55 1986. US$18.95.

Williams:1987:KFA

[Wil87] Robert Chadwell Williams. Klaus Fuchs, atom spy. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-50507-7. vi + 267 + 10 pp. LCCN UB271.R9 F838 1987.

Will:1988:HCJ

[Wil88] Clifford M. Will. Henry Cavendish, Johann von Soldner, and the deflection of light. American Journal of Physics, 56(5):413– 415, May 1988. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Wilson:1989:AKN

[Wil89] Curtis Wilson. Astronomy from Kepler to Newton: historical studies, volume 290 of Collected studies series. Variorum Reprints, London, UK, 1989. ISBN 0-86078-238-7. x + 292 pp. LCCN QB43.2 .W55 1989. URL http://www.gbv.de/dms/

bowker/toc/9780860782384.pdf.

Will:1990:GRH

[Wil90a] Clifford M. Will. General Relativity at 75: How right was Einstein? Science, 250(4982):770–776, November 9, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/250/
4982/770.full.pdf.

Will:1990:WER

[Wil90b] Clifford M. Will. Was Einstein Right?: Putting General Relativity to the Test. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-282203-9 (paperback). xii + 274 pp. LCCN QC173.6 .W55 1990.

Will:1993:GCC

[Wil93a] Clifford M. Will. Is the gravitational constant constant? In Was Einstein Right?: Putting General Relativity to the Test [Wil93c], chapter 9, pages 160–180. ISBN 0-465-09086-9 (paperback). LCCN QC173.6 .W55 1993. URL http://www.loc.gov/

catdir/enhancements/fy0832/92053250-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0832/92053250-d.
html.

Will:1993:TEG

[Wil93b] Clifford M. Will. Theory and experiment in gravitational physics. Cambridge University Press, Cambridge, UK, revised edition, 1993. ISBN 0-521-43973-6 (paperback). xvi + 380 pp. LCCN QC178 .W47 1993. URL http://www.loc.

gov/catdir/description/cam025/92029555.html;   http:
//www.loc.gov/catdir/toc/cam028/92029555.html.

Will:1993:WER

[Wil93c] Clifford M. Will. Was Einstein Right?: Putting General Relativity to the Test. Basic Books, New York, NY, USA, second edition, 1993. ISBN 0-465-09086-9 (paperback). xii + 290 pp. LCCN QC173.6 .W55 1993. URL http://www.loc.gov/

catdir/enhancements/fy0832/92053250-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0832/92053250-d.
html.

Wilson:1996:HSS

[Wil96] Robert R. Wilson. Hiroshima: The scientists’ social and political reaction. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 140(3):350– 357, ???? 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Will:2000:ERE

[Wil00] Clifford M. Will. Einstein’s relativity and everyday life.

Web site, 2000. URL http://physicscentral.com/explore/
writers/will.cfm.

Will:2001:CBG

[Wil01] Clifford M. Will. The confrontation between General Relativity and experiment. Living Reviews in Relativity, 4:4–97, 2001. CODEN ???? ISSN 1433-8351. URL http://relativity.

livingreviews.org/Articles/lrr-2001-4/.

Wilczek:2005:NLA

[Wil05a] Frank Wilczek. Nobel Lecture: Asymptotic freedom: From paradox to paradigm. Reviews of Modern Physics, 77(3):857– 870, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.77.857;   http://rmp.aps.
org/abstract/RMP/v77/i3/p857_1.

Will:2005:RC

[Wil05b] Clifford M. Will. Relativity at the centenary. Physics World, 18(1):27–32, January 2005. CODEN PHWOEW.

ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a27.
pdf;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/1993/.

Will:2006:CBG

[Wil06] Clifford M. Will. The confrontation between General Relativity and experiment. Living Reviews in Relativity, 9:3–100, 2006. CODEN ???? ISSN 1433-8351. URL http://relativity.

livingreviews.org/Articles/lrr-2006-3/.

Williams:2011:MHB

[Wil11a] Hill Williams. Made in Hanford: the bomb that changed the world. Washington State University Press, Pullman, WA, USA, 2011. ISBN 0-87422-307-5. xvi + 190 pp. LCCN QC773.3.U5 W55 2011.

Wilson:2011:ERB

[Wil11b] R. Mark Wilson. Experiments reveal a Bose–Einstein condensate of photons. Physics Today, 64(2):10–11, February 1, 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.3554306.

Wilford:2013:BBD

[Wil13a] John Noble Wilford. Big Bang’s defenders weigh “fudge factor,” a blunder of Einstein’s, as fix for new crisis. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Williams:2013:MNJ

[Wil13b] Jody Williams. My Name Is Jody Williams: a Vermont Girl’s Winding Path to the Nobel Peace Prize. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 2013. ISBN 0-520-27025-8 (print), 0-520-27025-8. LCCN ???? Foreword by Eve Ensler.

Will:2014:CBG

[Wil14] Clifford M. Will. The confrontation between general relativity and experiment. Living Reviews in Relativity, 17(1):1–117, June 2014. ISSN 1433-8351.

Wiley:2016:BRP

[Wil16a] Daniel Wiley. Book review: The Physicist and the Philosopher: Einstein, Bergson, and the Debate that Changed Our Understanding of Time, by Jimena Canales. Configurations: a Journal of Literature, Science, and Technology, 24(3):406–408, Summer 2016. CODEN ???? ISSN 1063-1801 (print), 1080-6520 (electronic).

Williams:2016:SCA

[Wil16b] Susan Williams. Spies in the Congo: America’s atomic mission in World War II. Public Affairs, New York, NY, USa, 2016. ISBN 1-61039-654-5 (hardcover), 1-61039-655-3 (ebook). xxiv + 369 + 32 pp. LCCN ???? URL https:/

/www.theguardian.com/books/2016/sep/17/spies-in-the-
congo-susan-williams-review.

Will:2017:DEG

[Wil17] Clifford M. Will. Did Einstein get it right? A centennial assessment. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 161(1):18–30, March 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Will.pdf.

Winau:1969:BAE

[Win69] Rolf Winau. Buchbesprechungen: Albert Einstein/Arnold Sommerfeld Briefwechsel. Sechzig Briefe aus dem goldenen Zeitalter der modernen Physik. Herausgegeben und kommentiert von Armin Hermann. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 53(1):106, ???? 1969. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.org/

stable/20775744.

Winkler:1993:LUC

[Win93] Allan M. Winkler. Life under a cloud: American anxiety about the atom. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-507821-7. 282 pp. LCCN UA23 W485 1993.

Winkler:1999:LUC

[Win99] Allan M. Winkler. Life under a cloud: American anxiety about the atom. University of Illinois Press, Urbana, IL, USA, 1999.

ISBN 0-252-06773-8 (paperback). x + 290 pp. LCCN UA23 .W485 1999.

Wineland:2013:NLS

[Win13a] David J. Wineland. Nobel Lecture: Superposition, entanglement, and raising Schrödinger’s cat. Reviews of Modern Physics, 85(3):1103–1114, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1103;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1103_1.

Wineland:2013:SER

[Win13b] David J. Wineland. Superposition, entanglement, and raising Schrödinger’s cat. Annalen der Physik (Berlin), 525(10–11): 739–752, 2013. CODEN ANPYA2. ISSN 0003-3804 (print),

1521-3889 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/andp.201300736/abstract.

Wires:2008:PNP

[Wir08] Richard Wires. The politics of the Nobel Prize in literature: how the laureates were selected, 1901–2007. Edwin Mellen Press, Lewiston, NY, USA, 2008. ISBN 0-7734-4957-4. v + 269 pp. LCCN PN171.P75 W57 2008.

Wisseroth:1979:MWI

[Wis79] K. Wisseroth. Matter — what is it really — 100th anniversary of Albert Einstein March 14, 1979. Chemiker-Zeitung, 103(3): 83–98, ???? 1979. ISSN 0009-2894.

Wise:1982:MLB

[Wis82] M. Norton Wise. The Maxwell literature and British dynamical theory. Historical Studies in the Physical Sciences, 13(1):175– 205, ???? 1982. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757511.

Wise:2021:BRM

[Wis21] M. Norton Wise. Book review: Michael D. Gordin. Einstein in Bohemia. Isis, 112(1):203–204, ???? 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Witten:1962:GIC

[Wit62] Louis Witten, editor. Gravitation: an introduction to current research. Wiley, New York, NY, USA, 1962. x + 481 pp. LCCN QC178 .W829; QC178 .W58; QC178 .W5; QC178 .W78g.

Witkop:1982:PEH

[Wit82] Bernhard Witkop. Paul Ehrlich: His ideas and his legacy. In Bernhard et al. [BCS82], pages 146–166. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500153.

Witten:2005:EEF

[Wit05] Edward Witten. Einstein, Egypt and former glories. Physics World, 18(7):16, July 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/7/phwv18i7a23.
pdf.

Wrinch:1921:RBG

[WJ21] Dorothy Wrinch and Harold Jeffreys. The relation between geometry and Einstein’s theory of gravitation. Nature, 106(2677): 806–809, February 17, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v106/n2677/pdf/106806a0.pdf.

Wilson:2012:MLA

[WK12] John D. Wilson and Steven G. Kellman, editors. Magill’s literary annual, 2012: essay-reviews of 200 outstanding books published in the United States during 2011, with an annotated list of titles. Salem Press, Pasadena, CA, USA, 2012. ISBN 1-58765-967-0 (e-book). ???? pp. LCCN Z1035.A1. URL http:/

/literature.salempress.com/doi/book/10.3331/MLA2012.

Witelson:1999:EBAb

[WKH99a] S. F. Witelson, D. L. Kigar, and T. Harvey. The exceptional brain of Albert Einstein. Society for Neuroscience Abstracts, 25 (1–2):1140, ???? 1999. ISSN 0190-5295.

Witelson:1999:AEB

[WKH99b] Sandra F. Witelson, Debra L. Kigar, and Thomas Harvey. Albert Einstein’s brain. The Lancet (London, England), 354 (9192):1822, November 20, 1999. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547X (electronic).

Witelson:1999:EBAa

[WKH99c] Sandra F. Witelson, Debra L. Kigar, and Thomas Harvey. The exceptional brain of Albert Einstein. The Lancet (London, England), 353(9170):2149–2153, June 19, 1999. CODEN LANCAO.

ISSN 0140-6736 (print), 1474-547X (electronic). See erratum in Lancet, 1999 Jul 17, 354(9174):258.

Wilson:1912:STM

[WL12] E. B. Wilson and G. N. Lewis. The space–time manifold of relativity. Proc. Amer. Soc. Arts & Science, XLVIII(11):??, 1912.

Weber:1980:PSN

[WL80] Robert L. Weber and John Lenihan, editors. Pioneers of science: Nobel Prize winners in physics. The Institute of Physics, Bristol, UK, 1980. ISBN 0-85498-036-9. xviii + 272 pp. LCCN ????

Wang:2012:IFI

[WL12] Weize Wang and Xinwang Liu. Intuitionistic fuzzy information aggregation using Einstein operations. IEEE Transactions on Fuzzy Systems, 20(5):923–938, October 2012. CODEN IEFSEV. ISSN 1063-6706 (print), 1941-0034 (electronic). URL https://

dblp.org/db/journals/tfs/tfs20.html#WangL12.

Wallis:2008:BRA

[WM08] Max K. Wallis and Trevor W. Marshall. Book review: Aant Elzinga, Einstein’s Nobel Prize: A Glimpse behind Closed Doors. Sagamore Beach, MA: Science History Publications/ USA, 2006. Pp xii + 128. ISBN 0-88135-283-7. $39.95 (hardcover). British Journal for the History of Science, 41(1):148– 149, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weiler:2008:SVF

[WNS^{+}08] Chad N. Weiler, Tyler W. Neely, David R. Scherer, Ashton S. Bradley, Matthew J. Davis, Brian P. Anderson, et al. Spontaneous vortices in the formation of Bose–Einstein condensates. Nature, 455(7215):948–951, October 16, 2008. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v455/n7215/
full/nature07334.html.

Wofford:2008:HEP

[Wof08] Thomas Wofford. Hertz, Einstein, and the photoelectric effect. Physics Today, 61(5):10, May 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v61/i5/p10.

Wohlberg:1976:ARA

[Woh76] Margaret Wohlberg. A Los Alamos reader: a Los Alamos bicentennial presentation. Los Alamos County Museum of History, Los Alamos, NM, USA, 1976. 64 pp. LCCN F804.L6 W64 1976.

Woit:2006:EWF

[Woi06] Peter Woit. Not even wrong: the failure of string theory and the search for unity in physical law. Basic Books, New York, NY, USA, 2006. ISBN 0-465-09275-6. xix + 291 pp. LCCN

QC794.6.S85 W65 2006. URL http://www.loc.gov/catdir/
toc/ecip0613/2006013933.html. See critical review [Ber06b].

Wolk:1962:SPB

[Wol62] Herman S. Wolk. Scientists, politics, and the bomb. Air Force magazine, 45(10):44–47, October 1962. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://

www.airforce-magazine.com/MagazineArchive/Pages/1962/
October%201962/1062bomb.aspx.

Wolff:1967:TFY

[Wol67] Anthony Wolff. Twenty-five years with the bomb: A conversation with Nobel Prize-winner Eugene P. Wigner. Look Magazine, ??(??):58, 60–61, December 26, 1967. URL http://library.

ucsd.edu/dc/object/bb5564714p.

Wolters:1987:MMI

[Wol87] Gereon Wolters. Mach I, Mach II, Einstein und die Relativitätstheorie: eine Fälschung und ihre Folgen. (German) [Mach I, Mach II, Einstein and the Theory of Relativity: a falsification and its consequences]. Walter de Gruyter, Berlin, Germany, 1987. ISBN 3-11-010825-9. xv + 474 pp. LCCN Q143.M22 W65x 1987. URL 01; http://www.

zentralblatt-math.org/zmath/en/search/?an=0643.01029.

Wolfson:1993:NCC

[Wol93] Richard Wolfson. Nuclear choices: a citizen’s guide to nuclear technology. New liberal arts series. MIT Press, Cambridge, MA, USA, revised edition, 1993. ISBN 0-262-73108-8. xv + 467 pp. LCCN TK9145 .W59 1993.

Wolfson:1995:ERQ

[Wol95] Richard Wolfson. Einstein’s relativity and the quantum revolution: Modern physics for non-scientists. The SuperStar teachers series. Teaching Company, Springfield, VA, USA, 1995. LCCN QC173.65 .R64 1995. 4 videocassettes (720 minutes).

Wolke:2000:WET

[Wol00] Robert L. Wolke. What Einstein told his barber: more scientific answers to everyday questions. Dell Publishing, New York, NY, USA, 2000. ISBN 0-440-50879-7. xviii + 269 pp. LCCN Q173 .W7875 2000.

Wolfson:2003:SER

[Wol03] Richard Wolfson. Simply Einstein: Relativity demystified. W. W. Norton & Co., New York, NY, USA, 2003. ISBN 0-393-05154-4 (hardcover). x + 261 pp. LCCN QC173.57

.W65 2003. URL http://www.loc.gov/catdir/toc/fy035/
2002002984.html.

Wolk:2009:MHB

[Wol09] Herman S. Wolk. Making the H-bomb. Air Force magazine, 95 (3):66–69, March 2009. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.airforce-

magazine.com/MagazineArchive/Pages/2009/March%202009/
0309H-Bomb.aspx.

Wolff:2016:NPP

[Wol16] Barbara Wolff. The Nobel Prize in Physics 1921 — what happened to the prize money? Web site, March 2016. URL http://

www.einstein-website.de/z_information/nobelprizemoney.
html.

Wong:2005:WMT

[Won05] Samuel Y. S. Wong. Are we more than the sum of our parts? Should we listen to Albert Einstein? Australian family physician, 34(4):291–292, April 2005. ISSN 0300-8495.

Wu:1925:CEC

[Woo25] Yui Hsun Woo. The Compton effect. Thesis, University of Chicago, Chicago, IL, USA, 1925. 17 pp.

Woolard:1928:RPR

[Woo28] Edgar W. Woolard. Recent publications: Reviews: Investigations on the Theory of the Brownian Movement, by Albert Einstein. Edited with notes by R. Fürth. Translated by A. D. Cowper. New York, E. P. Dutton and Co., 1927. viii + 124 pages. American Mathematical Monthly, 35(6):318–320, 1928. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL 2-S;   http://links.jstor.org/sici?sici=
0002-9890(192806/07)35:6<318:IOTTOT>2.0.CO.

Woolf:1980:AEE

[Woo80a] Harry Woolf. Albert Einstein: encounter with America. In Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein [Woo80b], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Woolf:1980:SSP

[Woo80b] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Woodhouse:2003:SR

[Woo03] Nicholas M. J. Woodhouse. Special Relativity. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-426-6, 1-4471-0083-2 (e-book). ???? pp.

LCCN T57-57.97. URL http://d-nb.info/1011126907/34;
http://nbn-resolving.de/urn:nbn:de:1111-20110419318;
 http://www.springerlink.com/content/978-1-4471-0083-
6.

Woodward:2006:BRB

[Woo06] James F. Woodward. Book review: World Without Time: The Forgotten Legacy of Gödel and Einstein. By Palle Yourgrau, Basic Books, New York, New York, USA, 2005, viii + 210 pp., $24 (hard cover). ISBN 0-465-09293-4. Foundations of Physics, 36 (2):321–325, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-9018-8.

Worek:2008:NCP

[Wor08] Michael Worek, editor. Nobel: a century of Prize winners. Firefly Books, Buffalo, NY, USA, 2008. ISBN 1-55407-411-8, 1-55407-416-9 (paperback). 320 pp. LCCN ????

Worek:2010:NPS

[Wor10] Michael Worek. The Nobel Prize: the story of Alfred Nobel and the most famous prize in the world. Firefly Books, Richmond Hill, ON, Canada, 2010. ISBN 1-55407-711-7. 79 pp. LCCN 001.44 W67 2010.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Wenger:2004:EF

[WP04] Win Wenger and Richard Poe. The Einstein Factor. Gramercy, New York, NY, USA, 2004. ISBN 0-517-22320-1. 352 (est.) pp. LCCN BF431 .W455 2004.

Wrede:1966:BRA

[Wre66] R. C. Wrede. Book review: Albert Einstein and the Cosmic World Order, by Cornelius Lanczos. vi + 139 pp. 1956 (Wiley). Mathematics Magazine, 39(1):65, January 1966. CODEN MAMGA8. ISSN 0025-570X. URL http://www.jstor.org/

stable/2688999.

Weart:1978:LSH

[WS78] Spencer R. Weart and Gertrud Weiss Szilard, editors. Leo Szilard, his version of the facts: selected recollections and correspondence, volume 2 of The Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-19168-7 (hardcover). xxii + 244 pp. LCCN QC3 .S97 vol. 2; QC16.S95.

Wilson:1988:SMD

[WS88] Jane Wilson and Charlotte Serber, editors. Standing by and making do: women of wartime Los Alamos. Los Alamos Historical Society, Los Alamos, NM, USA, 1988. ISBN 0-941232-08-5 (paperback). xi + 130 pp. LCCN QC773.A1 S82 1988. US$8.95.

Walker:1989:DEE

[WS89] Evan Harris Walker and John Stachl. Did Einstein espouse his spouse’s ideas? Physics Today, 42(2):9, 11, 13, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weyl:2001:HWR

[WS^{+}01] Hermann Weyl, Erhard Scholz, et al., editors. Hermann Weyl’s Raum–Zeit–Materie and a general introduction to his scientific work, volume 30 of DMV Seminar. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2001. ISBN 3-7643-6476-9, 0-8176-6476-9. vi + 403 pp. LCCN QC173.58 .H37 2001.

Wilson:2008:SMD

[WS08] Jane Wilson and Charlotte Serber, editors. Standing by and making do: women of wartime Los Alamos. Los Alamos Historical Society, Los Alamos, NM, USA, second edition, 2008. ISBN 0-941232-08-5 (softcover). 152 pp. LCCN QC773.A1 S82 2008.

Wolper:2000:LIS

[WSG00] David L. Wolper, Roy Scheider, and Robert Guenette. Legends, icons & superstars of the 20th century. Warner Home Video, 2000.

Williams:2009:LLR

[WTB09] James G. Williams, Slava G. Turyshev, and Dale H. Boggs. Lunar laser ranging tests of the Equivalence Principle with the Earth and Moon. International journal of modern physics. D, Gravitation, astrophysics, cosmology, 18(7):1129–1175, July 2009. CODEN IMPDEO. ISSN 0218-2718. URL http://www.

worldscinet.com/ijmpd/18/1807/S021827180901500X.html.

Weisberg:1981:GWO

[WTF81] J. M. Weisberg, J. H. Taylor, and L. A. Fowler. Gravitational waves from an orbiting pulsar. Scientific American, 245(4):66– 74, October 1981. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wuensch:2005:ZWK

[Wue05] Daniela Wuensch. “Zwei wirkliche Kerle”: neues zur Entdeckung der Gravitationsgleichungen der allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert. (German) [“Two real guys”: news of the discovery by Albert Einstein and David Hilbert of the gravitational equations of General Relativity]. Termessos, Göttingen, Germany, 2005. ISBN 3-938016-04-3. 125 pp. LCCN QC173.6 .W84 2005.

Warburg:1918:MPS

[WvLuAE18] Emil Warburg, Max von Laue, and Arnold Sommerfeld und Albert Einstein. Zu Max Plancks sechzigstem Geburtstag: Ansprachen, gehalten am 26. April 1918 in der Deutschen Physikalischen Gesellschaft, nebst einer Erwiderung von M. Planck. C. F. Müllerschen Hofbuchhandlung m.b.H., Karlsruhe, Germany, 1918. ISSN 0005-9021. 36 pp.

Weber:1957:RCG

[WW57] Joseph Weber and John A. Wheeler. Reality of the cylindrical gravitational waves of Einstein and Rosen. Reviews of Modern Physics, 29(3):509–515, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.29.509;
http://rmp.aps.org/abstract/RMP/v29/i3/p509_1.

Winnacker:1975:UWK

[WW75] Karl Winnacker and Karl Wirtz. Das unverstandene Wunder: Kernenergie in Deutschland. (German) [The misunderstood wonder: nuclear energy in Germany]. Econ Verlag, Düsseldorf, Germany, 1975. ISBN 3-430-19792-9. 412 pp. LCCN QC792.78.G3 W56.

Wigner:1986:NLE

[WW86] Eugene P. Wigner and A. M. Weinberg. New light on the Einstein letter. The Oak Ridger, 37(1):7–??, ???? 1986. CODEN ???? ISSN 0890-6009.

Walker:2005:KHA

[WY05] Stephen Walker and Hiroaki Yokoyama. Kauntodaun Hiroshima: 08:15 August 6 1945. Hayakawa Shobō, Tōkyō, Japan, 2005. ISBN 4-15-208654-8. 438 pp. LCCN ????

Wyden:1984:DOB

[Wyd84] Peter Wyden. Day One: Before Hiroshima and After. Simon and Schuster, New York, NY, USA, 1984. ISBN 0-671-46142-7 (hardcover). xxxii + 412 pp. LCCN D767.25.H6 W93 1984.

Wyder:2015:ECN

[Wyd15] Margrit Wyder. Einstein and Co.: Nobelpreisträger in Zürich. (German) [Einstein and company: Nobel Prize winner in Zürich]. NZZ libro. Verlag Neue Zürcher Zeitung, Zürich, Switzerland, 2015. ISBN 3-03810-025-0 (hardcover). 256 pp. LCCN AS911.N9 W93 2015.

Wykstra:1976:ESP

[Wyk76] Steve Wykstra. On Einstein’s Second Postulate. The British Journal for the Philosophy of Science, 27(3):259–261, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/27/3/259.full.pdf+html.

Wheeler:1983:QTM

[WZ83] John Archibald Wheeler and Wojciech Hubert Zurek, editors. Quantum theory and measurement. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08315-0, 0-691-08316-9 (paperback). xxviii + 811 pp. LCCN QC174.125 .Q38 1983.

Xu:2009:AFS

[XMY^{+}09] Limei Xu, Francesco Mallamace, Zhenyu Yan, Francis W. Starr, Sergey V. Buldyrev, H. Eugene Stanley, et al. Appearance of a fractional Stokes–Einstein relation in water and a structural interpretation of its onset. Nature Physics, 5(8):565–569, July 5, 2009. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481

(electronic). URL http://www.nature.com/nphys/journal/
v5/n8/full/nphys1328.html.

Yaglom:1985:EPT

[Yag85] A. M. Yaglom. Einstein’s 1914 paper on the theory of randomly fluctuating series of observations. Problemy Peredachi Informatsii, 21(4):101–107, 1985. CODEN PPDIA5. ISSN 0555-2923.

Yam:2004:EE

[Yam04] Philip Yam. Everyday Einstein. Scientific American, 291 (3):50–55, September 2004. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v291/n3/
pdf/scientificamerican0904-50.pdf.

Yamazaki:2009:NEP

[Yam09] Masakatsu Yamazaki. Nuclear energy in postwar Japan and anti-nuclear movements in the 1950s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):132–145, December 2009. CODEN HISCDU. ISSN 0285-4821.

Yang:1980:EIT

[Yan80] Chen Ning Yang. Einstein’s impact on theoretical physics. Physics Today, 33(6):42–49, June 1980. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

scitation.aip.org/content/aip/magazine/physicstoday/
article/33/6/10.1063/1.2914117.

Yang:1987:BRC

[Yan87] C. N. Yang. Book review: The Collected Papers of Albert Einstein, Vol 1, the Early Years — 1879–1902 — J. Stachel. Nature, 330(6143):30–31, November 5, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Yang:2006:AEO

[Yan06] Chen Ning Yang. Albert Einstein: Opportunity and perception. International Journal of Modern Physics. A, Particles and Fields, Gravitation, Cosmology, 21(15):3031–3038, June 20, 2006. CODEN IMPAEF. ISSN 0217-751X.

Yang:2013:AEO

[Yan13a] Chen Ning Yang. Albert Einstein: Opportunity and perception. In Selected papers II, with commentaries [Yan13b], pages 272–279. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). LCCN QC21.3.

URL http://www.worldscientific.com/doi/abs/10.1142/
9789814449021_0036.

Yang:2013:SPI

[Yan13b] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.
com/worldscibooks/10.1142/8640.

Yang:2018:EPR

[YC18] Ying Yang and Huaixin Cao. Einstein–Podolsky–Rosen steering inequalities and applications. Entropy, 20(9):683, ???? 2018. CODEN ENTRFG. ISSN 1099-4300. URL https://dblp.org/

db/journals/entropy/entropy20.html#YangC18;   https:
//www.wikidata.org/entity/Q103830905.

Yoshioka:2011:TBE

[YCKG11] Kosuke Yoshioka, Eunmi Chae, and Makoto Kuwata-Gonokami. Transition to a Bose–Einstein condensate and relaxation explosion of excitons at sub-Kelvin temperatures open. Nature Communications, 2(5):328, May 31, 2011. CODEN NCAOBW. ISSN 2041-1723 (electronic). URL http://www.nature.com/

ncomms/journal/v2/n5/full/ncomms1335.html.

Yan:2020:ASP

[YCL20] Erjia Yan, Zheng Chen, and Kai Li. Authors’ status and the perceived quality of their work: Measuring citation sentiment change in Nobel articles. Journal of the Association for Information Science and Technology, 71(3):314–324, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yeang:2008:BRE

[Yea08] Chen-Pang Yeang. Book review: Einstein’s Jury: The Race to Test Relativity. University of Toronto Quarterly, 77(1):331– 332, 2008. CODEN ???? ISSN 0042-0247 (print), 1712-5278 (electronic).

Yerxa:2003:ECP

[Yer03] Donald A. Yerxa. Einstein’s clocks, Poincaré’s maps: An interview with Peter Galison, Part I. Historically Speaking, 5(2):5–9, 2003. CODEN ???? ISSN 1941-4188 (print), 1944-6438 (elec-

tronic). URL http://muse.jhu.edu/article/423117/pdf.

Yerxa:2004:ECP

[Yer04] Donald A. Yerxa. Einstein’s clocks, Poincaré’s maps: An interview with Peter Galison, Part II. Historically Speaking, 5(3):7–10, January 2004. CODEN ???? ISSN 1941-4188 (print), 1944-6438 (electronic). URL https://muse.jhu.edu/

article/423421/summary.

Yourgrau:2005:GEF

[YGKK05] Palle Yourgrau, Kurt Geginnen, and Susanne Kuhlmann-Krieg. Gödel, Einstein und die Folgen: Vermächtnis einer ungewöhnlichen Freundschaft. (German) [Gödel, Einstein, and the consequences. Legacy of an unusual friendship]. Beck, München, Germany, 2005. ISBN 3-406-52914-3. 234 pp. LCCN ???? URL http://www.gbv.de/

dms/faz-rez/F0C20050316107281.pdf;   http://www.gbv.
de/dms/hebis-darmstadt/toc/128002069.pdf.

Yilmaz:1975:DEF

[Yil75] Hüseyin Yilmaz. On the “derivation” of Einstein’s field equations. American Journal of Physics, 43(4):319–322, 1975.

CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/43/4/10.1119/1.10080.

Ying:2015:NEG

[Yin15] Leong Ying. From Newton, Einstein, to God: a poetic memoir. Greenleaf Book Group, Austin, TX, USA, 2015. ISBN 1-937110-71-0, 1-937110-72-9 (e-book). ???? pp. LCCN QC774.Y56 A3 2015.

Yamahata:1995:NJP

[YJ95] Yosuke Yamahata and Rupert Jenkins, editors. Nagasaki journey: the photographs of Yosuke Yamahata, August 10, 1945. Pomegranate Artbooks, San Francisco, CA, USA, 1995. ISBN 0-87654-360-3 (paperback). 127 pp. LCCN D767.25.N3 Y356 1995.

Yau:2010:SIS

[YN10] Shing-Tung Yau and Steven J. Nadis. The shape of inner space: string theory and the geometry of the universe’s hidden dimensions. Basic Books, New York, NY, USA, 2010. ISBN 0-465-02023-2. xix + 377 pp. LCCN QA691 .Y38 2010.

Yau:2011:STG

[YN11] Shing-Tung Yau and Steve Nadis. String theory and the geometry of the universe’s hidden dimensions. Notices of the American Mathematical Society, 58(8):1067–1076, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/
201108/rtx110801067p.pdf.

York:1970:ROP

[Yor70] Herbert F. (Herbert Frank) York. Race to oblivion; a participant’s view of the arms race. Simon and Schuster, New York, NY, USA, 1970. ISBN 0-671-20610-9. 256 pp.

LCCN UA23 .Y67 1970. URL http://www.learnworld.
com/ZNW/LWText.York.RaceToOblivion.html;   https://
en.wikipedia.org/wiki/Herbert_York.

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67 1989. URL http://www.loc.

gov/catdir/description/cam023/88062671.html;   http:
//www.loc.gov/catdir/toc/cam027/88062671.html.

York:1995:AP

[Yor95] Herbert F. (Herbert Frank) York. Arms and the physicist, volume 12 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-099-0. xiii + 294 pp. LCCN QC16.Y67 A3 1995.

Yost:1995:BPK

[Yos95] Res Yost. Boltzmann und Planck: Die Krise des Atomismus um die Jahrhundertwende und ihre Überwindung durch Einstein. (German) [Boltzmann and Planck: The crisis of atomism at the turn of the century and its surmounting by Einstein]. In Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers] [Jos95b], pages 35–51. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. LCCN QC16.J67 J67 1995. URL

http://www.springerlink.com/openurl.asp?genre=issue\
%26issn=0940-7677\%26volume=34. Edited posthumously by

Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Yourgrau:1999:GME

[YotY99] Palle Yourgrau and Palle. Disappearance of time Yourgrau. Gödel meets Einstein: time travel in the Gödel universe. Open Court, Chicago, IL, USA, 1999. ISBN 0-8126-9408-2. xxiv + 253 pp. LCCN BD638 .Y68 1999.

Yourgrau:1979:EVA

[You79] Wolfgang Yourgrau. Einstein— and the vanity of academia. In Aichelburg and Sexl [AS79], pages 213–220. CODEN LNCSD9. ISBN 3-528-08425-1 (print), 3-322-91080-6 (e-book). LCCN

QC16.E5 A66. DM48.00. URL http://link.springer.com/
chapter/10.1007/978-3-322-91080-6_16/.

Yourgrau:1991:DTK

[You91] Palle Yourgrau. The disappearance of time: Kurt Gödel and the idealistic tradition in philosophy. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41012-6 (hardcover). x + 182 pp. LCCN BD638 .Y68 1991. URL http://www.loc.gov/

catdir/description/cam024/91006646.html.

Young:1995:BRB

[You95] Marilyn Young. Book review: The Decision to Use the Atomic Bomb and the Architecture of an American Myth by Gar Alperovitz, Sanho Tree, Edward Rouse Winstead, Kathryn C. Morris, David J. Williams, Leo C. Maley III, Thad Williamson, Miranda Grieder. The American Historical Review, 100(5):1515– 1516, December 1995. CODEN ???? ISSN 0002-8762 (print),

1937-5239 (electronic). URL http://www.jstor.org/stable/
2169869.

Yourgrau:1998:CDG

[You98] Palle Yourgrau. Comments on “Did Gödel Surprise Einstein with a Rotating Universe and Time Travel?” by Giora Hon. Foundations of Physics, 28(11):1719–1727, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1018849827437. See [Hon96].

Yourgrau:2004:GEF

[You04] Palle Yourgrau. Gödel and Einstein: Friendship and Relativity. The Chronicle of Higher Education, ??(??):??, 2004. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://chronicle.

com/article/G-delEinstein-/4363/.

Yourgrau:2005:WTF

[You05] Palle Yourgrau. A World Without Time: the Forgotten Legacy of Gödel and Einstein. Basic Books, New York, NY, USA, 2005. ISBN 0-465-09293-4. viii + 210 + 4 pp. LCCN BD638 .Y73 2005.

Younger:2009:BNH

[You09] Stephen M. Younger. The Bomb: a New History. Ecco, New York, NY, USA, 2009. ISBN 0-06-153719-5. xiii + 238 pp. LCCN QC773.3.U5 Y68 2009.

Young:2011:ERC

[You11] Mervyn O. Young. Einstein’s relativity: a companion to Einstein’s The meaning of relativity. Mervyn O. Young, Millmerran, QLD, Australia, second edition, 2011. ISBN 0-646-56235-5. xviii + 804 pp. LCCN ????

Yukawa:1949:MMM

[Yuk49] Hideki Yukawa. Models and methods in the meson theory. Reviews of Modern Physics, 21(3):474–479, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.474;   http://rmp.aps.org/

abstract/RMP/v21/i3/p474_1. Special issue in honor of Albert Einstein on his 70th birthday.

Yunes:2015:EST

[Yun15] Nicolás Yunes. Einstein ‘?sigue teniendo razón?. (Spanish) [is Einstein still right?]. arXiv.org, ??(??):1–5, October 13, 2015.

URL https://arxiv.org/abs/1510.03845.

Zacek:1911:PRV

[Žáč11] Augustin Žáček. O principu relativity ve fysice. (Czech) [On the principle of relativity in physics]. Živa, 21(??):135–137, ???? 1911.

Zacek:1912:OEA

[Žáč12] Augustin Žáček. Odvození Einsteinova addičního theoremu pro skládání v případé rychlostí parallelních. (Czech) [Derivation of Einstein’s addition theorem for folding in the case of parallel speeds]. Časopis pro pěstování matematiky a fysiky, 41(??):538– 541, ???? 1912.

Zachary:1997:EFV

[Zac97] G. Pascal Zachary. Endless frontier: Vannevar Bush, engineer of the American Century. Free Press, New York, NY, USA, 1997. ISBN 0-684-82821-9. viii + 518 + 16 pp. LCCN TK140.B87 Z33 1997.

Zachary:1999:EFV

[Zac99a] G. Pascal Zachary. Endless frontier: Vannevar Bush, engineer of the American Century. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-74022-2 (paperback). viii + 518 pp. LCCN TK140.B87 Z33 1999.

Zackheim:1999:EDS

[Zac99b] Michele Zackheim. Einstein’s daughter: the search for Lieserl. Riverhead Books, New York, NY, USA, 1999. ISBN 1-57322-127-9. xiii + 301 pp. LCCN QC16.E5 Z33 1999.

Zahar:1973:WDEa

[Zah73a] Elie Zahar. Why did Einstein’s programme supersede Lorentz’s? (I). The British Journal for the Philosophy of Science, 24(2): 95–123, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/24/2/95.full.pdf+html.

Zahar:1973:WDEb

[Zah73b] Elie Zahar. Why did Einstein’s programme supersede Lorentz’s? (II). The British Journal for the Philosophy of Science, 24 (3):223–262, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/24/3/223.full.pdf+html;
http://links.jstor.org/sici?sici=0007-0882%28197309%
2924%3A3%3C223%3AWDEPSL%3E2.0.CO%3B2-Z.

Zahar:1976:WDE

[Zah76] Elie Zahar. Why did Einstein’s programme supersede Lorentz’s? In Howson [How76], pages 211–276. ISBN 0-521-21110-

7. LCCN Q175 .M5417. URL http://catdir.loc.gov/
catdir/enhancements/fy0906/75044580-d.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy0906/75044580-
t.html;   http://www.gbv.de/dms/bowker/toc/9780521211109.
pdf.

Zahar:1977:MER

[Zah77] Elie Zahar. Mach, Einstein, and the rise of modern science. The British Journal for the Philosophy of Science, 28(3):195–213, September 1977. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/28/3/195.full.pdf+html.

Zahar:1978:EDL

[Zah78] Elie Zahar. Einstein’s debt to Lorentz: a reply to Feyerabend and Miller. The British Journal for the Philosophy of Science, 29(1):49–60, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/29/1/49.full.pdf+html.

Zahar:1980:EMR

[Zah80] Elie Zahar. Einstein, Meyerson and the role of mathematics in physical discovery. The British Journal for the Philosophy of Science, 31(1):1–43, March 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
31/1/1.full.pdf+html;   http://links.jstor.org/sici?
sici=0007-0882%28198003%2931%3A1%3C1%3AEMATRO%3E2.0.
CO%3B2-B.

Zahar:1987:MRD

[Zah87] Elie Zahar. Meyerson’s ‘relativistic deduction’: Einstein versus Hegel. The British Journal for the Philosophy of Science, 38(1): 93–106, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/38/1/93.full.pdf+html.

Zahar:1989:ERS

[Zah89] Elie Zahar. Einstein’s revolution: a study in heuristic. Open Court, La Salle, IL, USA, 1989. ISBN 0-8126-9066-4, 0-8126-9067-2 (paperback). x + 373 pp. LCCN QC173.55 .Z34 1989.

Zaleski:1991:WOJ

[Zal91] C. Zaleski. Out of this world — otherworldly journeys from Gilgamesh to Albert Einstein — I. P. Couliano. Parabola — Myth Tradition and the Search for Meaning, 16(4):94, Winter 1991. ISSN 0362-1596.

Zandy:1993:LGE

[Zan93] Hassan F. Zandy. Letter: Galileo, Einstein, and the church. American Journal of Physics, 61(3):202, March 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://aapt.scitation.org/doi/10.1119/1.17289.

Zaun:2005:JSA

[Zau05] Jörg Zaun. Josef Scharl und Albert Einstein: Geschichte einer Freundschaft. (German) [Josef Scharl and Albert Einstein: History of a Friendship]. ????, ????, 2005. ???? pp. LCCN ????

Zavicka:1925:EUP

[Záv25] František Závička. Einsteinů princip relativnosti a teorie gravitační. (Czech) [Einstein’s principle of relativity and gravitational theory]. Jednoty československých matematiků a fysiků, Prague, Czechoslovakia, 1925. 166 pp.

Zych:2011:QIVa

[ZCPB11a] Magdalena Zych, Fabio Costa, Igor Pikovski, and Časlav Brukner. Quantum interferometric visibility as a witness of general relativistic proper time. arXiv, ??(??):??, May 23,

2011. URL http://arxiv.org/abs/1105.4531v1;   http:/
/www.rdmag.com/News/2011/10/General-Science-Physics-
One-clock-two-times-quantum-mechanics-meets-general-
relativity/.

Zych:2011:QIVb

[ZCPB11b] Magdalena Zych, Fabio Costa, Igor Pikovski, and Časlav Brukner. Quantum interferometric visibility as a witness of general relativistic proper time. Nature Communications, 2(10):??, September 5, 2011. CODEN NCAOBW. ISSN 2041-1723 (elec-

tronic). URL http://www.nature.com/ncomms/journal/v2/
n10/full/ncomms1498.html.

Zhao:2008:EES

[ZDM08] Shuang Zhao, Hans De Raedt, and Kristel Michielsen. Eventby-event simulation of Einstein–Podolsky–Rosen–Bohm experiments. Foundations of Physics, 38(4):322–347, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-008-9205-5.pdf.

Zimmer:2011:CPB

[ZDSF11] Karl Günter Zimmer, Max Delbrück, Phillip R. Sloan, and D. Brandon Fogel, editors. Creating a physical biology: the Three-Man Paper and early molecular biology. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-76782-5 (hardcover), 0-226-76783-3 (paperback). 319 pp. LCCN QH506 .C73 2011.

Zichichi:1991:GMC

[ZdSS91] Antonino Zichichi, Venzo de Sabbata, and Norma Sánchez, editors. Gravitation and modern cosmology: the cosmological constant problem, volume 56 of Ettore Majorana international science series. Physical sciences. Plenum Press, New York, NY, USA; London, UK, 1991. ISBN 0-306-44054-7. xiii + 228 pp.

LCCN QC178 .G63 1991. URL http://www.loc.gov/catdir/
enhancements/fy0820/91040369-t.html.

Zee:1989:OMT

[Zee89] Anthony Zee. An old man’s toy: gravity at work and play in Einstein’s universe. Macmillan Publishing Company, New York, NY, USA, 1989. ISBN 0-02-633440-2. xxxii + 272 pp. LCCN QC178 .Z44 1989.

Zee:1990:OMT

[Zee90] Anthony Zee. An old man’s toy: gravity at work and play in Einstein’s universe. Macmillan Publishing Company, New York, NY, USA, 1990. ISBN 0-02-040915-X. xxxii + 272 pp. LCCN QC178 .Z44 1990.

Zee:2001:EUG

[Zee01] Anthony Zee. Einstein’s universe: gravity at work and play. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514285-3 (paperback). xxxii + 282 pp.

LCCN QC178 .Z44 2001. URL http://www.loc.gov/catdir/
enhancements/fy0610/00068458-d.html.

Zee:2003:QFT

[Zee03] A. Zee. Quantum field theory in a nutshell. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-01019-6. xv + 518 pp. LCCN QC174.45 .Z44 2003. URL http://www.loc.gov/

catdir/description/prin031/2002031743.html;   http:/
/www.loc.gov/catdir/samples/prin031/2002031743.html;
 http://www.loc.gov/catdir/toc/prin032/2002031743.html.

Zee:2010:QFT

[Zee10] Anthony Zee. Quantum field theory in a nutshell. Princeton University Press, Princeton, NJ, USA, second edition, 2010. ISBN 0-691-14034-0. xxvi + 576 pp. LCCN QC174.45 .Z44 2010.

Zee:2013:EGN

[Zee13] Anthony Zee. Einstein Gravity in a Nutshell. In a nutshell. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-14558-X (hardcover). xxii + 866 pp. LCCN QC173.6 .Z44 2013.

Zeh:2007:PBD

[Zeh07] H. D. Zeh. The Physical Basis of the Direction of Time. Springer London, Guildford, Surrey, UK, fifth edition, 2007. ISBN 3-540-68000-4, 3-540-68001-2 (e-book). LCCN QC173.59.S65Z38

2007. URL http://catdir.loc.gov/catdir/toc/fy0714/
2007922924.html;   http://www.springerlink.com/openurl.
asp?genre=book&isbn=978-3-540-68000-0.

Zeilinger:1999:AEP

[Zei99] A. Zeilinger. Albert Einstein: Philosopher-scientist. Nature, 398(6724):210–211, March 18, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Zeilinger:2010:DPE

[Zei10] Anton Zeilinger. Dance of the photons: from Einstein to quantum teleportation. Farrar, Straus and Giroux, New York, NY, USA, 2010. ISBN 0-374-23966-5 (hardcover). 305 pp. LCCN QC174.12 .Z4513 2010.

Zellner:1990:THN

[Zel90] Arnold Zellner. Trygve Haalvelmo, 1989 Nobel Laureate in Economic Science. Chance, 3(1):33–35, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zenneck:2007:G

[Zen07] Jonathan Zenneck. Gravitation. In Renn et al. [RSS^{+}07a], pages 77–112. ISBN 1-4020-3999-9. ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https://link.

springer.com/chapter/10.1007/978-1-4020-4000-9_12.

Zewail:2002:VTT

[Zew02] Ahmed H. Zewail. Voyage through time: walks of life to the Nobel Prize. American University in Cairo Press, Cairo, Egypt, 2002. ISBN 977-424-677-2. xii + 287 pp. LCCN QD22.Z46 A3 2002.

Zhou:1985:AEC

[Zho85] C. Z. Zhou. Albert Einstein and contemporary Western philosophy of science. Chinese Studies in Philosophy, 17(1):70–90, ???? 1985. ISSN 0023-8627.

Ziadat:1994:ERE

[Zia94] Adel A. Ziadat. Early reception of Einstein’s Relativity in the Arab periodical press. Annals of Science, 51(1):17–35, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zichichi:2000:DEP

[Zic00] Antonino Zichichi. Dirac, Einstein and physics. Physics World, 13(3):17–18, March 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://physicsworld.

com/cws/article/print/2000/mar/02/dirac-einstein-and-
physics;   http://physicsworldarchive.iop.org/full/
pwa-pdf/13/3/phwv13i3a17.pdf.

Ziman:1982:SRV

[Zim82] John Ziman. Social responsibility in Victorian science. In Bernhard et al. [BCS82], pages 21–43. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.). URL http://www.

sciencedirect.com/science/article/pii/B9780080279398500074.

Zimmermann:1984:AEV

[Zim84] Rainer E. Zimmermann. Albert Einstein—Versuch einer totalisierenden Würdigung. (German) [Albert Einstein — attempt at a totalizing appreciation]. Philosophia Naturalis, 21(1):126–138, 1984. CODEN ???? ISSN 0031-8027.

Zindel:2001:G

[Zin01] Paul Zindel. The Gadget. HarperCollins, New York, NY, USA and London, UK, 2001. ISBN 0-06-027812-9, 0-06-028255-X. 184 pp. LCCN PZ7.Z647 Gad 2001. URL http://www.loc.

gov/catdir/description/hc041/00038903.html.

Zoellner:2009:UWEa

[Zoe09a] Tom Zoellner. Uranium: war, energy, and the rock that shaped the world. Viking, New York, NY, USA, 2009. ISBN 0-670-02064-8. xii + 337 pp. LCCN QD181.U7

Z64 2009. URL http://www.loc.gov/catdir/enhancements/
fy0906/2008029023-d.html.

Zoellner:2009:UWEb

[Zoe09b] Tom Zoellner. Uranium: War, Energy, and the Rock That Shaped the World. Penguin, New York, NY, USA, 2009. ISBN 0-14-311672-X (paperback). xii + 354 pp. LCCN QD181.U7 Z64 2009.

Zoller:2002:BEC

[Zol02] Peter Zoller. Bose–Einstein condensation: Making it with molecules. Nature, 417(6888):493–494, May 30, 2002. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v417/n6888/
full/417493a.html.

Zipkes:2010:TSI

[ZPSK10] Christoph Zipkes, Stefan Palzer, Carlo Sias, and Michael Köhl. A trapped single ion inside a Bose–Einstein condensate. Nature, 464(7287):388–391, March 18, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v464/n7287/
full/nature08865.html.

Zuckerman:1967:SNP

[Zuc67] Harriet Zuckerman. The sociology of the Nobel Prizes. Scientific American, 217(5):25–33, November 1967. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v217/n5/pdf/scientificamerican1167-25.pdf.

Zuckerman:1977:SEN

[Zuc77] Harriet Ann Zuckerman. Scientific Elite: Nobel Laureates in the United States. Free Press, New York, NY, USA, 1977. ISBN 0-02-935760-8, 0-02-935880-9 (paperback). xv + 335 pp. LCCN Q149.U5 Z8.

Zuckerman:1982:NIR

[Zuc82] Baron Solly Zuckerman. Nuclear illusion and reality. Viking, New York, NY, USA, 1982. ISBN 0-670-51822-0. 154 pp. LCCN UF767 .Z8.

Zuckerman:1983:NIR

[Zuc83] Baron Solly Zuckerman. Nuclear illusion and reality. Vintage Books, New York, NY, USA, 1983. ISBN 0-394-71363-X (paperback). xiv + 154 pp. LCCN UF767 .Z8 1983.

Zuckerman:1996:SEN

[Zuc96] Harriet Zuckerman. Scientific Elite: Nobel Laureates in the United States. Foundations of higher education. Transaction Publishers, New Brunswick, NJ, USA, 1996. ISBN 1-56000-855-5 (paperback). xlv + 335 pp. LCCN Q149.U5 Z8 1996.

Zuckerman:2018:SEN

[Zuc18] Harriet Zuckerman. Scientific Elite: Nobel Laureates in the United States. Routledge, Taylor and Francis Group, London, UK, 2018. ISBN 1-351-30686-3. xv + 335 pp. LCCN Q149.U5

.Z8 1989. URL http://public.ebookcentral.proquest.com/
choice/publicfullrecord.aspx?p=5359479.

Zuoxiu:1996:EPR

[Zuo96] He Zuoxiu. On the Einstein, Podolsky and Rosen Paradox and the relevant philosophical problems. In Cohen et al. [CHR96], pages 299–305. ISBN 94-015-8638-1. ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-015-8638-2_21.

Zweifel:1980:SWA

[Zwe80] P. F. Zweifel. The scientific work of Albert Einstein. Annals of Nuclear Energy, 7(4–5):279–287, ???? 1980. CODEN ANENDJ. ISSN 0306-4549 (print), 1873-2100 (electronic).

Zwicky:1937:MNC

[Zwi37] Fritz Zwicky. On the masses of nebulae and of clusters of nebulae. Astrophysical Journal, 86(3):217–246, October 1937. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Zhen:2019:EPR

[ZXL^{+}19] Yi-Zheng Zhen, Xin-Yu Xu, Li Li, Nai-Le Liu, and Kai Chen. The Einstein–Podolsky–Rosen steering and its certification. Entropy, 21(4):422, ???? 2019. CODEN EN-TRFG. ISSN 1099-4300. URL https://dblp.org/db/

journals/entropy/entropy21.html#ZhenXLLC19;   https:
//www.wikidata.org/entity/Q103833194.

Zhou:2014:LPW

[ZXLX14] Zhiwei Zhou, Rui Xing, Jing Liu, and Feiyue Xing. Landmark papers written by the Nobelists in physics from 1901 to 2012: a bibliometric analysis of their citations and journals. Scientometrics, 100(2):329–338, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1306-7.

Zhang:2023:MST

[ZZ23] Lingche Zhang and Qiuju Zhang. Mapping the scientific and technological landscape: an analysis of Nobel Prize-producing institutions. Scientometrics, 128(11):6129–6145, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL https://link.springer.com/article/10.
1007/s11192-023-04831-1.

Zhang:2019:TSG

[ZZY19] Helena H. Zhang, Alesia A. Zuccala, and Fred Y. Ye. Tracing the ‘swan groups’ of physics and economics in the key publications of Nobel Laureates. Scientometrics, 119(1):425–436, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-019-03036-9.